A Complete Bibliography of Scientific American: 1890–1899

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.06

Title word cross-reference

1/4 [9499]. $4 [1908].

-inch [9499].

1/2 [4561, 10684]. 1/2-Knot [10684]. 10 [1621]. 10-Inch [9992]. 155th [592]. 16 [3572]. 16th [3832]. 1889 [102, 1401]. 1892 [17, 3036, 2593, 2671, 3194]. 1893 [3542, 4138, 4157, 4183, 3743, 3786, 3805, 3823, 3843, 3863, 3885, 3908, 3932, 3964, 3976, 4053, 4079, 4098, 4118, 3587]. 1894 [4971, 5404]. 1895 [6037, 6330]. 1896 [6732, 6837, 6622]. 1897 [6916]. 1898 [8794]. 1899 [10276]. 1900 [6162, 6899, 8124, 9678, 9833, 10624, 10694, 9885].

2 [4117, 7024]. 2-Knot [10684]. 250-Kilowatt [2591].

3 [6444, 8102]. 3.2-Inch [9161]. 31 [209]. 35.2-Knot [9304].

42 [6879].

500-Ton [10260].

8-inch [9499].

9-Inch [137]. ’98 [8858].

A. [4017]. A.A.A.S. [4897, 7829]. A.D. [1401]. Aard [3317]. Aard-Varks [3317]. Abandonment [7655]. Abbey [3850, 9776]. Abdomen [3163]. Abnormal [5007, 10055, 9830]. Abroad [9899]. Abrogated [7919]. Absolutely [7385]. Absorbent [8042]. Absorbing [8176]. Abstracts [5746]. Abutments [10497]. Academy [4497, 5678, 7672, 9954, 10691, 8869, 5620, 9994]. Acceleration [6398]. Accepted [10684]. Accessible [7025]. Accident [358, 478, 979, 1754, 1787, 3364, 4920, 6258, 7024, 7850, 8219, 9179, 9061, 9751]. Accidents [3960, 5511]. Account [2607]. Accountings [7310]. Accuracy [10364]. Accurate [2739, 6922, 10420]. Acetate [2503]. Acetylene [5553, 6285, 6591, 6523, 8338, 7385, 8450, 8696, 8933, 8847, 9155, 9493, 8858, 9233, 9570, 10516, 10649, 9903, 10676, 10599, 10710, 10145, 10348, 10397, 10592, 9845, 5581, 6366]. Acetylene-Apparatus [9570]. Acetylene-Gas [10676]. Acetylene-Generator [10516]. Achievements [1917, 3344]. Acid [2773, 8067, 9113, 4310]. Acoustic [2660, 5133]. Acoustics [1499, 2532]. Acquired [10442]. Acquisition [7705]. Across [8172, 8763, 428, 6201, 7104, 7920, 7981, 10536, 9950, 10545]. Act [7732]. Acting [9777]. Action [642, 130, 2356, 5828, 5725, 6706, 9198, 1424, 10126]. Activity [10299, 9702]. Actually [9115]. Actuated [4708]. Actuating [9823]. Adamson [2401]. Adapted [6347, 1238]. Adding [1747, 4302, 6769]. Additions [3284, 9508]. Adhesion [3849]. Adirondack [7345]. Adjunct [6905]. Adjustable [1040, 1203, 714, 293, 469, 1241, 2181, 1350, 3094, 2866, 2581, 4235, 4967, 5560, 5534, 6020, 6351, 6067, 6870, 7154, 6688, 6499, 8337, 7493, 8363, 9799]. Adjuster [472, 5831, 7639]. Adjusting [1592, 6382]. Admiral [8961, 9163, 9840, 10586, 10240, 9913, 10354, 10585, 9639]. Advance [2984, 7779, 10763]. Advanced [5663]. Advancement [1997, 2774, 5004, 4985, 6018, 6037, 7218, 8035, 10184, 8103, 7706, 871, 9263, 10394]. Advantages [6397]. Advertise [707]. Advertisements [1555, 1580, 1611]. Advertisers [6967]. Advertising [578, 5482, 5798, 9771, 10560]. Advice [8740]. Aerating [1013]. Aerial [182, 1796, 2642, 3134, 3905, 4489, 5968, 5917, 5552, 10608, 2495, 4538]. Aerodromic [4691]. Aerodynamics [2421]. Aeronautics [7756]. Aerophile [7184]. Aeroplane [5169, 10735, 7901]. Affect [10065]. Affected [1987]. Affiliated [6037]. Afghanistan [4498]. Afloat [42, 2278, 7731, 9747]. Africa [4684, 4867, 8345, 7559, 7840, 9505, 10539, 10723, 9436]. African [2291]. After [608, 6613]. Again [4877, 9424, 10255]. Against [387, 2560, 4758, 4077, 5325, 7899, 8356]. Age [8315]. Agency [9771]. Agent [9745]. Ages [2602, 2754, 6221]. Agricultural [958, 3494, 5895, 6948, 9326, 9027]. Agriculture [3944, 5331, 9513]. Aid [5909, 7877, 8927, 9373, 10468, 10668]. Air [133, 345, 595, 1000, 114, 2019, 1298, 1471, 2914, 2722, 2937, 3058, 2801, 4165, 4210, 4029, 3906, 4667, 4545, 4453, 4343, 4654, 6346, 5777, 6728, 6805, 6652, 6803, 7008, 8090, 7434, 7366, 7568, 8401, 9008, 8949, 8626, 9189, 9742, 10144, 10273, 9614, 10047, 9626, 10452, 10094, 9662, 10280, 10092, 10301, 10558, 10115, 10180, 4636, 1424, 10446, 9905, 9928, 8859, 8792, 10234, 10162, 6321]. Air-Bath [9905]. Air-Cushion [10144]. Air-Pipes [10047]. Airship [10735]. Ala [6324, 6286]. Alabama [4539, 8951, 10414, 9690]. Alarm [528, 312, 28, 314, 2201, 2108, 2711, 2813, 3911, 5316, 5327, 6605, 7514, 8173, 9800, 10178, 9930]. Alarm-Matting [9800]. Alarms [559, 2197, 4238]. Alaska [2161, 10796, 2565]. Alaskan [8172, 8058]. Albany [4216, 10611]. Albatross [4909]. Albert [1986]. Alcohol [2117, 7385]. Alessandro [10263]. Alexander [7578]. Alfred [6378, 6406, 8022]. Algebraic [7541]. Algeria [5480]. Alias [4829]. Aligning [5875]. Alkali [4336, 4359]. All-Metal [1518]. Alleged [8034, 1995]. Allegheny [10392]. Alley [4715]. Alligators [4908]. Allis [5044]. Alloy [4506]. Alloys [10232]. Alpine [5603]. Alps [10054]. Altar [2726]. Altels [6199]. Alternating [172, 2528, 8144, 3409]. Alternator [2591]. Aluminum [1146, 1767, 3057, 2660, 4630, 5158, 8475, 7665, 246]. Alvan [7906, 7875]. Amalgamating [4695]. Amalgamator [1697, 4664, 5379, 6583]. Amateur [895, 4182, 10537]. Amateurs [455, 585, 1305, 3389]. Ambulance [7153, 9120]. Amendment [9603]. Amendments [569, 7635, 7572, 7569, 8212, 7611]. America [10583, 496, 981, 235, 1273, 1561, 1725, 3182, 4125, 3636, 4898, 6076, 5563, 6881, 8145, 7376, 9900, 9807, 8694, 10218]. American [2774, 2929, 8194, 871, 100, 582, 129, 1010, 334, 27, 1997, 1447, 1761, 2032, 1294, 1271, 2196, 1321, 1539, 1696, 1292, 1282, 2819, 2335, 3122, 2403, 3124, 3820, 4195, 3429, 3383, 3283, 3998, 3982, 4323, 3540, 5004, 4896, 5024, 4594, 5252, 4952, 4985, 4984, 4550, 5235, 4564, 5219, 5065, 6018, 5395, 5366, 6037, 5519, 5722, 5637, 6057, 7096, 7179, 6675, 6944, 7333, 6956, 7291, 7180, 7060, 6486, 7197, 6459, 8035, 7918, 8358, 8380, 7803, 7895, 8258, 7684, 8121, 7755, 7530, 8084, 7662, 7852, 7797, 8192, 8405, 7764]. American [8575, 9414, 8908, 9326, 9243, 8600, 8823, 9401, 8889, 9446, 9090, 9548, 8627, 8970, 9282, 9368, 8521, 9030, 8781, 8553, 10299, 10695, 10786, 9769, 10637, 10714, 9952, 10203, 10296, 10561, 10343, 10184, 9978, 9775, 10093, 9951, 8103, 8049, 7706, 8998, 3954, 7077, 7098, 9263, 8827, 10394, 707, 187, 213, 261, 281, 303, 327, 351, 432, 507, 630, 652, 679, 699, 439, 800, 835, 513, 797, 831, 801, 836, 485, 492, 516, 1110, 1137, 1224, 1265, 1284, 1356, 1476, 1500, 1526, 1548, 1604, 1628, 1653, 1457, 1560, 1585]. American [1616, 1408, 1454, 1557, 1582, 1613, 1366, 1411, 1317, 1341, 1365, 1387, 1433, 1484, 1510, 1537, 1639, 1664, 1618, 1642, 1617, 2816, 4211, 5798, 6959, 9418, 8582, 8911, 10277, 10436]. American-Built [9414]. Americana [3013]. Amerigo [8749]. Ammonia [658, 2354, 7294]. Ammunition [9294, 10355]. Among [5802]. Amoskeag [2324]. Ampere [158, 1522, 5554]. Amphibious [10030, 8520]. Amphitheater [2169]. Amphitrite [7251]. Amsterdam [7886]. Amstutz [5599]. Anaconda [7386]. Anaglyphs [4776]. Analogies [158, 1522]. Analysis [5848, 9008, 8899, 9090, 8898, 8817]. Anchor [7296]. Ancient [1716, 1475, 7853, 9195, 8654, 10762, 10402, 4863, 5063, 4547, 8894, 9789, 7348, 5480, 7886]. Andiron [2692]. Andrée [7991]. ane [6046]. Angeles [7979]. Angerville [8897]. Angle [5420, 8835]. Angular [8511]. Animal [424, 5629, 8195, 10080, 10256]. Animals [3811, 6219, 5655, 8989, 8874, 10739, 10764, 8641, 8660, 4087, 10617]. Ankle [8174]. Annam [1834]. Annealing [6782, 5859]. Annexation [3325]. Anniversary [6726, 6971, 6951, 9229, 8827]. Annual [174, 27, 1761, 2774, 3343, 4606, 4746, 6037, 8396, 7501, 9447, 8600, 9770, 10272, 9653, 8869, 6436, 7487, 7522, 10118]. Annunciator [2197]. Answered [9894]. Ant [1329]. Ant-Eater [1329]. Antarctic [6217, 8061, 7409, 9820]. Antediluvian [3833]. Antelope [5220, 5786, 6487]. Antelopes [4515]. Anthony [5275]. Anthracite [4299]. Anthropometry [5236]. Anti [43, 6810, 6560, 7716]. Anti-Friction [43, 6810]. Anti-Rattler [7716]. Anti-Rattling [6560]. Antique [6185, 10775, 9832]. Antiquity [8145]. Antiseptic [1969, 10164]. Anton [3508]. Ants [4867, 8266, 8879]. Antwerp [4553, 4827]. Anvil [468]. Any [7789, 9413, 6851]. Ape [2921]. Apennino [8727]. Apparatus [759, 736, 755, 231, 451, 182, 1652, 1248, 2276, 2116, 1415, 1300, 1734, 3107, 3022, 3039, 2931, 4039, 3980, 3228, 3331, 3314, 3957, 4144, 3892, 4472, 5044, 4494, 4524, 5108, 4390, 4384, 4695, 4926, 4759, 6126, 5872, 6238, 6082, 5888, 5666, 6975, 6716, 7026, 6638, 6591, 6906, 7007, 6998, 6652, 6656, 8150, 8447, 7452, 7363, 7976, 8764, 8701, 9352, 8718, 9555, 9391, 9570, 10235, 10651, 10776, 9113, 10693, 1814, 1753, 5581, 6366, 2495, 9616, 694, 1855]. Apparent [5178]. Appeal [8032, 7574]. Appearing [7349]. Appellate [8489, 8377]. Apple [1215]. Appliance [781, 4196, 3329, 10044, 10302]. Appliances [4049, 4076, 8196, 10766]. Application [8354, 6946]. Applications [10615, 7317, 8446]. Applied [994, 1566, 3911, 7271]. Appointment [8573]. Appreciation [10380]. Apprenticeship [6880, 6763, 7428]. Approaches [10462]. Approaching [2882, 5036, 9152]. Appropriation [8802, 8887]. Appropriations [8574, 9772]. Approved [8256, 8489]. April [290, 1621, 1540, 2544, 3472, 3832, 3572, 5591, 7592, 6647, 8786]. Apron [39]. Aqua [5968]. Aquarium [5296]. Aquatic [3134, 8989]. Aqueduct [685, 698, 2265, 694, 1248]. Arbor [3545]. Arboretum [4583]. Arc [582, 2780, 6527, 7789, 361]. Arch [8, 3973, 4469, 8156, 8551, 8831, 10454, 10426, 10260, 9827, 7692, 5350].

Archæological [8345, 7488]. Archeological [6658]. Arches [4191]. Archimedean [6127, 6833]. Architect [4649]. Architects [3541, 9447]. Architectural [830, 1407, 1507]. Architecture [3057, 9655, 9436]. Arctic [3813, 7040]. Arcturus [1826]. Area [8744]. Arena [5010]. Argentine [8295]. Argon [5979]. Argonaut [8516]. Arithmetic [4188]. Arizona [3828, 9331, 2835]. Arles [5010]. Arm [1854, 5218, 7810]. Armadillo [212]. Armadillos [3317]. Armament [9246, 10427]. Armand [8808]. Armature [300, 3148]. Armed [9576]. Armenia [9199, 9034]. Armor [885, 423, 930, 2286, 2810, 2948, 3820, 3974, 4899, 4841, 6225, 7346, 6889, 6648, 7298, 7967, 7664, 8417, 8465, 7718, 10695, 10183, 10634, 10090, 10748, 9817, 5859, 9832, 10508]. Armor-Clad [423, 10748]. Armored [1132, 804, 921, 6152, 7330, 8969, 8906, 9100]. Armory [5273, 9964, 10053]. Armour [3929, 3444]. Arms [1985, 10486]. Armstrong [137, 9867]. Army [404, 2202, 3519, 3315, 6047, 6465, 7278, 7824, 7719, 9385, 9210, 9305, 9069, 9159, 9088, 8930, 9948, 9866, 10610]. Arranged [1177]. Arrangements [9327]. Arranging [2979]. Arrester [1730, 6280, 6477]. Arrival [8324, 10490]. Arrow [1784]. Arsenal [9610, 4745]. Arsenate [10492]. Art [1063, 2985, 4037, 4401, 6299, 5909, 6475, 7673, 8326, 8147, 8857, 9624, 10016, 10065, 8093, 4127]. Artesian [211, 1813]. Arthur [8616]. Article [1210, 4845]. Articles [623, 1150, 2135]. Artificial [886, 1201, 2014, 1688, 2561, 3804, 4144, 3235, 5126, 5700, 5397, 5932, 7023, 8507, 9245, 8809, 10438, 9635]. Artificially [5964]. Artillery [7971, 8334]. Artist [7164]. Artistic [3654, 6169, 6915, 7131]. Artists [8846]. Artograph [5599]. Arts [3078, 3243, 8545, 8834, 9143, 3917]. Ascended [8123, 9284]. Ascensions [2495]. Ascertain [8125]. Aschaffenburg [7884]. Ash [105, 311, 5827]. Ashore [2656]. Asia [10687]. Aspect [8648]. Aspects [6379]. Asphalt [8221]. Asphyxiation [4205]. Assaying [1836]. Assignments [4817]. Assistant [7808]. Association [1997, 2774, 5004, 4985, 6018, 6037, 8035, 8122, 8148, 8575, 10184, 8103, 8194, 7706, 871, 7077, 7098, 9263, 10394, 1294, 1437, 2053, 2929, 3982, 5024, 4952, 6057, 6172, 7060, 7900, 8274, 8049, 3954, 8827]. Astigmatic [68]. Astoria [9912, 8304]. Astronomer [4725]. Astronomical [1589, 5189, 5850, 6329, 6957]. Athbara [9947]. Athletes [5090]. Atkinson [1751]. Atlanta [6303, 5067, 5294, 5935, 6198, 6234, 6253, 6326, 6285, 5538, 6349, 5918]. Atlantic [4901, 5763, 5838, 6024, 8241, 9445, 3097]. Atlantis [4839]. Atmospheric [3352, 6541]. Atomizer [316]. Atoms [10667]. Attaching [346, 2341, 6003]. Attachment [468, 132, 1035, 575, 687, 295, 710, 734, 693, 1962, 1774, 1852, 2167, 2017, 1445, 3037, 2624, 3497, 4001, 3287, 3570, 4139, 3750, 3751, 3774, 4563, 5874, 6187, 6023, 6811, 6768, 6772, 6693, 7622, 7433, 9432, 9287, 8742, 8978, 9121, 9495, 9452, 10653, 9987, 10331, 7259]. Attachments [1495, 2740]. Attacked [3326]. Attained [9531]. Attaining [5715]. Attending [9279]. Attention [1199]. Attitudes [8641, 8660]. Attorney [8211]. Attorneys [8083, 8087, 7574]. Attractive [5597].

Auchmuty [3839]. Auditoriums [7413]. Audubon [9868]. August [6037, 6916, 731, 1927, 2845, 3869, 5936, 8016, 6622, 9135, 10275]. Augusta [159, 3112]. Augustus [8108]. Aurora [7066]. Austin [2991, 6988]. Australia [1950, 1696, 2818, 7826]. Australian [6217]. Authenticated [3042]. Authorship [6144]. Auto [5543]. Auto-Phoro-Optometer [5543]. Automata [9615]. Automatic [1127, 1202, 325, 1149, 429, 417, 672, 978, 646, 441, 427, 1747, 1788, 2267, 2287, 2247, 2271, 1597, 2739, 2486, 2551, 3217, 2693, 4300, 3313, 3498, 3436, 4049, 3568, 4099, 4304, 3248, 4957, 5007, 4818, 5054, 4857, 4340, 4729, 4797, 4548, 4390, 5831, 6345, 5878, 5533, 6151, 5477, 5478, 6237, 5891, 6126, 6186, 5646, 5419, 6322, 5595, 7198, 6804, 6884, 6924, 6671, 7293, 7219, 6991, 6520, 7150, 6827, 8291, 8422, 7377, 7929, 7639, 7575, 7778, 8382, 7671, 7617, 7839, 8696, 8933, 9471, 9071, 9155, 8664, 9570, 10516]. Automatic [10645, 9727, 10044, 9960, 10075, 9680, 9803, 10625, 10312, 10242]. Automatically [444, 584, 4748, 4383, 8832]. Automaton [2506, 4863]. Automobile [5898, 5810, 5705, 9016, 10738, 10188, 10761, 10211, 10295, 10444, 10119, 10025, 9653, 10593, 10009, 10285, 10029]. Automobiles [8036, 9568, 9709, 10670]. Automobilism [10006]. Autotruck [9748]. Auxiliary [1087, 2093, 4158, 6752, 3446]. Avalanche [6199]. Aven [8808]. Avenue [1707, 2128, 3917, 5088, 4936, 4876, 4773, 4511, 4569, 7008, 8063, 7358, 7981, 9429, 9568, 4673, 4705, 5246, 6461]. Average [10294]. Aviator [2424]. Avion [9216]. Avondale [5945]. Award [6293, 7836]. Awarded [3055, 9418]. Awards [3652]. Awning [547, 3427, 7042]. Ax [9047]. Axle [714, 3473, 4987, 7198, 6440, 8216]. Axles [320, 6565, 8298]. Ayres [10738].

B.C. [1401]. B.L. [6566]. Babcock [3882]. Baboon [2291]. Baby [6581]. Bacillus [1261]. Back [940, 10793]. Back-Pedaling [10793]. Backward [10020]. Bacteria [2773]. Bad [2204, 9529]. Badge [3024]. Bag [523, 116, 4632, 6714, 6673]. Baggage [596, 9119, 9779]. Bagging [2993]. Bags [531, 1765, 2398]. Bahama [4633]. Bahamas [2220]. Bailey [4797, 8453]. Bait [3309]. Baked [6657]. Baker [1943, 2854]. Baking [9199]. Baku [3003, 6504]. Balance [7859, 10236]. Balanced [2395, 2793, 3977, 3937, 7242, 8104]. Balancing [8566]. Baldwin [6688]. Baling [1673]. Ball [1304, 2748, 5868, 6810, 7332, 7100, 7155, 7668, 8491, 9250, 8914, 9094, 10193]. Ball-Bearing [9250, 9094]. Ballasting [1108]. Ballistic [6889]. Balloon [979, 1216, 5179, 5870, 7139, 8475, 8002, 7933, 9395, 9131]. Ballooning [2019, 3009, 9933]. Balloons [8701, 10689]. Ballot [2755]. Balls [3161, 4142]. Baltic [5253]. Baltimore [870, 3507, 3687, 4608, 9316]. Bamboo [2458]. Banana [9761]. Band [940, 83, 2932, 3889, 7135, 9958]. Band-Saw-Setting [9958]. Banff [2377]. Bank [3697, 4895, 10183]. Banks [5134, 4593]. Banyan [962, 1447, 5446]. Bapid [7254]. Baptist [5202]. Bar [3746, 4798, 5385, 6123, 6351, 6369, 6822, 7668, 8197, 7669, 9214, 8976, 3021]. Barbed [3115, 3601]. Barbel [6992]. Barbel-Forming [6992]. Barbette [6566]. Barges [7089]. Barn [1808]. Barometer [6505, 7514, 2492]. Barrel [1087, 248, 1807, 1794]. Barrels [1872, 4030, 9989]. Barrow [10381]. Bars [6851]. Baseball [662, 7909]. Basement [2905]. Basin [674, 4126, 10220]. Basins [5717]. Basis [8989, 9402, 9761]. Basket [10645]. Bassett [6438]. Batchelor [7622]. Bates [3551]. Bath [1700, 1461, 1489, 6443, 7648, 9905]. Bathers [2045]. Bathing [10140]. Baths [2832, 7958, 8509, 9963]. Bats [7707]. Batteries [1414, 2243, 2783, 5137, 8893, 8822]. Battery [4305, 4204, 4569, 5386, 6131, 6466, 7674, 10103, 7317, 8446]. Battle [245, 1641, 2965, 3741, 4549, 5080, 6184, 6225, 5348, 5755, 6367, 10547, 10078, 5364, 6019, 7863]. Battleship [2297, 6100, 7030, 6655, 6863, 7253, 6734, 7318, 7011, 8217, 7477, 7660, 7643, 7930, 9498, 8951, 8680, 9372, 9347, 10695, 10771, 9852, 10439, 10238, 9712]. Battleships [6304, 6973, 7718, 9148, 8785, 8581, 8810, 9732, 9897, 10427]. Bauxite [4539]. Bay [2078, 2344, 5897, 8156, 8925]. Bayonne [9874]. Bazar [7761]. Bazin [7086, 7657]. be [885, 2171, 3719, 6254, 5701, 8291, 8463, 8211, 9047, 9899]. Beach [7887, 2865, 6378, 6424, 6406, 9911]. Beacon [8586]. Beader [1180]. Beam [122]. Bean [10376]. Beans [1784]. Bearing [536, 3826, 7332, 7100, 7155, 6866, 7668, 8264, 9250, 8914, 9094, 10473, 10496]. Bearings [6810, 8491, 10193]. Bears [9630]. Beat [6382]. Beater [645]. Beating [4991]. Beatrice [5220]. Beautiful [4888, 4697, 9038, 1681]. Beaver [4887]. Becklight [10676]. Become [4137]. Bed [424, 2644, 4161, 6020, 6852, 8760, 9353]. Bedclothes [1871, 7535]. Beds [3497, 6811]. Bedstead [3866, 10052]. Bedsteads [229]. Bee [9900]. Beecher [1825]. Beelin [8475]. Bees [2165]. Beeswax [4170]. Beet [5757, 8358, 8380, 8353, 8021, 8101, 9609]. Beetaloo [1950]. Beetle [9470]. Before [1274, 2485, 8760]. Beginning [1321, 1539]. Beheader [4237]. Belated [2431]. Belgian [4780, 6515, 8061]. Belgium [9869]. Bell [2021, 6220, 6048, 6369, 6713, 7044, 8892, 10191, 2882, 3445, 5325, 5834]. Belt [1185, 1207, 64, 1352, 1495, 3286, 3261, 3595, 4634, 5072, 4865, 6345, 6225, 6651]. Belting [374, 121, 4341]. Belts [644, 2586, 4032]. Bench [3913]. Bender [869]. Bending [278, 2449, 7237]. Beni [258]. Benin [8585]. Benjamin [1986, 7663, 4105]. Bennett [1697, 6350]. Bent [5297]. Benzine [2269, 7316]. Bequest [7941, 9404]. Berkefeld [4842]. Berlin [627, 998, 1329, 5220, 5503, 5786, 6732, 7099, 6487]. Berliner [3363, 3425, 5734, 7780]. Bersier [4940]. Berths [5387]. Bertillon [7619]. Bessemer [2255, 6686, 6780, 8772]. Bessiere [6854]. Best [2286, 3042]. Bestowing [6219]. Bethlehem [3870]. Beton [3039]. Better [4607, 10160]. Bettini [402]. Between [765, 2077, 2422, 3096, 3663, 4935, 1039, 2114, 3010, 3661, 5253, 5464, 6648, 7469, 8428, 10572, 10689]. Bevel [3955, 5574]. Beveling [4147]. Beveridge [4099]. Bicycle [81, 343, 315, 1417, 1603, 1467, 2566, 2592, 3744, 3306, 3887, 3549, 4797, 4571, 4464, 4407, 4747, 4737, 4711, 5254, 4819, 4380, 5195, 4993, 5878, 6347, 5909, 5424, 6346, 6189, 5699, 5940, 5779, 6060, 6021, 6316, 5440, 5781, 5465, 6133, 5768, 5436, 5466, 6009, 5407, 5934, 6056, 5756, 5854, 5523, 6436, 6945, 6399, 6363, 7106, 7097, 7104, 6484, 6691, 6377, 6445, 6441, 6369, 7332, 6460, 6359, 6713, 6364, 7078, 6370, 6633, 6980, 6361, 6446, 6897, 7487, 7837, 7430, 8238, 7554, 8167, 8280, 8299, 7522, 7824, 7953, 7556]. Bicycle [7926, 7975, 8430, 7763, 7500, 7740, 7495, 7501, 8402, 7595, 7951, 8473, 7952, 7783, 8003, 7942, 8319, 7596, 9260, 8849, 9031, 8561, 8513, 9206, 9424, 8892, 8932, 8975, 8858, 9354, 9532, 10027, 10069, 10002, 10081, 10015, 10189, 10026, 10010, 10229, 10241, 9588, 10520, 9689, 10165, 10005, 10013, 9759, 3449, 10020, 7167]. Bicycle-Chain [10010]. Bicycle-Crank [10165]. Bicycle-Lamp [9532]. Bicycle-Lock [9354]. Bicycle-Pump [9031]. Bicycle-Support [10013]. Bicycler [5832]. Bicycles [734, 1831, 2167, 1666, 4404, 4563, 4648, 5983, 6351, 6266, 6718, 6465, 6692, 8338, 7401, 7593, 9071, 8512, 9346, 9121, 8633, 10793, 10012, 10496]. Bicyclist [6653]. Biding [2076]. Bids [7664, 10684, 10660]. Big [1145, 1775, 2423, 8164, 10465, 8090]. Biggest [9483]. Bilbao [3039]. Bill [387, 3303, 7632, 8782, 8801, 466, 415, 1322, 1274, 2720, 6580, 6687, 6824, 8802, 8887, 8800, 9006, 9722]. Billets [4570]. Billing [2607]. Billroth [4712]. Bills [1764, 1855, 2485, 8824]. Biltmore [4583, 6449]. Bin [5422]. Binder [2133, 3387]. Binding [371, 388, 425, 4006, 7903, 7383]. Bins [7220]. Biograph [9592]. Biological [5981]. Biology [659]. Birch [1785]. Bird [1694, 4350, 7770, 10108]. Birds [2853, 5542, 6112, 5924, 9348, 3509, 8223, 10215]. Birmingham [6324, 6286]. Birth [10120]. Birth-Rate [10120]. Birthplace [6315]. Bishop [7776]. Bismarck [1744]. Bit [2016, 2588, 3244, 3592, 6482, 8511]. Bitter [3508]. Biwa [7217]. Bizerte [5880]. Black [2819, 8093]. Blackboards [911]. Blacksmith [6307]. Blackwall [7878]. Blackwell [5475]. Blade [80, 2488]. Blades [2400, 4243]. Blake [42, 2304, 3708, 5835, 5468]. Blanc [4285, 5099]. Blank [7903]. Blanket [497, 2690]. Blast [6220, 5873, 5666, 8244]. Blasters [2337]. Blasting [715, 1727, 5292]. Bleacher [5159]. Bleachery [4170]. Bleaching [5959]. Blind [297, 782, 1591, 7690, 6700]. Blindfolded [8498]. Block [1174, 310, 716, 1848, 1793, 4028, 3503, 5048, 2399]. Blocks [943, 3039, 6306]. Blow [819]. Blow-Out [819]. Blower [2199, 2884]. Blowing [2935, 4097]. Blown [7364]. Blowpipe [2269, 6479]. Blue [9734]. Blunder [8552]. Boa [2007, 6526]. Board [259, 2093, 1626, 5173, 6608, 8256, 7138]. Boards [3234]. Boat [209, 465, 1108, 52, 999, 1981, 1468, 1325, 2854, 2311, 3128, 3940, 4630, 4877, 5158, 4571, 4989, 4993, 4561, 5015, 4736, 5089, 5266, 6222, 5704, 5702, 6318, 6246, 5785, 6696, 7799, 8341, 8292, 7542, 7787, 8427, 8130, 9336, 9304, 8640, 9323, 8664, 9115, 9058, 8804, 9017, 8813, 8516, 9437, 9684, 10076, 10648, 9925, 9931, 8453, 99, 10747, 8302, 7138]. Boat-Launching [8664]. Boating [3023, 7977]. Boats [442, 1847, 1773, 2863, 2728, 2359, 4286, 4254, 3765, 4366, 5783, 6083, 5501, 5342, 7330, 6542, 6405, 7468, 9365, 8747, 8599, 9864, 10375, 10573]. Bob [937, 879, 5913]. Bodies [2914]. Body [1495, 10308]. Boers [10562]. Bog [674]. Bohemian [3216]. Boiler [1149, 30, 914, 741, 739, 1034, 1103, 1488, 2201, 1291, 2486, 3217, 2827, 4260, 3523, 3308, 3882, 4040, 4881, 4716, 5212, 5156, 5310, 5068, 4801, 4613, 4780, 6085, 6218, 5737, 6041, 6148, 5632, 6400, 6820, 6729, 6830, 8154, 8262, 7615, 7752, 7838, 7379, 7359, 9446, 8756, 8632, 9073, 8659, 9929, 10566, 9659, 9985, 10388, 10758, 6838]. Boiler-End [4260]. Boiler-Tube [9985]. Boilers [28, 314, 2268, 2242, 2503, 2826, 2721, 2813, 3506, 3784, 4885, 5329, 6078, 7162, 6901, 7426, 9039, 9494, 8720, 10651, 10102, 9802, 10187]. Bollee [7171]. Bologna [8727]. Bolt [709, 1494, 2742, 2968, 5537, 8698]. Bolt-Threading [2968]. Bolts [7450, 10050]. Bombardment [8953]. Bombay [7813, 9473]. Bone [1121]. Bonnetable [4919]. Book [597, 1816, 2607, 3451, 3788, 4918, 5633, 7472, 8170, 9529, 1663, 8082, 1322]. Bookbinders [3234]. Bookcase [4239]. Bookholder [5963]. Books [830, 166, 237, 262, 282, 304, 328, 407, 433, 458, 561, 751, 772, 825, 880, 1111, 1138, 1161, 1193, 950, 968, 491, 514, 188, 215, 353, 379, 539, 654, 681, 700, 724, 1025, 1069, 1091, 1362, 1285, 1308, 1336, 1384, 1451, 1654, 1719, 1740, 1801, 1858, 1882, 1938, 1955, 2100, 2121, 2159, 2173, 2189, 2237, 2258, 1786, 1407, 1507, 1556, 1388, 1533, 1635, 1660, 1478, 1501, 1527, 1550, 1574, 1606, 1630, 1903, 1920, 2140, 1314, 1339, 1535, 1581, 1612, 1637, 1662, 2347, 2387, 3014, 3061]. Books [3082, 3100, 3117, 3135, 3151, 3169, 3188, 3206, 3219, 2406, 2425, 2441, 2459, 2478, 2496, 2517, 2539, 2573, 2612, 2631, 2648, 2663, 2678, 2697, 2730, 2749, 2786, 2802, 2821, 2878, 2906, 2923, 2942, 2368, 2714, 3236, 3255, 3852, 3875, 3898, 3922, 3948, 3969, 3988, 3994, 4042, 4066, 4088, 4112, 4129, 4148, 4174, 4198, 4221, 4268, 4290, 4311, 3276, 3295, 3318, 3336, 3355, 3376, 3395, 3418, 3437, 3464, 3487, 3511, 3532, 3556, 3577, 3624, 3646, 3666, 3689, 3711, 3734, 3778, 3796, 3814, 4351, 4372]. Books [4978, 5018, 5056, 5076, 5110, 5146, 5181, 5204, 5222, 5278, 5299, 5318, 5335, 4394, 4414, 4437, 4457, 4480, 4501, 4531, 4554, 4573, 4599, 4621, 4638, 4658, 4676, 4698, 4738, 4761, 4783, 4804, 4831, 4869, 4889, 4911, 4928, 5240, 5351, 5370, 6031, 6050, 6070, 6091, 6114, 6156, 6175, 6191, 6210, 6229, 6268, 6287, 6309, 6331, 5389, 5410, 5430, 5449, 5470, 5486, 5526, 5545, 5566, 5584, 5657, 5670, 5686, 5707, 5748, 5770, 5791, 5817, 5840, 5861, 5883, 5902, 5925, 5949, 6371, 6390, 7156, 7190, 7210]. Books [7244, 7263, 7282, 7304, 7321, 7350, 6409, 6430, 6468, 6489, 6507, 6528, 6593, 6614, 6639, 6758, 6773, 6795, 6814, 6839, 6857, 6982, 7000, 7015, 7032, 7051, 7069, 7108, 7124, 7141, 6848, 7370, 7387, 8006, 8026, 8051, 8077, 8095, 8115, 8137, 8158, 8182, 8206, 8226, 8247, 8268, 8285, 8306, 8346, 8369, 8411, 8435, 8455, 8480, 7402, 7420, 7439, 7479, 7502, 7523, 7543, 7562, 7605, 7624, 7649, 7675, 7700, 7724, 7744, 7791, 7816, 7842, 7865, 7890, 7911, 7935, 7961, 7984, 7903, 8193, 8499, 8523, 9240]. Books [9255, 9274, 9296, 9319, 9340, 9361, 9378, 9438, 9459, 9478, 9541, 8568, 8619, 8645, 8668, 8729, 8750, 8774, 8795, 8815, 8836, 8860, 8882, 8900, 8919, 8941, 8964, 9001, 9021, 9063, 9082, 9124, 9182, 9201, 9220, 9595, 9618, 10432, 10456, 10500, 10527, 10552, 10577, 10601, 10627, 10655, 10677, 10752, 10780, 9644, 9693, 9714, 9736, 9784, 9808, 9834, 9858, 9887, 9915, 9941, 9967, 9996, 10032, 10057, 10083, 10130, 10172, 10198, 10243, 10265, 10288, 10358, 10403, 10318]. Boom [2320]. Boomer [1204, 1673]. Boot [280]. Boots [2511, 6298, 5518]. Bordeaux [3897]. Bore [7189, 10720]. Boring [2016, 3129, 2362, 4260, 4185, 3938, 4544, 5484, 8384, 10589]. Borki [5013]. Boschert [1204, 1673]. Boston [6063, 209, 3433, 6226, 6026, 6276, 7058, 7576, 8179, 8109, 9627, 9743, 9590, 10620, 3203]. Botanical [10726]. Bottle [1014, 2608, 6254, 5346, 6711, 7105, 6994, 7200, 8106, 8153, 8912, 9118, 9708, 9685]. Bottle-Closure [9118]. Bottles [3635, 6886, 7257, 8129, 8493]. Boulevard [5406]. Bound [10286]. Boundaries [5358]. Boundary [1228]. Bourgogne [9089]. Bouton [7171, 10024]. Bow [1241, 2074, 2364, 2377]. Bowling [4715]. Box [1203, 250, 2090, 1283, 3178, 2755, 3473, 4319, 4987, 4614, 4732, 6400, 6440, 8337, 8044, 9137, 10566, 10517, 10692, 9682]. Boxing [10332]. Boy [1121]. Boynton [1538, 4464]. Brace [418, 3244, 3592, 4612, 8174]. Bracing [9261]. Bracket [691, 5939, 6460, 9729]. Bradley [9177]. Braided [3730, 4160]. Brake [1151, 77, 81, 201, 393, 260, 669, 641, 1788, 1832, 2172, 4301, 4003, 4667, 4797, 4450, 5831, 5419, 6322, 5892, 6021, 6316, 5651, 5649, 5407, 6669, 6625, 8090, 7926, 7975, 8430, 7495, 8003, 8512, 8561, 10793, 9706, 10415, 10028]. Brakes [1000, 2017, 3424, 7639]. Brambel [7436]. Branch [8387, 10167]. Branches [8628]. Branchidæ [8630]. Branding [10700]. Brass [6062, 7737]. Brasseys [10118]. Brazer [7952]. Brazil [1052, 7764]. Brazilian [4255, 4256]. Brazos [3273]. Bread [550, 2107, 3227, 9199]. Break [3685]. Breakage [2324, 3252]. Breakdown [9547]. Breaker [2759, 3430, 9053, 9902]. Breaks [5311]. Breast [1732, 3197]. Breech [494, 4778, 8876]. Breech-Loading [4778]. Breeders [10254]. Breeding [9667, 1984]. Breslau [5418, 5090]. Breviceps [3026]. Brick [34, 573, 1208, 1181, 2988, 5046, 5461, 6607, 8461, 9553]. Brickill [10125]. Brickmaking [2702]. Bricks [7364]. Bridge [131, 1136, 1179, 323, 1239, 428, 1789, 2074, 1295, 1465, 2031, 1724, 1787, 2062, 2458, 3214, 3089, 2383, 3733, 3470, 4512, 4825, 5306, 4876, 5265, 4755, 4942, 5270, 5311, 5157, 5421, 5475, 5541, 6267, 7270, 6705, 6807, 6883, 8050, 7740, 7741, 7920, 7591, 8320, 8551, 9317, 9947, 10786, 10294, 10426, 10147, 7114, 5439, 25, 288, 676, 3410, 3333, 3984, 3941, 4825, 4943, 4705, 4773, 5813, 6167, 6182, 7075, 7510, 9179, 9551, 8763, 8505, 8605, 9474, 10709, 10525, 9575, 9585, 10536, 7692, 7170]. Bridgeport [3132]. Bridges [1979, 3428, 3548, 6935, 7757, 9471, 10151, 10371, 9827, 10760]. Bridging [139]. Bridle [690, 933, 2588, 2669]. Bridles [297]. Brief [8618, 8628]. Brigands [9631]. Briner [5827]. Brinton [10369, 4973]. Bristol [4634, 6058, 5554]. Britain [6515, 9427, 10685, 8278]. British [6172, 970, 42, 989, 2629, 4254, 3766, 4366, 6339, 7235, 6960, 8122, 8148, 7750, 7511, 7992, 8427, 10300, 9897]. Broad [3092]. Broadsword [8996]. Broadway [9911, 1707, 2128, 3555, 3631, 4307, 5035, 7358, 9557]. Broken [3294]. Bromide [10665]. Bronze [7068, 8326, 7768, 9251]. Brooder [2583, 9906]. Brooding [3895]. Brook [674, 4213]. Brooklyn [2382, 3089, 6009, 7786, 7811, 8478, 8, 1132, 1825, 1791, 1248, 1511, 2928, 4074, 3243, 3554, 3696, 5044, 4897, 4952, 4985, 5032, 5276, 5265, 5350, 6152, 6167, 6524, 6764, 7510, 7851, 8102, 7469, 7499, 9179, 9561, 8710, 8905, 8916, 9169, 9551, 8763, 8505, 9100, 10462, 9594, 10525, 10536]. Brooks [743, 7031]. Broom [1393, 10164]. Brooms [4956, 10254]. Brougham [8046, 10188, 10333]. Broughton [3618]. Brownlow [1599]. Brush [38, 1490, 2669, 4102, 5479, 9935, 2676, 1213, 4166]. Brushes [476]. Brussels [2109, 7709, 8736]. Brux [6029]. Bubble [686]. Bubbles [2708, 2951, 9742, 3576]. Buchanan [8429]. Buckle [940]. Buckles [3476]. Budapest [6770]. Buechner [10040]. Buenos [10738]. Buffalo [7077, 7098, 2374, 7060]. Buffaloes [5724]. Buffing [4686]. Buffington [6566, 7974]. Bug [5998]. Bugs [5971]. Build [1586, 9865, 8144]. Builder [3174]. Builders [4497]. Building [802, 837, 1118, 187, 213, 261, 303, 327, 351, 432, 507, 630, 652, 679, 699, 942, 1170, 1412, 1640, 1665, 1865, 1620, 1265, 1284, 1356, 1476, 1500, 1526, 1548, 1604, 1628, 1653, 3124, 3225, 3919, 3829, 3883, 4025, 3830, 3831, 3987, 4984, 4822, 4381, 5314, 6101, 6303, 7007, 6522, 7364, 8219, 7876, 8200, 7897, 8074, 7855, 8673, 9624, 9732, 9756, 1698, 7561, 5125, 8134, 281, 1198, 1891, 2229, 2941, 3078, 3828, 4288, 4083, 3394, 3504, 3944, 4477, 8082]. Buildings [1159, 1792, 3068, 2673, 2328, 3526, 3963, 5133, 5287, 4816, 4866, 5119, 5264, 6236, 6864, 7147, 6578, 9396, 9560, 9261, 9815, 9895, 10216, 8471, 3077, 3434]. Built [1079, 1237, 7918, 8427, 9414, 8302]. Bulkhead [9039]. Bull [850, 5010, 4589]. Bullet [4896, 4882, 5267, 7949]. Bulletins [1323]. Bulletproof [5851]. Bullets [7202, 8175, 9349]. Bundy [4304, 5833, 9903]. Bunsen [4611, 8131, 10370]. Buoy [5915, 8450, 9472, 10240]. Buoy-Cables [9472]. Buoys [2954]. Bureau [89, 6342, 9384, 9806, 9795, 8506, 8672, 10634, 10688, 10669, 10791, 9725]. Bureaus [9720]. Buren [5684]. Burgess [1917]. Burglar [559, 2197, 3911, 5155, 5316, 5217, 8239, 8173]. Burner [133, 939, 1875, 1462, 2562, 4257, 3452, 4327, 5192, 4707, 4611, 4609, 6417, 8131, 8383, 10348]. Burnham [4670]. Burning [58, 5192, 8714]. Burns [9876]. Burr [4362]. Burrowing [677]. Burrows [3509]. Bursting [9561]. Burt [4328]. Bushmen [2890]. Business [798, 829, 1043, 1068, 1093, 20, 46, 680, 702, 723, 748, 769, 792, 822, 854, 881, 903, 923, 946, 963, 982, 1004, 1023, 1113, 1139, 1162, 1189, 1218, 1242, 69, 93, 123, 145, 167, 214, 238, 263, 283, 305, 329, 352, 381, 408, 434, 459, 486, 509, 538, 562, 586, 611, 631, 653, 1549, 1266, 1286, 1879, 1901, 1919, 1934, 1952, 1972, 1988, 2009, 2026, 2046, 2064, 2082, 2102, 2123, 2142, 2154, 2175, 2191, 2215, 2234, 2260, 2279, 2298, 1309, 1332, 1358, 1380, 1403, 1425]. Business [1477, 1503, 1529, 1573, 1605, 1629, 1656, 1682, 1701, 1720, 1737, 1757, 1776, 1798, 1819, 1839, 1859, 1325, 2367, 2556, 2875, 2315, 2329, 2822, 2838, 2857, 2892, 2924, 2943, 2959, 2981, 2997, 3016, 3031, 3047, 3063, 3083, 3101, 3118, 3136, 3153, 3171, 3189, 3207, 3220, 2349, 2388, 2407, 2427, 2442, 2461, 2480, 2498, 2518, 2536, 2574, 2596, 2614, 2633, 2650, 2664, 2680, 2699, 2716, 2731, 2751, 2765, 2787, 2804, 3237, 3257, 3797, 3816, 3834, 3854, 3877, 3900, 3924, 3949, 3966, 3989, 4018, 4043, 4067, 4090, 4113]. Business [4131, 4150, 4176, 4199, 4223, 4245, 4270, 4291, 4313, 4329, 3278, 3297, 3320, 3338, 3356, 3377, 3397, 3420, 3439, 3465, 3488, 3512, 3534, 3558, 3579, 3602, 3625, 3644, 3668, 3691, 3713, 3735, 3759, 3780, 4352, 4373, 4890, 4912, 4929, 4946, 4960, 4979, 4998, 5019, 5038, 5057, 5073, 5092, 5111, 5128, 5147, 5163, 5182, 5205, 5223, 5239, 5259, 5279, 5300, 5319, 5336, 4395, 4416, 4439, 4459, 4482, 4502, 4517, 4532, 4556, 4575, 4601, 4623, 4639, 4659, 4678, 4699, 4718, 4740, 4762, 4785, 4806, 4832, 4851, 4871, 5353]. Business [5371, 5863, 5884, 5903, 5926, 5950, 5972, 5992, 6013, 6032, 6052, 6071, 6092, 6115, 6136, 6157, 6176, 6192, 6211, 6230, 6248, 6270, 6289, 6310, 6332, 6352, 5390, 5411, 5431, 5450, 5471, 5488, 5505, 5527, 5546, 5567, 5585, 5606, 5622, 5640, 5659, 5671, 5687, 5708, 5727, 5750, 5771, 5792, 5819, 5841, 6372, 6391, 6892, 6908, 6927, 6963, 6983, 7002, 7017, 7033, 7053, 7071, 7090, 7109, 7125, 7142, 7157, 7173, 7191, 7211, 7226, 7246, 7264, 7283, 7305, 7323, 7339, 7351, 6411, 6431, 6451, 6469, 6490, 6509, 6529]. Business [6549, 6571, 6594, 6615, 6640, 6662, 6680, 6701, 6720, 6738, 6755, 6774, 6796, 6815, 6840, 6859, 6873, 6950, 7233, 8034, 7656, 7367, 7389, 7937, 7962, 7985, 8007, 8027, 8052, 8075, 8096, 8116, 8138, 8159, 8183, 8207, 8227, 8248, 8269, 8286, 8308, 8329, 8347, 8370, 8390, 8412, 8432, 8456, 8481, 7403, 7421, 7440, 7458, 7481, 7503, 7524, 7544, 7563, 7582, 7606, 7626, 7650, 7676, 7697, 7725, 7746, 7772, 7792, 7818, 7843, 7866, 7891, 7912, 7491, 8501, 8524, 9064, 9083, 9103, 9126, 9144, 9164, 9183, 9202, 9221, 9237]. Business [9256, 9275, 9297, 9320, 9341, 9362, 9379, 9398, 9420, 9439, 9460, 9480, 9502, 9522, 9542, 9563, 8546, 8569, 8587, 8620, 8642, 8665, 8687, 8706, 8730, 8751, 8775, 8796, 8816, 8837, 8861, 8883, 8901, 8921, 8942, 8965, 8983, 9002, 9023, 9040, 8842, 8654, 9625, 9578, 9597, 10199, 10224, 10245, 10267, 10289, 10316, 10338, 10359, 10383, 10404, 10433, 10457, 10481, 10501, 10528, 10549, 10578, 10603, 10629, 10656, 10678, 10702, 10730, 10753, 10781, 9619, 10801, 9645, 9669, 9694, 9715, 9737, 9764, 9785, 9810, 9835, 9859, 9889, 9916, 9942, 9968]. Business [9997, 10033, 10059, 10084, 10109, 10131, 10152, 10173]. Business-Patent [7491]. Bust [6904]. Busting [10646]. Busts [7398]. Butter [202, 3804, 10591]. Butterflies [8421]. Butterfly [9218]. Butterworth [7663, 8576]. Button [3567, 4473, 10310]. Buttonhole [4095, 6243]. By-Products [9407].

C [1980, 4120, 6025, 6235, 9854]. Cab [5214, 9829]. Cabaret [6545]. Cabinet [576, 2433, 2491, 8579, 10123]. Cabinets [6848, 6463]. Cable [199, 810, 776, 1707, 2128, 2276, 1276, 1542, 2339, 3089, 2571, 3555, 3631, 4307, 3960, 4538, 4408, 4391, 4901, 4511, 6350, 5721, 5838, 6497, 6940, 7721, 9557, 8490, 10609]. Cable-Conveying [2276]. Cables [5744, 9472, 10709]. Cableway [5015, 8607]. Cabs [8199]. Cactus [5428, 10600, 8894, 4127]. Cadets [8724]. Caiman [2311]. Caisson [4984, 4986, 5946]. Cake [5441, 6657]. Cal [211, 7601, 9416]. Calcining [6786]. Calcium [8125, 8927]. Calculating [4941]. Calculator [2569]. Caldera [1888]. Caledonia [5194]. Calendar [5647]. Calendarial [10471]. Caliber [4308, 5267, 5191]. California [2610, 818, 720, 1889, 2423, 2638, 4328, 4467, 4838, 4635, 4402, 5515, 5612, 5521, 6737, 7335, 6462, 6891, 8389, 8201, 7887, 7979, 9537, 8556, 8711, 8519, 10141, 10641, 10221, 10213, 10336, 9805, 9338, 8004, 10775, 10598, 10170]. Californian [9965]. Caligraph [9515]. Caliper [606, 122, 4858]. Calipers [8746]. Calking [3640]. Call [4487]. Called [9279]. Calls [186]. Caloric [7765]. Calorimeter [9959]. Calumet [7102, 7120]. Calves [1593]. Cam [7012, 2081, 2099, 2112]. Camera [324, 224, 2883, 2843, 2995, 2971, 2606, 2758, 3474, 4235, 4879, 4589, 5965, 7163, 6825, 7337, 6587, 6771, 7905, 7432, 7720, 8281, 8361, 9333, 9095, 10072, 10518, 2815, 5579]. Cameras [1187]. Camille [9428]. Camp [9262]. Campaign [8987, 9593]. Campaigning [9108]. Campania [3417]. Camphor [1762, 359, 8487]. Camping [7977]. Can [445, 808, 249, 3053, 10716]. Canada [765, 2077, 4454]. Canadian [3043, 3585, 10204]. Canal [556, 3177, 2359, 4286, 3984, 5253, 5015, 8397, 9049, 10757, 10116, 10375, 7217, 277, 686, 173, 1124, 1732, 2188, 1343, 2791, 3403, 4015, 3941, 5141, 4522, 5218, 5776, 5805, 5813, 5899, 5839, 5814, 7750, 7686, 7467, 8702, 9769, 10708, 10137, 9651, 9665, 9711, 9791, 9675, 9286]. Canals [3362, 8677, 6024, 8552, 9744, 9699]. Canceling [6666]. Candle [8586]. Candles [4263]. Candlestick [6546, 8240]. Candlesticks [111]. Candy [4692]. Cane [257, 1416, 3073, 3110, 5381, 8764]. Canet [1355]. Canned [3620]. Canning [2760]. Cannon [1304, 4059, 5701, 5425, 6204, 5937, 5947, 7719, 8738]. Cannot [6254]. Cans [1692, 3462, 6164, 9054]. Canyon [2747, 2873]. Cap [1041, 1256, 9934]. Capacity [5174, 5941, 5703, 9446]. Cape [6024, 8948, 2737, 8907]. Capewell [4368]. Capillarity [2564]. Cappadocia [7206]. Caps [737]. Capstan [897]. Capstans [4734]. Capture [7435]. Captured [9140]. Car [1127, 715, 260, 107, 135, 155, 203, 294, 338, 498, 600, 620, 625, 995, 1104, 41, 557, 558, 552, 783, 1009, 1788, 2267, 1764, 2004, 1832, 1812, 1971, 2002, 2038, 1728, 2056, 2172, 1566, 1300, 2017, 2379, 2777, 2848, 2912, 2622, 2365, 2355, 3127, 3473, 3499, 3528, 3808, 3448, 3500, 3424, 3659, 5155, 4488, 4987, 4688, 5027, 4665, 4937, 5135, 5064, 4466, 4668, 4569, 5533, 5511, 6245, 6102, 5651, 5596, 5649, 5808, 6244, 6259, 5479, 6266, 5829, 6105, 5595, 6804, 6884, 7096, 6497, 6506, 7116]. Car [7209, 6440, 6540, 7312, 7100, 7155, 7042, 6625, 7295, 6902, 6524, 8090, 7639, 7896, 7860, 7973, 8325, 8359, 8387, 7721, 9053, 8829, 8763, 9751, 10609, 10589, 10415, 10350, 9876]. Car-Boring [10589]. Car-Heating [41]. Car-Seal [9053]. Carbide [8125, 8927]. Carbon [2579, 5476]. Carbonic [4310]. Carborundum [4615]. Carbureters [8993]. Card [1626, 5815, 6995, 10479, 9181]. Card-Board [1626]. Cardboard [10515]. Cardinal [9932]. Care [1525, 9096, 9028]. Career [5592, 5378, 5417, 5492]. Cargoes [2070]. Carli [4655]. Carmel [5176]. Carmelite [10314]. Carnegie [2915]. Carolina [3372]. Carpenter [5102, 2535, 3827, 4693]. Carpet [601, 56, 4585]. Carr [2986]. Carrara [9880]. Carriage [1379, 1423, 3174, 2323, 2677, 3708, 4168, 5085, 4655, 5213, 4429, 5328, 4595, 5892, 5550, 6128, 5958, 5598, 5933, 5667, 6566, 6718, 7171, 7208, 7260, 6835, 6581, 7261, 7331, 7121, 7790, 7974, 7998, 9016, 8607, 9057, 8915, 9819, 10023, 9700]. Carriages [2863, 2795, 4450, 5098, 5898, 5804, 5560, 5810, 6486, 7709, 10649, 10021, 10018]. Carried [8856]. Carrier [251, 2585, 5230, 8849, 8607]. Carriers [1147, 8082]. Carry [6347]. Carrying [2656, 8891]. Cars [1151, 345, 1174, 471, 162, 259, 1817, 1650, 2743, 3481, 4472, 4408, 4391, 4687, 5011, 5100, 4428, 4468, 5891, 6238, 5851, 5721, 5463, 5387, 6699, 6448, 6869, 8693, 8634, 10047, 10158, 9824, 9706, 9790, 9707, 10167, 10779, 9882, 10390]. Cart [1750, 2524, 4824, 6478]. Carter [366, 6165]. Carthaginian [7954]. Cartridge [4590, 10770]. Cartridges [5614]. Carved [3406]. Carving [560, 865, 7315]. Casanova [1855]. Cascade [5008, 7558]. Case [289, 5940, 7902, 9540, 10125, 9830, 9656, 3175, 8032, 7589]. Cases [91, 5343, 10646]. Cash [251, 2585]. Casing [1515, 7714, 8788]. Casings [3552]. Casino [1846]. Cask [6205, 9990]. Casks [9097]. Cassadaga [7225]. Casseuil [5858]. Cast [5287, 5458, 9583]. Cast-Steel [9583]. Castine [4864]. Casting [1825, 2148, 3401, 6062, 5576, 8326, 9251]. Castings [8144]. Castle [4919, 7884]. Castner [10698]. Castrography [1717]. Cat [2921, 4530]. Catacombs [7134]. Catalina [10727, 10264]. Cataract [7693]. Cataracts [7828]. Catastrophe [6029]. Catch [2345, 5173]. Catcher [670, 2887, 8069, 8089, 8155, 8453]. Caterpillar [945, 5070]. Cathedral [7384, 9691]. Cats [258, 3312]. Cattle [689, 2523, 3867, 4007, 4145, 3912, 8740]. Caucasus [8336, 9832]. Cause [7862, 9884]. Caustic [2503]. Cave [1915, 6919, 3302]. Caves [7414, 7735]. Cavities [3509]. Caw [4005]. Cedar [4009]. Ceiling [807, 650, 5423]. Ceilings [4191]. Celebration [1321, 1539, 1588, 3575, 5824, 7188, 1621, 3036]. Celery [604]. Celestial [16, 244]. Cell [8388, 10103, 5137]. Cellars [7220]. Cells [6219, 3389]. Cellular [10725]. Cement [1210, 626]. Census [849]. Cent [6446]. Centenarian [1466]. Centennial [981, 1169, 1571, 1588, 1621, 2809, 6821, 6951, 7756, 7396, 6791]. Center [807, 2093, 5271, 5873, 9133, 10074]. Center-Piece [807]. Centerboard [10453, 10207]. Centering [5760]. Central [151, 1791, 1946, 2817, 2531, 4204, 4876, 6750, 7487, 8314, 8367, 7518, 8388, 8275, 9431, 9841, 1877, 5088, 5415, 6612, 8889]. Centrales [115]. Centralization [3785]. Centrifugal [6668, 8989, 4715, 1424]. Century [1321, 1539, 6204, 6961, 9515, 10787, 10750, 3942, 10471]. Cephalopods [9939]. Cerda [3543]. Ceremonies [2534]. Certain [3402]. Cervera [9087, 9142]. Ceylon [505, 4840]. Chacma [2291]. Chain [35, 3849, 4798, 6275, 6688, 6399, 10793, 10010, 9832]. Chain-Less [10793]. Chainless [8167, 8280, 8299, 9588, 10520]. Chains [5248, 5854, 6067]. Chair [177, 2846, 3588, 4101, 3246, 5534, 6081, 8043, 8044]. Chairs [548, 1852, 3037, 3722]. Chalcedony [2835]. Chalk [6419]. Chalker [4751, 5102]. Challenger [10150]. Chamfering [1712]. Champagne [1241]. Championship [4110]. Champlain [9306]. Chance [3998, 3541]. Chances [10441]. Chandler [9974]. Chang [7065, 7076]. Change [9429, 8658, 10778]. Change-Making [8658]. Changer [3496]. Changes [3718, 4402, 7656, 7401, 8293, 8416, 8234, 8653, 8284]. Channel [449, 313, 4782, 9950]. Channeling [44]. Channels [4280]. Chaos [2762]. Chapel [7989, 4128, 7884]. Chapman [10495]. Character [7197]. Characteristic [5515, 8905, 8916, 1951]. Characteristics [9818]. Charcoal [5619]. Charcot [4036]. Charges [7899]. Chariot [10547]. Charity [7761]. Charles [3125, 5037, 8615, 9913, 1708, 2554, 3291, 6109]. Charleston [1772]. Charlton [7597]. Charm [10002]. Charter [5082]. Charts [8496]. Chase [4691, 6502]. Che [5990]. Che-Foo [5990]. Cheap [184, 4100, 4161, 3407, 4258, 3866, 4004, 4465, 7313, 9959, 9905]. Cheaper [9507]. Check [596, 205, 347, 296, 4954, 6125, 5652, 6901, 6562]. Checking [7470]. Chelan [10509]. Chemical [2032, 1258, 2335, 2579, 3402, 4379, 9324, 10082, 2152]. Chemicals [2472]. Chemist [5592, 5492, 5346]. Chemistry [6942, 10724]. Chest [1231, 4163]. Chicago [3542, 4138, 4157, 4183, 3743, 3786, 3805, 3823, 3843, 3863, 3885, 3908, 3932, 3964, 3976, 4053, 4079, 4098, 4118, 3587, 5141, 660, 745, 719, 221, 2185, 1792, 3068, 2438, 3096, 3444, 3697, 3723, 3661, 3495, 3706, 3777, 3842, 3520, 3565, 3611, 3653, 3676, 3721, 3608, 3260, 4327, 4408, 4370, 4943, 4755, 5218, 6101, 5776, 5805, 6281, 6163, 6278, 5684, 5654, 5445, 6522, 7588, 7647, 9039, 10411, 10239, 3806]. Chickamauga [8940]. Chicken [4237, 9906]. Chicken-Brooder [9906]. Chicory [4471]. Chief [9806, 9528]. Chiefs [9795]. Chignecto [1596]. Children [179, 1694, 5890, 10565, 4450, 9705]. Chile [1280, 5157]. Chilean [1772]. Chilian [1280, 1444]. Chilton [7644]. Chime [2675]. Chimes [1160, 3028, 7989]. Chimney [3215, 3807, 3307, 5307, 5695, 5495, 6401, 6582, 6679, 7736, 7739, 10421, 372]. China [3724, 4935, 6188, 5464, 7076, 9511, 9051, 9405, 8558, 10296]. Chinch [5998]. Chinese [387, 3304, 4921, 5315, 5405, 8408, 10101]. Chisel [3129, 6151, 5565]. Chlamydosaurus [6636]. Chlorination [198]. Chloroplatinite [10042]. Choctaw [2901]. Choice [5378, 5417]. Cholera [3002, 3001, 3201]. Chopper [154, 1331, 2415]. Christian [873]. Christopher [1016, 2881, 3622]. Chromatophores [6219]. Chromes [1262]. Chromo [10072]. Chromo-Camera [10072]. Chronograph [2763]. Chronophotographic [7720]. Chuck [1961, 3475, 5631]. Chucks [4731]. Chupaderos [5799]. Church [1499, 5013, 5033, 5202, 6048]. Churn [395, 1498, 2227, 5555]. Cicada [4829]. Cicadas [5125]. Cigar [2547, 5309, 5196]. Cigarette [8618, 4689]. Cigarettes [6061, 8817, 8945, 8733]. Cincinnati [3123, 3506, 5051, 4905, 6048]. Cincture [1035]. Cinematograph [7905, 9127]. Circle [7982]. Circles [7233]. Circuit [394, 10103]. Circuits [144]. Circular [3938, 8930]. Circulating [4123, 4883, 7685]. Cistern [1002]. Cisterns [3975]. Cities [2022, 9030]. City [581, 592, 1966, 1272, 3983, 5044, 4444, 5032, 5276, 5480, 5890, 7886, 7993, 7827, 10188, 10599, 7122, 7348, 478, 528, 1148, 776, 698, 1039, 1825, 1771, 1468, 1251, 3159, 3131, 2382, 2399, 2531, 3884, 3370, 4984, 4799, 5273, 4593, 4673, 4705, 4773, 5295, 5176, 5406, 5896, 5539, 5541, 5500, 6436, 7254, 6387, 6612, 6461, 6809, 7039, 7147, 6970, 6516, 6745, 6415, 6496, 6501, 8304, 7394, 8111, 7897, 7447, 7717, 7455, 7476, 7798, 7501, 7741, 7518, 8260, 8425, 8085, 8015, 7981, 8071, 7612, 7769, 8628, 10637, 9626, 10452]. City [9829, 10661]. Civil [100, 1761, 3283, 8405, 8038, 8600, 8783, 8713, 9628]. Civilized [9270]. Clad [423, 10748]. Clair [765, 892, 2077, 1815, 2293]. Clam [9520]. Clamp [134, 470, 226, 1297, 5291, 5853, 6976, 6787, 8363, 8766, 8679]. Clamped [2171]. Clark [7906, 9467, 372, 7875]. Clasp [1544]. Class [2629, 6100, 7030, 7011, 7643, 9498, 9372, 9993, 10556, 10771]. Classes [7311, 8439]. Clay [7220, 9572]. Clay-Cutting [9572]. Claypipe [10564]. Cleaned [4750]. Cleaner [449, 1135, 913, 2246, 2661, 3807, 6005, 8262, 9139, 10010]. Cleaning [887, 601, 314, 2036, 1490, 2582, 3884, 7116, 6656, 8150, 7708, 7823, 8244, 9553, 9410, 9030]. Cleansing [4857]. Clepsydra [5485]. Cleveland [1079, 10119]. Clevis [716]. Climate [8233]. Climatic [1987]. Climax [4584, 5699, 7118]. Climbing [372]. Clip [37, 2018, 9114]. Clock [3028, 2938, 3844, 5967, 6382, 7989, 10376, 9986]. Clocks [2953, 5782, 6481]. Clockwork [5345]. Clogs [2321]. Close [3052, 6349, 7252]. Closed [5016]. Closer [394]. Closet [2274, 3054]. Closing [584, 1233, 4155, 5985, 8832, 9266, 9269]. Closure [9118]. Cloth [537, 1625, 4822, 9988]. Clothes [472, 761, 645, 2092, 2417, 3681, 5289, 6295, 8958]. Clothes-Drying [3681]. Clothing [6920]. Cloud [5502]. Clowes [2137]. Club [3529, 5142]. Clutch [43, 1018, 4103, 5193, 7183, 6380, 7473, 10569]. Clutch-Pulleys [10569]. Co [1791, 4027, 7218, 556, 5044]. Co-Operation [7218]. Co. [4073, 4634]. Coach [3349, 6978]. Coal [889, 556, 2245, 1675, 2487, 2338, 3894, 3701, 3247, 4299, 4472, 4391, 4799, 4790, 4953, 6245, 6078, 6066, 6098, 6611, 6748, 8867, 9814, 10156]. Coalbrookdale [1789]. Coaling [4156, 7681, 9272, 10160]. Coast [1145, 55, 1641, 2771, 3458, 7251, 6823, 6949, 6793, 9178, 9069, 10614, 9674, 10004, 8496, 8878, 9690]. Coast-Defense [9178, 9674]. Coaster [10625, 10028]. Coasts [5563]. Coastwise [6024]. Coat [39]. Cock [1418, 4667, 5200]. Cockerel [8705]. Cocoanut [4287]. Cocoons [4392]. Cod [1122, 6024]. Codonophone [1183]. Coffee [4671, 6059, 7185, 9757]. Coffin [2931]. Coffin-Lowering [2931]. Cohoes [4324]. Coil [3409]. Coin [3178, 4236, 9031]. Coin-Controlled [9031]. Coin-Taking [3178]. Coins [10393]. Cold [2972, 5981, 5869, 3824, 4470, 7613, 7450, 10517]. Coliseum [6101, 10411]. Collapse [6101, 6505, 9559, 9290, 9656]. Collar [624, 3197]. Collars [5170, 6067]. Collecting [8989]. Collection [1089, 1941]. Collections [7333]. Collector [1849, 4493]. Collectors [7602]. College [976, 4769, 5500, 8952, 2772, 3504]. Collipulli [1280]. Collision [1946, 1264, 2313, 6567, 7085]. Collisions [4039, 9150, 9172, 10777, 10186]. Colomb [9840, 10586]. Colomés [9197]. Colon [3543]. Colonel [3839, 9408, 9134]. Colonial [4478]. Colonies [9677]. Colonization [4684]. Colony [961]. Color [2703, 3930, 6219, 5483, 8284, 3416, 1987]. Color-Bestowing [6219]. Colorado [2747]. Colored [4492]. Coloring [10665, 4567]. Colors [1647, 4298, 7551, 9132, 8868, 10105]. Colossal [1016]. Columbia [2956, 2870, 4231, 4266, 4244, 4404, 5234, 5956, 6000, 7095, 8299, 7790, 7618, 8809, 8771, 8528, 10021, 10337, 10534, 10368, 10169, 10572, 10129, 9976, 10712, 9914]. Columbian [2297, 2975, 3059, 4155, 4173, 3844, 3962, 3756, 3520, 3542, 4138, 4157, 4183, 3565, 3611, 3653, 3676, 3699, 3721, 3743, 3767, 3786, 3805, 3823, 3843, 3863, 3908, 3932, 3964, 3976, 4000, 4026, 4053, 4079, 4098, 4118, 3608, 4145, 3677, 3633, 3698, 3305, 3564, 3609, 3636, 3882, 3917, 3944, 3963, 3973, 3984, 3985, 3987, 4015, 4035, 4064, 4109, 4120, 4123, 4126, 4193, 4217, 4370, 7647, 4127, 1459, 1837, 3036, 2314, 3068, 2863, 4288, 4264, 3599, 3622, 4267, 3824, 4262, 4229, 3610, 3544, 3885, 3755, 4136, 4181]. Columbian [3587, 3612, 4210, 3758, 3917, 3650, 4478, 4529, 4455, 10788]. Columbus [1016, 1691, 2881, 2696, 2327, 3089, 3042, 3027, 2862, 3575, 3622, 3809, 10419, 10445]. Column [1713]. Combat [1773]. Combination [1519, 2607, 2986, 4771, 5621, 6604, 6747, 8132, 10025, 9636, 9703]. Combination-Tool [9636]. Combined [154, 954, 1461, 1253, 1982, 3023, 3889, 4282, 4879, 5379, 5402, 5847, 6852, 6668, 7181, 8912, 9112, 10568, 10280]. Combs [6080]. Combustion [8539]. Comet [743, 3143, 6145, 7031, 4213, 4012]. Comets [7066]. Comfortable [2846, 3654]. Coming [3404, 3539, 5154, 5538]. Command [8841]. Commerce [1226, 3383, 5805, 8033, 8335, 8601, 8661, 9051, 9092, 8758, 9192, 10644, 4231]. Commercial [390, 3697, 5476, 5803, 7076, 7856, 10297, 10094, 10136, 9604]. Commission [6970, 8377, 9487, 10214, 10443, 8335, 9448, 10137, 10116, 9791]. Commissioner [7808, 8615, 174, 2413, 3343, 4251, 4606, 5493, 6746, 7633, 7682, 8573, 8652, 8762, 9770, 10272, 10635, 10638, 10276]. Commissionership [1887]. Commodious [8848]. Commodore [8995]. Common [901, 9457]. Communication [8945]. Communications [2998, 5148, 5164, 5183, 9007]. Communities [10080]. Community [570]. Commutator [4102]. Commutators [4447]. Compact [1297, 1944, 2306, 3660, 3264, 5386, 5557, 10567]. Companies [9111]. Companion [6370, 7281]. Company [2728, 4005, 4075, 4029, 4123, 4569, 5539, 7008, 9431, 961, 2542, 3417, 3554, 4166, 3959, 4259, 3870, 3844, 4061, 3268, 4210, 4192, 4608, 4984, 5032, 5276, 4923, 4822, 4803, 4370, 4645, 4825, 4338, 5219, 4847, 4511, 5366, 5945, 6143, 5733, 7171, 7758, 7880, 8356, 9417, 10204, 7715]. Comparative [6497, 9744, 10250, 8641, 8660]. Comparatively [4402]. Compared [8711, 10398, 10797]. Comparison [6339, 7235, 8313, 8793, 10158, 10609, 10337, 10478, 9699, 7066]. Compass [4940, 5269]. Compensation [3412]. Competition [6726, 6781, 6821, 6515, 7797, 8897, 6898, 7896, 7855]. Competitions [5933]. Competitive [7394]. Complete [10219]. Completion [173, 8179, 8060, 8710]. Composing [1240, 3865]. Composite [4895, 6396]. Composition [10258]. Compound [78, 456, 2211, 2430, 2579, 2691, 3845, 4210, 4826, 5593, 6010, 5986, 5347, 6479, 9853]. Compound-Nickel [2579]. Compressed [133, 345, 6805, 6803, 7008, 7434, 7366, 7568, 9626, 10452, 10715, 10512, 9662]. Compressor [2722, 4165, 4210]. Compressors [4654]. Compromise [4337, 7967]. Comstock [9752]. Concave [16]. Concentrating [8564]. Concentration [7508, 9191]. Concentrator [527, 1795, 4081, 5379, 8151, 8297]. Concerning [2522, 5377, 9754]. Concerts [1439]. Concord [1120]. Concrete [1109, 971, 3554]. Concurrent [1399]. Condensation [7294]. Condensed [1229, 675, 3958]. Condenser [785, 2361, 4186, 6279, 8994]. Condensers [2721]. Condenses [8126]. Condensing [2505]. Condition [2035, 1709, 2736, 2746, 8439, 8593]. Conditions [6880, 8533, 9484]. Conductor [4791]. Conductors [4651, 10619]. Conduit [162, 1019, 4724, 5896, 5408, 10619]. Conduit-Crossing [10619]. Cone [2974, 6651]. Cones [4392]. Coney [2021]. Conference [268, 8736]. Confidence [7233]. Congealing [3107]. Congo [5542]. Congress [9678, 8336, 10419, 10445, 8082, 1271, 1866, 1274, 2485, 4096, 3446, 3841, 6916, 7486, 8149, 7947, 7759, 9356, 9869, 10708, 9720, 9629, 10508, 9898, 8213]. Congressional [2522, 2413, 2601, 3883, 7232]. Conical [3571]. Conjoined [3075]. Conn [1291, 3132, 5033, 416]. Connected [6485]. Connecticut [1524]. Connecting [1544]. Connection [559, 4242]. Connections [3548]. Conning [8770]. Conquers [8040]. Conservation [10308]. Considerations [10098]. Considered [8641, 8660]. Consolidated [8537]. Constitution [1319]. Constrictor [6526]. Constrictors [2007]. Constructed [6127, 8662]. Constructing [1109, 9660]. Construction [696, 1159, 2002, 3493, 4191, 3700, 4284, 4822, 4938, 6820, 6418, 7273, 8166, 8388, 8551, 9336, 8531, 9929, 10321, 9984, 10169, 10388, 10161, 10193]. Constructor [9528]. Constructors [9899]. Consular [10763, 9775]. Consuls [10299, 10345, 9951]. Consume [3022]. Consumers [9851]. Consuming [7738]. Consumption [1078, 1099, 1119, 1158]. Contagion [2071, 4663]. Containing [6868]. Conte [2916]. Contending [4110]. Contents [1618, 1642]. Contest [1908, 4060, 6648, 6163]. Continent [299, 7104]. Contingent [4256]. Continuous [34, 2222, 3864, 4406, 5759]. Contract [9947]. Contraction [5556, 7521]. Contractors [3998]. Contrast [10424]. Control [5284, 10595]. Controlled [5807, 5498, 8420, 9031]. Controller [6517, 9706]. Controlling [4410, 5874]. Controversy [2]. Convenience [4034]. Convenient [1171, 887, 1600, 2207, 2706, 2433, 2950, 2811, 3744, 3913, 4080, 3978, 5085, 4751, 4757, 5574, 5460, 7902, 7410, 7781, 7517, 8676, 9450, 2815]. Conveniently [5534]. Convention [1761, 1272, 842, 3177, 7486]. Conventions [9552]. Conversion [7871]. Convertible [9935, 9957, 9571]. Conveyance [2165]. Conveyer [553, 3894, 3245]. Conveying [2276, 9778].

Convict [9783]. Cooker [815, 4099]. Cooking [4257, 3262]. Cool [1230, 1128, 1962]. Cooler [1692, 3501, 4031, 9138, 10517]. Coolgardie [7826]. Cooling [696, 8176, 10422]. Copan [7348]. Cope [7028, 7695]. Copenhagen [6263]. Copier [6237]. Copies [3821, 7272]. Copper [4027, 4033, 8893, 10393, 7102, 7120]. Coppered [10352]. Coppersmithing [6828]. Copy [1172, 3451, 8685]. Copy-Holder [8685]. Copyholder [866]. Copying [2883, 4918, 5633, 6827]. Copyright [7635, 8468, 8846, 9650, 466, 1274, 6580]. Copyrights [9405]. Cordite [8032, 7589, 10355]. Corean [4529]. Cork [1152, 7593, 10715]. Cork-Drawing [1152]. Corks [8129]. Corliss [4185]. Corn [79, 5330, 5462, 5536, 6677, 6630, 10465, 10643]. Cornell [976, 2152, 5121, 5919]. Corner [3094, 8132]. Cornwall [9317]. Corona [7066]. Coronation [6846]. Corporation [4903, 5108]. Correct [3955, 5715]. Corrected [1995]. Correlative [9369]. Correnpondence [6190]. Correspondence [15, 142, 872, 894, 919, 941, 960, 997, 1017, 1037, 1059, 1086, 1106, 1157, 1186, 1212, 1236, 160, 183, 210, 234, 252, 276, 298, 321, 373, 400, 453, 477, 503, 534, 555, 718, 742, 766, 788, 817, 1259, 1279, 1895, 1916, 1931, 1949, 1968, 2006, 2023, 2042, 2060, 2080, 2094, 2118, 2138, 2151, 2168, 2186, 2209, 2230, 2253, 2292, 1302, 1328, 1353, 1373, 1398, 1420, 1469, 1570, 1598, 1622, 1678, 1695, 1715, 1731, 1752, 1770, 1790, 1811, 1857, 1873, 2326, 2358, 2888, 2919, 2957, 2976, 2990, 3012, 3044, 3056, 3130]. Correspondence [3147, 3166, 3183, 3204, 2381, 2402, 2419, 2437, 2454, 2471, 2493, 2512, 2552, 2567, 2625, 2646, 2657, 2672, 2695, 2761, 2782, 2798, 2851, 2874, 3232, 3269, 4010, 4084, 4146, 4194, 4214, 4261, 4283, 4326, 3292, 3310, 3332, 3351, 3374, 3390, 3411, 3457, 3483, 3505, 3527, 3550, 3573, 3597, 3641, 3682, 3707, 3776, 3792, 3810, 3896, 3916, 3945, 4346, 4369, 5066, 5087, 5104, 5122, 5197, 5233, 5251, 5272, 5312, 5332, 4432, 4452, 4496, 4552, 4568, 4592, 4619, 4652, 4690, 4714, 4754, 4777, 4828, 4848, 4884, 4906, 4922, 4992, 5012, 5050, 6190, 5365]. Correspondence [5384, 5969, 5988, 6011, 6049, 6064, 6086, 6132, 6150, 6224, 6284, 6302, 6323, 5401, 5426, 5443, 5517, 5540, 5562, 5618, 5635, 5653, 5668, 5720, 5741, 5767, 5811, 5836, 5856, 5876, 5900, 5923, 5944, 6365, 6407, 7119, 7152, 7186, 7204, 7222, 7241, 7280, 7334, 6427, 6464, 6525, 6543, 6564, 6586, 6610, 6631, 6694, 6715, 6733, 6766, 6832, 6888, 6903, 6977, 7010, 7041, 7084, 7101, 7382, 7475, 7498, 7560, 7577, 7646, 7766, 7812, 7907, 7932, 8000, 8047, 8178, 8283, 8322, 8342, 8364, 8386, 8406, 8451, 8476, 8517, 8543, 9415, 9434, 9475, 9497]. Correspondence [9518, 9536, 9558, 8583, 8613, 8637, 8700, 8745, 8769, 8855, 8875, 8938, 8960, 9060, 9075, 9117, 9157, 9232, 9268, 9292, 9314, 9335, 9355, 9392, 9574, 9591, 10148, 10261, 10282, 10309, 10334, 10351, 10378, 10430, 10451, 10476, 10571, 10594, 10621, 10672, 10728, 10749, 10772, 9612, 9640, 9663, 9686, 9710, 9731, 9804, 9828, 9856, 9883, 9909, 9937, 9962, 9991, 10051, 10077, 10106, 10127, 9582, 10746]. Correspondents [1633]. Corridor [8388]. Corrosion [8154, 10232]. Corrugated [3592, 4490]. Cosmical [1818]. Cost [9773, 685, 346, 1050, 832, 9553, 8592, 9743, 10250, 10139]. Costly [8421].

Cotes [2058]. Cotta [8931]. Cottage [5161, 7888, 8109, 7122]. Cottages [8521]. Cotton [154, 981, 467, 342, 605, 1331, 2070, 3093, 3185, 3790, 3709, 4542, 5005, 5959, 8296, 8787, 9339, 9407, 1184, 6099, 6143]. Counter [767, 4729, 6058, 7293]. Counterbalance [8088]. Counterbalanced [1136]. Counterweighted [7270]. Counting [4001]. Countries [7613, 8889]. Country [26, 7299, 8059, 8323, 9062]. County [593, 2170, 4645, 8519, 9539]. Couple [5193]. Coupled [9451]. Coupler [1127, 1728, 2119, 4957, 5595, 6804, 8291]. Couplers [2267, 2777, 3424]. Coupling [9, 107, 135, 155, 203, 294, 338, 498, 600, 620, 625, 995, 1104, 275, 369, 809, 935, 811, 1105, 1812, 1971, 2038, 2110, 2379, 2848, 2912, 2757, 3499, 3528, 3808, 4688, 5027, 5533, 5596, 5942, 6105, 6884, 6671, 6560, 6540, 8325, 7599, 8829, 10047, 10279]. Couplings [6728, 8788]. Courage [9068]. Course [875]. Court [4477, 7683, 10522, 334, 4193, 5532, 8814]. Courtesy [7732]. Courthouse [10522]. Covent [10491]. Cover [974, 2223, 4030, 5695, 8381, 9054]. Covered [6484]. Covering [181, 6148, 9494]. Cowl [5307, 6582, 7736, 10421]. Coyote [10699]. Crackers [636, 4594, 5701]. Cracks [5459]. Cradle [9249]. Cramp [1345, 5324, 8603]. Crane [1033, 744, 2041, 2674, 3623, 4596, 5945, 9454]. Cranes [3696]. Cranford [5161]. Crank [10165, 3450, 4819, 7595, 8789]. Cranks [3886]. Crankshafts [10329]. Crate [1378]. Crater [7289]. Cream [4586, 4907]. Creamery [3980]. Crepitans [7223]. Creusot [6045, 10742]. Crew [2394, 5919]. Crib [6811]. Cricket [9181]. Crime [8233, 10750]. Criminal [4477]. Crimping [5893]. Crippled [2736]. Crisis [6338]. Cristobal [3543]. Criterion [9155]. Critic [9624, 9894]. Criticism [9301, 10412]. Critics [1345]. Crookes [6488, 7552, 6789]. Crop [7661, 8362, 8296, 10465]. Crosby [4073]. Cross [607, 2800, 4233, 7463, 9120, 10449]. Cross-Cut [10449]. Cross-Cutting [2800, 4233]. Crossing [1542, 3635, 5415, 7693, 8020, 10619]. Crossings [138]. Croton [674, 694, 1805, 2790, 5121, 7168, 8614, 5971]. Croup [5211]. Crowded [2746]. Crowding [9790]. Crowning [3344]. Croze [1084]. Crozier [6566, 7974]. Crucible [3959]. Crude [2562, 3452, 5192, 9667]. Cruiser [1132, 302, 42, 637, 996, 2278, 2285, 1249, 1773, 1745, 2629, 3123, 2647, 3506, 3507, 3461, 3334, 4453, 4347, 4476, 5051, 6152, 5956, 5342, 6000, 6577, 7448, 8969, 8603, 8853, 9315, 9116, 9100, 9594, 10003, 10239, 9873]. Cruisers [423, 4156, 7918, 9538, 8906, 10684, 10660, 10556, 10352, 9897, 10344, 10366, 10398, 10252]. Crush [5789]. Crushed [2187]. Crusher [2469, 4749, 5361, 6748]. Crushing [5379]. Crustacea [5747]. Crustaceans [8879]. Crystal [1217]. Crystallization [9973]. Crystallized [747]. Cuba [6645, 9062, 8866, 9231, 9018, 8747, 8981, 9248, 10463, 9613, 10143, 3353]. Cuban [6012, 5857, 9059, 8803]. Cuirass [4882]. Culinary [862]. Cultivating [8787]. Cultivator [154, 342, 106, 230, 1331, 2166, 2416, 4735, 5381, 7598, 10590]. Culture [520, 1926, 9466, 6725]. Cunard [3417, 3294, 4052]. Cunningham [2384]. Cup [6076, 10761, 9807, 4125, 10583, 10150]. Cupid [9218]. Cups [2783].

Cure [1078, 5211, 4607, 8555, 1158]. Curing [292]. Curios [8585]. Curiosities [5345]. Curious [1397, 2506, 3615, 3752, 5254, 4800, 4675, 5382, 6505, 6865, 6404, 8037, 8418, 9540, 9751, 9830, 9821, 9615, 9539]. Curling [2451, 5830]. Current [172, 3864, 4493, 4726, 6280, 7411, 8629, 9902, 10518, 6678, 8144]. Currents [2528, 5007, 5920, 165, 3409]. Curtain [272, 473, 846, 853, 6438, 6917]. Curtain-Pole [846]. Curtains [5939]. Curves [7721]. Cushion [6829, 8473, 10144, 6321]. Cushioned [5630, 8198]. Cushions [4160, 5387]. Custom [8048]. Cut [12, 1693, 2979, 2510, 4158, 3955, 6447, 10449, 2872]. Cut-Loose [2872]. Cut-Off [12, 1693, 2510, 4158]. Cut-Out [6447]. Cutlery [5816]. Cutter [1176, 228, 1180, 550, 1253, 3889, 3812, 3594, 6240, 5537, 6124, 6262, 6729, 6806, 9061, 8976, 9681, 10591, 9985]. Cutter-Stocks [9681]. Cutters [10729]. Cutting [575, 525, 573, 1080, 1886, 1474, 2198, 2584, 2800, 4233, 4159, 3229, 5086, 5461, 6241, 8196, 7617, 10645, 9572, 9758, 10259, 1870]. Cutting-off [3229]. Cyanide [7234]. Cycle [1060, 1751, 3368, 4407, 5214, 5416, 6688, 7316, 7048, 6654, 9709, 9653, 10536, 10257]. Cycles [4508]. Cycling [3023, 10180]. Cyclist [6387]. Cyclists [8419]. Cyclometer [4770, 5467, 6439]. Cyclometers [10011]. Cyclone [350, 9351]. Cyclones [3225]. Cyclopedia [10436]. Cyclorama [6502]. Cylinder [992, 2362, 3401, 4361, 8529, 8780]. Cylinders [4185, 7366]. Cylindrical [5758]. Cyphers [9412]. Cyrus [2833]. Czar [6846].

D [5656, 10369]. D. [6514, 9449]. D.C [8683, 8895]. D.C. [6584, 7232, 8582]. Daguerre [991, 842]. Dahlia [5428]. Daily [3540, 9547]. Daimler [7208, 6425]. Dairy [3980]. Daisies [920]. Daisy [3946]. Dakota [1038, 1813, 7049]. Dalmeny [10736]. Dam [1950, 1344, 2687, 4324, 5176, 5987, 8201, 8614, 9537, 1805, 2991, 2790, 5121, 7168, 6988]. Damage [9594]. Damages [3307]. Damper [1202]. Dana [5656]. Dance [6853, 10168, 10425]. Dancers [6907]. Danforth [3175]. Danger [10462, 10739]. Dangers [2180, 4726, 9279]. Daniel [4973]. Danish [4586]. Danube [1303, 7686]. Daring [10014]. Dark [8242]. Darning [10235]. Darragh [1627]. Dasher [395]. Dashers [5555]. Dashiell [8876, 9794]. Data [9402]. Date [1228]. Dater [5646]. Dating [429, 2932]. Daugherty [5761]. Davis [6025]. Davison [4918]. Davoust [4630]. Dawson [10717]. Day [4408, 6835, 7942, 8518, 10043]. Days [4901, 8532]. Dead [1594, 9821]. Deadly [3600, 4424]. Deal [5178]. Death [4205, 3673, 5275, 8037, 7875, 7695, 7922, 7994, 8062, 7943, 9408, 9449, 9485, 8576, 8772, 10586, 10369, 10142, 9870, 10650, 9793, 9629, 9794, 10040, 10370, 10613, 10717, 5178]. Debating [2015]. Debris [1889]. Debt [6314]. Decadence [6763]. Decapitated [7262]. Decauville [10023]. December [1123, 102, 2244, 3165, 5247, 6257, 7292, 8400, 9491, 10719]. Deceptions [1995]. Deceptive [8083]. Decided [1344]. Decimalization [8518]. Deciphered [8039]. Decision [334, 1009, 5325, 4817, 4629, 5532, 7992, 8073, 8595, 10511, 5734, 10792]. Decisions [91, 271, 594, 1064, 661, 1669, 2589, 2978, 4728, 4995, 5062, 5140, 5746, 5344, 5447, 5522, 5826, 5938, 6002, 6090, 6990, 7067, 7513, 7553, 9110]. Deck [2849, 6699, 6863, 7718, 10612]. Decoration [6835]. Decreasing [1050]. Dedication [3068, 2534]. Deep [2095, 2905, 9983]. Deep-Mine [9983]. Deepen [4280]. Deepening [7944]. Defective [1030, 2826, 6522, 10565, 7531]. Defender [5722, 9807]. Defense [1145, 1156, 443, 3342, 3458, 7251, 6823, 6949, 6793, 9178, 10583, 9674]. Defenseless [1709]. Defenses [805, 806, 9069, 9018, 9136, 10614, 10004]. Defensive [1643]. Defraud [8357]. Dehorning [1593]. Delaware [556]. Delay [3821, 7272, 8189]. Delayed [5343]. Delcroix [4171]. Delicacy [3402]. Deliciosa [4802]. Delivering [3178, 2338, 4236, 8974]. Delivery [3618, 8069, 8089, 8155, 8844, 10250, 10025, 9912]. Delphi [7068]. Delusion [10744]. Demand [2303, 7466]. Densmore [7668]. Dental [530, 667, 1676, 2448, 5780, 5497, 6731]. Dentistry [2137]. Dents [9989]. Denudation [7980]. Denver [10556]. Deodorizing [2020]. Department [2872, 7427, 7823, 8491, 9280, 10642, 8641, 8660, 1734, 8628, 9244]. Departure [870, 6848, 9009, 10625]. Depoele [2554]. Deposit [8654]. Depot [151, 1400]. Depths [4387, 2873]. Derelict [5067]. Derrick [1049, 1511, 5218, 6104]. Deseret [5628]. Deserted [5694]. Desertions [2163]. Desiccated [4287]. Design [1016, 698, 2226, 3024, 3097]. Designed [9359, 10240]. Designing [1030]. Desirable [8353]. Desk [1014, 1676, 2608, 3246, 3750, 7781]. Desroziers [3864]. Dessau [6448]. Destined [8747]. Destroyer [2618, 4231, 4736, 5089, 6246, 6831, 8427, 9017, 9684]. Destroyers [8747, 10648]. Destroying [6871]. Destruction [1376, 2472, 3080, 3824, 9087, 9131, 8814, 10599]. Destructive [1524, 6706, 7978, 10721, 6838]. Detachable [6146]. Detacher [6669]. Detaching [6476]. Details [8462, 9336]. Detected [7374]. Detecting [10640]. Detective [2887, 6788]. Determination [7189]. Determinations [6996]. Determining [2706, 8854]. Detroit [3965, 7829]. Deutschland [8002, 10329, 10196]. Developed [2713]. Developing [1231, 9391]. Development [606, 6957, 8221, 8759, 9830, 6547]. Developments [9489]. Device [1152, 346, 444, 1174, 1230, 1231, 1061, 910, 595, 1128, 760, 762, 756, 1101, 644, 314, 578, 1233, 1981, 2036, 1765, 1786, 1351, 1872, 2037, 2288, 1489, 1371, 1521, 1543, 1395, 3052, 2966, 2322, 3071, 2582, 2969, 2741, 2935, 2341, 3313, 4236, 3978, 3933, 3935, 3681, 3367, 3728, 3747, 4713, 4823, 6085, 5557, 5760, 5558, 5960, 6003, 6282, 5362, 5830, 6168, 6711, 6604, 6442, 6476, 6670, 6886, 6562, 8195, 8092, 7928, 8912, 8703, 8891, 8664, 8993, 9230, 9269, 8658, 8722, 8580, 9803, 9989, 10722, 10281, 9931, 10673]. Device [10779, 9661, 3684]. Devices [544, 2245, 2357, 3822, 8218, 8069, 8089, 8155, 9410]. Dewar [6652, 9652]. Dewey [9078, 10454, 10354, 9639]. Dials [6481]. Diamond [343, 1886, 1603, 2566, 3368, 4011, 4508, 9602, 1932, 5359]. Diego [255]. Dies [10447]. Diesel [8721]. Dietary [8419]. Differential [1831, 8513, 9032]. Difficulties [2902, 4540, 9313]. Difficulty [7252]. Diffraction [1681]. Digger [643, 1015]. Digging [1011, 4280]. Diligence [9606]. Dion [10024, 7171]. Diphtheria [5211, 7613]. Direct [3148, 6485, 7871]. Direct-Connected [6485]. Directions [7665]. Director [8737]. Directories [5963]. Dirigible [7139, 9395]. Dirigo [4550]. Dirt [6478]. Disabled [10480]. Disappearance [617]. Disappearing [2658, 2852, 4168, 6045, 6566, 7974, 9057]. Disaster [5437, 8314, 8671, 8926, 8723, 8799, 10147, 6883, 10350, 10367]. Disasters [9188, 9111]. Disbarment [8232]. Disbarred [8211]. Discharge [7169, 8697]. Discharger [5007]. Discipline [9188]. Disclosed [1319]. Discourage [3303]. Discovered [5924, 8749, 9097, 6904]. Discoveries [8345, 9324, 8655, 10623]. Discovery [3688, 5979, 7068, 6658, 6456, 6498, 7877, 7637, 9822]. Discrimination [2560]. Discussion [10690]. Disease [2523, 3285]. Diseases [9605]. Dish [3179]. Dishes [7600]. Dishonest [7374]. Disinfectant [4524, 9285]. Disinfecting [6539, 7138]. Dismal [8702]. Dispatch [1147]. Dispatches [2165]. Dispeller [9981]. Dispersion [3757]. Display [1014, 4262, 4474, 10686]. Disposal [8111]. Disposed [5762]. Disposes [9370]. Disposing [9821]. Disposition [1889]. Distance [1327, 1439, 2494, 3010, 3345, 4768, 5832, 5514, 6202, 5455, 7601, 7532, 8854, 9234]. Distances [3787]. Distillation [2738]. Distilling [4494, 9555]. Distinguished [2533, 6955]. Distress [10413]. Distribute [2966]. Distributer [2693]. Distributing [378, 344, 4799]. Distribution [3164, 3371, 4726, 7730, 172]. Distributor [4300]. Ditching [366, 4543, 10545]. Diver [4013, 9046]. Diversions [1679, 8066]. Divider [3657, 7982]. Divides [10545]. Dividing [10307]. Divination [3909]. Diving [8429]. Divining [2899]. Dixon [3959]. Dock [744, 1241, 2231, 3507, 4347, 5921, 7024, 6611, 6969, 6444, 7798, 7786, 8235, 7811, 8102, 8478, 9550, 9288, 10534, 256, 8265, 8444, 8710, 9394, 7130]. Docks [7499, 9749]. Doctors [9869]. Doctrine [618]. Document [8676]. Documents [2018]. Dodge [8357]. Does [10027, 1763]. Dog [163, 993, 1672, 3589, 6653, 6752, 1121]. Dog-Headed [163]. Dogs [5123, 8897]. doing [9951]. Doll [401]. Dollar [1851, 4488, 7740]. Dollars [10212]. Domestic [2725, 4530]. Domesticating [1762]. Dominik [3508]. Don [10699]. Donatello [8663]. Door [296, 295, 84, 480, 2288, 2622, 2355, 3162, 5851, 5874, 9413, 2941]. Doors [910, 584, 665, 6146, 7782, 7768, 9039, 8693, 9727, 9934]. Doorways [3394]. Dormouse [5948]. Double [1240, 642, 13, 1519, 3129, 4341, 5274, 6151, 5593, 5852, 5565, 6699, 6805, 6863, 8423, 7885, 10612, 9777, 6700]. Double-Acting [9777]. Double-Deck [10612]. Double-Deck-Turret [6863]. Doubles [10574]. Dough [339]. Douglas [6172]. Dove [2058]. Dovecotes [1876]. Dovetail [7579, 6700]. Dowe [4882]. Down [650, 5311]. Downtown [131]. Dr [1099, 6588, 1078, 2137]. Dr. [1980, 1710, 3195, 2421, 3249, 5577, 8324, 9283, 10369]. Drag [6222, 6305]. Dragon [3508]. Drain [2057, 9376]. Drainage [340, 2293, 2796, 5218, 5776, 5805, 9752]. Dramatic [6905]. Draught [6397, 6396, 8366, 9287]. Draw [1136, 2553, 4876]. Drawbridge [6362, 7518, 5858, 5684, 6461, 7981]. Drawbridges [9660]. Drawer [1602, 5739]. Drawing [1152, 405, 1716, 1855, 1363, 1869, 3637, 5787, 6833]. Drawings [10095]. Dredge [1697, 8828]. Dredges [9214, 10303]. Dredging [1125, 4, 2095, 3021, 6046, 8639]. Dress [8429]. Dresser [1766, 6403]. Dressers [1001]. Dressing [5219]. Dried [5612]. Drier [34, 63, 1211, 2092, 4406, 5289, 6059, 7185, 9011]. Driers [10618]. Driggs [9161]. Drill [273, 274, 2222, 1727, 1965, 4024, 4210, 3571, 4670, 4696, 5779, 5879, 6867, 7377, 7902, 8072, 8339, 7904, 8407, 9956, 10330]. Drilling [1932, 4260, 3347, 4885, 5719, 7474]. Drills [992, 1441, 2210, 4210, 7835]. Drinker [7028]. Drinking [3874, 10422, 175]. Drive [689, 1913, 2623, 8636]. Driver [3369, 8977]. Drives [6737]. Driving [185, 220, 2222, 1441, 2111, 1330, 1395, 3148, 3484, 4380, 4648, 6706, 10652, 9759]. Driving-Gear [9759]. Drop [4002, 4407, 4798, 9823]. Drop-Actuating [9823]. Drops [5777]. Druggists [8579]. Drum [5716, 10618]. Drum-Driers [10618]. Dry [2231, 3507, 4347, 6969, 6466, 8265, 8235, 7499, 8710, 9394, 9550, 10534, 9749, 7130, 256, 5137, 7024, 6444, 7786, 7811, 8102, 8478]. Drydocks [8531]. Drying [759, 2545, 3701, 3681, 9391]. Duck [2093]. Duct [6900]. Dudley [4216, 6698]. Due [9547]. Duell [8615, 8652, 10635, 10638]. Duhamel [4167]. Duke [3543]. Dumb [417, 571, 6629, 7183]. Dumb-Waiter [571]. Dumbwaiter [1065, 3524]. Dumont [9395]. Dump [2524, 3072, 3769, 4322, 4685]. Dumping [1750, 2365, 4472, 5173, 6245, 7823]. Duplex [3886, 3845, 4210, 3890]. Duplication [10541]. Duplicator [6924]. Durability [9977]. Durable [2273, 2000, 2778, 4143, 6691, 6708]. During [3089, 5035, 6953, 6946, 9007, 9150, 9552]. Duryea [7331]. Dust [450, 2935, 3486]. Duster [10794]. Dusting [9935]. Dusting-Brush [9935]. Dustless [8471]. Dutch [6783]. Dwelling [732, 4861, 10336]. Dyeing [5959]. Dyer [8488]. Dynamite [269, 908, 2285, 3385, 4253]. Dynamo [1213, 5, 666, 768, 1907, 1928, 2342, 3196, 3148, 2816, 3864, 3660, 4756, 8263, 8144]. Dynamograph [6628]. Dynamometer [457, 5236]. Dynamos [1173, 4493, 5108, 4447].

Eagle [6377]. Ear [1171, 5630]. Eariphone [2504]. Early [818, 3688, 6325, 8428, 10017, 10009, 3601]. Ears [8879]. Earth [2898, 6813]. Earthquake [2467, 10796]. Earthquakes [9020]. Easels [2226]. Easily [1207, 1275, 2581, 2847, 5103, 4750]. East [7075, 4288, 4923, 5267, 5475, 8050, 7742, 8605, 10709, 10089, 9575, 10322, 10195]. Eastern [5044, 5032, 5276, 10478, 6919]. Eastman [10257]. Eastward [8259]. Eater [1329]. Eaters [5604]. Eating [8760]. Eccentric [845]. Eccentricities [6567]. Echoes [7522, 9922]. Eclipse [102, 3832, 3572, 5036, 5404, 6329, 8794, 6330, 6622]. Eclipser [3496]. Eclipses [10575]. Economic [1294, 1941, 6379, 8648, 10671]. Economical [2711]. Economics [5889, 10066]. Economies [7508]. Economy [374, 2221, 6161, 6497, 8462, 8625, 10410]. Eddy [5053]. Edgar [2629]. Edge [4041]. Edible [9607]. Edison [401, 706, 1907, 1928, 1791, 1909, 1818, 2131, 3055, 3859, 3326, 3621, 4433, 4423, 6836, 6638, 8564, 8508]. Edition [802, 837, 187, 213, 261, 281, 303, 327, 351, 432, 507, 630, 652, 679, 699, 1412, 1640, 1665, 1265, 1284, 1356, 1476, 1500, 1526, 1548, 1604, 1628, 1653]. Educated [9899, 9236]. Education [4898, 5663, 9582, 10746, 10345, 7009]. Educational [781, 4196]. Edward [1710, 3157, 10613, 7028, 1917, 6025, 6038, 10356, 6791]. Effect [1755, 1496, 3521, 6932, 9132, 10428, 8744]. Effective [645, 1797, 5029, 5726, 8873, 10415]. Effects [4253, 5805, 165]. Efficiency [8756, 10081]. Efficient [2248, 1474, 2136, 2967, 3007, 4100, 4143, 3623, 4031, 4406, 4941, 4707, 4843, 4409, 5228, 6059, 5360, 6107, 6102, 6040, 5742, 7313, 6708, 7296, 6729, 6506, 6442, 6517, 6401, 6583, 7736, 7953, 7782, 8017, 7640, 8003, 7599, 8956, 8977, 9409, 8871, 8698, 9011, 9824]. Effigies [3754]. Egg [3599, 8990, 9953]. Eggs [8515]. Egypt [8759, 8651]. Egyptian [10762, 10402]. Egyptians [9293]. Eiffel [1713, 7336]. Eight [6566, 8532]. Eighteenth [6204]. Eighth [5088, 9678]. Ejector [4748, 8828]. Elastic [35, 5714, 5518]. Elbe [7686]. Electric [9773, 10231, 582, 1160, 455, 18, 875, 736, 62, 649, 559, 300, 164, 162, 1019, 61, 143, 75, 1399, 1376, 1495, 1791, 1909, 2223, 2212, 1817, 1833, 1437, 2053, 2183, 1960, 2091, 3028, 2863, 2323, 2675, 2609, 2721, 3051, 2847, 2611, 3104, 2834, 3055, 2939, 2783, 3079, 3164, 2989, 3554, 4166, 3859, 4232, 3613, 3481, 4229, 3725, 4286, 4162, 3661, 3634, 3610, 3861, 3871, 3595, 3228, 4187, 3700, 4014, 4192, 5007, 4967, 4464, 4655, 4726, 4540, 5316, 4429, 4734, 4990, 4903, 4435, 4454, 4411, 4959, 5108, 5160, 4724]. Electric [4696, 4423, 5015, 5256, 4651, 5557, 5550, 5780, 6319, 5713, 5897, 5408, 6327, 6343, 6125, 5573, 5920, 6122, 5740, 5598, 6181, 5617, 5966, 5879, 6083, 5664, 5725, 7058, 6502, 6497, 6445, 7061, 6849, 6836, 7116, 6441, 6537, 6538, 6993, 6423, 7169, 7082, 6946, 7329, 8422, 8199, 7953, 7851, 7557, 7945, 7989, 7848, 8133, 8245, 7361, 8404, 7810, 7411, 7712, 7900, 7474, 8242, 8464, 8166, 9549, 8530, 8844, 8768, 8554, 8684, 9393, 8586, 8980, 9532, 10158, 10609, 10250, 10333, 10410, 10722, 10314, 10795, 10491, 10008, 10636, 9728, 9706]. Electric [9956, 10544, 10100, 10007, 10129, 7715, 5178, 6678, 1305, 6386, 7347, 2515, 3272, 144, 1134, 5108, 9829]. Electrical [1202, 627, 491, 514, 929, 115, 777, 172, 1868, 2210, 1323, 2129, 2073, 1848, 1810, 2254, 2465, 2422, 2452, 4006, 3651, 4108, 3588, 4137, 4121, 4746, 4433, 4528, 4570, 5621, 5896, 5911, 5980, 5999, 6103, 5957, 6744, 6750, 6416, 7779, 7803, 8365, 8193, 7908, 8928, 9303, 8910, 8955, 8982, 9800, 10037, 10428, 10079, 10619, 7066]. Electrically [1075, 851, 185, 3726, 5807, 5498, 5823, 7256]. Electrically-Fired [1075]. Electrician [3616]. Electricians [895, 2533, 2792]. Electricity [386, 778, 779, 1362, 1969, 1388, 1533, 1635, 1660, 2359, 2913, 3517, 3484, 4058, 3874, 4811, 8354, 7871, 7587, 9557, 8594, 9871, 9745, 2476, 5859, 4083]. Electrification [5777]. Electro [193, 310, 1969, 2131, 3057, 3146, 2887, 5599, 10495, 10615, 10518, 10257]. Electro-Artograph [5599]. Electro-Cycle [10257]. Electro-Mechanical [3146]. Electro-Metallurgy [10615]. Electro-Microscopic [1969]. Electro-Mobile [10495]. Electro-Motor [2131]. Electro-Motors [193]. Electro-Photo-Detective [2887]. Electro-Plating [3057]. Electro-Pneumatic [310]. Electro-Static [10518]. Electrodes [2325]. Electrographs [10518]. Electroliers [816]. Electrolytic [5725]. Electromotive [8544]. Electroplating [768, 1594, 5403]. Electrostatic [4753, 9902]. Element [6915, 7131]. Elements [9189]. Elephant [7333, 6368, 9369]. Elevated [454, 1347, 2548, 3548, 3284, 4799, 5684, 5654, 6802, 6384, 7239, 7434, 8063, 8554, 9627, 10210, 3272, 776, 7512]. Elevating [6129]. Elevator [889, 557, 1153, 584, 665, 1475, 2288, 2170, 3052, 4162, 3245, 4207, 4685, 5874, 6864, 8255, 8219, 7876, 8200, 8461, 7399, 8190]. Elevators [756, 916, 1983, 2357, 3367, 4335, 6187, 6084, 7361, 9526, 4136]. Eleventh [4746]. Eli [2304]. Elias [8123]. Elizabeth [5415]. Ellesmere [4597]. Elliptic [80]. Elliptical [3938]. Elm [5274]. Elms [3702]. Elswick [2658]. Ely [6378, 6406]. Emanations [2472]. Emancipation [7586]. Embankments [8340]. Embossing [278]. Embroidery [956, 2866]. Emergency [10129]. Emery [1766]. Emil [4548]. Empire [6339]. Employees [3893]. Employers [1667]. Employes [1667, 3471]. Employment [1392]. Empty [10770]. Enabling [4713]. End [317, 4260, 6402, 9303, 8693, 10742, 10787, 9924]. Ending [10276]. Ends [5253, 8298]. Endurance [8467, 10005]. Enemies [8004]. Energy [617, 4646, 5573, 5676, 6946, 10308]. Enforced [8291]. Engage [9604]. Engagement [9163, 9107]. Engine [900, 78, 456, 114, 431, 738, 646, 29, 1751, 1263, 1835, 1442, 2213, 1899, 2232, 2272, 2211, 2900, 2621, 2526, 2934, 2688, 2707, 3141, 2709, 2704, 4185, 3977, 4241, 4055, 3386, 4206, 4234, 4163, 3365, 3460, 3677, 4029, 3660, 5082, 5216, 4803, 4514, 4342, 4988, 5175, 4796, 4772, 4939, 5237, 5780, 6010, 5367, 5682, 6328, 5702, 5347, 5539, 6261, 6668, 7027, 6974, 6826, 6561, 7297, 7079, 6925, 6879, 6697, 7436, 7860, 7825, 7857, 7471, 7722, 7765, 7670, 7379, 9451, 8491, 8563, 8611, 8806, 8913, 8562, 9015, 9033]. Engine [10567, 10787, 10278, 10280, 10799, 10237, 10497, 10475, 10307, 7614, 209]. Engineeer [4540]. Engineer [1053, 4137, 8648, 1000, 4816, 7896, 7830, 8898]. Engineering [698, 1431, 1726, 3131, 3651, 3841, 6182, 6915, 7131, 7009, 8463, 9402, 9464, 10584, 10400, 9313, 9928, 3504, 6342, 9806, 10669]. Engineer’s [9090, 2242, 2826, 4230, 4746, 5028, 6416, 7803, 8273, 9208, 9281, 9628, 10068, 10256, 100, 1761, 3283, 7852, 8405, 9447, 8600, 9978]. Engines [159, 1055, 955, 891, 936, 1422, 2395, 2600, 2505, 4158, 4301, 3845, 3679, 3918, 3794, 4181, 3868, 5044, 4903, 4803, 4507, 4925, 4860, 5740, 5986, 6913, 6457, 8126, 8217, 7386, 8127, 8529, 8697, 9208, 8656, 9533, 10771, 10396, 10450, 10701, 4547, 6386, 9789]. England [5919, 7027, 6719, 7312, 7658, 8443, 7970, 9451, 9952, 9841, 10230]. English [3182, 3723, 3495, 4782, 7096, 7179, 7260, 7588, 9243, 9446, 8648, 9548, 8898, 8553, 10343, 10355, 9950]. Engraving [7863]. Engravings [1690]. Enlarged [7163]. Enlarging [1231, 6998]. Ensilage [208, 2415]. Enterprises [9763, 9604]. Entertainment [6905]. Entrance [3944, 5088, 7968, 10064]. Entrances [3434]. Entz [4569]. Enumeration [2797]. Envelop [9175]. Envelope [529, 1601, 3144, 4474, 6301]. Envelopes [4345, 4672, 7536]. Environment [9018]. Eolian [4656]. Equalizer [1948]. Equations [7541]. Equatorial [5333]. Equestrian [8663]. Equipment [6803, 7779, 7908, 9557, 10637]. Equitable [3346]. Equivalent [8830]. Eraser [911, 7087]. Erection [7591]. Ericsson [860, 841, 2618, 6297, 9178, 112]. Erie [7270, 9286]. Ernst [3195]. Erosion [3414]. Error [10711]. Eruption [2831, 2412, 10401]. Escape [501, 1945, 2434, 3180, 3348, 5982, 6283, 6808, 7623, 10377, 9936, 10776]. Escutcheon [118]. Essay [6726, 6781, 6821, 6898, 6951]. Essen [1234, 1209]. Essex [7170]. Essick [7674]. Established [439, 800, 835, 1457, 1560, 1585, 1616]. Establishment [3934, 5669]. Establishments [818, 10140]. Estate [6449]. Estimate [1076]. Etaings [3710]. Etc [759, 626, 208, 1080, 2197, 2133, 1764, 1378, 1514, 1875, 1808, 1465, 1296, 2018, 1569, 2400, 4301, 3225, 4280, 4028, 4191, 4082, 3701, 3408, 4107, 4160, 3914, 4041, 4057, 4472, 4474, 4880, 5134, 5028, 4672, 5211, 5100, 4860, 5939, 6146, 5717, 6106, 5963, 6067, 7166, 8006, 8026, 8051, 8077, 8095, 8115, 8137, 8158, 8182, 8206, 8226, 8247, 8268, 8285, 8306, 8346, 8369, 8411, 8435, 8455, 8480, 7624, 7649, 7675, 7700, 7724, 7744, 7791, 7816, 7842, 7865, 7890, 7911, 7935, 7961, 7984, 7536, 7600, 7977, 8764]. Etc [8499, 8523, 9240, 9255, 9274, 9296, 9319, 9340, 9361, 9378, 9438, 9459, 9478, 9541, 8568, 8619, 8645, 8668, 8729, 8750, 8774, 8795, 8815, 8836, 8860, 8882, 8900, 8919, 8941, 9001, 9021, 9063, 9082, 9124, 9182, 9201, 9220, 8788, 8677, 8765, 8579, 8493, 9595, 9618, 10432, 10456, 10500, 10527, 10552, 10577, 10601, 10627, 10655, 10677, 10752, 10780, 9644, 9693, 9714, 9736, 9784, 9808, 9834, 9858, 9887, 9915, 9941, 9967, 9996, 10032, 10057, 10083, 10130, 10172, 10198, 10243, 10265, 10288, 10358, 10403, 10318, 133, 1964]. etc [1943, 1897, 1913]. Etching [479]. Ether [3168]. Ethical [6379]. Ethylene [2871]. Etna [2831]. Eucalyptus [1032]. Eudynamis [4927]. Euphrates [10487]. Eureka [4027, 5159]. Europe [1876, 2058, 3105, 5177, 6294, 9549, 8712, 9375, 10120, 10538]. European [153, 7162, 9030, 10712]. Euyuk [7206]. Evade [4812]. Evans [8070]. Evaporation [1753]. Evaporator [390, 3521]. Evening [929, 2392]. Event [7291]. Ever [8662]. Evolution [999, 3979, 4972, 6065, 6089, 7972, 8169, 7046]. Evolutions [1965]. Examination [6147]. Examinations [2523, 8038]. Examiners [3953]. Examining [10217, 1814]. Example [4747, 1681]. Excavating [366, 1248]. Excavations [8630]. Excavator [4880, 5134]. Excellent [8441, 8593]. Exchange [3146, 2432]. Exchanging [10538]. Exclusion [2432, 3304]. Excursion [8336]. Execution [777]. Exercise [1486, 3315, 8996]. Exercises [2202, 1970, 4737]. Exercising [5558]. Exerted [5800]. Exhaust [3198]. Exhaustive [6218]. Exhibit [4005, 4073, 4166, 3959, 4259, 3495, 3870, 4055, 3844, 4008, 4123, 4037, 4370, 6285, 5620, 4027, 3870, 3803, 3811, 4075, 4029, 4064, 4120, 4041, 4057, 4017, 4523, 9252]. Exhibiting [1378, 4384]. Exhibition [386, 17, 2619, 4838, 6162, 7447, 9326, 8928, 8910, 8955, 8982, 9709, 10037, 10510, 10079, 9796, 29, 1846, 2252, 2073, 1443, 2738, 3723, 4827, 6198, 5436, 6436, 7099, 7487, 7756, 7522, 7826, 7832, 8274, 9653]. Exhibitions [9872]. Exhibitor [56, 712, 204, 789]. Exhibits [1010, 4173, 4061, 3677, 4190, 4529, 4455, 7659, 3522, 3544]. Expander [1180, 869, 4716, 5481, 9985]. Expanding [8175]. Expansible [1040]. Expansion [1263, 1422, 2505, 4507, 5237, 5556, 7990, 10620]. Expedition [1782, 3860, 4633, 3813, 4597, 6217, 6559, 7029, 8061, 8366, 7833, 8254, 8534, 8725, 6622]. Expeditions [1806, 7040]. Experiment [1090, 1547, 3909, 4309, 5920, 7655, 8063, 8316, 9402, 9652, 9038, 973, 2532, 3200, 3939, 6087, 2564, 3275]. Experimental [1020, 1508, 7189, 8450, 7824, 8682, 9215, 10220, 515, 1342, 1367, 1390, 1410, 1435, 1456, 1536, 1559, 1584, 1615, 1638, 1663, 10353, 3]. Experiments [706, 1688, 2295, 2421, 2708, 2951, 2891, 2528, 2801, 2310, 4541, 6218, 6638, 6570, 8856, 9614, 10666, 9734, 10724, 9874, 8744, 4499, 6547, 2477, 3576, 10736]. Expert [8926, 2932]. Expiration [2882, 5268]. Explained [9884]. Exploits [4169]. Exploration [2984, 7409, 6919]. Explorations [9673, 10674]. Explorers [9820]. Explosion [609, 1291, 2078, 2097, 2910, 2603, 4361, 4436, 5636, 6865, 7576, 8002, 7838, 7379, 9331, 9316, 10721, 10022, 9831, 9992, 9866, 9884, 6838]. Explosive [1257, 7223, 10710, 10446, 10126]. Explosives [483, 8294, 10298, 10092]. Export [7755, 8593, 9505, 10489, 10494, 10324]. Exportations [10274]. Exporters [9511]. Exports [9401, 8627, 9382, 10065, 10765]. Exposition [7, 4195, 3517, 3872, 3520, 3542, 4138, 4157, 4183, 3565, 3611, 3653, 3676, 3721, 3743, 3786, 3805, 3823, 3843, 3863, 3885, 3908, 3932, 3964, 3976, 4053, 4079, 4098, 4118, 3794, 3784, 3587, 4467, 4971, 6732, 6770, 6744, 6899, 8124, 9833, 10624, 10489, 10494, 10283, 10694, 9885, 495, 2297, 3068, 2863, 3029, 2975, 3059, 4304, 4288, 3622, 4267, 4155, 4262, 4229, 3745, 4162, 3610, 4173, 3706, 3394, 3434, 3844, 3962, 3756, 3544, 3699, 3767, 4000, 4026, 3755, 4136, 4181, 4189, 3608, 4145, 3677, 3612, 3633, 3698, 3983]. Exposition [3722, 4210, 4220, 4160, 3758, 3305, 3564, 3609, 3636, 3882, 3917, 3944, 3963, 3973, 3984, 3985, 3987, 4015, 4035, 4064, 4109, 4120, 4123, 4126, 4193, 4217, 3650, 4553, 4478, 4529, 4455, 4370, 5935, 6234, 6253, 6326, 6285, 5538, 6349, 5578, 5918, 6099, 6143, 6750, 7659, 7709, 7396, 7754, 9252, 9356, 8962, 9217, 10714, 10093, 4127]. Exposition-General [4035]. Expositions [4264]. Express [3650, 5155, 5851, 7027, 7103, 7880, 9451, 10382]. Extensible [1807, 4232, 8610]. Extension [419, 3159, 3388, 4687, 7061, 7451, 8084, 8019, 9169, 9353]. Extensions [8649]. Extermination [129, 5998]. Extinct [3393, 10218]. Extinguisher [3891, 9311]. Extinguishers [10082]. Extinguishing [755, 1300, 4412]. Extinguishment [4078, 8165]. Extraction [7590]. Extractor [2090, 4590]. Extracts [3107]. Extraordinary [4254]. Extreme [10228]. Extrusion [6851]. Eye [4727, 9171, 4589]. Eyes [9529]. Eyesight [7945, 10565].

F [6025]. F.R.S. [6488]. Fabric [2270]. Fabrics [9339]. Face [7521]. Facilitate [1061, 1727, 1299, 2398, 3408, 5192, 4862, 5715, 6298]. Facilities [3159, 7798]. Fact [9444]. Factor [10321]. Factory [2636, 2404, 3484, 5234, 5495, 8101, 8319, 8683, 10513]. Facts [5806, 10471]. Fails [466]. Failure [2285, 4899, 9583]. Failures [7040]. Fair [1010, 2196, 1250, 2229, 3160, 2363, 3764, 4306, 3260, 4041, 4057, 3841, 4635, 4886, 7895, 8258, 660, 745, 51, 1031, 60, 221, 1, 195, 1459, 1837, 1620, 1792, 2941, 2825, 2328, 2776, 3090, 3980, 3958, 3824, 3942, 4002, 4083, 3959, 3979, 3870, 4165, 4055, 3981, 4061, 4016, 3803, 4058, 4032, 4086, 4076, 4024, 3563, 4141, 4104, 3943, 4008, 4107, 4122, 3652, 3997, 4085, 4542, 4653, 6208, 6735]. Fairy [9539]. Faithful [2024]. Falguiere [3706]. Fall [1415, 2914, 3214, 3861, 7876, 8200, 7897, 9317, 9109, 10411, 1377, 5160]. Fallacy [9928, 10162]. Falling [1513]. Falls [5837, 6632, 10647, 10262, 7715, 1107, 2054, 3043, 2374, 2473, 5139, 6600, 6423]. False [8625, 10178]. Families [4494]. Family [1834, 3227, 4389]. Famine [8803, 2744]. Famous [2423, 8463, 10400]. Fan [1676, 3037, 5967, 8044]. Fancy [9444]. Fans [1913, 5742]. Faraday [1458]. Fargo [1038]. Farm [1058, 3718, 4820, 7980, 7548, 8101]. Farmer [4966, 9792, 3673]. Farmers [10139]. Farragut [9531]. Farthest [7047]. Fast [1847, 1996, 2114, 2668, 2484, 5071, 7620]. Fastener [1041, 318, 398, 1871, 1517, 1592, 2307, 2690, 4473, 4865, 4709, 5512, 5961, 6402, 7535, 8871, 9824, 10146]. Fasteners [6067]. Fastening [737, 1829, 4345, 7133, 7199, 8703]. Fastenings [8214]. Faster [8553]. Fastest [42, 8064, 7731, 9403, 9234, 10196, 10003, 9747, 10439, 8278]. Fate [9359]. Father [1699, 10759]. Fatigue [7921]. Faucet [396, 5717, 7517, 8872]. Faulty [9325, 10352]. Faure [9428]. Favor [3303]. Favorably [8377]. Favored [221]. Faye [6145]. Fayette [5679]. Feat [3335]. Feather [80, 4849]. Feats [10014]. Features [698, 8955, 10710, 10352, 10277]. February [104, 1346, 2593, 2376, 3316, 5438, 7429, 2671, 6437, 8606, 9658]. Feed [116, 1597, 1963, 1808, 2486, 3217, 2549, 2470, 3569, 5029, 5290, 4792, 4632, 5477, 6714, 7117, 6827, 6630, 7377, 8956, 10125, 10651, 10102, 7347]. Feed-Water [3569, 10651, 10102]. Feeder [1149, 2201, 2289]. Feeding [4774, 7805, 9661]. Feedwater [4587, 5516]. Fees [1587]. Feet [1525, 1771, 1415, 6679, 8408]. Fell [5428]. Felling [2800]. Fellow [2438]. Felt [2955]. Fence [572, 66, 206, 1131, 2134, 1929, 2530, 2778, 2592, 3407, 3846, 8977, 10673, 3115, 3601]. Fence-Making [1131]. Fence-Wires [10673]. Fences [1040, 975]. Fender [4665, 4937, 5135, 5064, 4428, 4668, 5721, 5649, 5463, 5479, 6506]. Fennecs [5503]. Ferdinand [5188]. Fermentation [2738]. Fernandez [6060]. Fernkorn [3508]. Ferns [2420]. Ferrel [2139]. Ferret [5089]. Ferris [3985]. Ferry [2728, 8091, 8302]. Fertilizers [821, 7802]. Festival [2132, 3089, 5090]. Few [8202]. Fiber [467, 1926, 2150, 2759, 5759, 9701, 4918]. Fibers [1926]. Field [938, 2995, 5269, 4365, 7537, 7642, 9161, 9153, 9281, 2833, 4814, 7647]. Field-Gun [9161]. Fields [4299, 8058, 8714]. Fiend [8488]. Fifth [4936, 9568]. Fiftieth [6726, 6971]. Fifty [4488, 4745, 6957, 6941, 6950, 6959, 6953, 6946, 10519, 10546, 7435]. Fight [1888]. Fighter [3253, 10472]. Fighting [4424, 7956]. Fights [9096]. Figs [10213]. Figures [3486]. File [972, 1523, 8676, 10447]. Filing [6848, 10123]. Filipstad [6297]. Filler [523]. Filling [1765, 2398]. Films [4309, 10185]. Filter [1217, 1627, 3501, 4857, 4842, 4750, 6165, 6480, 7713]. Filtering [4926, 9850]. Filters [3996]. Filtration [8018]. Fin [3812, 10453]. Financial [1855, 10463]. Find [7853]. Finder [4014, 5052, 8854]. Finders [8128]. Fine [4054, 6088]. Finger [36, 1062, 5558]. Fingers [9373]. Finisher [1816]. Finishing [3936, 10447]. Fire [528, 636, 755, 754, 209, 608, 609, 730, 1945, 1468, 1300, 1734, 1258, 2711, 2872, 2434, 2900, 3180, 3824, 3331, 3348, 4238, 3732, 3526, 4078, 4594, 5035, 4412, 5327, 4385, 4988, 4565, 5701, 5604, 5982, 6283, 6188, 6400, 6808, 6697, 7318, 8337, 7860, 7623, 8399, 7471, 7452, 7971, 7862, 7599, 8407, 8165, 7839, 8927, 8697, 8491, 9128, 8490, 9311, 8939, 9294, 9535, 10566, 9815, 9895, 9746, 10082, 9930, 10377, 9936, 10776, 9762, 10642, 10472, 8488, 9922]. Fire-Alarm [9930]. Fire-Escape [10377, 9936, 10776]. Fire-Extinguisher [9311]. Fire-Extinguishers [10082]. Firearms [898, 5009, 5395]. Fireboxes [9269]. Fired [1075]. Fireplace [4888]. Fireproof [1567, 4822, 8461]. Fireproofing [10461]. Fires [444, 2070, 3568, 5045, 8321, 7666]. Fireworks [4262, 3270, 6905]. Firing [483, 494, 1355, 6078, 8135, 10298]. Firm [1079, 7899, 8211]. First [777, 2629, 2327, 3289, 3531, 4550, 5273, 5801, 7030, 6384, 7011, 7643, 9498, 9140, 9372, 10771, 9674, 2774]. First-Class [7643, 9498, 10771]. Fiscal [10323, 10276]. Fish [759, 821, 3309, 5869, 8515, 9960, 10538, 10214, 8004, 10264, 1546, 8878, 9690]. Fish-Hook [9960]. Fisheries [7095, 2220, 3434]. Fishes [8520, 10727]. Fishing [291]. Fitters [2811]. Five [3150, 6010, 6446, 7187]. Fixed [6255]. Fixing [8509]. Fixture [272, 473, 6438, 6917]. Flag [3392, 9140, 8741, 10206]. Flagler [9870]. Flagstaffs [10216]. Flames [8673, 9010]. Flanging [1488]. Flash [184, 1061, 11, 2021, 2605, 3892, 10722]. Flash-Light [11, 10722]. Flashes [10349]. Flashless [9746]. Flat [2205, 4468]. Flats [2059]. Flax [9466]. Fleet [9142, 152, 8164, 8907, 9087, 9231, 8948, 10648]. Flexible [3113]. Flies [7770]. Flight [2221, 7023, 7580, 8507, 6547]. Float [1763]. Floating [1049, 988, 1511, 6046, 9394, 9550, 9288, 10251, 8265]. Flood [6613]. Floods [1524, 7802, 7751]. Flooring [323, 1297, 7927, 9012]. Floors [4191, 4207, 7708]. Florence [2676]. Florida [1215, 1675, 1828, 6872, 9975]. Flour [2491, 10274]. Flow [1440, 4580, 8856]. Flower [1396, 3265, 4802, 4849, 7600]. Flowers [2737, 2974, 2979, 2703, 4697, 5964, 6975, 8305, 7320]. Flue [1135, 1176, 314, 1488, 2661, 3807, 5695, 5481, 6806, 9139, 10729, 10304]. Flue-Cleaner [9139]. Flue-Cutters [10729]. Flue-Stopper [10304]. Flues [1103]. Fluid [2255, 7674, 8934, 10144]. Fluid-Pressure [10144]. Fluids [1547]. Flume [255, 2020]. Fluorine [377, 7854]. Fluorometer [8635]. Fluoroscope [6638]. Flush [2039]. Flusher [7516]. Fly [2180, 8617, 9236].

Flying [1783, 1497, 2378, 4219, 5305, 4997, 5049, 5747, 5948, 6171, 6570, 7901, 8004, 9554, 5106]. Flywheel [2324]. Focus [6637]. Focusing [3684]. Fodder [546]. Foes [4424]. Fog [2021, 7831, 9150, 9981]. Foldable [2092]. Folder [4434]. Folding [1081, 691, 717, 847, 733, 2031, 4101, 3264, 6081, 6581, 7556, 7417, 8043, 9249, 10076]. Folio [3788]. Followed [609]. Fontainebleau [3354]. Fonthill [3850]. Foo [5990]. Food [815, 3620, 8067, 8989, 9279, 9774]. Foods [9245]. Foot [1489, 1295, 1445, 2324, 3041, 3768, 5171, 5308, 6318, 8238, 7435, 10064]. Footgear [8971]. Foothills [2490]. Force [5360, 5800, 5676, 6277, 7182, 8782, 8801, 8544]. Forced [6397]. Forces [9576]. Foreign [832, 334, 2147, 2560, 7990, 8653, 8926, 9092, 8650, 9244, 9872, 10323]. Foreigners [4422]. Forest [5045, 9047, 4583]. Forestry [1271, 2229, 5024, 4455, 10468, 8952]. Forests [8519]. Forethought [8531, 9280]. Forge [6307]. Forged [4798, 8738]. Forger [9953]. Forging [2148, 2972, 6149]. Forgings [4002, 4407]. Fork [546]. Forks [185]. Form [896, 9433, 9936, 10681, 9905, 784, 1085, 1036, 2091, 2075, 2706, 2703, 3705, 8131, 10776, 10124]. Formal [288]. Formed [6400]. Former [6221]. Formidable [42]. Forming [4321, 6992]. Forms [10017, 10009]. Fort [8854, 5108]. Forth [288, 25]. Forthcoming [9527, 8784]. Fortifications [7346, 9687, 6687]. Fortify [9062]. Forts [8998]. Forty [1033, 10064, 2774]. Forty-First [2774]. Forty-Foot [10064]. Forty-Ton [1033]. Forum [10623]. Fossil [1933, 5459]. Foster [2921]. Foundation [4984, 5946]. Foundations [3554, 5264, 8050, 8074]. Founder [1170, 3839]. Foundry [568, 6048]. Fountain [85, 1396, 5860, 7118, 7851, 9389, 10328, 3508]. Fountains [3288, 4492]. Four [3041, 4241, 4498, 8090, 7579, 7518, 9451, 8780, 8929, 10589]. Four-Coupled [9451]. Four-Foot [3041]. Four-Masted [4498]. Four-Spindle [10589]. Fourth [9429, 9653]. Fowler [9490]. Fox [3910]. Frame [2250, 2866, 5287, 5828, 6131, 5466, 6363, 7837, 8473, 7596, 9260, 7430]. Framed [9261]. Frames [418, 1623, 6441]. Framing [9409]. France [496, 3745, 3706, 5179, 8036, 8560, 10560, 10311, 8808]. Francis [5428]. Francisco [8156, 9456, 996, 4467, 6969, 9370, 10472, 9963]. Frank [10142]. Frankfort [2073, 2422]. Franklin [2363]. Fraud [8356, 9642, 10194]. Fraudulent [8034, 8257]. Frederic [9209]. Frederick [873]. Free [4811, 7730]. Freight [1728, 1650, 2355, 3128, 2546, 8367, 7519, 8634, 9176, 9778, 9879]. French [92, 804, 1735, 2738, 3870, 5229, 5578, 6258, 6523, 9728, 9864, 9896, 10748, 10670, 3115, 3601]. Frequency [2528]. Fresenius [7922]. Fresno [7601]. Frick [390]. Friction [43, 4103, 6810, 7473, 10569]. Frictionless [7135]. Friedburg [4611]. Frikart [5237]. Frilled [6636]. Frisket [2287]. Frog [1066, 5231]. Front [6809, 7798]. Frost [5459]. Froude [5275]. Frozen [3914, 8718]. Fruit [63, 663, 1600, 4802, 4794, 5853, 8555]. Fruits [8783, 10714]. Fry [10538]. Fuel [2268, 3784, 8063, 8604, 10512, 10160, 10335].

Full [2992]. Fuller [6591]. Fulton [3518, 4475, 10718, 9611]. Fumigator [912]. Funeral [4669]. Funnel [954]. Furnace [480, 108, 109, 2827, 2529, 3145, 3523, 3546, 4801, 5366, 5873, 6830, 6383, 6786, 7738, 8890, 9073, 8659, 9984]. Furnaces [1597, 2289, 1514, 3613, 5366, 6220]. Furnishing [7272]. Furniture [3654, 6185, 4127]. Fürst [1744]. Further [8442]. Fuse [1041, 737, 2933, 5292]. Fused [7202]. Future [3131, 6120, 7973, 10747, 10723].

G [3673, 7875, 10369]. G.C.B [4994]. Ga [518, 6303]. Gadey [4343]. Gage [5496, 9411, 8720, 9826]. Gage-Glass [9411]. Galapagas [2296]. Gallery [7685]. Galley [200, 341]. Gallons [5174]. Galloway [4037]. Galton [6172]. Galvanizing [6239]. Galvanometer [8812]. Galveston [4512, 6762]. Gamblers [10222]. Game [129, 662, 231, 1085, 4921, 6608, 8761]. Games [6558, 6754]. Gang [3038, 3664, 10590]. Ganges [2383]. Garbage [4327, 8071, 9370]. Garden [692, 1329, 5220, 5436, 5416, 6173, 6436, 6408, 10491, 10726, 1698, 998, 2921, 4013, 4781, 4515, 5503, 5786, 5634, 5685, 6487, 7136, 9633]. Gardening [9076]. Gardens [4908, 6999]. Garibaldi [3410]. Garment [5960]. Garrett [391]. Garrison [4973]. Garter [791]. Gas [819, 939, 579, 609, 14, 1751, 1462, 1370, 3022, 2659, 2621, 3141, 2709, 4205, 4055, 3891, 3386, 3370, 4303, 4181, 4242, 4361, 5082, 4707, 4923, 5216, 4803, 4609, 4342, 6285, 5366, 5740, 5367, 5702, 5600, 6675, 7025, 7294, 6668, 6417, 6448, 6826, 6561, 6519, 8338, 7385, 7576, 8450, 7825, 7447, 7857, 8279, 8383, 7858, 8696, 8933, 8847, 9547, 9559, 8656, 9080, 9493, 9312, 8858, 9233, 9903, 10676, 10599, 10721, 9831, 10348, 10397, 10073, 10592, 10564, 6386, 10335, 2477, 5539]. Gas-Generating [10073]. Gas-Governor [9312]. Gases [133, 2801]. Gasocution [4275]. Gasoline [2213, 2677, 5082, 4803, 4796, 5682, 5702, 5783, 7209, 6974, 6561, 7857, 7670, 8848, 8656, 8562, 8578]. Gate [317, 1153, 138, 1058, 2271, 1729, 2977, 3313, 3248, 3749, 3956, 4467, 4820, 4363, 5100, 6402, 8448, 9471, 8677, 9077, 9332, 9266]. Gatemann [10041]. Gates [6362, 10146, 1303]. Gathering [1814]. Gatling [1075, 2502, 5349, 9583]. Gatti [6123]. Gauge [1062, 1418, 1768, 2119, 2986, 2694, 3202, 2723, 4073, 5308]. Gauges [3350]. Gazette [3821, 3862]. Gear [525, 1831, 1330, 3150, 4289, 3887, 3868, 4383, 4380, 4430, 4939, 4648, 5498, 5380, 7106, 7240, 7433, 8513, 8873, 10027, 10189, 9589, 9759]. Gear-Cutting [525]. Geared [4382, 7187, 9098]. Gearing [3955, 9032]. Gelatine [2453]. Gems [7636]. General [2257, 3861, 4035, 1932, 2384, 5108, 9870]. Generals [9180]. Generating [10073, 7715]. Generation [2808, 6946]. Generator [4793, 6417, 8696, 8933, 8847, 9155, 9493, 10516, 10676, 10145, 10397]. Generators [9845]. Generous [10412]. Geneva [5860]. Genus [10671]. Geographical [9844]. Geological [27, 2033, 4402, 6916, 9678, 8213, 8336, 1273, 6881, 7376]. Geologists [1866]. Geology [9263]. Geometrical [4196]. George [3508, 764, 1978, 5245, 9806].

Germ [3501, 3996, 10254]. Germ-Proof [3501, 3996]. German [1995, 7758, 7880, 10511, 423, 921, 404, 2202, 3962, 3315, 5123, 5377, 6047, 6515, 10763, 10345, 9827]. Germanic [9724]. Germany [3733, 5395, 5824, 7115, 8598, 10740, 10417]. Get [2920]. Getting [2130]. Geyser [9076, 10221]. Geysers [5837]. Giant [6221, 5448, 7238, 8389, 9445, 10600, 9573]. Giants [10108]. Gibraltar [10790]. Giffard [786]. Gift [8441, 7539]. Gigantic [482, 2170, 4775, 6112, 6890, 9635]. Gin [3790]. Ginning [4542]. Gins [10652]. Ginseng [1278]. Giovanni [8727]. Giraffe [10562]. Girder [918, 2659]. Girl [2609]. Girth [1592]. Giving [13]. Glacier [2565, 6199]. Glaciers [7621]. Glafcke [3493]. Glasgow [10637]. Glass [253, 319, 537, 1283, 2967, 2995, 3088, 3553, 5269, 5238, 4523, 4348, 5519, 5806, 5745, 7364, 8196, 7537, 7642, 7694, 9153, 9411, 9666, 1736, 990]. Glass-Polishing [2967]. Glazed [9529]. Glimpse [9885]. Globe [9450, 4803]. Globe-Holder [9450]. Glove [398]. Gloves [6609, 6590]. Glue [6007]. Glycerine [2830, 5765]. Glyptodont [4674]. Gnu [5685]. Go [8168]. Goebel [4344]. Going [7643, 8760, 9214, 10522, 99]. Gold [198, 1836, 2977, 3435, 3934, 6082, 7049, 8058, 7489, 8236, 8639, 8556, 9170, 9122, 10091, 10640, 7826]. Golden [4467, 2941, 9229]. Goldsmith [3723]. Gone [8431]. Good [618, 2242, 3176, 5098, 7097, 7290, 7548, 7739]. Goodale [764]. Goods [4189]. Gordon [8429]. Gores [5518]. Gorge [6600, 10151]. Gorham [6020]. Gospels [3688]. Gould [3186]. Government [7924, 1941, 2889, 5284, 4628, 6603, 7730, 7659, 8417, 7634, 8427, 9252, 8602, 8842, 9771, 10468, 10560, 10639, 9604, 7130]. Governmental [10066]. Governor [1298, 2688, 2452, 4514, 4939, 9312]. Grace [10440, 7989]. Graceful [7732]. Grade [1174, 9122, 10128]. Grader [4664]. Grades [10390]. Graft [4526]. Grafting [1121, 8681]. Graham [7906]. Grain [2247, 2149, 2507, 3072, 4322, 5422, 6869, 8461, 8362, 10045]. Gramophone [697, 6751]. Grand [5784, 7815, 151, 2054, 4015, 4126, 5160, 6750, 7487, 2747, 2873]. Granite [5520]. Grant [10473, 2626, 7201, 7717, 7541]. Grape [9805]. Graphophone [402, 8426]. Graphoscope [4879]. Grapnel [3133]. Grass [1849]. Grate [1514, 4123, 4343, 4883, 8197]. Grater [877]. Gratings [695]. Graves [4335]. Graving [8444]. Gravity [668, 1547, 2108, 2136, 6996, 9194]. Gravity-Operated [2108]. Grays [4017]. Great [358, 1039, 608, 1950, 1512, 1856, 1771, 1344, 1250, 1252, 2034, 2910, 2810, 3030, 2832, 2771, 2436, 3158, 2528, 3160, 3517, 4165, 3844, 3764, 3962, 3777, 3872, 3794, 3784, 3362, 4583, 4866, 4878, 4900, 4366, 6199, 6066, 5987, 5837, 5693, 5812, 5437, 5645, 5541, 6611, 6584, 7232, 6794, 8365, 8399, 7872, 7813, 8296, 7862, 7666, 9537, 9019, 9264, 9214, 9027, 8497, 9464, 10650, 9702, 9630, 9703, 9833, 10624, 9676, 8303, 10617, 3273, 372, 1226, 288, 765, 892, 942, 2231, 3029, 2593, 3394, 3434, 4117, 3803, 5141, 5006, 5418]. Great [6960, 6515, 6462, 8601, 8661, 9451, 8917, 8528, 9427, 9841, 10478, 10685, 10019, 10545, 10622, 8278]. Great-Headed [3030]. Greatest [6182].

Greatly [9603]. Greeley [7808, 8616]. Green [7463]. Greenerd [3545]. Greenland [3114, 7621]. Greenville [8540]. Greenwich [6717]. Greiner [4548]. Gresham [9061]. Greyhound [8241]. Greytown [10595]. Gridley [5357]. Griffin [3981]. Grills [695]. Grimm [4801]. Grind [2400]. Grinder [1254, 3498, 4405, 10395]. Grinding [624, 762, 430, 499, 2248, 3983, 4546, 5941, 7255, 8423, 9877, 9661]. Grinding-Wheels [9661]. Grindstones [10190]. Grip [199, 843, 1276, 4391, 7721]. Grooving [293]. Gross [10711, 9842]. Grosse [8442]. Ground [3663, 8718]. Grounding [2543]. Grounds [3354, 3372, 9867]. Group [8905, 8916, 9162]. Grove [1215]. Growing [6207, 6421, 9805, 10353]. Growth [945, 3051, 3106, 3752, 4800, 4339, 8627, 10055, 10233, 9586]. Growths [3615]. Gruson [10090]. Guard [450, 6, 994, 574, 1817, 3912, 3659, 5011, 5891]. Guide [688, 599, 3286, 4028, 3848, 3408, 3773, 6870, 7135]. Guiley [1848]. Guinea [5924]. Guitar [8112]. Gulf [8833]. Gum [1601]. Gun [137, 1075, 786, 483, 404, 908, 2401, 2618, 2502, 2582, 2658, 2636, 2404, 4168, 3732, 5117, 4778, 4745, 5349, 6141, 6698, 7189, 7278, 6793, 7273, 7778, 7974, 7839, 8683, 8631, 8719, 9161, 9057, 9383, 9989, 10228, 10022, 9583, 9992, 9674, 9948, 9866, 9798, 10312, 9499, 9884]. Gun-Barrels [9989]. Gun-Sight [5117]. Gunboat [1120, 8770]. Gunboats [4864, 6396, 8366, 9593]. Gundalow [9254]. Gunner [1854]. Gunpowder [1051, 270, 335]. Guns [1087, 1145, 1051, 269, 494, 443, 1512, 2286, 1436, 1485, 2001, 2034, 2810, 2345, 5510, 6790, 9195, 9538, 9417, 8997, 8939, 9294, 8991, 9746, 10769, 9610, 9867]. Gutta [1959, 7661]. Gutta-Percha [7661]. Gymnasium [6265]. Gymnastic [1970]. Gypsy [708, 10696]. Gyrograph [6169].

H [4276, 4994, 5577, 5357, 8615]. H. [4061]. Habitation [1546]. Habits [1942, 4887, 1984]. Habitual [8641, 8660]. Hack [1894]. Hackensack [2542]. Hagenbeck [4598]. Hailstones [6794, 7088]. Hair [5830, 10055, 932]. Hairless [2343]. Half [2052, 3632, 6961]. Halifax [256]. Hall [1771, 5295, 7737, 9235]. Halles [115]. Halsted [4943]. Halter [1372]. Halters [740]. Hamburg [4908, 4515, 6173, 3201]. Hame [5809]. Hamilton [2884, 2344, 10698]. Hammer [254, 3710, 9878]. Hammered [6088]. Hammerless [642, 7063]. Hammers [4041, 8092]. Hammond [6535]. Hand [9, 713, 156, 224, 1187, 1886, 1441, 2182, 1419, 2971, 2435, 2606, 4470, 5893, 6588, 6587, 7377, 8281, 8263, 9095, 7347]. Handicraft [4972]. Handle [180, 1523, 6351, 6369, 7600, 7669]. Handler [9779]. Handling [556, 821, 2245, 2830, 4570, 6123, 8362]. Handsome [3867, 9453]. Handy [6825, 7472]. Hanger [292, 2181, 1623, 1590, 2885, 6043, 6849, 6447, 8958, 10190, 9849]. Hangers [1081]. Hanging [1698]. Hanover [5685]. Hansom [7557]. Happy [3719]. Harbor [3021, 5981, 6824, 7948, 1932, 3342, 3433, 3755, 4510, 6276, 5869, 6556, 8450, 7968, 8995, 9316, 9602, 10064, 10303, 6762]. Harbors [4880, 8496, 10598]. Hardening [4895]. Hardwood [7708]. Harkness [4038].

Harlem [277, 694, 5153, 4593, 4705, 4773, 5839, 6461, 7445, 7455, 7476, 7801, 7465, 7981]. Harmless [5138]. Harmonic [9853]. Harmony [4944]. Harness [497, 205, 574, 2581, 5201]. Harp [6028]. Harps [4656]. Harrison [372, 4362]. Harrow [1253, 2166, 3477, 7151, 6772]. Hartford [7908, 8245, 7852]. Hartmann [6976]. Harvard [7788]. Harvest [1368, 2132, 4727, 5536]. Harvester [2415, 5462, 7383, 3185]. Harvesters [4860]. Harvesting [9338]. Harvey [8108]. Hassan [258]. Haste [9149]. Hatch [3162]. Hatchery [5869]. Hatching [1122]. Hats [2955]. Haulage [4445]. Hauling [10768]. Haunted [4830]. Haute [3214]. Havana [8265, 8444, 9394, 9550, 10507]. Haven [6208, 7908, 7861]. Hawaii [9430]. Hawaiian [3325]. Hay [426, 208, 1520, 2604, 3656, 4365, 10045, 9904]. Hay-Stacker [9904]. Hayne [10498]. Hayward [8108]. Head [293, 992, 1082, 117, 19, 650, 3974, 5232, 4779, 7641, 7617, 9071, 8677, 10543, 10049]. Head-Light [1082]. Headed [163, 3030, 7300]. Header [669]. Heading [2742]. Headlight [7165]. Health [4050, 9700, 8628, 7138]. Heart [9530]. Heat [1230, 2799, 7022, 6422, 7871, 8721]. Heated [133, 851]. Heater [1205, 1675, 1963, 1462, 2205, 2224, 2549, 2470, 3481, 4186, 4469, 5029, 4587, 5290, 5478, 7117, 9802]. Heating [10231, 41, 696, 1034, 4257, 4707, 4609, 4881, 5713, 5716, 8403, 10158, 10314, 7715]. Heavens [5247, 5326, 5190, 5591, 5936, 6257, 5438, 5735, 5849, 5494, 5677, 6203, 8954, 7292, 7132, 7592, 8016, 8400, 7429, 7807, 7534, 7710, 8301, 8220, 8110, 8786, 9135, 9491, 8606, 9052, 8695, 8870, 9388, 9307, 9219, 8494, 10275, 10719, 9658, 10071, 10163, 9726, 9955, 10588, 10493, 10374]. Heavy [885, 2286, 1436, 1485, 2135, 4016, 6280, 7290, 7366, 7782, 7719, 8477]. Hecla [7102, 7120, 3685]. Hedge [206, 8894]. Height [1771, 1350]. Heights [8701, 1966]. Heilmann [8133]. Hektograph [5268]. Held [1624, 1809]. Helena [8770]. Heliochromoscope [3167]. Helium [5910]. Hell [2977]. Helmholtz [4106, 5188]. Helping [152]. Hemet [8201]. Hemming [687]. Henri [10748]. Henrion [3196]. Henry [873, 5877, 7550, 8772, 1825, 4344, 10798]. Her [2394, 5294, 10441]. Herald [6163, 6278, 5669]. Herbert [7343]. Herman [5188]. Heron [3095, 5881, 1348, 4128]. Herrmann [7578]. Hibernia [5071]. Hichborn [9528, 10240, 10585]. Hickory [3406, 7596]. High [483, 1351, 1966, 3871, 4361, 4710, 4654, 4618, 5912, 6077, 6457, 6851, 6679, 7734, 8464, 9531, 8835, 9417, 10567, 9871, 10298, 9896, 10092, 10128, 10426, 10049, 7066, 144, 8544, 2928, 5813, 6063]. High-Level [10426]. High-Service [2928]. High-Speed [10049]. Higher [9148]. Highest [3733]. Highways [2935]. Hill [3983, 9734]. Hilliard [4770]. Hilling [604]. Hillside [4645]. Hinge [559, 621, 3108, 6146, 7313, 8448, 8973]. Hinges [2039, 1591]. Hint [10229]. Hints [1633, 2920, 9511, 4567]. Hip [181]. Hippopotamus [6612, 7136]. Historic [4370]. Historical [4195]. History [1292, 4778, 6325, 5810, 7588, 7662, 7549, 9386, 10015, 10506, 3122, 6154, 7333]. Hitchcock [1817]. Hits [9190]. Hoadley [9662]. Hoboken [2728]. Hoche [804]. Hoe [207, 227, 8979]. Hog [2185, 3795]. Hogs [5499]. Hoist [668, 553, 446, 2135, 2136]. Hoisting [3071, 3679, 4799, 7857, 7386]. Hold [1765, 1786, 5940]. Holdback [1371]. Holder [523, 597, 1172, 531, 447, 348, 639, 533, 956, 319, 38, 1393, 2153, 1299, 2207, 2075, 2272, 1375, 1569, 2883, 2950, 3451, 3552, 4967, 4431, 4546, 5424, 5699, 6298, 7061, 6812, 6768, 7087, 6419, 7472, 8321, 7783, 8474, 9450, 9559, 9112, 8685, 9035, 8807, 9036, 8959]. Holding [576, 603, 865, 1692, 2624, 4236, 5172, 8195, 8218, 8092, 8679]. Holds [1375]. Hole [4544]. Holes [2016]. Holiday [8441]. Holidays [9552]. Holland [6696, 7799, 7603, 8813]. Hollow [2474, 5565]. Holly [9194]. Holman [5257]. Holtz [326]. Holystone [2849]. Home [695, 673, 2726, 3480, 4497, 8405, 9886, 10517, 112, 5983]. Home-Made [10517]. Homestake [7049]. Homestead [2902, 2843, 2915]. Homing [8584]. Homogenetic [2797]. Hon [7663]. Hon. [8615]. Honey [9900, 10215]. Honey-Birds [10215]. Honolulu [3021, 9676]. Honor [4193]. Honored [4725, 9974]. Hoo [3919]. Hoo-Den [3919]. Hood [5560]. Hoof [2171, 6670]. Hook [205, 640, 9960, 8822, 10022, 9992, 9884]. Hooks [6266]. Hoop [622, 10645]. Hopkinson [9283]. Horizontal [1263, 6624, 8188, 10589]. Horn [6080]. Horned [8705]. Hornet [4736, 9886]. Horological [3793]. Horology [9177]. Horse [2129, 1651, 2171, 2987, 2343, 2690, 2709, 2704, 4241, 4368, 6107, 5696, 5347, 6669, 6476, 8275, 7426, 7538, 7691, 9516, 10055, 10609, 10250, 9707, 10773, 932, 4017]. Horseback [6299]. Horseless [6128, 5958, 7208, 6835, 7261, 7121, 8046, 7709, 7471, 9700]. Horses [760, 690, 2096, 4445, 5804, 7533, 7437]. Horseshoe [3329, 5385, 5593, 8105]. Horticultural [9326, 1891]. Horton [7452]. Hose [1105, 2757, 5421, 5982, 6671, 6728, 7599, 9114, 10795]. Hospice [10314]. Hospital [9327, 9091, 9159, 5313]. Hostile [8854]. Hot [689, 1205, 114, 1034, 2019, 2224, 3371, 4029, 7613, 7674]. Hotel [8304, 9922, 5084, 5681, 9912]. Hour [3632, 4561, 5266]. House [227, 2311, 3554, 3335, 4649, 5597, 7172, 7500, 7977, 7555, 8048, 9155, 9019, 9138, 10773, 3806, 8134, 221, 2836, 5044, 5313, 5616, 8224]. House-Cooler [9138]. Household [2920, 6295, 7410, 10164]. Housekeeping [9514]. Houses [732, 7299]. Housing [6379, 7311]. Houts [7150]. Howard [8194]. Howling [998]. Hub [1057, 4243, 6003]. Hubs [7332]. Hudson [556, 2170, 4645, 1179, 101, 1039, 428, 336, 1368, 1252, 1724, 4825, 5306, 5541, 7345, 6705, 8376, 8060, 8237, 8340]. Huge [6204, 10278]. Hulls [4386, 5403]. Human [4779, 5800, 8467, 10308, 9774, 10574, 6547]. Humane [8557]. Humbert [8719]. Humboldt [8519]. Humming [8223]. Hundred [4241, 3632, 4745, 6010, 5347, 6679, 7740, 7942, 7691]. Hung [7076, 7065]. Hungarian [6770]. Hunger [10542]. Hunt [2513, 8932]. Hunting [3095, 5797, 6753, 8001]. Huntington [5245]. Hunts [6737]. Hura [7223]. Hurricane [6972, 10477]. Hut [5125]. Hut-Building [5125]. Huxley [5877]. Hybrids [3375]. Hydraulic [198, 602, 646, 897, 2722, 2867, 3074, 3664, 3591, 3748, 6161, 5946, 5368, 6001, 6424, 6854, 6969, 7297, 6632, 8542, 10122, 10280, 10647]. Hydrocarbon [1875]. Hydrogen [5476, 9652, 10668, 10521, 6087]. Hydrographic [8496]. Hydrostatic [2570]. Hygienic [2096, 3199]. Hygroscope [3639]. Hysteria [10557].

I.H.P. [6010]. I846 [1560]. Ice [638, 2037, 1368, 2525, 3413, 3235, 4056, 4451, 4907, 5101, 5615, 5501, 5441, 6747, 6584, 6882, 8640, 9634, 10399, 8714]. Ice-Skates [9634]. Iceberg [10099, 7085]. Iceland [6712, 9673]. Icy [5563]. Idaho [2061]. Idea [7713]. Ideal [9584]. Ideas [1170, 8420]. Identification [7619]. If [3719]. Igniter [2933]. Igniters [6386]. Iguanodons [2109]. II [4336, 6089, 8148, 10053, 10425, 10723, 8834, 9358, 2120]. III [4359, 9143, 9374]. Ill [5823]. Illinois [1620, 1946, 3718, 9372]. Illuminant [5645]. Illuminating [1791, 2477]. Illumination [2727, 2939, 4108, 4990, 6514, 8982, 10525]. Illusion [6997, 7203, 6735]. Illusions [1649, 8066, 1238, 1281]. Illusive [4513]. Illustrating [4850]. Image [13]. Images [7163]. Imaginary [2088]. Imaginations [5442]. Imitation [662, 648]. Immigration [2466]. Impact [7202]. Impending [8234]. Imperfections [4562]. Imperial [3915, 4918, 10511]. Implement [692, 82, 958, 1869, 10048]. Implements [5895]. Import [9654]. Importance [6989, 10740]. Important [885, 2241, 5532, 7569, 8073, 7853, 10671]. Importation [10739]. Importing [9348]. Improved [714, 424, 417, 300, 43, 293, 312, 469, 1210, 672, 105, 316, 81, 297, 940, 596, 715, 716, 323, 523, 497, 597, 690, 418, 83, 201, 393, 879, 248, 202, 30, 933, 260, 107, 135, 155, 203, 251, 294, 338, 342, 394, 449, 498, 600, 601, 620, 625, 995, 1104, 395, 56, 78, 456, 272, 199, 624, 41, 205, 576, 992, 1172, 275, 472, 257, 445, 808, 347, 473, 605, 846, 626, 79, 862, 106, 134, 230, 324, 1002, 1084, 296, 450, 732, 530, 295]. Improved [339, 405, 84, 4, 340, 667, 571, 457, 1065, 781, 35, 557, 317, 1153, 164, 116, 1041, 1135, 63, 480, 546, 11, 938, 36, 85, 572, 663, 691, 717, 847, 66, 108, 109, 912, 231, 939, 762, 668, 692, 430, 253, 138, 688, 114, 669, 553, 206, 1082, 602, 531, 117, 156, 207, 426, 622, 446, 447, 621, 451, 638, 318, 499, 890, 44, 913, 273, 1211, 603, 474, 664, 448, 113, 535, 598, 1100, 918, 232, 547, 584, 280, 249, 313, 558, 607, 378, 744]. Improved [396, 224, 604, 157, 406, 376, 915, 31, 670, 867, 1175, 176, 208, 877, 524, 1206, 527, 579, 1235, 37, 470, 552, 643, 1176, 200, 341, 82, 57, 362, 551, 348, 67, 367, 397, 500, 392, 226, 767, 599, 843, 227, 639, 86, 548, 177, 812, 844, 710, 58, 64, 536, 914, 274, 756, 1101, 431, 845, 815, 398, 640, 532, 711, 916, 420, 544, 665, 301, 363, 757, 993, 136, 349, 671, 848, 452, 891, 878, 1102, 10, 178, 368, 573, 179, 228, 369]. Improved [809, 935, 644, 180, 533, 863, 1022, 233, 554, 1129, 421, 364, 119, 583, 934, 1232, 712, 471, 813, 526, 120, 481, 1178, 641, 810, 1060, 110, 204, 475, 647, 419, 577, 365, 422, 868, 977, 1130, 1155, 696, 111, 1589, 1298, 1794, 1417, 1832, 1812, 1971, 2038, 2276, 1276, 1625, 1542, 1379, 1515, 1498, 1961, 2110, 2227, 1750, 1602, 1835, 1728, 1945, 2205, 1808, 2150, 1418, 1729, 1372, 1520, 1463, 2270, 1516, 1865, 2056, 1911, 1912, 1748, 1749, 1999, 2113, 1795, 1733, 1590, 2203, 1277]. Improved [1464, 2003, 1693, 1442, 2040, 1465, 2198, 2057, 1254, 2249, 2116, 1964, 1870, 2172, 1491, 2228, 1349, 1394, 2199, 1677, 1893, 2274, 2250, 1768, 1810, 1492, 1563, 1714, 1892, 2986, 3217, 2937, 2588, 2755, 3197, 2827, 2742, 2988, 2793, 2968, 2379, 2585, 2848, 2912, 2622, 2417, 2487, 2547, 2562, 3093, 2535, 2416, 2524, 2448, 3179, 3198, 2548, 3144, 2355, 2661, 2434, 2900, 3180, 2530, 2549, 2529, 2582, 3072, 2621, 3038, 2757, 2971, 2435, 2867, 3074, 2604, 2885, 2525, 2643, 2508, 3108, 2969, 2674, 2606, 2830, 3025]. Improved [2639, 3004, 2970, 2469, 3053, 2741, 2396, 2356, 2488, 2583, 2669, 2702, 2764, 2380, 2694, 2724, 2509, 3127, 2526, 2934, 2812, 3202, 2307, 2489, 2563, 2670, 2993, 2510, 2345, 2361, 3073, 2357, 2688, 2723, 3162, 3005, 2312, 2794, 2828, 2707, 2640, 2449, 2545, 3150, 2309, 2366, 2527, 2952, 3181, 2308, 2397, 2964, 3429, 4302, 3248, 3244, 3451, 3523, 3937, 3261, 4003, 3306, 3473, 3592, 4102, 3499, 3528, 3808, 4233, 3245, 3938, 4191, 3474, 3452, 3827, 3640, 3262, 4081, 3347, 3524, 3595, 3228, 4187, 3700, 3388, 3331]. Improved [4103, 3348, 3846, 3569, 3546, 3386, 3656, 3748, 3477, 3547, 3729, 3771, 3448, 3290, 3727, 3502, 3250, 3293, 3453, 3703, 4277, 3287, 3746, 3865, 4319, 3680, 4209, 3426, 3503, 3349, 3790, 4206, 4234, 3749, 3614, 3478, 4281, 3314, 3570, 3387, 3246, 3500, 4159, 4163, 3769, 4007, 3226, 4289, 3230, 3593, 3525, 3365, 3460, 3521, 3720, 3566, 3594, 3825, 3231, 4082, 3770, 3366, 3431, 3704, 4320, 4119, 4390, 4819, 4380, 4473, 4881, 5046, 4381, 4716, 4987, 4688, 5027, 5289, 4665, 4937, 5135, 4954, 4735, 4858, 4885, 4384]. Improved [4514, 4345, 4820, 4385, 5290, 4342, 5117, 4363, 4654, 5028, 4614, 5232, 5212, 5249, 4425, 4410, 5230, 4367, 4730, 4975, 5047, 4968, 5030, 4749, 4525, 5291, 4859, 4495, 4708, 4989, 5156, 5310, 5328, 4988, 5009, 4405, 4588, 5255, 5064, 5231, 5118, 5068, 4596, 4671, 4791, 4938, 4507, 4792, 4844, 4902, 4364, 4821, 4430, 4610, 4586, 4795, 5397, 6021, 6316, 5461, 5834, 5440, 5913, 5737, 5651, 5596, 5649, 5462, 5986, 6205, 5808, 6244, 5717, 5481, 5853, 5516, 5463, 5422, 5349, 5682, 5696, 5809, 6028, 6328, 5613, 5650, 5984]. Improved [6296, 6170, 6022, 5914, 5697, 5638, 5561, 5361, 5828, 6166, 6259, 5719, 5943, 6348, 6041, 6279, 5512, 5961, 6131, 5535, 6004, 5942, 6005, 6300, 6240, 5398, 5575, 5680, 5894, 5594, 6260, 7332, 6830, 6992, 6460, 6673, 6482, 6440, 6540, 6748, 7185, 6362, 6731, 6629, 7239, 6537, 6383, 6581, 7117, 6826, 6561, 7297, 6624, 6812, 7151, 6544, 7064, 6458, 7236, 6649, 7081, 7133, 6388, 6403, 6695, 6887, 6709, 6850, 6518, 6749, 6785, 6689, 7062, 7045, 7237, 7255, 6918, 7258, 6672, 7314, 6479, 8401, 8382, 7903, 7926, 7975, 8430]. Improved [7949, 8104, 8262, 7615, 8325, 8359, 8297, 8423, 8339, 8131, 7473, 8383, 8196, 8448, 8105, 7641, 8152, 8177, 7380, 7362, 8020, 7904, 7594, 8214, 7931, 7738, 7722, 8218, 7689, 7798, 7996, 7494, 8360, 7378, 7515, 7540, 7839, 8561, 9250, 9118, 8976, 8829, 8685, 9054, 8746, 8513, 9032, 8872, 8890, 9012, 9077, 8563, 8678, 9158, 9137, 8789, 8850, 9093, 8611, 8806, 8913, 8632, 9174, 9213, 8790, 9332, 9035, 8936, 9310, 8559, 9153, 8657, 9265, 8851, 8957, 9371, 10145, 10700, 9958, 9906, 9930, 10304, 10190, 9877, 9961, 10048, 9587]. Improved [9659, 10236, 10279, 9852, 10305, 10474, 9637, 10497, 10681]. Improvement [277, 361, 1063, 740, 888, 1056, 1098, 2002, 2039, 1711, 1414, 2059, 2979, 2777, 2641, 2780, 2446, 3621, 3493, 3755, 3476, 5201, 4386, 4673, 4705, 5246, 5215, 6276, 6222, 5875, 6461, 6809, 7221, 7837, 7968, 8216, 9410, 9533, 8541, 9572, 10082, 9681, 9984, 10569, 10729, 10507, 9634, 10492, 9801, 10618, 10396, 9851, 9825]. Improvements [955, 908, 2137, 1983, 2241, 1867, 2353, 2586, 3433, 4240, 3686, 4747, 4969, 4768, 5781, 6363, 7510, 7361, 9527, 10326, 10450, 6789]. Improves [2987, 2365]. Improvised [1295, 10122]. Inaccessible [6569]. Inaccuracy [7971]. Inaudi [2569]. Incandescent [1050, 2, 3055, 3326, 5617, 6527]. Inch [137, 2584, 4670, 6566, 7278, 6793, 9161, 8997, 9383, 9215, 10228, 10022, 9992, 9674, 9866, 9798, 9884, 9499]. Incident [4726]. Incline [4685]. Inclined [1521, 7399]. Inclosed [5446]. Incompetent [2826]. Increase [2004, 6601, 8334, 8782, 8801, 9505, 9207]. Increased [7233]. Increasing [2303, 4925, 9206, 8650]. Incubator [4367, 6319]. Incubators [9412]. Independent [3475, 8638, 9375]. Index [23, 49, 657, 684, 705, 729, 753, 774, 795, 796, 827, 828, 906, 927, 949, 967, 986, 1007, 1028, 1047, 1073, 1096, 1116, 1143, 1167, 1195, 1222, 1223, 73, 1245, 97, 127, 149, 171, 192, 219, 242, 267, 287, 309, 333, 357, 384, 412, 438, 463, 490, 512, 542, 566, 590, 615, 633, 474, 525, 634, 1246, 858, 1270, 1290, 1823, 1842, 1863, 1884, 1905, 1924, 1940, 1957, 1976, 1992, 2012, 2029, 2050, 2068, 2086, 2105, 2126, 2145, 2158, 2178, 2194, 2218, 2239, 2263, 2283]. Index [2302, 1313, 1338, 1361, 1386, 1405, 1406, 1430, 1453, 1463, 1482, 1506, 1532, 1554, 1579, 1610, 1634, 1659, 1686, 1705, 1722, 1742, 1759, 1780, 1803, 1844, 2319, 2333, 2769, 2789, 2807, 2824, 2841, 2860, 2880, 2896, 2908, 2927, 2947, 2963, 2983, 3000, 3017, 3035, 3050, 3066, 3086, 3102, 3121, 3140, 3156, 3173, 3192, 3210, 2352, 3223, 2372, 2391, 2411, 2429, 2445, 2464, 2483, 2501, 2521, 2541, 2559, 2578, 2599, 2617, 2635, 2653, 2667, 2683, 2701, 2719, 2735, 2753, 2770, 3241, 3258, 3801, 3818, 3837, 3857, 3880]. Index [3903, 3927, 3951, 3971, 3993, 4022, 4047, 4071, 4093, 4115, 4134, 4153, 4179, 4202, 4226, 4249, 4273, 4295, 4316, 4333, 3281, 3300, 3323, 3340, 3360, 3381, 3398, 3422, 3442, 3468, 3492, 3515, 3537, 3561, 3582, 3606, 3629, 3648, 3671, 3694, 3716, 3739, 3762, 3782, 4356, 4376, 4893, 4915, 4933, 4949, 4964, 4982, 5002, 5022, 5042, 5059, 5078, 5096, 5114, 5131, 5151, 5167, 5186, 5208, 5226, 5243, 5262, 5282, 5303, 5322, 5339, 4398, 4419, 4441, 4462, 4485, 4504, 4520, 4536, 4559, 4578, 4604, 4626, 4643, 4657, 4681]. Index [4703, 4722, 4743, 4765, 4788, 4808, 4836, 4855, 4874, 5340, 5356, 5375, 5820, 5844, 5865, 5887, 5906, 5930, 5954, 5976, 5995, 6016, 6035, 6053, 6074, 6096, 6118, 6137, 6159, 6179, 6196, 6215, 6232, 6251, 6272, 6291, 5393, 6313, 6336, 6356, 5414, 5434, 5454, 5473, 5491, 5508, 5530, 5549, 5570, 5588, 5609, 5626, 5643, 5661, 5673, 5691, 5711, 5731, 5752, 5774, 5795, 6357, 5845, 6967, 6376, 6395, 6877, 6895, 6930, 6966, 6986, 7003, 7020, 7037, 7056, 7072, 7093, 7112, 7128, 7145, 7161, 7177, 7193, 7215, 7230, 7249]. Index [7268, 7286, 7309, 7325, 7340, 7355, 6413, 7357, 6434, 6453, 6472, 6494, 6511, 6533, 6553, 6575, 6598, 6619, 6643, 6664, 6684, 6704, 6723, 6742, 6760, 6778, 6800, 6818, 6844, 6860, 7371, 7391, 7938, 7964, 7986, 8008, 8029, 8056, 8079, 8098, 8118, 8142, 8161, 8185, 8208, 8229, 8252, 8270, 8287, 8310, 8331, 8349, 8373, 8392, 8413, 8437, 8458, 8482, 7407, 7423, 7443, 7460, 7483, 7507, 7528, 7547, 7566, 7583, 7609, 7629, 7653, 7679, 7702, 7727, 7747, 7773, 7795, 7820, 7846, 7868, 7893, 7914, 8484, 7916, 8503, 8526]. Index [9085, 9105, 9147, 9187, 9205, 9225, 9242, 9259, 9276, 9299, 9321, 9345, 9364, 9381, 9400, 9421, 9440, 9463, 9482, 9504, 9523, 9546, 9566, 8550, 8571, 8589, 8624, 8647, 8670, 8690, 8708, 8734, 8753, 8779, 8798, 8821, 8840, 8864, 8885, 8902, 8924, 8946, 8968, 8984, 9004, 9025, 9044, 9045, 9579, 9600, 10201, 10225, 10247, 10270, 10291, 10317, 10339, 10361, 10386, 10408, 10434, 10458, 10482, 10504, 10532, 10554, 10581, 10606, 10632, 10657, 10682, 10704, 10731, 10754, 10784, 10805, 9622, 9647, 9670, 9696, 9716, 9739, 9767, 9788, 9811, 9838]. Index [9862, 9891, 9918, 9944, 9970, 9998, 10035, 10061, 10085, 10113, 10132, 10153, 10175, 10176]. Indexes [7357]. India [4288, 2265, 2458, 2383, 2475, 3873, 4065, 4840, 5987, 6207, 7813, 9079, 2744]. Indian [3974, 4217, 8192, 9356, 9822, 9070, 10222, 8999]. Indiana [4549, 6184, 5679, 5348, 7030, 7011, 7477, 990, 1915]. Indianapolis [871]. Indians [1282, 9270]. Indicating [5496, 6886]. Indications [3002]. Indicator [571, 916, 1001, 1899, 3005, 2973, 3708, 4184, 4528, 4862, 4510, 5603, 5514, 6261, 7689, 10692]. Indies [3027]. Indoor [1486, 2451]. Induced [6397, 7946]. Induction [4753, 4752, 3409]. Industrial [518, 268, 4422, 6732, 7099, 6989, 7115, 7684, 8736, 8651, 9487, 10685, 10074, 10313, 10443]. Industries [959, 6324, 6286, 5757, 6939, 9243, 8545, 8834, 9143]. Industry [3620, 3567, 5238, 5271, 5198, 5330, 5519, 5385, 5520, 6242, 5423, 5619, 5745, 7205, 8358, 8380, 7511, 8353, 7825, 8011, 8282, 8343, 8452, 8597, 10211, 9761, 10069, 10141, 3449, 8694, 2676, 3353, 9609, 10310]. Inexpensive [2074, 1393, 2134, 2834, 3109, 4031, 3427, 3772, 5327, 6105, 6401, 6906, 7299]. Inflater [5878]. Inflating [6006]. Inflation [3161]. Influence [4506]. Influenced [9382]. Influences [1987]. Information [1460]. Infringer [3303]. Ingenious [4771, 6768, 9555, 10189, 9706, 10395, 10056]. Ingenuity [9783]. Ingersoll [6439]. Ingleton [2643, 6328, 6730]. Inhabitants [7243]. Injector [5544]. Injurious [5971]. Injustice [8468]. Ink [2608, 4005, 4492, 5613, 10498]. Ink-Well [10498]. Inkstand [6921]. Inlet [6695]. Inner [2898]. Inquiries [2522]. Inquiry [8691]. Insanity [4812]. Inscription [2940, 8039]. Insect [6831, 7580, 10794, 8681, 9847, 9667, 9099]. Insect-Powder [10794]. Inserting [8129]. Inspection [7209, 7155]. Instance [3414]. Instantaneous [1187, 3663, 3657, 5478]. Instead [10564]. Institute [1010, 2196, 2745, 3929, 3444, 3243, 3915, 4746, 7803, 7895, 8258, 7797, 9326, 9177, 970, 989, 3106, 4074, 3793]. Institution [4582, 8674, 8510]. Instructing [10217]. Instruction [247, 8090]. Instrument [530, 737, 157, 711, 1275, 2346, 3006, 3547, 3787, 3263, 5028, 5914, 5683, 7259, 7881, 10700, 5069, 3271]. Instruments [68, 693, 2364, 3479, 4104, 5234, 7221, 10697, 9719]. Insulator [2721, 6538, 9154, 9881]. Insurance [9869]. Insurrection [6012, 5857]. Intaglio [3754]. Intake [6480]. Intelligence [2756]. Intensity [6975]. Interest [2792, 5781, 7750]. Interesting [1010, 1090, 1460, 2627, 3045, 3265, 5382, 5655, 6658, 7923, 7549, 7959, 8316, 9389, 8955, 8655, 9110, 9831, 10666, 9730, 4499, 1281, 9690]. Interests [10539]. Interior [3987, 8379]. Interlocking [151, 414, 6607]. Internal [7559]. International [1866, 2738, 3841, 5285, 6065, 6111, 6027, 6099, 7218, 7732, 8736, 9252, 9217, 9678, 8213, 8336, 89, 268, 2432, 4060, 4110, 4996, 4338, 4827, 6089, 5918, 5544, 6916, 7486, 8962, 10519, 10546, 10205]. Intersecting [2382]. Interstate [8335]. Interview [1158]. Intra [3872]. Intra-Mural [3872]. Intrenching [2587]. Introduced [4049]. Introducing [10617]. Introduction [2161]. Invalid [6081]. Invalids [8514, 10052]. Invasion [10685, 9059]. Invented [10475]. Invention [1245, 1401, 1735, 3055, 2949, 2897, 4966, 4446, 5693, 6686, 6780, 6953, 6847, 9455, 10159, 7046]. Inventions [23, 49, 657, 684, 705, 729, 753, 774, 795, 796, 827, 828, 906, 927, 949, 967, 986, 1007, 1028, 1047, 1073, 1096, 1116, 1143, 1167, 1195, 1222, 1223, 73, 97, 127, 149, 171, 192, 219, 242, 267, 287, 309, 333, 357, 384, 412, 438, 463, 490, 512, 542, 566, 590, 615, 633, 858, 45, 70, 656, 682, 701, 725, 749, 770, 794, 823, 855, 883, 905, 926, 948, 964, 983, 1006, 1027, 1044, 1070, 1092, 1112, 1140, 1163, 1190, 1221, 1244, 94, 124, 146]. Inventions [168, 189, 216, 240, 264, 284, 306, 330, 354, 380, 409, 437, 460, 487, 506, 508, 540, 564, 587, 612, 632, 1270, 1290, 1823, 1842, 1863, 1884, 1905, 1924, 1940, 1957, 1976, 1992, 2012, 2029, 2050, 2068, 2086, 2105, 2126, 2145, 2158, 2178, 2194, 2218, 2239, 2263, 2283, 2302, 1313, 1338, 1361, 1386, 1405, 1406, 1430, 1453, 1482, 1506, 1532, 1554, 1579, 1610, 1634, 1659, 1686, 1705, 1722, 1742, 1759, 1780, 1803, 1667, 1267, 1287, 1861, 1880, 1900, 1921, 1936, 1953, 1973, 1989, 2008, 2025]. Inventions [2047, 2065, 2083, 2101, 2122, 2141, 2155, 2174, 2190, 2214, 2235, 2259, 2280, 2299, 1307, 1333, 1357, 1381, 1402, 1426, 1448, 1479, 1502, 1528, 1551, 1575, 1607, 1631, 1655, 1683, 1702, 1718, 1739, 1756, 1777, 1799, 1820, 1838, 2319, 2333, 2769, 2789, 2807, 2824, 2841, 2860, 2880, 2896, 2908, 2927, 2947, 2963, 2983, 3000, 3017, 3035, 3050, 3066, 3086, 3102, 3121, 3140, 3156, 3173, 3192, 3210, 2352, 3223, 2372, 2391, 2411, 2429, 2445, 2464, 2483, 2501, 2521, 2541, 2559, 2578, 2599, 2617, 2635, 2653]. Inventions [2667, 2683, 2701, 2719, 2735, 2753, 2685, 2316, 2330, 2785, 2803, 2820, 2837, 2856, 2876, 2893, 2922, 2944, 2960, 2980, 2996, 3015, 3032, 3046, 3062, 3081, 3099, 3116, 3137, 3152, 3170, 3187, 3205, 3218, 2348, 2369, 2386, 2408, 2426, 2440, 2460, 2479, 2497, 2516, 2537, 2555, 2575, 2595, 2613, 2632, 2649, 2662, 2679, 2698, 2715, 2732, 2750, 2766, 4096, 3241, 3258, 3801, 3818, 3837, 3857, 3880, 3903, 3927, 3951, 3971, 3993, 4022, 4047, 4071, 4093, 4115, 4134, 4153, 4179, 4202, 4226, 4249, 4273]. Inventions [4295, 4316, 4333, 3281, 3300, 3323, 3340, 3360, 3381, 3398, 3422, 3442, 3468, 3492, 3515, 3537, 3561, 3582, 3606, 3629, 3648, 3671, 3694, 3716, 3739, 3762, 3782, 3884, 3238, 3256, 3798, 3815, 3853, 3876, 3899, 3923, 3947, 3967, 3990, 4019, 4044, 4068, 4089, 4111, 4130, 4149, 4175, 4197, 4222, 4246, 4269, 4292, 4312, 4330, 3277, 3296, 3319, 3337, 3357, 3378, 3396, 3419, 3438, 3463, 3489, 3510, 3533, 3557, 3578, 3603, 3626, 3643, 3667, 3690, 3712, 3736, 3760, 3779, 4356, 4376, 4893, 4915, 4933, 4949]. Inventions [4964, 4982, 5002, 5022, 5042, 5059, 5078, 5096, 5114, 5131, 5151, 5167, 5186, 5208, 5226, 5243, 5262, 5282, 5303, 5322, 5339, 4398, 4419, 4441, 4462, 4485, 4504, 4520, 4536, 4559, 4578, 4604, 4626, 4643, 4657, 4681, 4703, 4722, 4743, 4765, 4788, 4808, 4836, 4855, 4874, 5061, 4353, 4371, 4892, 4910, 4930, 4945, 4963, 4977, 5001, 5017, 5041, 5055, 5075, 5093, 5109, 5127, 5145, 5162, 5180, 5203, 5221, 5242, 5258, 5277, 5298, 5317, 5334, 4393, 4415, 4438, 4458, 4481, 4500, 4516, 4533, 4555, 4574, 4600]. Inventions [4622, 4640, 4660, 4677, 4700, 4719, 4739, 4760, 4784, 4805, 4833, 4852, 4870, 5356, 5375, 5820, 5844, 5865, 5887, 5906, 5930, 5954, 5976, 5995, 6016, 6035, 6053, 6074, 6096, 6118, 6137, 6159, 6179, 6196, 6215, 6232, 6251, 6272, 6291, 5393, 6313, 6336, 6356, 5414, 5434, 5454, 5473, 5491, 5508, 5530, 5549, 5570, 5588, 5609, 5626, 5643, 5661, 5673, 5691, 5711, 5731, 5752, 5774, 5795, 5352, 5373, 5862, 5882, 5901, 5928, 5951, 5975, 5991, 6015, 6030, 6051, 6069, 6095, 6113, 6135, 6155, 6174, 6195, 6212]. Inventions [6228, 6250, 6269, 6288, 6308, 6333, 6355, 5388, 5409, 5429, 5451, 5469, 5487, 5504, 5525, 5548, 5569, 5583, 5605, 5623, 5639, 5658, 5672, 5688, 5706, 5728, 5749, 5769, 5790, 5818, 5843, 6932, 6406, 6376, 6395, 6877, 6895, 6930, 6966, 6986, 7003, 7020, 7037, 7056, 7072, 7093, 7112, 7128, 7145, 7161, 7177, 7193, 7215, 7230, 7249, 7268, 7286, 7309, 7325, 7340, 7355, 6413, 6434, 6453, 6472, 6494, 6511, 6533, 6553, 6575, 6598, 6619, 6643, 6664, 6684, 6704, 6723, 6742, 6760, 6778, 6800, 6818, 6844, 6860, 6374]. Inventions [6392, 6894, 6909, 6926, 6981, 7001, 7016, 7034, 7052, 7070, 7092, 7107, 7123, 7140, 7158, 7174, 7192, 7212, 7229, 7245, 7267, 7284, 7306, 7322, 7338, 7352, 6410, 6429, 6450, 6467, 6491, 6508, 6530, 6550, 6572, 6595, 6616, 6642, 6660, 6683, 6703, 6722, 6739, 6757, 6776, 6797, 6817, 6842, 6858, 6876, 8202, 7371, 7391, 7938, 7964, 7986, 8008, 8029, 8056, 8079, 8098, 8118, 8142, 8161, 8185, 8208, 8229, 8252, 8270, 8287, 8310, 8331, 8349, 8373, 8392, 8413, 8437, 8458, 8482, 7407, 7423, 7443, 7460]. Inventions [7483, 7507, 7528, 7547, 7566, 7583, 7609, 7629, 7653, 7679, 7702, 7727, 7747, 7773, 7795, 7820, 7846, 7868, 7893, 7914, 7959, 8277, 7375, 7369, 7388, 7936, 7960, 7983, 8005, 8025, 8054, 8076, 8094, 8114, 8140, 8157, 8181, 8205, 8225, 8251, 8267, 8289, 8307, 8328, 8351, 8371, 8394, 8410, 8433, 8454, 7405, 8479, 7419, 7438, 7457, 7480, 7504, 7525, 7546, 7565, 7581, 7604, 7625, 7652, 7678, 7699, 7723, 7745, 7771, 7794, 7817, 7845, 7864, 7889, 7910, 8503, 8526, 9085, 9105, 9147, 9187, 9205, 9225, 9242]. Inventions [9259, 9276, 9299, 9321, 9345, 9364, 9381, 9400, 9421, 9440, 9463, 9482, 9504, 9523, 9546, 9566, 8550, 8571, 8589, 8624, 8647, 8670, 8690, 8708, 8734, 8753, 8779, 8798, 8821, 8840, 8864, 8885, 8902, 8924, 8946, 8968, 8984, 9004, 9025, 9044, 8500, 8522, 9067, 9081, 9102, 9125, 9146, 9166, 9186, 9200, 9224, 9238, 9258, 9273, 9295, 9318, 9344, 9360, 9377, 9397, 9423, 9442, 9458, 9479, 9501, 9525, 9543, 9562, 8547, 8567, 8591, 8621, 8643, 8666, 8686, 8704, 8728, 8755, 8776, 8797, 8818, 8838, 8862, 8881]. Inventions [8904, 8920, 8944, 8963, 8986, 9000, 9022, 9043, 9579, 9600, 10201, 10225, 10247, 10270, 10291, 10317, 10339, 10361, 10386, 10408, 10434, 10458, 10482, 10504, 10532, 10554, 10581, 10606, 10632, 10657, 10682, 10704, 10731, 10754, 10784, 10805, 9622, 9647, 9670, 9696, 9716, 9739, 9767, 9788, 9811, 9838, 9862, 9891, 9918, 9944, 9970, 9998, 10035, 10061, 10085, 10113, 10132, 10153, 10175, 9577, 9596, 10223, 10244, 10266, 10287, 10315, 10341, 10357, 10385, 10405, 10431, 10455, 10484, 10502, 10526, 10551, 10576, 10602, 10628, 10654, 10680, 10706, 10733, 10751]. Inventions [10783, 10804, 9617, 9643, 9668, 9692, 9713, 9735, 9766, 9786, 9809, 9836, 9857, 9888, 9920, 9940, 9966, 10000, 10031, 10058, 10087, 10112, 10134, 10155, 10171, 10197]. Inventor [3303, 5195, 5564, 7778, 10650, 19]. Inventors [833, 21, 462, 489, 511, 565, 589, 614, 703, 727, 752, 773, 793, 826, 857, 882, 904, 924, 48, 966, 985, 1005, 1024, 1046, 1071, 1094, 1114, 72, 1142, 1166, 1194, 1219, 1243, 96, 126, 148, 170, 191, 218, 241, 266, 286, 308, 332, 356, 383, 411, 435, 1269, 1289, 1802, 1822, 1840, 1862, 1883, 1902, 1923, 1939, 1956, 1975, 1991, 2010, 2027, 2049, 2067, 2085, 2103, 2124, 2144, 2157, 2177, 1312, 2192, 2217, 2238, 2261, 2282, 2301, 1337, 1359, 1385, 1429, 1452, 1481, 1504, 1530, 1553, 1578, 1609]. Inventors [1632, 1657, 1685, 1704, 1741, 1758, 1779, 2431, 2825, 2318, 2332, 2805, 2839, 2858, 2879, 2895, 2907, 2925, 2946, 2962, 3034, 3048, 3064, 3084, 3119, 3139, 3154, 3190, 3208, 3221, 2351, 2371, 2390, 2410, 2444, 2462, 2481, 2500, 2520, 2540, 2557, 2577, 2598, 2615, 2652, 2666, 2681, 2718, 2734, 2768, 2560, 3020, 3240, 3279, 3835, 3855, 3878, 3901, 3925, 3970, 3992, 4021, 4046, 4070, 4091, 4132, 4151, 4177, 4200, 4225, 4248, 4271, 3298, 4294, 4314, 4332, 3321, 3359, 3380, 3441, 3467, 3491, 3514, 3536, 3560, 3581, 3605, 3628, 3647, 3670]. Inventors [3693, 3715, 3738, 3800, 4628, 4355, 4375, 4835, 4854, 4872, 4913, 4932, 4947, 4961, 4980, 4999, 5020, 5039, 5077, 5095, 5112, 5129, 5149, 5165, 5184, 5206, 5224, 4396, 5260, 5280, 5301, 5320, 5337, 4417, 4461, 4483, 4518, 4535, 4557, 4576, 4602, 4624, 4642, 4679, 4702, 4721, 4741, 4763, 4786, 4807, 5354, 5374, 5904, 5929, 5953, 5973, 5993, 6033, 6072, 6093, 6116, 6134, 6177, 6193, 6214, 6271, 6290, 6311, 6335, 6353, 5391, 5413, 5432, 5453, 5489, 5506, 5529, 5586, 5607, 5625, 5641, 5690, 5709, 5730, 5751, 5772, 5793, 5864, 5885, 6955]. Inventors [6375, 6394, 6928, 6965, 6985, 7018, 7036, 7054, 7110, 7126, 7143, 7160, 7176, 7214, 7227, 7247, 7265, 7285, 7308, 7324, 7354, 6412, 6432, 6452, 6470, 6493, 6510, 6532, 6552, 6574, 6597, 6618, 6663, 6681, 6741, 6759, 6777, 6799, 6843, 6874, 6621, 7632, 7375, 7390, 7406, 8117, 8141, 8160, 8184, 8228, 8249, 8309, 8330, 8348, 8372, 8391, 8436, 8457, 7422, 7441, 7459, 7482, 7506, 7527, 7608, 7627, 7701, 7726, 7819, 7867, 7913, 7963, 8028, 8055, 8078, 8097, 7687, 8445, 8575, 8502, 8525, 9084, 9104, 9184, 9203, 9222, 9241, 9298, 9342]. Inventors [9461, 9545, 9564, 8549, 8570, 8588, 8623, 8646, 8669, 8688, 8707, 8731, 8752, 8778, 8819, 8884, 8922, 8943, 8966, 9003, 9024, 9041, 9065, 9598, 9620, 10703, 9646, 9695, 9738, 9861, 9890, 9917, 9943, 9969, 10060, 10110, 10246, 10268, 10290, 10360, 10407, 10529, 10553, 10579, 10604, 10630]. Inverting [6855]. Investigations [10537, 9664, 7883]. Investments [7277]. Invisible [8512]. Involuntary [8420]. Iowa [593, 1264, 3415, 6225, 6655, 6367, 7660, 7643, 4633]. Iridescent [1736]. Iron [532, 26, 970, 585, 989, 1624, 2205, 2266, 1763, 2673, 3215, 3870, 3526, 5287, 4470, 4588, 5297, 5385, 5403, 6088, 5941, 6324, 6286, 5802, 5855, 6969, 7088, 6885, 7344, 6515, 8121, 7797, 8403, 8823, 8593, 10294, 10765, 10100, 10564, 1303]. Ironclad [4666]. Irons [851]. Ironwork [8244]. Irrigating [451, 8201, 8677]. Irrigation [2162, 3873, 4868, 9313, 7412]. Isabella [7014]. Isafjord [6712]. Island [2021, 3157, 5981, 5475, 10524, 754, 3920, 4839, 6645, 7741]. Islands [988, 8908, 2712, 3325, 8888, 9721, 10096, 9995, 8972, 8833]. Isles [6200]. Isolated [10347]. Issue [8317]. Issued [8356]. Italian [4064, 4684, 9389, 10286]. Italy [8655]. Items [5911, 5980, 5999]. Itself [10506]. IV [10355, 9390].

J [194, 2137, 2554]. Jack [391, 248, 1711, 6259, 6650, 7451, 9534, 4212]. Jack-Screws [1711]. Jacket [6834]. Jacketed [8659]. Jackson [999]. Jaffa [3149]. Jamaica [1443, 3543]. James [9235, 5275, 5656, 5357, 8355, 10372]. Janeiro [140]. January [1227, 2305, 5326, 6340, 8794, 8494]. Japan [2467, 4935, 5464, 7918, 7704, 10205, 10371, 6622, 7217]. Japanese [853, 1512, 3919, 3756, 5782, 6481, 7573, 8603, 9797, 9684, 9986]. Jar [2760, 5853]. Jarecki [4259]. Jarrah [1320]. Jars [8493]. Jarvis [196]. Jaw [5368, 5595]. Jaws [5800]. Jay [3186]. Jean [4036]. Jeantaud [5598]. Jenner [6791]. Jeopardizing [8012]. Jerome [7039]. Jersey [776, 4822, 4825, 528, 581, 1039, 1966, 1251, 1877, 5161, 5120, 5415, 5766, 6408, 7464]. Jerusalem [3149]. Jet [874, 999, 5457, 7137]. Jet-Propelled [874, 5457]. Jets [2891]. Jetties [3273, 6762]. Jetty [8528]. Jeweler [3435, 8363]. Jewelers [7166]. Jewfish [4479]. Jews [1293]. Jig [10447]. Jig-File [10447]. Jigger [2396]. Job [1053, 7620]. John [4297, 9283, 9490, 10099, 8070, 841, 7804, 9485, 9793]. Johns [4061, 10280]. Joining [9846]. Joint [1040, 86, 811, 2203, 2812, 5054, 4859, 6166, 8264, 9587, 8534]. Joints [4959, 5458, 7668, 7395]. Jones [1978]. Jørgensen [9964, 10053]. Joseph [2457, 2916, 10798]. Joshua [9512]. Journal [1256, 4614, 5232, 7301, 10626]. Jowett [4105]. Juan [10595, 8937, 8953, 9136]. Jubilee [6914, 8556, 9474, 9152, 8827]. Judge [1229, 7808, 8616]. Judgment [7589]. Juggler [8498]. Juice [7026]. Jules [4630]. July [619, 1827, 2775, 3775, 5849, 9052, 10163]. Jumping [1784, 7521]. June [522, 1746, 2705, 3683, 5735, 10276, 8954, 7807, 10071]. Jungfrau [5665]. Jupiter [2043, 3019]. Justices [8489]. Justifying [7762]. Juxtaposition [3724].

Kaiser [8442, 8259]. Kalb [5823]. Kangaroo [3253, 6999, 10332]. Kansas [3811, 7835, 8466]. Katahdin [3342, 5143]. Katcinas [8409]. Kearsarge [6863, 7718, 8785, 8810, 10238, 9798]. Keating [5465]. Kedah [6368]. Keel [10453, 9712]. Keely [9485, 9664, 9642]. Kelvin [8276]. Kensington [7763]. Kentucky [7718, 8785, 8810]. Kenwood [2044]. Kerosene [6144, 8563]. Kettle [2692]. Kettles [8999]. Keuka [4075]. Kew [2836]. Key [318, 863, 1022, 1809, 5250, 6282, 9413, 10802, 8833]. Keyhole [118]. Keys [888]. Kid [6609, 6590]. Kiel [790]. Killing [898, 2185, 2890]. Kiln [34, 2553, 4119, 4525]. Kilowatt [2591]. Kind [2130, 4589, 2081, 2099, 2112]. Kindergarten [3935]. Kindred [9329, 9334, 9358, 9374, 9390, 9419]. Kinds [2472]. Kinetograph [1818]. Kinetoscope [7196]. King [65, 7014, 8204, 8562]. Kingi [6636]. Kingsley [7840]. Kioto [7217]. Kirby [3686]. Kirchner [4691]. Kit [405]. Kitchen [2433, 2491, 5762]. Kite [2656, 5053, 6570, 7417, 7933, 9477, 10725, 9734, 9874]. Kites [7877, 8560]. Klatte [5248]. Klondike [8534, 8711, 8725]. Knapp [9336]. Kneader [339, 3227]. Kneading [4164]. Knee [5255]. Knife [36, 499, 890, 994, 427, 2248, 3498, 3753, 8678]. Knife-Grinding [2248]. Knight [9662]. Knitted [2270]. Knitting [4058, 4389, 5065, 5650]. Knob [295]. Knockdown [9990]. Knot [9304, 10684, 8453]. Koch [1078, 1099, 1119, 1319, 1158]. Kokomo [990]. Kola [899]. Konzi [6487]. Koodoo [5786]. Krag [9134, 9964, 10053]. Kruesi [9793]. Krupp [1234, 1209, 3522, 3544, 3803, 9215, 10508, 9499].

L.I. [2078]. Label [8824]. Labels [8558, 9130]. Laboe [3141]. Labor [3369, 4767, 7148, 7776, 7586, 7704]. Labor-Saving [3369]. Laboratories [3105, 7972]. Laboratory [4318, 5981, 6426, 7800, 10026, 2152, 9664]. Laboring [8439]. Labors [89]. Labrador [2054]. Lac [6242]. Lace [956]. Laces [4474]. Lachine [8404]. Lacing [4634, 5072]. Lack [9547]. Lacquer [1077]. Ladder [469, 419, 3729, 4687, 4385, 7451, 7452]. Lading [1764]. Lady [944, 2063, 7349, 1055]. Ladysmith [10769]. Lafayette [2655, 5406]. Lag [3939]. Lake [886, 1805, 2399, 10644, 10617, 7289, 1079, 1237, 1805, 3128, 2336, 4075, 4456, 5908, 6611, 10509, 9286, 7217]. Lakes [5825, 6256, 1226, 942, 2231, 1325, 5006, 5805, 8397, 8601, 8661, 10622]. Lamb [6590]. Lambrecht [2275]. Lamp [184, 11, 232, 57, 14, 1909, 2223, 1446, 3055, 2605, 2780, 3326, 3979, 4215, 4967, 5085, 5309, 5697, 6445, 7061, 6849, 6523, 6460, 6527, 8338, 7789, 7953, 7712, 8242, 8041, 8858, 9532, 9903, 10592, 1326, 7347]. Lamps [4303, 3454, 5617, 7810]. Land [1100, 6060, 9820, 4597, 5908]. Landmark [8463, 9464, 10400]. Lands [1296, 7980, 7925, 8717]. Langley [2421, 5305]. Language [1282, 9171, 9430]. Languages [9247]. Lantern [1235, 1755, 1496, 3725, 4208, 4378, 5850, 6887, 1238, 3416, 3271, 4567, 4850, 2515]. Lanterns [6563, 7347]. Larder [3772]. Large [898, 2016, 2180, 2148, 1696, 2673, 3148, 3128, 2590, 3401, 3710, 4308, 4899, 4451, 4694, 4591, 5941, 5420, 5744, 8203, 7887, 8319, 9559, 4087, 10448, 165, 10149]. Larger [8574]. Largest [2231, 2290, 1815, 2687, 2514, 6108, 7519, 7448, 9288, 8662, 10206, 8278]. Lark [4976]. Lasso [8469]. Last [9674, 10323, 7934, 10401]. Latch [665, 7782]. Latches [1494]. Late [4403, 7490, 9512, 9709, 10796, 9854, 9689]. Lateen [8640]. Latest [3974, 5455, 8241, 7361, 8013, 8341, 8581, 9404, 9455, 9347, 10196, 9897, 10238, 10427, 10234]. Lath [8891]. Latham [1849]. Lathe [1870, 2584, 3886, 3475, 4159, 4139, 4686, 4382, 6485, 6827, 6606, 8177]. Lathes [1774, 1445, 2272, 1569, 6925]. Lathing [4490]. Latimer [9467]. Latitude [6255, 7711]. Lauffen [2422]. Launch [1132, 302, 141, 2211, 2870, 3123, 3079, 3741, 5071, 5235, 6152, 5539, 5921, 6425, 6655, 8001, 8562, 8951, 8785, 9372, 8810]. Launched [8603]. Launches [4229, 5783]. Launching [7799, 8664, 9931]. Laundry [3251, 5520]. Lausanne [6385]. Lava [2061]. Law [920, 839, 519, 3424, 7635, 8291, 7571, 8295, 8234]. Lawn [913, 448, 1849, 5984, 6296, 6626, 8152, 8344, 8742]. Lawns [4868]. Lawrence [10757, 4724, 7944, 7920]. Laws [569, 2466, 4767, 8293, 8416, 7531, 8212, 8653, 9382, 9603, 7777, 9448]. Lawton [4119]. Layard [4994]. Laying [1248, 3113, 2796, 4901]. Layman [5704]. Lead [5031, 5601, 7255, 9158, 10492]. Leading [5238]. Leads [26, 2266]. Leaf [406, 6544, 7784]. League [6475]. Leak [59, 7786, 7811, 8478]. Leaky [5329]. Leamy [6428]. Leaning [6679]. Learning [4422]. Leather [44, 4032, 5937, 5738, 6420]. Leathern [3635]. Leber [6283]. Lecornu [10725]. Lecture [2492]. Ledger [474, 10128]. Lefebvre [7182]. Lefthandedness [4717]. Legal [3471]. Legislation [6555, 6667, 6536, 7446]. Legislative [7849]. Legitimate [9844]. Legs [3150]. Leipsic [7832]. Leipsig [4781]. Leipzig [2921, 4013, 4737, 5634]. Leland [8194]. Lemon [9961]. Lemon-Squeezer [9961]. Lemonade [319]. Lemons [8765]. Lenepe [3812]. Lengthening [4864, 8376]. Leni [3812]. Lens [68, 9877]. Lens-Grinding [9877]. Leonardo [8835]. Leprosy [222, 2071, 2055]. Less [10793, 10241]. Lesson [502]. Lessons [4790, 6019, 5437, 7509, 8925, 9107]. Let [7219]. Let-Off [7219]. Letter [88, 4918, 6237, 5633, 6848, 6666, 10692]. Letter-Box [10692]. Lettering [1816]. Letters [1964, 10017]. Level [1126, 1982, 3619, 4528, 5514, 10426]. Leveler [936, 1296]. Leveling [3547, 6085]. Levensau [5813]. Lever [9309]. Levers [1419]. Li [7076, 7065]. Libbey [8040]. Liberal [3917, 3078]. Librarian [7947, 9629]. Libraries [3842]. Library [3883, 7232, 7855, 8193, 6226, 8065, 8082]. Licensing [8273]. Lick [5521, 8737, 6622]. Lieut [8470]. Lieut. [402, 9794, 10469]. Lieutenant [6142, 8443]. Life [598, 544, 1108, 1817, 1421, 2656, 2465, 3133, 3659, 5011, 4887, 5314, 5286, 5891, 6126, 5369, 5915, 5718, 6932, 9197, 10294, 9869, 10240, 10302, 10789, 1576]. Life-Saving [3133, 10302, 10789]. Lifeboat [874, 5457, 7137, 6503, 8542, 8992]. Lift [1700, 4755, 7270, 6969, 7182, 9265]. Lifter [9]. Lifting [918, 1732, 3039, 4545]. Light [9773, 1050, 1061, 361, 11, 1082, 2, 1235, 1929, 1437, 2053, 2183, 2021, 3051, 2778, 3859, 4232, 3725, 4086, 3957, 3330, 3892, 3757, 3907, 4903, 4683, 5052, 5897, 5398, 5576, 6836, 6396, 7993, 7945, 8245, 8366, 8404, 7900, 8852, 9346, 9424, 9050, 9228, 10368, 10722, 9781, 9293, 2455, 2515, 7013, 3228]. Light-Weight [9346, 9424]. Lighter [420, 3891, 3658, 5309, 9218]. Lighthouse [2727]. Lighthouses [2954, 10693]. Lighting [582, 444, 455, 61, 75, 1960, 4303, 5108, 5160, 5713, 5740, 7969, 10491]. Lightning [2992, 3307, 5033, 5034, 5495, 6280, 5788, 7103, 10349, 9608, 4612]. Lights [3962, 6569]. Lightship [6879]. Lightships [2954, 2889]. Lilliputian [9308]. Lily [2836]. Limbs [5397, 5932]. Lime [2553]. Limitation [4813].

Limitations [5396, 9323, 9748]. Limits [8467]. Lincoln [764]. Line [1228, 472, 761, 574, 1641, 2656, 2417, 2973, 4751, 5102, 4564, 8387, 9290, 10078, 2686, 5160, 4645, 5235, 7345, 6459, 9557]. Line-of-Battle [10078]. Linen [10128]. Liner [8259, 9141, 10329, 9688]. Liners [10196, 10233]. Lines [193, 186, 1299, 5511, 5718, 10609, 10169, 10167, 8340]. Lining [5518]. Link [275, 9589]. Links [9, 4798]. Linoleum [113]. Linotype [4358]. Lion [9965, 2890, 10170]. Liquefaction [7854, 8626]. Liquefied [2801]. Liquefying [6652]. Liquid [603, 2969, 4309, 10094, 10092, 10301, 10558, 10115, 9652, 10446, 9928, 8859, 8792, 10234, 10162]. Liquid-Measuring [2969]. Liquids [1230, 2322, 2693, 7274, 6868, 9850, 1753]. List [1407, 1507, 1556, 1388, 1533, 1635, 1660, 6911]. Lister [273]. Litigation [4337]. Litter [5249]. Little [1329, 5786]. Live [3309, 4774, 10716]. Liverpool [9214]. Living [8535, 10727]. Lizard [6636]. Lizards [8284, 10218]. Lloyd [7758, 7880]. Loa [3373, 10401]. Loaded [5701]. Loader [4365, 5363]. Loading [664, 3769, 4778, 7116, 8764]. Loads [7290]. Lobster [1122]. Lobsters [2561]. Lock [312, 84, 524, 1206, 362, 1519, 1516, 2249, 1809, 1797, 3108, 2639, 3004, 4277, 4968, 5424, 8107, 9413, 9354]. Locking [734, 5200, 6168, 7928, 9230, 9803, 10743]. Locks [888, 5253, 10402, 7558, 10757]. Locomotive [938, 220, 1291, 2078, 2097, 1730, 1815, 2603, 2674, 2430, 2484, 2814, 3429, 3871, 3430, 3391, 3598, 3650, 5257, 5212, 4826, 4509, 4780, 5308, 6120, 5966, 6944, 6991, 7242, 6865, 6706, 6805, 7290, 7165, 6477, 6993, 7187, 7755, 7662, 8133, 7549, 7418, 8367, 7800, 8088, 7764, 9492, 9331, 9453, 8830, 8662, 9308, 8682, 9309, 9272, 9454, 9728, 10382, 9879, 5063, 9914, 9789]. Locomotive-Tenders [9272]. Locomotives [1082, 2212, 1696, 1815, 1473, 2691, 2546, 3870, 3617, 4086, 4400, 5847, 6079, 7688, 9446, 9287, 9549, 8530, 8780, 9098, 8627, 8553, 10420, 10343, 10046, 10230]. Locust [1918, 4829]. Loeb [5582]. Loftiest [5295]. Lofty [7147, 6578]. Log [716, 664, 5363, 7743, 3408]. Logging [535, 720, 2628, 5015, 7335, 6650, 7187, 7418]. Logs [10768]. London [1134, 61, 3915, 4775, 4942, 4815, 6344, 6999, 8199, 8399, 7878, 8594, 9128, 8715]. Long [1327, 1439, 1306, 3010, 3345, 5099, 4768, 5981, 6202, 5455, 7601, 7741, 7532, 9234, 10563, 10697, 6170, 9510]. Long-Distance [6202]. Long-Running [5099]. Long-Scale [10697]. Longest [7757]. Looking [1392]. Looms [7219]. Looped [2904]. Loose [7450, 2872]. Lord [3111, 6978, 8276]. Lorillard [141, 2311]. Lorraine [7014]. Loschwitz [7114]. Loss [3158, 5564, 9089, 8680, 9468, 9506, 10219, 10039]. Losses [9151]. Lost [9167]. Lottery [7271, 7946]. Louis [6972, 676, 2684, 3661, 5235, 4951, 5801, 5784, 5763, 6913, 7456, 6904]. Louisa [2580]. Louise [4276]. Louisiana [6725, 7149]. Loup [2850]. Louse [10671]. Louvre [2817]. Lovell [343, 1603, 2566, 3368, 4508, 6351, 5652]. Lovering [2457]. Low [28, 314, 2813, 4647, 4491, 8426, 9122, 7066, 7767, 144]. Low-Priced [4491]. Lowe [9393, 5124]. Lowell [8261]. Lower [2377, 6891]. Lowering [2931, 10281, 10776]. Lozere [8808]. Lubricator [2508, 3771, 4447, 5477, 6022, 7198, 7949]. Lubricators [4076]. Lucida [3474]. Ludwig [5188]. Lugano [2339]. Lumber [1211, 2229, 6123, 5871]. Lumbermen [1897]. Lumière [7905]. Luminous [118, 3288, 4192]. Lunar [6329]. Lunch [7293]. Lungs [1231]. Lunken [3956]. Luther [7550]. Lymph [1319, 2523]. Lyon [1012]. Lyons [3745, 4971].

M [7804]. M. [635, 1016]. M.C.B. [2119]. M.D. [6791]. Macaroni [3662]. Maché [7398]. Machias [4864]. Machine [709, 366, 115, 601, 430, 622, 499, 44, 664, 113, 604, 376, 208, 301, 671, 344, 623, 1208, 972, 763, 1080, 88, 1067, 326, 550, 1131, 608, 1747, 2247, 1488, 1816, 1626, 1886, 2248, 1474, 1794, 1625, 1911, 2198, 2228, 1810, 1929, 1783, 2182, 1497, 2076, 3023, 3129, 2967, 3182, 2378, 2742, 2988, 2968, 2487, 2530, 2702, 2764, 2509, 2993, 2449, 2545, 2590, 2398, 2400, 2849, 2796, 2474, 2362, 4260, 4185, 3551, 3450, 4280, 4302, 3938, 3347, 4164, 4321, 3936, 4063, 3701, 3247, 3229, 4095, 4121, 4142]. Machine [3774, 4340, 4543, 4941, 4409, 4885, 4730, 5047, 4792, 4694, 4426, 4544, 5305, 5219, 4752, 4997, 5049, 6151, 6245, 5759, 5461, 5719, 5536, 6061, 5816, 5738, 6206, 6241, 6171, 5510, 5997, 5565, 5733, 6675, 6520, 6399, 6441, 6992, 6624, 7064, 6709, 7237, 6918, 7258, 6420, 7082, 6947, 6769, 6693, 7575, 7640, 8360, 8203, 7536, 8298, 7671, 7690, 7474, 7762, 7520, 9548, 9174, 9310, 8765, 9211, 8974, 9212, 10645, 9952, 10589, 9958, 9877, 10395, 10056, 9637, 9850, 10073, 9753, 9826, 10242, 7541, 7901, 5106, 10149]. Machinery [556, 2800, 4054, 4057, 4542, 5650, 5978, 6948, 7148, 7776, 7826, 8221, 9074, 9338, 4109]. Machines [866, 687, 300, 688, 525, 865, 1256, 1565, 2342, 2623, 4058, 4062, 3563, 4729, 4753, 4389, 5065, 4546, 7255, 6666, 7622, 8891, 8978, 10049, 9554]. Machinist [254]. Mackay [6350]. Made [695, 673, 1667, 6946, 7303, 9047, 8741, 10517, 8496, 10518, 6676, 885]. Madison [3982, 5436, 5416, 6436, 3954]. Madrid [4195]. Maestro [10653]. Magazine [224, 1680, 1985, 2606, 3519, 5177, 4581, 5294, 9333, 9095, 10038]. Magazines [2133]. Magic [250, 2713, 3658, 7105, 8066, 8023, 8170, 9797, 2515, 8305]. Magician [10548]. Magnesium [5602, 10668]. Magnet [6676]. Magnetic [1283, 3849, 4212, 5967, 6133, 7590, 8564, 2671, 3939]. Magneto [4843]. Magnetographs [6676]. Magnets [5425]. Magnifying [1283]. Magnitude [5025, 2455]. Mahogany [5048]. Mail [558, 670, 3193, 4454, 5851, 7948, 8425, 8015, 8069, 8089, 8155, 8844, 9169, 10391, 3549]. Mailing [4730]. Main [2382, 3252, 3328, 8279, 9561, 2422]. Maine [1132, 3461, 5080, 6100, 5342, 5785, 8671, 9498, 8926, 8680, 8723, 8691, 8799, 8814, 9160]. Mains [1962, 2022, 10564]. Maintenance [3674]. Majestic [360, 5681]. Majolica [648]. Major [9180]. Major-Generals [9180]. Make [821, 4487, 9149, 8735]. Maker [3327, 2225]. Makers [1181, 7166]. Makeshift [4693]. Making [623, 1131, 1626, 1794, 2116, 1299, 2938, 2475, 2310, 3997, 4451, 4732, 6633, 7809, 8658, 9959, 9757, 9805]. Malay [5053, 7417]. Malformation [9540]. Malleco [1280, 5157]. Mammoth [5048, 6110, 7570, 9492, 9431, 10661, 1915]. Man [3530, 5081, 8169, 7397, 9227, 10486, 10054, 2762, 4424, 8145, 9079].

Management [3002]. Manchester [2324, 2188, 3941, 4903, 4522]. Maneuvers [7509]. Manger [3293]. Mangle [3703]. Manhattan [4799, 5314, 6527]. Manhole [609]. Manila [9009, 9140, 8917, 8925, 9687]. Manipulating [910]. Manitou [1318]. Mannesmann [4142]. Manometer [1713]. Mantels [6044]. Mantidae [2729]. Manual [247]. Manufactory [608, 4907]. Manufacture [981, 675, 291, 299, 1182, 2241, 1785, 1436, 1485, 2955, 2783, 2684, 3088, 4287, 3385, 3662, 4263, 3234, 4594, 5072, 5217, 4471, 4349, 4692, 5014, 4620, 4956, 4845, 5932, 6306, 6007, 6080, 5617, 5601, 6044, 5888, 5572, 6164, 6130, 5559, 6790, 6535, 6609, 6623, 6500, 8021, 7398, 7616, 9080, 8748, 8865, 9357, 10492, 10399, 10128, 9964, 10053, 10515, 990, 246, 9719, 3184]. Manufacturers [833, 8575]. Manufactures [3917, 9401, 10065, 9775]. Manufacturing [3039, 4259, 4061, 4849, 8477]. Manuscript [8807]. Map [2864, 5638, 7570]. Maps [16, 9982, 8496]. Marble [5945, 3302]. Marbleized [6044]. March [350, 223, 1438, 2447, 3384, 5494, 7534, 8695, 9726]. Marconi [10490, 10535, 10662]. Mareorama [9780]. Maria [2580, 3040, 3943, 9468, 9328]. Marine [1214, 4756, 4993, 4883, 6060, 5702, 7923, 9447, 10617, 9111]. Marino [9375]. Maritime [2414, 8758, 6569]. Mark [661, 6536, 6990, 7067, 7992, 8293, 8012, 7513, 7553, 10355, 10511, 10373, 5448]. Marker [1154, 4844]. Market [3768]. Marking [9988]. Marks [4096, 4252, 5377, 8558, 10635, 10638, 9606, 9844, 10740]. Marquette [8091]. Marquis [3543]. Mars [2898, 2904, 2903, 2864, 5116, 5288, 6521]. Marsh [9854]. Marshall [8022]. Martha [7122]. Martin [4036]. Marvelously [4491]. Marvin [9956, 8045]. Mashonas [6923]. Mask [7954]. Masonic [4444]. Masonry [2687, 9749]. Mass [6037, 2676, 4120, 4861, 6057, 8109]. Massachusetts [708, 3741, 6734, 7318, 8217, 7930, 10788]. Massage [9386]. Masses [10667]. Massys [7497]. Mast [746, 7810]. Masted [4498]. Masthead [2939]. Mastodon [2385]. Mat [3825]. Matabeles [6923]. Match [4211]. Matcher [10543]. Matcher-Head [10543]. Material [1030]. Materials [6127, 7007, 7927, 8830]. Mathematical [3006]. Matter [10391, 7066]. Matting [9800]. Mattresses [4160]. Mauna [3373, 10401]. Mauser [1680, 10488]. Maxim [9417, 1783, 4997, 5049, 6141, 7839, 4636, 5106]. May [389, 3587, 4421, 9047, 10065, 1668, 2620, 3574, 5677, 6837, 7710, 6707, 8870, 9955]. Mayer [8022, 7994]. Mayflower [7743]. Maynard [1710]. Mayor [6978]. Mcalpine [196]. McBride [1670]. McClure [8062]. McConnell [3996]. McDougall [1845]. McFarlan [6514]. McGill [8856]. McKinnon [8664]. Mcleod [696]. Meadow [1296]. Meadows [7464]. Means [313, 378, 2855, 3635, 3290, 3874, 3906, 4631, 7106, 9821, 10422]. Measure [233, 2547, 5871, 6605]. Measurement [4944, 7619, 10104, 6678, 5442]. Measurements [6922, 1915]. Measures [1089, 7006, 7733, 8735, 89]. Measuring [396, 2247, 1869, 1671, 2969, 2418, 3787, 3637, 5960, 5362, 6975, 6716, 6711, 7274, 7432, 8701, 8912, 8610, 10697, 9719]. Mechanic [247].

Mechanical [915, 151, 1912, 1363, 1568, 2200, 3146, 3025, 2853, 4033, 5230, 6690, 7852, 7909, 8215, 9978, 10010, 10591, 10068, 5483, 4188]. Mechanically [10449]. Mechanically-Operated [10449]. Mechanics [1431, 10256]. Mechanism [494, 1441, 2111, 1493, 1564, 1473, 2001, 7219, 6362, 6458, 6542, 8876, 9727, 9823, 10056, 10523, 10652, 10743]. Medal [7489, 9418, 7, 10283]. Mediæval [4666]. Medical [5137]. Medicine [9127, 10301]. Medicines [7059, 8295]. Medium [5798, 8638]. Meeting [970, 892, 27, 1437, 2774, 2929, 3982, 4897, 4952, 4985, 4746, 6037, 7852, 9447, 10184, 9978, 8049, 10691, 871, 7077, 7098, 9263, 8122, 8148, 7829, 7830]. Meigs [2384]. Melicerta [5970]. Melted [10393]. Melting [2393, 7520]. Melville [9806]. Membranes [3486]. Memorial [842, 5350, 7717, 991]. Memory [1016, 5195]. Men [798, 829, 6938, 8996, 9367, 10716]. Men-of-War [8996]. Menagerie [4598, 6612]. Mending [1775]. Mental [8169, 9521]. Menu [10479]. Mercantile [10423]. Mercedes [10097, 10166]. Mercerization [9339]. Merchant [152, 676]. Mercurial [3412]. Mercury [5154]. Merits [7752, 9744]. Merrimac [7170, 7561]. Mersey [9214]. Mesa [8040]. Message [3636, 7328]. Metal [278, 1063, 547, 607, 867, 943, 1518, 2255, 1646, 5047, 5423, 5556, 7172, 7590, 7997, 8716, 9507, 9973]. Metallic [1175, 176, 811, 758, 1748, 1749, 1829, 2273, 2594, 3727, 4239, 4490, 4672, 6317, 6851, 7873, 8198]. Metallo [1262]. Metallurgy [10615]. Metals [4336, 4359, 4664, 4506, 5576, 7921]. Metempsychosis [8180]. Meteor [10686]. Meteorite [5799]. Meteorites [593]. Meteorograph [5099]. Meteorological [8560]. Meteors [9469, 5069, 3194, 10597]. Meter [3655, 4242, 5554, 5621]. Method [584, 185, 1712, 1541, 2830, 3502, 3370, 4410, 4695, 7432, 8074, 8477, 7416, 8989, 9959, 10185, 9778, 10640, 10217, 10011, 2360, 6678, 7767]. Methods [1716, 1865, 4422, 8083, 7374, 9325, 10693, 9199]. Metric [2147, 7006]. Metropolitan [5654, 3203, 5142, 4847, 5684, 9431]. Mexican [1784, 7134, 8367, 8931, 10376]. Mexico [3828, 839, 8439, 1967, 9436]. Miantonomoh [2403, 3665]. Michigan [1237, 2035]. Micro [402]. Micro-Graphophone [402]. Micrograph [5399]. Micromotoscope [8024]. Microphone [3363]. Microscope [6147, 9993]. Microscopic [1969, 7163, 1814]. Microscopical [2619]. Microscopy [1850]. Middlings [31]. Midland [10046]. Midwinter [4306, 4467, 4838, 4635, 4886, 4653]. Migrations [9847]. Mile [6747, 10241]. Mileage [10209]. Miles [3069, 3632, 4561, 5266, 7942, 9180]. Military [1239, 1876, 2058, 3009, 5293, 9933, 7454]. Militia [1965]. Milk [1229, 249, 675, 1692, 3958, 4031, 4007, 4663, 7205, 6765, 9054]. Milk-Cans [9054]. Mill [257, 280, 993, 1181, 289, 2255, 2150, 1672, 3073, 2869, 2505, 4241, 3918, 3773, 3710, 4616, 5379, 5941, 6850, 6630, 8151, 8423, 9638, 3981]. Millennial [6770]. Milling [1445, 7064, 7671]. Million [8586, 10212]. Mills [1416, 2324, 7049, 10543, 4362]. Millstone [784]. Millstones [3290]. Milouna [7863]. Mine [3448, 4391, 6805, 9727, 9983, 7386, 7102, 7120]. Mine-Doors [9727]. Miner [3454, 5697, 6546, 10330]. Mineral [2832, 7511, 9721]. Miners [2337, 6710]. Mines [1444, 4953, 7049, 9316, 9752]. Miniature [1833, 3270, 7765]. Mining [1889, 2212, 2056, 2487, 3434, 4539, 6098, 6084, 6961, 6993, 8151, 7826, 8699, 10539]. Minneapolis [4924]. Minnesota [2628]. Minor [5061]. Minute [10241]. Miraculous [9123]. Mirages [5126]. Mirror [1623, 1589, 4563, 8812]. Mirrors [4147, 9797]. Miscellaneous [1354, 1562]. Miss [7840]. Mission [818, 7295]. Missionary [6979]. Mississippi [7802, 7751, 9252, 8962, 9217, 10310]. Missouri [3302]. Miter [8766]. Mitering [376]. Mixed [6078, 7616]. Mixer [971, 1912, 3618, 7080, 8956]. Mixing [4164, 4524]. Mo [1967]. Mobile [10495]. Mode [5315]. Model [2297, 3598, 6181, 6523, 7301, 7063, 9453, 10220, 1698]. Models [3755, 10218]. Modern [746, 504, 3405, 3251, 4284, 4816, 4966, 6227, 6236, 6880, 6359, 6727, 6897, 7972, 7466, 8190, 8533, 9068, 9195, 9535, 9560, 10402, 10523, 7317, 7046]. Modified [6765]. Mohawk [4324]. Moistener [529, 1601, 5196, 6243, 9175]. Mokis [10168, 10425]. Mold [10258]. Moldau [7686]. Molecular [6782]. molecules [2455]. Molten [1763]. Monarch [121]. Monazite [5803]. Monceau [4167]. Moncton [10720]. Money [250, 3327, 3285, 5739, 10464]. Monitor [245, 3665, 7251, 7568, 9178, 9425, 4367]. Monitors [8929, 9435, 9486, 9732]. Monkey [7181, 8472, 9369]. Monkeys [998]. Monopoly [5325]. Monster [4011, 3391, 4063]. Monstera [4802]. Monsters [3833, 3393]. Mont [4285, 5099]. Montana [3829]. Montauk [8490]. Monterey [2600]. Montgomery [2384, 4476]. Months [7683]. Monument [1016, 4108, 5195, 6297, 7841, 10718, 2696, 2626, 5521, 7201, 8246, 9733]. Monumental [3288]. Moodus [8577]. Moon [322, 2898, 3019, 2903, 2568, 5036, 5736, 5404, 8809, 6330, 9635]. Moonshiner’s [4264]. Moore [6514, 7969]. Mop [1036]. Mordey [2591]. Moretus [7224]. Morley [6038]. Morris [1732, 9079]. Morse [4024]. Mortar [8835, 8822]. Mortars [4308, 7366]. Mortise [6700]. Mortiser [6275]. Mortising [1911, 3129, 6151, 5565]. Morton [6106]. Mosasaurs [10192]. Moscow [6846]. Moses [2865, 3673]. Moss [9975]. Mossambicus [3026]. Mossberg [10191]. Most [42, 4424, 7688]. Mostar [5270]. Moth [708, 10696]. Mother [2921]. Motion [1304, 2844, 6261, 7189, 10194, 9853]. Motive [3785, 6277, 9429]. Motocycle [6281, 6163, 6293, 6247, 6223]. Motocycles [6792]. Motor [602, 363, 757, 110, 1676, 1907, 1928, 1833, 2131, 1370, 1391, 2115, 2091, 2200, 2354, 2847, 2309, 2623, 3168, 4100, 4187, 3250, 3453, 3330, 3455, 3851, 4818, 4425, 4975, 5098, 4466, 5256, 4650, 6278, 6040, 5680, 5556, 6318, 5783, 5933, 6305, 5524, 7316, 6485, 7260, 6448, 7312, 6654, 6690, 6486, 6902, 7008, 7434, 7790, 7896, 7973, 8401, 8344, 7858, 8387, 7998, 8721, 8578, 9156, 9289, 8877, 9196, 8595, 8915, 10649, 10021, 10023, 10008, 9662, 10768, 9706, 10280, 9642, 10167, 10007, 10018, 1305, 4310]. Motor-Carriage [10023]. Motor-Controller [9706]. Motors [193, 1173, 2254, 2834, 6803, 6946, 8949, 10295, 9730]. Motthaven [4705]. Mould [1109]. Moulder [10591]. Moulding [6206]. Moulds [1208, 8385].

Mount [581, 2339, 5922, 8123, 9098, 9393, 9284, 9416]. Mountain [1148, 4435, 8367, 7696, 8327, 9331, 9393, 9313]. Mountains [322, 7335, 10545, 9832]. Mounted [2992, 3811]. Mounting [2658, 7333, 10011, 4087]. Mouse [6767]. Mousetrap [2206]. Mouth [8528, 3273]. Mouths [850]. Movable [744, 4207, 4215]. Movement [915, 3025]. Movements [8420]. Moving [1914, 3335, 3470, 4876, 4705, 7673, 8447, 7739, 8065, 10260]. Mower [448, 5984, 6296, 8344, 8742]. Mowers [1849]. Mowing [10395]. Mowing-Machine-Sickle [10395]. Mr [141]. Mr. [1213, 1345, 2528, 6449, 7760, 8576, 5106]. Mrs. [2580, 5428]. Mucilage [8381]. Muffler [2794]. Muir [2565]. Muller [4801]. Multiphase [8490]. Multiphone [9848]. Multiphotography [5105]. Multiple [2933, 5292, 8529]. Multiple-Cylinder [8529]. Multiplex [7276]. Multiplication [8294, 9373]. Multipolar [4793]. Mummies [9822]. Mummy [258]. Mungsten [10426]. Municipal [7924, 10140, 4401]. Mural [3872]. Murphy [10229]. Museum [3122, 7333, 7647, 9676, 4814, 5628, 7224, 7588, 7856]. Mushroom [4051, 9008]. Music [406, 2207, 3788, 5136, 6544, 7707, 7784, 8044]. Musical [157, 2364, 2346, 3263, 3479, 5914, 7257, 7259]. Mussuck [5199]. Mysterious [4128, 3789, 5485, 8880, 10646].

N.H. [2324]. N.J. [372, 581, 2019, 2169, 3091, 3894, 5415, 7888]. N.P. [225]. N.Y. [819, 1182, 628, 674, 592, 1965, 1791, 2382, 4324, 5981, 6009, 7060, 7851, 7077, 7098]. Nadrai [2265]. Nafew [5652]. Nahant [9178]. Nail [36, 8092]. Nails [4368]. Najork [6318]. Name [3518, 9193]. Names [9844, 9070]. Naming [9923]. Nansen [6559, 7029, 8324, 7489]. Naphtha [5539]. Naples [8918]. Nasmyth [3940]. Nassau [225]. Natal [9994]. Nathaniel [4403]. National [6, 3697, 5837, 6744, 10489, 10365, 10641, 1437, 2053, 3254, 5678, 5436, 5416, 5733, 6750, 7427, 7900, 9493, 10494, 10322, 9954, 10691]. Nations [9605]. Native [9900, 9436]. Natural [3122, 2703, 3372, 4526, 5600, 6154, 7333, 6767, 8972, 3211]. Naturalist [3126]. Nature [10105]. Nautical [5768, 7363]. Nautico [1188]. Nautico-Terrestrial [1188]. Navaho [7412, 10222]. Naval [1888, 1965, 7658, 9447, 8710, 9527, 8917, 9107, 8724, 10118, 9908, 1030, 1460, 1773, 2636, 2918, 4182, 3404, 3539, 3584, 3840, 4841, 6339, 6949, 7509, 8235, 7634, 8533, 8584, 8802, 8887, 8761, 8814, 9383, 9384, 9722, 9795, 10769, 10713, 10662, 9899, 9719]. Navies [8313, 8969, 8793, 8825, 9476, 9519, 9641, 9782, 9938, 10107, 10284, 10499]. Navigate [595]. Navigating [5026]. Navigation [2414, 3905, 4338, 5917, 6325, 5912, 6315, 6727, 10608]. Navy [1132, 1511, 3696, 7024, 6790, 6444, 7786, 7811, 8102, 8478, 7499, 8905, 8916, 8683, 9454, 8895, 10462, 9753, 1118, 1586, 2163, 1847, 1512, 1460, 1436, 1485, 1619, 2353, 3212, 4230, 4255, 3732, 4256, 4581, 5228, 4917, 5580, 6601, 8072, 7542, 9414, 9280, 9162, 8581, 8906, 9210, 9159, 8845, 9508, 9315, 9207, 9129, 9294, 8781, 8950, 9732, 9897, 10138, 10398, 10585, 9772, 9867, 10610, 99]. Néant [6545]. Near [1148, 568, 2339, 2977, 5813, 5725]. Neat [645]. Nebula [5427, 10019]. Neck [1999, 2113, 2970]. Neckar [2422]. Necktie [1255]. Neckties [6067]. Nectarophora [10671]. Need [2964, 5228, 8906, 8574, 10297]. Needed [443, 1414, 2353, 2685, 4230, 5633, 6555, 8334, 9603, 10159]. Needle [2182, 5030, 6442, 6443]. Needs [10585]. Neff [3750]. Negative [2453]. Negotiable [8843, 9048]. Neisse [5418]. Neostyle [9496, 7882]. Nernst [9781]. Nerves [10182]. Nest [3509]. Net [8321]. Neutral [2543]. Nevada [7335]. Newark [302, 1249, 1947, 9116]. Newcomb [9816]. Newcomen [5175]. Newfoundland [10099]. Newly [8749, 10442]. News [7488, 9404, 10444, 10469]. Newspaper [4434, 9547]. Newspapers [1690, 3540, 7470]. Newton [1932]. Ney [7642]. Niagara [10204, 1107, 3043, 2374, 2473, 5006, 5139, 5552, 6600, 6364, 6423, 6632, 9019, 9516, 8831, 10151, 10636, 10647, 10262, 10074, 9585, 7715, 7692]. Nicaragua [1124, 1343, 2791, 3177, 3403, 7750, 7467, 9744, 10708, 9699, 9711, 10116]. Nickel [953, 2579, 10388]. Nickel-Steel [953]. Nicolas [10524]. Nicotine [8176, 8042]. Niger [4927]. Night [4215, 5797]. Nikola [9455]. Nile [8366, 7828]. Nine [2584, 5347, 7740]. Nine-Inch [2584]. Ninety [5966, 8586]. Ninety-Six [5966]. Ninth [6027]. Nipper [5086]. Nipple [2153]. Nitrate [1444]. Nitric [2773]. Nitro [2830, 5765]. Nitro-Glycerine [2830, 5765]. Nivet [7007]. No [209, 768, 3678, 7024, 6879, 6444, 7664, 8102, 8478, 99]. Nobel [7941, 9404]. Noble [8536]. Noise [5210]. Noiseless [8719]. Noises [8577]. Non [814, 6994, 7200, 8106, 8153, 9708, 9685, 2079, 2120]. Non-Refillable [6994, 8106, 8153, 9708, 9685]. Non-Sinkable [814]. Non-Venomous [2079, 2120]. Nonpareil [370]. Noriega [4218]. Normannia [545]. North [1038, 10218, 1216, 2170, 1890, 1325, 4267, 3984, 4645, 5253, 5813, 5563, 5908, 5870, 5899, 5814, 6182, 7047, 6809, 7991, 7758, 7880, 8378, 10286]. Northeastern [7027]. Northern [9451, 4013, 8345]. Northwest [7956]. Norton [1108]. Norway [1125]. Norwich [5033]. Nos [6304, 6973]. Nose [6714]. Noses [10424]. Notable [1595, 2938, 4675, 4400, 7291, 8188, 7957, 9827]. Note [7472, 5970]. Notes [53, 858, 22, 47, 683, 704, 726, 750, 771, 824, 856, 902, 925, 947, 965, 984, 1003, 1026, 1045, 1072, 1095, 1115, 1141, 1164, 1165, 1191, 1192, 1220, 71, 95, 125, 147, 169, 190, 217, 239, 265, 285, 307, 331, 355, 382, 410, 436, 461, 488, 510, 541, 563, 588, 613, 655, 24, 50, 969, 987, 1008, 1029, 1048, 1074, 1097, 1117, 1144, 1168, 1196, 1225, 1247, 74, 98, 128, 150, 243, 385, 413, 440, 464, 493, 517, 543, 567, 591, 616, 728, 775, 803]. Notes [838, 859, 884, 907, 928, 951, 54, 76, 161, 197, 399, 521, 580, 678, 780, 820, 876, 893, 922, 1021, 1042, 1133, 1354, 1562, 1633, 1268, 1288, 1800, 1821, 1841, 1860, 1881, 1904, 1922, 1937, 1954, 1974, 1990, 2011, 2028, 2048, 2066, 2084, 2104, 2125, 2143, 2156, 2176, 2193, 2216, 2236, 2262, 2281, 1310, 2300, 1311, 1334, 1335, 1360, 1382, 1383, 1404, 1427, 1428, 1449, 1450, 1480, 1505, 1531, 1552, 1577, 1608, 1658, 1684, 1703, 1721, 1738, 1778, 1316, 1340, 1925, 1935, 1958, 1977, 1993, 2013]. Notes [2030, 2051, 2069, 2087, 2106, 2127, 2146, 2179, 2195, 2219, 2240, 2264, 2284, 1364, 1389, 1409, 1432, 1434, 1455, 1483, 1509, 1534, 1558, 1583, 1614, 1687, 1706, 1723, 1743, 1760, 1781, 1804, 1824, 1843, 1864, 1885, 1906, 1260, 1374, 1470, 1769, 1896, 1910, 2005, 2256, 3160, 2317, 2331, 2788, 2806, 2823, 2840, 2859, 2877, 2894, 2926, 2945, 2961, 2982, 2999, 3018, 3033, 3049, 3065, 3085, 3103, 3120, 3138, 3155, 3172, 3191, 3209, 3222, 2350, 2370, 2389, 2409, 2428, 2443, 2463, 2482, 2499, 2519, 2538]. Notes [2558, 2576, 2597, 2616, 2634, 2651, 2665, 2682, 2700, 2717, 2733, 2752, 2767, 2842, 2861, 2909, 3067, 3087, 2550, 2710, 2781, 3239, 3280, 3817, 3836, 3856, 3879, 3902, 3926, 3950, 3968, 3991, 4020, 4045, 4069, 4092, 4114, 4133, 4152, 4178, 4201, 4224, 4247, 4272, 4293, 4315, 4331, 3299, 3322, 3339, 3358, 3379, 3399, 3421, 3440, 3466, 3490, 3513, 3535, 3559, 3580, 3604, 3627, 3645, 3669, 3692, 3714, 3737, 3761, 3781, 3799, 3520, 3542, 4138, 4157, 4183, 3565, 3611, 3653, 3676, 3699, 3721, 3743, 3767, 3786, 3805]. Notes [3823, 3843, 3863, 3885, 3908, 3932, 3964, 3976, 4000, 4026, 4053, 4079, 4098, 4118, 3274, 3242, 3259, 3819, 3838, 3858, 3881, 3904, 3928, 3952, 3972, 3995, 4023, 4048, 4072, 4094, 4116, 4135, 4154, 4180, 4203, 4227, 4250, 4274, 4296, 4317, 4334, 3282, 3301, 3324, 3341, 3361, 3382, 3400, 3423, 3443, 3469, 3516, 3538, 3562, 3583, 3607, 3630, 3649, 3672, 3695, 3717, 3740, 3763, 3783, 3802, 4778, 4886, 4354, 4374, 4891, 4914, 4931, 4948, 4962, 4981, 5000, 5021, 5040, 5058, 5074, 5094, 5113, 5130]. Notes [5150, 5166, 5185, 5207, 5225, 5241, 5261, 5281, 5302, 5321, 5338, 4397, 4418, 4440, 4460, 4484, 4503, 4519, 4534, 4558, 4577, 4603, 4625, 4641, 4661, 4680, 4701, 4720, 4742, 4764, 4787, 4809, 4834, 4853, 4873, 4357, 4377, 4916, 4934, 4950, 4965, 4983, 5003, 5023, 5043, 5060, 5079, 5097, 5115, 5132, 5152, 5168, 5187, 5209, 5227, 5244, 5263, 5283, 5304, 5323, 5341, 4399, 4420, 4442, 4463, 4486, 4505, 4521, 4537, 4560, 4579, 4605, 4627, 4644, 4662, 4682, 4704, 4723, 4744, 4766, 4789, 4810, 4837, 4875]. Notes [4894, 5935, 5355, 5372, 5866, 5886, 5905, 5927, 5952, 5974, 5994, 6014, 6034, 6054, 6073, 6094, 6117, 6138, 6158, 6178, 6194, 6213, 6231, 6249, 6273, 6292, 6312, 6334, 6354, 5392, 5412, 5433, 5452, 5472, 5490, 5507, 5528, 5547, 5568, 5587, 5608, 5624, 5642, 5660, 5674, 5689, 5710, 5729, 5753, 5773, 5794, 5821, 5842, 5394, 5435, 6017, 6036, 6055, 6075, 6097, 6119, 6139, 6160, 6180, 6197, 6216, 6233, 6252, 6274, 6337, 6358, 5474, 5509, 5531, 5571, 5589, 5610, 5627, 5644, 5662, 5675, 5692, 5712, 5732]. Notes [5754, 5775, 5796, 5822, 5846, 5867, 5907, 5931, 5955, 5977, 5996, 6373, 6393, 6893, 6910, 6929, 6964, 6984, 7004, 7019, 7035, 7055, 7073, 7091, 7111, 7127, 7144, 7159, 7175, 7194, 7213, 7228, 7248, 7266, 7287, 7307, 7326, 7341, 7353, 6414, 6433, 6454, 6471, 6492, 6512, 6531, 6551, 6573, 6596, 6617, 6641, 6661, 6682, 6702, 6721, 6740, 6756, 6775, 6798, 6816, 6841, 6861, 6875, 6435, 6455, 6987, 7005, 7021, 7038, 7057, 7074, 7094, 7113, 7129, 7146, 7178, 7195, 7216, 7231, 7250, 7269, 7288, 7327, 7342]. Notes [7356, 6473, 6495, 6513, 6534, 6554, 6576, 6599, 6620, 6644, 6665, 6685, 6724, 6743, 6761, 6779, 6801, 6819, 6845, 6862, 6878, 6896, 6912, 6931, 6968, 7368, 7392, 7939, 7965, 7987, 8009, 8030, 8053, 8080, 8099, 8119, 8139, 8162, 8186, 8209, 8230, 8250, 8271, 8288, 8311, 8332, 8350, 8374, 8393, 8414, 8434, 8459, 8483, 7404, 7424, 7442, 7461, 7484, 7505, 7526, 7545, 7564, 7584, 7607, 7628, 7651, 7677, 7698, 7728, 7748, 7774, 7793, 7821, 7844, 7869, 7892, 7915, 7501, 7372, 7393, 7988, 8010, 8031, 8057, 8081]. Notes [8100, 8120, 8143, 8163, 8187, 8210, 8231, 8253, 8272, 8290, 8312, 8333, 8352, 8375, 8395, 8415, 8438, 8460, 8485, 7408, 7425, 7444, 7462, 7485, 7529, 7567, 7585, 7610, 7630, 7654, 7680, 7703, 7729, 7749, 7775, 7796, 7822, 7847, 7870, 7894, 7917, 7940, 7966, 8504, 8527, 9086, 9106, 9145, 9165, 9185, 9204, 9223, 9239, 9257, 9277, 9300, 9322, 9343, 9363, 9380, 9399, 9422, 9441, 9462, 9481, 9503, 9524, 9544, 9565, 8548, 8572, 8590, 8622, 8644, 8667, 8689, 8709, 8732, 8754, 8777, 8820, 8839, 8863, 8886]. Notes [8903, 8923, 8947, 8967, 8985, 9005, 9026, 9042, 9066, 9580, 9599, 10226, 10248, 10269, 10292, 10319, 10340, 10362, 10384, 10406, 10435, 10459, 10483, 10503, 10530, 10550, 10580, 10605, 10631, 10658, 10679, 10705, 10732, 10755, 10782, 10803, 9621, 9648, 9671, 9697, 9717, 9740, 9765, 9787, 9812, 9837, 9860, 9892, 9919, 9945, 9971, 9999, 10034, 10062, 10086, 10111, 10133, 10154, 10174, 10200, 9581, 9601, 10249, 10271, 10293, 10320, 10342, 10363, 10387, 10409, 10437, 10460, 10485, 10505, 10533, 10555, 10582, 10607, 10633, 10659, 10683, 10707, 10734, 10756, 10785]. Notes [10806, 9623, 9649, 9672, 9698, 9718, 9741, 9768, 9813, 9839, 9863, 9893, 9921, 9946, 9972, 10001, 10036, 10063, 10088, 10114, 10135, 10157, 10177, 10202, 10227, 10436, 10327, 10418, 10514, 3076, 7077, 7098, 3353, 4530, 1088]. Notice [8878]. Nova [8302]. Novel [578, 782, 761, 1062, 956, 738, 735, 957, 2935, 3024, 2346, 2630, 2645, 2779, 4161, 3229, 3455, 3263, 3308, 4065, 3888, 3311, 3848, 4696, 5466, 6241, 5513, 5743, 7100, 7337, 7042, 6808, 6608, 6587, 6690, 6771, 7199, 6807, 7079, 7207, 8173, 7720, 8043, 8281, 7784, 7416, 8472, 7363, 7950, 7453, 7859, 7595, 7431, 9175, 8934, 8973, 9138, 8766, 9333, 9072, 8978, 8608, 9266, 8993, 9036, 8959, 9376, 9013, 8974, 9196, 9472, 9230, 9055, 9433, 9929, 9985, 9588, 10281, 10428, 10377, 9881, 9778, 10479, 10543, 10076, 10012]. Novelties [9689]. Novelty [90, 4306, 10777]. November [1054, 2164, 3070, 5190, 6203, 9469, 8301, 3194, 10597, 9388, 10588]. Nozzle [3198, 5868, 8697]. Nubian [3753]. Nuisance [909]. Number [6971, 6914, 6951, 7542]. Numbering [3551]. Numbers [1283]. Numerical [7541]. Numerous [5499]. Nut [714, 524, 1206, 320, 1797, 2639, 3004, 4277, 4968, 5537, 7223, 8107, 899]. Nutley [7888]. Nutmeg [877]. Nutritive [9774]. Nuts [8703].

O [9854, 8194]. O. [5194]. Oarlock [5561]. Obesity [8014]. Objective [10392]. Objects [537, 1170, 6147, 7163, 1572, 1814]. Observations [6521, 10470]. Observatory [5445, 7275, 6622, 1670, 4285, 4216, 5124, 5099, 7850, 7631, 8737, 9443, 10392, 9734, 7788, 6717, 5069, 2044]. Observe [6622]. Observed [6837, 8726]. Obsolete [9383]. Obstacles [8970, 9282]. Obtaining [7767]. Ocean [1237, 3383, 3224, 4951, 4387, 6325, 6315, 10233, 6009]. Oceanic [10478, 10466, 9688]. October [4117, 931, 2072, 3036, 3008, 7132, 8220, 9307, 10493]. Octopus [10448]. Odd [4339, 2438]. Odometer [5834]. Off [1080, 12, 1513, 1693, 2510, 4158, 7219, 3229, 6241, 7122]. Office [87, 337, 1413, 2531, 3821, 4319, 3612, 3953, 3862, 4866, 5264, 6864, 6666, 8396, 7876, 9560, 1698, 5971, 1699, 2736, 2746, 2522, 3860, 3687, 4025, 3585, 5169, 5343, 6555, 8398, 7919, 8356, 8318, 8782, 8801, 9029, 8574, 9625, 10068, 10639, 10511]. Officers [7634]. Offices [9547, 8582]. Official [2033, 1250, 8087, 8799, 9253, 9948, 6143]. Ogden [4993]. Ohio [9286, 3831, 5600, 5733]. Oil [579, 1235, 1785, 2268, 1446, 2966, 2562, 3053, 3452, 3794, 3784, 5005, 6144, 6504, 7260, 6879, 6519, 8384, 8063, 7400, 9137, 10626, 416, 9586]. Oil-Retaining [9137, 10626]. Oiling [5044]. Oils [133]. Oklahoma [3828]. Okra [103]. Old [852, 2953, 2484, 2375, 3742, 4284, 3462, 5119, 6164, 6920, 8418, 10078, 9986, 961, 9687]. Oldest [3479]. Olds [4796, 7261]. Olive [1446, 10141, 8975]. Olsen [4062]. Olympia [4323, 5611, 6613, 6735, 6577]. Olympian [6784, 6558, 6381, 9101]. Olympic [6754]. Omaha [9356, 8962]. Omelette [10548]. Omnibus [6344]. One [1083, 1851, 3069, 3632, 4444, 4745, 6679, 8063, 7942, 7426, 10453, 8446, 2081, 2099, 2112]. One-Rater [10453]. O’Neil [9913]. Onions [8171]. Ontario [9317]. Onton [55]. Opaque [6147, 1572]. Open [1765, 1713, 10103]. Opened [7834]. Opener [445, 808, 480]. Opening [288, 1198, 1233, 1964, 2077, 3587, 3608, 4814, 4827, 5695, 6099, 5839, 5814, 8405, 7518, 7856, 8260, 9269, 10757, 10664, 9910]. Openings [974]. Opera [3553, 7815, 7694, 10773]. Operated [1207, 1275, 2108, 1419, 4058, 4085, 5823, 5556, 7256, 10449, 9682]. Operates [7789]. Operating [2288, 3313, 3728, 4748, 4383, 6362, 9727, 10250]. Operation [7218, 6946, 7383]. Operations [3089, 4182, 3585]. Opinion [8926]. Opium [8694]. Opossum [163]. Opportunities [2949, 9167]. Opportunity [10365]. Optical [68, 3011, 4189, 4378, 7203, 6563, 10693, 1238, 1281, 1572]. Optics [3961]. Optometer [5083, 5543]. Orang [4781]. Orang-Outang [4781]. Orange [4990, 6834]. Oranges [8765]. Orchard [7130]. Orchards [7598]. Orchestral [8112]. Orchid [944]. Order [1786, 8356]. Ordnance [8477, 10203, 9554, 9719, 9527, 9384, 10634, 10688]. Ore [3701, 9638, 527, 1795, 1733, 2469, 2741, 2396, 4749, 5379, 5361]. Ore-Stamping [9638]. Oregon [5376, 5755, 6764, 7618, 8907, 9208, 8528, 8878, 7289]. Ores [1836, 4695, 6042, 9122]. Organization [2033]. Organs [4141, 10653]. Oriental [1541]. Origin [7636, 8192, 7811]. Original [1691, 1561, 1725, 4423, 10253]. Orion [10019]. Orleans [2965, 8811, 8939, 10611]. Ornamental [816, 585]. Ornaments [747, 673, 610]. Orton [10613, 10356]. Oscillating [2852, 4823]. Ostrich [2277, 10800]. Other [811, 2226, 2801, 5031, 8891, 494]. Otis [4120, 4136, 6129]. Otocyon [2958]. Ottawa [4454]. Otto [2213, 4181]. Our [131, 26, 953, 930, 1513, 2290, 1689, 1619, 2034, 3212, 2754, 4266, 7076, 6726, 6971, 6914, 6781, 6821, 7180, 6951, 6601, 7990, 8084, 8164, 8296, 8193, 7705, 8170, 7468, 8212, 8627, 9505, 8906, 9069, 9159, 9088, 9051, 9382, 9373, 9092, 9325, 9027, 8650, 9347, 8758, 8853, 9281, 8997, 8757, 8911, 8845, 8940, 8712, 9028, 8950, 10352, 9701, 10065, 10326, 10641, 9775, 10765, 10439, 9674, 10323, 9654, 10238, 9899, 10648, 9898, 10004, 10344, 10737, 10442, 10767, 9951, 8496, 7735, 9586, 1076, 1709, 2602, 4422, 8334, 9246, 9148, 9348, 9457]. our [9244, 10427, 10389, 10143, 10139, 10187, 9070]. Outang [4781]. Outfit [1851, 2312, 5576]. Outing [5704]. Outlet [2057]. Outlook [8651]. Oven [7977]. Over-Supply [6416]. Overhead [3289, 4562]. Overhead-Propelled [3289]. Overland [10278]. Overshoe [2341]. Owners [9606]. Oxen [5724]. Oxide [2579, 4349, 8893]. Oxygen [2801, 6652]. Oyster [3372, 2078]. Ozotype [9979].

P [2421, 5305, 5577, 8616]. P. [5194]. PA [3088, 568, 4803, 5295]. Pacific [2771, 9264, 7414, 9939]. Pacification [8866]. Package [1464, 8849, 9757]. Packing [785, 40, 1600, 1748, 1908, 5120, 6061, 6649]. Packings [2090]. Pad [8105]. Paddle [8278]. Padlock [500]. Padua [8663]. Pagoda [5446]. Pail [1258]. Paint [2669, 7873]. Painting [3563, 5238, 10216]. Paintings [10095]. Paints [7616]. Pair [8188]. Pairs [1177]. Palace [3824, 3917, 6750, 7487, 10561, 3944, 4109]. Palacio [1016]. Palæontological [1052]. Palisades [8237]. Palladio [9101]. Palm [6872]. Palms [9613]. Pampas [5369]. Pan [10093]. Pan-American [10093]. Panama [686, 173, 9769, 9744, 9651, 9665, 9699, 9675]. Panoramic [5965]. Pans [624, 8994]. Paper [1081, 531, 37, 299, 3285, 4689, 4732, 6124, 6623, 9529, 8891, 8748, 8865, 9357, 8843, 9048, 9799, 10128, 10515]. Paper-Punch [9799]. Papers [2133, 2018, 10123, 5971]. Papier [7398]. Papua [7243]. Papyri [7853]. Parachute [2917]. Parade [3584]. Paradise [4350, 5924]. Paradox [2570]. Paragon [4208, 4184]. Parallel [1277, 5358, 6786]. Paralysis [9046]. Paralyzed [1700]. Parasite [5499]. Parasites [5070, 8266, 10663]. Parasol [712, 789]. Parcels [10417]. Parcels-Post [10417]. Paris [9678, 478, 115, 7, 495, 143, 29, 3371, 3768, 4279, 5400, 6162, 5705, 6484, 7136, 6899, 7862, 8124, 7761, 7754, 9007, 9443, 10714, 10006, 10219, 10251, 9833, 10624, 9573, 9763, 10283, 10694, 10773, 9885]. Parisian [9953]. Park [1545, 1878, 2377, 4467, 6612, 7039, 2835, 5088, 4673, 4705, 5246, 5837, 6461, 8256, 8940, 10322, 10664, 9927]. Parks [10641]. Parkway [6009]. Parlor [1396, 5387]. Parrots [2756]. Parseval [7933]. Parson [3793]. Parsons [10675]. Part [8179, 10165]. Parthenon [8039]. Partial [5404]. Parties [7977]. Partition [1567, 8337]. Parts [6897]. Pass [466, 9910]. Passage [6687, 10255]. Passed [8802]. Passenger [345, 552, 2170, 1595, 4162, 4815, 6102, 7027, 7872, 7758, 2208, 3097]. Passes [8172]. Passing [9425, 10207]. Past [6953, 6946]. Pasteboard [6124]. Pasteur [6140, 635, 7456]. Patagonia [6112, 10674]. Patent [1169, 91, 839, 1197, 569, 24, 50, 969, 987, 1008, 1029, 1048, 1074, 1097, 1117, 1144, 1168, 1196, 1225, 1247, 74, 98, 128, 150, 243, 385, 413, 440, 464, 493, 517, 543, 567, 591, 616, 728, 775, 803, 838, 859, 884, 907, 928, 951, 337, 289, 1009, 1322, 1321, 1539, 1571, 1588, 1316, 1340, 1925, 1935, 1958, 1977, 1993, 2013, 2030, 2051, 2069, 2087, 2106, 2127, 2146, 2179, 2195, 2219, 2240, 2264, 2284, 1364, 1389, 1409, 1432, 1434, 1455, 1483, 1509, 1534, 1558]. Patent [1583, 1614, 1687, 1706, 1723, 1743, 1760, 1781, 1804, 1824, 1843, 1864, 1885, 1906, 1413, 1587, 2882, 2485, 2842, 2861, 2909, 3067, 3087, 2375, 3585, 3242, 3259, 3819, 3838, 3858, 3881, 3904, 3928, 3952, 3972, 3995, 4023, 4048, 4072, 4094, 4116, 4135, 4154, 4180, 4203, 4227, 4250, 4274, 4296, 4317, 4334, 3282, 3301, 3324, 3341, 3361, 3382, 3400, 3423, 3443, 3469, 3516, 3538, 3562, 3583, 3607, 3630, 3649, 3672, 3695, 3717, 3740, 3763, 3783, 3802, 3953, 3447, 4634, 5268, 4337, 4357, 4377, 4916]. Patent [4934, 4950, 4965, 4983, 5003, 5023, 5043, 5060, 5079, 5097, 5115, 5132, 5152, 5168, 5187, 5209, 5227, 5244, 5263, 5283, 5304, 5323, 5341, 4399, 4420, 4442, 4463, 4486, 4505, 4521, 4537, 4560, 4579, 4605, 4627, 4644, 4662, 4682, 4704, 4723, 4744, 4766, 4789, 4810, 4837, 4875, 4894, 5532, 6283, 5582, 5394, 5435, 6017, 6036, 6055, 6075, 6097, 6119, 6139, 6160, 6180, 6197, 6216, 6233, 6252, 6274, 6337, 6358, 5474, 5509, 5531, 5571, 5589, 5610, 5627, 5644, 5662, 5675, 5692, 5712, 5732, 5754]. Patent [5775, 5796, 5822, 5846, 5867, 5907, 5931, 5955, 5977, 5996, 5590, 5396, 5963, 7310, 7271, 6827, 6746, 7059, 6435, 6455, 6987, 7005, 7021, 7038, 7057, 7074, 7094, 7113, 7129, 7146, 7178, 7195, 7216, 7231, 7250, 7269, 7288, 7327, 7342, 7356, 6473, 6495, 6513, 6534, 6554, 6576, 6599, 6620, 6644, 6665, 6685, 6724, 6743, 6761, 6779, 6801, 6819, 6845, 6862, 6878, 6896, 6912, 6931, 6968, 6667, 6822, 6933, 6990, 7067, 8034, 7571, 7572, 7780, 7899, 7486, 8293, 8416, 8083, 7531, 7374, 7569, 8234, 7491, 8087]. Patent [7574, 8211, 8212, 7372, 7393, 7988, 8010, 8031, 8057, 8081, 8100, 8120, 8143, 8163, 8187, 8210, 8231, 8253, 8272, 8290, 8312, 8333, 8352, 8375, 8395, 8415, 8438, 8460, 8485, 7408, 7425, 7444, 7462, 7485, 7529, 7567, 7585, 7610, 7630, 7654, 7680, 7703, 7729, 7749, 7775, 7796, 7822, 7847, 7870, 7894, 7917, 7940, 7966, 7446, 8317, 7513, 7553, 8257, 7611, 7777, 8653, 8843, 9048, 9382, 8595, 9387, 10017, 9792, 9603, 10205, 10212, 9581, 9601, 10249, 10271, 10293, 10320, 10342, 10363, 10387, 10409, 10437, 10460]. Patent [10485, 10505, 10533, 10555, 10582, 10607, 10633, 10659, 10683, 10707, 10734, 10756, 10785, 10806, 9623, 9649, 9672, 9698, 9718, 9741, 9768, 9813, 9839, 9863, 9893, 9921, 9946, 9972, 10001, 10036, 10063, 10088, 10114, 10135, 10157, 10177, 10202, 10227, 10511, 3115, 3601, 415, 1909, 1699, 1621, 2720, 2736, 2746, 2522, 3363, 3821, 3326, 3860, 4025, 3862, 3425, 5169, 5343, 6555, 8398, 7919, 8318, 8782, 8801, 9448, 8800, 9029, 8574, 8805, 9625, 10068]. Patentability [3596]. Patented [45, 70, 656, 682, 701, 725, 749, 770, 794, 823, 855, 883, 905, 926, 948, 964, 983, 1006, 1027, 1044, 1070, 1092, 1112, 1140, 1163, 1190, 1221, 1244, 94, 124, 146, 168, 189, 216, 240, 264, 284, 306, 330, 354, 380, 409, 437, 460, 487, 506, 508, 540, 564, 587, 612, 632, 1267, 1287, 1861, 1880, 1900, 1921, 1936, 1953, 1973, 1989, 2008, 2025, 2047, 2065, 2083, 2101, 2122, 2141, 2155, 2174, 2190, 2214, 2235, 2259, 2280, 2299, 1307, 1333, 1357, 1381]. Patented [1402, 1426, 1448, 1479, 1502, 1528, 1551, 1575, 1607, 1631, 1655, 1683, 1702, 1718, 1739, 1756, 1777, 1799, 1820, 1838, 2316, 2330, 2785, 2803, 2820, 2837, 2856, 2876, 2893, 2922, 2944, 2960, 2980, 2996, 3015, 3032, 3046, 3062, 3081, 3099, 3116, 3137, 3152, 3170, 3187, 3205, 3218, 2348, 2369, 2386, 2408, 2426, 2440, 2460, 2479, 2497, 2516, 2537, 2555, 2575, 2595, 2613, 2632, 2649, 2662, 2679, 2698, 2715, 2732, 2750, 2766, 3238, 3256, 3798, 3815, 3853, 3876, 3899, 3923, 3947, 3967, 3990, 4019]. Patented [4044, 4068, 4089, 4111, 4130, 4149, 4175, 4197, 4222, 4246, 4269, 4292, 4312, 4330, 3277, 3296, 3319, 3337, 3357, 3378, 3396, 3419, 3438, 3463, 3489, 3510, 3533, 3557, 3578, 3603, 3626, 3643, 3667, 3690, 3712, 3736, 3760, 3779, 4353, 4371, 4892, 4910, 4930, 4945, 4963, 4977, 5001, 5017, 5041, 5055, 5075, 5093, 5109, 5127, 5145, 5162, 5180, 5203, 5221, 5242, 5258, 5277, 5298, 5317, 5334, 4393, 4415, 4438, 4458, 4481, 4500, 4516, 4533, 4555, 4574, 4600, 4622, 4640, 4660, 4677, 4700, 4719]. Patented [4739, 4760, 4784, 4805, 4833, 4852, 4870, 5352, 5373, 5862, 5882, 5901, 5928, 5951, 5975, 5991, 6015, 6030, 6051, 6069, 6095, 6113, 6135, 6155, 6174, 6195, 6212, 6228, 6250, 6269, 6288, 6308, 6333, 6355, 5388, 5409, 5429, 5451, 5469, 5487, 5504, 5525, 5548, 5569, 5583, 5605, 5623, 5639, 5658, 5672, 5688, 5706, 5728, 5749, 5769, 5790, 5818, 5843, 6374, 6392, 6894, 6909, 6926, 6981, 7001, 7016, 7034, 7052, 7070, 7092, 7107, 7123, 7140, 7158, 7174, 7192, 7212, 7229, 7245, 7267, 7284, 7306]. Patented [7322, 7338, 7352, 6410, 6429, 6450, 6467, 6491, 6508, 6530, 6550, 6572, 6595, 6616, 6642, 6660, 6683, 6703, 6722, 6739, 6757, 6776, 6797, 6817, 6842, 6858, 6876, 7369, 7388, 7936, 7960, 7983, 8005, 8025, 8054, 8076, 8094, 8114, 8140, 8157, 8181, 8205, 8225, 8251, 8267, 8289, 8307, 8328, 8351, 8371, 8394, 8410, 8433, 8454, 7405, 8479, 7419, 7438, 7457, 7480, 7504, 7525, 7546, 7565, 7581, 7604, 7625, 7652, 7678, 7699, 7723, 7745, 7771, 7794, 7817, 7845, 7864, 7889, 7910, 8500, 8522, 9067, 9081]. Patented [9102, 9125, 9146, 9166, 9186, 9200, 9224, 9238, 9258, 9273, 9295, 9318, 9344, 9360, 9377, 9397, 9423, 9442, 9458, 9479, 9501, 9525, 9543, 9562, 8547, 8567, 8591, 8621, 8643, 8666, 8686, 8704, 8728, 8755, 8776, 8797, 8818, 8838, 8862, 8881, 8904, 8920, 8944, 8963, 8986, 9000, 9022, 9043, 9577, 9596, 10223, 10244, 10266, 10287, 10315, 10341, 10357, 10385, 10405, 10431, 10455, 10484, 10502, 10526, 10551, 10576, 10602, 10628, 10654, 10680, 10706, 10733, 10751, 10783, 10804, 9617, 9643, 9668, 9692, 9713, 9735, 9766]. Patented [9786, 9809, 9836, 9857, 9888, 9920, 9940, 9966, 10000, 10031, 10058, 10087, 10112, 10134, 10155, 10171, 10197]. Patentees [8357]. Patents [271, 594, 1064, 832, 618, 334, 24, 50, 969, 987, 1008, 1029, 1048, 1074, 1097, 1117, 1144, 1168, 1196, 1225, 1247, 74, 98, 128, 150, 243, 385, 413, 440, 464, 493, 517, 543, 567, 591, 616, 728, 775, 803, 838, 859, 884, 907, 928, 951, 1669, 1316, 1340, 1925, 1935, 1958, 1977, 1993, 2013, 2030, 2051, 2069, 2087, 2106, 2127, 2146, 2179, 2195, 2219, 2240, 2264, 2284, 1364, 1389, 1409, 1432, 1434, 1455, 1483, 1509, 1534, 1558, 1583, 1614, 1687, 1706, 1723, 1743]. Patents [1760, 1781, 1804, 1824, 1843, 1864, 1885, 1906, 1315, 1636, 1661, 2160, 2589, 2373, 2601, 2842, 2861, 2909, 3067, 3087, 2978, 3445, 4096, 3821, 3242, 3259, 3819, 3838, 3858, 3881, 3904, 3928, 3952, 3972, 3995, 4023, 4048, 4072, 4094, 4116, 4135, 4154, 4180, 4203, 4227, 4250, 4274, 4296, 4317, 4334, 3282, 3301, 3324, 3341, 3361, 3382, 3400, 3423, 3443, 3469, 3516, 3538, 3562, 3583, 3607, 3630, 3649, 3672, 3695, 3717, 3740, 3763, 3783, 3802, 4228, 4629, 4728, 4995, 5062, 5140, 4443, 5061, 4357, 4377]. Patents [4916, 4934, 4950, 4965, 4983, 5003, 5023, 5043, 5060, 5079, 5097, 5115, 5132, 5152, 5168, 5187, 5209, 5227, 5244, 5263, 5283, 5304, 5323, 5341, 4399, 4420, 4442, 4463, 4486, 4505, 4521, 4537, 4560, 4579, 4605, 4627, 4644, 4662, 4682, 4704, 4723, 4744, 4766, 4789, 4810, 4837, 4875, 4894, 5746, 5344, 5447, 5522, 5826, 5938, 6002, 6090, 5394, 5435, 6017, 6036, 6055, 6075, 6097, 6119, 6139, 6160, 6180, 6197, 6216, 6233, 6252, 6274, 6337, 6358, 5474, 5509, 5531, 5571, 5589, 5610, 5627, 5644]. Patents [5662, 5675, 5692, 5712, 5732, 5754, 5775, 5796, 5822, 5846, 5867, 5907, 5931, 5955, 5977, 5996, 7272, 6960, 6911, 6435, 6455, 6987, 7005, 7021, 7038, 7057, 7074, 7094, 7113, 7129, 7146, 7178, 7195, 7216, 7231, 7250, 7269, 7288, 7327, 7342, 7356, 6473, 6495, 6513, 6534, 6554, 6576, 6599, 6620, 6644, 6665, 6685, 6724, 6743, 6761, 6779, 6801, 6819, 6845, 6862, 6878, 6896, 6912, 6931, 6968, 7234, 7573, 7946, 7372, 7393, 7988, 8010, 8031, 8057, 8081, 8100, 8120, 8143, 8163, 8187, 8210, 8231, 8253]. Patents [8272, 8290, 8312, 8333, 8375, 8395, 8415, 8438, 8460, 8485, 7408, 7425, 7444, 7462, 7485, 7529, 7567, 7585, 7610, 7630, 7654, 7680, 7703, 7729, 7749, 7775, 7796, 7822, 7847, 7870, 7894, 7917, 7940, 7966, 7634, 8418, 8508, 8649, 9405, 9006, 10300, 10740, 9581, 9601, 10249, 10271, 10293, 10320, 10342, 10363, 10387, 10409, 10437, 10460, 10485, 10505, 10533, 10555, 10582, 10607, 10633, 10659, 10683, 10707, 10734, 10756, 10785, 10806, 9623, 9649, 9672, 9698, 9718, 9741, 9768, 9813, 9839, 9863, 9893, 9921, 9946, 9972, 10001]. Patents [10036, 10063, 10088, 10114, 10135, 10157, 10177, 10202, 10227, 1933, 7277, 174, 1887, 2413, 3343, 4251, 4606, 5493, 7808, 7633, 8573, 8615, 8762, 9770, 10272, 10276]. Paterson [2809]. Path [2904, 6009, 10536, 10477]. Pathway [6484]. Patience [6847]. Patrick [980, 9539]. Patriotic [9301]. Patrol [2502, 4847]. Patterns [9988]. Patterson [4509]. Paul [4951, 5637, 6264, 6913, 6459]. Pavement [6829, 9691]. Pavements [9553]. Paving [2204, 8221]. Pawnbroking [10346]. Payment [519]. Pays [707]. Pea [10671]. Peace [7733, 8735, 9898]. Peach [7303]. Peak [1318]. Peanut [5330]. Pear [7955]. Pearl [3567, 10310]. Pearl-Button [10310]. Peary [3813, 6142, 8443, 8470, 7833, 8254, 10469]. Pease [6418]. Peat [10512, 9569, 10515]. Pecos [3333]. Peculiar [7838, 10349]. Pedal [3287, 5195, 10012]. Pedaling [10793]. Pedals [6364]. Pedestrian [5916]. Pekin [7806]. Pelican [8303]. Pelton [2115]. Pen [62, 85, 1299, 4005, 4008, 7118, 9271]. Pencil [348, 9158]. Pendent [1809]. Pending [6536]. Penduline [2689]. Pendulum [3412]. Pendulums [6382, 4991]. Penmen [713]. Penn [3077]. Pennington [1471]. Pennsylvania [581, 1400, 1251, 3504, 4299, 5415, 6098, 7519, 10382, 10078, 9879]. People [6379, 6932, 7345]. Peoples [5802]. Peral [9058]. Perambulator [6106]. Percentage [9190]. Percha [1959, 7661]. Percolator [9757]. Père [8091]. Perfected [7331]. Perfecting [6170]. Perforating [367, 9174]. Performance [7290, 8013]. Performing [3312]. Perfume [3107, 6975]. Perfumed [5964]. Perfumes [4300, 4944]. Peril [9526, 10138]. Periodical [7031, 4829]. Perishable [1150]. Peristyle [3973]. Permanence [6236]. Permanent [1899, 8194]. Permeability [6813]. Permutation [362]. Perpetual [2844, 10194]. Perry [194]. Perseverance [6847]. Persistence [3275]. Personal [1043, 1068, 1093, 20, 46, 680, 702, 723, 748, 769, 792, 822, 854, 881, 903, 923, 946, 963, 982, 1004, 1023, 1113, 1139, 1162, 1189, 1218, 1242, 69, 93, 123, 145, 167, 214, 238, 263, 283, 305, 329, 352, 381, 408, 434, 459, 486, 509, 538, 562, 586, 611, 631, 653, 1549, 1266, 1286, 1879, 1901, 1919, 1934, 1952, 1972, 1988, 2009, 2026, 2046, 2064, 2082, 2102, 2123, 2142, 2154, 2175, 2191, 2215, 2234, 2260, 2279, 2298, 1309, 1332, 1358, 1380, 1403, 1425, 1477, 1503]. Personal [1529, 1573, 1605, 1629, 1656, 1682, 1701, 1720, 1737, 1757, 1776, 1798, 1819, 1839, 1859, 2367, 2556, 2875, 2315, 2329, 2822, 2838, 2857, 2892, 2924, 2943, 2959, 2981, 2997, 3016, 3031, 3047, 3063, 3083, 3101, 3118, 3136, 3153, 3171, 3189, 3207, 3220, 2349, 2388, 2407, 2427, 2442, 2461, 2480, 2498, 2518, 2536, 2574, 2596, 2614, 2633, 2650, 2664, 2680, 2699, 2716, 2731, 2751, 2765, 2787, 2804, 3237, 3257, 3797, 3816, 3834, 3854, 3877, 3900, 3924, 3949, 3966, 3989, 4018, 4043, 4067, 4090, 4113, 4131, 4150, 4176]. Personal [4199, 4223, 4245, 4270, 4291, 4313, 4329, 3278, 3297, 3320, 3338, 3356, 3377, 3397, 3420, 3439, 3465, 3488, 3512, 3534, 3558, 3579, 3602, 3625, 3644, 3668, 3691, 3713, 3735, 3759, 3780, 4352, 4373, 4890, 4912, 4929, 4946, 4960, 4979, 4998, 5019, 5038, 5057, 5073, 5092, 5111, 5128, 5147, 5163, 5182, 5205, 5223, 5239, 5259, 5279, 5300, 5319, 5336, 4395, 4416, 4439, 4459, 4482, 4502, 4517, 4532, 4556, 4575, 4601, 4623, 4639, 4659, 4678, 4699, 4718, 4740, 4762, 4785, 4806, 4832, 4851, 4871, 5353, 5371, 5863, 5884]. Personal [5903, 5926, 5950, 5972, 5992, 6013, 6032, 6052, 6071, 6092, 6115, 6136, 6157, 6176, 6192, 6211, 6230, 6248, 6270, 6289, 6310, 6332, 6352, 5390, 5411, 5431, 5450, 5471, 5488, 5505, 5527, 5546, 5567, 5606, 5622, 5640, 5659, 5671, 5687, 5708, 5727, 5750, 5771, 5792, 5819, 5841, 6372, 6391, 6892, 6908, 6927, 6963, 6983, 7002, 7017, 7033, 7053, 7071, 7090, 7109, 7125, 7142, 7157, 7173, 7191, 7211, 7226, 7246, 7264, 7283, 7305, 7323, 7339, 7351, 6411, 6431, 6451, 6469, 6490, 6509, 6529, 6549, 6571, 6594, 6615]. Personal [6640, 6662, 6680, 6701, 6720, 6738, 6755, 6774, 6796, 6815, 6840, 6859, 6873, 7367, 7389, 7937, 7962, 7985, 8007, 8027, 8052, 8075, 8096, 8116, 8138, 8159, 8183, 8207, 8227, 8248, 8269, 8286, 8308, 8329, 8347, 8370, 8390, 8412, 8432, 8456, 8481, 7403, 7421, 7440, 7458, 7481, 7503, 7524, 7544, 7563, 7582, 7606, 7626, 7650, 7676, 7697, 7725, 7746, 7772, 7792, 7818, 7843, 7866, 7891, 7912, 8501, 8524, 9064, 9083, 9103, 9126, 9144, 9164, 9183, 9202, 9221, 9237, 9256, 9275, 9297, 9320, 9341, 9362, 9379, 9398, 9420]. Personal [9439, 9460, 9480, 9502, 9522, 9542, 9563, 8546, 8569, 8587, 8620, 8642, 8665, 8687, 8706, 8730, 8751, 8775, 8796, 8816, 8837, 8861, 8883, 8901, 8921, 8942, 8965, 8983, 9002, 9023, 9040, 9578, 9597, 10199, 10224, 10245, 10267, 10289, 10316, 10338, 10359, 10383, 10404, 10433, 10457, 10481, 10501, 10528, 10549, 10578, 10603, 10629, 10656, 10678, 10702, 10730, 10753, 10781, 9619, 10801, 9645, 9669, 9694, 9715, 9737, 9764, 9785, 9810, 9835, 9859, 9889, 9916, 9942, 9968, 9997, 10033, 10059, 10084, 10109, 10131, 10152, 10173, 5585]. Persons [6347]. Pertinent [10179]. Peru [6865, 7788]. Peryar [5987]. Pestle [280]. Pests [2920]. Petch [3751]. Peter [9608]. Peters [873]. Petersburg [6916, 7806]. Petitions [2485]. Petroleum [1370, 1391, 2456, 3677, 5192, 5645, 5764, 6153, 8243, 7979, 8604, 10160, 9667]. Ph.D [8194]. Phenomena [18, 6972, 9329, 7066, 8792, 9334, 9358, 9374, 9390, 9419]. Phenomenal [6557, 9251]. Philadelphia [3088, 5295, 637, 1771, 3092, 3077, 4803, 4451, 5501, 7856, 7841, 9315, 10494, 10570]. Philanthropy [7311, 7539]. Philbrook [5497]. Philip [9528]. Philippine [2712, 8888, 9488, 9721, 9655, 10096, 8972]. Philippines [9247, 10091, 9631, 9270, 9020, 9037]. Philosophy [7365, 10126]. Phonendoscope [7540]. Phonograph [7276, 6943, 7645, 236, 375, 722, 787]. Phonographic [401, 3636]. Phonographs [10331]. Phonostethescope [7540]. Phoro [5543]. Phosphate [3426]. Phosphorescent [8041]. Photo [1690, 2887, 3553]. Photo-Engravings [1690]. Photo-Opera [3553]. Photograph [3663, 4266, 5788, 5404, 6588, 6998]. Photographer [842]. Photographers [7013]. Photographic [54, 76, 161, 197, 399, 521, 580, 678, 780, 820, 876, 893, 922, 1021, 1042, 1133, 1851, 1255, 1260, 1374, 1470, 1769, 1896, 1910, 2005, 2256, 3110, 2550, 2710, 2781, 3892, 3274, 5888, 6143, 5978, 5427, 9326, 9391, 10666, 10563, 10185, 10510]. Photographing [4779, 10727]. Photographs [1061, 2432, 2713, 2855, 7551, 9650]. Photography [1304, 2985, 2494, 3090, 3213, 3436, 3930, 3698, 3459, 4298, 4513, 4551, 4572, 5502, 5989, 5962, 6941, 7169, 6589, 6483, 6548, 7673, 7831, 9679, 10105, 10095, 10121, 10429, 5069, 2233, 9616]. Photometer [2936]. Photomicrography [4759]. Photoret [4325]. Physical [3105, 3213, 4769, 2044]. Physician [10188]. Physicians [8149]. Physics [5934, 6056, 6937, 7637, 10724]. Physiology [9530, 10054]. Piano [418, 2356, 5198, 6343, 5715, 5828, 8514]. Pianos [1465, 3502, 4141]. Pick [784]. Picker [605, 3093, 3709]. Picking [663]. Picks [9801]. Picolet [4752]. Picture [2181, 1623, 2921, 3042, 7685]. Pictures [279, 8447]. Picturesque [9337]. Piece [807, 5271, 5630, 8132, 9251]. Pier [4564, 9290, 1846]. Pierce [8473]. Pierced [7088]. Piers [4984, 5946, 10598]. Pig [1624]. Pigeons [1147, 8584]. Pike [2344, 1318, 2580]. Pile [4565]. Piles [3826, 4758, 9264]. Pillows [3730, 4160]. Pilot [33, 9365]. Pilots [9365]. Pin [846, 1943, 6168, 7644, 10402]. Pinch [3746]. Pinches [998]. Pine [1215, 4392]. Pintsch [9080]. Pioneer [7319]. Pioneers [6047]. Pipa [3013]. Pipe [470, 1176, 67, 397, 136, 349, 554, 370, 1180, 1474, 1248, 1590, 2203, 2057, 1966, 2251, 2811, 2659, 2380, 2724, 4448, 4617, 5695, 6046, 5716, 6166, 5601, 6148, 6262, 7220, 7181, 7493, 7737, 7362, 8176, 8042, 7516, 8282, 8343, 8452, 10279, 9758, 9849]. Pipe-Hanger [9849]. Pipes [1080, 2037, 1515, 3113, 2796, 3914, 5200, 5031, 5725, 5985, 8150, 8385, 9494, 10047, 10259]. Pipette [4548]. Pirates [9631]. Pistol [1129, 484, 5943, 8608, 10488]. Piston [1748]. Pit [7303]. Pitcher [978, 7909]. Pitchfork [3889]. Pitman [7380, 10049]. Pitman-Head [10049]. Pittsburg [568]. Place [4411]. Places [3027, 4339, 8151]. Placing [5683]. Plague [9133, 9406]. Plainfield [3091]. Plains [2061]. Plan [814, 3043, 6970]. Plane [638, 5615]. Planer [2435]. Planers [575, 1330, 1569]. Planet [2864, 8612]. Planets [290, 731, 1123, 104, 522, 619, 1227, 223, 389, 1054, 931, 840, 1540, 1927, 2244, 1346, 1746, 1827, 1438, 1668, 2164, 2072, 1998, 2544, 2845, 3165, 2376, 2305, 2705, 2775, 2447, 2620, 3070, 3008, 2930, 3472, 3869, 3316, 3683, 3775, 3384, 3574, 3986]. Planing [4063, 10543]. Plankton [8989]. Plans [6579]. Plant [1020, 482, 225, 1960, 2771, 3010, 2439, 4264, 4799, 4451, 4511, 4569, 5366, 5897, 5823, 5740, 6181, 6423, 6879, 8365, 7601, 7950, 8868, 10599, 10461, 10096, 10313, 10616, 6539, 10234, 2531, 3882, 6208, 6632, 8245, 8404, 8417, 10647, 10262]. Planter [2003, 3680]. Planters [8979, 9823]. Plantin [7224]. Planting [7801]. Plants [10347, 344, 1941, 2571, 4339, 7365, 8787, 8638, 9701, 10794, 7136]. Plaster [5271]. Plastering [1210]. Plastic [2702]. Plata [3126]. Plate [525, 1234, 326, 2810, 2948, 3820, 4033, 4490, 4899, 5519, 5368, 6889, 7967, 7664, 8465, 8264, 8597, 9215, 10695, 10183, 9817, 9499, 990]. Plates [953, 930, 2243, 2992, 4895, 4448, 7997, 10666, 9750, 5859]. Platform [403, 1650, 5100, 6674]. Plating [885, 5, 666, 3057]. Platinum [952, 8236]. Playing [6343, 10653]. Plaza [5088]. Pleasant [581]. Pleasure [6068, 10025]. Pleiades [5107]. Plenty [8466]. Pliers [5648]. Plimsoll [1644, 1645]. Plow [2525, 2643, 3751, 7081, 6730, 6772, 8850, 9093, 9957, 10568, 10590]. Plows [6023]. Plug [370, 5329]. Plugger [2448, 6731, 7516]. Plumb [1982]. Plumbago [40]. Plummet [9733]. Plymouth [961, 4920]. Pneumatic [310, 551, 696, 143, 908, 1497, 1877, 3193, 3709, 4278, 3334, 5219, 5704, 6317, 5407, 5387, 6006, 6698, 6921, 6585, 7453, 8260, 8240, 8425, 8473, 8424, 7976, 8015, 9169, 9072, 9013, 8675, 9779, 9912]. Pneumatically [9682]. Pocket [2694, 3705, 5309, 6604, 6481, 6771, 7642, 10100]. Pod [6677]. Point [7549, 8950, 2021]. Poises [1951]. Poison [344, 1487]. Poisoning [4205, 4051, 8826]. Poisonous [1828]. Polar [1782, 1806, 2984, 6559, 7029]. Polaris [10470]. Pole [473, 846, 4544, 6320, 8610, 1216, 5870, 8378, 10286]. Poles [6676]. Police [2502]. Policemen [6387]. Policy [9244, 7658]. Polisher [890, 2507]. Polishing [253, 671, 2967, 4686, 8722]. Politicians [10440]. Pollok [10789]. Polychrome [2855]. Polymeter [2275]. Polyphone [10331]. Polytechnic [9177]. Pompeian [852]. Pompeii [8224]. Pont [5439]. Pont-Y-Pridd [5439]. Poor [2242, 4706, 10139]. Pop [5330]. Pope [3636, 7490]. Popular [1314, 1339]. Population [1689, 7667]. Porcelain [5572, 10762]. Pork [5120]. Porous [2783]. Port [3039, 5880, 6263, 8033, 8362, 7130]. Porta [8246]. Portable [57, 319, 1295, 2134, 1960, 2135, 2224, 1396, 4301, 4322, 4745, 5219, 6283, 6104, 5879, 6083, 5666, 7082, 7083, 7474, 8243, 10076, 10592, 9707, 6539]. Portchester [628]. Portelectric [1020]. Porter [7787]. Porters [3768]. Portiere [853]. Portion [5084]. Portions [4506]. Portland [9506]. Porto [9484, 9192, 9818, 10413, 8937, 10477]. Portrait [1691, 8749]. Position [290, 731, 1123, 104, 522, 619, 1227, 223, 389, 1054, 931, 840, 1540, 1927, 2244, 1346, 1746, 1827, 1438, 1668, 2164, 2072, 1998, 1643, 2544, 2845, 3165, 2376, 2305, 2705, 2775, 2447, 2620, 3070, 3008, 3472, 3869, 3316, 3683, 3775, 3384, 3574, 3986, 7789]. Positions [2930]. Possessions [9281, 9701, 10143, 10767]. Possibilities [5757, 7587, 8464]. Possible [2844]. Post [1040, 572, 87, 975, 4006, 3846, 4319, 3687, 3612, 4544, 6666, 7199, 8977, 10417, 8136, 7876, 8356]. Postage [5285, 10324]. Postal [2432, 6524, 8260, 7759, 8675]. Postponement [6223]. Pot [2418]. Potassium [4336, 4359, 10042]. Potato [643, 1015, 3680, 3586, 6110, 6421, 8990]. Potatoes [3586]. Potentialities [4379]. Potomac [2059]. Pots [7600]. Potsdam [3233, 3268]. Potter [7776]. Pottery [4119, 4037, 4525]. Pouch [670]. Poulson [7172]. Poultry [1378, 2583]. Pounder [10312]. Pounding [900]. Powder [5698, 6698, 8175, 8539, 9210, 8988, 10794, 9584, 10513, 10067, 10098]. Power [378, 1056, 1107, 1969, 2129, 1894, 1913, 3043, 2422, 2709, 2704, 2473, 2886, 3554, 3785, 3621, 4241, 3826, 3907, 5044, 5032, 5276, 5216, 4570, 4366, 5139, 4953, 6346, 6107, 6161, 5696, 6202, 5347, 5539, 5703, 6001, 5765, 6752, 6423, 6632, 6380, 7240, 8315, 8365, 8245, 8404, 8215, 7512, 8275, 7426, 7538, 7691, 9191, 9019, 8789, 9431, 9516, 8586, 9429, 10204, 10636, 10661, 10647, 10262, 9878, 9679, 7715, 9928]. Power-Hammer [9878]. Powered [9417]. Powerful [1297, 2041, 5086, 7688, 9879, 10103]. Powers [4017]. Practical [1031, 2273, 2594, 4616, 7665, 8242, 8628, 6547]. Practice [90, 235, 790, 2918, 4862, 6047, 7162, 7271, 7919, 8164, 8257, 8653, 8724]. Pratt [1708, 3106, 6651]. Pre [10788]. Pre-Columbian [10788]. Precious [4664]. Preliminary [10572, 10414]. Prentiss [5647, 5664]. Preparation [6590]. Preparing [6920, 8038]. Prescott [1986, 9331]. Prescription [8579]. Present [3131, 5756, 7428, 9129, 10398]. Preservation [2703, 8237]. Preservative [2420]. Preservatives [9279]. Preserve [371, 388, 425]. Preserver [9047]. Preserving [2760, 8868, 9113]. President [6172, 7328]. Press [1600, 1647, 2867, 4282, 3545, 3664, 3748, 4033, 4670, 4341, 5046, 5779, 6170, 5575, 5633, 5513, 6043, 6633, 7620, 10122]. Presses [132, 1204, 2287, 1673, 4001, 5005, 4492]. Pressing [1625]. Pressure [198, 1416, 3202, 3352, 3619, 4361, 4580, 4654, 6186, 6165, 5496, 6728, 8934, 10144]. Pressures [9871]. Prestidigitation [2295, 3590]. Prestidigitator [7578]. Preston [2641]. Prevent [3586, 4812, 9150, 9790]. Preventing [3145, 4039]. Prevention [4607, 5210, 6183, 7344, 9172, 8596, 10186]. Preventive [8716]. Price [346]. Priced [4647, 4491, 8426]. Prices [3633]. Pridd [5439]. Primitive [5802, 9199]. Prince [3157, 10561, 10286]. Princess [7262, 6389]. Princeton [7188, 9396]. Principle [3054, 9842]. Print [8824]. Printer [7762, 200, 341, 3865]. Printing [132, 113, 2287, 1647, 4001, 3346, 4033, 3847, 4492, 6321, 6170, 5513, 5978, 6950, 7620, 7881, 9391, 10639, 9632]. Prints [3098, 10665, 6568, 9130]. Priscilla [5160, 6425]. Prison [3493, 8388, 8274]. Private [6314]. Prize [1908, 3253, 4145, 6726, 6781, 6951, 10302, 10789, 6898]. Prizes [5933]. Probable [7862, 8465, 9207]. Probably [3042]. Problem [1185, 1352, 6864, 6415, 6496, 8292, 8088, 10273, 10636, 10366, 6802, 8235]. Problems [4540, 5889, 10463, 10644]. Proceedings [6057, 7060]. Process [198, 2972, 4545, 5248, 5759, 6042, 6686, 6780, 7551, 10461, 9499, 1753]. Proclamation [1250]. Proctor [1671, 1472]. Produced [18]. Producing [7163, 7551]. Production [1201, 2014, 1688, 6121, 6851, 7815, 7666, 8867, 10094, 9814, 10668, 10325, 10156]. Products [2738, 8151, 9407, 10096, 10767, 8972]. Prof [1119, 2780, 6456, 9652, 9816, 7922, 10613]. Prof. [4297, 5305, 7695, 8062, 7552, 8040, 10040, 10370, 9854, 9974, 9235]. Profession [3651, 9367]. Professor [3125, 2457, 4106, 5037, 6652, 7994, 7943, 1158]. Profitable [7311]. Profits [495]. Prognostications [7995]. Programme [10713]. Progress [1146, 269, 61, 1039, 1459, 1252, 2162, 1890, 1726, 3124, 3001, 3212, 2328, 3540, 3461, 3345, 3931, 3260, 4446, 4360, 5848, 6938, 6953, 6946, 9243, 10694, 9633, 9927]. Project [173, 6204]. Projectile [1257, 3142]. Projectiles [4321, 5191, 6716, 7189, 9162]. Projecting [8447]. Projection [4208, 6563, 1572]. Projections [2855]. Projectors [5838]. Prolonged [10413]. Promoting [10558]. Promotion [7009]. Proof [3501, 3996, 4896, 5155, 4882, 5217, 7951, 8609]. Propaganda [7225]. Propagation [2561]. Propelled [988, 874, 3289, 5171, 5457, 7137, 9437]. Propeller [80, 848, 2488, 4219, 4209, 4243, 5758, 6083, 5915, 8146]. Propellers [8744]. Propelling [1981, 1493, 1564, 4631, 6458, 6542, 7416, 10375]. Proper [9028]. Properties [2660, 5133, 4506, 9233, 2420]. Property [8012, 9487, 8736]. Proportions [6457]. Proposal [9865]. Proposed [569, 387, 814, 428, 1415, 1724, 1805, 2114, 2601, 2560, 2776, 4025, 3284, 3765, 4840, 4782, 6024, 5541, 6182, 5665, 6384, 6803, 6601, 6667, 6793, 6822, 8212, 7920, 8060, 8417, 7570, 7446, 8085, 8397, 9508, 9429, 9435, 9486, 10660, 10352, 10344, 10427, 10536, 10195, 9711, 10398, 10790, 2208, 3097, 10043]. Propulsion [4286, 3591, 4488, 7089]. Proscribed [920]. Prospect [2668]. Prospectors [7977]. Prosperity [8086, 8191]. Protect [7632]. Protected [6577, 9315]. Protecting [3526, 5725, 6692, 9472]. Protection [313, 4077, 4758, 7873, 8736, 9815, 9895]. Protective [3822]. Protector [2640, 5652, 7554, 7950, 9197]. Prouty [1672]. Providence [2686, 7121]. Proving [3354, 9867]. Prunes [5612]. Pruning [82, 10048]. Prying [6123]. Psychology [2897]. Ptomaine [8826]. Ptychoraphis [3112]. Public [3077, 6226, 6236, 6314, 7925, 8717, 9700, 6475, 7855]. Publication [3821, 3862]. Publications [166, 237, 262, 282, 304, 328, 407, 433, 458, 561, 751, 772, 825, 880, 1111, 1138, 1161, 1193, 188, 215, 353, 379, 539, 654, 681, 700, 724, 1025, 1069, 1091, 513, 797, 831, 1285, 1308, 1336, 1384, 1451, 1654, 1719, 1740, 1801, 1858, 1882, 1938, 1955, 2100, 2121, 2173, 2189, 2237, 2258, 1478, 1501, 1527, 1550, 1574, 1606, 1630, 1903, 1920, 2140, 1408, 1454, 1557, 1582, 1613, 2347, 2387, 3014, 3061, 3082, 3100, 3117, 3135, 3151, 3169, 3188, 3206, 3219, 2406, 2425, 2441, 2459]. Publications [2478, 2496, 2517, 2539, 2573, 2612, 2631, 2648, 2663, 2678, 2697, 2730, 2749, 2786, 2802, 2821, 2878, 2906, 2923, 2942, 2368, 2714, 3236, 3255, 3852, 3875, 3898, 3922, 3948, 3969, 3988, 4042, 4066, 4088, 4112, 4129, 4148, 4174, 4198, 4221, 4268, 4290, 4311, 3276, 3295, 3318, 3336, 3355, 3376, 3395, 3437, 3464, 3487, 3511, 3532, 3556, 3577, 3624, 3646, 3666, 3689, 3711, 3734, 3778, 3796, 3814, 4351, 4372, 4978, 5018, 5056, 5076, 5110, 5146, 5181, 5204, 5222, 5278, 5299, 5318, 5335, 4394, 4414]. Publications [4437, 4457, 4480, 4501, 4531, 4554, 4573, 4599, 4621, 4638, 4658, 4676, 4698, 4738, 4761, 4783, 4804, 4831, 4869, 4889, 4911, 4928, 5240, 5351, 5370, 6031, 6050, 6070, 6091, 6114, 6156, 6175, 6191, 6210, 6229, 6268, 6287, 6309, 6331, 5389, 5410, 5430, 5449, 5470, 5486, 5526, 5545, 5566, 5584, 5657, 5670, 5686, 5707, 5748, 5770, 5791, 5817, 5840, 5861, 5883, 5902, 5925, 5949, 6371, 6390, 7156, 7190, 7210, 7244, 7263, 7282, 7304, 7321, 7350, 6409, 6430, 6468, 6489, 6507, 6528, 6593, 6614, 6639]. Publications [6758, 6773, 6795, 6814, 6839, 6857, 6982, 7000, 7015, 7032, 7051, 7069, 7108, 7124, 7141, 7370, 7387, 7402, 7420, 7439, 7479, 7502, 7523, 7543, 7562, 7605, 3418]. Published [514]. Puddler [4142]. Pueblo [8545, 9143, 8834]. Pug [1181]. Puget [7130, 7435]. Pulaski [819]. Pulitzer [1198]. Pull [2611]. Puller [2225, 2366, 5726]. Pulley [83, 2417]. Pulleys [374, 7449, 10569]. Pulling [4413]. Pullman [9704, 1595]. Pulp [623, 10515]. Pulpit [7884, 7384]. Puluj [8041]. Pulverized [5619]. Pulverizing [2166, 2869, 4409]. Pump [1013, 1868, 1298, 2551, 2900, 4143, 4975, 6346, 5360, 6107, 5418, 6063, 5380, 5703, 6668, 7236, 7182, 7594, 8424, 9031, 9983, 9777]. Pumping [694, 2312, 5498, 6974]. Pumps [4453]. Punch [347, 4470, 5402, 5368, 9799]. Punching [2764, 5047, 4694]. Puncture [7951, 8609]. Puncture-Proof [8609]. Punishment [5315]. Punnett [6370]. Pupin [4768, 6588]. Purchased [8757]. Purchases [10371]. Purdue [2655, 5679, 7800, 8682]. Pure [7385]. Purification [3874, 3975]. Purifier [31, 1963, 2470, 4469, 4587, 10102]. Purify [175]. Purifying [9352]. Puritan [1377]. Purposes [4881, 7364]. Purse [4733]. Pursues [1772]. Put [3568]. Putnam [9209]. Puzzle [3599, 4733, 6788]. Pyramidical [6068]. Pyrogallic [9113]. Pythons [2712].

Quadricycle [6697]. Quadruple [9853]. Quaint [9986]. Quaker [3983]. Qualities [7927]. Quality [1229]. Quarantine [2964]. Quarries [9880, 3233]. Quarry [10775]. Quarrying [1088]. Quartz [7636]. Quatre [10748]. Quebec [7920]. Queen [7014, 7320, 326, 7603]. Queer [8800]. Queries [858, 22, 47, 683, 704, 726, 750, 771, 824, 856, 902, 925, 947, 965, 984, 1003, 1026, 1045, 1072, 1095, 1115, 1141, 1164, 1165, 1191, 1192, 1220, 71, 95, 125, 147, 169, 190, 217, 239, 265, 285, 307, 331, 355, 382, 410, 436, 461, 488, 510, 541, 563, 588, 613, 655, 1633, 1268, 1288, 1800, 1821, 1841, 1860, 1881, 1904, 1922, 1937, 1954, 1974, 1990, 2011, 2028, 2048, 2066, 2084, 2104, 2125, 2143, 2156, 2176, 2193, 2216, 2236, 2262, 2281, 1310, 2300, 1311, 1334, 1335]. Queries [1360, 1382, 1383, 1404, 1427, 1428, 1449, 1450, 1480, 1505, 1531, 1552, 1577, 1608, 1658, 1684, 1703, 1721, 1738, 1778, 2317, 2331, 2788, 2806, 2823, 2840, 2859, 2877, 2894, 2926, 2945, 2961, 2982, 2999, 3018, 3033, 3049, 3065, 3085, 3103, 3120, 3138, 3155, 3172, 3191, 3209, 3222, 2350, 2370, 2389, 2409, 2428, 2443, 2463, 2482, 2499, 2519, 2538, 2558, 2576, 2597, 2616, 2634, 2651, 2665, 2682, 2700, 2717, 2733, 2752, 2767, 3239, 3280, 3817, 3836, 3856, 3879, 3902, 3926, 3950, 3968, 3991, 4020, 4045, 4069, 4092]. Queries [4114, 4133, 4152, 4178, 4201, 4224, 4247, 4272, 4293, 4315, 4331, 3299, 3322, 3339, 3358, 3379, 3399, 3421, 3440, 3466, 3490, 3513, 3535, 3559, 3580, 3604, 3627, 3645, 3669, 3692, 3714, 3737, 3761, 3781, 3799, 4354, 4374, 4891, 4914, 4931, 4948, 4962, 4981, 5000, 5021, 5040, 5058, 5074, 5094, 5113, 5130, 5150, 5166, 5185, 5207, 5225, 5241, 5261, 5281, 5302, 5321, 5338, 4397, 4418, 4440, 4460, 4484, 4503, 4519, 4534, 4558, 4577, 4603, 4625, 4641, 4661, 4680, 4701, 4720, 4742, 4764, 4787, 4809, 4834, 4853, 4873]. Queries [5355, 5372, 5866, 5886, 5905, 5927, 5952, 5974, 5994, 6014, 6034, 6054, 6073, 6094, 6117, 6138, 6158, 6178, 6194, 6213, 6231, 6249, 6273, 6292, 6312, 6334, 6354, 5392, 5412, 5433, 5452, 5472, 5490, 5507, 5528, 5547, 5568, 5587, 5608, 5624, 5642, 5660, 5674, 5689, 5710, 5729, 5753, 5773, 5794, 5821, 5842, 6373, 6393, 6893, 6910, 6929, 6964, 6984, 7004, 7019, 7035, 7055, 7073, 7091, 7111, 7127, 7144, 7159, 7175, 7194, 7213, 7228, 7248, 7266, 7287, 7307, 7326, 7341, 7353, 6414, 6433, 6454, 6471, 6492, 6512]. Queries [6531, 6551, 6573, 6596, 6617, 6641, 6661, 6682, 6702, 6721, 6740, 6756, 6775, 6798, 6816, 6841, 6861, 6875, 7368, 7392, 7939, 7965, 7987, 8009, 8030, 8053, 8080, 8099, 8119, 8139, 8162, 8186, 8209, 8230, 8250, 8271, 8288, 8311, 8332, 8350, 8374, 8393, 8414, 8434, 8459, 8483, 7404, 7424, 7442, 7461, 7484, 7505, 7526, 7545, 7564, 7584, 7607, 7628, 7651, 7677, 7698, 7728, 7748, 7774, 7793, 7821, 7844, 7869, 7892, 7915, 8504, 8527, 9086, 9106, 9145, 9165, 9185, 9204, 9223, 9239, 9257, 9277, 9300, 9322, 9343, 9363]. Queries [9380, 9399, 9422, 9441, 9462, 9481, 9503, 9524, 9544, 9565, 8548, 8572, 8590, 8622, 8644, 8667, 8689, 8709, 8732, 8754, 8777, 8820, 8839, 8863, 8886, 8903, 8923, 8947, 8967, 8985, 9005, 9026, 9042, 9066, 9580, 9599, 10226, 10248, 10269, 10292, 10319, 10340, 10362, 10384, 10406, 10435, 10459, 10483, 10503, 10530, 10550, 10580, 10605, 10631, 10658, 10679, 10705, 10732, 10755, 10782, 10803, 9621, 9648, 9671, 9697, 9717, 9740, 9765, 9787, 9812, 9837, 9860, 9892, 9919, 9945, 9971, 9999, 10034, 10062, 10086, 10111, 10133, 10154, 10174, 10200, 10436]. Quest [8378]. Question [7704, 8465, 10634, 10179, 10066, 9817]. Questionable [415, 10208]. Quick [2052, 3618, 4237, 4772]. Quicksand [4410]. Quiet [4727]. Quilter [392, 1565]. Quilting [2509]. Quint [6867]. Quintin [7497]. Quinto [7454].

Rabbit [6737]. Race [875, 1651, 2114, 2610, 3768, 3897, 5199, 5804, 6281, 6247, 6223, 7104, 6835, 7121, 8036, 4060, 6278, 6027]. Races [1399, 5958, 6111, 10467, 10596, 4996, 6076]. Racing [6189, 6065, 6089, 10557, 10593, 10519, 10546]. Rack [1808, 1852, 6295, 6205, 10045]. Racquet [3529]. Radial [5719]. Radiant [7066]. Radiations [18, 6676]. Radiography [6659, 9530]. Radiophone [10079]. Radley [1674]. Raft [598, 9264]. Rags [6920]. Raid [6802]. Rail [86, 1209, 2004, 1829, 1517, 1264, 2812, 2307, 2592, 3470, 4172, 5054, 4959, 4859, 5231, 5458, 5829, 6900, 7133, 7082, 7329, 8214, 8264, 7806, 8327, 7908, 7861, 7970, 7395, 8515, 8598, 9587, 10210, 9278]. Rail-Joint [9587]. Rail-Road [9278]. Rail-Way [1264]. Railroad [867, 177, 812, 844, 1066, 1177, 2271, 1749, 1646, 1517, 1877, 1754, 1979, 1897, 3127, 2307, 2489, 2594, 2911, 2571, 4049, 3471, 3548, 4284, 4172, 3589, 5044, 4799, 4691, 4435, 5032, 5276, 4454, 4878, 4900, 5231, 4724, 4815, 4511, 4569, 5415, 6208, 6201, 5684, 7179, 7270, 6567, 6384, 6807, 7295, 7008, 8020, 8198, 8300, 9507, 9121, 8692, 8693, 8634, 10777, 10389, 581, 1538, 1400, 1251, 4423, 5415, 8314, 7358, 7908, 7861, 10487, 10382, 9879, 10379]. Railroads [325, 6, 3822, 3999, 5284, 7235, 6935, 7779, 8354, 7434, 7848, 7956, 10044, 10229, 10540, 10687, 10210]. Rails [3589, 4711, 4411, 8214, 9750]. Railway [889, 1134, 765, 138, 535, 1175, 551, 471, 892, 1148, 776, 454, 758, 1039, 441, 235, 33, 403, 55, 1542, 1280, 2077, 2273, 1252, 1300, 1946, 1930, 1369, 3214, 2548, 2670, 2383, 2592, 3555, 3631, 4328, 3554, 3960, 3661, 3495, 3634, 3289, 3700, 3727, 3503, 3349, 3614, 3912, 3659, 4543, 4433, 4645, 5064, 5011, 4840, 4782, 5515, 5408, 6327, 6258, 6077, 6259, 6181, 5665, 6320, 5723, 5654, 6398, 6718, 7256, 7027, 7239, 6388, 7103, 6501, 7530, 7929, 7588, 7813, 8069, 8089, 8155, 7373, 8327, 8949, 8768, 9451]. Railway [9431, 10296, 10047, 9763, 9707, 9682, 10209, 10766, 9882, 10745, 3272, 1596, 2170, 1946, 1347, 3149, 3089, 2313, 4307, 3428, 4538, 4464, 4780, 6129, 6068, 6209, 8367, 7764, 9098, 9393, 9841, 10046, 1318, 9910]. Railway-Cars [10047, 9882]. Railway-Sand-Box [9682]. Railways [1200, 1848, 3494, 3284, 3931, 4172, 4014, 4651, 4905, 5725, 8166, 9366, 9226, 9593, 9896, 10797, 10743, 8594]. Rain [1201, 2014, 1688, 2510, 2475, 2310, 5777, 7713, 7666]. Rainfall [10758]. Rainier [5922]. Rainwater [1693]. Raised [3432]. Raising [2322, 3392, 3906, 5872, 9230, 10281, 9760, 9199]. Rake [156, 207]. Raleigh [2647]. Ralston [9056]. Ram [3074, 3342, 5143]. Rameses [8497]. Ramming [5067]. Ramps [9176]. Rams [1853]. Rand [4210]. Randall [4276]. Raney [4970]. Range [1830, 5052, 8128, 8854, 10228]. Range-Finder [8854]. Ranges [1514, 6541]. Ransome [971]. Rapid [494, 2071, 3159, 3732, 4926, 6579, 6516, 6745, 8377, 8189, 7948, 8085, 7591, 8320, 8256, 7469, 8166, 7839, 8489, 8939, 9294, 9391, 9746, 9865, 10440, 9988, 10760, 6970]. Rapid-Fire [8939, 9294, 9746]. Rapidity [7580]. Rapids [8404]. Rare [3541]. Rat [8557]. Ratchet [2222, 4584, 8935]. Rate [10120]. Rater [10453]. Rates [5266]. Rather [8842]. Ratios [8529]. Ratisbon [7415]. Rattler [7716]. Rattlesnake [9830]. Rattling [6560]. Ravenna [7279]. Raw [8830]. Rawhide [5947]. Ray [6588, 6638, 6906, 6637, 7302, 6483, 6548, 10744, 10149]. Rays [8048, 9132, 9632, 7883]. Razor [3910]. Reached [8701]. Reaching [1351]. Reaction [4547]. Read [4713, 5103]. Readers [570, 6341]. Readily [994, 1566, 1350, 3911, 4235, 5762, 6499]. Reading [2846, 4080]. Ready [799, 7616, 834]. Reagan [4883]. Real [2088]. Reaping [2993]. Rear [10240, 9913, 10585, 9639]. Rear-Admiral [10240, 9913, 10585, 9639]. Reason [10117]. Reasonable [3719]. Rebel [1772]. Rebuilding [6522]. Rebuke [10559]. Receipts [10436]. Received [2998, 5148, 5164, 5183]. Receiver [1764, 5630, 6125]. Receives [8470]. Receiving [2040, 8896]. Recently [514, 45, 70, 656, 682, 701, 725, 749, 770, 794, 823, 855, 883, 905, 926, 948, 964, 983, 1006, 1027, 1044, 1070, 1092, 1112, 1140, 1163, 1190, 1221, 1244, 94, 124, 146, 168, 189, 216, 240, 264, 284, 306, 330, 354, 380, 409, 437, 460, 487, 506, 508, 540, 564, 587, 612, 632, 1267, 1287, 1861, 1880, 1900, 1921, 1936, 1953, 1973, 1989, 2008, 2025, 2047, 2065, 2083, 2101, 2122, 2141, 2155, 2174, 2190, 2214, 2235, 2259, 2280, 2299, 1307, 1333, 1357]. Recently [1381, 1402, 1426, 1448, 1479, 1502, 1528, 1551, 1575, 1607, 1631, 1655, 1683, 1702, 1718, 1739, 1756, 1777, 1799, 1820, 1838, 2316, 2330, 2785, 2803, 2820, 2837, 2856, 2876, 2893, 2922, 2944, 2960, 2980, 2996, 3015, 3032, 3046, 3062, 3081, 3099, 3116, 3137, 3152, 3170, 3187, 3205, 3218, 2348, 2369, 2386, 2408, 2426, 2440, 2460, 2479, 2497, 2516, 2537, 2555, 2575, 2595, 2613, 2632, 2649, 2662, 2679, 2698, 2715, 2732, 2750, 2766, 4049, 3238, 3256, 3798, 3815, 3853, 3876, 3899, 3923, 3947, 3967]. Recently [3990, 4019, 4044, 4068, 4089, 4111, 4130, 4149, 4175, 4197, 4222, 4246, 4269, 4292, 4312, 4330, 3277, 3296, 3319, 3337, 3357, 3378, 3396, 3419, 3438, 3463, 3489, 3510, 3533, 3557, 3578, 3603, 3626, 3643, 3667, 3690, 3712, 3736, 3760, 3779, 4353, 4371, 4892, 4910, 4930, 4945, 4963, 4977, 5001, 5017, 5041, 5055, 5075, 5093, 5109, 5127, 5145, 5162, 5180, 5203, 5221, 5242, 5258, 5277, 5298, 5317, 5334, 4393, 4415, 4438, 4458, 4481, 4500, 4516, 4533, 4555, 4574, 4600, 4622, 4640, 4660, 4677]. Recently [4700, 4719, 4739, 4760, 4784, 4805, 4833, 4852, 4870, 5924, 5352, 5373, 5862, 5882, 5901, 5928, 5951, 5975, 5991, 6015, 6030, 6051, 6069, 6095, 6113, 6135, 6155, 6174, 6195, 6212, 6228, 6250, 6269, 6288, 6308, 6333, 6355, 5388, 5409, 5429, 5451, 5469, 5487, 5504, 5525, 5548, 5569, 5583, 5605, 5623, 5639, 5658, 5672, 5688, 5706, 5728, 5749, 5769, 5790, 5818, 5843, 6374, 6392, 6894, 6909, 6926, 6981, 7001, 7016, 7034, 7052, 7070, 7092, 7107, 7123, 7140, 7158, 7174, 7192, 7212, 7229, 7245]. Recently [7267, 7284, 7306, 7322, 7338, 7352, 6410, 6429, 6450, 6467, 6491, 6508, 6530, 6550, 6572, 6595, 6616, 6642, 6660, 6683, 6703, 6722, 6739, 6757, 6776, 6797, 6817, 6842, 6858, 6876, 7369, 7388, 7936, 7960, 7983, 8005, 8025, 8054, 8076, 8094, 8114, 8140, 8157, 8181, 8205, 8225, 8251, 8267, 8289, 8307, 8328, 8351, 8371, 8394, 8410, 8433, 8454, 7405, 8479, 7419, 7438, 7457, 7480, 7504, 7525, 7546, 7565, 7581, 7604, 7625, 7652, 7678, 7699, 7723, 7745, 7771, 7794, 7817, 7845, 7864, 7889, 7910, 8757]. Recently [8500, 8522, 9067, 9081, 9102, 9125, 9146, 9166, 9186, 9200, 9224, 9238, 9258, 9273, 9295, 9318, 9344, 9360, 9377, 9397, 9423, 9442, 9458, 9479, 9501, 9525, 9543, 9562, 8547, 8567, 8591, 8621, 8643, 8666, 8686, 8704, 8728, 8755, 8776, 8797, 8818, 8838, 8862, 8881, 8904, 8920, 8944, 8963, 8986, 9000, 9022, 9043, 8726, 10475, 9577, 9596, 10223, 10244, 10266, 10287, 10315, 10341, 10357, 10385, 10405, 10431, 10455, 10484, 10502, 10526, 10551, 10576, 10602, 10628, 10654, 10680, 10706, 10733, 10751, 10783, 10804, 9617]. Recently [9643, 9668, 9692, 9713, 9735, 9766, 9786, 9809, 9836, 9857, 9888, 9920, 9940, 9966, 10000, 10031, 10058, 10087, 10112, 10134, 10155, 10171, 10197, 6904]. Receptacle [5698]. Reception [8869]. Recesses [3571]. Reclamation [7464]. Reclining [4101]. Recognition [9639]. Recoil [7366]. Reconnaissance [8995]. Reconstructed [9116, 10239]. Reconstruction [8702, 9366]. Record [8259, 8317, 8538, 9925]. Recorder [917, 2739, 2989, 4304, 3619, 5832, 6627, 8936]. Recording [325, 1747, 3202, 4281, 4104, 3893, 5016, 5554, 6541, 9556]. Records [6077, 8442, 8909, 10005]. Recovered [9538]. Recovery [7980, 9122, 8509]. Recreations [3459]. Recruiting [8930]. Red [3233, 3268, 4009, 9120]. Redoutable [92]. Redrawing [5855]. Reduced [832, 10255]. Reduction [1587, 8937]. Redwood [720, 8389, 8519]. Reece [4095]. Reef [1932, 9602]. Reel [6348]. Refillable [6994, 7200, 8106, 8153, 9708, 9685]. Refilled [6254]. Refining [3435, 3934]. Reflecting [3957]. Reforged [6889]. Reform [2466, 8318, 8783]. Reformatory [8495]. Refraction [3757]. Refrigerating [2590, 8638]. Refrigeration [2022]. Refrigerator [1150, 7410]. Refuse [5762, 7993]. Regard [618, 8295]. Regia [5766, 8222]. Regiment [5273]. Regimental [9119]. Register [767, 421, 917, 1930, 2527, 4970, 9112]. Registering [5362, 5871, 7363, 9477]. Registration [4252, 8087, 8824]. Regular [9403]. Regulated [2847, 3726]. Regulating [6186, 7043]. Regulation [7381, 8824]. Regulator [1202, 672, 1173, 2201, 1416, 2551, 3569, 5516, 7411, 8934, 10144]. Reid [4586, 4612]. Rein [599, 843, 639, 4954]. Reina [10097, 10166]. Reindeer [2161]. Reinforce [7430]. Reinforced [6301]. Reinforcement [9078]. Rejected [8032]. Rejoinder [5592]. Relating [91, 271, 594, 1064, 334, 1669, 2589, 2601, 2978, 4817, 4629, 4728, 4995, 5062, 5140, 5746, 5344, 5447, 5522, 5826, 5938, 6002, 6090]. Relation [7097, 7548, 8101, 7634, 9775, 9898, 10758]. Relations [2147, 7076]. Relay [7104, 9983]. Reliable [1301, 10067]. Reliance [1049]. Relic [9160]. Relics [4478, 5857]. Relief [4495, 6728, 7570, 8534, 8725, 8803, 9029, 10121, 5313, 9091]. Relieve [10542]. Reloading [5614, 10770]. Rem [8449]. Rem-Sho [8449]. Remains [9142, 10788]. Remarkable [1773, 1754, 1264, 2579, 2801, 4561, 5636, 6001, 5956, 6421, 7885, 7532, 7955, 10147, 9762, 10221, 9569, 9267, 3414]. Remedy [1119]. Remember [10531]. Remembering [5911, 5980]. Removable [229, 1674, 10566]. Removal [1932, 2393, 9464, 10400]. Removed [8463]. Remover [6831]. Removing [6478, 8218, 7997, 8580, 9989]. Renaissance [7315]. René [7014]. Renewal [8320]. Reno [7399]. Repair [3252, 6350, 6980]. Repaired [3294, 3328]. Repairing [1241, 5779, 5405, 8478, 8146, 9712, 9608]. Repairs [8102]. Repeater [32]. Repeating [978, 5835, 5999, 5943, 5957, 8608]. Repeating-Electrical [5957]. Repeats [10506]. Replace [8594]. Replacer [1566]. Replies [1345]. Reply [8899]. Report [174, 1413, 2413, 3343, 4251, 4606, 5493, 6746, 6579, 7343, 8396, 8335, 8379, 9163, 8489, 8799, 9253, 8762, 8814, 9513, 8674, 9510, 9770, 10272, 10688, 10669, 10276, 10116, 10791, 10467, 8377]. Reports [1399, 10763]. Repoussé [502]. Representatives [221]. Represented [3958, 3997]. Reproduction [852]. Republic [9375]. Republics [7180]. Required [9606]. Research [3213, 4379, 10253]. Researches [10019]. Reservation [7412]. Reserve [1965]. Reserves [2918]. Reservoir [7039, 5176]. Reservoirs [674, 8381]. Resetting [7450]. Residence [852, 628, 1878, 3132]. Resistance [2914, 10273, 10180]. Resolution [2601]. Resonance [2706, 5133, 973]. Resources [6339, 9248, 9721, 7243, 8972, 3211]. Respiration [4144]. Respirator [5582]. Responsibilities [7682]. Responsibility [9111]. Rest [713, 117, 574, 1854, 4139, 6925, 8238]. Restoration [6381]. Resumption [3862]. Retaining [6282, 9137, 10626]. Retina [5989]. Retrospect [6360, 7360, 8486, 9567]. Return [6145, 6142, 7760, 8254, 9194, 9231, 10354]. Returned [7470]. Revenue [9061, 9006]. Reverse [10412]. Reversible [548, 1591, 8935]. Reversing [1330, 3868, 4939, 5986, 8127, 8739, 10056]. Review [8628, 3404, 3539, 3584]. Revise [9448]. Revision [9173]. Revival [9368]. Revoking [7234]. Revolutionized [6962, 9128]. Revolvers [642]. Revolves [7046]. Revolving [860, 227, 917, 2743, 5479, 5441, 6428, 10693]. Reward [4488]. Reynaud [3011]. Rheinfelden [8365]. Rheostat [1326]. Rhine [8365]. Rhinoceros [5634]. Ribbon [2950, 5854, 8878]. Ribbons [9977]. Rican [10477]. Rice [520, 6725]. Rich [4706]. Richard [7490]. Richardson [10372]. Rico [9484, 9818, 10413, 9192, 8937]. Rid [2920]. Ride [10229]. Rider [10014, 4029]. Riders [5781]. Ridge [181]. Riding [1035, 6299, 7454, 7999, 10016, 10020]. Rife [646]. Rifle [1087, 978, 1680, 1830, 2563, 3519, 5835, 6566, 7063, 8175]. Rifles [5177, 4581, 9964, 10053]. Righthandedness [4717]. Rights [1667, 8843, 9048, 9387]. Rigid [7114]. Riker [10544]. Riley [1980, 5037, 6109]. Rim [2873, 5440, 6787]. Rims [6633]. Ring [846, 1062, 3592, 7044, 8679, 3200]. Ringens [5970]. Rink [6584]. Rio [140, 3485]. Ripper [9213]. Ripping [7622, 8978]. Ritual [8134]. River [277, 101, 892, 694, 2077, 1890, 3333, 4923, 5134, 5160, 4825, 5153, 4705, 4773, 5157, 5541, 6182, 6461, 6809, 7075, 8050, 7445, 7455, 7476, 8060, 7618, 7981, 8614, 8528, 8605, 10709, 10089, 9575, 10195, 9286, 8340, 765, 1179, 1039, 428, 1098, 1377, 1368, 1252, 1724, 3410, 4324, 4669, 5306, 4943, 4593, 5364, 5475, 7095, 7345, 6705, 6824, 8376, 10595, 10310, 3273]. Rivers [3635, 4880]. Riverside [211, 1878, 8636]. Riveting [6399, 6420]. Rivett [6606]. Road [2339, 2446, 4161, 3872, 3749, 3941, 6602, 8715, 9267, 1707, 2128, 6600, 9627, 9278]. Roadbed [3614]. Roads [1521, 3176, 5098, 7096, 7097, 6802, 6803, 7548, 8059, 8323, 8554, 10139]. Roaring [4908]. Robbers [4077]. Robert [3518, 10718, 9611]. Robinson [6630]. Roche [6235]. Rochester [2929, 9556]. Rock [992, 2210, 4210, 4692, 4696, 8156, 7904, 9956]. Rocker [6106]. Rocket [3133]. Rocking [3037, 3588, 8044, 8197]. Rocks [8856]. Rocky [10545]. Rod [9608, 3200]. Rods [291, 2474, 2899]. Roentgen [6456, 6498, 6589, 6483]. Rogers [2972]. Roll [2474]. Rolled [5248]. Roller [1100, 289, 1416, 3981, 3790, 7086, 7657, 8341, 7495, 8292, 9336, 9115, 10473, 9987, 10496]. Roller-Bearing [10473, 10496]. Rollers [10258]. Rolling [1234, 1209, 2255, 1943, 1874, 3722, 3710]. Rolls [5232]. Roman [2169, 5270, 5480, 7958, 9097, 10623]. Rome [3010, 3410, 8654, 9608]. Rondout [556]. Roof [3681, 4444]. Roofing [3226]. Roofs [181]. Rookeries [10170]. Rookery [8303]. Room [115, 6265, 8242]. Rope [226, 2571, 4085, 5291, 7693]. Ropes [1306]. Ropp [6126]. Rordame [4012]. Rose [9512]. Rotary [955, 1442, 1647, 2526, 2934, 4206, 4234, 3655, 4434, 5555, 5852, 7436, 7594, 8611, 8806, 8913, 9533, 9015, 9033, 9496, 10618, 10396, 10450, 10799, 10237, 10497, 10475, 9638, 7614]. Rotated [1257]. Route [7806, 8090]. Row [6562]. Rowell [441]. Rowland [9980]. Royal [9613, 5616, 6154, 6717]. Rubber [2932, 2586, 3161, 4282, 3478, 6321, 7087, 9079, 3184, 121]. Rudolf [4801]. Ruins [7206]. Rule [3890, 4171, 7683]. Ruler [1277]. Rules [2089, 4050, 8398, 7919]. Run [2211, 5101, 7532, 6170]. Runabout [10129]. Runaway [760, 8255]. Runner [1674]. Running [259, 3290, 5099, 7433]. Running-Board [259]. Runs [9234]. Rupture [2704]. Ruric [3267]. Russia [5013, 6889, 9049, 10203, 9754]. Russian [4190, 3267, 6246, 9414, 10695, 5331, 9034]. Rust [6183, 7344, 8716].

S [2421, 2865, 3519, 5305]. S. [194]. Sacks [3906]. Sacramental [2726]. Sad [851, 532, 4588]. Saddle [391, 2581, 3663, 3570]. Saddles [1035, 3888, 8932]. Safe [1159, 4211, 8654]. Safes [5316]. Safety [756, 916, 665, 343, 441, 75, 33, 315, 2172, 1491, 1603, 2017, 1494, 1521, 2345, 2357, 2794, 2566, 4049, 3979, 3910, 4076, 3549, 3705, 3999, 3224, 3350, 4428, 6187, 5739, 7183, 6901, 6710, 7644, 8190, 9471, 9495, 9551, 9411, 8720, 10044, 10321, 10774, 10766, 10743]. Safety-Attachment [9495]. Safety-Gate [9471]. Sagacity [7533]. Sail [6592]. Sailing [1806, 2303, 4550, 6108, 6785, 9683]. Sailor [7945, 9350]. Saint [3508]. Salamanca [4649]. Salaries [337]. Sale [8295]. Salicylic [8067]. Salmon [336, 7095]. Saloon [5784, 5160]. Salt [1182, 1908, 2829, 10617, 2399]. Saltpeter [7735]. Salvatore [2339]. Samoan [9995]. Sampling [2741, 9211]. Sampson [8961, 9163]. Samsons [9099]. Samuel [1644, 1645]. San [2339, 4467, 6969, 8156, 255, 996, 9370, 8937, 9456, 8953, 9136, 9375, 10595, 10472, 10524, 9963]. Sand [1208, 5666, 8244, 9682, 3486]. Sanding [9310]. Sandstone [3233, 3268]. Sandy [8822, 10022, 9992, 9884]. Sanguir [3080]. Sanitary [2511, 4527]. Sanitation [9489]. Santa [3040, 3943, 10727]. Santiago [9163, 8995, 9109, 9195, 9107, 9088, 9018, 8987, 8981, 9168]. Santos [9395]. Santos-Dumont [9395]. Saracibar [4649]. Sargent [3145]. Sash [3686, 4709, 6244, 5512, 6164, 8218, 8790, 9265, 9036, 10236]. Sash-Balance [10236]. Sash-Holder [9036]. Satellite [9926]. Satisfactory [3675]. Saturn [9926]. Sausage [3552]. Savage [7063, 8175]. Saver [2465, 8017]. Saving [10347, 544, 1421, 3133, 3369, 5286, 6126, 6082, 6714, 9328, 10302, 10789]. Saw [83, 993, 10, 647, 12, 1894, 1672, 3978, 4233, 3408, 6275, 6004, 6870, 7135, 6850, 9958, 10449]. Saw-Mill [1672]. Sawdust [1597]. Sawing [721, 3408, 6206, 7082]. Sawmill [3038, 2795, 5328, 7083]. Saws [972, 2036, 3936, 3848, 4426]. Saxony [7114]. Scab [3586]. Scaffold [691, 3593, 8541, 9729]. Scaffold-Bracket [9729]. Scaffolds [7232, 9495]. Scale [3456, 7997, 10778, 10697, 10758]. Scales [4745, 4970]. Scarf [6168]. Scars [6568]. Scene [4065]. Scenes [7618]. Scheme [8256, 8377]. Schemes [7470, 10208]. Schenectady [10230]. Schley [8995]. Scholarship [2772]. School [929, 659, 1170, 3174, 3246, 3480, 6265, 8407, 9177, 9489, 9514, 10565, 4078, 8261, 7454, 8165, 9843]. Schools [2257, 3839, 5229, 6989, 8495]. Science [8368, 1508, 1324, 1898, 6037, 5909, 7218, 6474, 9465, 10327, 10418, 10514, 10181, 7539, 7319, 515, 1342, 1367, 1390, 1410, 1435, 1456, 1536, 1559, 1584, 1615, 1638, 1663, 8827, 8733, 3, 1997, 2774, 5004, 4985, 6018, 8035, 7941, 7427, 10184, 10038, 8103, 8194, 7706, 871, 9229, 9263, 10394]. Sciences [3243, 5678, 7672, 9954, 10691, 8869, 5620]. Scientific [901, 1679, 6570, 7972, 7659, 8066, 8147, 8628, 9552, 10364, 236, 375, 722, 787, 3076, 2899, 9152, 10419, 10445, 707, 187, 213, 261, 281, 303, 327, 351, 432, 507, 630, 652, 679, 699, 439, 800, 835, 513, 797, 831, 801, 836, 485, 492, 516, 1110, 1137, 1224, 1265, 1284, 1356, 1476, 1500, 1526, 1548, 1604, 1628, 1653, 1457, 1560, 1585, 1616, 1408, 1454, 1557, 1582, 1613, 1366, 1411, 1317, 1341, 1365, 1387, 1433, 1484, 1510, 1537, 1639, 1664, 1618, 1642, 1617, 2816, 4211, 5798, 6959, 9418, 8582]. Scientific [8911, 10277, 10436]. Scissors [1254, 1964, 5359]. Scoop [8657]. Score [1101]. Scorpion [3600]. Scotia [8302]. Scotte [5667]. Scottish [8302]. Scourer [2507]. Scourge [5331]. Scouring [1652]. Scout [4217]. Scow [1932, 7823]. Scraper [4, 2795, 9535]. Screen [710, 2952, 3247, 4298, 6146, 6692]. Screened [2187]. Screw [1870, 2172, 2211, 2098, 2817, 2992, 2584, 3417, 4159, 3369, 4244, 5172, 4448, 6127, 6833, 6827, 7617, 7758, 7880, 8146, 10242, 8744]. Screw-cutting [1870]. Screwdriver [5172]. Screws [1711, 2972]. Scrubbing [9212]. Sculptors [10778]. Sculpture [4122]. Sculpturing [4340]. Scythe [6499]. Sea [677, 1775, 2637, 3252, 4156, 3364, 3920, 5026, 5813, 5899, 5814, 5778, 8384, 7681, 8397, 7373, 7643, 8841, 9096, 9188, 9170, 9214, 9150, 9172, 9115, 9151, 10232, 9965, 10004, 10186, 99, 10170]. Sea-Coast [10004]. Sea-Going [7643, 9214, 99]. Seabury [494, 9161]. Seagoing [6655, 8091]. Seal [4672, 6868, 9053, 8634, 8565, 6960]. Sealer [529, 9175]. Seam [9213]. Seam-Ripper [9213]. Seamless [943, 4142]. Seaport [805, 806]. Seaports [1709]. Search [3962, 5052, 10415]. Searching [2977]. Searchlights [10642]. Seas [5253]. Seat [548, 2167, 3264, 5895, 6749]. Seated [4637, 6370]. Sebae [1184]. Second [27, 1321, 1539, 6100, 7827, 2965, 4569]. Second-Class [6100]. Secondary [1414, 2243]. Secrecy [10161]. Secretary [8194, 1460, 7343, 8379, 9513, 9510]. Secrets [7115, 10737]. Section [6851, 7611, 9141]. Sectional [6041, 5921, 5632, 7012]. Sections [5193]. Sectors [4753, 4752]. Securing [4243, 6565]. See [4713]. Seed [5005, 9823, 10616]. Seed-Planters [9823]. Seeds [7730]. Seeley [3125]. Segmental [7278]. Seismograph [8045]. Selection [9791]. Self [1108, 1592, 3769, 4857, 4970, 6343, 7116, 7043, 8403, 9266, 10375, 7383, 3844]. Self-Ballasting [1108]. Self-Binding [7383]. Self-Cleansing [4857]. Self-Closing [9266]. Self-Heating [8403]. Self-Loading [3769, 7116]. Self-Playing [6343]. Self-Propelling [10375]. Self-Register [4970]. Self-Regulating [7043]. Self-Winding [3844]. Selling [1197]. Sells [10252]. Semi [6821, 6951, 7670, 10312]. Semi-Automatic [10312]. Semi-Centennial [6821, 6951]. Semi-Traction [7670]. Senate [9817, 9772]. Sending [279]. Senegal [3832]. Senor [4649]. Sensation [4683]. Sensationalism [9465]. Sensitive [274]. Sentinel [9354]. Separable [1237, 864]. Separating [2338, 3701, 4664, 6519]. Separator [2507, 4586, 7400]. September [840, 1998, 2930, 3986, 5404, 6039, 6330, 8110, 9219, 10374]. Sereno [2572]. Serious [7197]. Serpent [505]. Serpollet [1423, 4613, 5705]. Servant [2024]. Service [889, 1966, 2502, 3193, 4204, 4543, 4454, 6063, 7103, 7898, 7948, 8260, 8038, 8015, 8387, 8783, 9445, 8763, 9775, 1421, 2928, 2748, 8584, 9120, 8713, 9829]. Serviceable [6674]. Sesquicentennial [7188]. Session [8600, 9628]. Set [972, 4565, 8679]. Setting [1103, 3978, 7450, 9958]. Settlement [7925]. Seven [2602, 2754, 7611]. Seventeen [5125, 4829]. Seventeen-Year [5125, 4829]. Seventh [1707, 2128, 4876, 7358, 4773, 8213, 8336]. Seventy [3150, 5273, 8275]. Seventy-First [5273]. Seventy-Five [3150]. Several [1083]. Severe [9535]. Seville [5084]. Sewer [143, 2057, 3022, 2910, 2382, 7025, 6695, 8684]. Sewerage [340, 2818, 3203]. Sewers [1109, 4279]. Sewing [687, 688, 301, 608, 1565, 6947, 7622, 8978, 9826]. Sewing-Machine [9826]. Sextant [5420]. Sextuplex [8068]. Shade [204, 5894, 6917, 6518, 7465, 7753, 7714, 9987]. Shade-Roller [9987]. Shades [5939]. Shadows [7696]. Shaft [536, 420, 1775, 3252, 3685, 3294, 3328, 5875, 7859]. Shafting [5193, 5170]. Shafts [917, 3052, 6084, 8788]. Shaker [6747, 8197]. Shaking [4883]. Shall [8846, 9604]. Shamrock [10150, 10337, 10534, 10368, 10572, 10480, 10441]. Shapers [4240]. Shaping [1063, 3450]. Sharks [7491]. Sharpener [890, 348, 4955, 5359, 8678]. Sharpening [3329, 5816, 6812]. Shattered [5034]. Shaver [6747]. Shaving [622]. Shear [3150]. Shearing [2764]. Shears [5402, 5359]. Sheathed [10352]. Shed [9290, 10260]. Sheeley [2760]. Sheet [278, 1063]. Sheets [10763]. Sheffield [4220, 6902]. Shelf [349, 3094]. Shell [2553, 4590, 10041]. Shells [9594, 10770]. Shelving [1351]. Shely [2759]. Shepherd [5123, 8897]. Sherman [65]. Shield [476, 2743, 4896, 5851, 6424]. Shields [892]. Shifter [6345, 6651]. Shifting [845, 3367]. Shingle [176]. Shingling [2723]. Ship [860, 92, 277, 245, 42, 804, 921, 942, 1170, 1596, 2188, 1471, 1772, 2642, 2937, 3158, 3058, 2394, 3741, 4323, 3458, 3267, 3766, 3675, 3362, 4244, 3638, 3941, 3731, 4877, 5324, 4550, 4498, 4522, 4549, 4924, 5080, 5067, 4497, 5294, 6184, 5812, 5899, 5785, 6225, 5348, 5755, 7086, 7050, 6367, 8470, 8397, 9197, 9091, 8724, 8854, 8896, 9227, 9614, 10078, 4636, 7561, 9286, 3758, 3921, 9931]. Shipbuilder [1345]. Shipbuilders [10344]. Shipbuilding [2771, 6956, 7291, 9368, 9427]. Shipman [2312]. Shipment [1696, 7719]. Shippers [8740]. Shipping [1378, 8538]. Ships [544, 814, 443, 1513, 2290, 2130, 1641, 2668, 3674, 3742, 4856, 5403, 5376, 5405, 5580, 7758, 8164, 9159]. Sho [8449]. Shoal [9602]. Shoals [6200]. Shock [3387, 5178]. Shoddy [6920]. Shoe [81, 393, 398, 2171, 3315, 8722]. Shoes [2511, 2321, 6298, 5518, 3184]. Shooting [2611]. Shop [2089, 9753]. Shore [8340]. Shorthand [7785]. Shot [6648, 7298]. Shovel [4468, 6869]. Show [4407, 7501, 9689, 5416]. Shower [6443, 7648]. Showing [3939]. Shown [3979, 4016, 4086, 4160]. Shutter [1187, 1102, 2758, 5961, 10104, 5579]. Shutters [7432]. Siam [8204]. Siberia [6201]. Siberian [4598, 9278, 10379]. Sibley [976, 2772]. Sick [1700]. Sickle [2400, 4405, 10395]. Side [131, 3226, 5921, 5684, 5654, 6425, 6370, 7197, 7718, 8693]. Sided [7579]. Sidewalk [974, 2334]. Siding [229]. Siege [9007, 9610]. Siegfried [921]. Siemens [3195, 3249]. Siena [9691]. Sierra [7335]. Sierras [2439]. Sifter [105, 5827]. Sight [2563, 5117, 5009, 4717]. Sign [1282, 4192]. Signal [325, 310, 164, 452, 151, 1214, 414, 2489, 2670, 3503, 4390, 6103, 6122, 6887, 7165, 8020]. Signaling [1877, 1473, 3228, 5557, 6991, 6879, 7358]. Signals [8980, 10745]. Signor [10490]. Signs [8980]. Sikkim [2458]. Silchester [9097]. Silesia [5418]. Silk [2450, 6207, 9401, 10438, 1184]. Sill [2074]. Silver [1836, 3435, 3934, 3809, 8509, 2453]. Silvering [4147]. Similar [6441]. Similigraph [1126]. Simonds [4846]. Simoon [390]. Simple [186, 1085, 1036, 1131, 994, 1929, 1796, 2207, 2018, 1944, 2000, 2136, 1797, 2091, 2075, 1445, 2200, 1301, 2758, 3007, 2690, 3407, 4258, 3330, 3660, 3264, 4121, 3705, 4242, 4757, 4750, 5256, 4794, 6127, 5742, 5726, 5830, 6105, 6583, 7783, 8361, 8003, 7599, 7600, 8107, 9532, 8873, 8812, 9095, 9141, 10013, 10123, 10421, 10652, 10673, 10724, 10146, 10258, 10422, 9758, 9849, 9729, 10701, 10259, 9882, 9907, 6678, 1814, 1326, 3639, 3496, 5579, 7767]. Simplex [9114]. Simplified [4759, 8994, 9616]. Simplon [7638]. Simpson [8355]. Sims [706]. Since [6939]. Sine [9949]. Sine-Wave [9949]. Singer [608]. Singing [2853, 9010]. Single [1496, 2076, 6980, 8327, 8738]. Singletrees [1543]. Singular [1787, 3364]. Sinkable [814]. Sinking [3766, 5946, 7253, 8444, 9724]. Sinks [626, 4877, 5717]. Siphon [694, 2361, 2725, 7303, 10075, 9520]. Sir [4994, 6172, 8772, 9490, 10717, 8070]. Sisal [225]. Site [51, 60, 5121]. Situation [8981]. Six [4050, 5966, 9215, 7683]. Six-Inch [9215]. Sixteen [6793, 10228, 9674].

Sixteen-Inch [10228, 9674]. Sixty [2709]. Size [6794, 10233]. Sizes [2251]. Sizing [6007, 6693]. Skate [9683, 1984]. Skate-Sailing [9683]. Skaters [8174]. Skates [6592, 9634]. Skating [3235, 6584]. Sketches [7501]. Sketching [3314]. Skiascope [5083]. Skill [2938]. Skimmer [5815, 10102]. Skins [5738, 6590]. Skirmishing [10572]. Skirt [6853, 6692, 7554]. Sky [2392, 6340, 9535, 6647, 6437, 6707]. Sky-Scraper [9535]. Skylight [368]. Slack [5831, 7639]. Slag [501]. Slate [5520, 6044, 10325, 9334, 9358, 9374, 9390, 9419, 9329]. Slates [887]. Slatting [7690]. Slaughtering [736]. Slave [7559]. Sled [879, 848, 4449]. Sleds [9452]. Sleep [8168]. Sleeper [1670, 2307]. Sleepers [1177]. Sleeping [3500, 4687]. Sleigh [937, 1674, 4209, 5255, 5913]. Sleigh-Runner [1674]. Sleighs [3266]. Slicing [550]. Slide [886, 891, 1969, 1870, 2045, 2232, 2793, 3890, 3937, 4139, 5535, 8104, 10701, 3496, 4850]. Slide-Valve [10701]. Slides [3416, 3271, 4567]. Sliding [3162, 9934, 8340]. Sling [426]. Slipper [944]. Slow [7721, 4991]. Slowly [9149]. Small [1676, 1985, 2623, 5216, 4506, 5267, 5191, 6542, 8638, 8842, 1305, 5333, 8144]. Smaller [10667]. Smallest [2514, 9375]. Smith [8658]. Smithsonian [4582, 8674, 8510]. Smoke [909, 1496, 2713, 3145, 5159, 5716, 7738, 8176, 8596, 9851]. Smoke-Consumers [9851]. Smokeless [270, 335, 8175, 8539, 9210, 8988, 9584, 10513, 10067, 10098]. Smoker [5138]. Smokers [4034]. Smokestacks [10281]. Smoking [4620]. Smuggling [1324]. Smyrna [10213]. Snake [791, 1828, 7300, 10168, 10425, 932]. Snakes [1942, 3895, 9457, 7602, 1951, 1987, 1487, 2079, 2120]. Snap [640]. Snow [1376, 1440, 3104, 2393, 2439, 4056, 3315, 4565, 5101, 6478, 7956, 7520]. Snows [5288]. Snowstorm [9509, 9723]. So-Called [9279]. Soap [38, 2708, 2951, 4527, 9285, 3576]. Social [7655]. Societies [6037]. Society [100, 27, 1761, 2032, 1273, 2335, 2619, 3283, 6881, 8219, 7830, 8200, 7376, 7852, 8405, 8600, 9978, 9755, 2015, 4401, 5313, 4974, 7009, 9447, 9120, 9868]. Sociological [8316]. Socket [9432, 1326]. Socket-Wrenches [9432]. Soda [311, 1444, 2503, 10328, 10542]. Sodium [4336, 4359]. Sodom [674]. Sofa [6852]. Soft [1675]. Soil [5848, 7802, 10663]. Soils [9500]. Sokol [6246]. Solar [18, 693, 3572, 5025, 6422, 8858, 8794, 6622, 7066]. Soldering [1067, 8385, 7665, 8403, 10100]. Soldier [8971, 9330]. Soldiers [2587, 9028]. Solicitors [6822]. Solid [943, 1763, 6141, 10521]. Solids [4580]. Solution [8465, 8088]. Solutions [9113]. Solved [1348]. Solving [7541]. Some [959, 629, 1734, 3002, 2920, 4795, 5065, 4400, 5806, 5655, 6486, 7299, 6972, 8418, 7957, 8628, 8521, 8955, 9457, 9233, 8726, 9615, 10471, 10009, 9827, 9874, 10800, 10019, 10328, 10670, 4547, 7414, 8998, 10598, 9939, 165, 1850, 2233, 1915, 10218, 3201, 10098]. Something [9074, 10531]. Song [4976, 9348]. Sons [4017, 9417]. Sorely [10159]. Sorghum [2241]. Sorting [6709]. Soudan [6719, 9327, 9593]. Sound [2706, 7435, 7130, 4850]. Sounder [5250]. Sounding [1465]. Source [4953, 8191, 9928]. South [1813, 4015, 7049, 5439, 518, 3372, 3920, 6646, 7180, 8084, 8889, 8970, 9282, 10723, 2291]. Southbury [416]. Southern [8201, 3185, 5515, 6112, 10221, 10598, 10170]. Southwestern [6725]. Spacer [8807]. Spaces [5016, 7801]. Spaghetti [7809]. Spain [55, 4649, 5084, 8313, 8898, 8793, 9167, 9151, 9873, 10252]. Span [4876]. Spanish [8427, 8899, 9090, 9140, 9538, 9062, 9253, 8948, 8747, 9058, 9017, 8781, 9349, 9840, 9594]. Spar [8852]. Spark [1730, 6477]. Sparkstötting [4449]. Speaking [2504, 3230, 5091, 5807]. Spearing [2344]. Special [4382, 7952, 8845]. Species [10671]. Specific [1547]. Specimens [3811, 4958]. Spectacle [2784]. Spectroscopic [9616]. Spectrum [5700]. Speech [2940]. Speed [637, 1513, 2088, 1619, 3005, 3871, 3674, 4254, 4184, 4366, 4917, 4772, 4561, 5912, 6077, 5956, 6716, 6517, 6605, 6457, 6764, 6734, 6380, 7401, 7734, 7689, 8467, 7432, 8464, 7721, 9531, 9148, 10567, 10233, 9896, 10039, 10104, 10049, 10414, 10005, 10180]. Speed-32 [4561]. Speeding [6079]. Speeds [6398, 7530]. Speedway [5153, 4593, 7455, 7476, 7445, 7801, 7465]. Spelling [3933]. Spent [10464]. Spider [8617]. Spiders [4388]. Spiegeleisen [5014]. Spindle [864, 1395, 2450, 5565, 10589]. Spinning [864, 2450, 8566]. Spiral [6484]. Spire [5033]. Spirit [9334, 9358, 9374, 9390, 9419]. Spiritualism [10181]. Spiritualistic [8113, 8136]. Splice [1306]. Spoke [2225]. Sponge [887, 2220]. Sponges [9666]. Spool [576, 6976]. Spooled [2184]. Spools [1856]. Sporting [1323, 5704, 5724]. Sportsmen [9796]. Spot [2436, 2593]. Spots [4591, 5787, 6837, 8726]. Sprague [8595]. Spray [6144]. Spraying [7976, 8993]. Spread [2071, 3285]. Spreading [6670]. Spree [4669]. Spring [5981, 5869, 8198, 7669, 363, 1518, 1379, 1494, 1543, 1948, 2644, 6749, 9960]. Spring-Cushioned [8198]. Spring-Supported [7669]. Springfield [6037, 1813, 4861, 6189, 6057, 9964, 10053]. Springless [2249]. Springs [6130]. Sprinkler [2829, 3568, 4412, 6327, 6626, 8152, 8768]. Sprocket [5215]. Squadron [9078]. Squadrons [8899, 8898]. Square [2535, 3827, 5574, 5436, 5416, 6436, 7154, 6689, 9409]. Squeezer [9961]. Squeezing [4142]. St [8123]. St. [765, 892, 676, 2077, 1815, 2293, 2684, 3661, 5235, 4951, 5801, 5784, 5637, 5763, 6264, 6913, 6916, 6972, 6459, 7944, 7920, 7806, 10099, 10757, 9608]. Stability [4856, 6578, 7477, 10611]. Stable [6656]. Stacker [1520, 2604, 3656, 6585, 9904]. Stacking [208]. Stadium [6381]. Staff [201, 1793]. Stage [504, 1651, 8066, 10523, 9267]. Stained [537, 5745]. Stains [2453]. Staircase [2468]. Stairways [4215]. Staking [5738, 5648]. Stalk [2198]. Stalks [467, 10643]. Stall [4007, 9707]. Stamp [2932, 4282, 3773, 5285, 4616, 6131, 7049, 8151, 10423]. Stamping [88, 7536, 9638]. Stamps [429, 6321]. Stand [663, 1966, 1396, 3744, 4080, 4235, 6825, 8633, 5333]. Standard [1089, 1467, 2119, 4362, 9034]. Standing [5536]. Stands [2226]. Star [3163, 1994, 9724, 9688]. Starling [4958]. Stars [6039, 2994]. Starter [77, 783]. Starting [6122]. Starts [7991]. State [1965, 1620, 1280, 3811, 3504, 3830, 3963, 4780, 5869, 9351, 8552, 8495, 9609, 7138, 4633, 4324, 5897, 7050, 8952]. States [1169, 1421, 5143, 6099, 6143, 5580, 6342, 7278, 6879, 8072, 7824, 7569, 8234, 7542, 7777, 9527, 8683, 8581, 8724, 9315, 9129, 9294, 8597, 8930, 8713, 10398, 9908, 10220, 9867, 99, 849, 765, 302, 387, 1089, 2278, 1762, 2077, 2149, 1932, 2266, 1643, 1293, 3931, 4308, 3415, 4476, 5051, 5067, 6339, 6314, 5956, 6294, 6225, 5348, 5755, 6000, 5456, 7235, 6475, 6734, 6939, 6367, 7011, 6577, 7273, 7568, 8416, 7825, 7872, 8313, 7570, 8011, 7925, 7718, 7643, 7754, 8899, 9448, 8898, 9466, 8810, 8793, 8717, 9226, 8599, 9100]. States [9814, 10090, 10391, 6919, 10353]. Static [1810, 2254, 8203, 9730, 10518, 10149]. Station [4204, 5515, 7813, 8275, 9431, 7715, 1791, 3092, 5032, 5276, 4705, 9590, 10727, 8918]. Stationary [2268]. Stations [2964, 5286, 5498, 7872, 8530, 10661, 10160, 5265]. Statistical [10541]. Statistics [1293, 5377, 7923, 10389, 10328]. Statue [13, 3077, 4267, 3706, 3809, 5295, 7068, 8931, 9306, 1825, 2363, 8663]. Statues [10778]. Stature [7879]. Status [3471, 5756, 7428, 9129]. Statutes [5396, 7572, 7611, 7569]. Steady [6925]. Steadying [2637]. Steak [4040]. Steam [1149, 1055, 988, 914, 431, 815, 757, 874, 1053, 28, 314, 741, 738, 785, 1011, 1034, 2093, 1515, 2224, 2114, 2041, 1726, 1899, 2088, 2232, 1423, 1440, 2503, 2395, 2677, 2643, 2674, 2688, 2828, 2405, 2800, 2654, 3977, 3401, 4073, 3664, 4163, 4289, 4076, 3251, 3308, 4124, 3284, 3530, 3619, 4107, 3710, 5044, 5071, 4708, 5156, 5310, 4988, 4756, 4595, 5175, 4616, 4772, 5015, 4939, 4907, 5330, 4468, 4846, 4613, 5049, 5833, 6325, 6218, 6279, 5457, 5551, 6315, 6261, 6344, 6148, 5667, 5847, 6342, 7162, 6730]. Steam [7079, 6979, 6422, 7779, 8354, 8126, 8017, 7848, 7722, 7823, 8387, 7957, 8440, 7448, 7691, 8529, 8592, 8594, 8632, 9073, 9156, 8739, 9454, 10787, 9806, 9659, 10675, 10669, 10167, 10701, 9802, 7614]. Steam-Boiler [9659]. Steam-Boilers [9802]. Steam-Engines [10701]. Steam-Propelled [988]. Steamboat [961, 4436, 7657, 8376, 9611]. Steamer [358, 478, 545, 961, 1237, 1377, 2098, 1994, 2817, 2403, 2901, 3622, 3531, 3328, 3518, 3485, 3364, 4920, 5160, 5194, 6350, 5897, 5784, 5908, 5637, 5763, 6264, 7345, 7519, 8146, 9724, 2208, 3097, 8278]. Steamers [1079, 896, 2336, 2686, 4052, 3224, 4951, 4338]. Steamship [1241, 1744, 1867, 1325, 3124, 3252, 3685, 3417, 3294, 3392, 3415, 4456, 4801, 6459, 6934, 7898, 10219, 9111]. Steamships [2685, 6727, 7880, 8428, 9445]. Stearic [3851]. Stearin [4263]. Steel [988, 26, 302, 970, 141, 953, 996, 989, 1995, 1845, 1511, 1773, 2915, 2673, 2686, 2901, 3485, 4008, 4634, 4798, 4550, 4512, 4570, 4865, 5157, 5264, 6131, 6141, 6149, 6633, 7344, 6954, 7449, 8461, 7797, 8388, 8282, 8343, 8452, 8059, 8551, 8593, 8831, 8991, 10294, 10561, 9583, 9702, 9703, 10388, 10765, 10303, 10737, 10375, 4220, 7692]. Steering [1981, 4289, 4940, 9071]. Stein [4597]. Stem [4647]. Stencil [9174]. Stenographic [62]. Stepped [403]. Stereopticon [2306, 7196, 10124]. Stereoscope [6855]. Stereoscopic [3098]. Sterilization [2799]. Stern [1241, 5611]. Sternwheelers [9359]. Sterry [2513]. Stevens [2745]. Stick [1240, 1970, 3865]. Sticks [8741]. Still [2320, 6426, 8357, 9056]. Stilt [3897]. Stirrup [10012]. Stirrup-Pedal [10012]. Stock [2185, 2251, 4173, 3861, 4584, 4774, 8511, 9332, 9824]. Stock-Cars [9824]. Stock-Yard [9332]. Stocker [6419]. Stocking [336]. Stockings [2641, 10235]. Stocks [1624, 9681]. Stockton [2610]. Stoker [1568]. Stomach [8535, 10429]. Stone [4362, 5219, 5202, 671, 550, 2187, 1712, 2405, 3430, 3227, 6222, 6403]. Stoner [4794]. Stones [9539]. Stop [760, 782, 441, 1602, 4667, 5200, 6261, 6084, 9234]. Stopper [59, 6711, 8279, 8493, 10304]. Stopping [900, 1473, 3747, 6122]. Storage [3824, 4305, 4204, 4569, 7317, 8446, 3389]. Store [4207]. Stored [9657]. Stories [10800]. Storing [6205]. Storm [39, 1376, 4117, 5560, 6557, 7514, 2671]. Storms [2808]. Story [5736]. Stove [579, 58, 136, 349, 8243, 7977, 10158, 10279]. Stove-Pipe [349, 10279]. Stovepipe [3109, 8474, 8871]. Stoves [2289, 1875, 4257, 9385]. Straight [2973]. Straightening [4426, 6679]. Straightway [549]. Strainer [954, 4671]. Straining [2418]. Strains [1979]. Straits [10790]. Stranded [4265]. Stranding [6459]. Strange [9755, 3375]. Strategic [9049]. Stratford [10350, 10367]. Straw [58, 5789, 6585, 7303]. Straw-Burning [58]. Straws [4212]. Street [1151, 162, 776, 592, 2246, 2022, 3092, 3289, 3884, 4303, 3284, 3822, 4172, 4488, 4943, 5064, 5118, 5011, 4466, 4428, 5198, 4651, 4905, 6327, 5463, 6266, 5829, 5684, 7116, 6524, 6501, 7597, 8166, 7823, 8949, 8768, 9431, 9366, 9030, 8763, 10158, 9790]. Streets [139, 131, 2913, 7753, 3201]. Strength [6603, 7575, 7492, 8505, 9630, 9705]. Strengthening [5031]. Stress [9973]. Stretcher [66, 475, 3407, 4585]. Strike [4790, 4878, 4900, 8648, 10742]. Strikes [2911, 6294]. Striking [8597]. String [418]. Stringed [2364, 7221]. Stringing [3502]. Strip [10645, 3616]. Strips [6206]. Strong [3866, 8852]. Strop [3910]. Struck [5033, 5495]. Structural [8244]. Structure [7239]. Structures [7344, 7873]. Students [1292]. Studies [1052]. Study [5890, 2994, 9500]. Stuffing [1203, 2090]. Stuffing-Box [2090]. Stump [5726]. Style [10364]. Suabian [2132]. Sub [3134]. Sub-Aquatic [3134]. Submarine [465, 2854, 2618, 3765, 4989, 5962, 6696, 6940, 7799, 9316, 8699, 9058, 8804, 8813, 8516, 9864, 10311, 10747]. Submerged [8074, 55]. Subscribers [1199, 6341, 8911]. Subsidiary [5044]. Subsoil [6023, 6772]. Substitute [1077, 1646, 4617, 4348, 5854]. Subterranean [10313]. Suburban [8109]. Suburbs [7979]. Subway [696, 609, 3413, 10570, 6026, 7058, 8179, 9743]. Subways [1109]. Success [7115, 8597, 10737, 937]. Successful [6515, 8088, 8813, 8675, 10583, 9499, 1576]. Successfully [8631]. Successor [9816]. Suckow [4142]. Suction [3525]. Sudden [2704]. Sufferers [8803]. Sugar [2241, 2116, 2117, 3073, 5757, 5381, 8358, 8380, 8353, 8021, 8101, 8011, 8979, 3353, 9609]. Suggested [4078]. Suggestions [3020, 3224, 8558, 9754, 3961, 1850, 2233]. Suits [3859, 5396, 7310]. Sulky [10568]. Sulphide [6087]. Sulphur [7149]. Sulzer [29]. Sumatra [482, 5634]. Summer [659, 5890, 6047]. Sun [2593, 6188, 2436, 2412, 4591, 5560, 5612, 5787, 6481, 6837, 8726, 10575].

Sunday [7805]. Sundials [3405]. Sunfish [898, 7281]. Sunk [1773]. Sunken [2977, 3432, 5872, 9760]. Sunshade [178]. Superimposed [10690]. Superior [6611]. Supersede [763]. Supplement [801, 836, 1366, 1411, 1618, 1642, 8629, 8911, 8845]. Supplemental [4492]. Supply [1128, 1344, 3370, 5176, 6161, 5823, 6001, 5847, 6416, 7466, 7769, 8989, 9294, 9262, 9302, 9426, 9819, 9675, 8614, 9456]. Supplying [1914]. Support [201, 813, 761, 2183, 1395, 2644, 3071, 2450, 4258, 4232, 3888, 3454, 4469, 4954, 5939, 5594, 8238, 8264, 7716, 10306, 10013]. Supported [7669]. Supporter [6518, 8790]. Supporting [8633, 9931]. Supports [4614]. Suppression [8257]. Supremacy [5778, 7684, 8440, 8823, 9027]. Supreme [334, 5532]. Sure [4237]. Surety [4984]. Surface [6, 2986, 2505, 3791, 6803]. Surgery [9127, 10301]. Surplusage [10115]. Surprise [9271]. Surrounding [3203]. Surveying [8300, 8496]. Surveying-How [8496]. Surveyor [711]. Surveyors [5028]. Susini [3168]. Suspended [5723, 10498]. Suspender [64]. Suspension [1239, 428, 1724, 5311, 5541, 8505, 8605, 7114, 9585]. Sustained [1909, 7780, 8805]. Sutro [9963]. Swamp [4515, 5136, 8702]. Sweden [870, 6297]. Swedish [4449, 8293]. Sweeper [3104, 5118, 5829, 7597]. Sweepings [3435]. Swept [10524]. Swift [642]. Swimmer [2740]. Swimming [2202, 4140]. Swindles [3804]. Swing [12, 1350, 4830, 5231, 7931]. Swinging [897, 5895, 7044, 8541]. Swiss [7078]. Switch [452, 878, 151, 414, 2223, 3728, 7256, 6537, 8422, 7929, 10743, 10779]. Switch-Locking [10743]. Switch-Throwing [10779]. Switzerland [675, 1767, 7667, 10741]. Sword [3163, 7478]. Swordfish [4169]. Sworn [8652]. Sydney [2818]. Synchronizer [5664]. Synthesis [5476]. Syracuse [1182]. System [310, 1020, 528, 41, 340, 551, 151, 696, 255, 172, 1355, 1848, 2147, 2293, 1876, 2058, 1793, 2543, 2414, 3726, 4303, 5044, 4562, 5025, 5237, 5663, 6046, 5896, 5400, 5647, 5654, 6397, 6880, 6763, 7150, 7271, 7006, 6933, 6514, 7619, 7568, 8019, 7473, 7969, 8425, 7428, 8388, 8257, 7358, 7861, 7970, 8068, 9194, 9128, 8980, 9556, 8598, 9792, 10399, 9901, 9912, 9980, 10389, 10210, 3272, 3203, 582, 1169, 414, 1321, 1539, 5234]. Systems [2818, 3164, 3610, 5723, 6497, 6480, 6524, 7394].

T [3291, 6689]. T-Square [6689]. Table [1081, 717, 573, 733, 865, 1618, 1642, 2866, 2829, 3388, 4779, 8023]. Tables [9572]. Tachocycle [4566]. Tacker [4585]. Tackle [291]. Tacks [6500]. Tacony [3088]. Tag [934, 8634]. Tail [574, 5173]. Tailed [5685]. Tailless [5053]. Take [2453]. Taking [1061, 3178, 7450]. Talking [3096, 3069]. Tall [5264, 6864, 8673, 9624, 9815, 9895, 10216]. Tallapoosa [518]. Tallest [3215, 9560]. Tamalpais [9098, 9416]. Tame [8223]. Tandem [3845, 7336]. Tank [5174, 6239, 6805, 10537, 9656]. Tape [233]. Taper [1774]. Tapper [5537]. Tappet [1349]. Tapping [1872, 6624]. Tar [6426]. Target [484, 790, 2918, 4862]. Tattooing [8857]. Taxation [7924, 9387]. Taylor [5963, 6503]. Tea [2692, 4671, 10353]. Teach [3933]. Teachers [5500]. Teaching [3935, 4140]. Team [3867]. Teasing [5010]. Technical [9720].

Technology [2745]. Tectorium [4348]. Telautograph [3482]. Telegraph [863, 1022, 32, 279, 1275, 2531, 3346, 4028, 3847, 5250, 5383, 6936, 7881, 8068, 9980]. Telegraphic [730, 8784]. Telegraphing [2340]. Telegraphs [7048]. Telegraphy [7734, 7874, 8791, 10136, 10535, 9901, 10662, 9949, 9875, 10070, 9950, 10610, 10689, 10467, 10741, 10596]. Telemeter [7062]. Telephone [186, 1818, 2040, 1394, 1439, 3146, 2868, 3344, 4218, 3425, 5234, 4843, 4487, 4904, 4757, 5734, 5630, 6300, 5455, 5400, 5963, 7150, 7207, 6958, 6463, 7780, 8559, 2325, 2882, 3445, 5325]. Telephones [4421, 4465, 4541, 4969]. Telephoning [4768]. Telephony [1327, 3345, 6634]. Telephotography [9855]. Telescope [2257, 6957, 7275, 9833, 10624, 10563]. Telescopes [3684, 5333]. Telescriptor [8492]. Tell [6717]. Telltale [9733]. Tellurian [583, 3480, 4902]. Tellurion [9680]. Temperature [5496, 6256, 6541, 7767]. Temperatures [5825, 6851]. Tempered [4027]. Temperer [7080]. Tempering [6239]. Temple [2438, 2399, 4444]. Temporary [4705]. Ten [7278, 10022, 9866, 9884]. Ten-Inch [10022, 9866, 9884]. Tenders [9272]. Tenement [4938]. Tennessee [1098, 7756, 7659, 7396]. Tennis [3529]. Tennyson [3111]. Tension [644, 1001, 4309, 3791, 6562, 144]. Tent [9385]. Tequendama [7693]. Teredo [4758, 8773]. Teresa [9468, 9328]. Terminal [8530, 9590]. Terminals [6167]. Terminus [1251]. Termites [4867]. Terms [1617]. Terra [8931]. Terracing [2490]. Terre [3214]. Terrestrial [1188, 5910]. Terrible [4253]. Terrific [5765]. Territories [10442]. Terror [3458, 7568, 9017]. Terry [4632]. Tesla [2528, 5564, 9455]. Test [3157, 3402, 3826, 6889, 7974, 7691, 9535, 8675, 9215, 9948, 5483, 9499, 8136]. Tested [8631, 10041]. Tester [7575]. Testimonial [496]. Testing [930, 1969, 2034, 4062, 5983, 7007, 6792, 6897, 7800, 7928, 9291, 10537, 10616, 9845, 9705, 9719, 9914]. Tests [885, 52, 1985, 1745, 2523, 3652, 4957, 6603, 6602, 7927, 8602, 10026, 10662, 9875, 10741]. Tether [8874]. Teutonic [1996, 1994]. Texas [2991, 4512, 2310, 5376, 7253, 7343, 9198, 9852]. Textile [6939, 8261]. Texts [3688]. Thames [7834]. Thanks [6995]. Thatcher [7804]. Thawing [2037, 3914, 8718, 9745]. Theater [1852, 1898, 3011, 4936, 7413, 9101, 10491]. Theatrical [1649]. Their [832, 2703, 7375, 9752, 10778, 2420, 1667, 3325, 8846]. Them [2920]. Theodor [4712]. Theodoric [7279]. Theoretically [3955]. Theory [8368, 4683, 10193]. Therapeutics [3060]. Thermo [5920]. Thermo-Electric [5920]. Thermometer [4281, 5016, 6541]. Thermometers [5103]. Thermophone [6635]. Thief [2887, 6788]. Thill [369, 809, 935, 4258, 5942, 6560, 7716]. Thimble [226]. Thin [9924]. Things [629, 901]. Third [7329, 8063, 7908, 7861, 7970, 8605, 8598, 10210, 4511, 7008, 7981]. Thirteen [8997, 9383, 9798]. Thirteen-Inch [8997, 9383, 9798]. Thirty [2324, 5266, 8453]. Thirty-Foot [2324]. Thirty-Three [5266]. Thistle [5331]. Thistles [920]. Thomas [2513, 4668, 6235, 5877, 9262]. Thomson [10142]. Thorne [5997]. Thoroughbred [4173]. Thoroughfare [8536]. Thought [7552, 10759, 9369]. Thousand [3069, 4488, 8275, 7426, 7538].

Thrashing [2993, 3774, 4792]. Thread [372, 576]. Threader [2182, 5030]. Threading [709, 1080, 2968, 6442, 9758]. Three [2543, 3164, 5266, 4958, 6973, 7595, 7839, 7691, 10427, 10312]. Three-Pounder [10312]. Three-Wire [2543]. Throttle [9309]. Throttle-Lever [9309]. Throw [5718]. Throwing [10779]. Thrust [3685]. Thunder [2808]. Thunderbolt [7978]. Thurston [2780]. Tiber [3410]. Ticket [3178, 9112]. Ticket-Delivering [3178]. Tickets [346]. Tidal [4510, 10720]. Tide [9289]. Tie [607, 867, 1175, 177, 844, 758, 1749, 1464, 1517, 2273, 2594, 3727, 3589, 6673, 8198, 9750, 1829]. Tie-Dog [3589]. Tie-Plates [9750]. Ties [1646, 1369, 9507, 8113]. Tiffany [4523]. Tiffin [5733]. Tiger [4598, 5797]. Tight [5518]. Tightener [1207, 810, 3286, 3261, 8743]. Tightening [10673]. Tilbury [6783]. Tile [573, 3226, 4381]. Tiles [9376]. Tiling [1210]. Timber [6603, 6444, 7499, 8602, 9819]. Timbers [1544]. Time [429, 421, 2108, 1930, 2739, 2484, 2989, 2748, 4304, 3726, 3893, 5646, 6125, 5647, 6188, 6627, 8422, 7733, 8936, 9556, 10423, 10255, 6717, 7735]. Time-Recording [9556]. Time-Stamp [10423]. Timely [10559]. Times [1978, 4284, 6163, 6278]. Tin [2684, 3462, 6164, 6121, 8597]. Tin-Plate [8597]. Tinner [4057]. Tip [7333]. Tipple [2056, 2338]. Tire [5878, 5714, 5440, 6317, 6006, 6691, 7593, 7453, 7951, 7450, 8698, 9072, 8609, 8580, 8743, 10474, 10191, 10050]. Tire-Bell [10191]. Tire-Bolts [10050]. Tireless [2024]. Tires [6980, 8424, 7450, 9013]. Titmouse [2689]. Tivoli [3010]. Tobacco [292, 554, 4620, 5196, 8042, 10326]. Toboggan [2045]. Tocci [2294]. Toeplee [326]. Toilet [1014, 5698]. Token [10380]. Tomb [3186, 7014, 7279, 7717, 7456]. Tombstone [1397]. Ton [1033, 3150, 3710, 4596, 4745, 5966, 7187, 10260]. Tongs [4617]. Tongue [5594, 10306]. Toning [8236]. Tonograph [7814]. Tool [575, 667, 228, 180, 533, 1103, 1593, 2272, 1375, 1569, 2811, 2587, 2337, 3640, 3637, 3571, 4771, 4955, 5219, 4431, 5893, 5497, 5614, 5985, 6812, 6768, 6546, 7166, 8851, 9636, 9758, 10259, 10770, 10256, 3231]. Tool-Users [10256]. Tools [1083, 3182, 4054, 4041, 4057, 4795, 4431, 8239, 9548, 9952, 10056]. Top [554, 813, 259, 1796, 2779, 6169, 10421]. Topaze [6389]. Topeka [8853]. Topographic [4171]. Topographical [9982]. Toponymy [9488]. Tops [6404]. Tor [7454]. Torch [5602]. Torfrida [1055]. Tornado [1038, 1967, 7835]. Torpedo [706, 442, 52, 790, 1773, 1745, 3134, 4254, 3940, 4877, 5158, 4475, 4736, 5089, 5266, 6246, 5342, 5785, 7330, 6696, 6405, 8135, 7468, 7542, 7787, 8427, 9304, 9323, 8747, 8699, 9058, 8804, 9017, 8599, 9864, 9684, 10648, 9925, 8453, 99]. Torpedo-Boat [9017]. Torpedoes [1355, 3840, 5683]. Torsion [3730, 4160]. Tortoises [2296]. Tortugas [8833]. Total [102, 1771, 3832, 3572, 6329, 8794, 6622, 10575]. Touch [5715]. Tour [5179]. Tournament [8784]. Towboat [8684]. Tower [419, 1771, 1713, 2928, 2776, 5295, 5174, 5421, 5982, 6885, 7989, 7623, 8770, 1472, 2314, 4942]. Towing [896, 2359, 3849, 5015]. Towns [2022]. Toxophilite [4974]. Toy [179, 1129, 735, 1694, 3045, 2693, 3311, 4059, 4824, 5852, 7164, 7770, 7645]. Toys [2630, 2645, 3478]. Trace [2624, 7179, 9035]. Trace-Holder [9035]. Trace-Holding [2624]. Track [812, 7418, 7518, 2610]. Tracks [689, 8059]. Tract [8219, 8200]. Traction [936, 2004, 2707, 2571, 3365, 3460, 4608, 4925, 4847, 4511, 4860, 4569, 5682, 6328, 7329, 7394, 7848, 8464, 7670, 7379, 8949, 8554, 9626, 10452, 10278, 10681]. Traction-Wheel [10681]. Trade [661, 2147, 3839, 4096, 4252, 5377, 6989, 7180, 6536, 6990, 7067, 8121, 7755, 7992, 8293, 7990, 8084, 7559, 8012, 7513, 7553, 8823, 8889, 8593, 9505, 9325, 8650, 8970, 9282, 8712, 8558, 8495, 10635, 10638, 9606, 9702, 9844, 10740, 10764, 10323, 9654, 10324, 9754, 10737, 10143, 10511, 10442, 9951, 10373, 9586]. Trade-Mark [10373]. Trade-Marks [9606]. Trademark [9448]. Trademarks [9006]. Trades [6515]. Traffic [4815]. Tragedy [7761, 9168]. Train [1595, 1930, 1793, 2691, 4077, 5200, 6122, 6634, 8063, 8064, 9403, 9308, 10260]. Train-Shed [10260]. Trainer [5983]. Training [4769, 5144, 7437, 8528]. Trainmen [10217]. Trains [33, 1914, 4039]. Trajan [7050]. Tramcars [4613]. Tramway [90, 4085, 5705, 6448, 5552]. Tramways [6385]. Trans [9252, 8962, 9217, 9278, 10379]. Trans-Mississippi [9252, 8962, 9217]. Trans-Siberian [9278, 10379]. Transandine [3428]. Transatlantic [6934, 7898, 8146, 8909, 10196, 10255]. Transcaucasia [9199]. Transcontinental [10285]. Transference [7552]. Transferring [10185]. Transformed [3907]. Transit [693, 3159, 5154, 7254, 6802, 6970, 6579, 6516, 6745, 8377, 8189, 8085, 8256, 7469, 8489, 9865, 10440, 10760]. Transmission [2129, 2422, 3010, 3621, 5573, 6202, 7601]. Transmissions [8215]. Transmitter [5468, 6103, 6300, 5455, 7780, 8559, 2325]. Transmitting [7240]. Transplantation [10182]. Transportation [2149, 2456, 3720, 4747, 5744, 6415, 6496, 6962, 7798, 8515, 9723, 2941, 3394]. Transported [6462]. Transposed [3098]. Transsiberian [6209]. Trap [649, 1232, 1301, 5833, 7025, 6767, 6753, 8557]. Trapeze [6428]. Travelers [10464]. Traveling [2052, 2334, 3623, 3696, 4327, 5229, 5293, 9176]. Travels [7840]. Tread [6317, 9310]. Tread-Sanding [9310]. Treadle [1417]. Treadmill [3199]. Treating [2150, 6042, 7026]. Treatment [1099, 3586, 9386]. Treaty [8540]. Trebert [10028]. Tree [962, 1447, 2800, 3615, 3752, 4800, 5274, 4526, 5034, 5448, 5446, 6267, 6872, 6834, 7885, 7955, 8487, 10336, 1184, 6462]. Tree-Felling [2800]. Trees [1762, 2423, 8389, 7465, 7753]. Trestle [7996]. Treves [7384]. Trial [637, 999, 1120, 3142, 3974, 3675, 3334, 4997, 5049, 5026, 4924, 6225, 6734, 7008, 7718, 8068, 7660, 8427, 7359, 10414, 9925]. Trials [2948, 4917, 4841, 6764, 7448, 8813]. Tribes [9037]. Tribune [6361]. Trick [504, 1197, 3163, 3265, 4572, 7694, 7999, 8990, 10016, 10014, 9846]. Trick-Rider [10014]. Tricks [2627, 2375, 3590, 8566]. Tricks-Old [2375]. Tricolor [4551]. Tricycle [119, 1391, 5026, 4637, 5482, 5367, 5551, 5764, 6153, 7171, 7238, 6388, 10030, 10024]. Tries [8063]. Trifling [7849]. Trigger [2001]. Trilby [6997]. Trimmer [79]. Trinity [7768]. Trio [7426]. Trip [1120, 1996, 2568, 4924, 7824, 7660, 8427, 8319, 10119]. Tripe [8171]. Triple [326, 1263, 1422, 2505, 3521, 4244, 4507, 4382, 4780, 5237]. Triple-Screw [4244]. Triumph [10584]. Triumphal [8, 3973]. Trojan [10773]. Trolley [2320, 3634, 4014, 4562, 4791, 4724, 5968, 5511, 5721, 6900, 6447, 6501, 7329, 6602, 8019, 7397, 8323, 7612, 9509, 9751, 10410, 10415, 10350, 10367, 10390, 9876]. Trolling [3309]. Troops [9009, 8940]. Troopships [9305]. Trophies [6784]. Tropical [6408, 10767]. Tropics [2107, 9108, 10716]. Trotting [2987]. Trousers [6043]. Trouve [2424]. Truant [9843]. Truck [248, 1012, 1807, 4187, 3729, 5808, 6079, 6866, 8359]. Trucks [7100, 10768]. True [9402]. Trumpet [1171]. Trumpeters [3642]. Trundler [2184]. Trunk [193, 2640, 3789, 3234]. Truss [918, 1599, 3593]. Trusts [5590]. Truth [8147]. Tub [1461, 5520]. Tube [869, 2706, 2504, 3230, 4142, 4716, 5807, 6005, 5632, 6262, 6820, 6729, 6637, 7302, 6980, 6806, 6514, 8154, 8262, 7969, 7752, 8260, 8425, 8015, 8282, 8343, 8452, 7359, 8756, 8675, 9985, 9912, 10187]. Tubercular [2523]. Tuberculosis [1261]. Tubes [785, 143, 943, 3193, 5329, 5855, 6400, 7496, 9169, 6789]. Tubular [785, 1648, 5068, 4782, 6418]. Tuck [4844]. Tucker [7550]. Tug [5809, 5369, 8428]. Tugboat [1422]. Tugrin [9981]. Tunicata [6677]. Tuning [185]. Tunnel [1200, 1252, 8085, 7469, 9429, 9865, 10790, 10208, 765, 101, 892, 2077, 1890, 1815, 2293, 3157, 4923, 7878, 7834, 8060, 7638, 9911, 10089, 10195, 10570]. Tunneling [2360]. Tunnels [1039, 10760, 5611]. Turbine [980, 4124, 6736, 8440, 7538, 7691, 8739, 8599, 10305, 10675]. Turbines [8188, 9516]. Turbinia [8013]. Turkish [7863]. Turn [8298, 9228]. Turner [406, 6544, 7784]. Turning [4260]. Turret [2852, 6045, 6863, 6867, 7318, 10690]. Turreted [860]. Turrets [5293, 10612, 9852, 9908]. Turtle [3030, 6890]. Turtles [6173]. Tusayan [8409]. Tuyères [501]. Twain [5448]. Twelve [2546, 3632, 4901, 7764]. Twelve-Wheeled [2546]. Twentieth [10471]. Twenty [6446, 7187, 8532]. Twenty-Eight [8532]. Twenty-Five [6446, 7187]. Twin [2211, 2098, 2468, 3417, 3946, 3702, 5595, 7758, 7880, 7555]. Twin-Jaw [5595]. Twine [8959]. Twins [2294, 3216, 3075]. Twist [4024]. Two [850, 111, 1806, 3362, 5218, 5308, 4637, 6347, 7300, 7942, 8128, 7538, 9234, 8655, 9110, 10381, 10165, 9822, 9099, 7413]. Two-Headed [7300]. Two-Part [10165]. Tyndall [4297]. Type [1240, 364, 220, 2031, 1944, 1735, 7668, 8402, 7762, 8631, 9153, 10237]. Type-Writer [364, 1735]. Types [6563, 7839, 10670]. Typesetting [763, 4360, 5997]. Typewriter [3007, 3254, 5252, 4713, 4427, 5761, 5383, 6535, 7668, 8449, 7785, 10101, 9977]. Typewriting [866, 2228, 4729, 6769]. Typical [10194, 9500, 7046]. Tyrotoxicon [861].

U. [3519]. U.S. [996, 1436, 1485, 1641, 1772, 4581, 5080, 5467]. U.S.S. [3965].

Umbrella [847, 712, 789, 8130, 8832]. Umbria [3252, 3294, 3328]. Unclean [3285]. Uncolored [2855]. Under-Trolley [9876]. Underground [1134, 2814, 4724, 4651, 6501, 7329, 6602, 8019, 7612, 8594, 8715, 9509, 10410, 10619, 9657]. Undertaking [6182]. Undeveloped [9248]. Unicycle [1538, 3566, 4364, 4821, 9433, 1347]. Uniform [9330, 9350]. Uniforms [9050]. Union [2531, 2748, 8736, 268, 6969, 10205]. Unit [6962]. United [6919, 849, 1169, 765, 302, 387, 1089, 2278, 1762, 2077, 2149, 1932, 1421, 2266, 1643, 1293, 3931, 4308, 3415, 4476, 5051, 5143, 5067, 6339, 6314, 5956, 5580, 6294, 6225, 5348, 5755, 6342, 6000, 5456, 7235, 7278, 6879, 6475, 6734, 6939, 6367, 7011, 6577, 7273, 7568, 8416, 8072, 7824, 7825, 7872, 7569, 8234, 8313, 7570, 8011, 7925, 7542, 7718, 7643, 7754, 7777, 8899, 9448, 8898, 9466, 9527, 8683, 8581, 8810, 8724, 8793, 9315, 8717, 9129, 9294, 9226, 8599, 8930, 9100, 8713, 9814, 10090, 10398, 9908, 10220, 9867, 10391]. United [99, 10353]. Unites [8597]. Universal [1523, 3456, 4971, 10242, 10043]. Universally [8363]. University [2655, 5679, 5445, 4633, 976, 2152, 7188, 7800, 8771, 9396, 8682, 8856, 9473, 9914]. unknown [6952]. Unloading [6238, 6869]. Unscientific [2899]. Unsuccessful [8063]. Unsuspected [9526]. Upheaval [10509]. Upon [1344, 1877, 4711, 4779, 9293]. Upton [7182]. Urban [1689]. Urchins [677]. Urgent [8318, 8574]. Urgently [8334]. Use [911, 292, 901, 1248, 1296, 1943, 2257, 2740, 2472, 2726, 2829, 3213, 4207, 4303, 5170, 5134, 4925, 5102, 4468, 6295, 5510, 7207, 7598, 8321, 7977, 7721, 8244, 9016, 9376, 10164, 10423, 236, 722, 1462, 4880, 5236, 6441, 375, 787]. Used [6833, 9719, 1734]. Useful [1083, 1535, 1581, 1612, 1637, 1662, 2868, 9470, 8446]. Usefulness [7638]. Users [10256]. Uses [811, 2725, 10715, 10643, 9975, 9569, 246, 2420]. Using [10465]. Utilization [1056, 2886, 3462, 4646, 6277, 8071]. Utilizing [1107, 3043, 2473, 5139, 7828].

V [1980, 9419]. Va [9539]. Vacant [7925]. Vaccination [5629, 6791]. Vacuum [6514, 7969, 8994, 8982, 6871, 7066]. Valentine [5037, 6109]. Valley [10487, 1897, 8300]. Valuable [8740, 9407]. Value [3861, 3447, 5803, 6329, 8125, 9193, 10136]. Values [7548, 9774]. Valve [1000, 431, 891, 481, 549, 1729, 1491, 2232, 2395, 3041, 2793, 2794, 3429, 3977, 4073, 3937, 4003, 4163, 3525, 3956, 3350, 4383, 4495, 4708, 4430, 6186, 5498, 5535, 6006, 6728, 7236, 8104, 7494, 8127, 7928, 9291, 9013, 8873, 9589, 10774, 10701]. Valve-Gear [9589]. Valves [4107]. Vander [5577]. Vanderbilt [6449]. Vanishing [2063]. Vanity [6735]. Vapor [7026, 8001, 8994]. Vapor-Condenser [8994]. Vaporization [7496]. Vaporizer [65]. Variable [4648, 6255, 5703, 6380, 9759]. Variation [7711]. Varied [2420]. Variety [3266, 4431]. Various [2251, 6563]. Vark [3795]. Varks [3317]. Varying [8744]. Vase [4009, 4527, 8880, 3942]. Vast [10389]. Vatican [9592]. Vault [974, 5398]. Vaulting [1541]. Vaults [4895, 5217, 10183]. Vegetable [3051]. Vegetation [5081, 6891]. Vehicle [813, 526, 320, 1150, 39, 1518, 3285, 5171, 5419, 6322, 5714, 7045, 6565, 8216, 7433, 9014, 8743, 10025, 9571, 10474, 10544, 9907, 10285]. Vehicle-Tire [10474]. Vehicles [77, 1371, 1493, 1564, 1948, 3769, 4823, 5173, 10008, 10007]. Vein [6066]. Velocipede [120, 1178, 1188, 1677, 1893, 1419, 3888, 4056, 4993, 5101, 6008, 6882, 6902, 7153, 7648]. Velocipedes [2111, 4082]. Velocity [2706, 10416]. Vending [6520, 7575]. Veneer [5575]. Venetian [5297]. Venezuelan [7252]. Venomous [505, 4388, 2079, 2120]. Ventilated [2511, 8042]. Ventilating [1205, 696, 2884, 3054]. Ventilator [807, 450, 471, 2199, 3109, 4278, 5307, 6102, 6854, 7042, 6463, 9138, 9882]. Venus [8612]. Veragua [3543]. Verde [8907, 8948, 10775]. Verdict [9656]. Vergne [6305, 6753]. Vermont [8896]. Version [7466]. Versus [5916, 10402, 9749, 10760]. Vertical [2362, 5555, 5565]. Very [4618, 7517, 5499]. Vespucci [8749]. Vessel [862, 603, 746, 1845, 1671, 3181, 2849, 3262, 3747, 4489, 4386, 6108, 6785, 7731, 9747, 10311]. Vessels [1076, 59, 953, 930, 2303, 2637, 2862, 2654, 3591, 3432, 3906, 4631, 5228, 5872, 5718, 6868, 7416, 8852, 9760, 9070, 9908]. Vestibuled [1009]. Vesuvius [2285, 1745, 3334, 10587]. Vettii [8224]. Viaduct [592, 1280, 4905, 7518, 8636, 2850]. Vibration [7580]. Vibrations [2654]. Vibratory [3060]. Vicenza [9101]. Vickers [9417]. Victimizing [6621]. Victor [768, 315, 1666, 5523, 10029]. Victoria [159, 2591, 3766, 5766, 6883, 8222, 9474, 3806]. Victorious [9231]. Victory [7298, 7930, 8917]. Vienna [1734, 3508, 3724, 9406]. View [3917, 3985, 4015, 4035, 4126, 8781]. Viewing [3271]. Views [9416]. Vigilant [4060, 4125]. Viking [3758, 3638, 3921]. Village [3920, 5694]. Ville [5084]. Villerville [4265]. Vinci [8835]. Vine [6240]. Vinegar [5559]. Vineyard [7122]. Violation [9842]. Violent [2412]. Violin [5444]. Virginia [959, 2098, 3211]. Viscosimeter [9517]. Vise [957, 3913, 7166]. Visible [5761]. Vision [3275]. Visit [989, 1834, 7065, 9134, 2835, 5620]. Visitations [1918]. Visited [3027]. Visiting [3921]. Vital [1293]. Vitrified [10564]. Viviscope [6856]. Vivisection [4813]. Vlacke [3795]. Volcano [2831, 3373]. Volcanoes [9020]. Volson [7943]. Volt [158, 1522]. Volta [10263]. Voltmeter [10100]. Voyage [2327, 5801]. vs [8368, 1401]. Vulcanizer [4282]. Vulnerable [8758].

W [2137, 2833, 4061, 4017, 9806]. Wabash [10599]. Wabble [647]. Waddell [4569]. Wage [10253]. Wagon [641, 2225, 4512, 4846, 5892, 5594, 6130, 6402, 8298, 8059, 8323, 9534, 10795, 10306, 10129]. Wagon-Jack [9534]. Wagon-Tongue [10306]. Wagons [716, 317, 4016, 3769, 4391, 4847, 10649, 10250]. Waiter [417, 571, 6629, 7183]. Waldorf [9912]. Wales [5439]. Walhalla [7415]. Walker [7478]. Walkers [3897]. Walking [650, 10568, 10264]. Wall [610, 8891]. Wallingford [1291]. Walls [1567, 6522]. Walmsley [6479]. Waltham [4370, 10373]. want [2130]. Wanted [544, 1458, 2465, 3884, 3953, 5005]. War [860, 92, 1076, 746, 42, 804, 921, 953, 930, 1513, 2290, 1643, 3158, 2403, 2394, 3531, 3674, 4323, 3458, 3518, 3765, 3267, 3766, 3675, 4244, 3731, 4877, 4366, 4856, 4924, 5067, 4935, 5294, 5812, 5369, 5376, 5405, 5785, 5464, 7048, 8396, 8164, 7397, 8428, 7733, 8996, 9131, 9303, 8761, 8735, 9070, 7735, 7742, 9006, 9244, 9840]. Warburton [3941]. Ward [1825, 9209]. Ware [4127]. Warfare [8533, 9068, 9190, 8699]. Waring [9408]. Warmers [2503]. Warner [1599]. Warning [7687, 8445]. Warnings [8672]. Warp [6693]. Warship [1772, 1853]. Warships [7681, 9414, 9246, 9090, 8757, 9253, 10690, 10039, 9923, 10187]. Wart [3795]. Was [3294]. Wash [739, 1461, 3247, 5346]. Washburne [6067]. Washer [320, 319, 1733, 3179, 3426]. Washing [2117, 2396, 3701, 6918, 7258, 8360, 8765]. Washington [7232, 8683, 2059, 2062, 1272, 2636, 2404, 4009, 5922, 6584, 6790, 7435, 8065, 7841, 8582, 9454, 8895, 9610, 7130, 9733]. Washtubs [626]. Waste [4646, 4953]. Wastes [8111]. Watch [4054, 4004, 4370, 5699, 6067, 9756, 10373]. Watches [4491, 4675, 5091, 5382]. Watchmaker [3231, 5648]. Watchman [2989, 6627]. Water [886, 1217, 1128, 689, 1205, 1056, 1060, 110, 419, 577, 28, 314, 1013, 980, 1034, 1107, 476, 1962, 1716, 1652, 1475, 1732, 2095, 1627, 1963, 2274, 1768, 1947, 1344, 2224, 2115, 1914, 2020, 2966, 2891, 3041, 2832, 2928, 2542, 2549, 2510, 2309, 2527, 3113, 2514, 2813, 2470, 3054, 3091, 2799, 2473, 2886, 3371, 3501, 3248, 3569, 3770, 3874, 3268, 3975, 3655, 3619, 3350, 4123, 3907, 4857, 4469, 5008, 5029, 4528, 5290, 4883, 5139, 5176, 5478, 6165, 5823, 5680, 6202, 6063, 5632, 6188, 5985, 6047, 5518, 6008, 5524]. Water [6833, 6400, 6820, 6480, 7117, 6809, 6885, 6519, 6736, 8315, 7623, 7752, 7713, 7769, 7359, 7381, 9352, 9561, 8614, 8756, 9555, 9196, 9056, 9456, 8720, 8659, 9262, 10125, 10232, 10651, 10102, 10517, 10305, 10280, 10422, 10328, 10542, 9661, 9802, 9675, 175, 10187, 8018, 2836]. Water-Distilling [9555]. Water-Feeding [9661]. Water-Heater [9802]. Water-Jacketed [8659]. Water-Still [9056]. Water-Tight [5518]. Water-Tube [10187]. Water-Ways [1652]. Water-Wheel [10305]. Watering [2913, 4774]. Waters [10712]. Waterspout [7122]. Watervliet [4745]. Waterworks [10620]. Watson [2572]. Wave [130, 7022, 8877, 9437, 9949, 10126]. Wave-Propelled [9437]. Waved [1299]. Waverly [10129]. Waves [5777, 4850, 2455]. Wax [3658, 9776]. Way [1264, 9272, 9472, 9660, 9988]. Wayne [5108]. Ways [900, 1652, 10465]. Weak [8950]. Wealth [9034]. Wearing [7927]. Weather [5456, 6602, 7995, 10368, 2015, 8506, 8672, 10791, 9725]. Weaver [651]. Weaving [5693]. Webb [4497]. Wedderburn [8356, 8232]. Wedge [9924]. Weed [1784, 2366, 4040, 5524]. Weeder [365]. Weeds [9607]. Weehawken [2019, 2169, 3894]. Week [3089]. Weekly [519]. Weighing [4970, 8657, 8974, 10420]. Weight [1913, 4650, 7274, 7879, 9206, 9346, 9424, 9050, 10391]. Weights [89, 1089, 6164, 7006]. Weir [8382]. Weird [2784]. Welded [7395]. Welding [4411, 4959, 5458, 6149, 6441]. Welds [7492]. Weldsless [5248]. Well [819, 1051, 1813, 1648, 5613, 5600, 7497, 10498, 416]. Wells [211, 3975, 4545, 6504, 8384, 7979, 7149, 9449]. Welsbach [8805, 10792]. Wenham [14]. Went [9115]. Werner [3195, 3249]. West [5684, 7826, 3027, 5654, 10802, 3211]. Westcott [4731]. Western [2531, 2748, 4192, 10218]. Westinghouse [172, 568, 414]. Westminster [9776]. Westphalica [8246]. Westville [6208]. Wethey [6786]. Weyde [5577]. Whale [1853, 4265, 8037, 7435, 7887]. Whaleback [1845, 2208]. Wharf [1033]. Wheat [6183, 7080, 8191, 9302, 9426, 10274, 9199]. Wheel [253, 367, 526, 577, 980, 1057, 1518, 1732, 2004, 1766, 2115, 1874, 2076, 2514, 3777, 3748, 4082, 3770, 3268, 5008, 4775, 5215, 4860, 5916, 6425, 7045, 7043, 6736, 7381, 9014, 8743, 9573, 10305, 9907, 10681, 7167, 3985]. Wheeled [2546, 7764, 9452]. Wheeler [4403, 9180]. Wheeling [6752]. Wheels [811, 220, 2180, 5714, 6706, 6787, 6565, 10381, 9661]. Which [5633, 9413, 10735, 3509]. Whiffletree [2624]. Whisk [4956]. Whistles [2828]. Whitaker [4650]. White [4120, 4867, 5685, 10716, 1994, 5559, 7255, 6368, 9724, 9688, 9910]. White-Tailed [5685]. Whitney [2304, 2304]. Widah [4350]. Widow [4709]. Wilcox [3882]. Wild [10764, 9037, 2093]. Wilhelm [8442, 8259]. Will [8125, 9413]. William [10717, 196, 2139, 3077, 4037, 6488, 8246, 4038]. Williams [6038, 5245]. Wilson [9290]. Wilzin [427]. Wimshurst [1810, 4752]. Winby [3650]. Winch [6672]. Wind [6040, 6277, 7043, 6557, 9261, 10416, 10524, 1325]. Wind-Swept [10524]. Winder [4647]. Winding [3150, 3679, 3844]. Windlass [422]. Windmill [977, 1213, 2308, 4710, 4618, 5852, 6260, 6708, 7240]. Windmills [5380]. Window [204, 1233, 476, 2250, 1490, 2952, 3427, 3772, 3686, 4348, 5851, 6244, 5894, 5745, 7781, 7451, 7714, 9265, 9230, 9803]. Window-Locking [9803]. Window-Sash-Lift [9265]. Windows [910, 710]. Windsor [9922]. Windt [7760]. Wine [4075, 3997, 7220, 9097, 9657, 9805]. Wineglasses [9123]. Wines [2638]. Winged [7930]. Wings [7580]. Winnebago [593]. Winter [1524, 3114, 9016]. Winton [7998, 8915, 10018]. Wiper [4447]. Wire [1040, 66, 810, 475, 975, 537, 629, 778, 779, 1856, 2184, 2543, 2694, 3088, 3407, 3825, 3730, 4160, 4014, 4085, 4822, 6976, 7278, 7273, 8991, 9703, 9948, 3115, 3601]. Wireless [7874, 8791, 10136, 9901, 9875, 10070, 9950, 10610, 10689, 10467, 10741, 10596]. Wires [143, 1376, 2340, 4028, 4303, 6538, 6562, 10673]. Wis [3954]. Wisconsin [2313, 10771, 3754]. Wise [9173]. Wish [3719, 10759]. Without [7044, 7275, 7450, 2472, 2340, 4303, 4753, 4752, 5804, 7924, 8535, 9234, 2325, 10518, 8144]. Withstand [3225]. Woehnelt [9902]. Wolf [2819]. Wolsey [9932]. Woman [3394, 3987]. Womans [6303]. Women [8408]. Won [4125]. Wonderful [5427]. Wonderfully [4004]. Wonders [386, 9521]. Wood [1154, 560, 2036, 1256, 1320, 1646, 2449, 2972, 3406, 7237, 7315, 9507, 9074, 9637, 10461, 9846, 7943]. Wood-Workers [1154]. Wood-Working [1256, 9074, 9637]. Wooden [945, 1369, 3181, 2321, 4328, 7418, 9349, 9749]. Woodpecker [5484]. Woods [9348, 10007]. Woodworking [7640, 8851]. Wool [864, 2545, 4413, 9011]. Wool-Drier [9011]. Woolen [5959, 6920].

Worcester [4120, 1258]. word [4729]. Worden [7596]. Words [4050]. Work [1076, 1210, 585, 1031, 895, 865, 1124, 1561, 1725, 3723, 3860, 3848, 3526, 3260, 4279, 5287, 4986, 4713, 4816, 4365, 4470, 4582, 5297, 4546, 5286, 5760, 6088, 6298, 7333, 6441, 7739, 8166, 7952, 8628, 8846, 8773, 9384, 10214, 10443, 9725, 7412, 10149]. Work-Holding [865]. Worker [202, 878, 1102]. Workers [153, 1154]. Working [1256, 2685, 5802, 6379, 7311, 9074, 9637]. Workingmen [2035]. Works [1234, 1209, 1767, 1947, 2915, 3131, 3091, 2405, 6236, 5733, 6915, 6969, 6474, 9776, 627, 372, 1995, 1444, 3710, 6063, 8564]. Workshop [4318]. World [244, 660, 745, 26, 1837, 2052, 1815, 2266, 2687, 2602, 2754, 3531, 3587, 3446, 5295, 4646, 6108, 6182, 8064, 7688, 7519, 7827, 7448, 8969, 8532, 9483, 9403, 9288, 8825, 9476, 9519, 9576, 9677, 10003, 9641, 9782, 9938, 10107, 10284, 10499, 10209, 10797, 51, 1031, 60, 221, 1, 195, 1620, 2297, 1792, 2229, 2941, 2825, 2328, 2776, 3090, 2975, 3059, 2363, 4155, 4173, 3870, 3844, 3962, 3756, 3520, 3542, 4138, 4157, 4183, 3565, 3611, 3653, 3676, 3699, 3721, 3743, 3767, 3786, 3805, 3823, 3843, 3863, 3908, 3932]. World [3964, 3976, 4000, 4026, 4053, 4079, 4098, 4118, 3608, 4145, 3677, 3563, 3633, 3698, 4104, 4122, 3652, 3305, 3564, 3609, 3636, 3882, 3917, 3944, 3963, 3973, 3984, 3985, 3987, 4015, 4035, 4064, 4109, 4120, 4123, 4126, 4193, 4217, 4041, 4057, 3841, 4370, 6121, 9302, 9426, 10156, 4127]. Worlds [3260, 3885]. Worms [3406, 8879]. Worth [5911, 5980, 5999, 5957]. Wound [7278]. Wounded [6588, 9096]. Wrapper [3144]. Wreck [4097, 4276, 9160]. Wrecked [9253]. Wrecking [8516]. Wrecking-Boat [8516]. Wrench [67, 397, 868, 1130, 1155, 1492, 1563, 1714, 1892, 2000, 2380, 2724, 2397, 3594, 3366, 3431, 3704, 4320, 5213, 4610, 5460, 5743, 7181, 7314, 7493, 7737, 7362, 7378, 7515, 8472, 7431, 8698, 8957, 9371, 9055, 8935, 10050]. Wrenches [9432, 9825]. Wrestling [6227]. Wringer [1036]. Wrinkle [721]. Writer [364, 1944, 1735]. Writing [9329, 10364, 9334, 9358, 9374, 9390, 9419]. Written [4713]. Wrought [3215, 5287, 5855, 7449]. Wurdemann [10307].

X [6588, 6638, 6906, 6637, 7302, 6483, 6548, 8048, 9132, 9632, 7883, 10744, 10149]. X-Ray [10744, 10149]. Xiphopages [10574]. XVII [6904].

Yacht [1055, 1108, 2093, 4060, 4996, 5722, 6065, 6111, 6027, 8848, 10321, 10519, 10546, 10557, 10169, 10161, 10467, 10596, 9976, 10712, 6089]. Yachting [10622]. Yachts [2114, 2088, 4110, 7957]. Yale [5069]. Yalu [5364, 6019]. Yard [9332, 556, 1132, 581, 1511, 3696, 7024, 6790, 6444, 7786, 7811, 8102, 8478, 7499, 8905, 8916, 8683, 8603, 9454, 8895, 10462, 9753]. Yards [2185, 5324]. Yarn [1001, 5959]. Yarrow [6820]. Yaryan [1753]. Year [4606, 6360, 7360, 8466, 8486, 10323, 10276, 9567, 5125, 4829, 3115, 8538]. Years [6957, 6941, 6950, 6959, 6953, 6946, 10519, 10546]. Yellowstone [5837]. Yerkes [3291, 5445, 7850, 7631]. Yillimani [9284]. Yoke [1999, 2113, 2970]. York [1148, 698, 17, 592, 1978, 1468, 1932, 3159, 3089, 2531, 3884, 3370, 4256, 3755, 4984, 4799, 4825, 5273, 4593, 4673, 4705, 4773, 5313, 4510, 5176, 5896, 5539, 5669, 5869, 5541, 5500, 6436, 7024, 7254, 6387, 6612, 6461, 6809, 7039, 7147, 6970, 6556, 6516, 6745, 6444, 6415, 6496, 6501, 8377, 8304, 7394, 8314, 8111, 8450, 7897, 7447, 7717, 7455, 7476, 7968, 7798, 7501, 7672, 7512, 7741, 7518, 8260, 8274, 8425, 8085, 8015, 7981, 7908, 8071, 7612, 7769, 8614, 8537, 8552, 8495, 9626, 10452, 9602, 9829, 10064, 10661, 9753]. York [9843, 10664, 10303, 8869, 5620, 7138, 139, 1156, 970, 151, 87, 51, 1039, 428, 60, 1200, 519, 2204, 2095, 1545, 1344, 1878, 2696, 2619, 3131, 2728, 2977, 3096, 4049, 3555, 3631, 3839, 3575, 3539, 3548, 3830, 3284, 3392, 3529, 3675, 3731, 4453, 4477, 4347, 4936, 4866, 5296, 5153, 5119, 5246, 4815, 4569, 5406, 5681, 5436, 5475, 5416, 6315, 6384, 6666, 6705, 6750, 6803, 6579, 6524, 7008, 6602, 7487, 8258, 7434, 8033, 8189, 8219, 7989, 8019, 7876, 8050, 8200, 8362, 7741, 7469, 8166, 7823, 7753, 8628, 9351, 9290]. York [8554, 8491, 9431, 8910, 8489, 9365, 8636, 9366, 9552, 8505, 10119, 10609, 10726, 10614, 10764, 9865, 9762, 10642, 10210, 9609]. Yorker [1468]. Young [9367, 10698]. Yukiahs [7934]. Yukon [9910, 8300, 9359, 8692].

Zapon [1077]. Zea [6677]. Zealand [10587, 4767, 9337]. Zebras [5144]. Zeiss [9153]. Zero [5456]. Zerograph [8767]. Zinc [4349]. Zoological [998, 2921, 4013, 4781, 4515, 5503, 5786, 5634, 5685, 6487, 10664, 10727, 1329, 5220, 4908, 6173, 6999, 9633, 9927, 8918]. Zoology [8641, 8660].

References

Anonymous:1890:WFa

[1] Anonymous. The World’s Fair of 1892. Scientific American, 62(1):2, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

       (electronic). URL http://www.nature.com/scientificamerican/
       journal/v62/n1/pdf/scientificamerican01041890-2.pdf.

Anonymous:1890:ILCa

[2] Anonymous. The incandescent light controversy. Scientific American, 62(1):2–3, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

       scientificamerican/journal/v62/n1/pdf/scientificamerican01041890-
       2a.pdf.

Anonymous:1890:ES

[3] Anonymous. “Experimental Science”. Scientific American, 62(1):3, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

       (electronic). URL http://www.nature.com/scientificamerican/
       journal/v62/n1/pdf/scientificamerican01041890-3.pdf.

Anonymous:1890:IDSa

[4] Anonymous. An improved dredging scraper. Scientific American, 62(1):4, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n1/pdf/scientificamerican01041890-
4a.pdf.

Anonymous:1890:NPDa

[5] Anonymous. New plating dynamo. Scientific American, 62(1):4, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-4b.pdf.

Anonymous:1890:NSG

[6] Anonymous. The national surface guard for railroads. Scientific American, 62(1):4, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n1/pdf/scientificamerican01041890-
4c.pdf.

Anonymous:1890:PEM

[7] Anonymous. The Paris Exposition Medal. Scientific American, 62(1):4, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-4.pdf.

Anonymous:1890:BTA

[8] Anonymous. Brooklyn’s triumphal arch. Scientific American, 62(1):5, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-5a.pdf.

Anonymous:1890:HLC

[9] Anonymous. A hand lifter for coupling links. Scientific American, 62(1):5, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n1/pdf/scientificamerican01041890-
5.pdf.

Anonymous:1890:ISa

[10] Anonymous. An improved saw. Scientific American, 62(1):5, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-5b.pdf.

Anonymous:1890:IFLa

[11] Anonymous. An improved flash-light lamp. Scientific American, 62(1):5–6, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n1/pdf/scientificamerican01041890-
5c.pdf.

Anonymous:1890:NSC

[12] Anonymous. New swing cut-off saw. Scientific American, 62(1):6, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-6b.pdf.

Anonymous:1890:SGD

[13] Anonymous. Statue giving a double image. Scientific American, 62(1):6, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-6.pdf.

Anonymous:1890:WGL

[14] Anonymous. The Wenham gas lamp. Scientific American, 62(1):6, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-6a.pdf.

Anonymous:1890:Ca

[15] Anonymous. Correspondence. Scientific American, 62(1):7, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-7.pdf.

Anonymous:1890:CCMa

[16] Anonymous. Concave celestial maps. Scientific American, 62(1):7, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-7a.pdf.

Anonymous:1890:NYE

[17] Anonymous. The New York Exhibition of 1892. Scientific American, 62(1):8, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n1/pdf/scientificamerican01041890-
8.pdf.

Anonymous:1890:EPPa

[18] Anonymous. Electric phenomena produced by solar radiations. Scientific American, 62(1):9, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n1/pdf/scientificamerican01041890-
9.pdf.

Anonymous:1890:IHb

[19] Anonymous. The Inventor’s head. Scientific American, 62(1):9, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-9a.pdf.

Anonymous:1890:BPa

[20] Anonymous. Business and personal. Scientific American, 62(1):10, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-10.pdf.

Anonymous:1890:Ia

[21] Anonymous. To inventors. Scientific American, 62(1):10, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-10b.pdf.

Anonymous:1890:NQa

[22] Anonymous. Notes and queries. Scientific American, 62(1):10, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-10a.pdf.

Anonymous:1890:IIa

[23] Anonymous. Index of inventions. Scientific American, 62(1):10–13, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-10c.pdf.

Anonymous:1890:PNPa

[24] Anonymous. Patent notes: Patents. Scientific American, 62(1):14–16, January 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n1/pdf/scientificamerican01041890-14.pdf.

Anonymous:1890:FB

[25] Anonymous. The Forth Bridge. Scientific American, 62(2):17, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n2/pdf/scientificamerican01111890-17.pdf.

Anonymous:1890:ISO

[26] Anonymous. In iron and steel our country now leads the world. Scientific American, 62(2):18, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
18.pdf.

Anonymous:1890:SAM

[27] Anonymous. Second annual meeting of the American geological society. Scientific American, 62(2):19, January 11, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n2/pdf/scientificamerican01111890-19.pdf.

Anonymous:1890:LWAa

[28] Anonymous. A low water alarm for steam boilers. Scientific American, 62(2):20, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
20.pdf.

Anonymous:1890:SEP

[29] Anonymous. The Sulzer engine at the Paris Exhibition. Scientific American, 62(2):20, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
20a.pdf.

Anonymous:1890:IBy

[30] Anonymous. An improved boiler. Scientific American, 62(2): 20–21, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
20b.pdf.

Anonymous:1890:IMPa

[31] Anonymous. An improved middlings purifier. Scientific American, 62(2):21, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
21.pdf.

Anonymous:1890:NTR

[32] Anonymous. A new telegraph repeater. Scientific American, 62 (2):21, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
21a.pdf.

Anonymous:1890:SPRa

[33] Anonymous. A safety pilot for railway trains. Scientific American, 62(2):21, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
21b.pdf.

Anonymous:1890:CBK

[34] Anonymous. A continuous brick kiln and drier. Scientific American, 62(2):21–22, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
21c.pdf.

Anonymous:1890:IECa

[35] Anonymous. An improved elastic chain. Scientific American, 62(2):22, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
22a.pdf.

Anonymous:1890:IFN

[36] Anonymous. An improved finger nail knife. Scientific American, 62(2):22, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
22c.pdf.

Anonymous:1890:IPCa

[37] Anonymous. An improved paper clip. Scientific American, 62 (2):22, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
22d.pdf.

Anonymous:1890:SBH

[38] Anonymous. A soap and brush holder. Scientific American, 62 (2):22, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
22.pdf.

Anonymous:1890:SCV

[39] Anonymous. A storm coat and vehicle apron. Scientific American, 62(2):22, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
22b.pdf.

Anonymous:1890:PP

[40] Anonymous. Plumbago packing. Scientific American, 62(2):22–23, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n2/pdf/scientificamerican01111890-22e.pdf.

Anonymous:1890:ICHa

[41] Anonymous. An improved car-heating system. Scientific American, 62(2):25, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
25.pdf.

Anonymous:1890:NBW

[42] Anonymous. The new British war ship Blake, the fastest and most formidable cruiser afloat. Scientific American, 62(2):25–26, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n2/pdf/scientificamerican01111890-25a.pdf.

Anonymous:1890:IAF

[43] Anonymous. An improved anti-friction clutch. Scientific American, 62(2):26, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
26.pdf.

Anonymous:1890:ILCb

[44] Anonymous. An improved leather channeling machine. Scientific American, 62(2):26, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
26a.pdf.

Anonymous:1890:RPIa

[45] Anonymous. Recently patented inventions. Scientific American, 62(2):27, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n2/pdf/scientificamerican01111890-
27.pdf.

Anonymous:1890:BPb

[46] Anonymous. Business and personal. Scientific American, 62(2):28, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n2/pdf/scientificamerican01111890-28.pdf.

Anonymous:1890:NQb

[47] Anonymous. Notes and queries. Scientific American, 62(2):28–29, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n2/pdf/scientificamerican01111890-28a.pdf.

Anonymous:1890:Ic

[48] Anonymous. To inventors. Scientific American, 62(2):29, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n2/pdf/scientificamerican01111890-29.pdf.

Anonymous:1890:IIb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n2/pdf/scientificamerican01111890-29a.pdf.

Anonymous:1890:PNPb

[50] Anonymous. Patent notes: Patents. Scientific American, 62(2):32, January 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1890:NYS

[51] Anonymous. New York as a site for the World’s Fair. Scientific American, 62(3):34, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
34.pdf.

Anonymous:1890:TBT

[52] Anonymous. Torpedo boat tests. Scientific American, 62(3):34, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-34a.pdf.

Anonymous:1890:N

[53] Anonymous. Notes. Scientific American, 62(3):34–35, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-34b.pdf.

Anonymous:1890:PNa

[54] Anonymous. Photographic notes. Scientific American, 62(3):35, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-35.pdf.

Anonymous:1890:SRO

[55] Anonymous. The Submerged Railway of Onton on the coast of Spain. Scientific American, 62(3):36, January 18, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n3/pdf/scientificamerican01181890-36.pdf.

Anonymous:1890:ICEa

[56] Anonymous. An improved carpet exhibitor. Scientific American, 62(3):37, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
37a.pdf.

Anonymous:1890:IPLa

[57] Anonymous. Improved portable lamp. Scientific American, 62 (3):37, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
37.pdf.

Anonymous:1890:ISBa

[58] Anonymous. An improved straw-burning stove. Scientific American, 62(3):37, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
37c.pdf.

Anonymous:1890:LSV

[59] Anonymous. A leak stopper for vessels. Scientific American, 62(3):37, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
37b.pdf.

Anonymous:1890:SWF

[60] Anonymous. The site for the World’s Fair in New York. Scientific American, 62(3):38–40, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
38.pdf.

Anonymous:1890:PEL

[61] Anonymous. Progress of electric lighting in London. Scientific American, 62(3):40–41, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
40.pdf.

Anonymous:1890:ESP

[62] Anonymous. An electric stenographic pen. Scientific American, 62(3):42, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
42b.pdf.

Anonymous:1890:IFDa

[63] Anonymous. An improved fruit drier. Scientific American, 62 (3):42, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
42.pdf.

Anonymous:1890:ISBb

[64] Anonymous. An improved suspender belt. Scientific American, 62(3):42, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
42a.pdf.

Anonymous:1890:SKV

[65] Anonymous. The Sherman “King” vaporizer. Scientific American, 62(3):42, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
42c.pdf.

Anonymous:1890:IFW

[66] Anonymous. An improved fence wire stretcher. Scientific American, 62(3):43, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
43.pdf.

Anonymous:1890:IPWa

[67] Anonymous. An improved pipe wrench. Scientific American, 62 (3):43, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
43a.pdf.

Anonymous:1890:NAL

[68] Anonymous. The new astigmatic lens for optical instruments. Scientific American, 62(3):43, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
43b.pdf.

Anonymous:1890:BPc

[69] Anonymous. Business and personal. Scientific American, 62(3):44, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-44a.pdf.

Anonymous:1890:RPIb

[70] Anonymous. Recently patented inventions. Scientific American, 62(3):44, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n3/pdf/scientificamerican01181890-
44.pdf.

Anonymous:1890:NQc

[71] Anonymous. Notes and queries. Scientific American, 62(3):44–45, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-44b.pdf.

Anonymous:1890:Id

[72] Anonymous. To inventors. Scientific American, 62(3):45, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-45.pdf.

Anonymous:1890:IIc

[73] Anonymous. Index of inventions. Scientific American, 62(3):45–46, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-45a.pdf.

Anonymous:1890:PNPc

[74] Anonymous. Patent notes: Patents. Scientific American, 62(3):48, January 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n3/pdf/scientificamerican01181890-48.pdf.

Anonymous:1890:SEL

[75] Anonymous. Safety in electric lighting. Scientific American, 62 (4):50–51, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
50.pdf.

Anonymous:1890:PNb

[76] Anonymous. Photographic notes. Scientific American, 62(4):51, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-51.pdf.

Anonymous:1890:BSV

[77] Anonymous. A brake and starter for vehicles. Scientific American, 62(4):52, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
52a.pdf.

Anonymous:1890:ICEb

[78] Anonymous. An improved compound engine. Scientific American, 62(4):52, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
52.pdf.

Anonymous:1890:ICT

[79] Anonymous. An improved corn trimmer. Scientific American, 62(4):52, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
52b.pdf.

Anonymous:1890:FBE

[80] Anonymous. The feather blade elliptic propeller. Scientific American, 62(4):53, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
53a.pdf.

Anonymous:1890:IBBa

[81] Anonymous. An improved bicycle brake shoe. Scientific American, 62(4):53, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
53b.pdf.

Anonymous:1890:IPI

[82] Anonymous. An improved pruning implement. Scientific American, 62(4):53, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
53.pdf.

Anonymous:1890:IBSa

[83] Anonymous. An improved band saw pulley. Scientific American, 62(4):54, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
54a.pdf.

Anonymous:1890:IDL

[84] Anonymous. An improved door lock. Scientific American, 62(4):54, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-54c.pdf.

Anonymous:1890:IFPa

[85] Anonymous. An improved fountain pen. Scientific American, 62(4):54, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
54.pdf.

Anonymous:1890:IRJ

[86] Anonymous. An improved rail joint. Scientific American, 62(4):54, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-54b.pdf.

Anonymous:1890:NYP

[87] Anonymous. The New York post office. Scientific American, 62 (4):55–56, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
55.pdf.

Anonymous:1890:NLS

[88] Anonymous. New letter stamping machine. Scientific American, 62(4):56, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
56.pdf.

Anonymous:1890:LIB

[89] Anonymous. Labors of the International Bureau of Weights and Measures. Scientific American, 62(4):57–58, January 25, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n4/pdf/scientificamerican01251890-57.pdf.

Anonymous:1890:NTPa

[90] Anonymous. A novelty in tramway practice. Scientific American, 62(4):58, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
58.pdf.

Anonymous:1890:DRPa

[91] Anonymous. Decisions relating to patent cases. Scientific American, 62(4):58–59, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
58a.pdf.

Anonymous:1890:FWS

[92] Anonymous. The French war ship “Redoutable”. Scientific American, 62(4):59, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
59.pdf.

Anonymous:1890:BPd

[93] Anonymous. Business and personal. Scientific American, 62(4):60, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-60a.pdf.

Anonymous:1890:RPIc

[94] Anonymous. Recently patented inventions. Scientific American, 62(4):60, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n4/pdf/scientificamerican01251890-
60.pdf.

Anonymous:1890:NQd

[95] Anonymous. Notes and queries. Scientific American, 62(4):60–61, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-60b.pdf.

Anonymous:1890:Ie

[96] Anonymous. To inventors. Scientific American, 62(4):61, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-61.pdf.

Anonymous:1890:IId

[97] Anonymous. Index of inventions. Scientific American, 62(4):61–62, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-61a.pdf.

Anonymous:1890:PNPd

[98] Anonymous. Patent notes: Patents. Scientific American, 62(4):64, January 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n4/pdf/scientificamerican01251890-64.pdf.

Drake:1890:SGT

[99] F. J. Drake. Sea-going torpedo boat no. 1, United States Navy. Scientific American, 62(5):65, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
65.pdf.

Anonymous:1890:ASCa

[100] Anonymous. American Society of Civil Engineers. Scientific American, 62(5):66, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
66b.pdf.

Anonymous:1890:HRT

[101] Anonymous. The Hudson River Tunnel. Scientific American, 62(5):66, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
66.pdf.

Anonymous:1890:TED

[102] Anonymous. The total eclipse of December 22, 1889. Scientific American, 62(5):66, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
66a.pdf.

Anonymous:1890:O

[103] Anonymous. Okra. Scientific American, 62(5):66–67, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n5/pdf/scientificamerican02011890-66c.pdf.

Anonymous:1890:PPF

[104] Anonymous. Position of the planets for February. Scientific American, 62(5):67, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
67.pdf.

Anonymous:1890:IASa

[105] Anonymous. An improved ash sifter. Scientific American, 62 (5):68, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
68d.pdf.

Anonymous:1890:ICa

[106] Anonymous. An improved cultivator. Scientific American, 62 (5):68, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
68b.pdf.

Anonymous:1890:ICCa

[107] Anonymous. An improved car coupling. Scientific American, 62 (5):68, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
68a.pdf.

Anonymous:1890:IFa

[108] Anonymous. An improved furnace. Scientific American, 62(5): 68, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
68.pdf.

Anonymous:1890:IFb

[109] Anonymous. An improved furnace. Scientific American, 62(5): 68, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
68c.pdf.

Anonymous:1890:IWM

[110] Anonymous. An improved water motor. Scientific American, 62 (5):69, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
69a.pdf.

Anonymous:1890:TIC

[111] Anonymous. Two improved candlesticks. Scientific American, 62(5):69, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
69.pdf.

Maccord:1890:EH

[112] C. W. Maccord. Ericsson’s home. Scientific American, 62(5): 69–70, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
69b.pdf.

Anonymous:1890:ILP

[113] Anonymous. Improved linoleum printing machine. Scientific American, 62(5):72, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
72.pdf.

Anonymous:1890:IHA

[114] Anonymous. An improved hot air engine. Scientific American, 62(5):73, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
73.pdf.

Anonymous:1890:EMR

[115] Anonymous. Electrical machine room of the Halles–Centrales, Paris. Scientific American, 62(5):73–74, February 1, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n5/pdf/scientificamerican02011890-73a.pdf.

Anonymous:1890:IFBa

[116] Anonymous. An improved feed bag. Scientific American, 62 (5):74, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
74b.pdf.

Anonymous:1890:IHRa

[117] Anonymous. An improved head rest. Scientific American, 62 (5):74, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
74.pdf.

Anonymous:1890:LKE

[118] Anonymous. A luminous keyhole escutcheon. Scientific American, 62(5):74, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
74a.pdf.

Anonymous:1890:ITa

[119] Anonymous. An improved tricycle. Scientific American, 62(5): 75, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
75a.pdf.

Anonymous:1890:IVa

[120] Anonymous. An improved velocipede. Scientific American, 62 (5):75, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
75c.pdf.

Anonymous:1890:MRB

[121] Anonymous. The Monarch Rubber belting. Scientific American, 62(5):75, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
75.pdf.

Anonymous:1890:NBC

[122] Anonymous. New beam caliper. Scientific American, 62(5):75, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n5/pdf/scientificamerican02011890-75b.pdf.

Anonymous:1890:BPe

[123] Anonymous. Business and personal. Scientific American, 62 (5):76, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
76a.pdf.

Anonymous:1890:RPId

[124] Anonymous. Recently patented inventions. Scientific American, 62(5):76, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
76.pdf.

Anonymous:1890:NQe

[125] Anonymous. Notes and queries. Scientific American, 62(5):76– 77, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
76b.pdf.

Anonymous:1890:Ig

[126] Anonymous. To inventors. Scientific American, 62(5):77, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n5/pdf/scientificamerican02011890-77.pdf.

Anonymous:1890:IIe

[127] Anonymous. Index of inventions. Scientific American, 62(5):77, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n5/pdf/scientificamerican02011890-77a.pdf.

Anonymous:1890:PNPe

[128] Anonymous. Patent notes: Patents. Scientific American, 62 (5):80, February 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n5/pdf/scientificamerican02011890-
80.pdf.

Anonymous:1890:EAG

[129] Anonymous. Extermination of American game. Scientific American, 62(6):82, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
82a.pdf.

Anonymous:1890:WA

[130] Anonymous. Wave action. Scientific American, 62(6):82, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n6/pdf/scientificamerican02081890-82.pdf.

Anonymous:1890:BOD

[131] Anonymous. Bridge our downtown side streets. Scientific American, 62(6):82–83, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
82b.pdf.

Anonymous:1890:APP

[132] Anonymous. An attachment for printing presses. Scientific American, 62(6):84, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
84c.pdf.

Anonymous:1890:BHC

[133] Anonymous. A burner for heated compressed air and oils gases etc. Scientific American, 62(6):84, February 8, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n6/pdf/scientificamerican02081890-84.pdf.

Anonymous:1890:ICb

[134] Anonymous. An improved clamp. Scientific American, 62(6):84, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n6/pdf/scientificamerican02081890-84a.pdf.

Anonymous:1890:ICCb

[135] Anonymous. An improved car coupling. Scientific American, 62 (6):84, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
84b.pdf.

Anonymous:1890:ISPa

[136] Anonymous. An improved stove pipe. Scientific American, 62 (6):84, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
84d.pdf.

Anonymous:1890:AIG

[137] Anonymous. The Armstrong 9-inch gun. Scientific American, 62 (6):85–86, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
85.pdf.

Anonymous:1890:IGR

[138] Anonymous. Improved gate for railway crossings. Scientific American, 62(6):86, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
86.pdf.

Anonymous:1890:BNY

[139] Anonymous. Bridging New York streets. Scientific American, 62(6):87, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
87.pdf.

Anonymous:1890:RJ

[140] Anonymous. Rio De Janeiro. Scientific American, 62(6):88, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n6/pdf/scientificamerican02081890-88.pdf.

Anonymous:1890:MLS

[141] Anonymous. Mr Lorillard’s steel launch. Scientific American, 62(6):89, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
89.pdf.

Anonymous:1890:Cb

[142] Anonymous. Correspondence. Scientific American, 62(6):90, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n6/pdf/scientificamerican02081890-90.pdf.

Anonymous:1890:PSE

[143] Anonymous. Paris sewer electric wires and pneumatic tubes. Scientific American, 62(6):91, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
91a.pdf.

Sloane:1890:HLT

[144] T. O’Conor Sloane. High and low tension electric circuits. Scientific American, 62(6):91, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
91.pdf.

Anonymous:1890:BPf

[145] Anonymous. Business and personal. Scientific American, 62 (6):92, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
92a.pdf.

Anonymous:1890:RPIe

[146] Anonymous. Recently patented inventions. Scientific American, 62(6):92, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
92.pdf.

Anonymous:1890:NQf

[147] Anonymous. Notes and queries. Scientific American, 62(6):92– 93, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
92b.pdf.

Anonymous:1890:Ij

[148] Anonymous. To inventors. Scientific American, 62(6):93, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n6/pdf/scientificamerican02081890-93.pdf.

Anonymous:1890:IIf

[149] Anonymous. Index of inventions. Scientific American, 62(6): 93–94, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
93a.pdf.

Anonymous:1890:PNPf

[150] Anonymous. Patent notes: Patents. Scientific American, 62 (6):96, February 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n6/pdf/scientificamerican02081890-
96.pdf.

Anonymous:1890:MIS

[151] Anonymous. The mechanical interlocking switch and signal system at the Grand Central Depot, New York. Scientific American, 62(7):97, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
97.pdf.

Anonymous:1890:HMF

[152] Anonymous. Helping the Merchant Fleet. Scientific American, 62(7):98, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
98.pdf.

Anonymous:1890:EW

[153] Anonymous. European workers. Scientific American, 62(7):98– 99, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
98a.pdf.

Anonymous:1890:CCC

[154] Anonymous. A combined cotton chopper and cultivator. Scientific American, 62(7):100, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
100.pdf.

Anonymous:1890:ICCc

[155] Anonymous. An improved car coupling. Scientific American, 62 (7):100, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
100c.pdf.

Anonymous:1890:IHRb

[156] Anonymous. An improved hand rake. Scientific American, 62 (7):100, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
100a.pdf.

Anonymous:1890:IMI

[157] Anonymous. An improved musical instrument. Scientific American, 62(7):100, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
100d.pdf.

Sloane:1890:VAA

[158] T. O’Conor Sloane. Volt and ampere analogies. Scientific American, 62(7):100, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
100b.pdf.

Anonymous:1890:EAV

[159] Anonymous. Engines of the Augusta Victoria. Scientific American, 62(7):101–102, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
101.pdf.

Anonymous:1890:Cc

[160] Anonymous. Correspondence. Scientific American, 62(7):103, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n7/pdf/scientificamerican02151890-103.pdf.

Anonymous:1890:PNc

[161] Anonymous. Photographic notes. Scientific American, 62(7): 103, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
103a.pdf.

Anonymous:1890:NECa

[162] Anonymous. A new electric conduit for street cars. Scientific American, 62(7):104, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
104.pdf.

Anonymous:1890:DHO

[163] Anonymous. The dog-headed opossum. Scientific American, 62 (7):105–106, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
105.pdf.

Anonymous:1890:IES

[164] Anonymous. An improved electric signal. Scientific American, 62(7):106, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
106.pdf.

Hopkins:1890:SEL

[165] Geo. M. Hopkins. Some effects of large currents. Scientific American, 62(7):107, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
107.pdf.

Anonymous:1890:BNBa

[166] Anonymous. Books: New books and publications. Scientific American, 62(7):108, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
108a.pdf.

Anonymous:1890:BPg

[167] Anonymous. Business and personal. Scientific American, 62(7): 108, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
108b.pdf.

Anonymous:1890:RPIf

[168] Anonymous. Recently patented inventions. Scientific American, 62(7):108, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
108.pdf.

Anonymous:1890:NQg

[169] Anonymous. Notes and queries. Scientific American, 62(7):108– 109, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
108c.pdf.

Anonymous:1890:Ik

[170] Anonymous. To inventors. Scientific American, 62(7):109, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n7/pdf/scientificamerican02151890-109.pdf.

Anonymous:1890:IIg

[171] Anonymous. Index of inventions. Scientific American, 62(7): 109–110, February 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n7/pdf/scientificamerican02151890-
109a.pdf.

Anonymous:1890:WAC

[172] Anonymous. The Westinghouse Alternating Current System of Electrical Distribution. Scientific American, 62(8):113, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n8/pdf/scientificamerican02221890-113.pdf.

Anonymous:1890:PCN

[173] Anonymous. The Panama Canal — a new project for its completion. Scientific American, 62(8):114, February 22, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n8/pdf/scientificamerican02221890-114.pdf.

Anonymous:1890:ARC

[174] Anonymous. Annual report of the Commissioner of Patents. Scientific American, 62(8):114–115, February 22, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n8/pdf/scientificamerican02221890-114a.pdf.

Austen:1890:HPD

[175] Peter T. Austen. How to Purify drinking water. Scientific American, 62(8):115, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
115.pdf.

Anonymous:1890:IMSa

[176] Anonymous. An improved metallic shingle. Scientific American, 62(8):116, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
116d.pdf.

Anonymous:1890:IRTa

[177] Anonymous. An improved railroad tie and chair. Scientific American, 62(8):116, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
116c.pdf.

Anonymous:1890:ISc

[178] Anonymous. An improved sunshade hat. Scientific American, 62(8):116, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
116b.pdf.

Anonymous:1890:ITCa

[179] Anonymous. An improved toy for children. Scientific American, 62(8):116, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
116a.pdf.

Anonymous:1890:ITHa

[180] Anonymous. An improved tool and handle. Scientific American, 62(8):116, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
116.pdf.

Anonymous:1890:RHC

[181] Anonymous. A ridge and hip covering for roofs. Scientific American, 62(8):116, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
116e.pdf.

Anonymous:1890:NAA

[182] Anonymous. A new aerial apparatus. Scientific American, 62 (8):117–118, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
117.pdf.

Anonymous:1890:Cd

[183] Anonymous. Correspondence. Scientific American, 62(8):119, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n8/pdf/scientificamerican02221890-119.pdf.

Anonymous:1890:CFL

[184] Anonymous. A cheap flash lamp. Scientific American, 62(8):123, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n8/pdf/scientificamerican02221890-123a.pdf.

Anonymous:1890:NME

[185] Anonymous. A new method of electrically driving tuning forks. Scientific American, 62(8):123, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
123b.pdf.

Anonymous:1890:SCT

[186] Anonymous. Simple calls for telephone lines. Scientific American, 62(8):123, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
123.pdf.

Anonymous:1890:SABa

[187] Anonymous. Scientific American building edition. Scientific American, 62(8):124, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
124b.pdf.

Anonymous:1890:NBPa

[188] Anonymous. New books and publications. Scientific American, 62(8):124, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
124a.pdf.

Anonymous:1890:RPIg

[189] Anonymous. Recently patented inventions. Scientific American, 62(8):124, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
124.pdf.

Anonymous:1890:NQh

[190] Anonymous. Notes and queries. Scientific American, 62(8):124– 125, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
124c.pdf.

Anonymous:1890:Il

[191] Anonymous. To inventors. Scientific American, 62(8):125, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n8/pdf/scientificamerican02221890-125.pdf.

Anonymous:1890:IIh

[192] Anonymous. Index of inventions. Scientific American, 62(8): 125–126, February 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n8/pdf/scientificamerican02221890-
125a.pdf.

Anonymous:1890:EMT

[193] Anonymous. Electro-motors on trunk lines. Scientific American, 62(9):130, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
130a.pdf.

Anonymous:1890:JP

[194] Anonymous. S. J. Perry. Scientific American, 62(9):130, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-130b.pdf.

Anonymous:1890:WFb

[195] Anonymous. The World’s Fair. Scientific American, 62(9):130, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-130.pdf.

Anonymous:1890:WJM

[196] Anonymous. William Jarvis Mcalpine. Scientific American, 62 (9):130–131, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
130c.pdf.

Anonymous:1890:PNd

[197] Anonymous. Photographic notes. Scientific American, 62(9):131, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-131.pdf.

Anonymous:1890:HPG

[198] Anonymous. Hydraulic pressure gold chlorination process. Scientific American, 62(9):132, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
132a.pdf.

Anonymous:1890:ICGa

[199] Anonymous. An improved cable grip. Scientific American, 62(9):132, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-132b.pdf.

Anonymous:1890:IPGa

[200] Anonymous. An improved Printer’s galley. Scientific American, 62(9):132, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
132.pdf.

Anonymous:1890:IBSb

[201] Anonymous. An improved brake staff support. Scientific American, 62(9):133, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
133.pdf.

Anonymous:1890:IBW

[202] Anonymous. An improved butter worker. Scientific American, 62(9):133, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
133c.pdf.

Anonymous:1890:ICCd

[203] Anonymous. An improved car coupling. Scientific American, 62(9):133, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-133b.pdf.

Anonymous:1890:IWSa

[204] Anonymous. An improved window shade exhibitor. Scientific American, 62(9):133, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
133a.pdf.

Anonymous:1890:ICHb

[205] Anonymous. An improved check hook for harness. Scientific American, 62(9):133–134, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
133d.pdf.

Anonymous:1890:IHF

[206] Anonymous. An improved hedge fence. Scientific American, 62(9):134, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-134b.pdf.

Anonymous:1890:IHRc

[207] Anonymous. An improved hoe or rake. Scientific American, 62(9):134, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-134a.pdf.

Anonymous:1890:IMSb

[208] Anonymous. An improved machine for stacking hay, ensilage, etc. Scientific American, 62(9):134, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
134.pdf.

Anonymous:1890:NBF

[209] Anonymous. The new Boston fire boat “Engine No. 31”. Scientific American, 62(9):135, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
135.pdf.

Anonymous:1890:Ce

[210] Anonymous. Correspondence. Scientific American, 62(9):135–136, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-135a.pdf.

Anonymous:1890:AWR

[211] Anonymous. The artesian wells of Riverside, Cal. Scientific American, 62(9):136, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
136.pdf.

Anonymous:1890:Ab

[212] Anonymous. The armadillo. Scientific American, 62(9):137–138, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-137.pdf.

Anonymous:1890:SABb

[213] Anonymous. Scientific American building edition. Scientific American, 62(9):139, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
139b.pdf.

Anonymous:1890:BPh

[214] Anonymous. Business and personal. Scientific American, 62(9):139, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-139c.pdf.

Anonymous:1890:NBPb

[215] Anonymous. New books and publications. Scientific American, 62(9):139, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
139a.pdf.

Anonymous:1890:RPIh

[216] Anonymous. Recently patented inventions. Scientific American, 62(9):139, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n9/pdf/scientificamerican03011890-
139.pdf.

Anonymous:1890:NQi

[217] Anonymous. Notes and queries. Scientific American, 62(9):139–141, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-139d.pdf.

Anonymous:1890:Im

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-141.pdf.

Anonymous:1890:IIi

[219] Anonymous. Index of inventions. Scientific American, 62(9):141–142, March 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n9/pdf/scientificamerican03011890-141a.pdf.

Anonymous:1890:NTL

[220] Anonymous. New type of locomotive driving wheels. Scientific American, 62(10):145–146, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
145.pdf.

Anonymous:1890:WFC

[221] Anonymous. The World’s Fair of 1892 — Chicago favored by the House of Representatives. Scientific American, 62(10):146–147, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-146.pdf.

Anonymous:1890:L

[222] Anonymous. Leprosy. Scientific American, 62(10):147, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-147.pdf.

Anonymous:1890:PPMa

[223] Anonymous. Position of the planets for March. Scientific American, 62(10):147, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
147a.pdf.

Anonymous:1890:IMHa

[224] Anonymous. An improved magazine hand camera. Scientific American, 62(10):148, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
148.pdf.

Anonymous:1890:SPN

[225] Anonymous. The sisal plant at Nassau, N.P. Scientific American, 62(10):148, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
148a.pdf.

Anonymous:1890:IRCa

[226] Anonymous. An improved rope clamp and thimble. Scientific American, 62(10):149, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
149.pdf.

Anonymous:1890:IRHa

[227] Anonymous. An improved revolving house hoe. Scientific American, 62(10):149, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
149b.pdf.

Anonymous:1890:ITCb

[228] Anonymous. An improved tool cutter. Scientific American, 62(10):149, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-149a.pdf.

Anonymous:1890:RSB

[229] Anonymous. A removable siding for bedsteads. Scientific American, 62(10):149, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
149c.pdf.

Anonymous:1890:ICd

[230] Anonymous. An improved cultivator. Scientific American, 62 (10):149–150, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
149d.pdf.

Anonymous:1890:IGA

[231] Anonymous. An improved game apparatus. Scientific American, 62(10):150, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
150.pdf.

Anonymous:1890:ILa

[232] Anonymous. An improved lamp. Scientific American, 62(10):150, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-150b.pdf.

Anonymous:1890:ITMa

[233] Anonymous. An improved tape measure. Scientific American, 62(10):150, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
150a.pdf.

Anonymous:1890:Cf

[234] Anonymous. Correspondence. Scientific American, 62(10):151, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-151.pdf.

Anonymous:1890:RPA

[235] Anonymous. Railway practice in America. Scientific American, 62(10):152–154, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
152.pdf.

Hopkins:1890:SUPa

[236] Geo. M. Hopkins. The scientific use of the phonograph. Scientific American, 62(10):155, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
155.pdf.

Anonymous:1890:BNBb

[237] Anonymous. Books: New books and publications. Scientific American, 62(10):156, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
156a.pdf.

Anonymous:1890:BPi

[238] Anonymous. Business and personal. Scientific American, 62(10):156, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-156b.pdf.

Anonymous:1890:NQj

[239] Anonymous. Notes and queries. Scientific American, 62(10):156, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-156c.pdf.

Anonymous:1890:RPIi

[240] Anonymous. Recently patented inventions. Scientific American, 62(10):156, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n10/pdf/scientificamerican03081890-
156.pdf.

Anonymous:1890:In

[241] Anonymous. To inventors. Scientific American, 62(10):157, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-157.pdf.

Anonymous:1890:IIj

[242] Anonymous. Index of inventions. Scientific American, 62(10):157–158, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-157a.pdf.

Anonymous:1890:PNPg

[243] Anonymous. Patent notes: Patents. Scientific American, 62(10):160, March 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n10/pdf/scientificamerican03081890-160.pdf.

Anonymous:1890:CW

[244] Anonymous. The celestial world. Scientific American, 62(11):162, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-162a.pdf.

Anonymous:1890:MBS

[245] Anonymous. The monitor as a battle ship. Scientific American, 62(11):162, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
162.pdf.

Hovey:1890:MUA

[246] H. C. Hovey. The manufacture and uses of aluminum. Scientific American, 62(11):163–164, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
163.pdf.

Anonymous:1890:MIM

[247] Anonymous. Manual instruction of the mechanic. Scientific American, 62(11):164, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
164.pdf.

Anonymous:1890:IBT

[248] Anonymous. An improved barrel truck and jack. Scientific American, 62(11):165, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
165b.pdf.

Anonymous:1890:IMCa

[249] Anonymous. An improved milk can. Scientific American, 62(11):165, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-165.pdf.

Anonymous:1890:MMB

[250] Anonymous. The magic money box. Scientific American, 62(11):165, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-165a.pdf.

Anonymous:1890:ICCe

[251] Anonymous. An improved cash carrier. Scientific American, 62 (11):165–166, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
165c.pdf.

Anonymous:1890:Cg

[252] Anonymous. Correspondence. Scientific American, 62(11):166, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-166b.pdf.

Anonymous:1890:IGP

[253] Anonymous. An improved glass polishing wheel. Scientific American, 62(11):166, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
166.pdf.

Anonymous:1890:NMH

[254] Anonymous. A new machinists’ hammer. Scientific American, 62(11):166, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
166a.pdf.

Anonymous:1890:SDF

[255] Anonymous. The San Diego flume system. Scientific American, 62(11):167, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
167.pdf.

Anonymous:1890:NDD

[256] Anonymous. The new Dry Dock at Halifax. Scientific American, 62(11):168–169, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
168.pdf.

Anonymous:1890:ICM

[257] Anonymous. An improved cane mill. Scientific American, 62(11):169, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-169b.pdf.

Anonymous:1890:MCB

[258] Anonymous. The mummy cats of Beni Hassan. Scientific American, 62(11):169, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
169a.pdf.

Anonymous:1890:RBT

[259] Anonymous. A running-board for the top of cars. Scientific American, 62(11):169, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
169.pdf.

Anonymous:1890:ICBa

[260] Anonymous. An improved car brake. Scientific American, 62 (11):169–170, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
169c.pdf.

Anonymous:1890:SABc

[261] Anonymous. Scientific American building edition. Scientific American, 62(11):171, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
171b.pdf.

Anonymous:1890:BNBc

[262] Anonymous. Books: New books and publications. Scientific American, 62(11):171, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
171a.pdf.

Anonymous:1890:BPj

[263] Anonymous. Business and personal. Scientific American, 62(11):171, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-171c.pdf.

Anonymous:1890:RPIj

[264] Anonymous. Recently patented inventions. Scientific American, 62(11):171, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n11/pdf/scientificamerican03151890-
171.pdf.

Anonymous:1890:NQk

[265] Anonymous. Notes and queries. Scientific American, 62(11):171–173, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-171d.pdf.

Anonymous:1890:Io

[266] Anonymous. To inventors. Scientific American, 62(11):173, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-173.pdf.

Anonymous:1890:IIk

[267] Anonymous. Index of inventions. Scientific American, 62(11):173–174, March 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n11/pdf/scientificamerican03151890-173a.pdf.

Anonymous:1890:NCI

[268] Anonymous. New conference of the International Industrial Union. Scientific American, 62(12):178, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
178.pdf.

Anonymous:1890:PDG

[269] Anonymous. Progress of dynamite guns. Scientific American, 62 (12):178–179, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
178a.pdf.

Anonymous:1890:SGa

[270] Anonymous. Smokeless gunpowder. Scientific American, 62(12): 180–181, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
180.pdf.

Anonymous:1890:DRPb

[271] Anonymous. Decisions relating to patents. Scientific American, 62(12):181, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
181a.pdf.

Anonymous:1890:ICF

[272] Anonymous. An improved curtain fixture. Scientific American, 62(12):181, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
181.pdf.

Anonymous:1890:ILDa

[273] Anonymous. An improved lister and drill. Scientific American, 62(12):181, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
181b.pdf.

Anonymous:1890:ISDa

[274] Anonymous. An improved sensitive drill. Scientific American, 62 (12):181–182, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
181c.pdf.

Anonymous:1890:ICLa

[275] Anonymous. An improved coupling link. Scientific American, 62(12):182, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
182.pdf.

Anonymous:1890:Ch

[276] Anonymous. Correspondence. Scientific American, 62(12):183, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n12/pdf/scientificamerican03221890-183a.pdf.

Anonymous:1890:HRI

[277] Anonymous. The Harlem River Improvement and Ship Canal. Scientific American, 62(12):183, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
183.pdf.

Anonymous:1890:BES

[278] Anonymous. Bending and embossing sheet metal. Scientific American, 62(12):184, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
184.pdf.

Anonymous:1890:SPT

[279] Anonymous. Sending pictures by telegraph. Scientific American, 62(12):184, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
184a.pdf.

Anonymous:1890:IMB

[280] Anonymous. An improved mill boot or pestle. Scientific American, 62(12):184–186, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
184b.pdf.

Anonymous:1890:SABd

[281] Anonymous. Scientific American building edition. Scientific American, 62(12):187, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
187b.pdf.

Anonymous:1890:BNBd

[282] Anonymous. Books: New books and publications. Scientific American, 62(12):187, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
187a.pdf.

Anonymous:1890:BPk

[283] Anonymous. Business and personal. Scientific American, 62(12):187, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n12/pdf/scientificamerican03221890-187c.pdf.

Anonymous:1890:RPIk

[284] Anonymous. Recently patented inventions. Scientific American, 62(12):187, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n12/pdf/scientificamerican03221890-
187.pdf.

Anonymous:1890:NQl

[285] Anonymous. Notes and queries. Scientific American, 62(12):187–188, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n12/pdf/scientificamerican03221890-187d.pdf.

Anonymous:1890:Ip

[286] Anonymous. To inventors. Scientific American, 62(12):188, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n12/pdf/scientificamerican03221890-188.pdf.

Anonymous:1890:IIl

[287] Anonymous. Index of inventions. Scientific American, 62(12):188–190, March 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n12/pdf/scientificamerican03221890-188a.pdf.

Anonymous:1890:FOG

[288] Anonymous. Formal opening of the Great Forth Bridge. Scientific American, 62(13):194, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
194.pdf.

Anonymous:1890:RMP

[289] Anonymous. Roller mill patent case. Scientific American, 62(13):194, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n13/pdf/scientificamerican03291890-194a.pdf.

Anonymous:1890:PPAa

[290] Anonymous. Position of the planets for April. Scientific American, 62(13):194–195, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
194b.pdf.

Anonymous:1890:MFR

[291] Anonymous. The manufacture of fishing rods and tackle. Scientific American, 62(13):196, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
196.pdf.

Anonymous:1890:HUC

[292] Anonymous. A hanger for use in curing tobacco. Scientific American, 62(13):197, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
197.pdf.

Anonymous:1890:IAG

[293] Anonymous. An improved adjustable grooving head. Scientific American, 62(13):197, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
197a.pdf.

Anonymous:1890:ICCf

[294] Anonymous. An improved car coupling. Scientific American, 62(13):197, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
197b.pdf.

Anonymous:1890:IDKa

[295] Anonymous. An improved door knob attachment. Scientific American, 62(13):197, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
197c.pdf.

Anonymous:1890:IDC

[296] Anonymous. An improved door check. Scientific American, 62 (13):197–198, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
197d.pdf.

Anonymous:1890:IBBb

[297] Anonymous. An improved blind for bridles. Scientific American, 62(13):198, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
198.pdf.

Anonymous:1890:Ci

[298] Anonymous. Correspondence. Scientific American, 62(13):199, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n13/pdf/scientificamerican03291890-199.pdf.

Anonymous:1890:MPC

[299] Anonymous. Manufacture of paper on the continent. Scientific American, 62(13):200–201, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
200.pdf.

Anonymous:1890:IAE

[300] Anonymous. An improved armature for electric machines. Scientific American, 62(13):201, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
201.pdf.

Anonymous:1890:ISMa

[301] Anonymous. An improved sewing machine. Scientific American, 62(13):201, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
201a.pdf.

Anonymous:1890:LUS

[302] Anonymous. Launch of the United States steel cruiser Newark. Scientific American, 62(13):202, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
202.pdf.

Anonymous:1890:SABe

[303] Anonymous. Scientific American building edition. Scientific American, 62(13):203, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
203a.pdf.

Anonymous:1890:BNBe

[304] Anonymous. Books: New books and publications. Scientific American, 62(13):203, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
203c.pdf.

Anonymous:1890:BPl

[305] Anonymous. Business and personal. Scientific American, 62(13):203, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n13/pdf/scientificamerican03291890-203b.pdf.

Anonymous:1890:RPIl

[306] Anonymous. Recently patented inventions. Scientific American, 62(13):203, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n13/pdf/scientificamerican03291890-
203.pdf.

Anonymous:1890:NQm

[307] Anonymous. Notes and queries. Scientific American, 62(13):204–205, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n13/pdf/scientificamerican03291890-204.pdf.

Anonymous:1890:Iq

[308] Anonymous. To inventors. Scientific American, 62(13):205, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n13/pdf/scientificamerican03291890-205.pdf.

Anonymous:1890:IIm

[309] Anonymous. Index of inventions. Scientific American, 62(13):205–206, March 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n13/pdf/scientificamerican03291890-205a.pdf.

Anonymous:1890:EPB

[310] Anonymous. Electro-pneumatic block signal system. Scientific American, 62(14):209, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
209.pdf.

Anonymous:1890:SAa

[311] Anonymous. Soda ash. Scientific American, 62(14):210–211, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-210.pdf.

Anonymous:1890:IALa

[312] Anonymous. An improved alarm lock. Scientific American, 62(14):212, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-212.pdf.

Anonymous:1890:IMCb

[313] Anonymous. An improved means of channel protection. Scientific American, 62(14):212, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
212c.pdf.

Anonymous:1890:LWAb

[314] Anonymous. Low water alarm and flue cleaning device for steam boilers. Scientific American, 62(14):212, April 5, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n14/pdf/scientificamerican04051890-212b.pdf.

Anonymous:1890:VSB

[315] Anonymous. The Victor safety bicycle. Scientific American, 62(14):212, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-212a.pdf.

Anonymous:1890:IAa

[316] Anonymous. An improved atomizer. Scientific American, 62(14):213, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-213.pdf.

Anonymous:1890:IEGa

[317] Anonymous. An improved end gate for wagons. Scientific American, 62(14):213, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
213b.pdf.

Anonymous:1890:IKF

[318] Anonymous. An improved key fastener. Scientific American, 62 (14):213, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
213a.pdf.

Anonymous:1890:PLH

[319] Anonymous. A portable lemonade holder and glass washer. Scientific American, 62(14):213, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
213c.pdf.

Anonymous:1890:NWV

[320] Anonymous. A nut and washer for vehicle axles. Scientific American, 62(14):213–214, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
213d.pdf.

Anonymous:1890:Cj

[321] Anonymous. Correspondence. Scientific American, 62(14):215, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-215.pdf.

Anonymous:1890:MM

[322] Anonymous. Mountains of the Moon. Scientific American, 62 (14):217–218, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
217.pdf.

Anonymous:1890:IBFa

[323] Anonymous. An improved bridge flooring. Scientific American, 62(14):218, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
218.pdf.

Anonymous:1890:ICi

[324] Anonymous. An improved camera. Scientific American, 62(14):218, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-218a.pdf.

Anonymous:1890:ARS

[325] Anonymous. Automatic recording signal for railroads. Scientific American, 62(14):219, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
219.pdf.

Anonymous:1890:QNT

[326] Anonymous. Queen’s new triple plate Toeplee–Holtz machine. Scientific American, 62(14):219, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
219a.pdf.

Anonymous:1890:SABf

[327] Anonymous. Scientific American building edition. Scientific American, 62(14):220, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
220b.pdf.

Anonymous:1890:BNBf

[328] Anonymous. Books: New books and publications. Scientific American, 62(14):220, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
220a.pdf.

Anonymous:1890:BPm

[329] Anonymous. Business and personal. Scientific American, 62(14):220, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-220c.pdf.

Anonymous:1890:RPIm

[330] Anonymous. Recently patented inventions. Scientific American, 62(14):220, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n14/pdf/scientificamerican04051890-
220.pdf.

Anonymous:1890:NQn

[331] Anonymous. Notes and queries. Scientific American, 62(14):220–221, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-220d.pdf.

Anonymous:1890:Ir

[332] Anonymous. To inventors. Scientific American, 62(14):221, April 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-221.pdf.

Anonymous:1890:IIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n14/pdf/scientificamerican04051890-221a.pdf.

Anonymous:1890:NDS

[334] Anonymous. A new decision by the Supreme Court relating to foreign and American patents. Scientific American, 62(15):226, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n15/pdf/scientificamerican04121890-226a.pdf.

Anonymous:1890:SGb

[335] Anonymous. Smokeless gunpowder. Scientific American, 62(15):226, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n15/pdf/scientificamerican04121890-226.pdf.

Anonymous:1890:SHS

[336] Anonymous. Stocking the Hudson with salmon. Scientific American, 62(15):226–227, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
226b.pdf.

Anonymous:1890:POS

[337] Anonymous. Patent office salaries. Scientific American, 62(15):227, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n15/pdf/scientificamerican04121890-227.pdf.

Anonymous:1890:ICCg

[338] Anonymous. An improved car coupling. Scientific American, 62 (15):228, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
228c.pdf.

Anonymous:1890:IDKb

[339] Anonymous. An improved dough kneader. Scientific American, 62(15):228, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
228a.pdf.

Anonymous:1890:IDSb

[340] Anonymous. An improved drainage and sewerage system. Scientific American, 62(15):228, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
228b.pdf.

Anonymous:1890:IPGb

[341] Anonymous. An improved Printer’s galley. Scientific American, 62(15):228, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
228.pdf.

Anonymous:1890:ICCh

[342] Anonymous. An improved cotton cultivator. Scientific American, 62(15):229, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
229.pdf.

Anonymous:1890:LDS

[343] Anonymous. The Lovell diamond safety bicycle. Scientific American, 62(15):229, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
229a.pdf.

Anonymous:1890:MDP

[344] Anonymous. A machine for distributing poison on plants. Scientific American, 62(15):229, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
229b.pdf.

Anonymous:1890:CAP

[345] Anonymous. Compressed air passenger cars. Scientific American, 62(15):229–230, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
229c.pdf.

Anonymous:1890:DAC

[346] Anonymous. A device for attaching cost or price tickets. Scientific American, 62(15):231, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
231c.pdf.

Anonymous:1890:ICPa

[347] Anonymous. An improved check punch. Scientific American, 62 (15):231, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
231a.pdf.

Anonymous:1890:IPS

[348] Anonymous. An improved pencil sharpener and holder. Scientific American, 62(15):231, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
231.pdf.

Anonymous:1890:ISPb

[349] Anonymous. An improved stove-pipe shelf. Scientific American, 62(15):231, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
231b.pdf.

Anonymous:1890:CM

[350] Anonymous. The Cyclone of March 26,27,and 28. Scientific American, 62(15):232–235, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
232.pdf.

Anonymous:1890:SABg

[351] Anonymous. Scientific American building edition. Scientific American, 62(15):236, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
236b.pdf.

Anonymous:1890:BPn

[352] Anonymous. Business and personal. Scientific American, 62(15):236, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n15/pdf/scientificamerican04121890-236c.pdf.

Anonymous:1890:NBPc

[353] Anonymous. New books and publications. Scientific American, 62(15):236, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
236a.pdf.

Anonymous:1890:RPIn

[354] Anonymous. Recently patented inventions. Scientific American, 62(15):236, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n15/pdf/scientificamerican04121890-
236.pdf.

Anonymous:1890:NQo

[355] Anonymous. Notes and queries. Scientific American, 62(15):236–237, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n15/pdf/scientificamerican04121890-236d.pdf.

Anonymous:1890:Iu

[356] Anonymous. To inventors. Scientific American, 62(15):237, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n15/pdf/scientificamerican04121890-237.pdf.

Anonymous:1890:IIo

[357] Anonymous. Index of inventions. Scientific American, 62(15):237–238, April 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n15/pdf/scientificamerican04121890-237a.pdf.

Anonymous:1890:AGS

[358] Anonymous. Accident to a great steamer. Scientific American, 62(16):242, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
242a.pdf.

Anonymous:1890:Ck

[359] Anonymous. Camphor. Scientific American, 62(16):242, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-242b.pdf.

Anonymous:1890:M

[360] Anonymous. The majestic. Scientific American, 62(16):242, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-242.pdf.

Anonymous:1890:IALb

[361] Anonymous. Improvement of the ARC light. Scientific American, 62(16):242–243, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
242c.pdf.

Anonymous:1890:IPLb

[362] Anonymous. An improved permutation lock. Scientific American, 62(16):244, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
244c.pdf.

Anonymous:1890:ISMb

[363] Anonymous. An improved Spring motor. Scientific American, 62(16):244, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
244.pdf.

Anonymous:1890:ITW

[364] Anonymous. An improved type-writer. Scientific American, 62(16):244, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-244a.pdf.

Anonymous:1890:IWa

[365] Anonymous. An improved weeder. Scientific American, 62(16):244, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-244b.pdf.

Anonymous:1890:CDE

[366] Anonymous. Carter’s ditching and excavating machine. Scientific American, 62(16):245, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
245c.pdf.

Anonymous:1890:IPWb

[367] Anonymous. An improved perforating wheel. Scientific American, 62(16):245, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
245.pdf.

Anonymous:1890:ISe

[368] Anonymous. An improved skylight. Scientific American, 62(16):245, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-245d.pdf.

Anonymous:1890:ITCc

[369] Anonymous. An improved thill coupling. Scientific American, 62(16):245, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
245a.pdf.

Anonymous:1890:NPP

[370] Anonymous. The Nonpareil pipe plug. Scientific American, 62(16):245, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-245b.pdf.

Anonymous:1890:PBb

[371] Anonymous. Preserve for binding. Scientific American, 62(16):245, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-245e.pdf.

Anonymous:1890:CGC

[372] Anonymous. Climbing the Great Chimney of the Clark Thread Works, at Harrison, N.J. Scientific American, 62(16):246, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-246.pdf.

Anonymous:1890:Cl

[373] Anonymous. Correspondence. Scientific American, 62(16):247, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-247.pdf.

Anonymous:1890:EBP

[374] Anonymous. Economy of belting and pulleys. Scientific American, 62(16):247, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
247a.pdf.

Hopkins:1890:SUPb

[375] Geo. M. Hopkins. The scientific use of the phonograph. Scientific American, 62(16):248, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
248.pdf.

Anonymous:1890:IMMa

[376] Anonymous. An improved mitering machine. Scientific American, 62(16):249, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
249.pdf.

Anonymous:1890:F

[377] Anonymous. Fluorine. Scientific American, 62(16):249–250, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-249a.pdf.

Anonymous:1890:IMDa

[378] Anonymous. Improved means of distributing power. Scientific American, 62(16):250, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
250.pdf.

Anonymous:1890:NBPd

[379] Anonymous. New books and publications. Scientific American, 62(16):251, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
251a.pdf.

Anonymous:1890:RPIo

[380] Anonymous. Recently patented inventions. Scientific American, 62(16):251, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n16/pdf/scientificamerican04191890-
251.pdf.

Anonymous:1890:BPo

[381] Anonymous. Business and personal. Scientific American, 62(16):252, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-252.pdf.

Anonymous:1890:NQp

[382] Anonymous. Notes and queries. Scientific American, 62(16):252–253, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-252a.pdf.

Anonymous:1890:Ix

[383] Anonymous. To inventors. Scientific American, 62(16):253, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-253.pdf.

Anonymous:1890:IIp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-253a.pdf.

Anonymous:1890:PNPh

[385] Anonymous. Patent notes: Patents. Scientific American, 62(16):256, April 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n16/pdf/scientificamerican04191890-256.pdf.

Anonymous:1890:EWE

[386] Anonymous. Exhibition of the wonders of electricity. Scientific American, 62(17):258, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
258.pdf.

Anonymous:1890:PBA

[387] Anonymous. A proposed bill against the Chinese in the United States. Scientific American, 62(17):258–259, April 26, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n17/pdf/scientificamerican04261890-258a.pdf.

Anonymous:1890:PBc

[388] Anonymous. Preserve for binding. Scientific American, 62(17):259, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-259a.pdf.

Anonymous:1890:PPMb

[389] Anonymous. Position of the planets for may. Scientific American, 62(17):259, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
259.pdf.

Anonymous:1890:FCE

[390] Anonymous. Frick’s commercial evaporator “Simoon”. Scientific American, 62(17):260, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
260.pdf.

Anonymous:1890:GSJ

[391] Anonymous. Garrett’s saddle jack. Scientific American, 62(17):260, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-260b.pdf.

Anonymous:1890:IQa

[392] Anonymous. An improved quilter. Scientific American, 62(17):260, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-260a.pdf.

Anonymous:1890:IBSc

[393] Anonymous. An improved brake shoe. Scientific American, 62(17):261, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-261.pdf.

Anonymous:1890:ICCi

[394] Anonymous. An improved circuit closer. Scientific American, 62(17):261, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
261e.pdf.

Anonymous:1890:ICDa

[395] Anonymous. An improved churn dasher. Scientific American, 62(17):261, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
261a.pdf.

Anonymous:1890:IMF

[396] Anonymous. An improved measuring faucet. Scientific American, 62(17):261, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
261d.pdf.

Anonymous:1890:IPWc

[397] Anonymous. An improved pipe wrench. Scientific American, 62 (17):261, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
261b.pdf.

Anonymous:1890:ISG

[398] Anonymous. An improved shoe or glove fastener. Scientific American, 62(17):261, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
261c.pdf.

Anonymous:1890:PNe

[399] Anonymous. Photographic notes. Scientific American, 62(17):262, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-262.pdf.

Anonymous:1890:Cn

[400] Anonymous. Correspondence. Scientific American, 62(17):263, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-263a.pdf.

Anonymous:1890:EPD

[401] Anonymous. Edison’s phonographic doll. Scientific American, 62 (17):263, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
263.pdf.

Anonymous:1890:LBN

[402] Anonymous. Lieut. Bettini’s new micro-graphophone. Scientific American, 62(17):264, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
264.pdf.

Anonymous:1890:SPRb

[403] Anonymous. The stepped platform railway. Scientific American, 62(17):265–266, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
265.pdf.

Anonymous:1890:NGG

[404] Anonymous. The new gun of the German Army. Scientific American, 62(17):266–267, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
266.pdf.

Anonymous:1890:IDKc

[405] Anonymous. An improved drawing kit. Scientific American, 62(17):267, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-267.pdf.

Anonymous:1890:IML

[406] Anonymous. An improved music leaf turner. Scientific American, 62(17):267, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
267a.pdf.

Anonymous:1890:BNBg

[407] Anonymous. Books: New books and publications. Scientific American, 62(17):268, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
268a.pdf.

Anonymous:1890:BPp

[408] Anonymous. Business and personal. Scientific American, 62(17):268, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-268b.pdf.

Anonymous:1890:RPIp

[409] Anonymous. Recently patented inventions. Scientific American, 62(17):268, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n17/pdf/scientificamerican04261890-
268.pdf.

Anonymous:1890:NQq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-268c.pdf.

Anonymous:1890:Iy

[411] Anonymous. To inventors. Scientific American, 62(17):269, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-269.pdf.

Anonymous:1890:IIq

[412] Anonymous. Index of inventions. Scientific American, 62(17):269–271, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-269a.pdf.

Anonymous:1890:PNPi

[413] Anonymous. Patent notes: Patents. Scientific American, 62(17):272, April 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n17/pdf/scientificamerican04261890-272.pdf.

Anonymous:1890:WIS

[414] Anonymous. The Westinghouse Interlocking Switch and Signal System. Scientific American, 62(18):273, May 3, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n18/pdf/scientificamerican05031890-273.pdf.

Anonymous:1890:QPB

[415] Anonymous. A questionable Patent Bill. Scientific American, 62(18):274, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
274.pdf.

Hovey:1890:OWS

[416] E. O. Hovey. The oil well at Southbury, Conn. Scientific American, 62(18):275, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
275.pdf.

Anonymous:1890:IAD

[417] Anonymous. An improved automatic dumb waiter. Scientific American, 62(18):276, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
276.pdf.

Anonymous:1890:IBP

[418] Anonymous. An improved brace for piano string frames. Scientific American, 62(18):276, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
276b.pdf.

Anonymous:1890:IWT

[419] Anonymous. An improved water tower and extension ladder. Scientific American, 62(18):276, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
276a.pdf.

Anonymous:1890:ISLa

[420] Anonymous. An improved shaft lighter. Scientific American, 62 (18):277, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
277c.pdf.

Anonymous:1890:ITR

[421] Anonymous. An improved time register. Scientific American, 62 (18):277, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
277.pdf.

Anonymous:1890:IWb

[422] Anonymous. An improved windlass. Scientific American, 62(18):277, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-277b.pdf.

Anonymous:1890:NGAa

[423] Anonymous. New German armor-clad cruisers. Scientific American, 62(18):277, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
277a.pdf.

Anonymous:1890:IAB

[424] Anonymous. An improved animal bed. Scientific American, 62 (18):277–278, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
277d.pdf.

Anonymous:1890:PBd

[425] Anonymous. Preserve for binding. Scientific American, 62(18):278, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-278.pdf.

Anonymous:1890:IHSa

[426] Anonymous. An improved hay sling. Scientific American, 62(18):279, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-279a.pdf.

Anonymous:1890:WAK

[427] Anonymous. The Wilzin automatic knife. Scientific American, 62(18):279, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
279.pdf.

Anonymous:1890:PSBa

[428] Anonymous. The proposed suspension bridge across the Hudson River at New York. Scientific American, 62(18):280–283, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-280.pdf.

Anonymous:1890:ATD

[429] Anonymous. Automatic time and dating stamps. Scientific American, 62(18):283, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
283b.pdf.

Anonymous:1890:IGM

[430] Anonymous. An improved grinding machine. Scientific American, 62(18):283, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
283.pdf.

Anonymous:1890:ISEa

[431] Anonymous. An improved steam engine valve. Scientific American, 62(18):283, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
283a.pdf.

Anonymous:1890:SABh

[432] Anonymous. Scientific American building edition. Scientific American, 62(18):284, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
284b.pdf.

Anonymous:1890:BNBh

[433] Anonymous. Books: New books and publications. Scientific American, 62(18):284, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
284a.pdf.

Anonymous:1890:BPq

[434] Anonymous. Business and personal. Scientific American, 62(18):284, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-284c.pdf.

Anonymous:1890:Iz

[435] Anonymous. To inventors. Scientific American, 62(18):284, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-284e.pdf.

Anonymous:1890:NQr

[436] Anonymous. Notes and queries. Scientific American, 62(18):284, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-284d.pdf.

Anonymous:1890:RPIq

[437] Anonymous. Recently patented inventions. Scientific American, 62(18):284, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
284.pdf.

Anonymous:1890:IIr

[438] Anonymous. Index of inventions. Scientific American, 62(18):284–286, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-284f.pdf.

Anonymous:1890:SAEa

[439] Anonymous. The Scientific American established 1846. Scientific American, 62(18):288, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n18/pdf/scientificamerican05031890-
288a.pdf.

Anonymous:1890:PNPj

[440] Anonymous. Patent notes: Patents. Scientific American, 62(18):288, May 3, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n18/pdf/scientificamerican05031890-288.pdf.

Anonymous:1890:RAR

[441] Anonymous. The Rowell automatic railway safety stop. Scientific American, 62(19):289, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
289.pdf.

Anonymous:1890:TB

[442] Anonymous. Torpedo boats. Scientific American, 62(19):290, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-290.pdf.

Anonymous:1890:SGN

[443] Anonymous. Ships and guns needed for defense. Scientific American, 62(19):290–291, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
290a.pdf.

Anonymous:1890:DAL

[444] Anonymous. A device for automatically lighting fires. Scientific American, 62(19):292, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
292a.pdf.

Anonymous:1890:ICOa

[445] Anonymous. An improved can opener. Scientific American, 62(19):292, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-292c.pdf.

Anonymous:1890:IHc

[446] Anonymous. An improved hoist. Scientific American, 62(19):292, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-292.pdf.

Anonymous:1890:IHd

[447] Anonymous. An improved hat holder. Scientific American, 62(19):292, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-292b.pdf.

Anonymous:1890:ILM

[448] Anonymous. An improved lawn mower. Scientific American, 62(19):292, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-292d.pdf.

Anonymous:1890:ICCj

[449] Anonymous. An improved channel cleaner. Scientific American, 62(19):293, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
293b.pdf.

Anonymous:1890:IDG

[450] Anonymous. An improved dust guard and ventilator. Scientific American, 62(19):293, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
293c.pdf.

Anonymous:1890:IIAa

[451] Anonymous. An improved irrigating apparatus. Scientific American, 62(19):293, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
293a.pdf.

Anonymous:1890:ISS

[452] Anonymous. An improved switch signal. Scientific American, 62 (19):293, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
293.pdf.

Anonymous:1890:Co

[453] Anonymous. Correspondence. Scientific American, 62(19):295, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-295.pdf.

Anonymous:1890:NER

[454] Anonymous. New elevated railway. Scientific American, 62(19):296, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-296.pdf.

Anonymous:1890:ELA

[455] Anonymous. Electric lighting for amateurs. Scientific American, 62(19):297–298, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
297.pdf.

Anonymous:1890:ICEc

[456] Anonymous. An improved compound engine. Scientific American, 62(19):299, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
299.pdf.

Anonymous:1890:IDa

[457] Anonymous. An improved dynamometer. Scientific American, 62 (19):299, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
299a.pdf.

Anonymous:1890:BNBi

[458] Anonymous. Books: New books and publications. Scientific American, 62(19):300, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
300a.pdf.

Anonymous:1890:BPr

[459] Anonymous. Business and personal. Scientific American, 62(19):300, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-300b.pdf.

Anonymous:1890:RPIr

[460] Anonymous. Recently patented inventions. Scientific American, 62(19):300, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n19/pdf/scientificamerican05101890-
300.pdf.

Anonymous:1890:NQs

[461] Anonymous. Notes and queries. Scientific American, 62(19):300–301, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-300c.pdf.

Anonymous:1890:Iba

[462] Anonymous. To inventors. Scientific American, 62(19):301, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-301.pdf.

Anonymous:1890:IIs

[463] Anonymous. Index of inventions. Scientific American, 62(19):301–302, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-301a.pdf.

Anonymous:1890:PNPk

[464] Anonymous. Patent notes: Patents. Scientific American, 62(19):304, May 10, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n19/pdf/scientificamerican05101890-304.pdf.

Anonymous:1890:NSBa

[465] Anonymous. A new submarine boat. Scientific American, 62(20):305, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-305.pdf.

Anonymous:1890:CBF

[466] Anonymous. The Copyright Bill fails to pass. Scientific American, 62(20):306, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
306.pdf.

Anonymous:1890:FCS

[467] Anonymous. A fiber from cotton stalks. Scientific American, 62 (20):306–307, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
306a.pdf.

Anonymous:1890:AA

[468] Anonymous. An anvil attachment. Scientific American, 62(20):308, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-308a.pdf.

Anonymous:1890:IALc

[469] Anonymous. An improved adjustable ladder. Scientific American, 62(20):308, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
308b.pdf.

Anonymous:1890:IPCb

[470] Anonymous. An improved pipe clamp. Scientific American, 62(20):308, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-308c.pdf.

Anonymous:1890:IVR

[471] Anonymous. An improved ventilator for railway cars. Scientific American, 62(20):308, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
308.pdf.

Anonymous:1890:ICLb

[472] Anonymous. An improved clothes line adjuster. Scientific American, 62(20):309, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
309c.pdf.

Anonymous:1890:ICPb

[473] Anonymous. An improved curtain pole fixture. Scientific American, 62(20):309, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
309.pdf.

Anonymous:1890:ILI

[474] Anonymous. An improved ledger index. Scientific American, 62 (20):309, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
309b.pdf.

Anonymous:1890:IWSb

[475] Anonymous. An improved wire stretcher. Scientific American, 62(20):309, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
309a.pdf.

Anonymous:1890:WSW

[476] Anonymous. A water shield for window brushes. Scientific American, 62(20):309, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
309d.pdf.

Anonymous:1890:Cp

[477] Anonymous. Correspondence. Scientific American, 62(20):311, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-311.pdf.

Anonymous:1890:ASCb

[478] Anonymous. The accident to the steamer City of Paris. Scientific American, 62(20):312–313, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
312.pdf.

Anonymous:1890:E

[479] Anonymous. Etching. Scientific American, 62(20):313, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-313a.pdf.

Anonymous:1890:IFDb

[480] Anonymous. An improved furnace door opener. Scientific American, 62(20):313, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
313b.pdf.

Anonymous:1890:IVb

[481] Anonymous. An improved valve. Scientific American, 62(20):313, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-313.pdf.

Anonymous:1890:GPS

[482] Anonymous. A gigantic plant from Sumatra. Scientific American, 62(20):314–315, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
314.pdf.

Anonymous:1890:NGF

[483] Anonymous. A new gun for firing high explosives. Scientific American, 62(20):315, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
315.pdf.

Anonymous:1890:NTPb

[484] Anonymous. A new target pistol. Scientific American, 62(20):315, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-315a.pdf.

Anonymous:1890:SAc

[485] Anonymous. Scientific American. Scientific American, 62(20):316, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-316a.pdf.

Anonymous:1890:BPs

[486] Anonymous. Business and personal. Scientific American, 62(20):316, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-316b.pdf.

Anonymous:1890:RPIs

[487] Anonymous. Recently patented inventions. Scientific American, 62(20):316, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n20/pdf/scientificamerican05171890-
316.pdf.

Anonymous:1890:NQt

[488] Anonymous. Notes and queries. Scientific American, 62(20):316–317, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-316c.pdf.

Anonymous:1890:Ibc

[489] Anonymous. To inventors. Scientific American, 62(20):317, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-317.pdf.

Anonymous:1890:IIt

[490] Anonymous. Index of inventions. Scientific American, 62(20):317–318, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-317a.pdf.

Anonymous:1890:EB

[491] Anonymous. Electrical books. Scientific American, 62(20):319, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-319.pdf.

Anonymous:1890:SAd

[492] Anonymous. The Scientific American. Scientific American, 62(20):320, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-320a.pdf.

Anonymous:1890:PNPl

[493] Anonymous. Patent notes: Patents! Scientific American, 62(20):320, May 17, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n20/pdf/scientificamerican05171890-320.pdf.

Anonymous:1890:SBM

[494] Anonymous. Seabury breech mechanism for rapid firing and other guns. Scientific American, 62(21):321, May 24, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n21/pdf/scientificamerican05241890-321.pdf.

Anonymous:1890:PPE

[495] Anonymous. Profits of the Paris Exposition. Scientific American, 62(21):322, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
322.pdf.

Anonymous:1890:ATF

[496] Anonymous. America’s testimonial to France. Scientific American, 62(21):322–323, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
322a.pdf.

Anonymous:1890:IBHa

[497] Anonymous. An improved blanket harness. Scientific American, 62(21):324, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
324b.pdf.

Anonymous:1890:ICCk

[498] Anonymous. An improved car coupling. Scientific American, 62 (21):324, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
324.pdf.

Anonymous:1890:IKG

[499] Anonymous. An improved knife grinding machine. Scientific American, 62(21):324, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
324c.pdf.

Anonymous:1890:IPa

[500] Anonymous. An improved padlock. Scientific American, 62(21):324, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-324a.pdf.

Anonymous:1890:SET

[501] Anonymous. A slag escape for Tuyères. Scientific American, 62(21):324, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-324d.pdf.

Anonymous:1890:LR

[502] Anonymous. A lesson in repoussé. Scientific American, 62(21):325–326, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-325.pdf.

Anonymous:1890:Cq

[503] Anonymous. Correspondence. Scientific American, 62(21):327, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-327.pdf.

Anonymous:1890:MSTa

[504] Anonymous. A modern stage trick. Scientific American, 62(21):329, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-329.pdf.

Anonymous:1890:VCSa

[505] Anonymous. A venomous Ceylon serpent. Scientific American, 62(21):329, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
330.pdf.

Anonymous:1890:RPIt

[506] Anonymous. Recently patented inventions. Scientific American, 62(21):330, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
330a.pdf.

Anonymous:1890:SABi

[507] Anonymous. Scientific American building edition. Scientific American, 62(21):331, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
331a.pdf.

Anonymous:1890:RPIu

[508] Anonymous. Recently patented inventions. Scientific American, 62(21):331, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
331.pdf.

Anonymous:1890:BPt

[509] Anonymous. Business and personal. Scientific American, 62(21):332, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-332.pdf.

Anonymous:1890:NQu

[510] Anonymous. Notes and queries. Scientific American, 62(21):332–333, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-332a.pdf.

Anonymous:1890:Ibd

[511] Anonymous. To inventors. Scientific American, 62(21):333, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-333.pdf.

Anonymous:1890:IIu

[512] Anonymous. Index of inventions. Scientific American, 62(21):333–334, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-333a.pdf.

Anonymous:1890:SAPa

[513] Anonymous. The Scientific American publications for 1890. Scientific American, 62(21):335, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
335b.pdf.

Anonymous:1890:EBR

[514] Anonymous. Electrical books recently published. Scientific American, 62(21):335, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
335.pdf.

Hopkins:1890:ES

[515] Geo. M. Hopkins. Experimental science. Scientific American, 62 (21):335, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n21/pdf/scientificamerican05241890-
335a.pdf.

Anonymous:1890:SAf

[516] Anonymous. The Scientific American. Scientific American, 62(21):336, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-336a.pdf.

Anonymous:1890:PNPm

[517] Anonymous. Patent notes: Patents! Scientific American, 62(21):336, May 24, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n21/pdf/scientificamerican05241890-336.pdf.

Anonymous:1890:IST

[518] Anonymous. The industrial South — Tallapoosa, Ga. Scientific American, 62(22):338, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
337.pdf.

Anonymous:1890:WPL

[519] Anonymous. A weekly payment law in New York. Scientific American, 62(22):338, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
338.pdf.

Anonymous:1890:RC

[520] Anonymous. Rice culture. Scientific American, 62(22):338–339, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-338a.pdf.

Anonymous:1890:PNf

[521] Anonymous. Photographic notes. Scientific American, 62(22):339, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-339a.pdf.

Anonymous:1890:PPJa

[522] Anonymous. Position of the planets for June. Scientific American, 62(22):339, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
339.pdf.

Anonymous:1890:IBFb

[523] Anonymous. An improved bag filler and holder. Scientific American, 62(22):340, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
340.pdf.

Anonymous:1890:INLa

[524] Anonymous. An improved nut lock. Scientific American, 62(22):340, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-340c.pdf.

Anonymous:1890:IPGc

[525] Anonymous. An index plate for gear-cutting machines. Scientific American, 62(22):340, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
340a.pdf.

Anonymous:1890:IVW

[526] Anonymous. An improved vehicle wheel. Scientific American, 62 (22):340, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
340b.pdf.

Anonymous:1890:IOC

[527] Anonymous. An improved Ore concentrator. Scientific American, 62(22):340–341, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
340d.pdf.

Anonymous:1890:FAS

[528] Anonymous. The fire alarm system of Jersey City. Scientific American, 62(22):341, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
341.pdf.

Anonymous:1890:EMS

[529] Anonymous. An envelope moistener and sealer. Scientific American, 62(22):342, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
342c.pdf.

Anonymous:1890:IDI

[530] Anonymous. An improved dental instrument. Scientific American, 62(22):342, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
342.pdf.

Anonymous:1890:IHP

[531] Anonymous. An improved holder for paper bags. Scientific American, 62(22):342, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
342b.pdf.

Anonymous:1890:ISIa

[532] Anonymous. An improved sad iron. Scientific American, 62(22):342, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-342d.pdf.

Anonymous:1890:ITHb

[533] Anonymous. An improved tool holder. Scientific American, 62(22):342, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-342a.pdf.

Anonymous:1890:Cr

[534] Anonymous. Correspondence. Scientific American, 62(22):343–344, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-343.pdf.

Anonymous:1890:ILRa

[535] Anonymous. An improved logging railway. Scientific American, 62(22):344, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
344.pdf.

Anonymous:1890:ISBc

[536] Anonymous. An improved shaft bearing. Scientific American, 62 (22):344, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
344a.pdf.

Anonymous:1890:SGO

[537] Anonymous. Stained glass and objects of wire cloth. Scientific American, 62(22):347, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
347.pdf.

Anonymous:1890:BPu

[538] Anonymous. Business and personal. Scientific American, 62(22):348, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-348b.pdf.

Anonymous:1890:NBPe

[539] Anonymous. New books and publications. Scientific American, 62(22):348, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
348a.pdf.

Anonymous:1890:RPIv

[540] Anonymous. Recently patented inventions. Scientific American, 62(22):348, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n22/pdf/scientificamerican05311890-
348.pdf.

Anonymous:1890:NQv

[541] Anonymous. Notes and queries. Scientific American, 62(22):348–349, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-348c.pdf.

Anonymous:1890:IIv

[542] Anonymous. Index of inventions. Scientific American, 62(22):349–350, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-349.pdf.

Anonymous:1890:PNPn

[543] Anonymous. Patent notes: Patents! Scientific American, 62(22):352, May 31, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n22/pdf/scientificamerican05311890-352.pdf.

Anonymous:1890:ISLb

[544] Anonymous. Improved ships and life saving devices wanted. Scientific American, 62(23):354, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
354.pdf.

Anonymous:1890:NSN

[545] Anonymous. The new steamer Normannia. Scientific American, 62(23):354–355, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
354a.pdf.

Anonymous:1890:IFFa

[546] Anonymous. An improved fodder fork. Scientific American, 62(23):356, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-356c.pdf.

Anonymous:1890:IMAa

[547] Anonymous. An improved metal awning. Scientific American, 62(23):356, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
356a.pdf.

Anonymous:1890:IRS

[548] Anonymous. An improved reversible seat for chairs. Scientific American, 62(23):356, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
356d.pdf.

Anonymous:1890:NSV

[549] Anonymous. A new straightway valve. Scientific American, 62(23):356, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-356b.pdf.

Anonymous:1890:SBC

[550] Anonymous. Stone’s bread cutter or slicing machine. Scientific American, 62(23):356, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
356.pdf.

Anonymous:1890:IPR

[551] Anonymous. An improved pneumatic railway system. Scientific American, 62(23):357, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
357.pdf.

Anonymous:1890:IPCc

[552] Anonymous. An improved passenger car. Scientific American, 62(23):357–358, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
357a.pdf.

Anonymous:1890:IHCa

[553] Anonymous. An improved hoist and conveyer. Scientific American, 62(23):358, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
358a.pdf.

Anonymous:1890:ITPa

[554] Anonymous. An improved tobacco pipe top. Scientific American, 62(23):358, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
358.pdf.

Anonymous:1890:Cs

[555] Anonymous. Correspondence. Scientific American, 62(23):359, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-359.pdf.

Anonymous:1890:CHM

[556] Anonymous. Coal handling machinery at the Rondout Yard of the Delaware and Hudson Canal Co. Scientific American, 62(23):360–361, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-360.pdf.

Anonymous:1890:IECb

[557] Anonymous. An improved elevator car. Scientific American, 62(23):361, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-361.pdf.

Anonymous:1890:IMCc

[558] Anonymous. An improved mail car. Scientific American, 62 (23):361–362, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
361a.pdf.

Anonymous:1890:HCE

[559] Anonymous. Hinge connection for electric burglar alarms. Scientific American, 62(23):363, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
363a.pdf.

Anonymous:1890:WC

[560] Anonymous. Wood carving. Scientific American, 62(23):363, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-363.pdf.

Anonymous:1890:BNBj

[561] Anonymous. Books: New books and publications. Scientific American, 62(23):364, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
364a.pdf.

Anonymous:1890:BPv

[562] Anonymous. Business and personal. Scientific American, 62(23):364, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-364b.pdf.

Anonymous:1890:NQw

[563] Anonymous. Notes and queries. Scientific American, 62(23):364, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-364c.pdf.

Anonymous:1890:RPIw

[564] Anonymous. Recently patented inventions. Scientific American, 62(23):364, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n23/pdf/scientificamerican06071890-
364.pdf.

Anonymous:1890:Ibe

[565] Anonymous. To inventors. Scientific American, 62(23):365, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-365.pdf.

Anonymous:1890:IIw

[566] Anonymous. Index of inventions. Scientific American, 62(23):365–366, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-365a.pdf.

Anonymous:1890:PNPo

[567] Anonymous. Patent notes: Patents. Scientific American, 62(23):368, June 7, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n23/pdf/scientificamerican06071890-368.pdf.

Anonymous:1890:WFN

[568] Anonymous. The Westinghouse Foundry, near Pittsburg, Pa. Scientific American, 62(24):369–370, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
369.pdf.

Anonymous:1890:PAP

[569] Anonymous. Proposed amendments to the patent laws. Scientific American, 62(24):370–371, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
370.pdf.

Anonymous:1890:CRa

[570] Anonymous. A community of readers. Scientific American, 62(24):371, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-371.pdf.

Anonymous:1890:IDW

[571] Anonymous. An improved dumb-waiter indicator. Scientific American, 62(24):372, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
372.pdf.

Anonymous:1890:IFPb

[572] Anonymous. An improved fence post. Scientific American, 62(24):372, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-372b.pdf.

Anonymous:1890:ITB

[573] Anonymous. An improved tile or brick cutting table. Scientific American, 62(24):372, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
372c.pdf.

Anonymous:1890:TGL

[574] Anonymous. A tail guard and line rest for harness. Scientific American, 62(24):372, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
372a.pdf.

Anonymous:1890:CTA

[575] Anonymous. A cutting tool attachment for planers. Scientific American, 62(24):373, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
373b.pdf.

Anonymous:1890:ICHc

[576] Anonymous. An improved cabinet for holding spool thread. Scientific American, 62(24):373, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
373.pdf.

Anonymous:1890:IWW

[577] Anonymous. An improved water wheel. Scientific American, 62(24):373, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-373c.pdf.

Anonymous:1890:NAD

[578] Anonymous. A novel advertising device. Scientific American, 62 (24):373, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
373a.pdf.

Anonymous:1890:IOG

[579] Anonymous. An improved oil or gas stove. Scientific American, 62(24):373–374, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
373d.pdf.

Anonymous:1890:PNg

[580] Anonymous. Photographic notes. Scientific American, 62(24):374–375, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-374.pdf.

Anonymous:1890:MPY

[581] Anonymous. The Mount Pleasant Yard of the Pennsylvania Railroad, at Jersey City, N.J. Scientific American, 62(24):376–377, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-376.pdf.

Anonymous:1890:ASE

[582] Anonymous. The “American System” of electric arc lighting. Scientific American, 62(24):377, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
377.pdf.

Anonymous:1890:ITd

[583] Anonymous. An improved Tellurian. Scientific American, 62(24):377– 378, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://en.wikipedia.org/wiki/Tellurion;
 http://www.nature.com/scientificamerican/journal/v62/n24/
pdf/scientificamerican06141890-377a.pdf.

Anonymous:1890:IMAb

[584] Anonymous. An improved method of automatically closing elevator doors. Scientific American, 62(24):378–379, June 14, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n24/pdf/scientificamerican06141890-378.pdf.

Anonymous:1890:OIW

[585] Anonymous. Ornamental iron work for amateurs. Scientific American, 62(24):379, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
379.pdf.

Anonymous:1890:BPw

[586] Anonymous. Business and personal. Scientific American, 62(24):380, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-380a.pdf.

Anonymous:1890:RPIx

[587] Anonymous. Recently patented inventions. Scientific American, 62(24):380, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n24/pdf/scientificamerican06141890-
380.pdf.

Anonymous:1890:NQx

[588] Anonymous. Notes and queries. Scientific American, 62(24):380–381, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-380b.pdf.

Anonymous:1890:Ibg

[589] Anonymous. To inventors. Scientific American, 62(24):381, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-381.pdf.

Anonymous:1890:IIx

[590] Anonymous. Index of inventions. Scientific American, 62(24):381–382, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-381a.pdf.

Anonymous:1890:PNPp

[591] Anonymous. Patent notes: Patents. Scientific American, 62(24):384, June 14, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n24/pdf/scientificamerican06141890-384.pdf.

Anonymous:1890:SVN

[592] Anonymous. The 155th Street Viaduct, New York City, N.Y. Scientific American, 62(25):385, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
385.pdf.

Anonymous:1890:WCI

[593] Anonymous. The Winnebago County (Iowa) meteorites. Scientific American, 62(25):386–387, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
386.pdf.

Anonymous:1890:DRPc

[594] Anonymous. Decisions relating to patents. Scientific American, 62(25):387, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
387.pdf.

Anonymous:1890:DNA

[595] Anonymous. A device to navigate the air. Scientific American, 62(25):388, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
388.pdf.

Anonymous:1890:IBCa

[596] Anonymous. An improved baggage check. Scientific American, 62(25):388, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
388b.pdf.

Anonymous:1890:IBHb

[597] Anonymous. An improved book holder. Scientific American, 62(25):388, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-388a.pdf.

Anonymous:1890:ILRb

[598] Anonymous. An improved life raft. Scientific American, 62(25):388, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-388c.pdf.

Anonymous:1890:IRGa

[599] Anonymous. An improved rein guide. Scientific American, 62(25):388, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-388d.pdf.

Anonymous:1890:ICCl

[600] Anonymous. An improved car coupling. Scientific American, 62 (25):389, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
389c.pdf.

Anonymous:1890:ICCm

[601] Anonymous. An improved carpet cleaning machine. Scientific American, 62(25):389, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
389d.pdf.

Anonymous:1890:IHM

[602] Anonymous. An improved hydraulic motor. Scientific American, 62(25):389, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
389.pdf.

Anonymous:1890:ILH

[603] Anonymous. An improved liquid holding vessel. Scientific American, 62(25):389, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
389b.pdf.

Anonymous:1890:IMHb

[604] Anonymous. An improved machine for hilling celery. Scientific American, 62(25):389, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
389a.pdf.

Anonymous:1890:ICPc

[605] Anonymous. An improved cotton picker. Scientific American, 62 (25):389–390, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
389e.pdf.

Anonymous:1890:DC

[606] Anonymous. The development of the caliper. Scientific American, 62(25):390, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
390.pdf.

Anonymous:1890:IMCd

[607] Anonymous. An improved metal cross tie. Scientific American, 62(25):390–391, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
390a.pdf.

Anonymous:1890:SGS

[608] Anonymous. Singer’s great sewing machine manufactory after the fire. Scientific American, 62(25):392, June 21, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n25/pdf/scientificamerican06211890-392.pdf.

Anonymous:1890:SME

[609] Anonymous. A subway manhole explosion followed by a gas fire. Scientific American, 62(25):393, June 21, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v62/
n25/pdf/scientificamerican06211890-393.pdf.

Anonymous:1890:WO

[610] Anonymous. Wall ornaments. Scientific American, 62(25):395, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-395.pdf.

Anonymous:1890:BPx

[611] Anonymous. Business and personal. Scientific American, 62(25):396, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-396a.pdf.

Anonymous:1890:RPIy

[612] Anonymous. Recently patented inventions. Scientific American, 62(25):396, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n25/pdf/scientificamerican06211890-
396.pdf.

Anonymous:1890:NQy

[613] Anonymous. Notes and queries. Scientific American, 62(25):396–397, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-396b.pdf.

Anonymous:1890:Ibi

[614] Anonymous. To inventors. Scientific American, 62(25):397, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-397.pdf.

Anonymous:1890:IIy

[615] Anonymous. Index of inventions. Scientific American, 62(25):397–398, June 21, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-397a.pdf.

Anonymous:1890:PNPq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n25/pdf/scientificamerican06211890-400.pdf.

Anonymous:1890:DE

[617] Anonymous. Disappearance of energy. Scientific American, 62(26):402, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n26/pdf/scientificamerican06281890-402.pdf.

Anonymous:1890:GDR

[618] Anonymous. Good doctrine in regard to patents. Scientific American, 62(26):402, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
402a.pdf.

Anonymous:1890:PPJb

[619] Anonymous. Position of the planets in July. Scientific American, 62(26):402–403, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
402b.pdf.

Anonymous:1890:ICCn

[620] Anonymous. An improved car coupling. Scientific American, 62 (26):404, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
404.pdf.

Anonymous:1890:IHe

[621] Anonymous. An improved hinge. Scientific American, 62(26):404, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n26/pdf/scientificamerican06281890-404b.pdf.

Anonymous:1890:IHSb

[622] Anonymous. An improved hoop shaving machine. Scientific American, 62(26):404, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
404a.pdf.

Anonymous:1890:MMA

[623] Anonymous. A machine for making articles from pulp. Scientific American, 62(26):404, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
404c.pdf.

Anonymous:1890:ICGb

[624] Anonymous. An improved collar for grinding pans. Scientific American, 62(26):404–405, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
404d.pdf.

Anonymous:1890:ICCo

[625] Anonymous. An improved car coupling. Scientific American, 62 (26):405, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
405.pdf.

Anonymous:1890:ICS

[626] Anonymous. Improved cement sinks, washtubs, etc. Scientific American, 62(26):405–406, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
405a.pdf.

Anonymous:1890:BEW

[627] Anonymous. The Berlin Electrical Works. Scientific American, 62(26):407, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
407.pdf.

Anonymous:1890:RP

[628] Anonymous. A residence at Portchester, N.Y. Scientific American, 62(26):408, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
408.pdf.

Anonymous:1890:STW

[629] Anonymous. Some things in wire. Scientific American, 62(26):409–410, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n26/pdf/scientificamerican06281890-409.pdf.

Anonymous:1890:SABj

[630] Anonymous. Scientific American building edition. Scientific American, 62(26):411, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
411b.pdf.

Anonymous:1890:BPy

[631] Anonymous. Business and personal. Scientific American, 62(26):411, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n26/pdf/scientificamerican06281890-411a.pdf.

Anonymous:1890:RPIz

[632] Anonymous. Recently patented inventions. Scientific American, 62(26):411, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v62/n26/pdf/scientificamerican06281890-
411.pdf.

Anonymous:1890:IIz

[633] Anonymous. Index of inventions. Scientific American, 62(26):411–412, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n26/pdf/scientificamerican06281890-411c.pdf.

Anonymous:1890:Ibj

[634] Anonymous. Index. Scientific American, 62(26):413–414, June 28, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v62/n26/pdf/scientificamerican06281890-413.pdf.

Anonymous:1890:MP

[635] Anonymous. M. Pasteur. Scientific American, 63(1):1–2, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-1.pdf.

Anonymous:1890:FC

[636] Anonymous. Fire crackers. Scientific American, 63(1):2–3, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-2.pdf.

Anonymous:1890:STC

[637] Anonymous. Speed trial of the cruiser Philadelphia. Scientific American, 63(1):3, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n1/pdf/scientificamerican07051890-
3.pdf.

Anonymous:1890:IIPa

[638] Anonymous. An improved ice plane. Scientific American, 63(1):4, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-4d.pdf.

Anonymous:1890:IRHb

[639] Anonymous. An improved rein holder. Scientific American, 63(1):4, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-4c.pdf.

Anonymous:1890:ISHa

[640] Anonymous. An improved snap hook. Scientific American, 63(1):4, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-4b.pdf.

Anonymous:1890:IWBa

[641] Anonymous. An improved wagon brake. Scientific American, 63(1):4, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-4.pdf.

Anonymous:1890:SDA

[642] Anonymous. Swift’s double action and hammerless revolvers. Scientific American, 63(1):4, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n1/pdf/scientificamerican07051890-
4a.pdf.

Anonymous:1890:IPD

[643] Anonymous. An improved potato digger. Scientific American, 63(1):5, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-5c.pdf.

Anonymous:1890:ITDa

[644] Anonymous. An improved tension device for belts. Scientific American, 63(1):5, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n1/pdf/scientificamerican07051890-
5a.pdf.

Anonymous:1890:NECb

[645] Anonymous. A neat and effective clothes beater. Scientific American, 63(1):5, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n1/pdf/scientificamerican07051890-
5.pdf.

Anonymous:1890:RAH

[646] Anonymous. Rife’s automatic hydraulic engine. Scientific American, 63(1):5, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n1/pdf/scientificamerican07051890-
5b.pdf.

Anonymous:1890:IWSc

[647] Anonymous. An improved wabble saw. Scientific American, 63(1):5–6, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-5d.pdf.

Anonymous:1890:IMe

[648] Anonymous. Imitation of Majolica. Scientific American, 63(1):6–7, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-6.pdf.

Anonymous:1890:ET

[649] Anonymous. An electric trap. Scientific American, 63(1):8, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-8a.pdf.

Anonymous:1890:WCH

[650] Anonymous. Walking on the ceiling head down. Scientific American, 63(1):8, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n1/pdf/scientificamerican07051890-
8.pdf.

Anonymous:1890:W

[651] Anonymous. The weaver. Scientific American, 63(1):9, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-9.pdf.

Anonymous:1890:SABk

[652] Anonymous. Scientific American building edition. Scientific American, 63(1):10, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n1/pdf/scientificamerican07051890-
10b.pdf.

Anonymous:1890:BPz

[653] Anonymous. Business and personal. Scientific American, 63(1):10, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-10c.pdf.

Anonymous:1890:NBPf

[654] Anonymous. New books and publications. Scientific American, 63(1): 10, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-10a.pdf.

Anonymous:1890:NQz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-10d.pdf.

Anonymous:1890:RPIba

[656] Anonymous. Recently patented inventions. Scientific American, 63(1): 10, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-10.pdf.

Anonymous:1890:IIba

[657] Anonymous. Index of inventions. Scientific American, 63(1):11–12, July 5, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n1/pdf/scientificamerican07051890-11.pdf.

Anonymous:1890:Ac

[658] Anonymous. Ammonia. Scientific American, 63(2):16, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-16.pdf.

Anonymous:1890:SSBa

[659] Anonymous. A summer school of biology. Scientific American, 63(2):16, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-16a.pdf.

Anonymous:1890:CWFa

[660] Anonymous. The Chicago World’s Fair. Scientific American, 63 (2):16–17, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
16b.pdf.

Anonymous:1890:TMD

[661] Anonymous. Trade mark decisions. Scientific American, 63(2):17, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-17.pdf.

Anonymous:1890:IBGa

[662] Anonymous. An imitation baseball game. Scientific American, 63(2):18, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-18b.pdf.

Anonymous:1890:IFPc

[663] Anonymous. An improved fruit picking stand. Scientific American, 63(2):18, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
18.pdf.

Anonymous:1890:ILL

[664] Anonymous. An improved log loading machine. Scientific American, 63(2):18, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
18c.pdf.

Anonymous:1890:ISLc

[665] Anonymous. An improved safety latch for elevator doors. Scientific American, 63(2):18, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
18a.pdf.

Anonymous:1890:NPDb

[666] Anonymous. A new plating dynamo. Scientific American, 63(2):18, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-18d.pdf.

Anonymous:1890:IDT

[667] Anonymous. An improved dental tool. Scientific American, 63(2):19, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-19b.pdf.

Anonymous:1890:IGH

[668] Anonymous. An improved gravity hoist. Scientific American, 63(2):19, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-19c.pdf.

Anonymous:1890:IHBa

[669] Anonymous. An improved header brake. Scientific American, 63(2):19, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-19.pdf.

Anonymous:1890:IMPb

[670] Anonymous. An improved mail pouch catcher. Scientific American, 63(2):19, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
19d.pdf.

Anonymous:1890:ISPc

[671] Anonymous. An improved Stone polishing machine. Scientific American, 63(2):19, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
19a.pdf.

Anonymous:1890:IAR

[672] Anonymous. An improved automatic regulator. Scientific American, 63(2):19–20, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
19e.pdf.

Anonymous:1890:HMO

[673] Anonymous. Home made ornaments. Scientific American, 63(2):20, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-20.pdf.

Anonymous:1890:SBB

[674] Anonymous. The Sodom and Bog Brook Reservoirs in the Croton Basin, N.Y. Scientific American, 63(2):21, July 12, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n2/pdf/scientificamerican07121890-21.pdf.

Anonymous:1890:MCM

[675] Anonymous. Manufacture of condensed milk in Switzerland. Scientific American, 63(2):22, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
22.pdf.

Anonymous:1890:NMB

[676] Anonymous. The new Merchants’ Bridge at St. Louis. Scientific American, 63(2):22–23, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
22a.pdf.

Anonymous:1890:BSU

[677] Anonymous. Burrowing sea urchins. Scientific American, 63(2):23, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-23.pdf.

Anonymous:1890:PNh

[678] Anonymous. Photographic notes. Scientific American, 63(2):24–25, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-24.pdf.

Anonymous:1890:SABl

[679] Anonymous. Scientific American building edition. Scientific American, 63(2):26, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
26b.pdf.

Anonymous:1890:BPba

[680] Anonymous. Business and personal. Scientific American, 63(2):26, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-26c.pdf.

Anonymous:1890:NBPg

[681] Anonymous. New books and publications. Scientific American, 63(2):26, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
26a.pdf.

Anonymous:1890:RPIbb

[682] Anonymous. Recently patented inventions. Scientific American, 63(2):26, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n2/pdf/scientificamerican07121890-
26.pdf.

Anonymous:1890:NQba

[683] Anonymous. Notes and queries. Scientific American, 63(2):26–27, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-26d.pdf.

Anonymous:1890:IIbb

[684] Anonymous. Index of inventions. Scientific American, 63(2):27–28, July 12, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n2/pdf/scientificamerican07121890-27.pdf.

Anonymous:1890:CNA

[685] Anonymous. Cost of the new aqueduct. Scientific American, 63(3):32, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n3/pdf/scientificamerican07191890-32.pdf.

Anonymous:1890:PCBa

[686] Anonymous. The Panama Canal bubble. Scientific American, 63(3):32–33, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
32a.pdf.

Anonymous:1890:HAS

[687] Anonymous. A hemming attachment for sewing machines. Scientific American, 63(3):34, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
34a.pdf.

Anonymous:1890:IGS

[688] Anonymous. An improved guide for sewing machines. Scientific American, 63(3):34, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
34.pdf.

Anonymous:1890:HWD

[689] Anonymous. Hot water to drive cattle from tracks. Scientific American, 63(3):35, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
35.pdf.

Anonymous:1890:IBHc

[690] Anonymous. An improved bridle for horses. Scientific American, 63(3):35, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
35d.pdf.

Anonymous:1890:IFS

[691] Anonymous. An improved folding scaffold bracket. Scientific American, 63(3):35, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
35a.pdf.

Anonymous:1890:IGI

[692] Anonymous. An improved garden implement. Scientific American, 63(3):35, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
35c.pdf.

Anonymous:1890:SAT

[693] Anonymous. A solar attachment for transit instruments. Scientific American, 63(3):35, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
35b.pdf.

Anonymous:1890:NCA

[694] Anonymous. The new Croton Aqueduct — the Harlem River Siphon and Pumping Apparatus. Scientific American, 63(3):36, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n3/pdf/scientificamerican07191890-36.pdf.

Anonymous:1890:HMG

[695] Anonymous. Home made grills and gratings. Scientific American, 63(3):37, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
37.pdf.

Anonymous:1890:MPS

[696] Anonymous. The Mcleod pneumatic system of heating, ventilating and cooling, and improved subway construction. Scientific American, 63(3):38, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
38.pdf.

Anonymous:1890:G

[697] Anonymous. The gramophone. Scientific American, 63(3):39, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n3/pdf/scientificamerican07191890-39.pdf.

Anonymous:1890:NYC

[698] Anonymous. New York City aqueduct — its engineering features and design. Scientific American, 63(3):40–41, July 19, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n3/pdf/scientificamerican07191890-40.pdf.

Anonymous:1890:SABm

[699] Anonymous. Scientific American building edition. Scientific American, 63(3):42, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
42b.pdf.

Anonymous:1890:NBPh

[700] Anonymous. New books and publications. Scientific American, 63(3):42, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
42a.pdf.

Anonymous:1890:RPIbc

[701] Anonymous. Recently patented inventions. Scientific American, 63(3):42, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n3/pdf/scientificamerican07191890-
42.pdf.

Anonymous:1890:BPbb

[702] Anonymous. Business and personal. Scientific American, 63(3):43, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n3/pdf/scientificamerican07191890-43.pdf.

Anonymous:1890:Ibk

[703] Anonymous. To inventors. Scientific American, 63(3):43, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n3/pdf/scientificamerican07191890-43b.pdf.

Anonymous:1890:NQbb

[704] Anonymous. Notes and queries. Scientific American, 63(3):43, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n3/pdf/scientificamerican07191890-43a.pdf.

Anonymous:1890:IIbc

[705] Anonymous. Index of inventions. Scientific American, 63(3):43–44, July 19, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n3/pdf/scientificamerican07191890-43c.pdf.

Anonymous:1890:ESE

[706] Anonymous. Experiments with the Sims–Edison torpedo. Scientific American, 63(4):47–48, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
47.pdf.

Anonymous:1890:IPAa

[707] Anonymous. It pays to advertise in the Scientific American. Scientific American, 63(4):48, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
48.pdf.

Anonymous:1890:GMM

[708] Anonymous. The gypsy moth in Massachusetts. Scientific American, 63(4):48–49, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
48a.pdf.

Anonymous:1890:BTMa

[709] Anonymous. Bolt threading machine. Scientific American, 63(4):50, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-50a.pdf.

Anonymous:1890:ISAa

[710] Anonymous. An improved screen attachment for Windows. Scientific American, 63(4):50, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
50c.pdf.

Anonymous:1890:ISIb

[711] Anonymous. An improved Surveyor’s instrument. Scientific American, 63(4):50, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
50b.pdf.

Anonymous:1890:IUP

[712] Anonymous. An improved umbrella or parasol exhibitor. Scientific American, 63(4):50, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
50.pdf.

Anonymous:1890:HRP

[713] Anonymous. A hand rest for penmen. Scientific American, 63(4):51, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-51.pdf.

Anonymous:1890:IAAa

[714] Anonymous. An improved adjustable axle nut. Scientific American, 63(4):51, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
51c.pdf.

Anonymous:1890:IBCb

[715] Anonymous. An improved blasting car. Scientific American, 63(4):51, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-51b.pdf.

Anonymous:1890:IBCc

[716] Anonymous. An improved block and clevis for log wagons. Scientific American, 63(4):51, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
51d.pdf.

Anonymous:1890:IFT

[717] Anonymous. An improved folding table. Scientific American, 63(4):51, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-51a.pdf.

Anonymous:1890:Ct

[718] Anonymous. Correspondence. Scientific American, 63(4):52–54, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-52.pdf.

Anonymous:1890:Cu

[719] Anonymous. The Chicago. Scientific American, 63(4):54, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-54.pdf.

Anonymous:1890:RLC

[720] Anonymous. Redwood logging in California. Scientific American, 63(4):55–56, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
55.pdf.

Anonymous:1890:WS

[721] Anonymous. A wrinkle in sawing. Scientific American, 63(4):57, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-57a.pdf.

Hopkins:1890:SUPc

[722] Geo. M. Hopkins. The scientific use of the phonograph. Scientific American, 63(4):57, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
57.pdf.

Anonymous:1890:BPbc

[723] Anonymous. Business and personal. Scientific American, 63(4):58, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-58b.pdf.

Anonymous:1890:NBPi

[724] Anonymous. New books and publications. Scientific American, 63(4):58, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
58a.pdf.

Anonymous:1890:RPIbd

[725] Anonymous. Recently patented inventions. Scientific American, 63(4):58, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n4/pdf/scientificamerican07261890-
58.pdf.

Anonymous:1890:NQbc

[726] Anonymous. Notes and queries. Scientific American, 63(4):58–59, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-58c.pdf.

Anonymous:1890:Ibl

[727] Anonymous. To inventors. Scientific American, 63(4):59, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-59a.pdf.

Anonymous:1890:PNPr

[728] Anonymous. Patent notes: Patents! Scientific American, 63(4):59, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-59.pdf.

Anonymous:1890:IIbd

[729] Anonymous. Index of inventions. Scientific American, 63(4):59–60, July 26, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n4/pdf/scientificamerican07261890-59b.pdf.

Anonymous:1890:TF

[730] Anonymous. A telegraphic fire. Scientific American, 63(5):64, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-64.pdf.

Anonymous:1890:PPAb

[731] Anonymous. Position of the planets in August. Scientific American, 63(5):65, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
65.pdf.

Anonymous:1890:IDH

[732] Anonymous. Improved dwelling houses. Scientific American, 63(5):66, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-66b.pdf.

Anonymous:1890:NFTa

[733] Anonymous. New folding table. Scientific American, 63(5):66, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-66a.pdf.

Anonymous:1890:NLA

[734] Anonymous. New locking attachment for bicycles. Scientific American, 63(5):66, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
66c.pdf.

Anonymous:1890:NT

[735] Anonymous. A novel toy. Scientific American, 63(5):66, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-66.pdf.

Anonymous:1890:ESA

[736] Anonymous. Electric slaughtering apparatus. Scientific American, 63(5):66–67, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
66d.pdf.

Anonymous:1890:IFFb

[737] Anonymous. Instrument for fastening fuse caps. Scientific American, 63(5):67, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
67.pdf.

Anonymous:1890:NSE

[738] Anonymous. Novel steam engine. Scientific American, 63(5):67, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-67b.pdf.

Anonymous:1890:NWB

[739] Anonymous. New wash boiler. Scientific American, 63(5):67, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-67a.pdf.

Anonymous:1890:IHg

[740] Anonymous. An improvement in halters. Scientific American, 63(5):67–68, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
67c.pdf.

Anonymous:1890:NSBb

[741] Anonymous. A new steam boiler. Scientific American, 63(5):68, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-68.pdf.

Anonymous:1890:Cv

[742] Anonymous. Correspondence. Scientific American, 63(5):69, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-69a.pdf.

Anonymous:1890:CBa

[743] Anonymous. Comet Brooks 1 of 1890. Scientific American, 63(5):69, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-69b.pdf.

Anonymous:1890:IMDb

[744] Anonymous. Improved movable dock crane. Scientific American, 63(5):69, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
69.pdf.

Anonymous:1890:CWFb

[745] Anonymous. The Chicago World’s Fair. Scientific American, 63 (5):70–71, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
70.pdf.

Anonymous:1890:MMW

[746] Anonymous. Mast of a modern war vessel. Scientific American, 63(5):71–73, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
71.pdf.

Anonymous:1890:COa

[747] Anonymous. Crystallized ornaments. Scientific American, 63(5):73, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-73.pdf.

Anonymous:1890:BPbd

[748] Anonymous. Business and personal. Scientific American, 63(5):74, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-74a.pdf.

Anonymous:1890:RPIbe

[749] Anonymous. Recently patented inventions. Scientific American, 63(5):74, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
74.pdf.

Anonymous:1890:NQbd

[750] Anonymous. Notes and queries. Scientific American, 63(5):74–75, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-74b.pdf.

Anonymous:1890:BNBk

[751] Anonymous. Books: New books and publications. Scientific American, 63(5):75, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n5/pdf/scientificamerican08021890-
75.pdf.

Anonymous:1890:Ibm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-75a.pdf.

Anonymous:1890:IIbe

[753] Anonymous. Index of inventions. Scientific American, 63(5):75–76, August 2, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n5/pdf/scientificamerican08021890-75b.pdf.

Anonymous:1890:FI

[754] Anonymous. Fire Island. Scientific American, 63(6):80–81, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-80.pdf.

Anonymous:1890:FEA

[755] Anonymous. Fire extinguishing apparatus. Scientific American, 63(6):82, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
82b.pdf.

Anonymous:1890:ISDb

[756] Anonymous. An improved safety device for elevators. Scientific American, 63(6):82, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
82c.pdf.

Anonymous:1890:ISMc

[757] Anonymous. An improved steam motor. Scientific American, 63(6):82, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-82a.pdf.

Anonymous:1890:NMR

[758] Anonymous. New metallic railway tie. Scientific American, 63(6):82, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-82.pdf.

Anonymous:1890:DAF

[759] Anonymous. A drying apparatus for fish, etc. Scientific American, 63(6):83, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
83b.pdf.

Anonymous:1890:DSR

[760] Anonymous. A device to stop runaway horses. Scientific American, 63(6):83, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
83.pdf.

Anonymous:1890:NCL

[761] Anonymous. Novel clothes line support. Scientific American, 63(6):83, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-83a.pdf.

Anonymous:1890:IGD

[762] Anonymous. An improved grinding device. Scientific American, 63(6):83–85, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
83c.pdf.

Anonymous:1890:MSTb

[763] Anonymous. A machine to supersede typesetting. Scientific American, 63(6):85, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
85.pdf.

Anonymous:1890:GLG

[764] Anonymous. George Lincoln Goodale. Scientific American, 63(6):86, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-86.pdf.

Anonymous:1890:GRT

[765] Anonymous. The Great Railway Tunnel under the St. Clair River, between the United States and Canada. Scientific American, 63(6):87–88, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-87.pdf.

Anonymous:1890:Cx

[766] Anonymous. Correspondence. Scientific American, 63(6):89, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-89.pdf.

Anonymous:1890:IRCb

[767] Anonymous. An improved register or counter. Scientific American, 63(6):89, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
89b.pdf.

Anonymous:1890:NVN

[768] Anonymous. New Victor no. 0 electroplating dynamo. Scientific American, 63(6):89, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
89a.pdf.

Anonymous:1890:BPbe

[769] Anonymous. Business and personal. Scientific American, 63(6):90, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-90a.pdf.

Anonymous:1890:RPIbf

[770] Anonymous. Recently patented inventions. Scientific American, 63(6):90, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
90.pdf.

Anonymous:1890:NQbe

[771] Anonymous. Notes and queries. Scientific American, 63(6):90–91, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-90b.pdf.

Anonymous:1890:BNBl

[772] Anonymous. Books: New books and publications. Scientific American, 63(6):91, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n6/pdf/scientificamerican08091890-
91.pdf.

Anonymous:1890:Ibn

[773] Anonymous. To inventors. Scientific American, 63(6):91, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-91a.pdf.

Anonymous:1890:IIbf

[774] Anonymous. Index of inventions. Scientific American, 63(6):91–92, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-91b.pdf.

Anonymous:1890:PNPs

[775] Anonymous. Patent notes: Patents. Scientific American, 63(6):94, August 9, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n6/pdf/scientificamerican08091890-94.pdf.

Anonymous:1890:NECc

[776] Anonymous. New Elevated Cable Street Railway, Jersey City. Scientific American, 63(7):95, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
95.pdf.

Anonymous:1890:FEE

[777] Anonymous. The first electrical execution. Scientific American, 63(7):96, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
96.pdf.

Anonymous:1890:WEa

[778] Anonymous. Wire and electricity. Scientific American, 63(7):96–97, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-96a.pdf.

Anonymous:1890:WEb

[779] Anonymous. Wire and electricity. Scientific American, 63(7):96–97, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-96b.pdf.

Anonymous:1890:PNi

[780] Anonymous. Photographic notes. Scientific American, 63(7):97, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-97.pdf.

Anonymous:1890:IEAa

[781] Anonymous. An improved educational appliance. Scientific American, 63(7):98, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
98c.pdf.

Anonymous:1890:NBS

[782] Anonymous. Novel blind stop. Scientific American, 63(7):98, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-98.pdf.

Anonymous:1890:NCS

[783] Anonymous. A new car starter. Scientific American, 63(7):98, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-98b.pdf.

Anonymous:1890:NFM

[784] Anonymous. New form of millstone pick. Scientific American, 63(7):98, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
98a.pdf.

Anonymous:1890:NTPc

[785] Anonymous. New tubular packing for steam condenser tubes. Scientific American, 63(7):98, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
98d.pdf.

Anonymous:1890:GG

[786] Anonymous. The Giffard gun. Scientific American, 63(7):99, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-99.pdf.

Hopkins:1890:SUPd

[787] Geo. M. Hopkins. The scientific use of the phonograph. Scientific American, 63(7):100, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
100.pdf.

Anonymous:1890:Cy

[788] Anonymous. Correspondence. Scientific American, 63(7):101, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-101.pdf.

Anonymous:1890:NUP

[789] Anonymous. New umbrella or parasol exhibitor. Scientific American, 63(7):101, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
101a.pdf.

Anonymous:1890:TTP

[790] Anonymous. Torpedo target practice at Kiel. Scientific American, 63(7):102–104, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
102.pdf.

Anonymous:1890:GS

[791] Anonymous. The garter snake. Scientific American, 63(7):105–106, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-105.pdf.

Anonymous:1890:BPbf

[792] Anonymous. Business and personal. Scientific American, 63(7):107, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-107a.pdf.

Anonymous:1890:Ibo

[793] Anonymous. To inventors. Scientific American, 63(7):107, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-107b.pdf.

Anonymous:1890:RPIbg

[794] Anonymous. Recently patented inventions. Scientific American, 63(7):107, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
107.pdf.

Anonymous:1890:IIbg

[795] Anonymous. Index of inventions. Scientific American, 63(7): 107–108, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
107d.pdf.

Anonymous:1890:IIbh

[796] Anonymous. Index of inventions. Scientific American, 63(7): 107–109, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
107c.pdf.

Anonymous:1890:SAPb

[797] Anonymous. The Scientific American publications for 1890. Scientific American, 63(7):109, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
109b.pdf.

Anonymous:1890:BMa

[798] Anonymous. To business men. Scientific American, 63(7):109, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-109.pdf.

Anonymous:1890:NR

[799] Anonymous. Now ready. Scientific American, 63(7):109, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-109a.pdf.

Anonymous:1890:SAEb

[800] Anonymous. The Scientific American established 1846. Scientific American, 63(7):110, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
110a.pdf.

Anonymous:1890:SASa

[801] Anonymous. The Scientific American supplement. Scientific American, 63(7):110, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n7/pdf/scientificamerican08161890-
110b.pdf.

Anonymous:1890:BEa

[802] Anonymous. Building edition. Scientific American, 63(7):110, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-110c.pdf.

Anonymous:1890:PNPt

[803] Anonymous. Patent notes: Patents. Scientific American, 63(7):110, August 16, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n7/pdf/scientificamerican08161890-110.pdf.

Anonymous:1890:NFA

[804] Anonymous. The new French armored war ship Hoche. Scientific American, 63(8):111–112, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
111.pdf.

Anonymous:1890:SDb

[805] Anonymous. Seaport defenses. Scientific American, 63(8):112–113, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-112.pdf.

Anonymous:1890:SDc

[806] Anonymous. Seaport defenses. Scientific American, 63(8):112–113, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-112a.pdf.

Anonymous:1890:CCP

[807] Anonymous. A ceiling center-piece and ventilator. Scientific American, 63(8):114, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
114b.pdf.

Anonymous:1890:ICOb

[808] Anonymous. An improved can opener. Scientific American, 63 (8):114, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
114d.pdf.

Anonymous:1890:ITCd

[809] Anonymous. Improved thill coupling. Scientific American, 63 (8):114, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
114.pdf.

Anonymous:1890:IWCa

[810] Anonymous. An improved wire or cable tightener. Scientific American, 63(8):114, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
114c.pdf.

Anonymous:1890:MJC

[811] Anonymous. A metallic joint or coupling for wheels and other uses. Scientific American, 63(8):114, August 23, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n8/pdf/scientificamerican08231890-114a.pdf.

Anonymous:1890:IRTb

[812] Anonymous. An improved railroad track. Scientific American, 63(8):115, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
115.pdf.

Anonymous:1890:IVT

[813] Anonymous. An improved vehicle top support. Scientific American, 63(8):115, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
115b.pdf.

Anonymous:1890:PPNa

[814] Anonymous. A proposed plan for non-sinkable ships. Scientific American, 63(8):115, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
115a.pdf.

Anonymous:1890:ISF

[815] Anonymous. Improved steam food cooker. Scientific American, 63 (8):115–116, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
115c.pdf.

Anonymous:1890:OE

[816] Anonymous. Ornamental electroliers. Scientific American, 63 (8):116, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
116.pdf.

Anonymous:1890:Cz

[817] Anonymous. Correspondence. Scientific American, 63(8):117, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-117.pdf.

Anonymous:1890:EME

[818] Anonymous. The early mission establishments in California. Scientific American, 63(8):118, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
118.pdf.

Anonymous:1890:GWB

[819] Anonymous. A gas well blow-out at Pulaski, N.Y. Scientific American, 63(8):119, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
119.pdf.

Anonymous:1890:PNj

[820] Anonymous. Photographic notes. Scientific American, 63(8):120, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-120.pdf.

Anonymous:1890:HFM

[821] Anonymous. Handling fish to make fertilizers. Scientific American, 63(8):121, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
121.pdf.

Anonymous:1890:BPbg

[822] Anonymous. Business and personal. Scientific American, 63(8):122, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-122a.pdf.

Anonymous:1890:RPIbh 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
122.pdf.

Anonymous:1890:NQbf

[824] Anonymous. Notes and queries. Scientific American, 63(8):122– 123, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
122b.pdf.

Anonymous:1890:BNBm

[825] Anonymous. Books: New books and publications. Scientific American, 63(8):123, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
123.pdf.

Anonymous:1890:Ibq

[826] Anonymous. To inventors. Scientific American, 63(8):123, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-123a.pdf.

Anonymous:1890:IIbi

[827] Anonymous. Index of inventions. Scientific American, 63(8): 123–124, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
123c.pdf.

Anonymous:1890:IIbj

[828] Anonymous. Index of inventions. Scientific American, 63(8): 123–125, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
123b.pdf.

Anonymous:1890:BMb

[829] Anonymous. To business men. Scientific American, 63(8):124, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-124.pdf.

Anonymous:1890:ABa

[830] Anonymous. Architectural books. Scientific American, 63(8):125, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-125.pdf.

Anonymous:1890:SAPc

[831] Anonymous. The Scientific American publications for 1890. Scientific American, 63(8):125, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
125d.pdf.

Anonymous:1890:FPT

[832] Anonymous. Foreign patents their cost reduced. Scientific American, 63(8):125, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
125b.pdf.

Anonymous:1890:IMg

[833] Anonymous. To inventors and manufacturers. Scientific American, 63(8):125, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
125a.pdf.

Hopkins:1890:NR

[834] Geo. M. Hopkins. Now ready. Scientific American, 63(8):125, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-125c.pdf.

Anonymous:1890:SAEc

[835] Anonymous. The Scientific American established 1846. Scientific American, 63(8):126, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
126a.pdf.

Anonymous:1890:SASb

[836] Anonymous. The Scientific American supplement. Scientific American, 63(8):126, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n8/pdf/scientificamerican08231890-
126b.pdf.

Anonymous:1890:BEb

[837] Anonymous. Building edition. Scientific American, 63(8):126, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-126c.pdf.

Anonymous:1890:PNPu

[838] Anonymous. Patent notes: Patents. Scientific American, 63(8):126, August 23, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n8/pdf/scientificamerican08231890-126.pdf.

Anonymous:1890:NPL

[839] Anonymous. The new patent law of Mexico. Scientific American, 63(9):128, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
128.pdf.

Anonymous:1890:PPS

[840] Anonymous. Position of the planets in September. Scientific American, 63(9):128, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
128a.pdf.

Anonymous:1890:JE

[841] Anonymous. John Ericsson. Scientific American, 63(9):128–129, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n9/pdf/scientificamerican08301890-128b.pdf.

Anonymous:1890:PCM

[842] Anonymous. The Photographers’ Convention — a memorial to Daguerre. Scientific American, 63(9):129, August 30, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n9/pdf/scientificamerican08301890-129.pdf.

Anonymous:1890:IRGb

[843] Anonymous. An improved rein grip. Scientific American, 63(9):130, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n9/pdf/scientificamerican08301890-130b.pdf.

Anonymous:1890:IRTc

[844] Anonymous. An improved railroad tie. Scientific American, 63 (9):130, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
130a.pdf.

Anonymous:1890:ISEb

[845] Anonymous. An improved shifting eccentric. Scientific American, 63(9):130, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
130.pdf.

Anonymous:1890:ICPd

[846] Anonymous. An improved curtain-pole ring and pin. Scientific American, 63(9):130–131, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
130c.pdf.

Anonymous:1890:IFU

[847] Anonymous. An improved folding umbrella. Scientific American, 63(9):131, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
131.pdf.

Anonymous:1890:ISPd

[848] Anonymous. An improved sled propeller. Scientific American, 63(9):131, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
131a.pdf.

Anonymous:1890:CUSa

[849] Anonymous. The census of the United States. Scientific American, 63(9):132–133, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
132.pdf.

Anonymous:1890:BTMb

[850] Anonymous. A bull with two mouths. Scientific American, 63 (9):134, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
134.pdf.

Anonymous:1890:EHS

[851] Anonymous. Electrically heated sad Irons. Scientific American, 63(9):134, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
134a.pdf.

Anonymous:1890:ROP

[852] Anonymous. Reproduction of an old Pompeian residence. Scientific American, 63(9):135–137, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
135.pdf.

Anonymous:1890:JPC

[853] Anonymous. Japanese portiere or curtain. Scientific American, 63(9):137, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
137.pdf.

Anonymous:1890:BPbh

[854] Anonymous. Business and personal. Scientific American, 63(9):138, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n9/pdf/scientificamerican08301890-138a.pdf.

Anonymous:1890:RPIbi

[855] Anonymous. Recently patented inventions. Scientific American, 63(9):138, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
138.pdf.

Anonymous:1890:NQbg

[856] Anonymous. Notes and queries. Scientific American, 63(9):138– 139, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
138b.pdf.

Anonymous:1890:Ibr

[857] Anonymous. To inventors. Scientific American, 63(9):139, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n9/pdf/scientificamerican08301890-139.pdf.

Anonymous:1890:NQIa

[858] Anonymous. Notes and queries: Index of inventions. Scientific American, 63(9):139–140, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n9/pdf/scientificamerican08301890-
139a.pdf.

Anonymous:1890:PNPv

[859] Anonymous. Patent notes: Patents. Scientific American, 63(9):142, August 30, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n9/pdf/scientificamerican08301890-142.pdf.

Anonymous:1890:ERT

[860] Anonymous. Ericsson’s revolving turreted war ship. Scientific American, 63(10):144–145, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
144.pdf.

Anonymous:1890:T

[861] Anonymous. Tyrotoxicon. Scientific American, 63(10):145–146, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n10/pdf/scientificamerican09061890-145.pdf.

Anonymous:1890:ICV

[862] Anonymous. An improved culinary vessel. Scientific American, 63(10):146, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
146b.pdf.

Anonymous:1890:ITKa

[863] Anonymous. An improved telegraph key. Scientific American, 63 (10):146, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
146.pdf.

Anonymous:1890:SSS

[864] Anonymous. A separable spindle for spinning wool. Scientific American, 63(10):146, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
146c.pdf.

Anonymous:1890:WHT

[865] Anonymous. A work-holding table for carving machines. Scientific American, 63(10):146, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
146a.pdf.

Anonymous:1890:CTM

[866] Anonymous. A copyholder for typewriting machines. Scientific American, 63(10):147, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
147a.pdf.

Anonymous:1890:IMRa

[867] Anonymous. An improved metal railroad tie. Scientific American, 63(10):147, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
147b.pdf.

Anonymous:1890:IWd

[868] Anonymous. An improved wrench. Scientific American, 63(10): 147, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
147c.pdf.

Anonymous:1890:TEB

[869] Anonymous. A tube expander and bender. Scientific American, 63(10):147, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
147.pdf.

Anonymous:1890:DBS

[870] Anonymous. Departure of the Baltimore for Sweden. Scientific American, 63(10):148, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
148.pdf.

Hovey:1890:IMA

[871] H. Hovey. Indianapolis Meeting of the American Association for the Advancement of Science. Scientific American, 63(10):148–149, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n10/pdf/scientificamerican09061890-148a.pdf.

Anonymous:1890:Cbb

[872] Anonymous. Correspondence. Scientific American, 63(10):149, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n10/pdf/scientificamerican09061890-149.pdf.

Anonymous:1890:CHF

[873] Anonymous. Christian Henry Frederick Peters. Scientific American, 63(10):150, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
150.pdf.

Anonymous:1890:JPS

[874] Anonymous. A jet-propelled steam lifeboat. Scientific American, 63 (10):150–151, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
150a.pdf.

Anonymous:1890:ERC

[875] Anonymous. The electric race course. Scientific American, 63 (10):151–152, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
151.pdf.

Anonymous:1890:PNk

[876] Anonymous. Photographic notes. Scientific American, 63(10): 152–153, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
152.pdf.

Anonymous:1890:ING

[877] Anonymous. An improved nutmeg grater. Scientific American, 63(10):153, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
153a.pdf.

Anonymous:1890:ISWa

[878] Anonymous. An improved switch worker. Scientific American, 63 (10):153, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
153.pdf.

Anonymous:1890:IBSd

[879] Anonymous. An improved Bob sled. Scientific American, 63(10): 153–154, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
153b.pdf.

Anonymous:1890:BNBn

[880] Anonymous. Books: New books and publications. Scientific American, 63(10):155, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
155a.pdf.

Anonymous:1890:BPbi

[881] Anonymous. Business and personal. Scientific American, 63 (10):155, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
155b.pdf.

Anonymous:1890:Ibu

[882] Anonymous. To inventors. Scientific American, 63(10):155, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n10/pdf/scientificamerican09061890-155c.pdf.

Anonymous:1890:RPIbj

[883] Anonymous. Recently patented inventions. Scientific American, 63(10):155, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
155.pdf.

Anonymous:1890:PNPw

[884] Anonymous. Patent notes: Patents. Scientific American, 63 (10):158, September 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n10/pdf/scientificamerican09061890-
158.pdf.

Anonymous:1890:ITMb

[885] Anonymous. Important tests to be made of heavy armor plating. Scientific American, 63(11):160–162, September 13, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n11/pdf/scientificamerican09131890-160.pdf.

Anonymous:1890:ALW

[886] Anonymous. An artificial lake and water slide. Scientific American, 63(11):162, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
162.pdf.

Anonymous:1890:CSC

[887] Anonymous. A convenient sponge for cleaning slates. Scientific American, 63(11):162, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
162b.pdf.

Anonymous:1890:ILK

[888] Anonymous. An improvement in locks and keys. Scientific American, 63(11):162, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
162a.pdf.

Anonymous:1890:CER

[889] Anonymous. A coal elevator for railway service. Scientific American, 63(11):163, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
163a.pdf.

Anonymous:1890:IKS

[890] Anonymous. An improved knife sharpener and polisher. Scientific American, 63(11):163, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
163b.pdf.

Anonymous:1890:ISV

[891] Anonymous. An improved slide valve for engines. Scientific American, 63(11):163, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
163.pdf.

Anonymous:1890:MGS

[892] Anonymous. The meeting of the Great Shields of the St. Clair River Railway Tunnel. Scientific American, 63(11):164, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n11/pdf/scientificamerican09131890-164.pdf.

Anonymous:1890:PNl

[893] Anonymous. Photographic notes. Scientific American, 63(11): 164, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
164a.pdf.

Anonymous:1890:Cbc

[894] Anonymous. Correspondence. Scientific American, 63(11):165, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n11/pdf/scientificamerican09131890-165.pdf.

Anonymous:1890:WAE

[895] Anonymous. Work of amateur electricians. Scientific American, 63(11):165, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
165a.pdf.

Anonymous:1890:NFTb

[896] Anonymous. New form of towing steamers. Scientific American, 63(11):166, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
166a.pdf.

Anonymous:1890:SHC

[897] Anonymous. A swinging hydraulic capstan. Scientific American, 63(11):166, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
166.pdf.

Anonymous:1890:KLS

[898] Anonymous. Killing large sunfish with firearms. Scientific American, 63(11):167, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
167a.pdf.

Pike:1890:KN

[899] Nicolas Pike. The kola nut. Scientific American, 63(11):167, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n11/pdf/scientificamerican09131890-167.pdf.

Anonymous:1890:EPW

[900] Anonymous. Engine pounding — ways of stopping it. Scientific American, 63(11):168, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
168.pdf.

Anonymous:1890:SUC

[901] Anonymous. The scientific use of common things. Scientific American, 63(11):169–170, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
169.pdf.

Anonymous:1890:NQbh

[902] Anonymous. Notes and queries. Scientific American, 63(11): 170, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
170.pdf.

Anonymous:1890:BPbj

[903] Anonymous. Business and personal. Scientific American, 63 (11):171, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
171b.pdf.

Anonymous:1890:Ibv

[904] Anonymous. To inventors. Scientific American, 63(11):171, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n11/pdf/scientificamerican09131890-171a.pdf.

Anonymous:1890:RPIbk

[905] Anonymous. Recently patented inventions. Scientific American, 63 (11):171, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
171.pdf.

Anonymous:1890:IIbk

[906] Anonymous. Index of inventions. Scientific American, 63(11): 171–172, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
171c.pdf.

Anonymous:1890:PNPx

[907] Anonymous. Patent notes: Patents. Scientific American, 63 (11):174, September 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n11/pdf/scientificamerican09131890-
174.pdf.

Anonymous:1890:RIP

[908] Anonymous. Recent improvements in the pneumatic dynamite gun. Scientific American, 63(12):175, September 20, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n12/pdf/scientificamerican09201890-175.pdf.

Anonymous:1890:SN

[909] Anonymous. The smoke nuisance. Scientific American, 63(12): 176–177, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
176.pdf.

Anonymous:1890:DMD

[910] Anonymous. A device for manipulating doors and Windows. Scientific American, 63(12):178, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
178.pdf.

Anonymous:1890:EUB

[911] Anonymous. An eraser for use on blackboards. Scientific American, 63(12):178, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
178c.pdf.

Anonymous:1890:IFc

[912] Anonymous. An improved fumigator. Scientific American, 63 (12):178, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
178a.pdf.

Anonymous:1890:ILCc

[913] Anonymous. An improved lawn cleaner. Scientific American, 63 (12):178, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
178b.pdf.

Anonymous:1890:ISBd

[914] Anonymous. Improved steam boiler. Scientific American, 63 (12):178, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
178d.pdf.

Anonymous:1890:IMMb

[915] Anonymous. An improved mechanical movement. Scientific American, 63(12):179, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
179a.pdf.

Anonymous:1890:ISIc

[916] Anonymous. An improved safety indicator for elevators. Scientific American, 63(12):179, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
179b.pdf.

Anonymous:1890:RRR

[917] Anonymous. A register and recorder for revolving shafts. Scientific American, 63(12):179, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
179.pdf.

Anonymous:1890:ILT

[918] Anonymous. An improved lifting truss girder. Scientific American, 63 (12):179–180, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
179c.pdf.

Anonymous:1890:Cbd

[919] Anonymous. Correspondence. Scientific American, 63(12):181, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n12/pdf/scientificamerican09201890-181.pdf.

Anonymous:1890:DTP

[920] Anonymous. Daisies and thistles proscribed by law. Scientific American, 63(12):183, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
183.pdf.

Anonymous:1890:NGAb

[921] Anonymous. The new German armored war ship Siegfried. Scientific American, 63(12):185, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
185.pdf.

Anonymous:1890:PNm

[922] Anonymous. Photographic notes. Scientific American, 63(12): 186, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
186.pdf.

Anonymous:1890:BPbk

[923] Anonymous. Business and personal. Scientific American, 63 (12):187, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
187a.pdf.

Anonymous:1890:Ibx

[924] Anonymous. To inventors. Scientific American, 63(12):187, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n12/pdf/scientificamerican09201890-187c.pdf.

Anonymous:1890:NQbi

[925] Anonymous. Notes and queries. Scientific American, 63(12): 187, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
187b.pdf.

Anonymous:1890:RPIbl

[926] Anonymous. Recently patented inventions. Scientific American, 63 (12):187, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
187.pdf.

Anonymous:1890:IIbl

[927] Anonymous. Index of inventions. Scientific American, 63(12): 187–189, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
187d.pdf.

Anonymous:1890:PNPy

[928] Anonymous. Patent notes: Patents. Scientific American, 63 (12):190, September 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n12/pdf/scientificamerican09201890-
190.pdf.

Anonymous:1890:EES

[929] Anonymous. An evening electrical school. Scientific American, 63 (13):192, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
192.pdf.

Anonymous:1890:TAP

[930] Anonymous. Testing armor plates for our war vessels. Scientific American, 63(13):192, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
192a.pdf.

Anonymous:1890:PPO

[931] Anonymous. Position of the planets in October. Scientific American, 63(13):193, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
193.pdf.

Pike:1890:HHS

[932] Nicolas Pike. The horse hair snake. Scientific American, 63 (13):193, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
193a.pdf.

Anonymous:1890:IBz

[933] Anonymous. An improved bridle. Scientific American, 63(13): 194, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
194d.pdf.

Anonymous:1890:ITe

[934] Anonymous. An improved hat tag. Scientific American, 63 (13):194, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
194.pdf.

Anonymous:1890:ITCe

[935] Anonymous. An improved thill coupling. Scientific American, 63 (13):194, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
194c.pdf.

Anonymous:1890:LTE

[936] Anonymous. Leveler for traction engines. Scientific American, 63 (13):194, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
194b.pdf.

Anonymous:1890:SBS

[937] Anonymous. The “Success” bob sleigh. Scientific American, 63 (13):194, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
194a.pdf.

Anonymous:1890:IFLb

[938] Anonymous. Improved field locomotive. Scientific American, 63 (13):195, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
195a.pdf.

Anonymous:1890:IGB

[939] Anonymous. An improved gas burner. Scientific American, 63 (13):195, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
195.pdf.

Anonymous:1890:IBBc

[940] Anonymous. An improved back band buckle. Scientific American, 63 (13):195–196, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
195b.pdf.

Anonymous:1890:Cbe

[941] Anonymous. Correspondence. Scientific American, 63(13):197, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n13/pdf/scientificamerican09271890-197a.pdf.

Anonymous:1890:SBG

[942] Anonymous. Ship building on the Great Lakes. Scientific American, 63(13):197, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
197.pdf.

Anonymous:1890:STS

[943] Anonymous. Seamless tubes from solid blocks of metal. Scientific American, 63(13):198–199, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
198.pdf.

Anonymous:1890:LSO

[944] Anonymous. A lady slipper orchid. Scientific American, 63 (13):199, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
199.pdf.

Anonymous:1890:GWC

[945] Anonymous. Growth of a wooden caterpillar. Scientific American, 63 (13):201–202, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
201.pdf.

Anonymous:1890:BPbl

[946] Anonymous. Business and personal. Scientific American, 63 (13):203, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
203a.pdf.

Anonymous:1890:NQbj

[947] Anonymous. Notes and queries. Scientific American, 63(13): 203, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
203b.pdf.

Anonymous:1890:RPIbm

[948] Anonymous. Recently patented inventions. Scientific American, 63 (13):203, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
203.pdf.

Anonymous:1890:IIbm

[949] Anonymous. Index of inventions. Scientific American, 63(13): 204, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
204.pdf.

Anonymous:1890:Ba

[950] Anonymous. Books. Scientific American, 63(13):205, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n13/pdf/scientificamerican09271890-205.pdf.

Anonymous:1890:PNPz

[951] Anonymous. Patent notes: Patents. Scientific American, 63 (13):206, September 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n13/pdf/scientificamerican09271890-
206.pdf.

Anonymous:1890:P

[952] Anonymous. Platinum. Scientific American, 63(14):208, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-208.pdf.

Anonymous:1890:NSPa

[953] Anonymous. Nickel-steel plates for our war vessels. Scientific American, 63(14):208–210, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n14/pdf/scientificamerican10041890-
208a.pdf.

Anonymous:1890:CSF

[954] Anonymous. Combined strainer and funnel. Scientific American, 63(14):210, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n14/pdf/scientificamerican10041890-
210a.pdf.

Anonymous:1890:IRE

[955] Anonymous. Improvements in rotary engines. Scientific American, 63(14):210, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n14/pdf/scientificamerican10041890-
210b.pdf.

Anonymous:1890:NLE

[956] Anonymous. Novel lace or embroidery holder. Scientific American, 63(14):210, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n14/pdf/scientificamerican10041890-
210.pdf.

Anonymous:1890:NV

[957] Anonymous. A novel vise. Scientific American, 63(14):210, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-210c.pdf.

Anonymous:1890:NAI

[958] Anonymous. A new agricultural implement. Scientific American, 63(14):211, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n14/pdf/scientificamerican10041890-
211.pdf.

Anonymous:1890:SIV

[959] Anonymous. Some of the industries of Virginia. Scientific American, 63(14):211–212, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n14/pdf/scientificamerican10041890-
211a.pdf.

Anonymous:1890:Cbg

[960] Anonymous. Correspondence. Scientific American, 63(14):213, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-213.pdf.

Anonymous:1890:NSPb

[961] Anonymous. The new steamer Plymouth of the Old Colony Steamboat Company. Scientific American, 63(14):214–215, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-214.pdf.

Anonymous:1890:BT

[962] Anonymous. The banyan tree. Scientific American, 63(14):215–217, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-215.pdf.

Anonymous:1890:BPbm

[963] Anonymous. Business and personal. Scientific American, 63(14):218, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-218a.pdf.

Anonymous:1890:RPIbn

[964] Anonymous. Recently patented inventions. Scientific American, 63(14):218, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n14/pdf/scientificamerican10041890-
218.pdf.

Anonymous:1890:NQbk

[965] Anonymous. Notes and queries. Scientific American, 63(14):218–219, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-218c.pdf.

Anonymous:1890:Icj

[966] Anonymous. To inventors. Scientific American, 63(14):219, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-219.pdf.

Anonymous:1890:IIbn

[967] Anonymous. Index of inventions. Scientific American, 63(14):219–220, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-219a.pdf.

Anonymous:1890:Bb

[968] Anonymous. Books. Scientific American, 63(14):221, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-221.pdf.

Anonymous:1890:PNPba

[969] Anonymous. Patent notes: Patents. Scientific American, 63(14):222, October 4, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n14/pdf/scientificamerican10041890-222.pdf.

Anonymous:1890:MBI

[970] Anonymous. Meeting of the British Iron and Steel Institute in New York. Scientific American, 63(15):224–225, October 11, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n15/pdf/scientificamerican10111890-224.pdf.

Anonymous:1890:RCM

[971] Anonymous. The Ransome concrete mixer. Scientific American, 63(15):226, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
226.pdf.

Anonymous:1890:MSF

[972] Anonymous. A machine to set and file saws. Scientific American, 63(15):227, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
227a.pdf.

Hopkins:1890:ER

[973] Geo M. Hopkins. An experiment in resonance. Scientific American, 63(15):227, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
227.pdf.

Anonymous:1890:VCSb

[974] Anonymous. A vault cover sidewalk openings. Scientific American, 63(15):227–228, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
227b.pdf.

Anonymous:1890:PWF

[975] Anonymous. Post for wire fences. Scientific American, 63(15):228, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n15/pdf/scientificamerican10111890-228.pdf.

Anonymous:1890:SCC

[976] Anonymous. Sibley College Cornell University. Scientific American, 63(15):229, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
229.pdf.

Anonymous:1890:IWe

[977] Anonymous. An improved windmill. Scientific American, 63 (15):230, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
230a.pdf.

Anonymous:1890:PAR

[978] Anonymous. The pitcher automatic repeating rifle. Scientific American, 63(15):230, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
230.pdf.

Anonymous:1890:BAa

[979] Anonymous. A balloon accident. Scientific American, 63(15): 231–232, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
231.pdf.

Anonymous:1890:NPT

[980] Anonymous. The new “Patrick” turbine water wheel. Scientific American, 63(15):232, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
232.pdf.

Anonymous:1890:CCMb

[981] Anonymous. Centennial of the cotton manufacture in America. Scientific American, 63(15):233, October 11, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n15/pdf/scientificamerican10111890-233.pdf.

Anonymous:1890:BPbn

[982] Anonymous. Business and personal. Scientific American, 63 (15):234, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
234a.pdf.

Anonymous:1890:RPIbo

[983] Anonymous. Recently patented inventions. Scientific American, 63(15):234, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
234.pdf.

Anonymous:1890:NQbl

[984] Anonymous. Notes and queries. Scientific American, 63(15): 234–235, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
234b.pdf.

Anonymous:1890:Ick

[985] Anonymous. To inventors. Scientific American, 63(15):235, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n15/pdf/scientificamerican10111890-235.pdf.

Anonymous:1890:IIbo

[986] Anonymous. Index of inventions. Scientific American, 63(15): 235–237, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
235a.pdf.

Anonymous:1890:PNPbb

[987] Anonymous. Patent notes: Patents. Scientific American, 63 (15):238, October 11, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n15/pdf/scientificamerican10111890-
238.pdf.

Anonymous:1890:FSP

[988] Anonymous. Floating steam-propelled islands of steel. Scientific American, 63(16):240, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
240.pdf.

Anonymous:1890:VBI

[989] Anonymous. The visit of the British Iron and Steel Institute. Scientific American, 63(16):240–241, October 18, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n16/pdf/scientificamerican10181890-240a.pdf.

Hovey:1890:MPG

[990] H. C. Hovey. Manufacture of plate glass at Kokomo, Indiana. Scientific American, 63(16):241, October 18, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n16/pdf/scientificamerican10181890-241.pdf.

Anonymous:1890:DM

[991] Anonymous. The Daguerre Memorial. Scientific American, 63 (16):242–243, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
242.pdf.

Anonymous:1890:ICHd

[992] Anonymous. An improved cylinder head for rock drills. Scientific American, 63(16):243, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
243.pdf.

Anonymous:1890:ISMd

[993] Anonymous. An improved saw mill dog. Scientific American, 63(16):243, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
243a.pdf.

Anonymous:1890:SRA

[994] Anonymous. A simple and readily applied knife guard. Scientific American, 63(16):243, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
243b.pdf.

Anonymous:1890:ICCp

[995] Anonymous. An improved car coupling. Scientific American, 63 (16):243–244, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
243c.pdf.

Anonymous:1890:USCa

[996] Anonymous. The U.S. steel cruiser San Francisco. Scientific American, 63(16):244, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
244.pdf.

Anonymous:1890:Cbh

[997] Anonymous. Correspondence. Scientific American, 63(16):245, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n16/pdf/scientificamerican10181890-245.pdf.

Anonymous:1890:HMP

[998] Anonymous. The howling monkeys and the pinches in the Berlin Zoological Garden. Scientific American, 63(16):246, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n16/pdf/scientificamerican10181890-246.pdf.

Anonymous:1890:TNJ

[999] Anonymous. Trial of the new Jackson jet boat evolution. Scientific American, 63(16):247–249, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
247.pdf.

Anonymous:1890:EVA

[1000] Anonymous. An Engineer’s valve for air brakes. Scientific American, 63(16):249, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
249a.pdf.

Anonymous:1890:TIY

[1001] Anonymous. A tension indicator for yarn dressers. Scientific American, 63(16):249, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
249.pdf.

Anonymous:1890:ICn

[1002] Anonymous. An improved cistern. Scientific American, 63(16): 249–250, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
249b.pdf.

Anonymous:1890:NQbm

[1003] Anonymous. Notes and queries. Scientific American, 63(16):250, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n16/pdf/scientificamerican10181890-250.pdf.

Anonymous:1890:BPbo

[1004] Anonymous. Business and personal. Scientific American, 63 (16):251, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
251a.pdf.

Anonymous:1890:Icq

[1005] Anonymous. To inventors. Scientific American, 63(16):251, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n16/pdf/scientificamerican10181890-251b.pdf.

Anonymous:1890:RPIbp

[1006] Anonymous. Recently patented inventions. Scientific American, 63(16):251, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
251.pdf.

Anonymous:1890:IIbp

[1007] Anonymous. Index of inventions. Scientific American, 63(16): 251–253, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
251c.pdf.

Anonymous:1890:PNPbc

[1008] Anonymous. Patent notes: Patents! Scientific American, 63 (16):254, October 18, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n16/pdf/scientificamerican10181890-
254.pdf.

Anonymous:1890:VCP

[1009] Anonymous. The vestibuled car patent decision. Scientific American, 63(17):256, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
256.pdf.

Anonymous:1890:IEAb

[1010] Anonymous. Interesting exhibits at the American institute fair. Scientific American, 63(17):256–258, October 25, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n17/pdf/scientificamerican10251890-256a.pdf.

Anonymous:1890:SDd

[1011] Anonymous. Steam digging. Scientific American, 63(17):258, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-258.pdf.

Anonymous:1890:LT

[1012] Anonymous. Lyon’s truck. Scientific American, 63(17):259, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-259b.pdf.

Anonymous:1890:NAW

[1013] Anonymous. A new aerating and water pump. Scientific American, 63(17):259, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
259a.pdf.

Anonymous:1890:NDT

[1014] Anonymous. New desk, toilet and display bottle. Scientific American, 63(17):259, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
259c.pdf.

Anonymous:1890:NPDc

[1015] Anonymous. A new potato digger. Scientific American, 63(17):259, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-259.pdf.

Anonymous:1890:MPD

[1016] Anonymous. M. Palacio’s design for a colossal monument in memory of Christopher Columbus. Scientific American, 63(17):260, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-260.pdf.

Anonymous:1890:Cbi

[1017] Anonymous. Correspondence. Scientific American, 63(17):261, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-261.pdf.

Anonymous:1890:NC

[1018] Anonymous. A new clutch. Scientific American, 63(17):262, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-262.pdf.

Anonymous:1890:NECd

[1019] Anonymous. New electric conduit. Scientific American, 63(17):262, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-262a.pdf.

Anonymous:1890:EPPb

[1020] Anonymous. Experimental plant of the Portelectric system. Scientific American, 63(17):263–264, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
263.pdf.

Anonymous:1890:PNn

[1021] Anonymous. Photographic notes. Scientific American, 63(17): 264–265, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
264.pdf.

Anonymous:1890:ITKb

[1022] Anonymous. Improved telegraph key. Scientific American, 63 (17):265–266, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
265.pdf.

Anonymous:1890:BPbp

[1023] Anonymous. Business and personal. Scientific American, 63 (17):267, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
267a.pdf.

Anonymous:1890:Icr

[1024] Anonymous. To inventors. Scientific American, 63(17):267, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-267d.pdf.

Anonymous:1890:NBPj

[1025] Anonymous. New books and publications. Scientific American, 63(17):267, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
267c.pdf.

Anonymous:1890:NQbn

[1026] Anonymous. Notes and queries. Scientific American, 63(17):267, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n17/pdf/scientificamerican10251890-267b.pdf.

Anonymous:1890:RPIbq

[1027] Anonymous. Recently patented inventions. Scientific American, 63(17):267, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
267.pdf.

Anonymous:1890:IIbq

[1028] Anonymous. Index of inventions. Scientific American, 63(17): 267–269, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
267e.pdf.

Anonymous:1890:PNPbd

[1029] Anonymous. Patent notes: Patents! Scientific American, 63 (17):269, October 25, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n17/pdf/scientificamerican10251890-
270.pdf.

Anonymous:1890:DND

[1030] Anonymous. Defective Naval designing and material. Scientific American, 63(18):272, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
272a.pdf.

Anonymous:1890:PWW

[1031] Anonymous. Practical work for the World’s Fair. Scientific American, 63(18):272, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
272.pdf.

Pike:1890:E

[1032] Nicolas Pike. The eucalyptus. Scientific American, 63(18):273– 274, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
273.pdf.

Anonymous:1890:FTW

[1033] Anonymous. A forty-ton wharf crane. Scientific American, 63 (18):274–275, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
274.pdf.

Anonymous:1890:SHW

[1034] Anonymous. A steam or hot water heating boiler. Scientific American, 63(18):275–276, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
275.pdf.

Anonymous:1890:CAR

[1035] Anonymous. A cincture attachment for riding saddles. Scientific American, 63(18):276, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
276.pdf.

Anonymous:1890:SFMa

[1036] Anonymous. A simple form of mop wringer. Scientific American, 63(18):276, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
276a.pdf.

Anonymous:1890:Cbj

[1037] Anonymous. Correspondence. Scientific American, 63(18):277, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n18/pdf/scientificamerican11011890-277.pdf.

Anonymous:1890:TFN

[1038] Anonymous. The tornado at Fargo North Dakota. Scientific American, 63(18):278, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
278.pdf.

Anonymous:1890:PGR

[1039] Anonymous. Progress of the great railway tunnels under the Hudson River between New York and Jersey City. Scientific American, 63(18):279–280, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
279.pdf.

Anonymous:1890:APE

[1040] Anonymous. An adjustable post and expansible joint for wire fences. Scientific American, 63(18):281, November 1, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n18/pdf/scientificamerican11011890-281b.pdf.

Anonymous:1890:IFCa

[1041] Anonymous. An improved fuse cap fastener. Scientific American, 63(18):281, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
281a.pdf.

Anonymous:1890:PNo

[1042] Anonymous. Photographic notes. Scientific American, 63(18): 281, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
281.pdf.

Anonymous:1890:BPBq

[1043] Anonymous. Business and personal: Business and personal. Scientific American, 63(18):282, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
282a.pdf.

Anonymous:1890:RPIbr

[1044] Anonymous. Recently patented inventions. Scientific American, 63(18):282, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
282.pdf.

Anonymous:1890:NQbo

[1045] Anonymous. Notes and queries. Scientific American, 63(18): 282–283, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
282b.pdf.

Anonymous:1890:Icu

[1046] Anonymous. To inventors. Scientific American, 63(18):283, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n18/pdf/scientificamerican11011890-283.pdf.

Anonymous:1890:IIbr

[1047] Anonymous. Index of inventions. Scientific American, 63(18): 283–285, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
283a.pdf.

Anonymous:1890:PNPbe

[1048] Anonymous. Patent notes: Patents. Scientific American, 63 (18):286, November 1, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n18/pdf/scientificamerican11011890-
286.pdf.

Anonymous:1890:FDR

[1049] Anonymous. The floating derrick Reliance. Scientific American, 63(19):287, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
287.pdf.

Anonymous:1890:DCI

[1050] Anonymous. Decreasing cost of incandescent light. Scientific American, 63(19):288, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
288.pdf.

Anonymous:1890:NGW

[1051] Anonymous. New gunpowder as well as new guns. Scientific American, 63(19):288, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
288a.pdf.

Anonymous:1890:PSBb

[1052] Anonymous. Palæontological studies in Brazil. Scientific American, 63(19):288, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
288b.pdf.

Anonymous:1890:JSE

[1053] Anonymous. Job as a steam engineer. Scientific American, 63 (19):288–289, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
288c.pdf.

Anonymous:1890:PPNb

[1054] Anonymous. Position of the planets in November. Scientific American, 63(19):289, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
289.pdf.

Anonymous:1890:ESY

[1055] Anonymous. Engines of the steam yacht Lady Torfrida. Scientific American, 63(19):290–291, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
290.pdf.

Anonymous:1890:IUW

[1056] Anonymous. Improvement in the utilization of water power. Scientific American, 63(19):291, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
291a.pdf.

Anonymous:1890:NWH

[1057] Anonymous. A new wheel hub. Scientific American, 63(19):291, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n19/pdf/scientificamerican11081890-291.pdf.

Anonymous:1890:NFG

[1058] Anonymous. A new farm gate. Scientific American, 63(19): 291–292, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
291b.pdf.

Anonymous:1890:Cbl

[1059] Anonymous. Correspondence. Scientific American, 63(19):293– 294, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
293.pdf.

Anonymous:1890:IWCb

[1060] Anonymous. An improved water cycle. Scientific American, 63 (19):294, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
294.pdf.

Anonymous:1890:DFT

[1061] Anonymous. A device to facilitate taking photographs by flash light. Scientific American, 63(19):294–295, November 8, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n19/pdf/scientificamerican11081890-294a.pdf.

Anonymous:1890:NFR

[1062] Anonymous. Novel finger ring gauge. Scientific American, 63 (19):295, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
295.pdf.

Anonymous:1890:IASb

[1063] Anonymous. Improvement in the art of shaping sheet metal. Scientific American, 63(19):295–296, November 8, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n19/pdf/scientificamerican11081890-295a.pdf.

Anonymous:1890:DRPd

[1064] Anonymous. Decisions relating to patents. Scientific American, 63(19):297, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
297b.pdf.

Anonymous:1890:IDb

[1065] Anonymous. An improved dumbwaiter. Scientific American, 63 (19):297, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
297.pdf.

Anonymous:1890:NRF

[1066] Anonymous. New railroad frog. Scientific American, 63(19): 297, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
297c.pdf.

Anonymous:1890:NSM

[1067] Anonymous. New soldering machine. Scientific American, 63 (19):297, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
297a.pdf.

Anonymous:1890:BPBr

[1068] Anonymous. Business and personal: Business and personal. Scientific American, 63(19):298, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
298a.pdf.

Anonymous:1890:NBPk

[1069] Anonymous. New books and publications. Scientific American, 63(19):298, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
298b.pdf.

Anonymous:1890:RPIbs

[1070] Anonymous. Recently patented inventions. Scientific American, 63(19):298, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
298.pdf.

Anonymous:1890:Icw

[1071] Anonymous. To inventors. Scientific American, 63(19):299, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n19/pdf/scientificamerican11081890-299a.pdf.

Anonymous:1890:NQbp

[1072] Anonymous. Notes and queries. Scientific American, 63(19): 299, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
299.pdf.

Anonymous:1890:IIbs

[1073] Anonymous. Index of inventions. Scientific American, 63(19): 299–300, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
299b.pdf.

Anonymous:1890:PNPbf

[1074] Anonymous. Patent notes: Patents! Scientific American, 63 (19):301–302, November 8, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n19/pdf/scientificamerican11081890-
301.pdf.

Anonymous:1890:EFG

[1075] Anonymous. Electrically-fired Gatling gun. Scientific American, 63(20):303, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
303.pdf.

Anonymous:1890:HEO

[1076] Anonymous. How to estimate our work on war vessels. Scientific American, 63(20):304, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
304a.pdf.

Anonymous:1890:ZSL

[1077] Anonymous. Zapon a substitute for lacquer. Scientific American, 63(20):304, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
304.pdf.

Anonymous:1890:DKC

[1078] Anonymous. Dr Koch’s cure for consumption. Scientific American, 63(20):304–305, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
304b.pdf.

Anonymous:1890:LSB

[1079] Anonymous. Lake steamers built by a Cleveland firm. Scientific American, 63(20):306, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
306.pdf.

Anonymous:1890:MTC

[1080] Anonymous. A machine for threading and cutting off pipes etc. Scientific American, 63(20):306, November 15, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n20/pdf/scientificamerican11151890-306a.pdf.

Anonymous:1890:FTP

[1081] Anonymous. A folding table for paper hangers. Scientific American, 63(20):307, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
307.pdf.

Anonymous:1890:IHL

[1082] Anonymous. An improved head-light for locomotives. Scientific American, 63(20):307, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
307b.pdf.

Anonymous:1890:SUT

[1083] Anonymous. Several useful tools in one. Scientific American, 63 (20):307, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
307a.pdf.

Anonymous:1890:ICx

[1084] Anonymous. Improved Croze. Scientific American, 63(20):307– 308, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
307c.pdf.

Anonymous:1890:SFG

[1085] Anonymous. A simple form of game. Scientific American, 63 (20):308, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
308.pdf.

Anonymous:1890:Cbm

[1086] Anonymous. Correspondence. Scientific American, 63(20):309, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n20/pdf/scientificamerican11151890-309.pdf.

Anonymous:1890:ARB

[1087] Anonymous. Auxiliary rifle barrel for guns. Scientific American, 63(20):310, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
310.pdf.

Saunders:1890:NQ

[1088] Wm. L. Saunders. Notes on quarrying. Scientific American, 63 (20):310, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
310a.pdf.

Anonymous:1890:USCb

[1089] Anonymous. The United States collection of standard weights and measures. Scientific American, 63(20):311–313, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n20/pdf/scientificamerican11151890-311.pdf.

Anonymous:1890:IEb

[1090] Anonymous. An interesting experiment. Scientific American, 63 (20):313, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
313.pdf.

Anonymous:1890:NBPl

[1091] Anonymous. New books and publications. Scientific American, 63 (20):314, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
314a.pdf.

Anonymous:1890:RPIbt

[1092] Anonymous. Recently patented inventions. Scientific American, 63(20):314, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
314.pdf.

Anonymous:1890:BPBs

[1093] Anonymous. Business and personal: Business and personal. Scientific American, 63(20):315, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
315.pdf.

Anonymous:1890:Icy

[1094] Anonymous. To inventors. Scientific American, 63(20):315, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n20/pdf/scientificamerican11151890-315b.pdf.

Anonymous:1890:NQbq

[1095] Anonymous. Notes and queries. Scientific American, 63(20): 315, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
315a.pdf.

Anonymous:1890:IIbt

[1096] Anonymous. Index of inventions. Scientific American, 63(20): 315–317, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
315c.pdf.

Anonymous:1890:PNPbg

[1097] Anonymous. Patent notes: Patents. Scientific American, 63 (20):318, November 15, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n20/pdf/scientificamerican11151890-
318.pdf.

Anonymous:1890:TRI

[1098] Anonymous. The Tennessee River improvement. Scientific American, 63(21):320, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
320.pdf.

Anonymous:1890:DKT

[1099] Anonymous. Dr. Koch’s treatment for consumption. Scientific American, 63(21):320–322, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
320a.pdf.

Anonymous:1890:ILRc

[1100] Anonymous. An improved land roller. Scientific American, 63 (21):323, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
323a.pdf.

Anonymous:1890:ISDc

[1101] Anonymous. An improved score device. Scientific American, 63 (21):323, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
323.pdf.

Anonymous:1890:ISWb

[1102] Anonymous. An improved shutter worker. Scientific American, 63 (21):323, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
323b.pdf.

Anonymous:1890:TSB

[1103] Anonymous. A tool for setting boiler flues. Scientific American, 63(21):323, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
323c.pdf.

Anonymous:1890:ICCq

[1104] Anonymous. An improved car coupling. Scientific American, 63 (21):323–324, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
323d.pdf.

Anonymous:1890:NHC

[1105] Anonymous. New hose coupling. Scientific American, 63(21): 324, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
324.pdf.

Anonymous:1890:Cbn

[1106] Anonymous. Correspondence. Scientific American, 63(21):325, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n21/pdf/scientificamerican11221890-325.pdf.

Anonymous:1890:UNF

[1107] Anonymous. Utilizing Niagara Falls water power. Scientific American, 63(21):326, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
326.pdf.

Anonymous:1890:NSBc

[1108] Anonymous. The Norton self-ballasting life boat or yacht. Scientific American, 63(21):327–328, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
327.pdf.

Anonymous:1890:MCC

[1109] Anonymous. Mould for constructing concrete sewers and subways. Scientific American, 63(21):329, November 22, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n21/pdf/scientificamerican11221890-329.pdf.

Anonymous:1890:SAg

[1110] Anonymous. Scientific American. Scientific American, 63(21): 330, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
330b.pdf.

Anonymous:1890:BNBo

[1111] Anonymous. Books: New books and publications. Scientific American, 63(21):330, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
330a.pdf.

Anonymous:1890:RPIbu

[1112] Anonymous. Recently patented inventions. Scientific American, 63(21):330, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
330.pdf.

Anonymous:1890:BPbt

[1113] Anonymous. Business and personal. Scientific American, 63 (21):331, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
331.pdf.

Anonymous:1890:Icz

[1114] Anonymous. To inventors. Scientific American, 63(21):331, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n21/pdf/scientificamerican11221890-331b.pdf.

Anonymous:1890:NQbr

[1115] Anonymous. Notes and queries. Scientific American, 63(21): 331, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
331a.pdf.

Anonymous:1890:IIbu

[1116] Anonymous. Index of inventions. Scientific American, 63(21): 331–332, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
331c.pdf.

Anonymous:1890:PNPbh

[1117] Anonymous. Patent notes: Patents. Scientific American, 63(21): 333–334, November 22, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n21/pdf/scientificamerican11221890-
333.pdf.

Anonymous:1890:BNN

[1118] Anonymous. Building the new Navy. Scientific American, 63 (22):336, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
336.pdf.

Anonymous:1890:PKC

[1119] Anonymous. Prof. Koch’s consumption Remedy. Scientific American, 63(22):336, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
336a.pdf.

Anonymous:1890:TTG

[1120] Anonymous. Trial trip of the gunboat concord. Scientific American, 63(22):336–337, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
336b.pdf.

Anonymous:1890:GDB

[1121] Anonymous. Grafting a Dog’s bone on a boy. Scientific American, 63(22):337, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
337.pdf.

Anonymous:1890:HCL

[1122] Anonymous. Hatching cod and lobster. Scientific American, 63 (22):337, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
337b.pdf.

Anonymous:1890:PPD

[1123] Anonymous. Position of the planets in December. Scientific American, 63(22):337, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
337c.pdf.

Anonymous:1890:WNC

[1124] Anonymous. Work on the Nicaragua Canal. Scientific American, 63(22):337, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
337a.pdf.

Anonymous:1890:DN

[1125] Anonymous. Dredging in Norway. Scientific American, 63(22): 338, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
338a.pdf.

Anonymous:1890:SL

[1126] Anonymous. The Similigraph level. Scientific American, 63(22): 338, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
338.pdf.

Anonymous:1890:ACC

[1127] Anonymous. An automatic car coupler. Scientific American, 63 (22):339, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
339.pdf.

Anonymous:1890:DSC

[1128] Anonymous. A device to supply cool water. Scientific American, 63(22):339, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
339a.pdf.

Anonymous:1890:ITPb

[1129] Anonymous. An improved toy pistol. Scientific American, 63 (22):339, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
339c.pdf.

Anonymous:1890:IWf

[1130] Anonymous. An improved wrench. Scientific American, 63(22): 339, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
339d.pdf.

Anonymous:1890:SFMb

[1131] Anonymous. A simple fence-making machine. Scientific American, 63(22):339, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
339b.pdf.

Anonymous:1890:LAC

[1132] Anonymous. The launch of the armored cruiser Maine at the Brooklyn Navy Yard. Scientific American, 63(22):340, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n22/pdf/scientificamerican11291890-340.pdf.

Anonymous:1890:PNq

[1133] Anonymous. Photographic notes. Scientific American, 63(22): 340–341, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
340a.pdf.

Anonymous:1890:EUR

[1134] Anonymous. The Electric Underground Railway London. Scientific American, 63(22):342–344, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
342.pdf.

Anonymous:1890:IFCb

[1135] Anonymous. An improved flue cleaner. Scientific American, 63 (22):344, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
344.pdf.

Anonymous:1890:CDB

[1136] Anonymous. A counterbalanced draw bridge. Scientific American, 63(22):345, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
345.pdf.

Anonymous:1890:SAh

[1137] Anonymous. Scientific American. Scientific American, 63(22): 346, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
346b.pdf.

Anonymous:1890:BNBp

[1138] Anonymous. Books: New books and publications. Scientific American, 63(22):346, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
346a.pdf.

Anonymous:1890:BPbu

[1139] Anonymous. Business and personal. Scientific American, 63 (22):346, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
346c.pdf.

Anonymous:1890:RPIbv

[1140] Anonymous. Recently patented inventions. Scientific American, 63(22):346, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
346.pdf.

Anonymous:1890:NQbs

[1141] Anonymous. Notes and queries. Scientific American, 63(22): 346–347, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
346d.pdf.

Anonymous:1890:Idd

[1142] Anonymous. To inventors. Scientific American, 63(22):347, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n22/pdf/scientificamerican11291890-347.pdf.

Anonymous:1890:IIbv

[1143] Anonymous. Index of inventions. Scientific American, 63(22): 347–349, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
347a.pdf.

Anonymous:1890:PNPbi

[1144] Anonymous. Patent notes: Patents. Scientific American, 63 (22):350, November 29, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n22/pdf/scientificamerican11291890-
350.pdf.

Anonymous:1890:BGC

[1145] Anonymous. Big guns for coast defense. Scientific American, 63 (23):352, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
352.pdf.

Anonymous:1890:PA

[1146] Anonymous. Progress of aluminum. Scientific American, 63(23): 352, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
352a.pdf.

Anonymous:1890:PDC

[1147] Anonymous. Pigeons as dispatch carriers. Scientific American, 63 (23):352–353, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
352b.pdf.

Anonymous:1890:MRN

[1148] Anonymous. A mountain railway near New York City. Scientific American, 63(23):353–354, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
353.pdf.

Anonymous:1890:ASBa

[1149] Anonymous. An automatic steam boiler feeder. Scientific American, 63(23):354, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
354a.pdf.

Anonymous:1890:RVP

[1150] Anonymous. A refrigerator vehicle for perishable articles. Scientific American, 63(23):354, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
354.pdf.

Anonymous:1890:BSC

[1151] Anonymous. A brake for street cars. Scientific American, 63 (23):355, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
355.pdf.

Anonymous:1890:CDD

[1152] Anonymous. A cork-drawing device. Scientific American, 63 (23):355, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
355b.pdf.

Anonymous:1890:IEGb

[1153] Anonymous. An improved elevator gate. Scientific American, 63 (23):355, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
355a.pdf.

Anonymous:1890:MWW

[1154] Anonymous. A marker for wood-workers. Scientific American, 63(23):355, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
355c.pdf.

Anonymous:1890:IWg

[1155] Anonymous. An improved wrench. Scientific American, 63(23): 355–356, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
355d.pdf.

Anonymous:1890:DNY

[1156] Anonymous. The defense of New York. Scientific American, 63 (23):356, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
356.pdf.

Anonymous:1890:Cbo

[1157] Anonymous. Correspondence. Scientific American, 63(23):357, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n23/pdf/scientificamerican12061890-357.pdf.

Hacks:1890:CCI

[1158] Charles Hacks. Cure of consumption — an interview with Professor Koch. Scientific American, 63(23):358–359, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n23/pdf/scientificamerican12061890-358.pdf.

Anonymous:1890:SCB

[1159] Anonymous. Safe construction of buildings. Scientific American, 63 (23):360–361, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
360.pdf.

Anonymous:1890:EC

[1160] Anonymous. Electric chimes. Scientific American, 63(23):361, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n23/pdf/scientificamerican12061890-361.pdf.

Anonymous:1890:BNBq

[1161] Anonymous. Books: New books and publications. Scientific American, 63(23):362, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
362a.pdf.

Anonymous:1890:BPbv

[1162] Anonymous. Business and personal. Scientific American, 63 (23):362, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
362b.pdf.

Anonymous:1890:RPIbw

[1163] Anonymous. Recently patented inventions. Scientific American, 63(23):362, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
362.pdf.

Anonymous:1890:NQbt

[1164] Anonymous. Notes & queries. Scientific American, 63(23):362–363, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n23/pdf/scientificamerican12061890-362d.pdf.

Anonymous:1890:NQbu

[1165] Anonymous. Notes and queries. Scientific American, 63(23): 362–363, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
362c.pdf.

Anonymous:1890:Ide

[1166] Anonymous. To inventors. Scientific American, 63(23):363, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n23/pdf/scientificamerican12061890-363.pdf.

Anonymous:1890:IIbw

[1167] Anonymous. Index of inventions. Scientific American, 63(23): 363–365, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
363a.pdf.

Anonymous:1890:PNPbj

[1168] Anonymous. Patent notes: Patents. Scientific American, 63 (23):366, December 6, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n23/pdf/scientificamerican12061890-
366.pdf.

Anonymous:1890:CUSb

[1169] Anonymous. The centennial of the United States Patent System. Scientific American, 63(24):368, December 13, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n24/pdf/scientificamerican12131890-368.pdf.

Anonymous:1890:SSBb

[1170] Anonymous. A school of ship building — the objects and ideas of its founder. Scientific American, 63(24):368–370, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n24/pdf/scientificamerican12131890-368a.pdf.

Anonymous:1890:CET

[1171] Anonymous. A convenient ear trumpet. Scientific American, 63 (24):370, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
370.pdf.

Anonymous:1890:ICHe

[1172] Anonymous. An improved copy holder. Scientific American, 63 (24):370, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
370a.pdf.

Anonymous:1890:RDM

[1173] Anonymous. A regulator for dynamos and motors. Scientific American, 63(24):370, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
370b.pdf.

Anonymous:1890:DBC

[1174] Anonymous. A device to block cars on a grade. Scientific American, 63(24):371, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
371c.pdf.

Anonymous:1890:IMRb

[1175] Anonymous. An improved metallic railway tie. Scientific American, 63(24):371, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
371b.pdf.

Anonymous:1890:IPF

[1176] Anonymous. An improved pipe or flue cutter. Scientific American, 63(24):371, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
371a.pdf.

Anonymous:1890:RSA

[1177] Anonymous. Railroad sleepers arranged in pairs. Scientific American, 63(24):371, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
371.pdf.

Anonymous:1890:IVc

[1178] Anonymous. An improved velocipede. Scientific American, 63(24): 371–372, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
371d.pdf.

Anonymous:1890:HRB

[1179] Anonymous. The Hudson River bridge. Scientific American, 63 (24):372, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
372.pdf.

Anonymous:1890:PCBb

[1180] Anonymous. A pipe cutter, beader, and expander. Scientific American, 63(24):372, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
372a.pdf.

Anonymous:1890:PMB

[1181] Anonymous. A pug mill for brick makers. Scientific American, 63(24):372, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
372b.pdf.

Anonymous:1890:MSS

[1182] Anonymous. The manufacture of salt at Syracuse, N.Y. Scientific American, 63(24):373, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
373.pdf.

Anonymous:1890:Cbp

[1183] Anonymous. The Codonophone. Scientific American, 63(24): 374, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
374.pdf.

Coonley:1890:SSC

[1184] J. F. Coonley. The Sebae or silk cotton tree. Scientific American, 63(24):374–375, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
374a.pdf.

Anonymous:1890:BPbw

[1185] Anonymous. The belt problem. Scientific American, 63(24): 376, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
376a.pdf.

Anonymous:1890:Cbq

[1186] Anonymous. Correspondence. Scientific American, 63(24):376, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n24/pdf/scientificamerican12131890-376.pdf.

Anonymous:1890:ISHb

[1187] Anonymous. An instantaneous shutter for hand cameras. Scientific American, 63(24):376, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
376b.pdf.

Anonymous:1890:NTV

[1188] Anonymous. A nautico-terrestrial velocipede. Scientific American, 63(24):377, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
377.pdf.

Anonymous:1890:BPbx

[1189] Anonymous. Business and personal. Scientific American, 63 (24):378, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
378a.pdf.

Anonymous:1890:RPIbx

[1190] Anonymous. Recently patented inventions. Scientific American, 63(24):378, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
378.pdf.

Anonymous:1890:NQbv

[1191] Anonymous. Notes & queries. Scientific American, 63(24):378– 379, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
378c.pdf.

Anonymous:1890:NQbw

[1192] Anonymous. Notes and queries. Scientific American, 63(24): 378–379, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
378b.pdf.

Anonymous:1890:BNBr

[1193] Anonymous. Books: New books and publications. Scientific American, 63(24):379, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
379.pdf.

Anonymous:1890:Idf

[1194] Anonymous. To inventors. Scientific American, 63(24):379, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n24/pdf/scientificamerican12131890-379a.pdf.

Anonymous:1890:IIbx

[1195] Anonymous. Index of inventions. Scientific American, 63(24): 379–381, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
379b.pdf.

Anonymous:1890:PNPbk

[1196] Anonymous. Patent notes: Patents. Scientific American, 63 (24):382, December 13, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n24/pdf/scientificamerican12131890-
382.pdf.

Anonymous:1890:NPS

[1197] Anonymous. New patent selling trick. Scientific American, 63 (25):384, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
384a.pdf.

Anonymous:1890:ONP

[1198] Anonymous. Opening of the new Pulitzer Building. Scientific American, 63(25):384, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
384c.pdf.

Anonymous:1890:SAi

[1199] Anonymous. Subscribers, attention! Scientific American, 63 (25):384, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
384.pdf.

Anonymous:1890:TRN

[1200] Anonymous. Tunnel railways in New York. Scientific American, 63(25):384, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
384b.pdf.

Anonymous:1890:APR

[1201] Anonymous. The artificial production of rain. Scientific American, 63(25):384–385, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
384d.pdf.

Anonymous:1890:AED

[1202] Anonymous. An automatic electrical damper regulator. Scientific American, 63(25):386, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
386b.pdf.

Anonymous:1890:ASBb

[1203] Anonymous. An adjustable stuffing box. Scientific American, 63 (25):386, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
386d.pdf.

Anonymous:1890:BBP

[1204] Anonymous. Boomer & Boschert’s presses. Scientific American, 63(25):386, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
386a.pdf.

Anonymous:1890:HWV

[1205] Anonymous. A hot water ventilating heater. Scientific American, 63(25):386, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
386c.pdf.

Anonymous:1890:INLb

[1206] Anonymous. An improved nut lock. Scientific American, 63 (25):386, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
386.pdf.

Anonymous:1890:EOB

[1207] Anonymous. An easily operated belt tightener. Scientific American, 63(25):387, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
387b.pdf.

Anonymous:1890:MSBa

[1208] Anonymous. A machine to sand brick moulds. Scientific American, 63(25):387, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
387.pdf.

Anonymous:1890:RRK

[1209] Anonymous. Rail rolling at the Krupp Works, Essen. Scientific American, 63(25):387, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
387a.pdf.

Anonymous:1890:IAP

[1210] Anonymous. An improved article for plastering, tiling, and cement work. Scientific American, 63(25):388, December 20, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n25/pdf/scientificamerican12201890-388.pdf.

Anonymous:1890:ILDb

[1211] Anonymous. An improved lumber drier. Scientific American, 63 (25):388, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
388a.pdf.

Anonymous:1890:Cbr

[1212] Anonymous. Correspondence. Scientific American, 63(25):389, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n25/pdf/scientificamerican12201890-389a.pdf.

Anonymous:1890:MBW

[1213] Anonymous. Mr. Brush’s windmill dynamo. Scientific American, 63(25):389, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
389.pdf.

Anonymous:1890:NMS

[1214] Anonymous. New marine signal. Scientific American, 63(25): 390, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
390.pdf.

Anonymous:1890:PAG

[1215] Anonymous. A pine apple grove in Florida. Scientific American, 63(25):390, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
390a.pdf.

Anonymous:1890:NPB

[1216] Anonymous. To the North Pole in a balloon. Scientific American, 63 (25):391–392, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
391.pdf.

Anonymous:1890:CWFc

[1217] Anonymous. The “Crystal” water filter. Scientific American, 63 (25):392, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
392.pdf.

Anonymous:1890:BPby

[1218] Anonymous. Business and personal. Scientific American, 63 (25):393, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
393a.pdf.

Anonymous:1890:Idj

[1219] Anonymous. To inventors. Scientific American, 63(25):393, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n25/pdf/scientificamerican12201890-393c.pdf.

Anonymous:1890:NQbx

[1220] Anonymous. Notes and queries. Scientific American, 63(25): 393, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
393b.pdf.

Anonymous:1890:RPIby

[1221] Anonymous. Recently patented inventions. Scientific American, 63(25):393, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
393.pdf.

Anonymous:1890:IIby

[1222] Anonymous. Index of inventions. Scientific American, 63(25): 393–394, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
393e.pdf.

Anonymous:1890:IIbz

[1223] Anonymous. Index of inventions. Scientific American, 63(25): 393–397, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
393d.pdf.

Anonymous:1890:SAj

[1224] Anonymous. The Scientific American. Scientific American, 63(25): 395–397, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
395.pdf.

Anonymous:1890:PNPbl

[1225] Anonymous. Patent notes: Patents. Scientific American, 63 (25):398, December 20, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n25/pdf/scientificamerican12201890-
398.pdf.

Anonymous:1890:CGL

[1226] Anonymous. Commerce of the Great Lakes. Scientific American, 63(26):400, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
400.pdf.

Anonymous:1890:PPJc

[1227] Anonymous. Position of the planets in January. Scientific American, 63(26):400–401, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
400a.pdf.

Anonymous:1890:DBL

[1228] Anonymous. Date boundary line. Scientific American, 63(26): 401, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
401.pdf.

Anonymous:1890:HJQ

[1229] Anonymous. How to judge the quality of condensed milk. Scientific American, 63(26):401, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
401a.pdf.

Anonymous:1890:DCH

[1230] Anonymous. A device to cool or heat liquids. Scientific American, 63(26):402, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
402b.pdf.

Anonymous:1890:DDL

[1231] Anonymous. A device for developing the lungs and enlarging the chest. Scientific American, 63(26):402, December 27, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n26/pdf/scientificamerican12271890-402.pdf.

Anonymous:1890:ITf

[1232] Anonymous. An improved trap. Scientific American, 63(26): 402, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
402c.pdf.

Anonymous:1890:WOC

[1233] Anonymous. A window opening and closing device. Scientific American, 63(26):402, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
402a.pdf.

Anonymous:1890:PRK

[1234] Anonymous. Plate rolling at the Krupp Works, Essen. Scientific American, 63(26):403–404, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
403.pdf.

Anonymous:1890:IOL

[1235] Anonymous. Improved oil light lantern. Scientific American, 63 (26):404, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
404.pdf.

Anonymous:1890:Cbs

[1236] Anonymous. Correspondence. Scientific American, 63(26):405, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n26/pdf/scientificamerican12271890-405a.pdf.

Anonymous:1890:SOS

[1237] Anonymous. A separable ocean steamer built on Lake Michigan. Scientific American, 63(26):405, December 27, 1890. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v63/
n26/pdf/scientificamerican12271890-405.pdf.

Hopkins:1890:OIA

[1238] Geo. M. Hopkins. Optical illusions adapted to the lantern. Scientific American, 63(26):406, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
406.pdf.

Anonymous:1890:MSBb

[1239] Anonymous. A military suspension bridge. Scientific American, 63 (26):406–408, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
406a.pdf.

Anonymous:1890:DTC

[1240] Anonymous. A double type composing stick. Scientific American, 63(26):408, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
408a.pdf.

Anonymous:1890:RSC

[1241] Anonymous. Repairing the steamship La Champagne with an adjustable bow and stern dock. Scientific American, 63(26):408, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n26/pdf/scientificamerican12271890-408.pdf.

Anonymous:1890:BPbz

[1242] Anonymous. Business and personal. Scientific American, 63 (26):409, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
409a.pdf.

Anonymous:1890:Idm

[1243] Anonymous. To inventors. Scientific American, 63(26):409, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n26/pdf/scientificamerican12271890-409b.pdf.

Anonymous:1890:RPIbz

[1244] Anonymous. Recently patented inventions. Scientific American, 63(26):409, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
409.pdf.

Anonymous:1890:IIca

[1245] Anonymous. Index of invention. Scientific American, 63(26): 409–410, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
409c.pdf.

Anonymous:1890:Idn

[1246] Anonymous. Index. Scientific American, 63(26):411–413, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v63/n26/pdf/scientificamerican12271890-411.pdf.

Anonymous:1890:PNPbm

[1247] Anonymous. Patent notes: Patents. Scientific American, 63 (26):414, December 27, 1890. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v63/n26/pdf/scientificamerican12271890-
414.pdf.

Anonymous:1891:EPL

[1248] Anonymous. Excavating and pipe laying apparatus in use on the Brooklyn Aqueduct. Scientific American, 64(1):1–2, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-1.pdf.

Anonymous:1891:NCN

[1249] Anonymous. The new cruiser Newark. Scientific American, 64(1):2, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-2.pdf.

Anonymous:1891:OPG

[1250] Anonymous. Official proclamation of the great fair. Scientific American, 64(1):2–3, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
2a.pdf.

Anonymous:1891:NTP

[1251] Anonymous. New terminus of Pennsylvania Railroad in Jersey City. Scientific American, 64(1):3, January 3, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n1/pdf/scientificamerican01031891-3a.pdf.

Anonymous:1891:PGR

[1252] Anonymous. Progress of the great railway tunnel under the Hudson River. Scientific American, 64(1):3, January 3, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n1/pdf/scientificamerican01031891-3.pdf.

Anonymous:1891:CHC

[1253] Anonymous. A combined harrow and cutter. Scientific American, 64(1):4, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
4a.pdf.

Anonymous:1891:ISG

[1254] Anonymous. An improved scissors grinder. Scientific American, 64(1):4, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-4b.pdf.

Anonymous:1891:PNa

[1255] Anonymous. The photographic necktie. Scientific American, 64(1):4, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-4.pdf.

Anonymous:1891:JCW

[1256] Anonymous. A journal cap for wood-working machines. Scientific American, 64(1):5, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
5b.pdf.

Anonymous:1891:PRE

[1257] Anonymous. A projectile rotated by the explosive. Scientific American, 64(1):5, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
5a.pdf.

Anonymous:1891:WCF

[1258] Anonymous. The Worcester chemical fire pail. Scientific American, 64(1):5, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
5.pdf.

Anonymous:1891:Ca

[1259] Anonymous. Correspondence. Scientific American, 64(1):6, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-6.pdf.

Anonymous:1891:PNb

[1260] Anonymous. Photographic notes. Scientific American, 64(1):6–7, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-6a.pdf.

Anonymous:1891:BTa

[1261] Anonymous. Bacillus of tuberculosis. Scientific American, 64(1):7, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-7.pdf.

Hopkins:1891:MC

[1262] Geo. M. Hopkins. Metallo chromes. Scientific American, 64(1):8, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-8.pdf.

Anonymous:1891:HTE

[1263] Anonymous. Horizontal triple expansion engine. Scientific American, 64(1):8–9, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
8a.pdf.

Anonymous:1891:RRW

[1264] Anonymous. Remarkable rail-way collision in Iowa. Scientific American, 64(1):9, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
9.pdf.

Anonymous:1891:SABa

[1265] Anonymous. Scientific American building edition. Scientific American, 64(1):10, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
10a.pdf.

Anonymous:1891:BPa

[1266] Anonymous. Business and personal. Scientific American, 64(1):10, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-10b.pdf.

Anonymous:1891:RPIa

[1267] Anonymous. Recently patented inventions. Scientific American, 64(1):10, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n1/pdf/scientificamerican01031891-
10.pdf.

Anonymous:1891:NQa

[1268] Anonymous. Notes and queries. Scientific American, 64(1):10–11, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-10c.pdf.

Anonymous:1891:Ia

[1269] Anonymous. To inventors. Scientific American, 64(1):11, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-11a.pdf.

Anonymous:1891:IIa

[1270] Anonymous. Index of inventions. Scientific American, 64(1):11, January 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n1/pdf/scientificamerican01031891-11.pdf.

Anonymous:1891:AFC

[1271] Anonymous. The American Forestry Congress. Scientific American, 64(2):16, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
16a.pdf.

Anonymous:1891:WCC

[1272] Anonymous. Washington as a convention city. Scientific American, 64(2):16, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
16.pdf.

Anonymous:1891:GSA

[1273] Anonymous. The Geological Society of America. Scientific American, 64(2):16–17, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
16b.pdf.

Anonymous:1891:NCB

[1274] Anonymous. The new Copyright Bill now before Congress. Scientific American, 64(2):17–18, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
17.pdf.

Anonymous:1891:EOT

[1275] Anonymous. An easily operated telegraph instrument. Scientific American, 64(2):18, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
18.pdf.

Anonymous:1891:ICGa

[1276] Anonymous. An improved cable grip. Scientific American, 64 (2):18, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
18b.pdf.

Anonymous:1891:IPR

[1277] Anonymous. An improved parallel ruler. Scientific American, 64(2):18, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
18a.pdf.

Pike:1891:G

[1278] Nicolas Pike. The ginseng. Scientific American, 64(2):19–20, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n2/pdf/scientificamerican01101891-19.pdf.

Anonymous:1891:Cb

[1279] Anonymous. Correspondence. Scientific American, 64(2):21, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n2/pdf/scientificamerican01101891-21a.pdf.

Anonymous:1891:MVC

[1280] Anonymous. The Malleco Viaduct, Chilian State Railway, Collipulli, Chile. Scientific American, 64(2):21, January 10, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n2/pdf/scientificamerican01101891-21.pdf.

Hopkins:1891:IOI

[1281] Geo. M. Hopkins. Interesting optical illusions. Scientific American, 64(2):22, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
22.pdf.

Anonymous:1891:SLA

[1282] Anonymous. The sign language of the American Indians. Scientific American, 64(2):23–25, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
23.pdf.

Anonymous:1891:MMG

[1283] Anonymous. The magnetic magnifying glass and the box of numbers. Scientific American, 64(2):25, January 10, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n2/pdf/scientificamerican01101891-25.pdf.

Anonymous:1891:SABb

[1284] Anonymous. Scientific American building edition. Scientific American, 64(2):26, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
26b.pdf.

Anonymous:1891:BNBa

[1285] Anonymous. Books: New books and publications. Scientific American, 64(2):26, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
26a.pdf.

Anonymous:1891:BPb

[1286] Anonymous. Business and personal. Scientific American, 64(2):26, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n2/pdf/scientificamerican01101891-26c.pdf.

Anonymous:1891:RPIb

[1287] Anonymous. Recently patented inventions. Scientific American, 64(2):26, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n2/pdf/scientificamerican01101891-
26.pdf.

Anonymous:1891:NQb

[1288] Anonymous. Notes and queries. Scientific American, 64(2):26–27, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n2/pdf/scientificamerican01101891-26d.pdf.

Anonymous:1891:Ib

[1289] Anonymous. To inventors. Scientific American, 64(2):27, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n2/pdf/scientificamerican01101891-27.pdf.

Anonymous:1891:IIb

[1290] Anonymous. Index of inventions. Scientific American, 64(2):27–28, January 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n2/pdf/scientificamerican01101891-27a.pdf.

Anonymous:1891:LBEa

[1291] Anonymous. Locomotive boiler explosion, Wallingford, Conn. Scientific American, 64(3):31, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
31.pdf.

Anonymous:1891:SAH

[1292] Anonymous. Students of American history. Scientific American, 64(3):32, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
32.pdf.

Anonymous:1891:VSJ

[1293] Anonymous. Vital statistics of Jews in the United States. Scientific American, 64(3):32–33, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
32a.pdf.

Anonymous:1891:AEA

[1294] Anonymous. The American Economic Association. Scientific American, 64(3):33, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
33.pdf.

Anonymous:1891:IPFa

[1295] Anonymous. An improvised portable foot bridge. Scientific American, 64(3):34, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
34.pdf.

Anonymous:1891:LUM

[1296] Anonymous. A leveler for use on meadow lands, etc. Scientific American, 64(3):34, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
34a.pdf.

Anonymous:1891:CPF

[1297] Anonymous. A compact and powerful flooring clamp. Scientific American, 64(3):35, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
35b.pdf.

Anonymous:1891:IAP

[1298] Anonymous. An improved air pump governor. Scientific American, 64(3):35, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
35a.pdf.

Anonymous:1891:PHF

[1299] Anonymous. A pen holder to facilitate making waved lines. Scientific American, 64(3):35, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
35c.pdf.

Anonymous:1891:RCF

[1300] Anonymous. A railway car fire extinguishing apparatus. Scientific American, 64(3):35, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
35.pdf.

Anonymous:1891:SRT

[1301] Anonymous. A simple and reliable trap. Scientific American, 64 (3):35–36, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
35d.pdf.

Anonymous:1891:Cc

[1302] Anonymous. Correspondence. Scientific American, 64(3):37, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-37.pdf.

Anonymous:1891:IGD

[1303] Anonymous. The “Iron Gates” of the Danube. Scientific American, 64(3):38, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
38a.pdf.

Anonymous:1891:PCB

[1304] Anonymous. Photography of a cannon ball in motion. Scientific American, 64(3):38, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
38.pdf.

Hopkins:1891:SEM

[1305] Geo. M. Hopkins. Small electric motor for amateurs. Scientific American, 64(3):39–40, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
39.pdf.

Anonymous:1891:LSR

[1306] Anonymous. Long splice for ropes. Scientific American, 64(3): 40–41, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
40.pdf.

Anonymous:1891:RPIc

[1307] Anonymous. Recently patented inventions. Scientific American, 64(3):41, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
41.pdf.

Anonymous:1891:BNBb

[1308] Anonymous. Books: New books and publications. Scientific American, 64(3):42, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
42.pdf.

Anonymous:1891:BPc

[1309] Anonymous. Business and personal. Scientific American, 64(3):42, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-42a.pdf.

Anonymous:1891:NQc

[1310] Anonymous. Notes and queries. Scientific American, 64(3):42–43, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-42b.pdf.

Anonymous:1891:NQd

[1311] Anonymous. Notes & queries. Scientific American, 64(3):42–43, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-42c.pdf.

Anonymous:1891:Ic

[1312] Anonymous. To inventors. Scientific American, 64(3):43, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-43.pdf.

Anonymous:1891:IIc

[1313] Anonymous. Index of inventions. Scientific American, 64(3):43–44, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-43a.pdf.

Anonymous:1891:PBa

[1314] Anonymous. Popular books. Scientific American, 64(3):44, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-44.pdf.

Anonymous:1891:Pa

[1315] Anonymous. Patents! Scientific American, 64(3):45, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-45a.pdf.

Anonymous:1891:PNPa

[1316] Anonymous. Patent notes: Patents! Scientific American, 64(3):45, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n3/pdf/scientificamerican01171891-45.pdf.

Anonymous:1891:SAa

[1317] Anonymous. The Scientific American. Scientific American, 64 (3):46, January 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n3/pdf/scientificamerican01171891-
46.pdf.

Hastings:1891:MPP

[1318] Fred. R. Hastings. The Manitou and Pike’s Peak Railway. Scientific American, 64(4):47, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
47.pdf.

Anonymous:1891:CKL

[1319] Anonymous. The constitution of the Koch lymph disclosed. Scientific American, 64(4):48, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
48.pdf.

Anonymous:1891:JW

[1320] Anonymous. Jarrah wood. Scientific American, 64(4):48, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-48a.pdf.

Anonymous:1891:CBSa

[1321] Anonymous. Celebration of the beginning of the second century of the American Patent System. Scientific American, 64(4):48–49, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-48b.pdf.

Anonymous:1891:BPBa

[1322] Anonymous. The Book Patent Bill. Scientific American, 64(4):49, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-49.pdf.

Anonymous:1891:ESBa

[1323] Anonymous. Electrical sporting bulletins. Scientific American, 64(4):50, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
50.pdf.

Anonymous:1891:SS

[1324] Anonymous. The science of smuggling. Scientific American, 64 (4):50, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
50a.pdf.

Anonymous:1891:SNW

[1325] Anonymous. The steamship North Wind — the business boat of the Lakes. Scientific American, 64(4):51, January 24, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n4/pdf/scientificamerican01241891-51a.pdf.

Hopkins:1891:SLS

[1326] Geo. M. Hopkins. Simple lamp socket and rheostat. Scientific American, 64(4):51, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
51.pdf.

Anonymous:1891:LDTa

[1327] Anonymous. Long distance telephony. Scientific American, 64 (4):51–52, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
51b.pdf.

Anonymous:1891:Cd

[1328] Anonymous. Correspondence. Scientific American, 64(4):53, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-53.pdf.

Anonymous:1891:LAE

[1329] Anonymous. The little ant-eater in the Zoological Garden at Berlin. Scientific American, 64(4):54, January 24, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n4/pdf/scientificamerican01241891-54a.pdf.

Anonymous:1891:RDG

[1330] Anonymous. A reversing driving gear for planers. Scientific American, 64(4):54, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
54.pdf.

Anonymous:1891:CCC

[1331] Anonymous. A cotton chopper and cultivator. Scientific American, 64(4):57, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
57.pdf.

Anonymous:1891:BPd

[1332] Anonymous. Business and personal. Scientific American, 64(4):58, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-58a.pdf.

Anonymous:1891:RPId

[1333] Anonymous. Recently patented inventions. Scientific American, 64(4):58, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
58.pdf.

Anonymous:1891:NQe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-58b.pdf.

Anonymous:1891:NQf

[1335] Anonymous. Notes & queries. Scientific American, 64(4):58–59, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-58c.pdf.

Anonymous:1891:BNBc

[1336] Anonymous. Books: New books and publications. Scientific American, 64(4):59, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
59.pdf.

Anonymous:1891:Id

[1337] Anonymous. To inventors. Scientific American, 64(4):59, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-59a.pdf.

Anonymous:1891:IId

[1338] Anonymous. Index of inventions. Scientific American, 64(4):59–60, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-59b.pdf.

Anonymous:1891:PBb

[1339] Anonymous. Popular books. Scientific American, 64(4):60, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-60.pdf.

Anonymous:1891:PNPb

[1340] Anonymous. Patent notes: Patents! Scientific American, 64(4):61, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n4/pdf/scientificamerican01241891-61.pdf.

Anonymous:1891:SAc

[1341] Anonymous. The Scientific American. Scientific American, 64 (4):62, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
62a.pdf.

Hopkins:1891:ESa

[1342] Geo. M. Hopkins. Experimental science. Scientific American, 64(4):62, January 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n4/pdf/scientificamerican01241891-
62.pdf.

Anonymous:1891:NC

[1343] Anonymous. The Nicaragua Canal. Scientific American, 64(5):63, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-63.pdf.

Anonymous:1891:NYW

[1344] Anonymous. The New York water supply — a great dam decided upon. Scientific American, 64(5):64, January 31, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n5/pdf/scientificamerican01311891-64.pdf.

Anonymous:1891:MCS

[1345] Anonymous. Mr. Cramp the shipbuilder, replies to his critics. Scientific American, 64(5):64–65, January 31, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n5/pdf/scientificamerican01311891-64a.pdf.

Anonymous:1891:PPFa

[1346] Anonymous. Position of the planets in February. Scientific American, 64(5):65, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
65.pdf.

Anonymous:1891:UER

[1347] Anonymous. The Unicycle Elevated Railway. Scientific American, 64(5):66, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
66.pdf.

Anonymous:1891:HHS

[1348] Anonymous. How Heron solved it. Scientific American, 64(5):67, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-67a.pdf.

Anonymous:1891:ITa

[1349] Anonymous. An improved tappet. Scientific American, 64(5):67, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-67.pdf.

Anonymous:1891:SRA

[1350] Anonymous. A swing readily adjustable for height. Scientific American, 64(5):67–68, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
67b.pdf.

Anonymous:1891:DRH

[1351] Anonymous. A device for reaching high shelving. Scientific American, 64(5):68, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
68.pdf.

Anonymous:1891:BPe

[1352] Anonymous. The belt problem. Scientific American, 64(5):69, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-69a.pdf.

Anonymous:1891:Ce

[1353] Anonymous. Correspondence. Scientific American, 64(5):69, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-69.pdf.

Anonymous:1891:MNa

[1354] Anonymous. Miscellaneous notes. Scientific American, 64(5):69, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-69b.pdf.

Anonymous:1891:CSF

[1355] Anonymous. The Canet system of firing torpedoes. Scientific American, 64(5):70–71, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
70.pdf.

Anonymous:1891:SABc

[1356] Anonymous. Scientific American building edition. Scientific American, 64(5):74, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
74a.pdf.

Anonymous:1891:RPIe

[1357] Anonymous. Recently patented inventions. Scientific American, 64(5):74, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
74.pdf.

Anonymous:1891:BPf

[1358] Anonymous. Business and personal. Scientific American, 64(5):75, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-75.pdf.

Anonymous:1891:Ie

[1359] Anonymous. To inventors. Scientific American, 64(5):75, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-75b.pdf.

Anonymous:1891:NQg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-75a.pdf.

Anonymous:1891:IIe

[1361] Anonymous. Index of inventions. Scientific American, 64(5):75–76, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-75c.pdf.

Anonymous:1891:BEa

[1362] Anonymous. Books on electricity. Scientific American, 64(5):76, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-76.pdf.

Anonymous:1891:MD

[1363] Anonymous. Mechanical drawing. Scientific American, 64(5):77, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-77a.pdf.

Anonymous:1891:PNPc

[1364] Anonymous. Patent notes: Patents. Scientific American, 64(5):77, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n5/pdf/scientificamerican01311891-77.pdf.

Anonymous:1891:SAd

[1365] Anonymous. The Scientific American. Scientific American, 64 (5):78, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
78a.pdf.

Anonymous:1891:SASa

[1366] Anonymous. The Scientific American supplement. Scientific American, 64(5):78, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
78b.pdf.

Hopkins:1891:ESb

[1367] Geo. M. Hopkins. Experimental science. Scientific American, 64(5):78, January 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n5/pdf/scientificamerican01311891-
78.pdf.

Anonymous:1891:HRI

[1368] Anonymous. The Hudson River ice harvest. Scientific American, 64(6):80, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
80.pdf.

Anonymous:1891:WRT

[1369] Anonymous. Wooden railway ties. Scientific American, 64(6): 80–81, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
80a.pdf.

Anonymous:1891:NGP

[1370] Anonymous. New gas and petroleum motor. Scientific American, 64(6):82–84, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
82.pdf.

Anonymous:1891:HDV

[1371] Anonymous. A holdback device for vehicles. Scientific American, 64(6):84, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
84a.pdf.

Anonymous:1891:IH

[1372] Anonymous. An improved halter. Scientific American, 64(6):84, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n6/pdf/scientificamerican02071891-84.pdf.

Anonymous:1891:Cf

[1373] Anonymous. Correspondence. Scientific American, 64(6):85, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n6/pdf/scientificamerican02071891-85a.pdf.

Anonymous:1891:PNc

[1374] Anonymous. Photographic notes. Scientific American, 64(6):85, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n6/pdf/scientificamerican02071891-85.pdf.

Anonymous:1891:THH

[1375] Anonymous. A tool holder that holds. Scientific American, 64 (6):85, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
85b.pdf.

Anonymous:1891:DEW

[1376] Anonymous. Destruction of electric wires by a snow storm. Scientific American, 64(6):86, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
86.pdf.

Anonymous:1891:FRS

[1377] Anonymous. The Fall River steamer Puritan. Scientific American, 64(6):87–89, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
87.pdf.

Anonymous:1891:CSE

[1378] Anonymous. A crate for shipping and exhibiting poultry, etc. Scientific American, 64(6):89, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
89a.pdf.

Anonymous:1891:ICSa

[1379] Anonymous. An improved carriage Spring. Scientific American, 64(6):89, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
89.pdf.

Anonymous:1891:BPg

[1380] Anonymous. Business and personal. Scientific American, 64 (6):90, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
90a.pdf.

Anonymous:1891:RPIf

[1381] Anonymous. Recently patented inventions. Scientific American, 64(6):90, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
90.pdf.

Anonymous:1891:NQh

[1382] Anonymous. Notes and queries. Scientific American, 64(6):90– 91, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
90b.pdf.

Anonymous:1891:NQi

[1383] Anonymous. Notes & queries. Scientific American, 64(6):90–91, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n6/pdf/scientificamerican02071891-90c.pdf.

Anonymous:1891:BNBd

[1384] Anonymous. Books: New books and publications. Scientific American, 64(6):91, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
91.pdf.

Anonymous:1891:If

[1385] Anonymous. To inventors. Scientific American, 64(6):91, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n6/pdf/scientificamerican02071891-91a.pdf.

Anonymous:1891:IIf

[1386] Anonymous. Index of inventions. Scientific American, 64(6): 91–92, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
91b.pdf.

Anonymous:1891:SAe

[1387] Anonymous. The Scientific American. Scientific American, 64 (6):92, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
92a.pdf.

Anonymous:1891:LBEb

[1388] Anonymous. List of books on electricity. Scientific American, 64(6):92, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
92.pdf.

Anonymous:1891:PNPd

[1389] Anonymous. Patent notes: Patents. Scientific American, 64 (6):94, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
94.pdf.

Hopkins:1891:ESc

[1390] Geo. M. Hopkins. Experimental science. Scientific American, 64(6):94, February 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n6/pdf/scientificamerican02071891-
94a.pdf.

Anonymous:1891:PMT

[1391] Anonymous. A petroleum motor tricycle. Scientific American, 64(7):95, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
95.pdf.

Anonymous:1891:LEa

[1392] Anonymous. Looking for employment. Scientific American, 64 (7):96–97, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
96.pdf.

Anonymous:1891:IBH

[1393] Anonymous. An inexpensive broom holder. Scientific American, 64(7):98, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
98a.pdf.

Anonymous:1891:ITb

[1394] Anonymous. An improved telephone. Scientific American, 64 (7):98, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
98.pdf.

Anonymous:1891:SSD

[1395] Anonymous. A spindle support and driving device. Scientific American, 64(7):98–99, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
98b.pdf.

Anonymous:1891:PPFb

[1396] Anonymous. Portable parlor fountain and flower stand. Scientific American, 64(7):99, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
99.pdf.

Anonymous:1891:CT

[1397] Anonymous. A curious tombstone. Scientific American, 64(7): 99–100, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
99a.pdf.

Anonymous:1891:Cg

[1398] Anonymous. Correspondence. Scientific American, 64(7):101, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n7/pdf/scientificamerican02141891-101.pdf.

Anonymous:1891:CER

[1399] Anonymous. Concurrent electric reports of races. Scientific American, 64(7):102, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
102.pdf.

Anonymous:1891:NDP

[1400] Anonymous. New depot of the Pennsylvania Railroad. Scientific American, 64(7):102–105, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
102a.pdf.

Anonymous:1891:IDV

[1401] Anonymous. Invention in 1889 A.D. vs. invention B.C. Scientific American, 64(7):105–106, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
105.pdf.

Anonymous:1891:RPIg

[1402] Anonymous. Recently patented inventions. Scientific American, 64 (7):106–107, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
106.pdf.

Anonymous:1891:BPh

[1403] Anonymous. Business and personal. Scientific American, 64(7): 107, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
107.pdf.

Anonymous:1891:NQj

[1404] Anonymous. Notes and queries. Scientific American, 64(7):107, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n7/pdf/scientificamerican02141891-107a.pdf.

Anonymous:1891:IIg

[1405] Anonymous. Index of inventions. Scientific American, 64(7): 107–108, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
107c.pdf.

Anonymous:1891:IIh

[1406] Anonymous. Index of inventions. Scientific American, 64(7): 107–108, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
107b.pdf.

Anonymous:1891:LABa

[1407] Anonymous. List of architectural books. Scientific American, 64(7):108, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
108.pdf.

Anonymous:1891:SAPa

[1408] Anonymous. The Scientific American publications for 1891. Scientific American, 64(7):109, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
109b.pdf.

Anonymous:1891:PNPe

[1409] Anonymous. Patent notes: Patents! Scientific American, 64 (7):109, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
109.pdf.

Hopkins:1891:ESd

[1410] Geo. M. Hopkins. Experimental science. Scientific American, 64 (7):109, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
109a.pdf.

Anonymous:1891:SASb

[1411] Anonymous. The Scientific American supplement. Scientific American, 64(7):110, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n7/pdf/scientificamerican02141891-
110.pdf.

Anonymous:1891:BEb

[1412] Anonymous. Building edition. Scientific American, 64(7):110, February 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n7/pdf/scientificamerican02141891-110a.pdf.

Anonymous:1891:POR

[1413] Anonymous. Patent office report for 1890. Scientific American, 64(8):112, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
112.pdf.

Anonymous:1891:INS

[1414] Anonymous. Improvement needed in secondary batteries. Scientific American, 64(8):112–113, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
112a.pdf.

Anonymous:1891:PAF

[1415] Anonymous. A proposed apparatus for a fall of 1,000 feet. Scientific American, 64(8):114, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
114a.pdf.

Anonymous:1891:RPR

[1416] Anonymous. A roller pressure regulator for cane mills. Scientific American, 64(8):114, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
114.pdf.

Anonymous:1891:IBT

[1417] Anonymous. An improved bicycle treadle. Scientific American, 64(8):115, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
115a.pdf.

Anonymous:1891:IGC

[1418] Anonymous. An improved gauge cock. Scientific American, 64 (8):115, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
115.pdf.

Anonymous:1891:VOH

[1419] Anonymous. A velocipede operated by hand levers. Scientific American, 64(8):115, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
115b.pdf.

Anonymous:1891:Ch

[1420] Anonymous. Correspondence. Scientific American, 64(8):116, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n8/pdf/scientificamerican02211891-116.pdf.

Anonymous:1891:USLa

[1421] Anonymous. The United States Life Saving Service. Scientific American, 64(8):117, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
117.pdf.

Anonymous:1891:TEE

[1422] Anonymous. Triple expansion engines for a tugboat. Scientific American, 64(8):118, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
118.pdf.

Anonymous:1891:SSC

[1423] Anonymous. Serpollet’s steam carriage. Scientific American, 64 (8):119–121, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
119.pdf.

Hopkins:1891:CAA

[1424] Geo. M. Hopkins. Centrifugal action of air. Scientific American, 64(8):121, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
121.pdf.

Anonymous:1891:BPi

[1425] Anonymous. Business and personal. Scientific American, 64(8): 122, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
122a.pdf.

Anonymous:1891:RPIh

[1426] Anonymous. Recently patented inventions. Scientific American, 64(8):122, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
122.pdf.

Anonymous:1891:NQk

[1427] Anonymous. Notes and queries. Scientific American, 64(8):122– 123, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
122b.pdf.

Anonymous:1891:NQl

[1428] Anonymous. Notes & queries. Scientific American, 64(8):122– 123, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
122c.pdf.

Anonymous:1891:Ig

[1429] Anonymous. To inventors. Scientific American, 64(8):123, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n8/pdf/scientificamerican02211891-123.pdf.

Anonymous:1891:IIi

[1430] Anonymous. Index of inventions. Scientific American, 64(8): 123–124, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
123a.pdf.

Anonymous:1891:EM

[1431] Anonymous. Engineering and mechanics. Scientific American, 64 (8):124, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
124.pdf.

Anonymous:1891:PNPf

[1432] Anonymous. Patent notes: Patents! Scientific American, 64 (8):125, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
125.pdf.

Anonymous:1891:SAf

[1433] Anonymous. The Scientific American. Scientific American, 64 (8):126, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
126b.pdf.

Anonymous:1891:PNPg

[1434] Anonymous. Patent notes: Patents. Scientific American, 64 (8):126, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
126a.pdf.

Hopkins:1891:ESe

[1435] Geo. M. Hopkins. Experimental science. Scientific American, 64 (8):126, February 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n8/pdf/scientificamerican02211891-
126.pdf.

Anonymous:1891:MHGa

[1436] Anonymous. Manufacture of heavy guns for U.S. Navy. Scientific American, 64(9):127, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
127.pdf.

Anonymous:1891:MNE

[1437] Anonymous. Meeting of the National Electric Light Association. Scientific American, 64(9):128, February 28, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n9/pdf/scientificamerican02281891-128.pdf.

Anonymous:1891:PPMa

[1438] Anonymous. Position of the planets in March. Scientific American, 64(9):129, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
129.pdf.

Anonymous:1891:LDTb

[1439] Anonymous. Long distance telephone concerts. Scientific American, 64(9):130, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
130a.pdf.

Anonymous:1891:SSF

[1440] Anonymous. A steam snow flow. Scientific American, 64(9):130, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n9/pdf/scientificamerican02281891-130.pdf.

Anonymous:1891:DMH

[1441] Anonymous. A driving mechanism for hand drills. Scientific American, 64(9):130–131, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
130b.pdf.

Anonymous:1891:IRE

[1442] Anonymous. An improved rotary engine. Scientific American, 64(9):131, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
131.pdf.

Anonymous:1891:JE

[1443] Anonymous. The Jamaica Exhibition. Scientific American, 64 (9):131–132, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
131a.pdf.

Anonymous:1891:CNS

[1444] Anonymous. The Chilian Nitrate of Soda Mines and Works. Scientific American, 64(9):135, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
135.pdf.

Anonymous:1891:SMA

[1445] Anonymous. Simple milling attachment for foot lathes. Scientific American, 64(9):135, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
135a.pdf.

Anonymous:1891:OOL

[1446] Anonymous. Olive oil lamp. Scientific American, 64(9):135–137, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n9/pdf/scientificamerican02281891-135b.pdf.

Anonymous:1891:ABT

[1447] Anonymous. An American banyan tree. Scientific American, 64 (9):137, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
137.pdf.

Anonymous:1891:RPIi

[1448] Anonymous. Recently patented inventions. Scientific American, 64(9):138, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
138.pdf.

Anonymous:1891:NQm

[1449] Anonymous. Notes and queries. Scientific American, 64(9):138– 139, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
138a.pdf.

Anonymous:1891:NQn

[1450] Anonymous. Notes & queries. Scientific American, 64(9):138– 139, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
138b.pdf.

Anonymous:1891:BNBe

[1451] Anonymous. Books: New books and publications. Scientific American, 64(9):139, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
139.pdf.

Anonymous:1891:Ij

[1452] Anonymous. To inventors. Scientific American, 64(9):139, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n9/pdf/scientificamerican02281891-139a.pdf.

Anonymous:1891:IIj

[1453] Anonymous. Index of inventions. Scientific American, 64(9): 139–140, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
139b.pdf.

Anonymous:1891:SAPb

[1454] Anonymous. The Scientific American publications for 1891. Scientific American, 64(9):141, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
141b.pdf.

Anonymous:1891:PNPh

[1455] Anonymous. Patent notes: Patents! Scientific American, 64 (9):141, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
141.pdf.

Hopkins:1891:ESf

[1456] Geo. M. Hopkins. Experimental science. Scientific American, 64 (9):141, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
141a.pdf.

Anonymous:1891:SAEa

[1457] Anonymous. The Scientific American established 1846. Scientific American, 64(9):142, February 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n9/pdf/scientificamerican02281891-
142.pdf.

Anonymous:1891:AFW

[1458] Anonymous. Another Faraday wanted. Scientific American, 64(10):144, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-144a.pdf.

Anonymous:1891:CFPa

[1459] Anonymous. Columbian Fair progress. Scientific American, 64(10):144, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-144.pdf.

Anonymous:1891:INI

[1460] Anonymous. Interesting Naval information by the Secretary of the Navy. Scientific American, 64(10):144–145, March 7, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n10/pdf/scientificamerican03071891-144b.pdf.

Anonymous:1891:CBW

[1461] Anonymous. A combined bath and wash tub. Scientific American, 64(10):146, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
146a.pdf.

Anonymous:1891:HUG

[1462] Anonymous. A heater for use with a gas burner. Scientific American, 64(10):146, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
146.pdf.

Anonymous:1891:IIk

[1463] Anonymous. An improved index. Scientific American, 64(10):146, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-146b.pdf.

Anonymous:1891:IPT

[1464] Anonymous. An improved package tie. Scientific American, 64(10):146, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-146c.pdf.

Anonymous:1891:ISBa

[1465] Anonymous. Improved sounding bridge for pianos, etc. Scientific American, 64(10):146–147, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
146d.pdf.

Anonymous:1891:Ci

[1466] Anonymous. A centenarian. Scientific American, 64(10):147, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-147.pdf.

Anonymous:1891:SB

[1467] Anonymous. A standard bicycle. Scientific American, 64(10):147, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-147a.pdf.

Anonymous:1891:NFB

[1468] Anonymous. New fire boat for New York City — The New Yorker. Scientific American, 64(10):148, March 7, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n10/pdf/scientificamerican03071891-148.pdf.

Anonymous:1891:Cj

[1469] Anonymous. Correspondence. Scientific American, 64(10):149, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-149.pdf.

Anonymous:1891:PNd

[1470] Anonymous. Photographic notes. Scientific American, 64(10):149, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-149a.pdf.

Anonymous:1891:PAS

[1471] Anonymous. The Pennington air ship. Scientific American, 64(10):150, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-150.pdf.

Anonymous:1891:PT

[1472] Anonymous. The Proctor Tower. Scientific American, 64(10):151, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-151.pdf.

Anonymous:1891:SSM

[1473] Anonymous. A stopping and signaling mechanism for locomotives. Scientific American, 64(10):151, March 7, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n10/pdf/scientificamerican03071891-151a.pdf.

Anonymous:1891:EPC

[1474] Anonymous. An efficient pipe cutting machine. Scientific American, 64(10):151–152, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
151b.pdf.

Anonymous:1891:AWE

[1475] Anonymous. An ancient water elevator. Scientific American, 64 (10):153, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
153.pdf.

Anonymous:1891:SABd

[1476] Anonymous. Scientific American building edition. Scientific American, 64(10):154, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
154b.pdf.

Anonymous:1891:BPj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-154c.pdf.

Anonymous:1891:NBPa

[1478] Anonymous. New books and publications. Scientific American, 64(10):154, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
154a.pdf.

Anonymous:1891:RPIj

[1479] Anonymous. Recently patented inventions. Scientific American, 64(10):154, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n10/pdf/scientificamerican03071891-
154.pdf.

Anonymous:1891:NQo

[1480] Anonymous. Notes and queries. Scientific American, 64(10):154–155, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-154d.pdf.

Anonymous:1891:Ik

[1481] Anonymous. To inventors. Scientific American, 64(10):155, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-155.pdf.

Anonymous:1891:IIl

[1482] Anonymous. Index of inventions. Scientific American, 64(10):155–156, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-155a.pdf.

Anonymous:1891:PNPi

[1483] Anonymous. Patent notes: Patents. Scientific American, 64(10):157, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-157.pdf.

Anonymous:1891:SAg

[1484] Anonymous. The Scientific American. Scientific American, 64(10):158, March 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n10/pdf/scientificamerican03071891-158.pdf.

Anonymous:1891:MHGb

[1485] Anonymous. Manufacture of heavy guns for U.S. Navy. Scientific American, 64(11):159, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
159.pdf.

Anonymous:1891:IEa

[1486] Anonymous. Indoor exercise. Scientific American, 64(11):160, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-160.pdf.

Pike:1891:PS

[1487] Nicolas Pike. Poison of snakes. Scientific American, 64(11):161, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-161.pdf.

Anonymous:1891:BFF

[1488] Anonymous. Boiler flue flanging machine. Scientific American, 64(11):162, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
162a.pdf.

Anonymous:1891:FBD

[1489] Anonymous. A foot bath device. Scientific American, 64(11):162, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-162b.pdf.

Anonymous:1891:WCB

[1490] Anonymous. A window cleaning brush. Scientific American, 64(11):162, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-162.pdf.

Anonymous:1891:ISV

[1491] Anonymous. An improved safety valve. Scientific American, 64(11):163, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-163.pdf.

Anonymous:1891:IWa

[1492] Anonymous. An improved wrench. Scientific American, 64(11):163, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-163b.pdf.

Anonymous:1891:PMVa

[1493] Anonymous. A propelling mechanism for vehicles. Scientific American, 64(11):163, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
163a.pdf.

Anonymous:1891:SBS

[1494] Anonymous. A safety bolt for Spring latches. Scientific American, 64(11):163, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
163c.pdf.

Anonymous:1891:EBB

[1495] Anonymous. An electric belt and body attachments. Scientific American, 64(11):163–165, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
163d.pdf.

Anonymous:1891:SESa

[1496] Anonymous. A smoke effect with a single lantern. Scientific American, 64(11):165, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
165.pdf.

Anonymous:1891:PFM

[1497] Anonymous. A pneumatic flying machine. Scientific American, 64 (11):167–168, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
167.pdf.

Anonymous:1891:ICa

[1498] Anonymous. An improved churn. Scientific American, 64(11):168, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-168.pdf.

Anonymous:1891:CAa

[1499] Anonymous. Church acoustics. Scientific American, 64(11):169, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-169.pdf.

Anonymous:1891:SABe

[1500] Anonymous. Scientific American building edition. Scientific American, 64(11):170, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
170b.pdf.

Anonymous:1891:NBPb

[1501] Anonymous. New books and publications. Scientific American, 64(11):170, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
170a.pdf.

Anonymous:1891:RPIk

[1502] Anonymous. Recently patented inventions. Scientific American, 64(11):170, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
170.pdf.

Anonymous:1891:BPk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-171.pdf.

Anonymous:1891:Il

[1504] Anonymous. To inventors. Scientific American, 64(11):171, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-171b.pdf.

Anonymous:1891:NQp

[1505] Anonymous. Notes and queries. Scientific American, 64(11):171, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-171a.pdf.

Anonymous:1891:IIm

[1506] Anonymous. Index of inventions. Scientific American, 64(11):171–172, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-171c.pdf.

Anonymous:1891:LABb

[1507] Anonymous. List of architectural books. Scientific American, 64(11):172, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n11/pdf/scientificamerican03141891-
172.pdf.

Anonymous:1891:ES

[1508] Anonymous. Experimental science. Scientific American, 64(11):173, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-173a.pdf.

Anonymous:1891:PNPj

[1509] Anonymous. Patent notes: Patents. Scientific American, 64(11):173, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-173.pdf.

Anonymous:1891:SAi

[1510] Anonymous. The Scientific American. Scientific American, 64(11):174, March 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n11/pdf/scientificamerican03141891-174.pdf.

Anonymous:1891:NSFa

[1511] Anonymous. The new steel floating derrick at the Brooklyn Navy Yard. Scientific American, 64(12):175–176, March 21, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n12/pdf/scientificamerican03211891-175.pdf.

Anonymous:1891:GGJ

[1512] Anonymous. The great guns of the Japanese Navy. Scientific American, 64(12):176, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
176.pdf.

Anonymous:1891:FSO

[1513] Anonymous. The falling off in speed of our war ships. Scientific American, 64(12):176–177, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
176a.pdf.

Anonymous:1891:GFR

[1514] Anonymous. A grate for furnaces, ranges, etc. Scientific American, 64(12):178, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
178.pdf.

Anonymous:1891:ICSb

[1515] Anonymous. An improved casing for steam pipes. Scientific American, 64(12):178, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
178a.pdf.

Anonymous:1891:ILa

[1516] Anonymous. An improved lock. Scientific American, 64(12):178, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-178b.pdf.

Anonymous:1891:NRT

[1517] Anonymous. A new railroad tie, rail fastener, and rail. Scientific American, 64(12):178–179, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
178c.pdf.

Anonymous:1891:AMS

[1518] Anonymous. An all-metal Spring vehicle wheel. Scientific American, 64(12):179, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
179.pdf.

Anonymous:1891:DCL

[1519] Anonymous. A double combination lock. Scientific American, 64(12):179, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
179b.pdf.

Anonymous:1891:IHS

[1520] Anonymous. An improved hay stacker. Scientific American, 64(12):179, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-179a.pdf.

Anonymous:1891:SDI

[1521] Anonymous. A safety device for inclined roads. Scientific American, 64(12):179–180, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
179c.pdf.

Sloane:1891:AVA

[1522] T. O’conor Sloane. Ampere and volt analogies. Scientific American, 64(12):180–181, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
180.pdf.

Anonymous:1891:UFH

[1523] Anonymous. Universal file handle. Scientific American, 64(12):181, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-181.pdf.

Anonymous:1891:DWFa

[1524] Anonymous. Destructive winter floods in Connecticut. Scientific American, 64(12):182–184, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
182.pdf.

Anonymous:1891:CFa

[1525] Anonymous. The care of the feet. Scientific American, 64(12):185, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-185.pdf.

Anonymous:1891:SABf

[1526] Anonymous. Scientific American building edition. Scientific American, 64(12):186, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
186b.pdf.

Anonymous:1891:NBPc

[1527] Anonymous. New books and publications. Scientific American, 64(12):186, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
186a.pdf.

Anonymous:1891:RPIl

[1528] Anonymous. Recently patented inventions. Scientific American, 64(12):186, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
186.pdf.

Anonymous:1891:BPl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-187.pdf.

Anonymous:1891:Im

[1530] Anonymous. To inventors. Scientific American, 64(12):187, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-187b.pdf.

Anonymous:1891:NQq

[1531] Anonymous. Notes and queries. Scientific American, 64(12):187, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-187a.pdf.

Anonymous:1891:IIn

[1532] Anonymous. Index of inventions. Scientific American, 64(12):187–188, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-187c.pdf.

Anonymous:1891:LBEc

[1533] Anonymous. List of books on electricity. Scientific American, 64(12):188, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
188.pdf.

Anonymous:1891:PNPk

[1534] Anonymous. Patent notes: Patents. Scientific American, 64(12):189, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-189.pdf.

Anonymous:1891:UBa

[1535] Anonymous. Useful books. Scientific American, 64(12):189, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-189a.pdf.

Hopkins:1891:ESg

[1536] Geo. M. Hopkins. Experimental science. Scientific American, 64(12):189, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n12/pdf/scientificamerican03211891-
189b.pdf.

Anonymous:1891:SAj

[1537] Anonymous. The Scientific American. Scientific American, 64(12):190, March 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n12/pdf/scientificamerican03211891-190.pdf.

Anonymous:1891:BUR

[1538] Anonymous. The Boynton Unicycle Railroad. Scientific American, 64(13):191–192, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
191.pdf.

Anonymous:1891:CBSb

[1539] Anonymous. The celebration of the beginning of the second century of the American Patent System. Scientific American, 64(13):192–193, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-192.pdf.

Anonymous:1891:PPAa

[1540] Anonymous. Position of the planets in April. Scientific American, 64(13):193–194, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
193.pdf.

Anonymous:1891:OMV

[1541] Anonymous. Oriental method of vaulting. Scientific American, 64(13):194, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
194.pdf.

Anonymous:1891:ICR

[1542] Anonymous. An improved cable railway crossing. Scientific American, 64(13):195, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
195.pdf.

Anonymous:1891:SDS

[1543] Anonymous. A Spring device for singletrees. Scientific American, 64(13):195, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
195a.pdf.

Anonymous:1891:CCT

[1544] Anonymous. A clasp for connecting timbers. Scientific American, 64(13):195–197, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
195b.pdf.

Anonymous:1891:NPN

[1545] Anonymous. A new park for New York. Scientific American, 64(13):198, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-198.pdf.

Seal:1891:FH

[1546] Wm. P. Seal. A fish habitation. Scientific American, 64(13):199–201, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-199.pdf.

Anonymous:1891:ESG

[1547] Anonymous. Experiment in specific gravity of fluids. Scientific American, 64(13):201, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
201.pdf.

Anonymous:1891:SABg

[1548] Anonymous. Scientific American building edition. Scientific American, 64(13):202, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
202b.pdf.

Anonymous:1891:BPBb

[1549] Anonymous. Business and personal: Business and personal. Scientific American, 64(13):202, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
202c.pdf.

Anonymous:1891:NBPd

[1550] Anonymous. New books and publications. Scientific American, 64(13):202, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
202a.pdf.

Anonymous:1891:RPIm

[1551] Anonymous. Recently patented inventions. Scientific American, 64(13):202, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
202.pdf.

Anonymous:1891:NQr

[1552] Anonymous. Notes & queries. Scientific American, 64(13):202–203, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-202d.pdf.

Anonymous:1891:In

[1553] Anonymous. To inventors. Scientific American, 64(13):203, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-203.pdf.

Anonymous:1891:IIo

[1554] Anonymous. Index of inventions. Scientific American, 64(13):203–204, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-203a.pdf.

Anonymous:1891:Aa

[1555] Anonymous. Advertisements. Scientific American, 64(13):204, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-204.pdf.

Anonymous:1891:LB

[1556] Anonymous. List of books. Scientific American, 64(13):204, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-204a.pdf.

Anonymous:1891:SAPc

[1557] Anonymous. The Scientific American publications for 1891. Scientific American, 64(13):205, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
205b.pdf.

Anonymous:1891:PNPl

[1558] Anonymous. Patent notes: Patents. Scientific American, 64(13):205, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n13/pdf/scientificamerican03281891-205.pdf.

Hopkins:1891:ESh

[1559] Geo. M. Hopkins. Experimental science. Scientific American, 64(13):205, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
205a.pdf.

Anonymous:1891:SAEb

[1560] Anonymous. The Scientific American established I846. Scientific American, 64(13):206, March 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n13/pdf/scientificamerican03281891-
206.pdf.

Anonymous:1891:OWAa

[1561] Anonymous. Original work in America. Scientific American, 64 (14):208–209, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
208.pdf.

Anonymous:1891:MNb

[1562] Anonymous. Miscellaneous notes. Scientific American, 64(14):209–210, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-209.pdf.

Anonymous:1891:IWb

[1563] Anonymous. An improved wrench. Scientific American, 64(14):210, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-210c.pdf.

Anonymous:1891:PMVb

[1564] Anonymous. A propelling mechanism for vehicles. Scientific American, 64(14):210, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
210b.pdf.

Anonymous:1891:QSM

[1565] Anonymous. A quilter for sewing machines. Scientific American, 64(14):210, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
210a.pdf.

Anonymous:1891:RAC

[1566] Anonymous. A readily applied car replacer. Scientific American, 64(14):210, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
210.pdf.

Anonymous:1891:FPW

[1567] Anonymous. Fireproof partition walls. Scientific American, 64(14):211, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-211a.pdf.

Anonymous:1891:MS

[1568] Anonymous. A mechanical stoker. Scientific American, 64(14):211, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-211b.pdf.

Anonymous:1891:THL

[1569] Anonymous. A tool holder for lathes, planers etc. Scientific American, 64(14):211, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
211.pdf.

Anonymous:1891:Ck

[1570] Anonymous. Correspondence. Scientific American, 64(14):212, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-212.pdf.

Anonymous:1891:PC

[1571] Anonymous. The patent centennial. Scientific American, 64 (14):213–215, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
213.pdf.

Hopkins:1891:OPO

[1572] Geo. M. Hopkins. Optical projection of opaque objects. Scientific American, 64(14):216–217, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
216.pdf.

Anonymous:1891:BPm

[1573] Anonymous. Business and personal. Scientific American, 64(14):218, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-218c.pdf.

Anonymous:1891:NBPe

[1574] Anonymous. New books and publications. Scientific American, 64(14):218, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
218a.pdf.

Anonymous:1891:RPIn

[1575] Anonymous. Recently patented inventions. Scientific American, 64(14):218, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
218.pdf.

Lovell:1891:SL

[1576] John P. Lovell. A successful life. Scientific American, 64(14):218, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-218b.pdf.

Anonymous:1891:NQs

[1577] Anonymous. Notes and queries. Scientific American, 64(14):218–219, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-218d.pdf.

Anonymous:1891:Io

[1578] Anonymous. To inventors. Scientific American, 64(14):219, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-219.pdf.

Anonymous:1891:IIp

[1579] Anonymous. Index of inventions. Scientific American, 64(14):219, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-219a.pdf.

Anonymous:1891:Ab

[1580] Anonymous. Advertisements. Scientific American, 64(14):220, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-220.pdf.

Anonymous:1891:UBb

[1581] Anonymous. Useful books. Scientific American, 64(14):220, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-220a.pdf.

Anonymous:1891:SAPd

[1582] Anonymous. The Scientific American publications for 1891. Scientific American, 64(14):221, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
221b.pdf.

Anonymous:1891:PNPm

[1583] Anonymous. Patent notes: Patents. Scientific American, 64(14):221, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n14/pdf/scientificamerican04041891-221.pdf.

Hopkins:1891:ESi

[1584] Geo. M. Hopkins. Experimental science. Scientific American, 64 (14):221, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
221a.pdf.

Anonymous:1891:SAEc

[1585] Anonymous. The Scientific American established 1846. Scientific American, 64(14):222, April 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n14/pdf/scientificamerican04041891-
222.pdf.

Anonymous:1891:BN

[1586] Anonymous. Build up the Navy. Scientific American, 64(15):224, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-224a.pdf.

Anonymous:1891:RPF

[1587] Anonymous. The reduction of patent fees. Scientific American, 64(15):224, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
224.pdf.

Anonymous:1891:PCCa

[1588] Anonymous. The patent centennial celebration. Scientific American, 64(15):224–225, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
224b.pdf.

Anonymous:1891:IAM

[1589] Anonymous. An improved astronomical mirror. Scientific American, 64(15):226, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
226c.pdf.

Anonymous:1891:IPH

[1590] Anonymous. An improved pipe hanger. Scientific American, 64(15):226, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-226a.pdf.

Anonymous:1891:RBH

[1591] Anonymous. Reversible blind hinges. Scientific American, 64(15):226, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-226d.pdf.

Anonymous:1891:SAGa

[1592] Anonymous. A self adjusting girth fastener. Scientific American, 64(15):226, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
226b.pdf.

Anonymous:1891:TDC

[1593] Anonymous. A tool for dehorning calves. Scientific American, 64(15):226, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
226.pdf.

Anonymous:1891:ED

[1594] Anonymous. Electroplating the dead. Scientific American, 64 (15):227–229, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
227.pdf.

Anonymous:1891:NNP

[1595] Anonymous. A notable new Pullman passenger train. Scientific American, 64(15):229, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
229.pdf.

Anonymous:1891:CSR

[1596] Anonymous. The Chignecto Ship Railway. Scientific American, 64(15):230, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
230.pdf.

Anonymous:1891:ASF

[1597] Anonymous. Automatic sawdust feed for furnaces. Scientific American, 64(15):231, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
231.pdf.

Anonymous:1891:Cl

[1598] Anonymous. Correspondence. Scientific American, 64(15):232, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-232.pdf.

Anonymous:1891:BWT

[1599] Anonymous. Brownlow & Warner’s truss. Scientific American, 64(15):233, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
233d.pdf.

Anonymous:1891:CFPb

[1600] Anonymous. A convenient fruit packing press. Scientific American, 64(15):233, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
233a.pdf.

Anonymous:1891:EGM

[1601] Anonymous. An envelope gum moistener. Scientific American, 64(15):233, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
233c.pdf.

Anonymous:1891:IDS

[1602] Anonymous. An improved drawer stop. Scientific American, 64(15):233, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-233.pdf.

Anonymous:1891:LDS

[1603] Anonymous. The Lovell diamond safety bicycle. Scientific American, 64(15):233, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
233b.pdf.

Anonymous:1891:SABh

[1604] Anonymous. Scientific American building edition. Scientific American, 64(15):234, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
234b.pdf.

Anonymous:1891:BPn

[1605] Anonymous. Business and personal. Scientific American, 64(15):234, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-234c.pdf.

Anonymous:1891:NBPf

[1606] Anonymous. New books and publications. Scientific American, 64(15):234, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
234a.pdf.

Anonymous:1891:RPIo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
234.pdf.

Anonymous:1891:NQt

[1608] Anonymous. Notes and queries. Scientific American, 64(15):234–235, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-234d.pdf.

Anonymous:1891:Ip

[1609] Anonymous. To inventors. Scientific American, 64(15):235, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-235.pdf.

Anonymous:1891:IIq

[1610] Anonymous. Index of inventions. Scientific American, 64(15):235–236, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-235a.pdf.

Anonymous:1891:Ac

[1611] Anonymous. Advertisements. Scientific American, 64(15):236, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-236.pdf.

Anonymous:1891:UBc

[1612] Anonymous. Useful books. Scientific American, 64(15):236, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-236a.pdf.

Anonymous:1891:SAPe

[1613] Anonymous. The Scientific American publications for 1891. Scientific American, 64(15):237, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
237b.pdf.

Anonymous:1891:PNPn

[1614] Anonymous. Patent notes: Patents. Scientific American, 64(15):237, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n15/pdf/scientificamerican04111891-237.pdf.

Hopkins:1891:ESj

[1615] Geo. M. Hopkins. Experimental science. Scientific American, 64 (15):237, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
237a.pdf.

Anonymous:1891:SAEd

[1616] Anonymous. The Scientific American established 1846. Scientific American, 64(15):238, April 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n15/pdf/scientificamerican04111891-
238.pdf.

Anonymous:1891:TSA

[1617] Anonymous. Terms for the Scientific American. Scientific American, 64(16):240, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
240.pdf.

Anonymous:1891:TCSa

[1618] Anonymous. Table of contents of Scientific American supplement. Scientific American, 64(16):240, April 18, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n16/pdf/scientificamerican04181891-240a.pdf.

Anonymous:1891:SON

[1619] Anonymous. Speed in our new Navy. Scientific American, 64 (16):240–241, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
240b.pdf.

Anonymous:1891:ISBb

[1620] Anonymous. The Illinois state building at the World’s Fair. Scientific American, 64(16):242, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
242.pdf.

Anonymous:1891:PCCb

[1621] Anonymous. The Patent Centennial Celebration — April 8, 9, and 10. Scientific American, 64(16):243–245, April 18, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n16/pdf/scientificamerican04181891-243.pdf.

Anonymous:1891:Cm

[1622] Anonymous. Correspondence. Scientific American, 64(16):245–246, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-245.pdf.

Anonymous:1891:HPM

[1623] Anonymous. A hanger for picture or mirror frames. Scientific American, 64(16):246, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
246a.pdf.

Anonymous:1891:HSP

[1624] Anonymous. How stocks of pig iron are held. Scientific American, 64(16):246, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
246.pdf.

Anonymous:1891:ICP

[1625] Anonymous. Improved cloth pressing machine. Scientific American, 64(16):247–249, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
247.pdf.

Anonymous:1891:CBM

[1626] Anonymous. A card-board making machine. Scientific American, 64(16):249, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
249a.pdf.

Anonymous:1891:DWFb

[1627] Anonymous. The Darragh water filter. Scientific American, 64(16):249, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-249.pdf.

Anonymous:1891:SABi

[1628] Anonymous. Scientific American building edition. Scientific American, 64(16):250, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
250a.pdf.

Anonymous:1891:BPo

[1629] Anonymous. Business and personal. Scientific American, 64(16):250, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-250b.pdf.

Anonymous:1891:NBPg

[1630] Anonymous. New books and publications. Scientific American, 64(16):250, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
250c.pdf.

Anonymous:1891:RPIp

[1631] Anonymous. Recently patented inventions. Scientific American, 64(16):250, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
250.pdf.

Anonymous:1891:Iq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-251a.pdf.

Anonymous:1891:NQHa

[1633] Anonymous. Notes and queries: Hints to correspondents. Scientific American, 64(16):251, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
251.pdf.

Anonymous:1891:IIr

[1634] Anonymous. Index of inventions. Scientific American, 64(16):251–252, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-251b.pdf.

Anonymous:1891:LBEd

[1635] Anonymous. List of books on electricity. Scientific American, 64(16):252, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
252.pdf.

Anonymous:1891:Pd

[1636] Anonymous. Patents! Scientific American, 64(16):253, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-253.pdf.

Anonymous:1891:UBd

[1637] Anonymous. Useful books! Scientific American, 64(16):253, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-253a.pdf.

Hopkins:1891:ESk

[1638] Geo. M. Hopkins. Experimental science. Scientific American, 64 (16):253, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n16/pdf/scientificamerican04181891-
253b.pdf.

Anonymous:1891:SAk

[1639] Anonymous. The Scientific American. Scientific American, 64(16):254, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-254.pdf.

Anonymous:1891:BEc

[1640] Anonymous. Building edition. Scientific American, 64(16):254, April 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n16/pdf/scientificamerican04181891-254a.pdf.

Anonymous:1891:UCL

[1641] Anonymous. U.S. coast line of battle ships. Scientific American, 64(17):255, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
255.pdf.

Anonymous:1891:TCSb

[1642] Anonymous. Table of contents of Scientific American supplement. Scientific American, 64(17):256, April 25, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n17/pdf/scientificamerican04251891-256a.pdf.

Anonymous:1891:USP

[1643] Anonymous. The United States’ position in a defensive war. Scientific American, 64(17):256, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
256.pdf.

Anonymous:1891:SPa

[1644] Anonymous. Samuel Plimsoll. Scientific American, 64(17):256–257, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-256b.pdf.

Anonymous:1891:SPb

[1645] Anonymous. Samuel Plimsoll. Scientific American, 64(17):256–257, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-256c.pdf.

Anonymous:1891:MSW

[1646] Anonymous. Metal as a substitute for wood in railroad ties. Scientific American, 64(17):257, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
257.pdf.

Anonymous:1891:RPP

[1647] Anonymous. Rotary press for printing in colors. Scientific American, 64(17):258, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
258.pdf.

Anonymous:1891:NTW

[1648] Anonymous. New tubular well. Scientific American, 64(17):259, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-259.pdf.

Anonymous:1891:TI

[1649] Anonymous. Theatrical illusions. Scientific American, 64(17):259, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-259a.pdf.

Anonymous:1891:PFC

[1650] Anonymous. Platform for freight cars. Scientific American, 64 (17):259–261, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
259b.pdf.

Anonymous:1891:HRS

[1651] Anonymous. The horse race on the stage. Scientific American, 64(17):263–264, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
263.pdf.

Anonymous:1891:ASW

[1652] Anonymous. Apparatus for scouring water-ways. Scientific American, 64(17):265–266, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
265.pdf.

Anonymous:1891:SABj

[1653] Anonymous. Scientific American building edition. Scientific American, 64(17):266, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
266b.pdf.

Anonymous:1891:BNBf

[1654] Anonymous. Books: New books and publications. Scientific American, 64(17):266, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
266a.pdf.

Anonymous:1891:RPIq

[1655] Anonymous. Recently patented inventions. Scientific American, 64(17):266, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
266.pdf.

Anonymous:1891:BPp

[1656] Anonymous. Business and personal. Scientific American, 64(17):267, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-267.pdf.

Anonymous:1891:Ir

[1657] Anonymous. To inventors. Scientific American, 64(17):267, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-267b.pdf.

Anonymous:1891:NQu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-267a.pdf.

Anonymous:1891:IIs

[1659] Anonymous. Index of inventions. Scientific American, 64(17):267–268, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-267c.pdf.

Anonymous:1891:LBEe

[1660] Anonymous. List of books on electricity. Scientific American, 64(17):268, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
268.pdf.

Anonymous:1891:Pe

[1661] Anonymous. Patents! Scientific American, 64(17):269, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-269.pdf.

Anonymous:1891:UBe

[1662] Anonymous. Useful books! Scientific American, 64(17):269, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-269a.pdf.

Hopkins:1891:NBE

[1663] Geo. M. Hopkins. The new book experimental science. Scientific American, 64(17):269, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n17/pdf/scientificamerican04251891-
269b.pdf.

Anonymous:1891:SAl

[1664] Anonymous. The Scientific American. Scientific American, 64(17):270, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-270.pdf.

Anonymous:1891:BEd

[1665] Anonymous. Building edition. Scientific American, 64(17):270, April 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n17/pdf/scientificamerican04251891-270a.pdf.

Anonymous:1891:VB

[1666] Anonymous. Victor bicycles. Scientific American, 64(18):271, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-271.pdf.

Anonymous:1891:REI

[1667] Anonymous. Rights of employers to inventions made by their employes. Scientific American, 64(18):272, May 2, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n18/pdf/scientificamerican05021891-272.pdf.

Anonymous:1891:PPMb

[1668] Anonymous. Position of the planets in May. Scientific American, 64(18):272–273, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
272a.pdf.

Anonymous:1891:DRP

[1669] Anonymous. Decisions relating to patents. Scientific American, 64(18):273, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
273.pdf.

Anonymous:1891:MOS

[1670] Anonymous. McBride’s Observatory sleeper. Scientific American, 64(18):274, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
274a.pdf.

Anonymous:1891:PMVc

[1671] Anonymous. Proctor’s measuring vessel. Scientific American, 64 (18):274, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
274.pdf.

Anonymous:1891:PSM

[1672] Anonymous. Prouty’s saw-mill dog. Scientific American, 64(18):274, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-274b.pdf.

Anonymous:1891:BBB

[1673] Anonymous. Boomer & Boschert baling presses. Scientific American, 64(18):275, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
275.pdf.

Anonymous:1891:RRS

[1674] Anonymous. Radley’s removable sleigh-runner. Scientific American, 64(18):275, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
275a.pdf.

Anonymous:1891:FHS

[1675] Anonymous. The “Florida” heater for soft coal. Scientific American, 64(18):275–276, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
275b.pdf.

Anonymous:1891:DFS

[1676] Anonymous. Desk fan or small dental motor. Scientific American, 64(18):276, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
276a.pdf.

Anonymous:1891:IVa

[1677] Anonymous. An improved velocipede. Scientific American, 64(18):276, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-276.pdf.

Anonymous:1891:Cn

[1678] Anonymous. Correspondence. Scientific American, 64(18):277, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-277.pdf.

Anonymous:1891:SD

[1679] Anonymous. Scientific diversions. Scientific American, 64(18):279, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-279.pdf.

Anonymous:1891:MMR

[1680] Anonymous. The Mauser magazine rifle. Scientific American, 64(18):279–280, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
279a.pdf.

Hopkins:1891:BED

[1681] M. Hopkins. Beautiful example of diffraction. Scientific American, 64(18):281, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
281.pdf.

Anonymous:1891:BPq

[1682] Anonymous. Business and personal. Scientific American, 64(18):282, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-282a.pdf.

Anonymous:1891:RPIr

[1683] Anonymous. Recently patented inventions. Scientific American, 64(18):282, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n18/pdf/scientificamerican05021891-
282.pdf.

Anonymous:1891:NQv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-282b.pdf.

Anonymous:1891:Is

[1685] Anonymous. To inventors. Scientific American, 64(18):283, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-283.pdf.

Anonymous:1891:IIt

[1686] Anonymous. Index of inventions. Scientific American, 64(18):283, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-283a.pdf.

Anonymous:1891:PNPo

[1687] Anonymous. Patent notes: Patents. Scientific American, 64(18):285, May 2, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n18/pdf/scientificamerican05021891-285.pdf.

Anonymous:1891:EAP

[1688] Anonymous. Experiments for the artificial production of rain. Scientific American, 64(19):288, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
288a.pdf.

Anonymous:1891:OUP

[1689] Anonymous. Our urban population. Scientific American, 64(19):288, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-288.pdf.

Anonymous:1891:PEN

[1690] Anonymous. Photo-engravings for newspapers. Scientific American, 64(19):288–290, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
288b.pdf.

Anonymous:1891:OPC

[1691] Anonymous. An original portrait of Columbus. Scientific American, 64(19):290, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
290.pdf.

Anonymous:1891:CHM

[1692] Anonymous. A cooler for holding milk cans. Scientific American, 64(19):291, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
291b.pdf.

Anonymous:1891:IRC

[1693] Anonymous. An improved rainwater cut-off. Scientific American, 64(19):291, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
291.pdf.

Anonymous:1891:TBC

[1694] Anonymous. A toy bird for children. Scientific American, 64(19):291, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-291a.pdf.

Anonymous:1891:Co

[1695] Anonymous. Correspondence. Scientific American, 64(19):292, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-292.pdf.

Anonymous:1891:LSAa

[1696] Anonymous. Large shipment of American locomotives to Australia. Scientific American, 64(19):293, May 9, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n19/pdf/scientificamerican05091891-293.pdf.

Anonymous:1891:BDA

[1697] Anonymous. The Bennett dredge and amalgamator. Scientific American, 64(19):294, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
294.pdf.

Hovey:1891:HGM

[1698] H. C. Hovey. A hanging garden and a model office building. Scientific American, 64(19):295, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
295.pdf.

Anonymous:1891:FPO

[1699] Anonymous. The father of the Patent Office. Scientific American, 64(19):295–297, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
295a.pdf.

Anonymous:1891:BLS

[1700] Anonymous. Bath lift for the sick and paralyzed. Scientific American, 64(19):297, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
297.pdf.

Anonymous:1891:BPr

[1701] Anonymous. Business and personal. Scientific American, 64(19):298, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-298a.pdf.

Anonymous:1891:RPIs

[1702] Anonymous. Recently patented inventions. Scientific American, 64(19):298, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n19/pdf/scientificamerican05091891-
298.pdf.

Anonymous:1891:NQw

[1703] Anonymous. Notes and queries. Scientific American, 64(19):298–299, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-298b.pdf.

Anonymous:1891:Iu

[1704] Anonymous. To inventors. Scientific American, 64(19):299, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-299.pdf.

Anonymous:1891:IIu

[1705] Anonymous. Index of inventions. Scientific American, 64(19):299–300, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-299a.pdf.

Anonymous:1891:PNPp

[1706] Anonymous. Patent notes: Patents. Scientific American, 64(19):301, May 9, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n19/pdf/scientificamerican05091891-301.pdf.

Anonymous:1891:BSAa

[1707] Anonymous. The Broadway and Seventh Avenue Cable Road. Scientific American, 64(20):303, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
303.pdf.

Anonymous:1891:CPa

[1708] Anonymous. Charles Pratt. Scientific American, 64(20):304, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-304a.pdf.

Anonymous:1891:DCO

[1709] Anonymous. Defenseless condition of our seaports. Scientific American, 64(20):304, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
304.pdf.

Anonymous:1891:DEM

[1710] Anonymous. Dr. Edward Maynard. Scientific American, 64(20): 304–305, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
304b.pdf.

Anonymous:1891:IJS

[1711] Anonymous. Improvement in jack-screws. Scientific American, 64 (20):306, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
306a.pdf.

Anonymous:1891:NMC

[1712] Anonymous. New method of chamfering Stone. Scientific American, 64(20):306, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
306.pdf.

Anonymous:1891:OCM

[1713] Anonymous. Open column manometer on the Eiffel tower. Scientific American, 64(20):306, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
306b.pdf.

Anonymous:1891:IWc

[1714] Anonymous. An improved wrench. Scientific American, 64(20):306–308, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-306c.pdf.

Anonymous:1891:Cq

[1715] Anonymous. Correspondence. Scientific American, 64(20):309, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-309.pdf.

Anonymous:1891:AMD

[1716] Anonymous. Ancient methods of drawing water. Scientific American, 64(20):311–313, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
311.pdf.

Anonymous:1891:Cr

[1717] Anonymous. Castrography. Scientific American, 64(20):313, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-313.pdf.

Anonymous:1891:RPIt

[1718] Anonymous. Recently patented inventions. Scientific American, 64(20):314, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
314.pdf.

Anonymous:1891:BNBg

[1719] Anonymous. Books: New books and publications. Scientific American, 64(20):315, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n20/pdf/scientificamerican05161891-
315b.pdf.

Anonymous:1891:BPs

[1720] Anonymous. Business and personal. Scientific American, 64(20):315, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-315.pdf.

Anonymous:1891:NQx

[1721] Anonymous. Notes and queries. Scientific American, 64(20):315, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-315a.pdf.

Anonymous:1891:IIv

[1722] Anonymous. Index of inventions. Scientific American, 64(20):315–316, May 16, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-315c.pdf.

Anonymous:1891:PNPq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n20/pdf/scientificamerican05161891-317.pdf.

Anonymous:1891:PHR

[1724] Anonymous. The proposed Hudson River suspension bridge. Scientific American, 64(21):319, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
319.pdf.

Anonymous:1891:OWAb

[1725] Anonymous. Original work in America. Scientific American, 64(21):320, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-320.pdf.

Anonymous:1891:PSEa

[1726] Anonymous. Progress of steam engineering. Scientific American, 64(21):320–321, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
320a.pdf.

Anonymous:1891:DFBa

[1727] Anonymous. A drill to facilitate blasting. Scientific American, 64(21):322, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
322.pdf.

Anonymous:1891:IFC

[1728] Anonymous. An improved freight car coupler. Scientific American, 64(21):322, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
322c.pdf.

Anonymous:1891:IGV

[1729] Anonymous. An improved gate valve. Scientific American, 64(21):322, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-322b.pdf.

Anonymous:1891:LSAb

[1730] Anonymous. A locomotive Spark arrester. Scientific American, 64(21):322, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
322a.pdf.

Anonymous:1891:Cs

[1731] Anonymous. Correspondence. Scientific American, 64(21):325, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-325.pdf.

Anonymous:1891:BWL

[1732] Anonymous. Breast wheel for lifting water into the Morris Canal. Scientific American, 64(21):326, May 23, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n21/pdf/scientificamerican05231891-326a.pdf.

Anonymous:1891:IOW

[1733] Anonymous. An improved Ore washer. Scientific American, 64(21):326, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-326.pdf.

Anonymous:1891:SAU

[1734] Anonymous. Some apparatus used by the Vienna Fire Department. Scientific American, 64(21):327–328, May 23, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n21/pdf/scientificamerican05231891-327.pdf.

Anonymous:1891:TWF

[1735] Anonymous. The type-writer a French invention. Scientific American, 64(21):328–329, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
328.pdf.

Hopkins:1891:IG

[1736] M. Hopkins. Iridescent glass. Scientific American, 64(21):329, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-329.pdf.

Anonymous:1891:BPt

[1737] Anonymous. Business and personal. Scientific American, 64(21):330, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-330a.pdf.

Anonymous:1891:NQy

[1738] Anonymous. Notes and queries. Scientific American, 64(21):330, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-330b.pdf.

Anonymous:1891:RPIu

[1739] Anonymous. Recently patented inventions. Scientific American, 64(21):330, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
330.pdf.

Anonymous:1891:BNBh

[1740] Anonymous. Books: New books and publications. Scientific American, 64(21):331, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n21/pdf/scientificamerican05231891-
331.pdf.

Anonymous:1891:Ix

[1741] Anonymous. To inventors. Scientific American, 64(21):331, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-331a.pdf.

Anonymous:1891:IIw

[1742] Anonymous. Index of inventions. Scientific American, 64(21):331–332, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-331b.pdf.

Anonymous:1891:PNPr

[1743] Anonymous. Patent notes: Patents. Scientific American, 64(21):332, May 23, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n21/pdf/scientificamerican05231891-332.pdf.

Anonymous:1891:NSFb

[1744] Anonymous. The new steamship Fürst Bismarck. Scientific American, 64(22):335, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
335.pdf.

Anonymous:1891:TTC

[1745] Anonymous. Tests of the torpedo cruiser Vesuvius. Scientific American, 64(22):336, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
336.pdf.

Anonymous:1891:PPJa

[1746] Anonymous. Position of the planets in June. Scientific American, 64(22):337, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
337.pdf.

Anonymous:1891:AAR

[1747] Anonymous. An automatic adding and recording machine. Scientific American, 64(22):338, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
338a.pdf.

Anonymous:1891:IMP

[1748] Anonymous. An improved metallic piston packing. Scientific American, 64(22):338, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
338.pdf.

Anonymous:1891:IMR

[1749] Anonymous. An improved metallic railroad tie. Scientific American, 64(22):338, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
338b.pdf.

Anonymous:1891:IDC

[1750] Anonymous. An improved dumping cart. Scientific American, 64 (22):339, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
339.pdf.

Anonymous:1891:ACG

[1751] Anonymous. Atkinson’s cycle gas engine. Scientific American, 64 (22):339–340, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
339a.pdf.

Anonymous:1891:Cu

[1752] Anonymous. Correspondence. Scientific American, 64(22):341, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n22/pdf/scientificamerican05301891-341.pdf.

Peirce:1891:YPA

[1753] H. H. D. Peirce. The Yaryan process and apparatus for evaporation of liquids. Scientific American, 64(22):342–343, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n22/pdf/scientificamerican05301891-342.pdf.

Anonymous:1891:RRA

[1754] Anonymous. A remarkable railroad accident. Scientific American, 64(22):343–345, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
343.pdf.

Anonymous:1891:NLE

[1755] Anonymous. New lantern effect. Scientific American, 64(22):345–346, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n22/pdf/scientificamerican05301891-345.pdf.

Anonymous:1891:RPIv

[1756] Anonymous. Recently patented inventions. Scientific American, 64(22):346–347, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n22/pdf/scientificamerican05301891-
346.pdf.

Anonymous:1891:BPu

[1757] Anonymous. Business and personal. Scientific American, 64(22):347, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n22/pdf/scientificamerican05301891-347.pdf.

Anonymous:1891:Iy

[1758] Anonymous. To inventors. Scientific American, 64(22):347, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n22/pdf/scientificamerican05301891-347a.pdf.

Anonymous:1891:IIx

[1759] Anonymous. Index of inventions. Scientific American, 64(22):347–348, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n22/pdf/scientificamerican05301891-347b.pdf.

Anonymous:1891:PNPs

[1760] Anonymous. Patent notes: Patents. Scientific American, 64(22):349, May 30, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n22/pdf/scientificamerican05301891-349.pdf.

Anonymous:1891:ACA

[1761] Anonymous. Annual convention of the American Society of Civil Engineers. Scientific American, 64(23):352–353, June 6, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n23/pdf/scientificamerican06061891-352.pdf.

Anonymous:1891:DCT

[1762] Anonymous. Domesticating camphor trees in the United States. Scientific American, 64(23):353, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
353.pdf.

Anonymous:1891:WDSa

[1763] Anonymous. Why does solid iron float on molten iron? Scientific American, 64(23):353, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
353a.pdf.

Anonymous:1891:CRB

[1764] Anonymous. A car receiver for bills of lading, etc. Scientific American, 64(23):354, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
354a.pdf.

Anonymous:1891:DHBa

[1765] Anonymous. A device to hold bags open for filling. Scientific American, 64(23):354, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
354b.pdf.

Anonymous:1891:NEW

[1766] Anonymous. New emery wheel dresser. Scientific American, 64(23):354, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-354.pdf.

Anonymous:1891:AWSa

[1767] Anonymous. Aluminum works in Switzerland. Scientific American, 64(23):355, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
355.pdf.

Anonymous:1891:IWG

[1768] Anonymous. An improved water gauge. Scientific American, 64 (23):355, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
355a.pdf.

Anonymous:1891:PNe

[1769] Anonymous. Photographic notes. Scientific American, 64(23):356, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-356.pdf.

Anonymous:1891:Cv

[1770] Anonymous. Correspondence. Scientific American, 64(23):357, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-357a.pdf.

Anonymous:1891:GTC

[1771] Anonymous. Great tower of the City Hall, Philadelphia — total height, 547 feet. Scientific American, 64(23):357, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-357.pdf.

Anonymous:1891:UWC

[1772] Anonymous. The U.S. warship Charleston pursues a Chilean rebel ship. Scientific American, 64(23):358, June 6, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n23/pdf/scientificamerican06061891-358.pdf.

Anonymous:1891:RNC

[1773] Anonymous. Remarkable Naval combat — a steel cruiser sunk by torpedo boats. Scientific American, 64(23):359–360, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-359.pdf.

Anonymous:1891:NTA

[1774] Anonymous. New taper attachment for lathes. Scientific American, 64(23):360–361, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
360.pdf.

Anonymous:1891:MBS

[1775] Anonymous. Mending a big shaft at sea. Scientific American, 64(23):361, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
361.pdf.

Anonymous:1891:BPv

[1776] Anonymous. Business and personal. Scientific American, 64(23):362, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-362a.pdf.

Anonymous:1891:RPIw

[1777] Anonymous. Recently patented inventions. Scientific American, 64(23):362, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n23/pdf/scientificamerican06061891-
362.pdf.

Anonymous:1891:NQz

[1778] Anonymous. Notes and queries. Scientific American, 64(23):362–363, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-362b.pdf.

Anonymous:1891:Iz

[1779] Anonymous. To inventors. Scientific American, 64(23):363, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-363.pdf.

Anonymous:1891:IIy

[1780] Anonymous. Index of inventions. Scientific American, 64(23):363–364, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-363a.pdf.

Anonymous:1891:PNPt

[1781] Anonymous. Patent notes: Patents. Scientific American, 64(23):364, June 6, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n23/pdf/scientificamerican06061891-364.pdf.

Anonymous:1891:APE

[1782] Anonymous. Another polar expedition. Scientific American, 64(24):368, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-368a.pdf.

Anonymous:1891:MFM

[1783] Anonymous. Maxim’s flying machine. Scientific American, 64(24):368, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-368.pdf.

Anonymous:1891:AWM

[1784] Anonymous. The “arrow weed” and Mexican “jumping beans”. Scientific American, 64(24):369, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
369.pdf.

Anonymous:1891:MBO

[1785] Anonymous. The manufacture of Birch oil. Scientific American, 64(24):369–370, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
369a.pdf.

Anonymous:1891:DHBb

[1786] Anonymous. A device to hold books in order. Scientific American, 64(24):370, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
370.pdf.

Anonymous:1891:SBA

[1787] Anonymous. A singular bridge accident. Scientific American, 64 (24):370, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
370a.pdf.

Anonymous:1891:ACBa

[1788] Anonymous. An automatic car brake. Scientific American, 64 (24):370–371, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
370b.pdf.

Anonymous:1891:CBa

[1789] Anonymous. Coalbrookdale bridge. Scientific American, 64(24):371, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-371.pdf.

Anonymous:1891:Cw

[1790] Anonymous. Correspondence. Scientific American, 64(24):372, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-372.pdf.

Anonymous:1891:EEI

[1791] Anonymous. The Edison Electric Illuminating Co.’s Central Station in Brooklyn, N.Y. Scientific American, 64(24):373, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-373.pdf.

Anonymous:1891:WFB

[1792] Anonymous. The World’s Fair buildings at Chicago. Scientific American, 64(24):374–375, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
374.pdf.

Anonymous:1891:TSB

[1793] Anonymous. The train staff block system. Scientific American, 64(24):375–376, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
375.pdf.

Anonymous:1891:IBM

[1794] Anonymous. An improved barrel making machine. Scientific American, 64(24):376–377, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
376.pdf.

Anonymous:1891:IOC

[1795] Anonymous. An improved Ore concentrator. Scientific American, 64(24):377, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
377.pdf.

Anonymous:1891:SAT

[1796] Anonymous. Simple aerial top. Scientific American, 64(24):377, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-377b.pdf.

Anonymous:1891:SEN

[1797] Anonymous. A simple and effective nut lock. Scientific American, 64(24):377, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
377a.pdf.

Anonymous:1891:BPw

[1798] Anonymous. Business and personal. Scientific American, 64(24):378, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-378a.pdf.

Anonymous:1891:RPIx

[1799] Anonymous. Recently patented inventions. Scientific American, 64(24):378, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
378.pdf.

Anonymous:1891:NQba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-378b.pdf.

Anonymous:1891:BNBi

[1801] Anonymous. Books: New books and publications. Scientific American, 64(24):379, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n24/pdf/scientificamerican06131891-
379.pdf.

Anonymous:1891:Iba

[1802] Anonymous. To inventors. Scientific American, 64(24):379, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-379a.pdf.

Anonymous:1891:IIz

[1803] Anonymous. Index of inventions. Scientific American, 64(24):379, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-379b.pdf.

Anonymous:1891:PNPu

[1804] Anonymous. Patent notes: Patents. Scientific American, 64(24):380, June 13, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n24/pdf/scientificamerican06131891-380.pdf.

Anonymous:1891:PNCa

[1805] Anonymous. The proposed new Croton Lake Dam and the new Croton Lake. Scientific American, 64(25):383, June 20, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n25/pdf/scientificamerican06201891-383.pdf.

Anonymous:1891:STP

[1806] Anonymous. The sailing of two polar expeditions. Scientific American, 64(25):384–385, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
384.pdf.

Anonymous:1891:EBT

[1807] Anonymous. An extensible barrel truck. Scientific American, 64 (25):386, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
386b.pdf.

Anonymous:1891:IFR

[1808] Anonymous. An improved feed rack, barn, etc. Scientific American, 64(25):386, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
386.pdf.

Anonymous:1891:LKH

[1809] Anonymous. A lock with key held pendent. Scientific American, 64(25):386, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
386a.pdf.

Anonymous:1891:IWS

[1810] Anonymous. Improved Wimshurst static electrical machine. Scientific American, 64(25):387, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
387.pdf.

Anonymous:1891:Cx

[1811] Anonymous. Correspondence. Scientific American, 64(25):388, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n25/pdf/scientificamerican06201891-388.pdf.

Anonymous:1891:ICCa

[1812] Anonymous. An improved car coupling. Scientific American, 64 (25):388, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
388a.pdf.

Anonymous:1891:AWSb

[1813] Anonymous. The artesian well at Springfield, South Dakota. Scientific American, 64(25):390, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
390.pdf.

Hopkins:1891:SAG

[1814] M. Hopkins. Simple apparatus for gathering and examining microscopic objects. Scientific American, 64(25):391, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v64/
n25/pdf/scientificamerican06201891-391.pdf.

Anonymous:1891:SCTa

[1815] Anonymous. The St. Clair Tunnel locomotives — the largest locomotive in the world. Scientific American, 64(25):391–392, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n25/pdf/scientificamerican06201891-391a.pdf.

Anonymous:1891:BFL

[1816] Anonymous. A book Finisher’s lettering machine. Scientific American, 64(25):392, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
392.pdf.

Anonymous:1891:HLG

[1817] Anonymous. The Hitchcock life guard for electric cars. Scientific American, 64(25):392, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
392a.pdf.

Anonymous:1891:EKC

[1818] Anonymous. Edison’s kinetograph and cosmical telephone. Scientific American, 64(25):393, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
393.pdf.

Anonymous:1891:BPx

[1819] Anonymous. Business and personal. Scientific American, 64(25):394, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n25/pdf/scientificamerican06201891-394a.pdf.

Anonymous:1891:RPIy

[1820] Anonymous. Recently patented inventions. Scientific American, 64(25):394, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n25/pdf/scientificamerican06201891-
394.pdf.

Anonymous:1891:NQbb

[1821] Anonymous. Notes and queries. Scientific American, 64(25):394–395, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n25/pdf/scientificamerican06201891-394b.pdf.

Anonymous:1891:Ibb

[1822] Anonymous. To inventors. Scientific American, 64(25):395, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n25/pdf/scientificamerican06201891-395.pdf.

Anonymous:1891:IIba

[1823] Anonymous. Index of inventions. Scientific American, 64(25):395–396, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n25/pdf/scientificamerican06201891-395a.pdf.

Anonymous:1891:PNPv

[1824] Anonymous. Patent notes: Patents. Scientific American, 64(25):397, June 20, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n25/pdf/scientificamerican06201891-397.pdf.

Anonymous:1891:CHW

[1825] Anonymous. Casting the Henry Ward Beecher Statue for the City of Brooklyn. Scientific American, 64(26):399–400, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-399.pdf.

Anonymous:1891:Ad

[1826] Anonymous. Arcturus. Scientific American, 64(26):400–401, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-400.pdf.

Anonymous:1891:PPJb

[1827] Anonymous. Position of the planets in July. Scientific American, 64(26):401, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
401.pdf.

Anonymous:1891:PSF

[1828] Anonymous. The poisonous snake of Florida. Scientific American, 64(26):401, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
401a.pdf.

Anonymous:1891:MTR

[1829] Anonymous. A metallic tie and rail fastening. Scientific American, 64(26):402, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
402.pdf.

Anonymous:1891:RR

[1830] Anonymous. Rifle range. Scientific American, 64(26):402, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-402a.pdf.

Anonymous:1891:DGB

[1831] Anonymous. A differential gear for bicycles. Scientific American, 64(26):403, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
403a.pdf.

Anonymous:1891:ICB

[1832] Anonymous. An improved car brake. Scientific American, 64(26):403, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-403.pdf.

Anonymous:1891:MEM

[1833] Anonymous. A miniature electric motor. Scientific American, 64 (26):403, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
403b.pdf.

Anonymous:1891:VFA

[1834] Anonymous. A visit to a family in annam. Scientific American, 64(26):403, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
403c.pdf.

Anonymous:1891:IEb

[1835] Anonymous. An improved engine. Scientific American, 64(26):403–404, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-403d.pdf.

Anonymous:1891:AGS

[1836] Anonymous. The assaying of gold and silver ores. Scientific American, 64(26):404, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
404.pdf.

Anonymous:1891:CWF

[1837] Anonymous. The Columbian World’s Fair, 1893. Scientific American, 64(26):405–408, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
405.pdf.

Anonymous:1891:RPIz

[1838] Anonymous. Recently patented inventions. Scientific American, 64(26):408–409, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v64/n26/pdf/scientificamerican06271891-
408.pdf.

Anonymous:1891:BPy

[1839] Anonymous. Business and personal. Scientific American, 64(26):409, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-409.pdf.

Anonymous:1891:Ibc

[1840] Anonymous. To inventors. Scientific American, 64(26):409, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-409b.pdf.

Anonymous:1891:NQbc

[1841] Anonymous. Notes and queries. Scientific American, 64(26):409, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-409a.pdf.

Anonymous:1891:IIbb

[1842] Anonymous. Index of inventions. Scientific American, 64(26):409–410, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-409c.pdf.

Anonymous:1891:PNPw

[1843] Anonymous. Patent notes: Patents. Scientific American, 64(26):410, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-410.pdf.

Anonymous:1891:Ibd

[1844] Anonymous. Index. Scientific American, 64(26):411–412, June 27, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v64/n26/pdf/scientificamerican06271891-411.pdf.

Anonymous:1891:MWS

[1845] Anonymous. The McDougall whaleback steel vessel. Scientific American, 65(1):1–2, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
1.pdf.

Anonymous:1891:CPE

[1846] Anonymous. The Casino and Pier at the Exhibition. Scientific American, 65(1):2, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
2.pdf.

Anonymous:1891:FBN

[1847] Anonymous. Fast boats for the Navy. Scientific American, 65(1):2–4, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-2a.pdf.

Anonymous:1891:GEB

[1848] Anonymous. Guiley’s electrical block system for railways. Scientific American, 65(1):4, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
4.pdf.

Anonymous:1891:LGC

[1849] Anonymous. Latham’s grass collector for lawn mowers. Scientific American, 65(1):5, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
5.pdf.

Hopkins:1891:SSM

[1850] M. Hopkins. Some suggestions in microscopy. Scientific American, 65(1):5, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
5a.pdf.

Anonymous:1891:ODP

[1851] Anonymous. A one dollar photographic outfit. Scientific American, 65(1):5–6, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
5b.pdf.

Anonymous:1891:RAT

[1852] Anonymous. Rack attachment for theater chairs. Scientific American, 65(1):6, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
6.pdf.

Anonymous:1891:WRW

[1853] Anonymous. A warship rams a whale. Scientific American, 65(1):6, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-6a.pdf.

Anonymous:1891:GAR

[1854] Anonymous. Gunner’s arm rest. Scientific American, 65(1):6–8, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-6b.pdf.

Anonymous:1891:DFBb

[1855] Anonymous. Drawing of financial bills by the Casanova Apparatus. Scientific American, 65(1):8–9, July 4, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n1/pdf/scientificamerican07041891-8.pdf.

Anonymous:1891:GSW

[1856] Anonymous. Great spools of wire. Scientific American, 65(1):9, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-9.pdf.

Anonymous:1891:Cy

[1857] Anonymous. Correspondence. Scientific American, 65(1):10, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-10.pdf.

Anonymous:1891:BNBj

[1858] Anonymous. Books: New books and publications. Scientific American, 65(1):11, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n1/pdf/scientificamerican07041891-
11c.pdf.

Anonymous:1891:BPz

[1859] Anonymous. Business and personal. Scientific American, 65(1):11, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-11a.pdf.

Anonymous:1891:NQbd

[1860] Anonymous. Notes and queries. Scientific American, 65(1):11, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-11b.pdf.

Anonymous:1891:RPIba

[1861] Anonymous. Recently patented inventions. Scientific American, 65(1): 11, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-11.pdf.

Anonymous:1891:Ibe

[1862] Anonymous. To inventors. Scientific American, 65(1):12, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-12.pdf.

Anonymous:1891:IIbc

[1863] Anonymous. Index of inventions. Scientific American, 65(1):12, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-12a.pdf.

Anonymous:1891:PNPx

[1864] Anonymous. Patent notes: Patents. Scientific American, 65(1):14, July 4, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n1/pdf/scientificamerican07041891-14.pdf.

Anonymous:1891:IMB

[1865] Anonymous. Improved methods of building. Scientific American, 65(2):15–16, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
15.pdf.

Anonymous:1891:ICGb

[1866] Anonymous. International Congress of geologists. Scientific American, 65(2):16, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
16.pdf.

Anonymous:1891:SI

[1867] Anonymous. Steamship improvements. Scientific American, 65 (2):16–18, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
16a.pdf.

Anonymous:1891:EPa

[1868] Anonymous. Electrical pump. Scientific American, 65(2):18, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-18.pdf.

Anonymous:1891:MDI

[1869] Anonymous. A measuring and drawing implement. Scientific American, 65(2):18, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
18a.pdf.

Anonymous:1891:ISSa

[1870] Anonymous. Improved slide and screw-cutting lathe. Scientific American, 65(2):19–20, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
19.pdf.

Anonymous:1891:BF

[1871] Anonymous. A bedclothes fastener. Scientific American, 65(2):20, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-20.pdf.

Anonymous:1891:DTB

[1872] Anonymous. A device for tapping barrels. Scientific American, 65(2):20, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-20a.pdf.

Anonymous:1891:Cz

[1873] Anonymous. Correspondence. Scientific American, 65(2):21, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-21.pdf.

Anonymous:1891:RW

[1874] Anonymous. The rolling wheel. Scientific American, 65(2):21, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-21a.pdf.

Anonymous:1891:HBS

[1875] Anonymous. A hydrocarbon burner for stoves, etc. Scientific American, 65(2):22, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
22.pdf.

Anonymous:1891:SMDa

[1876] Anonymous. The system of military dovecotes in Europe. Scientific American, 65(2):22–23, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
22a.pdf.

Anonymous:1891:PSU

[1877] Anonymous. Pneumatic signaling upon the Central Railroad of New Jersey. Scientific American, 65(2):23–24, July 11, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n2/pdf/scientificamerican07111891-23.pdf.

Anonymous:1891:RRP

[1878] Anonymous. Residence on Riverside Park, New York. Scientific American, 65(2):25, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
25.pdf.

Anonymous:1891:BPbc

[1879] Anonymous. Business and personal. Scientific American, 65(2):26, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-26a.pdf.

Anonymous:1891:RPIbb

[1880] Anonymous. Recently patented inventions. Scientific American, 65(2):26, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
26.pdf.

Anonymous:1891:NQbe

[1881] Anonymous. Notes and queries. Scientific American, 65(2):26–27, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-26b.pdf.

Anonymous:1891:BNBk

[1882] Anonymous. Books: New books and publications. Scientific American, 65(2):27, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n2/pdf/scientificamerican07111891-
27.pdf.

Anonymous:1891:Ibf

[1883] Anonymous. To inventors. Scientific American, 65(2):27, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-27a.pdf.

Anonymous:1891:IIbd

[1884] Anonymous. Index of inventions. Scientific American, 65(2):27–28, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-27b.pdf.

Anonymous:1891:PNPy

[1885] Anonymous. Patent notes: Patents. Scientific American, 65(2):29, July 11, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n2/pdf/scientificamerican07111891-29.pdf.

Anonymous:1891:DCH

[1886] Anonymous. Diamond cutting by hand and machine. Scientific American, 65(3):31–32, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
31.pdf.

Anonymous:1891:CPb

[1887] Anonymous. The Commissionership of Patents. Scientific American, 65(3):32, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
32.pdf.

Anonymous:1891:CNF

[1888] Anonymous. The Caldera Naval Fight. Scientific American, 65 (3):32–33, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
32a.pdf.

Anonymous:1891:DMD

[1889] Anonymous. The disposition of mining debris in California. Scientific American, 65(3):33–34, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
33.pdf.

Anonymous:1891:PNR

[1890] Anonymous. Progress of the North River Tunnel. Scientific American, 65(3):34, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
34.pdf.

Anonymous:1891:HB

[1891] Anonymous. The Horticultural Building. Scientific American, 65(3):35, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-35a.pdf.

Anonymous:1891:IWd

[1892] Anonymous. An improved wrench. Scientific American, 65(3):35, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-35.pdf.

Anonymous:1891:IVb

[1893] Anonymous. An improved velocipede. Scientific American, 65(3):35–36, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-35b.pdf.

Anonymous:1891:PHS

[1894] Anonymous. A power hack saw. Scientific American, 65(3):36, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-36.pdf.

Anonymous:1891:Cbb

[1895] Anonymous. Correspondence. Scientific American, 65(3):37, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-37.pdf.

Anonymous:1891:PNf

[1896] Anonymous. Photographic notes. Scientific American, 65(3):37, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-37a.pdf.

Anonymous:1891:VRL

[1897] Anonymous. Valley’s railroad for lumbermen etc. Scientific American, 65(3):38, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
38.pdf.

Anonymous:1891:ST

[1898] Anonymous. Science in the theater. Scientific American, 65(3):39–40, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-39.pdf.

Anonymous:1891:PSEb

[1899] Anonymous. Permanent steam engine indicator. Scientific American, 65(3):41, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
41.pdf.

Anonymous:1891:RPIbc

[1900] Anonymous. Recently patented inventions. Scientific American, 65(3):42, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
42.pdf.

Anonymous:1891:BPbd

[1901] Anonymous. Business and personal. Scientific American, 65(3):43, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-43.pdf.

Anonymous:1891:Ibg

[1902] Anonymous. To inventors. Scientific American, 65(3):43, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-43c.pdf.

Anonymous:1891:NBPh

[1903] Anonymous. New books and publications. Scientific American, 65(3):43, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n3/pdf/scientificamerican07181891-
43b.pdf.

Anonymous:1891:NQbf

[1904] Anonymous. Notes and queries. Scientific American, 65(3):43, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-43a.pdf.

Anonymous:1891:IIbe

[1905] Anonymous. Index of inventions. Scientific American, 65(3):43–44, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-43d.pdf.

Anonymous:1891:PNPz

[1906] Anonymous. Patent notes: Patents! Scientific American, 65(3):45, July 18, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n3/pdf/scientificamerican07181891-45.pdf.

Anonymous:1891:EDMa

[1907] Anonymous. Edison dynamo and motor. Scientific American, 65(4):47, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-47.pdf.

Anonymous:1891:SPP

[1908] Anonymous. Salt packing — a $4,000 prize contest. Scientific American, 65(4):48, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
48.pdf.

Anonymous:1891:EEL

[1909] Anonymous. The Edison Electric Lamp Patent sustained. Scientific American, 65(4):48–49, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
48a.pdf.

Anonymous:1891:PNg

[1910] Anonymous. Photographic notes. Scientific American, 65(4):49–50, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-49.pdf.

Anonymous:1891:IMMa

[1911] Anonymous. An improved mortising machine. Scientific American, 65(4):50, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
50.pdf.

Anonymous:1891:IMMb

[1912] Anonymous. Improved mechanical mixer. Scientific American, 65(4):50, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-50a.pdf.

Anonymous:1891:WPD

[1913] Anonymous. A weight power to drive fans etc. Scientific American, 65(4):51, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
51.pdf.

Anonymous:1891:SMT

[1914] Anonymous. Supplying moving trains with water. Scientific American, 65(4):51–52, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
51a.pdf.

Hovey:1891:SMM

[1915] H. C. Hovey. Some measurements in the Mammoth Cave of Indiana. Scientific American, 65(4):52, July 25, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n4/pdf/scientificamerican07251891-52.pdf.

Anonymous:1891:Cbc

[1916] Anonymous. Correspondence. Scientific American, 65(4):53, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-53.pdf.

Anonymous:1891:EBH

[1917] Anonymous. Edward Burgess and his achievements. Scientific American, 65(4):55–56, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
55.pdf.

Anonymous:1891:LV

[1918] Anonymous. Locust visitations. Scientific American, 65(4):57, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-57.pdf.

Anonymous:1891:BPbe

[1919] Anonymous. Business and personal. Scientific American, 65(4):58, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-58b.pdf.

Anonymous:1891:NBPi

[1920] Anonymous. New books and publications. Scientific American, 65(4):58, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
58a.pdf.

Anonymous:1891:RPIbd

[1921] Anonymous. Recently patented inventions. Scientific American, 65(4):58, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n4/pdf/scientificamerican07251891-
58.pdf.

Anonymous:1891:NQbg

[1922] Anonymous. Notes and queries. Scientific American, 65(4):58–59, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-58c.pdf.

Anonymous:1891:Ibi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-59.pdf.

Anonymous:1891:IIbf

[1924] Anonymous. Index of inventions. Scientific American, 65(4):59–60, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-59a.pdf.

Anonymous:1891:PNPba

[1925] Anonymous. Patent notes: Patents! Scientific American, 65(4):61, July 25, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n4/pdf/scientificamerican07251891-61.pdf.

Anonymous:1891:FFC

[1926] Anonymous. Fibers and fiber culture. Scientific American, 65(5):64, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-64.pdf.

Anonymous:1891:PPAb

[1927] Anonymous. Position of the planets in August. Scientific American, 65(5):65, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n5/pdf/scientificamerican08011891-
65.pdf.

Anonymous:1891:EDMb

[1928] Anonymous. Edison dynamo and motor. Scientific American, 65(5):66–68, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n5/pdf/scientificamerican08011891-
66.pdf.

Anonymous:1891:LSF

[1929] Anonymous. A light and simple fence machine. Scientific American, 65(5):68, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n5/pdf/scientificamerican08011891-
68.pdf.

Anonymous:1891:RTT

[1930] Anonymous. A railway train time register. Scientific American, 65(5):68, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n5/pdf/scientificamerican08011891-
68a.pdf.

Anonymous:1891:Cbd

[1931] Anonymous. Correspondence. Scientific American, 65(5):69, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-69.pdf.

Anonymous:1891:USD

[1932] Anonymous. The United States drilling scow General Newton and the removal of Diamond Reef, New York Harbor. Scientific American, 65(5):70, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n5/pdf/scientificamerican08011891-
70.pdf.

Gribble:1891:FP

[1933] T. Graham Gribble. Fossil patents. Scientific American, 65(5):71–73, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-71.pdf.

Anonymous:1891:BPbf

[1934] Anonymous. Business and personal. Scientific American, 65(5):74, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-74b.pdf.

Anonymous:1891:PNPbb

[1935] Anonymous. Patent notes: Patents. Scientific American, 65(5):74, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-74a.pdf.

Anonymous:1891:RPIbe (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n5/pdf/scientificamerican08011891-
74.pdf.

Anonymous:1891:NQbh

[1937] Anonymous. Notes and queries. Scientific American, 65(5):74–75, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-74c.pdf.

Anonymous:1891:BNBl

[1938] Anonymous. Books: New books and publications. Scientific American, 65(5):75, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n5/pdf/scientificamerican08011891-
75.pdf.

Anonymous:1891:Ibj

[1939] Anonymous. To inventors. Scientific American, 65(5):75, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-75a.pdf.

Anonymous:1891:IIbg

[1940] Anonymous. Index of inventions. Scientific American, 65(5):75–76, August 1, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n5/pdf/scientificamerican08011891-75b.pdf.

Anonymous:1891:GCE

[1941] Anonymous. The Government collection of economic plants. Scientific American, 65(6):80, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
80.pdf.

Anonymous:1891:HS

[1942] Anonymous. Habits of snakes. Scientific American, 65(6):81, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-81.pdf.

Anonymous:1891:RPE

[1943] Anonymous. A rolling pin, etc., for bakers’ use. Scientific American, 65(6):82, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
82.pdf.

Anonymous:1891:SCTb

[1944] Anonymous. A simple and compact type writer. Scientific American, 65(6):82, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
82a.pdf.

Anonymous:1891:IFE

[1945] Anonymous. An improved fire escape. Scientific American, 65(6):82–83, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-82b.pdf.

Anonymous:1891:RCI

[1946] Anonymous. Railway collision Illinois Central Railway. Scientific American, 65(6):83, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
83.pdf.

Anonymous:1891:NWW

[1947] Anonymous. The Newark water works. Scientific American, 65(6):84, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-84.pdf.

Anonymous:1891:SEV

[1948] Anonymous. A Spring equalizer for vehicles. Scientific American, 65(6):84, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
84a.pdf.

Anonymous:1891:Cbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-85.pdf.

Anonymous:1891:GDB

[1950] Anonymous. The great dam at Beetaloo, Australia. Scientific American, 65(6):86–87, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
86.pdf.

OReilly:1891:CPS

[1951] G. R. O’Reilly. Characteristic poises of snakes. Scientific American, 65(6):87–89, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
87.pdf.

Anonymous:1891:BPbg

[1952] Anonymous. Business and personal. Scientific American, 65(6):90, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-90a.pdf.

Anonymous:1891:RPIbf

[1953] Anonymous. Recently patented inventions. Scientific American, 65(6):90, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
90.pdf.

Anonymous:1891:NQbi

[1954] Anonymous. Notes and queries. Scientific American, 65(6):90–91, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-90b.pdf.

Anonymous:1891:BNBm

[1955] Anonymous. Books: New books and publications. Scientific American, 65(6):91, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n6/pdf/scientificamerican08081891-
91.pdf.

Anonymous:1891:Ibk

[1956] Anonymous. To inventors. Scientific American, 65(6):91, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-91a.pdf.

Anonymous:1891:IIbh

[1957] Anonymous. Index of inventions. Scientific American, 65(6):91–92, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-91b.pdf.

Anonymous:1891:PNPbc

[1958] Anonymous. Patent notes: Patents. Scientific American, 65(6):94, August 8, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n6/pdf/scientificamerican08081891-94.pdf.

Anonymous:1891:GP

[1959] Anonymous. Gutta percha. Scientific American, 65(7):96–98, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n7/pdf/scientificamerican08151891-96.pdf.

Anonymous:1891:PEL

[1960] Anonymous. Portable electric lighting plant. Scientific American, 65(7):98, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
98.pdf.

Anonymous:1891:ICd

[1961] Anonymous. An improved chuck. Scientific American, 65(7):98–99, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n7/pdf/scientificamerican08151891-98a.pdf.

Anonymous:1891:ACW

[1962] Anonymous. An attachment to cool water in mains. Scientific American, 65(7):99, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
99.pdf.

Anonymous:1891:FWH

[1963] Anonymous. A feed water heater and purifier. Scientific American, 65(7):99–100, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
99a.pdf.

Anonymous:1891:ISO

[1964] Anonymous. Improved scissors for opening letters, etc. Scientific American, 65(7):100, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
100.pdf.

Anonymous:1891:EDS

[1965] Anonymous. Evolutions and drill of the N.Y. State Naval Reserve Militia. Scientific American, 65(7):101, August 15, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n7/pdf/scientificamerican08151891-101.pdf.

Anonymous:1891:NHS

[1966] Anonymous. New high service stand pipe, Jersey City Heights. Scientific American, 65(7):102, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
102.pdf.

Baskett:1891:MMT

[1967] James Newton Baskett. The Mexico, Mo tornado. Scientific American, 65(7):103–104, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
103.pdf.

Anonymous:1891:Cbf

[1968] Anonymous. Correspondence. Scientific American, 65(7):104, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n7/pdf/scientificamerican08151891-104.pdf.

Anonymous:1891:EMS

[1969] Anonymous. Electro-microscopic slide for testing the antiseptic power of electricity. Scientific American, 65(7):104, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n7/pdf/scientificamerican08151891-104a.pdf.

Anonymous:1891:GES

[1970] Anonymous. Gymnastic exercises with the stick. Scientific American, 65(7):105, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
105.pdf.

Anonymous:1891:ICCb

[1971] Anonymous. An improved car coupling. Scientific American, 65(7):105, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
105a.pdf.

Anonymous:1891:BPbh

[1972] Anonymous. Business and personal. Scientific American, 65(7):106, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n7/pdf/scientificamerican08151891-106a.pdf.

Anonymous:1891:RPIbg

[1973] Anonymous. Recently patented inventions. Scientific American, 65(7):106, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
106.pdf.

Anonymous:1891:NQbj

[1974] Anonymous. Notes and queries. Scientific American, 65(7):106– 107, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
106b.pdf.

Anonymous:1891:Ibl

[1975] Anonymous. To inventors. Scientific American, 65(7):107, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n7/pdf/scientificamerican08151891-107.pdf.

Anonymous:1891:IIbi

[1976] Anonymous. Index of inventions. Scientific American, 65(7): 107–108, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n7/pdf/scientificamerican08151891-
107a.pdf.

Anonymous:1891:PNPbd

[1977] Anonymous. Patent notes: Patents. Scientific American, 65(7):110, August 15, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n7/pdf/scientificamerican08151891-110.pdf.

Anonymous:1891:GJN

[1978] Anonymous. George Jones of the New York “Times”. Scientific American, 65(8):112, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
112a.pdf.

Anonymous:1891:SRB

[1979] Anonymous. Strains on railroad bridges. Scientific American, 65(8):112, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
112.pdf.

Anonymous:1891:DCV

[1980] Anonymous. Dr. C. V. Riley. Scientific American, 65(8):112–114, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n8/pdf/scientificamerican08221891-112b.pdf.

Anonymous:1891:BPSa

[1981] Anonymous. A boat propelling and steering device. Scientific American, 65(8):114, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
114.pdf.

Anonymous:1891:CPL

[1982] Anonymous. A combined plumb and level. Scientific American, 65(8):114–115, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
114a.pdf.

Anonymous:1891:IEc

[1983] Anonymous. Improvements in elevators. Scientific American, 65 (8):115, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
115a.pdf.

Seal:1891:BHS

[1984] W. M. P. Seal. The breeding habits of the skate. Scientific American, 65(8):115, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
115.pdf.

Anonymous:1891:TMS

[1985] Anonymous. Tests of magazine small arms. Scientific American, 65(8):116–117, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
116.pdf.

Anonymous:1891:ABP

[1986] Anonymous. Albert Benjamin Prescott. Scientific American, 65 (8):119–120, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
119.pdf.

Oreilly:1891:CSA

[1987] G. R. O’reilly. Color of snakes as affected by climatic influences. Scientific American, 65(8):121, August 22, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n8/pdf/scientificamerican08221891-121.pdf.

Anonymous:1891:BPbi

[1988] Anonymous. Business and personal. Scientific American, 65(8):122, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n8/pdf/scientificamerican08221891-122a.pdf.

Anonymous:1891:RPIbh

[1989] Anonymous. Recently patented inventions. Scientific American, 65(8):122, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
122.pdf.

Anonymous:1891:NQbk

[1990] Anonymous. Notes and queries. Scientific American, 65(8):122– 123, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
122b.pdf.

Anonymous:1891:Ibn

[1991] Anonymous. To inventors. Scientific American, 65(8):123, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n8/pdf/scientificamerican08221891-123.pdf.

Anonymous:1891:IIbj

[1992] Anonymous. Index of inventions. Scientific American, 65(8): 123–124, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n8/pdf/scientificamerican08221891-
123a.pdf.

Anonymous:1891:PNPbe

[1993] Anonymous. Patent notes: Patents. Scientific American, 65(8):126, August 22, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n8/pdf/scientificamerican08221891-126.pdf.

Anonymous:1891:WSS

[1994] Anonymous. White Star steamer Teutonic. Scientific American, 65(9):127, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
127.pdf.

Anonymous:1891:ADG

[1995] Anonymous. “Alleged Deceptions in German Steel Works” corrected. Scientific American, 65(9):128, August 29, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n9/pdf/scientificamerican08291891-128a.pdf.

Anonymous:1891:FTT

[1996] Anonymous. The fast trip of the teutonic. Scientific American, 65(9):128, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
128.pdf.

Anonymous:1891:AAA

[1997] Anonymous. The American Association for the Advancement of Science. Scientific American, 65(9):128–129, August 29, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n9/pdf/scientificamerican08291891-128b.pdf.

Anonymous:1891:PPS

[1998] Anonymous. Position of the planets in September. Scientific American, 65(9):129–130, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
129.pdf.

Anonymous:1891:INYa

[1999] Anonymous. An improved neck yoke. Scientific American, 65 (9):130, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
130a.pdf.

Anonymous:1891:SDW

[2000] Anonymous. A simple and durable wrench. Scientific American, 65(9):130, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
130.pdf.

Anonymous:1891:TMG

[2001] Anonymous. A trigger mechanism for guns. Scientific American, 65(9):130, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
130b.pdf.

Anonymous:1891:ICCc

[2002] Anonymous. An improvement in car construction. Scientific American, 65(9):131, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
131.pdf.

Anonymous:1891:IPa

[2003] Anonymous. An improved planter. Scientific American, 65(9):131, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-131a.pdf.

Anonymous:1891:CWR

[2004] Anonymous. A car wheel and rail to increase traction. Scientific American, 65(9):131–132, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
131b.pdf.

Anonymous:1891:PNh

[2005] Anonymous. Photographic notes. Scientific American, 65(9):132, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-132.pdf.

Anonymous:1891:Cbg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-133.pdf.

Anonymous:1891:BC

[2007] Anonymous. Boa constrictors. Scientific American, 65(9):137–138, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-137.pdf.

Anonymous:1891:RPIbi

[2008] Anonymous. Recently patented inventions. Scientific American, 65 (9):138–139, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
138.pdf.

Anonymous:1891:BPbj

[2009] Anonymous. Business and personal. Scientific American, 65(9):139, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-139.pdf.

Anonymous:1891:Ibo

[2010] Anonymous. To inventors. Scientific American, 65(9):139, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-139b.pdf.

Anonymous:1891:NQbl

[2011] Anonymous. Notes and queries. Scientific American, 65(9):139, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-139a.pdf.

Anonymous:1891:IIbk

[2012] Anonymous. Index of inventions. Scientific American, 65(9): 139–140, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n9/pdf/scientificamerican08291891-
139c.pdf.

Anonymous:1891:PNPbf

[2013] Anonymous. Patent notes: Patents. Scientific American, 65(9):142, August 29, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n9/pdf/scientificamerican08291891-142.pdf.

Anonymous:1891:APR

[2014] Anonymous. The artificial production of rain. Scientific American, 65(10):144–145, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
144.pdf.

Anonymous:1891:WDSb

[2015] Anonymous. The Weather Debating Society. Scientific American, 65 (10):145–146, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
145.pdf.

Anonymous:1891:BBL

[2016] Anonymous. A bit for boring large holes. Scientific American, 65(10):146, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
146a.pdf.

Anonymous:1891:SACa

[2017] Anonymous. A safety attachment for car brakes. Scientific American, 65(10):146, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
146.pdf.

Anonymous:1891:SCP

[2018] Anonymous. A simple clip for papers, documents, etc. Scientific American, 65(10):146–147, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
146b.pdf.

Anonymous:1891:HAB

[2019] Anonymous. Hot air ballooning, Weehawken, N.J. Scientific American, 65(10):147–149, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
147.pdf.

Anonymous:1891:WDF

[2020] Anonymous. A water deodorizing flume. Scientific American, 65 (10):149, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
149.pdf.

Anonymous:1891:NFL

[2021] Anonymous. New flash light and fog bell at Coney Island Point. Scientific American, 65(10):150, September 5, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n10/pdf/scientificamerican09051891-150.pdf.

Anonymous:1891:RTC

[2022] Anonymous. Refrigeration for towns and cities by street mains. Scientific American, 65(10):151, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
151.pdf.

Anonymous:1891:Cbh

[2023] Anonymous. Correspondence. Scientific American, 65(10):152, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n10/pdf/scientificamerican09051891-152.pdf.

Anonymous:1891:FTS

[2024] Anonymous. A faithful and tireless servant. Scientific American, 65 (10):153–154, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
153.pdf.

Anonymous:1891:RPIbj

[2025] Anonymous. Recently patented inventions. Scientific American, 65(10):154, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
154.pdf.

Anonymous:1891:BPbk

[2026] Anonymous. Business and personal. Scientific American, 65 (10):155, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
155.pdf.

Anonymous:1891:Ibp

[2027] Anonymous. To inventors. Scientific American, 65(10):155, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n10/pdf/scientificamerican09051891-155b.pdf.

Anonymous:1891:NQbm

[2028] Anonymous. Notes and queries. Scientific American, 65(10): 155, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
155a.pdf.

Anonymous:1891:IIbl

[2029] Anonymous. Index of inventions. Scientific American, 65(10): 155–156, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
155c.pdf.

Anonymous:1891:PNPbg

[2030] Anonymous. Patent notes: Patents. Scientific American, 65 (10):158, September 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n10/pdf/scientificamerican09051891-
158.pdf.

Anonymous:1891:NTF

[2031] Anonymous. A new type of folding bridge. Scientific American, 65(11):159, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
159.pdf.

Anonymous:1891:ACS

[2032] Anonymous. The American chemical society. Scientific American, 65(11):160, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
160b.pdf.

Anonymous:1891:OGO

[2033] Anonymous. An official geological organization. Scientific American, 65(11):160, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
160a.pdf.

Anonymous:1891:TOG

[2034] Anonymous. Testing of our great guns. Scientific American, 65 (11):160, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
160.pdf.

Anonymous:1891:CWM

[2035] Anonymous. Condition of workingmen in Michigan. Scientific American, 65(11):161, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
161.pdf.

Anonymous:1891:DCW

[2036] Anonymous. A device for cleaning wood saws. Scientific American, 65(11):162, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
162d.pdf.

Anonymous:1891:DTI

[2037] Anonymous. A device for thawing ice from pipes. Scientific American, 65(11):162, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
162c.pdf.

Anonymous:1891:ICCd

[2038] Anonymous. An improved car coupling. Scientific American, 65 (11):162, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
162a.pdf.

Anonymous:1891:IFH

[2039] Anonymous. An improvement in flush hinges. Scientific American, 65(11):162, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
162.pdf.

Anonymous:1891:IRT

[2040] Anonymous. An improved receiving telephone. Scientific American, 65(11):162, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
162b.pdf.

Anonymous:1891:PSC

[2041] Anonymous. A powerful steam crane. Scientific American, 65(11): 163–165, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
163.pdf.

Anonymous:1891:Cbi

[2042] Anonymous. Correspondence. Scientific American, 65(11):165, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n11/pdf/scientificamerican09121891-165.pdf.

Anonymous:1891:J

[2043] Anonymous. Jupiter. Scientific American, 65(11):165–166, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n11/pdf/scientificamerican09121891-165a.pdf.

Hale:1891:KPO

[2044] George E. Hale. The Kenwood Physical Observatory. Scientific American, 65(11):166–168, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
166.pdf.

Anonymous:1891:NTS

[2045] Anonymous. New toboggan slide for bathers. Scientific American, 65(11):169, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
169.pdf.

Anonymous:1891:BPbl

[2046] Anonymous. Business and personal. Scientific American, 65 (11):170, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
170a.pdf.

Anonymous:1891:RPIbk

[2047] Anonymous. Recently patented inventions. Scientific American, 65 (11):170, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
170.pdf.

Anonymous:1891:NQbn

[2048] Anonymous. Notes and queries. Scientific American, 65(11): 170–171, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
170b.pdf.

Anonymous:1891:Ibq

[2049] Anonymous. To inventors. Scientific American, 65(11):171, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n11/pdf/scientificamerican09121891-171.pdf.

Anonymous:1891:IIbm

[2050] Anonymous. Index of inventions. Scientific American, 65(11): 171–172, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
171a.pdf.

Anonymous:1891:PNPbh

[2051] Anonymous. Patent notes: Patents. Scientific American, 65 (11):174, September 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n11/pdf/scientificamerican09121891-
174.pdf.

Anonymous:1891:QTH

[2052] Anonymous. Quick traveling half around the world. Scientific American, 65(12):176, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
176.pdf.

Anonymous:1891:NELa

[2053] Anonymous. National Electric Light Association. Scientific American, 65(12):176–177, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
176a.pdf.

Anonymous:1891:GFL

[2054] Anonymous. The Grand Falls of Labrador. Scientific American, 65(12):177, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
177.pdf.

Anonymous:1891:L

[2055] Anonymous. Leprosy. Scientific American, 65(12):177–178, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n12/pdf/scientificamerican09191891-177a.pdf.

Anonymous:1891:IMC

[2056] Anonymous. An improved mining car and tipple. Scientific American, 65(12):178, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
178a.pdf.

Anonymous:1891:ISD

[2057] Anonymous. Improved sewer and drain pipe outlet. Scientific American, 65(12):178, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
178.pdf.

Anonymous:1891:SMDb

[2058] Anonymous. The system of military dove Cotes in Europe. Scientific American, 65(12):179, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
179.pdf.

Anonymous:1891:IPFb

[2059] Anonymous. Improvement of Potomac Flats, Washington. Scientific American, 65(12):180–181, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
180.pdf.

Anonymous:1891:Cbj

[2060] Anonymous. Correspondence. Scientific American, 65(12):181, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n12/pdf/scientificamerican09191891-181.pdf.

Wright:1891:LPI

[2061] G. F. Wright. The lava plains of Idaho. Scientific American, 65 (12):182–183, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
182.pdf.

Anonymous:1891:WB

[2062] Anonymous. The Washington bridge. Scientific American, 65(12): 183–185, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
183.pdf.

Anonymous:1891:VL

[2063] Anonymous. The vanishing lady. Scientific American, 65(12): 185, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
185.pdf.

Anonymous:1891:BPbm

[2064] Anonymous. Business and personal. Scientific American, 65 (12):186, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
186a.pdf.

Anonymous:1891:RPIbl

[2065] Anonymous. Recently patented inventions. Scientific American, 65 (12):186, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
186.pdf.

Anonymous:1891:NQbo

[2066] Anonymous. Notes and queries. Scientific American, 65(12): 186–187, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
186b.pdf.

Anonymous:1891:Ibr

[2067] Anonymous. To inventors. Scientific American, 65(12):187, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n12/pdf/scientificamerican09191891-187.pdf.

Anonymous:1891:IIbn

[2068] Anonymous. Index of inventions. Scientific American, 65(12): 187–188, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
187a.pdf.

Anonymous:1891:PNPbi

[2069] Anonymous. Patent notes: Patents. Scientific American, 65 (12):190, September 19, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n12/pdf/scientificamerican09191891-
190.pdf.

Anonymous:1891:FCC

[2070] Anonymous. Fires in cotton cargoes. Scientific American, 65 (13):192, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
192.pdf.

Anonymous:1891:CRS

[2071] Anonymous. The contagion and rapid spread of leprosy. Scientific American, 65(13):192–193, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
192a.pdf.

Anonymous:1891:PPO

[2072] Anonymous. Position of the planets in October. Scientific American, 65(13):193, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
193.pdf.

Anonymous:1891:FEE

[2073] Anonymous. The Frankfort Electrical Exhibition. Scientific American, 65(13):194, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
194.pdf.

Anonymous:1891:IBBa

[2074] Anonymous. An inexpensive bow bridge sill. Scientific American, 65(13):195, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
195b.pdf.

Anonymous:1891:SFH

[2075] Anonymous. A simple form of hat holder. Scientific American, 65(13):195, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
195.pdf.

Anonymous:1891:SWB

[2076] Anonymous. A single wheel biding machine. Scientific American, 65(13):195, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
195a.pdf.

Anonymous:1891:OSC

[2077] Anonymous. Opening of the St. Clair River Railway Tunnel between the United States and Canada. Scientific American, 65(13): 196–197, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
196.pdf.

Anonymous:1891:LEO

[2078] Anonymous. The locomotive explosion at Oyster Bay, L.I. Scientific American, 65(13):198, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
198.pdf.

Seiss:1891:NVSa

[2079] C. Few Seiss. Non-venomous snakes. Scientific American, 65 (13):199, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
199.pdf.

Anonymous:1891:Cbk

[2080] Anonymous. Correspondence. Scientific American, 65(13):200, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n13/pdf/scientificamerican09261891-200.pdf.

Pentz:1891:OKCa

[2081] A. D. Pentz. One kind of cam. Scientific American, 65(13):201, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n13/pdf/scientificamerican09261891-201.pdf.

Anonymous:1891:BPbn

[2082] Anonymous. Business and personal. Scientific American, 65 (13):202, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
202a.pdf.

Anonymous:1891:RPIbm

[2083] Anonymous. Recently patented inventions. Scientific American, 65 (13):202, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
202.pdf.

Anonymous:1891:NQbp

[2084] Anonymous. Notes and queries. Scientific American, 65(13): 202–203, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
202b.pdf.

Anonymous:1891:Ibs

[2085] Anonymous. To inventors. Scientific American, 65(13):203, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n13/pdf/scientificamerican09261891-203.pdf.

Anonymous:1891:IIbo

[2086] Anonymous. Index of inventions. Scientific American, 65(13): 203–204, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
203a.pdf.

Anonymous:1891:PNPbj

[2087] Anonymous. Patent notes: Patents. Scientific American, 65 (13):206, September 26, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n13/pdf/scientificamerican09261891-
206.pdf.

Anonymous:1891:RIS

[2088] Anonymous. Real and imaginary speed of steam yachts. Scientific American, 65(14):208, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
208.pdf.

Anonymous:1891:SR

[2089] Anonymous. Shop rules. Scientific American, 65(14):208–209, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-208a.pdf.

Anonymous:1891:ESBb

[2090] Anonymous. An extractor for stuffing-box packings. Scientific American, 65(14):210, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
210a.pdf.

Anonymous:1891:SFE

[2091] Anonymous. A simple form of electric motor. Scientific American, 65(14):210, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
210.pdf.

Anonymous:1891:FCD

[2092] Anonymous. A foldable clothes drier. Scientific American, 65 (14):210–211, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
210b.pdf.

Anonymous:1891:ACBb

[2093] Anonymous. The auxiliary center board steam yacht Wild Duck. Scientific American, 65(14):211–212, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
211.pdf.

Anonymous:1891:Cbl

[2094] Anonymous. Correspondence. Scientific American, 65(14):213, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-213.pdf.

Anonymous:1891:DWD

[2095] Anonymous. Deep water dredging around New York. Scientific American, 65(14):214, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
214.pdf.

Anonymous:1891:HH

[2096] Anonymous. Hygienic horses. Scientific American, 65(14):214, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-214a.pdf.

Anonymous:1891:LEb

[2097] Anonymous. Locomotive explosion. Scientific American, 65(14):215, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-215.pdf.

Anonymous:1891:TSS

[2098] Anonymous. The twin screw steamer Virginia. Scientific American, 65(14):215–216, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
215a.pdf.

Pentz:1891:OKCb

[2099] A. D. Pentz. One kind of cam. Scientific American, 65(14):217, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-217.pdf.

Anonymous:1891:BNBn

[2100] Anonymous. Books: New books and publications. Scientific American, 65(14):218, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
218a.pdf.

Anonymous:1891:RPIbn

[2101] Anonymous. Recently patented inventions. Scientific American, 65(14):218, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n14/pdf/scientificamerican10031891-
218.pdf.

Anonymous:1891:BPbo

[2102] Anonymous. Business and personal. Scientific American, 65(14):219, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-219.pdf.

Anonymous:1891:Ibu

[2103] Anonymous. To inventors. Scientific American, 65(14):219, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-219b.pdf.

Anonymous:1891:NQbq

[2104] Anonymous. Notes and queries. Scientific American, 65(14):219, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-219a.pdf.

Anonymous:1891:IIbp

[2105] Anonymous. Index of inventions. Scientific American, 65(14):219–220, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-219c.pdf.

Anonymous:1891:PNPbk

[2106] Anonymous. Patent notes: Patents. Scientific American, 65(14):222, October 3, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n14/pdf/scientificamerican10031891-222.pdf.

Anonymous:1891:BTb

[2107] Anonymous. The bread of the tropics. Scientific American, 65 (15):224–225, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
224.pdf.

Anonymous:1891:GOT

[2108] Anonymous. A gravity-operated time alarm. Scientific American, 65(15):226, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
226.pdf.

Anonymous:1891:IBv

[2109] Anonymous. The iguanodons at Brussels. Scientific American, 65(15):226, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
226a.pdf.

Anonymous:1891:ICe

[2110] Anonymous. An improved coupling. Scientific American, 65 (15):226, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
226b.pdf.

Anonymous:1891:DMV

[2111] Anonymous. A driving mechanism for velocipedes. Scientific American, 65(15):227, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
227.pdf.

Pentz:1891:OKCc

[2112] A. D. Pentz. One kind of cam. Scientific American, 65(15):227–228, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n15/pdf/scientificamerican10101891-227a.pdf.

Anonymous:1891:INYb

[2113] Anonymous. An improved neck yoke. Scientific American, 65 (15):228–229, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
228.pdf.

Anonymous:1891:PRB

[2114] Anonymous. The proposed race between fast steam yachts. Scientific American, 65(15):229, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
229.pdf.

Anonymous:1891:PWW

[2115] Anonymous. The Pelton water wheel and motor. Scientific American, 65(15):230, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
230.pdf.

Anonymous:1891:ISM

[2116] Anonymous. An improved sugar making apparatus. Scientific American, 65(15):231, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
231a.pdf.

Anonymous:1891:WSA

[2117] Anonymous. Washing sugar with alcohol. Scientific American, 65(15):231, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
231.pdf.

Anonymous:1891:Cbn

[2118] Anonymous. Correspondence. Scientific American, 65(15):232, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n15/pdf/scientificamerican10101891-232.pdf.

Anonymous:1891:SGM

[2119] Anonymous. The standard gauge for the M.C.B. coupler. Scientific American, 65(15):232, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
232a.pdf.

Seiss:1891:NVSb

[2120] C. Few Seiss. Non-venomous snakes — II. Scientific American, 65(15):233, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
233.pdf.

Anonymous:1891:BNBo

[2121] Anonymous. Books: New books and publications. Scientific American, 65(15):234, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
234a.pdf.

Anonymous:1891:RPIbo

[2122] Anonymous. Recently patented inventions. Scientific American, 65(15):234, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
234.pdf.

Anonymous:1891:BPbp

[2123] Anonymous. Business and personal. Scientific American, 65 (15):235, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
235.pdf.

Anonymous:1891:Ibw

[2124] Anonymous. To inventors. Scientific American, 65(15):235, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n15/pdf/scientificamerican10101891-235b.pdf.

Anonymous:1891:NQbr

[2125] Anonymous. Notes and queries. Scientific American, 65(15):235, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n15/pdf/scientificamerican10101891-235a.pdf.

Anonymous:1891:IIbq

[2126] Anonymous. Index of inventions. Scientific American, 65(15): 235–236, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
235c.pdf.

Anonymous:1891:PNPbl

[2127] Anonymous. Patent notes: Patents. Scientific American, 65 (15):238, October 10, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n15/pdf/scientificamerican10101891-
238.pdf.

Anonymous:1891:BSAb

[2128] Anonymous. Broadway and Seventh Avenue Cable Road. Scientific American, 65(16):239, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
239.pdf.

Anonymous:1891:ETH

[2129] Anonymous. Electrical transmission of 300 horse power. Scientific American, 65(16):240, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
240a.pdf.

Anonymous:1891:SWW

[2130] Anonymous. The ships we want not the kind we are getting. Scientific American, 65(16):240, October 17, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n16/pdf/scientificamerican10171891-240.pdf.

Anonymous:1891:NEE

[2131] Anonymous. A new Edison electro-motor. Scientific American, 65 (16):240–241, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
240b.pdf.

Anonymous:1891:SHF

[2132] Anonymous. The Suabian Harvest Festival. Scientific American, 65(16):242, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
242.pdf.

Anonymous:1891:BPMa

[2133] Anonymous. A binder for papers, magazines, etc. Scientific American, 65(16):243, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
243c.pdf.

Anonymous:1891:IPFc

[2134] Anonymous. An inexpensive portable fence. Scientific American, 65(16):243, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
243a.pdf.

Anonymous:1891:PHH

[2135] Anonymous. A portable hoist for heavy articles. Scientific American, 65(16):243, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
243b.pdf.

Anonymous:1891:SEG

[2136] Anonymous. A simple and efficient gravity hoist. Scientific American, 65(16):243, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
243.pdf.

Anonymous:1891:DJW

[2137] Anonymous. Dr J. W. Clowes’ improvements in dentistry. Scientific American, 65(16):244, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
244.pdf.

Anonymous:1891:Cbo

[2138] Anonymous. Correspondence. Scientific American, 65(16):245, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n16/pdf/scientificamerican10171891-245.pdf.

Anonymous:1891:WF

[2139] Anonymous. William Ferrel. Scientific American, 65(16):247–248, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n16/pdf/scientificamerican10171891-247.pdf.

Anonymous:1891:NBPj

[2140] Anonymous. New books and publications. Scientific American, 65(16):249, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
249a.pdf.

Anonymous:1891:RPIbp

[2141] Anonymous. Recently patented inventions. Scientific American, 65(16):249, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
249.pdf.

Anonymous:1891:BPbq

[2142] Anonymous. Business and personal. Scientific American, 65 (16):250, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
250.pdf.

Anonymous:1891:NQbs

[2143] Anonymous. Notes and queries. Scientific American, 65(16): 250–251, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
250a.pdf.

Anonymous:1891:Ibx

[2144] Anonymous. To inventors. Scientific American, 65(16):251, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n16/pdf/scientificamerican10171891-251.pdf.

Anonymous:1891:IIbr

[2145] Anonymous. Index of inventions. Scientific American, 65(16): 251–252, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
251a.pdf.

Anonymous:1891:PNPbm

[2146] Anonymous. Patent notes: Patents! Scientific American, 65 (16):254, October 17, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n16/pdf/scientificamerican10171891-
254.pdf.

Anonymous:1891:RFT

[2147] Anonymous. Relations of foreign trade to the metric system. Scientific American, 65(17):256, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
256.pdf.

Anonymous:1891:LCL

[2148] Anonymous. Large casting and large forging. Scientific American, 65(17):256–257, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
256a.pdf.

Anonymous:1891:TGU

[2149] Anonymous. Transportation of grain in the United States. Scientific American, 65(17):258–259, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
258.pdf.

Anonymous:1891:IFT

[2150] Anonymous. An improved fiber treating mill. Scientific American, 65(17):259–260, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
259.pdf.

Anonymous:1891:Cbp

[2151] Anonymous. Correspondence. Scientific American, 65(17):261–262, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n17/pdf/scientificamerican10241891-261.pdf.

Anonymous:1891:NCL

[2152] Anonymous. The new Chemical Laboratory of Cornell University. Scientific American, 65(17):262–264, October 24, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n17/pdf/scientificamerican10241891-262.pdf.

Anonymous:1891:NH

[2153] Anonymous. Nipple holder. Scientific American, 65(17):264, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n17/pdf/scientificamerican10241891-264.pdf.

Anonymous:1891:BPbr

[2154] Anonymous. Business and personal. Scientific American, 65 (17):265, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
265a.pdf.

Anonymous:1891:RPIbq

[2155] Anonymous. Recently patented inventions. Scientific American, 65(17):265, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
265.pdf.

Anonymous:1891:NQbt

[2156] Anonymous. Notes and queries. Scientific American, 65(17): 265–267, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
265b.pdf.

Anonymous:1891:Iby

[2157] Anonymous. To inventors. Scientific American, 65(17):267, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n17/pdf/scientificamerican10241891-267.pdf.

Anonymous:1891:IIbs

[2158] Anonymous. Index of inventions. Scientific American, 65(17): 267–268, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n17/pdf/scientificamerican10241891-
267a.pdf.

Anonymous:1891:BNBp

[2159] Anonymous. Books: New books. Scientific American, 65(17):268, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n17/pdf/scientificamerican10241891-268.pdf.

Anonymous:1891:Pf

[2160] Anonymous. Patents! Scientific American, 65(17):270, October 24, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n17/pdf/scientificamerican10241891-270.pdf.

Anonymous:1891:IRA

[2161] Anonymous. The introduction of reindeer into Alaska. Scientific American, 65(18):272, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
272a.pdf.

Anonymous:1891:PI

[2162] Anonymous. Progress of irrigation. Scientific American, 65(18):272, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n18/pdf/scientificamerican10311891-272.pdf.

Anonymous:1891:DNN

[2163] Anonymous. Desertions from the new Navy. Scientific American, 65(18):272–273, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
272b.pdf.

Anonymous:1891:PPN

[2164] Anonymous. Position of the planets in November. Scientific American, 65(18):273, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
273.pdf.

Anonymous:1891:CDB

[2165] Anonymous. The conveyance of dispatches by bees. Scientific American, 65(18):274–275, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
274.pdf.

Anonymous:1891:PHC

[2166] Anonymous. A pulverizing harrow and cultivator. Scientific American, 65(18):275, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
275.pdf.

Anonymous:1891:SABk

[2167] Anonymous. A seat attachment for bicycles. Scientific American, 65(18):275–276, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
275a.pdf.

Anonymous:1891:Cbq

[2168] Anonymous. Correspondence. Scientific American, 65(18):277, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n18/pdf/scientificamerican10311891-277.pdf.

Anonymous:1891:RAW

[2169] Anonymous. Roman Amphitheater at Weehawken N.J. Scientific American, 65(18):278, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
278.pdf.

Anonymous:1891:GPE

[2170] Anonymous. Gigantic passenger elevator of the North Hudson County Railway. Scientific American, 65(18):279, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n18/pdf/scientificamerican10311891-279.pdf.

Anonymous:1891:HSCa

[2171] Anonymous. A horse shoe to be clamped on the hoof. Scientific American, 65(18):279–281, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
279a.pdf.

Anonymous:1891:ISSb

[2172] Anonymous. An improved screw safety car brake. Scientific American, 65(18):281, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
281.pdf.

Anonymous:1891:BNBq

[2173] Anonymous. Books: New books and publications. Scientific American, 65(18):282, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
282a.pdf.

Anonymous:1891:RPIbr

[2174] Anonymous. Recently patented inventions. Scientific American, 65(18):282, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
282.pdf.

Anonymous:1891:BPbs

[2175] Anonymous. Business and personal. Scientific American, 65 (18):283, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
283.pdf.

Anonymous:1891:NQbu

[2176] Anonymous. Notes and queries. Scientific American, 65(18):283, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n18/pdf/scientificamerican10311891-283a.pdf.

Anonymous:1891:Ibz

[2177] Anonymous. To inventors. Scientific American, 65(18):284, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n18/pdf/scientificamerican10311891-284.pdf.

Anonymous:1891:IIbt

[2178] Anonymous. Index of inventions. Scientific American, 65(18):284, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n18/pdf/scientificamerican10311891-284a.pdf.

Anonymous:1891:PNPbn

[2179] Anonymous. Patent notes: Patents! Scientific American, 65 (18):286, October 31, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n18/pdf/scientificamerican10311891-
286.pdf.

Anonymous:1891:DLF

[2180] Anonymous. Dangers of large fly wheels. Scientific American, 65 (19):288–289, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
288.pdf.

Anonymous:1891:APH

[2181] Anonymous. An adjustable picture hanger. Scientific American, 65(19):290, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
290.pdf.

Anonymous:1891:MHN

[2182] Anonymous. A machine and hand needle threader. Scientific American, 65(19):290, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
290b.pdf.

Anonymous:1891:NELb

[2183] Anonymous. A new electric light support. Scientific American, 65(19):290, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
290a.pdf.

Anonymous:1891:TSW

[2184] Anonymous. A trundler for spooled wire. Scientific American, 65(19):290, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
290c.pdf.

Anonymous:1891:HKC

[2185] Anonymous. Hog killing at the Chicago Stock Yards. Scientific American, 65(19):291–292, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
291.pdf.

Anonymous:1891:Cbr

[2186] Anonymous. Correspondence. Scientific American, 65(19):293, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n19/pdf/scientificamerican11071891-293.pdf.

Anonymous:1891:HSCb

[2187] Anonymous. How Stone is crushed and screened. Scientific American, 65(19):294–295, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
294.pdf.

Anonymous:1891:MSC

[2188] Anonymous. The Manchester Ship Canal. Scientific American, 65 (19):295–297, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
295.pdf.

Anonymous:1891:BNBr

[2189] Anonymous. Books: New books and publications. Scientific American, 65(19):298, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
298a.pdf.

Anonymous:1891:RPIbs

[2190] Anonymous. Recently patented inventions. Scientific American, 65(19):298, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
298.pdf.

Anonymous:1891:BPbt

[2191] Anonymous. Business and personal. Scientific American, 65 (19):299, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
299.pdf.

Anonymous:1891:Icf

[2192] Anonymous. To inventors. Scientific American, 65(19):299, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n19/pdf/scientificamerican11071891-299b.pdf.

Anonymous:1891:NQbv

[2193] Anonymous. Notes and queries. Scientific American, 65(19): 299, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
299a.pdf.

Anonymous:1891:IIbu

[2194] Anonymous. Index of inventions. Scientific American, 65(19): 299–300, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
299c.pdf.

Anonymous:1891:PNPbo

[2195] Anonymous. Patent notes: Patents! Scientific American, 65 (19):300, November 7, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n19/pdf/scientificamerican11071891-
302.pdf.

Anonymous:1891:AIF

[2196] Anonymous. American institute fair. Scientific American, 65(20): 304–305, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
304.pdf.

Anonymous:1891:ABA

[2197] Anonymous. An annunciator for burglar alarms, etc. Scientific American, 65(20):306, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
306a.pdf.

Anonymous:1891:ISC

[2198] Anonymous. An improved stalk cutting machine. Scientific American, 65(20):306, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
306b.pdf.

Anonymous:1891:IVBa

[2199] Anonymous. An improved ventilator or blower. Scientific American, 65(20):306, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
306c.pdf.

Anonymous:1891:SMM

[2200] Anonymous. A simple mechanical motor. Scientific American, 65 (20):306, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
306.pdf.

Anonymous:1891:BFR

[2201] Anonymous. A boiler feeder, regulator, and alarm. Scientific American, 65(20):307, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
307.pdf.

Anonymous:1891:GAS

[2202] Anonymous. The German Army swimming exercises. Scientific American, 65(20):307, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
307b.pdf.

Anonymous:1891:IPJ

[2203] Anonymous. An improved pipe joint. Scientific American, 65 (20):307, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
307a.pdf.

Anonymous:1891:BPNa

[2204] Anonymous. Bad paving in New York. Scientific American, 65 (20):308, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
308.pdf.

Anonymous:1891:IFI

[2205] Anonymous. An improved flat iron and heater. Scientific American, 65(20):308, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
308c.pdf.

Anonymous:1891:M

[2206] Anonymous. A mousetrap. Scientific American, 65(20):308, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n20/pdf/scientificamerican11141891-308a.pdf.

Anonymous:1891:SCM

[2207] Anonymous. A simple and convenient music holder. Scientific American, 65(20):308, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
308b.pdf.

Avery:1891:PWP

[2208] Harold Avery. A proposed “Whaleback” passenger steamer. Scientific American, 65(20):309, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
309.pdf.

Anonymous:1891:Cbt

[2209] Anonymous. Correspondence. Scientific American, 65(20):309– 310, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
309a.pdf.

Anonymous:1891:ERD

[2210] Anonymous. Electrical rock drills. Scientific American, 65(20): 310, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
310.pdf.

Anonymous:1891:TSL

[2211] Anonymous. A twin screw launch run by a compound engine. Scientific American, 65(20):311, November 14, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n20/pdf/scientificamerican11141891-311.pdf.

Anonymous:1891:EML

[2212] Anonymous. Electric mining locomotives. Scientific American, 65 (20):312, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
312.pdf.

Anonymous:1891:OGE

[2213] Anonymous. The “Otto” gasoline engine. Scientific American, 65 (20):312, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
312a.pdf.

Anonymous:1891:RPIbt

[2214] Anonymous. Recently patented inventions. Scientific American, 65(20):313, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
313.pdf.

Anonymous:1891:BPbu

[2215] Anonymous. Business and personal. Scientific American, 65 (20):314, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
314.pdf.

Anonymous:1891:NQbw

[2216] Anonymous. Notes and queries. Scientific American, 65(20): 314–315, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
314a.pdf.

Anonymous:1891:Ich

[2217] Anonymous. To inventors. Scientific American, 65(20):315, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n20/pdf/scientificamerican11141891-315.pdf.

Anonymous:1891:IIbv

[2218] Anonymous. Index of inventions. Scientific American, 65(20): 315–316, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
315a.pdf.

Anonymous:1891:PNPbp

[2219] Anonymous. Patent notes: Patents. Scientific American, 65 (20):316, November 14, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n20/pdf/scientificamerican11141891-
316.pdf.

Coonley:1891:SFB

[2220] J. F. Coonley. The sponge fisheries of the Bahamas. Scientific American, 65(21):319, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
319.pdf.

Anonymous:1891:EF

[2221] Anonymous. Economy of flight. Scientific American, 65(21): 320–321, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
320.pdf.

Anonymous:1891:CDR

[2222] Anonymous. A continuous driving ratchet drill. Scientific American, 65(21):322, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
322b.pdf.

Anonymous:1891:ELC

[2223] Anonymous. An electric lamp cover and switch. Scientific American, 65(21):322, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
322.pdf.

Anonymous:1891:PHW

[2224] Anonymous. A portable hot water or steam heater. Scientific American, 65(21):322, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
322a.pdf.

Anonymous:1891:WMS

[2225] Anonymous. A wagon Maker’s spoke puller. Scientific American, 65(21):322, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
322c.pdf.

Anonymous:1891:DEO

[2226] Anonymous. A design for easels and other stands. Scientific American, 65(21):323, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
323b.pdf.

Anonymous:1891:ICi

[2227] Anonymous. An improved churn. Scientific American, 65(21): 323, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
323.pdf.

Anonymous:1891:ITM

[2228] Anonymous. An improved typewriting machine. Scientific American, 65(21):323, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
323a.pdf.

Anonymous:1891:WFL

[2229] Anonymous. The World’s Fair Lumber and Forestry Building. Scientific American, 65(21):324, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
324.pdf.

Anonymous:1891:Cbu

[2230] Anonymous. Correspondence. Scientific American, 65(21):325, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n21/pdf/scientificamerican11211891-325.pdf.

Anonymous:1891:LDD

[2231] Anonymous. Largest dry dock on the Great Lakes. Scientific American, 65(21):327, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
327.pdf.

Anonymous:1891:SESb

[2232] Anonymous. A steam engine slide valve. Scientific American, 65 (21):327–328, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
327a.pdf.

Hopkins:1891:SSP

[2233] M. Hopkins. Some suggestions in photography. Scientific American, 65(21):329, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
329.pdf.

Anonymous:1891:BPbv

[2234] Anonymous. Business and personal. Scientific American, 65 (21):330, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
330a.pdf.

Anonymous:1891:RPIbu

[2235] Anonymous. Recently patented inventions. Scientific American, 65(21):330, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
330.pdf.

Anonymous:1891:NQbx

[2236] Anonymous. Notes and queries. Scientific American, 65(21): 330–331, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
330b.pdf.

Anonymous:1891:BNBs

[2237] Anonymous. Books: New books and publications. Scientific American, 65(21):331, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
331.pdf.

Anonymous:1891:Icj

[2238] Anonymous. To inventors. Scientific American, 65(21):331, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n21/pdf/scientificamerican11211891-331a.pdf.

Anonymous:1891:IIbw

[2239] Anonymous. Index of inventions. Scientific American, 65(21): 331–332, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
331b.pdf.

Anonymous:1891:PNPbq

[2240] Anonymous. Patent notes: Patents. Scientific American, 65 (21):333, November 21, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n21/pdf/scientificamerican11211891-
333.pdf.

Anonymous:1891:IISa

[2241] Anonymous. Important improvements in sorghum sugar manufacture. Scientific American, 65(22):336, November 28, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n22/pdf/scientificamerican11281891-336a.pdf.

Anonymous:1891:PEG

[2242] Anonymous. Poor engineers and good boilers. Scientific American, 65(22):336, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
336b.pdf.

Anonymous:1891:PSB

[2243] Anonymous. Plates for secondary batteries. Scientific American, 65(22):336, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
336.pdf.

Anonymous:1891:PPD

[2244] Anonymous. Position of the planets in December. Scientific American, 65(22):337, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
337.pdf.

Anonymous:1891:CHD

[2245] Anonymous. Coal handling devices. Scientific American, 65(22): 338, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
338a.pdf.

Anonymous:1891:NSC

[2246] Anonymous. A new street cleaner. Scientific American, 65(22): 338, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
338.pdf.

Anonymous:1891:AGM

[2247] Anonymous. An automatic grain measuring machine. Scientific American, 65(22):339, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
339c.pdf.

Anonymous:1891:EKG

[2248] Anonymous. An efficient knife-grinding machine. Scientific American, 65(22):339, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
339b.pdf.

Anonymous:1891:ISL

[2249] Anonymous. An improved springless lock. Scientific American, 65 (22):339, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
339a.pdf.

Anonymous:1891:IWF

[2250] Anonymous. An improved window frame. Scientific American, 65 (22):339, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
339d.pdf.

Anonymous:1891:PSV

[2251] Anonymous. A pipe die stock for various sizes of pipe. Scientific American, 65(22):339, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
339.pdf.

Anonymous:1891:E

[2252] Anonymous. The Exhibition of 1892–93. Scientific American, 65 (22):340, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
340.pdf.

Anonymous:1891:Cbv

[2253] Anonymous. Correspondence. Scientific American, 65(22):341, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n22/pdf/scientificamerican11281891-341.pdf.

Anonymous:1891:SEM

[2254] Anonymous. Static electrical motors. Scientific American, 65(22): 342–343, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
342.pdf.

Anonymous:1891:BFM

[2255] Anonymous. Bessemer’s fluid metal rolling mill. Scientific American, 65(22):343, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
343.pdf.

Anonymous:1891:PNi

[2256] Anonymous. Photographic notes. Scientific American, 65(22): 344, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
344.pdf.

Anonymous:1891:TSG

[2257] Anonymous. A telescope for schools and general use. Scientific American, 65(22):345, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
345.pdf.

Anonymous:1891:BNBt

[2258] Anonymous. Books: New books and publications. Scientific American, 65(22):346, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
346a.pdf.

Anonymous:1891:RPIbv

[2259] Anonymous. Recently patented inventions. Scientific American, 65(22):346, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
346.pdf.

Anonymous:1891:BPbw

[2260] Anonymous. Business and personal. Scientific American, 65 (22):347, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
347.pdf.

Anonymous:1891:Ick

[2261] Anonymous. To inventors. Scientific American, 65(22):347, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n22/pdf/scientificamerican11281891-347b.pdf.

Anonymous:1891:NQby

[2262] Anonymous. Notes and queries. Scientific American, 65(22): 347, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
347a.pdf.

Anonymous:1891:IIbx

[2263] Anonymous. Index of inventions. Scientific American, 65(22): 347–348, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
347c.pdf.

Anonymous:1891:PNPbr

[2264] Anonymous. Patent notes: Patents. Scientific American, 65 (22):349, November 28, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n22/pdf/scientificamerican11281891-
349.pdf.

Anonymous:1891:NAI

[2265] Anonymous. The Nadrai Aqueduct, India. Scientific American, 65 (23):351–352, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
351.pdf.

Anonymous:1891:USLb

[2266] Anonymous. The United States leads the world in iron. Scientific American, 65(23):352, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
352.pdf.

Anonymous:1891:ACC

[2267] Anonymous. Automatic car couplers. Scientific American, 65(23): 352–353, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
352a.pdf.

Anonymous:1891:OFU

[2268] Anonymous. Oil fuel under stationary boilers. Scientific American, 65(23):353, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
353.pdf.

Anonymous:1891:BB

[2269] Anonymous. Benzine blowpipe. Scientific American, 65(23):354, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n23/pdf/scientificamerican12051891-354.pdf.

Anonymous:1891:IKF

[2270] Anonymous. An improved knitted fabric. Scientific American, 65 (23):354–355, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
354a.pdf.

Anonymous:1891:ARG

[2271] Anonymous. An automatic railroad gate. Scientific American, 65(23):355, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
355.pdf.

Anonymous:1891:THE

[2272] Anonymous. A tool holder for engine lathes. Scientific American, 65(23):355, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
355a.pdf.

Anonymous:1891:PDM

[2273] Anonymous. A practical, durable, metallic railway tie. Scientific American, 65(23):355–356, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
355b.pdf.

Anonymous:1891:IWCa

[2274] Anonymous. An improved water closet. Scientific American, 65 (23):356, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
356.pdf.

Anonymous:1891:LP

[2275] Anonymous. Lambrecht’s polymeter. Scientific American, 65(23): 356–357, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
356a.pdf.

Anonymous:1891:ICCe

[2276] Anonymous. Improved cable-conveying apparatus. Scientific American, 65(23):358, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
358.pdf.

Anonymous:1891:O

[2277] Anonymous. The ostrich. Scientific American, 65(23):359–360, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n23/pdf/scientificamerican12051891-359.pdf.

Anonymous:1891:AUS

[2278] Anonymous. Another United States cruiser afloat. Scientific American, 65(23):361, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
361.pdf.

Anonymous:1891:BPbx

[2279] Anonymous. Business and personal. Scientific American, 65 (23):362, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
362a.pdf.

Anonymous:1891:RPIbw

[2280] Anonymous. Recently patented inventions. Scientific American, 65(23):362, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
362.pdf.

Anonymous:1891:NQbz

[2281] Anonymous. Notes and queries. Scientific American, 65(23): 362–363, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
362b.pdf.

Anonymous:1891:Icl

[2282] Anonymous. To inventors. Scientific American, 65(23):363, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n23/pdf/scientificamerican12051891-363.pdf.

Anonymous:1891:IIby

[2283] Anonymous. Index of inventions. Scientific American, 65(23): 363–364, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
363a.pdf.

Anonymous:1891:PNPbs

[2284] Anonymous. Patent notes: Patents. Scientific American, 65 (23):365, December 5, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n23/pdf/scientificamerican12051891-
365.pdf.

Anonymous:1891:FDC

[2285] Anonymous. Failure of the dynamite cruiser Vesuvius. Scientific American, 65(24):368, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
368.pdf.

Anonymous:1891:HGB

[2286] Anonymous. Heavy guns and the best armor. Scientific American, 65(24):368–370, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
368a.pdf.

Anonymous:1891:AFP

[2287] Anonymous. An automatic frisket for printing presses. Scientific American, 65(24):370, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
370b.pdf.

Anonymous:1891:EDO

[2288] Anonymous. An elevator door operating device. Scientific American, 65(24):370, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
370a.pdf.

Anonymous:1891:FSF

[2289] Anonymous. A feeder for stoves and furnaces. Scientific American, 65(24):370, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
370.pdf.

Anonymous:1891:LON

[2290] Anonymous. The largest of our new war ships. Scientific American, 65(24):371, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
371.pdf.

Pike:1891:CSA

[2291] Nicolas Pike. The Chacma or South African baboon. Scientific American, 65(24):371–372, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
371a.pdf.

Anonymous:1891:Cbx

[2292] Anonymous. Correspondence. Scientific American, 65(24):373, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n24/pdf/scientificamerican12121891-373a.pdf.

Anonymous:1891:SCTc

[2293] Anonymous. The St. Clair Tunnel drainage system. Scientific American, 65(24):373, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
373.pdf.

Anonymous:1891:TT

[2294] Anonymous. The Tocci twins. Scientific American, 65(24):374, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n24/pdf/scientificamerican12121891-374.pdf.

Anonymous:1891:EPb

[2295] Anonymous. Experiments in prestidigitation. Scientific American, 65 (24):374–375, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
374a.pdf.

Anonymous:1891:GT

[2296] Anonymous. The Galapagas tortoises. Scientific American, 65(24): 375–376, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
375.pdf.

Anonymous:1891:MBW

[2297] Anonymous. Model battleship at the World’s Columbian Exposition. Scientific American, 65(24):376, December 12, 1891. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v65/
n24/pdf/scientificamerican12121891-376.pdf.

Anonymous:1891:BPby

[2298] Anonymous. Business and personal. Scientific American, 65 (24):377, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
377a.pdf.

Anonymous:1891:RPIbx

[2299] Anonymous. Recently patented inventions. Scientific American, 65(24):377, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
377.pdf.

Anonymous:1891:NQca

[2300] Anonymous. Notes and queries. Scientific American, 65(24): 377–379, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
377b.pdf.

Anonymous:1891:Icm

[2301] Anonymous. To inventors. Scientific American, 65(24):379, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v65/n24/pdf/scientificamerican12121891-379.pdf.

Anonymous:1891:IIbz

[2302] Anonymous. Index of inventions. Scientific American, 65(24): 379, December 12, 1891. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v65/n24/pdf/scientificamerican12121891-
379a.pdf.

Anonymous:1892:DSV

[2303] Anonymous. Demand for sailing vessels increasing. Scientific American, 66(1):2, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n1/pdf/scientificamerican01021892-
2.pdf.

Anonymous:1892:EWE

[2304] Anonymous. Eli Whitney and Eli Whitney Blake. Scientific American, 66(1):2, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n1/pdf/scientificamerican01021892-
2a.pdf.

Anonymous:1892:PPJa

[2305] Anonymous. Position of the planets in January. Scientific American, 66(1):2–3, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n1/pdf/scientificamerican01021892-
2b.pdf.

Anonymous:1892:CS

[2306] Anonymous. A compact stereopticon. Scientific American, 66(1):4, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-4a.pdf.

Anonymous:1892:IRR

[2307] Anonymous. An improved railroad rail, sleeper and rail fastener. Scientific American, 66(1):4, January 2, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n1/pdf/scientificamerican01021892-4.pdf.

Anonymous:1892:IWa

[2308] Anonymous. An improved windmill. Scientific American, 66(1):4, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-4b.pdf.

Anonymous:1892:IWM

[2309] Anonymous. An improved water motor. Scientific American, 66(1):4, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-4c.pdf.

Anonymous:1892:TRM

[2310] Anonymous. The Texas rain making experiments. Scientific American, 66(1):5–6, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n1/pdf/scientificamerican01021892-
5.pdf.

Anonymous:1892:LHB

[2311] Anonymous. Lorillard house boat Caiman. Scientific American, 66(1):7, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-7.pdf.

Anonymous:1892:ISP

[2312] Anonymous. An improved Shipman pumping outfit. Scientific American, 66(1):8, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n1/pdf/scientificamerican01021892-
8.pdf.

Anonymous:1892:WRC

[2313] Anonymous. A Wisconsin Railway collision. Scientific American, 66(1):8, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n1/pdf/scientificamerican01021892-
8a.pdf.

Anonymous:1892:CTa

[2314] Anonymous. The Columbian Tower. Scientific American, 66(1):9, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-9.pdf.

Anonymous:1892:BPa

[2315] Anonymous. Business and personal. Scientific American, 66(1):10, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-10a.pdf.

Anonymous:1892:RPIa

[2316] Anonymous. Recently patented inventions. Scientific American, 66(1):10, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n1/pdf/scientificamerican01021892-
10.pdf.

Anonymous:1892:NQa

[2317] Anonymous. Notes and queries. Scientific American, 66(1):10–11, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-10b.pdf.

Anonymous:1892:Ia

[2318] Anonymous. To inventors. Scientific American, 66(1):11, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-11.pdf.

Anonymous:1892:IIa

[2319] Anonymous. Index of inventions. Scientific American, 66(1):11–14, January 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n1/pdf/scientificamerican01021892-11a.pdf.

Anonymous:1892:BST

[2320] Anonymous. Boom still for the trolley. Scientific American, 66(2):16, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-16.pdf.

Anonymous:1892:WSC

[2321] Anonymous. Wooden shoes and clogs. Scientific American, 66(2):16–17, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-16a.pdf.

Anonymous:1892:DRL

[2322] Anonymous. A device for raising liquids. Scientific American, 66(2):18, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-18.pdf.

Anonymous:1892:ECa

[2323] Anonymous. An electric carriage. Scientific American, 66(2):18, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-18a.pdf.

Anonymous:1892:BTF

[2324] Anonymous. Breakage of a thirty-foot flywheel at the Amoskeag Mills, Manchester, N.H. Scientific American, 66(2):19–20, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-19.pdf.

Cuttriss:1892:TTE

[2325] Chas Cuttriss. A telephone transmitter without electrodes. Scientific American, 66(2):20, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n2/pdf/scientificamerican01091892-
20.pdf.

Anonymous:1892:Ca

[2326] Anonymous. Correspondence. Scientific American, 66(2):21, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-21.pdf.

Anonymous:1892:FVC

[2327] Anonymous. The first voyage of Columbus. Scientific American, 66(2):22, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n2/pdf/scientificamerican01091892-
22.pdf.

Anonymous:1892:PWFa

[2328] Anonymous. Progress of the World’s Fair buildings. Scientific American, 66(2):23–25, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n2/pdf/scientificamerican01091892-
23.pdf.

Anonymous:1892:BPb

[2329] Anonymous. Business and personal. Scientific American, 66(2):26, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-26a.pdf.

Anonymous:1892:RPIb

[2330] Anonymous. Recently patented inventions. Scientific American, 66(2):26, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n2/pdf/scientificamerican01091892-
26.pdf.

Anonymous:1892:NQb

[2331] Anonymous. Notes and queries. Scientific American, 66(2):26–27, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-26b.pdf.

Anonymous:1892:Ib

[2332] Anonymous. To inventors. Scientific American, 66(2):27, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-27.pdf.

Anonymous:1892:IIb

[2333] Anonymous. Index of inventions. Scientific American, 66(2):27–28, January 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n2/pdf/scientificamerican01091892-27a.pdf.

Anonymous:1892:TSa

[2334] Anonymous. The traveling sidewalk. Scientific American, 66 (3):31–32, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
31.pdf.

Anonymous:1892:ACSa

[2335] Anonymous. The American chemical society. Scientific American, 66(3):32–33, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
32.pdf.

Anonymous:1892:NLS

[2336] Anonymous. New Lake steamers. Scientific American, 66(3):34, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n3/pdf/scientificamerican01161892-34.pdf.

Anonymous:1892:TMB

[2337] Anonymous. A tool for miners and blasters. Scientific American, 66(3):34, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
34b.pdf.

Anonymous:1892:TSD

[2338] Anonymous. A tipple for separating and delivering coal. Scientific American, 66(3):34, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
34a.pdf.

Anonymous:1892:CRM

[2339] Anonymous. The cable road on Mount San Salvatore, near Lugano. Scientific American, 66(3):35, January 16, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n3/pdf/scientificamerican01161892-35.pdf.

Anonymous:1892:TW

[2340] Anonymous. Telegraphing without wires. Scientific American, 66(3):36, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
36.pdf.

Anonymous:1892:OAD

[2341] Anonymous. An overshoe attaching device. Scientific American, 66(3):36–37, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
36a.pdf.

Anonymous:1892:DM

[2342] Anonymous. Dynamo machines. Scientific American, 66(3):38, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n3/pdf/scientificamerican01161892-38.pdf.

Anonymous:1892:HHa

[2343] Anonymous. The hairless horse. Scientific American, 66(3):38, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n3/pdf/scientificamerican01161892-38a.pdf.

Anonymous:1892:PSH

[2344] Anonymous. Pike spearing on Hamilton bay. Scientific American, 66(3):39, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
39.pdf.

Anonymous:1892:ISCa

[2345] Anonymous. An improved safety catch for guns. Scientific American, 66(3):40, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
40a.pdf.

Anonymous:1892:NMIa

[2346] Anonymous. A novel musical instrument. Scientific American, 66(3):40, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
40.pdf.

Anonymous:1892:BNBa 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
41a.pdf.

Anonymous:1892:RPIc

[2348] Anonymous. Recently patented inventions. Scientific American, 66(3):41, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n3/pdf/scientificamerican01161892-
41.pdf.

Anonymous:1892:BPc

[2349] Anonymous. Business and personal. Scientific American, 66(3):42, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n3/pdf/scientificamerican01161892-42.pdf.

Anonymous:1892:NQc

[2350] Anonymous. Notes and queries. Scientific American, 66(3):42–43, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n3/pdf/scientificamerican01161892-42a.pdf.

Anonymous:1892:Ic

[2351] Anonymous. To inventors. Scientific American, 66(3):43, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n3/pdf/scientificamerican01161892-43.pdf.

Anonymous:1892:IIc

[2352] Anonymous. Index of inventions. Scientific American, 66(3):43–44, January 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n3/pdf/scientificamerican01161892-43a.pdf.

Anonymous:1892:INN

[2353] Anonymous. Improvements needed in the new Navy. Scientific American, 66(4):48–50, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
48.pdf.

Anonymous:1892:AM

[2354] Anonymous. The ammonia motor. Scientific American, 66(4): 50–51, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
50.pdf.

Anonymous:1892:IFCa

[2355] Anonymous. An improved freight car door. Scientific American, 66(4):51, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
51a.pdf.

Anonymous:1892:IPA

[2356] Anonymous. An improved piano action. Scientific American, 66(4):51, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
51.pdf.

Anonymous:1892:ISD

[2357] Anonymous. Improved safety devices for elevators. Scientific American, 66(4):51, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
51b.pdf.

Anonymous:1892:Cb

[2358] Anonymous. Correspondence. Scientific American, 66(4):52, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n4/pdf/scientificamerican01231892-52.pdf.

Anonymous:1892:TCB

[2359] Anonymous. Towing canal boats by electricity. Scientific American, 66(4):52, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
52a.pdf.

Avery:1892:MT

[2360] Harold Avery. A method of tunneling. Scientific American, 66 (4):53–54, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
53.pdf.

Anonymous:1892:ISCb

[2361] Anonymous. Improved siphon condenser. Scientific American, 66(4):54, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
54a.pdf.

Anonymous:1892:VCB

[2362] Anonymous. Vertical cylinder boring machine. Scientific American, 66(4):54, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
54.pdf.

Anonymous:1892:WFF

[2363] Anonymous. The World’s Fair Franklin Statue. Scientific American, 66(4):55–57, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
55.pdf.

Anonymous:1892:BSM

[2364] Anonymous. A bow for stringed musical instruments. Scientific American, 66(4):57, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
57.pdf.

Anonymous:1892:IDCa

[2365] Anonymous. An improves dumping car. Scientific American, 66 (4):57, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
57a.pdf.

Anonymous:1892:IWP

[2366] Anonymous. An improved weed puller. Scientific American, 66 (4):57, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
57b.pdf.

Anonymous:1892:BPBa

[2367] Anonymous. Business and personal: Business and personal. Scientific American, 66(4):58, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
58b.pdf.

Anonymous:1892:NBPa

[2368] Anonymous. New books and publications. Scientific American, 66(4):58, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
58a.pdf.

Anonymous:1892:RPId

[2369] Anonymous. Recently patented inventions. Scientific American, 66(4):58, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n4/pdf/scientificamerican01231892-
58.pdf.

Anonymous:1892:NQd

[2370] Anonymous. Notes and queries. Scientific American, 66(4):58–59, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n4/pdf/scientificamerican01231892-58c.pdf.

Anonymous:1892:Id

[2371] Anonymous. To inventors. Scientific American, 66(4):59, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n4/pdf/scientificamerican01231892-59.pdf.

Anonymous:1892:IId

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n4/pdf/scientificamerican01231892-59a.pdf.

Anonymous:1892:P

[2373] Anonymous. Patents. Scientific American, 66(4):62, January 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n4/pdf/scientificamerican01231892-62.pdf.

Anonymous:1892:NFB

[2374] Anonymous. Niagara Falls and Buffalo. Scientific American, 66 (5):64, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
64.pdf.

Anonymous:1892:PTO

[2375] Anonymous. Patent tricks-old and new. Scientific American, 66 (5):64, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
64a.pdf.

Anonymous:1892:PPF

[2376] Anonymous. Position of the planets in February. Scientific American, 66(5):65–66, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
65.pdf.

Anonymous:1892:LBP

[2377] Anonymous. Lower Bow Park, Banff. Scientific American, 66 (5):66, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
66.pdf.

Anonymous:1892:FM

[2378] Anonymous. A flying machine. Scientific American, 66(5):67, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n5/pdf/scientificamerican01301892-67a.pdf.

Anonymous:1892:ICCa

[2379] Anonymous. An improved car coupling. Scientific American, 66 (5):67, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
67b.pdf.

Anonymous:1892:IPWa

[2380] Anonymous. An improved pipe wrench. Scientific American, 66 (5):67, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
67.pdf.

Anonymous:1892:Cc

[2381] Anonymous. Correspondence. Scientific American, 66(5):68, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n5/pdf/scientificamerican01301892-68.pdf.

Anonymous:1892:MIS

[2382] Anonymous. The main intersecting sewer of the City of Brooklyn, N.Y. Scientific American, 66(5):69, January 30, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n5/pdf/scientificamerican01301892-69.pdf.

Anonymous:1892:RBG

[2383] Anonymous. Railway bridge over the Ganges, India. Scientific American, 66(5):70, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
70.pdf.

Anonymous:1892:GMC

[2384] Anonymous. General Montgomery Cunningham Meigs. Scientific American, 66(5):71–72, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
71.pdf.

Anonymous:1892:Ma

[2385] Anonymous. The mastodon. Scientific American, 66(5):72, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n5/pdf/scientificamerican01301892-72.pdf.

Anonymous:1892:RPIe

[2386] Anonymous. Recently patented inventions. Scientific American, 66(5):73, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
73.pdf.

Anonymous:1892:BNBb

[2387] Anonymous. Books: New books and publications. Scientific American, 66(5):73–74, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n5/pdf/scientificamerican01301892-
73a.pdf.

Anonymous:1892:BPd

[2388] Anonymous. Business and personal. Scientific American, 66(5):74, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n5/pdf/scientificamerican01301892-74.pdf.

Anonymous:1892:NQe

[2389] Anonymous. Notes and queries. Scientific American, 66(5):74–75, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n5/pdf/scientificamerican01301892-74a.pdf.

Anonymous:1892:Ie

[2390] Anonymous. To inventors. Scientific American, 66(5):76, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n5/pdf/scientificamerican01301892-76.pdf.

Anonymous:1892:IIe

[2391] Anonymous. Index of inventions. Scientific American, 66(5):76, January 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n5/pdf/scientificamerican01301892-76a.pdf.

Anonymous:1892:ES

[2392] Anonymous. The evening sky. Scientific American, 66(6):80, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n6/pdf/scientificamerican02061892-80a.pdf.

Anonymous:1892:RSM

[2393] Anonymous. Removal of snow by melting. Scientific American, 66(6):80, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
80.pdf.

Anonymous:1892:WSH

[2394] Anonymous. The war ship and her crew. Scientific American, 66(6):80–81, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
80b.pdf.

Anonymous:1892:BVS

[2395] Anonymous. A balanced valve for steam engines. Scientific American, 66(6):82, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
82b.pdf.

Anonymous:1892:IOW

[2396] Anonymous. An improved Ore washing jigger. Scientific American, 66(6):82, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
82c.pdf.

Anonymous:1892:IWb

[2397] Anonymous. An improved wrench. Scientific American, 66(6): 82, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
82a.pdf.

Anonymous:1892:MFF

[2398] Anonymous. A machine to facilitate filling of bags. Scientific American, 66(6):82, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
82.pdf.

Anonymous:1892:TBS

[2399] Anonymous. The “Temple Block,” Salt Lake City. Scientific American, 66(6):83, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
83.pdf.

Anonymous:1892:MGS

[2400] Anonymous. A machine to grind sickle blades, etc. Scientific American, 66(6):83–84, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
83a.pdf.

Anonymous:1892:AG

[2401] Anonymous. The “Adamson” gun. Scientific American, 66(6): 84, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
84.pdf.

Anonymous:1892:Cd

[2402] Anonymous. Correspondence. Scientific American, 66(6):85, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n6/pdf/scientificamerican02061892-85a.pdf.

Anonymous:1892:NAW

[2403] Anonymous. The new American war steamer Miantonomoh. Scientific American, 66(6):85, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
85.pdf.

Anonymous:1892:WGF

[2404] Anonymous. The Washington gun factory. Scientific American, 66(6):86, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
86.pdf.

Anonymous:1892:SSW

[2405] Anonymous. Steam Stone works. Scientific American, 66(6): 87–89, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
87.pdf.

Anonymous:1892:BNBc

[2406] Anonymous. Books: New books and publications. Scientific American, 66(6):90, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
90a.pdf.

Anonymous:1892:BPe

[2407] Anonymous. Business and personal. Scientific American, 66 (6):90, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
90b.pdf.

Anonymous:1892:RPIf

[2408] Anonymous. Recently patented inventions. Scientific American, 66(6):90, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
90.pdf.

Anonymous:1892:NQf

[2409] Anonymous. Notes and queries. Scientific American, 66(6):90– 91, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
90c.pdf.

Anonymous:1892:If

[2410] Anonymous. To inventors. Scientific American, 66(6):91, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n6/pdf/scientificamerican02061892-91.pdf.

Anonymous:1892:IIf

[2411] Anonymous. Index of inventions. Scientific American, 66(6): 91–92, February 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n6/pdf/scientificamerican02061892-
91a.pdf.

Anonymous:1892:VES

[2412] Anonymous. A violent eruption of the Sun. Scientific American, 66(7):96, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
96.pdf.

Anonymous:1892:CRCa

[2413] Anonymous. The Congressional report of the Commissioner of Patents. Scientific American, 66(7):96–97, February 13, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n7/pdf/scientificamerican02131892-96a.pdf.

Anonymous:1892:NSM

[2414] Anonymous. New system of maritime navigation. Scientific American, 66(7):98, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
98.pdf.

Anonymous:1892:EHC

[2415] Anonymous. An ensilage harvester and chopper. Scientific American, 66(7):99, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
99b.pdf.

Anonymous:1892:ICa

[2416] Anonymous. An improved cultivator. Scientific American, 66 (7):99, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
99a.pdf.

Anonymous:1892:ICL

[2417] Anonymous. An improved clothes line pulley. Scientific American, 66(7):99, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
99c.pdf.

Anonymous:1892:SMP

[2418] Anonymous. A straining and measuring pot. Scientific American, 66(7):99, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
99.pdf.

Anonymous:1892:Ce

[2419] Anonymous. Correspondence. Scientific American, 66(7):100, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n7/pdf/scientificamerican02131892-100.pdf.

Pike:1892:FTP

[2420] N. Pike. Ferns: Their preservative properties and varied uses. Scientific American, 66(7):100, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
100a.pdf.

Anonymous:1892:DPL

[2421] Anonymous. Dr. S. P. Langley’s experiments in aerodynamics. Scientific American, 66(7):101, February 13, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n7/pdf/scientificamerican02131892-101.pdf.

Anonymous:1892:ETP

[2422] Anonymous. The electrical transmission of power between Lauffen on the Neckar and Frankfort on the Main. Scientific American, 66(7):102, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
102.pdf.

Anonymous:1892:CFB

[2423] Anonymous. California’s famous big trees. Scientific American, 66 (7):103–104, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
103.pdf.

Anonymous:1892:TA

[2424] Anonymous. Trouve’s aviator. Scientific American, 66(7):105, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n7/pdf/scientificamerican02131892-105.pdf.

Anonymous:1892:BNBd

[2425] Anonymous. Books: New books and publications. Scientific American, 66(7):106, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
106a.pdf.

Anonymous:1892:RPIg

[2426] Anonymous. Recently patented inventions. Scientific American, 66(7):106, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
106.pdf.

Anonymous:1892:BPf

[2427] Anonymous. Business and personal. Scientific American, 66(7): 107, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
107.pdf.

Anonymous:1892:NQg

[2428] Anonymous. Notes and queries. Scientific American, 66(7):107, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n7/pdf/scientificamerican02131892-107a.pdf.

Anonymous:1892:IIg

[2429] Anonymous. Index of inventions. Scientific American, 66(7): 107–108, February 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n7/pdf/scientificamerican02131892-
107b.pdf.

Anonymous:1892:NCL

[2430] Anonymous. New compound locomotive. Scientific American, 66 (8):111–112, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
111.pdf.

Anonymous:1892:BI

[2431] Anonymous. Belated inventors. Scientific American, 66(8):112– 113, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
112.pdf.

Anonymous:1892:EPI

[2432] Anonymous. Exclusion of photographs from the International Postal Exchange. Scientific American, 66(8):113–115, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n8/pdf/scientificamerican02201892-113.pdf.

Anonymous:1892:CKC

[2433] Anonymous. A convenient kitchen cabinet. Scientific American, 66(8):115, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
115.pdf.

Anonymous:1892:IFEa

[2434] Anonymous. An improved fire escape. Scientific American, 66 (8):115, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
115a.pdf.

Anonymous:1892:IHPa

[2435] Anonymous. An improved hand planer. Scientific American, 66 (8):115–116, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
115b.pdf.

Anonymous:1892:GSSa

[2436] Anonymous. A great spot on the Sun. Scientific American, 66 (8):116, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
116.pdf.

Anonymous:1892:Cf

[2437] Anonymous. Correspondence. Scientific American, 66(8):117, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n8/pdf/scientificamerican02201892-117.pdf.

Anonymous:1892:OFT

[2438] Anonymous. The Odd Fellows’ Temple, Chicago. Scientific American, 66(8):118–119, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
118.pdf.

Anonymous:1892:SPS

[2439] Anonymous. The snow plant of the Sierras. Scientific American, 66(8):119–120, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
119.pdf.

Anonymous:1892:RPIh

[2440] Anonymous. Recently patented inventions. Scientific American, 66(8):121, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
121.pdf.

Anonymous:1892:BNBe

[2441] Anonymous. Books: New books and publications. Scientific American, 66(8):121–122, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
121a.pdf.

Anonymous:1892:BPg

[2442] Anonymous. Business and personal. Scientific American, 66(8): 122, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
122.pdf.

Anonymous:1892:NQh

[2443] Anonymous. Notes and queries. Scientific American, 66(8):122– 123, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
122a.pdf.

Anonymous:1892:Ig

[2444] Anonymous. To inventors. Scientific American, 66(8):123, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n8/pdf/scientificamerican02201892-123.pdf.

Anonymous:1892:IIh

[2445] Anonymous. Index of inventions. Scientific American, 66(8): 123–124, February 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n8/pdf/scientificamerican02201892-
123a.pdf.

Anonymous:1892:RI

[2446] Anonymous. Road improvement. Scientific American, 66(9):128, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n9/pdf/scientificamerican02271892-128.pdf.

Anonymous:1892:PPMa

[2447] Anonymous. Position of the planets in March. Scientific American, 66(9):129, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
129.pdf.

Anonymous:1892:IDP

[2448] Anonymous. An improved dental plugger. Scientific American, 66(9):130, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
130.pdf.

Anonymous:1892:IWBa

[2449] Anonymous. An improved wood bending machine. Scientific American, 66(9):130, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
130a.pdf.

Anonymous:1892:SSSa

[2450] Anonymous. A silk spinning spindle support. Scientific American, 66(9):130, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
130b.pdf.

Anonymous:1892:ICb

[2451] Anonymous. Indoor curling. Scientific American, 66(9):131, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n9/pdf/scientificamerican02271892-131.pdf.

Anonymous:1892:NEG

[2452] Anonymous. New electrical governor. Scientific American, 66(9): 131–132, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
131a.pdf.

Drake:1892:HTS

[2453] J. V. Drake. How to take silver stains out of a gelatine negative. Scientific American, 66(9):132, February 27, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n9/pdf/scientificamerican02271892-132.pdf.

Anonymous:1892:Cg

[2454] Anonymous. Correspondence. Scientific American, 66(9):133, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n9/pdf/scientificamerican02271892-133.pdf.

Morton:1892:MML

[2455] Morton. Magnitude of molecules and light waves. Scientific American, 66(9):133, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
133a.pdf.

Anonymous:1892:TP

[2456] Anonymous. The transportation of petroleum. Scientific American, 66(9):134–135, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
134.pdf.

Anonymous:1892:PJL

[2457] Anonymous. Professor Joseph Lovering. Scientific American, 66 (9):136, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
136.pdf.

Anonymous:1892:BBS

[2458] Anonymous. A bamboo bridge in Sikkim, India. Scientific American, 66(9):137, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
137.pdf.

Anonymous:1892:BNBf

[2459] Anonymous. Books: New books and publications. Scientific American, 66(9):138, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
138a.pdf.

Anonymous:1892:RPIi

[2460] Anonymous. Recently patented inventions. Scientific American, 66(9):138, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
138.pdf.

Anonymous:1892:BPh

[2461] Anonymous. Business and personal. Scientific American, 66(9): 139, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
139.pdf.

Anonymous:1892:Ih

[2462] Anonymous. To inventors. Scientific American, 66(9):139, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n9/pdf/scientificamerican02271892-139b.pdf.

Anonymous:1892:NQi

[2463] Anonymous. Notes and queries. Scientific American, 66(9):139, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n9/pdf/scientificamerican02271892-139a.pdf.

Anonymous:1892:IIi

[2464] Anonymous. Index of inventions. Scientific American, 66(9): 139–140, February 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n9/pdf/scientificamerican02271892-
139c.pdf.

Anonymous:1892:ELS

[2465] Anonymous. An electrical life saver wanted. Scientific American, 66(10):144, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
144.pdf.

Anonymous:1892:RIL

[2466] Anonymous. Reform of the immigration laws. Scientific American, 66(10):144–145, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
144a.pdf.

Anonymous:1892:REJ

[2467] Anonymous. The recent earthquake in Japan. Scientific American, 66(10):146, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
146.pdf.

Anonymous:1892:TSb

[2468] Anonymous. A twin staircase. Scientific American, 66(10):146, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-146a.pdf.

Anonymous:1892:IOCa

[2469] Anonymous. An improved Ore crusher. Scientific American, 66(10):147, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-147.pdf.

Anonymous:1892:NFW

[2470] Anonymous. A new feed water heater and purifier. Scientific American, 66(10):147, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
147a.pdf.

Anonymous:1892:Ch

[2471] Anonymous. Correspondence. Scientific American, 66(10):148, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-148.pdf.

Anonymous:1892:DEA

[2472] Anonymous. Destruction of emanations of all kinds without the use of chemicals. Scientific American, 66(10):148, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n10/pdf/scientificamerican03051892-148a.pdf.

Anonymous:1892:UWPa

[2473] Anonymous. Utilizing the water power of Niagara Falls. Scientific American, 66(10):149, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
149.pdf.

Anonymous:1892:MRH

[2474] Anonymous. A machine to roll hollow rods. Scientific American, 66(10):150, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
150.pdf.

Anonymous:1892:RMI

[2475] Anonymous. Rain making in India. Scientific American, 66 (10):150–151, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
150a.pdf.

Bain:1892:WE

[2476] Foree Bain. What is electricity? Scientific American, 66(10):152–153, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-152.pdf.

Sloane:1892:EIG

[2477] T. O’Conor Sloane. Experiments with illuminating gas. Scientific American, 66(10):153, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
153.pdf.

Anonymous:1892:BNBg

[2478] Anonymous. Books: New books and publications. Scientific American, 66(10):154, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
154a.pdf.

Anonymous:1892:RPIj

[2479] Anonymous. Recently patented inventions. Scientific American, 66(10):154, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n10/pdf/scientificamerican03051892-
154.pdf.

Anonymous:1892:BPi

[2480] Anonymous. Business and personal. Scientific American, 66(10):155, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-155.pdf.

Anonymous:1892:Ij

[2481] Anonymous. To inventors. Scientific American, 66(10):155, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-155b.pdf.

Anonymous:1892:NQj

[2482] Anonymous. Notes and queries. Scientific American, 66(10):155, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-155a.pdf.

Anonymous:1892:IIj

[2483] Anonymous. Index of inventions. Scientific American, 66(10):155–156, March 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n10/pdf/scientificamerican03051892-155c.pdf.

Anonymous:1892:OTF

[2484] Anonymous. An old time fast locomotive. Scientific American, 66 (11):159–160, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
159.pdf.

Anonymous:1892:PBP

[2485] Anonymous. Patent bills and petitions now before Congress. Scientific American, 66(11):160–161, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
160.pdf.

Anonymous:1892:ABF

[2486] Anonymous. An automatic boiler feed. Scientific American, 66(11):162, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-162a.pdf.

Anonymous:1892:ICMa

[2487] Anonymous. An improved coal mining machine. Scientific American, 66(11):162, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
162b.pdf.

Anonymous:1892:IPBa

[2488] Anonymous. An improved propeller blade. Scientific American, 66(11):162, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
162.pdf.

Anonymous:1892:IRSa

[2489] Anonymous. An improved railroad signal. Scientific American, 66(11):163, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
163.pdf.

Anonymous:1892:TF

[2490] Anonymous. Terracing in the foothills. Scientific American, 66 (11):163–164, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
163a.pdf.

Anonymous:1892:FKC

[2491] Anonymous. A flour or kitchen cabinet. Scientific American, 66(11):164, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-164.pdf.

Sloane:1892:LB

[2492] T. O’conor Sloane. A lecture barometer. Scientific American, 66(11):164, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
164a.pdf.

Anonymous:1892:Ci

[2493] Anonymous. Correspondence. Scientific American, 66(11):165, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-165.pdf.

Anonymous:1892:DP

[2494] Anonymous. Distance photography. Scientific American, 66(11):166, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-166.pdf.

Smith:1892:AAA

[2495] J. Henry Smith. Apparatus for aerial ascensions. Scientific American, 66(11):167–168, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
167.pdf.

Anonymous:1892:BNBh

[2496] Anonymous. Books: New books and publications. Scientific American, 66(11):169, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
169a.pdf.

Anonymous:1892:RPIk 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n11/pdf/scientificamerican03121892-
169.pdf.

Anonymous:1892:BPj

[2498] Anonymous. Business and personal. Scientific American, 66(11):170, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-170.pdf.

Anonymous:1892:NQk

[2499] Anonymous. Notes and queries. Scientific American, 66(11):170–171, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-170a.pdf.

Anonymous:1892:Ik

[2500] Anonymous. To inventors. Scientific American, 66(11):171, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-171.pdf.

Anonymous:1892:IIk

[2501] Anonymous. Index of inventions. Scientific American, 66(11):171–172, March 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n11/pdf/scientificamerican03121892-171a.pdf.

Anonymous:1892:GGPa

[2502] Anonymous. The Gatling gun for police patrol service. Scientific American, 66(12):175, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
175.pdf.

Anonymous:1892:ASW

[2503] Anonymous. Acetate of soda warmers and caustic soda steam boilers. Scientific American, 66(12):176–177, March 19, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n12/pdf/scientificamerican03191892-176.pdf.

Anonymous:1892:STE

[2504] Anonymous. A speaking tube and eariphone. Scientific American, 66(12):178, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
178.pdf.

Anonymous:1892:SCT

[2505] Anonymous. Surface condensing triple expansion mill engines. Scientific American, 66(12):178–179, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
178a.pdf.

Anonymous:1892:CAa

[2506] Anonymous. A curious automaton. Scientific American, 66(12):179, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-179a.pdf.

Anonymous:1892:GSP

[2507] Anonymous. A grain scourer, polisher, and separator. Scientific American, 66(12):179, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
179.pdf.

Anonymous:1892:IL

[2508] Anonymous. An improved lubricator. Scientific American, 66(12):179, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-179b.pdf.

Anonymous:1892:IQM

[2509] Anonymous. An improved quilting machine. Scientific American, 66(12):180, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
180a.pdf.

Anonymous:1892:IRW

[2510] Anonymous. An improved rain water cut-off. Scientific American, 66(12):180, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
180.pdf.

Anonymous:1892:SVB

[2511] Anonymous. Sanitary, ventilated boots and shoes. Scientific American, 66(12):180, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
180b.pdf.

Anonymous:1892:Cj

[2512] Anonymous. Correspondence. Scientific American, 66(12):181, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-181.pdf.

Anonymous:1892:TSH

[2513] Anonymous. Thomas sterry hunt. Scientific American, 66(12):182, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-182.pdf.

Anonymous:1892:LSW

[2514] Anonymous. The largest and the smallest water wheel. Scientific American, 66(12):183, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
183.pdf.

Peckham:1892:ELM

[2515] W. C. Peckham. Electric light for magic lantern. Scientific American, 66(12):183–184, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
183a.pdf.

Anonymous:1892:RPIl

[2516] Anonymous. Recently patented inventions. Scientific American, 66(12):185, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
185.pdf.

Anonymous:1892:BNBi

[2517] Anonymous. Books: New books and publications. Scientific American, 66(12):185–186, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n12/pdf/scientificamerican03191892-
185a.pdf.

Anonymous:1892:BPk

[2518] Anonymous. Business and personal. Scientific American, 66(12):186, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-186.pdf.

Anonymous:1892:NQl

[2519] Anonymous. Notes and queries. Scientific American, 66(12):186–187, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-186a.pdf.

Anonymous:1892:Im

[2520] Anonymous. To inventors. Scientific American, 66(12):187, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-187.pdf.

Anonymous:1892:IIl

[2521] Anonymous. Index of inventions. Scientific American, 66(12):187–188, March 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n12/pdf/scientificamerican03191892-187a.pdf.

Anonymous:1892:CICa

[2522] Anonymous. Congressional inquiries concerning the Patent Office. Scientific American, 66(13):192, March 26, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n13/pdf/scientificamerican03261892-192.pdf.

Anonymous:1892:RLT

[2523] Anonymous. Recent lymph tests and examinations for tubercular disease in cattle. Scientific American, 66(13):192–194, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-192a.pdf.

Anonymous:1892:IDCb

[2524] Anonymous. An improved dump cart. Scientific American, 66(13):194, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-194b.pdf.

Anonymous:1892:IIP

[2525] Anonymous. An improved ice plow. Scientific American, 66(13):194, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-194.pdf.

Anonymous:1892:IREa

[2526] Anonymous. An improved rotary engine. Scientific American, 66(13):194, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
194a.pdf.

Anonymous:1892:IWR

[2527] Anonymous. An improved water register. Scientific American, 66(13):195, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
195.pdf.

Anonymous:1892:MTE

[2528] Anonymous. Mr. Tesla’s experiments on alternating currents of great frequency. Scientific American, 66(13):195–197, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-195a.pdf.

Anonymous:1892:IFa

[2529] Anonymous. An improved furnace. Scientific American, 66(13):197, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-197a.pdf.

Anonymous:1892:IFM

[2530] Anonymous. An improved fence machine. Scientific American, 66(13):197, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
197.pdf.

Anonymous:1892:WUC

[2531] Anonymous. The Western Union Central Telegraph Office and Plant in New York City. Scientific American, 66(13):198–199, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-198.pdf.

Hopkins:1892:EA

[2532] Geo. M. Hopkins. An experiment in acoustics. Scientific American, 66(13):199, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
199.pdf.

Anonymous:1892:DE

[2533] Anonymous. Distinguished electricians. Scientific American, 66 (13):199–200, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
199a.pdf.

Anonymous:1892:DCa

[2534] Anonymous. The dedication ceremonies. Scientific American, 66(13):200, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
200a.pdf.

Anonymous:1892:ICS

[2535] Anonymous. An improved Carpenter’s square. Scientific American, 66(13):200, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
200.pdf.

Anonymous:1892:BPl

[2536] Anonymous. Business and personal. Scientific American, 66(13):201, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-201a.pdf.

Anonymous:1892:RPIm

[2537] Anonymous. Recently patented inventions. Scientific American, 66(13):201, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
201.pdf.

Anonymous:1892:NQm

[2538] Anonymous. Notes and queries. Scientific American, 66(13):201–203, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-201b.pdf.

Anonymous:1892:BNBj

[2539] Anonymous. Books: New books and publications. Scientific American, 66(13):203, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n13/pdf/scientificamerican03261892-
203.pdf.

Anonymous:1892:In

[2540] Anonymous. To inventors. Scientific American, 66(13):203, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-203a.pdf.

Anonymous:1892:IIm

[2541] Anonymous. Index of inventions. Scientific American, 66(13):203–204, March 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n13/pdf/scientificamerican03261892-203b.pdf.

Anonymous:1892:HWC

[2542] Anonymous. The Hackensack Water Company. Scientific American, 66(14):207, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
207.pdf.

Anonymous:1892:GNW

[2543] Anonymous. The grounding of the neutral wire in the threewire system. Scientific American, 66(14):208, April 2, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n14/pdf/scientificamerican04021892-208.pdf.

Anonymous:1892:PPAa

[2544] Anonymous. Position of the planets in April. Scientific American, 66(14):209, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
209.pdf.

Anonymous:1892:IWD

[2545] Anonymous. An improved wool drying machine. Scientific American, 66(14):210, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
210.pdf.

Anonymous:1892:TWF

[2546] Anonymous. Twelve-wheeled freight locomotives. Scientific American, 66(14):210, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
210a.pdf.

Anonymous:1892:ICMb

[2547] Anonymous. An improved cigar measure. Scientific American, 66(14):211, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
211.pdf.

Anonymous:1892:IER

[2548] Anonymous. An improved elevated railway. Scientific American, 66(14):211, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
211b.pdf.

Anonymous:1892:IFW

[2549] Anonymous. An improved feed water heater. Scientific American, 66(14):211, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
211a.pdf.

Anonymous:1892:PNa

[2550] Anonymous. Photographic notes. Scientific American, 66(14):211, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n14/pdf/scientificamerican04021892-211c.pdf.

Anonymous:1892:APR

[2551] Anonymous. An automatic pump regulator. Scientific American, 66(14):211–212, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
211d.pdf.

Anonymous:1892:Ck

[2552] Anonymous. Correspondence. Scientific American, 66(14):213, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n14/pdf/scientificamerican04021892-213.pdf.

Anonymous:1892:DSL

[2553] Anonymous. A draw shell lime kiln. Scientific American, 66 (14):215–217, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
215.pdf.

Anonymous:1892:CJV

[2554] Anonymous. Charles J. Van Depoele. Scientific American, 66(14):217, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n14/pdf/scientificamerican04021892-217.pdf.

Anonymous:1892:RPIn

[2555] Anonymous. Recently patented inventions. Scientific American, 66(14):218, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
218.pdf.

Anonymous:1892:BPBb

[2556] Anonymous. Business and personal: Business and personal. Scientific American, 66(14):219, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n14/pdf/scientificamerican04021892-
219.pdf.

Anonymous:1892:Io

[2557] Anonymous. To inventors. Scientific American, 66(14):219, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n14/pdf/scientificamerican04021892-219b.pdf.

Anonymous:1892:NQn

[2558] Anonymous. Notes and queries. Scientific American, 66(14):219, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n14/pdf/scientificamerican04021892-219a.pdf.

Anonymous:1892:IIn

[2559] Anonymous. Index of inventions. Scientific American, 66(14):219–220, April 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n14/pdf/scientificamerican04021892-219c.pdf.

Anonymous:1892:PDA

[2560] Anonymous. Proposed discrimination against foreign inventors. Scientific American, 66(15):224, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
224.pdf.

Anonymous:1892:APL

[2561] Anonymous. Artificial propagation of lobsters. Scientific American, 66(15):224–225, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
224a.pdf.

Anonymous:1892:ICO

[2562] Anonymous. An improved crude oil burner. Scientific American, 66(15):226, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
226b.pdf.

Anonymous:1892:IRSb

[2563] Anonymous. An improved rifle sight. Scientific American, 66(15):226, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-226.pdf.

Sloane:1892:EC

[2564] T. O’conor Sloane. Experiment in capillarity. Scientific American, 66(15):226, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
226a.pdf.

Baldwin:1892:MGA

[2565] S. P. Baldwin. Muir Glacier, Alaska. Scientific American, 66 (15):227–228, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
227.pdf.

Anonymous:1892:LDS

[2566] Anonymous. The Lovell diamond safety bicycle. Scientific American, 66(15):228, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
228.pdf.

Anonymous:1892:Cl

[2567] Anonymous. Correspondence. Scientific American, 66(15):229, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-229a.pdf.

Anonymous:1892:TM

[2568] Anonymous. “A trip to the Moon”. Scientific American, 66(15):229, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-229.pdf.

Anonymous:1892:ICd

[2569] Anonymous. Inaudi, the calculator. Scientific American, 66(15):230, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-230.pdf.

Wood:1892:HP

[2570] R. W. Wood. A hydrostatic paradox. Scientific American, 66 (15):230–231, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
230a.pdf.

Anonymous:1892:RTC

[2571] Anonymous. Rope traction for cable railroad plants. Scientific American, 66(15):231–232, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
231.pdf.

Anonymous:1892:SW

[2572] Anonymous. Sereno Watson. Scientific American, 66(15):233, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-233.pdf.

Anonymous:1892:BNBk

[2573] Anonymous. Books: New books and publications. Scientific American, 66(15):234, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
234a.pdf.

Anonymous:1892:BPm

[2574] Anonymous. Business and personal. Scientific American, 66(15):234, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-234b.pdf.

Anonymous:1892:RPIo

[2575] Anonymous. Recently patented inventions. Scientific American, 66(15):234, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n15/pdf/scientificamerican04091892-
234.pdf.

Anonymous:1892:NQo

[2576] Anonymous. Notes and queries. Scientific American, 66(15):234–235, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-234c.pdf.

Anonymous:1892:Ip

[2577] Anonymous. To inventors. Scientific American, 66(15):235, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-235.pdf.

Anonymous:1892:IIo

[2578] Anonymous. Index of inventions. Scientific American, 66(15):235–236, April 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n15/pdf/scientificamerican04091892-235a.pdf.

Anonymous:1892:RCC

[2579] Anonymous. A remarkable chemical compound-nickel carbon oxide. Scientific American, 66(16):240, April 16, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n16/pdf/scientificamerican04161892-240.pdf.

Anonymous:1892:MML

[2580] Anonymous. Mrs. Maria Louisa Pike. Scientific American, 66 (16):240–241, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
240a.pdf.

Anonymous:1892:EAH

[2581] Anonymous. An easily adjustable harness saddle. Scientific American, 66(16):242, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
242a.pdf.

Anonymous:1892:IGC

[2582] Anonymous. An improved gun cleaning device. Scientific American, 66(16):242, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
242c.pdf.

Anonymous:1892:IPBb

[2583] Anonymous. An improved poultry brooder. Scientific American, 66(16):242, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
242.pdf.

Anonymous:1892:NIS

[2584] Anonymous. A nine-inch screw cutting lathe. Scientific American, 66(16):242, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
242b.pdf.

Anonymous:1892:ICCb

[2585] Anonymous. An improved cash carrier. Scientific American, 66(16):243, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n16/pdf/scientificamerican04161892-243b.pdf.

Anonymous:1892:IRBa

[2586] Anonymous. Improvements in rubber belts. Scientific American, 66(16):243, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
243.pdf.

Anonymous:1892:ITS

[2587] Anonymous. An intrenching tool for soldiers. Scientific American, 66(16):243, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
243a.pdf.

Anonymous:1892:IBBa

[2588] Anonymous. An improved bridle bit. Scientific American, 66 (16):243–244, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
243c.pdf.

Anonymous:1892:DRP

[2589] Anonymous. Decisions relating to patents. Scientific American, 66(16):244, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
244.pdf.

Anonymous:1892:LRM

[2590] Anonymous. A large refrigerating machine. Scientific American, 66(16):245, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
245.pdf.

Anonymous:1892:KMV

[2591] Anonymous. 250-kilowatt “Mordey–Victoria” alternator. Scientific American, 66(16):246–247, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
246.pdf.

Anonymous:1892:RFB

[2592] Anonymous. A rail fence bicycle railway. Scientific American, 66(16):247, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
247.pdf.

Anonymous:1892:GSSb

[2593] Anonymous. The Great Sun Spot of February, 1892. Scientific American, 66(16):247–248, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
247a.pdf.

Anonymous:1892:PMR

[2594] Anonymous. A practical metallic railroad tie. Scientific American, 66(16):248, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
248.pdf.

Anonymous:1892:RPIp

[2595] Anonymous. Recently patented inventions. Scientific American, 66(16):249, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n16/pdf/scientificamerican04161892-
249.pdf.

Anonymous:1892:BPn

[2596] Anonymous. Business and personal. Scientific American, 66(16):250, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n16/pdf/scientificamerican04161892-250.pdf.

Anonymous:1892:NQp

[2597] Anonymous. Notes and queries. Scientific American, 66(16):250–251, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n16/pdf/scientificamerican04161892-250a.pdf.

Anonymous:1892:Iq

[2598] Anonymous. To inventors. Scientific American, 66(16):251, April 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n16/pdf/scientificamerican04161892-251.pdf.

Anonymous:1892:IIq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n16/pdf/scientificamerican04161892-251a.pdf.

Anonymous:1892:EM

[2600] Anonymous. Engines of the Monterey. Scientific American, 66(17):255, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-255.pdf.

Anonymous:1892:PCR

[2601] Anonymous. A proposed Congressional resolution relating to patents. Scientific American, 66(17):256, April 23, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n17/pdf/scientificamerican04231892-256.pdf.

Anonymous:1892:SAOa

[2602] Anonymous. The seven ages of our world. Scientific American, 66(17):256–257, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n17/pdf/scientificamerican04231892-
256a.pdf.

Anonymous:1892:EL

[2603] Anonymous. Explosion of a locomotive. Scientific American, 66 (17):258, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n17/pdf/scientificamerican04231892-
258a.pdf.

Anonymous:1892:IHS

[2604] Anonymous. An improved hay stacker. Scientific American, 66(17):258, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-258.pdf.

Anonymous:1892:NFL

[2605] Anonymous. New flash lamp. Scientific American, 66(17):258, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-258b.pdf.

Anonymous:1892:IMHa

[2606] Anonymous. An improved magazine hand camera. Scientific American, 66(17):259–260, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n17/pdf/scientificamerican04231892-
259.pdf.

Anonymous:1892:CAB

[2607] Anonymous. A combination account and billing book. Scientific American, 66(17):260, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n17/pdf/scientificamerican04231892-
260.pdf.

Anonymous:1892:DIB

[2608] Anonymous. A desk ink bottle. Scientific American, 66(17):260–261, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-260a.pdf.

Anonymous:1892:EG

[2609] Anonymous. The “electric” girl. Scientific American, 66(17):262, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-262.pdf.

Anonymous:1892:SCR

[2610] Anonymous. The Stockton, California, Race Track. Scientific American, 66(17):263–264, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n17/pdf/scientificamerican04231892-
263.pdf.

Anonymous:1892:ESP

[2611] Anonymous. Electric shooting pull. Scientific American, 66(17):265, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-265.pdf.

Anonymous:1892:BNBl 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n17/pdf/scientificamerican04231892-
266a.pdf.

Anonymous:1892:RPIq

[2613] Anonymous. Recently patented inventions. Scientific American, 66(17):266, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n17/pdf/scientificamerican04231892-
266.pdf.

Anonymous:1892:BPo

[2614] Anonymous. Business and personal. Scientific American, 66(17):267, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-267.pdf.

Anonymous:1892:Ir

[2615] Anonymous. To inventors. Scientific American, 66(17):267, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-267b.pdf.

Anonymous:1892:NQq

[2616] Anonymous. Notes and queries. Scientific American, 66(17):267, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-267a.pdf.

Anonymous:1892:IIr

[2617] Anonymous. Index of inventions. Scientific American, 66(17):267–268, April 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n17/pdf/scientificamerican04231892-267c.pdf.

Anonymous:1892:EDS

[2618] Anonymous. Ericsson’s destroyer and its submarine gun. Scientific American, 66(18):272, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n18/pdf/scientificamerican04301892-
272a.pdf.

Anonymous:1892:ENY

[2619] Anonymous. Exhibition of the New York microscopical society. Scientific American, 66(18):272, April 30, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n18/pdf/scientificamerican04301892-272.pdf.

Anonymous:1892:PPMb

[2620] Anonymous. Position of the planets in May. Scientific American, 66(18):272–273, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n18/pdf/scientificamerican04301892-
272b.pdf.

Anonymous:1892:IGE

[2621] Anonymous. An improved gas engine. Scientific American, 66(18):274, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-274.pdf.

Anonymous:1892:ICDa

[2622] Anonymous. An improved car door. Scientific American, 66(18):275, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-275.pdf.

Anonymous:1892:MDS

[2623] Anonymous. A motor to drive small machines. Scientific American, 66(18):275, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n18/pdf/scientificamerican04301892-
275b.pdf.

Anonymous:1892:THW

[2624] Anonymous. A trace-holding whiffletree attachment. Scientific American, 66(18):275, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n18/pdf/scientificamerican04301892-
275a.pdf.

Anonymous:1892:Cm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-276.pdf.

Anonymous:1892:GM

[2626] Anonymous. The Grant Monument. Scientific American, 66(18):277, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-277.pdf.

Anonymous:1892:ITa

[2627] Anonymous. Interesting tricks. Scientific American, 66(18):279, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-279.pdf.

Anonymous:1892:LM

[2628] Anonymous. Logging in Minnesota. Scientific American, 66(18):279, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-279a.pdf.

Anonymous:1892:BFC

[2629] Anonymous. The British first class cruiser Edgar. Scientific American, 66(18):279–281, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n18/pdf/scientificamerican04301892-
279b.pdf.

Anonymous:1892:NTa

[2630] Anonymous. Novel toys. Scientific American, 66(18):281, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-281.pdf.

Anonymous:1892:BNBm

[2631] Anonymous. Books: New books and publications. Scientific American, 66(18):282, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n18/pdf/scientificamerican04301892-
282a.pdf.

Anonymous:1892:RPIr

[2632] Anonymous. Recently patented inventions. Scientific American, 66(18):282, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n18/pdf/scientificamerican04301892-
282.pdf.

Anonymous:1892:BPp

[2633] Anonymous. Business and personal. Scientific American, 66(18):283, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-283.pdf.

Anonymous:1892:NQr

[2634] Anonymous. Notes and queries. Scientific American, 66(18):283, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-283a.pdf.

Anonymous:1892:IIs

[2635] Anonymous. Index of inventions. Scientific American, 66(18):283–284, April 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n18/pdf/scientificamerican04301892-283b.pdf.

Anonymous:1892:NGFa

[2636] Anonymous. The Naval Gun Factory, Washington. Scientific American, 66(19):287, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
287.pdf.

Anonymous:1892:SVS

[2637] Anonymous. Steadying vessels at sea. Scientific American, 66(19):288, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-288.pdf.

Anonymous:1892:CW

[2638] Anonymous. California wines. Scientific American, 66(19):288–289, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-288a.pdf.

Anonymous:1892:INLa

[2639] Anonymous. An improved nut lock. Scientific American, 66(19):290, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-290a.pdf.

Anonymous:1892:ITP

[2640] Anonymous. An improved trunk protector. Scientific American, 66(19):290, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
290b.pdf.

Anonymous:1892:PIS

[2641] Anonymous. Preston’s improvement in stockings. Scientific American, 66(19):290, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
290.pdf.

Anonymous:1892:AS

[2642] Anonymous. An aerial ship. Scientific American, 66(19):291, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-291.pdf.

Anonymous:1892:IISa

[2643] Anonymous. Ingleton’s improved steam plow. Scientific American, 66(19):291–292, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
291a.pdf.

Anonymous:1892:BSS

[2644] Anonymous. A bed Spring and support. Scientific American, 66(19):292, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
292.pdf.

Anonymous:1892:NTb

[2645] Anonymous. Novel toys. Scientific American, 66(19):292, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-292a.pdf.

Anonymous:1892:Cn

[2646] Anonymous. Correspondence. Scientific American, 66(19):293, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-293.pdf.

Anonymous:1892:NCR

[2647] Anonymous. The new cruiser Raleigh. Scientific American, 66 (19):295–296, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
295.pdf.

Anonymous:1892:BNBn

[2648] Anonymous. Books: New books and publications. Scientific American, 66(19):297, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
297a.pdf.

Anonymous:1892:RPIs

[2649] Anonymous. Recently patented inventions. Scientific American, 66(19):297, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n19/pdf/scientificamerican05071892-
297.pdf.

Anonymous:1892:BPq

[2650] Anonymous. Business and personal. Scientific American, 66(19):298, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-298.pdf.

Anonymous:1892:NQs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-298a.pdf.

Anonymous:1892:Is

[2652] Anonymous. To inventors. Scientific American, 66(19):299, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-299.pdf.

Anonymous:1892:IIt

[2653] Anonymous. Index of inventions. Scientific American, 66(19):299–300, May 7, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n19/pdf/scientificamerican05071892-299a.pdf.

Anonymous:1892:VSV

[2654] Anonymous. Vibrations of steam vessels. Scientific American, 66 (20):304–305, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n20/pdf/scientificamerican05141892-
304.pdf.

Anonymous:1892:PUL

[2655] Anonymous. Purdue University, Lafayette, Ind. Scientific American, 66(20):306, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n20/pdf/scientificamerican05141892-
306.pdf.

Anonymous:1892:CLL

[2656] Anonymous. Carrying a life line ashore by a kite. Scientific American, 66(20):307–308, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n20/pdf/scientificamerican05141892-
307.pdf.

Anonymous:1892:Co

[2657] Anonymous. Correspondence. Scientific American, 66(20):309, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n20/pdf/scientificamerican05141892-309.pdf.

Anonymous:1892:NDE

[2658] Anonymous. New disappearing Elswick gun mounting. Scientific American, 66(20):310–311, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n20/pdf/scientificamerican05141892-
310.pdf.

Anonymous:1892:GGPb

[2659] Anonymous. A girder gas pipe. Scientific American, 66(20):311, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n20/pdf/scientificamerican05141892-311.pdf.

Anonymous:1892:APA

[2660] Anonymous. The acoustic properties of aluminum. Scientific American, 66(20):312, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n20/pdf/scientificamerican05141892-
312.pdf.

Anonymous:1892:IFCb

[2661] Anonymous. An improved flue cleaner. Scientific American, 66(20):312, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n20/pdf/scientificamerican05141892-312a.pdf.

Anonymous:1892:RPIt

[2662] Anonymous. Recently patented inventions. Scientific American, 66(20):313, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n20/pdf/scientificamerican05141892-
313.pdf.

Anonymous:1892:BNBo

[2663] Anonymous. Books: New books and publications. Scientific American, 66(20):313–314, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n20/pdf/scientificamerican05141892-
313a.pdf.

Anonymous:1892:BPr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n20/pdf/scientificamerican05141892-314.pdf.

Anonymous:1892:NQt

[2665] Anonymous. Notes and queries. Scientific American, 66(20):314–315, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n20/pdf/scientificamerican05141892-314a.pdf.

Anonymous:1892:Iu

[2666] Anonymous. To inventors. Scientific American, 66(20):315, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n20/pdf/scientificamerican05141892-315.pdf.

Anonymous:1892:IIu

[2667] Anonymous. Index of inventions. Scientific American, 66(20):315–316, May 14, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n20/pdf/scientificamerican05141892-315a.pdf.

Anonymous:1892:FSP

[2668] Anonymous. Fast ships in prospect. Scientific American, 66 (21):320–322, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
320.pdf.

Anonymous:1892:IPBc

[2669] Anonymous. An improved paint brush bridle. Scientific American, 66(21):322, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
322.pdf.

Anonymous:1892:IRSc

[2670] Anonymous. An improved railway signal. Scientific American, 66(21):322, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
322a.pdf.

Brown:1892:MSF

[2671] S. J. Brown and J. A. Hoogewerff. The magnetic storm of February 13, 1892. Scientific American, 66(21):323–324, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n21/pdf/scientificamerican05211892-323.pdf.

Anonymous:1892:Cp

[2672] Anonymous. Correspondence. Scientific American, 66(21):325, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n21/pdf/scientificamerican05211892-325a.pdf.

Anonymous:1892:ISL

[2673] Anonymous. Iron and steel in large buildings. Scientific American, 66(21):325, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
325.pdf.

Anonymous:1892:ILS

[2674] Anonymous. Improved locomotive steam crane. Scientific American, 66(21):326, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
326.pdf.

Anonymous:1892:ECb

[2675] Anonymous. An electric Chime. Scientific American, 66(21):327–328, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n21/pdf/scientificamerican05211892-327.pdf.

Hovey:1892:FMB

[2676] H. C. Hovey. The Florence (Mass.) brush industry. Scientific American, 66(21):328, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
328.pdf.

Anonymous:1892:GSC

[2677] Anonymous. A gasoline steam carriage. Scientific American, 66(21):329, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n21/pdf/scientificamerican05211892-329.pdf.

Anonymous:1892:BNBp

[2678] Anonymous. Books: New books and publications. Scientific American, 66(21):330, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
330a.pdf.

Anonymous:1892:RPIu

[2679] Anonymous. Recently patented inventions. Scientific American, 66(21):330, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n21/pdf/scientificamerican05211892-
330.pdf.

Anonymous:1892:BPs

[2680] Anonymous. Business and personal. Scientific American, 66(21):331, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n21/pdf/scientificamerican05211892-331.pdf.

Anonymous:1892:Iv

[2681] Anonymous. To inventors. Scientific American, 66(21):331, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n21/pdf/scientificamerican05211892-331b.pdf.

Anonymous:1892:NQu

[2682] Anonymous. Notes and queries. Scientific American, 66(21):331, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n21/pdf/scientificamerican05211892-331a.pdf.

Anonymous:1892:IIv

[2683] Anonymous. Index of inventions. Scientific American, 66(21):331–332, May 21, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n21/pdf/scientificamerican05211892-331c.pdf.

Anonymous:1892:MTS

[2684] Anonymous. The manufacture of tin at St. Louis. Scientific American, 66(22):335, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
335.pdf.

Anonymous:1892:NIN

[2685] Anonymous. New inventions needed for working steamships. Scientific American, 66(22):336, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
336.pdf.

Anonymous:1892:NSSa

[2686] Anonymous. The new steel steamers of the Providence Line. Scientific American, 66(22):336, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
336a.pdf.

Anonymous:1892:LMD

[2687] Anonymous. The largest masonry dam in the world. Scientific American, 66(22):336–337, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
336b.pdf.

Anonymous:1892:ISE

[2688] Anonymous. An improved steam engine governor. Scientific American, 66(22):338, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
338.pdf.

Anonymous:1892:PT

[2689] Anonymous. The penduline titmouse. Scientific American, 66(22):338, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n22/pdf/scientificamerican05281892-338a.pdf.

Anonymous:1892:SHB

[2690] Anonymous. A simple horse blanket fastener. Scientific American, 66(22):338, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
338b.pdf.

Anonymous:1892:TCL

[2691] Anonymous. A train of compound locomotives. Scientific American, 66(22):339, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
339a.pdf.

Anonymous:1892:TKA

[2692] Anonymous. Tea kettle andiron. Scientific American, 66(22):339, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n22/pdf/scientificamerican05281892-339.pdf.

Anonymous:1892:TAD

[2693] Anonymous. Toy automatic distributer of liquids. Scientific American, 66(22):339–340, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
339b.pdf.

Anonymous:1892:IPWb

[2694] Anonymous. An improved pocket wire gauge. Scientific American, 66(22):340, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
340.pdf.

Anonymous:1892:Cq

[2695] Anonymous. Correspondence. Scientific American, 66(22):341–342, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n22/pdf/scientificamerican05281892-341.pdf.

Anonymous:1892:CMN

[2696] Anonymous. Columbus Monument for New York. Scientific American, 66(22):342–345, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
342.pdf.

Anonymous:1892:BNBq

[2697] Anonymous. Books: New books and publications. Scientific American, 66(22):346, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
346a.pdf.

Anonymous:1892:RPIv

[2698] Anonymous. Recently patented inventions. Scientific American, 66(22):346, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n22/pdf/scientificamerican05281892-
346.pdf.

Anonymous:1892:BPt

[2699] Anonymous. Business and personal. Scientific American, 66(22):347, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n22/pdf/scientificamerican05281892-347.pdf.

Anonymous:1892:NQv

[2700] Anonymous. Notes and queries. Scientific American, 66(22):347, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n22/pdf/scientificamerican05281892-347a.pdf.

Anonymous:1892:IIw

[2701] Anonymous. Index of inventions. Scientific American, 66(22):347–348, May 28, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n22/pdf/scientificamerican05281892-347b.pdf.

Anonymous:1892:IPBd

[2702] Anonymous. Improved plastic brickmaking machine. Scientific American, 66(23):351, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
351.pdf.

Anonymous:1892:PFT

[2703] Anonymous. Preservation of flowers in their natural form and color. Scientific American, 66(23):352, June 4, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n23/pdf/scientificamerican06041892-352a.pdf.

Anonymous:1892:SRH

[2704] Anonymous. Sudden rupture of a 1500 horse power engine. Scientific American, 66(23):352, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
352.pdf.

Anonymous:1892:PPJb

[2705] Anonymous. Position of the planets in June. Scientific American, 66(23):353, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
353.pdf.

Anonymous:1892:CFR

[2706] Anonymous. A convenient form of resonance tube for determining the velocity of sound. Scientific American, 66(23):354, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n23/pdf/scientificamerican06041892-354.pdf.

Anonymous:1892:ITE

[2707] Anonymous. An improved traction engine. Scientific American, 66(23):354, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
354a.pdf.

Anonymous:1892:ESBa

[2708] Anonymous. Experiments with soap bubbles. Scientific American, 66(23):355, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
355.pdf.

Anonymous:1892:SHP

[2709] Anonymous. A sixty horse power gas engine. Scientific American, 66(23):355, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
355a.pdf.

Anonymous:1892:PNb

[2710] Anonymous. Photographic notes. Scientific American, 66(23):356–357, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n23/pdf/scientificamerican06041892-356.pdf.

Anonymous:1892:EFA

[2711] Anonymous. An economical fire alarm. Scientific American, 66(23):358, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n23/pdf/scientificamerican06041892-358.pdf.

Anonymous:1892:PPI

[2712] Anonymous. The pythons of the Philippine Islands. Scientific American, 66(23):359, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
359.pdf.

Anonymous:1892:MPD

[2713] Anonymous. Magic photographs developed by smoke. Scientific American, 66(23):359–360, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
359a.pdf.

Anonymous:1892:NBPb

[2714] Anonymous. New books and publications. Scientific American, 66(23):361, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
361a.pdf.

Anonymous:1892:RPIw (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n23/pdf/scientificamerican06041892-
361.pdf.

Anonymous:1892:BPu

[2716] Anonymous. Business and personal. Scientific American, 66(23):362, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n23/pdf/scientificamerican06041892-362.pdf.

Anonymous:1892:NQw

[2717] Anonymous. Notes and queries. Scientific American, 66(23):362–363, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n23/pdf/scientificamerican06041892-362a.pdf.

Anonymous:1892:Ix

[2718] Anonymous. To inventors. Scientific American, 66(23):363, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n23/pdf/scientificamerican06041892-363.pdf.

Anonymous:1892:IIx

[2719] Anonymous. Index of inventions. Scientific American, 66(23):363–364, June 4, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n23/pdf/scientificamerican06041892-363a.pdf.

Anonymous:1892:ANP

[2720] Anonymous. Another new Patent Bill. Scientific American, 66 (24):368–369, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
368.pdf.

Anonymous:1892:EIB

[2721] Anonymous. An electric insulator for boilers and condensers. Scientific American, 66(24):370, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
370a.pdf.

Anonymous:1892:HAC

[2722] Anonymous. An hydraulic air compressor. Scientific American, 66(24):370, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
370.pdf.

Anonymous:1892:ISG

[2723] Anonymous. An improved shingling gauge. Scientific American, 66(24):370, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
370b.pdf.

Anonymous:1892:IPWc

[2724] Anonymous. An improved pipe wrench. Scientific American, 66 (24):371, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
371.pdf.

Anonymous:1892:SDU

[2725] Anonymous. Siphon for domestic uses. Scientific American, 66(24):371, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n24/pdf/scientificamerican06111892-371a.pdf.

Anonymous:1892:SAH

[2726] Anonymous. A sacramental altar for home use. Scientific American, 66(24):371–372, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
371b.pdf.

Anonymous:1892:LI

[2727] Anonymous. Lighthouse illumination. Scientific American, 66(24):373, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n24/pdf/scientificamerican06111892-373.pdf.

Anonymous:1892:NBH

[2728] Anonymous. The new boats of the Hoboken Ferry Company, New York. Scientific American, 66(24):374, June 11, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v66/
n24/pdf/scientificamerican06111892-374.pdf.

Anonymous:1892:Mb

[2729] Anonymous. Mantidae. Scientific American, 66(24):375–377, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n24/pdf/scientificamerican06111892-375.pdf.

Anonymous:1892:BNBr

[2730] Anonymous. Books: New books and publications. Scientific American, 66(24):378, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
378a.pdf.

Anonymous:1892:BPv

[2731] Anonymous. Business and personal. Scientific American, 66(24):378, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n24/pdf/scientificamerican06111892-378b.pdf.

Anonymous:1892:RPIx

[2732] Anonymous. Recently patented inventions. Scientific American, 66(24):378, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n24/pdf/scientificamerican06111892-
378.pdf.

Anonymous:1892:NQx

[2733] Anonymous. Notes and queries. Scientific American, 66(24):378–379, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n24/pdf/scientificamerican06111892-378c.pdf.

Anonymous:1892:Iy

[2734] Anonymous. To inventors. Scientific American, 66(24):379, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n24/pdf/scientificamerican06111892-379.pdf.

Anonymous:1892:IIy

[2735] Anonymous. Index of inventions. Scientific American, 66(24):379–380, June 11, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n24/pdf/scientificamerican06111892-379a.pdf.

Anonymous:1892:CCPa

[2736] Anonymous. The crippled condition of the Patent Office. Scientific American, 66(25):384, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
384.pdf.

Anonymous:1892:CFa

[2737] Anonymous. Cape flowers. Scientific American, 66(25):384, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n25/pdf/scientificamerican06181892-384a.pdf.

Anonymous:1892:FIE

[2738] Anonymous. French International Exhibition of the products of fermentation and distillation. Scientific American, 66(25):384–385, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n25/pdf/scientificamerican06181892-384b.pdf.

Anonymous:1892:AAT

[2739] Anonymous. An accurate automatic time recorder. Scientific American, 66(25):386, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
386.pdf.

Anonymous:1892:ASU

[2740] Anonymous. Attachments for swimmers’ use. Scientific American, 66(25):386, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
386b.pdf.

Anonymous:1892:IOS

[2741] Anonymous. An improved Ore sampling device. Scientific American, 66(25):386, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
386a.pdf.

Anonymous:1892:IBH

[2742] Anonymous. An improved bolt heading machine. Scientific American, 66(25):387, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
387.pdf.

Anonymous:1892:RSC

[2743] Anonymous. A revolving shield for cars. Scientific American, 66(25):387, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
387a.pdf.

vanAllen:1892:FI

[2744] Frank van Allen. The famine in India. Scientific American, 66 (25):387–388, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
387b.pdf.

Anonymous:1892:SIT

[2745] Anonymous. Stevens institute of technology. Scientific American, 66(25):388, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
388.pdf.

Anonymous:1892:CCPb

[2746] Anonymous. The crowded condition of the Patent Office. Scientific American, 66(25):389, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
389.pdf.

Hovey:1892:GCC

[2747] H. C. Hovey. The Grand Canyon of the Colorado. Scientific American, 66(25):392–393, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
392.pdf.

Anonymous:1892:WUT

[2748] Anonymous. The Western Union Time Ball Service. Scientific American, 66(25):393, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
393.pdf.

Anonymous:1892:BNBs

[2749] Anonymous. Books: New books and publications. Scientific American, 66(25):394, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
394a.pdf.

Anonymous:1892:RPIy

[2750] Anonymous. Recently patented inventions. Scientific American, 66(25):394, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n25/pdf/scientificamerican06181892-
394.pdf.

Anonymous:1892:BPw

[2751] Anonymous. Business and personal. Scientific American, 66(25):395, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n25/pdf/scientificamerican06181892-395.pdf.

Anonymous:1892:NQy

[2752] Anonymous. Notes and queries. Scientific American, 66(25):395, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n25/pdf/scientificamerican06181892-395a.pdf.

Anonymous:1892:IIz

[2753] Anonymous. Index of inventions. Scientific American, 66(25):395–396, June 18, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n25/pdf/scientificamerican06181892-395b.pdf.

Anonymous:1892:SAOb

[2754] Anonymous. The seven ages of our world. Scientific American, 66(26):400–401, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n26/pdf/scientificamerican06251892-
400.pdf.

Anonymous:1892:IBBb

[2755] Anonymous. An improved ballot box. Scientific American, 66(26):402, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-402.pdf.

Anonymous:1892:IPc

[2756] Anonymous. Intelligence of parrots. Scientific American, 66(26):402, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-402a.pdf.

Anonymous:1892:IHCa

[2757] Anonymous. An improved hose coupling. Scientific American, 66 (26):402–403, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n26/pdf/scientificamerican06251892-
402b.pdf.

Anonymous:1892:SCS

[2758] Anonymous. A simple camera shutter. Scientific American, 66(26):403, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-403.pdf.

Anonymous:1892:SFB

[2759] Anonymous. The Shely fiber breaker. Scientific American, 66 (26):403–404, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n26/pdf/scientificamerican06251892-
403a.pdf.

Anonymous:1892:SCP

[2760] Anonymous. Sheeley’s canning or preserving jar. Scientific American, 66(26):404, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n26/pdf/scientificamerican06251892-
404.pdf.

Anonymous:1892:Cr

[2761] Anonymous. Correspondence. Scientific American, 66(26):405, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-405.pdf.

Serviss:1892:CM

[2762] Garrett P. Serviss. From chaos to man. Scientific American, 66 (26):405–407, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n26/pdf/scientificamerican06251892-
405a.pdf.

Anonymous:1892:Cu

[2763] Anonymous. The chronograph. Scientific American, 66(26):407, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-407.pdf.

Anonymous:1892:IPS

[2764] Anonymous. Improved punching and shearing machine. Scientific American, 66(26):407, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n26/pdf/scientificamerican06251892-
407a.pdf.

Anonymous:1892:BPx

[2765] Anonymous. Business and personal. Scientific American, 66(26):408, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-408a.pdf.

Anonymous:1892:RPIz

[2766] Anonymous. Recently patented inventions. Scientific American, 66(26):408, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v66/n26/pdf/scientificamerican06251892-
408.pdf.

Anonymous:1892:NQz

[2767] Anonymous. Notes and queries. Scientific American, 66(26):408–409, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-408b.pdf.

Anonymous:1892:Iz

[2768] Anonymous. To inventors. Scientific American, 66(26):409, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-409.pdf.

Anonymous:1892:IIba

[2769] Anonymous. Index of inventions. Scientific American, 66(26):409–410, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-409a.pdf.

Anonymous:1892:Iba

[2770] Anonymous. Index. Scientific American, 66(26):411–412, June 25, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v66/n26/pdf/scientificamerican06251892-411.pdf.

Anonymous:1892:GPC

[2771] Anonymous. A great Pacific coast shipbuilding plant. Scientific American, 67(1):1–2, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n1/pdf/scientificamerican07021892-
1.pdf.

Anonymous:1892:NSSb

[2772] Anonymous. A new scholarship at Sibley College. Scientific American, 67(1):2, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n1/pdf/scientificamerican07021892-
2.pdf.

Anonymous:1892:NAB

[2773] Anonymous. Nitric acid bacteria. Scientific American, 67(1):2–3, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-2a.pdf.

Anonymous:1892:FFA

[2774] Anonymous. The Forty-First Annual Meeting of the American Association for the Advancement of Science. Scientific American, 67(1):3, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-3.pdf.

Anonymous:1892:PPJc

[2775] Anonymous. Position of the planets in July. Scientific American, 67(1): 3, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-3a.pdf.

Anonymous:1892:PWFb

[2776] Anonymous. A proposed World’s Fair tower. Scientific American, 67(1): 4, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-4.pdf.

Anonymous:1892:ICCc

[2777] Anonymous. An improvement in car couplers. Scientific American, 67(1):5, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n1/pdf/scientificamerican07021892-
5c.pdf.

Anonymous:1892:LDF

[2778] Anonymous. A light and durable fence. Scientific American, 67(1):5, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-5b.pdf.

Anonymous:1892:NTc

[2779] Anonymous. A novel top. Scientific American, 67(1):5, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-5a.pdf.

Anonymous:1892:PTA

[2780] Anonymous. Prof. Thurston’s arc lamp improvement. Scientific American, 67(1):5, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n1/pdf/scientificamerican07021892-
5.pdf.

Anonymous:1892:PNc

[2781] Anonymous. Photographic notes. Scientific American, 67(1):6, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-6.pdf.

Anonymous:1892:Cv

[2782] Anonymous. Correspondence. Scientific American, 67(1):7, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-7.pdf.

Anonymous:1892:MPC

[2783] Anonymous. Manufacture of porous cups for electric batteries. Scientific American, 67(1):8, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n1/pdf/scientificamerican07021892-
8.pdf.

Anonymous:1892:WS

[2784] Anonymous. A weird spectacle. Scientific American, 67(1):9, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-9.pdf.

Anonymous:1892:RPIba

[2785] Anonymous. Recently patented inventions. Scientific American, 67(1): 10, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-10.pdf.

Anonymous:1892:BNBt

[2786] Anonymous. Books: New books and publications. Scientific American, 67(1):11, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n1/pdf/scientificamerican07021892-
11b.pdf.

Anonymous:1892:BPy

[2787] Anonymous. Business and personal. Scientific American, 67(1):11, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-11.pdf.

Anonymous:1892:NQba

[2788] Anonymous. Notes and queries. Scientific American, 67(1):11, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-11a.pdf.

Anonymous:1892:IIbb

[2789] Anonymous. Index of inventions. Scientific American, 67(1):11–12, July 2, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n1/pdf/scientificamerican07021892-11c.pdf.

Anonymous:1892:NCD

[2790] Anonymous. The new Croton Dam. Scientific American, 67(2):15–16, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-15.pdf.

Anonymous:1892:NC

[2791] Anonymous. The Nicaragua Canal. Scientific American, 67(2):16, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-16.pdf.

Anonymous:1892:IEa

[2792] Anonymous. Of interest to electricians. Scientific American, 67(2):16– 17, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-16a.pdf.

Anonymous:1892:IBS

[2793] Anonymous. An improved balanced slide valve. Scientific American, 67(2):18, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n2/pdf/scientificamerican07091892-
18.pdf.

Anonymous:1892:ISVa

[2794] Anonymous. An improved safety valve and muffler. Scientific American, 67(2):18, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n2/pdf/scientificamerican07091892-
18b.pdf.

Anonymous:1892:SSCa

[2795] Anonymous. A scraper for Sawmill carriages. Scientific American, 67(2):18, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n2/pdf/scientificamerican07091892-
18a.pdf.

Anonymous:1892:MLD

[2796] Anonymous. Machine for laying drainage pipes. Scientific American, 67(2):18–19, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n2/pdf/scientificamerican07091892-
18c.pdf.

Anonymous:1892:HE

[2797] Anonymous. Homogenetic enumeration. Scientific American, 67(2):19– 20, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-19.pdf.

Anonymous:1892:Cx

[2798] Anonymous. Correspondence. Scientific American, 67(2):21, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-21.pdf.

Anonymous:1892:SWH

[2799] Anonymous. Sterilization of water by heat. Scientific American, 67(2): 22, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-22.pdf.

Anonymous:1892:STF

[2800] Anonymous. Steam tree-felling and cross-cutting machinery. Scientific American, 67(2):23–25, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n2/pdf/scientificamerican07091892-
23.pdf.

Anonymous:1892:REL

[2801] Anonymous. Remarkable experiments with liquefied air and liquefied oxygen and other liquefied gases. Scientific American, 67(2):25, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-25.pdf.

Anonymous:1892:BNBu

[2802] Anonymous. Books: New books and publications. Scientific American, 67(2):26, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n2/pdf/scientificamerican07091892-
26a.pdf.

Anonymous:1892:RPIbb

[2803] Anonymous. Recently patented inventions. Scientific American, 67(2): 26, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-26.pdf.

Anonymous:1892:BPz

[2804] Anonymous. Business and personal. Scientific American, 67(2):27, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-27.pdf.

Anonymous:1892:Ibc

[2805] Anonymous. To inventors. Scientific American, 67(2):27, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-27b.pdf.

Anonymous:1892:NQbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-27a.pdf.

Anonymous:1892:IIbc

[2807] Anonymous. Index of inventions. Scientific American, 67(2):27–28, July 9, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n2/pdf/scientificamerican07091892-27c.pdf.

Anonymous:1892:GTS

[2808] Anonymous. The generation of thunder storms. Scientific American, 67(3):32, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
32.pdf.

Anonymous:1892:PCa

[2809] Anonymous. Paterson’s centennial. Scientific American, 67(3):32, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n3/pdf/scientificamerican07161892-32a.pdf.

Anonymous:1892:GGA

[2810] Anonymous. Great guns and armor plate. Scientific American, 67(3):32–33, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
32b.pdf.

Anonymous:1892:CTP

[2811] Anonymous. A convenient tool for pipe fitters. Scientific American, 67(3):34, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
34c.pdf.

Anonymous:1892:IRJ

[2812] Anonymous. An improved rail joint. Scientific American, 67(3):34, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n3/pdf/scientificamerican07161892-34b.pdf.

Anonymous:1892:LWA

[2813] Anonymous. A low water alarm for boilers. Scientific American, 67(3):34, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
34.pdf.

Anonymous:1892:UL

[2814] Anonymous. Underground locomotive. Scientific American, 67(3):34, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n3/pdf/scientificamerican07161892-34a.pdf.

Hopkins:1892:CC

[2815] Geo. M. Hopkins. A convenient camera. Scientific American, 67(3):35–36, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
35.pdf.

Anonymous:1892:SAD

[2816] Anonymous. The Scientific American dynamo. Scientific American, 67(3):37, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
37.pdf.

Anonymous:1892:CSS

[2817] Anonymous. The central screw steamer Le Louvre. Scientific American, 67(3):38, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
38.pdf.

Anonymous:1892:SSSb

[2818] Anonymous. The sewerage systems of Sydney, Australia. Scientific American, 67(3):39–40, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
39.pdf.

Anonymous:1892:ABW

[2819] Anonymous. The American black wolf. Scientific American, 67(3):41, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n3/pdf/scientificamerican07161892-41.pdf.

Anonymous:1892:RPIbc

[2820] Anonymous. Recently patented inventions. Scientific American, 67(3):42, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
42.pdf.

Anonymous:1892:BNBv

[2821] Anonymous. Books: New books and publications. Scientific American, 67(3):42–43, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n3/pdf/scientificamerican07161892-
42a.pdf.

Anonymous:1892:BPbc

[2822] Anonymous. Business and personal. Scientific American, 67(3):43, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n3/pdf/scientificamerican07161892-43.pdf.

Anonymous:1892:NQbc

[2823] Anonymous. Notes and queries. Scientific American, 67(3):43, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n3/pdf/scientificamerican07161892-43a.pdf.

Anonymous:1892:IIbd

[2824] Anonymous. Index of inventions. Scientific American, 67(3):43–44, July 16, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n3/pdf/scientificamerican07161892-43b.pdf.

Anonymous:1892:IWF

[2825] Anonymous. Inventors at the World’s Fair. Scientific American, 67(4):48, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
48.pdf.

Anonymous:1892:DBI

[2826] Anonymous. Defective boilers and incompetent engineers. Scientific American, 67(4):48–49, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
48a.pdf.

Anonymous:1892:IBF

[2827] Anonymous. An improved boiler furnace. Scientific American, 67(4):50, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-50.pdf.

Anonymous:1892:ISW

[2828] Anonymous. Improved steam whistles. Scientific American, 67(4):50, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-50b.pdf.

Anonymous:1892:SST

[2829] Anonymous. A salt sprinkler for table use. Scientific American, 67(4):50, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
50a.pdf.

Anonymous:1892:IMHb

[2830] Anonymous. Improved method of handling nitro-glycerine. Scientific American, 67(4):50–51, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
50c.pdf.

Anonymous:1892:EVE

[2831] Anonymous. Eruption of the volcano of Etna. Scientific American, 67(4):51, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
51.pdf.

Anonymous:1892:GMW

[2832] Anonymous. Great mineral water baths. Scientific American, 67(4):52, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-52.pdf.

Anonymous:1892:CWF

[2833] Anonymous. Cyrus W. Field. Scientific American, 67(4):53, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-53.pdf.

Anonymous:1892:IEM

[2834] Anonymous. Inexpensive electric motors. Scientific American, 67(4):54, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-54.pdf.

Hovey:1892:VCP

[2835] H. C. Hovey. A visit to Chalcedony Park, Arizona. Scientific American, 67(4):55–56, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
55.pdf.

Anonymous:1892:WLH

[2836] Anonymous. The Water Lily House at Kew. Scientific American, 67(4):57, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
57.pdf.

Anonymous:1892:RPIbd

[2837] Anonymous. Recently patented inventions. Scientific American, 67(4):58, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n4/pdf/scientificamerican07231892-
58.pdf.

Anonymous:1892:BPbd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-59.pdf.

Anonymous:1892:Ibd

[2839] Anonymous. To inventors. Scientific American, 67(4):59, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-59b.pdf.

Anonymous:1892:NQbd

[2840] Anonymous. Notes and queries. Scientific American, 67(4):59, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-59a.pdf.

Anonymous:1892:IIbe

[2841] Anonymous. Index of inventions. Scientific American, 67(4):59–60, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-59c.pdf.

Anonymous:1892:PNPa

[2842] Anonymous. Patent notes: Patents. Scientific American, 67(4):62, July 23, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n4/pdf/scientificamerican07231892-62.pdf.

Anonymous:1892:CHa

[2843] Anonymous. The camera at Homestead. Scientific American, 67(5):64, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-64a.pdf.

Anonymous:1892:PMP

[2844] Anonymous. Is perpetual motion possible? Scientific American, 67(5):64, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
64.pdf.

Anonymous:1892:PPAb

[2845] Anonymous. Position of the planets in August. Scientific American, 67(5):65, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
65.pdf.

Anonymous:1892:CRCb

[2846] Anonymous. A comfortable reading chair. Scientific American, 67(5):66, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
66a.pdf.

Anonymous:1892:ERE

[2847] Anonymous. An easily regulated electric motor. Scientific American, 67(5):66, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
66.pdf.

Anonymous:1892:ICCd

[2848] Anonymous. An improved car coupling. Scientific American, 67(5):66, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-66b.pdf.

Anonymous:1892:MHV

[2849] Anonymous. A machine to holystone a vessel’s deck. Scientific American, 67(5):66–67, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
66c.pdf.

Anonymous:1892:VDL

[2850] Anonymous. The Viaduct Du Loup. Scientific American, 67(5):67, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-67.pdf.

Anonymous:1892:Cy

[2851] Anonymous. Correspondence. Scientific American, 67(5):68, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-68.pdf.

Anonymous:1892:ODT

[2852] Anonymous. The oscillating disappearing turret. Scientific American, 67(5):69, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
69.pdf.

Anonymous:1892:MSB

[2853] Anonymous. Mechanical singing birds. Scientific American, 67(5):70, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-70.pdf.

Anonymous:1892:BSB

[2854] Anonymous. The Baker submarine boat. Scientific American, 67(5):71–72, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
71.pdf.

Anonymous:1892:PPMc

[2855] Anonymous. Polychrome projections by means of uncolored photographs. Scientific American, 67(5):72–74, July 30, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n5/pdf/scientificamerican07301892-72.pdf.

Anonymous:1892:RPIbe

[2856] Anonymous. Recently patented inventions. Scientific American, 67(5):74–75, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n5/pdf/scientificamerican07301892-
74.pdf.

Anonymous:1892:BPbe

[2857] Anonymous. Business and personal. Scientific American, 67(5):75, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-75.pdf.

Anonymous:1892:Ibe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-75b.pdf.

Anonymous:1892:NQbe

[2859] Anonymous. Notes and queries. Scientific American, 67(5):75, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-75a.pdf.

Anonymous:1892:IIbf

[2860] Anonymous. Index of inventions. Scientific American, 67(5):75–76, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-75c.pdf.

Anonymous:1892:PNPb

[2861] Anonymous. Patent notes: Patents. Scientific American, 67(5):78, July 30, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n5/pdf/scientificamerican07301892-78.pdf.

Anonymous:1892:VC

[2862] Anonymous. The vessels of Columbus. Scientific American, 67(6):79, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-79.pdf.

Anonymous:1892:EBC

[2863] Anonymous. Electric boats and carriages at the Columbian Exposition. Scientific American, 67(6):80, August 6, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n6/pdf/scientificamerican08061892-80.pdf.

Anonymous:1892:NMPa

[2864] Anonymous. A new map of the Planet Mars. Scientific American, 67(6):80, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n6/pdf/scientificamerican08061892-
80a.pdf.

Anonymous:1892:MBa

[2865] Anonymous. Moses S. Beach. Scientific American, 67(6):80–81, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-80b.pdf.

Anonymous:1892:AEF

[2866] Anonymous. Adjustable embroidery frame and table. Scientific American, 67(6):82, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n6/pdf/scientificamerican08061892-
82a.pdf.

Anonymous:1892:IHPb

[2867] Anonymous. An improved hydraulic press. Scientific American, 67(6):82, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n6/pdf/scientificamerican08061892-
82b.pdf.

Anonymous:1892:NUT

[2868] Anonymous. A new and useful telephone. Scientific American, 67(6):82, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-82.pdf.

Anonymous:1892:NMPb

[2869] Anonymous. A new mill for pulverizing. Scientific American, 67(6):83, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-83.pdf.

Anonymous:1892:LC

[2870] Anonymous. Launch of the Columbia. Scientific American, 67(6):84, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-84.pdf.

Anonymous:1892:E

[2871] Anonymous. Ethylene. Scientific American, 67(6):84–85, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-84a.pdf.

Anonymous:1892:FDC

[2872] Anonymous. Fire department “Cut-Loose”. Scientific American, 67(6):86, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n6/pdf/scientificamerican08061892-
86.pdf.

Hovey:1892:RDG

[2873] H. C. Hovey. On the rim and in the depths of the Grand Canyon. Scientific American, 67(6):87–89, August 6, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n6/pdf/scientificamerican08061892-87.pdf.

Anonymous:1892:Cz

[2874] Anonymous. Correspondence. Scientific American, 67(6):89, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-89.pdf.

Anonymous:1892:BPBf

[2875] Anonymous. Business and personal: Business and personal. Scientific American, 67(6):90, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n6/pdf/scientificamerican08061892-
90a.pdf.

Anonymous:1892:RPIbf

[2876] Anonymous. Recently patented inventions. Scientific American, 67(6):90, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n6/pdf/scientificamerican08061892-
90.pdf.

Anonymous:1892:NQbf

[2877] Anonymous. Notes and queries. Scientific American, 67(6):90–91, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-90b.pdf.

Anonymous:1892:BNBw

[2878] Anonymous. Books: New books and publications. Scientific American, 67(6):91, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n6/pdf/scientificamerican08061892-
91.pdf.

Anonymous:1892:Ibg

[2879] Anonymous. To inventors. Scientific American, 67(6):91, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-91a.pdf.

Anonymous:1892:IIbg

[2880] Anonymous. Index of inventions. Scientific American, 67(6):91, August 6, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n6/pdf/scientificamerican08061892-91b.pdf.

Anonymous:1892:CCa

[2881] Anonymous. Christopher Columbus. Scientific American, 67(7):95, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n7/pdf/scientificamerican08131892-95.pdf.

Anonymous:1892:AEB

[2882] Anonymous. The approaching expiration of the Bell Telephone patent. Scientific American, 67(7):96–97, August 13, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n7/pdf/scientificamerican08131892-96.pdf.

Anonymous:1892:CCH

[2883] Anonymous. Copying camera holder. Scientific American, 67(7):98, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n7/pdf/scientificamerican08131892-98a.pdf.

Anonymous:1892:HVB

[2884] Anonymous. Hamilton’s ventilating blower. Scientific American, 67(7):98, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
98c.pdf.

Anonymous:1892:IHa

[2885] Anonymous. An improved hanger. Scientific American, 67(7):98, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n7/pdf/scientificamerican08131892-98b.pdf.

Anonymous:1892:UWPb

[2886] Anonymous. The utilization of water power. Scientific American, 67(7):98, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
98.pdf.

Anonymous:1892:EPD

[2887] Anonymous. An electro-photo-detective thief catcher. Scientific American, 67(7):99, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
99.pdf.

Anonymous:1892:Cba

[2888] Anonymous. Correspondence. Scientific American, 67(7):100, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n7/pdf/scientificamerican08131892-100.pdf.

Anonymous:1892:NGL

[2889] Anonymous. New Government lightships. Scientific American, 67(7):102, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
102.pdf.

Gillmore:1892:BKL

[2890] Parker Gillmore and (“Ubique”). Bushmen killing a lion. Scientific American, 67(7):103–104, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
103.pdf.

Anonymous:1892:EWJ

[2891] Anonymous. Experiments with water jets. Scientific American, 67(7):105, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
105.pdf.

Anonymous:1892:BPbg

[2892] Anonymous. Business and personal. Scientific American, 67(7):106, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n7/pdf/scientificamerican08131892-106a.pdf.

Anonymous:1892:RPIbg

[2893] Anonymous. Recently patented inventions. Scientific American, 67(7):106, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
106.pdf.

Anonymous:1892:NQbg

[2894] Anonymous. Notes and queries. Scientific American, 67(7):106– 107, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
106b.pdf.

Anonymous:1892:Ibi

[2895] Anonymous. To inventors. Scientific American, 67(7):107, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n7/pdf/scientificamerican08131892-107.pdf.

Anonymous:1892:IIbh

[2896] Anonymous. Index of inventions. Scientific American, 67(7): 107–108, August 13, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n7/pdf/scientificamerican08131892-
107a.pdf.

Anonymous:1892:PI

[2897] Anonymous. Psychology of invention. Scientific American, 67 (8):112, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
112.pdf.

Anonymous:1892:EIM

[2898] Anonymous. The Earth and the inner Moon of Mars. Scientific American, 67(8):112–113, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
112a.pdf.

Hopkins:1892:USD

[2899] M. Hopkins. Unscientific and scientific divining rods. Scientific American, 67(8):114–115, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
114.pdf.

Anonymous:1892:IFEb

[2900] Anonymous. An improved fire engine pump. Scientific American, 67(8):115, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
115a.pdf.

Anonymous:1892:SSCb

[2901] Anonymous. The steel steamer Choctaw. Scientific American, 67(8):115, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
115.pdf.

Anonymous:1892:HDa

[2902] Anonymous. The homestead difficulties. Scientific American, 67 (8):116, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
116.pdf.

Anonymous:1892:MM

[2903] Anonymous. The Moon and Mars. Scientific American, 67(8):116, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n8/pdf/scientificamerican08201892-116a.pdf.

Anonymous:1892:LPM

[2904] Anonymous. The looped path of Mars. Scientific American, 67 (8):116–118, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
116b.pdf.

Anonymous:1892:DB

[2905] Anonymous. A deep basement. Scientific American, 67(8):118, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n8/pdf/scientificamerican08201892-118.pdf.

Anonymous:1892:BNBx

[2906] Anonymous. Books: New books and publications. Scientific American, 67(8):123, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n8/pdf/scientificamerican08201892-
123.pdf.

Anonymous:1892:Ibj

[2907] Anonymous. To inventors. Scientific American, 67(8):123, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n8/pdf/scientificamerican08201892-123a.pdf.

Anonymous:1892:IIbi

[2908] Anonymous. Index of inventions. Scientific American, 67(8):123, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n8/pdf/scientificamerican08201892-123b.pdf.

Anonymous:1892:PNPc

[2909] Anonymous. Patent notes: Patents. Scientific American, 67(8):126, August 20, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n8/pdf/scientificamerican08201892-126.pdf.

Anonymous:1892:EGS

[2910] Anonymous. Explosion in a great sewer. Scientific American, 67(9):128, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
128.pdf.

Anonymous:1892:RS

[2911] Anonymous. The railroad strikes. Scientific American, 67(9): 128–130, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
128a.pdf.

Anonymous:1892:ICCe

[2912] Anonymous. An improved car coupling. Scientific American, 67(9):130, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
130a.pdf.

Anonymous:1892:WSE

[2913] Anonymous. Watering streets by electricity. Scientific American, 67(9):130, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
130.pdf.

Anonymous:1892:FBR

[2914] Anonymous. The fall of bodies and the resistance of the air. Scientific American, 67(9):131, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
131.pdf.

Anonymous:1892:CSW

[2915] Anonymous. The Carnegie Steel Works, Homestead. Scientific American, 67(9):132, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
132.pdf.

Benjamin:1892:JC

[2916] Marcus Benjamin. Joseph Le Conte. Scientific American, 67 (9):133–134, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
133.pdf.

Anonymous:1892:NPa

[2917] Anonymous. A new parachute. Scientific American, 67(9):134–135, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n9/pdf/scientificamerican08271892-134.pdf.

Anonymous:1892:NRT

[2918] Anonymous. Naval Reserves at target practice. Scientific American, 67(9):135, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
135.pdf.

Anonymous:1892:Cbb

[2919] Anonymous. Correspondence. Scientific American, 67(9):136, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n9/pdf/scientificamerican08271892-136.pdf.

Anonymous:1892:HPS

[2920] Anonymous. Household pests — with some hints how to get rid of them. Scientific American, 67(9):136–137, August 27, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n9/pdf/scientificamerican08271892-136a.pdf.

Anonymous:1892:CFM

[2921] Anonymous. A cat as foster mother of an ape — a picture from the Leipzig Zoological Garden. Scientific American, 67(9):137, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n9/pdf/scientificamerican08271892-137.pdf.

Anonymous:1892:RPIbh 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
138.pdf.

Anonymous:1892:BNBy

[2923] Anonymous. Books: New books and publications. Scientific American, 67(9):138–139, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
138a.pdf.

Anonymous:1892:BPbh

[2924] Anonymous. Business and personal. Scientific American, 67(9):139, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n9/pdf/scientificamerican08271892-139.pdf.

Anonymous:1892:Ibk

[2925] Anonymous. To inventors. Scientific American, 67(9):139, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n9/pdf/scientificamerican08271892-139b.pdf.

Anonymous:1892:NQbh

[2926] Anonymous. Notes and queries. Scientific American, 67(9):139, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n9/pdf/scientificamerican08271892-139a.pdf.

Anonymous:1892:IIbj

[2927] Anonymous. Index of inventions. Scientific American, 67(9): 139–140, August 27, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n9/pdf/scientificamerican08271892-
139c.pdf.

Anonymous:1892:HSW

[2928] Anonymous. High-Service Water Tower, Brooklyn. Scientific American, 67(10):143, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
143.pdf.

Anonymous:1892:RMA

[2929] Anonymous. The Rochester Meeting of the American Association. Scientific American, 67(10):144–145, September 3, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n10/pdf/scientificamerican09031892-144.pdf.

Anonymous:1892:PPS

[2930] Anonymous. Positions of the planets in September. Scientific American, 67(10):145, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
145.pdf.

Anonymous:1892:CLA

[2931] Anonymous. A coffin-lowering apparatus. Scientific American, 67 (10):146, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
146a.pdf.

Anonymous:1892:ERB

[2932] Anonymous. “The Expert” rubber band dating stamp. Scientific American, 67(10):146, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
146b.pdf.

Anonymous:1892:MFI

[2933] Anonymous. A multiple fuse igniter. Scientific American, 67 (10):146, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
146.pdf.

Anonymous:1892:IREb

[2934] Anonymous. An improved rotary engine. Scientific American, 67 (10):146–147, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
146c.pdf.

Anonymous:1892:NDBa

[2935] Anonymous. A novel device for blowing dust from highways. Scientific American, 67(10):147, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
147.pdf.

Anonymous:1892:NPb

[2936] Anonymous. A new photometer. Scientific American, 67(10): 148, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
148.pdf.

Anonymous:1892:IAS

[2937] Anonymous. An improved air ship. Scientific American, 67 (10):150, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
150.pdf.

Anonymous:1892:NSC

[2938] Anonymous. Notable skill in clock making. Scientific American, 67(10):150, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
150a.pdf.

Anonymous:1892:MEI

[2939] Anonymous. Masthead electric illumination. Scientific American, 67(10):151, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
151.pdf.

Anonymous:1892:ISa

[2940] Anonymous. The inscription of speech. Scientific American, 67 (10):151–153, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
151a.pdf.

Anonymous:1892:GDT

[2941] Anonymous. The Golden Door of the Transportation Building of the World’s Fair. Scientific American, 67(10):153, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n10/pdf/scientificamerican09031892-153.pdf.

Anonymous:1892:BNBz

[2942] Anonymous. Books: New books and publications. Scientific American, 67(10):154, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
154a.pdf.

Anonymous:1892:BPbi

[2943] Anonymous. Business and personal. Scientific American, 67 (10):154, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
154b.pdf.

Anonymous:1892:RPIbi

[2944] Anonymous. Recently patented inventions. Scientific American, 67(10):154, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
154.pdf.

Anonymous:1892:NQbi

[2945] Anonymous. Notes and queries. Scientific American, 67(10): 154–155, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
154c.pdf.

Anonymous:1892:Ibl

[2946] Anonymous. To inventors. Scientific American, 67(10):155, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n10/pdf/scientificamerican09031892-155.pdf.

Anonymous:1892:IIbk

[2947] Anonymous. Index of inventions. Scientific American, 67(10): 155–156, September 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n10/pdf/scientificamerican09031892-
155a.pdf.

Anonymous:1892:RAP

[2948] Anonymous. Recent armor plate trials. Scientific American, 67 (11):160, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
160.pdf.

Anonymous:1892:OI

[2949] Anonymous. Opportunities for invention. Scientific American, 67 (11):160–161, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
160a.pdf.

Anonymous:1892:CRH

[2950] Anonymous. A convenient ribbon holder. Scientific American, 67 (11):162, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
162b.pdf.

Anonymous:1892:ESBb

[2951] Anonymous. Experiments with soap bubbles. Scientific American, 67(11):162, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
162a.pdf.

Anonymous:1892:IWS

[2952] Anonymous. An improved window screen. Scientific American, 67 (11):162, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
162.pdf.

Anonymous:1892:OC

[2953] Anonymous. Old clocks. Scientific American, 67(11):163, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n11/pdf/scientificamerican09101892-163.pdf.

Anonymous:1892:LLB

[2954] Anonymous. Lighthouses, lightships, and buoys. Scientific American, 67(11):164–165, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
164.pdf.

Anonymous:1892:MFH

[2955] Anonymous. Manufacture of felt hats. Scientific American, 67 (11):166, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
166.pdf.

Anonymous:1892:Cbc

[2956] Anonymous. The Columbia. Scientific American, 67(11):167, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n11/pdf/scientificamerican09101892-167.pdf.

Anonymous:1892:Cbd

[2957] Anonymous. Correspondence. Scientific American, 67(11):168, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n11/pdf/scientificamerican09101892-168.pdf.

Anonymous:1892:O

[2958] Anonymous. The Otocyon. Scientific American, 67(11):169, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n11/pdf/scientificamerican09101892-169.pdf.

Anonymous:1892:BPbj

[2959] Anonymous. Business and personal. Scientific American, 67 (11):170, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
170a.pdf.

Anonymous:1892:RPIbj

[2960] Anonymous. Recently patented inventions. Scientific American, 67 (11):170, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
170.pdf.

Anonymous:1892:NQbj

[2961] Anonymous. Notes and queries. Scientific American, 67(11): 170–171, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
170b.pdf.

Anonymous:1892:Ibm

[2962] Anonymous. To inventors. Scientific American, 67(11):171, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n11/pdf/scientificamerican09101892-171.pdf.

Anonymous:1892:IIbl

[2963] Anonymous. Index of inventions. Scientific American, 67(11): 171–172, September 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n11/pdf/scientificamerican09101892-
171a.pdf.

Anonymous:1892:NIQ

[2964] Anonymous. The need of improved quarantine stations. Scientific American, 67(12):176, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
176.pdf.

Anonymous:1892:SBN

[2965] Anonymous. The Second Battle of New Orleans. Scientific American, 67(12):176–177, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
176a.pdf.

Anonymous:1892:DDO

[2966] Anonymous. A device to distribute oil on water. Scientific American, 67(12):178, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
178d.pdf.

Anonymous:1892:EGP

[2967] Anonymous. An efficient glass-polishing machine. Scientific American, 67(12):178, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
178b.pdf.

Anonymous:1892:IBT

[2968] Anonymous. An improved bolt-threading machine. Scientific American, 67(12):178, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
178c.pdf.

Anonymous:1892:ILM

[2969] Anonymous. An improved liquid-measuring device. Scientific American, 67(12):178, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
178a.pdf.

Anonymous:1892:INY

[2970] Anonymous. An improved neck yoke. Scientific American, 67 (12):178, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
178.pdf.

Anonymous:1892:IHCb

[2971] Anonymous. Improved hand camera. Scientific American, 67 (12):179, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
179.pdf.

Anonymous:1892:RCF

[2972] Anonymous. Rogers’ cold forging process for wood screws. Scientific American, 67(12):179, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
179a.pdf.

Anonymous:1892:SLI

[2973] Anonymous. Straight line indicator. Scientific American, 67 (12):179, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
179b.pdf.

Anonymous:1892:CFc

[2974] Anonymous. The cone of flowers. Scientific American, 67(12): 179–180, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
179c.pdf.

Anonymous:1892:WCEa

[2975] Anonymous. The World’s Columbian Exposition. Scientific American, 67(12):180, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
180.pdf.

Anonymous:1892:Cbe

[2976] Anonymous. Correspondence. Scientific American, 67(12):181, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n12/pdf/scientificamerican09171892-181.pdf.

Anonymous:1892:SSG

[2977] Anonymous. Searching for sunken gold near Hell gate, New York. Scientific American, 67(12):181, September 17, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n12/pdf/scientificamerican09171892-181a.pdf.

Anonymous:1892:RDR

[2978] Anonymous. Recent decisions relating to patents. Scientific American, 67(12):184, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
184.pdf.

Anonymous:1892:IAC

[2979] Anonymous. Improvement in arranging cut flowers. Scientific American, 67(12):185–186, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
185.pdf.

Anonymous:1892:RPIbk

[2980] Anonymous. Recently patented inventions. Scientific American, 67 (12):186–187, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
186.pdf.

Anonymous:1892:BPbk

[2981] Anonymous. Business and personal. Scientific American, 67 (12):187, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
187.pdf.

Anonymous:1892:NQbk

[2982] Anonymous. Notes and queries. Scientific American, 67(12): 187, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
187a.pdf.

Anonymous:1892:IIbm

[2983] Anonymous. Index of inventions. Scientific American, 67(12): 187–188, September 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n12/pdf/scientificamerican09171892-
187b.pdf.

Anonymous:1892:APE

[2984] Anonymous. Advance of polar exploration. Scientific American, 67 (13):192–193, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
192.pdf.

Anonymous:1892:AP

[2985] Anonymous. Art and photography. Scientific American, 67(13): 194, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
194.pdf.

Anonymous:1892:CICb

[2986] Anonymous. Carr’s improved combination surface gauge. Scientific American, 67(13):194, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
194a.pdf.

Anonymous:1892:HHT

[2987] Anonymous. How horse trotting improves. Scientific American, 67 (13):195, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
195.pdf.

Anonymous:1892:IBMa

[2988] Anonymous. An improved brick machine. Scientific American, 67 (13):195, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
195a.pdf.

Anonymous:1892:WET

[2989] Anonymous. A Watchman’s electric time recorder. Scientific American, 67(13):195, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
195b.pdf.

Anonymous:1892:Cbf

[2990] Anonymous. Correspondence. Scientific American, 67(13):196, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n13/pdf/scientificamerican09241892-196.pdf.

Anonymous:1892:ATD

[2991] Anonymous. The Austin, Texas, Dam. Scientific American, 67 (13):197, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
197.pdf.

Anonymous:1892:LFM

[2992] Anonymous. Lightning full mounted screw plates. Scientific American, 67(13):197, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
197a.pdf.

Anonymous:1892:IRT

[2993] Anonymous. Improved reaping, thrashing, and bagging machine. Scientific American, 67(13):198, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
198.pdf.

Beach:1892:SS

[2994] A. E. Beach. The study of the stars. Scientific American, 67(13): 199–200, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
199.pdf.

Anonymous:1892:FGC

[2995] Anonymous. A field glass camera. Scientific American, 67(13): 201–202, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
201.pdf.

Anonymous:1892:RPIbl

[2996] Anonymous. Recently patented inventions. Scientific American, 67 (13):202, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
202.pdf.

Anonymous:1892:BPbl

[2997] Anonymous. Business and personal. Scientific American, 67 (13):203, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
203.pdf.

Anonymous:1892:CRa

[2998] Anonymous. Communications received. Scientific American, 67 (13):203, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
203b.pdf.

Anonymous:1892:NQbl

[2999] Anonymous. Notes and queries. Scientific American, 67(13): 203, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
203a.pdf.

Anonymous:1892:IIbn

[3000] Anonymous. Index of inventions. Scientific American, 67(13): 203–204, September 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n13/pdf/scientificamerican09241892-
203c.pdf.

Anonymous:1892:PCb

[3001] Anonymous. Progress of the Cholera. Scientific American, 67(14):208, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-208.pdf.

Anonymous:1892:CSI

[3002] Anonymous. Cholera and some indications for its management. Scientific American, 67(14):208–209, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
208a.pdf.

Anonymous:1892:B

[3003] Anonymous. Baku. Scientific American, 67(14):210, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-210.pdf.

Anonymous:1892:INLb

[3004] Anonymous. An improved nut lock. Scientific American, 67(14):211, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-211.pdf.

Anonymous:1892:ISI

[3005] Anonymous. An improved speed indicator. Scientific American, 67(14):211, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
211b.pdf.

Anonymous:1892:NMIb

[3006] Anonymous. New mathematical instrument. Scientific American, 67(14):211, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
211c.pdf.

Anonymous:1892:SET

[3007] Anonymous. A simple and efficient typewriter. Scientific American, 67(14):211, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
211a.pdf.

Anonymous:1892:PPO

[3008] Anonymous. Position of the planets in October. Scientific American, 67(14):212, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
212.pdf.

Anonymous:1892:MBb

[3009] Anonymous. Military ballooning. Scientific American, 67(14):213, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-213.pdf.

Anonymous:1892:LDT

[3010] Anonymous. A long distance transmission plant between Tivoli and Rome. Scientific American, 67(14):214–215, October 1, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n14/pdf/scientificamerican10011892-214.pdf.

Anonymous:1892:ROT

[3011] Anonymous. Reynaud’s optical theater. Scientific American, 67 (14):215, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
215.pdf.

Anonymous:1892:Cbg

[3012] Anonymous. Correspondence. Scientific American, 67(14):216–217, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-216.pdf.

Anonymous:1892:PAa

[3013] Anonymous. The Pipa Americana. Scientific American, 67(14):217, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-217.pdf.

Anonymous:1892:BNBba

[3014] Anonymous. Books: New books and publications. Scientific American, 67(14):218, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
218a.pdf.

Anonymous:1892:RPIbm

[3015] Anonymous. Recently patented inventions. Scientific American, 67(14):218, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n14/pdf/scientificamerican10011892-
218.pdf.

Anonymous:1892:BPbm

[3016] Anonymous. Business and personal. Scientific American, 67(14):219, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-219.pdf.

Anonymous:1892:IIbo

[3017] Anonymous. Index of inventions. Scientific American, 67(14):219, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-219b.pdf.

Anonymous:1892:NQbm

[3018] Anonymous. Notes and queries. Scientific American, 67(14):219, October 1, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n14/pdf/scientificamerican10011892-219a.pdf.

Anonymous:1892:JNM

[3019] Anonymous. Jupiter’s new Moon. Scientific American, 67(15):224, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n15/pdf/scientificamerican10081892-224a.pdf.

Anonymous:1892:SI

[3020] Anonymous. Suggestions for inventors. Scientific American, 67(15):224, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n15/pdf/scientificamerican10081892-224.pdf.

Anonymous:1892:DHH

[3021] Anonymous. Dredging the Honolulu Harbor Bar. Scientific American, 67(15):224–225, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
224b.pdf.

Anonymous:1892:ACSb

[3022] Anonymous. An apparatus to consume sewer gas. Scientific American, 67(15):226, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
226a.pdf.

Anonymous:1892:CCB

[3023] Anonymous. Combined cycling and boating machine. Scientific American, 67(15):226, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
226.pdf.

Anonymous:1892:NDBb

[3024] Anonymous. A novel design for a badge. Scientific American, 67 (15):226–227, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
226b.pdf.

Anonymous:1892:IMM

[3025] Anonymous. An improved mechanical movement. Scientific American, 67(15):227, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
227.pdf.

Anonymous:1892:BM

[3026] Anonymous. Breviceps Mossambicus. Scientific American, 67 (15):227–228, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
227a.pdf.

Anonymous:1892:PVC

[3027] Anonymous. Places visited by Columbus in the West Indies. Scientific American, 67(15):229–230, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
229.pdf.

Anonymous:1892:CEC

[3028] Anonymous. Clock with electric chimes. Scientific American, 67(15):230, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
230.pdf.

Anonymous:1892:GE

[3029] Anonymous. The Great Exposition. Scientific American, 67(15): 231–233, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
231.pdf.

Anonymous:1892:GHT

[3030] Anonymous. The great-headed turtle. Scientific American, 67(15):233, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n15/pdf/scientificamerican10081892-233.pdf.

Anonymous:1892:BPbn

[3031] Anonymous. Business and personal. Scientific American, 67(15):234, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n15/pdf/scientificamerican10081892-234a.pdf.

Anonymous:1892:RPIbn

[3032] Anonymous. Recently patented inventions. Scientific American, 67(15):234, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n15/pdf/scientificamerican10081892-
234.pdf.

Anonymous:1892:NQbn

[3033] Anonymous. Notes and queries. Scientific American, 67(15):234–235, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n15/pdf/scientificamerican10081892-234b.pdf.

Anonymous:1892:Ibn

[3034] Anonymous. To inventors. Scientific American, 67(15):235, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n15/pdf/scientificamerican10081892-235.pdf.

Anonymous:1892:IIbp

[3035] Anonymous. Index of inventions. Scientific American, 67(15):235, October 8, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n15/pdf/scientificamerican10081892-235a.pdf.

Anonymous:1892:CCO

[3036] Anonymous. The Columbian Celebration of October, 1892. Scientific American, 67(16):240–242, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n16/pdf/scientificamerican10151892-
240.pdf.

Anonymous:1892:FAR

[3037] Anonymous. A fan attachment for rocking chairs. Scientific American, 67(16):242, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n16/pdf/scientificamerican10151892-
242.pdf.

Anonymous:1892:IGS

[3038] Anonymous. An improved gang Sawmill. Scientific American, 67 (16):242–243, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n16/pdf/scientificamerican10151892-
242a.pdf.

Anonymous:1892:AML

[3039] Anonymous. Apparatus for manufacturing and lifting blocks of beton at the Port of Bilbao. Scientific American, 67(16):243–244, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n16/pdf/scientificamerican10151892-243.pdf.

Anonymous:1892:SM

[3040] Anonymous. The Santa Maria. Scientific American, 67(16):245, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n16/pdf/scientificamerican10151892-245.pdf.

Anonymous:1892:FFW

[3041] Anonymous. A four-foot water valve. Scientific American, 67 (16):246, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n16/pdf/scientificamerican10151892-
246a.pdf.

Anonymous:1892:PBA

[3042] Anonymous. The probably best authenticated picture of Columbus. Scientific American, 67(16):246, October 15, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n16/pdf/scientificamerican10151892-246.pdf.

Anonymous:1892:CPU

[3043] Anonymous. The Canadian plan for utilizing the power of Niagara Falls. Scientific American, 67(16):247, October 15, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n16/pdf/scientificamerican10151892-247.pdf.

Anonymous:1892:Cbh

[3044] Anonymous. Correspondence. Scientific American, 67(16):248–249, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n16/pdf/scientificamerican10151892-248.pdf.

Anonymous:1892:ITb

[3045] Anonymous. An interesting toy. Scientific American, 67(16):249, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n16/pdf/scientificamerican10151892-249.pdf.

Anonymous:1892:RPIbo

[3046] Anonymous. Recently patented inventions. Scientific American, 67(16):250, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n16/pdf/scientificamerican10151892-
250.pdf.

Anonymous:1892:BPbo

[3047] Anonymous. Business and personal. Scientific American, 67 (16):251, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n16/pdf/scientificamerican10151892-
251.pdf.

Anonymous:1892:Ibo

[3048] Anonymous. To inventors. Scientific American, 67(16):251, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n16/pdf/scientificamerican10151892-251b.pdf.

Anonymous:1892:NQbo

[3049] Anonymous. Notes and queries. Scientific American, 67(16):251, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n16/pdf/scientificamerican10151892-251a.pdf.

Anonymous:1892:IIbq

[3050] Anonymous. Index of inventions. Scientific American, 67(16): 251–252, October 15, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n16/pdf/scientificamerican10151892-
251c.pdf.

Anonymous:1892:ELV

[3051] Anonymous. Electric light and vegetable growth. Scientific American, 67(17):256–257, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
256.pdf.

Anonymous:1892:DCEa

[3052] Anonymous. A device to close elevator shafts. Scientific American, 67(17):258, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
258a.pdf.

Anonymous:1892:IOCb

[3053] Anonymous. An improved oil can. Scientific American, 67(17):258, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n17/pdf/scientificamerican10221892-258.pdf.

Anonymous:1892:NPV

[3054] Anonymous. New principle ventilating water closet. Scientific American, 67(17):258, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
258b.pdf.

Anonymous:1892:IIEa

[3055] Anonymous. The invention of the incandescent electric lamp awarded to Edison. Scientific American, 67(17):259, October 22, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n17/pdf/scientificamerican10221892-259.pdf.

Anonymous:1892:Cbi

[3056] Anonymous. Correspondence. Scientific American, 67(17):260–261, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n17/pdf/scientificamerican10221892-260.pdf.

Anonymous:1892:AEP

[3057] Anonymous. Aluminum electro-plating in architecture. Scientific American, 67(17):261, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
261.pdf.

Anonymous:1892:NAS

[3058] Anonymous. A new air ship. Scientific American, 67(17):262, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n17/pdf/scientificamerican10221892-262.pdf.

Anonymous:1892:WCEb

[3059] Anonymous. The World’s Columbian Exposition. Scientific American, 67(17):262–265, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
262a.pdf.

Anonymous:1892:VT

[3060] Anonymous. Vibratory therapeutics. Scientific American, 67 (17):265, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
265.pdf.

Anonymous:1892:BNBbb

[3061] Anonymous. Books: New books and publications. Scientific American, 67(17):266, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
266a.pdf.

Anonymous:1892:RPIbp

[3062] Anonymous. Recently patented inventions. Scientific American, 67(17):266, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
266.pdf.

Anonymous:1892:BPbp

[3063] Anonymous. Business and personal. Scientific American, 67 (17):267, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
267.pdf.

Anonymous:1892:Ibp

[3064] Anonymous. To inventors. Scientific American, 67(17):267, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n17/pdf/scientificamerican10221892-267b.pdf.

Anonymous:1892:NQbp

[3065] Anonymous. Notes and queries. Scientific American, 67(17):267, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n17/pdf/scientificamerican10221892-267a.pdf.

Anonymous:1892:IIbr

[3066] Anonymous. Index of inventions. Scientific American, 67(17): 267–268, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
267c.pdf.

Anonymous:1892:PNPd

[3067] Anonymous. Patent notes: Patents! Scientific American, 67 (17):270, October 22, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n17/pdf/scientificamerican10221892-
270.pdf.

Anonymous:1892:DCEb

[3068] Anonymous. Dedication of the Columbian Exposition buildings at Chicago. Scientific American, 67(18):272, October 29, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n18/pdf/scientificamerican10291892-272.pdf.

Anonymous:1892:TOT

[3069] Anonymous. Talking one thousand miles. Scientific American, 67 (18):272–273, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
272a.pdf.

Anonymous:1892:PPN

[3070] Anonymous. Position of the planets in November. Scientific American, 67(18):273, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
273.pdf.

Anonymous:1892:HDS

[3071] Anonymous. A hoisting device and support. Scientific American, 67(18):274, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
274a.pdf.

Anonymous:1892:IGD

[3072] Anonymous. An improved grain dump. Scientific American, 67 (18):274, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
274b.pdf.

Anonymous:1892:ISCc

[3073] Anonymous. An improved sugar cane mill. Scientific American, 67(18):274, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
274.pdf.

Anonymous:1892:IHR

[3074] Anonymous. An improved hydraulic ram. Scientific American, 67 (18):274–275, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
274c.pdf.

Anonymous:1892:CTb

[3075] Anonymous. Conjoined twins. Scientific American, 67(18):275–276, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n18/pdf/scientificamerican10291892-275.pdf.

Hopkins:1892:SN

[3076] Geo. Hopkins. Scientific notes. Scientific American, 67(18):276, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n18/pdf/scientificamerican10291892-276.pdf.

Anonymous:1892:SWP

[3077] Anonymous. The statue of William Penn for the Philadelphia Public Buildings. Scientific American, 67(18):277, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n18/pdf/scientificamerican10291892-277.pdf.

Anonymous:1892:LAB

[3078] Anonymous. The Liberal Arts Building. Scientific American, 67 (18):278, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
278.pdf.

Anonymous:1892:NEL

[3079] Anonymous. A new electric launch. Scientific American, 67(18): 279–281, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
279.pdf.

Anonymous:1892:DS

[3080] Anonymous. The destruction of Sanguir. Scientific American, 67 (18):281–282, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
281.pdf.

Anonymous:1892:RPIbq

[3081] Anonymous. Recently patented inventions. Scientific American, 67(18):282, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
282.pdf.

Anonymous:1892:BNBbc

[3082] Anonymous. Books: New books and publications. Scientific American, 67(18):283, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
283b.pdf.

Anonymous:1892:BPbq

[3083] Anonymous. Business and personal. Scientific American, 67 (18):283, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
283.pdf.

Anonymous:1892:Ibq

[3084] Anonymous. To inventors. Scientific American, 67(18):283, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n18/pdf/scientificamerican10291892-283c.pdf.

Anonymous:1892:NQbq

[3085] Anonymous. Notes and queries. Scientific American, 67(18):283, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n18/pdf/scientificamerican10291892-283a.pdf.

Anonymous:1892:IIbs

[3086] Anonymous. Index of inventions. Scientific American, 67(18): 283–284, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
283d.pdf.

Anonymous:1892:PNPe

[3087] Anonymous. Patent notes: Patents! Scientific American, 67 (18):286, October 29, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n18/pdf/scientificamerican10291892-
286.pdf.

Anonymous:1892:MWG

[3088] Anonymous. The manufacture of wire glass at Tacony, Philadelphia, PA. Scientific American, 67(19):287–288, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n19/pdf/scientificamerican11051892-287.pdf.

Anonymous:1892:NYB

[3089] Anonymous. New York and Brooklyn Bridge Cable Railway operations during the week of the Columbus Festival. Scientific American, 67(19):288, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
288.pdf.

Anonymous:1892:PWFc

[3090] Anonymous. Photography at the World’s Fair. Scientific American, 67(19):288–289, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
288a.pdf.

Anonymous:1892:PJW

[3091] Anonymous. Plainfield, N.J., water works. Scientific American, 67 (19):290–291, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
290.pdf.

Anonymous:1892:NBS

[3092] Anonymous. New Broad Street Station at Philadelphia. Scientific American, 67(19):291, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
291.pdf.

Anonymous:1892:ICP

[3093] Anonymous. An improved cotton picker. Scientific American, 67 (19):291–292, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
291a.pdf.

Anonymous:1892:ACSc

[3094] Anonymous. An adjustable corner shelf. Scientific American, 67 (19):292–293, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
292.pdf.

Anonymous:1892:HHb

[3095] Anonymous. Heron hunting. Scientific American, 67(19):294, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n19/pdf/scientificamerican11051892-294.pdf.

Anonymous:1892:TBN

[3096] Anonymous. Talking between New York and Chicago. Scientific American, 67(19):295, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
295.pdf.

Graham:1892:PDA

[3097] James Graham. Proposed design for an Atlantic passenger steamer. Scientific American, 67(19):295–296, November 5, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n19/pdf/scientificamerican11051892-295a.pdf.

Anonymous:1892:WSP

[3098] Anonymous. Why are stereoscopic prints transposed? Scientific American, 67(19):297–298, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
297.pdf.

Anonymous:1892:RPIbr

[3099] Anonymous. Recently patented inventions. Scientific American, 67(19):298, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
298.pdf.

Anonymous:1892:BNBbd

[3100] Anonymous. Books: New books and publications. Scientific American, 67(19):299, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
299b.pdf.

Anonymous:1892:BPbr

[3101] Anonymous. Business and personal. Scientific American, 67 (19):299, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
299.pdf.

Anonymous:1892:IIbt

[3102] Anonymous. Index of inventions. Scientific American, 67(19): 299, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
299c.pdf.

Anonymous:1892:NQbr

[3103] Anonymous. Notes and queries. Scientific American, 67(19): 299, November 5, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n19/pdf/scientificamerican11051892-
299a.pdf.

Anonymous:1892:ESS

[3104] Anonymous. Electric snow sweeper. Scientific American, 67 (20):303, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
303.pdf.

Anonymous:1892:PLE

[3105] Anonymous. The physical laboratories of Europe. Scientific American, 67(20):304, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
304.pdf.

Anonymous:1892:GPI

[3106] Anonymous. Growth of the Pratt Institute. Scientific American, 67 (20):304–305, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
304a.pdf.

Anonymous:1892:ACP

[3107] Anonymous. Apparatus for congealing perfume extracts. Scientific American, 67(20):306, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
306a.pdf.

Anonymous:1892:ILH

[3108] Anonymous. An improved lock hinge. Scientific American, 67 (20):306, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
306b.pdf.

Anonymous:1892:ISVb

[3109] Anonymous. An inexpensive stovepipe ventilator. Scientific American, 67(20):306, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
306.pdf.

Anonymous:1892:PCc

[3110] Anonymous. Photographic cane. Scientific American, 67(20): 307, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
307.pdf.

Anonymous:1892:LT

[3111] Anonymous. Lord Tennyson. Scientific American, 67(20):307– 308, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
307a.pdf.

Anonymous:1892:PAb

[3112] Anonymous. Ptychoraphis Augusta. Scientific American, 67(20): 308–309, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
308.pdf.

Anonymous:1892:LFW

[3113] Anonymous. Laying of flexible water pipes. Scientific American, 67(20):310, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
310.pdf.

Anonymous:1892:WG

[3114] Anonymous. A winter in Greenland. Scientific American, 67(20): 311–313, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
311.pdf.

Tanner:1892:FPY

[3115] A. M. Tanner. A French patent of the year 1860 for a barbed wire fence. Scientific American, 67(20):313–314, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n20/pdf/scientificamerican11121892-313.pdf.

Anonymous:1892:RPIbs

[3116] Anonymous. Recently patented inventions. Scientific American, 67(20):314, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
314.pdf.

Anonymous:1892:BNBbe

[3117] Anonymous. Books: New books and publications. Scientific American, 67(20):315, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
315b.pdf.

Anonymous:1892:BPbs

[3118] Anonymous. Business and personal. Scientific American, 67 (20):315, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
315.pdf.

Anonymous:1892:Ibr

[3119] Anonymous. To inventors. Scientific American, 67(20):315, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n20/pdf/scientificamerican11121892-315c.pdf.

Anonymous:1892:NQbs

[3120] Anonymous. Notes and queries. Scientific American, 67(20): 315, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
315a.pdf.

Anonymous:1892:IIbu

[3121] Anonymous. Index of inventions. Scientific American, 67(20): 315–316, November 12, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n20/pdf/scientificamerican11121892-
315d.pdf.

Anonymous:1892:AMN

[3122] Anonymous. The American Museum of Natural History. Scientific American, 67(21):320, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
320a.pdf.

Anonymous:1892:LCC

[3123] Anonymous. Launch of the cruiser Cincinnati. Scientific American, 67(21):320, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
320.pdf.

Anonymous:1892:PAS

[3124] Anonymous. Progress of American steamship building. Scientific American, 67(21):320, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
320b.pdf.

Anonymous:1892:PCS

[3125] Anonymous. Professor Charles A. Seeley. Scientific American, 67 (21):320, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
320c.pdf.

Anonymous:1892:NPc

[3126] Anonymous. A naturalist in La plata. Scientific American, 67 (21):321, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
321.pdf.

Anonymous:1892:IRC

[3127] Anonymous. An improved railroad car. Scientific American, 67 (21):322, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
322.pdf.

Anonymous:1892:LLF

[3128] Anonymous. A large Lake freight boat. Scientific American, 67 (21):322–323, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
322a.pdf.

Anonymous:1892:DCM

[3129] Anonymous. A double chisel mortising and boring machine. Scientific American, 67(21):323–324, November 19, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n21/pdf/scientificamerican11191892-323.pdf.

Anonymous:1892:Cbj

[3130] Anonymous. Correspondence. Scientific American, 67(21):325, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n21/pdf/scientificamerican11191892-325a.pdf.

Anonymous:1892:EWP

[3131] Anonymous. Engineering works, present and future, in the City of New York. Scientific American, 67(21):325, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n21/pdf/scientificamerican11191892-325.pdf.

Anonymous:1892:RBC

[3132] Anonymous. Residence at Bridgeport, Conn. Scientific American, 67 (21):326–327, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
326.pdf.

Anonymous:1892:LSR

[3133] Anonymous. The life-saving rocket grapnel. Scientific American, 67 (21):327–329, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
327.pdf.

Anonymous:1892:ASA

[3134] Anonymous. An aerial and sub-aquatic torpedo. Scientific American, 67(21):329, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
329.pdf.

Anonymous:1892:BNBbf

[3135] Anonymous. Books: New books and publications. Scientific American, 67(21):330, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
330a.pdf.

Anonymous:1892:BPbt

[3136] Anonymous. Business and personal. Scientific American, 67 (21):330, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
330b.pdf.

Anonymous:1892:RPIbt

[3137] Anonymous. Recently patented inventions. Scientific American, 67(21):330, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
330.pdf.

Anonymous:1892:NQbt

[3138] Anonymous. Notes and queries. Scientific American, 67(21): 330–331, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
330c.pdf.

Anonymous:1892:Ibu

[3139] Anonymous. To inventors. Scientific American, 67(21):331, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n21/pdf/scientificamerican11191892-331.pdf.

Anonymous:1892:IIbv

[3140] Anonymous. Index of inventions. Scientific American, 67(21): 331, November 19, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n21/pdf/scientificamerican11191892-
331a.pdf.

Anonymous:1892:LGE

[3141] Anonymous. A Laboe gas engine. Scientific American, 67(22): 335, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
335.pdf.

Anonymous:1892:RPT

[3142] Anonymous. A recent projectile trial. Scientific American, 67 (22):336, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
336.pdf.

Anonymous:1892:Cbk

[3143] Anonymous. The comet. Scientific American, 67(22):336–337, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n22/pdf/scientificamerican11261892-336a.pdf.

Anonymous:1892:IEW

[3144] Anonymous. An improved envelope and wrapper. Scientific American, 67(22):338, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
338a.pdf.

Anonymous:1892:SSP

[3145] Anonymous. Sargent’s smoke preventing furnace. Scientific American, 67(22):338, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
338.pdf.

Anonymous:1892:EMTa

[3146] Anonymous. Electro-mechanical telephone exchange. Scientific American, 67(22):339, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
339.pdf.

Anonymous:1892:Cbl

[3147] Anonymous. Correspondence. Scientific American, 67(22):340– 341, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
340.pdf.

Anonymous:1892:LDA

[3148] Anonymous. Large dynamo armature for direct driving. Scientific American, 67(22):342, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
342.pdf.

Anonymous:1892:JJR

[3149] Anonymous. The Jaffa and Jerusalem Railway. Scientific American, 67(22):342–343, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
342a.pdf.

Anonymous:1892:IWG

[3150] Anonymous. Improved winding gear for seventy-five ton shear legs. Scientific American, 67(22):345, November 26, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n22/pdf/scientificamerican11261892-345.pdf.

Anonymous:1892:BNBbg

[3151] Anonymous. Books: New books and publications. Scientific American, 67(22):346, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
346a.pdf.

Anonymous:1892:RPIbu

[3152] Anonymous. Recently patented inventions. Scientific American, 67(22):346, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
346.pdf.

Anonymous:1892:BPbu

[3153] Anonymous. Business and personal. Scientific American, 67 (22):347, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
347.pdf.

Anonymous:1892:Ibv

[3154] Anonymous. To inventors. Scientific American, 67(22):347, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n22/pdf/scientificamerican11261892-347b.pdf.

Anonymous:1892:NQbu

[3155] Anonymous. Notes and queries. Scientific American, 67(22): 347, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
347a.pdf.

Anonymous:1892:IIbw

[3156] Anonymous. Index of inventions. Scientific American, 67(22): 347–348, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
347c.pdf.

Anonymous:1892:PEI

[3157] Anonymous. Prince Edward Island Tunnel test. Scientific American, 67(22):443–444, November 26, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n22/pdf/scientificamerican11261892-
343.pdf.

Anonymous:1892:LGS

[3158] Anonymous. Loss of a great ship of war. Scientific American, 67 (23):351–352, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
351.pdf.

Anonymous:1892:ERT

[3159] Anonymous. Extension of rapid transit facilities in New York City. Scientific American, 67(23):352–353, December 3, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n23/pdf/scientificamerican12031892-352.pdf.

Anonymous:1892:NGFb

[3160] Anonymous. Notes from the great fair. Scientific American, 67 (23):353–354, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
353.pdf.

Anonymous:1892:IRBb

[3161] Anonymous. Inflation of rubber balls. Scientific American, 67 (23):354, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
354b.pdf.

Anonymous:1892:ISH

[3162] Anonymous. An improved sliding hatch door. Scientific American, 67(23):354, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
354a.pdf.

Anonymous:1892:STS

[3163] Anonymous. Sword trick — a star through the abdomen. Scientific American, 67(23):354, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
354.pdf.

Anonymous:1892:TSE

[3164] Anonymous. Three systems of electric distribution. Scientific American, 67(23):355, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
355.pdf.

Anonymous:1892:PPD

[3165] Anonymous. Position of the planets in December. Scientific American, 67(23):356, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
356.pdf.

Anonymous:1892:Cbm

[3166] Anonymous. Correspondence. Scientific American, 67(23):357, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n23/pdf/scientificamerican12031892-357.pdf.

Anonymous:1892:H

[3167] Anonymous. The heliochromoscope. Scientific American, 67(23): 358, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
358.pdf.

Anonymous:1892:SEM

[3168] Anonymous. De Susini’s ether motor. Scientific American, 67 (23):359–361, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
359.pdf.

Anonymous:1892:BNBbh

[3169] Anonymous. Books: New books and publications. Scientific American, 67(23):362, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
362a.pdf.

Anonymous:1892:RPIbv

[3170] Anonymous. Recently patented inventions. Scientific American, 67(23):362, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
362.pdf.

Anonymous:1892:BPbv

[3171] Anonymous. Business and personal. Scientific American, 67 (23):363, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
363.pdf.

Anonymous:1892:NQbv

[3172] Anonymous. Notes and queries. Scientific American, 67(23):363, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n23/pdf/scientificamerican12031892-363a.pdf.

Anonymous:1892:IIbx

[3173] Anonymous. Index of inventions. Scientific American, 67(23): 363–364, December 3, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n23/pdf/scientificamerican12031892-
363b.pdf.

Anonymous:1892:CBS

[3174] Anonymous. A carriage builders’ school. Scientific American, 67 (24):368, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
368b.pdf.

Anonymous:1892:DCb

[3175] Anonymous. The Danforth Case. Scientific American, 67(24): 368, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
368.pdf.

Anonymous:1892:GR

[3176] Anonymous. Good roads. Scientific American, 67(24):368, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n24/pdf/scientificamerican12101892-368a.pdf.

Anonymous:1892:NCC

[3177] Anonymous. The Nicaragua Canal Convention. Scientific American, 67(24):369, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
369.pdf.

Anonymous:1892:CTT

[3178] Anonymous. A coin-taking and ticket-delivering box. Scientific American, 67(24):370, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
370c.pdf.

Anonymous:1892:IDW

[3179] Anonymous. An improved dish washer. Scientific American, 67 (24):370, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
370.pdf.

Anonymous:1892:IFEc

[3180] Anonymous. An improved fire escape. Scientific American, 67 (24):370, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
370a.pdf.

Anonymous:1892:IWV

[3181] Anonymous. An improved wooden vessel. Scientific American, 67 (24):370, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
370b.pdf.

Anonymous:1892:EMTb

[3182] Anonymous. English machine tools for America. Scientific American, 67(24):371–373, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
371.pdf.

Anonymous:1892:Cbn

[3183] Anonymous. Correspondence. Scientific American, 67(24):373, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n24/pdf/scientificamerican12101892-373.pdf.

Underwood:1892:MRS

[3184] B. G. Underwood. Manufacture of rubber shoes. Scientific American, 67(24):374–375, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
374.pdf.

Anonymous:1892:SCH

[3185] Anonymous. The Southern Cotton Harvester. Scientific American, 67 (24):375–376, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
375.pdf.

Anonymous:1892:JGT

[3186] Anonymous. Jay Gould’s tomb. Scientific American, 67(24): 377–378, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
377.pdf.

Anonymous:1892:RPIbw

[3187] Anonymous. Recently patented inventions. Scientific American, 67(24):378, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
378.pdf.

Anonymous:1892:BNBbi

[3188] Anonymous. Books: New books and publications. Scientific American, 67(24):378–379, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
378a.pdf.

Anonymous:1892:BPbw

[3189] Anonymous. Business and personal. Scientific American, 67 (24):379, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
379.pdf.

Anonymous:1892:Ibw

[3190] Anonymous. To inventors. Scientific American, 67(24):379, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n24/pdf/scientificamerican12101892-379b.pdf.

Anonymous:1892:NQbw

[3191] Anonymous. Notes and queries. Scientific American, 67(24): 379, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
379a.pdf.

Anonymous:1892:IIby

[3192] Anonymous. Index of inventions. Scientific American, 67(24): 379–380, December 10, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n24/pdf/scientificamerican12101892-
379c.pdf.

Anonymous:1892:PTM

[3193] Anonymous. Pneumatic tubes for mail service. Scientific American, 67(25):384, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
384.pdf.

Kirkwood:1892:MN

[3194] Daniel Kirkwood. The meteors of November, 1892. Scientific American, 67(25):384, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
384a.pdf.

Anonymous:1892:DEW

[3195] Anonymous. Dr. Ernst Werner Siemens. Scientific American, 67 (25):384–385, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
384b.pdf.

Anonymous:1892:HDb

[3196] Anonymous. The Henrion dynamo. Scientific American, 67(25): 386, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
386a.pdf.

Anonymous:1892:IBCa

[3197] Anonymous. An improved breast collar. Scientific American, 67 (25):386, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
386b.pdf.

Anonymous:1892:IEN

[3198] Anonymous. An improved exhaust nozzle. Scientific American, 67(25):386, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
386.pdf.

Anonymous:1892:HT

[3199] Anonymous. The hygienic treadmill. Scientific American, 67 (25):387, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
387.pdf.

Hopkins:1892:RRE

[3200] Geo. M. Hopkins. Rod and ring experiment. Scientific American, 67(25):387, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
387a.pdf.

Kenealy:1892:SCS

[3201] Annesley Kenealy. Some of the “Cholera Streets” in Hamburg. Scientific American, 67(25):387–388, December 17, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n25/pdf/scientificamerican12171892-387b.pdf.

Anonymous:1892:IRP

[3202] Anonymous. Improved recording pressure gauge. Scientific American, 67(25):388–389, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
388.pdf.

Underwood:1892:MSS

[3203] B. G. Underwood. Metropolitan sewerage system surrounding Boston. Scientific American, 67(25):390–391, December 17, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n25/pdf/scientificamerican12171892-390.pdf.

Anonymous:1892:Cbo

[3204] Anonymous. Correspondence. Scientific American, 67(25):392, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n25/pdf/scientificamerican12171892-392.pdf.

Anonymous:1892:RPIbx

[3205] Anonymous. Recently patented inventions. Scientific American, 67 (25):392–393, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
392a.pdf.

Anonymous:1892:BNBbj

[3206] Anonymous. Books: New books and publications. Scientific American, 67(25):393, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
393.pdf.

Anonymous:1892:BPbx

[3207] Anonymous. Business and personal. Scientific American, 67 (25):393, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
393a.pdf.

Anonymous:1892:Ibx

[3208] Anonymous. To inventors. Scientific American, 67(25):393, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n25/pdf/scientificamerican12171892-393c.pdf.

Anonymous:1892:NQbx

[3209] Anonymous. Notes and queries. Scientific American, 67(25): 393, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
393b.pdf.

Anonymous:1892:IIbz

[3210] Anonymous. Index of inventions. Scientific American, 67(25): 393–394, December 17, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n25/pdf/scientificamerican12171892-
393d.pdf.

Underwood:1892:NRV

[3211] B. G. Underwood. Natural resources of Virginia and West Virginia. Scientific American, 67(26):399, December 24, 1892. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v67/
n26/pdf/scientificamerican12241892-399.pdf.

Anonymous:1892:PON

[3212] Anonymous. Progress of our Navy. Scientific American, 67 (26):400, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
400.pdf.

Anonymous:1892:UPP

[3213] Anonymous. The use of photography in physical research. Scientific American, 67(26):400–402, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
400a.pdf.

Anonymous:1892:FRB

[3214] Anonymous. Fall of a railway bridge at Terre Haute. Scientific American, 67(26):402, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
402.pdf.

Anonymous:1892:TWI

[3215] Anonymous. The tallest wrought iron chimney. Scientific American, 67(26):402–403, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
402a.pdf.

Anonymous:1892:BT

[3216] Anonymous. The Bohemian twins. Scientific American, 67(26): 403, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
403.pdf.

Anonymous:1892:IAB

[3217] Anonymous. An improved automatic boiler feed. Scientific American, 67(26):403, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
403a.pdf.

Anonymous:1892:RPIby

[3218] Anonymous. Recently patented inventions. Scientific American, 67 (26):408–409, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
408.pdf.

Anonymous:1892:BNBbk

[3219] Anonymous. Books: New books and publications. Scientific American, 67(26):409, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
409.pdf.

Anonymous:1892:BPby

[3220] Anonymous. Business and personal. Scientific American, 67 (26):409, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
409a.pdf.

Anonymous:1892:Iby

[3221] Anonymous. To inventors. Scientific American, 67(26):409, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v67/n26/pdf/scientificamerican12241892-409c.pdf.

Anonymous:1892:NQby

[3222] Anonymous. Notes and queries. Scientific American, 67(26): 409, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
409b.pdf.

Anonymous:1892:IIca

[3223] Anonymous. Index of inventions. Scientific American, 67(26): 409–410, December 24, 1892. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v67/n26/pdf/scientificamerican12241892-
409d.pdf.

Anonymous:1893:SSO

[3224] Anonymous. Safety suggestions for ocean steamers. Scientific American, 68(1):2–3, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
2.pdf.

Anonymous:1893:BWC

[3225] Anonymous. A building to withstand cyclones, etc. Scientific American, 68(1):4, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
4a.pdf.

Anonymous:1893:ISR

[3226] Anonymous. An improved side and roofing tile. Scientific American, 68(1):4, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
4b.pdf.

Anonymous:1893:SFB

[3227] Anonymous. Stone’s family bread kneader. Scientific American, 68(1):4, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-4.pdf.

Anonymous:1893:IEL

[3228] Anonymous. An improved electric light signaling apparatus. Scientific American, 68(1):5, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
5.pdf.

Anonymous:1893:NCM

[3229] Anonymous. A novel cutting-off machine. Scientific American, 68(1):5, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-5a.pdf.

Anonymous:1893:ISTa

[3230] Anonymous. An improved speaking tube. Scientific American, 68(1):6, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-6.pdf.

Anonymous:1893:IWT

[3231] Anonymous. An improved Watchmaker’s tool. Scientific American, 68(1):6, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
6a.pdf.

Anonymous:1893:Ca

[3232] Anonymous. Correspondence. Scientific American, 68(1):7, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-7.pdf.

Anonymous:1893:PRSa

[3233] Anonymous. The Potsdam Red Sandstone Quarries. Scientific American, 68(1):8–10, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
8.pdf.

Anonymous:1893:MTB

[3234] Anonymous. Manufacture of trunk and bookbinders boards. Scientific American, 68(1):10, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
10.pdf.

Anonymous:1893:SAI

[3235] Anonymous. Skating on artificial ice. Scientific American, 68(1):11, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-11.pdf.

Anonymous:1893:BNBa

[3236] Anonymous. Books: New books and publications. Scientific American, 68(1):12, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
12a.pdf.

Anonymous:1893:BPa

[3237] Anonymous. Business and personal. Scientific American, 68(1):12, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-12b.pdf.

Anonymous:1893:RPIa

[3238] Anonymous. Recently patented inventions. Scientific American, 68(1):12, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n1/pdf/scientificamerican01071893-
12.pdf.

Anonymous:1893:NQa

[3239] Anonymous. Notes and queries. Scientific American, 68(1):12–13, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-12c.pdf.

Anonymous:1893:Ia

[3240] Anonymous. To inventors. Scientific American, 68(1):13, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-13.pdf.

Anonymous:1893:IIa

[3241] Anonymous. Index of inventions. Scientific American, 68(1):13, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-13a.pdf.

Anonymous:1893:PNPa

[3242] Anonymous. Patent notes: Patents. Scientific American, 68(1):16, January 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n1/pdf/scientificamerican01071893-16.pdf.

Anonymous:1893:BIA

[3243] Anonymous. The Brooklyn Institute of Arts and Sciences. Scientific American, 68(2):18–19, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
18.pdf.

Anonymous:1893:IBB

[3244] Anonymous. An improved bit brace. Scientific American, 68(2):20, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n2/pdf/scientificamerican01141893-20a.pdf.

Anonymous:1893:ICEa

[3245] Anonymous. An improved conveyer or elevator. Scientific American, 68(2):20, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
20b.pdf.

Anonymous:1893:ISCa

[3246] Anonymous. An improved school chair and desk. Scientific American, 68(2):20, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
20.pdf.

Anonymous:1893:MWS

[3247] Anonymous. A machine to wash and screen coal. Scientific American, 68(2):20, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
20c.pdf.

Anonymous:1893:IAW

[3248] Anonymous. An improved automatic water gate. Scientific American, 68(2):21, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
21.pdf.

Anonymous:1893:DWS

[3249] Anonymous. Dr. Werner Von Siemens. Scientific American, 68 (2):21–22, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
21a.pdf.

Anonymous:1893:IMa

[3250] Anonymous. An improved motor. Scientific American, 68(2):22, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n2/pdf/scientificamerican01141893-22.pdf.

Anonymous:1893:MSL

[3251] Anonymous. The modern steam laundry. Scientific American, 68(2):23, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
23.pdf.

Anonymous:1893:BRS

[3252] Anonymous. Breakage and repair at sea of the main shaft of the steamship Umbria. Scientific American, 68(2):24, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n2/pdf/scientificamerican01141893-24.pdf.

Anonymous:1893:KPF

[3253] Anonymous. The kangaroo as a prize fighter. Scientific American, 68(2):25–26, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
25.pdf.

Anonymous:1893:NTa

[3254] Anonymous. The “National” typewriter. Scientific American, 68 (2):27, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
27.pdf.

Anonymous:1893:BNBb

[3255] Anonymous. Books: New books and publications. Scientific American, 68(2):28, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
28a.pdf.

Anonymous:1893:RPIb

[3256] Anonymous. Recently patented inventions. Scientific American, 68(2):28, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n2/pdf/scientificamerican01141893-
28.pdf.

Anonymous:1893:BPb

[3257] Anonymous. Business and personal. Scientific American, 68(2):29, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n2/pdf/scientificamerican01141893-29.pdf.

Anonymous:1893:IIb

[3258] Anonymous. Index of inventions. Scientific American, 68(2):29, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n2/pdf/scientificamerican01141893-29a.pdf.

Anonymous:1893:PNPb

[3259] Anonymous. Patent notes: Patents. Scientific American, 68(2):32, January 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n2/pdf/scientificamerican01141893-32.pdf.

Anonymous:1893:PWF

[3260] Anonymous. Progress of worlds fair work at Chicago. Scientific American, 68(3):34–35, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
34.pdf.

Anonymous:1893:IBT

[3261] Anonymous. An improved belt tightener. Scientific American, 68(3):36, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
36a.pdf.

Anonymous:1893:ICV

[3262] Anonymous. An improved cooking vessel. Scientific American, 68(3):36, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
36b.pdf.

Anonymous:1893:NMI

[3263] Anonymous. A novel musical instrument. Scientific American, 68(3):36, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
36c.pdf.

Anonymous:1893:SCF

[3264] Anonymous. A simple, compact folding seat. Scientific American, 68(3):36, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
36.pdf.

Anonymous:1893:IFT

[3265] Anonymous. An interesting flower trick. Scientific American, 68 (3):36–37, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
36d.pdf.

Anonymous:1893:VSa

[3266] Anonymous. Variety in sleighs. Scientific American, 68(3):37, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n3/pdf/scientificamerican01211893-37.pdf.

Anonymous:1893:RWS

[3267] Anonymous. The Russian war ship Ruric. Scientific American, 68(3):37–38, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
37a.pdf.

Anonymous:1893:PRSb

[3268] Anonymous. The Potsdam Red Sandstone Company’s water wheel. Scientific American, 68(3):38, January 21, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n3/pdf/scientificamerican01211893-38.pdf.

Anonymous:1893:Cb

[3269] Anonymous. Correspondence. Scientific American, 68(3):39, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n3/pdf/scientificamerican01211893-39.pdf.

Anonymous:1893:FM

[3270] Anonymous. Fireworks in miniature. Scientific American, 68 (3):40, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
40a.pdf.

Hopkins:1893:IVL

[3271] Geo. M. Hopkins. Instrument for viewing lantern slides. Scientific American, 68(3):40, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
40.pdf.

Pruyn:1893:NEE

[3272] Henry S. Pruyn. A new elevated electric railway system. Scientific American, 68(3):40–41, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
40b.pdf.

Wilson:1893:GJM

[3273] Walfred W. Wilson. The great jetties at the mouth of the Brazos River. Scientific American, 68(3):41–42, January 21, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n3/pdf/scientificamerican01211893-41.pdf.

Anonymous:1893:PN

[3274] Anonymous. Photographic notes. Scientific American, 68(3): 42–43, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
42.pdf.

Sloane:1893:EPV

[3275] T. O’conor Sloane. Experiment in persistence of vision. Scientific American, 68(3):43, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
43.pdf.

Anonymous:1893:BNBc

[3276] Anonymous. Books: New books and publications. Scientific American, 68(3):44, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
44a.pdf.

Anonymous:1893:RPIc

[3277] Anonymous. Recently patented inventions. Scientific American, 68(3):44, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n3/pdf/scientificamerican01211893-
44.pdf.

Anonymous:1893:BPc

[3278] Anonymous. Business and personal. Scientific American, 68(3):45, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n3/pdf/scientificamerican01211893-45.pdf.

Anonymous:1893:Ib

[3279] Anonymous. To inventors. Scientific American, 68(3):45, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n3/pdf/scientificamerican01211893-45b.pdf.

Anonymous:1893:NQb

[3280] Anonymous. Notes and queries. Scientific American, 68(3):45, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n3/pdf/scientificamerican01211893-45a.pdf.

Anonymous:1893:IIc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n3/pdf/scientificamerican01211893-45c.pdf.

Anonymous:1893:PNPc

[3282] Anonymous. Patent notes: Patents. Scientific American, 68(3):48, January 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n3/pdf/scientificamerican01211893-48.pdf.

Anonymous:1893:ASC

[3283] Anonymous. American Society of Civil Engineers. Scientific American, 68(4):50, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
50.pdf.

Anonymous:1893:PAEa

[3284] Anonymous. Proposed additions to the elevated steam street railways in New York. Scientific American, 68(4):50, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-50a.pdf.

Anonymous:1893:UPM

[3285] Anonymous. Unclean paper money a vehicle for the spread of disease. Scientific American, 68(4):50–51, January 28, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n4/pdf/scientificamerican01281893-50b.pdf.

Anonymous:1893:BGT

[3286] Anonymous. A belt guide and tightener. Scientific American, 68(4):52, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
52a.pdf.

Anonymous:1893:IPA

[3287] Anonymous. An improved pedal attachment. Scientific American, 68(4):52, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
52.pdf.

Anonymous:1893:MLF

[3288] Anonymous. Monumental luminous fountains. Scientific American, 68(4):52, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
52b.pdf.

Anonymous:1893:FOP

[3289] Anonymous. The first overhead-propelled street railway. Scientific American, 68(4):53, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
53.pdf.

Anonymous:1893:IMRa

[3290] Anonymous. Improved means of running millstones. Scientific American, 68(4):53, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
53a.pdf.

Anonymous:1893:CY

[3291] Anonymous. Charles T. Yerkes. Scientific American, 68(4):54, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-54.pdf.

Anonymous:1893:Cc

[3292] Anonymous. Correspondence. Scientific American, 68(4):55, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-55.pdf.

Anonymous:1893:IMb

[3293] Anonymous. An improved manger. Scientific American, 68(4):55, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-55a.pdf.

Anonymous:1893:HBS

[3294] Anonymous. How the broken shaft of the Cunard steamship Umbria was repaired. Scientific American, 68(4):56–59, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-56.pdf.

Anonymous:1893:BNBd

[3295] Anonymous. Books: New books and publications. Scientific American, 68(4):60, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
60a.pdf.

Anonymous:1893:RPId

[3296] Anonymous. Recently patented inventions. Scientific American, 68(4):60, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n4/pdf/scientificamerican01281893-
60.pdf.

Anonymous:1893:BPd

[3297] Anonymous. Business and personal. Scientific American, 68(4):61, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-61.pdf.

Anonymous:1893:Ic

[3298] Anonymous. To inventors. Scientific American, 68(4):61, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-61b.pdf.

Anonymous:1893:NQc

[3299] Anonymous. Notes and queries. Scientific American, 68(4):61, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-61a.pdf.

Anonymous:1893:IId

[3300] Anonymous. Index of inventions. Scientific American, 68(4):61–63, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-61c.pdf.

Anonymous:1893:PNPd

[3301] Anonymous. Patent notes: Patents. Scientific American, 68(4):64, January 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n4/pdf/scientificamerican01281893-64.pdf.

Hovey:1893:MCM

[3302] E. O. Hovey. The Marble Cave of Missouri. Scientific American, 68(5):65, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
65.pdf.

Anonymous:1893:ABD

[3303] Anonymous. Another bill to discourage the inventor and favor the infringer. Scientific American, 68(5):66, February 4, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n5/pdf/scientificamerican02041893-66.pdf.

Anonymous:1893:EC

[3304] Anonymous. The exclusion of Chinese. Scientific American, 68 (5):66–67, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
66a.pdf.

Anonymous:1893:WCEa

[3305] Anonymous. The World’s Columbian Exposition. Scientific American, 68(5):68, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
68.pdf.

Anonymous:1893:IBa

[3306] Anonymous. An improved bicycle. Scientific American, 68(5):69, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n5/pdf/scientificamerican02041893-69.pdf.

Anonymous:1893:LDCa

[3307] Anonymous. Lightning damages a chimney. Scientific American, 68(5):69, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
69a.pdf.

Anonymous:1893:NSBa

[3308] Anonymous. A novel steam boiler. Scientific American, 68(5):69, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n5/pdf/scientificamerican02041893-69b.pdf.

Anonymous:1893:TLF

[3309] Anonymous. Trolling with live fish bait. Scientific American, 68(5):71, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
71.pdf.

Anonymous:1893:Cd

[3310] Anonymous. Correspondence. Scientific American, 68(5):71–72, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n5/pdf/scientificamerican02041893-71a.pdf.

Anonymous:1893:NTb

[3311] Anonymous. A novel toy. Scientific American, 68(5):72, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n5/pdf/scientificamerican02041893-72a.pdf.

Anonymous:1893:PC

[3312] Anonymous. Performing cats. Scientific American, 68(5):72, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n5/pdf/scientificamerican02041893-72.pdf.

Anonymous:1893:AGO

[3313] Anonymous. An automatic gate operating device. Scientific American, 68(5):73, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
73.pdf.

Anonymous:1893:ISAa

[3314] Anonymous. An improved sketching apparatus. Scientific American, 68(5):73, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
73b.pdf.

Anonymous:1893:SSEa

[3315] Anonymous. Snow shoe exercise in the German Army. Scientific American, 68(5):73, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
73a.pdf.

Anonymous:1893:PPF

[3316] Anonymous. Position of the planets in February. Scientific American, 68(5):74, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
74.pdf.

Lydekker:1893:AAV

[3317] R. Lydekker. Armadillos and aard-varks. Scientific American, 68 (5):74–75, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
74a.pdf.

Anonymous:1893:BNBe

[3318] Anonymous. Books: New books and publications. Scientific American, 68(5):76, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
76a.pdf.

Anonymous:1893:RPIe

[3319] Anonymous. Recently patented inventions. Scientific American, 68(5):76, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
76.pdf.

Anonymous:1893:BPe

[3320] Anonymous. Business and personal. Scientific American, 68 (5):77, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
77.pdf.

Anonymous:1893:Id

[3321] Anonymous. To inventors. Scientific American, 68(5):77, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n5/pdf/scientificamerican02041893-77b.pdf.

Anonymous:1893:NQd

[3322] Anonymous. Notes and queries. Scientific American, 68(5):77, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n5/pdf/scientificamerican02041893-77a.pdf.

Anonymous:1893:IIe

[3323] Anonymous. Index of inventions. Scientific American, 68(5): 77–79, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
77c.pdf.

Anonymous:1893:PNPe

[3324] Anonymous. Patent notes: Patents. Scientific American, 68 (5):80, February 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n5/pdf/scientificamerican02041893-
80.pdf.

Anonymous:1893:HIT

[3325] Anonymous. The Hawaiian Islands and their annexation. Scientific American, 68(6):82, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
82.pdf.

Anonymous:1893:EIL

[3326] Anonymous. The Edison Incandescent Lamp Patent attacked. Scientific American, 68(6):82–83, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
82a.pdf.

Anonymous:1893:MMa

[3327] Anonymous. The money maker. Scientific American, 68(6):84, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n6/pdf/scientificamerican02111893-84.pdf.

Anonymous:1893:MSS

[3328] Anonymous. The main shaft of the steamer Umbria as repaired. Scientific American, 68(6):84, February 11, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n6/pdf/scientificamerican02111893-84a.pdf.

Anonymous:1893:HSA

[3329] Anonymous. A horseshoe sharpening appliance. Scientific American, 68(6):85, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
85a.pdf.

Anonymous:1893:LSM

[3330] Anonymous. A light and simple motor. Scientific American, 68(6):85, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
85.pdf.

Anonymous:1893:IFA

[3331] Anonymous. An improved fire apparatus. Scientific American, 68 (6):85–86, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
85b.pdf.

Anonymous:1893:Ce

[3332] Anonymous. Correspondence. Scientific American, 68(6):87, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n6/pdf/scientificamerican02111893-87a.pdf.

Anonymous:1893:PRB

[3333] Anonymous. The Pecos River Bridge. Scientific American, 68 (6):87, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
87.pdf.

Anonymous:1893:TPC

[3334] Anonymous. Trial of the pneumatic cruiser Vesuvius. Scientific American, 68(6):88–91, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
88.pdf.

Anonymous:1893:FHM

[3335] Anonymous. A feat in house moving. Scientific American, 68 (6):91, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
91.pdf.

Anonymous:1893:BNBf

[3336] Anonymous. Books: New books and publications. Scientific American, 68(6):92, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
92a.pdf.

Anonymous:1893:RPIf

[3337] Anonymous. Recently patented inventions. Scientific American, 68(6):92, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
92.pdf.

Anonymous:1893:BPf

[3338] Anonymous. Business and personal. Scientific American, 68(6): 93, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
93.pdf.

Anonymous:1893:NQe

[3339] Anonymous. Notes and queries. Scientific American, 68(6):93, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n6/pdf/scientificamerican02111893-93a.pdf.

Anonymous:1893:IIf

[3340] Anonymous. Index of inventions. Scientific American, 68(6):93– 95, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
93b.pdf.

Anonymous:1893:PNPf

[3341] Anonymous. Patent notes: Patents. Scientific American, 68(6): 96, February 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n6/pdf/scientificamerican02111893-
96.pdf.

Anonymous:1893:HDR

[3342] Anonymous. The Harbor Defense Ram Katahdin. Scientific American, 68(7):97, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
97.pdf.

Anonymous:1893:ARC

[3343] Anonymous. The annual report of the Commissioner of Patents. Scientific American, 68(7):98, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
98a.pdf.

Anonymous:1893:CAT

[3344] Anonymous. The crowning achievements of the telephone. Scientific American, 68(7):98, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
98.pdf.

Anonymous:1893:PLD

[3345] Anonymous. Progress of long distance telephony. Scientific American, 68(7):98–99, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
98b.pdf.

Anonymous:1893:EPT

[3346] Anonymous. The equitable printing telegraph. Scientific American, 68(7):99, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
99.pdf.

Anonymous:1893:IDM

[3347] Anonymous. An improved drilling machine. Scientific American, 68(7):101, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
101.pdf.

Anonymous:1893:IFE

[3348] Anonymous. An improved fire escape. Scientific American, 68 (7):101, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
101c.pdf.

Anonymous:1893:IRCa

[3349] Anonymous. An improved railway coach. Scientific American, 68(7):101, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
101a.pdf.

Anonymous:1893:SVW

[3350] Anonymous. A safety valve for water gauges. Scientific American, 68(7):101, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
101b.pdf.

Anonymous:1893:Cf

[3351] Anonymous. Correspondence. Scientific American, 68(7):102, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n7/pdf/scientificamerican02181893-102.pdf.

Anonymous:1893:APa

[3352] Anonymous. Atmospheric pressure. Scientific American, 68(7): 102–104, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
102a.pdf.

Kiely:1893:NSI

[3353] Humphrey J. Kiely. Notes on the sugar industry in Cuba. Scientific American, 68(7):104–106, February 18, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n7/pdf/scientificamerican02181893-104.pdf.

Anonymous:1893:FPG

[3354] Anonymous. The Fontainebleau proving grounds. Scientific American, 68(7):106–107, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
106.pdf.

Anonymous:1893:BNBg

[3355] Anonymous. Books: New books and publications. Scientific American, 68(7):108, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
108a.pdf.

Anonymous:1893:BPg

[3356] Anonymous. Business and personal. Scientific American, 68(7): 108, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
108b.pdf.

Anonymous:1893:RPIg

[3357] Anonymous. Recently patented inventions. Scientific American, 68(7):108, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
108.pdf.

Anonymous:1893:NQf

[3358] Anonymous. Notes and queries. Scientific American, 68(7):108– 109, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
108c.pdf.

Anonymous:1893:Ie

[3359] Anonymous. To inventors. Scientific American, 68(7):109, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n7/pdf/scientificamerican02181893-109.pdf.

Anonymous:1893:IIg

[3360] Anonymous. Index of inventions. Scientific American, 68(7): 109–111, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
109a.pdf.

Anonymous:1893:PNPg

[3361] Anonymous. Patent notes: Patents. Scientific American, 68 (7):112, February 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n7/pdf/scientificamerican02181893-
112.pdf.

Anonymous:1893:TGS

[3362] Anonymous. The two great ship canals. Scientific American, 68(8):114, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
114.pdf.

Anonymous:1893:BMP

[3363] Anonymous. The Berliner Microphone Patent. Scientific American, 68(8):114–115, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
114a.pdf.

Anonymous:1893:SAS

[3364] Anonymous. Singular accident to a steamer at sea. Scientific American, 68(8):116, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
116.pdf.

Anonymous:1893:ITEa

[3365] Anonymous. An improved traction engine. Scientific American, 68(8):117, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
117a.pdf.

Anonymous:1893:IWb

[3366] Anonymous. An improved wrench. Scientific American, 68(8): 117, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
117.pdf.

Anonymous:1893:SDE

[3367] Anonymous. A shifting device for elevators. Scientific American, 68(8):117, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
117b.pdf.

Anonymous:1893:LDCb

[3368] Anonymous. The Lovell diamond cycle. Scientific American, 68 (8):117–118, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
117c.pdf.

Anonymous:1893:LSSa

[3369] Anonymous. A labor-saving screw driver. Scientific American, 68(8):118, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
118.pdf.

Anonymous:1893:NMG

[3370] Anonymous. New method of gas supply for New York City. Scientific American, 68(8):119, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
119.pdf.

Anonymous:1893:DHW

[3371] Anonymous. Distribution of hot water at Paris. Scientific American, 68(8):120, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
120.pdf.

Anonymous:1893:NOG

[3372] Anonymous. Natural oyster grounds of South Carolina. Scientific American, 68(8):121, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
121.pdf.

Anonymous:1893:VML

[3373] Anonymous. The volcano Mauna Loa. Scientific American, 68 (8):121, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
121a.pdf.

Anonymous:1893:Cg

[3374] Anonymous. Correspondence. Scientific American, 68(8):122, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n8/pdf/scientificamerican02251893-122.pdf.

Oswald:1893:SH

[3375] F. L. Oswald. Strange hybrids. Scientific American, 68(8):123, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n8/pdf/scientificamerican02251893-123.pdf.

Anonymous:1893:BNBh

[3376] Anonymous. Books: New books and publications. Scientific American, 68(8):124, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
124a.pdf.

Anonymous:1893:BPh

[3377] Anonymous. Business and personal. Scientific American, 68(8): 124, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
124b.pdf.

Anonymous:1893:RPIh

[3378] Anonymous. Recently patented inventions. Scientific American, 68(8):124, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
124.pdf.

Anonymous:1893:NQg

[3379] Anonymous. Notes and queries. Scientific American, 68(8):124– 125, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
124c.pdf.

Anonymous:1893:If

[3380] Anonymous. To inventors. Scientific American, 68(8):125, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n8/pdf/scientificamerican02251893-125.pdf.

Anonymous:1893:IIh

[3381] Anonymous. Index of inventions. Scientific American, 68(8): 125–126, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
125a.pdf.

Anonymous:1893:PNPh

[3382] Anonymous. Patent notes: Patents. Scientific American, 68 (8):128, February 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n8/pdf/scientificamerican02251893-
128.pdf.

Anonymous:1893:AOC

[3383] Anonymous. American ocean commerce. Scientific American, 68 (9):130, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
130.pdf.

Anonymous:1893:PPMa

[3384] Anonymous. Position of the planets in March. Scientific American, 68(9):131, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
131.pdf.

Anonymous:1893:MDa

[3385] Anonymous. The manufacture of dynamite. Scientific American, 68(9):132–133, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
132.pdf.

Anonymous:1893:IGE

[3386] Anonymous. An improved gas engine. Scientific American, 68(9):133, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-133.pdf.

Anonymous:1893:ISB

[3387] Anonymous. An improved shock binder. Scientific American, 68 (9):133–134, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
133a.pdf.

Anonymous:1893:IET

[3388] Anonymous. An improved extension table. Scientific American, 68(9):134, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
134.pdf.

Woolley:1893:SCA

[3389] C. L. Woolley. Storage cells for amateurs. Scientific American, 68(9):134, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
134a.pdf.

Anonymous:1893:Ch

[3390] Anonymous. Correspondence. Scientific American, 68(9):135, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-135a.pdf.

Anonymous:1893:MLa

[3391] Anonymous. A monster locomotive. Scientific American, 68(9):135, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-135.pdf.

Anonymous:1893:RFS

[3392] Anonymous. Raising the flag on the steamship New York. Scientific American, 68(9):136–137, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
136.pdf.

Anonymous:1893:EM

[3393] Anonymous. Extinct monsters. Scientific American, 68(9):137–138, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-137.pdf.

Anonymous:1893:GED

[3394] Anonymous. The Great Exposition — doorways to the Transportation Building and the Woman’s Building. Scientific American, 68(9):138–139, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
138.pdf.

Anonymous:1893:BNBi

[3395] Anonymous. Books: New books and publications. Scientific American, 68(9):140, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
140a.pdf.

Anonymous:1893:RPIi

[3396] Anonymous. Recently patented inventions. Scientific American, 68(9):140, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n9/pdf/scientificamerican03041893-
140.pdf.

Anonymous:1893:BPi

[3397] Anonymous. Business and personal. Scientific American, 68(9):141, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-141.pdf.

Anonymous:1893:IIi

[3398] Anonymous. Index of inventions. Scientific American, 68(9):141, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-141b.pdf.

Anonymous:1893:NQh

[3399] Anonymous. Notes and queries. Scientific American, 68(9):141, March 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-141a.pdf.

Anonymous:1893:PNPi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n9/pdf/scientificamerican03041893-144.pdf.

Anonymous:1893:CLS

[3401] Anonymous. The casting of a large steam cylinder. Scientific American, 68(10):145–146, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
145.pdf.

Anonymous:1893:DCCa

[3402] Anonymous. Delicacy of a certain chemical test. Scientific American, 68(10):146, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
146.pdf.

Anonymous:1893:NC

[3403] Anonymous. The Nicaragua Canal. Scientific American, 68(10): 146–147, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
146a.pdf.

Anonymous:1893:CNRa

[3404] Anonymous. The coming Naval Review. Scientific American, 68(10):147, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
147.pdf.

Anonymous:1893:MS

[3405] Anonymous. Modern sundials. Scientific American, 68(10):148, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n10/pdf/scientificamerican03111893-148.pdf.

Anonymous:1893:HWC

[3406] Anonymous. Hickory wood carved by worms. Scientific American, 68(10):148–149, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
148a.pdf.

Anonymous:1893:CSF

[3407] Anonymous. A cheap and simple fence wire stretcher. Scientific American, 68(10):149, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
149a.pdf.

Anonymous:1893:SGF

[3408] Anonymous. A saw guide to facilitate log sawing, etc. Scientific American, 68(10):149, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
149.pdf.

Wood:1893:ICA

[3409] R. W. Wood. An induction coil for alternating currents. Scientific American, 68(10):149–150, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
149b.pdf.

Anonymous:1893:GBR

[3410] Anonymous. The Garibaldi Bridge over the River Tiber at Rome. Scientific American, 68(10):150, March 11, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n10/pdf/scientificamerican03111893-150.pdf.

Anonymous:1893:Ci

[3411] Anonymous. Correspondence. Scientific American, 68(10):151, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n10/pdf/scientificamerican03111893-151.pdf.

Anonymous:1893:NMC

[3412] Anonymous. New mercurial compensation pendulum. Scientific American, 68(10):151, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
151a.pdf.

Anonymous:1893:ISd

[3413] Anonymous. Ice in the subway. Scientific American, 68(10):152, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n10/pdf/scientificamerican03111893-152.pdf.

Hovey:1893:RIR

[3414] H. C. Hovey. A remarkable instance of recent erosion. Scientific American, 68(10):152, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
152a.pdf.

Anonymous:1893:USS

[3415] Anonymous. United States steamship Iowa. Scientific American, 68(10):153, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
153.pdf.

Hopkins:1893:HCL

[3416] Geo. M. Hopkins. How to color lantern slides. Scientific American, 68(10):153–154, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
153a.pdf.

Anonymous:1893:CCNa

[3417] Anonymous. The Cunard Company’s new twin screw steamship Campania. Scientific American, 68(10):155, March 11, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n10/pdf/scientificamerican03111893-155.pdf.

Anonymous:1893:BNBj

[3418] Anonymous. Books: New books and publications. Scientific American, 68(10):156, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
156a.pdf.

Anonymous:1893:RPIj

[3419] Anonymous. Recently patented inventions. Scientific American, 68(10):156, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n10/pdf/scientificamerican03111893-
156.pdf.

Anonymous:1893:BPj

[3420] Anonymous. Business and personal. Scientific American, 68(10):157, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n10/pdf/scientificamerican03111893-157.pdf.

Anonymous:1893:NQi

[3421] Anonymous. Notes and queries. Scientific American, 68(10):157, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n10/pdf/scientificamerican03111893-157a.pdf.

Anonymous:1893:IIj

[3422] Anonymous. Index of inventions. Scientific American, 68(10):157–158, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n10/pdf/scientificamerican03111893-157b.pdf.

Anonymous:1893:PNPj

[3423] Anonymous. Patent notes: Patents. Scientific American, 68(10):160, March 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n10/pdf/scientificamerican03111893-160.pdf.

Anonymous:1893:NLB

[3424] Anonymous. The new law for brakes and car couplers. Scientific American, 68(11):162, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
162.pdf.

Anonymous:1893:TBPa

[3425] Anonymous. The telephone and the Berliner Patent. Scientific American, 68(11):162–163, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
162a.pdf.

Anonymous:1893:IPW

[3426] Anonymous. An improved phosphate washer. Scientific American, 68(11):164, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
164a.pdf.

Anonymous:1893:IWA

[3427] Anonymous. An inexpensive window awning. Scientific American, 68(11):164, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
164.pdf.

Anonymous:1893:BTR

[3428] Anonymous. Bridges on the Transandine Railway. Scientific American, 68(11):164–165, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
164b.pdf.

Anonymous:1893:AIL

[3429] Anonymous. American improved locomotive valve. Scientific American, 68(11):165, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
165.pdf.

Anonymous:1893:LSB

[3430] Anonymous. A locomotive Stone breaker. Scientific American, 68(11):165, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
165a.pdf.

Anonymous:1893:IWc

[3431] Anonymous. An improved wrench. Scientific American, 68(11):165–166, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-165b.pdf.

Anonymous:1893:HSV

[3432] Anonymous. How sunken vessels are raised. Scientific American, 68(11):166, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
166.pdf.

Anonymous:1893:IBH

[3433] Anonymous. Improvements in Boston Harbor. Scientific American, 68(11):167, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
167.pdf.

Anonymous:1893:GEEa

[3434] Anonymous. The Great Exposition entrances to the Fisheries and the Mining Buildings. Scientific American, 68(11):168, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-168.pdf.

Anonymous:1893:GSRa

[3435] Anonymous. Gold and silver refining from jewelers’ sweepings. Scientific American, 68(11):169–170, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
169.pdf.

Anonymous:1893:APb

[3436] Anonymous. Automatic photography. Scientific American, 68(11):170, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-170.pdf.

Anonymous:1893:BNBk

[3437] Anonymous. Books: New books and publications. Scientific American, 68(11):171, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
171a.pdf.

Anonymous:1893:RPIk

[3438] Anonymous. Recently patented inventions. Scientific American, 68(11):171, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n11/pdf/scientificamerican03181893-
171.pdf.

Anonymous:1893:BPk

[3439] Anonymous. Business and personal. Scientific American, 68(11):172, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-172.pdf.

Anonymous:1893:NQj

[3440] Anonymous. Notes and queries. Scientific American, 68(11):172–173, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-172a.pdf.

Anonymous:1893:Ig

[3441] Anonymous. To inventors. Scientific American, 68(11):173, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-173.pdf.

Anonymous:1893:IIk

[3442] Anonymous. Index of inventions. Scientific American, 68(11):173, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-173a.pdf.

Anonymous:1893:PNPk

[3443] Anonymous. Patent notes: Patents. Scientific American, 68(11):176, March 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n11/pdf/scientificamerican03181893-176.pdf.

Anonymous:1893:AIC

[3444] Anonymous. The armour institute, Chicago. Scientific American, 68(12):178, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
178.pdf.

Anonymous:1893:BTP

[3445] Anonymous. The Bell Telephone patents. Scientific American, 68(12):178, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
178a.pdf.

Anonymous:1893:WCA

[3446] Anonymous. The world’s Congress Auxiliary. Scientific American, 68(12):178, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
178b.pdf.

Anonymous:1893:VP

[3447] Anonymous. The value of a patent. Scientific American, 68 (12):178–179, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
178c.pdf.

Anonymous:1893:IMC

[3448] Anonymous. An improved mine car. Scientific American, 68(12):180, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-180a.pdf.

Bolder:1893:BI

[3449] Ben Bolder. The bicycle industry. Scientific American, 68(12):180, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-180.pdf.

Anonymous:1893:CSM

[3450] Anonymous. Crank shaping machine. Scientific American, 68 (12):180–181, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
180b.pdf.

Anonymous:1893:IBC

[3451] Anonymous. An improved book or copy holder. Scientific American, 68(12):181, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
181b.pdf.

Anonymous:1893:ICO

[3452] Anonymous. An improved crude oil burner. Scientific American, 68(12):181, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
181a.pdf.

Anonymous:1893:IMd

[3453] Anonymous. An improved motor. Scientific American, 68(12):181, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-181.pdf.

Anonymous:1893:SML

[3454] Anonymous. A hat support for miners’ lamps. Scientific American, 68(12):181, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
181c.pdf.

Anonymous:1893:NM

[3455] Anonymous. A novel motor. Scientific American, 68(12):182, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-182.pdf.

Anonymous:1893:US

[3456] Anonymous. The universal scale. Scientific American, 68(12):182, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-182a.pdf.

Anonymous:1893:Cj

[3457] Anonymous. Correspondence. Scientific American, 68(12):183, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-183a.pdf.

Anonymous:1893:NCD

[3458] Anonymous. The new coast defense war ship Terror. Scientific American, 68(12):183, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
183.pdf.

Anonymous:1893:RP

[3459] Anonymous. Recreations in photography. Scientific American, 68(12):184, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
184.pdf.

Anonymous:1893:ITEb

[3460] Anonymous. An improved traction engine. Scientific American, 68(12):185, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
185.pdf.

Anonymous:1893:PCM

[3461] Anonymous. Progress of the cruiser Maine. Scientific American, 68(12):185, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
185a.pdf.

Anonymous:1893:UOT

[3462] Anonymous. The utilization of old tin cans. Scientific American, 68(12):186, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
186.pdf.

Anonymous:1893:RPIl

[3463] Anonymous. Recently patented inventions. Scientific American, 68(12):187, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
187.pdf.

Anonymous:1893:BNBl

[3464] Anonymous. Books: New books and publications. Scientific American, 68(12):187–188, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n12/pdf/scientificamerican03251893-
187a.pdf.

Anonymous:1893:BPl

[3465] Anonymous. Business and personal. Scientific American, 68(12):188, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-188.pdf.

Anonymous:1893:NQk

[3466] Anonymous. Notes and queries. Scientific American, 68(12):188–189, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-188a.pdf.

Anonymous:1893:Ih

[3467] Anonymous. To inventors. Scientific American, 68(12):189, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-189.pdf.

Anonymous:1893:IIl

[3468] Anonymous. Index of inventions. Scientific American, 68(12):189–191, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-189a.pdf.

Anonymous:1893:PNPl

[3469] Anonymous. Patent notes: Patents. Scientific American, 68(12):192, March 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n12/pdf/scientificamerican03251893-192.pdf.

Anonymous:1893:MBR

[3470] Anonymous. Moving a bridge by rail. Scientific American, 68(13):193, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-193.pdf.

Anonymous:1893:LSRa

[3471] Anonymous. The legal status of railroad employes. Scientific American, 68(13):194, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
194.pdf.

Anonymous:1893:PPAa

[3472] Anonymous. Position of the planets in April. Scientific American, 68(13):195, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
195.pdf.

Anonymous:1893:ICA

[3473] Anonymous. An improved car axle box. Scientific American, 68 (13):196, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
196b.pdf.

Anonymous:1893:ICL

[3474] Anonymous. An improved Camera Lucida. Scientific American, 68(13):196, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
196.pdf.

Anonymous:1893:ILC

[3475] Anonymous. An independent lathe chuck. Scientific American, 68(13):196, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
196c.pdf.

Anonymous:1893:RIB

[3476] Anonymous. A recent improvement in buckles. Scientific American, 68(13):196, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
196a.pdf.

Anonymous:1893:IHa

[3477] Anonymous. An improved harrow. Scientific American, 68(13):197, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-197b.pdf.

Anonymous:1893:IRTa

[3478] Anonymous. Improved rubber toys. Scientific American, 68(13):197, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-197.pdf.

Anonymous:1893:OMI

[3479] Anonymous. The oldest musical instruments. Scientific American, 68(13):197, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
197a.pdf.

Anonymous:1893:THS

[3480] Anonymous. A Tellurian for the home and school. Scientific American, 68(13):197–198, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
197c.pdf.

Anonymous:1893:EHC

[3481] Anonymous. An electric heater for cars. Scientific American, 68 (13):198, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
198.pdf.

Anonymous:1893:Ta

[3482] Anonymous. The telautograph. Scientific American, 68(13):198, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-198a.pdf.

Anonymous:1893:Ck

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-199.pdf.

Anonymous:1893:FDE

[3484] Anonymous. Factory driving by electricity. Scientific American, 68(13):200–202, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
200.pdf.

Anonymous:1893:NSS

[3485] Anonymous. The new steel steamer El Rio. Scientific American, 68(13):202, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
202a.pdf.

Sloane:1893:DSF

[3486] T. O’conor Sloane. Dust and sand figures on membranes. Scientific American, 68(13):202, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
202.pdf.

Anonymous:1893:BNBm

[3487] Anonymous. Books: New books and publications. Scientific American, 68(13):203, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
203a.pdf.

Anonymous:1893:BPm

[3488] Anonymous. Business and personal. Scientific American, 68(13):203, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-203b.pdf.

Anonymous:1893:RPIm

[3489] Anonymous. Recently patented inventions. Scientific American, 68(13):203, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n13/pdf/scientificamerican04011893-
203.pdf.

Anonymous:1893:NQl

[3490] Anonymous. Notes and queries. Scientific American, 68(13):203–205, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-203c.pdf.

Anonymous:1893:Ij

[3491] Anonymous. To inventors. Scientific American, 68(13):205, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-205.pdf.

Anonymous:1893:IIm

[3492] Anonymous. Index of inventions. Scientific American, 68(13):205, April 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n13/pdf/scientificamerican04011893-205a.pdf.

Anonymous:1893:GIP

[3493] Anonymous. Glafcke’s improvement in prison construction. Scientific American, 68(14):209–210, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
209.pdf.

Anonymous:1893:AR

[3494] Anonymous. Agricultural railways. Scientific American, 68(14): 210–211, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
210.pdf.

Anonymous:1893:ERE

[3495] Anonymous. English railway exhibit for Chicago. Scientific American, 68(14):212, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
212a.pdf.

Hopkins:1893:SSC

[3496] Geo. M. Hopkins. Simple slide changer and eclipser. Scientific American, 68(14):212, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
212.pdf.

Anonymous:1893:AB

[3497] Anonymous. An attachment for beds. Scientific American, 68(14):213, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-213a.pdf.

Anonymous:1893:AKG

[3498] Anonymous. An automatic knife grinder. Scientific American, 68(14):213, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
213.pdf.

Anonymous:1893:ICCa

[3499] Anonymous. An improved car coupling. Scientific American, 68 (14):213, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
213c.pdf.

Anonymous:1893:ISCb

[3500] Anonymous. An improved sleeping car. Scientific American, 68(14):213, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-213b.pdf.

Anonymous:1893:GPW

[3501] Anonymous. Germ-proof water filter and cooler. Scientific American, 68(14):214, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
214.pdf.

Anonymous:1893:IMS

[3502] Anonymous. Improved method of stringing pianos. Scientific American, 68(14):214, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
214c.pdf.

Anonymous:1893:IRB

[3503] Anonymous. An improved railway block signal. Scientific American, 68(14):214, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
214b.pdf.

Anonymous:1893:NEB

[3504] Anonymous. The new Engineering Building of Pennsylvania State College. Scientific American, 68(14):214, April 8, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n14/pdf/scientificamerican04081893-214a.pdf.

Anonymous:1893:Cl

[3505] Anonymous. Correspondence. Scientific American, 68(14):215, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-215.pdf.

Anonymous:1893:BNC

[3506] Anonymous. Boilers for the new cruiser Cincinnati. Scientific American, 68(14):216–217, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
216.pdf.

Anonymous:1893:CBD

[3507] Anonymous. The cruiser Baltimore in dry dock. Scientific American, 68(14):216–217, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
216a.pdf.

Anonymous:1893:FSG

[3508] Anonymous. Fountain of Saint George and the Dragon, Vienna, by Anton Dominik Bitter Von Fernkorn. Scientific American, 68(14):217, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-217.pdf.

Gibbs:1893:BWN

[3509] Morris Gibbs. Birds which nest in cavities and Burrows. Scientific American, 68(14):218, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
218.pdf.

Anonymous:1893:RPIn

[3510] Anonymous. Recently patented inventions. Scientific American, 68(14):219, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
219.pdf.

Anonymous:1893:BNBn

[3511] Anonymous. Books: New books and publications. Scientific American, 68(14):219–220, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n14/pdf/scientificamerican04081893-
219a.pdf.

Anonymous:1893:BPn

[3512] Anonymous. Business and personal. Scientific American, 68(14):220, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-220.pdf.

Anonymous:1893:NQm

[3513] Anonymous. Notes and queries. Scientific American, 68(14):220–221, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-220a.pdf.

Anonymous:1893:Ik

[3514] Anonymous. To inventors. Scientific American, 68(14):221, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-221.pdf.

Anonymous:1893:IIn

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-221a.pdf.

Anonymous:1893:PNPm

[3516] Anonymous. Patent notes: Patents. Scientific American, 68(14):224, April 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n14/pdf/scientificamerican04081893-224.pdf.

Anonymous:1893:EGE

[3517] Anonymous. Electricity at the great exposition. Scientific American, 68(15):226, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
226.pdf.

Anonymous:1893:NRFa

[3518] Anonymous. The name of Robert Fulton for a war steamer. Scientific American, 68(15):226, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
226a.pdf.

Anonymous:1893:NUA

[3519] Anonymous. The new U. S. Army magazine rifle. Scientific American, 68(15):226, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
226b.pdf.

Anonymous:1893:NWCa

[3520] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(15):227–228, April 15, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n15/pdf/scientificamerican04151893-227.pdf.

Anonymous:1893:ITEc

[3521] Anonymous. Improved triple effect evaporator. Scientific American, 68(15):228, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
228a.pdf.

Anonymous:1893:KE

[3522] Anonymous. The Krupp Exhibits. Scientific American, 68(15):228, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-228.pdf.

Anonymous:1893:IBF

[3523] Anonymous. An improved boiler furnace. Scientific American, 68(15):229, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
229b.pdf.

Anonymous:1893:IDa

[3524] Anonymous. An improved dumbwaiter. Scientific American, 68 (15):229, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
229c.pdf.

Anonymous:1893:ISV

[3525] Anonymous. An improved suction valve. Scientific American, 68(15):229, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
229.pdf.

Anonymous:1893:PIW

[3526] Anonymous. Protecting iron work in buildings from fire. Scientific American, 68(15):229, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
229a.pdf.

Anonymous:1893:Cm

[3527] Anonymous. Correspondence. Scientific American, 68(15):230, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-230.pdf.

Anonymous:1893:ICCb

[3528] Anonymous. An improved car coupling. Scientific American, 68 (15):230–231, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
230a.pdf.

Anonymous:1893:RTC

[3529] Anonymous. The Racquet and Tennis Club of New York. Scientific American, 68(15):232, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
232.pdf.

Anonymous:1893:SMa

[3530] Anonymous. The steam man. Scientific American, 68(15):233, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-233.pdf.

Anonymous:1893:FWS

[3531] Anonymous. The first war steamer of the world. Scientific American, 68(15):233–234, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
233a.pdf.

Anonymous:1893:BNBo

[3532] Anonymous. Books: New books and publications. Scientific American, 68(15):235, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
235a.pdf.

Anonymous:1893:RPIo

[3533] Anonymous. Recently patented inventions. Scientific American, 68(15):235, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n15/pdf/scientificamerican04151893-
235.pdf.

Anonymous:1893:BPo

[3534] Anonymous. Business and personal. Scientific American, 68(15):236, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-236.pdf.

Anonymous:1893:NQn

[3535] Anonymous. Notes and queries. Scientific American, 68(15):236–237, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-236a.pdf.

Anonymous:1893:Il

[3536] Anonymous. To inventors. Scientific American, 68(15):237, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-237.pdf.

Anonymous:1893:IIo

[3537] Anonymous. Index of inventions. Scientific American, 68(15):237–238, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-237a.pdf.

Anonymous:1893:PNPn

[3538] Anonymous. Patent notes: Patents. Scientific American, 68(15):240, April 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n15/pdf/scientificamerican04151893-240.pdf.

Anonymous:1893:CNRb

[3539] Anonymous. The coming Naval Review at New York. Scientific American, 68(16):242, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
242b.pdf.

Anonymous:1893:PAD

[3540] Anonymous. Progress of American daily newspapers. Scientific American, 68(16):242, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
242.pdf.

Anonymous:1893:RCA

[3541] Anonymous. Rare chance for architects. Scientific American, 68 (16):242, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
242a.pdf.

Anonymous:1893:NWCb

[3542] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(16):243–244, April 22, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n16/pdf/scientificamerican04221893-243.pdf.

Anonymous:1893:CCC

[3543] Anonymous. Cristobal Colon De La Cerda, Duke of Veragua, Marquis of Jamaica. Scientific American, 68(16):244, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-244a.pdf.

Anonymous:1893:KEC

[3544] Anonymous. The Krupp Exhibits for the Columbian Exposition. Scientific American, 68(16):244, April 22, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n16/pdf/scientificamerican04221893-244.pdf.

Anonymous:1893:GAP

[3545] Anonymous. The Greenerd arbor press. Scientific American, 68 (16):245, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
245a.pdf.

Anonymous:1893:IFb

[3546] Anonymous. An improved furnace. Scientific American, 68(16):245, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-245b.pdf.

Anonymous:1893:ILI

[3547] Anonymous. An improved leveling instrument. Scientific American, 68(16):245, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
245.pdf.

Anonymous:1893:NBE

[3548] Anonymous. New bridges and elevated railroad connections in New York. Scientific American, 68(16):245, April 22, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n16/pdf/scientificamerican04221893-245c.pdf.

Anonymous:1893:NMS

[3549] Anonymous. The new Mail safety bicycle. Scientific American, 68(16):245, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
245d.pdf.

Anonymous:1893:Cn

[3550] Anonymous. Correspondence. Scientific American, 68(16):246, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-246.pdf.

Anonymous:1893:BNM

[3551] Anonymous. The Bates’ numbering machine. Scientific American, 68(16):247, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
247.pdf.

Anonymous:1893:SCH

[3552] Anonymous. A sausage casings holder. Scientific American, 68(16):247, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-247a.pdf.

Anonymous:1893:POG

[3553] Anonymous. The photo-opera glass. Scientific American, 68(16):248, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-248.pdf.

Anonymous:1893:CFE

[3554] Anonymous. The concrete foundations for the electric power house of the Brooklyn Railway Company. Scientific American, 68(16):248–249, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-248a.pdf.

Anonymous:1893:BCRa

[3555] Anonymous. The Broadway Cable Railway, New York. Scientific American, 68(16):249–250, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
249.pdf.

Anonymous:1893:BNBp

[3556] Anonymous. Books: New books and publications. Scientific American, 68(16):251, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
251a.pdf.

Anonymous:1893:RPIp

[3557] Anonymous. Recently patented inventions. Scientific American, 68(16):251, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n16/pdf/scientificamerican04221893-
251.pdf.

Anonymous:1893:BPp

[3558] Anonymous. Business and personal. Scientific American, 68(16):252, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-252.pdf.

Anonymous:1893:NQo

[3559] Anonymous. Notes and queries. Scientific American, 68(16):252–253, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-252a.pdf.

Anonymous:1893:In

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-253.pdf.

Anonymous:1893:IIp

[3561] Anonymous. Index of inventions. Scientific American, 68(16):253, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-253a.pdf.

Anonymous:1893:PNPo

[3562] Anonymous. Patent notes: Patents. Scientific American, 68(16):256, April 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n16/pdf/scientificamerican04221893-256.pdf.

Anonymous:1893:PMW

[3563] Anonymous. The painting machines at the World’s Fair. Scientific American, 68(17):257, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
257.pdf.

Anonymous:1893:WCEb

[3564] Anonymous. The World’s Columbian Exposition. Scientific American, 68(17):258–259, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
258.pdf.

Anonymous:1893:NWCc

[3565] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(17):259–260, April 29, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n17/pdf/scientificamerican04291893-259.pdf.

Anonymous:1893:IU

[3566] Anonymous. An improved unicycle. Scientific American, 68(17):260, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n17/pdf/scientificamerican04291893-260.pdf.

Anonymous:1893:PBI

[3567] Anonymous. The pearl button industry. Scientific American, 68 (17):260–261, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
260a.pdf.

Anonymous:1893:ASP

[3568] Anonymous. An automatic sprinkler to put out fires. Scientific American, 68(17):261, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
261b.pdf.

Anonymous:1893:IFW

[3569] Anonymous. An improved feed-water regulator. Scientific American, 68(17):261, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
261c.pdf.

Anonymous:1893:ISAb

[3570] Anonymous. An improved saddle attachment. Scientific American, 68(17):261, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
261.pdf.

Anonymous:1893:TDC

[3571] Anonymous. A tool to drill conical recesses. Scientific American, 68(17):261, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
261a.pdf.

Anonymous:1893:TSE

[3572] Anonymous. The total solar eclipse of April 16. Scientific American, 68(17):262, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
262.pdf.

Anonymous:1893:Co

[3573] Anonymous. Correspondence. Scientific American, 68(17):263, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n17/pdf/scientificamerican04291893-263a.pdf.

Anonymous:1893:PPMb

[3574] Anonymous. Position of the planets in May. Scientific American, 68(17):263, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
263.pdf.

Anonymous:1893:CCNb

[3575] Anonymous. The Columbus Celebration, New York. Scientific American, 68(17):264–265, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
264.pdf.

Sloane:1893:ESB

[3576] T. O’conor Sloane. Experiments with soap bubbles. Scientific American, 68(17):266, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
266.pdf.

Anonymous:1893:BNBq

[3577] Anonymous. Books: New books and publications. Scientific American, 68(17):267, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
267a.pdf.

Anonymous:1893:RPIq

[3578] Anonymous. Recently patented inventions. Scientific American, 68(17):267, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n17/pdf/scientificamerican04291893-
267.pdf.

Anonymous:1893:BPq

[3579] Anonymous. Business and personal. Scientific American, 68(17):268, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n17/pdf/scientificamerican04291893-268.pdf.

Anonymous:1893:NQp

[3580] Anonymous. Notes and queries. Scientific American, 68(17):268–269, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n17/pdf/scientificamerican04291893-268a.pdf.

Anonymous:1893:Io

[3581] Anonymous. To inventors. Scientific American, 68(17):269, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n17/pdf/scientificamerican04291893-269.pdf.

Anonymous:1893:IIq

[3582] Anonymous. Index of inventions. Scientific American, 68(17):269, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n17/pdf/scientificamerican04291893-269a.pdf.

Anonymous:1893:PNPp

[3583] Anonymous. Patent notes: Patents. Scientific American, 68(17):272, April 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n17/pdf/scientificamerican04291893-272.pdf.

Anonymous:1893:NPR

[3584] Anonymous. The Naval Parade and Review. Scientific American, 68(18):273, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
273.pdf.

Anonymous:1893:OCP

[3585] Anonymous. Operations of the Canadian Patent Office. Scientific American, 68(18):274, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
274.pdf.

Anonymous:1893:TPP

[3586] Anonymous. Treatment of potatoes to prevent potato scab. Scientific American, 68(18):274, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
274a.pdf.

Anonymous:1893:OWCa

[3587] Anonymous. Opening of the world’s Columbian Exposition, Chicago, May 1, 1893. Scientific American, 68(18):274–275, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n18/pdf/scientificamerican05061893-274b.pdf.

Anonymous:1893:ERC

[3588] Anonymous. An electrical rocking chair. Scientific American, 68 (18):276, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
276.pdf.

Anonymous:1893:TDR

[3589] Anonymous. A tie-dog for railroad rails. Scientific American, 68(18):276, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
276b.pdf.

Anonymous:1893:TP

[3590] Anonymous. Tricks in prestidigitation. Scientific American, 68(18):276, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-276a.pdf.

Anonymous:1893:HPV

[3591] Anonymous. Hydraulic propulsion of vessels. Scientific American, 68(18):277, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
277b.pdf.

Anonymous:1893:ICBa

[3592] Anonymous. An improved corrugated bit brace ring. Scientific American, 68(18):277, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
277a.pdf.

Anonymous:1893:ISTb

[3593] Anonymous. An improved scaffold truss. Scientific American, 68(18):277, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
277c.pdf.

Anonymous:1893:IWCa

[3594] Anonymous. An improved wrench and cutter. Scientific American, 68(18):277, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
277.pdf.

Anonymous:1893:IEB

[3595] Anonymous. An improved electric belt. Scientific American, 68 (18):277–278, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
277d.pdf.

Anonymous:1893:Pa

[3596] Anonymous. Patentability. Scientific American, 68(18):278, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-278.pdf.

Anonymous:1893:Cp

[3597] Anonymous. Correspondence. Scientific American, 68(18):279–281, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-279.pdf.

Anonymous:1893:MLb

[3598] Anonymous. A model locomotive. Scientific American, 68(18):281, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-281.pdf.

Anonymous:1893:CEP

[3599] Anonymous. A Columbian egg puzzle. Scientific American, 68 (18):281–283, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
281a.pdf.

Anonymous:1893:DS

[3600] Anonymous. The deadly scorpion. Scientific American, 68(18):283, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-283a.pdf.

Tanner:1893:AEF

[3601] A. M. Tanner. Another early French patent for a barbed wire fence. Scientific American, 68(18):283, May 6, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n18/pdf/scientificamerican05061893-283.pdf.

Anonymous:1893:BPr

[3602] Anonymous. Business and personal. Scientific American, 68(18):284, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-284a.pdf.

Anonymous:1893:RPIr

[3603] Anonymous. Recently patented inventions. Scientific American, 68(18):284, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n18/pdf/scientificamerican05061893-
284.pdf.

Anonymous:1893:NQq

[3604] Anonymous. Notes and queries. Scientific American, 68(18):284–285, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-284b.pdf.

Anonymous:1893:Ip

[3605] Anonymous. To inventors. Scientific American, 68(18):285, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-285.pdf.

Anonymous:1893:IIr

[3606] Anonymous. Index of inventions. Scientific American, 68(18):285, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-285a.pdf.

Anonymous:1893:PNPq

[3607] Anonymous. Patent notes: Patents. Scientific American, 68(18):287, May 6, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n18/pdf/scientificamerican05061893-287.pdf.

Anonymous:1893:OWCb

[3608] Anonymous. Opening of the World’s Columbian Exposition at Chicago. Scientific American, 68(19):290, May 13, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n19/pdf/scientificamerican05131893-290.pdf.

Anonymous:1893:WCEc

[3609] Anonymous. The World’s Columbian Exposition. Scientific American, 68(19):290, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
290a.pdf.

Anonymous:1893:ESC

[3610] Anonymous. The electric systems at the Columbian Exposition. Scientific American, 68(19):290–291, May 13, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n19/pdf/scientificamerican05131893-290b.pdf.

Anonymous:1893:NWCd

[3611] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(19):291, May 13, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n19/pdf/scientificamerican05131893-291a.pdf.

Anonymous:1893:POC

[3612] Anonymous. The post office of the Columbian Exposition. Scientific American, 68(19):291, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
291.pdf.

Anonymous:1893:EF

[3613] Anonymous. Electric furnaces. Scientific American, 68(19):292, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-292.pdf.

Anonymous:1893:IRR

[3614] Anonymous. An improved railway roadbed. Scientific American, 68(19):292, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
292a.pdf.

Anonymous:1893:CTGa

[3615] Anonymous. Curious tree growths. Scientific American, 68(19):293, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-293.pdf.

Anonymous:1893:SEa

[3616] Anonymous. “Strip, the electrician”. Scientific American, 68(19):293, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-293a.pdf.

Anonymous:1893:L

[3617] Anonymous. Locomotives of 1831 and 1893. Scientific American, 68(19):294–295, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
294.pdf.

Anonymous:1893:BQD

[3618] Anonymous. The Broughton quick delivery mixer. Scientific American, 68(19):295, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
295.pdf.

Anonymous:1893:SPW

[3619] Anonymous. A steam pressure and water level recorder. Scientific American, 68(19):295, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
295a.pdf.

Anonymous:1893:CFI

[3620] Anonymous. The canned food industry. Scientific American, 68 (19):296, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
296a.pdf.

Anonymous:1893:EIT

[3621] Anonymous. Edison’s improvement in the transmission of power. Scientific American, 68(19):296, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
296.pdf.

Anonymous:1893:CESa

[3622] Anonymous. The Columbian Exposition steamer Christopher Columbus. Scientific American, 68(19):297–298, May 13, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n19/pdf/scientificamerican05131893-297.pdf.

Anonymous:1893:ETCa

[3623] Anonymous. An efficient traveling crane. Scientific American, 68 (19):298, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
298.pdf.

Anonymous:1893:BNBr

[3624] Anonymous. Books: New books and publications. Scientific American, 68(19):299, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
299a.pdf.

Anonymous:1893:BPs

[3625] Anonymous. Business and personal. Scientific American, 68(19):299, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-299b.pdf.

Anonymous:1893:RPIs

[3626] Anonymous. Recently patented inventions. Scientific American, 68(19):299, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n19/pdf/scientificamerican05131893-
299.pdf.

Anonymous:1893:NQr

[3627] Anonymous. Notes and queries. Scientific American, 68(19):299–301, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-299c.pdf.

Anonymous:1893:Iq

[3628] Anonymous. To inventors. Scientific American, 68(19):301, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-301.pdf.

Anonymous:1893:IIs

[3629] Anonymous. Index of inventions. Scientific American, 68(19):301, May 13, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-301a.pdf.

Anonymous:1893:PNPr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n19/pdf/scientificamerican05131893-304.pdf.

Anonymous:1893:BCRb

[3631] Anonymous. The Broadway Cable Railway, New York. Scientific American, 68(20):305, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n20/pdf/scientificamerican05201893-
305.pdf.

Anonymous:1893:OHT

[3632] Anonymous. One hundred and twelve and one half miles per hour. Scientific American, 68(20):306, May 20, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n20/pdf/scientificamerican05201893-306a.pdf.

Anonymous:1893:PWCa

[3633] Anonymous. Prices at the World’s Columbian Exposition. Scientific American, 68(20):306, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n20/pdf/scientificamerican05201893-
306.pdf.

Anonymous:1893:ERTa

[3634] Anonymous. The electric railway trolley. Scientific American, 68 (20):306–307, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n20/pdf/scientificamerican05201893-
306b.pdf.

Anonymous:1893:CRM

[3635] Anonymous. The crossing of rivers by means of leathern bottles. Scientific American, 68(20):308, May 20, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n20/pdf/scientificamerican05201893-308a.pdf.

Anonymous:1893:WCEd

[3636] Anonymous. The World’s Columbian Exposition — the Pope’s phonographic message to America. Scientific American, 68(20):308, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-308.pdf.

Anonymous:1893:MDT

[3637] Anonymous. A measuring and drawing tool. Scientific American, 68(20):309, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n20/pdf/scientificamerican05201893-
309.pdf.

Anonymous:1893:VSb

[3638] Anonymous. A Viking ship. Scientific American, 68(20):309, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-309a.pdf.

Hopkins:1893:SH

[3639] M. Hopkins. Simple hygroscope. Scientific American, 68(20):309–310, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-309b.pdf.

Anonymous:1893:ICT

[3640] Anonymous. An improved calking tool. Scientific American, 68(20):310, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-310.pdf.

Anonymous:1893:Cq

[3641] Anonymous. Correspondence. Scientific American, 68(20):311, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-311.pdf.

Anonymous:1893:Tb

[3642] Anonymous. The trumpeters. Scientific American, 68(20):314–315, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-314.pdf.

Anonymous:1893:RPIt

[3643] Anonymous. Recently patented inventions. Scientific American, 68(20):315–316, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n20/pdf/scientificamerican05201893-
315.pdf.

Anonymous:1893:BPt

[3644] Anonymous. Business and personal. Scientific American, 68(20):316, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-316.pdf.

Anonymous:1893:NQs

[3645] Anonymous. Notes and queries. Scientific American, 68(20):316–317, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-316a.pdf.

Anonymous:1893:BNBs

[3646] Anonymous. Books: New books and publications. Scientific American, 68(20):317, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n20/pdf/scientificamerican05201893-
317.pdf.

Anonymous:1893:Ir

[3647] Anonymous. To inventors. Scientific American, 68(20):317, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-317a.pdf.

Anonymous:1893:IIt

[3648] Anonymous. Index of inventions. Scientific American, 68(20):317, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-317b.pdf.

Anonymous:1893:PNPs

[3649] Anonymous. Patent notes: Patents. Scientific American, 68(20):320, May 20, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n20/pdf/scientificamerican05201893-320.pdf.

Anonymous:1893:WEL

[3650] Anonymous. Winby’s express locomotive at the Columbian Exposition. Scientific American, 68(21):321, May 27, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n21/pdf/scientificamerican05271893-321.pdf.

Anonymous:1893:EEP

[3651] Anonymous. Electrical engineering as a profession. Scientific American, 68(21):322, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
322a.pdf.

Anonymous:1893:TAW

[3652] Anonymous. Tests and awards at the World’s Fair. Scientific American, 68(21):322, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
322.pdf.

Anonymous:1893:NWCe

[3653] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(21):323, May 27, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n21/pdf/scientificamerican05271893-323.pdf.

Anonymous:1893:ACF

[3654] Anonymous. Artistic and comfortable furniture. Scientific American, 68(21):324, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
324a.pdf.

Anonymous:1893:RWM

[3655] Anonymous. A rotary water meter. Scientific American, 68(21):324, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-324.pdf.

Anonymous:1893:IHS

[3656] Anonymous. An improved hay stacker. Scientific American, 68 (21):324–325, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
324b.pdf.

Anonymous:1893:IDb

[3657] Anonymous. The instantaneous divider. Scientific American, 68 (21):325, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
325a.pdf.

Anonymous:1893:MWL

[3658] Anonymous. The magic wax lighter. Scientific American, 68(21):325, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-325b.pdf.

Anonymous:1893:RCL

[3659] Anonymous. A railway car life guard. Scientific American, 68(21):325, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-325.pdf.

Anonymous:1893:SCE

[3660] Anonymous. A simple and compact engine and dynamo. Scientific American, 68(21):325–326, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
325c.pdf.

Anonymous:1893:ERB

[3661] Anonymous. The electric railway between Chicago and St. Louis. Scientific American, 68(21):328–329, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
328.pdf.

Anonymous:1893:MMb

[3662] Anonymous. Manufacture of macaroni. Scientific American, 68(21):329, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-329.pdf.

Anonymous:1893:BSG

[3663] Anonymous. Between saddle and ground — an instantaneous photograph. Scientific American, 68(21):330, May 27, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n21/pdf/scientificamerican05271893-330b.pdf.

Anonymous:1893:GHS

[3664] Anonymous. A gang hydraulic steam press. Scientific American, 68(21):330, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
330.pdf.

Anonymous:1893:MMc

[3665] Anonymous. The monitor Miantonomoh. Scientific American, 68 (21):330, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
330a.pdf.

Anonymous:1893:BNBt

[3666] Anonymous. Books: New books and publications. Scientific American, 68(21):331, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
331a.pdf.

Anonymous:1893:RPIu

[3667] Anonymous. Recently patented inventions. Scientific American, 68(21):331, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n21/pdf/scientificamerican05271893-
331.pdf.

Anonymous:1893:BPu

[3668] Anonymous. Business and personal. Scientific American, 68(21):332, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-332.pdf.

Anonymous:1893:NQt

[3669] Anonymous. Notes and queries. Scientific American, 68(21):332–333, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-332a.pdf.

Anonymous:1893:It

[3670] Anonymous. To inventors. Scientific American, 68(21):333, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-333.pdf.

Anonymous:1893:IIu

[3671] Anonymous. Index of inventions. Scientific American, 68(21):333, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-333a.pdf.

Anonymous:1893:PNPt

[3672] Anonymous. Patent notes: Patents. Scientific American, 68(21):336, May 27, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n21/pdf/scientificamerican05271893-336.pdf.

Anonymous:1893:DMG

[3673] Anonymous. Death of Moses G. Farmer. Scientific American, 68(22):338, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
338b.pdf.

Anonymous:1893:MSW

[3674] Anonymous. The maintenance of the speed of war ships. Scientific American, 68(22):338, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
338a.pdf.

Anonymous:1893:STW

[3675] Anonymous. Satisfactory trial of the war ship New York. Scientific American, 68(22):338, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
338.pdf.

Anonymous:1893:NWCf

[3676] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(22):339, June 3, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n22/pdf/scientificamerican06031893-339.pdf.

Anonymous:1893:PEE

[3677] Anonymous. Petroleum engine exhibits at the World’s Columbian Exposition. Scientific American, 68(22):340, June 3, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n22/pdf/scientificamerican06031893-340.pdf.

Anonymous:1893:N

[3678] Anonymous. No. 999. Scientific American, 68(22):341, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n22/pdf/scientificamerican06031893-341.pdf.

Anonymous:1893:HWE

[3679] Anonymous. Hoisting and winding engines. Scientific American, 68(22):341–342, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
341a.pdf.

Anonymous:1893:IPP

[3680] Anonymous. An improved potato planter. Scientific American, 68(22):342, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
342a.pdf.

Anonymous:1893:RCD

[3681] Anonymous. A roof clothes-drying device. Scientific American, 68(22):342, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
342.pdf.

Anonymous:1893:Cr

[3682] Anonymous. Correspondence. Scientific American, 68(22):343, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n22/pdf/scientificamerican06031893-343a.pdf.

Anonymous:1893:PPJa

[3683] Anonymous. Position of the planets in June. Scientific American, 68(22):343, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
343.pdf.

Hopkins:1893:FDT

[3684] M. Hopkins. Focusing device for telescopes. Scientific American, 68(22):344, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
344.pdf.

Anonymous:1893:BTS

[3685] Anonymous. Break of thrust shaft of steamship Hecla. Scientific American, 68(22):345, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
345a.pdf.

Anonymous:1893:KWS

[3686] Anonymous. Kirby’s window sash improvements. Scientific American, 68(22):345, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
345.pdf.

Anonymous:1893:NBP

[3687] Anonymous. The new Baltimore Post Office. Scientific American, 68(22):346, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
346a.pdf.

Anonymous:1893:RDE

[3688] Anonymous. Recent discovery of early texts of the Gospels. Scientific American, 68(22):346, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
346.pdf.

Anonymous:1893:BNBu

[3689] Anonymous. Books: New books and publications. Scientific American, 68(22):347, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
347a.pdf.

Anonymous:1893:RPIv

[3690] Anonymous. Recently patented inventions. Scientific American, 68(22):347, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n22/pdf/scientificamerican06031893-
347.pdf.

Anonymous:1893:BPv

[3691] Anonymous. Business and personal. Scientific American, 68(22):348, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n22/pdf/scientificamerican06031893-348.pdf.

Anonymous:1893:NQu

[3692] Anonymous. Notes and queries. Scientific American, 68(22):348–349, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n22/pdf/scientificamerican06031893-348a.pdf.

Anonymous:1893:Iv

[3693] Anonymous. To inventors. Scientific American, 68(22):349, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n22/pdf/scientificamerican06031893-349.pdf.

Anonymous:1893:IIv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n22/pdf/scientificamerican06031893-349a.pdf.

Anonymous:1893:PNPu

[3695] Anonymous. Patent notes: Patents. Scientific American, 68(22):352, June 3, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n22/pdf/scientificamerican06031893-352.pdf.

Anonymous:1893:TCB

[3696] Anonymous. Traveling cranes at the Brooklyn Navy Yard. Scientific American, 68(23):353–354, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v68/n23/pdf/scientificamerican06101893-
353.pdf.

Anonymous:1893:CNB

[3697] Anonymous. The commercial national bank of Chicago. Scientific American, 68(23):354, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n23/pdf/scientificamerican06101893-
354.pdf.

Anonymous:1893:PWCb

[3698] Anonymous. Photography at the World’s Columbian Exposition. Scientific American, 68(23):354, June 10, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n23/pdf/scientificamerican06101893-354a.pdf.

Anonymous:1893:NWCg

[3699] Anonymous. Notes from the World’s Columbian Exposition. Scientific American, 68(23):355, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n23/pdf/scientificamerican06101893-
355.pdf.

Anonymous:1893:IER

[3700] Anonymous. Improved electric railway construction. Scientific American, 68(23):356, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n23/pdf/scientificamerican06101893-
356.pdf.

Anonymous:1893:MWD

[3701] Anonymous. A machine for washing, drying and separating coal, ore, etc. Scientific American, 68(23):356, June 10, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n23/pdf/scientificamerican06101893-356a.pdf.

Anonymous:1893:TE

[3702] Anonymous. The twin elms. Scientific American, 68(23):356, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-356b.pdf.

Anonymous:1893:IMe

[3703] Anonymous. An improved mangle. Scientific American, 68(23):357, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-357a.pdf.

Anonymous:1893:IWd

[3704] Anonymous. An improved wrench. Scientific American, 68(23):357, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-357.pdf.

Anonymous:1893:SFS

[3705] Anonymous. A simple form of safety pocket. Scientific American, 68(23):357–358, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n23/pdf/scientificamerican06101893-
357b.pdf.

Anonymous:1893:FSF

[3706] Anonymous. Falguiere’s statue “La France” at the Chicago Exposition. Scientific American, 68(23):358, June 10, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n23/pdf/scientificamerican06101893-358.pdf.

Anonymous:1893:Cs

[3707] Anonymous. Correspondence. Scientific American, 68(23):359–360, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-359.pdf.

Anonymous:1893:BCI

[3708] Anonymous. Blake’s carriage indicator. Scientific American, 68(23):361, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-361.pdf.

Anonymous:1893:PCP

[3709] Anonymous. Pneumatic cotton picker. Scientific American, 68(23):361, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-361a.pdf.

Anonymous:1893:TSH

[3710] Anonymous. The 100 ton steam hammer and the large rolling mill of the Etaings Works. Scientific American, 68(23):361–362, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-361b.pdf.

Anonymous:1893:BNBv

[3711] Anonymous. Books: New books and publications. Scientific American, 68(23):363, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n23/pdf/scientificamerican06101893-
363a.pdf.

Anonymous:1893:RPIw

[3712] Anonymous. Recently patented inventions. Scientific American, 68(23):363, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n23/pdf/scientificamerican06101893-
363.pdf.

Anonymous:1893:BPw

[3713] Anonymous. Business and personal. Scientific American, 68(23):364, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-364.pdf.

Anonymous:1893:NQv

[3714] Anonymous. Notes and queries. Scientific American, 68(23):364, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-364a.pdf.

Anonymous:1893:Ix

[3715] Anonymous. To inventors. Scientific American, 68(23):365, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-365.pdf.

Anonymous:1893:IIw

[3716] Anonymous. Index of inventions. Scientific American, 68(23):365–366, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-365a.pdf.

Anonymous:1893:PNPv

[3717] Anonymous. Patent notes: Patents. Scientific American, 68(23):368, June 10, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n23/pdf/scientificamerican06101893-368.pdf.

Anonymous:1893:FCI

[3718] Anonymous. Farm changes in Illinois. Scientific American, 68(24):370, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-370a.pdf.

Anonymous:1893:IYW

[3719] Anonymous. If you wish to be happy, be reasonable. Scientific American, 68(24):370, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
370.pdf.

Anonymous:1893:ITa

[3720] Anonymous. Improved transportation. Scientific American, 68 (24):370–371, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
370b.pdf.

Anonymous:1893:NWCh

[3721] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(24):371, June 17, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n24/pdf/scientificamerican06171893-371a.pdf.

Anonymous:1893:RCE

[3722] Anonymous. Rolling chairs at the Exposition. Scientific American, 68(24):371, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
371.pdf.

Anonymous:1893:EGW

[3723] Anonymous. English goldsmiths’ work for the Chicago Exhibition. Scientific American, 68(24):372, June 17, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n24/pdf/scientificamerican06171893-372.pdf.

Anonymous:1893:VCJ

[3724] Anonymous. Vienna and China in juxtaposition. Scientific American, 68(24):372, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
372a.pdf.

Anonymous:1893:ELL

[3725] Anonymous. An electric light lantern. Scientific American, 68(24):373, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-373b.pdf.

Anonymous:1893:ERTb

[3726] Anonymous. An electrically regulated time system. Scientific American, 68(24):373, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
373c.pdf.

Anonymous:1893:IMRb

[3727] Anonymous. An improved metallic railway tie. Scientific American, 68(24):373, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
373.pdf.

Anonymous:1893:SOD

[3728] Anonymous. A switch operating device. Scientific American, 68 (24):373, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
373a.pdf.

Anonymous:1893:ILT

[3729] Anonymous. An improved ladder and truck. Scientific American, 68(24):375, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
375b.pdf.

Anonymous:1893:TBWa

[3730] Anonymous. Torsion braided wire pillows. Scientific American, 68(24):375, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
375a.pdf.

Anonymous:1893:WSN

[3731] Anonymous. The war ship New York. Scientific American, 68(24):375, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-375.pdf.

Anonymous:1893:NRFb

[3732] Anonymous. The Navy rapid fire gun. Scientific American, 68 (24):376–378, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
376.pdf.

Anonymous:1893:GHB

[3733] Anonymous. Germany’s highest bridge. Scientific American, 68 (24):379, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
379.pdf.

Anonymous:1893:BNBw

[3734] Anonymous. Books: New books and publications. Scientific American, 68(24):380, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
380a.pdf.

Anonymous:1893:BPx

[3735] Anonymous. Business and personal. Scientific American, 68(24):380, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-380b.pdf.

Anonymous:1893:RPIx

[3736] Anonymous. Recently patented inventions. Scientific American, 68(24):380, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n24/pdf/scientificamerican06171893-
380.pdf.

Anonymous:1893:NQw

[3737] Anonymous. Notes and queries. Scientific American, 68(24):380–381, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-380c.pdf.

Anonymous:1893:Iy

[3738] Anonymous. To inventors. Scientific American, 68(24):381, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-381.pdf.

Anonymous:1893:IIx

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-381a.pdf.

Anonymous:1893:PNPw

[3740] Anonymous. Patent notes: Patents. Scientific American, 68(24):384, June 17, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n24/pdf/scientificamerican06171893-384.pdf.

Anonymous:1893:LBS

[3741] Anonymous. Launch of the battle ship Massachusetts. Scientific American, 68(25):386, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
386a.pdf.

Anonymous:1893:ONS

[3742] Anonymous. The old and the new ships. Scientific American, 68(25):386, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
386.pdf.

Anonymous:1893:NWCi

[3743] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 68(25):387–388, June 24, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n25/pdf/scientificamerican06241893-387.pdf.

Anonymous:1893:CBS

[3744] Anonymous. A convenient bicycle stand. Scientific American, 68 (25):388, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
388c.pdf.

Anonymous:1893:ELF

[3745] Anonymous. Exposition at Lyons, France, in 1894. Scientific American, 68(25):388, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
388a.pdf.

Anonymous:1893:IPB

[3746] Anonymous. An improved pinch bar. Scientific American, 68(25):388, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-388b.pdf.

Anonymous:1893:VSD

[3747] Anonymous. A vessel stopping device. Scientific American, 68(25):388, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-388.pdf.

Anonymous:1893:IHW

[3748] Anonymous. An improved hydraulic wheel press. Scientific American, 68(25):389, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
389a.pdf.

Anonymous:1893:IRG

[3749] Anonymous. An improved road gate. Scientific American, 68(25):389, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-389c.pdf.

Anonymous:1893:NDA

[3750] Anonymous. Neff’s desk attachment. Scientific American, 68(25):389, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-389b.pdf.

Anonymous:1893:PPAb

[3751] Anonymous. Petch’s plow attachment. Scientific American, 68(25):389, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-389.pdf.

Anonymous:1893:CTGb

[3752] Anonymous. A curious tree growth. Scientific American, 68(25):390, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-390b.pdf.

Anonymous:1893:NK

[3753] Anonymous. A Nubian knife. Scientific American, 68(25):390, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-390.pdf.

Lewis:1893:IEW

[3754] T. H. Lewis. Intaglio effigies of Wisconsin. Scientific American, 68(25):390, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
390a.pdf.

Anonymous:1893:NYH

[3755] Anonymous. New York Harbor improvement models at the Columbian Exposition. Scientific American, 68(25):391–392, June 24, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v68/
n25/pdf/scientificamerican06241893-391.pdf.

Anonymous:1893:JWC

[3756] Anonymous. Japanese at the World’s Columbian Exposition. Scientific American, 68(25):392, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
392.pdf.

Anonymous:1893:RDL

[3757] Anonymous. Refraction and dispersion of light. Scientific American, 68(25):392, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
392a.pdf.

Anonymous:1893:VSC

[3758] Anonymous. The Viking Ship for the Columbian Exposition. Scientific American, 68(25):393, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
393.pdf.

Anonymous:1893:BPy

[3759] Anonymous. Business and personal. Scientific American, 68(25):394, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-394a.pdf.

Anonymous:1893:RPIy

[3760] Anonymous. Recently patented inventions. Scientific American, 68(25):394, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v68/n25/pdf/scientificamerican06241893-
394.pdf.

Anonymous:1893:NQx

[3761] Anonymous. Notes and queries. Scientific American, 68(25):394–395, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-394b.pdf.

Anonymous:1893:IIy

[3762] Anonymous. Index of inventions. Scientific American, 68(25):395, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-395.pdf.

Anonymous:1893:PNPx

[3763] Anonymous. Patent notes: Patents. Scientific American, 68(25):400, June 24, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v68/n25/pdf/scientificamerican06241893-400.pdf.

Anonymous:1893:GF

[3764] Anonymous. The great fair. Scientific American, 69(1):2, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-2a.pdf.

Anonymous:1893:PSW

[3765] Anonymous. Proposed submarine war boats. Scientific American, 69(1):2, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
2b.pdf.

Anonymous:1893:SBW

[3766] Anonymous. Sinking of the British war ship Victoria. Scientific American, 69(1):2, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
2.pdf.

Anonymous:1893:NWCj

[3767] Anonymous. Notes from the World’s Columbian Exposition. Scientific American, 69(1):3, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
3.pdf.

Anonymous:1893:FRM

[3768] Anonymous. Foot race of market porters at Paris. Scientific American, 69(1):4, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
4b.pdf.

Anonymous:1893:ISL

[3769] Anonymous. Improved self-loading vehicles and dump wagons. Scientific American, 69(1):4, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
4a.pdf.

Anonymous:1893:IWW

[3770] Anonymous. An improved water wheel. Scientific American, 69(1):4, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-4.pdf.

Anonymous:1893:ILa

[3771] Anonymous. An improved lubricator. Scientific American, 69(1):5, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-5b.pdf.

Anonymous:1893:IWL

[3772] Anonymous. An inexpensive window larder. Scientific American, 69(1): 5, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-5.pdf.

Anonymous:1893:SMG

[3773] Anonymous. A stamp mill guide. Scientific American, 69(1):5, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-5c.pdf.

Anonymous:1893:TMA

[3774] Anonymous. A thrashing machine attachment. Scientific American, 69(1):5, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
5a.pdf.

Anonymous:1893:PPJb

[3775] Anonymous. Position of the planets in July. Scientific American, 69(1):6–7, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
6.pdf.

Anonymous:1893:Ct

[3776] Anonymous. Correspondence. Scientific American, 69(1):7–8, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-7.pdf.

Anonymous:1893:GWC

[3777] Anonymous. The great wheel at Chicago. Scientific American, 69(1):8– 11, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-8.pdf.

Anonymous:1893:BNBx

[3778] Anonymous. Books: New books and publications. Scientific American, 69(1):12, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n1/pdf/scientificamerican07011893-
12a.pdf.

Anonymous:1893:RPIz

[3779] Anonymous. Recently patented inventions. Scientific American, 69(1): 12, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-12.pdf.

Anonymous:1893:BPz

[3780] Anonymous. Business and personal. Scientific American, 69(1):13, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-13.pdf.

Anonymous:1893:NQy

[3781] Anonymous. Notes and queries. Scientific American, 69(1):13, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-13a.pdf.

Anonymous:1893:IIz

[3782] Anonymous. Index of inventions. Scientific American, 69(1):13–14, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-13b.pdf.

Anonymous:1893:PNPy

[3783] Anonymous. Patent notes: Patents. Scientific American, 69(1):16, July 1, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n1/pdf/scientificamerican07011893-16.pdf.

Anonymous:1893:OFB

[3784] Anonymous. Oil fuel and boilers at the great exposition. Scientific American, 69(2):17, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
17.pdf.

Anonymous:1893:CMP

[3785] Anonymous. The centralization of motive power. Scientific American, 69(2):18, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
18.pdf.

Anonymous:1893:NWCk

[3786] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(2):19, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
19.pdf.

Anonymous:1893:IMDa

[3787] Anonymous. An instrument for measuring distances. Scientific American, 69(2):20, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
20b.pdf.

Anonymous:1893:MBF

[3788] Anonymous. A music book of folio. Scientific American, 69(2):20, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-20.pdf.

Anonymous:1893:MT

[3789] Anonymous. The mysterious trunk. Scientific American, 69(2):20, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-20a.pdf.

Anonymous:1893:IRCb

[3790] Anonymous. An improved roller cotton gin. Scientific American, 69(2):20–21, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
20c.pdf.

Anonymous:1893:ST

[3791] Anonymous. Surface tension. Scientific American, 69(2):21, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-21.pdf.

Anonymous:1893:Cu

[3792] Anonymous. Correspondence. Scientific American, 69(2):22, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-22.pdf.

Anonymous:1893:PHI

[3793] Anonymous. Parsons’ Horological Institute. Scientific American, 69(2):23–24, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
23.pdf.

Anonymous:1893:OEG

[3794] Anonymous. Oil engines at the great exposition. Scientific American, 69(2):24–25, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
24.pdf.

Anonymous:1893:WHV

[3795] Anonymous. The wart hog, or vlacke vark. Scientific American, 69(2): 27, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-27.pdf.

Anonymous:1893:BNBy

[3796] Anonymous. Books: New books and publications. Scientific American, 69(2):28, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n2/pdf/scientificamerican07081893-
28a.pdf.

Anonymous:1893:BPba

[3797] Anonymous. Business and personal. Scientific American, 69(2):28, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-28b.pdf.

Anonymous:1893:RPIba

[3798] Anonymous. Recently patented inventions. Scientific American, 69(2): 28, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-28.pdf.

Anonymous:1893:NQz

[3799] Anonymous. Notes and queries. Scientific American, 69(2):28–29, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-28c.pdf.

Anonymous:1893:Iz

[3800] Anonymous. To inventors. Scientific American, 69(2):29, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-29.pdf.

Anonymous:1893:IIba

[3801] Anonymous. Index of inventions. Scientific American, 69(2):29–30, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-29a.pdf.

Anonymous:1893:PNPz

[3802] Anonymous. Patent notes: Patents. Scientific American, 69(2):32, July 8, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n2/pdf/scientificamerican07081893-32.pdf.

Anonymous:1893:KEG

[3803] Anonymous. The Krupp Exhibit at the Great Fair. Scientific American, 69(3):33, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
33.pdf.

Anonymous:1893:ABB

[3804] Anonymous. Artificial butter and butter swindles. Scientific American, 69(3):34, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
34.pdf.

Anonymous:1893:NWCl

[3805] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(3):35, July 15, 1893. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
35.pdf.

Edis:1893:VHC

[3806] Robert W. Edis. The Victoria House, Chicago. Scientific American, 69(3):36–37, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
36.pdf.

Anonymous:1893:FCC

[3807] Anonymous. A flue and chimney cleaner. Scientific American, 69(3):37, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-37b.pdf.

Anonymous:1893:ICCc

[3808] Anonymous. An improved car coupling. Scientific American, 69(3):37, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-37.pdf.

Anonymous:1893:SSC

[3809] Anonymous. The silver statue of Columbus. Scientific American, 69(3):37, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
37a.pdf.

Anonymous:1893:Cv

[3810] Anonymous. Correspondence. Scientific American, 69(3):39, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-39.pdf.

Anonymous:1893:KEM

[3811] Anonymous. The Kansas Exhibit of mounted specimens of the animals of the state. Scientific American, 69(3):41–43, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-41.pdf.

Anonymous:1893:FCL

[3812] Anonymous. The fin cutter Leni Lenepe. Scientific American, 69(3):43, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-43.pdf.

Anonymous:1893:PAEb

[3813] Anonymous. The Peary Arctic Expedition. Scientific American, 69(3):43, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
43a.pdf.

Anonymous:1893:BNBz

[3814] Anonymous. Books: New books and publications. Scientific American, 69(3):44, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
44a.pdf.

Anonymous:1893:RPIbb

[3815] Anonymous. Recently patented inventions. Scientific American, 69(3):44, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n3/pdf/scientificamerican07151893-
44.pdf.

Anonymous:1893:BPbb

[3816] Anonymous. Business and personal. Scientific American, 69(3):45, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-45.pdf.

Anonymous:1893:NQba

[3817] Anonymous. Notes and queries. Scientific American, 69(3):45, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-45a.pdf.

Anonymous:1893:IIbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-45b.pdf.

Anonymous:1893:PNPba

[3819] Anonymous. Patent notes: Patents. Scientific American, 69(3):48, July 15, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n3/pdf/scientificamerican07151893-48.pdf.

Anonymous:1893:AAP

[3820] Anonymous. American armor plate. Scientific American, 69(4):50, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-50a.pdf.

Anonymous:1893:DPP

[3821] Anonymous. The delay in the publication of the Patent Office Gazette and of copies of patents. Scientific American, 69(4):50, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-50.pdf.

Anonymous:1893:PDS

[3822] Anonymous. Protective devices on street railroads. Scientific American, 69(4):50, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n4/pdf/scientificamerican07221893-
50b.pdf.

Anonymous:1893:NWCm

[3823] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(4):51, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n4/pdf/scientificamerican07221893-
51.pdf.

Anonymous:1893:CSP

[3824] Anonymous. The Cold storage palace at the Columbian Fair and its destruction by fire. Scientific American, 69(4):52–53, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-52.pdf.

Anonymous:1893:IWM

[3825] Anonymous. An improved wire mat. Scientific American, 69(4):53, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-53a.pdf.

Anonymous:1893:TBPb

[3826] Anonymous. Test of bearing power of piles. Scientific American, 69(4):53, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n4/pdf/scientificamerican07221893-
53.pdf.

Anonymous:1893:ICS

[3827] Anonymous. An improved Carpenter’s square. Scientific American, 69(4):54–55, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n4/pdf/scientificamerican07221893-
54.pdf.

Anonymous:1893:AON

[3828] Anonymous. The Arizona, Oklahoma and New Mexico Building. Scientific American, 69(4):56, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n4/pdf/scientificamerican07221893-
56c.pdf.

Anonymous:1893:MBa

[3829] Anonymous. The Montana building. Scientific American, 69(4):56, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-56b.pdf.

Anonymous:1893:NYS

[3830] Anonymous. New York state building. Scientific American, 69(4):56, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-56.pdf.

Anonymous:1893:OB

[3831] Anonymous. The Ohio building. Scientific American, 69(4):56, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-56a.pdf.

Anonymous:1893:TEA

[3832] Anonymous. The total eclipse of the 16th of April in Senegal. Scientific American, 69(4):57–58, July 22, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n4/pdf/scientificamerican07221893-57.pdf.

Anonymous:1893:AM

[3833] Anonymous. Antediluvian monsters. Scientific American, 69(4):59–60, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-59.pdf.

Anonymous:1893:BPbc

[3834] Anonymous. Business and personal. Scientific American, 69(4):61, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-61.pdf.

Anonymous:1893:Ibd

[3835] Anonymous. To inventors. Scientific American, 69(4):61, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-61b.pdf.

Anonymous:1893:NQbb

[3836] Anonymous. Notes and queries. Scientific American, 69(4):61, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-61a.pdf.

Anonymous:1893:IIbc

[3837] Anonymous. Index of inventions. Scientific American, 69(4):61–62, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-61c.pdf.

Anonymous:1893:PNPbb

[3838] Anonymous. Patent notes: Patents. Scientific American, 69(4):64, July 22, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n4/pdf/scientificamerican07221893-64.pdf.

Anonymous:1893:CAF

[3839] Anonymous. Colonel Auchmuty, founder of the New York trade schools. Scientific American, 69(5):66, July 29, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n5/pdf/scientificamerican07291893-66b.pdf.

Anonymous:1893:NTc

[3840] Anonymous. Naval torpedoes. Scientific American, 69(5):66, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-66a.pdf.

Anonymous:1893:WFI

[3841] Anonymous. The World’s Fair International Engineering Congress. Scientific American, 69(5):66, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n5/pdf/scientificamerican07291893-
66.pdf.

Anonymous:1893:LC

[3842] Anonymous. Libraries in Chicago. Scientific American, 69(5):66–67, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-66c.pdf.

Anonymous:1893:NWCn

[3843] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(5):67, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n5/pdf/scientificamerican07291893-
67.pdf.

Anonymous:1893:GES

[3844] Anonymous. The great exhibit of the Self-Winding Clock Company at the World’s Columbian Exposition. Scientific American, 69(5):68–69, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-68.pdf.

Anonymous:1893:DTC

[3845] Anonymous. Duplex tandem compound engines. Scientific American, 69(5):69, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n5/pdf/scientificamerican07291893-
69.pdf.

Anonymous:1893:IFP

[3846] Anonymous. An improved fence post. Scientific American, 69(5):72, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-72.pdf.

Anonymous:1893:NPT

[3847] Anonymous. A new printing telegraph. Scientific American, 69(5):72, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-72a.pdf.

Anonymous:1893:NWG

[3848] Anonymous. A novel work guide for saws. Scientific American, 69(5):72, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-72b.pdf.

Anonymous:1893:CTM

[3849] Anonymous. Chain towing by magnetic adhesion. Scientific American, 69(5):73–74, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n5/pdf/scientificamerican07291893-
73.pdf.

Anonymous:1893:FA

[3850] Anonymous. Fonthill abbey. Scientific American, 69(5):75, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-75.pdf.

Anonymous:1893:SMb

[3851] Anonymous. The stearic motor. Scientific American, 69(5):75, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-75a.pdf.

Anonymous:1893:BNBba

[3852] Anonymous. Books: New books and publications. Scientific American, 69(5):76, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n5/pdf/scientificamerican07291893-
76a.pdf.

Anonymous:1893:RPIbc

[3853] Anonymous. Recently patented inventions. Scientific American, 69(5):76, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n5/pdf/scientificamerican07291893-
76.pdf.

Anonymous:1893:BPbd

[3854] Anonymous. Business and personal. Scientific American, 69(5):77, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-77.pdf.

Anonymous:1893:Ibe

[3855] Anonymous. To inventors. Scientific American, 69(5):77, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-77b.pdf.

Anonymous:1893:NQbc

[3856] Anonymous. Notes and queries. Scientific American, 69(5):77, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-77a.pdf.

Anonymous:1893:IIbd

[3857] Anonymous. Index of inventions. Scientific American, 69(5):77–78, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-77c.pdf.

Anonymous:1893:PNPbc

[3858] Anonymous. Patent notes: Patents. Scientific American, 69(5):80, July 29, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n5/pdf/scientificamerican07291893-80.pdf.

Anonymous:1893:EELa

[3859] Anonymous. The Edison, electric light suits. Scientific American, 69(6):82, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
82b.pdf.

Anonymous:1893:EWP

[3860] Anonymous. The expedition of work in the Patent Office. Scientific American, 69(6):82, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
82.pdf.

Anonymous:1893:FVG

[3861] Anonymous. The fall in the value of general electric stock. Scientific American, 69(6):82, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
82c.pdf.

Anonymous:1893:RPP

[3862] Anonymous. Resumption of publication of the Patent Office Gazette. Scientific American, 69(6):82, August 5, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n6/pdf/scientificamerican08051893-82a.pdf.

Anonymous:1893:NWCo

[3863] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(6):83, August 5, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n6/pdf/scientificamerican08051893-83.pdf.

Anonymous:1893:DCCb

[3864] Anonymous. The Desroziers continuous current dynamo. Scientific American, 69(6):84, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
84a.pdf.

Anonymous:1893:IPC

[3865] Anonymous. An improved Printer’s composing stick. Scientific American, 69(6):84, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
84.pdf.

Anonymous:1893:SCB

[3866] Anonymous. A strong and cheap bedstead. Scientific American, 69(6):84, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
84b.pdf.

Anonymous:1893:HTC

[3867] Anonymous. A handsome team of cattle. Scientific American, 69(6):85, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-85a.pdf.

Anonymous:1893:RGE

[3868] Anonymous. A reversing gear for engines. Scientific American, 69(6):85, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-85.pdf.

Anonymous:1893:PPAc

[3869] Anonymous. Position of the planets in August. Scientific American, 69(6):86–87, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
86.pdf.

Anonymous:1893:FEL

[3870] Anonymous. The French Exhibit of locomotives. — the Bethlehem Iron Company’s exhibit at the World’s Fair. Scientific American, 69(6):87, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-87.pdf.

Anonymous:1893:HSE

[3871] Anonymous. A high speed electric locomotive. Scientific American, 69(6):88, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
88.pdf.

Anonymous:1893:IMRc

[3872] Anonymous. The intra-mural road at the great exposition. Scientific American, 69(6):88–89, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
88a.pdf.

Anonymous:1893:IIbe

[3873] Anonymous. Irrigation in India. Scientific American, 69(6):89–90, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-89.pdf.

Anonymous:1893:PDW

[3874] Anonymous. The purification of drinking water by means of electricity. Scientific American, 69(6):91, August 5, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n6/pdf/scientificamerican08051893-91.pdf.

Anonymous:1893:BNBbb

[3875] Anonymous. Books: New books and publications. Scientific American, 69(6):92, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
92a.pdf.

Anonymous:1893:RPIbd (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n6/pdf/scientificamerican08051893-
92.pdf.

Anonymous:1893:BPbe

[3877] Anonymous. Business and personal. Scientific American, 69(6):93, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-93.pdf.

Anonymous:1893:Ibg

[3878] Anonymous. To inventors. Scientific American, 69(6):93, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-93b.pdf.

Anonymous:1893:NQbd

[3879] Anonymous. Notes and queries. Scientific American, 69(6):93, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-93a.pdf.

Anonymous:1893:IIbf

[3880] Anonymous. Index of inventions. Scientific American, 69(6):93–94, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-93c.pdf.

Anonymous:1893:PNPbd

[3881] Anonymous. Patent notes: Patents. Scientific American, 69(6):96, August 5, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n6/pdf/scientificamerican08051893-96.pdf.

Anonymous:1893:WCEe

[3882] Anonymous. The World’s Columbian Exposition — the Babcock & Wilcox Boiler Plant. Scientific American, 69(7):97, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-97.pdf.

Anonymous:1893:NCL

[3883] Anonymous. The new Congressional library building. Scientific American, 69(7):98, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
98a.pdf.

Anonymous:1893:NIS

[3884] Anonymous. New inventions for street cleaning wanted in New York City. Scientific American, 69(7):98, August 12, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n7/pdf/scientificamerican08121893-98.pdf.

Anonymous:1893:NWCp

[3885] Anonymous. Notes from the worlds Columbian Exposition Chicago 1893. Scientific American, 69(7):99, August 12, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n7/pdf/scientificamerican08121893-99.pdf.

Anonymous:1893:DLC

[3886] Anonymous. Duplex lathe for cranks. Scientific American, 69 (7):100, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
100a.pdf.

Anonymous:1893:NBG

[3887] Anonymous. A new bicycle gear. Scientific American, 69(7):100, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-100.pdf.

Anonymous:1893:NSV

[3888] Anonymous. A novel support for velocipede saddles. Scientific American, 69(7):100, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
100b.pdf.

Anonymous:1893:CPB

[3889] Anonymous. A combined pitchfork and band cutter. Scientific American, 69(7):101, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
101b.pdf.

Anonymous:1893:DSR

[3890] Anonymous. The “Duplex” slide rule. Scientific American, 69 (7):101, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
101c.pdf.

Anonymous:1893:GLE

[3891] Anonymous. A gas lighter or extinguisher. Scientific American, 69(7):101, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
101d.pdf.

Anonymous:1893:PFL

[3892] Anonymous. A photographic flash light apparatus. Scientific American, 69(7):101, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
101a.pdf.

Anonymous:1893:RTE

[3893] Anonymous. Recording time of employees. Scientific American, 69(7):101, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
101.pdf.

Anonymous:1893:CCW

[3894] Anonymous. Coal conveyer at Weehawken, N.J. Scientific American, 69(7):104, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
104.pdf.

Anonymous:1893:BS

[3895] Anonymous. Brooding snakes. Scientific American, 69(7):105, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-105.pdf.

Anonymous:1893:Cw

[3896] Anonymous. Correspondence. Scientific American, 69(7):106, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-106.pdf.

Anonymous:1893:RSW

[3897] Anonymous. The race of stilt walkers at Bordeaux. Scientific American, 69(7):107, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
107.pdf.

Anonymous:1893:BNBbc

[3898] Anonymous. Books: New books and publications. Scientific American, 69(7):108, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
108a.pdf.

Anonymous:1893:RPIbe

[3899] Anonymous. Recently patented inventions. Scientific American, 69(7):108, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
108.pdf.

Anonymous:1893:BPbf

[3900] Anonymous. Business and personal. Scientific American, 69(7):109, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-109.pdf.

Anonymous:1893:Ibi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-109b.pdf.

Anonymous:1893:NQbe

[3902] Anonymous. Notes and queries. Scientific American, 69(7):109, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-109a.pdf.

Anonymous:1893:IIbg

[3903] Anonymous. Index of inventions. Scientific American, 69(7): 109–110, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n7/pdf/scientificamerican08121893-
109c.pdf.

Anonymous:1893:PNPbe

[3904] Anonymous. Patent notes: Patents. Scientific American, 69(7):112, August 12, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n7/pdf/scientificamerican08121893-112.pdf.

Anonymous:1893:AN

[3905] Anonymous. Aerial navigation. Scientific American, 69(8):114, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-114b.pdf.

Anonymous:1893:RVM

[3906] Anonymous. Raising vessels by means of air sacks. Scientific American, 69(8):114, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
114.pdf.

Anonymous:1893:WPT

[3907] Anonymous. Water power transformed into light. Scientific American, 69(8):114, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
114a.pdf.

Anonymous:1893:NWCq

[3908] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(8):115, August 19, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n8/pdf/scientificamerican08191893-115.pdf.

Anonymous:1893:ED

[3909] Anonymous. An experiment in divination. Scientific American, 69(8):116, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
116b.pdf.

Anonymous:1893:FSR

[3910] Anonymous. The Fox safety razor and strop. Scientific American, 69(8):116, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
116a.pdf.

Anonymous:1893:RAB

[3911] Anonymous. A readily applied burglar alarm. Scientific American, 69(8):116, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
116.pdf.

Anonymous:1893:RCG

[3912] Anonymous. A railway cattle guard. Scientific American, 69(8):116, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-116c.pdf.

Anonymous:1893:CBV

[3913] Anonymous. A convenient bench vise. Scientific American, 69 (8):117, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
117.pdf.

Anonymous:1893:TFP

[3914] Anonymous. Thawing out frozen pipes, etc. Scientific American, 69(8):117, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
117a.pdf.

Anonymous:1893:IILa

[3915] Anonymous. The Imperial Institute, London. Scientific American, 69(8):117–118, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
117b.pdf.

Anonymous:1893:Cx

[3916] Anonymous. Correspondence. Scientific American, 69(8):118, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-118.pdf.

Anonymous:1893:WCEf

[3917] Anonymous. The World’s Columbian Exposition — a view in Columbian Avenue, Palace of Manufactures and Liberal Arts. Scientific American, 69(8):119–120, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
119.pdf.

Anonymous:1893:ME

[3918] Anonymous. Mill engines. Scientific American, 69(8):120, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-120.pdf.

Anonymous:1893:HJB

[3919] Anonymous. Hoo-Den, or Japanese building. Scientific American, 69(8):121, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
121.pdf.

Anonymous:1893:SSI

[3920] Anonymous. South Sea Island village. Scientific American, 69 (8):121–123, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
121a.pdf.

Anonymous:1893:VVS

[3921] Anonymous. Visiting the Viking Ship. Scientific American, 69 (8):123, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
123.pdf.

Anonymous:1893:BNBbd

[3922] Anonymous. Books: New books and publications. Scientific American, 69(8):124, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
124a.pdf.

Anonymous:1893:RPIbf

[3923] Anonymous. Recently patented inventions. Scientific American, 69(8):124, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
124.pdf.

Anonymous:1893:BPbg

[3924] Anonymous. Business and personal. Scientific American, 69(8):125, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-125.pdf.

Anonymous:1893:Ibj

[3925] Anonymous. To inventors. Scientific American, 69(8):125, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-125b.pdf.

Anonymous:1893:NQbf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-125a.pdf.

Anonymous:1893:IIbh

[3927] Anonymous. Index of inventions. Scientific American, 69(8): 125–126, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n8/pdf/scientificamerican08191893-
125c.pdf.

Anonymous:1893:PNPbf

[3928] Anonymous. Patent notes: Patents. Scientific American, 69(8):128, August 19, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n8/pdf/scientificamerican08191893-128.pdf.

Anonymous:1893:AI

[3929] Anonymous. Armour institute. Scientific American, 69(9):130, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-130a.pdf.

Anonymous:1893:CPa

[3930] Anonymous. Color photography. Scientific American, 69(9):130, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-130b.pdf.

Anonymous:1893:PRU

[3931] Anonymous. Progress of railways in the United States. Scientific American, 69(9):130, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
130.pdf.

Anonymous:1893:NWCr

[3932] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(9):131, August 26, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n9/pdf/scientificamerican08261893-131.pdf.

Anonymous:1893:DTS

[3933] Anonymous. A device to teach spelling. Scientific American, 69(9):132, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
132c.pdf.

Anonymous:1893:GSRb

[3934] Anonymous. A gold and silver refining establishment. Scientific American, 69(9):132, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
132b.pdf.

Anonymous:1893:KTD

[3935] Anonymous. A kindergarten teaching device. Scientific American, 69(9):132, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
132.pdf.

Anonymous:1893:MFS

[3936] Anonymous. A machine for finishing saws. Scientific American, 69(9):132, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
132a.pdf.

Anonymous:1893:IBS

[3937] Anonymous. An improved balanced slide valve. Scientific American, 69(9):132–133, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
132d.pdf.

Anonymous:1893:ICEb

[3938] Anonymous. Improved circular and elliptical boring machine. Scientific American, 69(9):133–134, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
133.pdf.

Hopkins:1893:ESM

[3939] M. Hopkins. Experiment showing magnetic lag. Scientific American, 69(9):134, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
134.pdf.

Anonymous:1893:NTBa

[3940] Anonymous. Nasmyth’s torpedo boat of 1853. Scientific American, 69(9):135, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
135.pdf.

Anonymous:1893:WRB

[3941] Anonymous. Warburton Road Bridge — Manchester Ship Canal. Scientific American, 69(9):136, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
136.pdf.

Anonymous:1893:CVF

[3942] Anonymous. The Century Vase at the Fair. Scientific American, 69(9):137, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
137.pdf.

Anonymous:1893:SMF

[3943] Anonymous. The Santa Maria at the Fair. Scientific American, 69(9):137, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
137a.pdf.

Anonymous:1893:WCEg

[3944] Anonymous. World’s Columbian Exposition — entrance to the Palace of Agriculture Building. Scientific American, 69(9):138, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-138.pdf.

Anonymous:1893:Cz

[3945] Anonymous. Correspondence. Scientific American, 69(9):138–139, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-138a.pdf.

Anonymous:1893:TD

[3946] Anonymous. A twin daisy. Scientific American, 69(9):139–140, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-139.pdf.

Anonymous:1893:RPIbg

[3947] Anonymous. Recently patented inventions. Scientific American, 69(9):140, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
140.pdf.

Anonymous:1893:BNBbe

[3948] Anonymous. Books: New books and publications. Scientific American, 69(9):141, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
141.pdf.

Anonymous:1893:BPbh

[3949] Anonymous. Business and personal. Scientific American, 69(9):141, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-141a.pdf.

Anonymous:1893:NQbg

[3950] Anonymous. Notes and queries. Scientific American, 69(9):141, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-141b.pdf.

Anonymous:1893:IIbi

[3951] Anonymous. Index of inventions. Scientific American, 69(9): 141–142, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n9/pdf/scientificamerican08261893-
141c.pdf.

Anonymous:1893:PNPbg

[3952] Anonymous. Patent notes: Patents. Scientific American, 69(9):144, August 26, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n9/pdf/scientificamerican08261893-144.pdf.

Anonymous:1893:POE

[3953] Anonymous. Patent office examiners wanted. Scientific American, 69(10):146, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
146.pdf.

Hovey:1893:AAM

[3954] H. C. Hovey. The American Association at Madison, Wis. Scientific American, 69(10):146–147, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
146a.pdf.

Anonymous:1893:BGC

[3955] Anonymous. Bevel gearing cut theoretically correct. Scientific American, 69(10):148, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
148a.pdf.

Anonymous:1893:LGV

[3956] Anonymous. The “Lunken” gate valve. Scientific American, 69 (10):148, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
148.pdf.

Anonymous:1893:LRA

[3957] Anonymous. A light reflecting apparatus. Scientific American, 69 (10):148, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
148b.pdf.

Anonymous:1893:CMR

[3958] Anonymous. Condensed milk represented at the Fair. Scientific American, 69(10):149, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
149.pdf.

Anonymous:1893:EDC

[3959] Anonymous. Exhibit of the Dixon Crucible Company at the Fair. Scientific American, 69(10):149–150, September 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n10/pdf/scientificamerican09021893-149a.pdf.

Anonymous:1893:CRA

[3960] Anonymous. Cable railway accidents. Scientific American, 69 (10):150, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
150.pdf.

Hopking:1893:SO

[3961] M. Hopking. Suggestions in optics. Scientific American, 69 (10):151, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
151.pdf.

Anonymous:1893:GGS

[3962] Anonymous. The great German search lights at the World’s Columbian Exposition. Scientific American, 69(10):152, September 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n10/pdf/scientificamerican09021893-152.pdf.

Anonymous:1893:WCEh

[3963] Anonymous. The World’s Columbian Exposition — state buildings. Scientific American, 69(10):153–154, September 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n10/pdf/scientificamerican09021893-153.pdf.

Anonymous:1893:NWCs

[3964] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(10):154, September 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n10/pdf/scientificamerican09021893-154.pdf.

Anonymous:1893:NUD

[3965] Anonymous. The new U.S.S. Detroit. Scientific American, 69 (10):155, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
155.pdf.

Anonymous:1893:BPbi

[3966] Anonymous. Business and personal. Scientific American, 69 (10):156, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
156a.pdf.

Anonymous:1893:RPIbh

[3967] Anonymous. Recently patented inventions. Scientific American, 69(10):156, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
156.pdf.

Anonymous:1893:NQbh

[3968] Anonymous. Notes and queries. Scientific American, 69(10): 156–157, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
156b.pdf.

Anonymous:1893:BNBbf

[3969] Anonymous. Books: New books and publications. Scientific American, 69(10):157, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
157.pdf.

Anonymous:1893:Ibk

[3970] Anonymous. To inventors. Scientific American, 69(10):157, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n10/pdf/scientificamerican09021893-157a.pdf.

Anonymous:1893:IIbj

[3971] Anonymous. Index of inventions. Scientific American, 69(10): 157–158, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
157b.pdf.

Anonymous:1893:PNPbh

[3972] Anonymous. Patent notes: Patents. Scientific American, 69 (10):160, September 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n10/pdf/scientificamerican09021893-
160.pdf.

Anonymous:1893:WCEi

[3973] Anonymous. The World’s Columbian Exposition — triumphal arch and peristyle. Scientific American, 69(11):161, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n11/pdf/scientificamerican09091893-161.pdf.

Anonymous:1893:LAT

[3974] Anonymous. The latest armor trial at Indian head. Scientific American, 69(11):162, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
162.pdf.

Anonymous:1893:PWW

[3975] Anonymous. The purification of water in wells and cisterns. Scientific American, 69(11):162–163, September 9, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n11/pdf/scientificamerican09091893-162a.pdf.

Anonymous:1893:NWCt

[3976] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(11):163, September 9, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n11/pdf/scientificamerican09091893-163.pdf.

Anonymous:1893:BSE

[3977] Anonymous. A balanced steam engine valve. Scientific American, 69(11):164, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
164a.pdf.

Anonymous:1893:CSS

[3978] Anonymous. A convenient saw setting device. Scientific American, 69(11):164, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
164b.pdf.

Anonymous:1893:ESL

[3979] Anonymous. Evolution of the safety lamp as shown at the Fair. Scientific American, 69(11):164, September 9, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n11/pdf/scientificamerican09091893-164.pdf.

Anonymous:1893:CDA

[3980] Anonymous. Creamery and dairy apparatus at the Fair. Scientific American, 69(11):165, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
165a.pdf.

Anonymous:1893:GRM

[3981] Anonymous. The Griffin Roller Mill at the Fair. Scientific American, 69(11):165, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
165.pdf.

Anonymous:1893:MAA

[3982] Anonymous. Meeting of the American Association at Madison. Scientific American, 69(11):166–167, September 9, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n11/pdf/scientificamerican09091893-166.pdf.

Anonymous:1893:QCG

[3983] Anonymous. The Quaker city grinding hill at the Exposition. Scientific American, 69(11):167, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
167.pdf.

Anonymous:1893:WCEj

[3984] Anonymous. The World’s Columbian Exposition — the North Canal and Bridge. Scientific American, 69(11):168, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n11/pdf/scientificamerican09091893-168.pdf.

Anonymous:1893:WCEk

[3985] Anonymous. The World’s Columbian Exposition — a view from the Ferris Wheel. Scientific American, 69(11):169–170, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n11/pdf/scientificamerican09091893-169.pdf.

Anonymous:1893:PPS

[3986] Anonymous. Position of the planets in September. Scientific American, 69(11):170, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
170.pdf.

Anonymous:1893:WCEl

[3987] Anonymous. The World’s Columbian Exposition — interior of Woman’s building. Scientific American, 69(11):171, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n11/pdf/scientificamerican09091893-171.pdf.

Anonymous:1893:BNBbg

[3988] Anonymous. Books: New books and publications. Scientific American, 69(11):172, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
172a.pdf.

Anonymous:1893:BPbj

[3989] Anonymous. Business and personal. Scientific American, 69 (11):172, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
172b.pdf.

Anonymous:1893:RPIbi

[3990] Anonymous. Recently patented inventions. Scientific American, 69(11):172, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
172.pdf.

Anonymous:1893:NQbi

[3991] Anonymous. Notes and queries. Scientific American, 69(11): 172–173, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
172c.pdf.

Anonymous:1893:Ibl

[3992] Anonymous. To inventors. Scientific American, 69(11):173, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n11/pdf/scientificamerican09091893-173.pdf.

Anonymous:1893:IIbk

[3993] Anonymous. Index of inventions. Scientific American, 69(11): 173–174, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
173a.pdf.

Anonymous:1893:BNBbh

[3994] Anonymous. Books: New books. Scientific American, 69(11): 175, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
175.pdf.

Anonymous:1893:PNPbi

[3995] Anonymous. Patent notes: Patents. Scientific American, 69 (11):176, September 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n11/pdf/scientificamerican09091893-
176.pdf.

Anonymous:1893:MGP

[3996] Anonymous. The McConnell germ-proof filters. Scientific American, 69(12):177, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
177.pdf.

Anonymous:1893:WMR

[3997] Anonymous. Wine making represented at the Fair. Scientific American, 69(12):177–178, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
177a.pdf.

Anonymous:1893:CAC

[3998] Anonymous. A chance for American contractors. Scientific American, 69(12):178, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
178.pdf.

Anonymous:1893:SR

[3999] Anonymous. Safety on railroads. Scientific American, 69(12): 178–179, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
178a.pdf.

Anonymous:1893:NWCu

[4000] Anonymous. Notes from the World’s Columbian Exposition. Scientific American, 69(12):179–180, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
179.pdf.

Anonymous:1893:CAP

[4001] Anonymous. A counting attachment for printing presses. Scientific American, 69(12):180, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
180.pdf.

Anonymous:1893:DFF

[4002] Anonymous. Drop forgings at the Fair. Scientific American, 69 (12):180, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
180a.pdf.

Anonymous:1893:IBV

[4003] Anonymous. An improved brake valve. Scientific American, 69 (12):180, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
180c.pdf.

Anonymous:1893:WCW

[4004] Anonymous. A wonderfully cheap watch. Scientific American, 69 (12):180, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
180b.pdf.

Anonymous:1893:CPI

[4005] Anonymous. The Caw’s pen and ink company exhibit. Scientific American, 69(12):181, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
181b.pdf.

Anonymous:1893:EBP

[4006] Anonymous. An electrical binding post. Scientific American, 69 (12):181, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
181.pdf.

Anonymous:1893:ISM

[4007] Anonymous. An improved stall for milk cattle. Scientific American, 69(12):181, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
181c.pdf.

Anonymous:1893:SPE

[4008] Anonymous. A steel pen exhibit at the Fair. Scientific American, 69(12):181, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
181a.pdf.

Anonymous:1893:WRC

[4009] Anonymous. Washington’s red Cedar vase. Scientific American, 69(12):182, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
182.pdf.

Anonymous:1893:Cba

[4010] Anonymous. Correspondence. Scientific American, 69(12):183, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n12/pdf/scientificamerican09161893-183.pdf.

Anonymous:1893:MDb

[4011] Anonymous. A monster diamond. Scientific American, 69(12): 183, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
183b.pdf.

Anonymous:1893:RC

[4012] Anonymous. The Rordame Comet. Scientific American, 69(12): 183, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
183a.pdf.

Anonymous:1893:NDL

[4013] Anonymous. The Northern Diver at the Leipzig Zoological Garden. Scientific American, 69(12):184, September 16, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n12/pdf/scientificamerican09161893-184.pdf.

Anonymous:1893:TWF

[4014] Anonymous. A trolley wire finder for electric railways. Scientific American, 69(12):184, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
184a.pdf.

Anonymous:1893:WCEm

[4015] Anonymous. The World’s Columbian Exposition — a view on the Grand South Canal. Scientific American, 69(12):185–186, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n12/pdf/scientificamerican09161893-185.pdf.

Anonymous:1893:HWS

[4016] Anonymous. Heavy wagons shown at the Fair. Scientific American, 69(12):187–188, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
187.pdf.

Anonymous:1893:WGS

[4017] Anonymous. A. W. Grays Sons Exhibit of “Horse Powers”. Scientific American, 69(12):187–188, September 16, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n12/pdf/scientificamerican09161893-187a.pdf.

Anonymous:1893:BPbk

[4018] Anonymous. Business and personal. Scientific American, 69 (12):189, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
189a.pdf.

Anonymous:1893:RPIbj

[4019] Anonymous. Recently patented inventions. Scientific American, 69 (12):189, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
189.pdf.

Anonymous:1893:NQbj

[4020] Anonymous. Notes and queries. Scientific American, 69(12): 189–190, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
189b.pdf.

Anonymous:1893:Ibm

[4021] Anonymous. To inventors. Scientific American, 69(12):190, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n12/pdf/scientificamerican09161893-190.pdf.

Anonymous:1893:IIbl

[4022] Anonymous. Index of inventions. Scientific American, 69(12): 190, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
190a.pdf.

Anonymous:1893:PNPbj

[4023] Anonymous. Patent notes: Patents. Scientific American, 69 (12):192, September 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n12/pdf/scientificamerican09161893-
192.pdf.

Anonymous:1893:MTD

[4024] Anonymous. The Morse twist drill at the Fair. Scientific American, 69(13):193, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
193.pdf.

Anonymous:1893:NPO

[4025] Anonymous. A new Patent Office building proposed. Scientific American, 69(13):194, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
194.pdf.

Anonymous:1893:NWCv

[4026] Anonymous. Notes from the World’s Columbian Exposition. Scientific American, 69(13):195, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
195.pdf.

Anonymous:1893:ETCb

[4027] Anonymous. The Eureka Tempered Copper Co.’s Exhibit. Scientific American, 69(13):196, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
196a.pdf.

Anonymous:1893:GBT

[4028] Anonymous. A guide block for telegraph wires, etc. Scientific American, 69(13):196, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
196b.pdf.

Anonymous:1893:RHA

[4029] Anonymous. The Rider Hot Air Engine Company Exhibit. Scientific American, 69(13):196, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
196.pdf.

Anonymous:1893:ICBb

[4030] Anonymous. An inside cover for barrels. Scientific American, 69 (13):197, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
197b.pdf.

Anonymous:1893:IEMa

[4031] Anonymous. An inexpensive and efficient milk cooler. Scientific American, 69(13):197, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
197.pdf.

Anonymous:1893:LBF

[4032] Anonymous. Leather belts at the Fair. Scientific American, 69 (13):197, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
197a.pdf.

Anonymous:1893:MPC

[4033] Anonymous. Mechanical press for copper plate printing. Scientific American, 69(13):197–198, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
197c.pdf.

Anonymous:1893:CSa

[4034] Anonymous. A convenience for smokers. Scientific American, 69 (13):198, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
198.pdf.

Anonymous:1893:WCEn

[4035] Anonymous. The World’s Columbian Exposition-general view. Scientific American, 69(13):199, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
199.pdf.

Anonymous:1893:JMC

[4036] Anonymous. Jean Martin Charcot. Scientific American, 69(13): 200, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
200.pdf.

Anonymous:1893:WGE

[4037] Anonymous. William Galloway’s exhibit of art pottery. Scientific American, 69(13):200, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
200a.pdf.

Benjamin:1893:WH

[4038] Marcus Benjamin. William Harkness. Scientific American, 69 (13):201, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
201.pdf.

Anonymous:1893:APC

[4039] Anonymous. Apparatus for preventing collisions of trains. Scientific American, 69(13):201–202, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
201a.pdf.

Anonymous:1893:WSB

[4040] Anonymous. The Weed steak boiler. Scientific American, 69(13): 202–203, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
202.pdf.

Anonymous:1893:WFEa

[4041] Anonymous. A World’s Fair Exhibit of hammers, edge tools, etc. Scientific American, 69(13):203–204, September 23, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n13/pdf/scientificamerican09231893-203.pdf.

Anonymous:1893:BNBbi

[4042] Anonymous. Books: New books and publications. Scientific American, 69(13):205, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
205a.pdf.

Anonymous:1893:BPbl

[4043] Anonymous. Business and personal. Scientific American, 69 (13):205, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
205b.pdf.

Anonymous:1893:RPIbk

[4044] Anonymous. Recently patented inventions. Scientific American, 69 (13):205, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
205.pdf.

Anonymous:1893:NQbk

[4045] Anonymous. Notes and queries. Scientific American, 69(13): 205–206, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
205c.pdf.

Anonymous:1893:Ibn

[4046] Anonymous. To inventors. Scientific American, 69(13):206, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n13/pdf/scientificamerican09231893-206.pdf.

Anonymous:1893:IIbm

[4047] Anonymous. Index of inventions. Scientific American, 69(13): 206, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
206a.pdf.

Anonymous:1893:PNPbk

[4048] Anonymous. Patent notes: Patents. Scientific American, 69 (13):208, September 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n13/pdf/scientificamerican09231893-
208.pdf.

Anonymous:1893:ASA

[4049] Anonymous. Automatic safety appliances recently introduced on a New York railroad. Scientific American, 69(14):210, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n14/pdf/scientificamerican09301893-210c.pdf.

Anonymous:1893:HRS

[4050] Anonymous. Health rules in six words. Scientific American, 69 (14):210, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
210b.pdf.

Anonymous:1893:MP

[4051] Anonymous. Mushroom poisoning. Scientific American, 69(14): 210, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
210a.pdf.

Anonymous:1893:NCS

[4052] Anonymous. The new Cunard steamers. Scientific American, 69 (14):210, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
210.pdf.

Anonymous:1893:NWCw

[4053] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(14):211–212, September 30, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n14/pdf/scientificamerican09301893-211.pdf.

Anonymous:1893:FWT

[4054] Anonymous. Fine watch tools and machinery. Scientific American, 69(14):212, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
212.pdf.

Anonymous:1893:GEEb

[4055] Anonymous. A gas engine exhibit at the Fair. Scientific American, 69(14):212, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
212a.pdf.

Anonymous:1893:SIV

[4056] Anonymous. A snow and ice velocipede. Scientific American, 69 (14):213, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
213.pdf.

Anonymous:1893:WFEb

[4057] Anonymous. A World’s Fair Exhibit of tinners’ tools, machinery, etc. Scientific American, 69(14):213, September 30, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n14/pdf/scientificamerican09301893-213a.pdf.

Anonymous:1893:KMO

[4058] Anonymous. Knitting machines operated by electricity at the Fair. Scientific American, 69(14):213–214, September 30, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n14/pdf/scientificamerican09301893-213b.pdf.

Anonymous:1893:TC

[4059] Anonymous. A toy cannon. Scientific American, 69(14):214, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n14/pdf/scientificamerican09301893-214a.pdf.

Anonymous:1893:VCI

[4060] Anonymous. The vigilant to contest the International Yacht Race. Scientific American, 69(14):214, September 30, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n14/pdf/scientificamerican09301893-214.pdf.

Anonymous:1893:HWJ

[4061] Anonymous. The H. W. Johns Manufacturing Company’s exhibits at the Fair. Scientific American, 69(14):215, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n14/pdf/scientificamerican09301893-215.pdf.

Anonymous:1893:OTM

[4062] Anonymous. The Olsen testing machines. Scientific American, 69 (14):216, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
216.pdf.

Anonymous:1893:MPM

[4063] Anonymous. A monster planing machine. Scientific American, 69 (14):216–217, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
216a.pdf.

Anonymous:1893:WCEo

[4064] Anonymous. The World’s Columbian Exposition — the Italian Exhibit. Scientific American, 69(14):217–219, September 30, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n14/pdf/scientificamerican09301893-217.pdf.

Anonymous:1893:NSI

[4065] Anonymous. Novel scene in India. Scientific American, 69(14): 219, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
219.pdf.

Anonymous:1893:BNBbj

[4066] Anonymous. Books: New books and publications. Scientific American, 69(14):220, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
220a.pdf.

Anonymous:1893:BPbm

[4067] Anonymous. Business and personal. Scientific American, 69 (14):220, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
220b.pdf.

Anonymous:1893:RPIbl

[4068] Anonymous. Recently patented inventions. Scientific American, 69 (14):220, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
220.pdf.

Anonymous:1893:NQbl

[4069] Anonymous. Notes and queries. Scientific American, 69(14): 220–221, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
220c.pdf.

Anonymous:1893:Ibo

[4070] Anonymous. To inventors. Scientific American, 69(14):221, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n14/pdf/scientificamerican09301893-221.pdf.

Anonymous:1893:IIbn

[4071] Anonymous. Index of inventions. Scientific American, 69(14): 221–222, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
221a.pdf.

Anonymous:1893:PNPbl

[4072] Anonymous. Patent notes: Patents. Scientific American, 69 (14):224, September 30, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n14/pdf/scientificamerican09301893-
224.pdf.

Anonymous:1893:CSG

[4073] Anonymous. Crosby steam gauge and valve Co.’s exhibit. Scientific American, 69(15):225–226, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
225.pdf.

Anonymous:1893:BI

[4074] Anonymous. The Brooklyn Institute. Scientific American, 69(15):226, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n15/pdf/scientificamerican10071893-226c.pdf.

Anonymous:1893:LKW

[4075] Anonymous. The Lake Keuka Wine Company Exhibit. Scientific American, 69(15):226, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
226a.pdf.

Anonymous:1893:LSSb

[4076] Anonymous. Lubricators and safety steam appliances at the Fair. Scientific American, 69(15):226, October 7, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n15/pdf/scientificamerican10071893-226.pdf.

Anonymous:1893:PAT

[4077] Anonymous. Protection against train robbers. Scientific American, 69(15):226, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
226b.pdf.

Anonymous:1893:SSF

[4078] Anonymous. A suggested School of Fire Extinguishment. Scientific American, 69(15):226, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
226d.pdf.

Anonymous:1893:NWCx

[4079] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(15):227, October 7, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n15/pdf/scientificamerican10071893-227.pdf.

Anonymous:1893:CRSa

[4080] Anonymous. A convenient reading stand. Scientific American, 69(15):228, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
228.pdf.

Anonymous:1893:ICd

[4081] Anonymous. An improved concentrator. Scientific American, 69 (15):228, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
228b.pdf.

Anonymous:1893:IWV

[4082] Anonymous. An improved wheel for velocipedes, etc. Scientific American, 69(15):228, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
228a.pdf.

Anonymous:1893:EBF

[4083] Anonymous. The Electricity Building at the Fair. Scientific American, 69(15):229, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
229.pdf.

Anonymous:1893:Cbb

[4084] Anonymous. Correspondence. Scientific American, 69(15):231, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n15/pdf/scientificamerican10071893-231.pdf.

Anonymous:1893:WRT

[4085] Anonymous. A wire rope tramway operated at the Fair. Scientific American, 69(15):232–233, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
232.pdf.

Anonymous:1893:LLS

[4086] Anonymous. Light locomotives shown at the Fair. Scientific American, 69(15):233–234, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
233.pdf.

Dyche:1893:MLA

[4087] L. L. Dyche. Mounting of large animals. Scientific American, 69(15):234, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
234.pdf.

Anonymous:1893:BNBbk

[4088] Anonymous. Books: New books and publications. Scientific American, 69(15):236, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
236a.pdf.

Anonymous:1893:RPIbm

[4089] Anonymous. Recently patented inventions. Scientific American, 69(15):236, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n15/pdf/scientificamerican10071893-
236.pdf.

Anonymous:1893:BPbn

[4090] Anonymous. Business and personal. Scientific American, 69(15):237, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n15/pdf/scientificamerican10071893-237.pdf.

Anonymous:1893:Ibp

[4091] Anonymous. To inventors. Scientific American, 69(15):237, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n15/pdf/scientificamerican10071893-237b.pdf.

Anonymous:1893:NQbm

[4092] Anonymous. Notes and queries. Scientific American, 69(15):237, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n15/pdf/scientificamerican10071893-237a.pdf.

Anonymous:1893:IIbo

[4093] Anonymous. Index of inventions. Scientific American, 69(15):237–238, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n15/pdf/scientificamerican10071893-237c.pdf.

Anonymous:1893:PNPbm

[4094] Anonymous. Patent notes: Patents. Scientific American, 69(15):240, October 7, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n15/pdf/scientificamerican10071893-240.pdf.

Anonymous:1893:RBM

[4095] Anonymous. The Reece buttonhole machine. Scientific American, 69(16):241, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
241.pdf.

Anonymous:1893:CPT

[4096] Anonymous. Congress on patents, trade marks and inventions. Scientific American, 69(16):242, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
242.pdf.

Anonymous:1893:BW

[4097] Anonymous. Blowing up a wreck. Scientific American, 69(16): 242–243, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
242a.pdf.

Anonymous:1893:NWCy

[4098] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(16):243, October 14, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n16/pdf/scientificamerican10141893-243.pdf.

Anonymous:1893:BAC

[4099] Anonymous. The Beveridge automatic cooker. Scientific American, 69(16):244, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
244b.pdf.

Anonymous:1893:CEMa

[4100] Anonymous. A cheap and efficient motor. Scientific American, 69(16):244, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
244c.pdf.

Anonymous:1893:FRC

[4101] Anonymous. A folding and reclining chair. Scientific American, 69(16):244, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
244.pdf.

Anonymous:1893:ICBc

[4102] Anonymous. An improved commutator brush. Scientific American, 69(16):244, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
244d.pdf.

Anonymous:1893:IFC

[4103] Anonymous. An improved friction clutch. Scientific American, 69(16):244, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
244a.pdf.

Anonymous:1893:RIW

[4104] Anonymous. Recording instruments at the World’s Fair. Scientific American, 69(16):245, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
245.pdf.

Anonymous:1893:BJ

[4105] Anonymous. Benjamin Jowett. Scientific American, 69(16):247, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n16/pdf/scientificamerican10141893-247.pdf.

Anonymous:1893:PH

[4106] Anonymous. Professor Helmholtz. Scientific American, 69(16):247, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n16/pdf/scientificamerican10141893-247b.pdf.

Anonymous:1893:SVE

[4107] Anonymous. Steam valves, etc., at the Fair. Scientific American, 69(16):247, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
247a.pdf.

Anonymous:1893:EIM

[4108] Anonymous. Electrical illumination of a monument. Scientific American, 69(16):248, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
248.pdf.

Anonymous:1893:WCEp

[4109] Anonymous. The World’s Columbian Exposition — the Palace of Machinery. Scientific American, 69(16):249–250, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n16/pdf/scientificamerican10141893-249.pdf.

Anonymous:1893:YCI

[4110] Anonymous. The yachts contending for the International Championship. Scientific American, 69(16):250–251, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n16/pdf/scientificamerican10141893-250.pdf.

Anonymous:1893:RPIbn

[4111] Anonymous. Recently patented inventions. Scientific American, 69(16):252, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
252.pdf.

Anonymous:1893:BNBbl

[4112] Anonymous. Books: New books and publications. Scientific American, 69(16):253, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
253b.pdf.

Anonymous:1893:BPbo

[4113] Anonymous. Business and personal. Scientific American, 69 (16):253, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
253.pdf.

Anonymous:1893:NQbn

[4114] Anonymous. Notes and queries. Scientific American, 69(16):253, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n16/pdf/scientificamerican10141893-253a.pdf.

Anonymous:1893:IIbp

[4115] Anonymous. Index of inventions. Scientific American, 69(16): 253–254, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
253c.pdf.

Anonymous:1893:PNPbn

[4116] Anonymous. Patent notes: Patents. Scientific American, 69 (16):256, October 14, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n16/pdf/scientificamerican10141893-
256.pdf.

Anonymous:1893:GSO

[4117] Anonymous. The Great Storm of October 2. Scientific American, 69(17):258, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
258.pdf.

Anonymous:1893:NWCz

[4118] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(17):259, October 21, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n17/pdf/scientificamerican10211893-259.pdf.

Anonymous:1893:LIP

[4119] Anonymous. Lawton’s improved pottery kiln. Scientific American, 69(17):260, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
260a.pdf.

Anonymous:1893:WCEq

[4120] Anonymous. The World’s Columbian Exposition — Exhibit of Otis C. White, Worcester, Mass. Scientific American, 69(17):260, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n17/pdf/scientificamerican10211893-260.pdf.

Anonymous:1893:SEM

[4121] Anonymous. A simple electrical machine. Scientific American, 69 (17):260–261, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
260b.pdf.

Anonymous:1893:SWF

[4122] Anonymous. Sculpture at the World’s Fair. Scientific American, 69(17):261, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
261.pdf.

Anonymous:1893:WCEr

[4123] Anonymous. The World’s Columbian Exposition — exhibit of the water circulating grate company. Scientific American, 69(17):261, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n17/pdf/scientificamerican10211893-261a.pdf.

Anonymous:1893:NST

[4124] Anonymous. A new steam turbine. Scientific American, 69(17):263, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n17/pdf/scientificamerican10211893-263.pdf.

Anonymous:1893:ACW

[4125] Anonymous. The America’s Cup won by the Vigilant. Scientific American, 69(17):264, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
264.pdf.

Anonymous:1893:WCEs

[4126] Anonymous. The World’s Columbian Exposition — a view on the Grand Basin. Scientific American, 69(17):265, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n17/pdf/scientificamerican10211893-265.pdf.

Hovey:1893:CFA

[4127] H. C. Hovey. Cactus furniture and art ware at the World’s Columbian Exposition. Scientific American, 69(17):266–267, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n17/pdf/scientificamerican10211893-266.pdf.

Anonymous:1893:MCH

[4128] Anonymous. The mysterious Chapel of Heron. Scientific American, 69(17):267, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
267.pdf.

Anonymous:1893:BNBbm

[4129] Anonymous. Books: New books and publications. Scientific American, 69(17):268, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
268a.pdf.

Anonymous:1893:RPIbo

[4130] Anonymous. Recently patented inventions. Scientific American, 69(17):268, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
268.pdf.

Anonymous:1893:BPbp

[4131] Anonymous. Business and personal. Scientific American, 69 (17):269, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
269.pdf.

Anonymous:1893:Ibq

[4132] Anonymous. To inventors. Scientific American, 69(17):269, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n17/pdf/scientificamerican10211893-269b.pdf.

Anonymous:1893:NQbo

[4133] Anonymous. Notes and queries. Scientific American, 69(17):269, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n17/pdf/scientificamerican10211893-269a.pdf.

Anonymous:1893:IIbq 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
269c.pdf.

Anonymous:1893:PNPbo

[4135] Anonymous. Patent notes: Patents. Scientific American, 69 (17):272, October 21, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n17/pdf/scientificamerican10211893-
272.pdf.

Anonymous:1893:OEC

[4136] Anonymous. The Otis Elevators at the Columbian Exposition. Scientific American, 69(18):273–274, October 28, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n18/pdf/scientificamerican10281893-273.pdf.

Anonymous:1893:HBE

[4137] Anonymous. How to become an electrical engineer. Scientific American, 69(18):274, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
274.pdf.

Anonymous:1893:NWCba

[4138] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(18):275, October 28, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n18/pdf/scientificamerican10281893-275.pdf.

Anonymous:1893:LSRb

[4139] Anonymous. Lathe slide rest attachment. Scientific American, 69(18):276, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
276a.pdf.

Anonymous:1893:TS

[4140] Anonymous. The teaching of swimming. Scientific American, 69 (18):276, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
276.pdf.

Anonymous:1893:POF

[4141] Anonymous. Pianos and organs at the Fair. Scientific American, 69(18):277, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
277a.pdf.

Anonymous:1893:SMS

[4142] Anonymous. Suckow’s machine for squeezing puddlers’ balls and the Mannesmann seamless tube machine. Scientific American, 69 (18):277, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
277.pdf.

Anonymous:1893:DEP

[4143] Anonymous. A durable and efficient pump. Scientific American, 69 (18):277–278, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
277b.pdf.

Anonymous:1893:NAA

[4144] Anonymous. New apparatus for artificial respiration. Scientific American, 69(18):280, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
280.pdf.

Anonymous:1893:PCW

[4145] Anonymous. Prize cattle at the World’s Columbian Exposition. Scientific American, 69(18):281, October 28, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n18/pdf/scientificamerican10281893-281.pdf.

Anonymous:1893:Cbc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n18/pdf/scientificamerican10281893-282.pdf.

Anonymous:1893:BSM

[4147] Anonymous. Beveling and silvering mirrors. Scientific American, 69(18):283, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
283.pdf.

Anonymous:1893:BNBbn

[4148] Anonymous. Books: New books and publications. Scientific American, 69(18):284, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
284a.pdf.

Anonymous:1893:RPIbp

[4149] Anonymous. Recently patented inventions. Scientific American, 69(18):284, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
284.pdf.

Anonymous:1893:BPbq

[4150] Anonymous. Business and personal. Scientific American, 69 (18):285, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
285.pdf.

Anonymous:1893:Ibr

[4151] Anonymous. To inventors. Scientific American, 69(18):285, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n18/pdf/scientificamerican10281893-285b.pdf.

Anonymous:1893:NQbp

[4152] Anonymous. Notes and queries. Scientific American, 69(18):285, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n18/pdf/scientificamerican10281893-285a.pdf.

Anonymous:1893:IIbr

[4153] Anonymous. Index of inventions. Scientific American, 69(18): 285–286, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
285c.pdf.

Anonymous:1893:PNPbp

[4154] Anonymous. Patent notes: Patents. Scientific American, 69 (18):288, October 28, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n18/pdf/scientificamerican10281893-
288.pdf.

Anonymous:1893:CWC

[4155] Anonymous. Closing of the World’s Columbian Exposition. Scientific American, 69(19):290, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
290.pdf.

Anonymous:1893:CCS

[4156] Anonymous. Coaling cruisers at sea. Scientific American, 69(19): 290–291, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
290a.pdf.

Anonymous:1893:NWCbb

[4157] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(19):291, November 4, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n19/pdf/scientificamerican11041893-291.pdf.

Anonymous:1893:ACE

[4158] Anonymous. An auxiliary cut-off for engines. Scientific American, 69(19):292, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
292b.pdf.

Anonymous:1893:ISCc

[4159] Anonymous. An improved screw cutting lathe. Scientific American, 69(19):292, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
292.pdf.

Anonymous:1893:TBWb

[4160] Anonymous. Torsion braided wire mattresses, pillows, cushions, etc., shown at the Exposition. Scientific American, 69(19):292, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n19/pdf/scientificamerican11041893-292a.pdf.

Anonymous:1893:CNRc

[4161] Anonymous. A cheap and novel road bed. Scientific American, 69(19):293, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
293a.pdf.

Anonymous:1893:EPE

[4162] Anonymous. An electric passenger elevator at the Exposition. Scientific American, 69(19):293, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
293c.pdf.

Anonymous:1893:ISE

[4163] Anonymous. An improved steam engine valve chest. Scientific American, 69(19):293, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
293b.pdf.

Anonymous:1893:KMM

[4164] Anonymous. A kneading or mixing machine. Scientific American, 69(19):293, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
293.pdf.

Anonymous:1893:GAC

[4165] Anonymous. A great air compressor at the Fair. Scientific American, 69(19):294, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
294.pdf.

Anonymous:1893:EBE

[4166] Anonymous. Exhibit of Brush Electric Company. Scientific American, 69(19):295–296, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
295.pdf.

Anonymous:1893:DDM

[4167] Anonymous. Duhamel Du Monceau. Scientific American, 69 (19):296, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
296a.pdf.

Anonymous:1893:DGC

[4168] Anonymous. Disappearing gun carriage. Scientific American, 69 (19):296, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
296.pdf.

Anonymous:1893:SEb

[4169] Anonymous. Swordfish exploits. Scientific American, 69(19): 297, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
297.pdf.

Anonymous:1893:BB

[4170] Anonymous. A beeswax bleachery. Scientific American, 69(19): 297–298, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
297a.pdf.

Anonymous:1893:DTR

[4171] Anonymous. Delcroix’s topographic rule. Scientific American, 69 (19):298, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
298a.pdf.

Anonymous:1893:RRS

[4172] Anonymous. A railroad rail for street railways. Scientific American, 69(19):298, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
298.pdf.

Anonymous:1893:ETS

[4173] Anonymous. Exhibits of thoroughbred stock at the World’s Columbian Exposition. Scientific American, 69(19):299, November 4, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n19/pdf/scientificamerican11041893-299.pdf.

Anonymous:1893:BNBbo

[4174] Anonymous. Books: New books and publications. Scientific American, 69(19):300, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
300a.pdf.

Anonymous:1893:RPIbq

[4175] Anonymous. Recently patented inventions. Scientific American, 69(19):300, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
300.pdf.

Anonymous:1893:BPbr

[4176] Anonymous. Business and personal. Scientific American, 69 (19):301, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
301.pdf.

Anonymous:1893:Ibu

[4177] Anonymous. To inventors. Scientific American, 69(19):301, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n19/pdf/scientificamerican11041893-301b.pdf.

Anonymous:1893:NQbq

[4178] Anonymous. Notes and queries. Scientific American, 69(19): 301, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
301a.pdf.

Anonymous:1893:IIbs

[4179] Anonymous. Index of inventions. Scientific American, 69(19): 301–302, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
301c.pdf.

Anonymous:1893:PNPbq

[4180] Anonymous. Patent notes: Patents. Scientific American, 69 (19):304, November 4, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n19/pdf/scientificamerican11041893-
304.pdf.

Anonymous:1893:OGEa

[4181] Anonymous. The Otto gas engines at the Columbian Exposition. Scientific American, 69(20):305–306, November 11, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n20/pdf/scientificamerican11111893-305.pdf.

Anonymous:1893:ANO

[4182] Anonymous. Amateur Naval operations. Scientific American, 69 (20):306–307, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
306.pdf.

Anonymous:1893:NWCbc

[4183] Anonymous. Notes from the World’s Columbian Exposition Chicago 1893. Scientific American, 69(20):307, November 11, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n20/pdf/scientificamerican11111893-307.pdf.

Anonymous:1893:PSI

[4184] Anonymous. The “Paragon” speed indicator. Scientific American, 69(20):308, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
308.pdf.

Anonymous:1893:BMC

[4185] Anonymous. Boring machine for Corliss engine cylinders. Scientific American, 69(20):308–309, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
308a.pdf.

Anonymous:1893:HC

[4186] Anonymous. A heater and condenser. Scientific American, 69 (20):309, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
309a.pdf.

Anonymous:1893:IEMb

[4187] Anonymous. An improved electric motor truck. Scientific American, 69(20):309, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
309.pdf.

Felt:1893:MA

[4188] Dorr E. Felt. Mechanical arithmetic. Scientific American, 69(20): 309–310, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
309b.pdf.

Anonymous:1893:OGEb

[4189] Anonymous. Optical goods at the Exposition. Scientific American, 69(20):310, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
310.pdf.

Anonymous:1893:RE

[4190] Anonymous. Russian exhibits. Scientific American, 69(20):311– 312, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
311.pdf.

Anonymous:1893:ICF

[4191] Anonymous. Improved construction of floors, ceilings, arches, etc. Scientific American, 69(20):312, November 11, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n20/pdf/scientificamerican11111893-312.pdf.

Anonymous:1893:WEC

[4192] Anonymous. Western Electric Company’s luminous sign. Scientific American, 69(20):312–313, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
312a.pdf.

Anonymous:1893:WCEt

[4193] Anonymous. The World’s Columbian Exposition — the Court of Honor. Scientific American, 69(20):313, November 11, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n20/pdf/scientificamerican11111893-313.pdf.

Anonymous:1893:Cbe

[4194] Anonymous. Correspondence. Scientific American, 69(20):314– 315, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
314.pdf.

Anonymous:1893:AHE

[4195] Anonymous. American Historical Exposition in Madrid. Scientific American, 69(20):315, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
315.pdf.

Anonymous:1893:GEA

[4196] Anonymous. A geometrical educational appliance. Scientific American, 69(20):315, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
315a.pdf.

Anonymous:1893:RPIbr

[4197] Anonymous. Recently patented inventions. Scientific American, 69(20):316, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
316.pdf.

Anonymous:1893:BNBbp

[4198] Anonymous. Books: New books and publications. Scientific American, 69(20):316–317, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
316a.pdf.

Anonymous:1893:BPbs

[4199] Anonymous. Business and personal. Scientific American, 69 (20):317, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
317.pdf.

Anonymous:1893:Ibw

[4200] Anonymous. To inventors. Scientific American, 69(20):317, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n20/pdf/scientificamerican11111893-317b.pdf.

Anonymous:1893:NQbr

[4201] Anonymous. Notes and queries. Scientific American, 69(20): 317, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
317a.pdf.

Anonymous:1893:IIbt

[4202] Anonymous. Index of inventions. Scientific American, 69(20): 317–318, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
317c.pdf.

Anonymous:1893:PNPbr

[4203] Anonymous. Patent notes: Patents. Scientific American, 69 (20):319, November 11, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n20/pdf/scientificamerican11111893-
319.pdf.

Anonymous:1893:SBC

[4204] Anonymous. A storage battery for central station service. Scientific American, 69(21):321–322, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
321.pdf.

Anonymous:1893:DGA

[4205] Anonymous. Death by gas asphyxiation and poisoning. Scientific American, 69(21):322–323, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
322.pdf.

Anonymous:1893:IREa

[4206] Anonymous. An improved rotary engine. Scientific American, 69 (21):324, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
324a.pdf.

Anonymous:1893:MEU

[4207] Anonymous. A movable elevator for use on store floors. Scientific American, 69(21):324, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
324b.pdf.

Anonymous:1893:PPL

[4208] Anonymous. The “Paragon” projection lantern. Scientific American, 69(21):324, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
324.pdf.

Anonymous:1893:IPS

[4209] Anonymous. An improved propeller sleigh. Scientific American, 69 (21):324–325, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
324c.pdf.

Anonymous:1893:RDC

[4210] Anonymous. The Rand Drill Company’s compound duplex air compressor and rock drills at the Columbian Exposition. Scientific American, 69(21):325–326, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
325.pdf.

Anonymous:1893:SAM

[4211] Anonymous. The Scientific American match safe. Scientific American, 69(21):326, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
326.pdf.

Anonymous:1893:MJS

[4212] Anonymous. Magnetic Jack straws. Scientific American, 69(21): 326–327, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
326a.pdf.

Anonymous:1893:BC

[4213] Anonymous. Brooks’ Comet of 1893. Scientific American, 69 (21):327, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
327.pdf.

Anonymous:1893:Cbf

[4214] Anonymous. Correspondence. Scientific American, 69(21):327, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n21/pdf/scientificamerican11181893-327a.pdf.

Anonymous:1893:MNL

[4215] Anonymous. Movable night lamp for stairways. Scientific American, 69(21):328, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
328.pdf.

Anonymous:1893:NDO

[4216] Anonymous. The new Dudley Observatory, in Albany. Scientific American, 69(21):328, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
328a.pdf.

Anonymous:1893:WCEu

[4217] Anonymous. The World’s Columbian Exposition — the Indian scout. Scientific American, 69(21):329–330, November 18, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n21/pdf/scientificamerican11181893-329.pdf.

Anonymous:1893:NTd

[4218] Anonymous. Noriega’s telephone. Scientific American, 69(21): 330–331, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
330.pdf.

Anonymous:1893:FP

[4219] Anonymous. The flying propeller. Scientific American, 69(21): 331, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
331a.pdf.

Anonymous:1893:SSEb

[4220] Anonymous. Sheffield Steel at the Exposition. Scientific American, 69(21):331, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
331.pdf.

Anonymous:1893:BNBbq

[4221] Anonymous. Books: New books and publications. Scientific American, 69(21):332, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
332a.pdf.

Anonymous:1893:RPIbs

[4222] Anonymous. Recently patented inventions. Scientific American, 69(21):332, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
332.pdf.

Anonymous:1893:BPbt

[4223] Anonymous. Business and personal. Scientific American, 69 (21):333, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
333.pdf.

Anonymous:1893:NQbs

[4224] Anonymous. Notes and queries. Scientific American, 69(21): 333, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
333a.pdf.

Anonymous:1893:Ibx

[4225] Anonymous. To inventors. Scientific American, 69(21):334, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n21/pdf/scientificamerican11181893-334.pdf.

Anonymous:1893:IIbu

[4226] Anonymous. Index of inventions. Scientific American, 69(21): 334, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
334a.pdf.

Anonymous:1893:PNPbs

[4227] Anonymous. Patent notes: Patents. Scientific American, 69 (21):335, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n21/pdf/scientificamerican11181893-
335.pdf.

Anonymous:1893:Pb

[4228] Anonymous. Patents. Scientific American, 69(21):336, November 18, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1893:ELC

[4229] Anonymous. Electric launches at the Columbian Exposition. Scientific American, 69(22):337, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
337.pdf.

Anonymous:1893:MEN

[4230] Anonymous. More engineers needed for the Navy. Scientific American, 69(22):338, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
338.pdf.

Anonymous:1893:CNC

[4231] Anonymous. The Columbia, the new Commerce destroyer. Scientific American, 69(22):338–339, November 25, 1893. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
338a.pdf.

Anonymous:1893:EELb

[4232] Anonymous. An extensible electric light support. Scientific American, 69(22):340, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
340c.pdf.

Anonymous:1893:ICCd

[4233] Anonymous. Improved cross-cutting saw. Scientific American, 69 (22):340, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
340a.pdf.

Anonymous:1893:IREb

[4234] Anonymous. An improved rotary engine. Scientific American, 69 (22):340, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
340.pdf.

Anonymous:1893:RAC

[4235] Anonymous. A readily adjustable camera stand. Scientific American, 69(22):340, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
340b.pdf.

Anonymous:1893:CHD

[4236] Anonymous. A coin holding and delivering device. Scientific American, 69(22):341, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
341.pdf.

Anonymous:1893:QSC

[4237] Anonymous. A quick and sure chicken beheader. Scientific American, 69(22):341, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
341a.pdf.

Anonymous:1893:NFA

[4238] Anonymous. New fire alarms. Scientific American, 69(22):341– 342, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
341b.pdf.

Anonymous:1893:MBb

[4239] Anonymous. A metallic bookcase. Scientific American, 69(22): 342, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
342.pdf.

Anonymous:1893:ISf

[4240] Anonymous. Improvements in shapers. Scientific American, 69 (22):344, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
344.pdf.

Anonymous:1893:FHH

[4241] Anonymous. Four hundred horse power mill engine. Scientific American, 69(22):345–346, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
345.pdf.

Anonymous:1893:SGM

[4242] Anonymous. A simple gas meter connection. Scientific American, 69(22):346, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
346.pdf.

Anonymous:1893:SPB

[4243] Anonymous. Securing propeller blades to the hub. Scientific American, 69(22):346, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
346a.pdf.

Anonymous:1893:TSW

[4244] Anonymous. The triple-screw war ship Columbia. Scientific American, 69(22):347, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
347.pdf.

Anonymous:1893:BPbu

[4245] Anonymous. Business and personal. Scientific American, 69 (22):348, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
348a.pdf.

Anonymous:1893:RPIbt

[4246] Anonymous. Recently patented inventions. Scientific American, 69(22):348, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
348.pdf.

Anonymous:1893:NQbt

[4247] Anonymous. Notes and queries. Scientific American, 69(22): 348–349, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
348b.pdf.

Anonymous:1893:Iby

[4248] Anonymous. To inventors. Scientific American, 69(22):349, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n22/pdf/scientificamerican11251893-349.pdf.

Anonymous:1893:IIbv

[4249] Anonymous. Index of inventions. Scientific American, 69(22): 349–350, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
349a.pdf.

Anonymous:1893:PNPbt

[4250] Anonymous. Patent notes: Patents. Scientific American, 69 (22):352, November 25, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n22/pdf/scientificamerican11251893-
352.pdf.

Anonymous:1893:RCP

[4251] Anonymous. Report of the Commissioner of Patents. Scientific American, 69(23):354, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
354.pdf.

Anonymous:1893:RTM

[4252] Anonymous. The registration of trade marks. Scientific American, 69(23):354, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
354a.pdf.

Anonymous:1893:TED

[4253] Anonymous. Terrible effects of dynamite. Scientific American, 69(23):354, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
354b.pdf.

Anonymous:1893:NBT

[4254] Anonymous. New British torpedo boats of extraordinary speed. Scientific American, 69(23):354–355, December 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n23/pdf/scientificamerican12021893-354c.pdf.

Anonymous:1893:NBN

[4255] Anonymous. The new Brazilian Navy. Scientific American, 69 (23):355, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
355.pdf.

Anonymous:1893:NYC

[4256] Anonymous. The New York Contingent of the Brazilian Navy. Scientific American, 69(23):355, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
355a.pdf.

Anonymous:1893:BCH

[4257] Anonymous. A burner for cooking or heating stoves. Scientific American, 69(23):356, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
356.pdf.

Anonymous:1893:CST

[4258] Anonymous. A cheap and simple thill support. Scientific American, 69(23):356, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
356b.pdf.

Anonymous:1893:EJM

[4259] Anonymous. Exhibit of Jarecki Manufacturing Company. Scientific American, 69(23):356, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
356a.pdf.

Anonymous:1893:BET

[4260] Anonymous. Boiler-end turning, boring, and drilling machine. Scientific American, 69(23):357–358, December 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n23/pdf/scientificamerican12021893-357.pdf.

Anonymous:1893:Cbg

[4261] Anonymous. Correspondence. Scientific American, 69(23):359, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n23/pdf/scientificamerican12021893-359a.pdf.

Anonymous:1893:DFC

[4262] Anonymous. Display of fireworks at the Columbian Exposition. Scientific American, 69(23):359, December 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n23/pdf/scientificamerican12021893-359.pdf.

Anonymous:1893:MSC

[4263] Anonymous. Manufacture of stearin candles. Scientific American, 69(23):360, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
360.pdf.

Anonymous:1893:CEMb

[4264] Anonymous. The Columbian Expositions — a “Moonshiner’s” plant. Scientific American, 69(23):360–361, December 2, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n23/pdf/scientificamerican12021893-360a.pdf.

Anonymous:1893:WSV

[4265] Anonymous. A whale stranded at Villerville. Scientific American, 69(23):361, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
361.pdf.

Anonymous:1893:OPC

[4266] Anonymous. Our photograph of the Columbia. Scientific American, 69(23):361–363, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
361a.pdf.

Anonymous:1893:CESb

[4267] Anonymous. The Columbian Exposition — statue of “The North”. Scientific American, 69(23):363, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
363.pdf.

Anonymous:1893:BNBbr

[4268] Anonymous. Books: New books and publications. Scientific American, 69(23):364, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
364a.pdf.

Anonymous:1893:RPIbu

[4269] Anonymous. Recently patented inventions. Scientific American, 69(23):364, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
364.pdf.

Anonymous:1893:BPbv

[4270] Anonymous. Business and personal. Scientific American, 69 (23):365, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
365.pdf.

Anonymous:1893:Ibz

[4271] Anonymous. To inventors. Scientific American, 69(23):365, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n23/pdf/scientificamerican12021893-365b.pdf.

Anonymous:1893:NQbu

[4272] Anonymous. Notes and queries. Scientific American, 69(23):365, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n23/pdf/scientificamerican12021893-365a.pdf.

Anonymous:1893:IIbw

[4273] Anonymous. Index of inventions. Scientific American, 69(23): 365–366, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
365c.pdf.

Anonymous:1893:PNPbu

[4274] Anonymous. Patent notes: Patents. Scientific American, 69 (23):368, December 2, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n23/pdf/scientificamerican12021893-
368.pdf.

Anonymous:1893:G

[4275] Anonymous. “Gasocution”. Scientific American, 69(24):370, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n24/pdf/scientificamerican12091893-370.pdf.

Anonymous:1893:WLH

[4276] Anonymous. The wreck of the Louise H. Randall. Scientific American, 69(24):370–371, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
370a.pdf.

Anonymous:1893:INL

[4277] Anonymous. An improved nut lock. Scientific American, 69 (24):372, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
372.pdf.

Anonymous:1893:PV

[4278] Anonymous. A pneumatic ventilator. Scientific American, 69 (24):372, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
372a.pdf.

Anonymous:1893:WSP

[4279] Anonymous. Work on the sewers of Paris. Scientific American, 69 (24):372–373, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
372b.pdf.

Anonymous:1893:DMD

[4280] Anonymous. A digging machine to deepen channels, etc. Scientific American, 69(24):373, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
373.pdf.

Anonymous:1893:IRTb

[4281] Anonymous. An improved recording thermometer. Scientific American, 69(24):373, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
373a.pdf.

Anonymous:1893:CRSb

[4282] Anonymous. A combined rubber stamp vulcanizer and press. Scientific American, 69(24):373–374, December 9, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n24/pdf/scientificamerican12091893-373b.pdf.

Anonymous:1893:Cbh

[4283] Anonymous. Correspondence. Scientific American, 69(24):375, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n24/pdf/scientificamerican12091893-375a.pdf.

Anonymous:1893:RCO

[4284] Anonymous. Railroad construction in old and modern times. Scientific American, 69(24):375, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
375.pdf.

Anonymous:1893:MBO

[4285] Anonymous. The Mont Blanc Observatory. Scientific American, 69(24):376, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
376.pdf.

Anonymous:1893:EPC

[4286] Anonymous. Electric propulsion of canal boats. Scientific American, 69(24):376–377, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
376a.pdf.

Anonymous:1893:MDC

[4287] Anonymous. The manufacture of desiccated cocoanut. Scientific American, 69(24):377–378, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
377.pdf.

Anonymous:1893:CEE

[4288] Anonymous. The Columbian Exposition — East India Building. Scientific American, 69(24):379, December 9, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n24/pdf/scientificamerican12091893-379.pdf.

Anonymous:1893:ISS

[4289] Anonymous. Improved steam steering gear. Scientific American, 69(24):379, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
379a.pdf.

Anonymous:1893:BNBbs

[4290] Anonymous. Books: New books and publications. Scientific American, 69(24):380, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
380a.pdf.

Anonymous:1893:BPbw

[4291] Anonymous. Business and personal. Scientific American, 69 (24):380, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
380b.pdf.

Anonymous:1893:RPIbv

[4292] Anonymous. Recently patented inventions. Scientific American, 69(24):380, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
380.pdf.

Anonymous:1893:NQbv

[4293] Anonymous. Notes and queries. Scientific American, 69(24): 380–381, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
380c.pdf.

Anonymous:1893:Icg

[4294] Anonymous. To inventors. Scientific American, 69(24):381, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n24/pdf/scientificamerican12091893-381.pdf.

Anonymous:1893:IIbx

[4295] Anonymous. Index of inventions. Scientific American, 69(24): 381–382, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
381a.pdf.

Anonymous:1893:PNPbv

[4296] Anonymous. Patent notes: Patents. Scientific American, 69 (24):384, December 9, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n24/pdf/scientificamerican12091893-
384.pdf.

Anonymous:1893:PJT

[4297] Anonymous. Prof. John Tyndall. Scientific American, 69(25): 386, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
386.pdf.

Anonymous:1893:SPC

[4298] Anonymous. Screen photography in colors. Scientific American, 69(25):386, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
386a.pdf.

Anonymous:1893:PAC

[4299] Anonymous. The Pennsylvania anthracite coal fields. Scientific American, 69(25):386–387, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
386b.pdf.

Anonymous:1893:ADP

[4300] Anonymous. Automatic distributor of perfumes. Scientific American, 69(25):388, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
388b.pdf.

Anonymous:1893:BPEa

[4301] Anonymous. A brake for portable engines, etc. Scientific American, 69(25):388, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
388c.pdf.

Anonymous:1893:IAM

[4302] Anonymous. An improved adding machine. Scientific American, 69(25):388, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
388.pdf.

Anonymous:1893:NSL

[4303] Anonymous. A new system of lighting street gas lamps without the use of wires. Scientific American, 69(25):388, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n25/pdf/scientificamerican12161893-388a.pdf.

Anonymous:1893:BAT

[4304] Anonymous. The Bundy automatic time recorder at the Exposition. Scientific American, 69(25):389, December 16, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n25/pdf/scientificamerican12161893-389.pdf.

Anonymous:1893:NSBb

[4305] Anonymous. A new storage battery. Scientific American, 69(25): 389–390, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
389a.pdf.

Anonymous:1893:NMF

[4306] Anonymous. A novelty for the midwinter fair. Scientific American, 69(25):390–391, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
390.pdf.

Anonymous:1893:BCRc

[4307] Anonymous. The Broadway Cable Railway. Scientific American, 69(25):392, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
392.pdf.

Anonymous:1893:USL

[4308] Anonymous. United States large caliber mortars. Scientific American, 69(25):393–395, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
393.pdf.

Anonymous:1893:ETL

[4309] Anonymous. Experiment on the tension of liquid films. Scientific American, 69(25):395, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
395a.pdf.

Michaud:1893:CAM

[4310] Gustave Michaud. A carbonic acid motor. Scientific American, 69(25):395, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
395.pdf.

Anonymous:1893:BNBbt

[4311] Anonymous. Books: New books and publications. Scientific American, 69(25):396, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
396a.pdf.

Anonymous:1893:RPIbw

[4312] Anonymous. Recently patented inventions. Scientific American, 69(25):396, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
396.pdf.

Anonymous:1893:BPbx

[4313] Anonymous. Business and personal. Scientific American, 69 (25):397, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
397.pdf.

Anonymous:1893:Ich

[4314] Anonymous. To inventors. Scientific American, 69(25):397, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n25/pdf/scientificamerican12161893-397b.pdf.

Anonymous:1893:NQbw

[4315] Anonymous. Notes and queries. Scientific American, 69(25): 397, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
397a.pdf.

Anonymous:1893:IIby

[4316] Anonymous. Index of inventions. Scientific American, 69(25): 397–398, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
397c.pdf.

Anonymous:1893:PNPbw

[4317] Anonymous. Patent notes: Patents. Scientific American, 69 (25):400, December 16, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n25/pdf/scientificamerican12161893-
400.pdf.

Anonymous:1893:LW

[4318] Anonymous. The laboratory and the workshop. Scientific American, 69(26):402–403, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
402.pdf.

Anonymous:1893:IPO

[4319] Anonymous. An improved post office box. Scientific American, 69(26):404, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
404b.pdf.

Anonymous:1893:IWe

[4320] Anonymous. An improved wrench. Scientific American, 69(26): 404, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
404c.pdf.

Anonymous:1893:MFP

[4321] Anonymous. A machine for forming projectiles. Scientific American, 69(26):404, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
404.pdf.

Anonymous:1893:PGD

[4322] Anonymous. A portable grain dump. Scientific American, 69 (26):404, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
404a.pdf.

Anonymous:1893:NAW

[4323] Anonymous. The new American war ship Olympia. Scientific American, 69(26):405, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
405.pdf.

Anonymous:1893:SDM

[4324] Anonymous. The State Dam, Mohawk River, at Cohoes, N.Y. Scientific American, 69(26):405–406, December 23, 1893. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v69/
n26/pdf/scientificamerican12231893-405a.pdf.

Anonymous:1893:Pd

[4325] Anonymous. The Photoret. Scientific American, 69(26):406, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n26/pdf/scientificamerican12231893-406.pdf.

Anonymous:1893:Cbi

[4326] Anonymous. Correspondence. Scientific American, 69(26):407, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n26/pdf/scientificamerican12231893-407.pdf.

Anonymous:1893:TGB

[4327] Anonymous. The traveling garbage burner of Chicago. Scientific American, 69(26):408–409, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
408.pdf.

Anonymous:1893:BWR

[4328] Anonymous. The Burt Wooden Railway, California. Scientific American, 69(26):409–410, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
409.pdf.

Anonymous:1893:BPby 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
411a.pdf.

Anonymous:1893:RPIbx

[4330] Anonymous. Recently patented inventions. Scientific American, 69(26):411, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
411.pdf.

Anonymous:1893:NQbx

[4331] Anonymous. Notes and queries. Scientific American, 69(26): 411–413, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
411b.pdf.

Anonymous:1893:Ici

[4332] Anonymous. To inventors. Scientific American, 69(26):413, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v69/n26/pdf/scientificamerican12231893-413.pdf.

Anonymous:1893:IIbz

[4333] Anonymous. Index of inventions. Scientific American, 69(26): 413–414, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
413a.pdf.

Anonymous:1893:PNPbx

[4334] Anonymous. Patent notes: Patents. Scientific American, 69 (26):415, December 23, 1893. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v69/n26/pdf/scientificamerican12231893-
415.pdf.

Anonymous:1894:GE

[4335] Anonymous. The Graves elevators. Scientific American, 70(1):1, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-1.pdf.

Anonymous:1894:AMPa

[4336] Anonymous. The alkali metals potassium and sodium — II. Scientific American, 70(1):2, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
2a.pdf.

Anonymous:1894:PLCa

[4337] Anonymous. Patent litigation and compromise. Scientific American, 70(1):2, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
2.pdf.

Anonymous:1894:NSI

[4338] Anonymous. The new steamers of the International Navigation Company. Scientific American, 70(1):2–3, January 6, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n1/pdf/scientificamerican01061894-2b.pdf.

Anonymous:1894:GPO

[4339] Anonymous. The growth of plants in odd places. Scientific American, 70(1):4, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
4.pdf.

Anonymous:1894:ASM

[4340] Anonymous. An automatic sculpturing machine. Scientific American, 70(1):4–5, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
4a.pdf.

Anonymous:1894:DBPa

[4341] Anonymous. A double belting press. Scientific American, 70(1):5, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-5.pdf.

Anonymous:1894:IGE

[4342] Anonymous. An improved gas engine. Scientific American, 70(1):5, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-5a.pdf.

Anonymous:1894:GAG

[4343] Anonymous. The Gadey air grate. Scientific American, 70(1):5–6, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-5b.pdf.

Anonymous:1894:HG

[4344] Anonymous. Henry Goebel. Scientific American, 70(1):6, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-6a.pdf.

Anonymous:1894:IFE

[4345] Anonymous. An improved fastening for envelopes. Scientific American, 70(1):6, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
6.pdf.

Anonymous:1894:Ca

[4346] Anonymous. Correspondence. Scientific American, 70(1):7, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-7.pdf.

Anonymous:1894:CNY

[4347] Anonymous. The cruiser New York in dry dock. Scientific American, 70(1):8, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
8.pdf.

Anonymous:1894:TSW

[4348] Anonymous. Tectorium, substitute for window glass. Scientific American, 70(1):9, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
9.pdf.

Anonymous:1894:MOZ

[4349] Anonymous. Manufacture of oxide of zinc. Scientific American, 70(1):9–10, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
9a.pdf.

Anonymous:1894:WBP

[4350] Anonymous. The Widah bird of paradise. Scientific American, 70(1):11, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-11.pdf.

Anonymous:1894:BNBa

[4351] Anonymous. Books: New books and publications. Scientific American, 70(1):12, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
12a.pdf.

Anonymous:1894:BPa

[4352] Anonymous. Business and personal. Scientific American, 70(1):12, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-12b.pdf.

Anonymous:1894:RPIa

[4353] Anonymous. Recently patented inventions. Scientific American, 70(1):12, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n1/pdf/scientificamerican01061894-
12.pdf.

Anonymous:1894:NQa

[4354] Anonymous. Notes and queries. Scientific American, 70(1):12–13, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-12c.pdf.

Anonymous:1894:Ia

[4355] Anonymous. To inventors. Scientific American, 70(1):13, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-13.pdf.

Anonymous:1894:IIa

[4356] Anonymous. Index of inventions. Scientific American, 70(1):13–14, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-13a.pdf.

Anonymous:1894:PNPa

[4357] Anonymous. Patent notes: Patents. Scientific American, 70(1):15, January 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n1/pdf/scientificamerican01061894-15.pdf.

Anonymous:1894:L

[4358] Anonymous. The “Linotype”. Scientific American, 70(2):17, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-17.pdf.

Anonymous:1894:AMPb

[4359] Anonymous. The alkali metals potassium and sodium — III. Scientific American, 70(2):18, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
18.pdf.

Anonymous:1894:PT

[4360] Anonymous. Progress in typesetting. Scientific American, 70 (2):18, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
18a.pdf.

Anonymous:1894:AEH

[4361] Anonymous. Another explosion of a high pressure gas cylinder. Scientific American, 70(2):18–19, January 13, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n2/pdf/scientificamerican01131894-18b.pdf.

Anonymous:1894:HSBa

[4362] Anonymous. The Harrison Standard Burr Stone Mills. Scientific American, 70(2):20, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
20a.pdf.

Anonymous:1894:IGa

[4363] Anonymous. An improved gate. Scientific American, 70(2):20, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-20b.pdf.

Anonymous:1894:IUa

[4364] Anonymous. An improved unicycle. Scientific American, 70(2):20, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-20.pdf.

Anonymous:1894:HLFa

[4365] Anonymous. A hay loader for field work. Scientific American, 70(2):21, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
21.pdf.

Anonymous:1894:NBW

[4366] Anonymous. New British war boats of great power and speed. Scientific American, 70(2):21, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
21a.pdf.

Anonymous:1894:IMI

[4367] Anonymous. The improved “Monitor” incubator. Scientific American, 70(2):21–22, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
21b.pdf.

Anonymous:1894:CHN

[4368] Anonymous. The Capewell horse nails. Scientific American, 70 (2):22–23, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
22.pdf.

Anonymous:1894:Cb

[4369] Anonymous. Correspondence. Scientific American, 70(2):23, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-23.pdf.

Anonymous:1894:HEW

[4370] Anonymous. Historic exhibit of the Waltham Watch Company at the World’s Columbian Exposition, at Chicago. Scientific American, 70(2):25–26, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
25.pdf.

Anonymous:1894:RPIb

[4371] Anonymous. Recently patented inventions. Scientific American, 70(2):27, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
27.pdf.

Anonymous:1894:BNBb

[4372] Anonymous. Books: New books and publications. Scientific American, 70(2):27–28, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n2/pdf/scientificamerican01131894-
27a.pdf.

Anonymous:1894:BPb

[4373] Anonymous. Business and personal. Scientific American, 70(2):28, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-28.pdf.

Anonymous:1894:NQb

[4374] Anonymous. Notes and queries. Scientific American, 70(2):28–29, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-28a.pdf.

Anonymous:1894:Ib

[4375] Anonymous. To inventors. Scientific American, 70(2):29, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-29.pdf.

Anonymous:1894:IIb

[4376] Anonymous. Index of inventions. Scientific American, 70(2):29–30, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-29a.pdf.

Anonymous:1894:PNPb

[4377] Anonymous. Patent notes: Patents. Scientific American, 70(2):31, January 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n2/pdf/scientificamerican01131894-31.pdf.

Anonymous:1894:NOL

[4378] Anonymous. A new optical lantern. Scientific American, 70(3):33, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-33.pdf.

Anonymous:1894:PCR

[4379] Anonymous. The potentialities of chemical research. Scientific American, 70(3):34–35, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
34.pdf.

Anonymous:1894:IBD

[4380] Anonymous. An improved bicycle driving gear. Scientific American, 70(3):36, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
36b.pdf.

Anonymous:1894:IBTa

[4381] Anonymous. An improved building tile. Scientific American, 70 (3):36, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
36.pdf.

Anonymous:1894:STG

[4382] Anonymous. Special triple geared lathe. Scientific American, 70(3):36, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
36a.pdf.

Anonymous:1894:AOV

[4383] Anonymous. An automatically operating valve gear. Scientific American, 70(3):37, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
37a.pdf.

Anonymous:1894:IEA

[4384] Anonymous. An improved exhibiting apparatus. Scientific American, 70(3):37, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
37b.pdf.

Anonymous:1894:IFL

[4385] Anonymous. An improved fire ladder. Scientific American, 70 (3):37, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
37.pdf.

Anonymous:1894:IVH

[4386] Anonymous. Improvement in vessels’ hulls. Scientific American, 70(3):37–38, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
37c.pdf.

Anonymous:1894:OD

[4387] Anonymous. In ocean depths. Scientific American, 70(3):38, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-38.pdf.

Pike:1894:VS

[4388] Nicolas Pike. Venomous spiders. Scientific American, 70(3):39, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-39.pdf.

Anonymous:1894:FKM

[4389] Anonymous. Family knitting machines. Scientific American, 70 (3):40, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
40.pdf.

Anonymous:1894:IAS

[4390] Anonymous. Improved automatic signal apparatus. Scientific American, 70(3):41–42, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
41.pdf.

Anonymous:1894:CGC

[4391] Anonymous. A cable grip for coal cars or mine wagons. Scientific American, 70(3):43, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
43a.pdf.

Anonymous:1894:PCC

[4392] Anonymous. Pine cones, or cocoons? Scientific American, 70 (3):43, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
43.pdf.

Anonymous:1894:RPIc 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
44.pdf.

Anonymous:1894:BNBc

[4394] Anonymous. Books: New books and publications. Scientific American, 70(3):44–45, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n3/pdf/scientificamerican01201894-
44a.pdf.

Anonymous:1894:BPc

[4395] Anonymous. Business and personal. Scientific American, 70(3):45, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-45.pdf.

Anonymous:1894:Ic

[4396] Anonymous. To inventors. Scientific American, 70(3):45, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-45b.pdf.

Anonymous:1894:NQc

[4397] Anonymous. Notes and queries. Scientific American, 70(3):45, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-45a.pdf.

Anonymous:1894:IIc

[4398] Anonymous. Index of inventions. Scientific American, 70(3):45–46, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-45c.pdf.

Anonymous:1894:PNPc

[4399] Anonymous. Patent notes: Patents. Scientific American, 70(3):47, January 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n3/pdf/scientificamerican01201894-47.pdf.

Anonymous:1894:SNL

[4400] Anonymous. Some notable locomotives. Scientific American, 70 (4):49–50, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
49.pdf.

Anonymous:1894:MAS

[4401] Anonymous. The Municipal Art Society. Scientific American, 70 (4):50–51, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
50.pdf.

Anonymous:1894:CRG

[4402] Anonymous. Comparatively recent geological changes in California. Scientific American, 70(4):52, January 27, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n4/pdf/scientificamerican01271894-52.pdf.

Anonymous:1894:LNW

[4403] Anonymous. The late Nathaniel Wheeler. Scientific American, 70(4):52–53, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
52a.pdf.

Anonymous:1894:CBa

[4404] Anonymous. Columbia bicycles for 1894. Scientific American, 70(4):53, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
53.pdf.

Anonymous:1894:ISG

[4405] Anonymous. An improved sickle grinder. Scientific American, 70(4):53, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
53a.pdf.

Anonymous:1894:ECD

[4406] Anonymous. An efficient and continuous drier. Scientific American, 70(4):53–55, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
53b.pdf.

Anonymous:1894:BDFa

[4407] Anonymous. Bicycle drop forgings at the cycle show. Scientific American, 70(4):55, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
55.pdf.

Anonymous:1894:CCC

[4408] Anonymous. Chicago cable cars on Chicago day. Scientific American, 70(4):56, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
56a.pdf.

Anonymous:1894:EPM

[4409] Anonymous. An efficient pulverizing machine. Scientific American, 70(4):56, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
56b.pdf.

Anonymous:1894:IMCa

[4410] Anonymous. Improved method of controlling quicksand. Scientific American, 70(4):56, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
56.pdf.

Anonymous:1894:EWRa

[4411] Anonymous. Electric welding of rails in place. Scientific American, 70(4):57–58, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
57.pdf.

Anonymous:1894:FES

[4412] Anonymous. A fire extinguishing sprinkler. Scientific American, 70(4):59, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
59a.pdf.

Anonymous:1894:WP

[4413] Anonymous. Wool pulling. Scientific American, 70(4):59, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n4/pdf/scientificamerican01271894-59.pdf.

Anonymous:1894:BNBd

[4414] Anonymous. Books: New books and publications. Scientific American, 70(4):60, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
60a.pdf.

Anonymous:1894:RPId

[4415] Anonymous. Recently patented inventions. Scientific American, 70(4):60, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n4/pdf/scientificamerican01271894-
60.pdf.

Anonymous:1894:BPd

[4416] Anonymous. Business and personal. Scientific American, 70(4):61, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n4/pdf/scientificamerican01271894-61.pdf.

Anonymous:1894:Id

[4417] Anonymous. To inventors. Scientific American, 70(4):61, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n4/pdf/scientificamerican01271894-61b.pdf.

Anonymous:1894:NQd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n4/pdf/scientificamerican01271894-61a.pdf.

Anonymous:1894:IId

[4419] Anonymous. Index of inventions. Scientific American, 70(4):61–62, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n4/pdf/scientificamerican01271894-61c.pdf.

Anonymous:1894:PNPd

[4420] Anonymous. Patent notes: Patents. Scientific American, 70(4):63, January 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n4/pdf/scientificamerican01271894-63.pdf.

Anonymous:1894:AMT

[4421] Anonymous. All may have telephones. Scientific American, 70 (5):64, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
65.pdf.

Anonymous:1894:FLO

[4422] Anonymous. Foreigners are learning our industrial methods. Scientific American, 70(5):66, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
66.pdf.

Anonymous:1894:OEE

[4423] Anonymous. The original Edison Electric Railroad. Scientific American, 70(5):66, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
66a.pdf.

Anonymous:1894:FMM

[4424] Anonymous. Fighting Man’s most deadly foes. Scientific American, 70(5):66–67, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
66b.pdf.

Anonymous:1894:IM

[4425] Anonymous. An improved motor. Scientific American, 70(5):68, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n5/pdf/scientificamerican02031894-68c.pdf.

Anonymous:1894:MSSa

[4426] Anonymous. A machine for straightening saws. Scientific American, 70(5):68, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
68a.pdf.

Anonymous:1894:NTa

[4427] Anonymous. A new typewriter. Scientific American, 70(5):68, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n5/pdf/scientificamerican02031894-68b.pdf.

Anonymous:1894:SFSa

[4428] Anonymous. A safety fender for street cars. Scientific American, 70(5):68, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
68.pdf.

Anonymous:1894:ECa

[4429] Anonymous. Electric carriage. Scientific American, 70(5):69, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n5/pdf/scientificamerican02031894-69.pdf.

Anonymous:1894:IVG

[4430] Anonymous. An improved valve gear. Scientific American, 70 (5):69–70, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
69a.pdf.

Anonymous:1894:THV

[4431] Anonymous. A tool holder with a variety of tools. Scientific American, 70(5):70, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
70.pdf.

Anonymous:1894:Cc

[4432] Anonymous. Correspondence. Scientific American, 70(5):71, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n5/pdf/scientificamerican02031894-71.pdf.

Anonymous:1894:ENE

[4433] Anonymous. Edison’s new electrical railway. Scientific American, 70(5):72, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
72.pdf.

Anonymous:1894:RNF

[4434] Anonymous. A rotary newspaper folder. Scientific American, 70(5):72, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
72a.pdf.

Anonymous:1894:EMR

[4435] Anonymous. An electric mountain railroad. Scientific American, 70(5):73–74, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
73.pdf.

Anonymous:1894:SE

[4436] Anonymous. A steamboat explosion. Scientific American, 70 (5):75, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
75.pdf.

Anonymous:1894:BNBe

[4437] Anonymous. Books: New books and publications. Scientific American, 70(5):76, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
76a.pdf.

Anonymous:1894:RPIe

[4438] Anonymous. Recently patented inventions. Scientific American, 70(5):76, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
76.pdf.

Anonymous:1894:BPe

[4439] Anonymous. Business and personal. Scientific American, 70 (5):77, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
77.pdf.

Anonymous:1894:NQe

[4440] Anonymous. Notes and queries. Scientific American, 70(5):77, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n5/pdf/scientificamerican02031894-77a.pdf.

Anonymous:1894:IIe

[4441] Anonymous. Index of inventions. Scientific American, 70(5): 77–78, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
77b.pdf.

Anonymous:1894:PNPe

[4442] Anonymous. Patent notes: Patents. Scientific American, 70 (5):79, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n5/pdf/scientificamerican02031894-
79.pdf.

Anonymous:1894:P

[4443] Anonymous. Patents. Scientific American, 70(5):80, February 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n5/pdf/scientificamerican02031894-80.pdf.

Anonymous:1894:CUO

[4444] Anonymous. A city under one roof — the masonic Temple. Scientific American, 70(6):81–82, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
81.pdf.

Anonymous:1894:HH

[4445] Anonymous. Haulage by horses. Scientific American, 70(6):82– 83, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
82.pdf.

Anonymous:1894:PI

[4446] Anonymous. Progress and invention. Scientific American, 70 (6):83, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
83.pdf.

Anonymous:1894:LWC

[4447] Anonymous. A lubricator and wiper for the commutators of dynamos. Scientific American, 70(6):84, February 10, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n6/pdf/scientificamerican02101894-84b.pdf.

Anonymous:1894:SPPa

[4448] Anonymous. Screw plates for pipe. Scientific American, 70(6): 84, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
84.pdf.

Anonymous:1894:SSS

[4449] Anonymous. A Swedish sled — the Sparkstötting. Scientific American, 70(6):84, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
84a.pdf.

Anonymous:1894:BCC

[4450] Anonymous. Brake for Children’s carriages. Scientific American, 70(6):84–85, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
84c.pdf.

Anonymous:1894:LPI

[4451] Anonymous. A large Philadelphia ice making plant. Scientific American, 70(6):85–86, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
85.pdf.

Anonymous:1894:Cd

[4452] Anonymous. Correspondence. Scientific American, 70(6):87, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n6/pdf/scientificamerican02101894-87.pdf.

Anonymous:1894:APC

[4453] Anonymous. Air pumps of the cruiser New York. Scientific American, 70(6):88, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
88.pdf.

Anonymous:1894:ERM

[4454] Anonymous. Electric railroad mail service in Ottawa, Canada. Scientific American, 70(6):88, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
88a.pdf.

Anonymous:1894:CEF

[4455] Anonymous. The Columbian Exposition — forestry exhibits. Scientific American, 70(6):89, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
89.pdf.

Anonymous:1894:NLS

[4456] Anonymous. A new Lake steamship. Scientific American, 70 (6):90, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
90.pdf.

Anonymous:1894:BNBf

[4457] Anonymous. Books: New books and publications. Scientific American, 70(6):91, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
91a.pdf.

Anonymous:1894:RPIf

[4458] Anonymous. Recently patented inventions. Scientific American, 70(6):91, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
91.pdf.

Anonymous:1894:BPf

[4459] Anonymous. Business and personal. Scientific American, 70(6): 92, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
92.pdf.

Anonymous:1894:NQf

[4460] Anonymous. Notes and queries. Scientific American, 70(6):92– 93, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
92a.pdf.

Anonymous:1894:Ie

[4461] Anonymous. To inventors. Scientific American, 70(6):93, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n6/pdf/scientificamerican02101894-93.pdf.

Anonymous:1894:IIf

[4462] Anonymous. Index of inventions. Scientific American, 70(6):93– 94, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
93a.pdf.

Anonymous:1894:PNPf

[4463] Anonymous. Patent notes: Patents. Scientific American, 70(6): 96, February 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n6/pdf/scientificamerican02101894-
96.pdf.

Anonymous:1894:BBE

[4464] Anonymous. The Boynton Bicycle Electric Railway. Scientific American, 70(7):97, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
97.pdf.

Anonymous:1894:CT

[4465] Anonymous. Cheap telephones. Scientific American, 70(7):98, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n7/pdf/scientificamerican02171894-98.pdf.

Anonymous:1894:NSCa

[4466] Anonymous. A new street car motor. Scientific American, 70 (7):98, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
98a.pdf.

Anonymous:1894:CMEa

[4467] Anonymous. The California Midwinter Exposition in Golden Gate Park, San Francisco. Scientific American, 70(7):98–99, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n7/pdf/scientificamerican02171894-98b.pdf.

Anonymous:1894:SSU

[4468] Anonymous. A steam shovel for use on flat cars. Scientific American, 70(7):100, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
100.pdf.

Anonymous:1894:ASW

[4469] Anonymous. An arch support, water purifier and heater. Scientific American, 70(7):101, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
101a.pdf.

Anonymous:1894:HPW

[4470] Anonymous. A hand punch for work on Cold iron. Scientific American, 70(7):101, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
101.pdf.

Anonymous:1894:MCa

[4471] Anonymous. Manufacture of chicory. Scientific American, 70(7): 101–102, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
101b.pdf.

Anonymous:1894:ADC

[4472] Anonymous. Apparatus for dumping coal cars, etc. Scientific American, 70(7):102, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
102.pdf.

Anonymous:1894:IBF

[4473] Anonymous. An improved button fastener. Scientific American, 70(7):102, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
102a.pdf.

Anonymous:1894:EDL

[4474] Anonymous. An envelope for the display of laces, etc. Scientific American, 70(7):102–103, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
102b.pdf.

Anonymous:1894:FT

[4475] Anonymous. Fulton’s torpedo. Scientific American, 70(7):104– 105, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
104.pdf.

Anonymous:1894:NUSa

[4476] Anonymous. The new United States cruiser Montgomery. Scientific American, 70(7):105, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
105.pdf.

Anonymous:1894:CCB

[4477] Anonymous. The Criminal Court Building of New York. Scientific American, 70(7):105–106, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
105a.pdf.

Anonymous:1894:CECa

[4478] Anonymous. The Columbian Exposition — colonial relics. Scientific American, 70(7):107, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
107.pdf.

Anonymous:1894:J

[4479] Anonymous. The jewfish. Scientific American, 70(7):107, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n7/pdf/scientificamerican02171894-107a.pdf.

Anonymous:1894:BNBg

[4480] Anonymous. Books: New books and publications. Scientific American, 70(7):108, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
108a.pdf.

Anonymous:1894:RPIg

[4481] Anonymous. Recently patented inventions. Scientific American, 70(7):108, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
108.pdf.

Anonymous:1894:BPg

[4482] Anonymous. Business and personal. Scientific American, 70(7): 109, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
109.pdf.

Anonymous:1894:If

[4483] Anonymous. To inventors. Scientific American, 70(7):109, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n7/pdf/scientificamerican02171894-109b.pdf.

Anonymous:1894:NQg

[4484] Anonymous. Notes and queries. Scientific American, 70(7):109, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n7/pdf/scientificamerican02171894-109a.pdf.

Anonymous:1894:IIg

[4485] Anonymous. Index of inventions. Scientific American, 70(7): 109–110, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
109c.pdf.

Anonymous:1894:PNPg

[4486] Anonymous. Patent notes: Patents. Scientific American, 70 (7):111, February 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n7/pdf/scientificamerican02171894-
111.pdf.

Anonymous:1894:HMT

[4487] Anonymous. How to make a telephone call. Scientific American, 70(8):113, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
113.pdf.

Anonymous:1894:FTD

[4488] Anonymous. The fifty thousand dollar reward for street car propulsion. Scientific American, 70(8):114–115, February 24, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n8/pdf/scientificamerican02241894-114.pdf.

Anonymous:1894:AV

[4489] Anonymous. An aerial vessel. Scientific American, 70(8):116, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n8/pdf/scientificamerican02241894-116.pdf.

Anonymous:1894:CPM

[4490] Anonymous. Corrugated plate metallic lathing. Scientific American, 70(8):116, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
116c.pdf.

Anonymous:1894:MLP

[4491] Anonymous. Marvelously low-priced watches. Scientific American, 70(8):116, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
116a.pdf.

Anonymous:1894:SCI

[4492] Anonymous. Supplemental colored ink fountains for printing presses. Scientific American, 70(8):116, February 24, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n8/pdf/scientificamerican02241894-116b.pdf.

Anonymous:1894:CCD

[4493] Anonymous. A current collector for dynamos. Scientific American, 70(8):117, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
117b.pdf.

Anonymous:1894:DAF

[4494] Anonymous. A distilling apparatus for families. Scientific American, 70(8):117, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
117.pdf.

Anonymous:1894:IRV

[4495] Anonymous. An improved relief valve. Scientific American, 70 (8):117, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
117a.pdf.

Anonymous:1894:Ce

[4496] Anonymous. Correspondence. Scientific American, 70(8):119, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n8/pdf/scientificamerican02241894-119.pdf.

Anonymous:1894:WAH

[4497] Anonymous. The Webb Academy and Home for Ship Builders. Scientific American, 70(8):120–121, February 24, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n8/pdf/scientificamerican02241894-120.pdf.

Anonymous:1894:FMS

[4498] Anonymous. The four-masted ship Afghanistan. Scientific American, 70(8):121–122, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
121.pdf.

Hillig:1894:IE

[4499] F. J. Hillig. Interesting experiments. Scientific American, 70(8): 122–123, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
122.pdf.

Anonymous:1894:RPIh

[4500] Anonymous. Recently patented inventions. Scientific American, 70 (8):123–124, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
123.pdf.

Anonymous:1894:BNBh

[4501] Anonymous. Books: New books and publications. Scientific American, 70(8):124, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
124.pdf.

Anonymous:1894:BPh

[4502] Anonymous. Business and personal. Scientific American, 70(8): 124, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
124a.pdf.

Anonymous:1894:NQh

[4503] Anonymous. Notes and queries. Scientific American, 70(8):124– 125, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
124b.pdf.

Anonymous:1894:IIh

[4504] Anonymous. Index of inventions. Scientific American, 70(8): 125–126, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
125.pdf.

Anonymous:1894:PNPh

[4505] Anonymous. Patent notes: Patents. Scientific American, 70 (8):127, February 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n8/pdf/scientificamerican02241894-
127.pdf.

Anonymous:1894:ISP

[4506] Anonymous. Influence of small portions of alloy on the properties of metals. Scientific American, 70(9):130–132, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-130.pdf.

Anonymous:1894:ITE

[4507] Anonymous. Improved triple expansion engines. Scientific American, 70(9):132, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n9/pdf/scientificamerican03031894-
132.pdf.

Anonymous:1894:LDC

[4508] Anonymous. The Lovell diamond cycles. Scientific American, 70(9):133, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-133.pdf.

Anonymous:1894:PL

[4509] Anonymous. The Patterson locomotive. Scientific American, 70(9):133, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-133a.pdf.

Anonymous:1894:TIN

[4510] Anonymous. Tidal indicator, New York Harbor. Scientific American, 70(9):133–134, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n9/pdf/scientificamerican03031894-
133b.pdf.

Anonymous:1894:TAR

[4511] Anonymous. The Third Avenue Railroad Company’s new cable traction plant. Scientific American, 70(9):135, March 3, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n9/pdf/scientificamerican03031894-135.pdf.

Anonymous:1894:GTS

[4512] Anonymous. The Galveston (Texas) steel wagon bridge. Scientific American, 70(9):136, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n9/pdf/scientificamerican03031894-
136.pdf.

Anonymous:1894:IP

[4513] Anonymous. Illusive photography. Scientific American, 70(9):137, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-137.pdf.

Anonymous:1894:IEG

[4514] Anonymous. An improved engine governor. Scientific American, 70(9):138, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n9/pdf/scientificamerican03031894-
138.pdf.

Anonymous:1894:SAH

[4515] Anonymous. Swamp antelopes in the Hamburg Zoological Garden. Scientific American, 70(9):139, March 3, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n9/pdf/scientificamerican03031894-139.pdf.

Anonymous:1894:RPIi

[4516] Anonymous. Recently patented inventions. Scientific American, 70(9):140, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n9/pdf/scientificamerican03031894-
140.pdf.

Anonymous:1894:BPi

[4517] Anonymous. Business and personal. Scientific American, 70(9):141, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-141.pdf.

Anonymous:1894:Ih

[4518] Anonymous. To inventors. Scientific American, 70(9):141, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-141b.pdf.

Anonymous:1894:NQi

[4519] Anonymous. Notes and queries. Scientific American, 70(9):141, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-141a.pdf.

Anonymous:1894:IIi

[4520] Anonymous. Index of inventions. Scientific American, 70(9):141–142, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-141c.pdf.

Anonymous:1894:PNPi

[4521] Anonymous. Patent notes: Patents. Scientific American, 70(9):143, March 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n9/pdf/scientificamerican03031894-143.pdf.

Anonymous:1894:MSC

[4522] Anonymous. The Manchester Ship Canal. Scientific American, 70(10):145, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
145.pdf.

Anonymous:1894:TGE

[4523] Anonymous. The Tiffany Glass Exhibit. Scientific American, 70 (10):146–147, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
146.pdf.

Anonymous:1894:DMA

[4524] Anonymous. A disinfectant mixing apparatus. Scientific American, 70(10):148, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
148b.pdf.

Anonymous:1894:IPK

[4525] Anonymous. An improved pottery kiln. Scientific American, 70(10):148, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-148.pdf.

Anonymous:1894:NTG

[4526] Anonymous. A natural tree graft. Scientific American, 70(10):148, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-148a.pdf.

Anonymous:1894:SSV

[4527] Anonymous. The sanitary soap vase. Scientific American, 70(10):148, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-148c.pdf.

Anonymous:1894:EWL

[4528] Anonymous. Electrical water level indicator. Scientific American, 70(10):149–151, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
149.pdf.

Anonymous:1894:CECb

[4529] Anonymous. The Columbian Exposition — Corean exhibits. Scientific American, 70(10):151, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
151.pdf.

Pike:1894:NDC

[4530] Nicolas Pike. Notes on the domestic cat. Scientific American, 70(10):155, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
155.pdf.

Anonymous:1894:BNBi

[4531] Anonymous. Books: New books and publications. Scientific American, 70(10):156, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
156a.pdf.

Anonymous:1894:BPj

[4532] Anonymous. Business and personal. Scientific American, 70(10):156, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-156b.pdf.

Anonymous:1894:RPIj

[4533] Anonymous. Recently patented inventions. Scientific American, 70(10):156, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n10/pdf/scientificamerican03101894-
156.pdf.

Anonymous:1894:NQj

[4534] Anonymous. Notes and queries. Scientific American, 70(10):156–157, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-156c.pdf.

Anonymous:1894:Ij

[4535] Anonymous. To inventors. Scientific American, 70(10):157, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-157.pdf.

Anonymous:1894:IIj

[4536] Anonymous. Index of inventions. Scientific American, 70(10):157–158, March 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-157a.pdf.

Anonymous:1894:PNPj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n10/pdf/scientificamerican03101894-159.pdf.

Anonymous:1894:ACR

[4538] Anonymous. The Aerial Cable Railway. Scientific American, 70(11):161, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
161.pdf.

Anonymous:1894:BMAa

[4539] Anonymous. Bauxite mining in Alabama. Scientific American, 70(11):162, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
162a.pdf.

Anonymous:1894:DPE

[4540] Anonymous. The difficulties and problems of the electric engineeer. Scientific American, 70(11):162, March 17, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n11/pdf/scientificamerican03171894-162.pdf.

Anonymous:1894:ET

[4541] Anonymous. Experiments with telephones. Scientific American, 70 (11):162–163, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
162b.pdf.

Anonymous:1894:CGM

[4542] Anonymous. Cotton ginning machinery at the Fair. Scientific American, 70(11):164, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
164a.pdf.

Anonymous:1894:DMR

[4543] Anonymous. A ditching machine for railway service. Scientific American, 70(11):164, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
164b.pdf.

Anonymous:1894:PPH

[4544] Anonymous. A post or pole hole boring machine. Scientific American, 70(11):164, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
164.pdf.

Anonymous:1894:ALP

[4545] Anonymous. An air lifting process for wells. Scientific American, 70(11):165, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
165.pdf.

Anonymous:1894:WHG

[4546] Anonymous. A work holder for grinding machines. Scientific American, 70(11):165, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
165a.pdf.

Durfee:1894:SAR

[4547] W. F. Durfee. Some ancient reaction engines. Scientific American, 70(11):165–167, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
165b.pdf.

Anonymous:1894:EGA

[4548] Anonymous. Emil Greiner’s automatic pipette. Scientific American, 70(11):167, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
167a.pdf.

Anonymous:1894:NBS

[4549] Anonymous. The new battle ship Indiana. Scientific American, 70(11):167, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
167.pdf.

Anonymous:1894:DFA

[4550] Anonymous. The Dirigo — the first American steel sailing ship. Scientific American, 70(11):168, March 17, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n11/pdf/scientificamerican03171894-168.pdf.

Anonymous:1894:TPa

[4551] Anonymous. Tricolor photography. Scientific American, 70(11):169, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n11/pdf/scientificamerican03171894-169.pdf.

Anonymous:1894:Cf

[4552] Anonymous. Correspondence. Scientific American, 70(11):170, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n11/pdf/scientificamerican03171894-170.pdf.

Anonymous:1894:AE

[4553] Anonymous. The Antwerp Exposition. Scientific American, 70 (11):170–171, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
170a.pdf.

Anonymous:1894:BNBj

[4554] Anonymous. Books: New books and publications. Scientific American, 70(11):172, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
172a.pdf.

Anonymous:1894:RPIk

[4555] Anonymous. Recently patented inventions. Scientific American, 70(11):172, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n11/pdf/scientificamerican03171894-
172.pdf.

Anonymous:1894:BPk

[4556] Anonymous. Business and personal. Scientific American, 70(11):173, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n11/pdf/scientificamerican03171894-173.pdf.

Anonymous:1894:Ik

[4557] Anonymous. To inventors. Scientific American, 70(11):173, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n11/pdf/scientificamerican03171894-173b.pdf.

Anonymous:1894:NQk

[4558] Anonymous. Notes and queries. Scientific American, 70(11):173, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n11/pdf/scientificamerican03171894-173a.pdf.

Anonymous:1894:IIk

[4559] Anonymous. Index of inventions. Scientific American, 70(11):173–174, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n11/pdf/scientificamerican03171894-173c.pdf.

Anonymous:1894:PNPk

[4560] Anonymous. Patent notes: Patents. Scientific American, 70(11):175, March 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n11/pdf/scientificamerican03171894-175.pdf.

Anonymous:1894:RBS

[4561] Anonymous. Remarkable boat speed-32 1/2 miles per hour. Scientific American, 70(12):178, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
178.pdf.

Anonymous:1894:IOT

[4562] Anonymous. The imperfections of the overhead trolley system. Scientific American, 70(12):178–180, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
178a.pdf.

Anonymous:1894:MAB

[4563] Anonymous. A mirror attachment for bicycles. Scientific American, 70(12):180, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
180b.pdf.

Anonymous:1894:NPA

[4564] Anonymous. New pier for the American line. Scientific American, 70(12):180, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
180a.pdf.

Anonymous:1894:SFP

[4565] Anonymous. To set fire to a pile of snow. Scientific American, 70(12):180, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
180.pdf.

Anonymous:1894:T

[4566] Anonymous. The tachocycle. Scientific American, 70(12):181, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-181.pdf.

Hopkins:1894:HCL

[4567] M. Hopkins. Hints on coloring lantern slides. Scientific American, 70(12):181–182, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
181a.pdf.

Anonymous:1894:Cg

[4568] Anonymous. Correspondence. Scientific American, 70(12):183, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-183.pdf.

Anonymous:1894:WES

[4569] Anonymous. The Waddell–Entz storage battery car traction plant of the Second Avenue Railroad Company, New York. Scientific American, 70(12):184, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
184.pdf.

Anonymous:1894:HSBb

[4570] Anonymous. Handling steel billets by electrical power. Scientific American, 70(12):185–187, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
185.pdf.

Anonymous:1894:BB

[4571] Anonymous. Bicycle boat. Scientific American, 70(12):187, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-187.pdf.

Anonymous:1894:TPb

[4572] Anonymous. Trick photography. Scientific American, 70(12):187, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-187a.pdf.

Anonymous:1894:BNBk

[4573] Anonymous. Books: New books and publications. Scientific American, 70(12):188, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
188a.pdf.

Anonymous:1894:RPIl

[4574] Anonymous. Recently patented inventions. Scientific American, 70(12):188, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n12/pdf/scientificamerican03241894-
188.pdf.

Anonymous:1894:BPl

[4575] Anonymous. Business and personal. Scientific American, 70(12):189, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-189.pdf.

Anonymous:1894:Il

[4576] Anonymous. To inventors. Scientific American, 70(12):189, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-189b.pdf.

Anonymous:1894:NQl

[4577] Anonymous. Notes and queries. Scientific American, 70(12):189, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-189a.pdf.

Anonymous:1894:IIl

[4578] Anonymous. Index of inventions. Scientific American, 70(12):189–190, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-189c.pdf.

Anonymous:1894:PNPl

[4579] Anonymous. Patent notes: Patents. Scientific American, 70(12):191, March 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n12/pdf/scientificamerican03241894-191.pdf.

Anonymous:1894:FSU

[4580] Anonymous. The flow of solids under pressure. Scientific American, 70(13):194, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
194b.pdf.

Anonymous:1894:MRU

[4581] Anonymous. Magazine rifles for the U.S. Navy. Scientific American, 70(13):194, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
194.pdf.

Anonymous:1894:RWS

[4582] Anonymous. Recent work at the Smithsonian Institution. Scientific American, 70(13):194, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
194a.pdf.

Anonymous:1894:BFG

[4583] Anonymous. Biltmore Forest — a great arboretum. Scientific American, 70(13):194–195, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
194c.pdf.

Anonymous:1894:CRS

[4584] Anonymous. The “Climax” ratchet stock. Scientific American, 70(13):196, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
196b.pdf.

Anonymous:1894:CST

[4585] Anonymous. A carpet stretcher and tacker. Scientific American, 70(13):196, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
196.pdf.

Anonymous:1894:RID

[4586] Anonymous. Reid’s improved Danish cream separator. Scientific American, 70(13):196, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
196a.pdf.

Anonymous:1894:FHP

[4587] Anonymous. A feedwater heater and purifier. Scientific American, 70(13):197, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
197b.pdf.

Anonymous:1894:ISI

[4588] Anonymous. An improved sad iron. Scientific American, 70(13):197, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-197.pdf.

Anonymous:1894:NKC

[4589] Anonymous. A new kind of camera — “The Bull’s Eye”. Scientific American, 70(13):197, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
197a.pdf.

Anonymous:1894:CSE

[4590] Anonymous. A cartridge shell extractor. Scientific American, 70(13):198, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
198a.pdf.

Anonymous:1894:LRS

[4591] Anonymous. Large recent Sun spots. Scientific American, 70(13):198, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-198.pdf.

Anonymous:1894:Ch

[4592] Anonymous. Correspondence. Scientific American, 70(13):199, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-199a.pdf.

Anonymous:1894:NSBa

[4593] Anonymous. The new speedway on the banks of the Harlem River, in New York City. Scientific American, 70(13):199, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-199.pdf.

Anonymous:1894:AMF

[4594] Anonymous. American manufacture of fire crackers. Scientific American, 70(13):200, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
200.pdf.

Anonymous:1894:NSCb

[4595] Anonymous. New steam carriage. Scientific American, 70(13):200–201, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-200a.pdf.

Anonymous:1894:ITCa

[4596] Anonymous. Improved 160 ton crane. Scientific American, 70(13):201, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-201.pdf.

Anonymous:1894:SEE

[4597] Anonymous. The Stein Expedition to Ellesmere Land. Scientific American, 70(13):202, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
202.pdf.

Anonymous:1894:STH

[4598] Anonymous. The Siberian tiger in Hagenbeck’s menagerie. Scientific American, 70(13):203, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
203.pdf.

Anonymous:1894:BNBl

[4599] Anonymous. Books: New books and publications. Scientific American, 70(13):204, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
204a.pdf.

Anonymous:1894:RPIm

[4600] Anonymous. Recently patented inventions. Scientific American, 70(13):204, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n13/pdf/scientificamerican03311894-
204.pdf.

Anonymous:1894:BPm

[4601] Anonymous. Business and personal. Scientific American, 70(13):205, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-205.pdf.

Anonymous:1894:In

[4602] Anonymous. To inventors. Scientific American, 70(13):205, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-205b.pdf.

Anonymous:1894:NQm

[4603] Anonymous. Notes and queries. Scientific American, 70(13):205, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-205a.pdf.

Anonymous:1894:IIm

[4604] Anonymous. Index of inventions. Scientific American, 70(13):205–206, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-205c.pdf.

Anonymous:1894:PNPm

[4605] Anonymous. Patent notes: Patents. Scientific American, 70(13):207, March 31, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n13/pdf/scientificamerican03311894-207.pdf.

Anonymous:1894:ARC

[4606] Anonymous. The annual report of the Commissioner of Patents for the year 1893. Scientific American, 70(14):210, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n14/pdf/scientificamerican04071894-210.pdf.

Anonymous:1894:PBT

[4607] Anonymous. Prevention better than cure. Scientific American, 70(14):210–211, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
210a.pdf.

Anonymous:1894:BTC

[4608] Anonymous. The Baltimore Traction Company. Scientific American, 70(14):212, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
212a.pdf.

Anonymous:1894:HGB

[4609] Anonymous. A heating gas burner. Scientific American, 70(14):212, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-212b.pdf.

Anonymous:1894:IW

[4610] Anonymous. An improved wrench. Scientific American, 70(14):212, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-212.pdf.

Anonymous:1894:FBB

[4611] Anonymous. Friedburg’s Bunsen burner. Scientific American, 70 (14):213, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
213.pdf.

Anonymous:1894:RLB

[4612] Anonymous. Reid’s “Lightning” brace. Scientific American, 70(14):213, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-213a.pdf.

Anonymous:1894:STS

[4613] Anonymous. Steam tramcars with Serpollet’s boiler. Scientific American, 70(14):213–214, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
213b.pdf.

Anonymous:1894:IJB

[4614] Anonymous. Improved journal box supports. Scientific American, 70(14):214, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
214.pdf.

Anonymous:1894:Ci

[4615] Anonymous. Carborundum. Scientific American, 70(14):215, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-215.pdf.

Anonymous:1894:PSS

[4616] Anonymous. A practical steam stamp mill. Scientific American, 70(14):216, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
216a.pdf.

Anonymous:1894:SPT

[4617] Anonymous. Substitute for pipe tongs. Scientific American, 70(14):216, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-216.pdf.

Anonymous:1894:VHW

[4618] Anonymous. Very high windmill. Scientific American, 70(14):217, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-217.pdf.

Anonymous:1894:Cj

[4619] Anonymous. Correspondence. Scientific American, 70(14):218, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-218.pdf.

Anonymous:1894:MST

[4620] Anonymous. Manufacture of smoking tobacco. Scientific American, 70(14):219, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
219.pdf.

Anonymous:1894:BNBm

[4621] Anonymous. Books: New books and publications. Scientific American, 70(14):220, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
220a.pdf.

Anonymous:1894:RPIn

[4622] Anonymous. Recently patented inventions. Scientific American, 70(14):220, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n14/pdf/scientificamerican04071894-
220.pdf.

Anonymous:1894:BPn

[4623] Anonymous. Business and personal. Scientific American, 70(14):221, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-221.pdf.

Anonymous:1894:Io

[4624] Anonymous. To inventors. Scientific American, 70(14):221, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-221b.pdf.

Anonymous:1894:NQn

[4625] Anonymous. Notes and queries. Scientific American, 70(14):221, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-221a.pdf.

Anonymous:1894:IIn

[4626] Anonymous. Index of inventions. Scientific American, 70(14):221–222, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-221c.pdf.

Anonymous:1894:PNPn

[4627] Anonymous. Patent notes: Patents. Scientific American, 70(14):223, April 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n14/pdf/scientificamerican04071894-223.pdf.

Anonymous:1894:IGb

[4628] Anonymous. Inventors and the Government. Scientific American, 70(15):226–227, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
226.pdf.

Anonymous:1894:DRPa

[4629] Anonymous. Decision relating to patents. Scientific American, 70(15):227, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
227.pdf.

Anonymous:1894:ABJ

[4630] Anonymous. Aluminum boat — the Jules Davoust. Scientific American, 70(15):228, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
228a.pdf.

Anonymous:1894:MPV

[4631] Anonymous. Means for propelling vessels. Scientific American, 70(15):228, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
228b.pdf.

Anonymous:1894:TFBa

[4632] Anonymous. Terry’s feed bag. Scientific American, 70(15):228, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n15/pdf/scientificamerican04141894-228.pdf.

Wilson:1894:BES

[4633] B. M. Wilson. The Bahama Expedition of the State University of Iowa. Scientific American, 70(15):229–230, April 14, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n15/pdf/scientificamerican04141894-229.pdf.

Anonymous:1894:BCP

[4634] Anonymous. The Bristol Co.’s patent steel belt lacing. Scientific American, 70(15):230–231, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
230.pdf.

Anonymous:1894:CMF

[4635] Anonymous. The California midwinter fair. Scientific American, 70(15):232, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
232.pdf.

Dam:1894:MAS

[4636] H. J. W. Dam. The Maxim air ship. Scientific American, 70 (15):233–235, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
233.pdf.

Anonymous:1894:TST

[4637] Anonymous. A two seated tricycle. Scientific American, 70(15):235, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n15/pdf/scientificamerican04141894-235.pdf.

Anonymous:1894:BNBn

[4638] Anonymous. Books: New books and publications. Scientific American, 70(15):236, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
236a.pdf.

Anonymous:1894:BPo

[4639] Anonymous. Business and personal. Scientific American, 70(15):236, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n15/pdf/scientificamerican04141894-236b.pdf.

Anonymous:1894:RPIo

[4640] Anonymous. Recently patented inventions. Scientific American, 70(15):236, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n15/pdf/scientificamerican04141894-
236.pdf.

Anonymous:1894:NQo

[4641] Anonymous. Notes and queries. Scientific American, 70(15):236–237, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n15/pdf/scientificamerican04141894-236c.pdf.

Anonymous:1894:Iq

[4642] Anonymous. To inventors. Scientific American, 70(15):237, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n15/pdf/scientificamerican04141894-237.pdf.

Anonymous:1894:IIo

[4643] Anonymous. Index of inventions. Scientific American, 70(15):237–238, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n15/pdf/scientificamerican04141894-237a.pdf.

Anonymous:1894:PNPo

[4644] Anonymous. Patent notes: Patents. Scientific American, 70(15):239, April 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n15/pdf/scientificamerican04141894-239.pdf.

Anonymous:1894:HLN

[4645] Anonymous. The Hillside Line of the North Hudson County Railway Company. Scientific American, 70(16):241–242, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n16/pdf/scientificamerican04211894-241.pdf.

Anonymous:1894:UWE

[4646] Anonymous. The utilization of the waste energy of the world. Scientific American, 70(16):242–243, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n16/pdf/scientificamerican04211894-
242.pdf.

Anonymous:1894:LPS

[4647] Anonymous. A low priced stem winder. Scientific American, 70(16):244, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-244a.pdf.

Anonymous:1894:VDG

[4648] Anonymous. A variable driving gear for bicycles. Scientific American, 70(16):244, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n16/pdf/scientificamerican04211894-
244.pdf.

Anonymous:1894:HSS

[4649] Anonymous. House of Senor Saracibar, architect, Salamanca, Spain. Scientific American, 70(16):244–245, April 21, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n16/pdf/scientificamerican04211894-244b.pdf.

Anonymous:1894:WNW

[4650] Anonymous. Whitaker’s new weight motor. Scientific American, 70(16):245, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n16/pdf/scientificamerican04211894-
245.pdf.

Anonymous:1894:UCE

[4651] Anonymous. Underground conductors for electric street railways. Scientific American, 70(16):245–246, April 21, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n16/pdf/scientificamerican04211894-245a.pdf.

Anonymous:1894:Ck

[4652] Anonymous. Correspondence. Scientific American, 70(16):247–248, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-247.pdf.

Anonymous:1894:MF

[4653] Anonymous. The Midwinter Fair. Scientific American, 70(16):248–249, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-248.pdf.

Anonymous:1894:IHP

[4654] Anonymous. Improved high pressure air compressors. Scientific American, 70(16):249–251, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n16/pdf/scientificamerican04211894-
249.pdf.

Anonymous:1894:CECc

[4655] Anonymous. Carli’s electric carriage. Scientific American, 70(16):251, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-251a.pdf.

Anonymous:1894:EHa

[4656] Anonymous. Eolian harps. Scientific American, 70(16):251, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-251.pdf.

Anonymous:1894:IIp

[4657] Anonymous. Index of inventions. Scientific American, 70(16):251–253, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-253.pdf.

Anonymous:1894:BNBo

[4658] Anonymous. Books: New books and publications. Scientific American, 70(16):252, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n16/pdf/scientificamerican04211894-
252a.pdf.

Anonymous:1894:BPp

[4659] Anonymous. Business and personal. Scientific American, 70(16):252, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-252b.pdf.

Anonymous:1894:RPIp

[4660] Anonymous. Recently patented inventions. Scientific American, 70(16):252, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n16/pdf/scientificamerican04211894-
252.pdf.

Anonymous:1894:NQp

[4661] Anonymous. Notes and queries. Scientific American, 70(16):252–253, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-252c.pdf.

Anonymous:1894:PNPp

[4662] Anonymous. Patent notes: Patents. Scientific American, 70(16):255, April 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n16/pdf/scientificamerican04211894-255.pdf.

Anonymous:1894:CMa

[4663] Anonymous. Contagion in milk. Scientific American, 70(17):258–259, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-258.pdf.

Anonymous:1894:GAS

[4664] Anonymous. A grader and amalgamator for separating precious metals. Scientific American, 70(17):260, April 28, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n17/pdf/scientificamerican04281894-260.pdf.

Anonymous:1894:ICFa

[4665] Anonymous. An improved car fender. Scientific American, 70(17):260, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-260b.pdf.

Anonymous:1894:MI

[4666] Anonymous. A Mediæval ironclad. Scientific American, 70(17):260, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-260a.pdf.

Anonymous:1894:ABS

[4667] Anonymous. An air brake stop cock. Scientific American, 70 (17):260–261, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
260c.pdf.

Anonymous:1894:TCF

[4668] Anonymous. Thomas’ car fender. Scientific American, 70(17):261, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-261.pdf.

Anonymous:1894:FRS

[4669] Anonymous. A funeral on the River Spree. Scientific American, 70(17):261–262, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
261a.pdf.

Anonymous:1894:BID

[4670] Anonymous. Burnham’s 13 inch drill press. Scientific American, 70(17):262, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
262a.pdf.

Anonymous:1894:ITCb

[4671] Anonymous. Improved tea and coffee strainer. Scientific American, 70(17):262, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
262b.pdf.

Anonymous:1894:MSE

[4672] Anonymous. A metallic seal for envelopes, etc. Scientific American, 70(17):262, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
262.pdf.

Anonymous:1894:PAIa

[4673] Anonymous. The Park Avenue improvement in New York City. Scientific American, 70(17):263, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
263.pdf.

Anonymous:1894:G

[4674] Anonymous. The Glyptodont. Scientific American, 70(17):264–265, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-264.pdf.

Anonymous:1894:NCW

[4675] Anonymous. Notable and curious watches. Scientific American, 70(17):265–267, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
265.pdf.

Anonymous:1894:BNBp

[4676] Anonymous. Books: New books and publications. Scientific American, 70(17):268, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
268a.pdf.

Anonymous:1894:RPIq

[4677] Anonymous. Recently patented inventions. Scientific American, 70(17):268, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n17/pdf/scientificamerican04281894-
268.pdf.

Anonymous:1894:BPq

[4678] Anonymous. Business and personal. Scientific American, 70(17):269, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-269.pdf.

Anonymous:1894:Ir

[4679] Anonymous. To inventors. Scientific American, 70(17):269, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-269b.pdf.

Anonymous:1894:NQq

[4680] Anonymous. Notes and queries. Scientific American, 70(17):269, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-269a.pdf.

Anonymous:1894:IIq

[4681] Anonymous. Index of inventions. Scientific American, 70(17):269–270, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-269c.pdf.

Anonymous:1894:PNPq

[4682] Anonymous. Patent notes: Patents. Scientific American, 70(17):271, April 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n17/pdf/scientificamerican04281894-271.pdf.

Anonymous:1894:NTL

[4683] Anonymous. A new theory of light sensation. Scientific American, 70(18):274, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
274.pdf.

Anonymous:1894:ICAa

[4684] Anonymous. Italian colonization in Africa. Scientific American, 70(18):274–275, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
274a.pdf.

Anonymous:1894:IED

[4685] Anonymous. An incline elevator and dump. Scientific American, 70(18):276, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
276.pdf.

Anonymous:1894:PBL

[4686] Anonymous. Polishing and buffing lathe. Scientific American, 70(18):276, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
276a.pdf.

Anonymous:1894:ELS

[4687] Anonymous. An extension ladder for sleeping cars. Scientific American, 70(18):276–277, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
276b.pdf.

Anonymous:1894:ICCa

[4688] Anonymous. An improved car coupling. Scientific American, 70 (18):277, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
277.pdf.

Anonymous:1894:CP

[4689] Anonymous. Cigarette paper. Scientific American, 70(18):277–278, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-277a.pdf.

Anonymous:1894:Cl

[4690] Anonymous. Correspondence. Scientific American, 70(18):278, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-278.pdf.

Anonymous:1894:CKA

[4691] Anonymous. The Chase–Kirchner aerodromic railroad. Scientific American, 70(18):279, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
279.pdf.

Anonymous:1894:MRC

[4692] Anonymous. Manufacture of rock candy. Scientific American, 70(18):280, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
280.pdf.

Anonymous:1894:CMb

[4693] Anonymous. A Carpenter’s makeshift. Scientific American, 70(18):281, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-281.pdf.

Anonymous:1894:LPM

[4694] Anonymous. A large punching machine. Scientific American, 70 (18):281, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
281a.pdf.

Anonymous:1894:MAA

[4695] Anonymous. A method of and apparatus for amalgamating ores. Scientific American, 70(18):281–282, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
281b.pdf.

Anonymous:1894:NER

[4696] Anonymous. Novel electric rock drill. Scientific American, 70(18):282, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-282.pdf.

Anonymous:1894:WFB

[4697] Anonymous. Why flowers are beautiful. Scientific American, 70 (18):283, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
283.pdf.

Anonymous:1894:BNBq

[4698] Anonymous. Books: New books and publications. Scientific American, 70(18):284, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
284a.pdf.

Anonymous:1894:BPr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-284b.pdf.

Anonymous:1894:RPIr

[4700] Anonymous. Recently patented inventions. Scientific American, 70(18):284, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n18/pdf/scientificamerican05051894-
284.pdf.

Anonymous:1894:NQr

[4701] Anonymous. Notes and queries. Scientific American, 70(18):284–285, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-284c.pdf.

Anonymous:1894:Is

[4702] Anonymous. To inventors. Scientific American, 70(18):285, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-285.pdf.

Anonymous:1894:IIr

[4703] Anonymous. Index of inventions. Scientific American, 70(18):285–286, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-285a.pdf.

Anonymous:1894:PNPr

[4704] Anonymous. Patent notes: Patents. Scientific American, 70(18):287, May 5, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n18/pdf/scientificamerican05051894-287.pdf.

Anonymous:1894:PAIb

[4705] Anonymous. The Park Avenue improvement in New York City — the temporary Harlem River Bridge and the moving of the Motthaven Station. Scientific American, 70(19):289, May 12, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n19/pdf/scientificamerican05121894-289.pdf.

Anonymous:1894:RP

[4706] Anonymous. The rich and the poor. Scientific American, 70 (19):290–291, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n19/pdf/scientificamerican05121894-
290.pdf.

Anonymous:1894:EHG

[4707] Anonymous. An efficient heating gas burner. Scientific American, 70(19):292, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n19/pdf/scientificamerican05121894-
292b.pdf.

Anonymous:1894:ISA

[4708] Anonymous. An improved steam actuated valve. Scientific American, 70(19):292, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n19/pdf/scientificamerican05121894-
292.pdf.

Anonymous:1894:WSF

[4709] Anonymous. A widow sash fastener. Scientific American, 70(19):292, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-292a.pdf.

Anonymous:1894:HW

[4710] Anonymous. A high windmill. Scientific American, 70(19):292–293, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-292c.pdf.

Anonymous:1894:BUR

[4711] Anonymous. The bicycle upon rails. Scientific American, 70(19):293, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-293.pdf.

Anonymous:1894:TB

[4712] Anonymous. Theodor Billroth. Scientific American, 70(19):293–294, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-293a.pdf.

Anonymous:1894:DETa

[4713] Anonymous. A device enabling the typewriter to see and read work as it is written. Scientific American, 70(19):294, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-294.pdf.

Anonymous:1894:Cn

[4714] Anonymous. Correspondence. Scientific American, 70(19):295, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-295.pdf.

Beach:1894:CBA

[4715] A. E. Beach. The centrifugal bowling alley. Scientific American, 70(19):297–298, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n19/pdf/scientificamerican05121894-
297.pdf.

Anonymous:1894:IBTb

[4716] Anonymous. An improved boiler tube expander. Scientific American, 70(19):298, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n19/pdf/scientificamerican05121894-
298.pdf.

Anonymous:1894:RLS

[4717] Anonymous. Righthandedness and lefthandedness of sight. Scientific American, 70(19):298–299, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n19/pdf/scientificamerican05121894-
298a.pdf.

Anonymous:1894:BPs

[4718] Anonymous. Business and personal. Scientific American, 70(19):300, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-300a.pdf.

Anonymous:1894:RPIs

[4719] Anonymous. Recently patented inventions. Scientific American, 70(19):300, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n19/pdf/scientificamerican05121894-
300.pdf.

Anonymous:1894:NQs

[4720] Anonymous. Notes and queries. Scientific American, 70(19):300–301, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-300b.pdf.

Anonymous:1894:It

[4721] Anonymous. To inventors. Scientific American, 70(19):301, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-301.pdf.

Anonymous:1894:IIs

[4722] Anonymous. Index of inventions. Scientific American, 70(19):301–302, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-301a.pdf.

Anonymous:1894:PNPs

[4723] Anonymous. Patent notes: Patents. Scientific American, 70(19):303, May 12, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n19/pdf/scientificamerican05121894-303.pdf.

Anonymous:1894:LUC

[4724] Anonymous. The Lawrence underground conduit electric trolley railroad. Scientific American, 70(20):305, May 19, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n20/pdf/scientificamerican05191894-305.pdf.

Anonymous:1894:AH

[4725] Anonymous. An astronomer honored. Scientific American, 70(20):306, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-306.pdf.

Anonymous:1894:DIE

[4726] Anonymous. Dangers incident to electric current distribution. Scientific American, 70(20):306, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
306a.pdf.

Anonymous:1894:HQE

[4727] Anonymous. “The harvest of a quiet eye”. Scientific American, 70(20):306–307, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
306b.pdf.

Anonymous:1894:DRPb

[4728] Anonymous. Decisions relating to patents. Scientific American, 70(20):307, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
307.pdf.

Anonymous:1894:AWC

[4729] Anonymous. An automatic word counter for typewriting machines. Scientific American, 70(20):308, May 19, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n20/pdf/scientificamerican05191894-308b.pdf.

Anonymous:1894:IMM

[4730] Anonymous. An improved mailing machine. Scientific American, 70(20):308, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
308.pdf.

Anonymous:1894:WC

[4731] Anonymous. The Westcott chucks. Scientific American, 70(20):308, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-308a.pdf.

Anonymous:1894:PBM

[4732] Anonymous. Paper box making. Scientific American, 70(20):309–310, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-309.pdf.

Anonymous:1894:PP

[4733] Anonymous. A puzzle purse. Scientific American, 70(20):310, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-310.pdf.

Anonymous:1894:ECb

[4734] Anonymous. Electric capstans. Scientific American, 70(20):312, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-312.pdf.

Anonymous:1894:ICb

[4735] Anonymous. An improved cultivator. Scientific American, 70(20):313, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-313.pdf.

Anonymous:1894:TBDa

[4736] Anonymous. The torpedo boat destroyer hornet. Scientific American, 70(20):313–314, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
313a.pdf.

Anonymous:1894:BEL

[4737] Anonymous. Bicycle exercises at Leipzig. Scientific American, 70 (20):315, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
315.pdf.

Anonymous:1894:BNBr

[4738] Anonymous. Books: New books and publications. Scientific American, 70(20):316, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
316a.pdf.

Anonymous:1894:RPIt

[4739] Anonymous. Recently patented inventions. Scientific American, 70(20):316, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n20/pdf/scientificamerican05191894-
316.pdf.

Anonymous:1894:BPt

[4740] Anonymous. Business and personal. Scientific American, 70(20):317, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-317.pdf.

Anonymous:1894:Iv

[4741] Anonymous. To inventors. Scientific American, 70(20):317, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-317b.pdf.

Anonymous:1894:NQt

[4742] Anonymous. Notes and queries. Scientific American, 70(20):317, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-317a.pdf.

Anonymous:1894:IIt

[4743] Anonymous. Index of inventions. Scientific American, 70(20):317–318, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-317c.pdf.

Anonymous:1894:PNPt

[4744] Anonymous. Patent notes: Patents. Scientific American, 70(20):319, May 19, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n20/pdf/scientificamerican05191894-319.pdf.

Anonymous:1894:OHF

[4745] Anonymous. The one hundred and fifty ton portable gun scales of the Watervliet Arsenal. Scientific American, 70(21):321–322, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-321.pdf.

Anonymous:1894:EAM

[4746] Anonymous. Eleventh annual meeting of the institute of electrical engineers. Scientific American, 70(21):322, May 26, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n21/pdf/scientificamerican05261894-322.pdf.

Anonymous:1894:BEI

[4747] Anonymous. The bicycle as an example for improvements in transportation. Scientific American, 70(21):322–323, May 26, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n21/pdf/scientificamerican05261894-322a.pdf.

Anonymous:1894:AOE

[4748] Anonymous. An automatically operating ejector. Scientific American, 70(21):324, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
324a.pdf.

Anonymous:1894:IOC

[4749] Anonymous. An improved Ore crusher. Scientific American, 70(21):324, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-324.pdf.

Anonymous:1894:SECa

[4750] Anonymous. A simple and easily cleaned filter. Scientific American, 70(21):324, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
324b.pdf.

Anonymous:1894:CLC

[4751] Anonymous. A convenient line chalker. Scientific American, 70 (21):324–325, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
324c.pdf.

Anonymous:1894:PWI

[4752] Anonymous. Picolet’s Wimshurst induction machine without sectors. Scientific American, 70(21):325, May 26, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n21/pdf/scientificamerican05261894-325.pdf.

Anonymous:1894:EIM

[4753] Anonymous. Electrostatic induction machines without sectors. Scientific American, 70(21):325–326, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
325a.pdf.

Anonymous:1894:Co

[4754] Anonymous. Correspondence. Scientific American, 70(21):327–328, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-327.pdf.

Anonymous:1894:NLB

[4755] Anonymous. New lift bridge, Chicago. Scientific American, 70(21):328, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-328.pdf.

Anonymous:1894:MSD

[4756] Anonymous. Marine steam dynamo. Scientific American, 70(21):329, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-329a.pdf.

Anonymous:1894:SCT

[4757] Anonymous. A simple and convenient telephone. Scientific American, 70(21):329, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
329.pdf.

Anonymous:1894:PPA

[4758] Anonymous. Protection of piles against the teredo. Scientific American, 70(21):330, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
330.pdf.

Anonymous:1894:SAP

[4759] Anonymous. Simplified apparatus for photomicrography. Scientific American, 70(21):331–332, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
331.pdf.

Anonymous:1894:RPIu

[4760] Anonymous. Recently patented inventions. Scientific American, 70(21):332, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
332.pdf.

Anonymous:1894:BNBs

[4761] Anonymous. Books: New books and publications. Scientific American, 70(21):332–333, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n21/pdf/scientificamerican05261894-
332a.pdf.

Anonymous:1894:BPu

[4762] Anonymous. Business and personal. Scientific American, 70(21):333, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-333.pdf.

Anonymous:1894:Ix

[4763] Anonymous. To inventors. Scientific American, 70(21):333, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-333b.pdf.

Anonymous:1894:NQu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-333a.pdf.

Anonymous:1894:IIu

[4765] Anonymous. Index of inventions. Scientific American, 70(21):333–334, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-333c.pdf.

Anonymous:1894:PNPu

[4766] Anonymous. Patent notes: Patents. Scientific American, 70(21):335, May 26, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n21/pdf/scientificamerican05261894-335.pdf.

Anonymous:1894:NZL

[4767] Anonymous. New Zealand labor laws. Scientific American, 70(22):338, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-338a.pdf.

Anonymous:1894:PIL

[4768] Anonymous. Pupin’s improvements in long distance telephoning. Scientific American, 70(22):338, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
338.pdf.

Anonymous:1894:PTC

[4769] Anonymous. Physical training in college. Scientific American, 70(22):338–339, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
338b.pdf.

Anonymous:1894:HC

[4770] Anonymous. The Hilliard cyclometer. Scientific American, 70(22):340, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-340a.pdf.

Anonymous:1894:ICT

[4771] Anonymous. An ingenious combination tool. Scientific American, 70(22):340, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
340.pdf.

Anonymous:1894:QSS

[4772] Anonymous. A quick speed steam engine. Scientific American, 70(22):340, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
340c.pdf.

Anonymous:1894:SAB

[4773] Anonymous. The Seventh Avenue Bridge over the Harlem River, New York City. Scientific American, 70(22):340, June 2, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n22/pdf/scientificamerican06021894-340b.pdf.

Anonymous:1894:FWL

[4774] Anonymous. Feeding and watering live stock. Scientific American, 70(22):341, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
341.pdf.

Anonymous:1894:LGW

[4775] Anonymous. The London gigantic wheel. Scientific American, 70(22):341, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
341a.pdf.

Anonymous:1894:Aa

[4776] Anonymous. Anaglyphs. Scientific American, 70(22):341–342, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-341b.pdf.

Anonymous:1894:Cq

[4777] Anonymous. Correspondence. Scientific American, 70(22):342, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-342.pdf.

Anonymous:1894:NHB

[4778] Anonymous. Notes on the history of the breech-loading gun. Scientific American, 70(22):343, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
343.pdf.

Anonymous:1894:PHH

[4779] Anonymous. Photographing a human head upon a table. Scientific American, 70(22):344, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
344.pdf.

Anonymous:1894:TBL

[4780] Anonymous. Triple boiler locomotive, Belgian State Railway. Scientific American, 70(22):344–345, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
344a.pdf.

Anonymous:1894:OOL

[4781] Anonymous. The orang-outang in the Leipsig Zoological Garden. Scientific American, 70(22):345–346, June 2, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n22/pdf/scientificamerican06021894-345.pdf.

Anonymous:1894:PTR

[4782] Anonymous. Proposed tubular railway under the English channel. Scientific American, 70(22):347, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
347.pdf.

Anonymous:1894:BNBt

[4783] Anonymous. Books: New books and publications. Scientific American, 70(22):348, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
348a.pdf.

Anonymous:1894:RPIv

[4784] Anonymous. Recently patented inventions. Scientific American, 70(22):348, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n22/pdf/scientificamerican06021894-
348.pdf.

Anonymous:1894:BPv

[4785] Anonymous. Business and personal. Scientific American, 70(22):349, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-349.pdf.

Anonymous:1894:Iy

[4786] Anonymous. To inventors. Scientific American, 70(22):349, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-349b.pdf.

Anonymous:1894:NQv

[4787] Anonymous. Notes and queries. Scientific American, 70(22):349, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-349a.pdf.

Anonymous:1894:IIv

[4788] Anonymous. Index of inventions. Scientific American, 70(22):349–350, June 2, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-349c.pdf.

Anonymous:1894:PNPv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n22/pdf/scientificamerican06021894-351.pdf.

Anonymous:1894:CSL

[4790] Anonymous. The coal strike and its lessons. Scientific American, 70(23):354–355, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
354.pdf.

Anonymous:1894:ITCc

[4791] Anonymous. An improved trolley conductor. Scientific American, 70(23):356, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
356b.pdf.

Anonymous:1894:ITMa

[4792] Anonymous. An improved thrashing machine feed. Scientific American, 70(23):356, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
356.pdf.

Anonymous:1894:NMG

[4793] Anonymous. New multipolar generator. Scientific American, 70 (23):356, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
356a.pdf.

Anonymous:1894:SFSb

[4794] Anonymous. A simple fruit Stoner. Scientific American, 70(23):357, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-357a.pdf.

Anonymous:1894:SIT

[4795] Anonymous. Some improved tools. Scientific American, 70(23):357, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-357.pdf.

Anonymous:1894:OGE

[4796] Anonymous. The Olds gasoline engine. Scientific American, 70 (23):357–358, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
357b.pdf.

Anonymous:1894:BAB

[4797] Anonymous. The Bailey automatic bicycle brake. Scientific American, 70(23):358, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
358a.pdf.

Anonymous:1894:CLD

[4798] Anonymous. Chain links drop forged from bar steel. Scientific American, 70(23):358, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
358.pdf.

Anonymous:1894:CHD

[4799] Anonymous. The coal hoisting and distributing plant of the Manhattan elevated railroad, of New York City. Scientific American, 70(23):359, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-359.pdf.

Anonymous:1894:CTG

[4800] Anonymous. A curious tree growth. Scientific American, 70(23):360, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-360.pdf.

Anonymous:1894:RMB

[4801] Anonymous. The Rudolf Muller boiler furnace on the steamship Grimm. Scientific American, 70(23):360, June 9, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n23/pdf/scientificamerican06091894-360a.pdf.

Anonymous:1894:FFMa

[4802] Anonymous. Flower and fruit of the Monstera Deliciosa. Scientific American, 70(23):361–362, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
361.pdf.

Anonymous:1894:GGE

[4803] Anonymous. The gas and gasoline engines of the Globe Gas Engine Company, of Philadelphia, Pa. Scientific American, 70(23):363, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-363.pdf.

Anonymous:1894:BNBu

[4804] Anonymous. Books: New books and publications. Scientific American, 70(23):364, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
364a.pdf.

Anonymous:1894:RPIw

[4805] Anonymous. Recently patented inventions. Scientific American, 70(23):364, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n23/pdf/scientificamerican06091894-
364.pdf.

Anonymous:1894:BPw

[4806] Anonymous. Business and personal. Scientific American, 70(23):365, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-365.pdf.

Anonymous:1894:Iz

[4807] Anonymous. To inventors. Scientific American, 70(23):365, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-365b.pdf.

Anonymous:1894:IIw

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-365c.pdf.

Anonymous:1894:NQw

[4809] Anonymous. Notes and queries. Scientific American, 70(23):365, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-365a.pdf.

Anonymous:1894:PNPw

[4810] Anonymous. Patent notes: Patents. Scientific American, 70(23):367, June 9, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n23/pdf/scientificamerican06091894-367.pdf.

Anonymous:1894:FE

[4811] Anonymous. Free electricity. Scientific American, 70(24):370, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-370c.pdf.

Anonymous:1894:HPE

[4812] Anonymous. How to prevent and evade insanity. Scientific American, 70(24):370, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
370a.pdf.

Anonymous:1894:LV

[4813] Anonymous. The limitation of vivisection. Scientific American, 70(24):370, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
370d.pdf.

Anonymous:1894:OFM

[4814] Anonymous. Opening of the Field Museum. Scientific American, 70(24):370, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
370.pdf.

Anonymous:1894:PRT

[4815] Anonymous. Passenger railroad traffic in New York and London. Scientific American, 70(24):370, June 16, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n24/pdf/scientificamerican06161894-370b.pdf.

Anonymous:1894:EWM

[4816] Anonymous. The Engineer’s work in modern buildings. Scientific American, 70(24):370–371, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
370e.pdf.

Anonymous:1894:DRA

[4817] Anonymous. A decision relating to assignments. Scientific American, 70(24):371, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
371.pdf.

Anonymous:1894:AMa

[4818] Anonymous. An automatic motor. Scientific American, 70(24):372, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-372.pdf.

Anonymous:1894:IBC

[4819] Anonymous. An improved bicycle Crank. Scientific American, 70(24):372, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
372b.pdf.

Anonymous:1894:IFG

[4820] Anonymous. An improved farm gate. Scientific American, 70(24):372, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-372c.pdf.

Anonymous:1894:IUb

[4821] Anonymous. An improved unicycle. Scientific American, 70(24):372, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-372a.pdf.

Anonymous:1894:FBC

[4822] Anonymous. The fireproof building construction of the New Jersey Wire Cloth Company. Scientific American, 70(24):373–374, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-373.pdf.

Anonymous:1894:ODV

[4823] Anonymous. An oscillating device for vehicles. Scientific American, 70(24):374, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
374a.pdf.

Anonymous:1894:TC

[4824] Anonymous. A toy cart. Scientific American, 70(24):374, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-374.pdf.

Anonymous:1894:HRBa

[4825] Anonymous. The Hudson River Bridge of the New York and New Jersey Bridge Company. Scientific American, 70(24):375, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-375.pdf.

Anonymous:1894:NCL

[4826] Anonymous. A new compound locomotive. Scientific American, 70(24):376, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
376.pdf.

Anonymous:1894:OIE

[4827] Anonymous. Opening of the International Exhibition at Antwerp. Scientific American, 70(24):377, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
377.pdf.

Anonymous:1894:Cr

[4828] Anonymous. Correspondence. Scientific American, 70(24):378, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-378.pdf.

Riley:1894:PCA

[4829] C. V. Riley. The periodical cicada, alias seventeen-year locust. Scientific American, 70(24):378–379, June 16, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n24/pdf/scientificamerican06161894-378a.pdf.

Anonymous:1894:HSa

[4830] Anonymous. The haunted swing. Scientific American, 70(24):379, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-379.pdf.

Anonymous:1894:BNBv

[4831] Anonymous. Books: New books and publications. Scientific American, 70(24):380, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
380a.pdf.

Anonymous:1894:BPx

[4832] Anonymous. Business and personal. Scientific American, 70(24):380, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-380b.pdf.

Anonymous:1894:RPIx

[4833] Anonymous. Recently patented inventions. Scientific American, 70(24):380, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n24/pdf/scientificamerican06161894-
380.pdf.

Anonymous:1894:NQx

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-380c.pdf.

Anonymous:1894:Iba

[4835] Anonymous. To inventors. Scientific American, 70(24):381, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-381.pdf.

Anonymous:1894:IIx

[4836] Anonymous. Index of inventions. Scientific American, 70(24):381–382, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-381a.pdf.

Anonymous:1894:PNPx

[4837] Anonymous. Patent notes: Patents. Scientific American, 70(24):383, June 16, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n24/pdf/scientificamerican06161894-383.pdf.

Anonymous:1894:CMEb

[4838] Anonymous. The California midwinter exhibition. Scientific American, 70(25):385, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
385.pdf.

Anonymous:1894:IAa

[4839] Anonymous. The Island of Atlantis. Scientific American, 70(25):386, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n25/pdf/scientificamerican06231894-386a.pdf.

Anonymous:1894:PRI

[4840] Anonymous. Proposed railway from India to Ceylon. Scientific American, 70(25):386, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
386.pdf.

Anonymous:1894:RTN

[4841] Anonymous. Recent trials of Naval armor. Scientific American, 70(25):386–387, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
386b.pdf.

Anonymous:1894:BFa

[4842] Anonymous. The Berkefeld filter. Scientific American, 70(25):388, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n25/pdf/scientificamerican06231894-388a.pdf.

Anonymous:1894:EMTa

[4843] Anonymous. An efficient magneto telephone. Scientific American, 70(25):388, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
388.pdf.

Anonymous:1894:ITMb

[4844] Anonymous. An improved tuck marker. Scientific American, 70 (25):388, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
388b.pdf.

Anonymous:1894:NAM

[4845] Anonymous. A new article of manufacture. Scientific American, 70(25):388, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
388c.pdf.

Anonymous:1894:SSW

[4846] Anonymous. The Simonds steam wagon. Scientific American, 70 (25):389, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
389.pdf.

Anonymous:1894:PWM

[4847] Anonymous. The patrol wagons of the Metropolitan Traction Company. Scientific American, 70(25):389–390, June 23, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v70/
n25/pdf/scientificamerican06231894-389a.pdf.

Anonymous:1894:Cs

[4848] Anonymous. Correspondence. Scientific American, 70(25):391, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n25/pdf/scientificamerican06231894-391.pdf.

Anonymous:1894:FFMb

[4849] Anonymous. Flower and feather manufacturing. Scientific American, 70(25):392, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
392.pdf.

Hopkins:1894:LSI

[4850] M. Hopkins. Lantern slide illustrating sound waves. Scientific American, 70(25):395, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
395.pdf.

Anonymous:1894:BPy

[4851] Anonymous. Business and personal. Scientific American, 70(25):396, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n25/pdf/scientificamerican06231894-396a.pdf.

Anonymous:1894:RPIy

[4852] Anonymous. Recently patented inventions. Scientific American, 70(25):396, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v70/n25/pdf/scientificamerican06231894-
396.pdf.

Anonymous:1894:NQy

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n25/pdf/scientificamerican06231894-396b.pdf.

Anonymous:1894:Ibb

[4854] Anonymous. To inventors. Scientific American, 70(25):397, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n25/pdf/scientificamerican06231894-397.pdf.

Anonymous:1894:IIy

[4855] Anonymous. Index of inventions. Scientific American, 70(25):397–398, June 23, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v70/n25/pdf/scientificamerican06231894-397a.pdf.

Anonymous:1894:SWS

[4856] Anonymous. The stability of war ships. Scientific American, 71(1):2–3, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-2.pdf.

Anonymous:1894:ASC

[4857] Anonymous. An automatic self-cleansing water filter. Scientific American, 71(1):4, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
4a.pdf.

Anonymous:1894:ICd

[4858] Anonymous. An improved caliper. Scientific American, 71(1):4, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-4b.pdf.

Anonymous:1894:IRJ

[4859] Anonymous. An improved rail joint. Scientific American, 71(1):4, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-4c.pdf.

Anonymous:1894:TWH

[4860] Anonymous. A traction wheel for harvesters, traction engines, etc. Scientific American, 71(1):4, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
4.pdf.

Anonymous:1894:DSM

[4861] Anonymous. A dwelling at Springfield, Mass. Scientific American, 71(1):5, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
5.pdf.

Anonymous:1894:IFT

[4862] Anonymous. An indicator to facilitate target practice. Scientific American, 71(1):5–6, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
5a.pdf.

Durfee:1894:AA

[4863] W. F. Durfee. An ancient automaton. Scientific American, 71(1):6, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-6.pdf.

Anonymous:1894:LGM

[4864] Anonymous. The lengthening of the gunboats Machias and Castine. Scientific American, 71(1):7, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
7.pdf.

Anonymous:1894:NSBb

[4865] Anonymous. New steel belt fastener. Scientific American, 71(1):7, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-7a.pdf.

Anonymous:1894:GOB

[4866] Anonymous. Great office buildings of New York. Scientific American, 71(1):8–9, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
8.pdf.

Anonymous:1894:TWA

[4867] Anonymous. The termites or white ants of Africa. Scientific American, 71(1):9–11, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
9.pdf.

Anonymous:1894:ILa

[4868] Anonymous. The irrigation of lawns. Scientific American, 71(1):11, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-11.pdf.

Anonymous:1894:BNBw

[4869] Anonymous. Books: New books and publications. Scientific American, 71(1):12, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n1/pdf/scientificamerican07071894-
12a.pdf.

Anonymous:1894:RPIz

[4870] Anonymous. Recently patented inventions. Scientific American, 71(1): 12, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-12.pdf.

Anonymous:1894:BPz

[4871] Anonymous. Business and personal. Scientific American, 71(1):13, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-13.pdf.

Anonymous:1894:Ibe

[4872] Anonymous. To inventors. Scientific American, 71(1):13, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-13b.pdf.

Anonymous:1894:NQz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-13a.pdf.

Anonymous:1894:IIz

[4874] Anonymous. Index of inventions. Scientific American, 71(1):13–14, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-13c.pdf.

Anonymous:1894:PNPy

[4875] Anonymous. Patent notes: Patents. Scientific American, 71(1):15, July 7, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n1/pdf/scientificamerican07071894-15.pdf.

Anonymous:1894:MCD

[4876] Anonymous. Moving the central draw span of seventh avenue bridge. Scientific American, 71(2):17–18, July 14, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n2/pdf/scientificamerican07141894-17.pdf.

Anonymous:1894:ATBa

[4877] Anonymous. Again a torpedo boat sinks a war ship. Scientific American, 71(2):18, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
18.pdf.

Anonymous:1894:GRSa

[4878] Anonymous. The great railroad strike. Scientific American, 71 (2):18–19, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
18a.pdf.

Anonymous:1894:CCG

[4879] Anonymous. A combined camera and graphoscope. Scientific American, 71(2):20, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
20a.pdf.

Anonymous:1894:EUH

[4880] Anonymous. An excavator for use in harbors, rivers, etc. Scientific American, 71(2):20, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
20b.pdf.

Anonymous:1894:IBH

[4881] Anonymous. An improved boiler for heating purposes. Scientific American, 71(2):20, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
20.pdf.

Anonymous:1894:DBPb

[4882] Anonymous. The Dowe bullet proof cuirass. Scientific American, 71(2):21, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
21.pdf.

Anonymous:1894:RMW

[4883] Anonymous. The Reagan marine water circulating and shaking grate. Scientific American, 71(2):21, July 14, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n2/pdf/scientificamerican07141894-21a.pdf.

Anonymous:1894:Cu

[4884] Anonymous. Correspondence. Scientific American, 71(2):22, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n2/pdf/scientificamerican07141894-22.pdf.

Anonymous:1894:IDM

[4885] Anonymous. Improved drilling machine for boilers. Scientific American, 71(2):23–24, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
23.pdf.

Anonymous:1894:NMF

[4886] Anonymous. Notes on the midwinter fair. Scientific American, 71(2):24, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n2/pdf/scientificamerican07141894-24.pdf.

Anonymous:1894:LHB

[4887] Anonymous. The life and habits of the beaver. Scientific American, 71(2):25–26, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
25.pdf.

Anonymous:1894:BFb

[4888] Anonymous. A beautiful fireplace. Scientific American, 71(2):27–28, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n2/pdf/scientificamerican07141894-27.pdf.

Anonymous:1894:BNBx

[4889] Anonymous. Books: New books and publications. Scientific American, 71(2):29, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
29a.pdf.

Anonymous:1894:BPba

[4890] Anonymous. Business and personal. Scientific American, 71(2):29, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n2/pdf/scientificamerican07141894-29b.pdf.

Anonymous:1894:NQba

[4891] Anonymous. Notes and queries. Scientific American, 71(2):29, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n2/pdf/scientificamerican07141894-29c.pdf.

Anonymous:1894:RPIba

[4892] Anonymous. Recently patented inventions. Scientific American, 71(2):29, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n2/pdf/scientificamerican07141894-
29.pdf.

Anonymous:1894:IIba

[4893] Anonymous. Index of inventions. Scientific American, 71(2):29–30, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n2/pdf/scientificamerican07141894-29d.pdf.

Anonymous:1894:PNPz

[4894] Anonymous. Patent notes: Patents. Scientific American, 71(2):31, July 14, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n2/pdf/scientificamerican07141894-31.pdf.

Anonymous:1894:HCP

[4895] Anonymous. Hardening composite plates for bank vaults. Scientific American, 71(3):33, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
33.pdf.

Anonymous:1894:ABP

[4896] Anonymous. An American bullet proof shield. Scientific American, 71(3):34, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
34.pdf.

Anonymous:1894:BM

[4897] Anonymous. Brooklyn meeting of the A.A.A.S. Scientific American, 71(3):34, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
34c.pdf.

Anonymous:1894:EA

[4898] Anonymous. Education in America. Scientific American, 71(3):34, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-34b.pdf.

Anonymous:1894:FLA

[4899] Anonymous. Failure of a large armor plate. Scientific American, 71(3):34, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
34a.pdf.

Anonymous:1894:GRSb

[4900] Anonymous. The great railroad strike. Scientific American, 71 (3):34–35, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
34d.pdf.

Anonymous:1894:LAC

[4901] Anonymous. Laying an Atlantic cable in twelve days. Scientific American, 71(3):35–36, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
35.pdf.

Anonymous:1894:ITa

[4902] Anonymous. An improved Tellurian. Scientific American, 71(3):36, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-36.pdf.

Anonymous:1894:ELE

[4903] Anonymous. Electric light engines, Manchester Corporation. Scientific American, 71(3):36–37, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
36a.pdf.

Anonymous:1894:NTb

[4904] Anonymous. A new telephone. Scientific American, 71(3):37, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-37a.pdf.

Anonymous:1894:VSR

[4905] Anonymous. Viaduct for street railways, Cincinnati. Scientific American, 71(3):37, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
37.pdf.

Anonymous:1894:Cv

[4906] Anonymous. Correspondence. Scientific American, 71(3):38–39, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-38.pdf.

Anonymous:1894:SIC

[4907] Anonymous. A steam ice cream manufactory. Scientific American, 71(3):40, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
40.pdf.

Anonymous:1894:RAZ

[4908] Anonymous. Roaring alligators in the Zoological Gardens at Hamburg. Scientific American, 71(3):41–43, July 21, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n3/pdf/scientificamerican07211894-41.pdf.

Anonymous:1894:Ab

[4909] Anonymous. The albatross. Scientific American, 71(3):43–44, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-43.pdf.

Anonymous:1894:RPIbb

[4910] Anonymous. Recently patented inventions. Scientific American, 71(3):44, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
44.pdf.

Anonymous:1894:BNBy

[4911] Anonymous. Books: New books and publications. Scientific American, 71(3):45, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n3/pdf/scientificamerican07211894-
45.pdf.

Anonymous:1894:BPbb

[4912] Anonymous. Business and personal. Scientific American, 71(3):45, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-45a.pdf.

Anonymous:1894:Ibg

[4913] Anonymous. To inventors. Scientific American, 71(3):45, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-45c.pdf.

Anonymous:1894:NQbb

[4914] Anonymous. Notes and queries. Scientific American, 71(3):45, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-45b.pdf.

Anonymous:1894:IIbb

[4915] Anonymous. Index of inventions. Scientific American, 71(3):45–46, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-45d.pdf.

Anonymous:1894:PNPba

[4916] Anonymous. Patent notes: Patents. Scientific American, 71(3):47, July 21, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n3/pdf/scientificamerican07211894-47.pdf.

Anonymous:1894:NST

[4917] Anonymous. Navy speed trials. Scientific American, 71(4):50–51, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-50.pdf.

Anonymous:1894:DIF

[4918] Anonymous. Davison’s “Imperial Fiber” letter copying book. Scientific American, 71(4):52, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
52.pdf.

Anonymous:1894:CBb

[4919] Anonymous. The Castle of Bonnetable. Scientific American, 71 (4):52–53, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
52a.pdf.

Anonymous:1894:ASP

[4920] Anonymous. The accident to the steamer Plymouth. Scientific American, 71(4):53, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
53.pdf.

Anonymous:1894:CGa

[4921] Anonymous. A Chinese game. Scientific American, 71(4):53, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-53a.pdf.

Anonymous:1894:Cw

[4922] Anonymous. Correspondence. Scientific American, 71(4):54, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-54.pdf.

Anonymous:1894:ERTa

[4923] Anonymous. The East River Tunnel of the East River Gas Company. Scientific American, 71(4):55, July 28, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n4/pdf/scientificamerican07281894-55.pdf.

Anonymous:1894:TTN

[4924] Anonymous. Trial trip of the new war ship Minneapolis. Scientific American, 71(4):56–57, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
56.pdf.

Anonymous:1894:IUT

[4925] Anonymous. Increasing use of traction engines. Scientific American, 71(4):57, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
57.pdf.

Anonymous:1894:RFA

[4926] Anonymous. Rapid filtering apparatus. Scientific American, 71 (4):57–58, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
57a.pdf.

Anonymous:1894:EN

[4927] Anonymous. Eudynamis Niger. Scientific American, 71(4):59–60, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-59.pdf.

Anonymous:1894:BNBz

[4928] Anonymous. Books: New books and publications. Scientific American, 71(4):61, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
61a.pdf.

Anonymous:1894:BPbc

[4929] Anonymous. Business and personal. Scientific American, 71(4):61, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-61b.pdf.

Anonymous:1894:RPIbc

[4930] Anonymous. Recently patented inventions. Scientific American, 71(4):61, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n4/pdf/scientificamerican07281894-
61.pdf.

Anonymous:1894:NQbc

[4931] Anonymous. Notes and queries. Scientific American, 71(4):61–62, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-61c.pdf.

Anonymous:1894:Ibi

[4932] Anonymous. To inventors. Scientific American, 71(4):62, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-62.pdf.

Anonymous:1894:IIbc

[4933] Anonymous. Index of inventions. Scientific American, 71(4):62, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-62a.pdf.

Anonymous:1894:PNPbb

[4934] Anonymous. Patent notes: Patents. Scientific American, 71(4):63, July 28, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n4/pdf/scientificamerican07281894-63.pdf.

Anonymous:1894:WBJ

[4935] Anonymous. War between Japan and China. Scientific American, 71(5):66–67, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
66.pdf.

Anonymous:1894:FAT

[4936] Anonymous. The fifth avenue theater, New York. Scientific American, 71(5):68, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
68.pdf.

Anonymous:1894:ICFb

[4937] Anonymous. An improved car fender. Scientific American, 71(5):68, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n5/pdf/scientificamerican08041894-68a.pdf.

Anonymous:1894:ITCd

[4938] Anonymous. Improved tenement construction. Scientific American, 71(5):69, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
69.pdf.

Anonymous:1894:SEG

[4939] Anonymous. A steam engine governor and reversing gear. Scientific American, 71(5):69, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
69a.pdf.

Anonymous:1894:BSC

[4940] Anonymous. Bersier’s steering compass. Scientific American, 71 (5):69–70, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
69b.pdf.

Anonymous:1894:ECM

[4941] Anonymous. An efficient calculating machine. Scientific American, 71(5):71, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
71.pdf.

Anonymous:1894:NTBa

[4942] Anonymous. The new Tower Bridge, London. Scientific American, 71(5):72–73, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
72.pdf.

Anonymous:1894:HSBc

[4943] Anonymous. The Halsted Street Bridge over the Chicago River. Scientific American, 71(5):73–75, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
73.pdf.

Anonymous:1894:HMP

[4944] Anonymous. The harmony and measurement of perfumes. Scientific American, 71(5):75–76, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
75.pdf.

Anonymous:1894:RPIbd

[4945] Anonymous. Recently patented inventions. Scientific American, 71(5):76–77, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n5/pdf/scientificamerican08041894-
76.pdf.

Anonymous:1894:BPbd

[4946] Anonymous. Business and personal. Scientific American, 71(5):77, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n5/pdf/scientificamerican08041894-77.pdf.

Anonymous:1894:Ibj

[4947] Anonymous. To inventors. Scientific American, 71(5):77, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n5/pdf/scientificamerican08041894-77b.pdf.

Anonymous:1894:NQbd

[4948] Anonymous. Notes and queries. Scientific American, 71(5):77, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n5/pdf/scientificamerican08041894-77a.pdf.

Anonymous:1894:IIbd

[4949] Anonymous. Index of inventions. Scientific American, 71(5):77–78, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n5/pdf/scientificamerican08041894-77c.pdf.

Anonymous:1894:PNPbc

[4950] Anonymous. Patent notes: Patents. Scientific American, 71(5):79, August 4, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n5/pdf/scientificamerican08041894-79.pdf.

Anonymous:1894:NOS

[4951] Anonymous. The new ocean steamers St. Louis and St. Paul. Scientific American, 71(6):81–82, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
81.pdf.

Anonymous:1894:BMAb

[4952] Anonymous. The Brooklyn meeting of the American Association. Scientific American, 71(6):82, August 11, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n6/pdf/scientificamerican08111894-82.pdf.

Anonymous:1894:WCM

[4953] Anonymous. The waste of coal mines as a source of power. Scientific American, 71(6):82–84, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
82a.pdf.

Anonymous:1894:ICR

[4954] Anonymous. An improved check rein support. Scientific American, 71(6):84, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
84a.pdf.

Anonymous:1894:NTS

[4955] Anonymous. New tool sharpener. Scientific American, 71(6):84, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n6/pdf/scientificamerican08111894-84.pdf.

Anonymous:1894:MWB

[4956] Anonymous. Manufacture of whisk brooms. Scientific American, 71(6):85–87, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
85.pdf.

Anonymous:1894:ACT

[4957] Anonymous. Automatic coupler tests. Scientific American, 71 (6):88, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
88.pdf.

Anonymous:1894:TSS

[4958] Anonymous. Three specimens of starling. Scientific American, 71(6):89–90, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
89.pdf.

Anonymous:1894:EWRb

[4959] Anonymous. The electric welding of rail joints. Scientific American, 71(6):91–92, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
91.pdf.

Anonymous:1894:BPbe

[4960] Anonymous. Business and personal. Scientific American, 71(6):93, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n6/pdf/scientificamerican08111894-93a.pdf.

Anonymous:1894:Ibk

[4961] Anonymous. To inventors. Scientific American, 71(6):93, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n6/pdf/scientificamerican08111894-93c.pdf.

Anonymous:1894:NQbe

[4962] Anonymous. Notes and queries. Scientific American, 71(6):93, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n6/pdf/scientificamerican08111894-93b.pdf.

Anonymous:1894:RPIbe

[4963] Anonymous. Recently patented inventions. Scientific American, 71(6):93, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n6/pdf/scientificamerican08111894-
93.pdf.

Anonymous:1894:IIbe

[4964] Anonymous. Index of inventions. Scientific American, 71(6):93–94, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n6/pdf/scientificamerican08111894-93d.pdf.

Anonymous:1894:PNPbd

[4965] Anonymous. Patent notes: Patents. Scientific American, 71(6):95, August 11, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n6/pdf/scientificamerican08111894-95.pdf.

Anonymous:1894:FMI

[4966] Anonymous. The farmer and modern invention. Scientific American, 71(7):98–99, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
98.pdf.

Anonymous:1894:AEL

[4967] Anonymous. An adjustable electric lamp holder. Scientific American, 71(7):100, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
100a.pdf.

Anonymous:1894:INL

[4968] Anonymous. An improved nut lock. Scientific American, 71(7):100, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n7/pdf/scientificamerican08181894-100b.pdf.

Anonymous:1894:ITb

[4969] Anonymous. Improvements in telephones. Scientific American, 71(7):100, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
100.pdf.

Anonymous:1894:RSR

[4970] Anonymous. Raney’s self-register for weighing scales. Scientific American, 71(7):100–101, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
100c.pdf.

Anonymous:1894:LUE

[4971] Anonymous. The Lyons Universal Exposition of 1894. Scientific American, 71(7):101–102, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
101.pdf.

Anonymous:1894:EHb

[4972] Anonymous. The evolution of handicraft. Scientific American, 71(7):103, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
103.pdf.

Benjamin:1894:DGB

[4973] Marcus Benjamin. Daniel Garrison Brinton. Scientific American, 71(7):104–105, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
104.pdf.

Anonymous:1894:TS

[4974] Anonymous. The Toxophilite Society. Scientific American, 71 (7):105, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
105.pdf.

Anonymous:1894:IMPa

[4975] Anonymous. An improved motor or pump. Scientific American, 71(7):105–106, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
105a.pdf.

Anonymous:1894:SL

[4976] Anonymous. The song of the lark. Scientific American, 71(7): 107–108, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
107.pdf.

Anonymous:1894:RPIbf

[4977] Anonymous. Recently patented inventions. Scientific American, 71 (7):108–109, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
108.pdf.

Anonymous:1894:BNBba

[4978] Anonymous. Books: New books and publications. Scientific American, 71(7):109, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
109.pdf.

Anonymous:1894:BPbf

[4979] Anonymous. Business and personal. Scientific American, 71(7):109, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n7/pdf/scientificamerican08181894-109a.pdf.

Anonymous:1894:Ibl

[4980] Anonymous. To inventors. Scientific American, 71(7):109, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n7/pdf/scientificamerican08181894-109c.pdf.

Anonymous:1894:NQbf

[4981] Anonymous. Notes and queries. Scientific American, 71(7):109, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n7/pdf/scientificamerican08181894-109b.pdf.

Anonymous:1894:IIbf

[4982] Anonymous. Index of inventions. Scientific American, 71(7): 109–110, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n7/pdf/scientificamerican08181894-
109d.pdf.

Anonymous:1894:PNPbe

[4983] Anonymous. Patent notes: Patents. Scientific American, 71(7):111, August 18, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n7/pdf/scientificamerican08181894-111.pdf.

Anonymous:1894:CFP

[4984] Anonymous. Caisson foundation piers of the American Surety Company’s building in New York City. Scientific American, 71(8):113, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-113.pdf.

Anonymous:1894:BMAc

[4985] Anonymous. The Brooklyn meeting of the American Association for the Advancement of Science. Scientific American, 71(8):114–115, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-114.pdf.

Anonymous:1894:CWa

[4986] Anonymous. Caisson work. Scientific American, 71(8):115, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-115.pdf.

Anonymous:1894:ICAb

[4987] Anonymous. An improved car axle box. Scientific American, 71(8):116, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
116.pdf.

Anonymous:1894:ISFa

[4988] Anonymous. An improved steam fire engine. Scientific American, 71(8):116, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
116a.pdf.

Anonymous:1894:ISBa

[4989] Anonymous. An improved submarine boat. Scientific American, 71(8):116–117, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
116b.pdf.

Anonymous:1894:EIO

[4990] Anonymous. Electric illumination of an orange. Scientific American, 71(8):117–118, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
117.pdf.

Summers:1894:SBP

[4991] C. R. Summers. Slow beating pendulums. Scientific American, 71(8):118, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
118.pdf.

Anonymous:1894:Cx

[4992] Anonymous. Correspondence. Scientific American, 71(8):119, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-119.pdf.

Anonymous:1894:OMV

[4993] Anonymous. Ogden’s marine velocipede or bicycle boat. Scientific American, 71(8):121, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
121.pdf.

Anonymous:1894:SHL

[4994] Anonymous. Sir A. H. Layard, G.C.B. Scientific American, 71 (8):121, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
121a.pdf.

Anonymous:1894:DRPc

[4995] Anonymous. Decisions relating to patents. Scientific American, 71(8):122, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
122.pdf.

Anonymous:1894:IYR

[4996] Anonymous. The International Yacht Races. Scientific American, 71(8):123, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
123.pdf.

Anonymous:1894:TMF

[4997] Anonymous. A trial of Maxim’s flying machine. Scientific American, 71(8):123–124, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
123a.pdf.

Anonymous:1894:BPbg

[4998] Anonymous. Business and personal. Scientific American, 71(8):125, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-125a.pdf.

Anonymous:1894:Ibm

[4999] Anonymous. To inventors. Scientific American, 71(8):125, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-125c.pdf.

Anonymous:1894:NQbg

[5000] Anonymous. Notes and queries. Scientific American, 71(8):125, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-125b.pdf.

Anonymous:1894:RPIbg

[5001] Anonymous. Recently patented inventions. Scientific American, 71(8):125, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
125.pdf.

Anonymous:1894:IIbg

[5002] Anonymous. Index of inventions. Scientific American, 71(8): 125–126, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n8/pdf/scientificamerican08251894-
125d.pdf.

Anonymous:1894:PNPbf

[5003] Anonymous. Patent notes: Patents. Scientific American, 71(8):127, August 25, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n8/pdf/scientificamerican08251894-127.pdf.

Anonymous:1894:AAA

[5004] Anonymous. The American Association for the Advancement of Science. Scientific American, 71(9):130–131, September 1, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n9/pdf/scientificamerican09011894-130.pdf.

Anonymous:1894:CSO

[5005] Anonymous. Cotton seed oil presses wanted. Scientific American, 71(9):131, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
131a.pdf.

Anonymous:1894:NGL

[5006] Anonymous. Niagara and the Great Lakes. Scientific American, 71(9):131, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
131.pdf.

Anonymous:1894:ADA

[5007] Anonymous. An automatic discharger of abnormal electric currents. Scientific American, 71(9):132, September 1, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n9/pdf/scientificamerican09011894-132b.pdf.

Anonymous:1894:CWW

[5008] Anonymous. The cascade water wheel. Scientific American, 71 (9):132, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
132a.pdf.

Anonymous:1894:ISFb

[5009] Anonymous. An improved sight for firearms. Scientific American, 71(9):132, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
132.pdf.

Anonymous:1894:BTA

[5010] Anonymous. Bull teasing in the Arena of Arles. Scientific American, 71(9):133, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
133a.pdf.

Anonymous:1894:LGS

[5011] Anonymous. A life guard for street railway cars. Scientific American, 71(9):133, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
133.pdf.

Anonymous:1894:Cy

[5012] Anonymous. Correspondence. Scientific American, 71(9):134, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n9/pdf/scientificamerican09011894-134.pdf.

Anonymous:1894:CBR

[5013] Anonymous. Church of Borki, Russia. Scientific American, 71 (9):135, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
135.pdf.

Anonymous:1894:MS

[5014] Anonymous. Manufacture of Spiegeleisen. Scientific American, 71 (9):136–137, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
136.pdf.

Anonymous:1894:SECb

[5015] Anonymous. Steam and electric cableway for logging and canal boat towing. Scientific American, 71(9):137–139, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n9/pdf/scientificamerican09011894-137.pdf.

Anonymous:1894:RTC

[5016] Anonymous. A recording thermometer for closed spaces. Scientific American, 71(9):139–140, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
139.pdf.

Anonymous:1894:RPIbh

[5017] Anonymous. Recently patented inventions. Scientific American, 71(9):140, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
140.pdf.

Anonymous:1894:BNBbb

[5018] Anonymous. Books: New books and publications. Scientific American, 71(9):141, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
141.pdf.

Anonymous:1894:BPbh

[5019] Anonymous. Business and personal. Scientific American, 71(9): 141, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
141a.pdf.

Anonymous:1894:Ibn

[5020] Anonymous. To inventors. Scientific American, 71(9):141, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n9/pdf/scientificamerican09011894-141c.pdf.

Anonymous:1894:NQbh

[5021] Anonymous. Notes and queries. Scientific American, 71(9):141, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n9/pdf/scientificamerican09011894-141b.pdf.

Anonymous:1894:IIbh

[5022] Anonymous. Index of inventions. Scientific American, 71(9): 141–142, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
141d.pdf.

Anonymous:1894:PNPbg

[5023] Anonymous. Patent notes: Patents. Scientific American, 71 (9):143, September 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n9/pdf/scientificamerican09011894-
143.pdf.

Anonymous:1894:AFA

[5024] Anonymous. American Forestry Association. Scientific American, 71(10):146, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
146.pdf.

Anonymous:1894:MSSb

[5025] Anonymous. The magnitude of the solar system. Scientific American, 71(10):146–147, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
146a.pdf.

Anonymous:1894:TSN

[5026] Anonymous. Trial at sea of a navigating tricycle. Scientific American, 71(10):148, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
148.pdf.

Anonymous:1894:ICCb

[5027] Anonymous. An improved car coupling. Scientific American, 71 (10):149, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
149a.pdf.

Anonymous:1894:IIEa

[5028] Anonymous. An improved instrument for engineers, surveyors, etc. Scientific American, 71(10):149, September 8, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n10/pdf/scientificamerican09081894-149.pdf.

Anonymous:1894:EFW

[5029] Anonymous. An effective feed water heater. Scientific American, 71 (10):149–150, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
149b.pdf.

Anonymous:1894:INT

[5030] Anonymous. An improved needle threader. Scientific American, 71(10):150, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
150.pdf.

Anonymous:1894:SLO

[5031] Anonymous. Strengthening lead and other pipes. Scientific American, 71(10):150, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
150a.pdf.

Anonymous:1894:EPSa

[5032] Anonymous. The Eastern Power Station of the Brooklyn City Railroad Company. Scientific American, 71(10):151–152, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n10/pdf/scientificamerican09081894-151.pdf.

Anonymous:1894:CSN

[5033] Anonymous. Church spire, Norwich, Conn., struck by lightning. Scientific American, 71(10):152, September 8, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n10/pdf/scientificamerican09081894-152a.pdf.

Anonymous:1894:TSL

[5034] Anonymous. A tree shattered by lightning. Scientific American, 71(10):152, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
152.pdf.

Anonymous:1894:BDFb

[5035] Anonymous. Broadway during a fire. Scientific American, 71(10): 153–154, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
153.pdf.

Anonymous:1894:AEM

[5036] Anonymous. Approaching eclipse of the Moon. Scientific American, 71(10):154, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
154.pdf.

Anonymous:1894:PCV

[5037] Anonymous. Professor Charles Valentine Riley. Scientific American, 71(10):155–156, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
155.pdf.

Anonymous:1894:BPbi

[5038] Anonymous. Business and personal. Scientific American, 71 (10):157, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
157a.pdf.

Anonymous:1894:Ibo

[5039] Anonymous. To inventors. Scientific American, 71(10):157, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n10/pdf/scientificamerican09081894-157c.pdf.

Anonymous:1894:NQbi

[5040] Anonymous. Notes and queries. Scientific American, 71(10): 157, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
157b.pdf.

Anonymous:1894:RPIbi

[5041] Anonymous. Recently patented inventions. Scientific American, 71(10):157, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
157.pdf.

Anonymous:1894:IIbi

[5042] Anonymous. Index of inventions. Scientific American, 71(10): 157–158, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
157d.pdf.

Anonymous:1894:PNPbh

[5043] Anonymous. Patent notes: Patents. Scientific American, 71 (10):159, September 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n10/pdf/scientificamerican09081894-
159.pdf.

Anonymous:1894:ASE

[5044] Anonymous. Allis steam engines, oiling system and subsidiary apparatus in the Eastern Power House of the Brooklyn City Railroad Co. Scientific American, 71(11):161, September 15, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n11/pdf/scientificamerican09151894-161.pdf.

Anonymous:1894:FF

[5045] Anonymous. The forest fires. Scientific American, 71(11):162– 163, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
162.pdf.

Anonymous:1894:IBP

[5046] Anonymous. An improved brick press. Scientific American, 71 (11):164, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
164b.pdf.

Anonymous:1894:IMPb

[5047] Anonymous. An improved metal punching machine. Scientific American, 71(11):164, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
164.pdf.

Anonymous:1894:MBM

[5048] Anonymous. Mammoth block of mahogany. Scientific American, 71(11):164, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
164a.pdf.

Anonymous:1894:TMS

[5049] Anonymous. Trial of Maxim’s steam flying machine. Scientific American, 71(11):165–166, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
165.pdf.

Anonymous:1894:Cz

[5050] Anonymous. Correspondence. Scientific American, 71(11):167– 168, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
167.pdf.

Anonymous:1894:NUSb

[5051] Anonymous. The new United States cruiser Cincinnati. Scientific American, 71(11):168, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
168.pdf.

Anonymous:1894:SLR

[5052] Anonymous. A search light range finder. Scientific American, 71 (11):168, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
168a.pdf.

Anonymous:1894:EMTb

[5053] Anonymous. The Eddy Malay tailless kite. Scientific American, 71 (11):169–170, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
169.pdf.

Anonymous:1894:ARJ

[5054] Anonymous. An automatic rail joint. Scientific American, 71(11): 171–172, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
171.pdf.

Anonymous:1894:RPIbj

[5055] Anonymous. Recently patented inventions. Scientific American, 71 (11):172–173, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
172.pdf.

Anonymous:1894:BNBbc

[5056] Anonymous. Books: New books and publications. Scientific American, 71(11):173, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
173b.pdf.

Anonymous:1894:BPbj

[5057] Anonymous. Business and personal. Scientific American, 71 (11):173, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
173.pdf.

Anonymous:1894:NQbj

[5058] Anonymous. Notes and queries. Scientific American, 71(11): 173, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
173a.pdf.

Anonymous:1894:IIbj

[5059] Anonymous. Index of inventions. Scientific American, 71(11): 173–174, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
173c.pdf.

Anonymous:1894:PNPbi

[5060] Anonymous. Patent notes: Patents. Scientific American, 71 (11):175, September 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n11/pdf/scientificamerican09151894-
175.pdf.

Anonymous:1894:PMI

[5061] Anonymous. Patents for minor inventions. Scientific American, 71 (12):178–179, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
178.pdf.

Anonymous:1894:DRPd

[5062] Anonymous. Decisions relating to patents. Scientific American, 71 (12):179, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
179.pdf.

Durfee:1894:AL

[5063] W. F. Durfee. An ancient locomotive. Scientific American, 71 (12):180–181, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
180.pdf.

Anonymous:1894:ISRa

[5064] Anonymous. An improved street railway car fender. Scientific American, 71(12):181, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
181a.pdf.

Anonymous:1894:SNA

[5065] Anonymous. Some new American knitting machines. Scientific American, 71(12):181, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
181.pdf.

Anonymous:1894:Cbc

[5066] Anonymous. Correspondence. Scientific American, 71(12):182, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n12/pdf/scientificamerican09221894-182.pdf.

Anonymous:1894:USW

[5067] Anonymous. The United States war ship Atlanta ramming a derelict. Scientific American, 71(12):183–184, September 22, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n12/pdf/scientificamerican09221894-183.pdf.

Anonymous:1894:ITBa

[5068] Anonymous. An improved tubular boiler. Scientific American, 71 (12):184, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
184a.pdf.

Elkin:1894:IPM

[5069] W. L. Elkin. Instrument for the photography of meteors for the Yale Observatory. Scientific American, 71(12):184, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n12/pdf/scientificamerican09221894-184.pdf.

Anonymous:1894:PC

[5070] Anonymous. Parasites on a caterpillar. Scientific American, 71 (12):185, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
185.pdf.

Anonymous:1894:HFS

[5071] Anonymous. The Hibernia — a fast steam launch. Scientific American, 71(12):185–186, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
185a.pdf.

Anonymous:1894:MBL

[5072] Anonymous. Manufacture of belt lacing. Scientific American, 71 (12):187–188, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
187.pdf.

Anonymous:1894:BPbk

[5073] Anonymous. Business and personal. Scientific American, 71 (12):189, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
189a.pdf.

Anonymous:1894:NQbk

[5074] Anonymous. Notes and queries. Scientific American, 71(12): 189, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
189b.pdf.

Anonymous:1894:RPIbk

[5075] Anonymous. Recently patented inventions. Scientific American, 71 (12):189, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
189.pdf.

Anonymous:1894:BNBbd

[5076] Anonymous. Books: New books and publications. Scientific American, 71(12):190, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
190.pdf.

Anonymous:1894:Ibq

[5077] Anonymous. To inventors. Scientific American, 71(12):190, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n12/pdf/scientificamerican09221894-190a.pdf.

Anonymous:1894:IIbk

[5078] Anonymous. Index of inventions. Scientific American, 71(12): 190, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
190b.pdf.

Anonymous:1894:PNPbj

[5079] Anonymous. Patent notes: Patents. Scientific American, 71 (12):191, September 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n12/pdf/scientificamerican09221894-
191.pdf.

Anonymous:1894:UBS

[5080] Anonymous. The U.S. battle ship Maine. Scientific American, 71 (13):193, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
193.pdf.

Anonymous:1894:MV

[5081] Anonymous. Man and vegetation. Scientific American, 71(13): 194–195, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
194.pdf.

Anonymous:1894:CGG

[5082] Anonymous. The charter gas and gasoline engine. Scientific American, 71(13):196, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
196a.pdf.

Anonymous:1894:OS

[5083] Anonymous. The optometer Skiascope. Scientific American, 71 (13):196, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
196.pdf.

Anonymous:1894:PHV

[5084] Anonymous. A portion of the Hotel De Ville, Seville, Spain. Scientific American, 71(13):197–198, September 29, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n13/pdf/scientificamerican09291894-197.pdf.

Anonymous:1894:CCL

[5085] Anonymous. A convenient carriage lamp. Scientific American, 71 (13):198, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
198a.pdf.

Anonymous:1894:PCN

[5086] Anonymous. A powerful cutting nipper. Scientific American, 71 (13):198, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
198.pdf.

Anonymous:1894:Cbd

[5087] Anonymous. Correspondence. Scientific American, 71(13):199, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n13/pdf/scientificamerican09291894-199.pdf.

Anonymous:1894:EAP

[5088] Anonymous. The Eighth Avenue Plaza at the entrance to Central Park. Scientific American, 71(13):200, September 29, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n13/pdf/scientificamerican09291894-200.pdf.

Anonymous:1894:TBDb

[5089] Anonymous. The torpedo boat destroyer Ferret. Scientific American, 71(13):201, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
201.pdf.

Anonymous:1894:FAB

[5090] Anonymous. The Festival of Athletes in Breslau. Scientific American, 71(13):201–203, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
201a.pdf.

Anonymous:1894:SW

[5091] Anonymous. Speaking watches. Scientific American, 71(13):203– 204, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
203.pdf.

Anonymous:1894:BPbl

[5092] Anonymous. Business and personal. Scientific American, 71 (13):205, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
205a.pdf.

Anonymous:1894:RPIbl

[5093] Anonymous. Recently patented inventions. Scientific American, 71 (13):205, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
205.pdf.

Anonymous:1894:NQbl

[5094] Anonymous. Notes and queries. Scientific American, 71(13): 205–206, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
205b.pdf.

Anonymous:1894:Ibr

[5095] Anonymous. To inventors. Scientific American, 71(13):206, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n13/pdf/scientificamerican09291894-206.pdf.

Anonymous:1894:IIbl

[5096] Anonymous. Index of inventions. Scientific American, 71(13): 206, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
206a.pdf.

Anonymous:1894:PNPbk

[5097] Anonymous. Patent notes: Patents. Scientific American, 71 (13):207, September 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n13/pdf/scientificamerican09291894-
207.pdf.

Anonymous:1894:MCG

[5098] Anonymous. Motor carriages and good roads. Scientific American, 71(14):210–211, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
210.pdf.

Anonymous:1894:LRM

[5099] Anonymous. The long-running meteorograph of the Mont Blanc Observatory. Scientific American, 71(14):212–213, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n14/pdf/scientificamerican10061894-212.pdf.

Anonymous:1894:PGC

[5100] Anonymous. A platform gate for cars, etc. Scientific American, 71(14):213, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
213.pdf.

Anonymous:1894:VRS

[5101] Anonymous. A velocipede to run on snow and ice. Scientific American, 71(14):213, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
213a.pdf.

Anonymous:1894:LCC

[5102] Anonymous. A line chalker for carpenters’ use. Scientific American, 71(14):213–214, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
213b.pdf.

Anonymous:1894:ERTb

[5103] Anonymous. Easily read thermometers. Scientific American, 71 (14):214, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
214.pdf.

Anonymous:1894:Cbe

[5104] Anonymous. Correspondence. Scientific American, 71(14):215, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-215.pdf.

Anonymous:1894:Ma

[5105] Anonymous. Multiphotography. Scientific American, 71(14):216, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-216.pdf.

Riley:1894:MMF

[5106] C. V. Riley. Mr. Maxim’s flying machine. Scientific American, 71(14):217, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
217.pdf.

Payne:1894:P

[5107] Wm. W. Payne. The Pleiades. Scientific American, 71(14):218, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-218.pdf.

Anonymous:1894:FWE

[5108] Anonymous. The Fort Wayne Electric Corporation — its dynamos and General Electric lighting apparatus. Scientific American, 71(14):219, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-219.pdf.

Anonymous:1894:RPIbm

[5109] Anonymous. Recently patented inventions. Scientific American, 71(14):220, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
220.pdf.

Anonymous:1894:BNBbe

[5110] Anonymous. Books: New books and publications. Scientific American, 71(14):220–221, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n14/pdf/scientificamerican10061894-
220a.pdf.

Anonymous:1894:BPbm

[5111] Anonymous. Business and personal. Scientific American, 71(14):221, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-221.pdf.

Anonymous:1894:Ibs

[5112] Anonymous. To inventors. Scientific American, 71(14):221, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-221b.pdf.

Anonymous:1894:NQbm

[5113] Anonymous. Notes and queries. Scientific American, 71(14):221, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-221a.pdf.

Anonymous:1894:IIbm

[5114] Anonymous. Index of inventions. Scientific American, 71(14):221–222, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-221c.pdf.

Anonymous:1894:PNPbl

[5115] Anonymous. Patent notes: Patents. Scientific American, 71(14):223, October 6, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n14/pdf/scientificamerican10061894-223.pdf.

Anonymous:1894:Mb

[5116] Anonymous. Mars. Scientific American, 71(15):226–227, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n15/pdf/scientificamerican10131894-226.pdf.

Anonymous:1894:IGS

[5117] Anonymous. An improved gun-sight. Scientific American, 71 (15):228, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
228a.pdf.

Anonymous:1894:ISS

[5118] Anonymous. An improved street sweeper. Scientific American, 71(15):228, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
228.pdf.

Anonymous:1894:ONB

[5119] Anonymous. Old and new buildings of New York. Scientific American, 71(15):228–229, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
228b.pdf.

Anonymous:1894:PPN

[5120] Anonymous. Pork packing in New Jersey. Scientific American, 71 (15):229–230, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
229.pdf.

Anonymous:1894:NCDa

[5121] Anonymous. The new Croton Dam on the Cornell site. Scientific American, 71(15):230, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
230.pdf.

Anonymous:1894:Cbf

[5122] Anonymous. Correspondence. Scientific American, 71(15):231, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n15/pdf/scientificamerican10131894-231.pdf.

Anonymous:1894:GSD

[5123] Anonymous. German shepherd dogs. Scientific American, 71 (15):232, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
232.pdf.

Anonymous:1894:LO

[5124] Anonymous. The Lowe Observatory. Scientific American, 71 (15):233, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
233.pdf.

Lander:1894:HBS

[5125] Benjamin Lander. Hut-building seventeen-year cicadas. Scientific American, 71(15):233–235, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
233a.pdf.

Anonymous:1894:AMb

[5126] Anonymous. Artificial mirages. Scientific American, 71(15):235– 236, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
235.pdf.

Anonymous:1894:RPIbn

[5127] Anonymous. Recently patented inventions. Scientific American, 71 (15):236–237, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
236.pdf.

Anonymous:1894:BPbn

[5128] Anonymous. Business and personal. Scientific American, 71 (15):237, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
237.pdf.

Anonymous:1894:Ibu

[5129] Anonymous. To inventors. Scientific American, 71(15):237, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n15/pdf/scientificamerican10131894-237b.pdf.

Anonymous:1894:NQbn

[5130] Anonymous. Notes and queries. Scientific American, 71(15):237, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n15/pdf/scientificamerican10131894-237a.pdf.

Anonymous:1894:IIbn

[5131] Anonymous. Index of inventions. Scientific American, 71(15): 237–238, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
237c.pdf.

Anonymous:1894:PNPbm

[5132] Anonymous. Patent notes: Patents. Scientific American, 71 (15):239, October 13, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n15/pdf/scientificamerican10131894-
239.pdf.

Anonymous:1894:APR

[5133] Anonymous. Acoustic properties and resonance of buildings. Scientific American, 71(16):242–243, October 20, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n16/pdf/scientificamerican10201894-242.pdf.

Anonymous:1894:EUR

[5134] Anonymous. An excavator for use on river banks, etc. Scientific American, 71(16):244, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
244a.pdf.

Anonymous:1894:ICFc

[5135] Anonymous. An improved car fender. Scientific American, 71 (16):244, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
244.pdf.

Anonymous:1894:SMa

[5136] Anonymous. Swamp music. Scientific American, 71(16):244–245, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n16/pdf/scientificamerican10201894-244b.pdf.

Anonymous:1894:DCM

[5137] Anonymous. “Dry Cell” medical batteries. Scientific American, 71(16):245, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
245.pdf.

Anonymous:1894:HSc

[5138] Anonymous. The harmless smoker. Scientific American, 71(16):245, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n16/pdf/scientificamerican10201894-245a.pdf.

Anonymous:1894:UWP

[5139] Anonymous. Utilizing the water power of Niagara Falls. Scientific American, 71(16):245–246, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
245b.pdf.

Anonymous:1894:DRPe

[5140] Anonymous. Decisions relating to patents. Scientific American, 71(16):246, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
246.pdf.

Anonymous:1894:GCC

[5141] Anonymous. The Great Chicago Canal. Scientific American, 71 (16):247, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
247.pdf.

Anonymous:1894:MCb

[5142] Anonymous. The Metropolitan Club. Scientific American, 71 (16):248–249, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
248.pdf.

Anonymous:1894:USR

[5143] Anonymous. The United States Ram Katahdin. Scientific American, 71(16):249–250, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
249.pdf.

Anonymous:1894:TZ

[5144] Anonymous. The training of zebras. Scientific American, 71 (16):251–252, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
251.pdf.

Anonymous:1894:RPIbo

[5145] Anonymous. Recently patented inventions. Scientific American, 71 (16):252–253, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
252.pdf.

Anonymous:1894:BNBbf

[5146] Anonymous. Books: New books and publications. Scientific American, 71(16):253, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
253.pdf.

Anonymous:1894:BPbo

[5147] Anonymous. Business and personal. Scientific American, 71 (16):253, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
253a.pdf.

Anonymous:1894:CRa

[5148] Anonymous. Communications received. Scientific American, 71 (16):253, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
253c.pdf.

Anonymous:1894:Ibv

[5149] Anonymous. To inventors. Scientific American, 71(16):253, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n16/pdf/scientificamerican10201894-253d.pdf.

Anonymous:1894:NQbo

[5150] Anonymous. Notes and queries. Scientific American, 71(16):253, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n16/pdf/scientificamerican10201894-253b.pdf.

Anonymous:1894:IIbo

[5151] Anonymous. Index of inventions. Scientific American, 71(16): 253–254, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
253e.pdf.

Anonymous:1894:PNPbn

[5152] Anonymous. Patent notes: Patents. Scientific American, 71 (16):255, October 20, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n16/pdf/scientificamerican10201894-
255.pdf.

Anonymous:1894:NHR

[5153] Anonymous. The new Harlem River Speedway, New York. Scientific American, 71(17):257, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
257.pdf.

Anonymous:1894:CTM

[5154] Anonymous. The coming transit of Mercury. Scientific American, 71(17):258–259, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
258.pdf.

Anonymous:1894:BPEa

[5155] Anonymous. A burglar proof express car. Scientific American, 71(17):260, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
260.pdf.

Anonymous:1894:ISBb

[5156] Anonymous. An improved steam boiler. Scientific American, 71 (17):260, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
260b.pdf.

Anonymous:1894:SBM

[5157] Anonymous. Steel bridge, Malleco River, Chile. Scientific American, 71(17):260, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
260a.pdf.

Anonymous:1894:ATBb

[5158] Anonymous. An aluminum torpedo boat. Scientific American, 71(17):261, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
261.pdf.

Anonymous:1894:ESB

[5159] Anonymous. The “Eureka” smoke bleacher. Scientific American, 71 (17):261–262, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
261a.pdf.

Anonymous:1894:GSE

[5160] Anonymous. The Grand Saloon and electric lighting of the steamer Priscilla, of the Fall River Line. Scientific American, 71(17): 264–266, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
264.pdf.

Anonymous:1894:CCN

[5161] Anonymous. A cottage at Cranford, New Jersey. Scientific American, 71(17):267–268, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
267.pdf.

Anonymous:1894:RPIbp

[5162] Anonymous. Recently patented inventions. Scientific American, 71 (17):268–269, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
268.pdf.

Anonymous:1894:BPbp

[5163] Anonymous. Business and personal. Scientific American, 71 (17):269, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
269.pdf.

Anonymous:1894:CRb

[5164] Anonymous. Communications received. Scientific American, 71 (17):269, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
269b.pdf.

Anonymous:1894:Ibw

[5165] Anonymous. To inventors. Scientific American, 71(17):269, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n17/pdf/scientificamerican10271894-269c.pdf.

Anonymous:1894:NQbp

[5166] Anonymous. Notes and queries. Scientific American, 71(17):269, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n17/pdf/scientificamerican10271894-269a.pdf.

Anonymous:1894:IIbp

[5167] Anonymous. Index of inventions. Scientific American, 71(17): 269–270, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
269d.pdf.

Anonymous:1894:PNPbo

[5168] Anonymous. Patent notes: Patents. Scientific American, 71 (17):271, October 27, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n17/pdf/scientificamerican10271894-
271.pdf.

Anonymous:1894:APO

[5169] Anonymous. The aeroplane in the Patent Office. Scientific American, 71(18):274–275, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
274.pdf.

Anonymous:1894:CUS

[5170] Anonymous. Collars for use on shafting. Scientific American, 71 (18):276, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
276b.pdf.

Anonymous:1894:FPV

[5171] Anonymous. A foot propelled vehicle. Scientific American, 71 (18):276, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
276.pdf.

Anonymous:1894:SHS

[5172] Anonymous. A screw holding screwdriver. Scientific American, 71(18):276, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
276d.pdf.

Anonymous:1894:TBC

[5173] Anonymous. A tail board catch for dumping vehicles. Scientific American, 71(18):276, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
276a.pdf.

Anonymous:1894:TTG

[5174] Anonymous. A tower tank — 30,000 gallons capacity. Scientific American, 71(18):276, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
276c.pdf.

Anonymous:1894:NSE

[5175] Anonymous. A Newcomen steam engine. Scientific American, 71 (18):277–278, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
277.pdf.

Anonymous:1894:WSN

[5176] Anonymous. The water supply of New York City — the new Carmel Dam and Reservoir. Scientific American, 71(18):279, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n18/pdf/scientificamerican11031894-279.pdf.

Anonymous:1894:MRE

[5177] Anonymous. The magazine rifles of Europe. Scientific American, 71 (18):280–281, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
280.pdf.

Goelet:1894:HDA

[5178] Augustin H. Goelet. How to deal with apparent death from electric shock. Scientific American, 71(18):281–282, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n18/pdf/scientificamerican11031894-281.pdf.

Anonymous:1894:TFBb

[5179] Anonymous. A tour of France in a balloon. Scientific American, 71 (18):283–284, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
283.pdf.

Anonymous:1894:RPIbq

[5180] Anonymous. Recently patented inventions. Scientific American, 71(18):284, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
284.pdf.

Anonymous:1894:BNBbg

[5181] Anonymous. Books: New books and publications. Scientific American, 71(18):285, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
285.pdf.

Anonymous:1894:BPbq

[5182] Anonymous. Business and personal. Scientific American, 71 (18):285, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
285a.pdf.

Anonymous:1894:CRc

[5183] Anonymous. Communications received. Scientific American, 71 (18):285, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
285c.pdf.

Anonymous:1894:Ibx

[5184] Anonymous. To inventors. Scientific American, 71(18):285, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n18/pdf/scientificamerican11031894-285d.pdf.

Anonymous:1894:NQbq

[5185] Anonymous. Notes and queries. Scientific American, 71(18): 285, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
285b.pdf.

Anonymous:1894:IIbq

[5186] Anonymous. Index of inventions. Scientific American, 71(18): 285–286, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
285e.pdf.

Anonymous:1894:PNPbp

[5187] Anonymous. Patent notes: Patents. Scientific American, 71 (18):287, November 3, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n18/pdf/scientificamerican11031894-
287.pdf.

Anonymous:1894:HLFb

[5188] Anonymous. Herman Ludwig Ferdinand Von Helmholtz. Scientific American, 71(19):289–290, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
289.pdf.

Anonymous:1894:Ac

[5189] Anonymous.  Astronomical.  Scientific American, 71(19):290–

291, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
290.pdf.

Anonymous:1894:HN

[5190] Anonymous. The heavens in November. Scientific American, 71 (19):291, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
291.pdf.

Anonymous:1894:SCP

[5191] Anonymous. Small caliber projectiles. Scientific American, 71 (19):291, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
291a.pdf.

Anonymous:1894:BFBa

[5192] Anonymous. A burner to facilitate the burning of crude petroleum. Scientific American, 71(19):292, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
292b.pdf.

Anonymous:1894:CCS

[5193] Anonymous. A clutch to couple sections of shafting. Scientific American, 71(19):292, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
292.pdf.

Anonymous:1894:NPS

[5194] Anonymous. The new P. and O. steamer Caledonia. Scientific American, 71(19):292, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
292a.pdf.

Anonymous:1894:MMI

[5195] Anonymous. Monument to the memory of the inventor of the bicycle pedal. Scientific American, 71(19):293, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n19/pdf/scientificamerican11101894-293.pdf.

Anonymous:1894:TCM

[5196] Anonymous. A tobacco or cigar moistener. Scientific American, 71 (19):293–294, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
293a.pdf.

Anonymous:1894:Cbg

[5197] Anonymous. Correspondence. Scientific American, 71(19):295, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n19/pdf/scientificamerican11101894-295.pdf.

Anonymous:1894:SPI

[5198] Anonymous. Street piano industry. Scientific American, 71(19): 296–297, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
296.pdf.

Anonymous:1894:MR

[5199] Anonymous. A Mussuck race. Scientific American, 71(19):297, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n19/pdf/scientificamerican11101894-297.pdf.

Anonymous:1894:LSC

[5200] Anonymous. A locking stop cock for train pipes. Scientific American, 71(19):297–298, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
297a.pdf.

Anonymous:1894:IHa

[5201] Anonymous. An improvement in harness. Scientific American, 71 (19):298, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
298.pdf.

Anonymous:1894:SBC

[5202] Anonymous. A stone Baptist church. Scientific American, 71(19): 299–300, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
299.pdf.

Anonymous:1894:RPIbr

[5203] Anonymous. Recently patented inventions. Scientific American, 71(19):300, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
300.pdf.

Anonymous:1894:BNBbh

[5204] Anonymous. Books: New books and publications. Scientific American, 71(19):301, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
301.pdf.

Anonymous:1894:BPbr

[5205] Anonymous. Business and personal. Scientific American, 71 (19):301, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
301a.pdf.

Anonymous:1894:Iby

[5206] Anonymous. To inventors. Scientific American, 71(19):301, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n19/pdf/scientificamerican11101894-301c.pdf.

Anonymous:1894:NQbr

[5207] Anonymous. Notes and queries. Scientific American, 71(19): 301, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
301b.pdf.

Anonymous:1894:IIbr

[5208] Anonymous. Index of inventions. Scientific American, 71(19): 301–302, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
301d.pdf.

Anonymous:1894:PNPbq

[5209] Anonymous. Patent notes: Patents. Scientific American, 71 (19):303, November 10, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n19/pdf/scientificamerican11101894-
303.pdf.

Anonymous:1894:PN

[5210] Anonymous. The prevention of noise. Scientific American, 71(20): 306–308, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
306.pdf.

Anonymous:1894:NCDb

[5211] Anonymous. The new cure for diphtheria, croup, etc. Scientific American, 71(20):308–309, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
308.pdf.

Anonymous:1894:ILB

[5212] Anonymous. An improved locomotive boiler. Scientific American, 71(20):309, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
309.pdf.

Anonymous:1894:CWb

[5213] Anonymous. A carriage wrench. Scientific American, 71(20): 309–310, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
309a.pdf.

Anonymous:1894:CCa

[5214] Anonymous. A cycle cab. Scientific American, 71(20):310, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n20/pdf/scientificamerican11171894-310.pdf.

Anonymous:1894:SWI

[5215] Anonymous. A sprocket wheel improvement. Scientific American, 71 (20):310–311, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
310a.pdf.

Anonymous:1894:GES

[5216] Anonymous. A gas engine for small power. Scientific American, 71(20):312, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
312a.pdf.

Anonymous:1894:MBP

[5217] Anonymous. The manufacture of burglar proof vaults. Scientific American, 71(20):312, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
312.pdf.

Anonymous:1894:TAD

[5218] Anonymous. Two arm derrick — Chicago Drainage Canal. Scientific American, 71(20):313, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
313.pdf.

Anonymous:1894:PSD

[5219] Anonymous. Portable stone dressing machine of the American Pneumatic Tool Company. Scientific American, 71(20):313–314, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n20/pdf/scientificamerican11171894-313a.pdf.

Anonymous:1894:BAZ

[5220] Anonymous. The Beatrice antelope in the Zoological Garden in Berlin. Scientific American, 71(20):315–316, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n20/pdf/scientificamerican11171894-315.pdf.

Anonymous:1894:RPIbs

[5221] Anonymous. Recently patented inventions. Scientific American, 71(20):316, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
316.pdf.

Anonymous:1894:BNBbi

[5222] Anonymous. Books: New books and publications. Scientific American, 71(20):317, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
317.pdf.

Anonymous:1894:BPbs

[5223] Anonymous. Business and personal. Scientific American, 71 (20):317, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
317a.pdf.

Anonymous:1894:Ibz

[5224] Anonymous. To inventors. Scientific American, 71(20):317, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n20/pdf/scientificamerican11171894-317c.pdf.

Anonymous:1894:NQbs

[5225] Anonymous. Notes and queries. Scientific American, 71(20): 317, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
317b.pdf.

Anonymous:1894:IIbs

[5226] Anonymous. Index of inventions. Scientific American, 71(20): 317–318, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
317d.pdf.

Anonymous:1894:PNPbr

[5227] Anonymous. Patent notes: Patents. Scientific American, 71 (20):319, November 17, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n20/pdf/scientificamerican11171894-
319.pdf.

Anonymous:1894:NEV

[5228] Anonymous. The need of efficient vessels for the Navy. Scientific American, 71(21):322–323, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
322.pdf.

Anonymous:1894:FTS

[5229] Anonymous. The French traveling schools. Scientific American, 71(21):324, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
324.pdf.

Anonymous:1894:IMCb

[5230] Anonymous. An improved mechanical carrier. Scientific American, 71(21):325, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
325a.pdf.

Anonymous:1894:ISRb

[5231] Anonymous. An improved swing rail railroad frog. Scientific American, 71(21):325, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
325.pdf.

Anonymous:1894:IJH

[5232] Anonymous. An improved journal head for rolls. Scientific American, 71(21):325–326, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
325b.pdf.

Anonymous:1894:Cbh

[5233] Anonymous. Correspondence. Scientific American, 71(21):327, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n21/pdf/scientificamerican11241894-327a.pdf.

Anonymous:1894:CTS

[5234] Anonymous. The Columbia Telephone System — its factory and instruments. Scientific American, 71(21):327, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n21/pdf/scientificamerican11241894-327.pdf.

Anonymous:1894:LSL

[5235] Anonymous. The launch of the St. Louis of the American Line. Scientific American, 71(21):328–329, November 24, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n21/pdf/scientificamerican11241894-328.pdf.

Anonymous:1894:NDU

[5236] Anonymous. New dynamometer for use in anthropometry. Scientific American, 71(21):329, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
329.pdf.

Anonymous:1894:TEE

[5237] Anonymous. Triple expansion engine — Frikart’s system. Scientific American, 71(21):329–330, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
329a.pdf.

Anonymous:1894:LGP

[5238] Anonymous. The leading and glass painting industry. Scientific American, 71(21):331–332, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
331.pdf.

Anonymous:1894:BPbt

[5239] Anonymous. Business and personal. Scientific American, 71 (21):333, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
333b.pdf.

Anonymous:1894:NBP

[5240] Anonymous. New books and publications. Scientific American, 71 (21):333, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
333a.pdf.

Anonymous:1894:NQbt

[5241] Anonymous. Notes and queries. Scientific American, 71(21): 333, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
333c.pdf.

Anonymous:1894:RPIbt

[5242] Anonymous. Recently patented inventions. Scientific American, 71(21):333, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
333.pdf.

Anonymous:1894:IIbt

[5243] Anonymous. Index of inventions. Scientific American, 71(21): 334, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
334.pdf.

Anonymous:1894:PNPbs

[5244] Anonymous. Patent notes: Patents. Scientific American, 71 (21):335, November 24, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n21/pdf/scientificamerican11241894-
335.pdf.

Anonymous:1894:GHW

[5245] Anonymous. George Huntington Williams. Scientific American, 71(22):337, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
337a.pdf.

Anonymous:1894:PAIc

[5246] Anonymous. The Park Avenue improvement in New York. Scientific American, 71(22):337, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
337.pdf.

Anonymous:1894:HD

[5247] Anonymous. The heavens in December. Scientific American, 71 (22):338–339, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
338.pdf.

Anonymous:1894:RWC

[5248] Anonymous. Rolled weldsless chains — Klatte’s process. Scientific American, 71(22):340–341, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
340.pdf.

Anonymous:1894:ILc

[5249] Anonymous. An improved litter. Scientific American, 71(22):341, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n22/pdf/scientificamerican12011894-341a.pdf.

Anonymous:1894:TKS

[5250] Anonymous. A telegraph key and sounder. Scientific American, 71(22):341, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
341.pdf.

Anonymous:1894:Cbi

[5251] Anonymous. Correspondence. Scientific American, 71(22):342, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n22/pdf/scientificamerican12011894-342.pdf.

Anonymous:1894:AT

[5252] Anonymous. The American typewriter. Scientific American, 71 (22):343, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
343.pdf.

Anonymous:1894:LEC

[5253] Anonymous. The locks at the ends of the canal between the North and Baltic Seas. Scientific American, 71(22):345–346, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n22/pdf/scientificamerican12011894-345.pdf.

Anonymous:1894:CBj

[5254] Anonymous. A curious bicycle. Scientific American, 71(22):346, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n22/pdf/scientificamerican12011894-346a.pdf.

Anonymous:1894:ISK

[5255] Anonymous. An improved sleigh knee. Scientific American, 71 (22):346, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
346.pdf.

Anonymous:1894:SEM

[5256] Anonymous. Simple electric motor. Scientific American, 71(22): 346–347, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
346b.pdf.

Anonymous:1894:HL

[5257] Anonymous. The Holman locomotive. Scientific American, 71 (22):347–348, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
347.pdf.

Anonymous:1894:RPIbu

[5258] Anonymous. Recently patented inventions. Scientific American, 71 (22):348–349, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
348.pdf.

Anonymous:1894:BPbu

[5259] Anonymous. Business and personal. Scientific American, 71 (22):349, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
349.pdf.

Anonymous:1894:Ice

[5260] Anonymous. To inventors. Scientific American, 71(22):349, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n22/pdf/scientificamerican12011894-349b.pdf.

Anonymous:1894:NQbu

[5261] Anonymous. Notes and queries. Scientific American, 71(22):349, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n22/pdf/scientificamerican12011894-349a.pdf.

Anonymous:1894:IIbu

[5262] Anonymous. Index of inventions. Scientific American, 71(22): 349–350, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
349c.pdf.

Anonymous:1894:PNPbt

[5263] Anonymous. Patent notes: Patents. Scientific American, 71 (22):351, December 1, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n22/pdf/scientificamerican12011894-
351.pdf.

Anonymous:1894:SFT

[5264] Anonymous. Steel foundations of tall office buildings. Scientific American, 71(23):353, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
353.pdf.

Anonymous:1894:NBB

[5265] Anonymous. The new Brooklyn Bridge Stations. Scientific American, 71(23):354, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
354a.pdf.

Anonymous:1894:TBR

[5266] Anonymous. A torpedo boat that rates over thirty-three miles per hour. Scientific American, 71(23):354, December 8, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n23/pdf/scientificamerican12081894-354.pdf.

Anonymous:1894:SCB

[5267] Anonymous. The small caliber bullet in the East. Scientific American, 71(23):354–355, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
354b.pdf.

Anonymous:1894:EHP

[5268] Anonymous. Expiration of the hektograph patent. Scientific American, 71(23):355, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
355.pdf.

Anonymous:1894:CFG

[5269] Anonymous. The compass field glass. Scientific American, 71 (23):356, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
356.pdf.

Anonymous:1894:RBM

[5270] Anonymous. The Roman bridge of Mostar. Scientific American, 71 (23):356–357, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
356a.pdf.

Anonymous:1894:PCP

[5271] Anonymous. The plaster center piece industry. Scientific American, 71(23):357–358, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
357.pdf.

Anonymous:1894:Cbk

[5272] Anonymous. Correspondence. Scientific American, 71(23):358, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n23/pdf/scientificamerican12081894-358.pdf.

Anonymous:1894:NAS

[5273] Anonymous. The new Armory of the Seventy-First Regiment, New York City. Scientific American, 71(23):360–361, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n23/pdf/scientificamerican12081894-360.pdf.

Anonymous:1894:DETb

[5274] Anonymous. A double elm tree. Scientific American, 71(23):361, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n23/pdf/scientificamerican12081894-361.pdf.

Anonymous:1894:DJA

[5275] Anonymous. Death of James Anthony Froude. Scientific American, 71(23):361–362, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
361a.pdf.

Anonymous:1894:EPSb

[5276] Anonymous. The Eastern Power Station of the Brooklyn City Railroad Company. Scientific American, 71(23):363–364, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n23/pdf/scientificamerican12081894-363.pdf.

Anonymous:1894:RPIbv

[5277] Anonymous. Recently patented inventions. Scientific American, 71 (23):364–365, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
364.pdf.

Anonymous:1894:BNBbj

[5278] Anonymous. Books: New books and publications. Scientific American, 71(23):365, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
365.pdf.

Anonymous:1894:BPbv

[5279] Anonymous. Business and personal. Scientific American, 71 (23):365, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
365a.pdf.

Anonymous:1894:Icg

[5280] Anonymous. To inventors. Scientific American, 71(23):365, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n23/pdf/scientificamerican12081894-365c.pdf.

Anonymous:1894:NQbv

[5281] Anonymous. Notes and queries. Scientific American, 71(23):365, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n23/pdf/scientificamerican12081894-365b.pdf.

Anonymous:1894:IIbv

[5282] Anonymous. Index of inventions. Scientific American, 71(23): 365–366, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
365d.pdf.

Anonymous:1894:PNPbu

[5283] Anonymous. Patent notes: Patents. Scientific American, 71 (23):367, December 8, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n23/pdf/scientificamerican12081894-
367.pdf.

Anonymous:1894:GCR

[5284] Anonymous. Government control of railroads. Scientific American, 71(24):370, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
370b.pdf.

Anonymous:1894:IPS

[5285] Anonymous. An international postage stamp. Scientific American, 71(24):370, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
370a.pdf.

Anonymous:1894:WLS

[5286] Anonymous. Work of the life saving stations. Scientific American, 71(24):370, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
370.pdf.

Anonymous:1894:CWI

[5287] Anonymous. Cast and wrought iron for frame work of buildings. Scientific American, 71(24):370–371, December 15, 1894. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v71/
n24/pdf/scientificamerican12151894-370c.pdf.

Anonymous:1894:SMb

[5288] Anonymous. The snows of Mars. Scientific American, 71(24): 371, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
371.pdf.

Anonymous:1894:ICDa

[5289] Anonymous. An improved clothes drier. Scientific American, 71 (24):372, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
372a.pdf.

Anonymous:1894:IFW

[5290] Anonymous. An improved feed water heater. Scientific American, 71(24):372, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
372b.pdf.

Anonymous:1894:IRC

[5291] Anonymous. An improved rope clamp. Scientific American, 71 (24):372, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
372c.pdf.

Anonymous:1894:MBF

[5292] Anonymous. A multiple blasting fuse. Scientific American, 71 (24):372, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
372.pdf.

Anonymous:1894:TMT

[5293] Anonymous. Traveling military turrets. Scientific American, 71 (24):373–374, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
373.pdf.

Anonymous:1894:WSA

[5294] Anonymous. The war ship Atlanta and her magazine. Scientific American, 71(24):375, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
375.pdf.

Anonymous:1894:TNC

[5295] Anonymous. The tower of the new City Hall, Philadelphia, Pa — the loftiest statue in the world. Scientific American, 71(24):376, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n24/pdf/scientificamerican12151894-376.pdf.

Anonymous:1894:NAN

[5296] Anonymous. New aquarium for New York. Scientific American, 71 (24):377–379, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
377.pdf.

Anonymous:1894:VBI

[5297] Anonymous. Venetian or bent iron work. Scientific American, 71 (24):379, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
379.pdf.

Anonymous:1894:RPIbw

[5298] Anonymous. Recently patented inventions. Scientific American, 71(24):380, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
380.pdf.

Anonymous:1894:BNBbk

[5299] Anonymous. Books: New books and publications. Scientific American, 71(24):380–381, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
380a.pdf.

Anonymous:1894:BPbw

[5300] Anonymous. Business and personal. Scientific American, 71 (24):381, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
381.pdf.

Anonymous:1894:Ich

[5301] Anonymous. To inventors. Scientific American, 71(24):381, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n24/pdf/scientificamerican12151894-381b.pdf.

Anonymous:1894:NQbw

[5302] Anonymous. Notes and queries. Scientific American, 71(24): 381, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
381a.pdf.

Anonymous:1894:IIbw

[5303] Anonymous. Index of inventions. Scientific American, 71(24): 381–382, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
381c.pdf.

Anonymous:1894:PNPbv

[5304] Anonymous. Patent notes: Patents. Scientific American, 71 (24):383, December 15, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n24/pdf/scientificamerican12151894-
383.pdf.

Anonymous:1894:PPL

[5305] Anonymous. Prof. S. P. Langley’s flying machine. Scientific American, 71(25):386, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
386.pdf.

Anonymous:1894:HRBb

[5306] Anonymous. The Hudson River bridge. Scientific American, 71 (25):386–387, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
386a.pdf.

Anonymous:1894:CCV

[5307] Anonymous. A chimney cowl or ventilator. Scientific American, 71(25):388, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
388.pdf.

Anonymous:1894:TFG

[5308] Anonymous. A two foot gauge locomotive. Scientific American, 71(25):388, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
388a.pdf.

Anonymous:1894:PLCb

[5309] Anonymous. A pocket lamp and cigar lighter. Scientific American, 71(25):388–389, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
388b.pdf.

Anonymous:1894:ISBc

[5310] Anonymous. An improved steam boiler. Scientific American, 71 (25):389, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
389.pdf.

Anonymous:1894:SBB

[5311] Anonymous. A suspension bridge breaks down. Scientific American, 71(25):389, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
389a.pdf.

Anonymous:1894:Cbl

[5312] Anonymous. Correspondence. Scientific American, 71(25):390, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n25/pdf/scientificamerican12221894-390.pdf.

Anonymous:1894:SNY

[5313] Anonymous. The Society of the New York Hospital and its new House of Relief. Scientific American, 71(25):391–392, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n25/pdf/scientificamerican12221894-391.pdf.

Anonymous:1894:MLB

[5314] Anonymous. The Manhattan life building. Scientific American, 71 (25):392–395, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
392.pdf.

Anonymous:1894:CMP

[5315] Anonymous. A Chinese mode of punishment. Scientific American, 71(25):395, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
395.pdf.

Anonymous:1894:EBA

[5316] Anonymous. An electric burglar alarm for safes. Scientific American, 71(25):395, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
395a.pdf.

Anonymous:1894:RPIbx

[5317] Anonymous. Recently patented inventions. Scientific American, 71(25):396, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
396.pdf.

Anonymous:1894:BNBbl

[5318] Anonymous. Books: New books and publications. Scientific American, 71(25):396–397, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
396a.pdf.

Anonymous:1894:BPbx

[5319] Anonymous. Business and personal. Scientific American, 71 (25):397, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
397.pdf.

Anonymous:1894:Ici

[5320] Anonymous. To inventors. Scientific American, 71(25):397, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n25/pdf/scientificamerican12221894-397b.pdf.

Anonymous:1894:NQbx

[5321] Anonymous. Notes and queries. Scientific American, 71(25): 397, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
397a.pdf.

Anonymous:1894:IIbx

[5322] Anonymous. Index of inventions. Scientific American, 71(25): 397–398, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
397c.pdf.

Anonymous:1894:PNPbw

[5323] Anonymous. Patent notes: Patents. Scientific American, 71 (25):399, December 22, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n25/pdf/scientificamerican12221894-
399.pdf.

Anonymous:1894:CSY

[5324] Anonymous. The Cramp Ship Yards. Scientific American, 71 (26):401, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
401.pdf.

Anonymous:1894:DAB

[5325] Anonymous. Decision against the Bell Telephone monopoly. Scientific American, 71(26):402, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
402.pdf.

Anonymous:1894:HJ

[5326] Anonymous. The heavens in January. Scientific American, 71(26): 402–403, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
402a.pdf.

Anonymous:1894:IFA

[5327] Anonymous. An inexpensive fire alarm. Scientific American, 71 (26):404, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
404.pdf.

Anonymous:1894:ISC

[5328] Anonymous. An improved Sawmill carriage. Scientific American, 71(26):404, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
404a.pdf.

Anonymous:1894:PLT

[5329] Anonymous. A plug for leaky tubes in boilers. Scientific American, 71(26):404, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
404b.pdf.

Anonymous:1894:SPPb

[5330] Anonymous. The steam peanut and pop corn industry. Scientific American, 71(26):405, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
405.pdf.

Hofer:1894:RTS

[5331] E. Hofer. The Russian thistle — a scourge to agriculture. Scientific American, 71(26):406, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
406.pdf.

Anonymous:1894:Cbm

[5332] Anonymous. Correspondence. Scientific American, 71(26):407, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n26/pdf/scientificamerican12291894-407.pdf.

Hopkins:1894:ESS

[5333] M. Hopkins. Equatorial stand for small telescopes. Scientific American, 71(26):409–410, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
409.pdf.

Anonymous:1894:RPIby

[5334] Anonymous. Recently patented inventions. Scientific American, 71 (26):410–411, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
410.pdf.

Anonymous:1894:BNBbm

[5335] Anonymous. Books: New books and publications. Scientific American, 71(26):411, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
411.pdf.

Anonymous:1894:BPby

[5336] Anonymous. Business and personal. Scientific American, 71 (26):411, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
411a.pdf.

Anonymous:1894:Icj

[5337] Anonymous. To inventors. Scientific American, 71(26):411, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n26/pdf/scientificamerican12291894-411c.pdf.

Anonymous:1894:NQby (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
411b.pdf.

Anonymous:1894:IIby

[5339] Anonymous. Index of inventions. Scientific American, 71(26): 411–412, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
411d.pdf.

Anonymous:1894:Ick

[5340] Anonymous. Index. Scientific American, 71(26):413–414, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v71/n26/pdf/scientificamerican12291894-413.pdf.

Anonymous:1894:PNPbx

[5341] Anonymous. Patent notes: Patents. Scientific American, 71 (26):415, December 29, 1894. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v71/n26/pdf/scientificamerican12291894-
415.pdf.

Anonymous:1895:TBC

[5342] Anonymous. Torpedo boats for the cruiser Maine. Scientific American, 72(1):1–2, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n1/pdf/scientificamerican01051895-
1.pdf.

Anonymous:1895:DCP

[5343] Anonymous. Delayed cases in the Patent Office. Scientific American, 72(1):2, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n1/pdf/scientificamerican01051895-
2.pdf.

Anonymous:1895:DRPa

[5344] Anonymous. Decisions relating to patents. Scientific American, 72(1):3, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-3.pdf.

Anonymous:1895:CCa

[5345] Anonymous. The curiosities of clockwork. Scientific American, 72(1):4–5, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n1/pdf/scientificamerican01051895-
4.pdf.

Anonymous:1895:NCW

[5346] Anonymous. New Chemist’s wash bottle. Scientific American, 72(1):5, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-5.pdf.

Anonymous:1895:NHH

[5347] Anonymous. Nine hundred horse power compound engine. Scientific American, 72(1):5–7, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n1/pdf/scientificamerican01051895-
5a.pdf.

Anonymous:1895:USBa

[5348] Anonymous. The United States battle ship Indiana. Scientific American, 72(1):8, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n1/pdf/scientificamerican01051895-
8.pdf.

Anonymous:1895:IGG

[5349] Anonymous. Improved Gatling gun. Scientific American, 72(1):9–10, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-9.pdf.

Anonymous:1895:BMA

[5350] Anonymous. Brooklyn’s Memorial Arch. Scientific American, 72(1):11, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-11.pdf.

Anonymous:1895:BNBa 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n1/pdf/scientificamerican01051895-
12a.pdf.

Anonymous:1895:RPIa

[5352] Anonymous. Recently patented inventions. Scientific American, 72(1):12, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n1/pdf/scientificamerican01051895-
12.pdf.

Anonymous:1895:BPa

[5353] Anonymous. Business and personal. Scientific American, 72(1):13, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-13.pdf.

Anonymous:1895:Ia

[5354] Anonymous. To inventors. Scientific American, 72(1):13, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-13b.pdf.

Anonymous:1895:NQa

[5355] Anonymous. Notes and queries. Scientific American, 72(1):13, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-13a.pdf.

Anonymous:1895:IIa

[5356] Anonymous. Index of inventions. Scientific American, 72(1):13–14, January 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n1/pdf/scientificamerican01051895-13c.pdf.

Anonymous:1895:JHG

[5357] Anonymous. James H. Gridley. Scientific American, 72(2):18, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-18.pdf.

Anonymous:1895:PBa

[5358] Anonymous. Parallel boundaries. Scientific American, 72(2):18– 19, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
18a.pdf.

Anonymous:1895:DSSa

[5359] Anonymous. The “Diamond” shears and scissors sharpener. Scientific American, 72(2):20, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
20b.pdf.

Anonymous:1895:EFP

[5360] Anonymous. An efficient force pump. Scientific American, 72 (2):20, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
20a.pdf.

Anonymous:1895:IOC

[5361] Anonymous. An improved Ore crusher. Scientific American, 72 (2):20, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
20c.pdf.

Anonymous:1895:NMR

[5362] Anonymous. A new measuring and registering device. Scientific American, 72(2):20, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
20.pdf.

Anonymous:1895:NLL

[5363] Anonymous. A new log loader. Scientific American, 72(2):21, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-21.pdf.

Anonymous:1895:BYR

[5364] Anonymous. The Battle of the Yalu River. Scientific American, 72(2):21–23, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
21a.pdf.

Anonymous:1895:Ca

[5365] Anonymous. Correspondence. Scientific American, 72(2):23, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-23.pdf.

Anonymous:1895:AGF

[5366] Anonymous. The American Gas Furnace Company’s gas plant and furnaces. Scientific American, 72(2):24–25, January 12, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n2/pdf/scientificamerican01121895-24.pdf.

Anonymous:1895:GET

[5367] Anonymous. Gas engine tricycle. Scientific American, 72(2):25, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-25.pdf.

Anonymous:1895:HJP

[5368] Anonymous. Hydraulic jaw plate punch. Scientific American, 72 (2):25–26, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
25a.pdf.

Anonymous:1895:LPT

[5369] Anonymous. Life in the Pampas — a tug of war. Scientific American, 72(2):27–28, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
27.pdf.

Anonymous:1895:BNBb 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
29a.pdf.

Anonymous:1895:BPb

[5371] Anonymous. Business and personal. Scientific American, 72(2):29, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-29b.pdf.

Anonymous:1895:NQb

[5372] Anonymous. Notes and queries. Scientific American, 72(2):29, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-29c.pdf.

Anonymous:1895:RPIb

[5373] Anonymous. Recently patented inventions. Scientific American, 72(2):29, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n2/pdf/scientificamerican01121895-
29.pdf.

Anonymous:1895:Ib

[5374] Anonymous. To inventors. Scientific American, 72(2):30, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-30.pdf.

Anonymous:1895:IIb

[5375] Anonymous. Index of inventions. Scientific American, 72(2):30, January 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n2/pdf/scientificamerican01121895-30a.pdf.

Anonymous:1895:NWS

[5376] Anonymous. The new war ships Texas and Oregon. Scientific American, 72(3):33, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
33.pdf.

Anonymous:1895:SCG

[5377] Anonymous. Statistics concerning German trade marks. Scientific American, 72(3):34, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
34.pdf.

Anonymous:1895:CCb

[5378] Anonymous. On the choice of a career. Scientific American, 72 (3):34–36, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
34a.pdf.

Anonymous:1895:CCM

[5379] Anonymous. A combined crushing mill, amalgamator and Ore concentrator. Scientific American, 72(3):36, January 19, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n3/pdf/scientificamerican01191895-36.pdf.

Anonymous:1895:PGW

[5380] Anonymous. A pump gear for windmills. Scientific American, 72(3):36, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
36a.pdf.

Anonymous:1895:SCC

[5381] Anonymous. A sugar cane cultivator. Scientific American, 72 (3):36, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
36b.pdf.

Anonymous:1895:CIW

[5382] Anonymous. Curious and interesting watches. Scientific American, 72(3):37, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
37.pdf.

Anonymous:1895:TT

[5383] Anonymous. The typewriter telegraph. Scientific American, 72 (3):37–38, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
37a.pdf.

Anonymous:1895:Cb

[5384] Anonymous. Correspondence. Scientific American, 72(3):39, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n3/pdf/scientificamerican01191895-39.pdf.

Anonymous:1895:BHI

[5385] Anonymous. The bar and horseshoe iron industry. Scientific American, 72(3):41–43, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
41.pdf.

Anonymous:1895:CBa

[5386] Anonymous. A compact battery. Scientific American, 72(3):43, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n3/pdf/scientificamerican01191895-43a.pdf.

Anonymous:1895:PBC

[5387] Anonymous. Pneumatic berths and cushions in parlor cars. Scientific American, 72(3):43, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
43.pdf.

Anonymous:1895:RPIc

[5388] Anonymous. Recently patented inventions. Scientific American, 72(3):44, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
44.pdf.

Anonymous:1895:BNBc

[5389] Anonymous. Books: New books and publications. Scientific American, 72(3):44–45, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n3/pdf/scientificamerican01191895-
44a.pdf.

Anonymous:1895:BPc

[5390] Anonymous. Business and personal. Scientific American, 72(3):45, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n3/pdf/scientificamerican01191895-45.pdf.

Anonymous:1895:Ic

[5391] Anonymous. To inventors. Scientific American, 72(3):45, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n3/pdf/scientificamerican01191895-45b.pdf.

Anonymous:1895:NQc

[5392] Anonymous. Notes and queries. Scientific American, 72(3):45, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n3/pdf/scientificamerican01191895-45a.pdf.

Anonymous:1895:IIc

[5393] Anonymous. Index of inventions. Scientific American, 72(3):45–46, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n3/pdf/scientificamerican01191895-45c.pdf.

Anonymous:1895:PNPa

[5394] Anonymous. Patent notes: Patents. Scientific American, 72(3):47, January 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n3/pdf/scientificamerican01191895-47.pdf.

Anonymous:1895:AFG

[5395] Anonymous. American firearms in Germany. Scientific American, 72(4):50, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
50.pdf.

Anonymous:1895:SLP

[5396] Anonymous. The statutes of limitations in patent suits. Scientific American, 72(4):50–51, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
50a.pdf.

Anonymous:1895:IAL

[5397] Anonymous. Improved artificial limbs. Scientific American, 72 (4):52, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
52.pdf.

Anonymous:1895:IVL

[5398] Anonymous. An improved vault light. Scientific American, 72 (4):52, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
52b.pdf.

Anonymous:1895:M

[5399] Anonymous. The micrograph. Scientific American, 72(4):52, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n4/pdf/scientificamerican01261895-52a.pdf.

Anonymous:1895:NTS

[5400] Anonymous. The new telephone system of Paris. Scientific American, 72(4):53–55, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
53.pdf.

Anonymous:1895:Cd

[5401] Anonymous. Correspondence. Scientific American, 72(4):55, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n4/pdf/scientificamerican01261895-55.pdf.

Anonymous:1895:CPS

[5402] Anonymous. Combined punch and shears. Scientific American, 72(4):55, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
55a.pdf.

Anonymous:1895:EHI

[5403] Anonymous. The electroplating of the hulls of iron ships. Scientific American, 72(4):56, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
56.pdf.

Anonymous:1895:PPE

[5404] Anonymous. Photograph of the partial eclipse of the Moon, September 14, 1894. Scientific American, 72(4):57, January 26, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n4/pdf/scientificamerican01261895-57.pdf.

Anonymous:1895:RCW

[5405] Anonymous. Repairing Chinese war ships. Scientific American, 72(4):57, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
57a.pdf.

Anonymous:1895:BLCa

[5406] Anonymous. The Boulevard Lafayette, in the City of New York. Scientific American, 72(4):58–59, January 26, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n4/pdf/scientificamerican01261895-58.pdf.

Anonymous:1895:PBB

[5407] Anonymous. A pneumatic bicycle brake. Scientific American, 72(4):59, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
59.pdf.

Anonymous:1895:ERC

[5408] Anonymous. An electric railway conduit. Scientific American, 72 (4):59–60, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
59a.pdf.

Anonymous:1895:RPId

[5409] Anonymous. Recently patented inventions. Scientific American, 72(4):60–61, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
60.pdf.

Anonymous:1895:BNBd

[5410] Anonymous. Books: New books and publications. Scientific American, 72(4):61, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n4/pdf/scientificamerican01261895-
61.pdf.

Anonymous:1895:BPd

[5411] Anonymous. Business and personal. Scientific American, 72(4):61, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n4/pdf/scientificamerican01261895-61a.pdf.

Anonymous:1895:NQd

[5412] Anonymous. Notes and queries. Scientific American, 72(4):61–62, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n4/pdf/scientificamerican01261895-61b.pdf.

Anonymous:1895:Id

[5413] Anonymous. To inventors. Scientific American, 72(4):62, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n4/pdf/scientificamerican01261895-62.pdf.

Anonymous:1895:IId

[5414] Anonymous. Index of inventions. Scientific American, 72(4):62, January 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n4/pdf/scientificamerican01261895-62a.pdf.

Anonymous:1895:CCRa

[5415] Anonymous. The crossing of the Central Railroad of New Jersey and the Pennsylvania Railroad in Elizabeth, N.J. Scientific American, 72(5):65, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
65.pdf.

Anonymous:1895:NCS

[5416] Anonymous. The National Cycle Show at Madison Square Garden, New York. Scientific American, 72(5):66, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n5/pdf/scientificamerican02021895-66.pdf.

Anonymous:1895:CCc

[5417] Anonymous. On the choice of a career. Scientific American, 72 (5):66–67, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
66a.pdf.

Anonymous:1895:GBP

[5418] Anonymous. The Great Breslau Pump, Neisse, Silesia. Scientific American, 72(5):68, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
68.pdf.

Anonymous:1895:AVBa

[5419] Anonymous. An automatic vehicle brake. Scientific American, 72(5):69, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
69.pdf.

Anonymous:1895:LAS

[5420] Anonymous. A large angle sextant. Scientific American, 72 (5):69, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
69a.pdf.

Anonymous:1895:HBT

[5421] Anonymous. A hose bridge and tower. Scientific American, 72 (5):69–70, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
69b.pdf.

Anonymous:1895:IG

[5422] Anonymous. An improved grain bin. Scientific American, 72 (5):70–71, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
70.pdf.

Anonymous:1895:MCI

[5423] Anonymous. The metal ceiling industry. Scientific American, 72 (5):72–73, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
72.pdf.

Anonymous:1895:BHL

[5424] Anonymous. A bicycle holder and lock. Scientific American, 72(5):73, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
73a.pdf.

Anonymous:1895:CMa

[5425] Anonymous. Cannon magnets. Scientific American, 72(5):73, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n5/pdf/scientificamerican02021895-73.pdf.

Anonymous:1895:Ce

[5426] Anonymous. Correspondence. Scientific American, 72(5):74, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n5/pdf/scientificamerican02021895-74.pdf.

Anonymous:1895:WPN

[5427] Anonymous. A wonderful photographic nebula. Scientific American, 72(5):74–75, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
74a.pdf.

Anonymous:1895:CDM

[5428] Anonymous. Cactus dahlia, Mrs. Francis Fell. Scientific American, 72(5):75–76, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
75.pdf.

Anonymous:1895:RPIe

[5429] Anonymous. Recently patented inventions. Scientific American, 72(5):76, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
76.pdf.

Anonymous:1895:BNBe

[5430] Anonymous. Books: New books and publications. Scientific American, 72(5):76–77, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
76a.pdf.

Anonymous:1895:BPe

[5431] Anonymous. Business and personal. Scientific American, 72 (5):77, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
77.pdf.

Anonymous:1895:Ie

[5432] Anonymous. To inventors. Scientific American, 72(5):77, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n5/pdf/scientificamerican02021895-77b.pdf.

Anonymous:1895:NQe

[5433] Anonymous. Notes and queries. Scientific American, 72(5):77, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n5/pdf/scientificamerican02021895-77a.pdf.

Anonymous:1895:IIe

[5434] Anonymous. Index of inventions. Scientific American, 72(5): 77–79, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
77c.pdf.

Anonymous:1895:PNPb

[5435] Anonymous. Patent notes: Patents. Scientific American, 72 (5):79, February 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n5/pdf/scientificamerican02021895-
79.pdf.

Anonymous:1895:NBEa

[5436] Anonymous. The National Bicycle Exhibition in Madison Square Garden, New York. Scientific American, 72(6):81–87, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n6/pdf/scientificamerican02091895-81.pdf.

Anonymous:1895:LGD

[5437] Anonymous. Lessons of a great disaster. Scientific American, 72(6):82, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
82.pdf.

Anonymous:1895:HF

[5438] Anonymous. The heavens in February. Scientific American, 72 (6):82–83, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
82a.pdf.

Williams:1895:PPB

[5439] Llewelyn Williams. Pont-Y-Pridd Bridge, South Wales. Scientific American, 72(6):84, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
84.pdf.

Anonymous:1895:IBR

[5440] Anonymous. An improved bicycle Rim and tire. Scientific American, 72(6):84–85, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
84a.pdf.

Anonymous:1895:RIC

[5441] Anonymous. A revolving ice cake. Scientific American, 72(6):85, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n6/pdf/scientificamerican02091895-85.pdf.

Scripture:1895:MI

[5442] E. W. Scripture. On the measurement of imaginations. Scientific American, 72(6):85, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
85a.pdf.

Anonymous:1895:Cf

[5443] Anonymous. Correspondence. Scientific American, 72(6):87, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n6/pdf/scientificamerican02091895-87.pdf.

Anonymous:1895:NV

[5444] Anonymous. A new violin. Scientific American, 72(6):88, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n6/pdf/scientificamerican02091895-88.pdf.

Anonymous:1895:YOU

[5445] Anonymous. The Yerkes Observatory — University of Chicago. Scientific American, 72(6):88, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
88a.pdf.

Anonymous:1895:PIBa

[5446] Anonymous. A pagoda inclosed by a banyan tree. Scientific American, 72(6):89–90, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
89.pdf.

Anonymous:1895:DRPb

[5447] Anonymous. Decisions relating to patents. Scientific American, 72(6):90–91, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
90.pdf.

Anonymous:1895:GTM

[5448] Anonymous. The giant tree Mark Twain. Scientific American, 72(6):91–92, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
91.pdf.

Anonymous:1895:BNBf

[5449] Anonymous. Books: New books and publications. Scientific American, 72(6):93, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
93a.pdf.

Anonymous:1895:BPf

[5450] Anonymous. Business and personal. Scientific American, 72 (6):93, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
93b.pdf.

Anonymous:1895:RPIf

[5451] Anonymous. Recently patented inventions. Scientific American, 72(6):93, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
93.pdf.

Anonymous:1895:NQf

[5452] Anonymous. Notes and queries. Scientific American, 72(6):93– 94, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n6/pdf/scientificamerican02091895-
93c.pdf.

Anonymous:1895:If

[5453] Anonymous. To inventors. Scientific American, 72(6):94, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n6/pdf/scientificamerican02091895-94.pdf.

Anonymous:1895:IIf

[5454] Anonymous. Index of inventions. Scientific American, 72(6):94, February 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n6/pdf/scientificamerican02091895-94a.pdf.

Anonymous:1895:LLD

[5455] Anonymous. The latest long distance telephone transmitter. Scientific American, 72(7):97, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
97.pdf.

Anonymous:1895:ZWU

[5456] Anonymous. Zero weather over the United States. Scientific American, 72(7):98, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
98.pdf.

Anonymous:1895:NJP

[5457] Anonymous. A new jet-propelled steam lifeboat. Scientific American, 72(7):98–99, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
98a.pdf.

Anonymous:1895:CWR

[5458] Anonymous. Cast welding rail joints. Scientific American, 72 (7):100–102, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
100.pdf.

Udden:1895:FFC

[5459] J. A. Udden. Fossil frost cracks. Scientific American, 72(7): 102–103, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
102.pdf.

Anonymous:1895:CW

[5460] Anonymous. A convenient wrench. Scientific American, 72(7): 103, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
103.pdf.

Anonymous:1895:IBC

[5461] Anonymous. An improved brick cutting machine. Scientific American, 72(7):104, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
104.pdf.

Anonymous:1895:ICH

[5462] Anonymous. An improved corn harvester. Scientific American, 72(7):104, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
104a.pdf.

Anonymous:1895:IFS

[5463] Anonymous. An improved fender for street cars. Scientific American, 72(7):104, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
104b.pdf.

Anonymous:1895:WBJ

[5464] Anonymous. The war between Japan and China. Scientific American, 72(7):105–106, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
105.pdf.

Anonymous:1895:KB

[5465] Anonymous. Keating bicycle. Scientific American, 72(7):106, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n7/pdf/scientificamerican02161895-106a.pdf.

Anonymous:1895:NBF

[5466] Anonymous. A novel bicycle frame. Scientific American, 72 (7):106, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
106.pdf.

Anonymous:1895:UC

[5467] Anonymous. The U.S. cyclometer. Scientific American, 72(7): 106, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
106b.pdf.

Anonymous:1895:BT

[5468] Anonymous. The Blake transmitter. Scientific American, 72(7): 107–108, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
107.pdf.

Anonymous:1895:RPIg

[5469] Anonymous. Recently patented inventions. Scientific American, 72(7):108, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
108.pdf.

Anonymous:1895:BNBg

[5470] Anonymous. Books: New books and publications. Scientific American, 72(7):109, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
109a.pdf.

Anonymous:1895:BPg

[5471] Anonymous. Business and personal. Scientific American, 72(7): 109, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
109.pdf.

Anonymous:1895:NQg

[5472] Anonymous. Notes and queries. Scientific American, 72(7):109, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n7/pdf/scientificamerican02161895-109b.pdf.

Anonymous:1895:IIg

[5473] Anonymous. Index of inventions. Scientific American, 72(7): 109–110, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
109c.pdf.

Anonymous:1895:PNPc

[5474] Anonymous. Patent notes: Patents. Scientific American, 72 (7):111, February 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n7/pdf/scientificamerican02161895-
111.pdf.

Anonymous:1895:NBEb

[5475] Anonymous. The new bridge over the East River at Blackwell’s Island, New York. Scientific American, 72(8):113, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n8/pdf/scientificamerican02231895-113.pdf.

Anonymous:1895:CSC

[5476] Anonymous. The commercial synthesis of carbon and hydrogen. Scientific American, 72(8):114–115, February 23, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n8/pdf/scientificamerican02231895-114.pdf.

Anonymous:1895:AFL

[5477] Anonymous. An automatic feed lubricator. Scientific American, 72(8):116, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
116.pdf.

Anonymous:1895:AIW

[5478] Anonymous. An automatic instantaneous water heater. Scientific American, 72(8):116, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
116a.pdf.

Anonymous:1895:RBC

[5479] Anonymous. A revolving brush car fender. Scientific American, 72(8):116, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
116b.pdf.

Anonymous:1895:ARCa

[5480] Anonymous. An Ancient Roman city in Algeria. Scientific American, 72(8):117, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
117.pdf.

Anonymous:1895:IFE

[5481] Anonymous. An improved flue expander. Scientific American, 72 (8):117–119, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
117a.pdf.

Anonymous:1895:AT

[5482] Anonymous. The advertising tricycle. Scientific American, 72 (8):120, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
120.pdf.

Benjamin:1895:MCT

[5483] Marcus Benjamin. A mechanical color test. Scientific American, 72 (8):121–122, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
121.pdf.

Anonymous:1895:BW

[5484] Anonymous. The boring woodpecker. Scientific American, 72(8): 122–123, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
122.pdf.

Anonymous:1895:MC

[5485] Anonymous. The mysterious Clepsydra. Scientific American, 72 (8):123, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
123.pdf.

Anonymous:1895:BNBh

[5486] Anonymous. Books: New books and publications. Scientific American, 72(8):124, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
124a.pdf.

Anonymous:1895:RPIh

[5487] Anonymous. Recently patented inventions. Scientific American, 72(8):124, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
124.pdf.

Anonymous:1895:BPh

[5488] Anonymous. Business and personal. Scientific American, 72(8): 125, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
125.pdf.

Anonymous:1895:Ih

[5489] Anonymous. To inventors. Scientific American, 72(8):125, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n8/pdf/scientificamerican02231895-125b.pdf.

Anonymous:1895:NQh

[5490] Anonymous. Notes and queries. Scientific American, 72(8):125, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n8/pdf/scientificamerican02231895-125a.pdf.

Anonymous:1895:IIh

[5491] Anonymous. Index of inventions. Scientific American, 72(8): 125–126, February 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n8/pdf/scientificamerican02231895-
125c.pdf.

Anonymous:1895:CCd

[5492] Anonymous. The career of a chemist. Scientific American, 72(9):130, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-130a.pdf.

Anonymous:1895:RCP

[5493] Anonymous. Report of the Commissioner of Patents for 1894. Scientific American, 72(9):130, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
130.pdf.

Anonymous:1895:HMa

[5494] Anonymous. The heavens in March. Scientific American, 72(9):130–131, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-130b.pdf.

Anonymous:1895:FCS

[5495] Anonymous. A factory chimney struck by lightning. Scientific American, 72(9):132, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
132.pdf.

Anonymous:1895:GIP

[5496] Anonymous. A gage indicating pressure and temperature. Scientific American, 72(9):133, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
133a.pdf.

Anonymous:1895:PDT

[5497] Anonymous. Philbrook’s dental tool. Scientific American, 72(9):133, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-133.pdf.

Anonymous:1895:ECV

[5498] Anonymous. Electrically controlled valve gear for pumping stations. Scientific American, 72(9):133–134, March 2, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n9/pdf/scientificamerican03021895-133b.pdf.

Michels:1895:PVN

[5499] John Michels. A parasite very numerous in hogs. Scientific American, 72(9):134, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
134.pdf.

Anonymous:1895:TCN

[5500] Anonymous. The Teachers College of New York City. Scientific American, 72(9):135, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
135.pdf.

Anonymous:1895:PIBb

[5501] Anonymous. Philadelphia ice boats. Scientific American, 72(9): 136–137, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
136.pdf.

Anonymous:1895:CPa

[5502] Anonymous. Cloud photography. Scientific American, 72(9):137–139, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-137.pdf.

Anonymous:1895:FBZ

[5503] Anonymous. Fennecs in the Berlin Zoological Garden. Scientific American, 72(9):139–140, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
139.pdf.

Anonymous:1895:RPIi

[5504] Anonymous. Recently patented inventions. Scientific American, 72(9):140–141, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n9/pdf/scientificamerican03021895-
140.pdf.

Anonymous:1895:BPi

[5505] Anonymous. Business and personal. Scientific American, 72(9):141, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-141.pdf.

Anonymous:1895:Ij

[5506] Anonymous. To inventors. Scientific American, 72(9):141, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-141b.pdf.

Anonymous:1895:NQi

[5507] Anonymous. Notes and queries. Scientific American, 72(9):141, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-141a.pdf.

Anonymous:1895:IIi

[5508] Anonymous. Index of inventions. Scientific American, 72(9):141–142, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-141c.pdf.

Anonymous:1895:PNPd

[5509] Anonymous. Patent notes: Patents. Scientific American, 72(9):143, March 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n9/pdf/scientificamerican03021895-143.pdf.

Anonymous:1895:NUM

[5510] Anonymous. A new use for machine guns. Scientific American, 72(10):146, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
146.pdf.

Anonymous:1895:ATC

[5511] Anonymous. Accidents on trolley car lines. Scientific American, 72(10):146–147, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
146a.pdf.

Anonymous:1895:ISFa

[5512] Anonymous. An improved sash fastener. Scientific American, 72(10):148, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
148.pdf.

Anonymous:1895:NPP

[5513] Anonymous. A novel printing press. Scientific American, 72(10):148, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-148a.pdf.

Anonymous:1895:DLI

[5514] Anonymous. A distance level indicator. Scientific American, 72 (10):148–149, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
148b.pdf.

Anonymous:1895:CRSa

[5515] Anonymous. Characteristic railway station in Southern California. Scientific American, 72(10):149, March 9, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n10/pdf/scientificamerican03091895-149.pdf.

Anonymous:1895:IFR

[5516] Anonymous. An improved feedwater regulator. Scientific American, 72(10):149, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
149a.pdf.

Anonymous:1895:Cg

[5517] Anonymous. Correspondence. Scientific American, 72(10):150, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-150.pdf.

Anonymous:1895:WTL

[5518] Anonymous. Water-tight lining for the elastic gores of boots and shoes. Scientific American, 72(10):150, March 9, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n10/pdf/scientificamerican03091895-150a.pdf.

Anonymous:1895:APG

[5519] Anonymous. The American plate glass industry. Scientific American, 72(10):151, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
151.pdf.

Anonymous:1895:GSL

[5520] Anonymous. The granite and slate laundry tub industry. Scientific American, 72(10):152, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
152.pdf.

Anonymous:1895:LMC

[5521] Anonymous. The Lick Monument to California. Scientific American, 72(10):153, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
153.pdf.

Anonymous:1895:DRPc

[5522] Anonymous. Decisions relating to patents. Scientific American, 72(10):154, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
154.pdf.

Anonymous:1895:VB

[5523] Anonymous. The Victor bicycle. Scientific American, 72(10):155, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-155.pdf.

Anonymous:1895:WWM

[5524] Anonymous. The Weed water motor. Scientific American, 72 (10):155–156, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
155a.pdf.

Anonymous:1895:RPIj

[5525] Anonymous. Recently patented inventions. Scientific American, 72(10):156, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
156.pdf.

Anonymous:1895:BNBi

[5526] Anonymous. Books: New books and publications. Scientific American, 72(10):156–157, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n10/pdf/scientificamerican03091895-
156a.pdf.

Anonymous:1895:BPj

[5527] Anonymous. Business and personal. Scientific American, 72(10):157, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-157.pdf.

Anonymous:1895:NQj

[5528] Anonymous. Notes and queries. Scientific American, 72(10):157–158, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-157a.pdf.

Anonymous:1895:Ik

[5529] Anonymous. To inventors. Scientific American, 72(10):158, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-158.pdf.

Anonymous:1895:IIj

[5530] Anonymous. Index of inventions. Scientific American, 72(10):158–160, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-158a.pdf.

Anonymous:1895:PNPe

[5531] Anonymous. Patent notes: Patents. Scientific American, 72(10):160, March 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n10/pdf/scientificamerican03091895-160.pdf.

Anonymous:1895:IPD

[5532] Anonymous. An important patent decision by the Supreme Court. Scientific American, 72(11):162–163, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
162.pdf.

Anonymous:1895:ACC

[5533] Anonymous. An automatic car coupling. Scientific American, 72(11):164, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
164b.pdf.

Anonymous:1895:CAC

[5534] Anonymous. A conveniently adjustable chair. Scientific American, 72(11):164, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
164c.pdf.

Anonymous:1895:ISV

[5535] Anonymous. An improved slide valve. Scientific American, 72(11):164, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n11/pdf/scientificamerican03161895-164.pdf.

Anonymous:1895:MHS

[5536] Anonymous. A machine to harvest standing corn. Scientific American, 72(11):164, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
164a.pdf.

Anonymous:1895:NBC

[5537] Anonymous. A new bolt cutter and nut tapper. Scientific American, 72(11):165, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
165.pdf.

Anonymous:1895:CAEa

[5538] Anonymous. The coming Atlanta Exposition. Scientific American, 72(11):165–166, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
165a.pdf.

Anonymous:1895:NLG

[5539] Anonymous. The naphtha launch of the Gas Engine and Power Company of New York City. Scientific American, 72(11):167, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n11/pdf/scientificamerican03161895-167.pdf.

Anonymous:1895:Ch

[5540] Anonymous. Correspondence. Scientific American, 72(11):167–168, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n11/pdf/scientificamerican03161895-167a.pdf.

Anonymous:1895:PGS

[5541] Anonymous. The proposed great suspension bridge over the Hudson River, New York City. Scientific American, 72(11):169, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n11/pdf/scientificamerican03161895-169.pdf.

Anonymous:1895:BCa

[5542] Anonymous. The birds of the Congo. Scientific American, 72 (11):169–170, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
169a.pdf.

Anonymous:1895:APO

[5543] Anonymous. The auto-phoro-optometer. Scientific American, 72 (11):171, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
171.pdf.

Anonymous:1895:IIk

[5544] Anonymous. The International injector. Scientific American, 72 (11):171–172, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
171a.pdf.

Anonymous:1895:BNBj

[5545] Anonymous. Books: New books and publications. Scientific American, 72(11):173, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
173a.pdf.

Anonymous:1895:BPk

[5546] Anonymous. Business and personal. Scientific American, 72(11):173, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n11/pdf/scientificamerican03161895-173b.pdf.

Anonymous:1895:NQk

[5547] Anonymous. Notes and queries. Scientific American, 72(11):173, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n11/pdf/scientificamerican03161895-173c.pdf.

Anonymous:1895:RPIk

[5548] Anonymous. Recently patented inventions. Scientific American, 72(11):173, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n11/pdf/scientificamerican03161895-
173.pdf.

Anonymous:1895:IIl

[5549] Anonymous. Index of inventions. Scientific American, 72(11):173–174, March 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n11/pdf/scientificamerican03161895-173d.pdf.

Anonymous:1895:EC

[5550] Anonymous. An electric carriage. Scientific American, 72(12):177, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-177a.pdf.

Anonymous:1895:NSTa

[5551] Anonymous. A new steam tricycle. Scientific American, 72(12):177, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-177.pdf.

Anonymous:1895:NAT

[5552] Anonymous. The Niagara Aerial Tramway. Scientific American, 72(12):178, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
178.pdf.

Anonymous:1895:A

[5553] Anonymous. Acetylene. Scientific American, 72(12):178–179, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-178a.pdf.

Anonymous:1895:BRA

[5554] Anonymous. Bristol’s recording Ampère meter. Scientific American, 72(12):180, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
180a.pdf.

Anonymous:1895:CVRa

[5555] Anonymous. A churn with vertical and rotary dashers. Scientific American, 72(12):180, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
180b.pdf.

Anonymous:1895:MOE

[5556] Anonymous. A motor operated by the expansion and contraction of metal. Scientific American, 72(12):180, March 23, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n12/pdf/scientificamerican03231895-180.pdf.

Anonymous:1895:CES

[5557] Anonymous. A compact electric signaling device. Scientific American, 72(12):181, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
181.pdf.

Anonymous:1895:FED

[5558] Anonymous. A finger exercising device. Scientific American, 72(12):181, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-181a.pdf.

Anonymous:1895:MWV

[5559] Anonymous. Manufacture of White vinegar. Scientific American, 72(12):181–182, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
181b.pdf.

Anonymous:1895:ASS

[5560] Anonymous. An adjustable storm and Sun hood for carriages. Scientific American, 72(12):182, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
182.pdf.

Anonymous:1895:IO

[5561] Anonymous. An improved oarlock. Scientific American, 72(12):182, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-182a.pdf.

Anonymous:1895:Ci

[5562] Anonymous. Correspondence. Scientific American, 72(12):182–184, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-182b.pdf.

Anonymous:1895:ICN

[5563] Anonymous. The icy coasts of North America. Scientific American, 72(12):184, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
184.pdf.

Anonymous:1895:ITL

[5564] Anonymous. Inventor Tesla’s loss. Scientific American, 72(12):185–186, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-185.pdf.

Anonymous:1895:VDS

[5565] Anonymous. Vertical double spindle hollow chisel mortising machine. Scientific American, 72(12):187–188, March 23, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n12/pdf/scientificamerican03231895-187.pdf.

Anonymous:1895:BNBk

[5566] Anonymous. Books: New books and publications. Scientific American, 72(12):189, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
189b.pdf.

Anonymous:1895:BPl

[5567] Anonymous. Business and personal. Scientific American, 72(12):189, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-189a.pdf.

Anonymous:1895:NQl

[5568] Anonymous. Notes and queries. Scientific American, 72(12):189, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-189c.pdf.

Anonymous:1895:RPIl

[5569] Anonymous. Recently patented inventions. Scientific American, 72(12):189, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n12/pdf/scientificamerican03231895-
189.pdf.

Anonymous:1895:IIm

[5570] Anonymous. Index of inventions. Scientific American, 72(12):189–190, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-189d.pdf.

Anonymous:1895:PNPf

[5571] Anonymous. Patent notes: Patents. Scientific American, 72(12):191, March 23, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n12/pdf/scientificamerican03231895-191.pdf.

Anonymous:1895:MPb

[5572] Anonymous. The manufacture of porcelain. Scientific American, 72(13):193–194, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
193.pdf.

Anonymous:1895:ETEa

[5573] Anonymous. The electric transmission of energy. Scientific American, 72(13):194–195, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
194.pdf.

Anonymous:1895:CSB

[5574] Anonymous. A convenient square and bevel. Scientific American, 72(13):196, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
196.pdf.

Anonymous:1895:IVP

[5575] Anonymous. An improved veneer press. Scientific American, 72(13):196, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
196b.pdf.

Anonymous:1895:OCL

[5576] Anonymous. Outfit for casting light metals. Scientific American, 72(13):196, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
196a.pdf.

Anonymous:1895:DPH

[5577] Anonymous. Dr. P. H. Vander Weyde. Scientific American, 72(13):197, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n13/pdf/scientificamerican03301895-197.pdf.

Anonymous:1895:FEa

[5578] Anonymous. The French Exposition of 1900. Scientific American, 72(13):197–198, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
197a.pdf.

Hopkins:1895:SCS

[5579] George M. Hopkins. A simple camera shutter. Scientific American, 72(13):198–199, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
198.pdf.

Anonymous:1895:SNU

[5580] Anonymous. Ships of the new United States Navy. Scientific American, 72(13):199–202, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
199.pdf.

Sloane:1895:AA

[5581] T. O’Conor Sloane. Acetylene apparatus. Scientific American, 72(13):202, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
202.pdf.

Anonymous:1895:LPR

[5582] Anonymous. The Loeb patent respirator. Scientific American, 72 (13):203–204, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
203.pdf.

Anonymous:1895:RPIm

[5583] Anonymous. Recently patented inventions. Scientific American, 72 (13):204–205, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
204.pdf.

Anonymous:1895:BNBl

[5584] Anonymous. Books: New books and publications. Scientific American, 72(13):205, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n13/pdf/scientificamerican03301895-
205.pdf.

Anonymous:1895:BPm

[5585] Anonymous. Business and personal. Scientific American, 72(13):205, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n13/pdf/scientificamerican03301895-205a.pdf.

Anonymous:1895:Il

[5586] Anonymous. To inventors. Scientific American, 72(13):205, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n13/pdf/scientificamerican03301895-205c.pdf.

Anonymous:1895:NQm

[5587] Anonymous. Notes and queries. Scientific American, 72(13):205, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n13/pdf/scientificamerican03301895-205b.pdf.

Anonymous:1895:IIn

[5588] Anonymous. Index of inventions. Scientific American, 72(13):205–206, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n13/pdf/scientificamerican03301895-205d.pdf.

Anonymous:1895:PNPg

[5589] Anonymous. Patent notes: Patents. Scientific American, 72(13):207, March 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n13/pdf/scientificamerican03301895-207.pdf.

Anonymous:1895:PTa

[5590] Anonymous. Patent trusts. Scientific American, 72(14):210, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-210.pdf.

Anonymous:1895:HAa

[5591] Anonymous. The heavens in April. Scientific American, 72(14):210–211, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-210a.pdf.

Anonymous:1895:CCRb

[5592] Anonymous. The career of a chemist — a rejoinder. Scientific American, 72(14):211, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
211.pdf.

Anonymous:1895:CDH

[5593] Anonymous. A compound or double horseshoe. Scientific American, 72(14):212, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
212.pdf.

Anonymous:1895:IWT

[5594] Anonymous. An improved wagon tongue support. Scientific American, 72(14):212, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
212b.pdf.

Anonymous:1895:TJA

[5595] Anonymous. A twin-jaw, automatic car coupler. Scientific American, 72(14):212, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
212a.pdf.

Anonymous:1895:ICC

[5596] Anonymous. An improved car coupling. Scientific American, 72 (14):212–213, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
212c.pdf.

Anonymous:1895:AH

[5597] Anonymous. An attractive house. Scientific American, 72(14):213, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-213.pdf.

Anonymous:1895:JEC

[5598] Anonymous. Jeantaud’s electric carriage. Scientific American, 72 (14):213–214, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
213a.pdf.

Anonymous:1895:AEA

[5599] Anonymous. The Amstutz electro-artograph. Scientific American, 72(14):215, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
215.pdf.

Anonymous:1895:ONG

[5600] Anonymous. An Ohio natural gas well. Scientific American, 72(14):216, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-216.pdf.

Anonymous:1895:MLP

[5601] Anonymous. Manufacture of lead pipe. Scientific American, 72 (14):217–218, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
217.pdf.

Anonymous:1895:MT

[5602] Anonymous. A magnesium torch. Scientific American, 72(14):218, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-218.pdf.

Anonymous:1895:AI

[5603] Anonymous. The Alpine indicator. Scientific American, 72(14):219, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-219.pdf.

Anonymous:1895:FEb

[5604] Anonymous. Fire eaters. Scientific American, 72(14):219–220, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-219a.pdf.

Anonymous:1895:RPIn

[5605] Anonymous. Recently patented inventions. Scientific American, 72(14):220–221, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n14/pdf/scientificamerican04061895-
220.pdf.

Anonymous:1895:BPn

[5606] Anonymous. Business and personal. Scientific American, 72(14):221, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-221.pdf.

Anonymous:1895:Im

[5607] Anonymous. To inventors. Scientific American, 72(14):221, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-221b.pdf.

Anonymous:1895:NQn

[5608] Anonymous. Notes and queries. Scientific American, 72(14):221, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-221a.pdf.

Anonymous:1895:IIo

[5609] Anonymous. Index of inventions. Scientific American, 72(14):221–222, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-221c.pdf.

Anonymous:1895:PNPh

[5610] Anonymous. Patent notes: Patents. Scientific American, 72(14):223, April 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n14/pdf/scientificamerican04061895-223.pdf.

Anonymous:1895:STO

[5611] Anonymous. The Stern Tunnels of the Olympia. Scientific American, 72(15):226–227, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n15/pdf/scientificamerican04131895-
226.pdf.

Anonymous:1895:CSD

[5612] Anonymous. California Sun dried prunes. Scientific American, 72(15):228, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n15/pdf/scientificamerican04131895-
228a.pdf.

Anonymous:1895:IIW

[5613] Anonymous. An improved ink well. Scientific American, 72(15):228, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-228.pdf.

Anonymous:1895:RTC

[5614] Anonymous. A reloading tool for cartridges. Scientific American, 72(15):228, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n15/pdf/scientificamerican04131895-
228b.pdf.

Anonymous:1895:IP

[5615] Anonymous. The ice plane. Scientific American, 72(15):229, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-229.pdf.

Anonymous:1895:RH

[5616] Anonymous. Royal E. House. Scientific American, 72(15):229–230, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-229a.pdf.

Anonymous:1895:MIE

[5617] Anonymous. The manufacture of incandescent electric lamps. Scientific American, 72(15):230, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n15/pdf/scientificamerican04131895-
230.pdf.

Anonymous:1895:Cj

[5618] Anonymous. Correspondence. Scientific American, 72(15):231, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-231.pdf.

Anonymous:1895:PCI

[5619] Anonymous. The pulverized charcoal industry. Scientific American, 72(15):232, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n15/pdf/scientificamerican04131895-
232.pdf.

Osborn:1895:VEN

[5620] H. F. Osborn. A visit to the exhibit of the New York Academy of Sciences. Scientific American, 72(15):233–235, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n15/pdf/scientificamerican04131895-233.pdf.

Anonymous:1895:CEM

[5621] Anonymous. A combination electrical meter. Scientific American, 72(15):235, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n15/pdf/scientificamerican04131895-
235.pdf.

Anonymous:1895:BPo

[5622] Anonymous. Business and personal. Scientific American, 72(15):236, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-236a.pdf.

Anonymous:1895:RPIo

[5623] Anonymous. Recently patented inventions. Scientific American, 72(15):236, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n15/pdf/scientificamerican04131895-
236.pdf.

Anonymous:1895:NQo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-236b.pdf.

Anonymous:1895:In

[5625] Anonymous. To inventors. Scientific American, 72(15):237, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-237.pdf.

Anonymous:1895:IIp

[5626] Anonymous. Index of inventions. Scientific American, 72(15):237–238, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-237a.pdf.

Anonymous:1895:PNPi

[5627] Anonymous. Patent notes: Patents. Scientific American, 72(15):239, April 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n15/pdf/scientificamerican04131895-239.pdf.

Anonymous:1895:DM

[5628] Anonymous. The Deseret Museum. Scientific American, 72(16):241, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-241.pdf.

Anonymous:1895:AV

[5629] Anonymous. Animal vaccination. Scientific American, 72(16):242–244, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-242.pdf.

Anonymous:1895:CEP

[5630] Anonymous. A cushioned ear piece for the telephone receiver. Scientific American, 72(16):244, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n16/pdf/scientificamerican04201895-
244.pdf.

Anonymous:1895:NC

[5631] Anonymous. A new chuck. Scientific American, 72(16):244, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-244a.pdf.

Anonymous:1895:SWT

[5632] Anonymous. A sectional water tube boiler. Scientific American, 72(16):244–245, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n16/pdf/scientificamerican04201895-
244b.pdf.

Anonymous:1895:LCB

[5633] Anonymous. A letter copying book with which a press is not needed. Scientific American, 72(16):245, April 20, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n16/pdf/scientificamerican04201895-245a.pdf.

Anonymous:1895:SRL

[5634] Anonymous. The Sumatra rhinoceros in the Leipzig Zoological Garden. Scientific American, 72(16):245, April 20, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n16/pdf/scientificamerican04201895-245.pdf.

Anonymous:1895:Ck

[5635] Anonymous. Correspondence. Scientific American, 72(16):246, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-246.pdf.

Anonymous:1895:REa

[5636] Anonymous. A remarkable explosion. Scientific American, 72(16):248, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-248.pdf.

Anonymous:1895:NASa

[5637] Anonymous. The new American steamer St. Paul. Scientific American, 72(16):249–250, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n16/pdf/scientificamerican04201895-
249.pdf.

Anonymous:1895:IMa

[5638] Anonymous. An improved map. Scientific American, 72(16):251–252, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-251.pdf.

Anonymous:1895:RPIp

[5639] Anonymous. Recently patented inventions. Scientific American, 72(16):252–253, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n16/pdf/scientificamerican04201895-
252.pdf.

Anonymous:1895:BPp

[5640] Anonymous. Business and personal. Scientific American, 72(16):253, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-253.pdf.

Anonymous:1895:Iq

[5641] Anonymous. To inventors. Scientific American, 72(16):253, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-253b.pdf.

Anonymous:1895:NQp

[5642] Anonymous. Notes and queries. Scientific American, 72(16):253, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-253a.pdf.

Anonymous:1895:IIq

[5643] Anonymous. Index of inventions. Scientific American, 72(16):253–254, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-253c.pdf.

Anonymous:1895:PNPj

[5644] Anonymous. Patent notes: Patents. Scientific American, 72(16):255, April 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n16/pdf/scientificamerican04201895-255.pdf.

Anonymous:1895:PGI

[5645] Anonymous. Petroleum, the great illuminant. Scientific American, 72(17):258–259, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
258.pdf.

Anonymous:1895:ATD

[5646] Anonymous. An automatic time dater. Scientific American, 72(17):260, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n17/pdf/scientificamerican04271895-260a.pdf.

Anonymous:1895:PCT

[5647] Anonymous. The Prentiss calendar and time system. Scientific American, 72(17):260, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
260.pdf.

Anonymous:1895:WSP

[5648] Anonymous. A Watchmaker’s staking pliers. Scientific American, 72(17):260, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
260b.pdf.

Anonymous:1895:ICF

[5649] Anonymous. An improved car fender and brake. Scientific American, 72(17):261, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
261.pdf.

Anonymous:1895:IKM

[5650] Anonymous. Improved knitting machinery. Scientific American, 72(17):261, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
261a.pdf.

Anonymous:1895:ICB

[5651] Anonymous. An improved car brake. Scientific American, 72 (17):261–262, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
261b.pdf.

Anonymous:1895:NLC

[5652] Anonymous. The Nafew–Lovell check protector. Scientific American, 72(17):262, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
262.pdf.

Anonymous:1895:Cl

[5653] Anonymous. Correspondence. Scientific American, 72(17):263, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n17/pdf/scientificamerican04271895-263.pdf.

Anonymous:1895:WSM

[5654] Anonymous. The West Side Metropolitan Elevated Railway System of Chicago. Scientific American, 72(17):264, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n17/pdf/scientificamerican04271895-264.pdf.

Anonymous:1895:SIA

[5655] Anonymous. Some interesting animals. Scientific American, 72 (17):265–266, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
265.pdf.

Anonymous:1895:JDD

[5656] Anonymous. James D. Dana. Scientific American, 72(17):267–268, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n17/pdf/scientificamerican04271895-267.pdf.

Anonymous:1895:BNBm 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
268a.pdf.

Anonymous:1895:RPIq

[5658] Anonymous. Recently patented inventions. Scientific American, 72(17):268, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n17/pdf/scientificamerican04271895-
268.pdf.

Anonymous:1895:BPq

[5659] Anonymous. Business and personal. Scientific American, 72(17):269, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n17/pdf/scientificamerican04271895-269.pdf.

Anonymous:1895:NQq

[5660] Anonymous. Notes and queries. Scientific American, 72(17):269, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n17/pdf/scientificamerican04271895-269a.pdf.

Anonymous:1895:IIr

[5661] Anonymous. Index of inventions. Scientific American, 72(17):269–270, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n17/pdf/scientificamerican04271895-269b.pdf.

Anonymous:1895:PNPk

[5662] Anonymous. Patent notes: Patents. Scientific American, 72(17):271, April 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n17/pdf/scientificamerican04271895-271.pdf.

Anonymous:1895:ASE

[5663] Anonymous. The advanced system of education. Scientific American, 72(18):274–275, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
274.pdf.

Anonymous:1895:PES

[5664] Anonymous. The Prentiss electric synchronizer. Scientific American, 72(18):276, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
276.pdf.

Anonymous:1895:PRJ

[5665] Anonymous. The proposed railway up the Jungfrau. Scientific American, 72(18):276, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
276a.pdf.

Anonymous:1895:PSBa

[5666] Anonymous. Portable sand blast apparatus. Scientific American, 72(18):277, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
277.pdf.

Anonymous:1895:SSCa

[5667] Anonymous. The Scotte steam carriage. Scientific American, 72 (18):277–278, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
277a.pdf.

Anonymous:1895:Cn

[5668] Anonymous. Correspondence. Scientific American, 72(18):279, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n18/pdf/scientificamerican05041895-279.pdf.

Anonymous:1895:NYH

[5669] Anonymous. The New York Herald establishment. Scientific American, 72(18):280–283, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
280.pdf.

Anonymous:1895:BNBn

[5670] Anonymous. Books: New books and publications. Scientific American, 72(18):284, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
284b.pdf.

Anonymous:1895:BPr

[5671] Anonymous. Business and personal. Scientific American, 72(18):284, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n18/pdf/scientificamerican05041895-284a.pdf.

Anonymous:1895:RPIr

[5672] Anonymous. Recently patented inventions. Scientific American, 72(18):284, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n18/pdf/scientificamerican05041895-
284.pdf.

Anonymous:1895:IIs

[5673] Anonymous. Index of inventions. Scientific American, 72(18):285, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n18/pdf/scientificamerican05041895-285a.pdf.

Anonymous:1895:NQr

[5674] Anonymous. Notes and queries. Scientific American, 72(18):285, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n18/pdf/scientificamerican05041895-285.pdf.

Anonymous:1895:PNPl

[5675] Anonymous. Patent notes: Patents. Scientific American, 72(18):287, May 4, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n18/pdf/scientificamerican05041895-287.pdf.

Anonymous:1895:FEc

[5676] Anonymous. Force and energy. Scientific American, 72(19):290, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n19/pdf/scientificamerican05111895-290.pdf.

Anonymous:1895:HMb

[5677] Anonymous. The heavens in May. Scientific American, 72(19):290–291, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n19/pdf/scientificamerican05111895-290a.pdf.

Anonymous:1895:NASb

[5678] Anonymous. The National Academy of Sciences. Scientific American, 72(19):291, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
291.pdf.

Anonymous:1895:PUF

[5679] Anonymous. Purdue University, La Fayette, Indiana. Scientific American, 72(19):292, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
292.pdf.

Anonymous:1895:IWM

[5680] Anonymous. An improved water motor. Scientific American, 72 (19):293, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
293.pdf.

Anonymous:1895:HMN

[5681] Anonymous. The Hotel Majestic, New York. Scientific American, 72(19):293–294, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
293a.pdf.

Anonymous:1895:IGT

[5682] Anonymous. An improved gasoline traction engine. Scientific American, 72(19):294–295, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
294.pdf.

Anonymous:1895:IPT

[5683] Anonymous. Instrument for placing torpedoes. Scientific American, 72(19):295, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
295.pdf.

Anonymous:1895:VBS

[5684] Anonymous. The Van Buren Street Drawbridge of the Metropolitan West Side Elevated Railroad of Chicago. Scientific American, 72(19):296–298, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
296.pdf.

Anonymous:1895:WTG

[5685] Anonymous. The white-tailed gnu in the Hanover Zoological Garden. Scientific American, 72(19):299, May 11, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n19/pdf/scientificamerican05111895-299.pdf.

Anonymous:1895:BNBo

[5686] Anonymous. Books: New books and publications. Scientific American, 72(19):300, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
300a.pdf.

Anonymous:1895:BPs

[5687] Anonymous. Business and personal. Scientific American, 72(19):300, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n19/pdf/scientificamerican05111895-300b.pdf.

Anonymous:1895:RPIs

[5688] Anonymous. Recently patented inventions. Scientific American, 72(19):300, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n19/pdf/scientificamerican05111895-
300.pdf.

Anonymous:1895:NQs

[5689] Anonymous. Notes and queries. Scientific American, 72(19):300–301, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n19/pdf/scientificamerican05111895-300c.pdf.

Anonymous:1895:Ir

[5690] Anonymous. To inventors. Scientific American, 72(19):301, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n19/pdf/scientificamerican05111895-301.pdf.

Anonymous:1895:IIt

[5691] Anonymous. Index of inventions. Scientific American, 72(19):301–302, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n19/pdf/scientificamerican05111895-301a.pdf.

Anonymous:1895:PNPm

[5692] Anonymous. Patent notes: Patents. Scientific American, 72(19):303, May 11, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n19/pdf/scientificamerican05111895-303.pdf.

Anonymous:1895:GIW

[5693] Anonymous. A great invention in weaving. Scientific American, 72(20):306, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
306.pdf.

Anonymous:1895:DV

[5694] Anonymous. The deserted village. Scientific American, 72(20):306–307, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n20/pdf/scientificamerican05181895-306a.pdf.

Anonymous:1895:CFP

[5695] Anonymous. A chimney flue pipe opening cover. Scientific American, 72(20):308, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
308c.pdf.

Anonymous:1895:IHP

[5696] Anonymous. An improved horse power. Scientific American, 72(20):308, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n20/pdf/scientificamerican05181895-308a.pdf.

Anonymous:1895:IML

[5697] Anonymous. An improved Miner’s lamp. Scientific American, 72 (20):308, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
308.pdf.

Anonymous:1895:TPR

[5698] Anonymous. A toilet powder receptacle. Scientific American, 72 (20):308, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
308b.pdf.

Anonymous:1895:CBW

[5699] Anonymous. The “Climax” bicycle watch and holder. Scientific American, 72(20):309, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
309.pdf.

Anonymous:1895:AS

[5700] Anonymous. An artificial spectrum. Scientific American, 72 (20):309–310, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
309a.pdf.

Anonymous:1895:CLF

[5701] Anonymous. A cannon to be loaded with fire crackers. Scientific American, 72(20):310, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
310.pdf.

Anonymous:1895:MGE

[5702] Anonymous. The marine gasoline engine and gas engine boat. Scientific American, 72(20):311, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
311.pdf.

Anonymous:1895:PPV

[5703] Anonymous. A power pump with variable capacity. Scientific American, 72(20):312, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
312.pdf.

Anonymous:1895:LPS

[5704] Anonymous. The Layman pneumatic sporting and outing boat. Scientific American, 72(20):313–314, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
313.pdf.

Anonymous:1895:SAT

[5705] Anonymous. The Serpollet automobile tramway in Paris. Scientific American, 72(20):315–316, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
315.pdf.

Anonymous:1895:RPIt

[5706] Anonymous. Recently patented inventions. Scientific American, 72(20):316–317, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
316.pdf.

Anonymous:1895:BNBp

[5707] Anonymous. Books: New books and publications. Scientific American, 72(20):317, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n20/pdf/scientificamerican05181895-
317.pdf.

Anonymous:1895:BPt

[5708] Anonymous. Business and personal. Scientific American, 72(20):317, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n20/pdf/scientificamerican05181895-317a.pdf.

Anonymous:1895:Is

[5709] Anonymous. To inventors. Scientific American, 72(20):317, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n20/pdf/scientificamerican05181895-317c.pdf.

Anonymous:1895:NQt

[5710] Anonymous. Notes and queries. Scientific American, 72(20):317, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n20/pdf/scientificamerican05181895-317b.pdf.

Anonymous:1895:IIu

[5711] Anonymous. Index of inventions. Scientific American, 72(20):317–318, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n20/pdf/scientificamerican05181895-317d.pdf.

Anonymous:1895:PNPn

[5712] Anonymous. Patent notes: Patents. Scientific American, 72(20):319, May 18, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n20/pdf/scientificamerican05181895-319.pdf.

Anonymous:1895:ELH

[5713] Anonymous. Electric lighting and heating. Scientific American, 72(21):322–323, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
322.pdf.

Anonymous:1895:ETV

[5714] Anonymous. An elastic tire for vehicle wheels. Scientific American, 72(21):324, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
324b.pdf.

Anonymous:1895:FAC

[5715] Anonymous. To facilitate attaining correct touch on the piano. Scientific American, 72(21):324, May 25, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n21/pdf/scientificamerican05251895-324a.pdf.

Anonymous:1895:HDS

[5716] Anonymous. A heating drum in the smoke pipe. Scientific American, 72(21):324, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
324.pdf.

Anonymous:1895:IFB

[5717] Anonymous. Improved faucet for basins, sinks, etc. Scientific American, 72(21):324, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
324c.pdf.

Anonymous:1895:TLL

[5718] Anonymous. To throw life lines from vessels. Scientific American, 72(21):325, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
325.pdf.

Anonymous:1895:IRD

[5719] Anonymous. Improved radial drilling machine. Scientific American, 72(21):325–326, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
325a.pdf.

Anonymous:1895:Co

[5720] Anonymous. Correspondence. Scientific American, 72(21):326, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n21/pdf/scientificamerican05251895-326a.pdf.

Anonymous:1895:FCT

[5721] Anonymous. A fender for cable or trolley cars. Scientific American, 72(21):326, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
326.pdf.

Anonymous:1895:AYD

[5722] Anonymous. The American yacht defender. Scientific American, 72(21):327, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
327.pdf.

Anonymous:1895:SRS

[5723] Anonymous. Suspended railway systems. Scientific American, 72 (21):328–329, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
328.pdf.

Anonymous:1895:SOB

[5724] Anonymous. Sporting oxen and buffaloes. Scientific American, 72(21):329–331, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
329.pdf.

Anonymous:1895:PPN

[5725] Anonymous. Protecting the pipes near electric railways from electrolytic action. Scientific American, 72(21):331, May 25, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n21/pdf/scientificamerican05251895-331.pdf.

Anonymous:1895:SES

[5726] Anonymous. A simple and effective stump puller. Scientific American, 72(21):331, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
331a.pdf.

Anonymous:1895:BPu

[5727] Anonymous. Business and personal. Scientific American, 72(21):332, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n21/pdf/scientificamerican05251895-332a.pdf.

Anonymous:1895:RPIu

[5728] Anonymous. Recently patented inventions. Scientific American, 72(21):332, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n21/pdf/scientificamerican05251895-
332.pdf.

Anonymous:1895:NQu

[5729] Anonymous. Notes and queries. Scientific American, 72(21):332–333, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n21/pdf/scientificamerican05251895-332b.pdf.

Anonymous:1895:It

[5730] Anonymous. To inventors. Scientific American, 72(21):333, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n21/pdf/scientificamerican05251895-333.pdf.

Anonymous:1895:IIv

[5731] Anonymous. Index of inventions. Scientific American, 72(21):333, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n21/pdf/scientificamerican05251895-333a.pdf.

Anonymous:1895:PNPo

[5732] Anonymous. Patent notes: Patents. Scientific American, 72(21):335, May 25, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n21/pdf/scientificamerican05251895-335.pdf.

Anonymous:1895:WNM

[5733] Anonymous. Works of the National Machine Company, of Tiffin, Ohio. Scientific American, 72(22):337, June 1, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n22/pdf/scientificamerican06011895-337.pdf.

Anonymous:1895:BTD

[5734] Anonymous. The Berliner Telephone Decision. Scientific American, 72(22):338, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
338.pdf.

Anonymous:1895:HJa

[5735] Anonymous. The heavens in June. Scientific American, 72(22):338, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-338a.pdf.

Anonymous:1895:MS

[5736] Anonymous. The Moon’s story. Scientific American, 72(22):338–339, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-338b.pdf.

Anonymous:1895:IBa

[5737] Anonymous. An improved boiler. Scientific American, 72(22):340, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-340a.pdf.

Anonymous:1895:MSL

[5738] Anonymous. A machine for staking leather or skins. Scientific American, 72(22):340, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
340.pdf.

Anonymous:1895:SMD

[5739] Anonymous. A safety money drawer. Scientific American, 72(22):340, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-340b.pdf.

Anonymous:1895:GEE

[5740] Anonymous. Gas engines for electric lighting plant. Scientific American, 72(22):341, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
341.pdf.

Anonymous:1895:Cq

[5741] Anonymous. Correspondence. Scientific American, 72(22):342, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-342.pdf.

Anonymous:1895:SEF

[5742] Anonymous. Simple and efficient fans. Scientific American, 72(22):342, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-342a.pdf.

Anonymous:1895:NW

[5743] Anonymous. A novel wrench. Scientific American, 72(22):342–343, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-342b.pdf.

Anonymous:1895:TLC

[5744] Anonymous. The transportation of large cables. Scientific American, 72(22):344, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
344.pdf.

Anonymous:1895:SGW

[5745] Anonymous. The stained glass window industry. Scientific American, 72(22):345, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
345.pdf.

Anonymous:1895:ADR

[5746] Anonymous. Abstracts of decisions relating to patents. Scientific American, 72(22):346, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
346.pdf.

Anonymous:1895:FC

[5747] Anonymous. Flying crustacea. Scientific American, 72(22):347, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-347.pdf.

Anonymous:1895:BNBq

[5748] Anonymous. Books: New books and publications. Scientific American, 72(22):348, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
348a.pdf.

Anonymous:1895:RPIv

[5749] Anonymous. Recently patented inventions. Scientific American, 72(22):348, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n22/pdf/scientificamerican06011895-
348.pdf.

Anonymous:1895:BPv

[5750] Anonymous. Business and personal. Scientific American, 72(22):349, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-349.pdf.

Anonymous:1895:Iu

[5751] Anonymous. To inventors. Scientific American, 72(22):349, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-349b.pdf.

Anonymous:1895:IIx

[5752] Anonymous. Index of inventions. Scientific American, 72(22):349, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-349c.pdf.

Anonymous:1895:NQv

[5753] Anonymous. Notes and queries. Scientific American, 72(22):349, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-349a.pdf.

Anonymous:1895:PNPp

[5754] Anonymous. Patent notes: Patents. Scientific American, 72(22):351, June 1, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n22/pdf/scientificamerican06011895-351.pdf.

Anonymous:1895:USBb

[5755] Anonymous. The United States battle ship Oregon. Scientific American, 72(23):353, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
353.pdf.

Anonymous:1895:PSBb

[5756] Anonymous. The present status of the bicycle. Scientific American, 72(23):354, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
354.pdf.

Anonymous:1895:PBS

[5757] Anonymous. Possibilities of beet sugar industries. Scientific American, 72(23):354–355, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
354a.pdf.

Anonymous:1895:CPb

[5758] Anonymous. A cylindrical propeller. Scientific American, 72(23):356, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-356.pdf.

Anonymous:1895:CPF

[5759] Anonymous. A continuous process fiber machine. Scientific American, 72(23):356, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
356a.pdf.

Anonymous:1895:DCW

[5760] Anonymous. A device for centering work. Scientific American, 72(23):356, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
356b.pdf.

Anonymous:1895:DVT

[5761] Anonymous. The “Daugherty Visible” typewriter. Scientific American, 72(23):357–358, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
357.pdf.

Anonymous:1895:KRR

[5762] Anonymous. Kitchen refuse readily disposed of. Scientific American, 72(23):358–359, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
358.pdf.

Anonymous:1895:NASc

[5763] Anonymous. The new Atlantic steamer St. Louis. Scientific American, 72(23):360, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
360a.pdf.

Anonymous:1895:PTb

[5764] Anonymous. A petroleum tricycle. Scientific American, 72(23):360, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-360.pdf.

Anonymous:1895:TPN

[5765] Anonymous. Terrific power of nitro-glycerine. Scientific American, 72(23):361, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
361.pdf.

Anonymous:1895:VRN

[5766] Anonymous. The Victoria Regia in New Jersey. Scientific American, 72(23):361, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
361a.pdf.

Anonymous:1895:Cr

[5767] Anonymous. Correspondence. Scientific American, 72(23):362–363, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-362.pdf.

Anonymous:1895:NB

[5768] Anonymous. A nautical bicycle. Scientific American, 72(23):363–364, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-363.pdf.

Anonymous:1895:RPIw

[5769] Anonymous. Recently patented inventions. Scientific American, 72(23):364, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
364.pdf.

Anonymous:1895:BNBr

[5770] Anonymous. Books: New books and publications. Scientific American, 72(23):364–365, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n23/pdf/scientificamerican06081895-
364a.pdf.

Anonymous:1895:BPw

[5771] Anonymous. Business and personal. Scientific American, 72(23):365, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-365.pdf.

Anonymous:1895:Iv

[5772] Anonymous. To inventors. Scientific American, 72(23):365, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-365b.pdf.

Anonymous:1895:NQw

[5773] Anonymous. Notes and queries. Scientific American, 72(23):365, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-365a.pdf.

Anonymous:1895:IIy

[5774] Anonymous. Index of inventions. Scientific American, 72(23):365–366, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-365c.pdf.

Anonymous:1895:PNPq

[5775] Anonymous. Patent notes: Patents. Scientific American, 72(23):367, June 8, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n23/pdf/scientificamerican06081895-367.pdf.

Anonymous:1895:CDCa

[5776] Anonymous. The Chicago Drainage Canal. Scientific American, 72(24):369–370, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
369.pdf.

Anonymous:1895:EAR

[5777] Anonymous. The electrification of the air by rain drops and waves. Scientific American, 72(24):370, June 15, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n24/pdf/scientificamerican06151895-370.pdf.

Anonymous:1895:SSb

[5778] Anonymous. Supremacy of the sea. Scientific American, 72 (24):370–371, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
370a.pdf.

Anonymous:1895:DPB

[5779] Anonymous. A drill press for bicycle repairing. Scientific American, 72(24):372, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
372a.pdf.

Anonymous:1895:EDE

[5780] Anonymous. An electric dental engine. Scientific American, 72(24):372, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-372.pdf.

Anonymous:1895:IIBa

[5781] Anonymous. Improvements of interest to bicycle riders. Scientific American, 72(24):373, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
373.pdf.

Anonymous:1895:JC

[5782] Anonymous. Japanese clocks. Scientific American, 72(24):373–375, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-373a.pdf.

Anonymous:1895:NGM

[5783] Anonymous. A new gasoline motor for boats and launches. Scientific American, 72(24):375, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
375.pdf.

Anonymous:1895:GSS

[5784] Anonymous. Grand saloon of the steamer St. Louis. Scientific American, 72(24):376, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
376.pdf.

Anonymous:1895:TBW

[5785] Anonymous. Torpedo boat for the war ship Maine. Scientific American, 72(24):376–377, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
376a.pdf.

Anonymous:1895:LKA

[5786] Anonymous. The little Koodoo antelope in the Berlin Zoological Garden. Scientific American, 72(24):377–378, June 15, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n24/pdf/scientificamerican06151895-377.pdf.

Anonymous:1895:DSSb

[5787] Anonymous. A drawing of Sun spots. Scientific American, 72 (24):378–379, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
378.pdf.

Anonymous:1895:LP

[5788] Anonymous. A lightning photograph. Scientific American, 72(24):379, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-379.pdf.

Anonymous:1895:SC

[5789] Anonymous. A straw crush hat. Scientific American, 72(24):379–380, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-379a.pdf.

Anonymous:1895:RPIx

[5790] Anonymous. Recently patented inventions. Scientific American, 72(24):380–381, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
380.pdf.

Anonymous:1895:BNBs

[5791] Anonymous. Books: New books and publications. Scientific American, 72(24):381, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n24/pdf/scientificamerican06151895-
381b.pdf.

Anonymous:1895:BPx

[5792] Anonymous. Business and personal. Scientific American, 72(24):381, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-381.pdf.

Anonymous:1895:Iw

[5793] Anonymous. To inventors. Scientific American, 72(24):381, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-381c.pdf.

Anonymous:1895:NQx

[5794] Anonymous. Notes and queries. Scientific American, 72(24):381, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-381a.pdf.

Anonymous:1895:IIz

[5795] Anonymous. Index of inventions. Scientific American, 72(24):381–382, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-381d.pdf.

Anonymous:1895:PNPr

[5796] Anonymous. Patent notes: Patents. Scientific American, 72(24):383, June 15, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n24/pdf/scientificamerican06151895-383.pdf.

Anonymous:1895:HTN

[5797] Anonymous. Hunting the tiger by night. Scientific American, 72(25):385, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
385.pdf.

Anonymous:1895:SAA

[5798] Anonymous. The Scientific American as an advertising medium. Scientific American, 72(25):386, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
386b.pdf.

Anonymous:1895:CMb

[5799] Anonymous. The Chupaderos meteorite. Scientific American, 72 (25):386, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
386a.pdf.

Anonymous:1895:FEH

[5800] Anonymous. Force exerted by the human jaws. Scientific American, 72(25):386, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
386c.pdf.

Anonymous:1895:FVS

[5801] Anonymous. First voyage of the St. Louis. Scientific American, 72(25):386, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
386.pdf.

Anonymous:1895:IWA

[5802] Anonymous. Iron working among primitive peoples. Scientific American, 72(25):386, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
386d.pdf.

Anonymous:1895:CVM

[5803] Anonymous. The commercial value of Monazite. Scientific American, 72(25):386–387, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
386e.pdf.

Anonymous:1895:ARCb

[5804] Anonymous. Another race of carriages without horses. Scientific American, 72(25):387, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
387b.pdf.

Anonymous:1895:CDCb

[5805] Anonymous. Chicago Drainage Canal — its effects on the commerce of the Lakes. Scientific American, 72(25):387, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-387a.pdf.

Anonymous:1895:SFA

[5806] Anonymous. Some facts about glass. Scientific American, 72(25):387, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-387.pdf.

Anonymous:1895:ECSa

[5807] Anonymous. An electrically controlled speaking tube. Scientific American, 72(25):388, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
388b.pdf.

Anonymous:1895:ICT

[5808] Anonymous. An improved car truck. Scientific American, 72(25):388, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-388.pdf.

Anonymous:1895:IHT

[5809] Anonymous. An improved Hame tug. Scientific American, 72(25):388, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-388a.pdf.

Anonymous:1895:HAC

[5810] Anonymous. History of automobile carriages. Scientific American, 72(25):389–390, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
389.pdf.

Anonymous:1895:Cs

[5811] Anonymous. Correspondence. Scientific American, 72(25):391, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-391.pdf.

Anonymous:1895:GWS

[5812] Anonymous. A great war ship. Scientific American, 72(25):392, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-392.pdf.

Anonymous:1895:HBN

[5813] Anonymous. The High Bridge over the North Sea Canal near Levensau. Scientific American, 72(25):392, June 22, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n25/pdf/scientificamerican06221895-392a.pdf.

Anonymous:1895:ONS

[5814] Anonymous. The opening of the North Sea Canal. Scientific American, 72(25):394, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
394.pdf.

Anonymous:1895:CSa

[5815] Anonymous. The card skimmer. Scientific American, 72(25):395, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-395.pdf.

Anonymous:1895:MSC

[5816] Anonymous. Machine for sharpening cutlery. Scientific American, 72(25):395, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
395a.pdf.

Anonymous:1895:BNBt

[5817] Anonymous. Books: New books and publications. Scientific American, 72(25):396, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
396a.pdf.

Anonymous:1895:RPIy

[5818] Anonymous. Recently patented inventions. Scientific American, 72(25):396, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n25/pdf/scientificamerican06221895-
396.pdf.

Anonymous:1895:BPy

[5819] Anonymous. Business and personal. Scientific American, 72(25):397, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-397.pdf.

Anonymous:1895:IIbb

[5820] Anonymous. Index of inventions. Scientific American, 72(25):397, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-397b.pdf.

Anonymous:1895:NQy

[5821] Anonymous. Notes and queries. Scientific American, 72(25):397, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-397a.pdf.

Anonymous:1895:PNPs

[5822] Anonymous. Patent notes: Patents. Scientific American, 72(25):399, June 22, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n25/pdf/scientificamerican06221895-399.pdf.

Anonymous:1895:EOW

[5823] Anonymous. Electrically operated water supply plant, De Kalb, Ill. Scientific American, 72(26):401–402, June 29, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n26/pdf/scientificamerican06291895-401.pdf.

Anonymous:1895:CGa

[5824] Anonymous. The celebration in Germany. Scientific American, 72(26):402, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
402.pdf.

Anonymous:1895:TLa

[5825] Anonymous. Temperatures of lakes. Scientific American, 72 (26):402–403, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
402a.pdf.

Anonymous:1895:DRPd

[5826] Anonymous. Decisions relating to patents. Scientific American, 72(26):403, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
403.pdf.

Anonymous:1895:BASa

[5827] Anonymous. Briner’s ash sifter. Scientific American, 72(26):404, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-404b.pdf.

Anonymous:1895:IPA

[5828] Anonymous. An improved piano action frame. Scientific American, 72(26):404, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
404.pdf.

Anonymous:1895:SCR

[5829] Anonymous. A street car rail sweeper. Scientific American, 72(26):404, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-404a.pdf.

Anonymous:1895:SHC

[5830] Anonymous. A simple hair curling device. Scientific American, 72(26):404, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
404c.pdf.

Anonymous:1895:ABSa

[5831] Anonymous. An automatic brake slack adjuster. Scientific American, 72(26):405, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
405.pdf.

Anonymous:1895:BDR

[5832] Anonymous. A Bicycler’s distance recorder. Scientific American, 72(26):405, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
405c.pdf.

Anonymous:1895:BST

[5833] Anonymous. The Bundy steam trap. Scientific American, 72(26):405, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-405a.pdf.

Anonymous:1895:IBO

[5834] Anonymous. The improved Bell odometer. Scientific American, 72(26):405, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
405b.pdf.

Anonymous:1895:BRR

[5835] Anonymous. The Blake repeating rifle. Scientific American, 72(26):406, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-406.pdf.

Anonymous:1895:Ct

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-407.pdf.

Anonymous:1895:GFG

[5837] Anonymous. Great falls and geysers — Yellowstone National Park. Scientific American, 72(26):408, June 29, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v72/
n26/pdf/scientificamerican06291895-408a.pdf.

Anonymous:1895:PACa

[5838] Anonymous. The projectors of the Atlantic cable. Scientific American, 72(26):408, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
408.pdf.

Anonymous:1895:OHC

[5839] Anonymous. The opening of the Harlem Canal. Scientific American, 72(26):409–410, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
409.pdf.

Anonymous:1895:BNBu

[5840] Anonymous. Books: New books and publications. Scientific American, 72(26):411, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
411a.pdf.

Anonymous:1895:BPz

[5841] Anonymous. Business and personal. Scientific American, 72(26):411, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-411c.pdf.

Anonymous:1895:NQz

[5842] Anonymous. Notes and queries. Scientific American, 72(26):411, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-411b.pdf.

Anonymous:1895:RPIz

[5843] Anonymous. Recently patented inventions. Scientific American, 72(26):411, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v72/n26/pdf/scientificamerican06291895-
411.pdf.

Anonymous:1895:IIbc

[5844] Anonymous. Index of inventions. Scientific American, 72(26):412, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-412.pdf.

Anonymous:1895:Ix

[5845] Anonymous. Index. Scientific American, 72(26):413–415, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-413.pdf.

Anonymous:1895:PNPt

[5846] Anonymous. Patent notes: Patents. Scientific American, 72(26):415, June 29, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v72/n26/pdf/scientificamerican06291895-415.pdf.

Anonymous:1895:SSCb

[5847] Anonymous. A steam supply from combined locomotives. Scientific American, 73(1):1, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n1/pdf/scientificamerican07061895-
1.pdf.

Anonymous:1895:RPS

[5848] Anonymous. Recent progress in soil analysis. Scientific American, 73(1): 2, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-2.pdf.

Anonymous:1895:HJb

[5849] Anonymous. The heavens in July. Scientific American, 73(1):3, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-3.pdf.

Anonymous:1895:AL

[5850] Anonymous. An astronomical lantern. Scientific American, 73(1):4, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-4b.pdf.

Anonymous:1895:BDW

[5851] Anonymous. A bulletproof door and window shield for mail and express cars. Scientific American, 73(1):4, July 6, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n1/pdf/scientificamerican07061895-4d.pdf.

Anonymous:1895:DRT

[5852] Anonymous. A double rotary toy windmill. Scientific American, 73(1): 4, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-4c.pdf.

Anonymous:1895:IFJ

[5853] Anonymous. An improved fruit jar clamp. Scientific American, 73(1): 4, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-4.pdf.

Anonymous:1895:RSB

[5854] Anonymous. A ribbon substitute for bicycle chains. Scientific American, 73(1):4, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n1/pdf/scientificamerican07061895-
4a.pdf.

Anonymous:1895:RWI

[5855] Anonymous. Redrawing of wrought iron tubes. Scientific American, 73(1):5–6, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n1/pdf/scientificamerican07061895-
5.pdf.

Anonymous:1895:Cu

[5856] Anonymous. Correspondence. Scientific American, 73(1):7, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-7.pdf.

Anonymous:1895:RCI

[5857] Anonymous. Relics of a Cuban Insurrection. Scientific American, 73(1): 8, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-8.pdf.

Anonymous:1895:CDa

[5858] Anonymous. The Casseuil Drawbridge. Scientific American, 73(1):9, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-9.pdf.

Hanscom:1895:AAP

[5859] W. W. Hanscom. The annealing of armor plates by electricity. Scientific American, 73(1):9–10, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n1/pdf/scientificamerican07061895-
9a.pdf.

Anonymous:1895:FG

[5860] Anonymous. The fountain at Geneva. Scientific American, 73(1):11, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-11.pdf.

Anonymous:1895:BNBv

[5861] Anonymous. Books: New books and publications. Scientific American, 73(1):12, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n1/pdf/scientificamerican07061895-
12a.pdf.

Anonymous:1895:RPIba

[5862] Anonymous. Recently patented inventions. Scientific American, 73(1): 12, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-12.pdf.

Anonymous:1895:BPba

[5863] Anonymous. Business and personal. Scientific American, 73(1):13, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-13.pdf.

Anonymous:1895:Iy

[5864] Anonymous. To inventors. Scientific American, 73(1):13, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-13b.pdf.

Anonymous:1895:IIbd

[5865] Anonymous. Index of inventions. Scientific American, 73(1):13, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-13c.pdf.

Anonymous:1895:NQba

[5866] Anonymous. Notes and queries. Scientific American, 73(1):13, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-13a.pdf.

Anonymous:1895:PNPu

[5867] Anonymous. Patent notes: Patents. Scientific American, 73(1):14, July 6, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n1/pdf/scientificamerican07061895-14.pdf.

Anonymous:1895:BNa

[5868] Anonymous. The ball nozzle. Scientific American, 73(2):17–18, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-17.pdf.

Anonymous:1895:NYS

[5869] Anonymous. The New York State Fish Hatchery, Cold Spring Harbor. Scientific American, 73(2):18–19, July 13, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n2/pdf/scientificamerican07131895-18.pdf.

Anonymous:1895:NPBa

[5870] Anonymous. To the North Pole by balloon. Scientific American, 73(2):20, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
20.pdf.

Anonymous:1895:RLM

[5871] Anonymous. A registering lumber measure. Scientific American, 73(2):20, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
20a.pdf.

Anonymous:1895:ARS

[5872] Anonymous. Apparatus for raising sunken vessels. Scientific American, 73(2):21, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
21b.pdf.

Anonymous:1895:BFC

[5873] Anonymous. A blast furnace with center blast. Scientific American, 73(2):21, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
21c.pdf.

Anonymous:1895:DCE

[5874] Anonymous. A door controlling elevator attachment. Scientific American, 73(2):21, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
21a.pdf.

Anonymous:1895:SAI

[5875] Anonymous. A shaft aligning improvement. Scientific American, 73(2):21, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
21.pdf.

Anonymous:1895:Cv

[5876] Anonymous. Correspondence. Scientific American, 73(2):22, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-22.pdf.

Anonymous:1895:THH

[5877] Anonymous. Thomas Henry Huxley. Scientific American, 73(2):23, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-23.pdf.

Anonymous:1895:ABT

[5878] Anonymous. An automatic bicycle tire inflater. Scientific American, 73(2):24, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
24a.pdf.

Anonymous:1895:PED

[5879] Anonymous. Portable electric drill. Scientific American, 73(2):24, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-24.pdf.

Anonymous:1895:NPBb

[5880] Anonymous. The new Port of Bizerte. Scientific American, 73(2):25–26, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-25.pdf.

Anonymous:1895:H

[5881] Anonymous. The heron. Scientific American, 73(2):27–28, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-27.pdf.

Anonymous:1895:RPIbb 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
28.pdf.

Anonymous:1895:BNBw

[5883] Anonymous. Books: New books and publications. Scientific American, 73(2):29, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n2/pdf/scientificamerican07131895-
29.pdf.

Anonymous:1895:BPbb

[5884] Anonymous. Business and personal. Scientific American, 73(2):29, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-29a.pdf.

Anonymous:1895:Iz

[5885] Anonymous. To inventors. Scientific American, 73(2):29, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-29c.pdf.

Anonymous:1895:NQbb

[5886] Anonymous. Notes and queries. Scientific American, 73(2):29, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-29b.pdf.

Anonymous:1895:IIbe

[5887] Anonymous. Index of inventions. Scientific American, 73(2):29–30, July 13, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n2/pdf/scientificamerican07131895-29d.pdf.

Anonymous:1895:MPA

[5888] Anonymous. The manufacture of photographic apparatus. Scientific American, 73(3):33, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
33.pdf.

Anonymous:1895:PEa

[5889] Anonymous. Problems in economics. Scientific American, 73(3):34, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-34.pdf.

Anonymous:1895:SSCc

[5890] Anonymous. Summer study for city children. Scientific American, 73(3):34–35, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
34a.pdf.

Anonymous:1895:ALG

[5891] Anonymous. An automatic life guard for cars. Scientific American, 73(3):36, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
36c.pdf.

Anonymous:1895:CWB

[5892] Anonymous. A carriage and wagon brake. Scientific American, 73(3):36, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-36b.pdf.

Anonymous:1895:HCT

[5893] Anonymous. A hand crimping tool. Scientific American, 73(3):36, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-36d.pdf.

Anonymous:1895:IWS

[5894] Anonymous. An improved window shade. Scientific American, 73(3):36, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-36.pdf.

Anonymous:1895:SSA

[5895] Anonymous. A swinging seat for agricultural implements. Scientific American, 73(3):36, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
36a.pdf.

Anonymous:1895:ECSb

[5896] Anonymous. The electrical conduit system in New York City. Scientific American, 73(3):38, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
38.pdf.

Anonymous:1895:ELP

[5897] Anonymous. Electric light plant steamer Bay State. Scientific American, 73(3):39, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
39.pdf.

Anonymous:1895:AC

[5898] Anonymous. Automobile carriages. Scientific American, 73(3):40, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-40.pdf.

Anonymous:1895:NSS

[5899] Anonymous. The North Sea Ship Canal. Scientific American, 73(3):41, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-41.pdf.

Anonymous:1895:Cx

[5900] Anonymous. Correspondence. Scientific American, 73(3):42–44, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-42.pdf.

Anonymous:1895:RPIbc

[5901] Anonymous. Recently patented inventions. Scientific American, 73(3):44–45, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
44.pdf.

Anonymous:1895:BNBx

[5902] Anonymous. Books: New books and publications. Scientific American, 73(3):45, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n3/pdf/scientificamerican07201895-
45.pdf.

Anonymous:1895:BPbc

[5903] Anonymous. Business and personal. Scientific American, 73(3):45, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-45a.pdf.

Anonymous:1895:Ibb

[5904] Anonymous. To inventors. Scientific American, 73(3):45, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-45c.pdf.

Anonymous:1895:NQbc

[5905] Anonymous. Notes and queries. Scientific American, 73(3):45, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-45b.pdf.

Anonymous:1895:IIbf

[5906] Anonymous. Index of inventions. Scientific American, 73(3):45–46, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-45d.pdf.

Anonymous:1895:PNPv

[5907] Anonymous. Patent notes: Patents. Scientific American, 73(3):47, July 20, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n3/pdf/scientificamerican07201895-47.pdf.

Anonymous:1895:LSN

[5908] Anonymous. The Lake steamer North Land. Scientific American, 73(4):49–50, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
49.pdf.

Anonymous:1895:BASb

[5909] Anonymous. Bicycle an aid to science and art. Scientific American, 73(4):50, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
50.pdf.

Anonymous:1895:TH

[5910] Anonymous. Terrestrial helium. Scientific American, 73(4):50–51, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-50a.pdf.

Anonymous:1895:EIWa

[5911] Anonymous. Electrical items worth remembering. Scientific American, 73(4):51, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
51.pdf.

Anonymous:1895:HSN

[5912] Anonymous. High speed navigation. Scientific American, 73(4):52, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-52a.pdf.

Anonymous:1895:IBS

[5913] Anonymous. An improved Bob sleigh. Scientific American, 73(4):52, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-52d.pdf.

Anonymous:1895:IMI

[5914] Anonymous. An improved musical instrument. Scientific American, 73(4):52, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
52.pdf.

Anonymous:1895:PLB

[5915] Anonymous. A propeller life buoy. Scientific American, 73(4):52, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-52b.pdf.

Anonymous:1895:WVP

[5916] Anonymous. The wheel versus the pedestrian. Scientific American, 73(4):52, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
52c.pdf.

Anonymous:1895:AN

[5917] Anonymous. Aerial navigation. Scientific American, 73(4):53, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-53.pdf.

Anonymous:1895:IEA

[5918] Anonymous. The International Exposition at Atlanta. Scientific American, 73(4):53–54, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
53a.pdf.

Anonymous:1895:CCE

[5919] Anonymous. The Cornell crew in England. Scientific American, 73(4):55, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
55a.pdf.

Anonymous:1895:ETEb

[5920] Anonymous. Experiment on thermo-electric currents. Scientific American, 73(4):55, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
55.pdf.

Anonymous:1895:SSL

[5921] Anonymous. A sectional side launch dock. Scientific American, 73(4):56, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
56.pdf.

Anonymous:1895:MRW

[5922] Anonymous. Mount Rainier, Washington. Scientific American, 73(4):57, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
57.pdf.

Anonymous:1895:Cy

[5923] Anonymous. Correspondence. Scientific American, 73(4):58, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-58.pdf.

Anonymous:1895:PBR

[5924] Anonymous. The Paradise birds recently discovered in New Guinea. Scientific American, 73(4):59–60, July 27, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n4/pdf/scientificamerican07271895-59.pdf.

Anonymous:1895:BNBy

[5925] Anonymous. Books: New books and publications. Scientific American, 73(4):61, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
61a.pdf.

Anonymous:1895:BPbd

[5926] Anonymous. Business and personal. Scientific American, 73(4):61, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-61b.pdf.

Anonymous:1895:NQbd

[5927] Anonymous. Notes and queries. Scientific American, 73(4):61, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-61c.pdf.

Anonymous:1895:RPIbd 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n4/pdf/scientificamerican07271895-
61.pdf.

Anonymous:1895:Ibd

[5929] Anonymous. To inventors. Scientific American, 73(4):62, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-62.pdf.

Anonymous:1895:IIbg

[5930] Anonymous. Index of inventions. Scientific American, 73(4):62, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-62a.pdf.

Anonymous:1895:PNPw

[5931] Anonymous. Patent notes: Patents. Scientific American, 73(4):63, July 27, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n4/pdf/scientificamerican07271895-63.pdf.

Anonymous:1895:MAL

[5932] Anonymous. The manufacture of artificial limbs. Scientific American, 73(5):65–66, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
65.pdf.

Anonymous:1895:NPM

[5933] Anonymous. New prizes for motor carriage competitions. Scientific American, 73(5):66, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
66a.pdf.

Anonymous:1895:PBd

[5934] Anonymous. The physics of the bicycle. Scientific American, 73(5):66, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-66.pdf.

Anonymous:1895:AEN

[5935] Anonymous. Atlanta Exposition notes. Scientific American, 73 (5):66–67, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
66b.pdf.

Anonymous:1895:HAb

[5936] Anonymous. The heavens in August. Scientific American, 73(5):67, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-67a.pdf.

Anonymous:1895:LC

[5937] Anonymous. Leather cannon. Scientific American, 73(5):67, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-67.pdf.

Anonymous:1895:DRPe

[5938] Anonymous. Decisions relating to patents. Scientific American, 73(5):67–68, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
67b.pdf.

Anonymous:1895:BSS

[5939] Anonymous. A bracket support for shades, curtains, etc. Scientific American, 73(5):68, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
68b.pdf.

Anonymous:1895:CHB

[5940] Anonymous. A case to hold a bicycle. Scientific American, 73(5):68, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-68.pdf.

Anonymous:1895:IGM

[5941] Anonymous. An iron grinding mill of large capacity. Scientific American, 73(5):68, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
68a.pdf.

Anonymous:1895:ITCa

[5942] Anonymous. An improved thill coupling. Scientific American, 73(5):68, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-68c.pdf.

Anonymous:1895:IRP

[5943] Anonymous. Improved repeating pistol. Scientific American, 73(5):69, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-69.pdf.

Anonymous:1895:Cz

[5944] Anonymous. Correspondence. Scientific American, 73(5):70, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-70.pdf.

Anonymous:1895:AMC

[5945] Anonymous. The Avondale Marble Company’s new crane. Scientific American, 73(5):71, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
71.pdf.

Anonymous:1895:HCS

[5946] Anonymous. Hydraulic caisson sinking for foundation piers. Scientific American, 73(5):72, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
72.pdf.

Anonymous:1895:RC

[5947] Anonymous. The rawhide cannon. Scientific American, 73(5):73–74, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-73.pdf.

Anonymous:1895:FD

[5948] Anonymous. A flying dormouse. Scientific American, 73(5):75, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-75.pdf.

Anonymous:1895:BNBz

[5949] Anonymous. Books: New books and publications. Scientific American, 73(5):76, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
76a.pdf.

Anonymous:1895:BPbe

[5950] Anonymous. Business and personal. Scientific American, 73(5):76, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-76b.pdf.

Anonymous:1895:RPIbe

[5951] Anonymous. Recently patented inventions. Scientific American, 73(5):76, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n5/pdf/scientificamerican08031895-
76.pdf.

Anonymous:1895:NQbe

[5952] Anonymous. Notes and queries. Scientific American, 73(5):76–77, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-76c.pdf.

Anonymous:1895:Ibe

[5953] Anonymous. To inventors. Scientific American, 73(5):77, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-77.pdf.

Anonymous:1895:IIbh

[5954] Anonymous. Index of inventions. Scientific American, 73(5):77, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-77a.pdf.

Anonymous:1895:PNPx

[5955] Anonymous. Patent notes: Patents. Scientific American, 73(5):79, August 3, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n5/pdf/scientificamerican08031895-79.pdf.

Anonymous:1895:RSU

[5956] Anonymous. Remarkable speed of the United States cruiser Columbia. Scientific American, 73(6):82, August 10, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n6/pdf/scientificamerican08101895-82.pdf.

Anonymous:1895:WRE

[5957] Anonymous. Worth repeating-electrical. Scientific American, 73 (6):82, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
82a.pdf.

Anonymous:1895:HCR

[5958] Anonymous. The horseless carriage races. Scientific American, 73(6):82–83, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
82b.pdf.

Anonymous:1895:DBC

[5959] Anonymous. Dyeing and bleaching cotton and woolen yarn. Scientific American, 73(6):84, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
84.pdf.

Anonymous:1895:GMD

[5960] Anonymous. A garment measuring device. Scientific American, 73(6):85, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
85b.pdf.

Anonymous:1895:ISFb

[5961] Anonymous. An improved shutter fastener. Scientific American, 73(6):85, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
85a.pdf.

Anonymous:1895:SP

[5962] Anonymous. Submarine photography. Scientific American, 73 (6):85, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
85c.pdf.

Anonymous:1895:TPB

[5963] Anonymous. Taylor’s patent bookholder for telephone directories, etc. Scientific American, 73(6):85, August 10, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n6/pdf/scientificamerican08101895-85.pdf.

Anonymous:1895:FPA

[5964] Anonymous. Flowers perfumed artificially. Scientific American, 73(6):85–86, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
85d.pdf.

Anonymous:1895:PC

[5965] Anonymous. A panoramic camera. Scientific American, 73(6):86, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-86.pdf.

Anonymous:1895:NSTb

[5966] Anonymous. A ninety-six ton electric locomotive. Scientific American, 73(6):87, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
87.pdf.

Anonymous:1895:FMC

[5967] Anonymous. The fan and magnetic clock. Scientific American, 73(6):88, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
88.pdf.

Anonymous:1895:AAT

[5968] Anonymous. The Aqua aerial trolley. Scientific American, 73(6):89, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-89.pdf.

Anonymous:1895:Cbb

[5969] Anonymous. Correspondence. Scientific American, 73(6):90, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-90.pdf.

Dallinger:1895:NMR

[5970] W. H. Dallinger. Note on Melicerta Ringens. Scientific American, 73(6):91–92, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
91.pdf.

Riley:1895:CBI

[5971] C. V. Riley. Croton bugs injurious to office papers. Scientific American, 73(6):92, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
92.pdf.

Anonymous:1895:BPbf

[5972] Anonymous. Business and personal. Scientific American, 73(6):93, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-93a.pdf.

Anonymous:1895:Ibf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-93c.pdf.

Anonymous:1895:NQbf

[5974] Anonymous. Notes and queries. Scientific American, 73(6):93, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-93b.pdf.

Anonymous:1895:RPIbf

[5975] Anonymous. Recently patented inventions. Scientific American, 73(6):93, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n6/pdf/scientificamerican08101895-
93.pdf.

Anonymous:1895:IIbi

[5976] Anonymous. Index of inventions. Scientific American, 73(6):93–94, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-93d.pdf.

Anonymous:1895:PNPy

[5977] Anonymous. Patent notes: Patents. Scientific American, 73(6):95, August 10, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n6/pdf/scientificamerican08101895-95.pdf.

Anonymous:1895:PPM

[5978] Anonymous. Photographic printing by machinery. Scientific American, 73(7):97, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
97.pdf.

Anonymous:1895:DA

[5979] Anonymous. The discovery of argon. Scientific American, 73(7):98, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n7/pdf/scientificamerican08171895-98.pdf.

Anonymous:1895:EIWb

[5980] Anonymous. Electrical items worth remembering. Scientific American, 73(7):98–99, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
98a.pdf.

Anonymous:1895:BLCb

[5981] Anonymous. The Biological Laboratory at Cold Spring Harbor, Long Island, N.Y. Scientific American, 73(7):99, August 17, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n7/pdf/scientificamerican08171895-99.pdf.

Anonymous:1895:HTF

[5982] Anonymous. A hose tower and fire escape. Scientific American, 73(7):100, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
100b.pdf.

Anonymous:1895:HTT

[5983] Anonymous. The “Home Trainer” for testing bicycles. Scientific American, 73(7):100, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
100a.pdf.

Anonymous:1895:ILMa

[5984] Anonymous. An improved lawn mower. Scientific American, 73 (7):100, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
100.pdf.

Anonymous:1895:TCW

[5985] Anonymous. A tool for closing water pipes. Scientific American, 73(7):100, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
100c.pdf.

Anonymous:1895:ICR

[5986] Anonymous. Improved compound reversing engines. Scientific American, 73(7):101–104, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
101.pdf.

Anonymous:1895:GDP

[5987] Anonymous. The great dam of the Peryar, India. Scientific American, 73(7):104–105, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
104.pdf.

Anonymous:1895:Cbc

[5988] Anonymous. Correspondence. Scientific American, 73(7):106, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n7/pdf/scientificamerican08171895-106.pdf.

Anonymous:1895:PRa

[5989] Anonymous. Photography of the retina. Scientific American, 73(7):106, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
106a.pdf.

Anonymous:1895:CFa

[5990] Anonymous. Che-Foo. Scientific American, 73(7):107–108, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n7/pdf/scientificamerican08171895-107.pdf.

Anonymous:1895:RPIbg

[5991] Anonymous. Recently patented inventions. Scientific American, 73 (7):108–109, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
108.pdf.

Anonymous:1895:BPbg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n7/pdf/scientificamerican08171895-109.pdf.

Anonymous:1895:Ibg

[5993] Anonymous. To inventors. Scientific American, 73(7):109, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n7/pdf/scientificamerican08171895-109b.pdf.

Anonymous:1895:NQbg

[5994] Anonymous. Notes and queries. Scientific American, 73(7):109, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n7/pdf/scientificamerican08171895-109a.pdf.

Anonymous:1895:IIbj

[5995] Anonymous. Index of inventions. Scientific American, 73(7): 109–110, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n7/pdf/scientificamerican08171895-
109c.pdf.

Anonymous:1895:PNPz

[5996] Anonymous. Patent notes: Patents. Scientific American, 73(7):111, August 17, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n7/pdf/scientificamerican08171895-111.pdf.

Anonymous:1895:TTM

[5997] Anonymous. The Thorne typesetting machine. Scientific American, 73(8):113, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
113.pdf.

Anonymous:1895:CBE

[5998] Anonymous. Chinch bug extermination. Scientific American, 73 (8):114, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
114a.pdf.

Anonymous:1895:EIWc

[5999] Anonymous. Electrical items worth repeating. Scientific American, 73(8):114, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
114b.pdf.

Anonymous:1895:USC

[6000] Anonymous. The United States cruiser Columbia. Scientific American, 73(8):114, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
114.pdf.

Anonymous:1895:RHP

[6001] Anonymous. A remarkable hydraulic power supply. Scientific American, 73(8):114–115, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
114c.pdf.

Anonymous:1895:DRPf

[6002] Anonymous. Decisions relating to patents. Scientific American, 73(8):115, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
115.pdf.

Anonymous:1895:HAD

[6003] Anonymous. A hub attaching device. Scientific American, 73 (8):116, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
116a.pdf.

Anonymous:1895:ISa

[6004] Anonymous. An improved saw. Scientific American, 73(8):116, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n8/pdf/scientificamerican08241895-116.pdf.

Anonymous:1895:ITCb

[6005] Anonymous. An improved tube cleaner. Scientific American, 73(8):116, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
116c.pdf.

Anonymous:1895:PTI

[6006] Anonymous. A pneumatic tire inflating valve. Scientific American, 73(8):116, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
116b.pdf.

Anonymous:1895:MGS

[6007] Anonymous. Manufacture of glue and sizing. Scientific American, 73(8):117–118, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
117.pdf.

Anonymous:1895:WV

[6008] Anonymous. Water velocipede. Scientific American, 73(8):119, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n8/pdf/scientificamerican08241895-119.pdf.

Anonymous:1895:NBP

[6009] Anonymous. The new bicycle path, Ocean Parkway, Brooklyn, N.Y. Scientific American, 73(8):120, August 24, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n8/pdf/scientificamerican08241895-120.pdf.

Anonymous:1895:FHH

[6010] Anonymous. Five hundred I.H.P. compound engine. Scientific American, 73(8):121, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
121.pdf.

Anonymous:1895:Cbd

[6011] Anonymous. Correspondence. Scientific American, 73(8):122, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n8/pdf/scientificamerican08241895-122.pdf.

Anonymous:1895:CIa

[6012] Anonymous. The Cuban Insurrection. Scientific American, 73 (8):123–124, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
123.pdf.

Anonymous:1895:BPbh

[6013] Anonymous. Business and personal. Scientific American, 73(8):125, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n8/pdf/scientificamerican08241895-125a.pdf.

Anonymous:1895:NQbh

[6014] Anonymous. Notes and queries. Scientific American, 73(8):125, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n8/pdf/scientificamerican08241895-125b.pdf.

Anonymous:1895:RPIbh

[6015] Anonymous. Recently patented inventions. Scientific American, 73(8):125, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
125.pdf.

Anonymous:1895:IIbk

[6016] Anonymous. Index of inventions. Scientific American, 73(8): 125–126, August 24, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n8/pdf/scientificamerican08241895-
125c.pdf.

Anonymous:1895:PNPba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n8/pdf/scientificamerican08241895-127.pdf.

Anonymous:1895:AAA

[6018] Anonymous. American Association for the Advancement of Science. Scientific American, 73(9):130, August 31, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n9/pdf/scientificamerican08311895-130.pdf.

Anonymous:1895:LBY

[6019] Anonymous. The lessons of the Battle of the Yalu. Scientific American, 73(9):130–131, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
130a.pdf.

Anonymous:1895:GAB

[6020] Anonymous. The Gorham adjustable bed. Scientific American, 73(9):132, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
132a.pdf.

Anonymous:1895:IBBa

[6021] Anonymous. An improved bicycle brake. Scientific American, 73(9):132, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
132.pdf.

Anonymous:1895:ILa

[6022] Anonymous. An improved lubricator. Scientific American, 73 (9):132, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
132c.pdf.

Anonymous:1895:SAP

[6023] Anonymous. A subsoil attachment for plows. Scientific American, 73(9):132, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
132b.pdf.

Anonymous:1895:PACb

[6024] Anonymous. The proposed Atlantic Coastwise and Cape Cod Canals. Scientific American, 73(9):133, August 31, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n9/pdf/scientificamerican08311895-133.pdf.

Anonymous:1895:ECF

[6025] Anonymous. Edward C. F. Davis. Scientific American, 73(9):134, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n9/pdf/scientificamerican08311895-134.pdf.

Anonymous:1895:BS

[6026] Anonymous. The Boston Subway. Scientific American, 73(9):135, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n9/pdf/scientificamerican08311895-135.pdf.

Anonymous:1895:NIY

[6027] Anonymous. The Ninth International Yacht Race. Scientific American, 73(9):136–138, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
136.pdf.

Anonymous:1895:IHa

[6028] Anonymous. An improved harp. Scientific American, 73(9):138, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n9/pdf/scientificamerican08311895-138.pdf.

Anonymous:1895:CBf

[6029] Anonymous. The catastrophe in Brux. Scientific American, 73 (9):139–140, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
139.pdf.

Anonymous:1895:RPIbi

[6030] Anonymous. Recently patented inventions. Scientific American, 73(9):140, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
140.pdf.

Anonymous:1895:BNBba

[6031] Anonymous. Books: New books and publications. Scientific American, 73(9):140–141, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
140a.pdf.

Anonymous:1895:BPbi

[6032] Anonymous. Business and personal. Scientific American, 73(9):141, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n9/pdf/scientificamerican08311895-141.pdf.

Anonymous:1895:Ibh

[6033] Anonymous. To inventors. Scientific American, 73(9):141, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n9/pdf/scientificamerican08311895-141b.pdf.

Anonymous:1895:NQbi

[6034] Anonymous. Notes and queries. Scientific American, 73(9):141, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n9/pdf/scientificamerican08311895-141a.pdf.

Anonymous:1895:IIbl

[6035] Anonymous. Index of inventions. Scientific American, 73(9): 141–142, August 31, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n9/pdf/scientificamerican08311895-
141c.pdf.

Anonymous:1895:PNPbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n9/pdf/scientificamerican08311895-143.pdf.

Anonymous:1895:AMA

[6037] Anonymous. Annual meeting of the American Association for the Advancement of Science, and Affiliated Societies, at Springfield, Mass, August 29, 1895. Scientific American, 73(10):146, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n10/pdf/scientificamerican09071895-146.pdf.

Anonymous:1895:EWMa

[6038] Anonymous. Edward Williams Morley. Scientific American, 73 (10):147, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
147a.pdf.

Anonymous:1895:SSd

[6039] Anonymous. The stars in September. Scientific American, 73 (10):147, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
147.pdf.

Anonymous:1895:EWMb

[6040] Anonymous. An efficient wind motor. Scientific American, 73 (10):148, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
148a.pdf.

Anonymous:1895:ISBa

[6041] Anonymous. An improved sectional boiler. Scientific American, 73(10):148, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
148.pdf.

Anonymous:1895:NPT

[6042] Anonymous. New process for treating ores. Scientific American, 73(10):148, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
148b.pdf.

Anonymous:1895:THP

[6043] Anonymous. A trousers hanger and press. Scientific American, 73(10):148, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
148c.pdf.

Anonymous:1895:MMS

[6044] Anonymous. Manufacture of marbleized slate mantels. Scientific American, 73(10):149, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
149.pdf.

Anonymous:1895:CDT

[6045] Anonymous. The Creusot disappearing turret. Scientific American, 73(10):150, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
150.pdf.

Anonymous:1895:DAF

[6046] Anonymous. Dredging ane floating pipe system. Scientific American, 73(10):151–152, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
151.pdf.

Anonymous:1895:WPS

[6047] Anonymous. Water practice of the “Summer Pioneers” of the German Army. Scientific American, 73(10):152–154, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n10/pdf/scientificamerican09071895-152.pdf.

Anonymous:1895:CBFa

[6048] Anonymous. Church Bell Foundry in Cincinnati. Scientific American, 73(10):154, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
154.pdf.

Anonymous:1895:Cbg

[6049] Anonymous. Correspondence. Scientific American, 73(10):155, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n10/pdf/scientificamerican09071895-155.pdf.

Anonymous:1895:BNBbb

[6050] Anonymous. Books: New books and publications. Scientific American, 73(10):156, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
156a.pdf.

Anonymous:1895:RPIbj

[6051] Anonymous. Recently patented inventions. Scientific American, 73(10):156, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
156.pdf.

Anonymous:1895:BPbj

[6052] Anonymous. Business and personal. Scientific American, 73 (10):157, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
157.pdf.

Anonymous:1895:IIbm

[6053] Anonymous. Index of inventions. Scientific American, 73(10): 157, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
157b.pdf.

Anonymous:1895:NQbj

[6054] Anonymous. Notes and queries. Scientific American, 73(10): 157, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
157a.pdf.

Anonymous:1895:PNPbc

[6055] Anonymous. Patent notes: Patents. Scientific American, 73 (10):159, September 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n10/pdf/scientificamerican09071895-
159.pdf.

Anonymous:1895:PBe

[6056] Anonymous. The physics of the bicycle. Scientific American, 73 (11):162, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
162.pdf.

Anonymous:1895:PAA

[6057] Anonymous. Proceedings of the American Association at Springfield, Mass. Scientific American, 73(11):162–163, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n11/pdf/scientificamerican09141895-162a.pdf.

Anonymous:1895:BCb

[6058] Anonymous. The “Bristol” counter. Scientific American, 73 (11):164, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
164b.pdf.

Anonymous:1895:ECD

[6059] Anonymous. An efficient coffee drier. Scientific American, 73 (11):164, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
164.pdf.

Anonymous:1895:FML

[6060] Anonymous. Fernandez’s marine and land bicycle. Scientific American, 73(11):164, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
164c.pdf.

Anonymous:1895:MPC

[6061] Anonymous. A machine for packing cigarettes. Scientific American, 73(11):164, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
164a.pdf.

Anonymous:1895:BCc

[6062] Anonymous. Brass casting. Scientific American, 73(11):165, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n11/pdf/scientificamerican09141895-165.pdf.

Anonymous:1895:NHS

[6063] Anonymous. New High Service Pump, Boston Water Works. Scientific American, 73(11):166, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
166.pdf.

Anonymous:1895:Cbh

[6064] Anonymous. Correspondence. Scientific American, 73(11):167, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n11/pdf/scientificamerican09141895-167.pdf.

Anonymous:1895:EIRa

[6065] Anonymous. The evolution of the international racing yacht — I. Scientific American, 73(11):168–169, September 14, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n11/pdf/scientificamerican09141895-168.pdf.

Anonymous:1895:GCV

[6066] Anonymous. A great coal vein. Scientific American, 73(11): 169, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
169.pdf.

Anonymous:1895:WAF

[6067] Anonymous. The Washburne adjustable fasteners for watch chains, neckties, collars, etc. Scientific American, 73(11):170, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n11/pdf/scientificamerican09141895-170.pdf.

Anonymous:1895:PPR

[6068] Anonymous. The Pyramidical Pleasure Railway. Scientific American, 73(11):171–172, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
171.pdf.

Anonymous:1895:RPIbk

[6069] Anonymous. Recently patented inventions. Scientific American, 73 (11):172, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
172.pdf.

Anonymous:1895:BNBbc

[6070] Anonymous. Books: New books and publications. Scientific American, 73(11):172–173, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
172a.pdf.

Anonymous:1895:BPbk

[6071] Anonymous. Business and personal. Scientific American, 73 (11):173, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
173.pdf.

Anonymous:1895:Ibi

[6072] Anonymous. To inventors. Scientific American, 73(11):173, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n11/pdf/scientificamerican09141895-173b.pdf.

Anonymous:1895:NQbk

[6073] Anonymous. Notes and queries. Scientific American, 73(11): 173, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
173a.pdf.

Anonymous:1895:IIbn

[6074] Anonymous. Index of inventions. Scientific American, 73(11): 173–174, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
173c.pdf.

Anonymous:1895:PNPbd

[6075] Anonymous. Patent notes: Patents. Scientific American, 73 (11):175, September 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n11/pdf/scientificamerican09141895-
175.pdf.

Anonymous:1895:ACRa

[6076] Anonymous. The America Cup Races of 1895. Scientific American, 73(12):178, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
178.pdf.

Anonymous:1895:HSR

[6077] Anonymous. High speed railway records. Scientific American, 73 (12):178, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
178a.pdf.

Anonymous:1895:FBM

[6078] Anonymous. Firing of boilers with mixed coal. Scientific American, 73(12):178–179, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
178b.pdf.

Anonymous:1895:STL

[6079] Anonymous. Speeding truck for locomotives. Scientific American, 73(12):179, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
179.pdf.

Anonymous:1895:MHC

[6080] Anonymous. The manufacture of Horn combs. Scientific American, 73(12):180, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
180.pdf.

Anonymous:1895:FIC

[6081] Anonymous. A folding invalid chair. Scientific American, 73 (12):181, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
181a.pdf.

Anonymous:1895:GSA

[6082] Anonymous. A gold saving apparatus. Scientific American, 73 (12):181, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
181c.pdf.

Anonymous:1895:PEP

[6083] Anonymous. A portable electric propeller for boats. Scientific American, 73(12):181, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
181b.pdf.

Anonymous:1895:SEM

[6084] Anonymous. A stop for elevators and mining shafts. Scientific American, 73(12):181, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
181.pdf.

Anonymous:1895:BLD

[6085] Anonymous. A boiler leveling device. Scientific American, 73(12): 181–182, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
181d.pdf.

Anonymous:1895:Cbi

[6086] Anonymous. Correspondence. Scientific American, 73(12):182, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n12/pdf/scientificamerican09211895-182a.pdf.

Michaud:1895:EHS

[6087] Gustave Michaud. An experiment with hydrogen sulphide. Scientific American, 73(12):182, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
182.pdf.

Anonymous:1895:FHI

[6088] Anonymous. Fine hammered iron work. Scientific American, 73 (12):183, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
183.pdf.

Anonymous:1895:EIRb

[6089] Anonymous. The evolution of the International Racing Yacht — II. Scientific American, 73(12):184–188, September 21, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n12/pdf/scientificamerican09211895-184.pdf.

Anonymous:1895:DRPg

[6090] Anonymous. Decisions relating to patents. Scientific American, 73 (12):188, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
188.pdf.

Anonymous:1895:BNBbd

[6091] Anonymous. Books: New books and publications. Scientific American, 73(12):189, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
189a.pdf.

Anonymous:1895:BPbl

[6092] Anonymous. Business and personal. Scientific American, 73 (12):189, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
189b.pdf.

Anonymous:1895:Ibj

[6093] Anonymous. To inventors. Scientific American, 73(12):189, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n12/pdf/scientificamerican09211895-189d.pdf.

Anonymous:1895:NQbl

[6094] Anonymous. Notes and queries. Scientific American, 73(12): 189, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
189c.pdf.

Anonymous:1895:RPIbl

[6095] Anonymous. Recently patented inventions. Scientific American, 73 (12):189, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
189.pdf.

Anonymous:1895:IIbo

[6096] Anonymous. Index of inventions. Scientific American, 73(12): 190, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
190.pdf.

Anonymous:1895:PNPbe

[6097] Anonymous. Patent notes: Patents. Scientific American, 73 (12):191, September 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n12/pdf/scientificamerican09211895-
191.pdf.

Anonymous:1895:PCM

[6098] Anonymous. Pennsylvania coal mining. Scientific American, 73 (13):193, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
193.pdf.

Anonymous:1895:OCS

[6099] Anonymous. The opening of the Cotton States and International Exposition. Scientific American, 73(13):194, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n13/pdf/scientificamerican09281895-194.pdf.

Anonymous:1895:SCB

[6100] Anonymous. The second-class battleship Maine. Scientific American, 73(13):194, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
194a.pdf.

Anonymous:1895:CCBa

[6101] Anonymous. The collapse of the Coliseum Building, Chicago. Scientific American, 73(13):194–195, September 28, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n13/pdf/scientificamerican09281895-194b.pdf.

Anonymous:1895:EPC

[6102] Anonymous. An efficient passenger car ventilator. Scientific American, 73(13):196, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
196c.pdf.

Anonymous:1895:EST

[6103] Anonymous. An electrical signal transmitter. Scientific American, 73(13):196, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
196a.pdf.

Anonymous:1895:PD

[6104] Anonymous. A portable Derrick. Scientific American, 73(13): 196, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
196b.pdf.

Anonymous:1895:SIC

[6105] Anonymous. A simple and inexpensive car coupling. Scientific American, 73(13):196, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
196.pdf.

Anonymous:1895:MPR

[6106] Anonymous. Morton’s perambulator, rocker, etc. Scientific American, 73(13):196–197, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
196d.pdf.

Anonymous:1895:EHPa

[6107] Anonymous. An efficient horse power and pump. Scientific American, 73(13):197, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
197a.pdf.

Anonymous:1895:LSV

[6108] Anonymous. The largest sailing vessel in the world. Scientific American, 73(13):197, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
197.pdf.

Anonymous:1895:CVRb

[6109] Anonymous. Charles Valentine Riley. Scientific American, 73(13): 198–199, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
198.pdf.

Anonymous:1895:MPd

[6110] Anonymous. A mammoth potato. Scientific American, 73(13): 199, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
199.pdf.

Anonymous:1895:IYR

[6111] Anonymous. International yacht races of 1895. Scientific American, 73(13):200, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
200.pdf.

Anonymous:1895:GBS

[6112] Anonymous. The gigantic birds of Southern Patagonia. Scientific American, 73(13):202–204, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
202.pdf.

Anonymous:1895:RPIbm

[6113] Anonymous. Recently patented inventions. Scientific American, 73 (13):204, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
204.pdf.

Anonymous:1895:BNBbe

[6114] Anonymous. Books: New books and publications. Scientific American, 73(13):204–205, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
204a.pdf.

Anonymous:1895:BPbm

[6115] Anonymous. Business and personal. Scientific American, 73 (13):205, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
205.pdf.

Anonymous:1895:Ibk

[6116] Anonymous. To inventors. Scientific American, 73(13):205, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n13/pdf/scientificamerican09281895-205b.pdf.

Anonymous:1895:NQbm

[6117] Anonymous. Notes and queries. Scientific American, 73(13): 205, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
205a.pdf.

Anonymous:1895:IIbp

[6118] Anonymous. Index of inventions. Scientific American, 73(13): 205–206, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
205c.pdf.

Anonymous:1895:PNPbf

[6119] Anonymous. Patent notes: Patents. Scientific American, 73 (13):207, September 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n13/pdf/scientificamerican09281895-
207.pdf.

Anonymous:1895:LF

[6120] Anonymous. The locomotive of the future. Scientific American, 73(14):210, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
210.pdf.

Anonymous:1895:WPT

[6121] Anonymous. The World’s production of tin. Scientific American, 73(14):210–211, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
210a.pdf.

Anonymous:1895:ETS

[6122] Anonymous. Electric train “stopping” and “starting” signal. Scientific American, 73(14):212, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
212.pdf.

Anonymous:1895:GPB

[6123] Anonymous. Gatti’s prying bar for handling lumber. Scientific American, 73(14):212, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
212b.pdf.

Anonymous:1895:PPC

[6124] Anonymous. A paper and pasteboard cutter. Scientific American, 73(14):212, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
212a.pdf.

Anonymous:1895:ETC

[6125] Anonymous. An electric time check receiver. Scientific American, 73(14):212–213, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
212c.pdf.

Anonymous:1895:ALS

[6126] Anonymous. The automatic life saving apparatus of De Ropp. Scientific American, 73(14):213, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
213.pdf.

Anonymous:1895:ASC

[6127] Anonymous. An Archimedean screw constructed from simple materials. Scientific American, 73(14):213, October 5, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n14/pdf/scientificamerican10051895-213a.pdf.

Anonymous:1895:HC

[6128] Anonymous. A horseless carriage of 1827. Scientific American, 73(14):214, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
214.pdf.

Anonymous:1895:OER

[6129] Anonymous. The Otis Elevating Railway. Scientific American, 73(14):215, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
215.pdf.

Anonymous:1895:MWS

[6130] Anonymous. Manufacture of wagon springs. Scientific American, 73(14):216, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
216.pdf.

Anonymous:1895:ISFc

[6131] Anonymous. Improved steel frame stamp battery. Scientific American, 73(14):217–218, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
217.pdf.

Anonymous:1895:Cbj

[6132] Anonymous. Correspondence. Scientific American, 73(14):218–219, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n14/pdf/scientificamerican10051895-218.pdf.

Anonymous:1895:MBa

[6133] Anonymous. The magnetic bicycle. Scientific American, 73(14): 219–220, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
219.pdf.

Anonymous:1895:Ibl

[6134] Anonymous. To inventors. Scientific American, 73(14):220, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n14/pdf/scientificamerican10051895-220a.pdf.

Anonymous:1895:RPIbn

[6135] Anonymous. Recently patented inventions. Scientific American, 73(14):220, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n14/pdf/scientificamerican10051895-
220.pdf.

Anonymous:1895:BPbn

[6136] Anonymous. Business and personal. Scientific American, 73(14):221, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n14/pdf/scientificamerican10051895-221.pdf.

Anonymous:1895:IIbq

[6137] Anonymous. Index of inventions. Scientific American, 73(14):221, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n14/pdf/scientificamerican10051895-221b.pdf.

Anonymous:1895:NQbn

[6138] Anonymous. Notes and queries. Scientific American, 73(14):221, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n14/pdf/scientificamerican10051895-221a.pdf.

Anonymous:1895:PNPbg

[6139] Anonymous. Patent notes: Patents. Scientific American, 73(14):223, October 5, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n14/pdf/scientificamerican10051895-223.pdf.

Anonymous:1895:P

[6140] Anonymous. Pasteur. Scientific American, 73(15):225, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n15/pdf/scientificamerican10121895-225.pdf.

Anonymous:1895:NMG

[6141] Anonymous. The new Maxim gun of solid steel. Scientific American, 73(15):226, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
226a.pdf.

Anonymous:1895:RLP

[6142] Anonymous. The return of Lieutenant Peary. Scientific American, 73(15):226, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
226.pdf.

Anonymous:1895:OPC

[6143] Anonymous. The Official Photographic Company of the Cotton States Exposition. Scientific American, 73(15):226–227, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n15/pdf/scientificamerican10121895-226b.pdf.

Anonymous:1895:AKO

[6144] Anonymous. The authorship of the kerosene oil spray. Scientific American, 73(15):227, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
227a.pdf.

Anonymous:1895:RFC

[6145] Anonymous. The return of Faye’s comet. Scientific American, 73(15):227, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
227.pdf.

Anonymous:1895:DHS

[6146] Anonymous. A detachable hinge for screen doors, etc. Scientific American, 73(15):228, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
228.pdf.

Anonymous:1895:MEO

[6147] Anonymous. Microscope for the examination of opaque objects. Scientific American, 73(15):228, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
228a.pdf.

Anonymous:1895:SPB

[6148] Anonymous. Steam pipe and boiler covering. Scientific American, 73(15):228, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
228b.pdf.

Anonymous:1895:WFS

[6149] Anonymous. Welding and forging steel. Scientific American, 73 (15):229–230, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
229.pdf.

Anonymous:1895:Cbk

[6150] Anonymous. Correspondence. Scientific American, 73(15):231, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n15/pdf/scientificamerican10121895-231.pdf.

Anonymous:1895:ADC

[6151] Anonymous. An automatic double chisel mortising machine. Scientific American, 73(15):232, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
232.pdf.

Anonymous:1895:LAC

[6152] Anonymous. The launch of the armored cruiser Brooklyn. Scientific American, 73(15):232–234, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
232a.pdf.

Anonymous:1895:PTc

[6153] Anonymous. A petroleum tricycle. Scientific American, 73(15):234, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n15/pdf/scientificamerican10121895-234.pdf.

Anonymous:1895:RNH

[6154] Anonymous. The Royal Natural History. Scientific American, 73 (15):235–236, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
235.pdf.

Anonymous:1895:RPIbo

[6155] Anonymous. Recently patented inventions. Scientific American, 73 (15):236–237, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
236.pdf.

Anonymous:1895:BNBbf

[6156] Anonymous. Books: New books and publications. Scientific American, 73(15):237, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
237.pdf.

Anonymous:1895:BPbo

[6157] Anonymous. Business and personal. Scientific American, 73 (15):237, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
237a.pdf.

Anonymous:1895:NQbo

[6158] Anonymous. Notes and queries. Scientific American, 73(15):237, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n15/pdf/scientificamerican10121895-237b.pdf.

Anonymous:1895:IIbr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
237c.pdf.

Anonymous:1895:PNPbh

[6160] Anonymous. Patent notes: Patents. Scientific American, 73 (15):239, October 12, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n15/pdf/scientificamerican10121895-
239.pdf.

Anonymous:1895:EHPb

[6161] Anonymous. The economy of hydraulic power supply. Scientific American, 73(16):242, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
242a.pdf.

Anonymous:1895:PEb

[6162] Anonymous. The Paris Exhibition of 1900. Scientific American, 73(16):242, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
242.pdf.

Anonymous:1895:CTHa

[6163] Anonymous. The Chicago Times–Herald Motocycle Contest. Scientific American, 73(16):242–243, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
242b.pdf.

Anonymous:1895:MSW

[6164] Anonymous. Manufacture of sash weights from old tin cans. Scientific American, 73(16):244–245, October 19, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n16/pdf/scientificamerican10191895-244.pdf.

Anonymous:1895:CPW

[6165] Anonymous. The Carter pressure water filter. Scientific American, 73(16):245, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
245.pdf.

Anonymous:1895:IPJ

[6166] Anonymous. An improved pipe joint. Scientific American, 73 (16):245, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
245a.pdf.

Anonymous:1895:NTB

[6167] Anonymous. The new terminals of the Brooklyn Bridge. Scientific American, 73(16):246, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
246.pdf.

Anonymous:1895:SPL

[6168] Anonymous. A scarf pin locking device. Scientific American, 73 (16):246–248, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
246a.pdf.

Anonymous:1895:GAT

[6169] Anonymous. Gyrograph or artistic top. Scientific American, 73 (16):248, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
248a.pdf.

Anonymous:1895:ILR

[6170] Anonymous. Improved “Long Run” perfecting printing press. Scientific American, 73(16):248, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
248.pdf.

Anonymous:1895:NFMa

[6171] Anonymous. New flying machine. Scientific American, 73(16): 249–250, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
249.pdf.

Anonymous:1895:SDG

[6172] Anonymous. Sir Douglas Galton, President of the British Association. Scientific American, 73(16):250, October 19, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n16/pdf/scientificamerican10191895-250.pdf.

Anonymous:1895:TZG

[6173] Anonymous. Turtles in the Zoological Garden at Hamburg. Scientific American, 73(16):251–252, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
251.pdf.

Anonymous:1895:RPIbp

[6174] Anonymous. Recently patented inventions. Scientific American, 73 (16):252–253, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
252.pdf.

Anonymous:1895:BNBbg

[6175] Anonymous. Books: New books and publications. Scientific American, 73(16):253, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
253.pdf.

Anonymous:1895:BPbp

[6176] Anonymous. Business and personal. Scientific American, 73 (16):253, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
253a.pdf.

Anonymous:1895:Ibm

[6177] Anonymous. To inventors. Scientific American, 73(16):253, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n16/pdf/scientificamerican10191895-253c.pdf.

Anonymous:1895:NQbp

[6178] Anonymous. Notes and queries. Scientific American, 73(16):253, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n16/pdf/scientificamerican10191895-253b.pdf.

Anonymous:1895:IIbs

[6179] Anonymous. Index of inventions. Scientific American, 73(16): 253–254, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
253d.pdf.

Anonymous:1895:PNPbi

[6180] Anonymous. Patent notes: Patents. Scientific American, 73 (16):255, October 19, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n16/pdf/scientificamerican10191895-
255.pdf.

Anonymous:1895:MER

[6181] Anonymous. A model electric railway plant. Scientific American, 73(17):257, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
257.pdf.

Anonymous:1895:PNR

[6182] Anonymous. The proposed North River Bridge — the greatest engineering undertaking in the world. Scientific American, 73(17):258, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n17/pdf/scientificamerican10261895-258a.pdf.

Anonymous:1895:PRW

[6183] Anonymous. Prevention of rust in wheat. Scientific American, 73(17):258, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
258.pdf.

Anonymous:1895:BSI

[6184] Anonymous. The battle ship Indiana. Scientific American, 73 (17):258–259, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
258b.pdf.

Anonymous:1895:AF

[6185] Anonymous. Antique furniture. Scientific American, 73(17):260, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n17/pdf/scientificamerican10261895-260a.pdf.

Anonymous:1895:APR

[6186] Anonymous. An automatic pressure regulating valve. Scientific American, 73(17):260, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
260b.pdf.

Anonymous:1895:SAE

[6187] Anonymous. A safety attachment for elevators. Scientific American, 73(17):260, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
260.pdf.

Anonymous:1895:TCS

[6188] Anonymous. The time in China by the sun, water, and fire. Scientific American, 73(17):261–262, October 26, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n17/pdf/scientificamerican10261895-261.pdf.

Anonymous:1895:BRS

[6189] Anonymous. Bicycle racing at Springfield. Scientific American, 73(17):262, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
262.pdf.

Anonymous:1895:CCf

[6190] Anonymous. Correspondence: Correnpondence. Scientific American, 73(17):263–264, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
263.pdf.

Anonymous:1895:BNBbh

[6191] Anonymous. Books: New books and publications. Scientific American, 73(17):269, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
269c.pdf.

Anonymous:1895:BPbq

[6192] Anonymous. Business and personal. Scientific American, 73 (17):269, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
269a.pdf.

Anonymous:1895:Ibn

[6193] Anonymous. To inventors. Scientific American, 73(17):269, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n17/pdf/scientificamerican10261895-269d.pdf.

Anonymous:1895:NQbq

[6194] Anonymous. Notes and queries. Scientific American, 73(17):269, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n17/pdf/scientificamerican10261895-269b.pdf.

Anonymous:1895:RPIbq

[6195] Anonymous. Recently patented inventions. Scientific American, 73(17):269, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
269.pdf.

Anonymous:1895:IIbt 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
269e.pdf.

Anonymous:1895:PNPbj

[6197] Anonymous. Patent notes: Patents. Scientific American, 73 (17):271, October 26, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n17/pdf/scientificamerican10261895-
271.pdf.

Anonymous:1895:AEb

[6198] Anonymous. The Atlanta Exhibition. Scientific American, 73 (18):273, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
273.pdf.

Anonymous:1895:GAA

[6199] Anonymous. The great avalanche of the Altels Glacier. Scientific American, 73(18):274, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
274b.pdf.

Anonymous:1895:ISc

[6200] Anonymous. The isles of shoals. Scientific American, 73(18): 274, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
274.pdf.

Anonymous:1895:RAS

[6201] Anonymous. The railroad across Siberia. Scientific American, 73 (18):274, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
274a.pdf.

Anonymous:1895:LDT

[6202] Anonymous. Long-distance transmission of water power. Scientific American, 73(18):274–275, November 2, 1895. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
274c.pdf.

Anonymous:1895:HN

[6203] Anonymous. The heavens in November. Scientific American, 73 (18):275, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
275.pdf.

Anonymous:1895:ECP

[6204] Anonymous. An eighteenth century project for a huge cannon. Scientific American, 73(18):276, November 2, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n18/pdf/scientificamerican11021895-276a.pdf.

Anonymous:1895:ICS

[6205] Anonymous. An improved cask storing rack. Scientific American, 73(18):276, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
276b.pdf.

Anonymous:1895:MSM

[6206] Anonymous. A machine for sawing moulding strips. Scientific American, 73(18):276, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
276.pdf.

Anonymous:1895:SGI

[6207] Anonymous. Silk growing in India. Scientific American, 73(18): 277–280, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
277.pdf.

Anonymous:1895:FHW

[6208] Anonymous. The Fair Haven and Westville Railroad Plant. Scientific American, 73(18):281–282, November 2, 1895. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
281.pdf.

Anonymous:1895:TR

[6209] Anonymous. The Transsiberian Railway. Scientific American, 73 (18):282–283, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
282.pdf.

Anonymous:1895:BNBbi

[6210] Anonymous. Books: New books and publications. Scientific American, 73(18):284, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
284a.pdf.

Anonymous:1895:BPbr

[6211] Anonymous. Business and personal. Scientific American, 73 (18):284, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
284b.pdf.

Anonymous:1895:RPIbr

[6212] Anonymous. Recently patented inventions. Scientific American, 73(18):284, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
284.pdf.

Anonymous:1895:NQbr

[6213] Anonymous. Notes and queries. Scientific American, 73(18): 284–285, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
284c.pdf.

Anonymous:1895:Ibp

[6214] Anonymous. To inventors. Scientific American, 73(18):285, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n18/pdf/scientificamerican11021895-285.pdf.

Anonymous:1895:IIbu

[6215] Anonymous. Index of inventions. Scientific American, 73(18): 285, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
285a.pdf.

Anonymous:1895:PNPbk

[6216] Anonymous. Patent notes: Patents. Scientific American, 73 (18):287, November 2, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n18/pdf/scientificamerican11021895-
287.pdf.

Anonymous:1895:AAE

[6217] Anonymous. The Australian Antarctic Expedition. Scientific American, 73(19):290, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
290.pdf.

Anonymous:1895:ESB

[6218] Anonymous. Exhaustive steam boiler experiments. Scientific American, 73(19):290, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
290a.pdf.

Anonymous:1895:CCBb

[6219] Anonymous. Chromatophores, or the color-bestowing cells of animals. Scientific American, 73(19):290–291, November 9, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n19/pdf/scientificamerican11091895-290b.pdf.

Anonymous:1895:BBF

[6220] Anonymous. A bell for blast furnaces. Scientific American, 73 (19):292, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
292.pdf.

Anonymous:1895:GFA

[6221] Anonymous. A giant of former ages. Scientific American, 73 (19):292, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
292a.pdf.

Anonymous:1895:ISBb

[6222] Anonymous. An improvement on the Stone boat or drag. Scientific American, 73(19):292, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
292b.pdf.

Anonymous:1895:PMR

[6223] Anonymous. Postponement of the motocycle race. Scientific American, 73(19):293–294, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
293.pdf.

Anonymous:1895:Cbl

[6224] Anonymous. Correspondence. Scientific American, 73(19):295, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n19/pdf/scientificamerican11091895-295.pdf.

Anonymous:1895:TUS

[6225] Anonymous. Trial of United States battle ship Iowa’s belt armor. Scientific American, 73(19):296, November 9, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n19/pdf/scientificamerican11091895-296.pdf.

Anonymous:1895:BPLa

[6226] Anonymous. The Boston Public Library. Scientific American, 73 (19):297–298, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
297.pdf.

Anonymous:1895:MW

[6227] Anonymous. Modern wrestling. Scientific American, 73(19):299– 300, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
299.pdf.

Anonymous:1895:RPIbs

[6228] Anonymous. Recently patented inventions. Scientific American, 73(19):300, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
300.pdf.

Anonymous:1895:BNBbj

[6229] Anonymous. Books: New books and publications. Scientific American, 73(19):301, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
301.pdf.

Anonymous:1895:BPbs

[6230] Anonymous. Business and personal. Scientific American, 73 (19):301, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
301a.pdf.

Anonymous:1895:NQbs

[6231] Anonymous. Notes and queries. Scientific American, 73(19): 301, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
301b.pdf.

Anonymous:1895:IIbv

[6232] Anonymous. Index of inventions. Scientific American, 73(19): 301–302, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
301c.pdf.

Anonymous:1895:PNPbl

[6233] Anonymous. Patent notes: Patents. Scientific American, 73 (19):303, November 9, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n19/pdf/scientificamerican11091895-
303.pdf.

Anonymous:1895:AEc

[6234] Anonymous. The Atlanta Exposition. Scientific American, 73(20): 305–306, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
305.pdf.

Anonymous:1895:TCR

[6235] Anonymous. Thomas C. Roche. Scientific American, 73(20): 306, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
306.pdf.

Anonymous:1895:PMB

[6236] Anonymous. The permanence of modern buildings and public works. Scientific American, 73(20):306–307, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n20/pdf/scientificamerican11161895-306a.pdf.

Anonymous:1895:ALC

[6237] Anonymous. An automatic letter copier. Scientific American, 73 (20):308, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
308b.pdf.

Anonymous:1895:AUC

[6238] Anonymous. An apparatus for unloading cars. Scientific American, 73(20):308, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
308a.pdf.

Anonymous:1895:GTT

[6239] Anonymous. A galvanizing and tempering tank. Scientific American, 73(20):308, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
308c.pdf.

Anonymous:1895:IVC

[6240] Anonymous. An improved vine cutter. Scientific American, 73 (20):308, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
308.pdf.

Anonymous:1895:NCM

[6241] Anonymous. A novel cutting off machine. Scientific American, 73(20):309, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
309.pdf.

Anonymous:1895:LI

[6242] Anonymous. The lac industry. Scientific American, 73(20):309– 310, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
309a.pdf.

Anonymous:1895:BM

[6243] Anonymous. A buttonhole moistener. Scientific American, 73 (20):310, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
310.pdf.

Anonymous:1895:ICW

[6244] Anonymous. An improved car window sash. Scientific American, 73 (20):310–311, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
310a.pdf.

Anonymous:1895:CCDa

[6245] Anonymous. A coal car dumping machine. Scientific American, 73(20):312, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
312.pdf.

Anonymous:1895:RTB

[6246] Anonymous. The Russian torpedo boat destroyer Sokol. Scientific American, 73(20):313–314, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
313.pdf.

Anonymous:1895:MR

[6247] Anonymous. The motocycle race. Scientific American, 73(20): 315–316, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
315.pdf.

Anonymous:1895:BPbt

[6248] Anonymous. Business and personal. Scientific American, 73 (20):317, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
317a.pdf.

Anonymous:1895:NQbt

[6249] Anonymous. Notes and queries. Scientific American, 73(20): 317, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
317b.pdf.

Anonymous:1895:RPIbt

[6250] Anonymous. Recently patented inventions. Scientific American, 73(20):317, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
317.pdf.

Anonymous:1895:IIbw

[6251] Anonymous. Index of inventions. Scientific American, 73(20): 318, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
318.pdf.

Anonymous:1895:PNPbm

[6252] Anonymous. Patent notes: Patents. Scientific American, 73 (20):319, November 16, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n20/pdf/scientificamerican11161895-
319.pdf.

Anonymous:1895:AEd

[6253] Anonymous. The Atlanta Exposition. Scientific American, 73 (22):337, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
337.pdf.

Anonymous:1895:BCR

[6254] Anonymous. The bottle that cannot be refilled. Scientific American, 73(22):338, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
338.pdf.

Anonymous:1895:LFV

[6255] Anonymous. Latitude not fixed, but variable. Scientific American, 73(22):338, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
338a.pdf.

Anonymous:1895:TLb

[6256] Anonymous. The temperature of lakes. Scientific American, 73 (22):338–339, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
338b.pdf.

Anonymous:1895:HD

[6257] Anonymous. The heavens in December. Scientific American, 73 (22):339, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
339.pdf.

Anonymous:1895:FRA

[6258] Anonymous. A French railway accident. Scientific American, 73 (22):340, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
340.pdf.

Anonymous:1895:IRC

[6259] Anonymous. An improved railway car jack. Scientific American, 73(22):341, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
341c.pdf.

Anonymous:1895:IWb

[6260] Anonymous. An improved windmill. Scientific American, 73 (22):341, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
341a.pdf.

Anonymous:1895:SEI

[6261] Anonymous. A steam engine indicator stop motion. Scientific American, 73(22):341, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
341.pdf.

Anonymous:1895:TPC

[6262] Anonymous. A tube or pipe cutter. Scientific American, 73 (22):341, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
341b.pdf.

Anonymous:1895:NPC

[6263] Anonymous. The new Port of Copenhagen. Scientific American, 73 (22):342–343, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
342.pdf.

Anonymous:1895:SSP

[6264] Anonymous. The steamer St. Paul. Scientific American, 73 (22):345, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
345.pdf.

Anonymous:1895:SRG

[6265] Anonymous. A school room gymnasium. Scientific American, 73 (22):345–346, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
345a.pdf.

Anonymous:1895:SCH

[6266] Anonymous. Street car hooks for bicycles. Scientific American, 73(22):347, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
347a.pdf.

Anonymous:1895:TB

[6267] Anonymous. A tree bridge. Scientific American, 73(22):347, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n22/pdf/scientificamerican11301895-347.pdf.

Anonymous:1895:BNBbk

[6268] Anonymous. Books: New books and publications. Scientific American, 73(22):348, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
348a.pdf.

Anonymous:1895:RPIbu

[6269] Anonymous. Recently patented inventions. Scientific American, 73(22):348, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
348.pdf.

Anonymous:1895:BPbu

[6270] Anonymous. Business and personal. Scientific American, 73 (22):349, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
349.pdf.

Anonymous:1895:Ibq

[6271] Anonymous. To inventors. Scientific American, 73(22):349, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n22/pdf/scientificamerican11301895-349b.pdf.

Anonymous:1895:IIbx

[6272] Anonymous. Index of inventions. Scientific American, 73(22): 349, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
349c.pdf.

Anonymous:1895:NQbu

[6273] Anonymous. Notes and queries. Scientific American, 73(22): 349, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
349a.pdf.

Anonymous:1895:PNPbn

[6274] Anonymous. Patent notes: Patents. Scientific American, 73 (22):351, November 30, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n22/pdf/scientificamerican11301895-
351.pdf.

Anonymous:1895:CSM

[6275] Anonymous. The chain saw mortiser. Scientific American, 73 (23):353, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
353.pdf.

Anonymous:1895:IBHa

[6276] Anonymous. Improvement of Boston Harbor. Scientific American, 73(23):354, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
354.pdf.

Anonymous:1895:UWM

[6277] Anonymous. The utilization of wind as a motive force. Scientific American, 73(23):354, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
354a.pdf.

Anonymous:1895:CTHb

[6278] Anonymous. The Chicago Times–Herald Motor Race. Scientific American, 73(23):354–355, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
354b.pdf.

Anonymous:1895:ISCa

[6279] Anonymous. An improved steam condenser. Scientific American, 73(23):357, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
357a.pdf.

Anonymous:1895:LHC

[6280] Anonymous. A lightning and heavy current arrester. Scientific American, 73(23):357, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
357.pdf.

Anonymous:1895:CMR

[6281] Anonymous. The Chicago motocycle race. Scientific American, 73 (23):357–358, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
357b.pdf.

Anonymous:1895:KRD

[6282] Anonymous. A key retaining device. Scientific American, 73 (23):358, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
358.pdf.

Anonymous:1895:LPP

[6283] Anonymous. The Leber patent portable fire escape. Scientific American, 73(23):358, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
358a.pdf.

Anonymous:1895:Cbm

[6284] Anonymous. Correspondence. Scientific American, 73(23):359, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n23/pdf/scientificamerican12071895-359.pdf.

Anonymous:1895:AGE

[6285] Anonymous. The acetylene gas exhibit at the Atlanta Exposition. Scientific American, 73(23):360, December 7, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n23/pdf/scientificamerican12071895-360.pdf.

Anonymous:1895:IIBy

[6286] Anonymous. The iron industries of Birmingham, Ala. Scientific American, 73(23):361–364, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
361.pdf.

Anonymous:1895:BNBbl

[6287] Anonymous. Books: New books and publications. Scientific American, 73(23):364, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
364a.pdf.

Anonymous:1895:RPIbv

[6288] Anonymous. Recently patented inventions. Scientific American, 73(23):364, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
364.pdf.

Anonymous:1895:BPbv

[6289] Anonymous. Business and personal. Scientific American, 73 (23):365, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n23/pdf/scientificamerican12071895-
365.pdf.

Anonymous:1895:Ibt

[6290] Anonymous. To inventors. Scientific American, 73(23):365, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n23/pdf/scientificamerican12071895-365b.pdf.

Anonymous:1895:IIbz

[6291] Anonymous. Index of inventions. Scientific American, 73(23):365, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n23/pdf/scientificamerican12071895-365c.pdf.

Anonymous:1895:NQbv

[6292] Anonymous. Notes and queries. Scientific American, 73(23):365, December 7, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n23/pdf/scientificamerican12071895-365a.pdf.

Anonymous:1895:MA

[6293] Anonymous. The motocycle award. Scientific American, 73(24): 370, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
370.pdf.

Anonymous:1895:SUS

[6294] Anonymous. Strikes in the United States and Europe. Scientific American, 73(24):370–371, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
370a.pdf.

Anonymous:1895:CRH

[6295] Anonymous. A clothes rack for household use. Scientific American, 73(24):372, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
372a.pdf.

Anonymous:1895:ILMb

[6296] Anonymous. An improved lawn mower. Scientific American, 73 (24):372, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
372b.pdf.

Anonymous:1895:MEF

[6297] Anonymous. A monument to Ericsson, Filipstad, Sweden. Scientific American, 73(24):372, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
372.pdf.

Anonymous:1895:HFW

[6298] Anonymous. A holder to facilitate work on boots or shoes. Scientific American, 73(24):373, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
373.pdf.

Anonymous:1895:AHR

[6299] Anonymous. The art of horseback riding. Scientific American, 73 (24):373–374, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
373a.pdf.

Anonymous:1895:ITT

[6300] Anonymous. An improved telephone transmitter. Scientific American, 73(24):374, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
374a.pdf.

Anonymous:1895:REb

[6301] Anonymous. A reinforced envelope. Scientific American, 73(24): 374, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
374.pdf.

Anonymous:1895:Cbn

[6302] Anonymous. Correspondence. Scientific American, 73(24):375, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n24/pdf/scientificamerican12141895-375a.pdf.

Anonymous:1895:WBA

[6303] Anonymous. The Womans Building, Atlanta, Ga. Scientific American, 73(24):375, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
375.pdf.

Anonymous:1895:BNc

[6304] Anonymous. Battleships nos. 5 and 8. Scientific American, 73 (24):376–377, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
376.pdf.

Anonymous:1895:VMD

[6305] Anonymous. The De La Vergne motor drag. Scientific American, 73(24):377, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
377.pdf.

Anonymous:1895:MBb

[6306] Anonymous. Manufacture of hat blocks. Scientific American, 73 (24):378, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
378.pdf.

Anonymous:1895:BHF

[6307] Anonymous. The blacksmith and his forge. Scientific American, 73 (24):379–380, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
379.pdf.

Anonymous:1895:RPIbw

[6308] Anonymous. Recently patented inventions. Scientific American, 73(24):380, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
380.pdf.

Anonymous:1895:BNBbm

[6309] Anonymous. Books: New books and publications. Scientific American, 73(24):380–381, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
380a.pdf.

Anonymous:1895:BPbw

[6310] Anonymous. Business and personal. Scientific American, 73 (24):381, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
381.pdf.

Anonymous:1895:Ibu

[6311] Anonymous. To inventors. Scientific American, 73(24):381, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n24/pdf/scientificamerican12141895-381b.pdf.

Anonymous:1895:NQbw

[6312] Anonymous. Notes and queries. Scientific American, 73(24): 381, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
381a.pdf.

Anonymous:1895:IIca

[6313] Anonymous. Index of inventions. Scientific American, 73(24): 381–382, December 14, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n24/pdf/scientificamerican12141895-
381c.pdf.

Anonymous:1895:PPD

[6314] Anonymous. Private and public debt in the United States. Scientific American, 73(25):386, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
386.pdf.

Anonymous:1895:NYB

[6315] Anonymous. New York the birthplace of ocean steam navigation. Scientific American, 73(25):386–387, December 21, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n25/pdf/scientificamerican12211895-386a.pdf.

Anonymous:1895:IBBb

[6316] Anonymous. An improved bicycle brake. Scientific American, 73 (25):388, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
388b.pdf.

Anonymous:1895:MTP

[6317] Anonymous. A metallic tread pneumatic tire. Scientific American, 73(25):388, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
388.pdf.

Anonymous:1895:NFMb

[6318] Anonymous. Najork’s foot motor boat. Scientific American, 73 (25):388, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
388a.pdf.

Anonymous:1895:EI

[6319] Anonymous. An electric incubator. Scientific American, 73(25): 389, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
389.pdf.

Anonymous:1895:PRb

[6320] Anonymous. A pole railway. Scientific American, 73(25):389– 390, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
389a.pdf.

Anonymous:1895:ACRb

[6321] Anonymous. “Air Cushion” rubber printing stamps. Scientific American, 73(25):390, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
390a.pdf.

Anonymous:1895:AVBb

[6322] Anonymous. An automatic vehicle brake. Scientific American, 73 (25):390, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
390.pdf.

Anonymous:1895:Cbo

[6323] Anonymous. Correspondence. Scientific American, 73(25):391, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n25/pdf/scientificamerican12211895-391a.pdf.

Anonymous:1895:IIBba

[6324] Anonymous. The iron industries of Birmingham, Ala. Scientific American, 73(25):391, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
391.pdf.

Anonymous:1895:EHO

[6325] Anonymous. The early history of ocean steam navigation. Scientific American, 73(25):392–393, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
392.pdf.

Anonymous:1895:AEe

[6326] Anonymous. The Atlanta Exposition. Scientific American, 73(25): 393–394, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
393.pdf.

Anonymous:1895:ERS

[6327] Anonymous. The electric railway street sprinkler. Scientific American, 73(25):394, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
394.pdf.

Anonymous:1895:IITa

[6328] Anonymous. Ingleton’s improved traction engine. Scientific American, 73(25):394, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
394a.pdf.

Anonymous:1895:TLE

[6329] Anonymous. The total lunar eclipse: Its astronomical value. Scientific American, 73(25):395, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
395a.pdf.

Brooks:1895:EMS

[6330] William R. Brooks. The eclipse of the Moon, September 3, 1895. Scientific American, 73(25):395, December 21, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n25/pdf/scientificamerican12211895-395.pdf.

Anonymous:1895:BNBbn

[6331] Anonymous. Books: New books and publications. Scientific American, 73(25):396, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
396a.pdf.

Anonymous:1895:BPbx

[6332] Anonymous. Business and personal. Scientific American, 73 (25):396, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
396b.pdf.

Anonymous:1895:RPIbx

[6333] Anonymous. Recently patented inventions. Scientific American, 73(25):396, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
396.pdf.

Anonymous:1895:NQbx

[6334] Anonymous. Notes and queries. Scientific American, 73(25): 396–397, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
396c.pdf.

Anonymous:1895:Ibv

[6335] Anonymous. To inventors. Scientific American, 73(25):397, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n25/pdf/scientificamerican12211895-397.pdf.

Anonymous:1895:IIcb

[6336] Anonymous. Index of inventions. Scientific American, 73(25): 397, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
397a.pdf.

Anonymous:1895:PNPbo

[6337] Anonymous. Patent notes: Patents. Scientific American, 73 (25):399, December 21, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n25/pdf/scientificamerican12211895-
399.pdf.

Anonymous:1895:Cbp

[6338] Anonymous. The crisis. Scientific American, 73(26):402, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n26/pdf/scientificamerican12281895-402.pdf.

Anonymous:1895:NRU

[6339] Anonymous. The Naval resources of the United States and the British Empire — a comparison. Scientific American, 73(26):402–403, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n26/pdf/scientificamerican12281895-402a.pdf.

Anonymous:1895:JS

[6340] Anonymous. The January sky. Scientific American, 73(26):403, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n26/pdf/scientificamerican12281895-403a.pdf.

Anonymous:1895:RS

[6341] Anonymous. To readers and subscribers. Scientific American, 73 (26):403, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
403b.pdf.

Anonymous:1895:USBc

[6342] Anonymous. The United States Bureau of Steam Engineering. Scientific American, 73(26):403, December 28, 1895. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v73/
n26/pdf/scientificamerican12281895-403.pdf.

Anonymous:1895:ESP

[6343] Anonymous. The electric self-playing piano. Scientific American, 73(26):404, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
404a.pdf.

Anonymous:1895:SOL

[6344] Anonymous. A steam omnibus in London, 1833. Scientific American, 73(26):404, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
404.pdf.

Anonymous:1895:ABSb

[6345] Anonymous. An automatic belt shifter. Scientific American, 73 (26):405, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
405a.pdf.

Anonymous:1895:BPAa

[6346] Anonymous. Bicycle power air pump. Scientific American, 73 (26):405, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
405.pdf.

Anonymous:1895:BAC

[6347] Anonymous. A bicycle adapted to carry two persons. Scientific American, 73(26):405–406, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
405b.pdf.

Anonymous:1895:IRa

[6348] Anonymous. An improved reel. Scientific American, 73(26): 406–407, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
406.pdf.

Anonymous:1895:CAEb

[6349] Anonymous. Close of the Atlanta Exposition. Scientific American, 73(26):408, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
408.pdf.

Anonymous:1895:CRSb

[6350] Anonymous. The cable repair steamer Mackay–Bennett. Scientific American, 73(26):409–410, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
409.pdf.

Anonymous:1895:LAH

[6351] Anonymous. The Lovell adjustable handle bar for bicycles. Scientific American, 73(26):410, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
410.pdf.

Anonymous:1895:BPby

[6352] Anonymous. Business and personal. Scientific American, 73 (26):411, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
411a.pdf.

Anonymous:1895:Ibw

[6353] Anonymous. To inventors. Scientific American, 73(26):411, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n26/pdf/scientificamerican12281895-411c.pdf.

Anonymous:1895:NQby

[6354] Anonymous. Notes and queries. Scientific American, 73(26): 411, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
411b.pdf.

Anonymous:1895:RPIby

[6355] Anonymous. Recently patented inventions. Scientific American, 73(26):411, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
411.pdf.

Anonymous:1895:IIcc

[6356] Anonymous. Index of inventions. Scientific American, 73(26): 411–412, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
411d.pdf.

Anonymous:1895:Ibx

[6357] Anonymous. Index. Scientific American, 73(26):413–414, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v73/n26/pdf/scientificamerican12281895-413.pdf.

Anonymous:1895:PNPbp

[6358] Anonymous. Patent notes: Patents. Scientific American, 73 (26):415, December 28, 1895. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v73/n26/pdf/scientificamerican12281895-
415.pdf.

Anonymous:1896:MBa

[6359] Anonymous. The modern bicycle. Scientific American, 74(1):2, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-2.pdf.

Anonymous:1896:RY

[6360] Anonymous. A retrospect of the year 1895. Scientific American, 74(1):2–3, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n1/pdf/scientificamerican01041896-
2a.pdf.

Anonymous:1896:TBa

[6361] Anonymous. The tribune bicycle. Scientific American, 74(1):4–5, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-4.pdf.

Anonymous:1896:IDG

[6362] Anonymous. Improved drawbridge gates operating mechanism. Scientific American, 74(1):5, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n1/pdf/scientificamerican01041896-
5.pdf.

Anonymous:1896:BFIa

[6363] Anonymous. Bicycle frame improvements. Scientific American, 74(1):6, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-6.pdf.

Anonymous:1896:NBP

[6364] Anonymous. “Niagara” bicycle pedals. Scientific American, 74(1):6–7, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-6a.pdf.

Anonymous:1896:Ca

[6365] Anonymous. Correspondence. Scientific American, 74(1):7, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-7.pdf.

Sloane:1896:AA

[6366] T. O’Conor Sloane. Acetylene apparatus. Scientific American, 74(1):8, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-8.pdf.

Anonymous:1896:USB

[6367] Anonymous. The United States battle ship Iowa. Scientific American, 74(1):9, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n1/pdf/scientificamerican01041896-
9.pdf.

Anonymous:1896:WEK

[6368] Anonymous. The White Elephant Kedah. Scientific American, 74(1):10–11, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n1/pdf/scientificamerican01041896-
10.pdf.

Anonymous:1896:HBB

[6369] Anonymous. A handle bar bicycle bell. Scientific American, 74(1):11, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-11a.pdf.

Anonymous:1896:PCS

[6370] Anonymous. The Punnett companion side seated bicycle. Scientific American, 74(1):11, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n1/pdf/scientificamerican01041896-
11.pdf.

Anonymous:1896:BNBa

[6371] Anonymous. Books: New books and publications. Scientific American, 74(1):12, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n1/pdf/scientificamerican01041896-
12a.pdf.

Anonymous:1896:BPa

[6372] Anonymous. Business and personal. Scientific American, 74(1):12, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-12b.pdf.

Anonymous:1896:NQa

[6373] Anonymous. Notes and queries. Scientific American, 74(1):12, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-12c.pdf.

Anonymous:1896:RPIa 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n1/pdf/scientificamerican01041896-
12.pdf.

Anonymous:1896:Ia

[6375] Anonymous. To inventors. Scientific American, 74(1):13, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-13.pdf.

Anonymous:1896:IIa

[6376] Anonymous. Index of inventions. Scientific American, 74(1):13, January 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n1/pdf/scientificamerican01041896-13a.pdf.

Anonymous:1896:EB

[6377] Anonymous. The Eagle bicycle. Scientific American, 74(2):17, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-17.pdf.

Anonymous:1896:AEB

[6378] Anonymous. Alfred Ely Beach. Scientific American, 74(2):18–19, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-18.pdf.

Anonymous:1896:HWP

[6379] Anonymous. The housing of the working people — its economic and ethical aspects. Scientific American, 74(2):19, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-19.pdf.

Anonymous:1896:VSP

[6380] Anonymous. A variable speed power clutch. Scientific American, 74(2):21, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
21.pdf.

Anonymous:1896:ROS

[6381] Anonymous. Restoration of the Olympian Stadium. Scientific American, 74(2):21–22, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
21a.pdf.

Anonymous:1896:ABC

[6382] Anonymous. Adjusting the beat of clock pendulums. Scientific American, 74(2):22, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
22.pdf.

Anonymous:1896:IF

[6383] Anonymous. An improved furnace. Scientific American, 74(2): 22–23, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
22a.pdf.

Anonymous:1896:FPE

[6384] Anonymous. The first proposed elevated railroad for New York. Scientific American, 74(2):24, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
24.pdf.

Anonymous:1896:TLa

[6385] Anonymous. The tramways of Lausanne. Scientific American, 74 (2):24–25, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
24a.pdf.

Hopkins:1896:EIG

[6386] George M. Hopkins. Electric igniters for gas engines. Scientific American, 74(2):25–26, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
25.pdf.

Anonymous:1896:CPN

[6387] Anonymous. Cyclist policemen in New York City. Scientific American, 74(2):26, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
26a.pdf.

Anonymous:1896:IRT

[6388] Anonymous. An improved railway tricycle. Scientific American, 74(2):26, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
26.pdf.

Anonymous:1896:PT

[6389] Anonymous. Princess Topaze. Scientific American, 74(2):27, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-27.pdf.

Anonymous:1896:BNBb

[6390] Anonymous. Books: New books and publications. Scientific American, 74(2):28, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
28a.pdf.

Anonymous:1896:BPb

[6391] Anonymous. Business and personal. Scientific American, 74(2):28, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-28b.pdf.

Anonymous:1896:RPIb

[6392] Anonymous. Recently patented inventions. Scientific American, 74(2):28, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n2/pdf/scientificamerican01111896-
28.pdf.

Anonymous:1896:NQb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-28c.pdf.

Anonymous:1896:Ib

[6394] Anonymous. To inventors. Scientific American, 74(2):29, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-29.pdf.

Anonymous:1896:IIb

[6395] Anonymous. Index of inventions. Scientific American, 74(2):29, January 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n2/pdf/scientificamerican01111896-29a.pdf.

Anonymous:1896:LDC

[6396] Anonymous. The light draught composite gunboats. Scientific American, 74(3):33, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
33.pdf.

Anonymous:1896:AIF

[6397] Anonymous. The advantages of the induced over the forced draught system. Scientific American, 74(3):34, January 18, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n3/pdf/scientificamerican01181896-34.pdf.

Anonymous:1896:ARS

[6398] Anonymous. The acceleration of railway speeds. Scientific American, 74(3):34–35, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
34a.pdf.

Anonymous:1896:BCR

[6399] Anonymous. A bicycle chain riveting machine. Scientific American, 74(3):36, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
36.pdf.

Anonymous:1896:BFB

[6400] Anonymous. A boiler fire box formed of water tubes. Scientific American, 74(3):36, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
36b.pdf.

Anonymous:1896:IEC

[6401] Anonymous. An inexpensive, efficient chimney. Scientific American, 74(3):36, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
36c.pdf.

Anonymous:1896:WEG

[6402] Anonymous. A wagon end gate fastener. Scientific American, 74(3):36, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
36a.pdf.

Anonymous:1896:ISD

[6403] Anonymous. An improved Stone dresser. Scientific American, 74(3):37, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
37.pdf.

Anonymous:1896:CTa

[6404] Anonymous. Curious tops. Scientific American, 74(3):37–38, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n3/pdf/scientificamerican01181896-37a.pdf.

Anonymous:1896:NTB

[6405] Anonymous. The new torpedo boats. Scientific American, 74 (3):39, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
39.pdf.

Anonymous:1896:IAE

[6406] Anonymous. The inventions of Alfred Ely Beach. Scientific American, 74(3):40–41, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
40.pdf.

Anonymous:1896:Cb

[6407] Anonymous. Correspondence. Scientific American, 74(3):42, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n3/pdf/scientificamerican01181896-42.pdf.

Anonymous:1896:TGN

[6408] Anonymous. A tropical garden in New Jersey. Scientific American, 74(3):43, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
43.pdf.

Anonymous:1896:BNBc

[6409] Anonymous. Books: New books and publications. Scientific American, 74(3):44, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
44a.pdf.

Anonymous:1896:RPIc

[6410] Anonymous. Recently patented inventions. Scientific American, 74(3):44, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n3/pdf/scientificamerican01181896-
44.pdf.

Anonymous:1896:BPc

[6411] Anonymous. Business and personal. Scientific American, 74(3):45, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n3/pdf/scientificamerican01181896-45.pdf.

Anonymous:1896:Ic

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n3/pdf/scientificamerican01181896-45b.pdf.

Anonymous:1896:IIc

[6413] Anonymous. Index of inventions. Scientific American, 74(3):45, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n3/pdf/scientificamerican01181896-45c.pdf.

Anonymous:1896:NQc

[6414] Anonymous. Notes and queries. Scientific American, 74(3):45, January 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n3/pdf/scientificamerican01181896-45a.pdf.

Anonymous:1896:TPNa

[6415] Anonymous. The transportation problem in New York City. Scientific American, 74(4):50, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
50.pdf.

Anonymous:1896:SEE

[6416] Anonymous. The over-supply of electrical engineers. Scientific American, 74(4):50–51, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
50a.pdf.

Anonymous:1896:GGB

[6417] Anonymous. A gas generator and burner. Scientific American, 74(4):52, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
52b.pdf.

Anonymous:1896:PTC

[6418] Anonymous. Pease’s tubular construction. Scientific American, 74(4):52, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
52.pdf.

Anonymous:1896:SCH

[6419] Anonymous. Stocker’s chalk holder. Scientific American, 74(4):52, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-52a.pdf.

Anonymous:1896:LRM

[6420] Anonymous. A leather riveting machine. Scientific American, 74(4):53, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
53.pdf.

Anonymous:1896:RPG

[6421] Anonymous. Remarkable potato growing. Scientific American, 74(4):53, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
53a.pdf.

Anonymous:1896:SSH

[6422] Anonymous. Steam and solar heat. Scientific American, 74 (4):53–54, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
53b.pdf.

Anonymous:1896:NFEa

[6423] Anonymous. The Niagara Falls electric power plant. Scientific American, 74(4):55, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
55.pdf.

Anonymous:1896:BHS

[6424] Anonymous. The Beach hydraulic shield. Scientific American, 74 (4):56–57, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
56.pdf.

Anonymous:1896:DSW

[6425] Anonymous. The Daimler side wheel launch Priscilla. Scientific American, 74(4):57, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
57a.pdf.

Anonymous:1896:LST

[6426] Anonymous. Laboratory still for tar. Scientific American, 74(4):57, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-57.pdf.

Anonymous:1896:Cc

[6427] Anonymous. Correspondence. Scientific American, 74(4):58–59, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-58.pdf.

Anonymous:1896:LRT

[6428] Anonymous. The Leamy revolving trapeze. Scientific American, 74(4):59, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
59.pdf.

Anonymous:1896:RPId

[6429] Anonymous. Recently patented inventions. Scientific American, 74(4):60, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
60.pdf.

Anonymous:1896:BNBd

[6430] Anonymous. Books: New books and publications. Scientific American, 74(4):60–61, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n4/pdf/scientificamerican01251896-
60a.pdf.

Anonymous:1896:BPd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-61.pdf.

Anonymous:1896:Id

[6432] Anonymous. To inventors. Scientific American, 74(4):61, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-61b.pdf.

Anonymous:1896:NQd

[6433] Anonymous. Notes and queries. Scientific American, 74(4):61, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-61a.pdf.

Anonymous:1896:IId

[6434] Anonymous. Index of inventions. Scientific American, 74(4):61–62, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-61c.pdf.

Anonymous:1896:PNPa

[6435] Anonymous. Patent notes: Patents. Scientific American, 74(4):63, January 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n4/pdf/scientificamerican01251896-63.pdf.

Anonymous:1896:ABE

[6436] Anonymous. The Annual Bicycle Exhibition at the Madison Square Garden, New York City. Scientific American, 74(5):66, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n5/pdf/scientificamerican02011896-66.pdf.

Serviss:1896:FS

[6437] Garrett P. Serviss. The February sky. Scientific American, 74 (5):66–67, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
66a.pdf.

Anonymous:1896:BCF

[6438] Anonymous. The Bassett curtain fixture. Scientific American, 74(5):68, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
68.pdf.

Anonymous:1896:ICa

[6439] Anonymous. The Ingersoll cyclometer. Scientific American, 74 (5):69, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
69a.pdf.

Anonymous:1896:ICAa

[6440] Anonymous. An improved car axle box. Scientific American, 74(5):69, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
69.pdf.

Anonymous:1896:EWM

[6441] Anonymous. An electric welding machine for use on bicycle frames and similar work. Scientific American, 74(5):69–70, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n5/pdf/scientificamerican02011896-69b.pdf.

Anonymous:1896:ENT

[6442] Anonymous. An efficient needle threading device. Scientific American, 74(5):70, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
70a.pdf.

Anonymous:1896:SNB

[6443] Anonymous. A shower or needle bath. Scientific American, 74 (5):70, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
70.pdf.

Anonymous:1896:TDD

[6444] Anonymous. The timber Dry Dock No. 3 at the New York Navy Yard. Scientific American, 74(5):71, February 1, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n5/pdf/scientificamerican02011896-71.pdf.

Anonymous:1896:EBL

[6445] Anonymous. An electric bicycle lamp. Scientific American, 74 (5):72, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
72b.pdf.

Anonymous:1896:TFCa

[6446] Anonymous. A twenty-five cent bicycle. Scientific American, 74 (5):72, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
72a.pdf.

Anonymous:1896:THC

[6447] Anonymous. Trolley hanger and cut-out. Scientific American, 74(5):72, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
72.pdf.

Anonymous:1896:GMC

[6448] Anonymous. Gas motor cars of the Dessau tramway. Scientific American, 74(5):73–74, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
73.pdf.

Anonymous:1896:MVE

[6449] Anonymous. Mr. Vanderbilt’s estate, Biltmore. Scientific American, 74(5):75–76, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
75.pdf.

Anonymous:1896:RPIe

[6450] Anonymous. Recently patented inventions. Scientific American, 74(5):76, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
76.pdf.

Anonymous:1896:BPe

[6451] Anonymous. Business and personal. Scientific American, 74 (5):77, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
77.pdf.

Anonymous:1896:Ie

[6452] Anonymous. To inventors. Scientific American, 74(5):77, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n5/pdf/scientificamerican02011896-77b.pdf.

Anonymous:1896:IIe

[6453] Anonymous. Index of inventions. Scientific American, 74(5):77, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n5/pdf/scientificamerican02011896-77c.pdf.

Anonymous:1896:NQe

[6454] Anonymous. Notes and queries. Scientific American, 74(5):77, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n5/pdf/scientificamerican02011896-77a.pdf.

Anonymous:1896:PNPb

[6455] Anonymous. Patent notes: Patents. Scientific American, 74 (5):79, February 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n5/pdf/scientificamerican02011896-
79.pdf.

Anonymous:1896:PRD

[6456] Anonymous. Prof. Roentgen’s discovery. Scientific American, 74 (6):82, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
82.pdf.

Anonymous:1896:PHS

[6457] Anonymous. The proportions of high speed engines. Scientific American, 74(6):82–83, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
82a.pdf.

Anonymous:1896:IPM

[6458] Anonymous. An improved propelling mechanism. Scientific American, 74(6):84, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
84.pdf.

Anonymous:1896:SSS

[6459] Anonymous. The stranding of the steamship St. Paul of the American Line. Scientific American, 74(6):84–86, February 8, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n6/pdf/scientificamerican02081896-84a.pdf.

Anonymous:1896:IBL

[6460] Anonymous. An improved bicycle lamp bracket. Scientific American, 74(6):86–87, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
86.pdf.

Anonymous:1896:HRD

[6461] Anonymous. The Harlem River Drawbridge and the Park Avenue improvement in New York City. Scientific American, 74(6):88–89, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n6/pdf/scientificamerican02081896-88.pdf.

Anonymous:1896:TCG

[6462] Anonymous. A transported California “Great Tree”. Scientific American, 74(6):89, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
89.pdf.

Anonymous:1896:VTC

[6463] Anonymous. A ventilator for telephone cabinets. Scientific American, 74(6):89–90, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
89a.pdf.

Anonymous:1896:Cd

[6464] Anonymous. Correspondence. Scientific American, 74(6):90–91, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n6/pdf/scientificamerican02081896-90.pdf.

Anonymous:1896:NAB

[6465] Anonymous. New Army bicycles. Scientific American, 74(6):91, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n6/pdf/scientificamerican02081896-91.pdf.

Anonymous:1896:NDBa

[6466] Anonymous. A new dry battery. Scientific American, 74(6): 91–92, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
91a.pdf.

Anonymous:1896:RPIf

[6467] Anonymous. Recently patented inventions. Scientific American, 74(6):92, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
92.pdf.

Anonymous:1896:BNBe

[6468] Anonymous. Books: New books and publications. Scientific American, 74(6):92–93, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
92a.pdf.

Anonymous:1896:BPf

[6469] Anonymous. Business and personal. Scientific American, 74 (6):93, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
93.pdf.

Anonymous:1896:Ig

[6470] Anonymous. To inventors. Scientific American, 74(6):93, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n6/pdf/scientificamerican02081896-93b.pdf.

Anonymous:1896:NQf

[6471] Anonymous. Notes and queries. Scientific American, 74(6):93, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n6/pdf/scientificamerican02081896-93a.pdf.

Anonymous:1896:IIf

[6472] Anonymous. Index of inventions. Scientific American, 74(6): 93–94, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
93c.pdf.

Anonymous:1896:PNPc

[6473] Anonymous. Patent notes: Patents. Scientific American, 74 (6):95, February 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n6/pdf/scientificamerican02081896-
95.pdf.

Anonymous:1896:NWS

[6474] Anonymous. The new works of science. Scientific American, 74(7):98, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
98.pdf.

Anonymous:1896:PAL

[6475] Anonymous. The Public Art League of the United States. Scientific American, 74(7):98–99, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
98a.pdf.

Anonymous:1896:HDD

[6476] Anonymous. A horse detaching device. Scientific American, 74 (7):100, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
100.pdf.

Anonymous:1896:LSAa

[6477] Anonymous. A locomotive Spark arrester. Scientific American, 74(7):100, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
100c.pdf.

Anonymous:1896:SDR

[6478] Anonymous. A snow and dirt removing cart. Scientific American, 74(7):100, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
100b.pdf.

Anonymous:1896:WIC

[6479] Anonymous. Walmsley’s improved compound blowpipe. Scientific American, 74(7):100, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
100a.pdf.

Anonymous:1896:IFWa

[6480] Anonymous. An intake and filter for water systems. Scientific American, 74(7):101, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
101.pdf.

Anonymous:1896:JCP

[6481] Anonymous. Japanese clocks and pocket Sun dials. Scientific American, 74(7):101–102, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
101a.pdf.

Anonymous:1896:IBa

[6482] Anonymous. An improved bit. Scientific American, 74(7):102, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n7/pdf/scientificamerican02151896-102.pdf.

Anonymous:1896:RXR

[6483] Anonymous. Roentgen or X ray photography. Scientific American, 74(7):103, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
103.pdf.

Anonymous:1896:CSB

[6484] Anonymous. A covered spiral bicycle pathway in Paris. Scientific American, 74(7):104, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
104a.pdf.

Anonymous:1896:DCM

[6485] Anonymous. A direct-connected motor and lathe. Scientific American, 74(7):104, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
104.pdf.

Anonymous:1896:SAM

[6486] Anonymous. Some American motor carriages. Scientific American, 74(7):105–106, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
105.pdf.

Anonymous:1896:KAB

[6487] Anonymous. The Konzi antelope in the Berlin Zoological Garden. Scientific American, 74(7):107, February 15, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n7/pdf/scientificamerican02151896-107a.pdf.

Anonymous:1896:WCF

[6488] Anonymous. William Crookes, F.R.S. Scientific American, 74 (7):107, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
107.pdf.

Anonymous:1896:BNBf

[6489] Anonymous. Books: New books and publications. Scientific American, 74(7):108, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
108a.pdf.

Anonymous:1896:BPg

[6490] Anonymous. Business and personal. Scientific American, 74(7): 108, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
108b.pdf.

Anonymous:1896:RPIg

[6491] Anonymous. Recently patented inventions. Scientific American, 74(7):108, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
108.pdf.

Anonymous:1896:NQg

[6492] Anonymous. Notes and queries. Scientific American, 74(7):108– 109, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
108c.pdf.

Anonymous:1896:Ih

[6493] Anonymous. To inventors. Scientific American, 74(7):109, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n7/pdf/scientificamerican02151896-109.pdf.

Anonymous:1896:IIg

[6494] Anonymous. Index of inventions. Scientific American, 74(7):109, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n7/pdf/scientificamerican02151896-109a.pdf.

Anonymous:1896:PNPd

[6495] Anonymous. Patent notes: Patents. Scientific American, 74 (7):111, February 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n7/pdf/scientificamerican02151896-
111.pdf.

Anonymous:1896:TPNb

[6496] Anonymous. The transportation problem in New York City. Scientific American, 74(8):114, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
114.pdf.

Anonymous:1896:CEE

[6497] Anonymous. The comparative economy of the electric and the cable car systems. Scientific American, 74(8):115, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n8/pdf/scientificamerican02221896-115a.pdf.

Anonymous:1896:RD

[6498] Anonymous. Roentgen’s discovery. Scientific American, 74(8): 115, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
115.pdf.

Anonymous:1896:RAS

[6499] Anonymous. A readily adjustable scythe. Scientific American, 74(8):116, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
116.pdf.

Anonymous:1896:MT

[6500] Anonymous. The manufacture of tacks. Scientific American, 74 (8):116–118, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
116a.pdf.

Anonymous:1896:UTS

[6501] Anonymous. The underground trolley street railway in New York City. Scientific American, 74(8):119, February 22, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n8/pdf/scientificamerican02221896-119.pdf.

Anonymous:1896:CEC

[6502] Anonymous. Chase’s electric cyclorama. Scientific American, 74 (8):120, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
120.pdf.

Anonymous:1896:TLb

[6503] Anonymous. Taylor’s lifeboat. Scientific American, 74(8):121, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n8/pdf/scientificamerican02221896-121.pdf.

Anonymous:1896:BOW

[6504] Anonymous. Baku and its oil wells. Scientific American, 74(8): 121–122, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
121a.pdf.

Anonymous:1896:CCB

[6505] Anonymous. Curious collapse of a barometer. Scientific American, 74(8):123, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
123a.pdf.

Anonymous:1896:ECF

[6506] Anonymous. An efficient car fender. Scientific American, 74 (8):123, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
123.pdf.

Anonymous:1896:BNBg

[6507] Anonymous. Books: New books and publications. Scientific American, 74(8):124, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
124a.pdf.

Anonymous:1896:RPIh

[6508] Anonymous. Recently patented inventions. Scientific American, 74(8):124, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
124.pdf.

Anonymous:1896:BPh

[6509] Anonymous. Business and personal. Scientific American, 74(8): 125, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
125.pdf.

Anonymous:1896:Ij

[6510] Anonymous. To inventors. Scientific American, 74(8):125, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n8/pdf/scientificamerican02221896-125b.pdf.

Anonymous:1896:IIh

[6511] Anonymous. Index of inventions. Scientific American, 74(8):125, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n8/pdf/scientificamerican02221896-125c.pdf.

Anonymous:1896:NQh

[6512] Anonymous. Notes and queries. Scientific American, 74(8):125, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n8/pdf/scientificamerican02221896-125a.pdf.

Anonymous:1896:PNPe

[6513] Anonymous. Patent notes: Patents. Scientific American, 74 (8):127, February 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n8/pdf/scientificamerican02221896-
127.pdf.

Anonymous:1896:VTI

[6514] Anonymous. Vacuum tube illumination by the D. McFarlan Moore system. Scientific American, 74(9):129, February 29, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n9/pdf/scientificamerican02291896-129.pdf.

Anonymous:1896:SGB

[6515] Anonymous. Successful German and Belgian competition with Great Britain in the iron trades. Scientific American, 74(9):130, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n9/pdf/scientificamerican02291896-130.pdf.

Anonymous:1896:RTNa

[6516] Anonymous. Rapid transit in New York City. Scientific American, 74(9):130–131, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
130a.pdf.

Anonymous:1896:ESC

[6517] Anonymous. An efficient speed controller. Scientific American, 74(9):132, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
132.pdf.

Anonymous:1896:ISSa

[6518] Anonymous. An improved shade supporter. Scientific American, 74(9):132, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
132b.pdf.

Anonymous:1896:SGW

[6519] Anonymous. Separating gas and water from oil. Scientific American, 74(9):132, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
132a.pdf.

Anonymous:1896:AVM

[6520] Anonymous. An automatic vending machine. Scientific American, 74(9):133, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
133.pdf.

Anonymous:1896:ROM

[6521] Anonymous. Recent observations of Mars. Scientific American, 74 (9):133–134, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
133a.pdf.

Anonymous:1896:RDW

[6522] Anonymous. Rebuilding defective walls of a Chicago building. Scientific American, 74(9):136, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
136.pdf.

Anonymous:1896:FMA

[6523] Anonymous. A French model acetylene lamp. Scientific American, 74(9):137, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
137.pdf.

Anonymous:1896:SPC

[6524] Anonymous. Street postal car systems of New York and Brooklyn. Scientific American, 74(9):137–138, February 29, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n9/pdf/scientificamerican02291896-137a.pdf.

Anonymous:1896:Ce

[6525] Anonymous. Correspondence. Scientific American, 74(9):138, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n9/pdf/scientificamerican02291896-138.pdf.

Anonymous:1896:BC

[6526] Anonymous. The boa constrictor. Scientific American, 74(9): 139, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
139.pdf.

Anonymous:1896:MIAa

[6527] Anonymous. The Manhattan incandescent arc lamp. Scientific American, 74(9):139, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
139a.pdf.

Anonymous:1896:BNBh

[6528] Anonymous. Books: New books and publications. Scientific American, 74(9):140, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
140b.pdf.

Anonymous:1896:BPi

[6529] Anonymous. Business and personal. Scientific American, 74(9): 140, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
140a.pdf.

Anonymous:1896:RPIi

[6530] Anonymous. Recently patented inventions. Scientific American, 74(9):140, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
140.pdf.

Anonymous:1896:NQi

[6531] Anonymous. Notes and queries. Scientific American, 74(9):140– 141, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
140c.pdf.

Anonymous:1896:Ik

[6532] Anonymous. To inventors. Scientific American, 74(9):141, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n9/pdf/scientificamerican02291896-141.pdf.

Anonymous:1896:IIi

[6533] Anonymous. Index of inventions. Scientific American, 74(9): 141–142, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
141a.pdf.

Anonymous:1896:PNPf

[6534] Anonymous. Patent notes: Patents. Scientific American, 74 (9):143, February 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n9/pdf/scientificamerican02291896-
143.pdf.

Anonymous:1896:MHT

[6535] Anonymous. The manufacture of the Hammond typewriter. Scientific American, 74(10):145, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
145.pdf.

Anonymous:1896:PTM

[6536] Anonymous. Pending trade mark legislation. Scientific American, 74(10):146–147, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
146.pdf.

Anonymous:1896:IES

[6537] Anonymous. An improved electric switch. Scientific American, 74(10):148, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
148.pdf.

Anonymous:1896:IEW

[6538] Anonymous. An insulator for electric wires. Scientific American, 74(10):148, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
148a.pdf.

Francis:1896:PDP

[6539] W. H. Francis. A portable disinfecting plant. Scientific American, 74(10):148–149, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
148b.pdf.

Anonymous:1896:ICCa

[6540] Anonymous. An improved car coupling. Scientific American, 74 (10):149, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
149.pdf.

Anonymous:1896:NRT

[6541] Anonymous. A new recording thermometer for atmospheric ranges of temperature. Scientific American, 74(10):149, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n10/pdf/scientificamerican03071896-149a.pdf.

Anonymous:1896:MPS

[6542] Anonymous. Mechanism for propelling small boats. Scientific American, 74(10):149–150, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
149b.pdf.

Anonymous:1896:Cf

[6543] Anonymous. Correspondence. Scientific American, 74(10):150, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n10/pdf/scientificamerican03071896-150.pdf.

Anonymous:1896:IML

[6544] Anonymous. An improved music leaf turner. Scientific American, 74(10):151, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
151.pdf.

Anonymous:1896:CDN

[6545] Anonymous. The Cabaret Du Néant. Scientific American, 74 (10):152–153, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
152.pdf.

Anonymous:1896:MTC

[6546] Anonymous. A Miner’s tool and candlestick. Scientific American, 74(10):153, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
153.pdf.

Lilienthal:1896:PED

[6547] Otto Lilienthal. Practical experiments for the development of human flight. Scientific American, 74(10):153–155, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n10/pdf/scientificamerican03071896-153a.pdf.

Anonymous:1896:XRP

[6548] Anonymous. X ray photography. Scientific American, 74(10):155, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n10/pdf/scientificamerican03071896-155.pdf.

Anonymous:1896:BPj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n10/pdf/scientificamerican03071896-156a.pdf.

Anonymous:1896:RPIj

[6550] Anonymous. Recently patented inventions. Scientific American, 74(10):156, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n10/pdf/scientificamerican03071896-
156.pdf.

Anonymous:1896:NQj

[6551] Anonymous. Notes and queries. Scientific American, 74(10):156–157, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n10/pdf/scientificamerican03071896-156b.pdf.

Anonymous:1896:Il

[6552] Anonymous. To inventors. Scientific American, 74(10):157, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n10/pdf/scientificamerican03071896-157.pdf.

Anonymous:1896:IIj

[6553] Anonymous. Index of inventions. Scientific American, 74(10):157, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n10/pdf/scientificamerican03071896-157a.pdf.

Anonymous:1896:PNPg

[6554] Anonymous. Patent notes: Patents. Scientific American, 74(10):159, March 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n10/pdf/scientificamerican03071896-159.pdf.

Anonymous:1896:NPO

[6555] Anonymous. Needed Patent Office legislation. Scientific American, 74(11):162, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
162.pdf.

Anonymous:1896:NYH

[6556] Anonymous. New York Harbor. Scientific American, 74(11):162, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-162b.pdf.

Anonymous:1896:PWS

[6557] Anonymous. A phenomenal wind storm. Scientific American, 74(11):162, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
162a.pdf.

Anonymous:1896:OGa

[6558] Anonymous. The Olympian Games. Scientific American, 74 (11):162–163, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
162c.pdf.

Anonymous:1896:NPEa

[6559] Anonymous. Nansen’s Polar Expedition. Scientific American, 74(11):163, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
163.pdf.

Anonymous:1896:ART

[6560] Anonymous. An anti-rattling thill coupling. Scientific American, 74(11):164, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
164.pdf.

Anonymous:1896:IGG

[6561] Anonymous. Improved gas and gasoline engine. Scientific American, 74(11):164, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
164a.pdf.

Anonymous:1896:TDC

[6562] Anonymous. Tension device for check row wires. Scientific American, 74(11):164, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
164b.pdf.

Anonymous:1896:VTO

[6563] Anonymous. Various types of optical projection lanterns. Scientific American, 74(11):165, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
165.pdf.

Anonymous:1896:Cg

[6564] Anonymous. Correspondence. Scientific American, 74(11):166, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-166.pdf.

Anonymous:1896:SVW

[6565] Anonymous. Securing vehicle wheels to axles. Scientific American, 74(11):166, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
166a.pdf.

Anonymous:1896:BCD

[6566] Anonymous. The Buffington–Crozier disappearing Barbette carriage for the eight inch B.L. rifle. Scientific American, 74(11):167–168, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-167.pdf.

Anonymous:1896:ERC

[6567] Anonymous. The eccentricities of a railroad collision. Scientific American, 74(11):168, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
168.pdf.

Galton:1896:PS

[6568] Francis Galton. Prints of scars. Scientific American, 74(11):168, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-168a.pdf.

Corthell:1896:IML

[6569] Elmer Laurence Corthell. Inaccessible maritime lights. Scientific American, 74(11):169–170, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
169.pdf.

Anonymous:1896:RES

[6570] Anonymous. Recent experiments in scientific kite flying. Scientific American, 74(11):171, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
171.pdf.

Anonymous:1896:BPk

[6571] Anonymous. Business and personal. Scientific American, 74(11):172, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-172a.pdf.

Anonymous:1896:RPIk

[6572] Anonymous. Recently patented inventions. Scientific American, 74(11):172, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n11/pdf/scientificamerican03141896-
172.pdf.

Anonymous:1896:NQk

[6573] Anonymous. Notes and queries. Scientific American, 74(11):172–173, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-172b.pdf.

Anonymous:1896:Im

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-173.pdf.

Anonymous:1896:IIk

[6575] Anonymous. Index of inventions. Scientific American, 74(11):173, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-173a.pdf.

Anonymous:1896:PNPh

[6576] Anonymous. Patent notes: Patents. Scientific American, 74(11):175, March 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n11/pdf/scientificamerican03141896-175.pdf.

Anonymous:1896:USP

[6577] Anonymous. The United States protected cruiser Olympia. Scientific American, 74(12):177–178, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
177.pdf.

Anonymous:1896:SLB

[6578] Anonymous. Stability of lofty buildings. Scientific American, 74(12):178, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
178.pdf.

Anonymous:1896:RPN

[6579] Anonymous. Report on the plans for New York rapid transit. Scientific American, 74(12):178–179, March 21, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n12/pdf/scientificamerican03211896-178a.pdf.

Anonymous:1896:NCBa

[6580] Anonymous. A new Copyright Bill. Scientific American, 74(12):179, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-179.pdf.

Anonymous:1896:IFB

[6581] Anonymous. An improved folding baby carriage. Scientific American, 74(12):180, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
180a.pdf.

Anonymous:1896:NCC

[6582] Anonymous. A new chimney cowl. Scientific American, 74(12):180, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-180.pdf.

Anonymous:1896:SEA

[6583] Anonymous. A simple and efficient amalgamator. Scientific American, 74(12):180, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
180b.pdf.

Anonymous:1896:GIS

[6584] Anonymous. A great ice skating rink, Washington, D.C. Scientific American, 74(12):181, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
181.pdf.

Anonymous:1896:PSS

[6585] Anonymous. A pneumatic straw stacker. Scientific American, 74(12):181, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
181a.pdf.

Anonymous:1896:Ch

[6586] Anonymous. Correspondence. Scientific American, 74(12):182, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-182.pdf.

Anonymous:1896:NHC

[6587] Anonymous. A novel hand camera. Scientific American, 74 (12):182–183, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
182a.pdf.

Anonymous:1896:DPX

[6588] Anonymous. Dr. Pupin’s X ray photograph of a wounded hand. Scientific American, 74(12):184, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
184a.pdf.

Anonymous:1896:RP

[6589] Anonymous. Roentgen photography. Scientific American, 74(12):184, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-184.pdf.

Anonymous:1896:PLK

[6590] Anonymous. Preparation of Lamb and kid skins for gloves. Scientific American, 74(12):185–186, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
185.pdf.

Anonymous:1896:FAA

[6591] Anonymous. The Fuller acetylene apparatus. Scientific American, 74(12):187, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
187a.pdf.

Anonymous:1896:HSS

[6592] Anonymous. How to sail on skates. Scientific American, 74(12):187, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-187.pdf.

Anonymous:1896:BNBi

[6593] Anonymous. Books: New books and publications. Scientific American, 74(12):188, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
188a.pdf.

Anonymous:1896:BPl

[6594] Anonymous. Business and personal. Scientific American, 74(12):188, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-188b.pdf.

Anonymous:1896:RPIl

[6595] Anonymous. Recently patented inventions. Scientific American, 74(12):188, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n12/pdf/scientificamerican03211896-
188.pdf.

Anonymous:1896:NQl

[6596] Anonymous. Notes and queries. Scientific American, 74(12):188–189, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-188c.pdf.

Anonymous:1896:In

[6597] Anonymous. To inventors. Scientific American, 74(12):189, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-189.pdf.

Anonymous:1896:IIl

[6598] Anonymous. Index of inventions. Scientific American, 74(12):189, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-189a.pdf.

Anonymous:1896:PNPi

[6599] Anonymous. Patent notes: Patents. Scientific American, 74(12):191, March 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n12/pdf/scientificamerican03211896-191.pdf.

Anonymous:1896:GRN

[6600] Anonymous. The Gorge Road at Niagara Falls. Scientific American, 74(13):193, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
193.pdf.

Anonymous:1896:PIO

[6601] Anonymous. The proposed increase in our Navy. Scientific American, 74(13):194, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
194.pdf.

Anonymous:1896:WTN

[6602] Anonymous. Weather tests on the New York underground trolley road. Scientific American, 74(13):194, March 28, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n13/pdf/scientificamerican03281896-194a.pdf.

Anonymous:1896:GTS

[6603] Anonymous. The Government tests of the strength of timber. Scientific American, 74(13):194–195, March 28, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n13/pdf/scientificamerican03281896-194b.pdf.

Anonymous:1896:CDP

[6604] Anonymous. A combination device for the pocket. Scientific American, 74(13):196, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
196.pdf.

Anonymous:1896:NSM

[6605] Anonymous. A new speed measure and alarm. Scientific American, 74(13):196, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
196b.pdf.

Anonymous:1896:RL

[6606] Anonymous. The Rivett lathe. Scientific American, 74(13):196, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-196a.pdf.

Anonymous:1896:IBb

[6607] Anonymous. Interlocking brick. Scientific American, 74(13):197, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-197a.pdf.

Anonymous:1896:NGB

[6608] Anonymous. A novel game board. Scientific American, 74(13):197, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-197.pdf.

Anonymous:1896:MKG

[6609] Anonymous. Manufacture of kid gloves. Scientific American, 74 (13):197–198, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
197b.pdf.

Anonymous:1896:Ci

[6610] Anonymous. Correspondence. Scientific American, 74(13):199, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-199.pdf.

Anonymous:1896:GCD

[6611] Anonymous. A great coal dock on Lake Superior. Scientific American, 74(13):200, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
200.pdf.

Anonymous:1896:HCP

[6612] Anonymous. The hippopotamus in the Central Park Menagerie, New York City. Scientific American, 74(13):201–202, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-201.pdf.

Anonymous:1896:AFO

[6613] Anonymous. “After the Flood” at the Olympia. Scientific American, 74(13):203, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
203.pdf.

Anonymous:1896:BNBj

[6614] Anonymous. Books: New books and publications. Scientific American, 74(13):204, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
204a.pdf.

Anonymous:1896:BPm

[6615] Anonymous. Business and personal. Scientific American, 74(13):204, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-204b.pdf.

Anonymous:1896:RPIm

[6616] Anonymous. Recently patented inventions. Scientific American, 74(13):204, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n13/pdf/scientificamerican03281896-
204.pdf.

Anonymous:1896:NQm

[6617] Anonymous. Notes and queries. Scientific American, 74(13):204–205, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-204c.pdf.

Anonymous:1896:Io

[6618] Anonymous. To inventors. Scientific American, 74(13):205, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-205.pdf.

Anonymous:1896:IIm

[6619] Anonymous. Index of inventions. Scientific American, 74(13):205, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-205a.pdf.

Anonymous:1896:PNPj

[6620] Anonymous. Patent notes: Patents. Scientific American, 74(13):207, March 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n13/pdf/scientificamerican03281896-207.pdf.

Anonymous:1896:VI

[6621] Anonymous. The victimizing of inventors. Scientific American, 74(14):210, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
210.pdf.

Holden:1896:LOE

[6622] Edward S. Holden. The Lick Observatory Expedition to observe the total solar eclipse of August, 1896, in Japan. Scientific American, 74(14):210–211, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
210a.pdf.

Anonymous:1896:MPa

[6623] Anonymous. The manufacture of paper. Scientific American, 74 (14):211, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
211.pdf.

Anonymous:1896:IHTa

[6624] Anonymous. Improved horizontal tapping machine. Scientific American, 74(14):212, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
212c.pdf.

Anonymous:1896:NCBb

[6625] Anonymous. A new car brake. Scientific American, 74(14):212, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-212.pdf.

Anonymous:1896:NLS

[6626] Anonymous. A new lawn sprinkler. Scientific American, 74(14):212, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-212b.pdf.

Anonymous:1896:WTR

[6627] Anonymous. Watchman’s time recorder. Scientific American, 74 (14):212, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
212a.pdf.

Anonymous:1896:Da

[6628] Anonymous. A dynamograph. Scientific American, 74(14):213, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-213a.pdf.

Anonymous:1896:IDW

[6629] Anonymous. An improved dumb waiter. Scientific American, 74 (14):213, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
213b.pdf.

Anonymous:1896:RCF

[6630] Anonymous. The Robinson corn and feed mill. Scientific American, 74(14):213, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
213.pdf.

Anonymous:1896:Cj

[6631] Anonymous. Correspondence. Scientific American, 74(14):214, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-214.pdf.

Anonymous:1896:NFH

[6632] Anonymous. Niagara Falls Hydraulic Power Plant. Scientific American, 74(14):215, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
215.pdf.

Anonymous:1896:PMS

[6633] Anonymous. Press for making steel bicycle rims. Scientific American, 74(14):216, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
216.pdf.

Anonymous:1896:TT

[6634] Anonymous. Train telephony. Scientific American, 74(14):216–217, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-216a.pdf.

Anonymous:1896:Ta

[6635] Anonymous. The thermophone. Scientific American, 74(14):217, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-217.pdf.

Anonymous:1896:FLC

[6636] Anonymous. The frilled lizard — Chlamydosaurus Kingi. Scientific American, 74(14):217–218, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
217a.pdf.

Anonymous:1896:NXRa

[6637] Anonymous. The new X ray “Focus” tube. Scientific American, 74(14):218, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
218.pdf.

Anonymous:1896:EXR

[6638] Anonymous. The Edison X ray experiments, apparatus and fluoroscope. Scientific American, 74(14):219, April 4, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n14/pdf/scientificamerican04041896-219.pdf.

Anonymous:1896:BNBk

[6639] Anonymous. Books: New books and publications. Scientific American, 74(14):220, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
220a.pdf.

Anonymous:1896:BPn

[6640] Anonymous. Business and personal. Scientific American, 74(14):220, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-220b.pdf.

Anonymous:1896:NQn

[6641] Anonymous. Notes and queries. Scientific American, 74(14):220, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-220c.pdf.

Anonymous:1896:RPIn

[6642] Anonymous. Recently patented inventions. Scientific American, 74(14):220, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n14/pdf/scientificamerican04041896-
220.pdf.

Anonymous:1896:IIn

[6643] Anonymous. Index of inventions. Scientific American, 74(14):221, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-221.pdf.

Anonymous:1896:PNPk

[6644] Anonymous. Patent notes: Patents. Scientific American, 74(14):222, April 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n14/pdf/scientificamerican04041896-222.pdf.

Anonymous:1896:ICb

[6645] Anonymous. The Island of Cuba. Scientific American, 74(15):225, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-225.pdf.

Anonymous:1896:NS

[6646] Anonymous. The New South. Scientific American, 74(15):226, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-226.pdf.

Serviss:1896:AS

[6647] Garrett P. Serviss. The April sky. Scientific American, 74(15):226–227, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-226a.pdf.

Anonymous:1896:CBSa

[6648] Anonymous. The contest between shot and armor. Scientific American, 74(15):227, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n15/pdf/scientificamerican04111896-
227.pdf.

Anonymous:1896:IPa

[6649] Anonymous. An improved packing. Scientific American, 74(15):228, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-228.pdf.

Anonymous:1896:NLJ

[6650] Anonymous. A new logging jack. Scientific American, 74(15):228, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-228c.pdf.

Anonymous:1896:PCB

[6651] Anonymous. Pratt’s cone belt shifter. Scientific American, 74(15):228, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-228b.pdf.

Anonymous:1896:PDA

[6652] Anonymous. Professor Dewar’s apparatus for liquefying air and oxygen. Scientific American, 74(15):228, April 11, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n15/pdf/scientificamerican04111896-228a.pdf.

Anonymous:1896:DB

[6653] Anonymous. A dog bicyclist. Scientific American, 74(15):229, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-229.pdf.

Anonymous:1896:NMC

[6654] Anonymous. A new motor cycle. Scientific American, 74(15):229–231, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-229a.pdf.

Anonymous:1896:LSB

[6655] Anonymous. The launch of the seagoing battleship Iowa. Scientific American, 74(15):233–234, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n15/pdf/scientificamerican04111896-
233.pdf.

Anonymous:1896:SCA

[6656] Anonymous. A stable cleaning apparatus. Scientific American, 74(15):234, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n15/pdf/scientificamerican04111896-
234.pdf.

Anonymous:1896:CBa

[6657] Anonymous. A cake baked in a hat. Scientific American, 74 (15):234–235, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n15/pdf/scientificamerican04111896-
234a.pdf.

Anonymous:1896:IAD

[6658] Anonymous. An interesting archeological discovery. Scientific American, 74(15):235, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n15/pdf/scientificamerican04111896-
235a.pdf.

Anonymous:1896:R

[6659] Anonymous. Radiography. Scientific American, 74(15):235, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-235.pdf.

Anonymous:1896:RPIo

[6660] Anonymous. Recently patented inventions. Scientific American, 74(15):236, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n15/pdf/scientificamerican04111896-
236.pdf.

Anonymous:1896:NQo

[6661] Anonymous. Notes and queries. Scientific American, 74(15):236–237, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-236a.pdf.

Anonymous:1896:BPo

[6662] Anonymous. Business and personal. Scientific American, 74(15):237, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-237.pdf.

Anonymous:1896:Iq

[6663] Anonymous. To inventors. Scientific American, 74(15):237, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-237a.pdf.

Anonymous:1896:IIo

[6664] Anonymous. Index of inventions. Scientific American, 74(15):237, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-237b.pdf.

Anonymous:1896:PNPl

[6665] Anonymous. Patent notes: Patents. Scientific American, 74(15):238, April 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n15/pdf/scientificamerican04111896-238.pdf.

Anonymous:1896:LCM

[6666] Anonymous. Letter canceling machines at the New York post office. Scientific American, 74(16):241–242, April 18, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n16/pdf/scientificamerican04181896-241.pdf.

Anonymous:1896:PPL

[6667] Anonymous. Proposed patent legislation. Scientific American, 74 (16):242–243, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n16/pdf/scientificamerican04181896-
242.pdf.

Anonymous:1896:CGE

[6668] Anonymous. Combined gas engine and centrifugal pump. Scientific American, 74(16):244, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n16/pdf/scientificamerican04181896-
244a.pdf.

Anonymous:1896:HDB

[6669] Anonymous. A horse detacher and brake. Scientific American, 74(16):244, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n16/pdf/scientificamerican04181896-
244.pdf.

Anonymous:1896:HSD

[6670] Anonymous. A hoof spreading device. Scientific American, 74(16):244, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-244b.pdf.

Anonymous:1896:AHC

[6671] Anonymous. An automatic hose coupling. Scientific American, 74(16):245, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n16/pdf/scientificamerican04181896-
245.pdf.

Anonymous:1896:IWa

[6672] Anonymous. An improved winch. Scientific American, 74(16):245, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-245a.pdf.

Anonymous:1896:IBT

[6673] Anonymous. An improved bag tie. Scientific American, 74(16):245–246, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-245b.pdf.

Anonymous:1896:SP

[6674] Anonymous. Serviceable platform. Scientific American, 74(16):246–247, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-246.pdf.

Anonymous:1896:AGM

[6675] Anonymous. The American gas machine. Scientific American, 74(16):248, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n16/pdf/scientificamerican04181896-
248.pdf.

McKay:1896:MMR

[6676] John S. McKay. Magnetographs made by radiations from the poles of a magnet. Scientific American, 74(16):249, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-249.pdf.

Anonymous:1896:PCZ

[6677] Anonymous. Pod corn — Zea Tunicata. Scientific American, 74(16):250, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n16/pdf/scientificamerican04181896-
250a.pdf.

Goto:1896:SMM

[6678] Ichiro Goto. Simple method for the measurement of electric current. Scientific American, 74(16):250, April 18, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n16/pdf/scientificamerican04181896-250.pdf.

Anonymous:1896:SLC

[6679] Anonymous. Straightening a leaning chimney one hundred feet high. Scientific American, 74(16):251, April 18, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n16/pdf/scientificamerican04181896-251.pdf.

Anonymous:1896:BPp

[6680] Anonymous. Business and personal. Scientific American, 74(16):252, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-252.pdf.

Anonymous:1896:Ir

[6681] Anonymous. To inventors. Scientific American, 74(16):252, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-252c.pdf.

Anonymous:1896:NQp

[6682] Anonymous. Notes and queries. Scientific American, 74(16):252, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-252b.pdf.

Anonymous:1896:RPIp

[6683] Anonymous. Recently patented inventions. Scientific American, 74(16):252, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n16/pdf/scientificamerican04181896-
252a.pdf.

Anonymous:1896:IIp

[6684] Anonymous. Index of inventions. Scientific American, 74(16):252–253, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-252d.pdf.

Anonymous:1896:PNPm

[6685] Anonymous. Patent notes: Patents. Scientific American, 74(16):255, April 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n16/pdf/scientificamerican04181896-255.pdf.

Anonymous:1896:IBPa

[6686] Anonymous. The invention of the Bessemer process. Scientific American, 74(17):258, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
258.pdf.

Anonymous:1896:PFB

[6687] Anonymous. Passage of the Fortifications Bill. Scientific American, 74(17):258–259, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
258a.pdf.

Anonymous:1896:BAC

[6688] Anonymous. The Baldwin adjustable cycle chain. Scientific American, 74(17):260, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
260a.pdf.

Anonymous:1896:ISa

[6689] Anonymous. An improved T-square. Scientific American, 74(17):260, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-260.pdf.

Anonymous:1896:NMM

[6690] Anonymous. A novel mechanical motor. Scientific American, 74 (17):260, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
260b.pdf.

Anonymous:1896:DBT

[6691] Anonymous. A durable bicycle tire. Scientific American, 74(17):261, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-261.pdf.

Anonymous:1896:SPSa

[6692] Anonymous. Skirt protecting screen for bicycles. Scientific American, 74(17):261, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
261a.pdf.

Anonymous:1896:WSM

[6693] Anonymous. A warp sizing machine attachment. Scientific American, 74(17):261, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
261b.pdf.

Anonymous:1896:Ck

[6694] Anonymous. Correspondence. Scientific American, 74(17):262, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-262.pdf.

Anonymous:1896:ISIa

[6695] Anonymous. An improved sewer inlet. Scientific American, 74(17):262, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-262a.pdf.

Anonymous:1896:HST

[6696] Anonymous. The Holland submarine torpedo boat. Scientific American, 74(17):263, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
263.pdf.

Anonymous:1896:QFE

[6697] Anonymous. A quadricycle fire engine. Scientific American, 74(17):264, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-264.pdf.

Anonymous:1896:DPP

[6698] Anonymous. The Dudley powder pneumatic gun. Scientific American, 74(17):265–266, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
265.pdf.

Anonymous:1896:DDC

[6699] Anonymous. Double deck cars. Scientific American, 74(17):267, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-267.pdf.

Williams:1896:DDB

[6700] Emery Leverett Williams. The double dovetail and blind mortise. Scientific American, 74(17):267, April 25, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n17/pdf/scientificamerican04251896-267a.pdf.

Anonymous:1896:BPq

[6701] Anonymous. Business and personal. Scientific American, 74(17):268, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-268.pdf.

Anonymous:1896:NQq

[6702] Anonymous. Notes and queries. Scientific American, 74(17):268, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-268b.pdf.

Anonymous:1896:RPIq

[6703] Anonymous. Recently patented inventions. Scientific American, 74(17):268, April 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n17/pdf/scientificamerican04251896-
268a.pdf.

Anonymous:1896:IIq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n17/pdf/scientificamerican04251896-268c.pdf.

Anonymous:1896:NBH

[6705] Anonymous. The new bridge over the Hudson River at New York. Scientific American, 74(18):273, May 2, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n18/pdf/scientificamerican05021896-273.pdf.

Anonymous:1896:DAL

[6706] Anonymous. The destructive action of locomotive driving wheels. Scientific American, 74(18):274, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
274.pdf.

Serviss:1896:SM

[6707] Garrett P. Serviss. The sky in May. Scientific American, 74 (18):275–276, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
275.pdf.

Anonymous:1896:DEW

[6708] Anonymous. A durable and efficient windmill. Scientific American, 74(18):276, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
276b.pdf.

Anonymous:1896:ISMa

[6709] Anonymous. An improved sorting machine. Scientific American, 74(18):276, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
276.pdf.

Anonymous:1896:SMa

[6710] Anonymous. A safety hat for miners. Scientific American, 74(18):276, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-276a.pdf.

Anonymous:1896:BSM

[6711] Anonymous. A bottle stopper and measuring device. Scientific American, 74(18):277, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
277.pdf.

Anonymous:1896:IIr

[6712] Anonymous. Isafjord, Iceland. Scientific American, 74(18):277, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-277a.pdf.

Anonymous:1896:NBBa

[6713] Anonymous. A new bicycle bell. Scientific American, 74(18):277–278, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-277b.pdf.

Anonymous:1896:FSN

[6714] Anonymous. Feed saving nose bag. Scientific American, 74(18):278, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-278.pdf.

Anonymous:1896:Cl

[6715] Anonymous. Correspondence. Scientific American, 74(18):279, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-279.pdf.

Anonymous:1896:AMS

[6716] Anonymous. Apparatus for measuring the speed of projectiles. Scientific American, 74(18):280–281, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
280.pdf.

Dunbar:1896:ROH

[6717] D. Dunbar. The Royal Observatory and how they tell the time at Greenwich. Scientific American, 74(18):281–283, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n18/pdf/scientificamerican05021896-281.pdf.

Anonymous:1896:CBR

[6718] Anonymous. The carriage of bicycles by railway. Scientific American, 74(18):283, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
283a.pdf.

Anonymous:1896:ES

[6719] Anonymous. England and the Soudan. Scientific American, 74(18):283, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-283.pdf.

Anonymous:1896:BPr

[6720] Anonymous. Business and personal. Scientific American, 74(18):284, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-284.pdf.

Anonymous:1896:NQr

[6721] Anonymous. Notes and queries. Scientific American, 74(18):284, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-284b.pdf.

Anonymous:1896:RPIr

[6722] Anonymous. Recently patented inventions. Scientific American, 74(18):284, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n18/pdf/scientificamerican05021896-
284a.pdf.

Anonymous:1896:IIs

[6723] Anonymous. Index of inventions. Scientific American, 74(18):284–285, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-284c.pdf.

Anonymous:1896:PNPn

[6724] Anonymous. Patent notes: Patents. Scientific American, 74(18):286, May 2, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n18/pdf/scientificamerican05021896-286.pdf.

Childers:1896:RCS

[6725] H. H. Childers. Rice culture in Southwestern Louisiana. Scientific American, 74(19):289, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
289.pdf.

Anonymous:1896:OFAa

[6726] Anonymous. Our fiftieth anniversary prize essay competition. Scientific American, 74(19):290, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
290.pdf.

Anonymous:1896:MSN

[6727] Anonymous. Modern steamships and navigation. Scientific American, 74(19):290–291, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
290a.pdf.

Anonymous:1896:APR

[6728] Anonymous. Air pressure relief valve for hose couplings. Scientific American, 74(19):292, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
292b.pdf.

Anonymous:1896:EBT

[6729] Anonymous. An efficient boiler tube cutter. Scientific American, 74(19):292, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
292.pdf.

Anonymous:1896:ISP

[6730] Anonymous. The Ingleton steam plow. Scientific American, 74(19):292, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-292a.pdf.

Anonymous:1896:IDP

[6731] Anonymous. An improved dental plugger. Scientific American, 74(19):293, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
293.pdf.

Anonymous:1896:BIEa

[6732] Anonymous. Berlin Industrial Exposition of 1896. Scientific American, 74(19):293–294, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
293a.pdf.

Anonymous:1896:Cm

[6733] Anonymous. Correspondence. Scientific American, 74(19):294, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-294.pdf.

Anonymous:1896:STU

[6734] Anonymous. The speed trial of the United States battleship Massachusetts. Scientific American, 74(19):296–297, May 9, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n19/pdf/scientificamerican05091896-296.pdf.

Anonymous:1896:IVF

[6735] Anonymous. The Illusion Vanity Fair at the Olympia. Scientific American, 74(19):297–298, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
297.pdf.

Anonymous:1896:TWW

[6736] Anonymous. A turbine water wheel. Scientific American, 74(19):298, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-298.pdf.

Anonymous:1896:CRD

[6737] Anonymous. California rabbit drives and hunts. Scientific American, 74(19):299, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
299.pdf.

Anonymous:1896:BPs

[6738] Anonymous. Business and personal. Scientific American, 74(19):300, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-300a.pdf.

Anonymous:1896:RPIs

[6739] Anonymous. Recently patented inventions. Scientific American, 74(19):300, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n19/pdf/scientificamerican05091896-
300.pdf.

Anonymous:1896:NQs

[6740] Anonymous. Notes and queries. Scientific American, 74(19):300–301, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-300b.pdf.

Anonymous:1896:It

[6741] Anonymous. To inventors. Scientific American, 74(19):301, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-301.pdf.

Anonymous:1896:IIt

[6742] Anonymous. Index of inventions. Scientific American, 74(19):301, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-301a.pdf.

Anonymous:1896:PNPo

[6743] Anonymous. Patent notes: Patents. Scientific American, 74(19):303, May 9, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n19/pdf/scientificamerican05091896-303.pdf.

Anonymous:1896:NEEa

[6744] Anonymous. The national electrical exposition. Scientific American, 74(20):306, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
306.pdf.

Anonymous:1896:RTNb

[6745] Anonymous. Rapid transit in New York City. Scientific American, 74(20):306, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
306a.pdf.

Anonymous:1896:PCR

[6746] Anonymous. The patent Commissioner’s report for 1895. Scientific American, 74(20):306–307, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
306b.pdf.

Anonymous:1896:CIS

[6747] Anonymous. A combination ice shaver and mile shaker. Scientific American, 74(20):308, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
308b.pdf.

Anonymous:1896:ICCb

[6748] Anonymous. An improved coal crusher. Scientific American, 74 (20):308, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
308a.pdf.

Anonymous:1896:ISSb

[6749] Anonymous. An improved Spring seat. Scientific American, 74(20):308, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-308.pdf.

Anonymous:1896:NEEb

[6750] Anonymous. The National Electrical Exposition at the Grand Central Palace, New York. Scientific American, 74(20):309–310, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-309.pdf.

Anonymous:1896:G

[6751] Anonymous. The gramophone. Scientific American, 74(20):311, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-311.pdf.

Anonymous:1896:DPA

[6752] Anonymous. Dog power as an auxiliary in wheeling. Scientific American, 74(20):312, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
312.pdf.

Anonymous:1896:VHT

[6753] Anonymous. The De La Vergne hunting trap. Scientific American, 74(20):312, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
312a.pdf.

Anonymous:1896:OGb

[6754] Anonymous. The Olympic Games. Scientific American, 74(20):313–314, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-313.pdf.

Anonymous:1896:BPt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-315a.pdf.

Anonymous:1896:NQt

[6756] Anonymous. Notes and queries. Scientific American, 74(20):315, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-315b.pdf.

Anonymous:1896:RPIt

[6757] Anonymous. Recently patented inventions. Scientific American, 74(20):315, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
315.pdf.

Anonymous:1896:BNBl

[6758] Anonymous. Books: New books and publications. Scientific American, 74(20):316, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n20/pdf/scientificamerican05161896-
316.pdf.

Anonymous:1896:Iu

[6759] Anonymous. To inventors. Scientific American, 74(20):316, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-316a.pdf.

Anonymous:1896:IIu

[6760] Anonymous. Index of inventions. Scientific American, 74(20):316, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-316b.pdf.

Anonymous:1896:PNPp

[6761] Anonymous. Patent notes: Patents. Scientific American, 74(20):319, May 16, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n20/pdf/scientificamerican05161896-319.pdf.

Wilson:1896:JGH

[6762] Walfred Wilson. The jetties at Galveston Harbor. Scientific American, 74(21):321, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
321.pdf.

Anonymous:1896:DAS

[6763] Anonymous. The decadence of the apprenticeship system. Scientific American, 74(21):322, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
322.pdf.

Anonymous:1896:STB

[6764] Anonymous. The speed trials of the Brooklyn and the Oregon. Scientific American, 74(21):322–323, May 23, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n21/pdf/scientificamerican05231896-322a.pdf.

Anonymous:1896:MM

[6765] Anonymous. Modified milk. Scientific American, 74(21):324–326, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-324.pdf.

Anonymous:1896:Cn

[6766] Anonymous. Correspondence. Scientific American, 74(21):326, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-326a.pdf.

Anonymous:1896:NMT

[6767] Anonymous. A natural mouse trap. Scientific American, 74(21):326, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-326.pdf.

Anonymous:1896:ITH

[6768] Anonymous. An ingenious tool holder attachment. Scientific American, 74(21):328, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
328a.pdf.

Anonymous:1896:TAM

[6769] Anonymous. Typewriting and adding machine. Scientific American, 74(21):328, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
328.pdf.

Anonymous:1896:HME

[6770] Anonymous. The Hungarian Millennial Exposition of Budapest. Scientific American, 74(21):329–330, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
329.pdf.

Anonymous:1896:NPC

[6771] Anonymous. A novel pocket camera. Scientific American, 74(21):331, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-331.pdf.

Anonymous:1896:SHP

[6772] Anonymous. A subsoil and harrow plow attachment. Scientific American, 74(21):331, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
331a.pdf.

Anonymous:1896:BNBm

[6773] Anonymous. Books: New books and publications. Scientific American, 74(21):332, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
332a.pdf.

Anonymous:1896:BPu

[6774] Anonymous. Business and personal. Scientific American, 74(21):332, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-332b.pdf.

Anonymous:1896:NQu

[6775] Anonymous. Notes and queries. Scientific American, 74(21):332, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-332c.pdf.

Anonymous:1896:RPIu

[6776] Anonymous. Recently patented inventions. Scientific American, 74(21):332, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n21/pdf/scientificamerican05231896-
332.pdf.

Anonymous:1896:Iv

[6777] Anonymous. To inventors. Scientific American, 74(21):333, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-333.pdf.

Anonymous:1896:IIv

[6778] Anonymous. Index of inventions. Scientific American, 74(21):333, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-333a.pdf.

Anonymous:1896:PNPq

[6779] Anonymous. Patent notes: Patents. Scientific American, 74(21):335, May 23, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n21/pdf/scientificamerican05231896-335.pdf.

Anonymous:1896:IBPb

[6780] Anonymous. The invention of the Bessemer process. Scientific American, 74(22):338, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
338a.pdf.

Anonymous:1896:OPE

[6781] Anonymous. Our prize essay competition. Scientific American, 74(22):338, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
338.pdf.

Anonymous:1896:MA

[6782] Anonymous. Molecular annealing. Scientific American, 74(22):338–339, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-338b.pdf.

Anonymous:1896:DT

[6783] Anonymous. A Dutch Tilbury. Scientific American, 74(22):340, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-340.pdf.

Anonymous:1896:OT

[6784] Anonymous. Olympian trophies. Scientific American, 74(22):340, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-340a.pdf.

Anonymous:1896:ISV

[6785] Anonymous. An improved sailing vessel. Scientific American, 74 (22):341, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
341a.pdf.

Anonymous:1896:WPC

[6786] Anonymous. The Wethey parallel calcining furnace. Scientific American, 74(22):341, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
341.pdf.

Anonymous:1896:RCW

[6787] Anonymous. A Rim clamp for wheels. Scientific American, 74 (22):341–342, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
341b.pdf.

Anonymous:1896:TDP

[6788] Anonymous. The thief and detective puzzle. Scientific American, 74(22):342, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
342.pdf.

Hutchins:1896:ICT

[6789] C. C. Hutchins and F. C. Robinson. Improvements in Crookes tubes. Scientific American, 74(22):342, May 30, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n22/pdf/scientificamerican05301896-342a.pdf.

Anonymous:1896:MGW

[6790] Anonymous. The manufacture of guns at the Washington Navy Yard. Scientific American, 74(22):343, May 30, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n22/pdf/scientificamerican05301896-343.pdf.

Hunt:1896:EJM

[6791] Joseph H. Hunt. Edward Jenner, M.D., and the vaccination centennial. Scientific American, 74(22):344, May 30, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n22/pdf/scientificamerican05301896-344.pdf.

Anonymous:1896:TM

[6792] Anonymous. Testing motocycles. Scientific American, 74(22):345, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-345.pdf.

Anonymous:1896:PSI

[6793] Anonymous. The proposed sixteen inch gun for coast defense. Scientific American, 74(22):346, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
346.pdf.

Anonymous:1896:HGS

[6794] Anonymous. Hailstones of great size. Scientific American, 74(22):347, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-347.pdf.

Anonymous:1896:BNBn

[6795] Anonymous. Books: New books and publications. Scientific American, 74(22):348, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
348a.pdf.

Anonymous:1896:BPv

[6796] Anonymous. Business and personal. Scientific American, 74(22):348, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-348b.pdf.

Anonymous:1896:RPIv

[6797] Anonymous. Recently patented inventions. Scientific American, 74(22):348, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n22/pdf/scientificamerican05301896-
348.pdf.

Anonymous:1896:NQv

[6798] Anonymous. Notes and queries. Scientific American, 74(22):348–349, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-348c.pdf.

Anonymous:1896:Ix

[6799] Anonymous. To inventors. Scientific American, 74(22):349, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-349.pdf.

Anonymous:1896:IIw

[6800] Anonymous. Index of inventions. Scientific American, 74(22):349, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-349a.pdf.

Anonymous:1896:PNPr

[6801] Anonymous. Patent notes: Patents. Scientific American, 74(22):351, May 30, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n22/pdf/scientificamerican05301896-351.pdf.

Anonymous:1896:ERR

[6802] Anonymous. The elevated roads and the Raid Transit Problem. Scientific American, 74(23):354, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
354.pdf.

Anonymous:1896:PEN

[6803] Anonymous. Proposed equipment of the New York surface roads with compressed air motors. Scientific American, 74(23):354–355, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-354a.pdf.

Anonymous:1896:ACCa

[6804] Anonymous. An automatic car coupler. Scientific American, 74(23):356, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-356.pdf.

Anonymous:1896:DTC

[6805] Anonymous. A double tank compressed air mine locomotive. Scientific American, 74(23):356, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
356a.pdf.

Anonymous:1896:TFCb

[6806] Anonymous. A tube and flue cutter. Scientific American, 74(23):356, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-356b.pdf.

Anonymous:1896:NRBa

[6807] Anonymous. A novel railroad bridge. Scientific American, 74 (23):357–358, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
357.pdf.

Anonymous:1896:NFEb

[6808] Anonymous. A novel fire escape. Scientific American, 74(23):358–359, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-358.pdf.

Anonymous:1896:INR

[6809] Anonymous. The improvement of the North River Water Front, New York City. Scientific American, 74(23):360, June 6, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n23/pdf/scientificamerican06061896-360.pdf.

Anonymous:1896:AFB

[6810] Anonymous. Anti-friction ball bearings. Scientific American, 74 (23):361–362, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
361.pdf.

Anonymous:1896:CAB

[6811] Anonymous. A crib attachment for beds. Scientific American, 74(23):362, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
362.pdf.

Anonymous:1896:IHTb

[6812] Anonymous. An improved holder for tool sharpening. Scientific American, 74(23):363, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
363.pdf.

Anonymous:1896:PEa

[6813] Anonymous. The permeability of the Earth. Scientific American, 74(23):363, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
363a.pdf.

Anonymous:1896:BNBo

[6814] Anonymous. Books: New books and publications. Scientific American, 74(23):364, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
364a.pdf.

Anonymous:1896:BPw

[6815] Anonymous. Business and personal. Scientific American, 74(23):364, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-364b.pdf.

Anonymous:1896:NQw

[6816] Anonymous. Notes and queries. Scientific American, 74(23):364, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-364c.pdf.

Anonymous:1896:RPIw

[6817] Anonymous. Recently patented inventions. Scientific American, 74(23):364, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n23/pdf/scientificamerican06061896-
364.pdf.

Anonymous:1896:IIx

[6818] Anonymous. Index of inventions. Scientific American, 74(23):364–365, June 6, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-364d.pdf.

Anonymous:1896:PNPs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n23/pdf/scientificamerican06061896-367.pdf.

Anonymous:1896:CYW

[6820] Anonymous. Construction of the Yarrow water tube boiler. Scientific American, 74(24):369, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
369.pdf.

Anonymous:1896:OSC

[6821] Anonymous. Our semi-centennial essay competition. Scientific American, 74(24):370, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
370.pdf.

Anonymous:1896:PSP

[6822] Anonymous. Patent solicitors and the proposed patent bar. Scientific American, 74(24):370, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
370a.pdf.

Anonymous:1896:CDa

[6823] Anonymous. Coast defense. Scientific American, 74(24):371, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-371a.pdf.

Anonymous:1896:RHB

[6824] Anonymous. The River and Harbor Bill. Scientific American, 74(24):371, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
371.pdf.

Anonymous:1896:HCS

[6825] Anonymous. A handy camera stand. Scientific American, 74(24):372, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-372.pdf.

Anonymous:1896:IGE

[6826] Anonymous. Improved gas engine. Scientific American, 74(24):372, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-372a.pdf.

Anonymous:1896:PAC

[6827] Anonymous. Patent automatic copying lathe with automatic screw feed. Scientific American, 74(24):372, June 13, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v74/
n24/pdf/scientificamerican06131896-372b.pdf.

Anonymous:1896:Co

[6828] Anonymous. Coppersmithing. Scientific American, 74(24):373, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-373a.pdf.

Anonymous:1896:CP

[6829] Anonymous. A cushion pavement. Scientific American, 74(24):373, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-373.pdf.

Anonymous:1896:IBFa

[6830] Anonymous. Improved boiler furnace. Scientific American, 74(24):374, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-374a.pdf.

Anonymous:1896:IRD

[6831] Anonymous. Insect remover and destroyer. Scientific American, 74(24):374, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
374.pdf.

Anonymous:1896:Cq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-375.pdf.

Anonymous:1896:ASUa

[6833] Anonymous. Archimedean screw used for drawing water. Scientific American, 74(24):376, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
376.pdf.

Anonymous:1896:OTJ

[6834] Anonymous. Orange tree jacket. Scientific American, 74(24):376, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-376a.pdf.

Anonymous:1896:HCR

[6835] Anonymous. The horseless carriage race on decoration day. Scientific American, 74(24):377–378, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
377.pdf.

Anonymous:1896:ENE

[6836] Anonymous. Edison’s new electric light. Scientific American, 74 (24):378, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
378.pdf.

Anonymous:1896:SSO

[6837] Anonymous. Sun spots observed May 31, 1896. Scientific American, 74(24):379, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
379.pdf.

Woodmancy:1896:DBE

[6838] Frank Woodmancy. A destructive boiler explosion. Scientific American, 74(24):379, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
379a.pdf.

Anonymous:1896:BNBp

[6839] Anonymous. Books: New books and publications. Scientific American, 74(24):380, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
380a.pdf.

Anonymous:1896:BPx

[6840] Anonymous. Business and personal. Scientific American, 74(24):380, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-380b.pdf.

Anonymous:1896:NQx

[6841] Anonymous. Notes and queries. Scientific American, 74(24):380, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-380c.pdf.

Anonymous:1896:RPIx

[6842] Anonymous. Recently patented inventions. Scientific American, 74(24):380, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n24/pdf/scientificamerican06131896-
380.pdf.

Anonymous:1896:Iy

[6843] Anonymous. To inventors. Scientific American, 74(24):381, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-381.pdf.

Anonymous:1896:IIy

[6844] Anonymous. Index of inventions. Scientific American, 74(24):381, June 13, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-381a.pdf.

Anonymous:1896:PNPt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n24/pdf/scientificamerican06131896-383.pdf.

Anonymous:1896:MCC

[6846] Anonymous. Moscow and the Coronation of the Czar. Scientific American, 74(25):385, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n25/pdf/scientificamerican06201896-
385.pdf.

Anonymous:1896:PPI

[6847] Anonymous. Patience and perseverance in invention. Scientific American, 74(25):386–387, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n25/pdf/scientificamerican06201896-
386.pdf.

Anonymous:1896:BND

[6848] Anonymous. Books: New departure in letter filing cabinets. Scientific American, 74(25):388, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n25/pdf/scientificamerican06201896-
388.pdf.

Anonymous:1896:ELH

[6849] Anonymous. An electric lamp hanger. Scientific American, 74(25):388, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-388b.pdf.

Anonymous:1896:ISMb

[6850] Anonymous. An improved saw mill. Scientific American, 74(25):388, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-388a.pdf.

Anonymous:1896:PMB

[6851] Anonymous. The production of metallic bars of any section by extrusion at high temperatures. Scientific American, 74(25):389–390, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-389.pdf.

Anonymous:1896:CBSb

[6852] Anonymous. A combined bed and sofa. Scientific American, 74 (25):390–391, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n25/pdf/scientificamerican06201896-
390.pdf.

Anonymous:1896:SD

[6853] Anonymous. The skirt dance. Scientific American, 74(25):392, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-392.pdf.

Anonymous:1896:BHV

[6854] Anonymous. Bessiere’s hydraulic ventilator. Scientific American, 74(25):393–394, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n25/pdf/scientificamerican06201896-
393.pdf.

Anonymous:1896:ISb

[6855] Anonymous. An inverting stereoscope. Scientific American, 74(25):395, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-395a.pdf.

Anonymous:1896:V

[6856] Anonymous. The viviscope. Scientific American, 74(25):395, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-395.pdf.

Anonymous:1896:BNBq

[6857] Anonymous. Books: New books and publications. Scientific American, 74(25):396, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n25/pdf/scientificamerican06201896-
396a.pdf.

Anonymous:1896:RPIy 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n25/pdf/scientificamerican06201896-
396.pdf.

Anonymous:1896:BPy

[6859] Anonymous. Business and personal. Scientific American, 74(25):397, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-397.pdf.

Anonymous:1896:IIz

[6860] Anonymous. Index of inventions. Scientific American, 74(25):397, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-397b.pdf.

Anonymous:1896:NQy

[6861] Anonymous. Notes and queries. Scientific American, 74(25):397, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-397a.pdf.

Anonymous:1896:PNPu

[6862] Anonymous. Patent notes: Patents. Scientific American, 74(25):398, June 20, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n25/pdf/scientificamerican06201896-398.pdf.

Anonymous:1896:NDD

[6863] Anonymous. The new double-deck-turret battleship Kearsarge. Scientific American, 74(26):401, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n26/pdf/scientificamerican06271896-
401.pdf.

Anonymous:1896:EPT

[6864] Anonymous. The elevator problem in tall office buildings. Scientific American, 74(26):402–404, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v74/n26/pdf/scientificamerican06271896-
402.pdf.

Anonymous:1896:CLE

[6865] Anonymous. Curious locomotive explosion in Peru. Scientific American, 74(26):404, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n26/pdf/scientificamerican06271896-
404.pdf.

Anonymous:1896:TBb

[6866] Anonymous. Truck bearing. Scientific American, 74(26):404, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-404a.pdf.

Anonymous:1896:QTD

[6867] Anonymous. Quint’s turret drill. Scientific American, 74(26):405, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-405.pdf.

Anonymous:1896:SVC

[6868] Anonymous. Seal for vessels containing liquids. Scientific American, 74(26):405, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n26/pdf/scientificamerican06271896-
405a.pdf.

Anonymous:1896:GSU

[6869] Anonymous. Grain shovel for unloading cars. Scientific American, 74(26):405–408, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n26/pdf/scientificamerican06271896-
405b.pdf.

Anonymous:1896:ASGa

[6870] Anonymous. An adjustable saw guide. Scientific American, 74(26):408, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-408.pdf.

Barrows:1896:DV

[6871] John C. Barrows. A destroying vacuum. Scientific American, 74 (26):408–410, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n26/pdf/scientificamerican06271896-
408a.pdf.

Anonymous:1896:FTP

[6872] Anonymous. A Florida tree palm. Scientific American, 74(26):410, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-410.pdf.

Anonymous:1896:BPz

[6873] Anonymous. Business and personal. Scientific American, 74(26):411, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-411a.pdf.

Anonymous:1896:Iz

[6874] Anonymous. To inventors. Scientific American, 74(26):411, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-411c.pdf.

Anonymous:1896:NQz

[6875] Anonymous. Notes and queries. Scientific American, 74(26):411, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-411b.pdf.

Anonymous:1896:RPIz

[6876] Anonymous. Recently patented inventions. Scientific American, 74(26):411, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v74/n26/pdf/scientificamerican06271896-
411.pdf.

Anonymous:1896:IIba

[6877] Anonymous. Index of inventions. Scientific American, 74(26):411–414, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-411d.pdf.

Anonymous:1896:PNPv

[6878] Anonymous. Patent notes: Patents. Scientific American, 74(26):415, June 27, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v74/n26/pdf/scientificamerican06271896-415.pdf.

Anonymous:1896:OES

[6879] Anonymous. Oil engine signaling plant on United States Lightship No. 42. Scientific American, 75(1):1–2, July 4, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n1/pdf/scientificamerican07041896-1.pdf.

Anonymous:1896:ASUb

[6880] Anonymous. The apprenticeship system under modern conditions. Scientific American, 75(1):2–3, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n1/pdf/scientificamerican07041896-
2.pdf.

Anonymous:1896:GSA

[6881] Anonymous. The Geological Society of America. Scientific American, 75(1):3, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n1/pdf/scientificamerican07041896-
3.pdf.

Anonymous:1896:IVa

[6882] Anonymous. An ice velocipede. Scientific American, 75(1):4, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-4.pdf.

Anonymous:1896:VBD

[6883] Anonymous. The Victoria Bridge Disaster. Scientific American, 75(1): 4, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-4a.pdf.

Anonymous:1896:ACCb

[6884] Anonymous. Automatic car coupling. Scientific American, 75(1):5, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-5a.pdf.

Anonymous:1896:IWT

[6885] Anonymous. An iron water tower. Scientific American, 75(1):5, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-5.pdf.

Anonymous:1896:IDB

[6886] Anonymous. Indicating device for bottles. Scientific American, 75(1):5– 6, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-5b.pdf.

Anonymous:1896:ISL

[6887] Anonymous. An improved signal lantern. Scientific American, 75(1):6, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-6.pdf.

Anonymous:1896:Cr

[6888] Anonymous. Correspondence. Scientific American, 75(1):7, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-7.pdf.

Anonymous:1896:BTR

[6889] Anonymous. Ballistic test of reforged armor plate for the Russia. Scientific American, 75(1):8, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n1/pdf/scientificamerican07041896-
8.pdf.

Anonymous:1896:GTa

[6890] Anonymous. A gigantic turtle. Scientific American, 75(1):9–10, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-9.pdf.

Anonymous:1896:VLC

[6891] Anonymous. The vegetation of Lower California. Scientific American, 75(1):10–11, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n1/pdf/scientificamerican07041896-
10.pdf.

Anonymous:1896:BPba

[6892] Anonymous. Business and personal. Scientific American, 75(1):12, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-12a.pdf.

Anonymous:1896:NQba

[6893] Anonymous. Notes and queries. Scientific American, 75(1):12, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-12b.pdf.

Anonymous:1896:RPIba

[6894] Anonymous. Recently patented inventions. Scientific American, 75(1): 12, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-12.pdf.

Anonymous:1896:IIbb

[6895] Anonymous. Index of inventions. Scientific American, 75(1):12–13, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-12c.pdf.

Anonymous:1896:PNPw

[6896] Anonymous. Patent notes: Patents. Scientific American, 75(1):14, July 4, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n1/pdf/scientificamerican07041896-14.pdf.

Anonymous:1896:TPM

[6897] Anonymous. Testing the parts of modern bicycle. Scientific American, 75(2):17, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n2/pdf/scientificamerican07111896-
17.pdf.

Anonymous:1896:PEC

[6898] Anonymous. The Prize Essay Competition. Scientific American, 75(2):18, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n2/pdf/scientificamerican07111896-
18.pdf.

Anonymous:1896:PEb

[6899] Anonymous. The Paris Exposition of 1900. Scientific American, 75(2):18–20, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n2/pdf/scientificamerican07111896-
18a.pdf.

Anonymous:1896:DTR

[6900] Anonymous. The Duct trolley rail. Scientific American, 75(2):20, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-20a.pdf.

Anonymous:1896:SCB

[6901] Anonymous. Safety check for boilers. Scientific American, 75(2):20, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-20.pdf.

Anonymous:1896:SMV

[6902] Anonymous. The Sheffield motor velocipede car. Scientific American, 75(2):21–22, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n2/pdf/scientificamerican07111896-
21.pdf.

Anonymous:1896:Cs

[6903] Anonymous. Correspondence. Scientific American, 75(2):22, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-22.pdf.

Morand:1896:RDB

[6904] Henri Morand. Recently discovered bust of Louis XVII. Scientific American, 75(2):24, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n2/pdf/scientificamerican07111896-
24.pdf.

Anonymous:1896:FAD

[6905] Anonymous. Fireworks as an adjunct to dramatic entertainment. Scientific American, 75(2):25–27, July 11, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n2/pdf/scientificamerican07111896-25.pdf.

Anonymous:1896:IXR

[6906] Anonymous. Inexpensive X ray apparatus. Scientific American, 75(2):27, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n2/pdf/scientificamerican07111896-
27.pdf.

Anonymous:1896:Dc

[6907] Anonymous. The dancers. Scientific American, 75(2):27–28, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-27a.pdf.

Anonymous:1896:BPbb

[6908] Anonymous. Business and personal. Scientific American, 75(2):29, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-29a.pdf.

Anonymous:1896:RPIbb

[6909] Anonymous. Recently patented inventions. Scientific American, 75(2):29, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n2/pdf/scientificamerican07111896-
29.pdf.

Anonymous:1896:NQbb

[6910] Anonymous. Notes and queries. Scientific American, 75(2):29–30, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-29b.pdf.

Anonymous:1896:LP

[6911] Anonymous. The list of patents. Scientific American, 75(2):30, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-30.pdf.

Anonymous:1896:PNPx

[6912] Anonymous. Patent notes: Patents. Scientific American, 75(2):31, July 11, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n2/pdf/scientificamerican07111896-31.pdf.

Anonymous:1896:ESLa

[6913] Anonymous. The engines of the St. Louis and St. Paul. Scientific American, 75(3):33–34, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n3/pdf/scientificamerican07181896-
33.pdf.

Anonymous:1896:OJN

[6914] Anonymous. Our jubilee number. Scientific American, 75(3):34, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-34.pdf.

Anonymous:1896:AEEa

[6915] Anonymous. The artistic element in engineering works. Scientific American, 75(3):34–35, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n3/pdf/scientificamerican07181896-
34a.pdf.

Anonymous:1896:IGC

[6916] Anonymous. The International Geological Congress at St. Petersburg, August, 1897. Scientific American, 75(3):35, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-35.pdf.

Anonymous:1896:CSF

[6917] Anonymous. Curtain or shade fixture. Scientific American, 75(3):36, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-36.pdf.

Anonymous:1896:IWMa

[6918] Anonymous. Improved washing machine. Scientific American, 75(3):36, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-36a.pdf.

Mercer:1896:CEE

[6919] Henry C. Mercer. Cave exploration in the Eastern United States. Scientific American, 75(3):36–37, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n3/pdf/scientificamerican07181896-
36b.pdf.

Anonymous:1896:POW

[6920] Anonymous. Preparing old woolen rags for shoddy clothing. Scientific American, 75(3):37–38, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n3/pdf/scientificamerican07181896-
37.pdf.

Anonymous:1896:PI

[6921] Anonymous. A pneumatic inkstand. Scientific American, 75(3):38–39, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-38.pdf.

Anonymous:1896:AMa

[6922] Anonymous. Accurate measurements. Scientific American, 75(3):40–41, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-40.pdf.

Lemosof:1896:MM

[6923] P. Lemosof. The Matabeles and Mashonas. Scientific American, 75(3):41–42, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n3/pdf/scientificamerican07181896-
41.pdf.

Anonymous:1896:AD

[6924] Anonymous. An automatic duplicator. Scientific American, 75(3):42, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-42a.pdf.

Anonymous:1896:NSRa

[6925] Anonymous. A new steady rest for engine lathes. Scientific American, 75(3):42, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n3/pdf/scientificamerican07181896-
42.pdf.

Anonymous:1896:RPIbc

[6926] Anonymous. Recently patented inventions. Scientific American, 75(3):43–44, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n3/pdf/scientificamerican07181896-
43.pdf.

Anonymous:1896:BPbc

[6927] Anonymous. Business and personal. Scientific American, 75(3):44, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-44.pdf.

Anonymous:1896:Ibc

[6928] Anonymous. To inventors. Scientific American, 75(3):44, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-44b.pdf.

Anonymous:1896:NQbc

[6929] Anonymous. Notes and queries. Scientific American, 75(3):44, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-44a.pdf.

Anonymous:1896:IIbc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-44c.pdf.

Anonymous:1896:PNPy

[6931] Anonymous. Patent notes: Patents. Scientific American, 75(3):47, July 18, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n3/pdf/scientificamerican07181896-47.pdf.

Anonymous:1896:EIP

[6932] Anonymous. The effect of inventions on the People’s life. Scientific American, 75(4):50–51, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
50.pdf.

Anonymous:1896:PSa

[6933] Anonymous. The patent system. Scientific American, 75(4):51, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-51.pdf.

Anonymous:1896:TS

[6934] Anonymous. The transatlantic steamship. Scientific American, 75(4):52–55, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
52.pdf.

Anonymous:1896:RB

[6935] Anonymous. Railroads and bridges. Scientific American, 75(4):55–58, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-55.pdf.

Anonymous:1896:Tc

[6936] Anonymous. The telegraph. Scientific American, 75(4):58–59, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-58.pdf.

Anonymous:1896:Pa

[6937] Anonymous. Physics. Scientific American, 75(4):59–61, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-59.pdf.

Anonymous:1896:MPb

[6938] Anonymous. Men of progress. Scientific American, 75(4):61, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-61.pdf.

Anonymous:1896:TIU

[6939] Anonymous. The textile industries of the United States since 1846. Scientific American, 75(4):61–62, July 25, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n4/pdf/scientificamerican07251896-61a.pdf.

Anonymous:1896:SCc

[6940] Anonymous. The submarine cable. Scientific American, 75(4):62–63, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-62.pdf.

Anonymous:1896:FYPa

[6941] Anonymous. Fifty years of photography. Scientific American, 75(4):63–64, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
63.pdf.

Anonymous:1896:Cu

[6942] Anonymous. Chemistry. Scientific American, 75(4):64–65, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-64.pdf.

Anonymous:1896:Pb

[6943] Anonymous. The phonograph. Scientific American, 75(4):65–66, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-65.pdf.

Anonymous:1896:AL

[6944] Anonymous. The American locomotive. Scientific American, 75 (4):66–68, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
66.pdf.

Anonymous:1896:B

[6945] Anonymous. The bicycle. Scientific American, 75(4):68–69, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-68.pdf.

Anonymous:1896:PMG

[6946] Anonymous. The progress made in the generation of electric energy and its application to the operation of motors during the past fifty years. Scientific American, 75(4):69–72, July 25, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n4/pdf/scientificamerican07251896-69.pdf.

Anonymous:1896:SMb

[6947] Anonymous. The sewing machine. Scientific American, 75(4):72–73, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-72.pdf.

Anonymous:1896:AMb

[6948] Anonymous. Agricultural machinery. Scientific American, 75(4):74–76, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-74.pdf.

Anonymous:1896:NCDa

[6949] Anonymous. Naval and coast defense. Scientific American, 75(4):76–80, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-76.pdf.

Anonymous:1896:FYPb

[6950] Anonymous. Fifty years in the printing business. Scientific American, 75(4):80–81, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
80.pdf.

Anonymous:1896:PEO

[6951] Anonymous. The prize essay of our semi-centennial anniversary number. Scientific American, 75(4):82, July 25, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n4/pdf/scientificamerican07251896-82.pdf.

Anonymous:1896:U

[6952] Anonymous. [unknown]. Scientific American, 75(4):82, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-82a.pdf.

Anonymous:1896:PID

[6953] Anonymous. The progress of invention during the past fifty years. Scientific American, 75(4):82–83, July 25, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n4/pdf/scientificamerican07251896-82b.pdf.

Anonymous:1896:S

[6954] Anonymous. Steel. Scientific American, 75(4):83–84, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-83.pdf.

Anonymous:1896:DI

[6955] Anonymous. Distinguished inventors. Scientific American, 75(4):84–86, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-84.pdf.

Anonymous:1896:ASa

[6956] Anonymous. American shipbuilding. Scientific American, 75(4):86–88, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-86.pdf.

Anonymous:1896:DAT

[6957] Anonymous. Development of the astronomical telescope in fifty years. Scientific American, 75(4):88–89, July 25, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n4/pdf/scientificamerican07251896-88.pdf.

Anonymous:1896:Td

[6958] Anonymous. The telephone. Scientific American, 75(4):89–91, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-89.pdf.

Anonymous:1896:FYS

[6959] Anonymous. Fifty years of the Scientific American. Scientific American, 75(4):91–93, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
91.pdf.

Anonymous:1896:GSB

[6960] Anonymous. The Great Seal on British patents. Scientific American, 75(4):96, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
96a.pdf.

Anonymous:1896:HCM

[6961] Anonymous. A half century of mining. Scientific American, 75(4):96, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-96.pdf.

Anonymous:1896:UTR

[6962] Anonymous. Unit transportation revolutionized. Scientific American, 75(4):97, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n4/pdf/scientificamerican07251896-
97.pdf.

Anonymous:1896:BPbd

[6963] Anonymous. Business and personal. Scientific American, 75(4):98, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-98a.pdf.

Anonymous:1896:NQbd

[6964] Anonymous. Notes and queries. Scientific American, 75(4):98, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-98.pdf.

Anonymous:1896:Ibd

[6965] Anonymous. To inventors. Scientific American, 75(4):99, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-99.pdf.

Anonymous:1896:IIbd

[6966] Anonymous. Index of inventions. Scientific American, 75(4):99–102, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-99a.pdf.

Anonymous:1896:IAa

[6967] Anonymous. Index to advertisers. Scientific American, 75(4):114, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-114.pdf.

Anonymous:1896:PNPz

[6968] Anonymous. Patent notes: Patents. Scientific American, 75(4):115, July 25, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n4/pdf/scientificamerican07251896-115.pdf.

Anonymous:1896:HLD

[6969] Anonymous. Hydraulic lift dry dock at the Union Iron Works, San Francisco. Scientific American, 75(5):117, August 1, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n5/pdf/scientificamerican08011896-117.pdf.

Anonymous:1896:NPR

[6970] Anonymous. The new plan of the Rapid Transit Commission of New York City. Scientific American, 75(5):118, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-118a.pdf.

Anonymous:1896:OFAb

[6971] Anonymous. Our fiftieth anniversary number. Scientific American, 75(5):118, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n5/pdf/scientificamerican08011896-
118.pdf.

Anonymous:1896:SPSb

[6972] Anonymous. Some phenomena of the St. Louis Hurricane. Scientific American, 75(5):119, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n5/pdf/scientificamerican08011896-
119.pdf.

Anonymous:1896:TNB

[6973] Anonymous. The three new battleships, nos. 7, 8, 9. Scientific American, 75(5):119, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n5/pdf/scientificamerican08011896-
119a.pdf.

Anonymous:1896:GPE

[6974] Anonymous. Gasoline pumping engine. Scientific American, 75(5):121, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-121.pdf.

Anonymous:1896:AMI

[6975] Anonymous. Apparatus for measuring the intensity of the perfume of flowers. Scientific American, 75(5):121–122, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-121a.pdf.

Anonymous:1896:HSW

[6976] Anonymous. Hartmann’s spool wire clamp. Scientific American, 75(5):122, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n5/pdf/scientificamerican08011896-
122.pdf.

Anonymous:1896:Cv

[6977] Anonymous. Correspondence. Scientific American, 75(5):123, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-123.pdf.

Anonymous:1896:LMC

[6978] Anonymous. Lord Mayor’s coach. Scientific American, 75(5):124, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-124.pdf.

Anonymous:1896:SMc

[6979] Anonymous. A steam missionary. Scientific American, 75(5):124, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-124a.pdf.

Anonymous:1896:RST

[6980] Anonymous. The repair of single tube bicycle tires. Scientific American, 75(5):125–126, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n5/pdf/scientificamerican08011896-
125.pdf.

Anonymous:1896:RPIbd

[6981] Anonymous. Recently patented inventions. Scientific American, 75(5):127–128, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n5/pdf/scientificamerican08011896-
127.pdf.

Anonymous:1896:BNBr

[6982] Anonymous. Books: New books and publications. Scientific American, 75(5):128, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n5/pdf/scientificamerican08011896-
128.pdf.

Anonymous:1896:BPbe

[6983] Anonymous. Business and personal. Scientific American, 75(5):128, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-128a.pdf.

Anonymous:1896:NQbe

[6984] Anonymous. Notes and queries. Scientific American, 75(5):128–129, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-128b.pdf.

Anonymous:1896:Ibe

[6985] Anonymous. To inventors. Scientific American, 75(5):129, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-129.pdf.

Anonymous:1896:IIbe

[6986] Anonymous. Index of inventions. Scientific American, 75(5):129, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-129a.pdf.

Anonymous:1896:PNPba

[6987] Anonymous. Patent notes: Patents. Scientific American, 75(5):130, August 1, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n5/pdf/scientificamerican08011896-130.pdf.

Childers:1896:AD

[6988] H. H. Childers. The Austin Dam. Scientific American, 75(6):133, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-133.pdf.

Anonymous:1896:IITa

[6989] Anonymous. The importance of industrial and trade schools. Scientific American, 75(6):134, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
134.pdf.

Anonymous:1896:RPTa

[6990] Anonymous. Recent patent and trade mark decisions. Scientific American, 75(6):134–135, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
134a.pdf.

Anonymous:1896:ALS

[6991] Anonymous. Automatic locomotive signaling. Scientific American, 75(6):136, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
136.pdf.

Anonymous:1896:IBFb

[6992] Anonymous. An improved barbel-forming machine. Scientific American, 75(6):136, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
136b.pdf.

Anonymous:1896:NEM

[6993] Anonymous. A new electric mining locomotive. Scientific American, 75(6):136, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
136a.pdf.

Anonymous:1896:NNR

[6994] Anonymous. A new non-refillable bottle. Scientific American, 75(6):138, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
138.pdf.

Anonymous:1896:CTb

[6995] Anonymous. A card of thanks. Scientific American, 75(6):139, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-139.pdf.

Anonymous:1896:DG

[6996] Anonymous. Determinations of gravity. Scientific American, 75(6):140, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-140.pdf.

Anonymous:1896:ITa

[6997] Anonymous. The illusion of Trilby. Scientific American, 75(6):141–142, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-141.pdf.

Anonymous:1896:NPEb

[6998] Anonymous. New photograph enlarging apparatus. Scientific American, 75(6):142, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
142.pdf.

Anonymous:1896:KZG

[6999] Anonymous. A kangaroo at the Zoological Gardens, London. Scientific American, 75(6):143, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
143.pdf.

Anonymous:1896:BNBs

[7000] Anonymous. Books: New books and publications. Scientific American, 75(6):144, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
144a.pdf.

Anonymous:1896:RPIbe

[7001] Anonymous. Recently patented inventions. Scientific American, 75(6):144, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n6/pdf/scientificamerican08081896-
144.pdf.

Anonymous:1896:BPbf

[7002] Anonymous. Business and personal. Scientific American, 75(6):145, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-145.pdf.

Anonymous:1896:IIbf

[7003] Anonymous. Index of inventions. Scientific American, 75(6):145, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-145b.pdf.

Anonymous:1896:NQbf

[7004] Anonymous. Notes and queries. Scientific American, 75(6):145, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-145a.pdf.

Anonymous:1896:PNPbb

[7005] Anonymous. Patent notes: Patents. Scientific American, 75(6):147, August 8, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n6/pdf/scientificamerican08081896-147.pdf.

Anonymous:1896:MSW

[7006] Anonymous. The metric system of weights and measures. Scientific American, 75(7):150–151, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n7/pdf/scientificamerican08151896-
150.pdf.

Anonymous:1896:NAT

[7007] Anonymous. The Nivet apparatus for testing building materials. Scientific American, 75(7):152–153, August 15, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n7/pdf/scientificamerican08151896-152.pdf.

Anonymous:1896:TCA

[7008] Anonymous. Trial of the compressed air motor by the Third Avenue Railroad Company, New York. Scientific American, 75(7):153–154, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n7/pdf/scientificamerican08151896-153.pdf.

Anonymous:1896:SPE

[7009] Anonymous. The Society for the Promotion of Engineering Education. Scientific American, 75(7):154, August 15, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n7/pdf/scientificamerican08151896-154.pdf.

Anonymous:1896:Cw

[7010] Anonymous. Correspondence. Scientific American, 75(7):155, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n7/pdf/scientificamerican08151896-155.pdf.

Anonymous:1896:USF

[7011] Anonymous. The United States first class battleship Indiana. Scientific American, 75(7):156–158, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n7/pdf/scientificamerican08151896-
156.pdf.

Anonymous:1896:SCd

[7012] Anonymous. A sectional cam. Scientific American, 75(7):158, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n7/pdf/scientificamerican08151896-158.pdf.

Townsend:1896:NLP

[7013] C. F. Townsend. A new light for photographers. Scientific American, 75(7):158, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n7/pdf/scientificamerican08151896-
158a.pdf.

Anonymous:1896:TKR

[7014] Anonymous. The tomb of King René and Queen Isabella of Lorraine. Scientific American, 75(7):159, August 15, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n7/pdf/scientificamerican08151896-159.pdf.

Anonymous:1896:BNBt

[7015] Anonymous. Books: New books and publications. Scientific American, 75(7):160, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n7/pdf/scientificamerican08151896-
160a.pdf.

Anonymous:1896:RPIbf

[7016] Anonymous. Recently patented inventions. Scientific American, 75(7):160, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n7/pdf/scientificamerican08151896-
160.pdf.

Anonymous:1896:BPbg

[7017] Anonymous. Business and personal. Scientific American, 75(7):161, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n7/pdf/scientificamerican08151896-161.pdf.

Anonymous:1896:Ibg

[7018] Anonymous. To inventors. Scientific American, 75(7):161, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n7/pdf/scientificamerican08151896-161b.pdf.

Anonymous:1896:NQbg

[7019] Anonymous. Notes and queries. Scientific American, 75(7):161, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n7/pdf/scientificamerican08151896-161a.pdf.

Anonymous:1896:IIbg

[7020] Anonymous. Index of inventions. Scientific American, 75(7): 161–162, August 15, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n7/pdf/scientificamerican08151896-
161c.pdf.

Anonymous:1896:PNPbc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n7/pdf/scientificamerican08151896-163.pdf.

Anonymous:1896:RHW

[7022] Anonymous. The recent heat wave. Scientific American, 75(8):166, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n8/pdf/scientificamerican08221896-166.pdf.

Anonymous:1896:AFa

[7023] Anonymous. Artificial flight. Scientific American, 75(8):166–167, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n8/pdf/scientificamerican08221896-166a.pdf.

Anonymous:1896:ADD

[7024] Anonymous. Accident to Dry Dock No. 2 — New York Navy Yard. Scientific American, 75(8):168, August 22, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n8/pdf/scientificamerican08221896-168b.pdf.

Anonymous:1896:ASGb

[7025] Anonymous. An accessible sewer gas trap. Scientific American, 75(8):168, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
168.pdf.

Anonymous:1896:ATJ

[7026] Anonymous. Apparatus for treating juice and vapor. Scientific American, 75(8):168, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
168a.pdf.

Anonymous:1896:EPE

[7027] Anonymous. Express passenger engine, Northeastern Railway England. Scientific American, 75(8):169, August 22, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n8/pdf/scientificamerican08221896-169.pdf.

Benjamin:1896:EDC

[7028] Marcus Benjamin. Edward drinker cope. Scientific American, 75(8):170, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
170.pdf.

Anonymous:1896:NPEc

[7029] Anonymous. Nansen’s Polar Expedition. Scientific American, 75 (8):171, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
171.pdf.

Anonymous:1896:FCB

[7030] Anonymous. The first class battleship Indiana. Scientific American, 75(8):172–174, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
172.pdf.

Brooks:1896:BPC

[7031] William R. Brooks. The Brooks periodical comet. Scientific American, 75(8):174–175, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
174.pdf.

Anonymous:1896:BNBu

[7032] Anonymous. Books: New books and publications. Scientific American, 75(8):176, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
176a.pdf.

Anonymous:1896:BPbh

[7033] Anonymous. Business and personal. Scientific American, 75(8):176, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n8/pdf/scientificamerican08221896-176b.pdf.

Anonymous:1896:RPIbg

[7034] Anonymous. Recently patented inventions. Scientific American, 75(8):176, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
176.pdf.

Anonymous:1896:NQbh

[7035] Anonymous. Notes and queries. Scientific American, 75(8):176– 177, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n8/pdf/scientificamerican08221896-
176c.pdf.

Anonymous:1896:Ibh

[7036] Anonymous. To inventors. Scientific American, 75(8):177, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n8/pdf/scientificamerican08221896-177.pdf.

Anonymous:1896:IIbh

[7037] Anonymous. Index of inventions. Scientific American, 75(8):177, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n8/pdf/scientificamerican08221896-177a.pdf.

Anonymous:1896:PNPbd

[7038] Anonymous. Patent notes: Patents. Scientific American, 75(8):178, August 22, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n8/pdf/scientificamerican08221896-178.pdf.

Anonymous:1896:JPR

[7039] Anonymous. The Jerome Park Reservoir, New York City. Scientific American, 75(9):181, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
181.pdf.

Anonymous:1896:RFA

[7040] Anonymous. The recent failures of Arctic Expeditions. Scientific American, 75(9):182–183, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
182.pdf.

Anonymous:1896:Cx

[7041] Anonymous. Correspondence. Scientific American, 75(9):184, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n9/pdf/scientificamerican08291896-184b.pdf.

Anonymous:1896:NCA

[7042] Anonymous. A novel car awning ventilator. Scientific American, 75(9):184, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
184.pdf.

Anonymous:1896:NSRb

[7043] Anonymous. A new self-regulating wind wheel. Scientific American, 75(9):184, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
184c.pdf.

Anonymous:1896:RBS

[7044] Anonymous. To ring a bell without swinging it. Scientific American, 75(9):184, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
184a.pdf.

Anonymous:1896:IVW

[7045] Anonymous. An improved vehicle wheel. Scientific American, 75 (9):184–185, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
184d.pdf.

Prindle:1896:MRE

[7046] E. J. Prindle. The modern revolves — the evolution of a typical invention. Scientific American, 75(9):185, August 29, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n9/pdf/scientificamerican08291896-185.pdf.

Anonymous:1896:FN

[7047] Anonymous. The Farthest North. Scientific American, 75(9):187, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n9/pdf/scientificamerican08291896-187.pdf.

Anonymous:1896:CTW

[7048] Anonymous. Cycle telegraphs in war. Scientific American, 75 (9):187–188, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
187a.pdf.

Anonymous:1896:HGM

[7049] Anonymous. The Homestake Gold Mines and Stamp Mills, South Dakota. Scientific American, 75(9):188–190, August 29, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n9/pdf/scientificamerican08291896-188.pdf.

Anonymous:1896:TSS

[7050] Anonymous. Trajan’s Ship of State. Scientific American, 75(9):191, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n9/pdf/scientificamerican08291896-191.pdf.

Anonymous:1896:BNBv

[7051] Anonymous. Books: New books and publications. Scientific American, 75(9):192, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
192a.pdf.

Anonymous:1896:RPIbh

[7052] Anonymous. Recently patented inventions. Scientific American, 75(9):192, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
192.pdf.

Anonymous:1896:BPbi

[7053] Anonymous. Business and personal. Scientific American, 75(9):193, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n9/pdf/scientificamerican08291896-193.pdf.

Anonymous:1896:Ibi

[7054] Anonymous. To inventors. Scientific American, 75(9):193, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n9/pdf/scientificamerican08291896-193b.pdf.

Anonymous:1896:NQbi

[7055] Anonymous. Notes and queries. Scientific American, 75(9):193, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n9/pdf/scientificamerican08291896-193a.pdf.

Anonymous:1896:IIbi

[7056] Anonymous. Index of inventions. Scientific American, 75(9): 193–194, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n9/pdf/scientificamerican08291896-
193c.pdf.

Anonymous:1896:PNPbe

[7057] Anonymous. Patent notes: Patents. Scientific American, 75(9):196, August 29, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n9/pdf/scientificamerican08291896-196.pdf.

Anonymous:1896:BES

[7058] Anonymous. The Boston Electric Subway. Scientific American, 75(10):197, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
197.pdf.

Anonymous:1896:PMa

[7059] Anonymous. Patent medicines. Scientific American, 75(10): 198, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
198.pdf.

Anonymous:1896:PAA

[7060] Anonymous. Proceedings of the American Association at Buffalo, N.Y. Scientific American, 75(10):198–199, September 5, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n10/pdf/scientificamerican09051896-198a.pdf.

Anonymous:1896:EEL

[7061] Anonymous. An extension electric lamp holder. Scientific American, 75(10):200, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
200a.pdf.

Anonymous:1896:ITb

[7062] Anonymous. An improved telemeter. Scientific American, 75 (10):200, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
200b.pdf.

Anonymous:1896:SHRa

[7063] Anonymous. The Savage hammerless rifle — model 1895. Scientific American, 75(10):200, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
200.pdf.

Anonymous:1896:IMM

[7064] Anonymous. An improved milling machine. Scientific American, 75(10):201, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
201.pdf.

Anonymous:1896:VLH

[7065] Anonymous. The visit of Li hung chang. Scientific American, 75(10):202, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
202.pdf.

Levison:1896:CLH

[7066] Wallace Goold Levison. On the comparison of low and high vacuum electrical and radiant matter phenomena with the aurora, the solar corona, and comets. Scientific American, 75(10):203, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n10/pdf/scientificamerican09051896-203.pdf.

Anonymous:1896:RPTb

[7067] Anonymous. Recent patent and trade mark decisions. Scientific American, 75(10):206–207, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
206.pdf.

Anonymous:1896:DBS

[7068] Anonymous. Discovery of a bronze statue at Delphi. Scientific American, 75(10):207, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
207.pdf.

Anonymous:1896:BNBw

[7069] Anonymous. Books: New books and publications. Scientific American, 75(10):208, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
208a.pdf.

Anonymous:1896:RPIbi

[7070] Anonymous. Recently patented inventions. Scientific American, 75(10):208, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
208.pdf.

Anonymous:1896:BPbj

[7071] Anonymous. Business and personal. Scientific American, 75 (10):209, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
209.pdf.

Anonymous:1896:IIbj

[7072] Anonymous. Index of inventions. Scientific American, 75(10): 209, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
209b.pdf.

Anonymous:1896:NQbj

[7073] Anonymous. Notes and queries. Scientific American, 75(10): 209, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
209a.pdf.

Anonymous:1896:PNPbf

[7074] Anonymous. Patent notes: Patents. Scientific American, 75 (10):211, September 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n10/pdf/scientificamerican09051896-
211.pdf.

Anonymous:1896:NER

[7075] Anonymous. The New East River Bridge. Scientific American, 75 (11):213, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
213.pdf.

Anonymous:1896:LHC

[7076] Anonymous. Li Hung Chang and our commercial relations with China. Scientific American, 75(11):214, September 12, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n11/pdf/scientificamerican09121896-214.pdf.

Hovey:1896:NAAa

[7077] Horace C. Hovey. Notes on the American Association Meeting, Buffalo, N.Y. Scientific American, 75(11):215, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n11/pdf/scientificamerican09121896-215.pdf.

Anonymous:1896:NSB

[7078] Anonymous. A new Swiss bicycle. Scientific American, 75(11): 216, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
216.pdf.

Anonymous:1896:NSE

[7079] Anonymous. A novel steam engine. Scientific American, 75 (11):216, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
216b.pdf.

Anonymous:1896:WMT

[7080] Anonymous. A wheat mixer and temperer. Scientific American, 75(11):216, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
216a.pdf.

Anonymous:1896:IPb

[7081] Anonymous. An improved plow. Scientific American, 75(11): 217, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
217b.pdf.

Anonymous:1896:PER

[7082] Anonymous. Portable electric rail sawing machine. Scientific American, 75(11):217, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
217a.pdf.

Anonymous:1896:PSb

[7083] Anonymous. A portable Sawmill. Scientific American, 75(11): 217, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
217.pdf.

Anonymous:1896:Cy

[7084] Anonymous. Correspondence. Scientific American, 75(11):219, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n11/pdf/scientificamerican09121896-219.pdf.

Murray:1896:CI

[7085] James Murray. Collision with an iceberg. Scientific American, 75 (11):220, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
220.pdf.

Anonymous:1896:BRS

[7086] Anonymous. The Bazin roller ship. Scientific American, 75(11): 221–222, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
221.pdf.

Anonymous:1896:REH

[7087] Anonymous. A rubber eraser holder. Scientific American, 75 (11):222, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
222.pdf.

Anonymous:1896:IPH

[7088] Anonymous. Iron pierced by hailstones. Scientific American, 75 (11):223, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
223.pdf.

Anonymous:1896:PBa

[7089] Anonymous. The propulsion of barges. Scientific American, 75 (11):223–224, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
223a.pdf.

Anonymous:1896:BPbk

[7090] Anonymous. Business and personal. Scientific American, 75 (11):225, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
225a.pdf.

Anonymous:1896:NQbk

[7091] Anonymous. Notes and queries. Scientific American, 75(11): 225, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
225b.pdf.

Anonymous:1896:RPIbj

[7092] Anonymous. Recently patented inventions. Scientific American, 75 (11):225, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
225.pdf.

Anonymous:1896:IIbk

[7093] Anonymous. Index of inventions. Scientific American, 75(11): 225–226, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
225c.pdf.

Anonymous:1896:PNPbg

[7094] Anonymous. Patent notes: Patents. Scientific American, 75 (11):228, September 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n11/pdf/scientificamerican09121896-
228.pdf.

Anonymous:1896:CRS

[7095] Anonymous. The Columbia River salmon fisheries. Scientific American, 75(12):229, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
229.pdf.

Anonymous:1896:ACE

[7096] Anonymous. The American car on English roads. Scientific American, 75(12):230, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
230a.pdf.

Anonymous:1896:BRG

[7097] Anonymous. The bicycle in its relation to good roads. Scientific American, 75(12):230, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
230.pdf.

Hovey:1896:NAAb

[7098] Horace C. Hovey. Notes on the American Association Meeting, Buffalo, N.Y. Scientific American, 75(12):231, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n12/pdf/scientificamerican09191896-231.pdf.

Anonymous:1896:BIEb

[7099] Anonymous. The Berlin Industrial Exhibition. Scientific American, 75(12):233–234, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
233.pdf.

Anonymous:1896:NBBb

[7100] Anonymous. A novel ball bearing for car trucks. Scientific American, 75(12):234, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
234.pdf.

Anonymous:1896:Cz

[7101] Anonymous. Correspondence. Scientific American, 75(12):235, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n12/pdf/scientificamerican09191896-235.pdf.

Kibbie:1896:CHCa

[7102] William P. Kibbie. The Calumet and Hecla Copper Mine. Scientific American, 75(12):236, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
236.pdf.

Anonymous:1896:LER

[7103] Anonymous. Lightning express railway service. Scientific American, 75(12):237–238, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
237.pdf.

Anonymous:1896:BRR

[7104] Anonymous. The bicycle relay race across the continent. Scientific American, 75(12):238, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
238.pdf.

Anonymous:1896:MBb

[7105] Anonymous. The magic bottle. Scientific American, 75(12): 239, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
239.pdf.

Anonymous:1896:BGW

[7106] Anonymous. Bicycle “gear” — what it means. Scientific American, 75(12):239–240, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
239a.pdf.

Anonymous:1896:RPIbk

[7107] Anonymous. Recently patented inventions. Scientific American, 75 (12):240–241, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
240.pdf.

Anonymous:1896:BNBx

[7108] Anonymous. Books: New books and publications. Scientific American, 75(12):241, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
241.pdf.

Anonymous:1896:BPbl

[7109] Anonymous. Business and personal. Scientific American, 75 (12):241, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
241a.pdf.

Anonymous:1896:Ibj

[7110] Anonymous. To inventors. Scientific American, 75(12):241, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n12/pdf/scientificamerican09191896-241c.pdf.

Anonymous:1896:NQbl

[7111] Anonymous. Notes and queries. Scientific American, 75(12): 241, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
241b.pdf.

Anonymous:1896:IIbl

[7112] Anonymous. Index of inventions. Scientific American, 75(12): 241–242, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
241d.pdf.

Anonymous:1896:PNPbh

[7113] Anonymous. Patent notes: Patents. Scientific American, 75 (12):243, September 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n12/pdf/scientificamerican09191896-
243.pdf.

Grimshaw:1896:RSB

[7114] Robert Grimshaw. The rigid suspension bridge at Loschwitz, Saxony. Scientific American, 75(13):245, September 26, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n13/pdf/scientificamerican09261896-245.pdf.

Anonymous:1896:SGI

[7115] Anonymous. The secrets of Germany’s industrial success. Scientific American, 75(13):246–247, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
246.pdf.

Anonymous:1896:ESLb

[7116] Anonymous. An electric self-loading car for street cleaning. Scientific American, 75(13):249, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
249.pdf.

Anonymous:1896:IFWb

[7117] Anonymous. An improved feed water heater. Scientific American, 75 (13):249–250, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
249a.pdf.

Anonymous:1896:CFP

[7118] Anonymous. The “Climax” fountain pen. Scientific American, 75 (13):250, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
250.pdf.

Anonymous:1896:Cbb

[7119] Anonymous. Correspondence. Scientific American, 75(13):251, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n13/pdf/scientificamerican09261896-251.pdf.

Kibbie:1896:CHCb

[7120] William P. Kibbie. The Calumet and Hecla Copper Mine. Scientific American, 75(13):252–253, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
252.pdf.

Anonymous:1896:PHC

[7121] Anonymous. The Providence horseless carriage race. Scientific American, 75(13):253–254, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
253.pdf.

Vomsaal:1896:WCC

[7122] F. C. V. H. Vomsaal. Waterspout off cottage city, Martha’s vineyard. Scientific American, 75(13):255–256, September 26, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n13/pdf/scientificamerican09261896-255.pdf.

Anonymous:1896:RPIbl

[7123] Anonymous. Recently patented inventions. Scientific American, 75 (13):256–257, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
256.pdf.

Anonymous:1896:BNBy

[7124] Anonymous. Books: New books and publications. Scientific American, 75(13):257, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
257.pdf.

Anonymous:1896:BPbm

[7125] Anonymous. Business and personal. Scientific American, 75 (13):257, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
257a.pdf.

Anonymous:1896:Ibk

[7126] Anonymous. To inventors. Scientific American, 75(13):257, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n13/pdf/scientificamerican09261896-257c.pdf.

Anonymous:1896:NQbm

[7127] Anonymous. Notes and queries. Scientific American, 75(13): 257, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
257b.pdf.

Anonymous:1896:IIbm

[7128] Anonymous. Index of inventions. Scientific American, 75(13): 257–258, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
257d.pdf.

Anonymous:1896:PNPbi

[7129] Anonymous. Patent notes: Patents. Scientific American, 75 (13):260, September 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n13/pdf/scientificamerican09261896-
260.pdf.

Hawks:1896:GDD

[7130] A. M’L. Hawks. Government Dry Dock at Port Orchard, Puget Sound, Washington. Scientific American, 75(14):261, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-261.pdf.

Anonymous:1896:AEEb

[7131] Anonymous. The artistic element in engineering. Scientific American, 75(14):262, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n14/pdf/scientificamerican10031896-
262.pdf.

Brooks:1896:HO

[7132] William R. Brooks. The heavens for October. Scientific American, 75(14):263, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n14/pdf/scientificamerican10031896-
263.pdf.

Anonymous:1896:IRF

[7133] Anonymous. An improved rail fastening. Scientific American, 75(14):264, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n14/pdf/scientificamerican10031896-
264.pdf.

Anonymous:1896:MC

[7134] Anonymous. In Mexican catacombs. Scientific American, 75(14):264, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-264a.pdf.

Anonymous:1896:FBS

[7135] Anonymous. A frictionless band saw guide. Scientific American, 75(14):264–265, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n14/pdf/scientificamerican10031896-
264b.pdf.

Anonymous:1896:NHG

[7136] Anonymous. The new hippopotamus of the Garden of Plants, of Paris. Scientific American, 75(14):265, October 3, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n14/pdf/scientificamerican10031896-265.pdf.

Anonymous:1896:JPL

[7137] Anonymous. A jet propelled lifeboat. Scientific American, 75(14):268, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-268.pdf.

Sloane:1896:DBN

[7138] T. O’conor Sloane. The disinfecting boat of the New York State Board of Health. Scientific American, 75(14):268–270, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-268a.pdf.

Anonymous:1896:NDBb

[7139] Anonymous. A new dirigible balloon. Scientific American, 75 (14):271–272, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n14/pdf/scientificamerican10031896-
271.pdf.

Anonymous:1896:RPIbm

[7140] Anonymous. Recently patented inventions. Scientific American, 75(14):272, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n14/pdf/scientificamerican10031896-
272.pdf.

Anonymous:1896:BNBz

[7141] Anonymous. Books: New books and publications. Scientific American, 75(14):273, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n14/pdf/scientificamerican10031896-
273.pdf.

Anonymous:1896:BPbn

[7142] Anonymous. Business and personal. Scientific American, 75(14):273, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-273a.pdf.

Anonymous:1896:Ibm

[7143] Anonymous. To inventors. Scientific American, 75(14):273, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-273c.pdf.

Anonymous:1896:NQbn

[7144] Anonymous. Notes and queries. Scientific American, 75(14):273, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-273b.pdf.

Anonymous:1896:IIbn

[7145] Anonymous. Index of inventions. Scientific American, 75(14):273–274, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-273d.pdf.

Anonymous:1896:PNPbj

[7146] Anonymous. Patent notes: Patents. Scientific American, 75(14):275, October 3, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n14/pdf/scientificamerican10031896-275.pdf.

Anonymous:1896:LBN

[7147] Anonymous. The lofty buildings of New York City. Scientific American, 75(15):277, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
277.pdf.

Anonymous:1896:LM

[7148] Anonymous. Labor and machinery. Scientific American, 75(15): 278–279, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
278.pdf.

Frauckhauser:1896:LSW

[7149] J. B. Frauckhauser. The Louisiana sulphur wells. Scientific American, 75(15):280–281, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
280.pdf.

Anonymous:1896:HAT

[7150] Anonymous. The houts automatic telephone system. Scientific American, 75(15):281–282, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
281.pdf.

Anonymous:1896:IHa

[7151] Anonymous. An improved harrow. Scientific American, 75(15):282, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n15/pdf/scientificamerican10101896-282.pdf.

Anonymous:1896:Cbc

[7152] Anonymous. Correspondence. Scientific American, 75(15):283, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n15/pdf/scientificamerican10101896-283.pdf.

Anonymous:1896:VA

[7153] Anonymous. A velocipede ambulance. Scientific American, 75 (15):284, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
284.pdf.

Anonymous:1896:ASb

[7154] Anonymous. An adjustable square. Scientific American, 75(15):287, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n15/pdf/scientificamerican10101896-287.pdf.

Anonymous:1896:NBBc

[7155] Anonymous. A new ball bearing inspection car. Scientific American, 75(15):287, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
287a.pdf.

Anonymous:1896:BNBba

[7156] Anonymous. Books: New books and publications. Scientific American, 75(15):288, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
288a.pdf.

Anonymous:1896:BPbo

[7157] Anonymous. Business and personal. Scientific American, 75 (15):288, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
288b.pdf.

Anonymous:1896:RPIbn

[7158] Anonymous. Recently patented inventions. Scientific American, 75(15):288, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
288.pdf.

Anonymous:1896:NQbo

[7159] Anonymous. Notes and queries. Scientific American, 75(15): 288–289, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n15/pdf/scientificamerican10101896-
288c.pdf.

Anonymous:1896:Ibn

[7160] Anonymous. To inventors. Scientific American, 75(15):289, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n15/pdf/scientificamerican10101896-289.pdf.

Anonymous:1896:IIbo

[7161] Anonymous. Index of inventions. Scientific American, 75(15):289, October 10, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n15/pdf/scientificamerican10101896-289a.pdf.

Anonymous:1896:EPS

[7162] Anonymous. European practice in steam boilers. Scientific American, 75(16):294–295, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
294.pdf.

Anonymous:1896:CPE

[7163] Anonymous. Camera for producing enlarged images of microscopic objects. Scientific American, 75(16):296, October 17, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n16/pdf/scientificamerican10171896-296a.pdf.

Anonymous:1896:TAa

[7164] Anonymous. The toy artist. Scientific American, 75(16):296, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n16/pdf/scientificamerican10171896-296.pdf.

Anonymous:1896:LHS

[7165] Anonymous. A locomotive headlight and signal. Scientific American, 75(16):297, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
297a.pdf.

Anonymous:1896:VJT

[7166] Anonymous. A vise for jewelers, tool makers, etc. Scientific American, 75(16):297, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
297.pdf.

Sewall:1896:BW

[7167] E. D. Sewall. The bicycle wheel. Scientific American, 75(16): 297–298, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
297b.pdf.

Anonymous:1896:NCDb

[7168] Anonymous. The new Croton Dam. Scientific American, 75 (16):299, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
299.pdf.

Anonymous:1896:PED

[7169] Anonymous. Photography by an electric discharge. Scientific American, 75(16):299, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
299a.pdf.

Hovey:1896:EMB

[7170] Horace C. Hovey. The Essex–Merrimac Bridge. Scientific American, 75(16):300–301, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
300.pdf.

Anonymous:1896:DBC

[7171] Anonymous. The Dion, Bouton & Company tricycle and the Bollee carriage. Scientific American, 75(16):301–302, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n16/pdf/scientificamerican10171896-301.pdf.

Anonymous:1896:PMH

[7172] Anonymous. The Poulson metal house. Scientific American, 75 (16):302–303, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
302.pdf.

Anonymous:1896:BPbp

[7173] Anonymous. Business and personal. Scientific American, 75 (16):304, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
304a.pdf.

Anonymous:1896:RPIbo

[7174] Anonymous. Recently patented inventions. Scientific American, 75(16):304, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
304.pdf.

Anonymous:1896:NQbp

[7175] Anonymous. Notes and queries. Scientific American, 75(16): 304–305, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
304b.pdf.

Anonymous:1896:Ibo

[7176] Anonymous. To inventors. Scientific American, 75(16):305, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n16/pdf/scientificamerican10171896-305.pdf.

Anonymous:1896:IIbp

[7177] Anonymous. Index of inventions. Scientific American, 75(16):305, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n16/pdf/scientificamerican10171896-305a.pdf.

Anonymous:1896:PNPbk

[7178] Anonymous. Patent notes: Patents. Scientific American, 75 (16):307, October 17, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n16/pdf/scientificamerican10171896-
307.pdf.

Anonymous:1896:AER

[7179] Anonymous. American and English railroad trace. Scientific American, 75(17):310, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
310.pdf.

Anonymous:1896:OTS

[7180] Anonymous. Our trade with the South American republics. Scientific American, 75(17):310–311, October 24, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n17/pdf/scientificamerican10241896-310a.pdf.

Anonymous:1896:CPM

[7181] Anonymous. A combined pipe and monkey wrench. Scientific American, 75(17):312, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
312a.pdf.

Anonymous:1896:LUF

[7182] Anonymous. Lefebvre and Upton’s force and lift pump. Scientific American, 75(17):312, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
312b.pdf.

Anonymous:1896:NDW

[7183] Anonymous. A new dumb waiter safety clutch. Scientific American, 75(17):312, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
312.pdf.

Anonymous:1896:A

[7184] Anonymous. The aerophile. Scientific American, 75(17):313, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n17/pdf/scientificamerican10241896-313.pdf.

Anonymous:1896:ICD

[7185] Anonymous. An improved coffee drier. Scientific American, 75 (17):313–314, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
313a.pdf.

Anonymous:1896:Cbd

[7186] Anonymous. Correspondence. Scientific American, 75(17):315, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n17/pdf/scientificamerican10241896-315.pdf.

Anonymous:1896:TFT

[7187] Anonymous. A twenty-five ton geared logging locomotive. Scientific American, 75(17):316, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
316.pdf.

Anonymous:1896:SCP

[7188] Anonymous. The sesquicentennial celebration of Princeton University. Scientific American, 75(17):317–318, October 24, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n17/pdf/scientificamerican10241896-317.pdf.

Anonymous:1896:EDM

[7189] Anonymous. Experimental determination of the motion of projectiles inside the bore of a gun. Scientific American, 75(17):318–319, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n17/pdf/scientificamerican10241896-318.pdf.

Anonymous:1896:BNBbb

[7190] Anonymous. Books: New books and publications. Scientific American, 75(17):320, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
320b.pdf.

Anonymous:1896:BPbq

[7191] Anonymous. Business and personal. Scientific American, 75 (17):320, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
320a.pdf.

Anonymous:1896:RPIbp

[7192] Anonymous. Recently patented inventions. Scientific American, 75(17):320, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
320.pdf.

Anonymous:1896:IIbq

[7193] Anonymous. Index of inventions. Scientific American, 75(17):321, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n17/pdf/scientificamerican10241896-321a.pdf.

Anonymous:1896:NQbq

[7194] Anonymous. Notes and queries. Scientific American, 75(17):321, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n17/pdf/scientificamerican10241896-321.pdf.

Anonymous:1896:PNPbl

[7195] Anonymous. Patent notes: Patents. Scientific American, 75 (17):323, October 24, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n17/pdf/scientificamerican10241896-
323.pdf.

Anonymous:1896:KS

[7196] Anonymous. The kinetoscope stereopticon. Scientific American, 75(18):325, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
325.pdf.

Anonymous:1896:SSA

[7197] Anonymous. The serious side of the American character. Scientific American, 75(18):326–327, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
326.pdf.

Anonymous:1896:AAL

[7198] Anonymous. An automatic axle lubricator. Scientific American, 75(18):328, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
328.pdf.

Anonymous:1896:NPF

[7199] Anonymous. A novel post fastening. Scientific American, 75 (18):328, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
328b.pdf.

Anonymous:1896:NRBb

[7200] Anonymous. A non refillable bottle. Scientific American, 75 (18):328, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
328a.pdf.

Anonymous:1896:GM

[7201] Anonymous. The Grant Monument. Scientific American, 75(18): 329–330, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
329.pdf.

Anonymous:1896:BFIb

[7202] Anonymous. Bullets fused by impact. Scientific American, 75 (18):330, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
330.pdf.

Anonymous:1896:OI

[7203] Anonymous. An optical illusion. Scientific American, 75(18):330, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n18/pdf/scientificamerican10311896-330a.pdf.

Anonymous:1896:Cbe

[7204] Anonymous. Correspondence. Scientific American, 75(18):331, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n18/pdf/scientificamerican10311896-331.pdf.

Anonymous:1896:MI

[7205] Anonymous. The milk industry. Scientific American, 75(18):332, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n18/pdf/scientificamerican10311896-332.pdf.

Anonymous:1896:REC

[7206] Anonymous. The ruins of Euyuk, Cappadocia. Scientific American, 75(18):333–334, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
333.pdf.

Anonymous:1896:NUT

[7207] Anonymous. Novel use of the telephone. Scientific American, 75(18):334, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
334.pdf.

Anonymous:1896:DHC

[7208] Anonymous. A Daimler horseless carriage. Scientific American, 75(18):335, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
335.pdf.

Anonymous:1896:GIC

[7209] Anonymous. A gasoline inspection car. Scientific American, 75 (18):335, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
335a.pdf.

Anonymous:1896:BNBbc

[7210] Anonymous. Books: New books and publications. Scientific American, 75(18):336, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
336a.pdf.

Anonymous:1896:BPbr

[7211] Anonymous. Business and personal. Scientific American, 75 (18):336, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
336b.pdf.

Anonymous:1896:RPIbq

[7212] Anonymous. Recently patented inventions. Scientific American, 75(18):336, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
336.pdf.

Anonymous:1896:NQbr

[7213] Anonymous. Notes and queries. Scientific American, 75(18): 336–337, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
336c.pdf.

Anonymous:1896:Ibq

[7214] Anonymous. To inventors. Scientific American, 75(18):337, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n18/pdf/scientificamerican10311896-337.pdf.

Anonymous:1896:IIbr

[7215] Anonymous. Index of inventions. Scientific American, 75(18): 337–338, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
337a.pdf.

Anonymous:1896:PNPbm

[7216] Anonymous. Patent notes: Patents. Scientific American, 75 (18):339, October 31, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n18/pdf/scientificamerican10311896-
339.pdf.

Tanabe:1896:LBK

[7217] Sakuro Tanabe. The Lake Biwa–Kioto Canal, Japan. Scientific American, 75(19):341, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
341.pdf.

Anonymous:1896:ICO

[7218] Anonymous. International co-operation in the advancement of science. Scientific American, 75(19):342–344, November 7, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n19/pdf/scientificamerican11071896-342.pdf.

Anonymous:1896:ALM

[7219] Anonymous. An automatic let-off mechanism for looms. Scientific American, 75(19):344, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
344a.pdf.

Anonymous:1896:CPB

[7220] Anonymous. Clay pipe bins for wine cellars. Scientific American, 75(19):344, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
344.pdf.

Anonymous:1896:ISIb

[7221] Anonymous. An improvement in stringed instruments. Scientific American, 75(19):346, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
346.pdf.

Anonymous:1896:Cbf

[7222] Anonymous. Correspondence. Scientific American, 75(19):347, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n19/pdf/scientificamerican11071896-347.pdf.

Anonymous:1896:ENH

[7223] Anonymous. An explosive nut (Hura Crepitans). Scientific American, 75(19):348, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
348.pdf.

Anonymous:1896:PMM

[7224] Anonymous. The Plantin–Moretus Museum. Scientific American, 75 (19):348–350, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
348a.pdf.

Benjamin:1896:CP

[7225] Wiliam Benjamin. Cassadaga propaganda. Scientific American, 75(19):351, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
351.pdf.

Anonymous:1896:BPbs

[7226] Anonymous. Business and personal. Scientific American, 75 (19):352, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
352a.pdf.

Anonymous:1896:Ibr

[7227] Anonymous. To inventors. Scientific American, 75(19):352, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n19/pdf/scientificamerican11071896-352c.pdf.

Anonymous:1896:NQbs

[7228] Anonymous. Notes and queries. Scientific American, 75(19): 352, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
352b.pdf.

Anonymous:1896:RPIbr

[7229] Anonymous. Recently patented inventions. Scientific American, 75(19):352, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
352.pdf.

Anonymous:1896:IIbs

[7230] Anonymous. Index of inventions. Scientific American, 75(19): 352–353, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
352d.pdf.

Anonymous:1896:PNPbn

[7231] Anonymous. Patent notes: Patents. Scientific American, 75 (19):354, November 7, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n19/pdf/scientificamerican11071896-
354.pdf.

Anonymous:1896:GSC

[7232] Anonymous. The great scaffolds of the Congressional Library, Washington, D.C. Scientific American, 75(20):357, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n20/pdf/scientificamerican11141896-357.pdf.

Anonymous:1896:ICBa

[7233] Anonymous. The increased confidence in business circles. Scientific American, 75(20):358, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
358.pdf.

Anonymous:1896:RCP

[7234] Anonymous. The revoking of the cyanide patents. Scientific American, 75(20):358, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
358a.pdf.

Anonymous:1896:ACU

[7235] Anonymous. Another comparison of United States and British railroads. Scientific American, 75(20):358–359, November 14, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n20/pdf/scientificamerican11141896-358b.pdf.

Anonymous:1896:IPV

[7236] Anonymous. An improved pump valve. Scientific American, 75 (20):360, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
360.pdf.

Anonymous:1896:IWB

[7237] Anonymous. An improved wood bending machine. Scientific American, 75(20):360, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
360a.pdf.

Anonymous:1896:GTb

[7238] Anonymous. A giant tricycle. Scientific American, 75(20):361, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n20/pdf/scientificamerican11141896-361a.pdf.

Anonymous:1896:IER

[7239] Anonymous. An improved elevated railway structure. Scientific American, 75(20):361, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
361.pdf.

Anonymous:1896:WPT

[7240] Anonymous. A windmill power transmitting gear. Scientific American, 75(20):361–362, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
361b.pdf.

Anonymous:1896:Cbg

[7241] Anonymous. Correspondence. Scientific American, 75(20):363, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n20/pdf/scientificamerican11141896-363.pdf.

Anonymous:1896:BL

[7242] Anonymous. A balanced locomotive. Scientific American, 75(20): 365–366, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
365.pdf.

Avery:1896:PIR

[7243] Harold Avery. Papua, its inhabitants and resources. Scientific American, 75(20):367, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
367.pdf.

Anonymous:1896:BNBbd

[7244] Anonymous. Books: New books and publications. Scientific American, 75(20):368, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
368a.pdf.

Anonymous:1896:RPIbs

[7245] Anonymous. Recently patented inventions. Scientific American, 75(20):368, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
368.pdf.

Anonymous:1896:BPbt

[7246] Anonymous. Business and personal. Scientific American, 75 (20):369, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
369.pdf.

Anonymous:1896:Ibs

[7247] Anonymous. To inventors. Scientific American, 75(20):369, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n20/pdf/scientificamerican11141896-369b.pdf.

Anonymous:1896:NQbt

[7248] Anonymous. Notes and queries. Scientific American, 75(20): 369, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
369a.pdf.

Anonymous:1896:IIbt

[7249] Anonymous. Index of inventions. Scientific American, 75(20): 369–370, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
369c.pdf.

Anonymous:1896:PNPbo

[7250] Anonymous. Patent notes: Patents. Scientific American, 75 (20):372, November 14, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n20/pdf/scientificamerican11141896-
372.pdf.

Anonymous:1896:CDM

[7251] Anonymous. The coast defense monitor amphitrite. Scientific American, 75(21):373, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
373.pdf.

Anonymous:1896:CVD

[7252] Anonymous. The close of the Venezuelan Difficulty. Scientific American, 75(21):374, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
374a.pdf.

Anonymous:1896:SBT

[7253] Anonymous. The sinking of the battleship Texas. Scientific American, 75(21):374, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
374.pdf.

Anonymous:1896:BTN

[7254] Anonymous. Bapid transit in New York City. Scientific American, 75(21):375, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
375.pdf.

Anonymous:1896:IWL

[7255] Anonymous. Improved White lead grinding machines. Scientific American, 75(21):376, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
376.pdf.

Anonymous:1896:EOR

[7256] Anonymous. An electrically operated railway switch. Scientific American, 75(21):377, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
377.pdf.

Anonymous:1896:MBc

[7257] Anonymous. The musical bottles. Scientific American, 75(21): 377, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
377a.pdf.

Anonymous:1896:IWMb

[7258] Anonymous. An improved washing machine. Scientific American, 75 (21):377–378, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
377b.pdf.

Anonymous:1896:MIAb

[7259] Anonymous. A musical instrument attachment. Scientific American, 75(21):378–379, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
378.pdf.

Anonymous:1896:EOM

[7260] Anonymous. An English oil motor carriage. Scientific American, 75(21):380, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
380a.pdf.

Anonymous:1896:OHC

[7261] Anonymous. The Olds horseless carriage. Scientific American, 75 (21):380, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
380.pdf.

Anonymous:1896:DP

[7262] Anonymous. The decapitated princess. Scientific American, 75 (21):383, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
383.pdf.

Anonymous:1896:BNBbe

[7263] Anonymous. Books: New books and publications. Scientific American, 75(21):384, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
384c.pdf.

Anonymous:1896:BPbu

[7264] Anonymous. Business and personal. Scientific American, 75 (21):384, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
384a.pdf.

Anonymous:1896:Ibu

[7265] Anonymous. To inventors. Scientific American, 75(21):384, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n21/pdf/scientificamerican11211896-384d.pdf.

Anonymous:1896:NQbu

[7266] Anonymous. Notes and queries. Scientific American, 75(21): 384, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
384b.pdf.

Anonymous:1896:RPIbt

[7267] Anonymous. Recently patented inventions. Scientific American, 75(21):384, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
384.pdf.

Anonymous:1896:IIbu

[7268] Anonymous. Index of inventions. Scientific American, 75(21): 384–385, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
384e.pdf.

Anonymous:1896:PNPbp

[7269] Anonymous. Patent notes: Patents. Scientific American, 75 (21):387, November 21, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n21/pdf/scientificamerican11211896-
387.pdf.

Anonymous:1896:CLB

[7270] Anonymous. Counterweighted lift bridge on the erie railroad. Scientific American, 75(22):389–390, November 28, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n22/pdf/scientificamerican11281896-389.pdf.

Anonymous:1896:LSAb

[7271] Anonymous. The lottery system as applied to patent practice. Scientific American, 75(22):390–391, November 28, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n22/pdf/scientificamerican11281896-390.pdf.

Anonymous:1896:DFC

[7272] Anonymous. Delay in furnishing copies of patents. Scientific American, 75(22):391, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
391a.pdf.

Anonymous:1896:WGC

[7273] Anonymous. Wire gun construction in the United States. Scientific American, 75(22):391, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
391.pdf.

Anonymous:1896:MLW

[7274] Anonymous. Measuring liquids by weight. Scientific American, 75 (22):392, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
392a.pdf.

Anonymous:1896:TO

[7275] Anonymous. A telescope without an observatory. Scientific American, 75(22):392, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
392.pdf.

Anonymous:1896:MPc

[7276] Anonymous. Multiplex phonograph. Scientific American, 75(22): 393, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
393.pdf.

Scheible:1896:PI

[7277] Albert Scheible. Patents as investments. Scientific American, 75 (22):394–395, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
394.pdf.

Anonymous:1896:NTI

[7278] Anonymous. The new ten inch wire wound, segmental gun for the United States Army. Scientific American, 75(22):396–397, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n22/pdf/scientificamerican11281896-396.pdf.

Anonymous:1896:TTR

[7279] Anonymous. The tomb of Theodoric at Ravenna. Scientific American, 75(22):397–398, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
397.pdf.

Anonymous:1896:Cbh

[7280] Anonymous. Correspondence. Scientific American, 75(22):398, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n22/pdf/scientificamerican11281896-398.pdf.

Holder:1896:CS

[7281] C. F. Holder. A companion of the sunfish. Scientific American, 75(22):399, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
399.pdf.

Anonymous:1896:BNBbf

[7282] Anonymous. Books: New books and publications. Scientific American, 75(22):400, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
400b.pdf.

Anonymous:1896:BPbv

[7283] Anonymous. Business and personal. Scientific American, 75 (22):400, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
400a.pdf.

Anonymous:1896:RPIbu

[7284] Anonymous. Recently patented inventions. Scientific American, 75(22):400, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
400.pdf.

Anonymous:1896:Ibv

[7285] Anonymous. To inventors. Scientific American, 75(22):401, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n22/pdf/scientificamerican11281896-401a.pdf.

Anonymous:1896:IIbv

[7286] Anonymous. Index of inventions. Scientific American, 75(22): 401, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
401b.pdf.

Anonymous:1896:NQbv

[7287] Anonymous. Notes and queries. Scientific American, 75(22): 401, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
401.pdf.

Anonymous:1896:PNPbq

[7288] Anonymous. Patent notes: Patents. Scientific American, 75 (22):404, November 28, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n22/pdf/scientificamerican11281896-
404.pdf.

Wilbur:1896:CLO

[7289] Earl Morse Wilbur. Crater lake, Oregon. Scientific American, 75(23):405, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
405.pdf.

Anonymous:1896:GLP

[7290] Anonymous. Good locomotive performance with heavy loads. Scientific American, 75(23):406, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
406a.pdf.

Anonymous:1896:NEA

[7291] Anonymous. A notable event in American shipbuilding. Scientific American, 75(23):406, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
406.pdf.

Brooks:1896:HD

[7292] William R. Brooks. The heavens for December. Scientific American, 75(23):407, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
407.pdf.

Anonymous:1896:ALC

[7293] Anonymous. The automatic lunch counter. Scientific American, 75(23):408, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
408a.pdf.

Anonymous:1896:CAG

[7294] Anonymous. The condensation of ammonia gas. Scientific American, 75(23):408, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
408.pdf.

Anonymous:1896:RMC

[7295] Anonymous. A railroad mission car. Scientific American, 75(23): 409–410, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
409.pdf.

Anonymous:1896:EA

[7296] Anonymous. An efficient anchor. Scientific American, 75(23):410, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n23/pdf/scientificamerican12051896-410.pdf.

Anonymous:1896:IHE

[7297] Anonymous. An improved hydraulic engine. Scientific American, 75 (23):410–411, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
410a.pdf.

Anonymous:1896:VSA

[7298] Anonymous. The victory of shot over armor. Scientific American, 75(23):412, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
412.pdf.

Anonymous:1896:SIC

[7299] Anonymous. Some inexpensive country houses. Scientific American, 75(23):413, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
413a.pdf.

Anonymous:1896:THS

[7300] Anonymous. A two-headed snake. Scientific American, 75(23):413, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n23/pdf/scientificamerican12051896-413.pdf.

Anonymous:1896:MJ

[7301] Anonymous. A model journal. Scientific American, 75(23):414, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n23/pdf/scientificamerican12051896-414a.pdf.

Anonymous:1896:NXRb

[7302] Anonymous. A new X ray tube. Scientific American, 75(23):414, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n23/pdf/scientificamerican12051896-414.pdf.

Anonymous:1896:SMS

[7303] Anonymous. Siphon made of a straw and a peach pit. Scientific American, 75(23):414, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
414b.pdf.

Anonymous:1896:BNBbg

[7304] Anonymous. Books: New books and publications. Scientific American, 75(23):415, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
415a.pdf.

Anonymous:1896:BPbw 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
415b.pdf.

Anonymous:1896:RPIbv

[7306] Anonymous. Recently patented inventions. Scientific American, 75(23):415, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
415.pdf.

Anonymous:1896:NQbw

[7307] Anonymous. Notes and queries. Scientific American, 75(23): 415–416, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n23/pdf/scientificamerican12051896-
415c.pdf.

Anonymous:1896:Ibw

[7308] Anonymous. To inventors. Scientific American, 75(23):416, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n23/pdf/scientificamerican12051896-416.pdf.

Anonymous:1896:IIbw

[7309] Anonymous. Index of inventions. Scientific American, 75(23):416, December 5, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n23/pdf/scientificamerican12051896-416a.pdf.

Anonymous:1896:APS

[7310] Anonymous. Accountings in patent suits. Scientific American, 75 (24):422, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
422.pdf.

Anonymous:1896:PPH

[7311] Anonymous. Profitable philanthropy in the housing of the working classes. Scientific American, 75(24):422–423, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n24/pdf/scientificamerican12121896-422a.pdf.

Anonymous:1896:MCE

[7312] Anonymous. The motor car in England. Scientific American, 75 (24):423, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
423.pdf.

Anonymous:1896:CEH

[7313] Anonymous. A cheap and efficient hinge. Scientific American, 75 (24):424, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
424b.pdf.

Anonymous:1896:IWc

[7314] Anonymous. An improved wrench. Scientific American, 75(24): 424, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
424.pdf.

Anonymous:1896:RWC

[7315] Anonymous. A renaissance wood carving. Scientific American, 75 (24):424, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
424a.pdf.

Anonymous:1896:BMC

[7316] Anonymous. Benzine motor cycle. Scientific American, 75(24): 425–426, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
425.pdf.

Baxter:1896:MAS

[7317] William Baxter. Modern applications of the storage battery. Scientific American, 75(24):426–427, December 12, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n24/pdf/scientificamerican12121896-426.pdf.

Anonymous:1896:TBM

[7318] Anonymous. Turret of the battleship Massachusetts under fire. Scientific American, 75(24):428–429, December 12, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n24/pdf/scientificamerican12121896-428.pdf.

Hale:1896:PS

[7319] W. H. Hale. A pioneer of science. Scientific American, 75 (24):430, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
430.pdf.

Phelps:1896:QF

[7320] Stowe Phelps. “The Queen of Flowers”. Scientific American, 75 (24):431, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
431.pdf.

Anonymous:1896:BNBbh

[7321] Anonymous. Books: New books and publications. Scientific American, 75(24):432, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
432a.pdf.

Anonymous:1896:RPIbw

[7322] Anonymous. Recently patented inventions. Scientific American, 75(24):432, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
432.pdf.

Anonymous:1896:BPbx

[7323] Anonymous. Business and personal. Scientific American, 75 (24):433, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
433.pdf.

Anonymous:1896:Ibx

[7324] Anonymous. To inventors. Scientific American, 75(24):433, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n24/pdf/scientificamerican12121896-433b.pdf.

Anonymous:1896:IIbx

[7325] Anonymous. Index of inventions. Scientific American, 75(24): 433, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
433c.pdf.

Anonymous:1896:NQbx

[7326] Anonymous. Notes and queries. Scientific American, 75(24): 433, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
433a.pdf.

Anonymous:1896:PNPbr

[7327] Anonymous. Patent notes: Patents. Scientific American, 75 (24):434, December 12, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n24/pdf/scientificamerican12121896-
434.pdf.

Anonymous:1896:PMc

[7328] Anonymous. The President’s message. Scientific American, 75 (25):438, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
438.pdf.

Anonymous:1896:UTT

[7329] Anonymous. The underground trolley and the third rail in electric traction. Scientific American, 75(25):438–439, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n25/pdf/scientificamerican12191896-438a.pdf.

Anonymous:1896:ATB

[7330] Anonymous. Armored torpedo boats. Scientific American, 75 (25):439, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
439.pdf.

Anonymous:1896:PDC

[7331] Anonymous. The perfected Duryea carriage. Scientific American, 75(25):440, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
440.pdf.

Anonymous:1896:IBBa

[7332] Anonymous. Improved ball bearing for bicycle hubs. Scientific American, 75(25):440–441, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
440a.pdf.

Anonymous:1896:AMN

[7333] Anonymous. The American Museum of Natural History — its collections and work — the mounting of the elephant “Tip”. Scientific American, 75(25):441–442, December 19, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n25/pdf/scientificamerican12191896-441.pdf.

Anonymous:1896:Cbi

[7334] Anonymous. Correspondence. Scientific American, 75(25):443, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n25/pdf/scientificamerican12191896-443.pdf.

Anonymous:1896:LSN

[7335] Anonymous. Logging in the Sierra Nevada Mountains, California. Scientific American, 75(25):444–446, December 19, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n25/pdf/scientificamerican12191896-444.pdf.

Anonymous:1896:ET

[7336] Anonymous. The Eiffel tandem. Scientific American, 75(25): 446–447, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
446.pdf.

Anonymous:1896:NC

[7337] Anonymous. A novel camera. Scientific American, 75(25):447, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n25/pdf/scientificamerican12191896-447.pdf.

Anonymous:1896:RPIbx

[7338] Anonymous. Recently patented inventions. Scientific American, 75(25):448, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
448.pdf.

Anonymous:1896:BPby

[7339] Anonymous. Business and personal. Scientific American, 75 (25):449, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
449.pdf.

Anonymous:1896:IIby

[7340] Anonymous. Index of inventions. Scientific American, 75(25): 449, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
449b.pdf.

Anonymous:1896:NQby

[7341] Anonymous. Notes and queries. Scientific American, 75(25): 449, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
449a.pdf.

Anonymous:1896:PNPbs

[7342] Anonymous. Patent notes: Patents. Scientific American, 75 (25):452, December 19, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n25/pdf/scientificamerican12191896-
452.pdf.

Anonymous:1896:SHRb

[7343] Anonymous. Secretary Herbert’s report on the Texas. Scientific American, 75(26):454, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
454.pdf.

Anonymous:1896:PRI

[7344] Anonymous. The prevention of rust in iron and steel structures. Scientific American, 75(26):454–455, December 26, 1896. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v75/
n26/pdf/scientificamerican12261896-454a.pdf.

Anonymous:1896:HRS

[7345] Anonymous. The Hudson River steamer Adirondack, of the People’s Line. Scientific American, 75(26):456–457, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n26/pdf/scientificamerican12261896-456.pdf.

Anonymous:1896:AFc

[7346] Anonymous. Armor for fortifications. Scientific American, 75(26): 457–458, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
457.pdf.

Hopkins:1896:HFE

[7347] George M. Hopkins. Hand feed electric lamp for lanterns. Scientific American, 75(26):458–459, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
458.pdf.

Willoughby:1896:ACC

[7348] C. C. Willoughby. The ancient city of Copan. Scientific American, 75(26):460–461, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
460.pdf.

Caulk:1896:AL

[7349] William B. Caulk. The appearing lady. Scientific American, 75 (26):461, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
461.pdf.

Anonymous:1896:BNBbi

[7350] Anonymous. Books: New books and publications. Scientific American, 75(26):462, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
462a.pdf.

Anonymous:1896:BPbz

[7351] Anonymous. Business and personal. Scientific American, 75 (26):462, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
462b.pdf.

Anonymous:1896:RPIby

[7352] Anonymous. Recently patented inventions. Scientific American, 75(26):462, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
462.pdf.

Anonymous:1896:NQbz

[7353] Anonymous. Notes and queries. Scientific American, 75(26): 462–463, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
462c.pdf.

Anonymous:1896:Iby

[7354] Anonymous. To inventors. Scientific American, 75(26):463, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v75/n26/pdf/scientificamerican12261896-463.pdf.

Anonymous:1896:IIbz

[7355] Anonymous. Index of inventions. Scientific American, 75(26): 463–464, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
463a.pdf.

Anonymous:1896:PNPbt

[7356] Anonymous. Patent notes: Patents. Scientific American, 75 (26):464, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
464.pdf.

Anonymous:1896:IIca

[7357] Anonymous. Index: Indexes. Scientific American, 75(26):465– 466, December 26, 1896. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v75/n26/pdf/scientificamerican12261896-
465.pdf.

Anonymous:1897:SSB

[7358] Anonymous. The signaling system of the Broadway and Seventh Avenue Railroad. Scientific American, 76(1):3–4, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n1/pdf/scientificamerican01021897-3.pdf.

Anonymous:1897:WTB

[7359] Anonymous. The water tube boiler on its trial. Scientific American, 76(1):4, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n1/pdf/scientificamerican01021897-
4.pdf.

Anonymous:1897:RY

[7360] Anonymous. A retrospect of the year 1896. Scientific American, 76(1):4–5, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n1/pdf/scientificamerican01021897-
4a.pdf.

Anonymous:1897:LIE

[7361] Anonymous. The latest improvements in electric elevators. Scientific American, 76(1):7–8, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n1/pdf/scientificamerican01021897-
7.pdf.

Anonymous:1897:IPW

[7362] Anonymous. An improved pipe wrench. Scientific American, 76(1):8, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n1/pdf/scientificamerican01021897-8.pdf.

Anonymous:1897:NNR

[7363] Anonymous. A novel nautical registering apparatus. Scientific American, 76(1):8–10, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n1/pdf/scientificamerican01021897-
8a.pdf.

Anonymous:1897:BGB

[7364] Anonymous. Blown glass bricks for building purposes. Scientific American, 76(1):10–11, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n1/pdf/scientificamerican01021897-
10.pdf.

Anonymous:1897:PP

[7365] Anonymous. The philosophy of plants. Scientific American, 76(1):11, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n1/pdf/scientificamerican01021897-11.pdf.

Anonymous:1897:CAR

[7366] Anonymous. Compressed air recoil cylinders for heavy mortars. Scientific American, 76(1):11–12, January 2, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n1/pdf/scientificamerican01021897-11a.pdf.

Anonymous:1897:BPa

[7367] Anonymous. Business and personal. Scientific American, 76(1):12, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n1/pdf/scientificamerican01021897-12a.pdf.

Anonymous:1897:NQa

[7368] Anonymous. Notes and queries. Scientific American, 76(1):12, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n1/pdf/scientificamerican01021897-12b.pdf.

Anonymous:1897:RPIa

[7369] Anonymous. Recently patented inventions. Scientific American, 76(1):12, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n1/pdf/scientificamerican01021897-
12.pdf.

Anonymous:1897:BNBa

[7370] Anonymous. Books: New books and publications. Scientific American, 76(1):13, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n1/pdf/scientificamerican01021897-
13.pdf.

Anonymous:1897:IIa

[7371] Anonymous. Index of inventions. Scientific American, 76(1):13, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n1/pdf/scientificamerican01021897-13a.pdf.

Anonymous:1897:PNPa

[7372] Anonymous. Patent notes: Patents. Scientific American, 76(1):14, January 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n1/pdf/scientificamerican01021897-14.pdf.

Anonymous:1897:RTSa

[7373] Anonymous. A railway through the sea. Scientific American, 76(2):17, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-17.pdf.

Anonymous:1897:DPMa

[7374] Anonymous. Dishonest patent methods detected. Scientific American, 76(2):18, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
18.pdf.

Anonymous:1897:ITI

[7375] Anonymous. Inventors and their inventions. Scientific American, 76(2):18–19, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
18a.pdf.

Anonymous:1897:GSA

[7376] Anonymous. The Geological Society of America. Scientific American, 76(2):19, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
19.pdf.

Anonymous:1897:AFH

[7377] Anonymous. An automatic feed hand drill. Scientific American, 76(2):20, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
20.pdf.

Anonymous:1897:IWa

[7378] Anonymous. An improved wrench. Scientific American, 76(2):20, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-20b.pdf.

Anonymous:1897:TEB

[7379] Anonymous. A traction engine boiler explosion. Scientific American, 76(2):20, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
20a.pdf.

Anonymous:1897:IP

[7380] Anonymous. An improved Pitman. Scientific American, 76(2):21, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-21a.pdf.

Anonymous:1897:WWR

[7381] Anonymous. Water wheel regulation. Scientific American, 76(2):21, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-21.pdf.

Anonymous:1897:Ca

[7382] Anonymous. Correspondence. Scientific American, 76(2):23–24, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-23.pdf.

Prindle:1897:OSB

[7383] E. J. Prindle. The operation of the self-binding harvester. Scientific American, 76(2):24–25, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
24.pdf.

Anonymous:1897:PCTa

[7384] Anonymous. The pulpit of the Cathedral of Treves. Scientific American, 76(2):25–26, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
25.pdf.

Anonymous:1897:APA

[7385] Anonymous. Absolutely pure alcohol from acetylene gas. Scientific American, 76(2):26, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
26.pdf.

Anonymous:1897:HEA

[7386] Anonymous. Hoisting engines for the Anaconda Mine. Scientific American, 76(2):27, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
27.pdf.

Anonymous:1897:BNBb

[7387] Anonymous. Books: New books and publications. Scientific American, 76(2):28, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
28a.pdf.

Anonymous:1897:RPIb

[7388] Anonymous. Recently patented inventions. Scientific American, 76(2):28, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n2/pdf/scientificamerican01091897-
28.pdf.

Anonymous:1897:BPb

[7389] Anonymous. Business and personal. Scientific American, 76(2):29, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-29.pdf.

Anonymous:1897:Ia

[7390] Anonymous. To inventors. Scientific American, 76(2):29, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-29b.pdf.

Anonymous:1897:IIb

[7391] Anonymous. Index of inventions. Scientific American, 76(2):29, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-29c.pdf.

Anonymous:1897:NQb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-29a.pdf.

Anonymous:1897:PNPb

[7393] Anonymous. Patent notes: Patents. Scientific American, 76(2):30, January 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n2/pdf/scientificamerican01091897-30.pdf.

Anonymous:1897:CST

[7394] Anonymous. Competitive systems of traction in New York City. Scientific American, 76(3):34, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
34.pdf.

Anonymous:1897:WRJ

[7395] Anonymous. Welded rail joints. Scientific American, 76(3):34–35, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n3/pdf/scientificamerican01161897-34a.pdf.

Anonymous:1897:TCE

[7396] Anonymous. The Tennessee Centennial Exposition. Scientific American, 76(3):36–37, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
36.pdf.

Anonymous:1897:TMW

[7397] Anonymous. A trolley man of war. Scientific American, 76 (3):37–39, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
37.pdf.

Anonymous:1897:MPM

[7398] Anonymous. Manufacture of papier maché busts. Scientific American, 76(3):40, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
40.pdf.

Anonymous:1897:RIE

[7399] Anonymous. The Reno inclined elevator. Scientific American, 76(3):41, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
41.pdf.

Anonymous:1897:NOS

[7400] Anonymous. A new oil separator. Scientific American, 76(3): 41–42, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
41a.pdf.

Anonymous:1897:CSB

[7401] Anonymous. Changes of speed for bicycles. Scientific American, 76(3):43, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
43.pdf.

Anonymous:1897:BNBc

[7402] Anonymous. Books: New books and publications. Scientific American, 76(3):44, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
44a.pdf.

Anonymous:1897:BPc

[7403] Anonymous. Business and personal. Scientific American, 76(3):44, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n3/pdf/scientificamerican01161897-44b.pdf.

Anonymous:1897:NQc

[7404] Anonymous. Notes and queries. Scientific American, 76(3):44, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n3/pdf/scientificamerican01161897-44c.pdf.

Anonymous:1897:RPIc 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n3/pdf/scientificamerican01161897-
44.pdf.

Anonymous:1897:Ib

[7406] Anonymous. To inventors. Scientific American, 76(3):45, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n3/pdf/scientificamerican01161897-45.pdf.

Anonymous:1897:IIc

[7407] Anonymous. Index of inventions. Scientific American, 76(3):45, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n3/pdf/scientificamerican01161897-45a.pdf.

Anonymous:1897:PNPc

[7408] Anonymous. Patent notes: Patents. Scientific American, 76(3):46, January 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n3/pdf/scientificamerican01161897-46.pdf.

Anonymous:1897:AE

[7409] Anonymous. Antarctic exploration. Scientific American, 76(4): 50–51, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
50.pdf.

Anonymous:1897:CHR

[7410] Anonymous. A convenient household refrigerator. Scientific American, 76(4):52, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
52a.pdf.

Anonymous:1897:NEC

[7411] Anonymous. A new electric current regulator. Scientific American, 76(4):52, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
52.pdf.

Mindeleff:1897:RIW

[7412] Cosmos Mindeleff. Recent irrigation work on the Navaho reservation. Scientific American, 76(4):52–55, January 23, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n4/pdf/scientificamerican01231897-52b.pdf.

Anonymous:1897:TTA

[7413] Anonymous. A theater with two auditoriums. Scientific American, 76(4):55, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
55.pdf.

Holder:1897:SPC

[7414] C. F. Holder. Some Pacific caves. Scientific American, 76(4):56, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n4/pdf/scientificamerican01231897-56.pdf.

Anonymous:1897:WR

[7415] Anonymous. The Walhalla of Ratisbon. Scientific American, 76 (4):57–58, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
57.pdf.

Anonymous:1897:NMP

[7416] Anonymous. A novel method of propelling vessels. Scientific American, 76(4):58, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
58.pdf.

Anonymous:1897:FMK

[7417] Anonymous. A folding Malay kite. Scientific American, 76(4):59, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n4/pdf/scientificamerican01231897-59.pdf.

Anonymous:1897:LLW

[7418] Anonymous. Logging locomotive for wooden track. Scientific American, 76(4):59–60, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
59a.pdf.

Anonymous:1897:RPId

[7419] Anonymous. Recently patented inventions. Scientific American, 76(4):60, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
60.pdf.

Anonymous:1897:BNBd

[7420] Anonymous. Books: New books and publications. Scientific American, 76(4):61, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n4/pdf/scientificamerican01231897-
61b.pdf.

Anonymous:1897:BPd

[7421] Anonymous. Business and personal. Scientific American, 76(4):61, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n4/pdf/scientificamerican01231897-61.pdf.

Anonymous:1897:Ic

[7422] Anonymous. To inventors. Scientific American, 76(4):61, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n4/pdf/scientificamerican01231897-61c.pdf.

Anonymous:1897:IId

[7423] Anonymous. Index of inventions. Scientific American, 76(4):61, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n4/pdf/scientificamerican01231897-61d.pdf.

Anonymous:1897:NQd

[7424] Anonymous. Notes and queries. Scientific American, 76(4):61, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n4/pdf/scientificamerican01231897-61a.pdf.

Anonymous:1897:PNPd

[7425] Anonymous. Patent notes: Patents. Scientific American, 76(4):62, January 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n4/pdf/scientificamerican01231897-62.pdf.

Anonymous:1897:TOT

[7426] Anonymous. A trio of one thousand horse power boilers. Scientific American, 76(5):65–66, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
65.pdf.

Anonymous:1897:NDS

[7427] Anonymous. A National Department of Science. Scientific American, 76(5):66, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
66a.pdf.

Anonymous:1897:PSAa

[7428] Anonymous. Present status of the apprenticeship system. Scientific American, 76(5):66, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
66.pdf.

Brooks:1897:HF

[7429] William R. Brooks. The heavens for February. Scientific American, 76(5):66–69, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
66b.pdf.

Anonymous:1897:BFR

[7430] Anonymous. A bicycle frame reinforce. Scientific American, 76 (5):69, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
69.pdf.

Anonymous:1897:NW

[7431] Anonymous. A novel wrench. Scientific American, 76(5):69, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n5/pdf/scientificamerican01301897-69a.pdf.

Anonymous:1897:MMS

[7432] Anonymous. Method of measuring the speed of camera shutters. Scientific American, 76(5):69–71, January 30, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n5/pdf/scientificamerican01301897-69b.pdf.

Anonymous:1897:VRG

[7433] Anonymous. A vehicle running gear attachment. Scientific American, 76(5):71–72, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
71.pdf.

Anonymous:1897:CAM

[7434] Anonymous. Compressed air motor on the elevated railroads, New York. Scientific American, 76(5):72–73, January 30, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n5/pdf/scientificamerican01301897-72.pdf.

Anonymous:1897:CFF

[7435] Anonymous. Capture of a fifty foot whale in Puget Sound, Washington. Scientific American, 76(5):73–74, January 30, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n5/pdf/scientificamerican01301897-73.pdf.

Anonymous:1897:BRE

[7436] Anonymous. The Brambel rotary engine. Scientific American, 76 (5):74–75, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
74.pdf.

Anonymous:1897:THa

[7437] Anonymous. The training of horses. Scientific American, 76 (5):75–76, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
75.pdf.

Anonymous:1897:RPIe

[7438] Anonymous. Recently patented inventions. Scientific American, 76(5):76, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
76.pdf.

Anonymous:1897:BNBe

[7439] Anonymous. Books: New books and publications. Scientific American, 76(5):76–77, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
76a.pdf.

Anonymous:1897:BPe

[7440] Anonymous. Business and personal. Scientific American, 76(5):77, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n5/pdf/scientificamerican01301897-77.pdf.

Anonymous:1897:Id

[7441] Anonymous. To inventors. Scientific American, 76(5):77, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n5/pdf/scientificamerican01301897-77b.pdf.

Anonymous:1897:NQe

[7442] Anonymous. Notes and queries. Scientific American, 76(5):77, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n5/pdf/scientificamerican01301897-77a.pdf.

Anonymous:1897:IIe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n5/pdf/scientificamerican01301897-77c.pdf.

Anonymous:1897:PNPe

[7444] Anonymous. Patent notes: Patents. Scientific American, 76 (5):78–80, January 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n5/pdf/scientificamerican01301897-
78.pdf.

Anonymous:1897:HRSa

[7445] Anonymous. The Harlem River Speedway. Scientific American, 76(6):82, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
82.pdf.

Anonymous:1897:PPLa

[7446] Anonymous. Proposed patent legislation. Scientific American, 76(6):82, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
82a.pdf.

Anonymous:1897:GEN

[7447] Anonymous. The gas exhibition in New York City. Scientific American, 76(6):82–83, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
82b.pdf.

Anonymous:1897:STL

[7448] Anonymous. The steam trials of the largest cruiser in the world. Scientific American, 76(6):83, February 6, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n6/pdf/scientificamerican02061897-83.pdf.

Anonymous:1897:AWS

[7449] Anonymous. All wrought steel pulleys. Scientific American, 76 (6):84, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
84.pdf.

Anonymous:1897:SNT

[7450] Anonymous. Setting new tires and resetting loose tires Cold without taking out the tire bolts. Scientific American, 76(6):84, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n6/pdf/scientificamerican02061897-84a.pdf.

Anonymous:1897:ELW

[7451] Anonymous. An extension ladder and window jack. Scientific American, 76(6):84–85, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
84b.pdf.

Anonymous:1897:HFL

[7452] Anonymous. The Horton fire ladder apparatus. Scientific American, 76(6):85, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
85.pdf.

Anonymous:1897:NPTa

[7453] Anonymous. A novel pneumatic tire. Scientific American, 76 (6):85–87, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
85a.pdf.

Anonymous:1897:MRS

[7454] Anonymous. The Military Riding School at Tor Di Quinto. Scientific American, 76(6):88, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
88.pdf.

Anonymous:1897:HRSb

[7455] Anonymous. The Harlem River Speedway, New York City. Scientific American, 76(6):89–90, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
89.pdf.

Anonymous:1897:TLP

[7456] Anonymous. The tomb of Louis Pasteur. Scientific American, 76(6):91, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
91.pdf.

Anonymous:1897:RPIf

[7457] Anonymous. Recently patented inventions. Scientific American, 76(6):92, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
92.pdf.

Anonymous:1897:BPf

[7458] Anonymous. Business and personal. Scientific American, 76 (6):93, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
93.pdf.

Anonymous:1897:Ie

[7459] Anonymous. To inventors. Scientific American, 76(6):93, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n6/pdf/scientificamerican02061897-93b.pdf.

Anonymous:1897:IIf

[7460] Anonymous. Index of inventions. Scientific American, 76(6):93, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n6/pdf/scientificamerican02061897-93c.pdf.

Anonymous:1897:NQf

[7461] Anonymous. Notes and queries. Scientific American, 76(6):93, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n6/pdf/scientificamerican02061897-93a.pdf.

Anonymous:1897:PNPf

[7462] Anonymous. Patent notes: Patents. Scientific American, 76 (6):94, February 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n6/pdf/scientificamerican02061897-
94.pdf.

Anonymous:1897:GC

[7463] Anonymous. The Green cross. Scientific American, 76(7):98, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n7/pdf/scientificamerican02131897-98b.pdf.

Anonymous:1897:RNJ

[7464] Anonymous. The reclamation of the New Jersey meadows. Scientific American, 76(7):98, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
98a.pdf.

Anonymous:1897:STHa

[7465] Anonymous. Shade trees for the Harlem Speedway. Scientific American, 76(7):98, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
98.pdf.

Anonymous:1897:MVS

[7466] Anonymous. The modern version of supply and demand. Scientific American, 76(7):99, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
99d.pdf.

Anonymous:1897:NC

[7467] Anonymous. The Nicaragua Canal. Scientific American, 76(7): 99, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
99.pdf.

Anonymous:1897:ONT

[7468] Anonymous. Our new torpedo boats. Scientific American, 76 (7):99, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
99c.pdf.

Anonymous:1897:RTT

[7469] Anonymous. A rapid transit tunnel between New York and Brooklyn. Scientific American, 76(7):99, February 13, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n7/pdf/scientificamerican02131897-99a.pdf.

Anonymous:1897:SCRa

[7470] Anonymous. Schemes for checking returned newspapers. Scientific American, 76(7):99, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
99b.pdf.

Anonymous:1897:HFE

[7471] Anonymous. A horseless fire engine. Scientific American, 76 (7):100, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
100a.pdf.

Anonymous:1897:HNB

[7472] Anonymous. A handy note book holder. Scientific American, 76(7):100, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
100.pdf.

Anonymous:1897:IFC

[7473] Anonymous. An improved friction clutch system. Scientific American, 76(7):100–101, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
100b.pdf.

Anonymous:1897:PEDa

[7474] Anonymous. A portable electric drilling machine. Scientific American, 76(7):101–102, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
101.pdf.

Anonymous:1897:Cb

[7475] Anonymous. Correspondence. Scientific American, 76(7):103, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n7/pdf/scientificamerican02131897-103.pdf.

Anonymous:1897:HRSc

[7476] Anonymous. The Harlem River Speedway, New York City. Scientific American, 76(7):104–105, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
104.pdf.

Anonymous:1897:SBI

[7477] Anonymous. The stability of the battleship Indiana. Scientific American, 76(7):105–106, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
105.pdf.

Caulk:1897:SW

[7478] William B. Caulk. The sword Walker. Scientific American, 76 (7):107, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
107.pdf.

Anonymous:1897:BNBf

[7479] Anonymous. Books: New books and publications. Scientific American, 76(7):108, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
108a.pdf.

Anonymous:1897:RPIg

[7480] Anonymous. Recently patented inventions. Scientific American, 76(7):108, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
108.pdf.

Anonymous:1897:BPg

[7481] Anonymous. Business and personal. Scientific American, 76(7): 109, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
109.pdf.

Anonymous:1897:If

[7482] Anonymous. To inventors. Scientific American, 76(7):109, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n7/pdf/scientificamerican02131897-109b.pdf.

Anonymous:1897:IIg

[7483] Anonymous. Index of inventions. Scientific American, 76(7):109, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n7/pdf/scientificamerican02131897-109c.pdf.

Anonymous:1897:NQg

[7484] Anonymous. Notes and queries. Scientific American, 76(7):109, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n7/pdf/scientificamerican02131897-109a.pdf.

Anonymous:1897:PNPg

[7485] Anonymous. Patent notes: Patents. Scientific American, 76 (7):110, February 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n7/pdf/scientificamerican02131897-
110.pdf.

Anonymous:1897:CIP

[7486] Anonymous. Congress and the International Patent Convention. Scientific American, 76(8):114, February 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n8/pdf/scientificamerican02201897-114.pdf.

Anonymous:1897:ABE

[7487] Anonymous. The Annual Bicycle Exhibition at the Grand Central Palace, New York. Scientific American, 76(8):114–115, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n8/pdf/scientificamerican02201897-114a.pdf.

Anonymous:1897:AN

[7488] Anonymous. Archæological news. Scientific American, 76(8): 115, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
115d.pdf.

Anonymous:1897:GMN

[7489] Anonymous. A gold medal for Nansen. Scientific American, 76 (8):115, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
115b.pdf.

Anonymous:1897:LRP

[7490] Anonymous. The late Richard Pope. Scientific American, 76 (8):115, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
115.pdf.

Anonymous:1897:NBP

[7491] Anonymous. A new business-patent sharks. Scientific American, 76(8):115, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
115a.pdf.

Anonymous:1897:SWa

[7492] Anonymous. Strength of welds. Scientific American, 76(8):115, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n8/pdf/scientificamerican02201897-115c.pdf.

Anonymous:1897:APW

[7493] Anonymous. An adjustable pipe wrench. Scientific American, 76(8):116, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
116b.pdf.

Anonymous:1897:IV

[7494] Anonymous. An improved valve. Scientific American, 76(8):116, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n8/pdf/scientificamerican02201897-116.pdf.

Anonymous:1897:NRBa

[7495] Anonymous. A new roller bicycle brake. Scientific American, 76(8):116, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
116a.pdf.

Anonymous:1897:VT

[7496] Anonymous. Vaporization in tubes. Scientific American, 76(8): 117, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
117.pdf.

Anonymous:1897:QMW

[7497] Anonymous. Quintin Massys’ well. Scientific American, 76(8): 117–118, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
117a.pdf.

Anonymous:1897:Cc

[7498] Anonymous. Correspondence. Scientific American, 76(8):119– 120, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
119.pdf.

Anonymous:1897:TDD

[7499] Anonymous. The timber dry docks at the Brooklyn Navy Yard. Scientific American, 76(8):120–121, February 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n8/pdf/scientificamerican02201897-120.pdf.

Anonymous:1897:NHB

[7500] Anonymous. A new house bicycle. Scientific American, 76(8): 122, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
122.pdf.

Anonymous:1897:NSA

[7501] Anonymous. Notes and sketches at the annual bicycle show, New York City. Scientific American, 76(8):122–123, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n8/pdf/scientificamerican02201897-122a.pdf.

Anonymous:1897:BNBg

[7502] Anonymous. Books: New books and publications. Scientific American, 76(8):124, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
124a.pdf.

Anonymous:1897:BPh

[7503] Anonymous. Business and personal. Scientific American, 76(8): 124, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
124b.pdf.

Anonymous:1897:RPIh

[7504] Anonymous. Recently patented inventions. Scientific American, 76(8):124, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
124.pdf.

Anonymous:1897:NQh

[7505] Anonymous. Notes and queries. Scientific American, 76(8):124– 125, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n8/pdf/scientificamerican02201897-
124c.pdf.

Anonymous:1897:Ig

[7506] Anonymous. To inventors. Scientific American, 76(8):125, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n8/pdf/scientificamerican02201897-125.pdf.

Anonymous:1897:IIh

[7507] Anonymous. Index of inventions. Scientific American, 76(8):125, February 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n8/pdf/scientificamerican02201897-125a.pdf.

Anonymous:1897:EC

[7508] Anonymous. The economies of concentration. Scientific American, 76(9):130, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
130.pdf.

Anonymous:1897:LRN

[7509] Anonymous. The lessons of the recent Naval maneuvers. Scientific American, 76(9):130–131, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
130a.pdf.

Anonymous:1897:BBI

[7510] Anonymous. The Brooklyn Bridge improvements. Scientific American, 76(9):131, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
131.pdf.

Anonymous:1897:BMI

[7511] Anonymous. British mineral industry. Scientific American, 76 (9):131, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
131c.pdf.

Anonymous:1897:NYE

[7512] Anonymous. The New York Elevated’s new power. Scientific American, 76(9):131, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
131a.pdf.

Anonymous:1897:RPTa

[7513] Anonymous. Recent patent and trade mark decisions. Scientific American, 76(9):131, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
131b.pdf.

Anonymous:1897:BSA

[7514] Anonymous. A barometer and storm alarm. Scientific American, 76(9):132, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
132.pdf.

Anonymous:1897:IWb

[7515] Anonymous. An improved wrench. Scientific American, 76(9): 132, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
132b.pdf.

Anonymous:1897:PPF

[7516] Anonymous. A pipe plugger and flusher. Scientific American, 76(9):132, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
132a.pdf.

Anonymous:1897:VCF

[7517] Anonymous. A very convenient faucet. Scientific American, 76 (9):132, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
132c.pdf.

Anonymous:1897:ONY

[7518] Anonymous. Opening of the New York Central Four Track Drawbridge and Viaduct in New York City. Scientific American, 76 (9):133–135, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
133.pdf.

Anonymous:1897:PLF

[7519] Anonymous. The Pennsylvania, the largest freight steamer in the world. Scientific American, 76(9):136, February 27, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n9/pdf/scientificamerican02271897-136.pdf.

Anonymous:1897:SMM

[7520] Anonymous. A snow melting machine. Scientific American, 76 (9):137–138, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
137.pdf.

Anonymous:1897:CFJ

[7521] Anonymous. Contraction of the face in jumping. Scientific American, 76(9):138, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
138.pdf.

Anonymous:1897:EAB

[7522] Anonymous. Echoes of the Annual Bicycle Exhibition. Scientific American, 76(9):138–139, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
138a.pdf.

Anonymous:1897:BNBh

[7523] Anonymous. Books: New books and publications. Scientific American, 76(9):140, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
140a.pdf.

Anonymous:1897:BPi

[7524] Anonymous. Business and personal. Scientific American, 76(9): 140, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
140b.pdf.

Anonymous:1897:RPIi

[7525] Anonymous. Recently patented inventions. Scientific American, 76(9):140, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
140.pdf.

Anonymous:1897:NQi

[7526] Anonymous. Notes and queries. Scientific American, 76(9):140– 141, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
140c.pdf.

Anonymous:1897:Ih

[7527] Anonymous. To inventors. Scientific American, 76(9):141, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n9/pdf/scientificamerican02271897-141.pdf.

Anonymous:1897:IIi

[7528] Anonymous. Index of inventions. Scientific American, 76(9):141, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n9/pdf/scientificamerican02271897-141a.pdf.

Anonymous:1897:PNPh

[7529] Anonymous. Patent notes: Patents. Scientific American, 76 (9):142, February 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n9/pdf/scientificamerican02271897-
142.pdf.

Anonymous:1897:ARSa

[7530] Anonymous. American railway speeds. Scientific American, 76(10):146, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-146.pdf.

Anonymous:1897:DPL

[7531] Anonymous. “Defective patent laws”. Scientific American, 76 (10):146–147, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
146a.pdf.

Anonymous:1897:RLDa

[7532] Anonymous. The remarkable long distance run. Scientific American, 76(10):147, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
147a.pdf.

Anonymous:1897:SHa

[7533] Anonymous. Sagacity of horses. Scientific American, 76(10):147, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-147b.pdf.

Brooks:1897:HMa

[7534] William R. Brooks. The heavens for March. Scientific American, 76(10):147, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
147.pdf.

Anonymous:1897:BF

[7535] Anonymous. A bedclothes fastener. Scientific American, 76(10):148, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-148b.pdf.

Anonymous:1897:MSEa

[7536] Anonymous. A machine for stamping envelopes, etc. Scientific American, 76(10):148, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
148.pdf.

Anonymous:1897:NFG

[7537] Anonymous. A new field glass. Scientific American, 76(10):148, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-148a.pdf.

Anonymous:1897:TTHa

[7538] Anonymous. A two thousand horse power turbine. Scientific American, 76(10):149–151, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
149.pdf.

Baldwin:1897:GPS

[7539] E. R. Baldwin. A gift of philanthropy to science. Scientific American, 76(10):152–154, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
152.pdf.

Anonymous:1897:PIP

[7540] Anonymous. A “phonendoscope” or improved phonostethescope. Scientific American, 76(10):154, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
154.pdf.

Grant:1897:GMS

[7541] George B. Grant. Grant’s machine for solving numerical algebraic equations. Scientific American, 76(10):154–155, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-154a.pdf.

Anonymous:1897:TBN

[7542] Anonymous. Torpedo boat number 6, for the United States Navy. Scientific American, 76(10):155, March 6, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n10/pdf/scientificamerican03061897-155.pdf.

Anonymous:1897:BNBi

[7543] Anonymous. Books: New books and publications. Scientific American, 76(10):156, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
156a.pdf.

Anonymous:1897:BPj

[7544] Anonymous. Business and personal. Scientific American, 76(10):156, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-156b.pdf.

Anonymous:1897:NQj

[7545] Anonymous. Notes and queries. Scientific American, 76(10):156, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-156c.pdf.

Anonymous:1897:RPIj

[7546] Anonymous. Recently patented inventions. Scientific American, 76(10):156, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n10/pdf/scientificamerican03061897-
156.pdf.

Anonymous:1897:IIj

[7547] Anonymous. Index of inventions. Scientific American, 76(10):157, March 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n10/pdf/scientificamerican03061897-157.pdf.

Anonymous:1897:GRR

[7548] Anonymous. Good roads in relation to farm values. Scientific American, 76(11):162, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
162.pdf.

Anonymous:1897:IPL

[7549] Anonymous. An interesting point in locomotive history. Scientific American, 76(11):162–163, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
162a.pdf.

Anonymous:1897:LHT

[7550] Anonymous. Luther Henry Tucker. Scientific American, 76(11):163, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-163c.pdf.

Anonymous:1897:NPPa

[7551] Anonymous. A new process of producing photographs in colors. Scientific American, 76(11):163, March 13, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n11/pdf/scientificamerican03131897-163.pdf.

Anonymous:1897:PCTb

[7552] Anonymous. Prof. Crookes on thought transference. Scientific American, 76(11):163, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
163a.pdf.

Anonymous:1897:RPTb

[7553] Anonymous. Recent patent and trade mark decisions. Scientific American, 76(11):163, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
163b.pdf.

Anonymous:1897:BSP

[7554] Anonymous. A bicycle skirt protector. Scientific American, 76(11):164, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-164a.pdf.

Anonymous:1897:THb

[7555] Anonymous. A twin house. Scientific American, 76(11):164, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-164.pdf.

Anonymous:1897:FB

[7556] Anonymous. A folding bicycle. Scientific American, 76(11):165, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-165.pdf.

Anonymous:1897:EH

[7557] Anonymous. An electric Hansom. Scientific American, 76(11):165–166, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-165a.pdf.

Bishop:1897:CL

[7558] Edw. K. Bishop. Cascade locks. Scientific American, 76(11):167, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-167.pdf.

Anonymous:1897:IST

[7559] Anonymous. The internal slave trade of Africa. Scientific American, 76(11):168–169, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
168.pdf.

Anonymous:1897:Cd

[7560] Anonymous. Correspondence. Scientific American, 76(11):170, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-170.pdf.

Hovey:1897:AMS

[7561] Horace C. Hovey. About Merrimac ship building. Scientific American, 76(11):170–171, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
170a.pdf.

Anonymous:1897:BNBj

[7562] Anonymous. Books: New books and publications. Scientific American, 76(11):172, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
172a.pdf.

Anonymous:1897:BPk

[7563] Anonymous. Business and personal. Scientific American, 76(11):172, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-172b.pdf.

Anonymous:1897:NQk

[7564] Anonymous. Notes and queries. Scientific American, 76(11):172, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-172c.pdf.

Anonymous:1897:RPIk

[7565] Anonymous. Recently patented inventions. Scientific American, 76(11):172, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n11/pdf/scientificamerican03131897-
172.pdf.

Anonymous:1897:IIk

[7566] Anonymous. Index of inventions. Scientific American, 76(11):172–173, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-172d.pdf.

Anonymous:1897:PNPi

[7567] Anonymous. Patent notes: Patents. Scientific American, 76(11):174, March 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n11/pdf/scientificamerican03131897-174.pdf.

Anonymous:1897:CAS

[7568] Anonymous. Compressed air system on the United States monitor Terror. Scientific American, 76(12):177, March 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n12/pdf/scientificamerican03201897-177.pdf.

Anonymous:1897:IAU

[7569] Anonymous. Important amendments to the United States Patent Statutes. Scientific American, 76(12):178, March 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n12/pdf/scientificamerican03201897-178.pdf.

Anonymous:1897:PMR

[7570] Anonymous. The proposed mammoth relief map of the United States. Scientific American, 76(12):178–179, March 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n12/pdf/scientificamerican03201897-178a.pdf.

Anonymous:1897:ANP

[7571] Anonymous. Another new patent law. Scientific American, 76(12):179, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-179b.pdf.

Anonymous:1897:APS

[7572] Anonymous. Amendments to the patent statutes. Scientific American, 76(12):179, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n12/pdf/scientificamerican03201897-
179a.pdf.

Anonymous:1897:JP

[7573] Anonymous. Japanese patents. Scientific American, 76(12):179, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-179.pdf.

Anonymous:1897:PAA

[7574] Anonymous. Patent attorneys appeal. Scientific American, 76 (12):179–180, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n12/pdf/scientificamerican03201897-
179c.pdf.

Anonymous:1897:AST

[7575] Anonymous. Automatic strength tester and vending machine. Scientific American, 76(12):180, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n12/pdf/scientificamerican03201897-
180.pdf.

Anonymous:1897:BGE

[7576] Anonymous. The Boston gas explosion. Scientific American, 76 (12):180–182, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n12/pdf/scientificamerican03201897-
180a.pdf.

Anonymous:1897:Ce

[7577] Anonymous. Correspondence. Scientific American, 76(12):183, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-183.pdf.

Robinson:1897:AHP

[7578] W. E. Robinson. Alexander Herrmann prestidigitator. Scientific American, 76(12):184, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n12/pdf/scientificamerican03201897-
184.pdf.

Anonymous:1897:FSD

[7579] Anonymous. A four sided dovetail. Scientific American, 76(12):186, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-186.pdf.

Anonymous:1897:IWR

[7580] Anonymous. Insects’ wings — rapidity of vibration in flight. Scientific American, 76(12):187–188, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n12/pdf/scientificamerican03201897-
187.pdf.

Anonymous:1897:RPIl 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n12/pdf/scientificamerican03201897-
188.pdf.

Anonymous:1897:BPl

[7582] Anonymous. Business and personal. Scientific American, 76(12):189, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-189.pdf.

Anonymous:1897:IIl

[7583] Anonymous. Index of inventions. Scientific American, 76(12):189, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-189b.pdf.

Anonymous:1897:NQl

[7584] Anonymous. Notes and queries. Scientific American, 76(12):189, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-189a.pdf.

Anonymous:1897:PNPj

[7585] Anonymous. Patent notes: Patents. Scientific American, 76(12):190, March 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n12/pdf/scientificamerican03201897-190.pdf.

Anonymous:1897:EL

[7586] Anonymous. The emancipation of labor. Scientific American, 76(13):194, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
194.pdf.

Anonymous:1897:NPEa

[7587] Anonymous. New possibilities of electricity. Scientific American, 76(13):194, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
194a.pdf.

Anonymous:1897:ERH

[7588] Anonymous. English railway history in the Chicago Museum. Scientific American, 76(13):194–195, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
194b.pdf.

Anonymous:1897:JCC

[7589] Anonymous. The judgment in the Cordite Case. Scientific American, 76(13):195, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
195.pdf.

Anonymous:1897:MME

[7590] Anonymous. Magnetic metal extraction. Scientific American, 76 (13):195, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
195b.pdf.

Anonymous:1897:RBE

[7591] Anonymous. Rapid bridge erection. Scientific American, 76(13):195, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-195c.pdf.

Brooks:1897:HAa

[7592] William R. Brooks. The heavens for April. Scientific American, 76(13):195, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
195a.pdf.

Anonymous:1897:CTB

[7593] Anonymous. A cork tire for bicycles. Scientific American, 76(13):196, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-196a.pdf.

Anonymous:1897:IRP

[7594] Anonymous. An improved rotary pump. Scientific American, 76(13):196, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
196.pdf.

Anonymous:1897:NTC

[7595] Anonymous. A novel three Crank bicycle. Scientific American, 76(13):196, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
196c.pdf.

Anonymous:1897:WHF

[7596] Anonymous. The Worden hickory frame bicycle. Scientific American, 76(13):196, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
196b.pdf.

Anonymous:1897:CSS

[7597] Anonymous. The Charlton street sweeper. Scientific American, 76(13):197, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
197.pdf.

Anonymous:1897:CUO

[7598] Anonymous. A cultivator for use in orchards. Scientific American, 76(13):197–198, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
197a.pdf.

Anonymous:1897:SEF

[7599] Anonymous. A simple and efficient fire hose coupling. Scientific American, 76(13):198, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
198.pdf.

Anonymous:1897:SHF

[7600] Anonymous. Simple handle for flower pots, dishes, etc. Scientific American, 76(13):198–199, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
198a.pdf.

Anonymous:1897:LDT

[7601] Anonymous. The long distance transmission plant at Fresno, Cal. Scientific American, 76(13):200–201, March 27, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n13/pdf/scientificamerican03271897-200.pdf.

Gratacap:1897:CS

[7602] L. P. Gratacap. Collectors of snakes. Scientific American, 76(13):201, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-201.pdf.

Anonymous:1897:QH

[7603] Anonymous. The Queen of Holland. Scientific American, 76 (13):202–203, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
202.pdf.

Anonymous:1897:RPIm

[7604] Anonymous. Recently patented inventions. Scientific American, 76 (13):203–204, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
203.pdf.

Anonymous:1897:BNBk

[7605] Anonymous. Books: New books and publications. Scientific American, 76(13):204, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n13/pdf/scientificamerican03271897-
204.pdf.

Anonymous:1897:BPm

[7606] Anonymous. Business and personal. Scientific American, 76(13):204, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-204a.pdf.

Anonymous:1897:NQm

[7607] Anonymous. Notes and queries. Scientific American, 76(13):204–205, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-204b.pdf.

Anonymous:1897:Ij

[7608] Anonymous. To inventors. Scientific American, 76(13):205, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-205.pdf.

Anonymous:1897:IIm

[7609] Anonymous. Index of inventions. Scientific American, 76(13):205, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-205a.pdf.

Anonymous:1897:PNPk

[7610] Anonymous. Patent notes: Patents. Scientific American, 76(13):206, March 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n13/pdf/scientificamerican03271897-206.pdf.

Anonymous:1897:SSA

[7611] Anonymous. Section seven of the amendments to the patent statutes. Scientific American, 76(14):210, April 3, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n14/pdf/scientificamerican04031897-210.pdf.

Anonymous:1897:UTN

[7612] Anonymous. The underground trolley in New York City. Scientific American, 76(14):210–211, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
210a.pdf.

Anonymous:1897:DCHa

[7613] Anonymous. Diphtheria in Cold and hot countries. Scientific American, 76(14):211, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
211a.pdf.

Thurston:1897:RSE

[7614] R. H. Thurston. The “Rotary” steam engine. Scientific American, 76(14):211, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
211.pdf.

Anonymous:1897:IBa

[7615] Anonymous. An improved boiler. Scientific American, 76(14):212, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n14/pdf/scientificamerican04031897-212b.pdf.

Anonymous:1897:MRM

[7616] Anonymous. The manufacture of ready mixed paints. Scientific American, 76(14):212, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
212.pdf.

Anonymous:1897:NASa

[7617] Anonymous. A new automatic screw cutting die head. Scientific American, 76(14):212, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
212a.pdf.

Anonymous:1897:SCRb

[7618] Anonymous. Scenes on the Columbia River, Oregon. Scientific American, 76(14):213, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
213.pdf.

Anonymous:1897:BSIa

[7619] Anonymous. The Bertillon system of identification by measurement. Scientific American, 76(14):214–215, April 3, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n14/pdf/scientificamerican04031897-214.pdf.

Anonymous:1897:NFJ

[7620] Anonymous. A new fast job printing press. Scientific American, 76(14):216, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
216.pdf.

Tarr:1897:GG

[7621] Ralph S. Tarr. The glaciers of Greenland. Scientific American, 76(14):216–218, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
216a.pdf.

Anonymous:1897:BRA

[7622] Anonymous. Batchelor’s ripping attachment for sewing machines. Scientific American, 76(14):218, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
218a.pdf.

Anonymous:1897:FEW

[7623] Anonymous. A fire escape and water tower. Scientific American, 76(14):218, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
218.pdf.

Anonymous:1897:BNBl

[7624] Anonymous. Books: New books, etc. Scientific American, 76(14):219, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n14/pdf/scientificamerican04031897-219a.pdf.

Anonymous:1897:RPIn

[7625] Anonymous. Recently patented inventions. Scientific American, 76(14):219, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n14/pdf/scientificamerican04031897-
219.pdf.

Anonymous:1897:BPn

[7626] Anonymous. Business and personal. Scientific American, 76(14):220, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n14/pdf/scientificamerican04031897-220.pdf.

Anonymous:1897:Ik

[7627] Anonymous. To inventors. Scientific American, 76(14):220, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n14/pdf/scientificamerican04031897-220b.pdf.

Anonymous:1897:NQn

[7628] Anonymous. Notes and queries. Scientific American, 76(14):220, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n14/pdf/scientificamerican04031897-220a.pdf.

Anonymous:1897:IIn

[7629] Anonymous. Index of inventions. Scientific American, 76(14):220–221, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n14/pdf/scientificamerican04031897-220c.pdf.

Anonymous:1897:PNPl

[7630] Anonymous. Patent notes: Patents. Scientific American, 76(14):222, April 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n14/pdf/scientificamerican04031897-222.pdf.

Anonymous:1897:YO

[7631] Anonymous. The Yerkes Observatory. Scientific American, 76(15):225, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-225.pdf.

Anonymous:1897:BPIa

[7632] Anonymous. A bill to protect inventors. Scientific American, 76(15):226, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
226.pdf.

Anonymous:1897:NCP

[7633] Anonymous. The new Commissioner of Patents. Scientific American, 76(15):226, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
226a.pdf.

Anonymous:1897:RGP

[7634] Anonymous. Relation of the Government to the patents of Naval officers. Scientific American, 76(15):226, April 10, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n15/pdf/scientificamerican04101897-226b.pdf.

Anonymous:1897:ACLa

[7635] Anonymous. Amendments to the copyright law. Scientific American, 76(15):226–227, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
226c.pdf.

Anonymous:1897:GQO

[7636] Anonymous. Gems of quartz origin. Scientific American, 76(15):227, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-227.pdf.

Anonymous:1897:NDP

[7637] Anonymous. A new discovery in physics. Scientific American, 76(15):227, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
227a.pdf.

Anonymous:1897:UST

[7638] Anonymous. Usefulness of the Simplon Tunnel. Scientific American, 76(15):227, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
227b.pdf.

Anonymous:1897:ASA

[7639] Anonymous. Automatic slack adjuster for car brakes. Scientific American, 76(15):228, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
228b.pdf.

Anonymous:1897:EWM

[7640] Anonymous. An efficient woodworking machine. Scientific American, 76(15):228, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
228a.pdf.

Anonymous:1897:IHa

[7641] Anonymous. An improved die head. Scientific American, 76(15):228, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-228.pdf.

Anonymous:1897:NPF

[7642] Anonymous. Ney’s pocket field glass. Scientific American, 76(15):229, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-229.pdf.

Anonymous:1897:USF

[7643] Anonymous. The United States first-class sea-going battleship Iowa. Scientific American, 76(15):229–230, April 10, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n15/pdf/scientificamerican04101897-229a.pdf.

Anonymous:1897:CSPa

[7644] Anonymous. Chilton’s safety pin. Scientific American, 76(15):230, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-230.pdf.

Anonymous:1897:TP

[7645] Anonymous. A toy phonograph. Scientific American, 76(15):230, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-230a.pdf.

Anonymous:1897:Cf

[7646] Anonymous. Correspondence. Scientific American, 76(15):231, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-231.pdf.

Anonymous:1897:FCM

[7647] Anonymous. The Field Columbian Museum, Chicago. Scientific American, 76(15):233–235, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
233.pdf.

Anonymous:1897:VSB

[7648] Anonymous. A velocipede shower bath. Scientific American, 76(15):235, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-235.pdf.

Anonymous:1897:BNBm

[7649] Anonymous. Books: New books, etc. Scientific American, 76(15):236, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-236a.pdf.

Anonymous:1897:BPo

[7650] Anonymous. Business and personal. Scientific American, 76(15):236, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-236b.pdf.

Anonymous:1897:NQo

[7651] Anonymous. Notes and queries. Scientific American, 76(15):236, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-236c.pdf.

Anonymous:1897:RPIo 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n15/pdf/scientificamerican04101897-
236.pdf.

Anonymous:1897:IIo

[7653] Anonymous. Index of inventions. Scientific American, 76(15):236–237, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-236d.pdf.

Anonymous:1897:PNPm

[7654] Anonymous. Patent notes: Patents. Scientific American, 76(15):238, April 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n15/pdf/scientificamerican04101897-238.pdf.

Anonymous:1897:ASE

[7655] Anonymous. Abandonment of a social experiment. Scientific American, 76(16):242, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
242b.pdf.

Anonymous:1897:BC

[7656] Anonymous. Business changes. Scientific American, 76(16):242, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-242d.pdf.

Anonymous:1897:BRS

[7657] Anonymous. Bazin’s roller steamboat. Scientific American, 76(16):242, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-242e.pdf.

Anonymous:1897:ENP

[7658] Anonymous. England’s Naval Policy. Scientific American, 76(16):242, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-242a.pdf.

Anonymous:1897:GSE

[7659] Anonymous. The Government’s scientific exhibits at the Tennessee Exposition. Scientific American, 76(16):242, April 17, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n16/pdf/scientificamerican04171897-242c.pdf.

Anonymous:1897:TTBa

[7660] Anonymous. Trial trip of the battleship Iowa. Scientific American, 76(16):242, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
242.pdf.

Anonymous:1897:GPC

[7661] Anonymous. The gutta-percha crop of 1896. Scientific American, 76(16):243, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
243e.pdf.

Anonymous:1897:HAL

[7662] Anonymous. History of the American locomotive. Scientific American, 76(16):243, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
243f.pdf.

Anonymous:1897:HBB

[7663] Anonymous. The Hon Benjamin Butterworth. Scientific American, 76(16):243, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
243.pdf.

Anonymous:1897:NBAa

[7664] Anonymous. No bids for armor plate. Scientific American, 76(16):243, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-243a.pdf.

Anonymous:1897:PDS

[7665] Anonymous. Practical directions for soldering aluminum. Scientific American, 76(16):243, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
243d.pdf.

Anonymous:1897:PRG

[7666] Anonymous. Production of rain by great fires. Scientific American, 76(16):243, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
243c.pdf.

Anonymous:1897:PS

[7667] Anonymous. The population of Switzerland. Scientific American, 76(16):243, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
243b.pdf.

Anonymous:1897:DTB

[7668] Anonymous. The Densmore typewriter ball bearing type bar joints. Scientific American, 76(16):244, April 17, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n16/pdf/scientificamerican04171897-244b.pdf.

Anonymous:1897:SSH

[7669] Anonymous. A spring-supported handle bar. Scientific American, 76(16):244, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
244.pdf.

Anonymous:1897:STG

[7670] Anonymous. A semi-traction gasoline engine. Scientific American, 76(16):244, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
244a.pdf.

Anonymous:1897:NAM

[7671] Anonymous. A new automatic milling machine. Scientific American, 76(16):245, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
245.pdf.

Anonymous:1897:NYA

[7672] Anonymous. New York Academy of Sciences. Scientific American, 76(16):246–247, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
246.pdf.

Anonymous:1897:AMP

[7673] Anonymous. The art of moving photography. Scientific American, 76(16):248–250, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
248.pdf.

Anonymous:1897:EHF

[7674] Anonymous. The Essick hot fluid battery. Scientific American, 76(16):250, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
250.pdf.

Anonymous:1897:BNBn

[7675] Anonymous. Books: New books, etc. Scientific American, 76(16):251, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-251a.pdf.

Anonymous:1897:BPp

[7676] Anonymous. Business and personal. Scientific American, 76(16):251, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-251b.pdf.

Anonymous:1897:NQp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-251c.pdf.

Anonymous:1897:RPIp

[7678] Anonymous. Recently patented inventions. Scientific American, 76(16):251, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n16/pdf/scientificamerican04171897-
251.pdf.

Anonymous:1897:IIp

[7679] Anonymous. Index of inventions. Scientific American, 76(16):251–252, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-251d.pdf.

Anonymous:1897:PNPn

[7680] Anonymous. Patent notes: Patents. Scientific American, 76(16):254, April 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n16/pdf/scientificamerican04171897-254.pdf.

Anonymous:1897:CWS

[7681] Anonymous. Coaling warships at sea. Scientific American, 76 (17):257–258, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
257.pdf.

Anonymous:1897:RNC

[7682] Anonymous. Responsibilities of the new Commissioner. Scientific American, 76(17):258, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
258.pdf.

Anonymous:1897:SMR

[7683] Anonymous. The Six Months Rule in court. Scientific American, 76(17):258–259, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
258a.pdf.

Anonymous:1897:AIS

[7684] Anonymous. American industrial supremacy. Scientific American, 76(17):259, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
259.pdf.

Anonymous:1897:CPG

[7685] Anonymous. A circulating picture gallery. Scientific American, 76(17):259, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
259a.pdf.

Anonymous:1897:DME

[7686] Anonymous. Danube–Moldau–Elbe Canal. Scientific American, 76(17):259, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
259b.pdf.

Anonymous:1897:WIa

[7687] Anonymous. A warning to inventors. Scientific American, 76(17):259, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-259c.pdf.

Anonymous:1897:MPL

[7688] Anonymous. The most powerful locomotives in the world. Scientific American, 76(17):260, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
260.pdf.

Anonymous:1897:ISIa

[7689] Anonymous. An improved speed indicator. Scientific American, 76(17):261, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
261.pdf.

Anonymous:1897:NBSa

[7690] Anonymous. A new blind slatting machine. Scientific American, 76(17):261, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
261a.pdf.

Anonymous:1897:TTHb

[7691] Anonymous. Test of a three hundred horse power steam turbine. Scientific American, 76(17):261–263, April 24, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n17/pdf/scientificamerican04241897-261b.pdf.

Dunlap:1897:NSA

[7692] Orrin E. Dunlap. New Steel Arch Bridge over the Niagara. Scientific American, 76(17):264, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
264.pdf.

Anonymous:1897:CCT

[7693] Anonymous. Crossing the Cataract of Tequendama on a rope. Scientific American, 76(17):265, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
265.pdf.

Anonymous:1897:TOG

[7694] Anonymous. A trick opera glass. Scientific American, 76(17):265, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-265a.pdf.

Anonymous:1897:DPC

[7695] Anonymous. Death of Prof. Cope. Scientific American, 76(17):266, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-266.pdf.

Anonymous:1897:MSa

[7696] Anonymous. Mountain shadows. Scientific American, 76(17):267, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-267.pdf.

Anonymous:1897:BPq

[7697] Anonymous. Business and personal. Scientific American, 76(17):268, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-268a.pdf.

Anonymous:1897:NQq

[7698] Anonymous. Notes and queries. Scientific American, 76(17):268, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-268b.pdf.

Anonymous:1897:RPIq

[7699] Anonymous. Recently patented inventions. Scientific American, 76(17):268, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n17/pdf/scientificamerican04241897-
268.pdf.

Anonymous:1897:BNBo

[7700] Anonymous. Books: New books, etc. Scientific American, 76(17):269, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-269.pdf.

Anonymous:1897:Il

[7701] Anonymous. To inventors. Scientific American, 76(17):269, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-269a.pdf.

Anonymous:1897:IIq

[7702] Anonymous. Index of inventions. Scientific American, 76(17):269, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-269b.pdf.

Anonymous:1897:PNPo

[7703] Anonymous. Patent notes: Patents. Scientific American, 76(17):270, April 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n17/pdf/scientificamerican04241897-270.pdf.

Anonymous:1897:LQJ

[7704] Anonymous. The labor question in Japan. Scientific American, 76(18):274, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
274a.pdf.

Anonymous:1897:ONA

[7705] Anonymous. Our new acquisition. Scientific American, 76(18):274, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-274.pdf.

Hale:1897:AAA

[7706] William H. Hale. The American Association for the Advancement of Science. Scientific American, 76(18):274–275, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-274b.pdf.

Anonymous:1897:BM

[7707] Anonymous. Bats and music. Scientific American, 76(18):275, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-275d.pdf.

Anonymous:1897:CHF

[7708] Anonymous. Cleaning hardwood floors. Scientific American, 76 (18):275, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
275b.pdf.

Anonymous:1897:HCB

[7709] Anonymous. Horseless carriages at the Brussels Exposition. Scientific American, 76(18):275, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
275c.pdf.

Brooks:1897:HMb

[7710] William R. Brooks. The heavens for May. Scientific American, 76(18):275, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
275.pdf.

Chandler:1897:VL

[7711] S. C. Chandler. On the variation of latitude. Scientific American, 76(18):275, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
275a.pdf.

Anonymous:1897:NELa

[7712] Anonymous. A new electric lamp. Scientific American, 76(18):276, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-276a.pdf.

Anonymous:1897:NIR

[7713] Anonymous. The “New Idea” rain water filter. Scientific American, 76(18):276, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
276b.pdf.

Anonymous:1897:WSC

[7714] Anonymous. A window shade casing. Scientific American, 76(18):276, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-276.pdf.

Dunlap:1897:EHG

[7715] Orrin E. Dunlap. The electric heating of the generating station of the Niagara Falls Power Company. Scientific American, 76(18):277–278, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-277.pdf.

Anonymous:1897:TSA

[7716] Anonymous. A thill support and anti-rattler. Scientific American, 76(18):278, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
278.pdf.

Anonymous:1897:GMT

[7717] Anonymous. The Grant Memorial Tomb, New York City. Scientific American, 76(18):279, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
279.pdf.

Anonymous:1897:TSD

[7718] Anonymous. Trial of side and deck armor for the United States battleships Kearsarge and Kentucky. Scientific American, 76(18):280, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-280.pdf.

Anonymous:1897:HSA

[7719] Anonymous. Heavy shipment of Army cannon. Scientific American, 76(18):280–281, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
280a.pdf.

Anonymous:1897:NCC

[7720] Anonymous. A novel chronophotographic camera. Scientific American, 76(18):281–282, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
281.pdf.

Anonymous:1897:SSC

[7721] Anonymous. Slow speed cable car grip for use on curves. Scientific American, 76(18):282, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
282.pdf.

Anonymous:1897:ISEa

[7722] Anonymous. An improved steam engine. Scientific American, 76(18):283, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
283.pdf.

Anonymous:1897:RPIr

[7723] Anonymous. Recently patented inventions. Scientific American, 76(18):283, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n18/pdf/scientificamerican05011897-
283a.pdf.

Anonymous:1897:BNBp

[7724] Anonymous. Books: New books, etc. Scientific American, 76(18):284, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-284b.pdf.

Anonymous:1897:BPr

[7725] Anonymous. Business and personal. Scientific American, 76(18):284, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-284.pdf.

Anonymous:1897:Im

[7726] Anonymous. To inventors. Scientific American, 76(18):284, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-284c.pdf.

Anonymous:1897:IIr

[7727] Anonymous. Index of inventions. Scientific American, 76(18):284, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-284d.pdf.

Anonymous:1897:NQr

[7728] Anonymous. Notes and queries. Scientific American, 76(18):284, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-284a.pdf.

Anonymous:1897:PNPp

[7729] Anonymous. Patent notes: Patents. Scientific American, 76(18):286, May 1, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n18/pdf/scientificamerican05011897-286.pdf.

Anonymous:1897:FDS

[7730] Anonymous. Free distribution of seeds by the Government. Scientific American, 76(19):290, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
290a.pdf.

Anonymous:1897:FVA

[7731] Anonymous. The fastest vessel afloat. Scientific American, 76(19):290, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-290b.pdf.

Anonymous:1897:GAI

[7732] Anonymous. A graceful act of international courtesy. Scientific American, 76(19):290, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
290.pdf.

Anonymous:1897:WMT

[7733] Anonymous. War measures in time of peace. Scientific American, 76(19):290, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
290c.pdf.

Anonymous:1897:HST

[7734] Anonymous. High speed telegraphy. Scientific American, 76 (19):290–291, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
290d.pdf.

Hovey:1897:OSC

[7735] Horace C. Hovey. Our saltpeter caves in time of war. Scientific American, 76(19):291, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
291.pdf.

Anonymous:1897:ECC

[7736] Anonymous. An efficient chimney cowl. Scientific American, 76(19):292, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-292a.pdf.

Anonymous:1897:HBP

[7737] Anonymous. The Hall brass pipe wrench. Scientific American, 76(19):292, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
292b.pdf.

Anonymous:1897:ISC

[7738] Anonymous. An improved smoke consuming furnace. Scientific American, 76(19):292, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
292.pdf.

Anonymous:1897:GWC

[7739] Anonymous. Good work in chimney moving. Scientific American, 76(19):293, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
293a.pdf.

Anonymous:1897:NHDa

[7740] Anonymous. A nine hundred dollar bicycle bridge. Scientific American, 76(19):293, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
293.pdf.

Anonymous:1897:NYL

[7741] Anonymous. The New York and long Island bridge, New York City. Scientific American, 76(19):294–295, May 8, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n19/pdf/scientificamerican05081897-294.pdf.

Anonymous:1897:WE

[7742] Anonymous. The War in the East. Scientific American, 76(19):297–299, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-297.pdf.

Anonymous:1897:LM

[7743] Anonymous. The log of the Mayflower. Scientific American, 76(19):299, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-299.pdf.

Anonymous:1897:BNBq

[7744] Anonymous. Books: New books, etc. Scientific American, 76(19):300, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-300a.pdf.

Anonymous:1897:RPIs

[7745] Anonymous. Recently patented inventions. Scientific American, 76(19):300, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n19/pdf/scientificamerican05081897-
300.pdf.

Anonymous:1897:BPs

[7746] Anonymous. Business and personal. Scientific American, 76(19):301, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-301.pdf.

Anonymous:1897:IIs

[7747] Anonymous. Index of inventions. Scientific American, 76(19):301, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-301b.pdf.

Anonymous:1897:NQs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-301a.pdf.

Anonymous:1897:PNPq

[7749] Anonymous. Patent notes: Patents. Scientific American, 76(19):302, May 8, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n19/pdf/scientificamerican05081897-302.pdf.

Anonymous:1897:BIN

[7750] Anonymous. British interest in the Nicaragua Canal. Scientific American, 76(20):306, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
306b.pdf.

Anonymous:1897:MFa

[7751] Anonymous. The Mississippi floods. Scientific American, 76(20):306, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-306c.pdf.

Anonymous:1897:MWT

[7752] Anonymous. The merits of the water tube boiler. Scientific American, 76(20):306, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
306d.pdf.

Anonymous:1897:STS

[7753] Anonymous. Shade trees for the streets of New York. Scientific American, 76(20):306, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
306a.pdf.

Anonymous:1897:USPa

[7754] Anonymous. The United States and the Paris Exposition. Scientific American, 76(20):306, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
306.pdf.

Anonymous:1897:ALE

[7755] Anonymous. The American locomotive export trade. Scientific American, 76(20):306–307, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
306e.pdf.

Anonymous:1897:ATC

[7756] Anonymous. Aeronautics at the Tennessee Centennial Exhibition. Scientific American, 76(20):307, May 15, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n20/pdf/scientificamerican05151897-307d.pdf.

Anonymous:1897:LB

[7757] Anonymous. The longest bridges. Scientific American, 76(20):307, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-307e.pdf.

Anonymous:1897:NTSa

[7758] Anonymous. The new twin screw passenger ships of the North German Lloyd Company. Scientific American, 76(20):307, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-307a.pdf.

Anonymous:1897:PCa

[7759] Anonymous. The postal Congress. Scientific American, 76(20):307, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-307c.pdf.

Anonymous:1897:RMW

[7760] Anonymous. The return of Mr. De Windt. Scientific American, 76(20):307, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
307b.pdf.

Anonymous:1897:TCB

[7761] Anonymous. The tragedy at the Charity Bazar, Paris. Scientific American, 76(20):307, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
307.pdf.

Anonymous:1897:PTJ

[7762] Anonymous. A printers’ type justifying machine. Scientific American, 76(20):308, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
308.pdf.

Anonymous:1897:KB

[7763] Anonymous. The Kensington bicycle. Scientific American, 76 (20):308–309, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
308a.pdf.

Anonymous:1897:TWA

[7764] Anonymous. Twelve wheeled American locomotive for the Brazil Railway. Scientific American, 76(20):309, May 15, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n20/pdf/scientificamerican05151897-309.pdf.

Anonymous:1897:MCE

[7765] Anonymous. A miniature caloric engine. Scientific American, 76 (20):309–311, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
309a.pdf.

Anonymous:1897:Cg

[7766] Anonymous. Correspondence. Scientific American, 76(20):311, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-311a.pdf.

Sage:1897:SMO

[7767] C. Edward Sage. A simple method for obtaining a low temperature. Scientific American, 76(20):311, May 15, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n20/pdf/scientificamerican05151897-311.pdf.

Anonymous:1897:TBD

[7768] Anonymous. Trinity’s bronze doors. Scientific American, 76(20):312, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-312.pdf.

Anonymous:1897:WSN

[7769] Anonymous. The water supply of New York City. Scientific American, 76(20):314–315, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
314.pdf.

Anonymous:1897:TBF

[7770] Anonymous. A toy bird that flies. Scientific American, 76(20):315, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-315.pdf.

Anonymous:1897:RPIt

[7771] Anonymous. Recently patented inventions. Scientific American, 76(20):315–316, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n20/pdf/scientificamerican05151897-
315a.pdf.

Anonymous:1897:BPt

[7772] Anonymous. Business and personal. Scientific American, 76(20):316, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-316a.pdf.

Anonymous:1897:IIt

[7773] Anonymous. Index of inventions. Scientific American, 76(20):316, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-316b.pdf.

Anonymous:1897:NQt

[7774] Anonymous. Notes and queries. Scientific American, 76(20):316, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-316.pdf.

Anonymous:1897:PNPr

[7775] Anonymous. Patent notes: Patents. Scientific American, 76(20):318, May 15, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n20/pdf/scientificamerican05151897-318.pdf.

Anonymous:1897:BPLa

[7776] Anonymous. Bishop Potter on labor and machinery. Scientific American, 76(21):322, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
322a.pdf.

Anonymous:1897:USPb

[7777] Anonymous. “United States Patent Laws”. Scientific American, 76(21):322, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
322.pdf.

Anonymous:1897:IAG

[7778] Anonymous. The inventor of the automatic gun. Scientific American, 76(21):322–323, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
322b.pdf.

Anonymous:1897:AEE

[7779] Anonymous. An advance in the electrical equipment of steam railroads. Scientific American, 76(21):323, May 22, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n21/pdf/scientificamerican05221897-323a.pdf.

Anonymous:1897:BTT

[7780] Anonymous. The Berliner telephone transmitter patent sustained. Scientific American, 76(21):323, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
323.pdf.

Anonymous:1897:CWD

[7781] Anonymous. A convenient window desk. Scientific American, 76 (21):324, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
324a.pdf.

Anonymous:1897:ELH

[7782] Anonymous. An efficient latch for heavy doors. Scientific American, 76(21):324, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
324.pdf.

Anonymous:1897:SBH

[7783] Anonymous. A simple bicycle holder. Scientific American, 76 (21):324–325, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
324b.pdf.

Anonymous:1897:NML

[7784] Anonymous. A novel music leaf turner. Scientific American, 76(21):325, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-325a.pdf.

Anonymous:1897:ST

[7785] Anonymous. A shorthand typewriter. Scientific American, 76(21):325, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-325.pdf.

Anonymous:1897:LND

[7786] Anonymous. The leak in the new Dry Dock, Brooklyn Navy Yard. Scientific American, 76(21):326–327, May 22, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n21/pdf/scientificamerican05221897-326.pdf.

Anonymous:1897:TBP

[7787] Anonymous. The torpedo boat Porter. Scientific American, 76(21):328, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-328.pdf.

Bailey:1897:HOP

[7788] S. I. Bailey. Harvard Observatory in Peru. Scientific American, 76(21):329–331, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
329.pdf.

Anonymous:1897:ALO

[7789] Anonymous. An arc lamp that operates in any position. Scientific American, 76(21):331, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
331a.pdf.

Anonymous:1897:CMC

[7790] Anonymous. The Columbia motor carriage. Scientific American, 76(21):331, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
331.pdf.

Anonymous:1897:BNBr

[7791] Anonymous. Books: New books, etc. Scientific American, 76(21):332, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-332c.pdf.

Anonymous:1897:BPu

[7792] Anonymous. Business and personal. Scientific American, 76(21):332, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-332b.pdf.

Anonymous:1897:NQu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-332a.pdf.

Anonymous:1897:RPIu

[7794] Anonymous. Recently patented inventions. Scientific American, 76(21):332, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n21/pdf/scientificamerican05221897-
332.pdf.

Anonymous:1897:IIu

[7795] Anonymous. Index of inventions. Scientific American, 76(21):332–333, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-332d.pdf.

Anonymous:1897:PNPs

[7796] Anonymous. Patent notes: Patents. Scientific American, 76(21):334, May 22, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n21/pdf/scientificamerican05221897-334.pdf.

Anonymous:1897:ISIb

[7797] Anonymous. The iron and steel institute on American competition. Scientific American, 76(22):338, May 29, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n22/pdf/scientificamerican05291897-338d.pdf.

Anonymous:1897:ITF

[7798] Anonymous. Improved transportation facilities on the dock front, New York City. Scientific American, 76(22):338, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n22/pdf/scientificamerican05291897-338c.pdf.

Anonymous:1897:LHSa

[7799] Anonymous. The launching of the Holland submarine boat. Scientific American, 76(22):338, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
338b.pdf.

Anonymous:1897:NLP

[7800] Anonymous. A new locomotive for the Purdue University testing laboratory. Scientific American, 76(22):338, May 29, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n22/pdf/scientificamerican05291897-338a.pdf.

Anonymous:1897:PSH

[7801] Anonymous. The planting spaces on the Harlem Speedway. Scientific American, 76(22):338, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
338.pdf.

Anonymous:1897:MFF

[7802] Anonymous. The Mississippi floods as fertilizers of the soil. Scientific American, 76(22):338–339, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
338e.pdf.

Anonymous:1897:AIE

[7803] Anonymous. American institute of electrical engineers. Scientific American, 76(22):339, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
339b.pdf.

Anonymous:1897:JMT

[7804] Anonymous. John M. Thatcher. Scientific American, 76(22):339, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-339a.pdf.

Anonymous:1897:SF

[7805] Anonymous. Sunday feeding. Scientific American, 76(22):339, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-339c.pdf.

Anonymous:1897:SPP

[7806] Anonymous. St. Petersburg–Pekin rail route. Scientific American, 76(22):339, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
339d.pdf.

Brooks:1897:HJ

[7807] William R. Brooks. The heavens for June. Scientific American, 76(22):339, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
339.pdf.

Anonymous:1897:JGA

[7808] Anonymous. Judge Greeley, Assistant Commissioner of Patents. Scientific American, 76(22):339–340, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
339e.pdf.

Anonymous:1897:MSb

[7809] Anonymous. Making spaghetti. Scientific American, 76(22):340–341, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-340.pdf.

Anonymous:1897:MAE

[7810] Anonymous. A mast arm for electric lamps. Scientific American, 76(22):341, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
341.pdf.

Anonymous:1897:OLD

[7811] Anonymous. Origin of leak in the Dry Dock, Brooklyn Navy Yard. Scientific American, 76(22):342, May 29, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n22/pdf/scientificamerican05291897-342.pdf.

Anonymous:1897:Ch

[7812] Anonymous. Correspondence. Scientific American, 76(22):343, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-343.pdf.

Anonymous:1897:GRS

[7813] Anonymous. The great railway station at Bombay, India. Scientific American, 76(22):344, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
344.pdf.

Curtis:1897:T

[7814] H. Holbrook Curtis. The tonograph. Scientific American, 76 (22):345–346, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
345.pdf.

Anonymous:1897:PGO

[7815] Anonymous. The production of grand opera. Scientific American, 76(22):346–347, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
346.pdf.

Anonymous:1897:BNBs

[7816] Anonymous. Books: New books, etc. Scientific American, 76(22):348, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-348a.pdf.

Anonymous:1897:RPIv

[7817] Anonymous. Recently patented inventions. Scientific American, 76(22):348, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n22/pdf/scientificamerican05291897-
348.pdf.

Anonymous:1897:BPv

[7818] Anonymous. Business and personal. Scientific American, 76(22):349, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-349.pdf.

Anonymous:1897:In

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-349b.pdf.

Anonymous:1897:IIv

[7820] Anonymous. Index of inventions. Scientific American, 76(22):349, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-349c.pdf.

Anonymous:1897:NQv

[7821] Anonymous. Notes and queries. Scientific American, 76(22):349, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-349a.pdf.

Anonymous:1897:PNPt

[7822] Anonymous. Patent notes: Patents. Scientific American, 76(22):350, May 29, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n22/pdf/scientificamerican05291897-350.pdf.

Anonymous:1897:SDS

[7823] Anonymous. Steam dumping scow for the street cleaning department of New York. Scientific American, 76(23):353, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-353.pdf.

Anonymous:1897:EBT

[7824] Anonymous. Experimental bicycle trip in the United States Army. Scientific American, 76(23):354, June 5, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n23/pdf/scientificamerican06051897-354c.pdf.

Anonymous:1897:GEI

[7825] Anonymous. The gas engine industry in the United States. Scientific American, 76(23):354, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
354.pdf.

Anonymous:1897:GMM

[7826] Anonymous. Gold Mining Machinery Exhibition at Coolgardie, West Australia. Scientific American, 76(23):354, June 5, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n23/pdf/scientificamerican06051897-354a.pdf.

Anonymous:1897:SCW

[7827] Anonymous. The second city in the world. Scientific American, 76(23):354, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
354b.pdf.

Anonymous:1897:UCN

[7828] Anonymous. Utilizing the Cataracts of the Nile. Scientific American, 76(23):354, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
354d.pdf.

Anonymous:1897:DM

[7829] Anonymous. The Detroit Meeting of the A.A.A.S. Scientific American, 76(23):355, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
355e.pdf.

Anonymous:1897:ESM

[7830] Anonymous. Engineers’ Society Meeting. Scientific American, 76 (23):355, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
355f.pdf.

Anonymous:1897:FP

[7831] Anonymous. Fog and photography. Scientific American, 76(23):355, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-355c.pdf.

Anonymous:1897:LEa

[7832] Anonymous. The Leipsic Exhibition. Scientific American, 76(23):355, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-355b.pdf.

Anonymous:1897:NPEb

[7833] Anonymous. The new Peary Expedition. Scientific American, 76(23):355, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
355.pdf.

Anonymous:1897:NTT

[7834] Anonymous. The new Thames Tunnel opened. Scientific American, 76(23):355, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
355d.pdf.

Anonymous:1897:TDK

[7835] Anonymous. Tornado drills in Kansas. Scientific American, 76(23):355, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-355a.pdf.

Anonymous:1897:A

[7836] Anonymous. Award of over. Scientific American, 76(23):356, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-356.pdf.

Anonymous:1897:BFI

[7837] Anonymous. A bicycle frame improvement. Scientific American, 76(23):356, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
356a.pdf.

Anonymous:1897:PBE

[7838] Anonymous. Peculiar boiler explosion. Scientific American, 76 (23):357–358, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
357.pdf.

Anonymous:1897:TIT

[7839] Anonymous. Three improved types of the Maxim automatic rapid fire gun. Scientific American, 76(23):360–361, June 5, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n23/pdf/scientificamerican06051897-360.pdf.

Anonymous:1897:MKT

[7840] Anonymous. Miss Kingsley’s travels in Africa. Scientific American, 76(23):361–362, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
361.pdf.

Anonymous:1897:WMPa

[7841] Anonymous. The Washington Monument, Philadelphia. Scientific American, 76(23):362–363, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
362.pdf.

Anonymous:1897:BNBt

[7842] Anonymous. Books: New books, etc. Scientific American, 76(23):364, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-364a.pdf.

Anonymous:1897:BPw

[7843] Anonymous. Business and personal. Scientific American, 76(23):364, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-364b.pdf.

Anonymous:1897:NQw

[7844] Anonymous. Notes and queries. Scientific American, 76(23):364, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-364c.pdf.

Anonymous:1897:RPIw

[7845] Anonymous. Recently patented inventions. Scientific American, 76(23):364, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n23/pdf/scientificamerican06051897-
364.pdf.

Anonymous:1897:IIw

[7846] Anonymous. Index of inventions. Scientific American, 76(23):364–365, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-364d.pdf.

Anonymous:1897:PNPu

[7847] Anonymous. Patent notes: Patents. Scientific American, 76(23):367, June 5, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n23/pdf/scientificamerican06051897-367.pdf.

Anonymous:1897:ETS

[7848] Anonymous. Electric traction on steam railroads. Scientific American, 76(24):370, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
370a.pdf.

Anonymous:1897:LT

[7849] Anonymous. Legislative trifling. Scientific American, 76(24):370, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-370.pdf.

Anonymous:1897:AYO

[7850] Anonymous. Accident at the Yerkes Observatory. Scientific American, 76(24):371, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
371c.pdf.

Anonymous:1897:EFB

[7851] Anonymous. An electric fountain for Brooklyn, N.Y. Scientific American, 76(24):371, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
371b.pdf.

Anonymous:1897:HMA

[7852] Anonymous. The Hartford meeting of the American Society of Mechanical Engineers. Scientific American, 76(24):371, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-371e.pdf.

Anonymous:1897:IFA

[7853] Anonymous. An important find of ancient papyri. Scientific American, 76(24):371, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
371d.pdf.

Anonymous:1897:LF

[7854] Anonymous. The liquefaction of fluorine. Scientific American, 76(24):371, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
371f.pdf.

Anonymous:1897:NPL

[7855] Anonymous. The new Public Library Building Competition. Scientific American, 76(24):371, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
371a.pdf.

Anonymous:1897:OPC

[7856] Anonymous. Opening of Philadelphia’s Commercial Museum. Scientific American, 76(24):371, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
371.pdf.

Anonymous:1897:GGH

[7857] Anonymous. Gasoline or gas hoisting engine. Scientific American, 76(24):372, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
372a.pdf.

Anonymous:1897:NGM

[7858] Anonymous. New gas motor. Scientific American, 76(24):372, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-372.pdf.

Anonymous:1897:NSB

[7859] Anonymous. A novel shaft balance. Scientific American, 76(24):372, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-372b.pdf.

Anonymous:1897:FEC

[7860] Anonymous. A fire engine car. Scientific American, 76(24):373–375, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-373.pdf.

Anonymous:1897:TRSa

[7861] Anonymous. The third rail system on the New Haven Railroad. Scientific American, 76(24):376–378, June 12, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n24/pdf/scientificamerican06121897-376.pdf.

Anonymous:1897:PCG

[7862] Anonymous. The probable cause of the great Paris fire. Scientific American, 76(24):378, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
378.pdf.

Anonymous:1897:BMT

[7863] Anonymous. The Battle at Milouna — a Turkish engraving. Scientific American, 76(24):379, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
379.pdf.

Anonymous:1897:RPIx

[7864] Anonymous. Recently patented inventions. Scientific American, 76(24):380, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
380.pdf.

Anonymous:1897:BNBu

[7865] Anonymous. Books: New books, etc. Scientific American, 76 (24):380–381, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n24/pdf/scientificamerican06121897-
380a.pdf.

Anonymous:1897:BPx

[7866] Anonymous. Business and personal. Scientific American, 76(24):381, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-381.pdf.

Anonymous:1897:Io

[7867] Anonymous. To inventors. Scientific American, 76(24):381, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-381b.pdf.

Anonymous:1897:IIx

[7868] Anonymous. Index of inventions. Scientific American, 76(24):381, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-381c.pdf.

Anonymous:1897:NQx

[7869] Anonymous. Notes and queries. Scientific American, 76(24):381, June 12, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-381a.pdf.

Anonymous:1897:PNPv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n24/pdf/scientificamerican06121897-383.pdf.

Anonymous:1897:DCHb

[7871] Anonymous. Direct conversion of heat into electricity. Scientific American, 76(25):386, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
386a.pdf.

Anonymous:1897:GPS

[7872] Anonymous. Great passenger stations of the United States. Scientific American, 76(25):386, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
386c.pdf.

Anonymous:1897:PPM

[7873] Anonymous. Paint as a protection to metallic structures. Scientific American, 76(25):386, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
386b.pdf.

Anonymous:1897:WT

[7874] Anonymous. Wireless telegraphy. Scientific American, 76(25):386, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-386.pdf.

Anonymous:1897:DAG

[7875] Anonymous. Death of Alvan G. Clark. Scientific American, 76(25):387, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-387b.pdf.

Anonymous:1897:FEP

[7876] Anonymous. Fall of an elevator in the Post Office Building, New York. Scientific American, 76(25):387, June 19, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v76/
n25/pdf/scientificamerican06191897-387a.pdf.

Anonymous:1897:KAD

[7877] Anonymous. Kites as an aid to discovery. Scientific American, 76(25):387, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
387c.pdf.

Anonymous:1897:LBT

[7878] Anonymous. The London Blackwall Tunnel. Scientific American, 76(25):387, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
387.pdf.

Anonymous:1897:SWb

[7879] Anonymous. Stature and weight. Scientific American, 76(25):387, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-387d.pdf.

Anonymous:1897:NTSb

[7880] Anonymous. The new twin screw express steamships of the North German Lloyd Company. Scientific American, 76(25):388, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-388.pdf.

Anonymous:1897:NPTb

[7881] Anonymous. A new printing telegraph instrument. Scientific American, 76(25):389, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
389.pdf.

Anonymous:1897:N

[7882] Anonymous. The “Neostyle”. Scientific American, 76(25):389–390, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-389a.pdf.

Hale:1897:RIX

[7883] William H. Hale. Recent investigations of X rays. Scientific American, 76(25):391, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
391.pdf.

Anonymous:1897:PCC

[7884] Anonymous. Pulpit in the Castle Chapel at Aschaffenburg. Scientific American, 76(25):392, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
392a.pdf.

Anonymous:1897:RDT

[7885] Anonymous. A remarkable double tree. Scientific American, 76(25):392, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-392.pdf.

Anonymous:1897:ACN

[7886] Anonymous. The Ancient City of New Amsterdam. Scientific American, 76(25):393, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
393.pdf.

Anonymous:1897:LWC

[7887] Anonymous. A large whale on a California beach. Scientific American, 76(25):394, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
394.pdf.

Anonymous:1897:CNJ

[7888] Anonymous. A cottage at Nutley N.J. Scientific American, 76(25):395, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-395.pdf.

Anonymous:1897:RPIy

[7889] Anonymous. Recently patented inventions. Scientific American, 76(25):396, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
396.pdf.

Anonymous:1897:BNBv

[7890] Anonymous. Books: New books, etc. Scientific American, 76 (25):396–397, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n25/pdf/scientificamerican06191897-
396a.pdf.

Anonymous:1897:BPy

[7891] Anonymous. Business and personal. Scientific American, 76(25):397, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-397.pdf.

Anonymous:1897:NQy

[7892] Anonymous. Notes and queries. Scientific American, 76(25):397, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-397a.pdf.

Anonymous:1897:IIy

[7893] Anonymous. Index of inventions. Scientific American, 76(25):397–399, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-397b.pdf.

Anonymous:1897:PNPw

[7894] Anonymous. Patent notes: Patents. Scientific American, 76(25):399, June 19, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n25/pdf/scientificamerican06191897-399.pdf.

Anonymous:1897:AIFa

[7895] Anonymous. American institute fair. Scientific American, 76(26):402, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-402.pdf.

Anonymous:1897:EMC

[7896] Anonymous. The Engineer Motor Car Competition. Scientific American, 76(26):402, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
402a.pdf.

Anonymous:1897:FNB

[7897] Anonymous. Fall of a new building in New York City. Scientific American, 76(26):402–403, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
402b.pdf.

Anonymous:1897:ATS

[7898] Anonymous. Another transatlantic steamship service. Scientific American, 76(26):403, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
403a.pdf.

Anonymous:1897:CAP

[7899] Anonymous. Charges against a patent firm. Scientific American, 76(26):403, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
403.pdf.

Anonymous:1897:NELb

[7900] Anonymous. National Electric Light Association. Scientific American, 76(26):403, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
403b.pdf.

Herring:1897:AFM

[7901] A. M. Herring. The aeroplane flying machine. Scientific American, 76(26):403–404, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
403c.pdf.

Anonymous:1897:CDC

[7902] Anonymous. A convenient drill case. Scientific American, 76(26):404, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-404a.pdf.

Anonymous:1897:IBBa

[7903] Anonymous. Improved binding for blank books. Scientific American, 76(26):404, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
404b.pdf.

Anonymous:1897:IRD

[7904] Anonymous. An improved rock drill. Scientific American, 76(26):404, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-404.pdf.

Anonymous:1897:LCC

[7905] Anonymous. The Lumière cinematograph camera. Scientific American, 76(26):405, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
405.pdf.

Anonymous:1897:AGC

[7906] Anonymous. Alvan Graham clark. Scientific American, 76(26):406, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-406.pdf.

Anonymous:1897:Ci

[7907] Anonymous. Correspondence. Scientific American, 76(26):407, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-407.pdf.

Anonymous:1897:TRE

[7908] Anonymous. Third rail electrical equipment of the New York, New Haven and Hartford Railroad. Scientific American, 76(26):408, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-408.pdf.

Anonymous:1897:MBP

[7909] Anonymous. The mechanical baseball pitcher. Scientific American, 76(26):409–410, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
409.pdf.

Anonymous:1897:RPIz

[7910] Anonymous. Recently patented inventions. Scientific American, 76(26):410, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v76/n26/pdf/scientificamerican06261897-
410.pdf.

Anonymous:1897:BNBw

[7911] Anonymous. Books: New books, etc. Scientific American, 76(26):411, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-411b.pdf.

Anonymous:1897:BPz

[7912] Anonymous. Business and personal. Scientific American, 76(26):411, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-411.pdf.

Anonymous:1897:Iq

[7913] Anonymous. To inventors. Scientific American, 76(26):411, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-411c.pdf.

Anonymous:1897:IIz

[7914] Anonymous. Index of inventions. Scientific American, 76(26):411, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-411d.pdf.

Anonymous:1897:NQz

[7915] Anonymous. Notes and queries. Scientific American, 76(26):411, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-411a.pdf.

Anonymous:1897:Ir

[7916] Anonymous. Index. Scientific American, 76(26):413–414, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-413.pdf.

Anonymous:1897:PNPx

[7917] Anonymous. Patent notes: Patents. Scientific American, 76(26):415, June 26, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v76/n26/pdf/scientificamerican06261897-415.pdf.

Anonymous:1897:ABC

[7918] Anonymous. American built cruisers for Japan. Scientific American, 77(1):3–4, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n1/pdf/scientificamerican07031897-
3.pdf.

Anonymous:1897:NRP

[7919] Anonymous. The new rules of practice of the Patent Office abrogated. Scientific American, 77(1):4, July 3, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n1/pdf/scientificamerican07031897-4a.pdf.

Anonymous:1897:PBA

[7920] Anonymous. Proposed bridge across the St. Lawrence at Quebec. Scientific American, 77(1):4, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n1/pdf/scientificamerican07031897-
4.pdf.

Anonymous:1897:FM

[7921] Anonymous. The fatigue of metals. Scientific American, 77(1):4–5, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-4b.pdf.

Anonymous:1897:DPF

[7922] Anonymous. Death of Prof Fresenius. Scientific American, 77(1):5, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-5c.pdf.

Anonymous:1897:IMS

[7923] Anonymous. Interesting marine statistics. Scientific American, 77(1): 5, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-5b.pdf.

Anonymous:1897:MGT

[7924] Anonymous. Municipal government without taxation. Scientific American, 77(1):5, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n1/pdf/scientificamerican07031897-
5.pdf.

Anonymous:1897:SVP

[7925] Anonymous. Settlement of the vacant public lands of the United States. Scientific American, 77(1):5, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n1/pdf/scientificamerican07031897-
5a.pdf.

Anonymous:1897:IBBb

[7926] Anonymous. An improved bicycle brake. Scientific American, 77(1):6, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-6.pdf.

Anonymous:1897:TWQ

[7927] Anonymous. Tests of the wearing qualities of flooring materials. Scientific American, 77(1):6, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n1/pdf/scientificamerican07031897-
6a.pdf.

Anonymous:1897:VTL

[7928] Anonymous. A valve testing and locking device. Scientific American, 77(1):6, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n1/pdf/scientificamerican07031897-
6b.pdf.

Anonymous:1897:ARSb

[7929] Anonymous. An automatic railway switch. Scientific American, 77(1): 7, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-7.pdf.

Anonymous:1897:WVB

[7930] Anonymous. The “Winged Victory” for the battleship Massachusetts. Scientific American, 77(1):7, July 3, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n1/pdf/scientificamerican07031897-7a.pdf.

Anonymous:1897:IS

[7931] Anonymous. An improved swing. Scientific American, 77(1):7–8, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-7b.pdf.

Anonymous:1897:Cj

[7932] Anonymous. Correspondence. Scientific American, 77(1):9, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-9.pdf.

Anonymous:1897:PKB

[7933] Anonymous. Parseval’s kite balloon. Scientific American, 77(1):10, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-10.pdf.

Brown:1897:LY

[7934] Enos Brown. The last of the Yukiahs. Scientific American, 77(1):11, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-11.pdf.

Anonymous:1897:BNBx

[7935] Anonymous. Books: New books, etc. Scientific American, 77(1):12, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-12a.pdf.

Anonymous:1897:RPIba

[7936] Anonymous. Recently patented inventions. Scientific American, 77(1): 12, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-12.pdf.

Anonymous:1897:BPba

[7937] Anonymous. Business and personal. Scientific American, 77(1):13, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-13.pdf.

Anonymous:1897:IIba

[7938] Anonymous. Index of inventions. Scientific American, 77(1):13, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-13b.pdf.

Anonymous:1897:NQba

[7939] Anonymous. Notes and queries. Scientific American, 77(1):13, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-13a.pdf.

Anonymous:1897:PNPy

[7940] Anonymous. Patent notes: Patents. Scientific American, 77(1):15, July 3, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n1/pdf/scientificamerican07031897-15.pdf.

Anonymous:1897:NBSb

[7941] Anonymous. The Nobel Bequest to Science. Scientific American, 77(2):18, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
18.pdf.

Anonymous:1897:THM

[7942] Anonymous. Two hundred miles on a bicycle in one day. Scientific American, 77(2):18–19, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
18a.pdf.

Anonymous:1897:DPV

[7943] Anonymous. Death of Professor De Volson Wood. Scientific American, 77(2):19, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
19c.pdf.

Anonymous:1897:DSL

[7944] Anonymous. Deepening the St. Lawrence. Scientific American, 77(2):19, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
19e.pdf.

Anonymous:1897:ELS

[7945] Anonymous. Electric light and sailors’ eyesight. Scientific American, 77(2):19, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
19d.pdf.

Anonymous:1897:LIP

[7946] Anonymous. Lottery induced patents. Scientific American, 77(2):19, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-19b.pdf.

Anonymous:1897:NLC

[7947] Anonymous. A new Librarian of Congress. Scientific American, 77(2):19, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
19.pdf.

Anonymous:1897:NRH

[7948] Anonymous. New rapid harbor mail service. Scientific American, 77(2):19, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
19a.pdf.

Anonymous:1897:IBL

[7949] Anonymous. An improved bullet lubricator. Scientific American, 77(2):20, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
20b.pdf.

Anonymous:1897:NPPb

[7950] Anonymous. A novel plant protector. Scientific American, 77(2):20, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-20c.pdf.

Anonymous:1897:PPB

[7951] Anonymous. A puncture proof bicycle tire. Scientific American, 77(2):20, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
20a.pdf.

Anonymous:1897:SBB

[7952] Anonymous. A special brazer for bicycle work. Scientific American, 77(2):20, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
20.pdf.

Anonymous:1897:EEB

[7953] Anonymous. An efficient electric bicycle lamp. Scientific American, 77(2):20–21, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
20d.pdf.

Anonymous:1897:CM

[7954] Anonymous. A Carthaginian mask. Scientific American, 77(2):21, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-21a.pdf.

Anonymous:1897:RPTc

[7955] Anonymous. A remarkable pear tree. Scientific American, 77(2):21, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-21.pdf.

Anonymous:1897:FSR

[7956] Anonymous. Fighting snow on the railroads of the Northwest. Scientific American, 77(2):22–23, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
22.pdf.

Anonymous:1897:SNS

[7957] Anonymous. Some notable steam yachts. Scientific American, 77(2):24–25, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
24.pdf.

Anonymous:1897:RB

[7958] Anonymous. Roman baths. Scientific American, 77(2):25–26, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-25.pdf.

Anonymous:1897:IRIa

[7959] Anonymous. Interesting recent inventions. Scientific American, 77(2):26–28, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
26.pdf.

Anonymous:1897:RPIbb

[7960] Anonymous. Recently patented inventions. Scientific American, 77(2):28–29, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n2/pdf/scientificamerican07101897-
28.pdf.

Anonymous:1897:BNBy

[7961] Anonymous. Books: New books, etc. Scientific American, 77(2):29, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-29b.pdf.

Anonymous:1897:BPbb

[7962] Anonymous. Business and personal. Scientific American, 77(2):29, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-29.pdf.

Anonymous:1897:It

[7963] Anonymous. To inventors. Scientific American, 77(2):29, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-29c.pdf.

Anonymous:1897:IIbb

[7964] Anonymous. Index of inventions. Scientific American, 77(2):29, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-29d.pdf.

Anonymous:1897:NQbb

[7965] Anonymous. Notes and queries. Scientific American, 77(2):29, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-29a.pdf.

Anonymous:1897:PNPz

[7966] Anonymous. Patent notes: Patents. Scientific American, 77(2):31, July 10, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n2/pdf/scientificamerican07101897-31.pdf.

Anonymous:1897:APC

[7967] Anonymous. The armor plate compromise. Scientific American, 77(3):34, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
34c.pdf.

Anonymous:1897:IEN

[7968] Anonymous. Improvement of the entrance to New York Harbor. Scientific American, 77(3):34, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
34.pdf.

Anonymous:1897:MSV

[7969] Anonymous. The Moore system of vacuum tube lighting. Scientific American, 77(3):34, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
34b.pdf.

Anonymous:1897:TRSb

[7970] Anonymous. The third rail system in England. Scientific American, 77(3):34, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
34a.pdf.

Anonymous:1897:IAF

[7971] Anonymous. The inaccuracy of artillery fire. Scientific American, 77(3):34–35, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
34d.pdf.

Anonymous:1897:EMS

[7972] Anonymous. The evolution of modern scientific laboratories. Scientific American, 77(3):35, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
35a.pdf.

Anonymous:1897:FMC

[7973] Anonymous. The future of the motor car. Scientific American, 77(3):35, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-35.pdf.

Anonymous:1897:TBC

[7974] Anonymous. Test of the Buffington–Crozier disappearing gun carriage. Scientific American, 77(3):35, July 17, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n3/pdf/scientificamerican07171897-35b.pdf.

Anonymous:1897:IBBc

[7975] Anonymous. An improved bicycle brake. Scientific American, 77(3):36, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-36.pdf.

Anonymous:1897:PSAb

[7976] Anonymous. A pneumatic spraying apparatus. Scientific American, 77(3):36, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
36a.pdf.

Anonymous:1897:SOC

[7977] Anonymous. A stove and oven for “camping out” parties, prospectors, boating, house use, etc. Scientific American, 77(3):36, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-36b.pdf.

Anonymous:1897:DT

[7978] Anonymous. A destructive thunderbolt. Scientific American, 77(3):37, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-37.pdf.

Anonymous:1897:PWS

[7979] Anonymous. Petroleum wells in the suburbs of Los Angeles, California. Scientific American, 77(3):37–39, July 17, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n3/pdf/scientificamerican07171897-37a.pdf.

Anonymous:1897:DRF

[7980] Anonymous. The denudation and recovery of farm lands. Scientific American, 77(3):39–40, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
39.pdf.

Anonymous:1897:TAD

[7981] Anonymous. Third Avenue Drawbridge across the Harlem River, New York City. Scientific American, 77(3):41–43, July 17, 1897.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n3/pdf/scientificamerican07171897-41.pdf.

Anonymous:1897:NCD

[7982] Anonymous. New circle divider. Scientific American, 77(3):43, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-43.pdf.

Anonymous:1897:RPIbc

[7983] Anonymous. Recently patented inventions. Scientific American, 77(3):44, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n3/pdf/scientificamerican07171897-
44.pdf.

Anonymous:1897:BNBz

[7984] Anonymous. Books: New books, etc. Scientific American, 77(3):44–45, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-44a.pdf.

Anonymous:1897:BPbc

[7985] Anonymous. Business and personal. Scientific American, 77(3):45, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-45a.pdf.

Anonymous:1897:IIbc

[7986] Anonymous. Index of inventions. Scientific American, 77(3):45, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-45b.pdf.

Anonymous:1897:NQbc

[7987] Anonymous. Notes and queries. Scientific American, 77(3):45, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-45.pdf.

Anonymous:1897:PNPba

[7988] Anonymous. Patent notes: Patents. Scientific American, 77(3):47, July 17, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n3/pdf/scientificamerican07171897-47.pdf.

Anonymous:1897:ETC

[7989] Anonymous. Electric tower clock and chimes, Grace Chapel, New York. Scientific American, 77(4):49–50, July 24, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n4/pdf/scientificamerican07241897-49.pdf.

Anonymous:1897:EOF

[7990] Anonymous. The expansion of our foreign trade. Scientific American, 77(4):50–51, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n4/pdf/scientificamerican07241897-
50.pdf.

Anonymous:1897:ASN

[7991] Anonymous. Andrée starts North. Scientific American, 77(4):51, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-51d.pdf.

Anonymous:1897:BTM

[7992] Anonymous. British trade mark decision. Scientific American, 77(4):51, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-51c.pdf.

Anonymous:1897:CLCa

[7993] Anonymous. City light from city refuse. Scientific American, 77(4):51, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-51.pdf.

Anonymous:1897:DPMb

[7994] Anonymous. Death of Professor Mayer. Scientific American, 77(4):51, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-51b.pdf.

Anonymous:1897:WP

[7995] Anonymous. Weather prognostications. Scientific American, 77(4):51, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-51a.pdf.

Anonymous:1897:ITa

[7996] Anonymous. An improved trestle. Scientific American, 77(4):52, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-52.pdf.

Anonymous:1897:RSM

[7997] Anonymous. Removing scale from metal plates. Scientific American, 77(4):52, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n4/pdf/scientificamerican07241897-
52b.pdf.

Anonymous:1897:WMC

[7998] Anonymous. The Winton motor carriage. Scientific American, 77(4):52, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-52a.pdf.

Anonymous:1897:TR

[7999] Anonymous. Trick riding. Scientific American, 77(4):53–54, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-53.pdf.

Anonymous:1897:Ck

[8000] Anonymous. Correspondence. Scientific American, 77(4):55, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-55.pdf.

Anonymous:1897:VHL

[8001] Anonymous. A vapor hunting launch. Scientific American, 77(4):56, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-56.pdf.

Anonymous:1897:EBD

[8002] Anonymous. The explosion of the balloon Deutschland. Scientific American, 77(4):57, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n4/pdf/scientificamerican07241897-
57.pdf.

Anonymous:1897:SEB

[8003] Anonymous. A simple and efficient bicycle brake. Scientific American, 77(4):58, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n4/pdf/scientificamerican07241897-
58.pdf.

Holder:1897:CFF

[8004] C. F. Holder. The California flying fish and its enemies. Scientific American, 77(4):59–60, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n4/pdf/scientificamerican07241897-
59.pdf.

Anonymous:1897:RPIbd

[8005] Anonymous. Recently patented inventions. Scientific American, 77(4):60, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n4/pdf/scientificamerican07241897-
60.pdf.

Anonymous:1897:BNBba

[8006] Anonymous. Books: New books, etc. Scientific American, 77(4):61, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-61.pdf.

Anonymous:1897:BPbd

[8007] Anonymous. Business and personal. Scientific American, 77(4):61, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-61a.pdf.

Anonymous:1897:IIbd

[8008] Anonymous. Index of inventions. Scientific American, 77(4):61, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-61c.pdf.

Anonymous:1897:NQbd

[8009] Anonymous. Notes and queries. Scientific American, 77(4):61, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-61b.pdf.

Anonymous:1897:PNPbb

[8010] Anonymous. Patent notes: Patents. Scientific American, 77(4):63, July 24, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n4/pdf/scientificamerican07241897-63.pdf.

Anonymous:1897:SIU

[8011] Anonymous. The sugar industry in the United States. Scientific American, 77(5):66, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
66.pdf.

Anonymous:1897:JTM

[8012] Anonymous. Jeopardizing trade mark property. Scientific American, 77(5):66–67, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
66a.pdf.

Anonymous:1897:LPT

[8013] Anonymous. The latest performance of the Turbinia. Scientific American, 77(5):67, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
67.pdf.

Anonymous:1897:O

[8014] Anonymous. Obesity. Scientific American, 77(5):67, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-67c.pdf.

Anonymous:1897:PTM

[8015] Anonymous. Pneumatic tube mail service for New York City. Scientific American, 77(5):67, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
67a.pdf.

Brooks:1897:HAb

[8016] William R. Brooks. The heavens for August. Scientific American, 77(5):67, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
67b.pdf.

Anonymous:1897:ESS

[8017] Anonymous. An efficient steam saver. Scientific American, 77(5):68, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-68.pdf.

Hungerford:1897:WF

[8018] Churchill Hungerford. Water filtration. Scientific American, 77 (5):68–69, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
68a.pdf.

Anonymous:1897:EUT

[8019] Anonymous. Extension of the underground trolley system in New York. Scientific American, 77(5):69–71, July 31, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n5/pdf/scientificamerican07311897-69.pdf.

Anonymous:1897:IRC

[8020] Anonymous. An improved railroad crossing signal. Scientific American, 77(5):71, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
71.pdf.

Anonymous:1897:MBS

[8021] Anonymous. The manufacture of beet sugar. Scientific American, 77(5):72–73, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
72.pdf.

Anonymous:1897:AMM

[8022] Anonymous. Alfred Marshall Mayer. Scientific American, 77(5):74, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-74.pdf.

Anonymous:1897:MT

[8023] Anonymous. The magic table. Scientific American, 77(5):75, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-75a.pdf.

StClair:1897:M

[8024] D. F. St. Clair. The micromotoscope. Scientific American, 77(5):75, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-75.pdf.

Anonymous:1897:RPIbe

[8025] Anonymous. Recently patented inventions. Scientific American, 77(5):76, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n5/pdf/scientificamerican07311897-
76.pdf.

Anonymous:1897:BNBbb

[8026] Anonymous. Books: New books, etc. Scientific American, 77(5):76–77, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-76a.pdf.

Anonymous:1897:BPbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-77.pdf.

Anonymous:1897:Iu

[8028] Anonymous. To inventors. Scientific American, 77(5):77, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-77b.pdf.

Anonymous:1897:IIbe

[8029] Anonymous. Index of inventions. Scientific American, 77(5):77, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-77c.pdf.

Anonymous:1897:NQbe

[8030] Anonymous. Notes and queries. Scientific American, 77(5):77, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-77a.pdf.

Anonymous:1897:PNPbc

[8031] Anonymous. Patent notes: Patents. Scientific American, 77(5):79, July 31, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n5/pdf/scientificamerican07311897-79.pdf.

Anonymous:1897:ACC

[8032] Anonymous. Appeal in the Cordite Case rejected. Scientific American, 77(6):82, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
82.pdf.

Anonymous:1897:CPN

[8033] Anonymous. The commerce of the Port of New York. Scientific American, 77(6):82, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
82a.pdf.

Anonymous:1897:AFP

[8034] Anonymous. Alleged fraudulent patent business. Scientific American, 77(6):82–83, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
82b.pdf.

Anonymous:1897:AAA

[8035] Anonymous. American Association for the Advancement of Science. Scientific American, 77(6):83, August 7, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n6/pdf/scientificamerican08071897-83.pdf.

Anonymous:1897:ARF

[8036] Anonymous. The automobiles race in France. Scientific American, 77(6):83, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
83b.pdf.

Anonymous:1897:CDW

[8037] Anonymous. The curious death of a whale. Scientific American, 77(6):83, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
83d.pdf.

Anonymous:1897:PCS

[8038] Anonymous. Preparing for civil service examinations. Scientific American, 77(6):83, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
83a.pdf.

Anonymous:1897:PID

[8039] Anonymous. The Parthenon inscription deciphered. Scientific American, 77(6):83, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
83f.pdf.

Anonymous:1897:PLC

[8040] Anonymous. Prof. Libbey conquers the Mesa. Scientific American, 77(6):83, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
83c.pdf.

Anonymous:1897:PPLb

[8041] Anonymous. Puluj’s phosphorescent lamp. Scientific American, 77(6):83, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
83e.pdf.

Anonymous:1897:NAV

[8042] Anonymous. A nicotine absorbent ventilated tobacco pipe. Scientific American, 77(6):84, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
84b.pdf.

Anonymous:1897:NFC

[8043] Anonymous. A novel folding chair. Scientific American, 77(6):84, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-84.pdf.

Anonymous:1897:RCF

[8044] Anonymous. A rocking chair with fan and music box. Scientific American, 77(6):84, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
84a.pdf.

Triepel:1897:MS

[8045] Emma V. Triepel. The Marvin seismograph. Scientific American, 77(6):85, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
85.pdf.

Anonymous:1897:HB

[8046] Anonymous. A horseless Brougham. Scientific American, 77(6):85–86, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-85a.pdf.

Anonymous:1897:Cl

[8047] Anonymous. Correspondence. Scientific American, 77(6):87, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-87.pdf.

Anonymous:1897:XRC

[8048] Anonymous. The X rays in the custom house. Scientific American, 77(6):88–89, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
88.pdf.

Benjamin:1897:MAA

[8049] Marcus Benjamin. Meeting of the American Association. Scientific American, 77(6):89–90, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
89.pdf.

Anonymous:1897:FER

[8050] Anonymous. The foundations of the East River Bridge, New York. Scientific American, 77(6):90–91, August 7, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n6/pdf/scientificamerican08071897-90.pdf.

Anonymous:1897:BNBbc

[8051] Anonymous. Books: New books, etc. Scientific American, 77(6):92, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-92c.pdf.

Anonymous:1897:BPbf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-92a.pdf.

Anonymous:1897:NQbf

[8053] Anonymous. Notes and queries. Scientific American, 77(6):92, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-92b.pdf.

Anonymous:1897:RPIbf

[8054] Anonymous. Recently patented inventions. Scientific American, 77(6):92, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n6/pdf/scientificamerican08071897-
92.pdf.

Anonymous:1897:Ix

[8055] Anonymous. To inventors. Scientific American, 77(6):93, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-93.pdf.

Anonymous:1897:IIbf

[8056] Anonymous. Index of inventions. Scientific American, 77(6):93, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-93a.pdf.

Anonymous:1897:PNPbd

[8057] Anonymous. Patent notes: Patents. Scientific American, 77(6):95, August 7, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n6/pdf/scientificamerican08071897-95.pdf.

Anonymous:1897:AGF

[8058] Anonymous. The Alaskan gold fields. Scientific American, 77(7):98, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-98.pdf.

Anonymous:1897:SWT

[8059] Anonymous. Steel wagon tracks on country roads. Scientific American, 77(7):98, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
98a.pdf.

Anonymous:1897:PCH

[8060] Anonymous. Proposed completion of the Hudson River Tunnel. Scientific American, 77(7):98–99, August 14, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n7/pdf/scientificamerican08141897-98b.pdf.

Anonymous:1897:BAE

[8061] Anonymous. The Belgian Antarctic Expedition. Scientific American, 77(7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
99d.pdf.

Anonymous:1897:DPMc

[8062] Anonymous. Death of Prof. McClure. Scientific American, 77 (7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
99a.pdf.

Anonymous:1897:ETF

[8063] Anonymous. Elevated tries fuel oil — unsuccessful experiment with one train on third avenue. Scientific American, 77(7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-99b.pdf.

Anonymous:1897:FTW

[8064] Anonymous. The fastest train in the world. Scientific American, 77(7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
99.pdf.

Anonymous:1897:MNL

[8065] Anonymous. Moving to the new Library, Washington. Scientific American, 77(7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
99f.pdf.

Anonymous:1897:MSI

[8066] Anonymous. Magic: Stage illusions and scientific diversions. Scientific American, 77(7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
99e.pdf.

Anonymous:1897:SAF

[8067] Anonymous. Salicylic acid in food. Scientific American, 77(7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-99g.pdf.

Anonymous:1897:TST

[8068] Anonymous. Trial of a sextuplex telegraph system. Scientific American, 77(7):99, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
99c.pdf.

Anonymous:1897:RMCa

[8069] Anonymous. Railway mail catcher and delivery devices. Scientific American, 77(7):100, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
100.pdf.

Benjamin:1897:SJE

[8070] Marcus Benjamin. Sir John Evans. Scientific American, 77(7):101, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-101.pdf.

Anonymous:1897:UNY

[8071] Anonymous. The utilization of New York City garbage. Scientific American, 77(7):102–103, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
102.pdf.

Anonymous:1897:DUS

[8072] Anonymous. Drill in the United States Navy. Scientific American, 77(7):104–106, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
104.pdf.

Anonymous:1897:IDa

[8073] Anonymous. An important decision. Scientific American, 77 (7):106–107, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
106.pdf.

Anonymous:1897:NMB

[8074] Anonymous. A new method of building submerged foundations. Scientific American, 77(7):107, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
107.pdf.

Anonymous:1897:BPbg

[8075] Anonymous. Business and personal. Scientific American, 77(7):108, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-108a.pdf.

Anonymous:1897:RPIbg

[8076] Anonymous. Recently patented inventions. Scientific American, 77(7):108, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
108.pdf.

Anonymous:1897:BNBbd

[8077] Anonymous. Books: New books, etc. Scientific American, 77 (7):108–109, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n7/pdf/scientificamerican08141897-
108b.pdf.

Anonymous:1897:Iy

[8078] Anonymous. To inventors. Scientific American, 77(7):109, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-109a.pdf.

Anonymous:1897:IIbg

[8079] Anonymous. Index of inventions. Scientific American, 77(7):109, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-109b.pdf.

Anonymous:1897:NQbg

[8080] Anonymous. Notes and queries. Scientific American, 77(7):109, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-109.pdf.

Anonymous:1897:PNPbe

[8081] Anonymous. Patent notes: Patents. Scientific American, 77(7):111, August 14, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n7/pdf/scientificamerican08141897-111.pdf.

Prindle:1897:BCN

[8082] E. J. Prindle. The book carriers of the new Library of Congress Building. Scientific American, 77(8):113, August 21, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n8/pdf/scientificamerican08211897-113.pdf.

Anonymous:1897:DMP

[8083] Anonymous. Deceptive methods of patent attorneys. Scientific American, 77(8):114, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
114b.pdf.

Anonymous:1897:EOS

[8084] Anonymous. Extension of our South American trade. Scientific American, 77(8):114, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
114.pdf.

Anonymous:1897:PRT

[8085] Anonymous. The proposed rapid transit tunnel, New York City. Scientific American, 77(8):114, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
114a.pdf.

Anonymous:1897:P

[8086] Anonymous. Prosperity. Scientific American, 77(8):114–115, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n8/pdf/scientificamerican08211897-114c.pdf.

Anonymous:1897:ORP

[8087] Anonymous. Official registration of patent attorneys. Scientific American, 77(8):115, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
115a.pdf.

Anonymous:1897:SSL

[8088] Anonymous. A successful solution of the locomotive counterbalance problem. Scientific American, 77(8):115, August 21, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n8/pdf/scientificamerican08211897-115.pdf.

Anonymous:1897:RMCb

[8089] Anonymous. Railway mail catcher and delivery devices. Scientific American, 77(8):116, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
116.pdf.

Anonymous:1897:ABI

[8090] Anonymous. Air brake instruction car on the “Big Four Route”. Scientific American, 77(8):117–119, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
117.pdf.

Anonymous:1897:SFP

[8091] Anonymous. The seagoing ferry “Père Marquette”. Scientific American, 77(8):121–122, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
121.pdf.

Anonymous:1897:NHDb

[8092] Anonymous. A nail holding device for hammers. Scientific American, 77(8):122, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
122.pdf.

Caulk:1897:BA

[8093] William B. Caulk. Black art. Scientific American, 77(8):123–124, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n8/pdf/scientificamerican08211897-123.pdf.

Anonymous:1897:RPIbh

[8094] Anonymous. Recently patented inventions. Scientific American, 77(8):124, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
124.pdf.

Anonymous:1897:BNBbe

[8095] Anonymous. Books: New books, etc. Scientific American, 77 (8):124–125, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n8/pdf/scientificamerican08211897-
124a.pdf.

Anonymous:1897:BPbh

[8096] Anonymous. Business and personal. Scientific American, 77(8):125, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n8/pdf/scientificamerican08211897-125.pdf.

Anonymous:1897:Iz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n8/pdf/scientificamerican08211897-125b.pdf.

Anonymous:1897:IIbh

[8098] Anonymous. Index of inventions. Scientific American, 77(8):125, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n8/pdf/scientificamerican08211897-125c.pdf.

Anonymous:1897:NQbh

[8099] Anonymous. Notes and queries. Scientific American, 77(8):125, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n8/pdf/scientificamerican08211897-125a.pdf.

Anonymous:1897:PNPbf

[8100] Anonymous. Patent notes: Patents. Scientific American, 77(8):127, August 21, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n8/pdf/scientificamerican08211897-127.pdf.

Anonymous:1897:RBS

[8101] Anonymous. The relation of the beet sugar farm to the factory. Scientific American, 77(9):130, August 28, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n9/pdf/scientificamerican08281897-130.pdf.

Anonymous:1897:RDD

[8102] Anonymous. Repairs to Dry Dock No. 3 at the Brooklyn Navy Yard. Scientific American, 77(9):130, August 28, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n9/pdf/scientificamerican08281897-130a.pdf.

Benjamin:1897:AAA

[8103] Marcus Benjamin. The American Association for the Advancement of Science. Scientific American, 77(9):130–131, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-130b.pdf.

Anonymous:1897:IBS

[8104] Anonymous. An improved balanced slide valve. Scientific American, 77(9):132, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
132a.pdf.

Anonymous:1897:IHP

[8105] Anonymous. As improved horseshoe pad. Scientific American, 77(9):132, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
132b.pdf.

Anonymous:1897:NRBb

[8106] Anonymous. A non-refillable bottle. Scientific American, 77(9):132, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-132c.pdf.

Anonymous:1897:SNL

[8107] Anonymous. A simple nut lock. Scientific American, 77(9):132, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-132.pdf.

Anonymous:1897:HAH

[8108] Anonymous. Hayward Augustus Harvey. Scientific American, 77(9):133, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
133.pdf.

Anonymous:1897:SCB

[8109] Anonymous. A suburban cottage at Boston, Mass. Scientific American, 77(9):134, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
134.pdf.

Brooks:1897:HS

[8110] William B. Brooks. The heavens for September. Scientific American, 77(9):135, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
135.pdf.

Anonymous:1897:DWN

[8111] Anonymous. Disposal of the wastes of New York City. Scientific American, 77(9):136–137, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
136.pdf.

Anonymous:1897:OG

[8112] Anonymous. The orchestral guitar. Scientific American, 77(9): 137–138, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
137.pdf.

Caulk:1897:ST

[8113] W. B. Caulk. Spiritualistic ties. Scientific American, 77(9):139–140, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-139.pdf.

Anonymous:1897:RPIbi

[8114] Anonymous. Recently patented inventions. Scientific American, 77(9):140, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
140.pdf.

Anonymous:1897:BNBbf

[8115] Anonymous. Books: New books, etc. Scientific American, 77 (9):140–141, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n9/pdf/scientificamerican08281897-
140a.pdf.

Anonymous:1897:BPbi

[8116] Anonymous. Business and personal. Scientific American, 77(9):141, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-141.pdf.

Anonymous:1897:Ibc

[8117] Anonymous. To inventors. Scientific American, 77(9):141, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-141b.pdf.

Anonymous:1897:IIbi

[8118] Anonymous. Index of inventions. Scientific American, 77(9):141, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-141c.pdf.

Anonymous:1897:NQbi

[8119] Anonymous. Notes and queries. Scientific American, 77(9):141, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-141a.pdf.

Anonymous:1897:PNPbg

[8120] Anonymous. Patent notes: Patents. Scientific American, 77(9):143, August 28, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n9/pdf/scientificamerican08281897-143.pdf.

Anonymous:1897:AIT

[8121] Anonymous. The American iron trade for 1896. Scientific American, 77(10):146, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
146.pdf.

Anonymous:1897:BAMa

[8122] Anonymous. The British Association Meeting. Scientific American, 77(10):146–147, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
146a.pdf.

Anonymous:1897:MSEb

[8123] Anonymous. Mount St. Elias ascended. Scientific American, 77 (10):147, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
147a.pdf.

Anonymous:1897:PE

[8124] Anonymous. The Paris Exposition of 1900. Scientific American, 77(10):147, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
147.pdf.

Anonymous:1897:WAV

[8125] Anonymous. Will ascertain the value of calcium carbide. Scientific American, 77(10):147, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
147b.pdf.

Anonymous:1897:CSE

[8126] Anonymous. A condenses for steam engines. Scientific American, 77(10):148, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
148.pdf.

Anonymous:1897:RVE

[8127] Anonymous. A reversing valve for engines. Scientific American, 77(10):148, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
148b.pdf.

Anonymous:1897:TNR

[8128] Anonymous. Two new range finders. Scientific American, 77 (10):148, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
148a.pdf.

Anonymous:1897:ICB

[8129] Anonymous. Inserting corks in bottles. Scientific American, 77 (10):148–149, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
148c.pdf.

Anonymous:1897:UB

[8130] Anonymous. The umbrella boat. Scientific American, 77(10): 149, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
149.pdf.

Anonymous:1897:IFB

[8131] Anonymous. An improved form of Bunsen burner. Scientific American, 77(10):149–150, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
149a.pdf.

Anonymous:1897:CCP

[8132] Anonymous. Combination corner piece. Scientific American, 77 (10):150–151, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
150.pdf.

Anonymous:1897:HELa

[8133] Anonymous. The Heilmann electric locomotive. Scientific American, 77(10):152, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
152.pdf.

Mindelerp:1897:HBR

[8134] Cosmos Mindelerp. House building by ritual. Scientific American, 77 (10):153–154, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
153.pdf.

Anonymous:1897:FT

[8135] Anonymous. Firing a torpedo. Scientific American, 77(10):155, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n10/pdf/scientificamerican09041897-155.pdf.

Caulk:1897:SPT

[8136] W. B. Caulk. The “Spiritualistic Post Test”. Scientific American, 77(10):155, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
155a.pdf.

Anonymous:1897:BNBbg

[8137] Anonymous. Books: New books, etc. Scientific American, 77 (10):156, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
156c.pdf.

Anonymous:1897:BPbj

[8138] Anonymous. Business and personal. Scientific American, 77 (10):156, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
156a.pdf.

Anonymous:1897:NQbj

[8139] Anonymous. Notes and queries. Scientific American, 77(10): 156, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
156b.pdf.

Anonymous:1897:RPIbj

[8140] Anonymous. Recently patented inventions. Scientific American, 77(10):156, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
156.pdf.

Anonymous:1897:Ibd

[8141] Anonymous. To inventors. Scientific American, 77(10):157, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n10/pdf/scientificamerican09041897-157.pdf.

Anonymous:1897:IIbj

[8142] Anonymous. Index of inventions. Scientific American, 77(10): 157, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
157a.pdf.

Anonymous:1897:PNPbh

[8143] Anonymous. Patent notes: Patents. Scientific American, 77 (10):159, September 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n10/pdf/scientificamerican09041897-
159.pdf.

Hopkins:1897:HBS

[8144] Nevil Monroe Hopkins. How to build a small alternating current dynamo without castings. Scientific American, 77(11):161, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n11/pdf/scientificamerican09111897-161.pdf.

Anonymous:1897:AMA

[8145] Anonymous. The antiquity of Man in America. Scientific American, 77(11):162, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
162.pdf.

Anonymous:1897:RSP

[8146] Anonymous. Repairing the screw propeller of a transatlantic steamer. Scientific American, 77(11):162, September 11, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n11/pdf/scientificamerican09111897-162a.pdf.

Anonymous:1897:STA

[8147] Anonymous. Scientific truth in art. Scientific American, 77 (11):162, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
162b.pdf.

Anonymous:1897:BAMb

[8148] Anonymous. The British Association Meeting — II. Scientific American, 77(11):162–163, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
162c.pdf.

Anonymous:1897:CP

[8149] Anonymous. A Congress of Physicians. Scientific American, 77 (11):163, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
163.pdf.

Anonymous:1897:ACP

[8150] Anonymous. An apparatus for cleaning pipes. Scientific American, 77(11):164, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
164b.pdf.

Anonymous:1897:CPM

[8151] Anonymous. A concentrator for places mining or stamp mill products. Scientific American, 77(11):164, September 11, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n11/pdf/scientificamerican09111897-164a.pdf.

Anonymous:1897:ILS

[8152] Anonymous. An improved lawn sprinkler. Scientific American, 77 (11):164, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
164.pdf.

Anonymous:1897:NRBc

[8153] Anonymous. A non-refillable bottle. Scientific American, 77 (11):164, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
164c.pdf.

Anonymous:1897:CBT

[8154] Anonymous. The corrosion of a boiler tube. Scientific American, 77(11):168, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
168.pdf.

Anonymous:1897:RMCc

[8155] Anonymous. Railway mail catcher and delivery devices. Scientific American, 77(11):168–170, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
168a.pdf.

Anonymous:1897:ARSc

[8156] Anonymous. “Arch Rock,” San Francisco Bay. Scientific American, 77(11):171–172, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
171.pdf.

Anonymous:1897:RPIbk

[8157] Anonymous. Recently patented inventions. Scientific American, 77 (11):172–173, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
172.pdf.

Anonymous:1897:BNBbh

[8158] Anonymous. Books: New books, etc. Scientific American, 77 (11):173, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
173.pdf.

Anonymous:1897:BPbk

[8159] Anonymous. Business and personal. Scientific American, 77 (11):173, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
173a.pdf.

Anonymous:1897:Ibe

[8160] Anonymous. To inventors. Scientific American, 77(11):173, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n11/pdf/scientificamerican09111897-173c.pdf.

Anonymous:1897:IIbk

[8161] Anonymous. Index of inventions. Scientific American, 77(11): 173, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
173d.pdf.

Anonymous:1897:NQbk

[8162] Anonymous. Notes and queries. Scientific American, 77(11): 173, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
173b.pdf.

Anonymous:1897:PNPbi

[8163] Anonymous. Patent notes: Patents. Scientific American, 77 (11):175, September 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n11/pdf/scientificamerican09111897-
175.pdf.

Anonymous:1897:OBW

[8164] Anonymous. Our big war ships in Fleet practice. Scientific American, 77(12):178, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
178.pdf.

Anonymous:1897:SFE

[8165] Anonymous. A School of Fire Extinguishment. Scientific American, 77(12):178, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
178a.pdf.

Anonymous:1897:RWC

[8166] Anonymous. Rapid work in the construction of New York electric street railways. Scientific American, 77(12):178–179, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n12/pdf/scientificamerican09181897-178b.pdf.

Anonymous:1897:CBa

[8167] Anonymous. The chainless bicycle. Scientific American, 77(12): 179, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
179.pdf.

Anonymous:1897:HGS

[8168] Anonymous. How to go to sleep. Scientific American, 77(12): 179, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
179b.pdf.

Anonymous:1897:MEM

[8169] Anonymous. The mental evolution of man. Scientific American, 77(12):179, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
179a.pdf.

Anonymous:1897:ONB

[8170] Anonymous. Our new book on magic. Scientific American, 77 (12):179, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
179c.pdf.

Anonymous:1897:TO

[8171] Anonymous. Tripe and onions. Scientific American, 77(12): 179, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
179d.pdf.

Anonymous:1897:AAP

[8172] Anonymous. Across Alaskan passes. Scientific American, 77 (12):180, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
180a.pdf.

Anonymous:1897:NBAb

[8173] Anonymous. A novel burglar alarm. Scientific American, 77 (12):180, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
180.pdf.

Anonymous:1897:ABSa

[8174] Anonymous. An ankle brace for skaters. Scientific American, 77 (12):181, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
181a.pdf.

Anonymous:1897:SRS

[8175] Anonymous. The Savage rifle — smokeless powder and expanding bullets. Scientific American, 77(12):181, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n12/pdf/scientificamerican09181897-181.pdf.

Anonymous:1897:NASb

[8176] Anonymous. A nicotine absorbing, smoke cooling pipe. Scientific American, 77(12):181–182, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
181b.pdf.

Anonymous:1897:ILa

[8177] Anonymous. An improved lathe. Scientific American, 77(12): 182, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
182.pdf.

Anonymous:1897:Cn

[8178] Anonymous. Correspondence. Scientific American, 77(12):183, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n12/pdf/scientificamerican09181897-183.pdf.

Anonymous:1897:CPB

[8179] Anonymous. Completion of part of the Boston Subway. Scientific American, 77(12):184–186, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
184.pdf.

Anonymous:1897:M

[8180] Anonymous. “Metempsychosis”. Scientific American, 77(12):187– 188, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
187.pdf.

Anonymous:1897:RPIbl

[8181] Anonymous. Recently patented inventions. Scientific American, 77 (12):188, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
188.pdf.

Anonymous:1897:BNBbi

[8182] Anonymous. Books: New books, etc. Scientific American, 77(12): 188–189, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
188a.pdf.

Anonymous:1897:BPbl

[8183] Anonymous. Business and personal. Scientific American, 77 (12):189, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
189.pdf.

Anonymous:1897:Ibf

[8184] Anonymous. To inventors. Scientific American, 77(12):189, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n12/pdf/scientificamerican09181897-189b.pdf.

Anonymous:1897:IIbl

[8185] Anonymous. Index of inventions. Scientific American, 77(12): 189, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
189c.pdf.

Anonymous:1897:NQbl

[8186] Anonymous. Notes and queries. Scientific American, 77(12): 189, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
189a.pdf.

Anonymous:1897:PNPbj

[8187] Anonymous. Patent notes: Patents. Scientific American, 77 (12):191, September 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n12/pdf/scientificamerican09181897-
191.pdf.

Anonymous:1897:NPH

[8188] Anonymous. A notable pair of horizontal turbines. Scientific American, 77(13):193, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
193.pdf.

Anonymous:1897:DNYa

[8189] Anonymous. Delay of New York rapid transit. Scientific American, 77(13):194, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
194.pdf.

Anonymous:1897:SME

[8190] Anonymous. The safety of the modern elevator. Scientific American, 77(13):194, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
194a.pdf.

Anonymous:1897:WSP

[8191] Anonymous. Wheat as a source of prosperity. Scientific American, 77 (13):194–195, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
194b.pdf.

Anonymous:1897:OAI

[8192] Anonymous. Origin of the American Indian. Scientific American, 77(13):195, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
195b.pdf.

Anonymous:1897:OLE

[8193] Anonymous. Our library of electrical books. Scientific American, 77(13):195, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
195.pdf.

Benjamin:1897:LHP

[8194] Marcus Benjamin. Leland O. Howard, Ph.D., Permanent Secretary American Association of Science. Scientific American, 77 (13):195, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
195a.pdf.

Anonymous:1897:AHD

[8195] Anonymous. An animal holding device. Scientific American, 77 (13):196, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
196c.pdf.

Anonymous:1897:IGC

[8196] Anonymous. Improved glass cutting appliances. Scientific American, 77(13):196, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
196b.pdf.

Anonymous:1897:RGB

[8197] Anonymous. A rocking grate bar and shaker. Scientific American, 77(13):196, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
196a.pdf.

Anonymous:1897:SCM

[8198] Anonymous. A spring-cushioned metallic railroad tie. Scientific American, 77(13):196, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
196.pdf.

Anonymous:1897:ECL

[8199] Anonymous. Electric cabs in London. Scientific American, 77 (13):197, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
197.pdf.

Anonymous:1897:FET

[8200] Anonymous. Fall of an elevator at the Tract Society Building, New York. Scientific American, 77(13):199, September 25, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n13/pdf/scientificamerican09251897-199.pdf.

Anonymous:1897:HID

[8201] Anonymous. The Hemet Irrigating Dam, Southern California. Scientific American, 77(13):200–201, September 25, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n13/pdf/scientificamerican09251897-200.pdf.

Anonymous:1897:FNI

[8202] Anonymous. A few new inventions. Scientific American, 77(13): 201–202, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
201.pdf.

Anonymous:1897:LSM

[8203] Anonymous. A large static machine. Scientific American, 77 (13):203, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
203.pdf.

Anonymous:1897:KS

[8204] Anonymous. The King of Siam. Scientific American, 77(13): 203–204, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
203a.pdf.

Anonymous:1897:RPIbm

[8205] Anonymous. Recently patented inventions. Scientific American, 77 (13):204–205, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
204.pdf.

Anonymous:1897:BNBbj

[8206] Anonymous. Books: New books, etc. Scientific American, 77 (13):205, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
205.pdf.

Anonymous:1897:BPbm

[8207] Anonymous. Business and personal. Scientific American, 77 (13):205, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
205a.pdf.

Anonymous:1897:IIbm

[8208] Anonymous. Index of inventions. Scientific American, 77(13): 205, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
205c.pdf.

Anonymous:1897:NQbm

[8209] Anonymous. Notes and queries. Scientific American, 77(13): 205, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
205b.pdf.

Anonymous:1897:PNPbk

[8210] Anonymous. Patent notes: Patents. Scientific American, 77 (13):206, September 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n13/pdf/scientificamerican09251897-
206.pdf.

Anonymous:1897:PAF

[8211] Anonymous. A patent attorney firm to be disbarred. Scientific American, 77(14):210, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
210.pdf.

Anonymous:1897:PAO

[8212] Anonymous. Proposed amendments to our patent laws. Scientific American, 77(14):210–211, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
210a.pdf.

Hovey:1897:SIGa

[8213] Horace C. Hovey. The Seventh International Geological Congress. Scientific American, 77(14):211–212, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
211.pdf.

Anonymous:1897:IRR

[8214] Anonymous. Improved rails and rail fastenings. Scientific American, 77(14):212, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
212.pdf.

Anonymous:1897:MPT

[8215] Anonymous. Mechanical power transmissions. Scientific American, 77(14):212, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
212a.pdf.

Anonymous:1897:VAI

[8216] Anonymous. A vehicle axle improvement. Scientific American, 77(14):212, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
212b.pdf.

Anonymous:1897:EBM

[8217] Anonymous. The engines of the battleship Massachusetts. Scientific American, 77(14):213–214, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
213.pdf.

Anonymous:1897:ISH

[8218] Anonymous. Improved sash holding and removing devices. Scientific American, 77(14):214, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
214.pdf.

Anonymous:1897:EAT

[8219] Anonymous. The elevator accident at the Tract Society Building, New York. Scientific American, 77(14):215, October 2, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n14/pdf/scientificamerican10021897-215a.pdf.

Brooks:1897:HO

[8220] William R. Brooks. The heavens for October. Scientific American, 77(14):215, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
215.pdf.

Anonymous:1897:NDM

[8221] Anonymous. A new development in machinery for asphalt paving. Scientific American, 77(14):216–217, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
216.pdf.

Anonymous:1897:VR

[8222] Anonymous. Victoria Regia. Scientific American, 77(14):217, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n14/pdf/scientificamerican10021897-217.pdf.

Holder:1897:THB

[8223] Charles Frederick Holder. Tame humming birds. Scientific American, 77(14):217–218, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
217a.pdf.

Anonymous:1897:HVP

[8224] Anonymous. The House of the Vettii at Pompeii. Scientific American, 77(14):219, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
219.pdf.

Anonymous:1897:RPIbn

[8225] Anonymous. Recently patented inventions. Scientific American, 77(14):220, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n14/pdf/scientificamerican10021897-
220.pdf.

Anonymous:1897:BNBbk

[8226] Anonymous. Books: New books, etc. Scientific American, 77(14):221, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n14/pdf/scientificamerican10021897-221b.pdf.

Anonymous:1897:BPbn

[8227] Anonymous. Business and personal. Scientific American, 77(14):221, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n14/pdf/scientificamerican10021897-221.pdf.

Anonymous:1897:Ibg

[8228] Anonymous. To inventors. Scientific American, 77(14):221, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n14/pdf/scientificamerican10021897-221c.pdf.

Anonymous:1897:IIbn

[8229] Anonymous. Index of inventions. Scientific American, 77(14):221, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n14/pdf/scientificamerican10021897-221d.pdf.

Anonymous:1897:NQbn

[8230] Anonymous. Notes and queries. Scientific American, 77(14):221, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n14/pdf/scientificamerican10021897-221a.pdf.

Anonymous:1897:PNPbl

[8231] Anonymous. Patent notes: Patents. Scientific American, 77(14):222, October 2, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n14/pdf/scientificamerican10021897-222.pdf.

Anonymous:1897:WD

[8232] Anonymous. The Wedderburn disbarment. Scientific American, 77 (15):226–227, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
226.pdf.

Anonymous:1897:CCa

[8233] Anonymous. Climate and crime. Scientific American, 77(15):227, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-227c.pdf.

Anonymous:1897:ICU

[8234] Anonymous. Impending changes in United States Patent Law. Scientific American, 77(15):227, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
227.pdf.

Anonymous:1897:NDD

[8235] Anonymous. The Naval Dry Dock Problem. Scientific American, 77(15):227, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
227b.pdf.

Anonymous:1897:PGT

[8236] Anonymous. Platinum and gold toning. Scientific American, 77(15):227, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
227d.pdf.

Anonymous:1897:PPH

[8237] Anonymous. Preservation of the Palisades of the Hudson. Scientific American, 77(15):227, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
227a.pdf.

Anonymous:1897:BSF

[8238] Anonymous. A bicycle support and foot rest. Scientific American, 77(15):228, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
228a.pdf.

Anonymous:1897:BT

[8239] Anonymous. Burglars’ tools. Scientific American, 77(15):228, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-228.pdf.

Anonymous:1897:PCb

[8240] Anonymous. Pneumatic candlestick. Scientific American, 77(15):228, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-228b.pdf.

Anonymous:1897:LGA

[8241] Anonymous. The latest greyhound of the Atlantic. Scientific American, 77(15):229–230, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
229.pdf.

Anonymous:1897:PEDb

[8242] Anonymous. A practical electric dark Room lamp. Scientific American, 77(15):230–231, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
230.pdf.

Anonymous:1897:PPS

[8243] Anonymous. Portable petroleum stove. Scientific American, 77(15):232, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-232.pdf.

Anonymous:1897:USB

[8244] Anonymous. Use of the sand blast for cleaning structural ironwork. Scientific American, 77(15):232, October 9, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n15/pdf/scientificamerican10091897-232a.pdf.

Anonymous:1897:HELb

[8245] Anonymous. The Hartford Electric Light and Power Plant. Scientific American, 77(15):233–234, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
233.pdf.

Anonymous:1897:WMPb

[8246] Anonymous. The William I Monument at the Porta Westphalica. Scientific American, 77(15):235, October 9, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n15/pdf/scientificamerican10091897-235.pdf.

Anonymous:1897:BNBbl

[8247] Anonymous. Books: New books, etc. Scientific American, 77(15):236, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-236a.pdf.

Anonymous:1897:BPbo

[8248] Anonymous. Business and personal. Scientific American, 77(15):236, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-236b.pdf.

Anonymous:1897:Ibh

[8249] Anonymous. To inventors. Scientific American, 77(15):236, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-236d.pdf.

Anonymous:1897:NQbo

[8250] Anonymous. Notes and queries. Scientific American, 77(15):236, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-236c.pdf.

Anonymous:1897:RPIbo

[8251] Anonymous. Recently patented inventions. Scientific American, 77(15):236, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n15/pdf/scientificamerican10091897-
236.pdf.

Anonymous:1897:IIbo

[8252] Anonymous. Index of inventions. Scientific American, 77(15):237, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-237.pdf.

Anonymous:1897:PNPbm

[8253] Anonymous. Patent notes: Patents. Scientific American, 77(15):238, October 9, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n15/pdf/scientificamerican10091897-238.pdf.

Anonymous:1897:RPE

[8254] Anonymous. The return of the Peary Expedition. Scientific American, 77(16):241, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
241.pdf.

Anonymous:1897:ARE

[8255] Anonymous. Another runaway elevator. Scientific American, 77 (16):242, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
242.pdf.

Anonymous:1897:RTSb

[8256] Anonymous. Rapid transit scheme approved by the Park board. Scientific American, 77(16):242, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
242a.pdf.

Anonymous:1897:SFS

[8257] Anonymous. The suppression of a fraudulent system of patent practice. Scientific American, 77(16):242–243, October 16, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n16/pdf/scientificamerican10161897-242b.pdf.

Anonymous:1897:AIFb

[8258] Anonymous. The American institute fair, New York. Scientific American, 77(16):243, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
243b.pdf.

Anonymous:1897:NER

[8259] Anonymous. New eastward record for the liner Kaiser Wilhelm. Scientific American, 77(16):243, October 16, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n16/pdf/scientificamerican10161897-243.pdf.

Anonymous:1897:OPP

[8260] Anonymous. Opening of the pneumatic postal tube service in New York City. Scientific American, 77(16):243, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n16/pdf/scientificamerican10161897-243a.pdf.

Anonymous:1897:LTS

[8261] Anonymous. The Lowell Textile School. Scientific American, 77 (16):244–245, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
244.pdf.

Anonymous:1897:IBT

[8262] Anonymous. An improved boiler tube cleaner. Scientific American, 77(16):245, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
245.pdf.

Anonymous:1897:NHDc

[8263] Anonymous. A new hand dynamo. Scientific American, 77(16): 245–246, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
245a.pdf.

Anonymous:1897:RJS

[8264] Anonymous. A rail joint support and bearing plate. Scientific American, 77(16):246–247, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
246.pdf.

Anonymous:1897:FDD

[8265] Anonymous. A Floating Dry Dock for Havana. Scientific American, 77(16):247–248, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
247.pdf.

Anonymous:1897:PA

[8266] Anonymous. The parasites of ants. Scientific American, 77(16):251, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n16/pdf/scientificamerican10161897-251.pdf.

Anonymous:1897:RPIbp

[8267] Anonymous. Recently patented inventions. Scientific American, 77(16):252, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
252.pdf.

Anonymous:1897:BNBbm

[8268] Anonymous. Books: New books, etc. Scientific American, 77 (16):252–253, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
252a.pdf.

Anonymous:1897:BPbp

[8269] Anonymous. Business and personal. Scientific American, 77 (16):253, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
253.pdf.

Anonymous:1897:IIbp

[8270] Anonymous. Index of inventions. Scientific American, 77(16):253, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n16/pdf/scientificamerican10161897-253b.pdf.

Anonymous:1897:NQbp

[8271] Anonymous. Notes and queries. Scientific American, 77(16):253, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n16/pdf/scientificamerican10161897-253a.pdf.

Anonymous:1897:PNPbn

[8272] Anonymous. Patent notes: Patents. Scientific American, 77 (16):255, October 16, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n16/pdf/scientificamerican10161897-
255.pdf.

Anonymous:1897:LEb

[8273] Anonymous. The licensing of engineers. Scientific American, 77 (17):258, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
258.pdf.

Anonymous:1897:PAN

[8274] Anonymous. Prison Association of New York Exhibition. Scientific American, 77(17):258, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
258a.pdf.

Anonymous:1897:STHb

[8275] Anonymous. A seventy thousand horse power central station. Scientific American, 77(17):258–259, October 23, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n17/pdf/scientificamerican10231897-258b.pdf.

Anonymous:1897:LK

[8276] Anonymous. Lord Kelvin. Scientific American, 77(17):259, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n17/pdf/scientificamerican10231897-259.pdf.

Anonymous:1897:IRIb

[8277] Anonymous. Inventions: Recent inventions. Scientific American, 77(17):260, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
260.pdf.

Sinclair:1897:LFP

[8278] A. J. Sinclair. The largest and fastest paddle steamer in Great Britain. Scientific American, 77(17):261–262, October 23, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n17/pdf/scientificamerican10231897-261.pdf.

Anonymous:1897:GMS

[8279] Anonymous. A gas main stopper. Scientific American, 77(17):262, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n17/pdf/scientificamerican10231897-262.pdf.

Anonymous:1897:CBb

[8280] Anonymous. The chainless bicycle. Scientific American, 77(17): 262–263, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
262a.pdf.

Anonymous:1897:NHC

[8281] Anonymous. A novel hand camera. Scientific American, 77(17):264, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n17/pdf/scientificamerican10231897-264.pdf.

Anonymous:1897:SPTa

[8282] Anonymous. The steel pipe and tube industry. Scientific American, 77(17):264–266, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
264a.pdf.

Anonymous:1897:Co

[8283] Anonymous. Correspondence. Scientific American, 77(17):267, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n17/pdf/scientificamerican10231897-267.pdf.

Holder:1897:CCL

[8284] C. F. Holder. The color changes in lizards. Scientific American, 77 (17):267–268, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
267a.pdf.

Anonymous:1897:BNBbn

[8285] Anonymous. Books: New books, etc. Scientific American, 77 (17):269, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
269a.pdf.

Anonymous:1897:BPbq

[8286] Anonymous. Business and personal. Scientific American, 77 (17):269, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
269b.pdf.

Anonymous:1897:IIbq

[8287] Anonymous. Index of inventions. Scientific American, 77(17):269, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n17/pdf/scientificamerican10231897-269d.pdf.

Anonymous:1897:NQbq

[8288] Anonymous. Notes and queries. Scientific American, 77(17):269, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n17/pdf/scientificamerican10231897-269c.pdf.

Anonymous:1897:RPIbq

[8289] Anonymous. Recently patented inventions. Scientific American, 77(17):269, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
269.pdf.

Anonymous:1897:PNPbo

[8290] Anonymous. Patent notes: Patents. Scientific American, 77 (17):270, October 23, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n17/pdf/scientificamerican10231897-
270.pdf.

Anonymous:1897:ACLb

[8291] Anonymous. The automatic coupler law to be enforced. Scientific American, 77(18):274, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
274.pdf.

Anonymous:1897:RBP

[8292] Anonymous. The roller boat problem. Scientific American, 77 (18):274–275, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
274a.pdf.

Anonymous:1897:CSPb

[8293] Anonymous. Changes in Swedish patent and trade mark laws. Scientific American, 77(18):275, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
275b.pdf.

Anonymous:1897:ME

[8294] Anonymous. Multiplication of explosives. Scientific American, 77(18):275, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
275c.pdf.

Anonymous:1897:NAL

[8295] Anonymous. New Argentine law in regard to the sale of medicines. Scientific American, 77(18):275, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
275a.pdf.

Anonymous:1897:OGC

[8296] Anonymous. Our great cotton crop. Scientific American, 77 (18):275, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
275.pdf.

Anonymous:1897:ICa

[8297] Anonymous. An improved concentrator. Scientific American, 77 (18):276, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
276.pdf.

Anonymous:1897:MTE

[8298] Anonymous. A machine to turn the ends of wagon axles. Scientific American, 77(18):276, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
276a.pdf.

Anonymous:1897:CCB

[8299] Anonymous. The Columbia chainless bicycle. Scientific American, 77(18):277–278, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
277.pdf.

Anonymous:1897:SRY

[8300] Anonymous. Surveying for a railroad to the Yukon Valley. Scientific American, 77(18):278, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
278.pdf.

Brooks:1897:HN

[8301] William R. Brooks. The heavens for November. Scientific American, 77(18):279, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
279.pdf.

Sinclair:1897:SBF

[8302] A. J. Sinclair. A Scottish built ferry boat for Nova Scotia. Scientific American, 77(18):279–280, October 30, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n18/pdf/scientificamerican10301897-279a.pdf.

Holder:1897:GPR

[8303] C. F. Holder. A great pelican rookery. Scientific American, 77 (18):280–281, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
280.pdf.

Anonymous:1897:AHN

[8304] Anonymous. The Astoria Hotel, New York City. Scientific American, 77(18):281–282, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
281.pdf.

Anonymous:1897:MFb

[8305] Anonymous. “Magic Flowers”. Scientific American, 77(18):283, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n18/pdf/scientificamerican10301897-283.pdf.

Anonymous:1897:BNBbo

[8306] Anonymous. Books: New books, etc. Scientific American, 77 (18):284, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
284a.pdf.

Anonymous:1897:RPIbr

[8307] Anonymous. Recently patented inventions. Scientific American, 77(18):284, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
284.pdf.

Anonymous:1897:BPbr

[8308] Anonymous. Business and personal. Scientific American, 77 (18):285, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
285.pdf.

Anonymous:1897:Ibi

[8309] Anonymous. To inventors. Scientific American, 77(18):285, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n18/pdf/scientificamerican10301897-285b.pdf.

Anonymous:1897:IIbr

[8310] Anonymous. Index of inventions. Scientific American, 77(18):285, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n18/pdf/scientificamerican10301897-285c.pdf.

Anonymous:1897:NQbr

[8311] Anonymous. Notes and queries. Scientific American, 77(18):285, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n18/pdf/scientificamerican10301897-285a.pdf.

Anonymous:1897:PNPbp

[8312] Anonymous. Patent notes: Patents. Scientific American, 77 (18):286, October 30, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n18/pdf/scientificamerican10301897-
286.pdf.

Anonymous:1897:NUS

[8313] Anonymous. The Navies of the United States and Spain — a comparison. Scientific American, 77(19):290, November 6, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n19/pdf/scientificamerican11061897-290.pdf.

Anonymous:1897:DNYb

[8314] Anonymous. The disaster on the New York Central Railroad. Scientific American, 77(19):290–291, November 6, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n19/pdf/scientificamerican11061897-290a.pdf.

Anonymous:1897:AWP

[8315] Anonymous. The age of water power. Scientific American, 77 (19):291, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
291c.pdf.

Anonymous:1897:ISEb

[8316] Anonymous. An interesting sociological experiment. Scientific American, 77(19):291, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
291.pdf.

Anonymous:1897:RPIbs

[8317] Anonymous. A record patent issue. Scientific American, 77 (19):291, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
291b.pdf.

Anonymous:1897:UPO

[8318] Anonymous. An urgent Patent Office reform. Scientific American, 77(19):291, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
291a.pdf.

Anonymous:1897:TTL

[8319] Anonymous. A trip through a large bicycle factory. Scientific American, 77(19):292–293, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
292.pdf.

Anonymous:1897:RBR

[8320] Anonymous. A rapid bridge renewal. Scientific American, 77 (19):293–294, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
293.pdf.

Anonymous:1897:NHU

[8321] Anonymous. A net holder for use at fires. Scientific American, 77(19):294, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
294.pdf.

Anonymous:1897:Cq

[8322] Anonymous. Correspondence. Scientific American, 77(19):295, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n19/pdf/scientificamerican11061897-295.pdf.

Anonymous:1897:TWC

[8323] Anonymous. A trolley wagon for country roads. Scientific American, 77(19):296, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
296.pdf.

Anonymous:1897:ADNa

[8324] Anonymous. Arrival of Dr. Nansen. Scientific American, 77(19): 297–298, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
297.pdf.

Anonymous:1897:ICC

[8325] Anonymous. An improved car coupling. Scientific American, 77 (19):298, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
298.pdf.

Anonymous:1897:CAB

[8326] Anonymous. The casting of art bronze. Scientific American, 77 (19):299, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
299.pdf.

Anonymous:1897:SRM

[8327] Anonymous. A single rail mountain railway. Scientific American, 77 (19):299–300, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
299a.pdf.

Anonymous:1897:RPIbt

[8328] Anonymous. Recently patented inventions. Scientific American, 77(19):300, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
300.pdf.

Anonymous:1897:BPbs

[8329] Anonymous. Business and personal. Scientific American, 77 (19):301, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
301.pdf.

Anonymous:1897:Ibj

[8330] Anonymous. To inventors. Scientific American, 77(19):301, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n19/pdf/scientificamerican11061897-301b.pdf.

Anonymous:1897:IIbs

[8331] Anonymous. Index of inventions. Scientific American, 77(19): 301, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
301c.pdf.

Anonymous:1897:NQbs

[8332] Anonymous. Notes and queries. Scientific American, 77(19): 301, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
301a.pdf.

Anonymous:1897:PNPbq

[8333] Anonymous. Patent notes: Patents. Scientific American, 77 (19):302, November 6, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n19/pdf/scientificamerican11061897-
302.pdf.

Anonymous:1897:IOA

[8334] Anonymous. Increase in our artillery urgently needed. Scientific American, 77(20):306, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
306.pdf.

Anonymous:1897:RIC

[8335] Anonymous. Report of the Interstate Commerce Commission. Scientific American, 77(20):306–307, November 13, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n20/pdf/scientificamerican11131897-306a.pdf.

Hovey:1897:SIGb

[8336] E. O. Hovey. The Seventh International Geological Congress — The Caucasus Excursion. Scientific American, 77(20):307, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n20/pdf/scientificamerican11131897-307.pdf.

Anonymous:1897:AFB

[8337] Anonymous. An adjustable fire box partition. Scientific American, 77(20):308, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
308b.pdf.

Anonymous:1897:AGL

[8338] Anonymous. An acetylene gas lamp for bicycles. Scientific American, 77(20):308, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
308a.pdf.

Anonymous:1897:IDb

[8339] Anonymous. An improved drill. Scientific American, 77(20): 308, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
308.pdf.

Schenck:1897:SEH

[8340] Archibald A. Schenck. The sliding embankments of the Hudson River Shore Lines. Scientific American, 77(20):309–310, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n20/pdf/scientificamerican11131897-309.pdf.

Anonymous:1897:LRB

[8341] Anonymous. The latest roller boat. Scientific American, 77(20): 310–311, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
310.pdf.

Anonymous:1897:Cr

[8342] Anonymous. Correspondence. Scientific American, 77(20):311, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n20/pdf/scientificamerican11131897-311.pdf.

Anonymous:1897:SPTb

[8343] Anonymous. The steel pipe and tube industry. Scientific American, 77(20):312–313, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
312.pdf.

Anonymous:1897:MLM

[8344] Anonymous. A motor lawn mower. Scientific American, 77 (20):314, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
314.pdf.

Anonymous:1897:ADNb

[8345] Anonymous. Archæological discoveries in Northern Africa. Scientific American, 77(20):315, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
315.pdf.

Anonymous:1897:BNBbp

[8346] Anonymous. Books: New books, etc. Scientific American, 77 (20):316, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
316a.pdf.

Anonymous:1897:BPbt

[8347] Anonymous. Business and personal. Scientific American, 77 (20):316, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
316b.pdf.

Anonymous:1897:Ibk

[8348] Anonymous. To inventors. Scientific American, 77(20):316, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n20/pdf/scientificamerican11131897-316d.pdf.

Anonymous:1897:IIbt

[8349] Anonymous. Index of inventions. Scientific American, 77(20): 316, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
316e.pdf.

Anonymous:1897:NQbt

[8350] Anonymous. Notes and queries. Scientific American, 77(20): 316, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
316c.pdf.

Anonymous:1897:RPIbu

[8351] Anonymous. Recently patented inventions. Scientific American, 77(20):316, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
316.pdf.

Anonymous:1897:PNPbr

[8352] Anonymous. Patent notes: Patent. Scientific American, 77(20): 317, November 13, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n20/pdf/scientificamerican11131897-
317.pdf.

Anonymous:1897:BSIb

[8353] Anonymous. Is the beet sugar industry desirable? Scientific American, 77(21):322, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
322.pdf.

Anonymous:1897:AES

[8354] Anonymous. The application of electricity to steam railroads. Scientific American, 77(21):322–323, November 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n21/pdf/scientificamerican11201897-322a.pdf.

Anonymous:1897:JS

[8355] Anonymous. James E. Simpson. Scientific American, 77(21): 323, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
323b.pdf.

Anonymous:1897:POF

[8356] Anonymous. Post Office fraud order issued against Wedderburn & Company. Scientific American, 77(21):323, November 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n21/pdf/scientificamerican11201897-323a.pdf.

Anonymous:1897:SAD

[8357] Anonymous. Still another dodge to defraud patentees. Scientific American, 77(21):323, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
323.pdf.

Anonymous:1897:ABSb

[8358] Anonymous. The American beet sugar industry. Scientific American, 77(21):323–324, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
323c.pdf.

Anonymous:1897:ICT

[8359] Anonymous. An improved car truck. Scientific American, 77 (21):324, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
324.pdf.

Anonymous:1897:IWM

[8360] Anonymous. An improved washing machine. Scientific American, 77(21):324, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
324b.pdf.

Anonymous:1897:SC

[8361] Anonymous. A simple camera. Scientific American, 77(21):324, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n21/pdf/scientificamerican11201897-324a.pdf.

Anonymous:1897:HGC

[8362] Anonymous. Handling the grain crop at the Port of New York. Scientific American, 77(21):325–326, November 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n21/pdf/scientificamerican11201897-325.pdf.

Anonymous:1897:UAJ

[8363] Anonymous. A universally adjustable Jeweler’s clamp. Scientific American, 77(21):326, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
326.pdf.

Anonymous:1897:Cs

[8364] Anonymous. Correspondence. Scientific American, 77(21):327, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n21/pdf/scientificamerican11201897-327.pdf.

Anonymous:1897:GEP

[8365] Anonymous. The great electrical power plant at Rheinfelden on the Rhine. Scientific American, 77(21):329–330, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n21/pdf/scientificamerican11201897-329.pdf.

Anonymous:1897:LDG

[8366] Anonymous. Light draught gunboats for the Nile Expedition. Scientific American, 77(21):330–331, November 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n21/pdf/scientificamerican11201897-330.pdf.

Anonymous:1897:MFL

[8367] Anonymous. Mountain freight locomotive for the Mexican Central Railway. Scientific American, 77(21):331, November 20, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n21/pdf/scientificamerican11201897-331.pdf.

Aldrich:1897:SVT

[8368] Cyrus Aldrich. Science vs. theory. Scientific American, 77(21): 332, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
332b.pdf.

Anonymous:1897:BNBbq

[8369] Anonymous. Books: New books, etc. Scientific American, 77 (21):332, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
332a.pdf.

Anonymous:1897:BPbu

[8370] Anonymous. Business and personal. Scientific American, 77 (21):332, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
332c.pdf.

Anonymous:1897:RPIbv

[8371] Anonymous. Recently patented inventions. Scientific American, 77(21):332, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
332.pdf.

Anonymous:1897:Ibm

[8372] Anonymous. To inventors. Scientific American, 77(21):333, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n21/pdf/scientificamerican11201897-333a.pdf.

Anonymous:1897:IIbu

[8373] Anonymous. Index of inventions. Scientific American, 77(21): 333, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
333b.pdf.

Anonymous:1897:NQbu

[8374] Anonymous. Notes and queries. Scientific American, 77(21): 333, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
333.pdf.

Anonymous:1897:PNPbs

[8375] Anonymous. Patent notes: Patents. Scientific American, 77 (21):334, November 20, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n21/pdf/scientificamerican11201897-
334.pdf.

Anonymous:1897:LHR

[8376] Anonymous. Lengthening a Hudson River steamboat. Scientific American, 77(22):337, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
337.pdf.

Anonymous:1897:ACR

[8377] Anonymous. Appellate Commission Report favorably on the New York Rapid Transit Scheme. Scientific American, 77(22):338, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n22/pdf/scientificamerican11271897-338.pdf.

Anonymous:1897:QNP

[8378] Anonymous. The quest of the North Pole. Scientific American, 77(22):338, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
338a.pdf.

Anonymous:1897:RSI

[8379] Anonymous. Report of the Secretary of the Interior. Scientific American, 77(22):338–339, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
338b.pdf.

Anonymous:1897:ABSc

[8380] Anonymous. The American beet sugar industry. Scientific American, 77(22):339, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
339.pdf.

Anonymous:1897:CMR

[8381] Anonymous. A cover for mucilage reservoirs. Scientific American, 77(22):340, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
340b.pdf.

Anonymous:1897:IAW

[8382] Anonymous. An improved automatic weir. Scientific American, 77(22):340, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
340a.pdf.

Anonymous:1897:IGB

[8383] Anonymous. An improved gas burner. Scientific American, 77 (22):340, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
340.pdf.

Anonymous:1897:BOW

[8384] Anonymous. Boring oil wells at sea. Scientific American, 77 (22):341, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
341a.pdf.

Anonymous:1897:MSP

[8385] Anonymous. Moulds for soldering pipes. Scientific American, 77 (22):341, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
341.pdf.

Anonymous:1897:Ct

[8386] Anonymous. Correspondence. Scientific American, 77(22):343, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n22/pdf/scientificamerican11271897-343.pdf.

Anonymous:1897:SMC

[8387] Anonymous. Steam motor car for branch line service. Scientific American, 77(22):344, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
344.pdf.

Anonymous:1897:SCC

[8388] Anonymous. The steel cell and central corridor system of prison construction. Scientific American, 77(22):345–346, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n22/pdf/scientificamerican11271897-345.pdf.

Anonymous:1897:GRT

[8389] Anonymous. The giant redwood trees of California. Scientific American, 77(22):347, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
347.pdf.

Anonymous:1897:BPbv

[8390] Anonymous. Business and personal. Scientific American, 77 (22):348, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
348a.pdf.

Anonymous:1897:Ibn

[8391] Anonymous. To inventors. Scientific American, 77(22):348, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n22/pdf/scientificamerican11271897-348c.pdf.

Anonymous:1897:IIbv

[8392] Anonymous. Index of inventions. Scientific American, 77(22): 348, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
348d.pdf.

Anonymous:1897:NQbv

[8393] Anonymous. Notes and queries. Scientific American, 77(22): 348, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
348b.pdf.

Anonymous:1897:RPIbw

[8394] Anonymous. Recently patented inventions. Scientific American, 77(22):348, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
348.pdf.

Anonymous:1897:PNPbt

[8395] Anonymous. Patent notes: Patents. Scientific American, 77 (22):351, November 27, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n22/pdf/scientificamerican11271897-
351.pdf.

Anonymous:1897:AWO

[8396] Anonymous. The annual war office report. Scientific American, 77(23):354, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
354a.pdf.

Anonymous:1897:PSC

[8397] Anonymous. The proposed ship canal from the Lakes to the sea. Scientific American, 77(23):354, December 4, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n23/pdf/scientificamerican12041897-354.pdf.

Anonymous:1897:NPO

[8398] Anonymous. New Patent Office rules. Scientific American, 77 (23):354–355, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
354b.pdf.

Anonymous:1897:GFL

[8399] Anonymous. A great fire in London. Scientific American, 77 (23):355, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
355.pdf.

Brooks:1897:HD

[8400] William R. Brooks. The heavens for December. Scientific American, 77(23):355, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
355a.pdf.

Anonymous:1897:IAM

[8401] Anonymous. An improved air motor. Scientific American, 77 (23):356, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
356a.pdf.

Anonymous:1897:NTB

[8402] Anonymous. A new type of bicycle. Scientific American, 77(23):356, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n23/pdf/scientificamerican12041897-356.pdf.

Anonymous:1897:SHS

[8403] Anonymous. A self-heating soldering iron. Scientific American, 77(23):356, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
356b.pdf.

Anonymous:1897:LRE

[8404] Anonymous. The Lachine Rapids Electric Light and Power Plant. Scientific American, 77(23):357–358, December 4, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n23/pdf/scientificamerican12041897-357.pdf.

Anonymous:1897:ONH

[8405] Anonymous. Opening of the new home of the American Society of Civil Engineers. Scientific American, 77(23):358, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n23/pdf/scientificamerican12041897-358.pdf.

Anonymous:1897:Cu

[8406] Anonymous. Correspondence. Scientific American, 77(23):359–360, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n23/pdf/scientificamerican12041897-359.pdf.

Anonymous:1897:SFD

[8407] Anonymous. A school fire drill. Scientific American, 77(23): 360–361, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
360.pdf.

Anonymous:1897:FCW

[8408] Anonymous. The feet of Chinese women. Scientific American, 77(23):361, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
361.pdf.

Anonymous:1897:TK

[8409] Anonymous. Tusayan Katcinas. Scientific American, 77(23):362– 363, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
362.pdf.

Anonymous:1897:RPIbx

[8410] Anonymous. Recently patented inventions. Scientific American, 77 (23):363–364, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
363.pdf.

Anonymous:1897:BNBbr

[8411] Anonymous. Books: New books, etc. Scientific American, 77 (23):364, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
364.pdf.

Anonymous:1897:BPbw

[8412] Anonymous. Business and personal. Scientific American, 77 (23):364, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
364a.pdf.

Anonymous:1897:IIbw

[8413] Anonymous. Index of inventions. Scientific American, 77(23):364, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n23/pdf/scientificamerican12041897-364c.pdf.

Anonymous:1897:NQbw

[8414] Anonymous. Notes and queries. Scientific American, 77(23):364, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n23/pdf/scientificamerican12041897-364b.pdf.

Anonymous:1897:PNPbu

[8415] Anonymous. Patent notes: Patents. Scientific American, 77 (23):367, December 4, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n23/pdf/scientificamerican12041897-
367.pdf.

Anonymous:1897:CUS

[8416] Anonymous. Changes in United States patent laws. Scientific American, 77(24):370, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
370.pdf.

Anonymous:1897:PGA

[8417] Anonymous. The proposed Government Armor Plant. Scientific American, 77(24):370, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
370a.pdf.

Anonymous:1897:SCO

[8418] Anonymous. Some curious old patents. Scientific American, 77 (24):370–371, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
370b.pdf.

Anonymous:1897:DC

[8419] Anonymous. The dietary of cyclists. Scientific American, 77 (24):371, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
371a.pdf.

Anonymous:1897:IMC

[8420] Anonymous. Involuntary movements as controlled by ideas. Scientific American, 77(24):371, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
371b.pdf.

Walsh:1897:CB

[8421] George E. Walsh. Costly butterflies. Scientific American, 77 (24):371, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
371.pdf.

Anonymous:1897:AET

[8422] Anonymous. An automatic electric time switch. Scientific American, 77(24):372, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
372.pdf.

Anonymous:1897:IDG

[8423] Anonymous. An improved double grinding mill. Scientific American, 77(24):372, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
372a.pdf.

Anonymous:1897:PPT

[8424] Anonymous. A pump for pneumatic tires. Scientific American, 77(24):372, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
372b.pdf.

Anonymous:1897:PMT

[8425] Anonymous. Pneumatic mail tube system, New York City. Scientific American, 77(24):373–375, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
373.pdf.

Anonymous:1897:LPG

[8426] Anonymous. A low priced graphophone. Scientific American, 77 (24):375, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
375.pdf.

Anonymous:1897:TTBb

[8427] Anonymous. Trial trip of a British built torpedo boat destroyer for the Spanish Government. Scientific American, 77(24):376, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n24/pdf/scientificamerican12111897-376a.pdf.

Anonymous:1897:TWB

[8428] Anonymous. A “tug of war” between early steamships. Scientific American, 77(24):376, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
376.pdf.

Anonymous:1897:BGD

[8429] Anonymous. The Buchanan–Gordon diving dress. Scientific American, 77(24):377, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
377.pdf.

Anonymous:1897:IBBd

[8430] Anonymous. An improved bicycle brake. Scientific American, 77 (24):378, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
378.pdf.

Anonymous:1897:G

[8431] Anonymous. “Gone”. Scientific American, 77(24):379, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n24/pdf/scientificamerican12111897-379.pdf.

Anonymous:1897:BPbx

[8432] Anonymous. Business and personal. Scientific American, 77 (24):380, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
380a.pdf.

Anonymous:1897:RPIby

[8433] Anonymous. Recently patented inventions. Scientific American, 77(24):380, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
380.pdf.

Anonymous:1897:NQbx

[8434] Anonymous. Notes and queries. Scientific American, 77(24): 380–381, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
380b.pdf.

Anonymous:1897:BNBbs

[8435] Anonymous. Books: New books, etc. Scientific American, 77 (24):381, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
381.pdf.

Anonymous:1897:Ibo

[8436] Anonymous. To inventors. Scientific American, 77(24):381, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n24/pdf/scientificamerican12111897-381a.pdf.

Anonymous:1897:IIbx

[8437] Anonymous. Index of inventions. Scientific American, 77(24): 381, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
381b.pdf.

Anonymous:1897:PNPbv

[8438] Anonymous. Patent notes: Patents. Scientific American, 77 (24):383, December 11, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n24/pdf/scientificamerican12111897-
383.pdf.

Anonymous:1897:CLCb

[8439] Anonymous. Condition of the laboring classes in Mexico. Scientific American, 77(25):386, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
386.pdf.

Anonymous:1897:SST

[8440] Anonymous. The supremacy of the steam turbine. Scientific American, 77(25):386–387, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
386a.pdf.

Anonymous:1897:EHG

[8441] Anonymous. An excellent holiday gift. Scientific American, 77 (25):387, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
387d.pdf.

Anonymous:1897:FRK

[8442] Anonymous. Further records for the Kaiser Wilhelm Der Grosse. Scientific American, 77(25):387, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
387c.pdf.

Anonymous:1897:LPE

[8443] Anonymous. Lieutenant Peary in England. Scientific American, 77(25):387, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
387a.pdf.

Anonymous:1897:SHG

[8444] Anonymous. Sinking of the Havana Graving Dock. Scientific American, 77(25):387, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
387e.pdf.

Anonymous:1897:WIb

[8445] Anonymous. Warning to inventors. Scientific American, 77(25): 387, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
387.pdf.

Baxter:1897:OUA

[8446] William Baxter, Jr. One of the useful applications of the storage battery. Scientific American, 77(25):387, December 18, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n25/pdf/scientificamerican12181897-387b.pdf.

Anonymous:1897:APM

[8447] Anonymous. An apparatus for projecting moving pictures. Scientific American, 77(25):388, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
388b.pdf.

Anonymous:1897:IGH

[8448] Anonymous. An improved gate hinge. Scientific American, 77 (25):388, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
388a.pdf.

Anonymous:1897:RST

[8449] Anonymous. The Rem-Sho typewriter. Scientific American, 77 (25):388, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
388.pdf.

Anonymous:1897:EAG

[8450] Anonymous. Experimental acetylene gas buoy for New York Harbor. Scientific American, 77(25):389–390, December 18, 1897. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v77/
n25/pdf/scientificamerican12181897-389.pdf.

Anonymous:1897:Cv

[8451] Anonymous. Correspondence. Scientific American, 77(25):391, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n25/pdf/scientificamerican12181897-391.pdf.

Anonymous:1897:SPTc

[8452] Anonymous. The steel pipe and tube industry. Scientific American, 77(25):392–394, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
392.pdf.

Carden:1897:TKT

[8453] G. L. Carden. Thirty knot torpedo boat catcher “Bailey”. Scientific American, 77(25):394–395, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
394.pdf.

Anonymous:1897:RPIbz

[8454] Anonymous. Recently patented inventions. Scientific American, 77 (25):395–396, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
395.pdf.

Anonymous:1897:BNBbt

[8455] Anonymous. Books: New books, etc. Scientific American, 77 (25):396, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
396b.pdf.

Anonymous:1897:BPby

[8456] Anonymous. Business and personal. Scientific American, 77 (25):396, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
396.pdf.

Anonymous:1897:Ibp

[8457] Anonymous. To inventors. Scientific American, 77(25):396, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n25/pdf/scientificamerican12181897-396c.pdf.

Anonymous:1897:IIby

[8458] Anonymous. Index of inventions. Scientific American, 77(25): 396, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
396d.pdf.

Anonymous:1897:NQby

[8459] Anonymous. Notes and queries. Scientific American, 77(25): 396, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
396a.pdf.

Anonymous:1897:PNPbw

[8460] Anonymous. Patent notes: Patents. Scientific American, 77 (25):398, December 18, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n25/pdf/scientificamerican12181897-
398.pdf.

Anonymous:1897:FSB

[8461] Anonymous. Fireproof steel and brick grain elevator. Scientific American, 77(26):401, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
401.pdf.

Anonymous:1897:ED

[8462] Anonymous. Economy in details. Scientific American, 77(26): 402, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
402d.pdf.

Anonymous:1897:FEL

[8463] Anonymous. Famous engineering landmark to be removed. Scientific American, 77(26):402, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
402.pdf.

Anonymous:1897:PHS

[8464] Anonymous. The possibilities of high speed electric traction. Scientific American, 77(26):402, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
402a.pdf.

Anonymous:1897:PSAc

[8465] Anonymous. Probable solution of the armor plate question. Scientific American, 77(26):402, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
402b.pdf.

Anonymous:1897:YPK

[8466] Anonymous. A year of plenty in Kansas. Scientific American, 77 (26):402, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
402c.pdf.

Anonymous:1897:LHSb

[8467] Anonymous. The limits of human speed and endurance. Scientific American, 77(26):402–403, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
402e.pdf.

Anonymous:1897:CIa

[8468] Anonymous. Copyright injustice. Scientific American, 77(26): 403, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
403.pdf.

Anonymous:1897:L

[8469] Anonymous. The lasso. Scientific American, 77(26):403, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n26/pdf/scientificamerican12251897-403b.pdf.

Anonymous:1897:LPR

[8470] Anonymous. Lieut Peary receives a ship. Scientific American, 77 (26):403, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
403a.pdf.

Woodbury:1897:DB

[8471] C. J. Woodbury. Dustless buildings. Scientific American, 77 (26):403, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
403c.pdf.

Anonymous:1897:NMW

[8472] Anonymous. A novel monkey wrench. Scientific American, 77 (26):404, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
404.pdf.

Anonymous:1897:PPC

[8473] Anonymous. The Pierce pneumatic cushion bicycle frame. Scientific American, 77(26):404, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
404a.pdf.

Anonymous:1897:SHb

[8474] Anonymous. A stovepipe holder. Scientific American, 77(26): 404–405, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
404b.pdf.

Anonymous:1897:BAB

[8475] Anonymous. The Beelin aluminum balloon. Scientific American, 77 (26):405–406, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
405.pdf.

Anonymous:1897:Cw

[8476] Anonymous. Correspondence. Scientific American, 77(26):407, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n26/pdf/scientificamerican12251897-407.pdf.

Anonymous:1897:NMM

[8477] Anonymous. New method of manufacturing heavy ordnance. Scientific American, 77(26):408, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
408.pdf.

Anonymous:1897:RLDb

[8478] Anonymous. Repairing the leak at Dry Dock No. 3, Brooklyn Navy Yard. Scientific American, 77(26):408–409, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n26/pdf/scientificamerican12251897-408a.pdf.

Anonymous:1897:RPIca

[8479] Anonymous. Recently patented inventions. Scientific American, 77(26):410, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
410.pdf.

Anonymous:1897:BNBbu

[8480] Anonymous. Books: New books, etc. Scientific American, 77(26): 410–411, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
410a.pdf.

Anonymous:1897:BPbz

[8481] Anonymous. Business and personal. Scientific American, 77 (26):411, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
411.pdf.

Anonymous:1897:IIbz

[8482] Anonymous. Index of inventions. Scientific American, 77(26): 411, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
411b.pdf.

Anonymous:1897:NQbz

[8483] Anonymous. Notes and queries. Scientific American, 77(26): 411, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
411a.pdf.

Anonymous:1897:Ibq

[8484] Anonymous. Index. Scientific American, 77(26):413–414, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v77/n26/pdf/scientificamerican12251897-413.pdf.

Anonymous:1897:PNPbx

[8485] Anonymous. Patent notes: Patents. Scientific American, 77 (26):415, December 25, 1897. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v77/n26/pdf/scientificamerican12251897-
415.pdf.

Anonymous:1898:RYa

[8486] Anonymous. A retrospect of the year 1897. Scientific American, 78(1):4–5, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n1/pdf/scientificamerican01011898-
4.pdf.

Anonymous:1898:CTa

[8487] Anonymous. The Camphor tree. Scientific American, 78(1):5, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-5a.pdf.

Anonymous:1898:FFD

[8488] Anonymous. The “Fire Fiend” as a dyer. Scientific American, 78(1):5, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-5b.pdf.

Anonymous:1898:NYR

[8489] Anonymous. New York rapid transit report approved by the appellate justices. Scientific American, 78(1):5, January 1, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n1/pdf/scientificamerican01011898-5.pdf.

Anonymous:1898:MMF

[8490] Anonymous. The Montauk multiphase fire cable. Scientific American, 78(1):6, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n1/pdf/scientificamerican01011898-
6.pdf.

Anonymous:1898:FEB

[8491] Anonymous. Fire engine ball bearings for the New York fire department. Scientific American, 78(1):7, January 1, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n1/pdf/scientificamerican01011898-7.pdf.

Anonymous:1898:T

[8492] Anonymous. The telescriptor. Scientific American, 78(1):7–8, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-7a.pdf.

Anonymous:1898:SBJ

[8493] Anonymous. A stopper for bottles, jars, etc. Scientific American, 78(1):8, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n1/pdf/scientificamerican01011898-
8.pdf.

Serviss:1898:JH

[8494] Garrett P. Serviss. The January heavens. Scientific American, 78(1):8, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-8a.pdf.

Anonymous:1898:TSN

[8495] Anonymous. The trade schools of the New York State Reformatory. Scientific American, 78(1):9, January 1, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n1/pdf/scientificamerican01011898-9.pdf.

Barris:1898:HSH

[8496] Thomas C. Barris. Hydrographic surveying-how the maps and charts of our coast and harbors are made. Scientific American, 78(1):10–11, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-10.pdf.

Anonymous:1898:RG

[8497] Anonymous. Rameses the great. Scientific American, 78(1):11, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-11a.pdf.

Caulk:1898:BJ

[8498] W. B. Caulk. The blindfolded juggler. Scientific American, 78(1):11, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-11.pdf.

Anonymous:1898:BNBa

[8499] Anonymous. Books: New books, etc. Scientific American, 78(1):12, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-12a.pdf.

Anonymous:1898:RPIa

[8500] Anonymous. Recently patented inventions. Scientific American, 78(1):12, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n1/pdf/scientificamerican01011898-
12.pdf.

Anonymous:1898:BPa

[8501] Anonymous. Business and personal. Scientific American, 78(1):13, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-13.pdf.

Anonymous:1898:Ia

[8502] Anonymous. To inventors. Scientific American, 78(1):13, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-13b.pdf.

Anonymous:1898:IIa

[8503] Anonymous. Index of inventions. Scientific American, 78(1):13, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-13c.pdf.

Anonymous:1898:NQa

[8504] Anonymous. Notes and queries. Scientific American, 78(1):13, January 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n1/pdf/scientificamerican01011898-13a.pdf.

Anonymous:1898:SNY

[8505] Anonymous. Strength of the New York and Brooklyn Suspension Bridge. Scientific American, 78(2):18, January 8, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n2/pdf/scientificamerican01081898-18.pdf.

Anonymous:1898:WB

[8506] Anonymous. The Weather Bureau in 1897. Scientific American, 78(2):18–19, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
18a.pdf.

Anonymous:1898:AFa

[8507] Anonymous. Artificial flight. Scientific American, 78(2):19, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-19.pdf.

Anonymous:1898:EPa

[8508] Anonymous. Edison on patents. Scientific American, 78(2):19, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-19a.pdf.

Anonymous:1898:RSF

[8509] Anonymous. Recovery of silver from fixing baths. Scientific American, 78(2):19, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
19c.pdf.

Anonymous:1898:SI

[8510] Anonymous. The Smithsonian Institution, 1846–1896. Scientific American, 78(2):19, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
19b.pdf.

Anonymous:1898:ABS

[8511] Anonymous. An angular bit stock. Scientific American, 78(2):20, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-20.pdf.

Anonymous:1898:IBBa

[8512] Anonymous. An invisible brake for bicycles. Scientific American, 78(2):20, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
20b.pdf.

Anonymous:1898:IDB

[8513] Anonymous. An improved differential bicycle gear. Scientific American, 78(2):20, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
20a.pdf.

Anonymous:1898:PIa

[8514] Anonymous. A piano for invalids. Scientific American, 78(2):21, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-21.pdf.

Anonymous:1898:TRF

[8515] Anonymous. The transportation by rail of fish and fish eggs. Scientific American, 78(2):21, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
21a.pdf.

Anonymous:1898:SWBa

[8516] Anonymous. The submarine wrecking-boat “Argonaut”. Scientific American, 78(2):22, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
22.pdf.

Anonymous:1898:Ca

[8517] Anonymous. Correspondence. Scientific American, 78(2):23, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-23.pdf.

Anonymous:1898:DD

[8518] Anonymous. Decimalization of the day. Scientific American, 78(2):24, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-24.pdf.

Anonymous:1898:RFH

[8519] Anonymous. The redwood forests of Humboldt County, California. Scientific American, 78(2):24–25, January 8, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n2/pdf/scientificamerican01081898-24a.pdf.

Holder:1898:AF

[8520] C. F. Holder. Amphibious fishes. Scientific American, 78(2):25–26, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-25.pdf.

Anonymous:1898:SACa

[8521] Anonymous. Some American cottages. Scientific American, 78 (2):27–28, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
27.pdf.

Anonymous:1898:RPIb

[8522] Anonymous. Recently patented inventions. Scientific American, 78(2):28, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n2/pdf/scientificamerican01081898-
28.pdf.

Anonymous:1898:BNBb

[8523] Anonymous. Books: New books, etc. Scientific American, 78(2):29, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-29a.pdf.

Anonymous:1898:BPb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-29.pdf.

Anonymous:1898:Ib

[8525] Anonymous. To inventors. Scientific American, 78(2):29, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-29c.pdf.

Anonymous:1898:IIb

[8526] Anonymous. Index of inventions. Scientific American, 78(2):29, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-29d.pdf.

Anonymous:1898:NQb

[8527] Anonymous. Notes and queries. Scientific American, 78(2):29, January 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n2/pdf/scientificamerican01081898-29b.pdf.

Anonymous:1898:GTJ

[8528] Anonymous. The Great Training Jetty at the mouth of the Columbia River, Oregon. Scientific American, 78(3):33, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-33.pdf.

Anonymous:1898:CRM

[8529] Anonymous. Cylinder ratios in multiple-cylinder steam engines. Scientific American, 78(3):34, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
34.pdf.

Anonymous:1898:ELT

[8530] Anonymous. Electric locomotives at terminal stations. Scientific American, 78(3):34, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
34a.pdf.

Anonymous:1898:FCD

[8531] Anonymous. Forethought in the construction of drydocks. Scientific American, 78(3):34, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
34b.pdf.

Anonymous:1898:AWT

[8532] Anonymous. Around the world in twenty-eight days. Scientific American, 78(3):34–35, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
34c.pdf.

Anonymous:1898:CMN

[8533] Anonymous. The conditions of modern Naval warfare. Scientific American, 78(3):35, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
35e.pdf.

Anonymous:1898:JKR

[8534] Anonymous. Joint Klondike Relief Expedition. Scientific American, 78(3):35, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
35c.pdf.

Anonymous:1898:LS

[8535] Anonymous. Living without a stomach. Scientific American, 78 (3):35, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
35b.pdf.

Anonymous:1898:NT

[8536] Anonymous. A noble thoroughfare. Scientific American, 78(3):35, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-35a.pdf.

Anonymous:1898:NYC

[8537] Anonymous. New York Consolidated. Scientific American, 78 (3):35, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
35.pdf.

Anonymous:1898:YSR

[8538] Anonymous. A Year’s shipping record. Scientific American, 78 (3):35, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
35d.pdf.

Anonymous:1898:CSP

[8539] Anonymous. The combustion of smokeless powder. Scientific American, 78(3):36, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
36a.pdf.

Anonymous:1898:GT

[8540] Anonymous. The Greenville Treaty. Scientific American, 78(3):36, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-36b.pdf.

Anonymous:1898:SSI

[8541] Anonymous. A swinging scaffold improvement. Scientific American, 78(3):36, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
36.pdf.

Anonymous:1898:HL

[8542] Anonymous. A hydraulic lifeboat. Scientific American, 78(3): 37–38, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
37.pdf.

Anonymous:1898:Cb

[8543] Anonymous. Correspondence. Scientific American, 78(3):39, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-39.pdf.

Trowbridge:1898:HEF

[8544] John Trowbridge. High electromotive force. Scientific American, 78(3):41–42, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
41.pdf.

Mindeleff:1898:PAIa

[8545] Cosmos Mindeleff. Pueblo arts and industries. Scientific American, 78(3):42–43, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
42.pdf.

Anonymous:1898:BPc

[8546] Anonymous. Business and personal. Scientific American, 78(3):44, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-44a.pdf.

Anonymous:1898:RPIc

[8547] Anonymous. Recently patented inventions. Scientific American, 78(3):44, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n3/pdf/scientificamerican01151898-
44.pdf.

Anonymous:1898:NQc

[8548] Anonymous. Notes and queries. Scientific American, 78(3):44–45, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-44b.pdf.

Anonymous:1898:Ic

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-45.pdf.

Anonymous:1898:IIc

[8550] Anonymous. Index of inventions. Scientific American, 78(3):45, January 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n3/pdf/scientificamerican01151898-45a.pdf.

Anonymous:1898:ASB

[8551] Anonymous. The arch in steel bridge construction. Scientific American, 78(4):50, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
50b.pdf.

Anonymous:1898:NYS

[8552] Anonymous. The New York State Canals blunder. Scientific American, 78(4):50, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
50.pdf.

Anonymous:1898:WAF

[8553] Anonymous. Why are American faster than English locomotives? Scientific American, 78(4):50, January 22, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n4/pdf/scientificamerican01221898-50a.pdf.

Anonymous:1898:ETN

[8554] Anonymous. Electric traction on the New York elevated roads. Scientific American, 78(4):50–51, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
50c.pdf.

Anonymous:1898:FCa

[8555] Anonymous. The fruit cure. Scientific American, 78(4):51, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n4/pdf/scientificamerican01221898-51b.pdf.

Anonymous:1898:GJC

[8556] Anonymous. Gold jubilee in California. Scientific American, 78 (4):51, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
51a.pdf.

Anonymous:1898:HRT

[8557] Anonymous. A humane rat trap. Scientific American, 78(4):51, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n4/pdf/scientificamerican01221898-51c.pdf.

Anonymous:1898:SLT

[8558] Anonymous. Suggestions for labels and trade marks in China. Scientific American, 78(4):51, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
51.pdf.

Anonymous:1898:ITT

[8559] Anonymous. An improved telephone transmitter. Scientific American, 78(4):52, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
52.pdf.

Anonymous:1898:MKF

[8560] Anonymous. Meteorological kites in France. Scientific American, 78(4):52, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
52a.pdf.

Anonymous:1898:IBBb

[8561] Anonymous. An improved bicycle brake. Scientific American, 78(4):53, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
53a.pdf.

Anonymous:1898:KGL

[8562] Anonymous. The King gasoline launch engine. Scientific American, 78(4):53, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
53.pdf.

Anonymous:1898:IKE

[8563] Anonymous. An improved kerosene engine. Scientific American, 78(4):53–55, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
53b.pdf.

Anonymous:1898:EMC

[8564] Anonymous. The Edison Magnetic Concentrating Works. Scientific American, 78(4):55–57, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
55.pdf.

Sharp:1898:S

[8565] Benjamin Sharp. The fur seal. Scientific American, 78(4):57–59, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n4/pdf/scientificamerican01221898-57.pdf.

Caulk:1898:SBT

[8566] W. B. Caulk. Spinning and balancing tricks. Scientific American, 78(4):59–60, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
59.pdf.

Anonymous:1898:RPId

[8567] Anonymous. Recently patented inventions. Scientific American, 78(4):60–61, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
60.pdf.

Anonymous:1898:BNBc

[8568] Anonymous. Books: New books, etc. Scientific American, 78 (4):61, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n4/pdf/scientificamerican01221898-
61b.pdf.

Anonymous:1898:BPd

[8569] Anonymous. Business and personal. Scientific American, 78(4):61, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n4/pdf/scientificamerican01221898-61.pdf.

Anonymous:1898:Id

[8570] Anonymous. To inventors. Scientific American, 78(4):61, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n4/pdf/scientificamerican01221898-61c.pdf.

Anonymous:1898:IId

[8571] Anonymous. Index of inventions. Scientific American, 78(4):61, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n4/pdf/scientificamerican01221898-61d.pdf.

Anonymous:1898:NQd

[8572] Anonymous. Notes and queries. Scientific American, 78(4):61, January 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n4/pdf/scientificamerican01221898-61a.pdf.

Anonymous:1898:ACP

[8573] Anonymous. The appointment of a Commissioner of Patents. Scientific American, 78(5):66, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n5/pdf/scientificamerican01291898-
66.pdf.

Anonymous:1898:UNL

[8574] Anonymous. Urgent need for larger Patent Office appropriations. Scientific American, 78(5):66–67, January 29, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n5/pdf/scientificamerican01291898-66a.pdf.

Anonymous:1898:AAI

[8575] Anonymous. American Association of Inventors and Manufacturers. Scientific American, 78(5):67, January 29, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n5/pdf/scientificamerican01291898-67a.pdf.

Anonymous:1898:DMB

[8576] Anonymous. Death of Mr. Butterworth. Scientific American, 78(5):67, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n5/pdf/scientificamerican01291898-
67.pdf.

Anonymous:1898:MN

[8577] Anonymous. The Moodus noises. Scientific American, 78(5):67, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-67b.pdf.

Anonymous:1898:NGM

[8578] Anonymous. A new gasoline motor. Scientific American, 78(5):68, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-68a.pdf.

Anonymous:1898:PCDa

[8579] Anonymous. Prescription cabinet for druggists, etc. Scientific American, 78(5):68, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n5/pdf/scientificamerican01291898-
68b.pdf.

Anonymous:1898:TRD

[8580] Anonymous. A tire removing device. Scientific American, 78(5):68, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-68.pdf.

Anonymous:1898:LBU

[8581] Anonymous. The latest battleships for the United States Navy. Scientific American, 78(5):69–70, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v78/n5/pdf/scientificamerican01291898-
69.pdf.

Anonymous:1898:NSAa

[8582] Anonymous. New Scientific American offices in Washington, D.C. Scientific American, 78(5):70, January 29, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n5/pdf/scientificamerican01291898-70.pdf.

Anonymous:1898:Cc

[8583] Anonymous. Correspondence. Scientific American, 78(5):71, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-71.pdf.

Anonymous:1898:HPN

[8584] Anonymous. Homing pigeons in Naval Service. Scientific American, 78(5):72, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n5/pdf/scientificamerican01291898-
72.pdf.

Anonymous:1898:CBa

[8585] Anonymous. Curios from Benin. Scientific American, 78(5):73–74, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-73.pdf.

Anonymous:1898:NMCa

[8586] Anonymous. A ninety million candle power electric beacon. Scientific American, 78(5):75–76, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n5/pdf/scientificamerican01291898-
75.pdf.

Anonymous:1898:BPe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-77a.pdf.

Anonymous:1898:Ie

[8588] Anonymous. To inventors. Scientific American, 78(5):77, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-77c.pdf.

Anonymous:1898:IIe

[8589] Anonymous. Index of inventions. Scientific American, 78(5):77, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-77d.pdf.

Anonymous:1898:NQe

[8590] Anonymous. Notes and queries. Scientific American, 78(5):77, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n5/pdf/scientificamerican01291898-77b.pdf.

Anonymous:1898:RPIe

[8591] Anonymous. Recently patented inventions. Scientific American, 78(5):77, January 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n5/pdf/scientificamerican01291898-
77.pdf.

Anonymous:1898:CSa

[8592] Anonymous. Cost of steam in 1870 and 1897. Scientific American, 78(6):82, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
82e.pdf.

Anonymous:1898:ECI

[8593] Anonymous. Excellent condition of the iron and steel export trade. Scientific American, 78(6):82, February 5, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n6/pdf/scientificamerican02051898-82f.pdf.

Anonymous:1898:ERS

[8594] Anonymous. Electricity to replace steam on the London Underground Railways. Scientific American, 78(6):82, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n6/pdf/scientificamerican02051898-82a.pdf.

Anonymous:1898:SMP

[8595] Anonymous. Sprague motor patent decision. Scientific American, 78(6):82, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
82d.pdf.

Anonymous:1898:SPa

[8596] Anonymous. Smoke and its prevention. Scientific American, 78 (6):82, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
82c.pdf.

Anonymous:1898:SSTa

[8597] Anonymous. Striking success of the tin-plate industry in the unites states. Scientific American, 78(6):82, February 5, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n6/pdf/scientificamerican02051898-82.pdf.

Anonymous:1898:TRS

[8598] Anonymous. The third rail system in Germany. Scientific American, 78(6):82, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
82b.pdf.

Anonymous:1898:TUS

[8599] Anonymous. The turbine for United States torpedo boats. Scientific American, 78(6):82–83, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
82g.pdf.

Anonymous:1898:ASA

[8600] Anonymous. Annual session of the American Society of Civil Engineers. Scientific American, 78(6):83, February 5, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n6/pdf/scientificamerican02051898-83d.pdf.

Anonymous:1898:CGLa

[8601] Anonymous. Commerce on the Great Lakes. Scientific American, 78(6):83, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
83e.pdf.

Anonymous:1898:GTT

[8602] Anonymous. Government tests of timber. Scientific American, 78(6):83, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
83b.pdf.

Anonymous:1898:JCL

[8603] Anonymous. Japanese cruiser launched at Cramps’ Yard. Scientific American, 78(6):83, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
83c.pdf.

Anonymous:1898:PF

[8604] Anonymous. Petroleum fuel. Scientific American, 78(6):83, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n6/pdf/scientificamerican02051898-83.pdf.

Anonymous:1898:TER

[8605] Anonymous. A third East River Suspension Bridge. Scientific American, 78(6):83, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
83a.pdf.

Serviss:1898:HF

[8606] Garrett P. Serviss. The heavens in February. Scientific American, 78(6):83, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
83f.pdf.

Anonymous:1898:CCC

[8607] Anonymous. A cableway carrier and carriage. Scientific American, 78(6):84, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
84b.pdf.

Anonymous:1898:NRP

[8608] Anonymous. A novel repeating pistol. Scientific American, 78 (6):84, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
84.pdf.

Anonymous:1898:PPT

[8609] Anonymous. A puncture-proof tire. Scientific American, 78(6): 84, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
84a.pdf.

Anonymous:1898:EMP

[8610] Anonymous. An extensible measuring pole. Scientific American, 78(6):84–85, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
84c.pdf.

Anonymous:1898:IREa

[8611] Anonymous. An improved rotary engine. Scientific American, 78(6):85, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
85.pdf.

Anonymous:1898:PVa

[8612] Anonymous. The Planet Venus. Scientific American, 78(6):85– 86, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
85a.pdf.

Anonymous:1898:Cd

[8613] Anonymous. Correspondence. Scientific American, 78(6):87, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n6/pdf/scientificamerican02051898-87.pdf.

Anonymous:1898:CRD

[8614] Anonymous. The Croton River Dam, New York Water Supply. Scientific American, 78(6):88–89, February 5, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n6/pdf/scientificamerican02051898-88.pdf.

Anonymous:1898:HCH

[8615] Anonymous. Hon. Charles H. Duell, Commissioner of Patents. Scientific American, 78(6):89, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
89.pdf.

Anonymous:1898:JAP

[8616] Anonymous. Judge Arthur P. Greeley. Scientific American, 78 (6):89–90, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
89a.pdf.

Anonymous:1898:SFa

[8617] Anonymous. The spider and the fly. Scientific American, 78 (6):91, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
91.pdf.

Aldrich:1898:BC

[8618] Cyrus Aldrich. “A brief for the cigarette”. Scientific American, 78(6):92, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
92b.pdf.

Anonymous:1898:BNBd

[8619] Anonymous. Books: New books, etc. Scientific American, 78 (6):92, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
92a.pdf.

Anonymous:1898:BPf

[8620] Anonymous. Business and personal. Scientific American, 78 (6):92, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
92c.pdf.

Anonymous:1898:RPIf

[8621] Anonymous. Recently patented inventions. Scientific American, 78(6):92, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
92.pdf.

Anonymous:1898:NQf

[8622] Anonymous. Notes and queries. Scientific American, 78(6):92– 93, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n6/pdf/scientificamerican02051898-
92d.pdf.

Anonymous:1898:If

[8623] Anonymous. To inventors. Scientific American, 78(6):93, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n6/pdf/scientificamerican02051898-93.pdf.

Anonymous:1898:IIf

[8624] Anonymous. Index of inventions. Scientific American, 78(6):93, February 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n6/pdf/scientificamerican02051898-93a.pdf.

Anonymous:1898:FE

[8625] Anonymous. False economy. Scientific American, 78(7):98, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n7/pdf/scientificamerican02121898-98a.pdf.

Anonymous:1898:LA

[8626] Anonymous. The liquefaction of air. Scientific American, 78 (7):98, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
98.pdf.

Anonymous:1898:GOE

[8627] Anonymous. Growth in our exports of American locomotives. Scientific American, 78(7):98–99, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
98b.pdf.

Anonymous:1898:BRS

[8628] Anonymous. A brief review of some branches of the work, scientific and practical, of the Health Department of the City of New York. Scientific American, 78(7):99, February 12, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n7/pdf/scientificamerican02121898-99a.pdf.

Anonymous:1898:CSb

[8629] Anonymous. The current supplement. Scientific American, 78 (7):99, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
99c.pdf.

Anonymous:1898:EB

[8630] Anonymous. Excavations at Branchidæ. Scientific American, 78 (7):99, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
99b.pdf.

Anonymous:1898:GNT

[8631] Anonymous. Gun of new type successfully tested. Scientific American, 78(7):99, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
99.pdf.

Anonymous:1898:ISBa

[8632] Anonymous. An improved steam boiler. Scientific American, 78 (7):100, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
100.pdf.

Anonymous:1898:SSB

[8633] Anonymous. A supporting stand for bicycles. Scientific American, 78(7):100, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
100a.pdf.

Anonymous:1898:STR

[8634] Anonymous. A seal and tag for railroad freight cars. Scientific American, 78(7):100, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
100b.pdf.

Anonymous:1898:F

[8635] Anonymous. The fluorometer. Scientific American, 78(7):101– 102, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
101.pdf.

Anonymous:1898:RDV

[8636] Anonymous. The Riverside Drive Viaduct, New York. Scientific American, 78(7):102, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
102.pdf.

Anonymous:1898:Ce

[8637] Anonymous. Correspondence. Scientific American, 78(7):103, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n7/pdf/scientificamerican02121898-103.pdf.

Anonymous:1898:MSI

[8638] Anonymous. Medium and small independent refrigerating plants. Scientific American, 78(7):104–105, February 12, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n7/pdf/scientificamerican02121898-104.pdf.

Anonymous:1898:DG

[8639] Anonymous. Dredging for gold. Scientific American, 78(7):105– 106, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
105.pdf.

Anonymous:1898:LIB

[8640] Anonymous. A lateen ice boat. Scientific American, 78(7):106, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n7/pdf/scientificamerican02121898-106.pdf.

Beard:1898:HAAa

[8641] J. Carter Beard. Habitual attitudes of animals considered as a department of comparative zoology. Scientific American, 78(7):107, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n7/pdf/scientificamerican02121898-107.pdf.

Anonymous:1898:BPg

[8642] Anonymous. Business and personal. Scientific American, 78(7): 108, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
108a.pdf.

Anonymous:1898:RPIg

[8643] Anonymous. Recently patented inventions. Scientific American, 78(7):108, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
108.pdf.

Anonymous:1898:NQg

[8644] Anonymous. Notes and queries. Scientific American, 78(7):108– 109, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
108b.pdf.

Anonymous:1898:BNBe

[8645] Anonymous. Books: New books, etc. Scientific American, 78 (7):109, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
109.pdf.

Anonymous:1898:Ig

[8646] Anonymous. To inventors. Scientific American, 78(7):109, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n7/pdf/scientificamerican02121898-109a.pdf.

Anonymous:1898:IIg

[8647] Anonymous. Index of inventions. Scientific American, 78(7): 109–110, February 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n7/pdf/scientificamerican02121898-
109b.pdf.

Anonymous:1898:EAE

[8648] Anonymous. Economic aspect of the English engineers’ strike. Scientific American, 78(8):114, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
114.pdf.

Anonymous:1898:EPb

[8649] Anonymous. Extensions of patents. Scientific American, 78(8): 114, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
114a.pdf.

Anonymous:1898:OIF

[8650] Anonymous. Our increasing foreign trade. Scientific American, 78(8):114, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
114b.pdf.

Anonymous:1898:IOEa

[8651] Anonymous. The industrial outlook in Egypt. Scientific American, 78(8):114–115, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
114c.pdf.

Anonymous:1898:CDSa

[8652] Anonymous. Commissioner Duell sworn in. Scientific American, 78(8):115, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
115.pdf.

Anonymous:1898:CFP

[8653] Anonymous. Changes in foreign patent laws and practice. Scientific American, 78(8):115, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
115a.pdf.

Anonymous:1898:SDB

[8654] Anonymous. The safe deposit business in ancient Rome. Scientific American, 78(8):115, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
115c.pdf.

Anonymous:1898:TIDa

[8655] Anonymous. Two interesting discoveries in Italy. Scientific American, 78(8):115, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
115b.pdf.

Anonymous:1898:GGE

[8656] Anonymous. Gas and gasoline engines. Scientific American, 78 (8):116, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
116b.pdf.

Anonymous:1898:IWSa

[8657] Anonymous. An improved weighing scoop. Scientific American, 78(8):116, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
116.pdf.

Anonymous:1898:SCM

[8658] Anonymous. Smith’s change-making device. Scientific American, 78(8):116, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
116a.pdf.

Anonymous:1898:WJF

[8659] Anonymous. A water-jacketed furnace and boiler. Scientific American, 78(8):116, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
116c.pdf.

Beard:1898:HAAb

[8660] J. Carter Beard. Habitual attitudes of animals considered as a department of comparative zoology. Scientific American, 78(8):

117–118, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
117.pdf.

Anonymous:1898:CGLb

[8661] Anonymous. Commerce of the Great Lakes. Scientific American, 78(8):119, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
119.pdf.

Anonymous:1898:LLE

[8662] Anonymous. The largest locomotive ever constructed. Scientific American, 78(8):120, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
120.pdf.

Anonymous:1898:DES

[8663] Anonymous. Donatello’s Equestrian Statue at Padua. Scientific American, 78(8):121–122, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
121.pdf.

Anonymous:1898:MAB

[8664] Anonymous. The McKinnon automatic boat-launching device. Scientific American, 78(8):122–123, February 19, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n8/pdf/scientificamerican02191898-122.pdf.

Anonymous:1898:BPh

[8665] Anonymous. Business and personal. Scientific American, 78(8): 124, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
124a.pdf.

Anonymous:1898:RPIh

[8666] Anonymous. Recently patented inventions. Scientific American, 78(8):124, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
124.pdf.

Anonymous:1898:NQh

[8667] Anonymous. Notes and queries. Scientific American, 78(8):124– 125, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
124b.pdf.

Anonymous:1898:BNBf

[8668] Anonymous. Books: New books, etc. Scientific American, 78 (8):125, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
125.pdf.

Anonymous:1898:Ih

[8669] Anonymous. To inventors. Scientific American, 78(8):125, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n8/pdf/scientificamerican02191898-125a.pdf.

Anonymous:1898:IIh

[8670] Anonymous. Index of inventions. Scientific American, 78(8): 125–126, February 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n8/pdf/scientificamerican02191898-
125b.pdf.

Anonymous:1898:DM

[8671] Anonymous. The disaster to the “Maine”. Scientific American, 78(9):130, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
130.pdf.

Anonymous:1898:WBW

[8672] Anonymous. Weather Bureau warnings. Scientific American, 78 (9):130–131, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
130a.pdf.

Anonymous:1898:FTB

[8673] Anonymous. Flames in a tall building. Scientific American, 78 (9):131, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
131a.pdf.

Anonymous:1898:RSI

[8674] Anonymous. Report of the Smithsonian Institution. Scientific American, 78(9):131, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
131.pdf.

Anonymous:1898:STP

[8675] Anonymous. Successful test of pneumatic postal tube. Scientific American, 78(9):131–132, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
131b.pdf.

Anonymous:1898:CDF

[8676] Anonymous. A convenient document file. Scientific American, 78(9):132, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
132a.pdf.

Anonymous:1898:HGI

[8677] Anonymous. A head gate for irrigating canals, etc. Scientific American, 78(9):132, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
132.pdf.

Anonymous:1898:IKS

[8678] Anonymous. An improved knife sharpener. Scientific American, 78(9):132, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
132b.pdf.

Anonymous:1898:RSH

[8679] Anonymous. A ring set holding clamp. Scientific American, 78 (9):132, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
132c.pdf.

Anonymous:1898:LBM

[8680] Anonymous. The loss of the battleship “Maine”. Scientific American, 78(9):133–134, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
133.pdf.

Hale:1898:IG

[8681] William H. Hale. Insect grafting. Scientific American, 78(9):135, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n9/pdf/scientificamerican02261898-135.pdf.

Anonymous:1898:NEL

[8682] Anonymous. The new experimental locomotive for Purdue University. Scientific American, 78(9):136, February 26, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n9/pdf/scientificamerican02261898-136.pdf.

Anonymous:1898:GFU

[8683] Anonymous. The gun factory at the United States Washington, D.C., Navy Yard. Scientific American, 78(9):136–139, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n9/pdf/scientificamerican02261898-136a.pdf.

Anonymous:1898:ETS

[8684] Anonymous. Electric towboat in a sewer. Scientific American, 78(9):139, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
139.pdf.

Anonymous:1898:ICH

[8685] Anonymous. An improved copy-holder. Scientific American, 78 (9):139–140, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
139a.pdf.

Anonymous:1898:RPIi

[8686] Anonymous. Recently patented inventions. Scientific American, 78 (9):140–141, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
140.pdf.

Anonymous:1898:BPi

[8687] Anonymous. Business and personal. Scientific American, 78(9): 141, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
141.pdf.

Anonymous:1898:Ij

[8688] Anonymous. To inventors. Scientific American, 78(9):141, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n9/pdf/scientificamerican02261898-141b.pdf.

Anonymous:1898:NQi

[8689] Anonymous. Notes and queries. Scientific American, 78(9):141, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n9/pdf/scientificamerican02261898-141a.pdf.

Anonymous:1898:IIi

[8690] Anonymous. Index of inventions. Scientific American, 78(9): 141–142, February 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n9/pdf/scientificamerican02261898-
141c.pdf.

Anonymous:1898:MI

[8691] Anonymous. The “Maine” inquiry. Scientific American, 78(10):146, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-146.pdf.

Anonymous:1898:RYb

[8692] Anonymous. Railroad to the Yukon. Scientific American, 78(10):146, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-146a.pdf.

Anonymous:1898:SDE

[8693] Anonymous. Side doors or end doors on railroad cars. Scientific American, 78(10):146–147, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
146b.pdf.

Holder:1898:OIA

[8694] C. P. Holder. The opium industry in America. Scientific American, 78(10):147, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
147a.pdf.

Serviss:1898:HMa

[8695] Garrett P. Serviss. The heavens in March. Scientific American, 78(10):147, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
147.pdf.

Anonymous:1898:AAGa

[8696] Anonymous. An automatic acetylene gas generator. Scientific American, 78(10):148, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
148a.pdf.

Anonymous:1898:DNF

[8697] Anonymous. Discharge nozzle for fire engines. Scientific American, 78(10):148, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
148b.pdf.

Anonymous:1898:ETBa

[8698] Anonymous. An efficient tire bolt wrench. Scientific American, 78(10):148, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
148.pdf.

Anonymous:1898:SMTa

[8699] Anonymous. Submarine mining and torpedo warfare. Scientific American, 78(10):149–150, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
149.pdf.

Anonymous:1898:Cf

[8700] Anonymous. Correspondence. Scientific American, 78(10):151, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-151.pdf.

Anonymous:1898:AMH

[8701] Anonymous. Apparatus for measuring the heights reached by balloons. Scientific American, 78(10):152, March 5, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n10/pdf/scientificamerican03051898-152.pdf.

Anonymous:1898:RDSa

[8702] Anonymous. Reconstruction of the Dismal Swamp Canal. Scientific American, 78(10):153–154, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
153.pdf.

Anonymous:1898:DFN

[8703] Anonymous. A device for fastening nuts. Scientific American, 78(10):155, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
155a.pdf.

Anonymous:1898:RPIj

[8704] Anonymous. Recently patented inventions. Scientific American, 78(10):155, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n10/pdf/scientificamerican03051898-
155b.pdf.

Dexter:1898:HC

[8705] Edwin G. Dexter. A horned cockerel. Scientific American, 78(10):155, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-155.pdf.

Anonymous:1898:BPj

[8706] Anonymous. Business and personal. Scientific American, 78(10):156, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-156.pdf.

Anonymous:1898:Ik

[8707] Anonymous. To inventors. Scientific American, 78(10):156, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-156b.pdf.

Anonymous:1898:IIj

[8708] Anonymous. Index of inventions. Scientific American, 78(10):156, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-156c.pdf.

Anonymous:1898:NQj

[8709] Anonymous. Notes and queries. Scientific American, 78(10):156, March 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n10/pdf/scientificamerican03051898-156a.pdf.

Anonymous:1898:CBN

[8710] Anonymous. The completion of the Brooklyn Naval Dry Dock. Scientific American, 78(11):162, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
162b.pdf.

Anonymous:1898:KCC

[8711] Anonymous. Klondike and California compared. Scientific American, 78(11):162, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
162.pdf.

Anonymous:1898:OTE

[8712] Anonymous. Our trade with Europe. Scientific American, 78(11):162, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-162c.pdf.

Anonymous:1898:USC

[8713] Anonymous. The United States Civil Service. Scientific American, 78(11):162, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
162a.pdf.

Knerr:1898:BFI

[8714] E. B. Knerr. Burning fields of ice. Scientific American, 78(11):162–163, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-162d.pdf.

Anonymous:1898:LUR

[8715] Anonymous. London’s underground road. Scientific American, 78(11):163, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
163a.pdf.

Anonymous:1898:MRP

[8716] Anonymous. A metal rust preventive of 1402. Scientific American, 78(11):163, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
163b.pdf.

Anonymous:1898:PLU

[8717] Anonymous. The public lands of the United States. Scientific American, 78(11):163, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
163.pdf.

Anonymous:1898:ATF

[8718] Anonymous. An apparatus for thawing frozen ground. Scientific American, 78(11):164, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
164b.pdf.

Anonymous:1898:HNG

[8719] Anonymous. Humbert’s noiseless gun. Scientific American, 78(11):164, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-164a.pdf.

Anonymous:1898:SWGa

[8720] Anonymous. A safety water gage for boilers. Scientific American, 78(11):164, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
164.pdf.

Anonymous:1898:DHM

[8721] Anonymous. Diesel’s heat motor. Scientific American, 78(11):165–166, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-165.pdf.

Anonymous:1898:SPD

[8722] Anonymous. A shoe polishing device. Scientific American, 78(11):166, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-166.pdf.

Anonymous:1898:MD

[8723] Anonymous. The “Maine” disaster. Scientific American, 78(11):167, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-167.pdf.

Anonymous:1898:NPS

[8724] Anonymous. New practice ship for the United States Naval Cadets. Scientific American, 78(11):168, March 12, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n11/pdf/scientificamerican03121898-168.pdf.

Anonymous:1898:KRE

[8725] Anonymous. The Klondike Relief Expedition. Scientific American, 78(11):169–170, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
169.pdf.

Anonymous:1898:SRO

[8726] Anonymous. Some recently observed Sun spots. Scientific American, 78(11):171, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
171.pdf.

Anonymous:1898:AGB

[8727] Anonymous. The “Apennino” of Giovanni Bologna. Scientific American, 78(11):171–172, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
171a.pdf.

Anonymous:1898:RPIk

[8728] Anonymous. Recently patented inventions. Scientific American, 78(11):172, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n11/pdf/scientificamerican03121898-
172.pdf.

Anonymous:1898:BNBg

[8729] Anonymous. Books: New books, etc. Scientific American, 78(11):173, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-173b.pdf.

Anonymous:1898:BPk

[8730] Anonymous. Business and personal. Scientific American, 78(11):173, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-173a.pdf.

Anonymous:1898:Il

[8731] Anonymous. To inventors. Scientific American, 78(11):173, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-173d.pdf.

Anonymous:1898:NQk

[8732] Anonymous. Notes and queries. Scientific American, 78(11):173, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-173c.pdf.

Wallace:1898:CS

[8733] John Wallace. Cigarettes and science. Scientific American, 78(11):173, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-173.pdf.

Anonymous:1898:IIk

[8734] Anonymous. Index of inventions. Scientific American, 78(11):173–174, March 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n11/pdf/scientificamerican03121898-173e.pdf.

Anonymous:1898:WMM

[8735] Anonymous. War measures that make for peace. Scientific American, 78(12):178, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
178.pdf.

Anonymous:1898:BCI

[8736] Anonymous. The Brussels Conference of the International Union for the Protection of Industrial Property. Scientific American, 78(12):178–179, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-178a.pdf.

Anonymous:1898:NDL

[8737] Anonymous. The new director of the Lick Observatory. Scientific American, 78(12):179, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
179c.pdf.

Anonymous:1898:NSF

[8738] Anonymous. The new single forged cannon. Scientific American, 78(12):179, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
179.pdf.

Anonymous:1898:RSTa

[8739] Anonymous. The reversing steam turbine. Scientific American, 78(12):179, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
179a.pdf.

Anonymous:1898:VAC

[8740] Anonymous. Valuable advice to cattle shippers. Scientific American, 78(12):179, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
179b.pdf.

Anonymous:1898:HFS

[8741] Anonymous. How flag sticks are made. Scientific American, 78(12):180, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-180a.pdf.

Anonymous:1898:LMA

[8742] Anonymous. A lawn mower attachment. Scientific American, 78(12):180, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
180.pdf.

Anonymous:1898:VWT

[8743] Anonymous. A vehicle wheel tire tightener. Scientific American, 78(12):180, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
180b.pdf.

Durand:1898:EEV

[8744] W. F. Durand. Experiments on the effect of varying the area of screw propellers. Scientific American, 78(12):181–182, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-181.pdf.

Anonymous:1898:Cg

[8745] Anonymous. Correspondence. Scientific American, 78(12):183, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-183.pdf.

Anonymous:1898:ICa

[8746] Anonymous. Improved calipers. Scientific American, 78(12):184, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-184.pdf.

Anonymous:1898:SDT

[8747] Anonymous. Spanish destroyers and torpedo boats destined for Cuba. Scientific American, 78(12):184, March 19, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n12/pdf/scientificamerican03191898-184a.pdf.

Anonymous:1898:MPa

[8748] Anonymous. The manufacture of paper. Scientific American, 78 (12):185–186, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
185.pdf.

Anonymous:1898:NDP

[8749] Anonymous. A newly discovered portrait of Amerigo Vespucci. Scientific American, 78(12):187, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
187.pdf.

Anonymous:1898:BNBh

[8750] Anonymous. Books: New books, etc. Scientific American, 78(12):188, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-188c.pdf.

Anonymous:1898:BPl

[8751] Anonymous. Business and personal. Scientific American, 78(12):188, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-188a.pdf.

Anonymous:1898:Im

[8752] Anonymous. To inventors. Scientific American, 78(12):188, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-188d.pdf.

Anonymous:1898:IIl

[8753] Anonymous. Index of inventions. Scientific American, 78(12):188, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-188e.pdf.

Anonymous:1898:NQl

[8754] Anonymous. Notes and queries. Scientific American, 78(12):188, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n12/pdf/scientificamerican03191898-188b.pdf.

Anonymous:1898:RPIl

[8755] Anonymous. Recently patented inventions. Scientific American, 78(12):188, March 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n12/pdf/scientificamerican03191898-
188.pdf.

Anonymous:1898:EWT

[8756] Anonymous. The efficiency of the water tube boiler. Scientific American, 78(13):194, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
194a.pdf.

Anonymous:1898:ORP

[8757] Anonymous. Our recently purchased warships. Scientific American, 78(13):194, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
194.pdf.

Anonymous:1898:OMC

[8758] Anonymous. Is our maritime commerce vulnerable? Scientific American, 78(13):194–195, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
194b.pdf.

Anonymous:1898:DEa

[8759] Anonymous. The development of Egypt. Scientific American, 78(13):195, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
195b.pdf.

Anonymous:1898:EBG

[8760] Anonymous. Eating before going to bed. Scientific American, 78(13):195, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
195d.pdf.

Anonymous:1898:NNW

[8761] Anonymous. The new Naval war game. Scientific American, 78(13):195, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-195a.pdf.

Anonymous:1898:RCP

[8762] Anonymous. Report of the Commissioner of Patents for 1897. Scientific American, 78(13):195, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
195.pdf.

Anonymous:1898:SCS

[8763] Anonymous. Street car service across the Brooklyn Bridge. Scientific American, 78(13):195, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
195c.pdf.

Anonymous:1898:ALC

[8764] Anonymous. An apparatus for loading cane, etc. Scientific American, 78(13):196, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
196a.pdf.

Anonymous:1898:MWO

[8765] Anonymous. A machine for washing oranges, lemons, etc. Scientific American, 78(13):196, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
196b.pdf.

Anonymous:1898:NMCb

[8766] Anonymous. A novel miter clamp. Scientific American, 78(13):196, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-196.pdf.

Anonymous:1898:Z

[8767] Anonymous. The Zerograph. Scientific American, 78(13):197, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-197.pdf.

Anonymous:1898:ESR

[8768] Anonymous. An electric street railway sprinkler. Scientific American, 78(13):197–198, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
197a.pdf.

Anonymous:1898:Ch

[8769] Anonymous. Correspondence. Scientific American, 78(13):199, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-199.pdf.

Anonymous:1898:CTG

[8770] Anonymous. The conning tower of the gunboat “Helena”. Scientific American, 78(13):200, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
200.pdf.

Anonymous:1898:CU

[8771] Anonymous. Columbia University. Scientific American, 78(13):200–202, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-200a.pdf.

Anonymous:1898:DSH

[8772] Anonymous. Death of Sir Henry Bessemer. Scientific American, 78(13):203, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
203a.pdf.

Anonymous:1898:TW

[8773] Anonymous. The teredo and its work. Scientific American, 78(13):203, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-203.pdf.

Anonymous:1898:BNBi

[8774] Anonymous. Books: New books, etc. Scientific American, 78(13):204, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-204a.pdf.

Anonymous:1898:BPm

[8775] Anonymous. Business and personal. Scientific American, 78(13):204, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-204b.pdf.

Anonymous:1898:RPIm

[8776] Anonymous. Recently patented inventions. Scientific American, 78(13):204, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n13/pdf/scientificamerican03261898-
204.pdf.

Anonymous:1898:NQm

[8777] Anonymous. Notes and queries. Scientific American, 78(13):204–205, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-204c.pdf.

Anonymous:1898:In

[8778] Anonymous. To inventors. Scientific American, 78(13):205, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-205.pdf.

Anonymous:1898:IIm

[8779] Anonymous. Index of inventions. Scientific American, 78(13):205, March 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n13/pdf/scientificamerican03261898-205a.pdf.

Anonymous:1898:FCL

[8780] Anonymous. Four cylinder locomotives. Scientific American, 78 (14):210, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
210a.pdf.

Anonymous:1898:SVA

[8781] Anonymous. A Spanish view of the American Navy. Scientific American, 78(14):210, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
210.pdf.

Anonymous:1898:BIPa

[8782] Anonymous. Bill to increase the Patent Office force. Scientific American, 78(14):210–211, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
210b.pdf.

Anonymous:1898:FCS

[8783] Anonymous. The fruits of civil service reform. Scientific American, 78(14):211, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
211.pdf.

Anonymous:1898:FTT

[8784] Anonymous. Forthcoming telegraphic tournament. Scientific American, 78(14):211, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
211b.pdf.

Anonymous:1898:LBK

[8785] Anonymous. Launch of the battleships “Kearsarge” and “Kentucky”. Scientific American, 78(14):211, April 2, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n14/pdf/scientificamerican04021898-211c.pdf.

Serviss:1898:HAa

[8786] Garrett P. Serviss. The heavens in April. Scientific American, 78(14):211, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
211a.pdf.

Anonymous:1898:CCPa

[8787] Anonymous. Cultivating cotton plants. Scientific American, 78(14):212, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n14/pdf/scientificamerican04021898-212c.pdf.

Anonymous:1898:CSC

[8788] Anonymous. A casing for shafts, couplings, etc. Scientific American, 78(14):212, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
212a.pdf.

Anonymous:1898:IPCa

[8789] Anonymous. An improved power Crank. Scientific American, 78 (14):212, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
212.pdf.

Anonymous:1898:ISS

[8790] Anonymous. An improved sash supporter. Scientific American, 78(14):212, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
212b.pdf.

Anonymous:1898:WT

[8791] Anonymous. Wireless telegraphy. Scientific American, 78(14):213–214, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n14/pdf/scientificamerican04021898-213.pdf.

Peckham:1898:LAP

[8792] W. C. Peckham. Liquid air and its phenomena. Scientific American, 78(14):214–215, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
214.pdf.

Anonymous:1898:NUS

[8793] Anonymous. The Navies of the United States and Spain — a comparison. Scientific American, 78(14):216–218, April 2, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n14/pdf/scientificamerican04021898-216.pdf.

Anonymous:1898:TSE

[8794] Anonymous. The total solar eclipse, January 22, 1898. Scientific American, 78(14):218–219, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
218.pdf.

Anonymous:1898:BNBj

[8795] Anonymous. Books: New books, etc. Scientific American, 78(14):219, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n14/pdf/scientificamerican04021898-219b.pdf.

Anonymous:1898:BPn

[8796] Anonymous. Business and personal. Scientific American, 78(14):219, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n14/pdf/scientificamerican04021898-219a.pdf.

Anonymous:1898:RPIn

[8797] Anonymous. Recently patented inventions. Scientific American, 78(14):219, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n14/pdf/scientificamerican04021898-
219.pdf.

Anonymous:1898:IIn

[8798] Anonymous. Index of inventions. Scientific American, 78(14):220, April 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n14/pdf/scientificamerican04021898-220.pdf.

Anonymous:1898:ORM

[8799] Anonymous. The official report of the “Maine” disaster. Scientific American, 78(15):226, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
226.pdf.

Anonymous:1898:QPB

[8800] Anonymous. A “queer” Patent Bill. Scientific American, 78 (15):226–227, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
226a.pdf.

Anonymous:1898:BIPb

[8801] Anonymous. The bill to increase the Patent Office force. Scientific American, 78(15):227, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
227b.pdf.

Anonymous:1898:NABa

[8802] Anonymous. Naval Appropriation Bill passed. Scientific American, 78(15):227, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
227c.pdf.

Anonymous:1898:RCFa

[8803] Anonymous. Relief for Cuban famine sufferers. Scientific American, 78(15):227, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
227d.pdf.

Anonymous:1898:STBa

[8804] Anonymous. The submarine torpedo boat. Scientific American, 78(15):227, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
227.pdf.

Anonymous:1898:WPS

[8805] Anonymous. The Welsbach Patent sustained. Scientific American, 78(15):227, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
227a.pdf.

Anonymous:1898:IREb

[8806] Anonymous. An improved rotary engine. Scientific American, 78(15):228, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
228.pdf.

Anonymous:1898:MHSa

[8807] Anonymous. A manuscript holder and spacer. Scientific American, 78(15):228, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
228a.pdf.

Hovey:1898:AAL

[8808] Horace C. Hovey. The Aven Armand, Lozere, France. Scientific American, 78(15):228–229, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
228b.pdf.

Anonymous:1898:CAM

[8809] Anonymous. Columbia’s artificial Moon. Scientific American, 78 (15):229–230, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
229.pdf.

Anonymous:1898:LUS

[8810] Anonymous. The launch of the United States battleships “Kentucky” and “Kearsarge”. Scientific American, 78(15):231, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n15/pdf/scientificamerican04091898-231a.pdf.

Anonymous:1898:NO

[8811] Anonymous. The “New Orleans”. Scientific American, 78(15):231, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n15/pdf/scientificamerican04091898-231.pdf.

Anonymous:1898:SMG

[8812] Anonymous. A simple mirror galvanometer. Scientific American, 78(15):232–233, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
232.pdf.

Anonymous:1898:STH

[8813] Anonymous. Successful trials of the Holland submarine boat. Scientific American, 78(15):233–234, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
233.pdf.

Anonymous:1898:RNC

[8814] Anonymous. Report of the Naval Court on the destruction of the “Maine”. Scientific American, 78(15):234–235, April 9, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n15/pdf/scientificamerican04091898-234.pdf.

Anonymous:1898:BNBk

[8815] Anonymous. Books: New books, etc. Scientific American, 78(15):236, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n15/pdf/scientificamerican04091898-236c.pdf.

Anonymous:1898:BPo

[8816] Anonymous. Business and personal. Scientific American, 78(15):236, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n15/pdf/scientificamerican04091898-236b.pdf.

Anonymous:1898:RAC

[8817] Anonymous. Recent analysis of cigarettes. Scientific American, 78(15):236, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
236a.pdf.

Anonymous:1898:RPIo

[8818] Anonymous. Recently patented inventions. Scientific American, 78(15):236, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n15/pdf/scientificamerican04091898-
236.pdf.

Anonymous:1898:Io

[8819] Anonymous. To inventors. Scientific American, 78(15):237, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n15/pdf/scientificamerican04091898-237a.pdf.

Anonymous:1898:NQn

[8820] Anonymous. Notes and queries. Scientific American, 78(15):237, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n15/pdf/scientificamerican04091898-237.pdf.

Anonymous:1898:IIo

[8821] Anonymous. Index of inventions. Scientific American, 78(15):237–238, April 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n15/pdf/scientificamerican04091898-237b.pdf.

Anonymous:1898:SHM

[8822] Anonymous. The Sandy Hook mortar batteries. Scientific American, 78(16):241, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
241.pdf.

Anonymous:1898:ASI

[8823] Anonymous. American supremacy in the iron trade. Scientific American, 78(16):242, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
242.pdf.

Anonymous:1898:NBR

[8824] Anonymous. New bills for the regulation of print and label registration. Scientific American, 78(16):242–243, April 16, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n16/pdf/scientificamerican04161898-242a.pdf.

Anonymous:1898:NWa

[8825] Anonymous. The Navies of the world. Scientific American, 78(16):243, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n16/pdf/scientificamerican04161898-243a.pdf.

Anonymous:1898:PP

[8826] Anonymous. Ptomaine poisoning. Scientific American, 78(16):243, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n16/pdf/scientificamerican04161898-243.pdf.

Hovey:1898:JAA

[8827] Horace O. Hovey. Jubilee anniversary of the American Science Association. Scientific American, 78(16):243, April 16, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n16/pdf/scientificamerican04161898-243b.pdf.

Anonymous:1898:DEb

[8828] Anonymous. A dredge ejector. Scientific American, 78(16):244, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n16/pdf/scientificamerican04161898-244.pdf.

Anonymous:1898:ICC

[8829] Anonymous. An improved car coupling. Scientific American, 78 (16):244, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
244b.pdf.

Anonymous:1898:LER

[8830] Anonymous. A locomotive and its equivalent in raw materials. Scientific American, 78(16):244, April 16, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n16/pdf/scientificamerican04161898-244a.pdf.

Anonymous:1898:NSAb

[8831] Anonymous. The new steel arch at Niagara. Scientific American, 78(16):245–246, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
245.pdf.

Anonymous:1898:ACU

[8832] Anonymous. An automatically closing umbrella. Scientific American, 78(16):246, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
246.pdf.

Holder:1898:KGT

[8833] C. F. Holder. The Key of the Gulf — the Tortugas Islands. Scientific American, 78(16):248–249, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
248.pdf.

Mindeleff:1898:PAIb

[8834] Cosmos Mindeleff. Pueblo arts and industries — II. Scientific American, 78(16):249–250, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
249.pdf.

Anonymous:1898:LVH

[8835] Anonymous. Leonardo Da Vinci’s high angle mortar. Scientific American, 78(16):251, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
251.pdf.

Anonymous:1898:BNBl

[8836] Anonymous. Books: New books, etc. Scientific American, 78(16):252, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n16/pdf/scientificamerican04161898-252a.pdf.

Anonymous:1898:BPp

[8837] Anonymous. Business and personal. Scientific American, 78(16):252, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n16/pdf/scientificamerican04161898-252b.pdf.

Anonymous:1898:RPIp 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n16/pdf/scientificamerican04161898-
252.pdf.

Anonymous:1898:NQo

[8839] Anonymous. Notes and queries. Scientific American, 78(16):252–253, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n16/pdf/scientificamerican04161898-252c.pdf.

Anonymous:1898:IIp

[8840] Anonymous. Index of inventions. Scientific American, 78(16):253–254, April 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n16/pdf/scientificamerican04161898-253.pdf.

Anonymous:1898:CSd

[8841] Anonymous. The command of the sea. Scientific American, 78(17):258, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-258.pdf.

Anonymous:1898:RSB

[8842] Anonymous. Rather small business for the Government. Scientific American, 78(17):258, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
258a.pdf.

Anonymous:1898:NPPa

[8843] Anonymous. Negotiable paper for patent rights. Scientific American, 78(17):258–259, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
258b.pdf.

Anonymous:1898:EMD

[8844] Anonymous. Electric mail delivery. Scientific American, 78(17):259, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-259b.pdf.

Anonymous:1898:OSN

[8845] Anonymous. Our special Navy supplement. Scientific American, 78(17):259, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
259a.pdf.

Anonymous:1898:SACb

[8846] Anonymous. Shall artists copyright their work? Scientific American, 78(17):259, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
259.pdf.

Anonymous:1898:AGG

[8847] Anonymous. Acetylene gas generator. Scientific American, 78(17):260, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-260.pdf.

Anonymous:1898:CGY

[8848] Anonymous. A commodious gasoline yacht. Scientific American, 78(17):260, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
260a.pdf.

Anonymous:1898:BPCa

[8849] Anonymous. A bicycle package carrier. Scientific American, 78(17):261, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-261.pdf.

Anonymous:1898:IPa

[8850] Anonymous. An improved plow. Scientific American, 78(17):261, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-261a.pdf.

Anonymous:1898:IWT

[8851] Anonymous. An improved woodworking tool. Scientific American, 78(17):261, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
261b.pdf.

Anonymous:1898:LSS

[8852] Anonymous. A light and strong spar for vessels. Scientific American, 78(17):261, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
261c.pdf.

Anonymous:1898:ONC

[8853] Anonymous. Our new cruiser “Topeka”. Scientific American, 78 (17):262, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
262.pdf.

Anonymous:1898:RFD

[8854] Anonymous. The range-finder for determining the distance of a hostile ship or fort. Scientific American, 78(17):262, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-262a.pdf.

Anonymous:1898:Ci

[8855] Anonymous. Correspondence. Scientific American, 78(17):263, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-263.pdf.

Anonymous:1898:REF

[8856] Anonymous. Recent experiments on the flow of rocks carried out at McGill University. Scientific American, 78(17):264–265, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-264.pdf.

Anonymous:1898:AT

[8857] Anonymous. The art of tattooing. Scientific American, 78(17):265–266, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-265.pdf.

Anonymous:1898:SAG

[8858] Anonymous. The ’98 solar acetylene gas bicycle lamp. Scientific American, 78(17):266, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
266.pdf.

Peckham:1898:LA

[8859] W. C. Peckham. Liquid air. Scientific American, 78(17):267, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-267.pdf.

Anonymous:1898:BNBm

[8860] Anonymous. Books: New books, etc. Scientific American, 78(17):268, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-268a.pdf.

Anonymous:1898:BPq

[8861] Anonymous. Business and personal. Scientific American, 78(17):268, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-268b.pdf.

Anonymous:1898:RPIq

[8862] Anonymous. Recently patented inventions. Scientific American, 78(17):268, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n17/pdf/scientificamerican04231898-
268.pdf.

Anonymous:1898:NQp

[8863] Anonymous. Notes and queries. Scientific American, 78(17):268–269, April 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-268c.pdf.

Anonymous:1898:IIq

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n17/pdf/scientificamerican04231898-269.pdf.

Anonymous:1898:MPb

[8865] Anonymous. The manufacture of paper. Scientific American, 78 (18):273, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
273.pdf.

Anonymous:1898:PCa

[8866] Anonymous. The pacification of Cuba. Scientific American, 78(18):274, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-274.pdf.

Anonymous:1898:PCb

[8867] Anonymous. Production of coal in 1897. Scientific American, 78(18):274, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
274a.pdf.

Anonymous:1898:PPC

[8868] Anonymous. Preserving plant colors. Scientific American, 78(18):275, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-275a.pdf.

Hovey:1898:ARN

[8869] Edmund Otis Hovey. Annual reception of the New York Academy of Sciences. Scientific American, 78(18):275, April 30, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n18/pdf/scientificamerican04301898-275.pdf.

Serviss:1898:HMb

[8870] Garrett P. Serviss. The heavens in May. Scientific American, 78(18):275, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
275b.pdf.

Anonymous:1898:ESF

[8871] Anonymous. An efficient stovepipe fastener. Scientific American, 78(18):276, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
276b.pdf.

Anonymous:1898:IFa

[8872] Anonymous. An improved faucet. Scientific American, 78(18):276, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-276c.pdf.

Anonymous:1898:SEV

[8873] Anonymous. A simple and effective valve gear. Scientific American, 78(18):276, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
276.pdf.

Anonymous:1898:TAa

[8874] Anonymous. A tether for animals. Scientific American, 78(18):276, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-276a.pdf.

Anonymous:1898:Cj

[8875] Anonymous. Correspondence. Scientific American, 78(18):279, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-279.pdf.

Anonymous:1898:DBM

[8876] Anonymous. The Dashiell breech mechanism. Scientific American, 78(18):280, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
280.pdf.

Anonymous:1898:NWMa

[8877] Anonymous. The new wave motor. Scientific American, 78(18):281, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-281a.pdf.

Smith:1898:NNR

[8878] Hugh M. Smith. Notice of a new ribbon fish from the coast of Oregon. Scientific American, 78(18):281, April 30, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n18/pdf/scientificamerican04301898-281.pdf.

Weir:1898:EWC

[8879] James Weir. The ears of worms, crustaceans and ants. Scientific American, 78(18):282, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
282.pdf.

Caulk:1898:MV

[8880] W. B. Caulk. Mysterious vase. Scientific American, 78(18):283, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-283.pdf.

Anonymous:1898:RPIr

[8881] Anonymous. Recently patented inventions. Scientific American, 78(18):283–284, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n18/pdf/scientificamerican04301898-
283a.pdf.

Anonymous:1898:BNBn

[8882] Anonymous. Books: New books, etc. Scientific American, 78(18):284, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-284.pdf.

Anonymous:1898:BPr

[8883] Anonymous. Business and personal. Scientific American, 78(18):284, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-284a.pdf.

Anonymous:1898:Iq

[8884] Anonymous. To inventors. Scientific American, 78(18):284, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-284c.pdf.

Anonymous:1898:IIr

[8885] Anonymous. Index of inventions. Scientific American, 78(18):284, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-284d.pdf.

Anonymous:1898:NQq

[8886] Anonymous. Notes and queries. Scientific American, 78(18):284, April 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n18/pdf/scientificamerican04301898-284b.pdf.

Anonymous:1898:NABb

[8887] Anonymous. Naval Appropriation Bill. Scientific American, 78(19):290, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-290.pdf.

Anonymous:1898:PIb

[8888] Anonymous. The Philippine Islands. Scientific American, 78 (19):290–291, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n19/pdf/scientificamerican05071898-
290a.pdf.

Anonymous:1898:ATC

[8889] Anonymous. American trade with Central and South American countries. Scientific American, 78(19):291, May 7, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n19/pdf/scientificamerican05071898-291.pdf.

Anonymous:1898:IFb

[8890] Anonymous. An improved furnace. Scientific American, 78(19):292, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-292a.pdf.

Anonymous:1898:LCD

[8891] Anonymous. A lath carrying device for wall paper and other machines. Scientific American, 78(19):292, May 7, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n19/pdf/scientificamerican05071898-292b.pdf.

Anonymous:1898:NBBa

[8892] Anonymous. A new bicycle bell. Scientific American, 78(19):292, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-292.pdf.

Anonymous:1898:OCB

[8893] Anonymous. Oxide of copper batteries. Scientific American, 78(19):293, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-293.pdf.

Holder:1898:ACH

[8894] C. F. Holder. An ancient cactus hedge. Scientific American, 78(19):293–294, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n19/pdf/scientificamerican05071898-
293a.pdf.

Anonymous:1898:WDC

[8895] Anonymous. The Washington, D.C., Navy Yard. Scientific American, 78(19):295, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n19/pdf/scientificamerican05071898-
295.pdf.

Anonymous:1898:RSV

[8896] Anonymous. Receiving ship “Vermont”. Scientific American, 78 (19):296, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n19/pdf/scientificamerican05071898-
296.pdf.

Anonymous:1898:CSDa

[8897] Anonymous. Competition of shepherds’ dogs at Angerville. Scientific American, 78(19):297, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n19/pdf/scientificamerican05071898-
297.pdf.

Anonymous:1898:EEA

[8898] Anonymous. English Engineer’s analysis of squadrons of Spain and United States. Scientific American, 78(19):298, May 7, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n19/pdf/scientificamerican05071898-298.pdf.

Anonymous:1898:ASU

[8899] Anonymous. Analysis of Spanish and United States squadrons — a reply. Scientific American, 78(19):298–299, May 7, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n19/pdf/scientificamerican05071898-298a.pdf.

Anonymous:1898:BNBo

[8900] Anonymous. Books: New books, etc. Scientific American, 78(19):300, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-300c.pdf.

Anonymous:1898:BPs

[8901] Anonymous. Business and personal. Scientific American, 78(19):300, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-300a.pdf.

Anonymous:1898:IIs

[8902] Anonymous. Index of inventions. Scientific American, 78(19):300, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-300d.pdf.

Anonymous:1898:NQr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n19/pdf/scientificamerican05071898-300b.pdf.

Anonymous:1898:RPIs

[8904] Anonymous. Recently patented inventions. Scientific American, 78(19):300, May 7, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n19/pdf/scientificamerican05071898-
300.pdf.

Anonymous:1898:CGBa

[8905] Anonymous. A characteristic group at the Brooklyn Navy Yard. Scientific American, 78(20):305, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
305.pdf.

Anonymous:1898:NMA

[8906] Anonymous. The need for more armored cruisers in our Navy. Scientific American, 78(20):306, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
306.pdf.

Anonymous:1898:CVF

[8907] Anonymous. The Cape Verde Fleet and the “Oregon”. Scientific American, 78(20):306–307, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
306a.pdf.

Anonymous:1898:AI

[8908] Anonymous. The American islands. Scientific American, 78(20):307, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n20/pdf/scientificamerican05141898-307b.pdf.

Anonymous:1898:NTR

[8909] Anonymous. New transatlantic records. Scientific American, 78 (20):307, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
307c.pdf.

Anonymous:1898:NYE

[8910] Anonymous. The New York electrical exhibition. Scientific American, 78(20):307, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
307a.pdf.

Anonymous:1898:OSAa

[8911] Anonymous. To our Scientific American supplement subscribers. Scientific American, 78(20):307, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
307.pdf.

Anonymous:1898:CBM

[8912] Anonymous. A combined bottle and measuring device. Scientific American, 78(20):308, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
308b.pdf.

Anonymous:1898:IREc

[8913] Anonymous. An improved rotary engine. Scientific American, 78 (20):308, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
308.pdf.

Anonymous:1898:NBBb

[8914] Anonymous. A new ball bearing. Scientific American, 78(20):308, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n20/pdf/scientificamerican05141898-308a.pdf.

Anonymous:1898:WMC

[8915] Anonymous. The Winton motor carriage. Scientific American, 78(20):309–310, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
309.pdf.

Anonymous:1898:CGBb

[8916] Anonymous. A characteristic group at the Brooklyn Navy Yard. Scientific American, 78(20):311–312, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
311.pdf.

Anonymous:1898:GNV

[8917] Anonymous. The Great Naval Victory at Manila. Scientific American, 78(20):313, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
313.pdf.

Hovey:1898:ZSN

[8918] E. O. Hovey. The Zoological Station at Naples. Scientific American, 78(20):314–315, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
314.pdf.

Anonymous:1898:BNBp

[8919] Anonymous. Books: New books, etc. Scientific American, 78(20):316, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n20/pdf/scientificamerican05141898-316a.pdf.

Anonymous:1898:RPIt

[8920] Anonymous. Recently patented inventions. Scientific American, 78(20):316, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n20/pdf/scientificamerican05141898-
316.pdf.

Anonymous:1898:BPt

[8921] Anonymous. Business and personal. Scientific American, 78(20):317, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n20/pdf/scientificamerican05141898-317.pdf.

Anonymous:1898:Ir

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n20/pdf/scientificamerican05141898-317b.pdf.

Anonymous:1898:NQs

[8923] Anonymous. Notes and queries. Scientific American, 78(20):317, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n20/pdf/scientificamerican05141898-317a.pdf.

Anonymous:1898:IIt

[8924] Anonymous. Index of inventions. Scientific American, 78(20):317–318, May 14, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n20/pdf/scientificamerican05141898-317c.pdf.

Anonymous:1898:LMB

[8925] Anonymous. The lessons of Manila Bay. Scientific American, 78 (21):322, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
322.pdf.

Anonymous:1898:FEO

[8926] Anonymous. Foreign expert opinion on the “Maine” disaster. Scientific American, 78(21):322–323, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
322a.pdf.

Anonymous:1898:CCA

[8927] Anonymous. Calcium carbide as an aid to fire. Scientific American, 78(21):323, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
323b.pdf.

Anonymous:1898:EE

[8928] Anonymous. Electrical exhibition. Scientific American, 78(21):323, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-323a.pdf.

Anonymous:1898:FNM

[8929] Anonymous. The four new monitors. Scientific American, 78(21):323, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-323.pdf.

Anonymous:1898:USAa

[8930] Anonymous. United States Army recruiting circular. Scientific American, 78(21):323, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
323c.pdf.

Anonymous:1898:MST

[8931] Anonymous. A Mexican statue of terra cotta. Scientific American, 78(21):324, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
324a.pdf.

Anonymous:1898:NHB

[8932] Anonymous. The new Hunt bicycle saddles. Scientific American, 78(21):324, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
324.pdf.

Anonymous:1898:AAGb

[8933] Anonymous. An automatic acetylene gas generator. Scientific American, 78(21):325, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
325a.pdf.

Anonymous:1898:NFP

[8934] Anonymous. A novel fluid pressure regulator. Scientific American, 78(21):325, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
325.pdf.

Anonymous:1898:RRW

[8935] Anonymous. A reversible ratchet wrench. Scientific American, 78(21):325, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
325b.pdf.

Anonymous:1898:ITR

[8936] Anonymous. An improved time recorder. Scientific American, 78 (21):325–326, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
325c.pdf.

Anonymous:1898:PRR

[8937] Anonymous. Porto Rico and the reduction of San Juan. Scientific American, 78(21):326–327, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
326.pdf.

Anonymous:1898:Ck

[8938] Anonymous. Correspondence. Scientific American, 78(21):327, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-327.pdf.

Anonymous:1898:RFGa

[8939] Anonymous. The rapid-fire guns of the “New Orleans”. Scientific American, 78(21):328–330, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
328.pdf.

Anonymous:1898:OTC

[8940] Anonymous. Our troops at Chickamauga Park. Scientific American, 78(21):330–331, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
330.pdf.

Anonymous:1898:BNBq

[8941] Anonymous. Books: New books, etc. Scientific American, 78(21):332, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-332a.pdf.

Anonymous:1898:BPu

[8942] Anonymous. Business and personal. Scientific American, 78(21):332, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-332c.pdf.

Anonymous:1898:Is

[8943] Anonymous. To inventors. Scientific American, 78(21):332, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-332d.pdf.

Anonymous:1898:RPIu

[8944] Anonymous. Recently patented inventions. Scientific American, 78(21):332, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
332.pdf.

Mallet:1898:CC

[8945] J. W. Mallet. Communication on cigarettes. Scientific American, 78(21):332, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n21/pdf/scientificamerican05211898-
332b.pdf.

Anonymous:1898:IIu

[8946] Anonymous. Index of inventions. Scientific American, 78(21):333, May 21, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-333a.pdf.

Anonymous:1898:NQt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n21/pdf/scientificamerican05211898-333.pdf.

Anonymous:1898:SCV

[8948] Anonymous. The Spanish “Cape Verde” Fleet. Scientific American, 78(22):337, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n22/pdf/scientificamerican05281898-
337.pdf.

Anonymous:1898:AMSa

[8949] Anonymous. Air motors for street railway traction. Scientific American, 78(22):338, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n22/pdf/scientificamerican05281898-
338a.pdf.

Anonymous:1898:WPO

[8950] Anonymous. The weak point in our Navy. Scientific American, 78(22):338, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n22/pdf/scientificamerican05281898-
338.pdf.

Anonymous:1898:LBA

[8951] Anonymous. Launch of the battleship “Alabama”. Scientific American, 78(22):338–339, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n22/pdf/scientificamerican05281898-
338b.pdf.

Anonymous:1898:SCF

[8952] Anonymous. State College of Forestry. Scientific American, 78(22):339, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-339b.pdf.

Anonymous:1898:SJB

[8953] Anonymous. San Juan’s bombardment. Scientific American, 78 (22):339, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n22/pdf/scientificamerican05281898-
339a.pdf.

Berviss:1898:HJ

[8954] Garrett P. Berviss. The heavens in June. Scientific American, 78(22):339, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n22/pdf/scientificamerican05281898-
339.pdf.

Anonymous:1898:SIF

[8955] Anonymous. Some interesting features of the electrical exhibition. Scientific American, 78(22):340, May 28, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n22/pdf/scientificamerican05281898-340.pdf.

Anonymous:1898:EFM

[8956] Anonymous. An efficient feed mixer. Scientific American, 78(22):341, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-341b.pdf.

Anonymous:1898:IWa

[8957] Anonymous. An improved wrench. Scientific American, 78(22):341, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-341.pdf.

Anonymous:1898:NCH

[8958] Anonymous. A new clothes hanger. Scientific American, 78(22):341, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-341c.pdf.

Anonymous:1898:NTH

[8959] Anonymous. A novel twine holder. Scientific American, 78(22):341, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-341a.pdf.

Anonymous:1898:Cl

[8960] Anonymous. Correspondence. Scientific American, 78(22):342, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-342.pdf.

Anonymous:1898:AS

[8961] Anonymous. Admiral Sampson. Scientific American, 78(22):343–344, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-343.pdf.

Anonymous:1898:TMIa

[8962] Anonymous. The Trans-Mississippi and International Exposition at Omaha. Scientific American, 78(22):346–347, May 28, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n22/pdf/scientificamerican05281898-346.pdf.

Anonymous:1898:RPIv

[8963] Anonymous. Recently patented inventions. Scientific American, 78(22):348, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n22/pdf/scientificamerican05281898-
348.pdf.

Anonymous:1898:BNBr

[8964] Anonymous. Books: New books. Scientific American, 78(22):348–349, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-348a.pdf.

Anonymous:1898:BPv

[8965] Anonymous. Business and personal. Scientific American, 78(22):349, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-349.pdf.

Anonymous:1898:It

[8966] Anonymous. To inventors. Scientific American, 78(22):349, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-349b.pdf.

Anonymous:1898:NQu

[8967] Anonymous. Notes and queries. Scientific American, 78(22):349, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-349a.pdf.

Anonymous:1898:IIv

[8968] Anonymous. Index of inventions. Scientific American, 78(22):349–350, May 28, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n22/pdf/scientificamerican05281898-349c.pdf.

Anonymous:1898:ACN

[8969] Anonymous. The armored cruiser in the Navies of the world. Scientific American, 78(23):354, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
354.pdf.

Anonymous:1898:OSAb

[8970] Anonymous. Obstacles to South American trade. Scientific American, 78(23):354–355, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
354a.pdf.

Anonymous:1898:FSa

[8971] Anonymous. Footgear for the soldier. Scientific American, 78(23):355, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n23/pdf/scientificamerican06041898-355a.pdf.

Harrington:1898:NPR

[8972] M. W. Harrington. Natural products and resources of the Philippine Islands. Scientific American, 78(23):355, June 4, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n23/pdf/scientificamerican06041898-355.pdf.

Anonymous:1898:NH

[8973] Anonymous. A novel hinge. Scientific American, 78(23):356, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n23/pdf/scientificamerican06041898-356.pdf.

Anonymous:1898:NWD

[8974] Anonymous. A novel weighing and delivering machine. Scientific American, 78(23):356, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
356a.pdf.

Anonymous:1898:OB

[8975] Anonymous. The “Olive” bicycle. Scientific American, 78(23):356, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n23/pdf/scientificamerican06041898-356b.pdf.

Anonymous:1898:ICB

[8976] Anonymous. An improved cutter bar. Scientific American, 78 (23):356–357, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
356c.pdf.

Anonymous:1898:EFP

[8977] Anonymous. An efficient fence post driver. Scientific American, 78(23):357, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
357.pdf.

Anonymous:1898:NRA

[8978] Anonymous. A novel ripping attachment for sewing machines. Scientific American, 78(23):357–358, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
357a.pdf.

Anonymous:1898:SPH

[8979] Anonymous. Sugar planters hoe. Scientific American, 78(23):358, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n23/pdf/scientificamerican06041898-358.pdf.

Anonymous:1898:NSE

[8980] Anonymous. New system of electric signs and signals. Scientific American, 78(23):359–360, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
359.pdf.

Anonymous:1898:SSC

[8981] Anonymous. The situation at Santiago De Cuba. Scientific American, 78(23):361–362, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
361.pdf.

Anonymous:1898:VIE

[8982] Anonymous. Vacuum illumination at the electrical exhibition. Scientific American, 78(23):363, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
363.pdf.

Anonymous:1898:BPw

[8983] Anonymous. Business and personal. Scientific American, 78(23):364, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n23/pdf/scientificamerican06041898-364a.pdf.

Anonymous:1898:IIw

[8984] Anonymous. Index of inventions. Scientific American, 78(23):364, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n23/pdf/scientificamerican06041898-364c.pdf.

Anonymous:1898:NQv

[8985] Anonymous. Notes and queries. Scientific American, 78(23):364, June 4, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n23/pdf/scientificamerican06041898-364b.pdf.

Anonymous:1898:RPIw (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n23/pdf/scientificamerican06041898-
364.pdf.

Anonymous:1898:SCa

[8987] Anonymous. The Santiago Campaign. Scientific American, 78(24):370, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-370.pdf.

Anonymous:1898:SPb

[8988] Anonymous. Smokeless powder. Scientific American, 78(24):370, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-370a.pdf.

Anonymous:1898:CMC

[8989] Anonymous. The centrifugal method of collecting plankton, the basis of food supply for aquatic animals. Scientific American, 78(24):370–371, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-370b.pdf.

Anonymous:1898:EPT

[8990] Anonymous. Egg and potato trick. Scientific American, 78(24):372, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-372.pdf.

Anonymous:1898:SWGb

[8991] Anonymous. Steel wire guns. Scientific American, 78(24):372, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-372a.pdf.

Anonymous:1898:NL

[8992] Anonymous. A new lifeboat. Scientific American, 78(24):373, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-373.pdf.

Anonymous:1898:NSD

[8993] Anonymous. A novel spraying device for carbureters. Scientific American, 78(24):373, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n24/pdf/scientificamerican06111898-
373a.pdf.

Anonymous:1898:SVC

[8994] Anonymous. A simplified vapor-condenser for vacuum pans. Scientific American, 78(24):373–374, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n24/pdf/scientificamerican06111898-
373b.pdf.

Anonymous:1898:CSR

[8995] Anonymous. Commodore Schley’s reconnaissance of Santiago Harbor. Scientific American, 78(24):375, June 11, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n24/pdf/scientificamerican06111898-375.pdf.

Anonymous:1898:BEM

[8996] Anonymous. Broadsword exercise on men-of-war. Scientific American, 78(24):376, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n24/pdf/scientificamerican06111898-
376.pdf.

Anonymous:1898:ONT

[8997] Anonymous. Our new thirteen-inch guns. Scientific American, 78 (24):376–377, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n24/pdf/scientificamerican06111898-
376a.pdf.

Holder:1898:SAF

[8998] C. F. Holder. Some American forts. Scientific American, 78 (24):377–378, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n24/pdf/scientificamerican06111898-
377.pdf.

Reynolds:1898:IK

[8999] Cuyler Reynolds. Indian kettles. Scientific American, 78(24):379–380, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-379.pdf.

Anonymous:1898:RPIx

[9000] Anonymous. Recently patented inventions. Scientific American, 78(24):380–381, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n24/pdf/scientificamerican06111898-
380.pdf.

Anonymous:1898:BNBs

[9001] Anonymous. Books: New books, etc. Scientific American, 78(24):381, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-381.pdf.

Anonymous:1898:BPx

[9002] Anonymous. Business and personal. Scientific American, 78(24):381, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-381b.pdf.

Anonymous:1898:Iu

[9003] Anonymous. To inventors. Scientific American, 78(24):381, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-381a.pdf.

Anonymous:1898:IIx

[9004] Anonymous. Index of inventions. Scientific American, 78(24):381, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-381d.pdf.

Anonymous:1898:NQw

[9005] Anonymous. Notes and queries. Scientific American, 78(24):381, June 11, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n24/pdf/scientificamerican06111898-381c.pdf.

Anonymous:1898:TPW

[9006] Anonymous. Trademarks and patents and the War Revenue Bill. Scientific American, 78(25):386, June 18, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n25/pdf/scientificamerican06181898-386.pdf.

Anonymous:1898:CDSb

[9007] Anonymous. Communications during the siege of Paris. Scientific American, 78(25):386–387, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
386a.pdf.

Anonymous:1898:AAM

[9008] Anonymous. Analysis of air by a mushroom. Scientific American, 78(25):387, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
387b.pdf.

Anonymous:1898:DTM

[9009] Anonymous. Departure of troops for Manila. Scientific American, 78(25):387, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
387.pdf.

Anonymous:1898:SFb

[9010] Anonymous. Singing flames. Scientific American, 78(25):387, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-387a.pdf.

Anonymous:1898:EWD

[9011] Anonymous. An efficient wool-drier. Scientific American, 78(25):388, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-388.pdf.

Anonymous:1898:IFc

[9012] Anonymous. An improved flooring. Scientific American, 78(25):388, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-388b.pdf.

Anonymous:1898:NVP

[9013] Anonymous. A novel valve for pneumatic tires. Scientific American, 78(25):388, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
388a.pdf.

Anonymous:1898:NVW

[9014] Anonymous. A new vehicle wheel. Scientific American, 78(25):388, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-388c.pdf.

Anonymous:1898:NREa

[9015] Anonymous. A new rotary engine. Scientific American, 78(25):389, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-389.pdf.

Anonymous:1898:ACW

[9016] Anonymous. Automobile carriage for winter use. Scientific American, 78(25):389–390, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
389a.pdf.

Anonymous:1898:STBb

[9017] Anonymous. The Spanish torpedo-boat destroyer “Terror”. Scientific American, 78(25):391, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
391.pdf.

Anonymous:1898:SCEa

[9018] Anonymous. Santiago De Cuba — its environment and defenses. Scientific American, 78(25):392, June 18, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v78/
n25/pdf/scientificamerican06181898-392.pdf.

Anonymous:1898:GPH

[9019] Anonymous. A great power house at Niagara. Scientific American, 78(25):393–394, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
393.pdf.

Steere:1898:VEP

[9020] J. B. Steere. Volcanoes and earthquakes in the Philippines. Scientific American, 78(25):395, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
395.pdf.

Anonymous:1898:BNBt

[9021] Anonymous. Books: New books, etc. Scientific American, 78(25):396, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-396a.pdf.

Anonymous:1898:RPIy

[9022] Anonymous. Recently patented inventions. Scientific American, 78(25):396, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n25/pdf/scientificamerican06181898-
396.pdf.

Anonymous:1898:BPy

[9023] Anonymous. Business and personal. Scientific American, 78(25):397, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-397.pdf.

Anonymous:1898:Iv

[9024] Anonymous. To inventors. Scientific American, 78(25):397, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-397b.pdf.

Anonymous:1898:IIy

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-397c.pdf.

Anonymous:1898:NQx

[9026] Anonymous. Notes and queries. Scientific American, 78(25):397, June 18, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n25/pdf/scientificamerican06181898-397a.pdf.

Anonymous:1898:OGA

[9027] Anonymous. Our great agricultural supremacy. Scientific American, 78(26):402, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
402.pdf.

Anonymous:1898:PCO

[9028] Anonymous. Proper care for our soldiers. Scientific American, 78(26):402–403, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
402a.pdf.

Anonymous:1898:RPO

[9029] Anonymous. Relief for the Patent Office. Scientific American, 78(26):403, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
403.pdf.

Anonymous:1898:SCEb

[9030] Anonymous. Street cleaning in European and American cities. Scientific American, 78(26):403–404, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
403a.pdf.

Anonymous:1898:CCBa

[9031] Anonymous. A coin-controlled bicycle-pump. Scientific American, 78(26):404, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
404b.pdf.

Anonymous:1898:IDG

[9032] Anonymous. An improved differential gearing. Scientific American, 78(26):404, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
404.pdf.

Anonymous:1898:NREb

[9033] Anonymous. A new rotary engine. Scientific American, 78(26):404, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-404a.pdf.

Hovey:1898:SWR

[9034] E. O. Hovey. A standard of wealth in Russian Armenia. Scientific American, 78(26):405–406, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
405.pdf.

Anonymous:1898:ITH

[9035] Anonymous. An improved trace-holder. Scientific American, 78 (26):406, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
406.pdf.

Anonymous:1898:NSH

[9036] Anonymous. A novel sash-holder. Scientific American, 78(26):406, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-406a.pdf.

Steere:1898:WTP

[9037] J. B. Steere. The wild tribes of the Philippines. Scientific American, 78(26):407, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
407.pdf.

Dailey:1898:BE

[9038] H. B. Dailey. A beautiful experiment. Scientific American, 78(26):408, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-408.pdf.

Anonymous:1898:BBD

[9039] Anonymous. The boilers and bulkhead doors of the “Chicago”. Scientific American, 78(26):408–410, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
408a.pdf.

Anonymous:1898:BPz

[9040] Anonymous. Business and personal. Scientific American, 78(26):411, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-411a.pdf.

Anonymous:1898:Ix

[9041] Anonymous. To inventors. Scientific American, 78(26):411, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-411c.pdf.

Anonymous:1898:NQy

[9042] Anonymous. Notes and queries. Scientific American, 78(26):411, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-411b.pdf.

Anonymous:1898:RPIz

[9043] Anonymous. Recently patented inventions. Scientific American, 78(26):411, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v78/n26/pdf/scientificamerican06251898-
411.pdf.

Anonymous:1898:IIz

[9044] Anonymous. Index of inventions. Scientific American, 78(26):412, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-412.pdf.

Anonymous:1898:Iy

[9045] Anonymous. Index. Scientific American, 78(26):413–414, June 25, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v78/n26/pdf/scientificamerican06251898-413.pdf.

Anonymous:1898:DP

[9046] Anonymous. Diver’s paralysis. Scientific American, 79(1):4, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-4b.pdf.

Anonymous:1898:HAM

[9047] Anonymous. How the ax may be made the preserver of the forest. Scientific American, 79(1):4, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n1/pdf/scientificamerican07021898-
4a.pdf.

Anonymous:1898:NPPb

[9048] Anonymous. Negotiable paper for patent rights. Scientific American, 79(1):4, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n1/pdf/scientificamerican07021898-
4.pdf.

Anonymous:1898:RSC

[9049] Anonymous. Russia’s strategic canal. Scientific American, 79(1):4–5, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-4c.pdf.

Anonymous:1898:NLW

[9050] Anonymous. New light weight uniforms. Scientific American, 79(1):5, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-5b.pdf.

Anonymous:1898:OCC

[9051] Anonymous. Our commerce with China. Scientific American, 79(1):5, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-5.pdf.

Serviss:1898:HJ

[9052] Garrett P. Serviss. The heavens in July. Scientific American, 79(1):5, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-5a.pdf.

Anonymous:1898:CSBa

[9053] Anonymous. Car-seal breaker. Scientific American, 79(1):6, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-6b.pdf.

Anonymous:1898:ICM

[9054] Anonymous. An improved cover for milk-cans. Scientific American, 79(1):6, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n1/pdf/scientificamerican07021898-
6c.pdf.

Anonymous:1898:NWb

[9055] Anonymous. A novel wrench. Scientific American, 79(1):6, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-6.pdf.

Anonymous:1898:RWS

[9056] Anonymous. The Ralston water-still. Scientific American, 79(1):6, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-6a.pdf.

Anonymous:1898:NDG

[9057] Anonymous. A new disappearing gun carriage. Scientific American, 79(1):7, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n1/pdf/scientificamerican07021898-
7a.pdf.

Anonymous:1898:SSTb

[9058] Anonymous. The Spanish submarine torpedo boat “Peral”. Scientific American, 79(1):7, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n1/pdf/scientificamerican07021898-
7.pdf.

Anonymous:1898:CIa

[9059] Anonymous. The Cuban Invasion. Scientific American, 79(1):8, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-8.pdf.

Anonymous:1898:Cm

[9060] Anonymous. Correspondence. Scientific American, 79(1):9, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-9.pdf.

Anonymous:1898:ARC

[9061] Anonymous. Accident to the revenue cutter “Gresham”. Scientific American, 79(1):10, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n1/pdf/scientificamerican07021898-
10.pdf.

Anonymous:1898:HSFa

[9062] Anonymous. How the Spanish fortify the country in Cuba. Scientific American, 79(1):11, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n1/pdf/scientificamerican07021898-
11.pdf.

Anonymous:1898:BNBu

[9063] Anonymous. Books: New books, etc. Scientific American, 79(1):12, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-12a.pdf.

Anonymous:1898:BPba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-12b.pdf.

Anonymous:1898:Iz

[9065] Anonymous. To inventors. Scientific American, 79(1):12, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-12d.pdf.

Anonymous:1898:NQz

[9066] Anonymous. Notes and queries. Scientific American, 79(1):12, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-12c.pdf.

Anonymous:1898:RPIba

[9067] Anonymous. Recently patented inventions. Scientific American, 79(1): 12, July 2, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n1/pdf/scientificamerican07021898-12.pdf.

Anonymous:1898:CMW

[9068] Anonymous. Courage in modern warfare. Scientific American, 79(2):18, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-18a.pdf.

Anonymous:1898:OAC

[9069] Anonymous. Our Army and coast defenses. Scientific American, 79(2): 18, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-18.pdf.

Gerard:1898:INO

[9070] W. R. Gerard. The Indian names of our war vessels. Scientific American, 79(2):19, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n2/pdf/scientificamerican07091898-
19.pdf.

Anonymous:1898:ASH

[9071] Anonymous. An automatic steering head for bicycles. Scientific American, 79(2):20, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n2/pdf/scientificamerican07091898-
20.pdf.

Anonymous:1898:NPT

[9072] Anonymous. A novel pneumatic tire. Scientific American, 79(2):20, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-20a.pdf.

Anonymous:1898:NSB

[9073] Anonymous. A new steam boiler and furnace. Scientific American, 79(2):20, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n2/pdf/scientificamerican07091898-
20b.pdf.

Anonymous:1898:SNW

[9074] Anonymous. Something new in wood-working machinery. Scientific American, 79(2):21–22, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n2/pdf/scientificamerican07091898-
21.pdf.

Anonymous:1898:Cn

[9075] Anonymous. Correspondence. Scientific American, 79(2):23, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-23.pdf.

Anonymous:1898:GG

[9076] Anonymous. Gardening over a geyser. Scientific American, 79(2):24, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-24a.pdf.

Anonymous:1898:IGa

[9077] Anonymous. An improved gate. Scientific American, 79(2):24, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-24.pdf.

Anonymous:1898:RDSb

[9078] Anonymous. Reinforcement of the Dewey squadron. Scientific American, 79(2):25–26, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n2/pdf/scientificamerican07091898-
25.pdf.

Anonymous:1898:MIR

[9079] Anonymous. Morris, the “India Rubber Man”. Scientific American, 79(2):26, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n2/pdf/scientificamerican07091898-
26.pdf.

Anonymous:1898:MPG

[9080] Anonymous. The manufacture of Pintsch gas. Scientific American, 79 (2):27–28, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic).  URL http://en.wikipedia.org/wiki/
Pintsch_gas;   http://www.nature.com/scientificamerican/journal/
v79/n2/pdf/scientificamerican07091898-27.pdf.

Anonymous:1898:RPIbb

[9081] Anonymous. Recently patented inventions. Scientific American, 79(2):28–29, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n2/pdf/scientificamerican07091898-
28.pdf.

Anonymous:1898:BNBv

[9082] Anonymous. Books: New books, etc. Scientific American, 79(2):29, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-29b.pdf.

Anonymous:1898:BPbb

[9083] Anonymous. Business and personal. Scientific American, 79(2):29, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-29.pdf.

Anonymous:1898:Iba

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-29c.pdf.

Anonymous:1898:IIba

[9085] Anonymous. Index of inventions. Scientific American, 79(2):29, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-29d.pdf.

Anonymous:1898:NQba

[9086] Anonymous. Notes and queries. Scientific American, 79(2):29, July 9, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n2/pdf/scientificamerican07091898-29a.pdf.

Anonymous:1898:DCF

[9087] Anonymous. The destruction of Cervera’s Fleet. Scientific American, 79(3):34, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
34.pdf.

Anonymous:1898:OAS

[9088] Anonymous. Our Army at Santiago. Scientific American, 79(3):34, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-34a.pdf.

Anonymous:1898:LB

[9089] Anonymous. The loss of the “Bourgogne”. Scientific American, 79(3):34–35, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
34b.pdf.

Anonymous:1898:EAA

[9090] Anonymous. “The engineer’s” analysis of American and Spanish warships. Scientific American, 79(3):35, July 16, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n3/pdf/scientificamerican07161898-35b.pdf.

Anonymous:1898:HSR

[9091] Anonymous. The hospital ship “Relief”. Scientific American, 79(3):35, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-35a.pdf.

Anonymous:1898:OFC

[9092] Anonymous. Our foreign commerce. Scientific American, 79(3):35, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-35.pdf.

Anonymous:1898:IPb

[9093] Anonymous. An improved plow. Scientific American, 79(3):36, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-36.pdf.

Anonymous:1898:NBBc

[9094] Anonymous. A new ball-bearing. Scientific American, 79(3):36, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-36a.pdf.

Anonymous:1898:SMH

[9095] Anonymous. A simple magazine hand camera. Scientific American, 79(3):36, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
36b.pdf.

Anonymous:1898:CWS

[9096] Anonymous. Care of the wounded in sea fights. Scientific American, 79(3):37, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
37.pdf.

Anonymous:1898:RWC

[9097] Anonymous. Roman wine casks discovered at Silchester. Scientific American, 79(3):37–38, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
37a.pdf.

Anonymous:1898:GLM

[9098] Anonymous. Geared locomotives on the Mount Tamalpais Railway. Scientific American, 79(3):39, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
39.pdf.

Weir:1898:TIS

[9099] James Weir, Jr. Two insect Samsons. Scientific American, 79(3):40, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-40.pdf.

Anonymous:1898:USAb

[9100] Anonymous. The United States armored cruiser “Brooklyn”. Scientific American, 79(3):40–42, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
40a.pdf.

Anonymous:1898:OTP

[9101] Anonymous. The Olympian Theater of Palladio at Vicenza. Scientific American, 79(3):43–44, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
43.pdf.

Anonymous:1898:RPIbc

[9102] Anonymous. Recently patented inventions. Scientific American, 79(3):44, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n3/pdf/scientificamerican07161898-
44.pdf.

Anonymous:1898:BPbc

[9103] Anonymous. Business and personal. Scientific American, 79(3):45, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-45.pdf.

Anonymous:1898:Ibc

[9104] Anonymous. To inventors. Scientific American, 79(3):45, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-45b.pdf.

Anonymous:1898:IIbb

[9105] Anonymous. Index of inventions. Scientific American, 79(3):45, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-45c.pdf.

Anonymous:1898:NQbb

[9106] Anonymous. Notes and queries. Scientific American, 79(3):45, July 16, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n3/pdf/scientificamerican07161898-45a.pdf.

Anonymous:1898:LSN

[9107] Anonymous. Lessons of the Santiago Naval Engagement. Scientific American, 79(4):50, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
50.pdf.

Anonymous:1898:CTb

[9108] Anonymous. Campaigning in the tropics. Scientific American, 79(4):50–51, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
50a.pdf.

Anonymous:1898:FSb

[9109] Anonymous. The fall of Santiago. Scientific American, 79(4):51, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-51.pdf.

Anonymous:1898:TIDb

[9110] Anonymous. Two interesting decisions. Scientific American, 79(4):51, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-51a.pdf.

Sajous:1898:RSC

[9111] M. Sajous. Responsibility of steamship companies in marine disasters. Scientific American, 79(4):51–52, July 23, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n4/pdf/scientificamerican07231898-51b.pdf.

Anonymous:1898:CTH

[9112] Anonymous. Combined ticket holder and register. Scientific American, 79(4):52, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
52.pdf.

Bolling:1898:APP

[9113] Randolph Bolling. Apparatus for preserving pyrogallic acid solutions. Scientific American, 79(4):52, July 23, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n4/pdf/scientificamerican07231898-52a.pdf.

Anonymous:1898:SHC

[9114] Anonymous. The “Simplex” hose clip. Scientific American, 79(4):52–53, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-52b.pdf.

Anonymous:1898:RBA

[9115] Anonymous. The roller boat that actually went to sea. Scientific American, 79(4):53, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
53.pdf.

Anonymous:1898:RCN

[9116] Anonymous. The reconstructed cruiser “Newark”. Scientific American, 79(4):54, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
54.pdf.

Anonymous:1898:Co

[9117] Anonymous. Correspondence. Scientific American, 79(4):55, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-55.pdf.

Anonymous:1898:IBCa

[9118] Anonymous. An improved bottle-closure. Scientific American, 79(4):55, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-55a.pdf.

Anonymous:1898:RB

[9119] Anonymous. Regimental baggage. Scientific American, 79(4):56, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-56.pdf.

Anonymous:1898:RCS

[9120] Anonymous. The Red Cross Society and the Ambulance Service. Scientific American, 79(4):56–57, July 23, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n4/pdf/scientificamerican07231898-56a.pdf.

Anonymous:1898:RAB

[9121] Anonymous. A railroad attachment for bicycles. Scientific American, 79(4):57, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
57.pdf.

Anonymous:1898:RGL

[9122] Anonymous. Recovery of gold from low grade ores. Scientific American, 79(4):57–59, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
57a.pdf.

Caulk:1898:MW

[9123] W. B. Caulk. The miraculous wineglasses. Scientific American, 79(4):59–60, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
59.pdf.

Anonymous:1898:BNBw

[9124] Anonymous. Books: New books, etc. Scientific American, 79(4):61, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-61a.pdf.

Anonymous:1898:RPIbd

[9125] Anonymous. Recently patented inventions. Scientific American, 79(4):61, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n4/pdf/scientificamerican07231898-
61.pdf.

Anonymous:1898:BPbd

[9126] Anonymous. Business and personal. Scientific American, 79(4):62, July 23, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n4/pdf/scientificamerican07231898-62.pdf.

Anonymous:1898:CMS

[9127] Anonymous. The cinematograph in medicine and surgery. Scientific American, 79(5):66, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
66b.pdf.

Anonymous:1898:LFS

[9128] Anonymous. London’s fire system revolutionized. Scientific American, 79(5):66, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
66a.pdf.

Anonymous:1898:PSU

[9129] Anonymous. The present status of the United States Navy. Scientific American, 79(5):66, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
66.pdf.

Turpin:1898:PL

[9130] Perry B. Turpin. Prints and labels. Scientific American, 79(5):66–67, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-66c.pdf.

Anonymous:1898:DWB

[9131] Anonymous. Destruction of a war balloon. Scientific American, 79(5):67, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
67c.pdf.

Anonymous:1898:EXR

[9132] Anonymous. Effect of X rays on colors. Scientific American, 79(5):67, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-67d.pdf.

Anonymous:1898:NCP

[9133] Anonymous. A new center of the plague. Scientific American, 79(5):67, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-67b.pdf.

Anonymous:1898:VCK

[9134] Anonymous. The visit of Colonel Krag. Scientific American, 79(5):67, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-67.pdf.

Serviss:1898:HAb

[9135] Garrett P. Serviss. The heavens in August. Scientific American, 79(5):67, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
67a.pdf.

Anonymous:1898:SJD

[9136] Anonymous. San Juan and its defenses. Scientific American, 79(5):68–69, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
68.pdf.

Anonymous:1898:IOR

[9137] Anonymous. An improved oil-retaining box. Scientific American, 79(5):69, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
69.pdf.

Anonymous:1898:NHC

[9138] Anonymous. A novel house-cooler and ventilator. Scientific American, 79(5):69, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
69a.pdf.

Anonymous:1898:NFC

[9139] Anonymous. A new flue-cleaner. Scientific American, 79(5):69–71, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-69b.pdf.

Anonymous:1898:FSF

[9140] Anonymous. The first Spanish flag captured at Manila. Scientific American, 79(5):71, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
71.pdf.

Anonymous:1898:SSL

[9141] Anonymous. Simple section liner. Scientific American, 79(5):72, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-72.pdf.

Anonymous:1898:RCFb

[9142] Anonymous. The remains of Cervera’s fleet. Scientific American, 79(5):72–74, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
72a.pdf.

Mindeleff:1898:PAIc

[9143] Cosmos Mindeleff. Pueblo arts and industries — III. Scientific American, 79(5):75–76, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
75.pdf.

Anonymous:1898:BPbe

[9144] Anonymous. Business and personal. Scientific American, 79(5):77, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-77a.pdf.

Anonymous:1898:NQbc

[9145] Anonymous. Notes and queries. Scientific American, 79(5):77, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-77b.pdf.

Anonymous:1898:RPIbe

[9146] Anonymous. Recently patented inventions. Scientific American, 79(5):77, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n5/pdf/scientificamerican07301898-
77.pdf.

Anonymous:1898:IIbc

[9147] Anonymous. Index of inventions. Scientific American, 79(5):77–78, July 30, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n5/pdf/scientificamerican07301898-77c.pdf.

Anonymous:1898:HSO

[9148] Anonymous. Higher speed for our new battleships. Scientific American, 79(6):82, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
82.pdf.

Anonymous:1898:MHSb

[9149] Anonymous. Make haste slowly. Scientific American, 79(6):82, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-82a.pdf.

Anonymous:1898:PCDb

[9150] Anonymous. To prevent collisions during a fog at sea. Scientific American, 79(6):82–83, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
82b.pdf.

Anonymous:1898:SLS

[9151] Anonymous. Spain’s losses on the sea. Scientific American, 79(6):83, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-83a.pdf.

Hovey:1898:ASJ

[9152] Horace C. Hovey. The approaching scientific jubilee. Scientific American, 79(6):83, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
83.pdf.

Anonymous:1898:ITZ

[9153] Anonymous. An improved type of Zeiss field glass. Scientific American, 79(6):84, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
84.pdf.

Anonymous:1898:NI

[9154] Anonymous. A new insulator. Scientific American, 79(6):84, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-84a.pdf.

Anonymous:1898:CAA

[9155] Anonymous. The Criterion automatic acetylene house generator. Scientific American, 79(6):84–85, August 6, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n6/pdf/scientificamerican08061898-84b.pdf.

Anonymous:1898:NSMa

[9156] Anonymous. A new steam motor. Scientific American, 79(6):85, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-85.pdf.

Anonymous:1898:Cp

[9157] Anonymous. Correspondence. Scientific American, 79(6):86, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-86.pdf.

Anonymous:1898:ILP

[9158] Anonymous. An improved lead pencil. Scientific American, 79(6):86–87, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-86a.pdf.

Anonymous:1898:OAN

[9159] Anonymous. Our Army and Navy hospital ships. Scientific American, 79(6):88, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
88.pdf.

Anonymous:1898:RWM

[9160] Anonymous. A relic from the wreck of the “Maine”. Scientific American, 79(6):88, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
88a.pdf.

Anonymous:1898:IDS

[9161] Anonymous. The 3.2-inch Driggs–Seabury field-gun. Scientific American, 79(6):89, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
89.pdf.

Anonymous:1898:GNP

[9162] Anonymous. A group of Navy projectiles. Scientific American, 79(6):89–90, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
89a.pdf.

Anonymous:1898:ASR

[9163] Anonymous. Admiral Sampson’s report of the Santiago Engagement. Scientific American, 79(6):90–91, August 6, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n6/pdf/scientificamerican08061898-90.pdf.

Anonymous:1898:BPbf

[9164] Anonymous. Business and personal. Scientific American, 79(6):92, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-92a.pdf.

Anonymous:1898:NQbd

[9165] Anonymous. Notes and queries. Scientific American, 79(6):92, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n6/pdf/scientificamerican08061898-92b.pdf.

Anonymous:1898:RPIbf

[9166] Anonymous. Recently patented inventions. Scientific American, 79(6):92, August 6, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n6/pdf/scientificamerican08061898-
92.pdf.

Anonymous:1898:SLO

[9167] Anonymous. Spain’s lost opportunities. Scientific American, 79 (7):98, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
98a.pdf.

Anonymous:1898:ST

[9168] Anonymous. The Santiago Tragedy. Scientific American, 79(7):98, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-98.pdf.

Anonymous:1898:EPM

[9169] Anonymous. Extension of pneumatic mail tubes to Brooklyn. Scientific American, 79(7):99, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
99c.pdf.

Anonymous:1898:GS

[9170] Anonymous. Gold from the sea. Scientific American, 79(7):99, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-99.pdf.

Anonymous:1898:LE

[9171] Anonymous. Language of the eye. Scientific American, 79(7):99, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-99a.pdf.

Anonymous:1898:PCS

[9172] Anonymous. Prevention of collisions at sea. Scientific American, 79(7):99, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
99d.pdf.

Anonymous:1898:WR

[9173] Anonymous. A wise revision. Scientific American, 79(7):99, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-99b.pdf.

Anonymous:1898:ISP

[9174] Anonymous. An improved stencil perforating machine. Scientific American, 79(7):100, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
100.pdf.

Anonymous:1898:NEM

[9175] Anonymous. A novel envelop moistener and sealer. Scientific American, 79(7):100, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
100b.pdf.

Anonymous:1898:TFR

[9176] Anonymous. Traveling freight ramps. Scientific American, 79 (7):100, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
100a.pdf.

Anonymous:1898:BPIa

[9177] Anonymous. The Bradley Polytechnic Institute and School of Horology. Scientific American, 79(7):101–103, August 13, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n7/pdf/scientificamerican08131898-101.pdf.

Anonymous:1898:ECD

[9178] Anonymous. Ericsson’s coast-defense monitor “Nahant”. Scientific American, 79(7):104–105, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
104.pdf.

Anonymous:1898:ABB

[9179] Anonymous. The accident to the Brooklyn Bridge. Scientific American, 79(7):106, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
106.pdf.

Anonymous:1898:MGM

[9180] Anonymous. Major-Generals Miles and Wheeler. Scientific American, 79(7):107, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
107.pdf.

Caulk:1898:CC

[9181] W. B. Caulk. Card Cricket. Scientific American, 79(7):107–108, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-107a.pdf.

Anonymous:1898:BNBx

[9182] Anonymous. Books: New books, etc. Scientific American, 79 (7):109, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
109a.pdf.

Anonymous:1898:BPbg

[9183] Anonymous. Business and personal. Scientific American, 79(7):109, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-109b.pdf.

Anonymous:1898:Ibd

[9184] Anonymous. To inventors. Scientific American, 79(7):109, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-109d.pdf.

Anonymous:1898:NQbe

[9185] Anonymous. Notes and queries. Scientific American, 79(7):109, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n7/pdf/scientificamerican08131898-109c.pdf.

Anonymous:1898:RPIbg

[9186] Anonymous. Recently patented inventions. Scientific American, 79(7):109, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
109.pdf.

Anonymous:1898:IIbd

[9187] Anonymous. Index of inventions. Scientific American, 79(7): 109–110, August 13, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n7/pdf/scientificamerican08131898-
109e.pdf.

Anonymous:1898:DDS

[9188] Anonymous. Discipline and disasters at sea. Scientific American, 79(8):114, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
114b.pdf.

Anonymous:1898:NEA

[9189] Anonymous. The new elements in the air. Scientific American, 79(8):114, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
114.pdf.

Anonymous:1898:PHW

[9190] Anonymous. The percentage of hits in warfare. Scientific American, 79(8):114, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
114a.pdf.

Anonymous:1898:CPa

[9191] Anonymous. Concentration of power. Scientific American, 79 (8):115, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
115.pdf.

Anonymous:1898:PRC

[9192] Anonymous. Porto Rico’s commerce. Scientific American, 79 (8):115, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
115b.pdf.

Anonymous:1898:VN

[9193] Anonymous. The value of a name. Scientific American, 79(8):115, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n8/pdf/scientificamerican08201898-115a.pdf.

Anonymous:1898:HGR

[9194] Anonymous. The Holly gravity return system. Scientific American, 79(8):116, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
116.pdf.

Anonymous:1898:GAM

[9195] Anonymous. Guns, ancient and modern, at Santiago. Scientific American, 79(8):117–118, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
117.pdf.

Anonymous:1898:NWMb

[9196] Anonymous. A novel water motor. Scientific American, 79(8):118, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n8/pdf/scientificamerican08201898-118.pdf.

Anonymous:1898:CLS

[9197] Anonymous. The Colomés life and ship protector. Scientific American, 79(8):119, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
119.pdf.

Anonymous:1898:TAb

[9198] Anonymous. The “Texas” in action. Scientific American, 79 (8):120–122, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
120.pdf.

Hovey:1898:PMR

[9199] E. O. Hovey. Primitive methods of raising wheat and baking bread in Transcaucasia and Armenia. Scientific American, 79(8):122–124, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n8/pdf/scientificamerican08201898-122.pdf.

Anonymous:1898:RPIbh

[9200] Anonymous. Recently patented inventions. Scientific American, 79 (8):124–125, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
124.pdf.

Anonymous:1898:BNBy

[9201] Anonymous. Books: New books, etc. Scientific American, 79 (8):125, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
125b.pdf.

Anonymous:1898:BPbh

[9202] Anonymous. Business and personal. Scientific American, 79(8):125, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n8/pdf/scientificamerican08201898-125.pdf.

Anonymous:1898:Ibe

[9203] Anonymous. To inventors. Scientific American, 79(8):125, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n8/pdf/scientificamerican08201898-125c.pdf.

Anonymous:1898:NQbf

[9204] Anonymous. Notes and queries. Scientific American, 79(8):125, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n8/pdf/scientificamerican08201898-125a.pdf.

Anonymous:1898:IIbe

[9205] Anonymous. Index of inventions. Scientific American, 79(8): 125–126, August 20, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n8/pdf/scientificamerican08201898-
125d.pdf.

Anonymous:1898:IWB

[9206] Anonymous. The increasing weight of the bicycle. Scientific American, 79(9):130, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
130.pdf.

Anonymous:1898:PIN

[9207] Anonymous. Probable increase of the Navy. Scientific American, 79(9):130, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
130a.pdf.

Anonymous:1898:EEO

[9208] Anonymous. Engines and engineers of the “Oregon”. Scientific American, 79(9):130–131, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
130b.pdf.

Benjamin:1898:FWP

[9209] Marcus Benjamin. Frederic Ward Putnam. Scientific American, 79(9):131, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
131.pdf.

Anonymous:1898:NSP

[9210] Anonymous. The new smokeless powder for Army and Navy. Scientific American, 79(9):131–132, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
131a.pdf.

Anonymous:1898:NSMb

[9211] Anonymous. A new sampling machine. Scientific American, 79 (9):132, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
132.pdf.

Anonymous:1898:SMa

[9212] Anonymous. A scrubbing machine. Scientific American, 79(9):132, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n9/pdf/scientificamerican08271898-132a.pdf.

Anonymous:1898:ISR

[9213] Anonymous. An improved seam-ripper. Scientific American, 79 (9):132–133, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
132b.pdf.

Anonymous:1898:GSG

[9214] Anonymous. The great sea-going dredges on the Mersey Bar, Liverpool. Scientific American, 79(9):133–135, August 27, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n9/pdf/scientificamerican08271898-133.pdf.

Anonymous:1898:TSI

[9215] Anonymous. Test of a six-inch experimental Krupp plate. Scientific American, 79(9):136–137, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
136.pdf.

Anonymous:1898:A

[9216] Anonymous. The “Avion”. Scientific American, 79(9):137–138, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n9/pdf/scientificamerican08271898-137.pdf.

Anonymous:1898:TMIb

[9217] Anonymous. The Trans-Mississippi and International Exposition. Scientific American, 79(9):138–139, August 27, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n9/pdf/scientificamerican08271898-138.pdf.

Robinson:1898:CLT

[9218] W. E. Robinson. “Cupid lighter than a butterfly”. Scientific American, 79(9):139, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
139.pdf.

Serviss:1898:HS

[9219] Garrett P. Serviss. The heavens in September. Scientific American, 79(9):140, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
140.pdf.

Anonymous:1898:BNBz

[9220] Anonymous. Books: New books, etc. Scientific American, 79 (9):141, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
141a.pdf.

Anonymous:1898:BPbi

[9221] Anonymous. Business and personal. Scientific American, 79(9):141, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n9/pdf/scientificamerican08271898-141b.pdf.

Anonymous:1898:Ibf

[9222] Anonymous. To inventors. Scientific American, 79(9):141, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n9/pdf/scientificamerican08271898-141d.pdf.

Anonymous:1898:NQbg

[9223] Anonymous. Notes and queries. Scientific American, 79(9):141, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n9/pdf/scientificamerican08271898-141c.pdf.

Anonymous:1898:RPIbi

[9224] Anonymous. Recently patented inventions. Scientific American, 79(9):141, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
141.pdf.

Anonymous:1898:IIbf

[9225] Anonymous. Index of inventions. Scientific American, 79(9): 141–142, August 27, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n9/pdf/scientificamerican08271898-
141e.pdf.

Anonymous:1898:RUS

[9226] Anonymous. The railways of the United States. Scientific American, 79(10):146, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
146b.pdf.

Anonymous:1898:SMb

[9227] Anonymous. The ship and the man. Scientific American, 79 (10):146, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
146a.pdf.

Anonymous:1898:TL

[9228] Anonymous. Turn on the light. Scientific American, 79(10): 146, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
146.pdf.

Hovey:1898:GAS

[9229] Horace C. Hovey. Golden Anniversary of Science. Scientific American, 79(10):147–148, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
147.pdf.

Anonymous:1898:NWR

[9230] Anonymous. A novel window raising and locking device. Scientific American, 79(10):148, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
148.pdf.

Anonymous:1898:RVF

[9231] Anonymous. Return of the victorious Fleet from Cuba. Scientific American, 79(10):148–150, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
148a.pdf.

Anonymous:1898:Cq

[9232] Anonymous. Correspondence. Scientific American, 79(10):151, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n10/pdf/scientificamerican09031898-151.pdf.

Anonymous:1898:SPAa

[9233] Anonymous. Some of the properties of acetylene gas. Scientific American, 79(10):151, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
151a.pdf.

Anonymous:1898:TFL

[9234] Anonymous. The two fastest long distance runs without stop. Scientific American, 79(10):152–153, September 3, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n10/pdf/scientificamerican09031898-152.pdf.

Reynolds:1898:PJH

[9235] Cuyler Reynolds. Prof. James Hall. Scientific American, 79 (10):154, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
154.pdf.

Anonymous:1898:EF

[9236] Anonymous. The “Educated Fly”. Scientific American, 79(10): 155, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
155.pdf.

Anonymous:1898:BPbj

[9237] Anonymous. Business and personal. Scientific American, 79 (10):156, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
156a.pdf.

Anonymous:1898:RPIbj

[9238] Anonymous. Recently patented inventions. Scientific American, 79(10):156, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
156.pdf.

Anonymous:1898:NQbh

[9239] Anonymous. Notes and queries. Scientific American, 79(10): 156–157, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
156b.pdf.

Anonymous:1898:BNBba

[9240] Anonymous. Books: New books, etc. Scientific American, 79 (10):157, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
157.pdf.

Anonymous:1898:Ibg

[9241] Anonymous. To inventors. Scientific American, 79(10):157, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n10/pdf/scientificamerican09031898-157a.pdf.

Anonymous:1898:IIbg

[9242] Anonymous. Index of inventions. Scientific American, 79(10): 157–158, September 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n10/pdf/scientificamerican09031898-
157b.pdf.

Anonymous:1898:APE

[9243] Anonymous. American progress in English industries. Scientific American, 79(11):162, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
162a.pdf.

Anonymous:1898:WDO

[9244] Anonymous. The War Department and our new foreign policy. Scientific American, 79(11):162, September 10, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n11/pdf/scientificamerican09101898-162.pdf.

Anonymous:1898:AFb

[9245] Anonymous. Artificial foods. Scientific American, 79(11):162– 163, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
162b.pdf.

Anonymous:1898:AON

[9246] Anonymous. The armament of our new warships. Scientific American, 79(11):163, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
163.pdf.

Anonymous:1898:LPa

[9247] Anonymous. Languages of the Philippines. Scientific American, 79(11):163, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
163a.pdf.

Anonymous:1898:URC

[9248] Anonymous. The undeveloped resources of Cuba. Scientific American, 79(11):163, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
163b.pdf.

Anonymous:1898:FCb

[9249] Anonymous. A folding cradle. Scientific American, 79(11):164, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n11/pdf/scientificamerican09101898-164a.pdf.

Anonymous:1898:IBBc

[9250] Anonymous. An improved ball-bearing. Scientific American, 79 (11):164, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
164.pdf.

Anonymous:1898:PPBa

[9251] Anonymous. A phenomenal piece of bronze casting. Scientific American, 79(11):164–167, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
164b.pdf.

Anonymous:1898:GET

[9252] Anonymous. The Government Exhibit at the Trans-Mississippi and International Exposition. Scientific American, 79(11):168–169, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n11/pdf/scientificamerican09101898-168.pdf.

Anonymous:1898:ORW

[9253] Anonymous. The official report on the wrecked Spanish warships. Scientific American, 79(11):170–171, September 10, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n11/pdf/scientificamerican09101898-170.pdf.

Anonymous:1898:G

[9254] Anonymous. The gundalow. Scientific American, 79(11):171–172, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://en.wikipedia.org/wiki/Gundalow;
 http://www.nature.com/scientificamerican/journal/v79/n11/
pdf/scientificamerican09101898-171.pdf.

Anonymous:1898:BNBbb

[9255] Anonymous. Books: New books, etc. Scientific American, 79 (11):173, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
173c.pdf.

Anonymous:1898:BPbk

[9256] Anonymous. Business and personal. Scientific American, 79 (11):173, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
173a.pdf.

Anonymous:1898:NQbi

[9257] Anonymous. Notes and queries. Scientific American, 79(11): 173, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
173b.pdf.

Anonymous:1898:RPIbk

[9258] Anonymous. Recently patented inventions. Scientific American, 79 (11):173, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
173.pdf.

Anonymous:1898:IIbh

[9259] Anonymous. Index of inventions. Scientific American, 79(11): 174, September 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n11/pdf/scientificamerican09101898-
174.pdf.

Anonymous:1898:BF

[9260] Anonymous. The bicycle frame. Scientific American, 79(12): 178, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
178a.pdf.

Anonymous:1898:WBF

[9261] Anonymous. Wind bracing in framed buildings. Scientific American, 79(12):178, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
178.pdf.

Anonymous:1898:WSC

[9262] Anonymous. The water supply at camp Thomas. Scientific American, 79(12):178–179, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
178b.pdf.

Hovey:1898:GMA

[9263] E. O. Hovey. Geology at the meeting of the American Association for the Advancement of Science. Scientific American, 79(12):179, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n12/pdf/scientificamerican09171898-179.pdf.

Anonymous:1898:GRP

[9264] Anonymous. A great raft of piles on the Pacific. Scientific American, 79(12):180, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
180a.pdf.

Anonymous:1898:IWSb

[9265] Anonymous. An improved window-sash-lift. Scientific American, 79(12):180, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
180b.pdf.

Anonymous:1898:NSC

[9266] Anonymous. A novel self-closing gate. Scientific American, 79 (12):180, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
180.pdf.

Holder:1898:RSR

[9267] C. F. Holder. A remarkable stage road. Scientific American, 79 (12):181–182, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
181.pdf.

Anonymous:1898:Cr

[9268] Anonymous. Correspondence. Scientific American, 79(12):182, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n12/pdf/scientificamerican09171898-182.pdf.

Anonymous:1898:OCD

[9269] Anonymous. An opening and closing device for fireboxes. Scientific American, 79(12):183, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
183.pdf.

Steere:1898:CIP

[9270] J. B. Steere. The civilized Indians of the Philippines. Scientific American, 79(12):184–187, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
184.pdf.

Anonymous:1898:SPc

[9271] Anonymous. The surprise pen. Scientific American, 79(12): 187, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
187.pdf.

Anonymous:1898:NWC

[9272] Anonymous. A new way of coaling locomotive-tenders. Scientific American, 79(12):187–188, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
187a.pdf.

Anonymous:1898:RPIbl

[9273] Anonymous. Recently patented inventions. Scientific American, 79 (12):188–189, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
188.pdf.

Anonymous:1898:BNBbc

[9274] Anonymous. Books: New books, etc. Scientific American, 79 (12):189, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
189.pdf.

Anonymous:1898:BPbl

[9275] Anonymous. Business and personal. Scientific American, 79 (12):189, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
189a.pdf.

Anonymous:1898:IIbi

[9276] Anonymous. Index of inventions. Scientific American, 79(12): 189, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
189c.pdf.

Anonymous:1898:NQbj

[9277] Anonymous. Notes and queries. Scientific American, 79(12): 189, September 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n12/pdf/scientificamerican09171898-
189b.pdf.

Hovey:1898:TSR

[9278] Horace C. Hovey. The Trans-Siberian Rail-Road. Scientific American, 79(13):193, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
193.pdf.

Anonymous:1898:DAC

[9279] Anonymous. Dangers attending so-called food preservatives. Scientific American, 79(13):194, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
194b.pdf.

Anonymous:1898:FND

[9280] Anonymous. Forethought in the Navy department. Scientific American, 79(13):194, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
194a.pdf.

Anonymous:1898:ONP

[9281] Anonymous. Our new possessions as a field for engineers. Scientific American, 79(13):194, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
194.pdf.

Anonymous:1898:OSAc

[9282] Anonymous. Obstacles to South American trade. Scientific American, 79(13):194, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
194c.pdf.

Anonymous:1898:DJH

[9283] Anonymous. Dr. John Hopkinson. Scientific American, 79(13): 195, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
195a.pdf.

Anonymous:1898:MYA

[9284] Anonymous. Mount Yillimani ascended. Scientific American, 79 (13):195, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
195b.pdf.

Anonymous:1898:SD

[9285] Anonymous. Soap as a disinfectant. Scientific American, 79 (13):195, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
195c.pdf.

Irwin:1898:LEO

[9286] William Gilbert Irwin. The Lake Erie and Ohio River Ship Canal. Scientific American, 79(13):195, September 24, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n13/pdf/scientificamerican09241898-195.pdf.

Anonymous:1898:DAL

[9287] Anonymous. Draught attachment for locomotives. Scientific American, 79(13):196, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
196b.pdf.

Anonymous:1898:LFD

[9288] Anonymous. The largest floating dock in the world. Scientific American, 79(13):196, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
196a.pdf.

Anonymous:1898:NTM

[9289] Anonymous. A new tide motor. Scientific American, 79(13): 196, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
196.pdf.

Anonymous:1898:CWL

[9290] Anonymous. Collapse of the Wilson line pier shed at New York. Scientific American, 79(13):197–198, September 24, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n13/pdf/scientificamerican09241898-197.pdf.

Anonymous:1898:NTV

[9291] Anonymous. A new testing valve. Scientific American, 79(13): 198, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
198.pdf.

Anonymous:1898:Cv

[9292] Anonymous. Correspondence. Scientific American, 79(13):199, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n13/pdf/scientificamerican09241898-199.pdf.

Beasley:1898:NLU

[9293] Walter L. Beasley. A new light upon the Egyptians. Scientific American, 79(13):200, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
200.pdf.

Anonymous:1898:RFGb

[9294] Anonymous. Rapid-fire guns and ammunition supply in the United States Navy. Scientific American, 79(13):203–204, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n13/pdf/scientificamerican09241898-203.pdf.

Anonymous:1898:RPIbm

[9295] Anonymous. Recently patented inventions. Scientific American, 79 (13):204–205, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
204.pdf.

Anonymous:1898:BNBbd

[9296] Anonymous. Books: New books, etc. Scientific American, 79 (13):205, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
205.pdf.

Anonymous:1898:BPbm

[9297] Anonymous. Business and personal. Scientific American, 79 (13):205, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
205a.pdf.

Anonymous:1898:Ibh

[9298] Anonymous. To inventors. Scientific American, 79(13):205, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n13/pdf/scientificamerican09241898-205c.pdf.

Anonymous:1898:IIbj

[9299] Anonymous. Index of inventions. Scientific American, 79(13): 205, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
205d.pdf.

Anonymous:1898:NQbk

[9300] Anonymous. Notes and queries. Scientific American, 79(13): 205, September 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n13/pdf/scientificamerican09241898-
205b.pdf.

Anonymous:1898:PCc

[9301] Anonymous. Patriotic criticism. Scientific American, 79(14):210, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-210.pdf.

Anonymous:1898:WSW

[9302] Anonymous. The World’s supply of wheat. Scientific American, 79(14):210–211, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
210a.pdf.

Anonymous:1898:EEW

[9303] Anonymous. End of an electrical war. Scientific American, 79(14):211, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-211b.pdf.

Anonymous:1898:KTB

[9304] Anonymous. A 35.2-knot torpedo boat. Scientific American, 79(14):211, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-211c.pdf.

Anonymous:1898:NTA

[9305] Anonymous. New troopships for the Army. Scientific American, 79(14):211, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
211a.pdf.

Anonymous:1898:SCb

[9306] Anonymous. A statue of Champlain. Scientific American, 79(14):211, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-211d.pdf.

Serviss:1898:HO

[9307] Garrett P. Serviss. The heavens in October. Scientific American, 79(14):211, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
211.pdf.

Anonymous:1898:LLT

[9308] Anonymous. A Lilliputian locomotive and train. Scientific American, 79(14):212, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
212a.pdf.

Anonymous:1898:NLT

[9309] Anonymous. A new locomotive throttle-lever. Scientific American, 79(14):212, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
212.pdf.

Anonymous:1898:ITS

[9310] Anonymous. An improved tread-sanding machine. Scientific American, 79(14):213, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
213a.pdf.

Anonymous:1898:NFE

[9311] Anonymous. A new fire-extinguisher. Scientific American, 79(14):213, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-213b.pdf.

Anonymous:1898:NGG

[9312] Anonymous. A new gas-governor. Scientific American, 79(14):213, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-213.pdf.

Crafts:1898:DMI

[9313] H. A. Crafts. Difficulties of mountain irrigation engineering. Scientific American, 79(14):214, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
214.pdf.

Anonymous:1898:Cw

[9314] Anonymous. Correspondence. Scientific American, 79(14):215, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-215.pdf.

Anonymous:1898:PCP

[9315] Anonymous. Protected cruiser “Philadelphia,” of the United States Navy. Scientific American, 79(14):216, October 1, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n14/pdf/scientificamerican10011898-216.pdf.

Anonymous:1898:ESM

[9316] Anonymous. Explosion of submarine mines in Baltimore Harbor. Scientific American, 79(14):217–218, October 1, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n14/pdf/scientificamerican10011898-217.pdf.

Anonymous:1898:FBC

[9317] Anonymous. The fall of a bridge at Cornwall, Ontario. Scientific American, 79(14):219–220, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
219.pdf.

Anonymous:1898:RPIbn

[9318] Anonymous. Recently patented inventions. Scientific American, 79(14):220, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n14/pdf/scientificamerican10011898-
220.pdf.

Anonymous:1898:BNBbe

[9319] Anonymous. Books: New books, etc. Scientific American, 79(14):221, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-221b.pdf.

Anonymous:1898:BPbn

[9320] Anonymous. Business and personal. Scientific American, 79(14):221, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-221.pdf.

Anonymous:1898:IIbk

[9321] Anonymous. Index of inventions. Scientific American, 79(14):221, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-221c.pdf.

Anonymous:1898:NQbl

[9322] Anonymous. Notes and queries. Scientific American, 79(14):221, October 1, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n14/pdf/scientificamerican10011898-221a.pdf.

Anonymous:1898:LTB

[9323] Anonymous. The limitations of the torpedo boat. Scientific American, 79(15):226, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
226a.pdf.

Anonymous:1898:NCD

[9324] Anonymous. New chemical discoveries. Scientific American, 79(15):226, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-226.pdf.

Anonymous:1898:OFT

[9325] Anonymous. Our faulty trade methods. Scientific American, 79 (15):226–227, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
226b.pdf.

Anonymous:1898:AIP

[9326] Anonymous. The American institute photographic, horticultural, and agricultural exhibition. Scientific American, 79(15):227, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-227a.pdf.

Anonymous:1898:HAS

[9327] Anonymous. Hospital arrangements in the Soudan. Scientific American, 79(15):227, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
227.pdf.

Anonymous:1898:SMTb

[9328] Anonymous. Saving the “Maria Teresa”. Scientific American, 79(15):227, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
227c.pdf.

Anonymous:1898:SWK

[9329] Anonymous. “Slate writing and kindred phenomena”. Scientific American, 79(15):227, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
227d.pdf.

Anonymous:1898:USb

[9330] Anonymous. The uniform of the soldier. Scientific American, 79(15):227, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
227b.pdf.

Anonymous:1898:EML

[9331] Anonymous. Explosion of a mountain locomotive at Prescott, Arizona. Scientific American, 79(15):228, October 8, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n15/pdf/scientificamerican10081898-228.pdf.

Anonymous:1898:ISY

[9332] Anonymous. An improved stock-yard gate. Scientific American, 79(15):228, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
228a.pdf.

Anonymous:1898:NMCc

[9333] Anonymous. A novel magazine camera. Scientific American, 79(15):229, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-229.pdf.

Robinson:1898:SSWa

[9334] W. E. Robinson. Spirit slate writing and kindred phenomena. — I. Scientific American, 79(15):229–230, October 8, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n15/pdf/scientificamerican10081898-229a.pdf.

Anonymous:1898:Cx

[9335] Anonymous. Correspondence. Scientific American, 79(15):231, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-231.pdf.

Anonymous:1898:CDK

[9336] Anonymous. Construction and details of the knapp roller boat. Scientific American, 79(15):232, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
232.pdf.

Dickinson:1898:PNZ

[9337] Sidney Dickinson. Picturesque New Zealand. Scientific American, 79(15):233–234, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
233.pdf.

Covey:1898:CHM

[9338] Frank Covey. California harvesting machinery. Scientific American, 79(15):235, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
235.pdf.

Anonymous:1898:MCF

[9339] Anonymous. The “Mercerization” of cotton fabrics. Scientific American, 79(15):236, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
236.pdf.

Anonymous:1898:BNBbf

[9340] Anonymous. Books: New books, etc. Scientific American, 79(15):237, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-237a.pdf.

Anonymous:1898:BPbo

[9341] Anonymous. Business and personal. Scientific American, 79(15):237, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-237b.pdf.

Anonymous:1898:Ibi

[9342] Anonymous. To inventors. Scientific American, 79(15):237, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-237d.pdf.

Anonymous:1898:NQbm

[9343] Anonymous. Notes and queries. Scientific American, 79(15):237, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-237c.pdf.

Anonymous:1898:RPIbo

[9344] Anonymous. Recently patented inventions. Scientific American, 79(15):237, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n15/pdf/scientificamerican10081898-
237.pdf.

Anonymous:1898:IIbl

[9345] Anonymous. Index of inventions. Scientific American, 79(15):238, October 8, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n15/pdf/scientificamerican10081898-238.pdf.

Anonymous:1898:LWBa

[9346] Anonymous. Light-weight bicycles. Scientific American, 79(16):242, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n16/pdf/scientificamerican10151898-242.pdf.

Anonymous:1898:OLB

[9347] Anonymous. Our latest battleship. Scientific American, 79(16):242, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n16/pdf/scientificamerican10151898-242a.pdf.

Anonymous:1898:ISBb

[9348] Anonymous. Importing song birds for our woods. Scientific American, 79(16):242–243, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
242b.pdf.

Anonymous:1898:SWBb

[9349] Anonymous. Spanish wooden bullets. Scientific American, 79 (16):243, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
243a.pdf.

Anonymous:1898:USd

[9350] Anonymous. The uniform of the sailor. Scientific American, 79 (16):243, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
243.pdf.

Anonymous:1898:CNY

[9351] Anonymous. A Cyclone in New York state. Scientific American, 79 (16):244–245, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
244.pdf.

Anonymous:1898:APW

[9352] Anonymous. An apparatus for purifying water. Scientific American, 79(16):245, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
245a.pdf.

Anonymous:1898:NEB

[9353] Anonymous. A new extension bed. Scientific American, 79(16):245, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n16/pdf/scientificamerican10151898-245.pdf.

Anonymous:1898:SBL

[9354] Anonymous. The “Sentinel” bicycle-lock. Scientific American, 79 (16):245–246, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
245b.pdf.

Anonymous:1898:Cy

[9355] Anonymous. Correspondence. Scientific American, 79(16):247, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n16/pdf/scientificamerican10151898-247.pdf.

Anonymous:1898:OEIa

[9356] Anonymous. The Omaha Exposition and the Indian Congress. Scientific American, 79(16):248–249, October 15, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n16/pdf/scientificamerican10151898-248.pdf.

Anonymous:1898:MPc

[9357] Anonymous. The manufacture of paper. Scientific American, 79 (16):249–250, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
249.pdf.

Robinson:1898:SSWb

[9358] W. E. Robinson. Spirit slate writing and kindred phenomena — II. Scientific American, 79(16):251, October 15, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n16/pdf/scientificamerican10151898-251.pdf.

Anonymous:1898:FSD

[9359] Anonymous. Fate of sternwheelers designed for the Yukon. Scientific American, 79(16):251–252, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
251a.pdf.

Anonymous:1898:RPIbp

[9360] Anonymous. Recently patented inventions. Scientific American, 79(16):252, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
252.pdf.

Anonymous:1898:BNBbg

[9361] Anonymous. Books: New books, etc. Scientific American, 79 (16):253, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
253.pdf.

Anonymous:1898:BPbp

[9362] Anonymous. Business and personal. Scientific American, 79 (16):253, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
253a.pdf.

Anonymous:1898:NQbn

[9363] Anonymous. Notes and queries. Scientific American, 79(16):253, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n16/pdf/scientificamerican10151898-253b.pdf.

Anonymous:1898:IIbm

[9364] Anonymous. Index of inventions. Scientific American, 79(16): 253–254, October 15, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n16/pdf/scientificamerican10151898-
253c.pdf.

Anonymous:1898:PPBb

[9365] Anonymous. The pilots and pilot boats of New York. Scientific American, 79(17):257, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
257.pdf.

Anonymous:1898:RNY

[9366] Anonymous. Reconstruction of New York street railways. Scientific American, 79(17):258, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
258.pdf.

Anonymous:1898:NPY

[9367] Anonymous. A new profession for young men. Scientific American, 79(17):258–259, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
258a.pdf.

Anonymous:1898:RAS

[9368] Anonymous. The revival of American shipbuilding. Scientific American, 79(17):259, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
259.pdf.

Weir:1898:CTM

[9369] James Weir. Correlative thought in the monkey and the elephant. Scientific American, 79(17):259, October 22, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n17/pdf/scientificamerican10221898-259a.pdf.

Anonymous:1898:HSFb

[9370] Anonymous. How San Francisco disposes of its garbage. Scientific American, 79(17):260–262, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
260.pdf.

Anonymous:1898:IWc

[9371] Anonymous. An improved wrench. Scientific American, 79(17):262, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n17/pdf/scientificamerican10221898-262.pdf.

Anonymous:1898:LFC

[9372] Anonymous. The launch of the first class battleship “Illinois”. Scientific American, 79(17):264, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
264.pdf.

Anonymous:1898:OFA

[9373] Anonymous. Our fingers as an aid in multiplication. Scientific American, 79(17):265–266, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
265.pdf.

Robinson:1898:SSWc

[9374] W. E. Robinson. Spirit slate writing and kindred phenomena — III. Scientific American, 79(17):266, October 22, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n17/pdf/scientificamerican10221898-266.pdf.

Anonymous:1898:SMS

[9375] Anonymous. San Marino, the smallest independent republic in Europe. Scientific American, 79(17):267, October 22, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n17/pdf/scientificamerican10221898-267.pdf.

Anonymous:1898:NUD

[9376] Anonymous. A novel use for drain tiles. Scientific American, 79 (17):267–268, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
267a.pdf.

Anonymous:1898:RPIbq

[9377] Anonymous. Recently patented inventions. Scientific American, 79(17):268, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
268.pdf.

Anonymous:1898:BNBbh

[9378] Anonymous. Books: New books, etc. Scientific American, 79 (17):269, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
269b.pdf.

Anonymous:1898:BPbq

[9379] Anonymous. Business and personal. Scientific American, 79 (17):269, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
269.pdf.

Anonymous:1898:NQbo

[9380] Anonymous. Notes and queries. Scientific American, 79(17):269, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n17/pdf/scientificamerican10221898-269a.pdf.

Anonymous:1898:IIbn

[9381] Anonymous. Index of inventions. Scientific American, 79(17): 269–270, October 22, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n17/pdf/scientificamerican10221898-
269c.pdf.

Anonymous:1898:OEIb

[9382] Anonymous. Our exports influenced by our patent laws. Scientific American, 79(18):274, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
274.pdf.

Anonymous:1898:TIN

[9383] Anonymous. Is the thirteen-inch Naval gun obsolete? Scientific American, 79(18):274, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
274a.pdf.

Anonymous:1898:WNB

[9384] Anonymous. Work of the Naval Bureau of Ordnance. Scientific American, 79(18):274–275, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
274b.pdf.

Anonymous:1898:ATS

[9385] Anonymous. Army tent stoves. Scientific American, 79(18):275, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n18/pdf/scientificamerican10291898-275c.pdf.

Anonymous:1898:HMT

[9386] Anonymous. The history of the massage treatment. Scientific American, 79(18):275, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
275b.pdf.

Anonymous:1898:TPR

[9387] Anonymous. Taxation of patent rights. Scientific American, 79 (18):275, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
275.pdf.

Serviss:1898:HN

[9388] Garrett P. Serviss. The heavens in November. Scientific American, 79(18):275, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
275a.pdf.

Anonymous:1898:IIFa

[9389] Anonymous. An interesting Italian fountain. Scientific American, 79(18):276, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
276a.pdf.

Robinson:1898:SSWd

[9390] W. E. Robinson. Spirit slate writing and kindred phenomena. — IV. Scientific American, 79(18):276, October 29, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n18/pdf/scientificamerican10291898-276.pdf.

Anonymous:1898:RPP

[9391] Anonymous. Rapid photographic printing, developing, and drying apparatus. Scientific American, 79(18):277, October 29, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n18/pdf/scientificamerican10291898-277.pdf.

Anonymous:1898:Cz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n18/pdf/scientificamerican10291898-278.pdf.

Anonymous:1898:MLE

[9393] Anonymous. Mount Lowe Electric Mountain Railway. Scientific American, 79(18):279, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
279.pdf.

Anonymous:1898:HFDa

[9394] Anonymous. The Havana Floating Dry Dock. Scientific American, 79(18):280, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
280.pdf.

Anonymous:1898:DBS

[9395] Anonymous. The dirigible balloon of De Santos-Dumont. Scientific American, 79(18):281, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
281.pdf.

Anonymous:1898:NBP

[9396] Anonymous. The new buildings of Princeton University. Scientific American, 79(18):282–284, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
282.pdf.

Anonymous:1898:RPIbr

[9397] Anonymous. Recently patented inventions. Scientific American, 79(18):284, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
284.pdf.

Anonymous:1898:BPbr

[9398] Anonymous. Business and personal. Scientific American, 79 (18):285, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
285.pdf.

Anonymous:1898:NQbp

[9399] Anonymous. Notes and queries. Scientific American, 79(18):285, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n18/pdf/scientificamerican10291898-285a.pdf.

Anonymous:1898:IIbo

[9400] Anonymous. Index of inventions. Scientific American, 79(18): 285–286, October 29, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n18/pdf/scientificamerican10291898-
285b.pdf.

Anonymous:1898:ASM

[9401] Anonymous. American silk manufactures and exports. Scientific American, 79(19):290, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
290b.pdf.

Anonymous:1898:ETBb

[9402] Anonymous. Experiment the true basis of engineering data. Scientific American, 79(19):290, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
290a.pdf.

Anonymous:1898:FRT

[9403] Anonymous. The fastest regular train in the world. Scientific American, 79(19):290, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
290.pdf.

Anonymous:1898:LNN

[9404] Anonymous. The latest news of the Nobel Bequest. Scientific American, 79(19):290, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
290c.pdf.

Anonymous:1898:PCCa

[9405] Anonymous. Patents and copyrights in China. Scientific American, 79(19):291, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
291b.pdf.

Anonymous:1898:PVb

[9406] Anonymous. The Plague in Vienna. Scientific American, 79 (19):291, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
291a.pdf.

Anonymous:1898:VPC

[9407] Anonymous. Valuable by-products of cotton. Scientific American, 79(19):291, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
291.pdf.

Anonymous:1898:DCW

[9408] Anonymous. Death of Colonel Waring. Scientific American, 79 (19):291–292, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
291c.pdf.

Anonymous:1898:EFS

[9409] Anonymous. An efficient framing square. Scientific American, 79 (19):292, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
292a.pdf.

Anonymous:1898:ICD

[9410] Anonymous. An improvement in cleaning devices. Scientific American, 79(19):292, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
292.pdf.

Anonymous:1898:SGG

[9411] Anonymous. A safety gage-glass. Scientific American, 79(19): 292, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
292b.pdf.

Anonymous:1898:CIb

[9412] Anonymous. Cyphers incubators. Scientific American, 79(19): 293, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
293.pdf.

Anonymous:1898:KWW

[9413] Anonymous. A key which will lock any door. Scientific American, 79(19):293, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
293a.pdf.

Anonymous:1898:ABW

[9414] Anonymous. American-built warships for the Russian Navy. Scientific American, 79(19):294, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
294.pdf.

Anonymous:1898:Cbb

[9415] Anonymous. Correspondence. Scientific American, 79(19):295, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n19/pdf/scientificamerican11051898-295.pdf.

Anonymous:1898:VMT

[9416] Anonymous. Views from Mount Tamalpais, Cal. Scientific American, 79(19):296, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
296.pdf.

Anonymous:1898:NHPa

[9417] Anonymous. The new high powered guns of the Vickers, Sons & Maxim Company. Scientific American, 79(19):297–298, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n19/pdf/scientificamerican11051898-297.pdf.

Anonymous:1898:MAS

[9418] Anonymous. A medal awarded to the Scientific American. Scientific American, 79(19):298–299, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
298.pdf.

Robinson:1898:SSWe

[9419] W. E. Robinson. Spirit slate writing and kindred phenomena — V. Scientific American, 79(19):299, November 5, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n19/pdf/scientificamerican11051898-299.pdf.

Anonymous:1898:BPbs

[9420] Anonymous. Business and personal. Scientific American, 79 (19):300, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
300a.pdf.

Anonymous:1898:IIbp

[9421] Anonymous. Index of inventions. Scientific American, 79(19): 300, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
300c.pdf.

Anonymous:1898:NQbq

[9422] Anonymous. Notes and queries. Scientific American, 79(19): 300, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
300b.pdf.

Anonymous:1898:RPIbs

[9423] Anonymous. Recently patented inventions. Scientific American, 79(19):300, November 5, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n19/pdf/scientificamerican11051898-
300.pdf.

Anonymous:1898:LWBb

[9424] Anonymous. The light-weight bicycle again. Scientific American, 79(20):306, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
306b.pdf.

Anonymous:1898:PM

[9425] Anonymous. The passing of the monitor. Scientific American, 79 (20):306, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
306a.pdf.

Anonymous:1898:WWS

[9426] Anonymous. The World’s wheat supply. Scientific American, 79 (20):306, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
306.pdf.

Anonymous:1898:SGB

[9427] Anonymous. Shipbuilding in Great Britain. Scientific American, 79 (20):306–307, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
306c.pdf.

Anonymous:1898:CFa

[9428] Anonymous. Camille A. Faure. Scientific American, 79(20):307, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n20/pdf/scientificamerican11121898-307b.pdf.

Anonymous:1898:PCM

[9429] Anonymous. Proposed change of motive power for the fourth avenue tunnel. Scientific American, 79(20):307, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n20/pdf/scientificamerican11121898-307.pdf.

Gerard:1898:LH

[9430] W. R. Gerard. The language of Hawaii. Scientific American, 79 (20):307, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
307a.pdf.

Anonymous:1898:MCP

[9431] Anonymous. The mammoth central power station of the Metropolitan Street Railway Company, New York. Scientific American, 79 (20):308–309, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
308.pdf.

Anonymous:1898:ASW

[9432] Anonymous. An attachment for socket-wrenches. Scientific American, 79(20):309, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
309.pdf.

Anonymous:1898:UNF

[9433] Anonymous. A unicycle of novel form. Scientific American, 79 (20):309–310, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
309a.pdf.

Anonymous:1898:Cbc

[9434] Anonymous. Correspondence. Scientific American, 79(20):311, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n20/pdf/scientificamerican11121898-311.pdf.

Anonymous:1898:PNMa

[9435] Anonymous. The proposed new monitors. Scientific American, 79 (20):312–313, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
312.pdf.

Mindeleff:1898:NAA

[9436] Cosmos Mindeleff. Native architecture in Africa and New Mexico. Scientific American, 79(20):313–314, November 12, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n20/pdf/scientificamerican11121898-313.pdf.

Anonymous:1898:WPB

[9437] Anonymous. A wave-propelled boat. Scientific American, 79 (20):315, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
315.pdf.

Anonymous:1898:BNBbi

[9438] Anonymous. Books: New books, etc. Scientific American, 79 (20):316, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
316a.pdf.

Anonymous:1898:BPbt

[9439] Anonymous. Business and personal. Scientific American, 79 (20):316, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
316b.pdf.

Anonymous:1898:IIbq

[9440] Anonymous. Index of inventions. Scientific American, 79(20): 316, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
316d.pdf.

Anonymous:1898:NQbr

[9441] Anonymous. Notes and queries. Scientific American, 79(20): 316, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
316c.pdf.

Anonymous:1898:RPIbt

[9442] Anonymous. Recently patented inventions. Scientific American, 79(20):316, November 12, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n20/pdf/scientificamerican11121898-
316.pdf.

Anonymous:1898:OP

[9443] Anonymous. The Observatory of Paris. Scientific American, 79 (21):321, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
321.pdf.

Anonymous:1898:FF

[9444] Anonymous. Fact and fancy. Scientific American, 79(21):322, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n21/pdf/scientificamerican11191898-322.pdf.

Anonymous:1898:GSA

[9445] Anonymous. Giant steamships for the Atlantic service. Scientific American, 79(21):322, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
322a.pdf.

Anonymous:1898:BCA

[9446] Anonymous. Boiler capacity in American and English locomotives. Scientific American, 79(21):322–323, November 19, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n21/pdf/scientificamerican11191898-322b.pdf.

Anonymous:1898:AMSb

[9447] Anonymous. Annual meeting of the Society of Naval Architects and Marine Engineers. Scientific American, 79(21):323, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n21/pdf/scientificamerican11191898-323.pdf.

Anonymous:1898:CRP

[9448] Anonymous. Commission to revise the Patent and Trademark Laws of the United States. Scientific American, 79(21):323, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n21/pdf/scientificamerican11191898-323a.pdf.

Anonymous:1898:DDW

[9449] Anonymous. Death of D. A. Wells. Scientific American, 79 (21):323, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
323b.pdf.

Anonymous:1898:CGH

[9450] Anonymous. A convenient globe-holder. Scientific American, 79 (21):324, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
324b.pdf.

Anonymous:1898:FCE

[9451] Anonymous. Four-coupled express engine for the Great Northern Railway, England. Scientific American, 79(21):324, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n21/pdf/scientificamerican11191898-324a.pdf.

Anonymous:1898:WAS

[9452] Anonymous. A wheeled attachment for sleds. Scientific American, 79(21):324, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
324.pdf.

Anonymous:1898:HML

[9453] Anonymous. Handsome model locomotive. Scientific American, 79 (21):325, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
325.pdf.

Anonymous:1898:SLC

[9454] Anonymous. A steam locomotive crane at the Washington Navy Yard. Scientific American, 79(21):325, November 19, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n21/pdf/scientificamerican11191898-325a.pdf.

Anonymous:1898:NTL

[9455] Anonymous. Nikola Tesla’s latest invention. Scientific American, 79(21):326, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
326.pdf.

Anonymous:1898:SFW

[9456] Anonymous. San Francisco Water Supply. Scientific American, 79 (21):328–330, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
328.pdf.

Anonymous:1898:SOC

[9457] Anonymous. Some of our common snakes. Scientific American, 79 (21):331–332, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
331.pdf.

Anonymous:1898:RPIbu

[9458] Anonymous. Recently patented inventions. Scientific American, 79(21):332, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
332.pdf.

Anonymous:1898:BNBbj

[9459] Anonymous. Books: New books, etc. Scientific American, 79 (21):333, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
333b.pdf.

Anonymous:1898:BPbu

[9460] Anonymous. Business and personal. Scientific American, 79 (21):333, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
333.pdf.

Anonymous:1898:Ibj

[9461] Anonymous. To inventors. Scientific American, 79(21):333, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n21/pdf/scientificamerican11191898-333c.pdf.

Anonymous:1898:NQbs

[9462] Anonymous. Notes and queries. Scientific American, 79(21): 333, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
333a.pdf.

Anonymous:1898:IIbr

[9463] Anonymous. Index of inventions. Scientific American, 79(21): 333–334, November 19, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n21/pdf/scientificamerican11191898-
333d.pdf.

Anonymous:1898:RGE

[9464] Anonymous. The removal of a great engineering landmark. Scientific American, 79(22):338, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
338a.pdf.

Anonymous:1898:SS

[9465] Anonymous. Science and sensationalism. Scientific American, 79 (22):338, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
338.pdf.

Anonymous:1898:FCU

[9466] Anonymous. Flax culture in the United States. Scientific American, 79(22):338–339, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
338b.pdf.

Anonymous:1898:LC

[9467] Anonymous. Latimer clark. Scientific American, 79(22):339, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n22/pdf/scientificamerican11261898-339b.pdf.

Anonymous:1898:LMT

[9468] Anonymous. The loss of the “Maria Teresa”. Scientific American, 79(22):339, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
339.pdf.

Anonymous:1898:NM

[9469] Anonymous. The November meteors. Scientific American, 79 (22):339, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
339a.pdf.

Anonymous:1898:UBa

[9470] Anonymous. A useful beetle. Scientific American, 79(22):339, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n22/pdf/scientificamerican11261898-339c.pdf.

Anonymous:1898:ASG

[9471] Anonymous. An automatic safety-gate for bridges. Scientific American, 79(22):340, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
340.pdf.

Anonymous:1898:NWP

[9472] Anonymous. A novel way of protecting buoy-cables. Scientific American, 79(22):340, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
340b.pdf.

Anonymous:1898:UBb

[9473] Anonymous. The University of Bombay. Scientific American, 79 (22):340, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
340a.pdf.

Anonymous:1898:VJB

[9474] Anonymous. The Victoria Jubilee Bridge. Scientific American, 79 (22):341–342, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
341.pdf.

Anonymous:1898:Cbd

[9475] Anonymous. Correspondence. Scientific American, 79(22):343, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n22/pdf/scientificamerican11261898-343.pdf.

Anonymous:1898:NWe

[9476] Anonymous. The Navies of the world. Scientific American, 79 (22):344–346, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
344.pdf.

Anonymous:1898:NRK

[9477] Anonymous. A new registering kite. Scientific American, 79 (22):347, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
347.pdf.

Anonymous:1898:BNBbk

[9478] Anonymous. Books: New books, etc. Scientific American, 79 (22):348, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
348a.pdf.

Anonymous:1898:RPIbv

[9479] Anonymous. Recently patented inventions. Scientific American, 79(22):348, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
348.pdf.

Anonymous:1898:BPbv

[9480] Anonymous. Business and personal. Scientific American, 79 (22):349, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
349.pdf.

Anonymous:1898:NQbt

[9481] Anonymous. Notes and queries. Scientific American, 79(22): 349, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
349a.pdf.

Anonymous:1898:IIbs

[9482] Anonymous. Index of inventions. Scientific American, 79(22): 349–350, November 26, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n22/pdf/scientificamerican11261898-
349b.pdf.

Anonymous:1898:BW

[9483] Anonymous. “The biggest in the world”. Scientific American, 79(23):354, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
354.pdf.

Anonymous:1898:CPR

[9484] Anonymous. Conditions in porto rico. Scientific American, 79 (23):354, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
354c.pdf.

Anonymous:1898:DJK

[9485] Anonymous. Death of John Keely. Scientific American, 79(23):354, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n23/pdf/scientificamerican12031898-354b.pdf.

Anonymous:1898:PNMb

[9486] Anonymous. The proposed new monitors. Scientific American, 79(23):354, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
354a.pdf.

Anonymous:1898:IPCb

[9487] Anonymous. Industrial property commission. Scientific American, 79(23):355, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
355.pdf.

Anonymous:1898:PT

[9488] Anonymous. Philippine toponymy. Scientific American, 79(23):355, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n23/pdf/scientificamerican12031898-355c.pdf.

Anonymous:1898:RDSc

[9489] Anonymous. Recent developments in school sanitation. Scientific American, 79(23):355, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
355b.pdf.

Anonymous:1898:SJF

[9490] Anonymous. Sir John Fowler. Scientific American, 79(23):355, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n23/pdf/scientificamerican12031898-355d.pdf.

Serviss:1898:HD

[9491] Garrett P. Serviss. The heavens in December. Scientific American, 79(23):355, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
355a.pdf.

Anonymous:1898:AML

[9492] Anonymous. Another mammoth locomotive. Scientific American, 79 (23):356–357, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
356.pdf.

Anonymous:1898:NAG

[9493] Anonymous. The National acetylene gas generator. Scientific American, 79(23):357, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
357a.pdf.

Anonymous:1898:NCB

[9494] Anonymous. A new covering for boilers and pipes. Scientific American, 79(23):357, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
357.pdf.

Anonymous:1898:SAS

[9495] Anonymous. A safety-attachment for scaffolds. Scientific American, 79(23):357, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
357b.pdf.

Anonymous:1898:RN

[9496] Anonymous. The rotary neostyle. Scientific American, 79(23):358, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n23/pdf/scientificamerican12031898-358.pdf.

Anonymous:1898:Cbe

[9497] Anonymous. Correspondence. Scientific American, 79(23):359, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n23/pdf/scientificamerican12031898-359.pdf.

Anonymous:1898:FCBa

[9498] Anonymous. The first-class battleship “Maine”. Scientific American, 79(23):361, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
361.pdf.

Carden:1898:STI

[9499] Lieut. G. L. Carden and Ordnance Officer. Successful test of 61/4inch Krupp process plate by 8-inch gun. Scientific American, 79 (23):361–362, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
361a.pdf.

Benjamin:1898:STS

[9500] Marcus Benjamin. The study of typical soils. Scientific American, 79(23):362–364, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
362.pdf.

Anonymous:1898:RPIbw

[9501] Anonymous. Recently patented inventions. Scientific American, 79(23):364, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
364.pdf.

Anonymous:1898:BPbw

[9502] Anonymous. Business and personal. Scientific American, 79 (23):365, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
365.pdf.

Anonymous:1898:NQbu

[9503] Anonymous. Notes and queries. Scientific American, 79(23):365, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n23/pdf/scientificamerican12031898-365a.pdf.

Anonymous:1898:IIbt

[9504] Anonymous. Index of inventions. Scientific American, 79(23): 365–366, December 3, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n23/pdf/scientificamerican12031898-
365b.pdf.

Anonymous:1898:IOEb

[9505] Anonymous. Increase of our export trade to Africa. Scientific American, 79(24):370, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
370b.pdf.

Anonymous:1898:LPb

[9506] Anonymous. The loss of the “Portland”. Scientific American, 79 (24):370, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
370a.pdf.

Anonymous:1898:MRT

[9507] Anonymous. Metal railroad ties cheaper than wood ties. Scientific American, 79(24):370, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
370d.pdf.

Anonymous:1898:PAN

[9508] Anonymous. The proposed additions to the Navy. Scientific American, 79(24):370, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
370c.pdf.

Anonymous:1898:UTR

[9509] Anonymous. The underground trolley in the recent snowstorm. Scientific American, 79(24):370, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
370.pdf.

Anonymous:1898:SLR

[9510] Anonymous. Secretary Long’s report. Scientific American, 79(24): 370–371, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
370e.pdf.

Anonymous:1898:HEC

[9511] Anonymous. Hints for exporters to China. Scientific American, 79(24):371, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
371b.pdf.

Anonymous:1898:LJR

[9512] Anonymous. The late Joshua Rose. Scientific American, 79 (24):371, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
371c.pdf.

Anonymous:1898:RSA

[9513] Anonymous. The report of the Secretary of Agriculture. Scientific American, 79(24):371, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
371.pdf.

Anonymous:1898:SH

[9514] Anonymous. A school of housekeeping. Scientific American, 79 (24):371, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
371a.pdf.

Anonymous:1898:NCC

[9515] Anonymous. The New Century Caligraph. Scientific American, 79(24):372, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
372.pdf.

Anonymous:1898:NHPb

[9516] Anonymous. The new 2,500 horse power turbines at Niagara. Scientific American, 79(24):373, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
373.pdf.

Anonymous:1898:NV

[9517] Anonymous. A new viscosimeter. Scientific American, 79(24): 373–374, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
373a.pdf.

Anonymous:1898:Cbf

[9518] Anonymous. Correspondence. Scientific American, 79(24):375, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n24/pdf/scientificamerican12101898-375.pdf.

Anonymous:1898:NWf

[9519] Anonymous. Navies of the world. Scientific American, 79(24): 376–378, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
376.pdf.

Holder:1898:SC

[9520] C. F. Holder. The siphon of the clam. Scientific American, 79 (24):378–379, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
378.pdf.

Anonymous:1898:MW

[9521] Anonymous. Mental wonders. Scientific American, 79(24):379, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n24/pdf/scientificamerican12101898-379.pdf.

Anonymous:1898:BPbx

[9522] Anonymous. Business and personal. Scientific American, 79 (24):380, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
380a.pdf.

Anonymous:1898:IIbu

[9523] Anonymous. Index of inventions. Scientific American, 79(24): 380, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
380c.pdf.

Anonymous:1898:NQbv

[9524] Anonymous. Notes and queries. Scientific American, 79(24): 380, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
380b.pdf.

Anonymous:1898:RPIbx

[9525] Anonymous. Recently patented inventions. Scientific American, 79(24):380, December 10, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n24/pdf/scientificamerican12101898-
380.pdf.

Anonymous:1898:UPE

[9526] Anonymous. Unsuspected peril in elevators. Scientific American, 79(25):386, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
386.pdf.

Anonymous:1898:FIU

[9527] Anonymous. Forthcoming improvements in United States Naval Ordnance. Scientific American, 79(25):386–387, December 17, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n25/pdf/scientificamerican12171898-386a.pdf.

Anonymous:1898:CCPb

[9528] Anonymous. Chief Constructor Philip Hichborn. Scientific American, 79(25):387, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
387.pdf.

Anonymous:1898:GBP

[9529] Anonymous. Glazed book paper bad for the eyes. Scientific American, 79(25):387, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
387a.pdf.

Anonymous:1898:RPH

[9530] Anonymous. Radiography and the physiology of the heart. Scientific American, 79(25):387, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
387b.pdf.

Anonymous:1898:HSA

[9531] Anonymous. High speed attained by the “Farragut”. Scientific American, 79(25):388, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
388a.pdf.

Anonymous:1898:SEB

[9532] Anonymous. A simple electric bicycle-lamp. Scientific American, 79(25):388, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
388.pdf.

Anonymous:1898:IREd

[9533] Anonymous. An improvement in rotary engines. Scientific American, 79(25):389, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
389.pdf.

Anonymous:1898:NWJ

[9534] Anonymous. A new wagon-jack. Scientific American, 79(25): 389, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
389a.pdf.

Anonymous:1898:SFT

[9535] Anonymous. Severe fire test of the modern “sky-scraper”. Scientific American, 79(25):389–390, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
389b.pdf.

Anonymous:1898:Cbg

[9536] Anonymous. Correspondence. Scientific American, 79(25):391, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n25/pdf/scientificamerican12171898-391.pdf.

Anonymous:1898:GCD

[9537] Anonymous. A great California dam. Scientific American, 79(25): 392–393, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
392.pdf.

Anonymous:1898:GRS

[9538] Anonymous. Guns recovered from the Spanish cruisers. Scientific American, 79(25):393–394, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
393.pdf.

Bouldin:1898:PCV

[9539] Powhatan Bouldin. Patrick County, Va., and its curious “Fairy Stones”. Scientific American, 79(25):394–395, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n25/pdf/scientificamerican12171898-394.pdf.

Anonymous:1898:CCM

[9540] Anonymous. A curious case of malformation. Scientific American, 79(25):395, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
395.pdf.

Anonymous:1898:BNBbl

[9541] Anonymous. Books: New books, etc. Scientific American, 79 (25):396, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
396a.pdf.

Anonymous:1898:BPby

[9542] Anonymous. Business and personal. Scientific American, 79 (25):396, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
396b.pdf.

Anonymous:1898:RPIby

[9543] Anonymous. Recently patented inventions. Scientific American, 79(25):396, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
396.pdf.

Anonymous:1898:NQbw

[9544] Anonymous. Notes and queries. Scientific American, 79(25): 396–397, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
396c.pdf.

Anonymous:1898:Ibk

[9545] Anonymous. To inventors. Scientific American, 79(25):397, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n25/pdf/scientificamerican12171898-397.pdf.

Anonymous:1898:IIbv

[9546] Anonymous. Index of inventions. Scientific American, 79(25): 397–398, December 17, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n25/pdf/scientificamerican12171898-
397a.pdf.

Anonymous:1898:BDN

[9547] Anonymous. Breakdown in daily newspaper offices due to lack of gas. Scientific American, 79(26):402, December 24, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n26/pdf/scientificamerican12241898-402c.pdf.

Anonymous:1898:EAM

[9548] Anonymous. English and American machine tools. Scientific American, 79(26):402, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
402.pdf.

Anonymous:1898:ELE

[9549] Anonymous. Electric locomotives for Europe. Scientific American, 79(26):402, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
402b.pdf.

Anonymous:1898:HFDb

[9550] Anonymous. Havana’s floating dry dock. Scientific American, 79 (26):402, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
402d.pdf.

Anonymous:1898:SBB

[9551] Anonymous. Safety of the Brooklyn Bridge. Scientific American, 79(26):402, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
402a.pdf.

Anonymous:1898:SCN

[9552] Anonymous. Scientific conventions in New York during the holidays. Scientific American, 79(26):402–403, December 24, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n26/pdf/scientificamerican12241898-402e.pdf.

Anonymous:1898:CCBb

[9553] Anonymous. Cost of cleaning brick pavements. Scientific American, 79(26):403, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
403a.pdf.

Maxim:1898:FMO

[9554] Hiram Maxim. Flying machines and ordnance. Scientific American, 79(26):403, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
403.pdf.

Anonymous:1898:IWD

[9555] Anonymous. An ingenious water-distilling apparatus. Scientific American, 79(26):404, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
404.pdf.

Anonymous:1898:RSTb

[9556] Anonymous. The “Rochester” system of time-recording. Scientific American, 79(26):404, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
404a.pdf.

Anonymous:1898:EBC

[9557] Anonymous. Equipment of the Broadway Cable Line with electricity. Scientific American, 79(26):405–406, December 24, 1898. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v79/
n26/pdf/scientificamerican12241898-405.pdf.

Anonymous:1898:Cbh

[9558] Anonymous. Correspondence. Scientific American, 79(26):407, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n26/pdf/scientificamerican12241898-407.pdf.

Anonymous:1898:CLG

[9559] Anonymous. Collapse of a large gas holder. Scientific American, 79(26):408, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
408.pdf.

Anonymous:1898:TMO

[9560] Anonymous. The tallest of modern office buildings. Scientific American, 79(26):409–411, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
409.pdf.

Anonymous:1898:BWM

[9561] Anonymous. The bursting of a water main in Brooklyn. Scientific American, 79(26):411–412, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
411.pdf.

Anonymous:1898:RPIbz

[9562] Anonymous. Recently patented inventions. Scientific American, 79(26):412, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
412.pdf.

Anonymous:1898:BPbz

[9563] Anonymous. Business and personal. Scientific American, 79 (26):413, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
413.pdf.

Anonymous:1898:Ibl

[9564] Anonymous. To inventors. Scientific American, 79(26):413, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v79/n26/pdf/scientificamerican12241898-413b.pdf.

Anonymous:1898:NQbx

[9565] Anonymous. Notes and queries. Scientific American, 79(26): 413, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
413a.pdf.

Anonymous:1898:IIbw

[9566] Anonymous. Index of inventions. Scientific American, 79(26): 413–414, December 24, 1898. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v79/n26/pdf/scientificamerican12241898-
413c.pdf.

Anonymous:1899:RY

[9567] Anonymous. A retrospect of the year 1898. Scientific American, 80(1):4–5, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n1/pdf/scientificamerican01071899-
4.pdf.

Anonymous:1899:AFA

[9568] Anonymous. Automobiles for fifth avenue. Scientific American, 80(1):5, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-5a.pdf.

Farrington:1899:RUP

[9569] Oliver C. Farrington. Remarkable uses of peat. Scientific American, 80(1):5, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n1/pdf/scientificamerican01071899-
5.pdf.

Anonymous:1899:AAA

[9570] Anonymous. An automatic acetylene-apparatus. Scientific American, 80(1):6, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n1/pdf/scientificamerican01071899-
6b.pdf.

Anonymous:1899:CV

[9571] Anonymous. A convertible vehicle. Scientific American, 80(1):6, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-6a.pdf.

Anonymous:1899:ICC

[9572] Anonymous. An improvement in clay-cutting tables. Scientific American, 80(1):6, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n1/pdf/scientificamerican01071899-
6.pdf.

Anonymous:1899:GWP

[9573] Anonymous. The giant wheel of Paris. Scientific American, 80(1):7–8, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-7.pdf.

Anonymous:1899:Ca

[9574] Anonymous. Correspondence. Scientific American, 80(1):9, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-9.pdf.

Anonymous:1899:NER

[9575] Anonymous. The new East River Bridge. Scientific American, 80(1):10–11, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n1/pdf/scientificamerican01071899-
10.pdf.

Anonymous:1899:AFW

[9576] Anonymous. The armed forces of the world. Scientific American, 80(1):11–12, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n1/pdf/scientificamerican01071899-
11.pdf.

Anonymous:1899:RPIa

[9577] Anonymous. Recently patented inventions. Scientific American, 80(1):12–13, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n1/pdf/scientificamerican01071899-
12.pdf.

Anonymous:1899:BPa

[9578] Anonymous. Business and personal. Scientific American, 80(1):13, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-13.pdf.

Anonymous:1899:IIa

[9579] Anonymous. Index of inventions. Scientific American, 80(1):13, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-13b.pdf.

Anonymous:1899:NQa

[9580] Anonymous. Notes and queries. Scientific American, 80(1):13, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-13a.pdf.

Anonymous:1899:PNPa

[9581] Anonymous. Patent notes: Patents. Scientific American, 80(1):14, January 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n1/pdf/scientificamerican01071899-14.pdf.

Anonymous:1899:ECa

[9582] Anonymous. Education by correspondence. Scientific American, 80(2):18, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
18b.pdf.

Anonymous:1899:FGC

[9583] Anonymous. Failure of the Gatling cast-steel gun. Scientific American, 80(2):18, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
18a.pdf.

Anonymous:1899:ISPa

[9584] Anonymous. The ideal smokeless powder. Scientific American, 80(2):18, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
18.pdf.

Anonymous:1899:NSBa

[9585] Anonymous. New Suspension Bridge at Niagara. Scientific American, 80(2):18–19, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
18c.pdf.

Irwin:1899:GOO

[9586] William Gilbert Irwin. The growth of our oil trade. Scientific American, 80(2):19, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
19.pdf.

Anonymous:1899:IRJ

[9587] Anonymous. An improved rail-joint. Scientific American, 80(2):20, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-20b.pdf.

Anonymous:1899:NCBa

[9588] Anonymous. A novel chainless bicycle. Scientific American, 80 (2):20, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
20a.pdf.

Anonymous:1899:NLV

[9589] Anonymous. A new link valve-gear. Scientific American, 80(2):20, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-20.pdf.

Anonymous:1899:NBTa

[9590] Anonymous. The new Boston Terminal Station. Scientific American, 80(2):21–22, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
21.pdf.

Anonymous:1899:Cb

[9591] Anonymous. Correspondence. Scientific American, 80(2):23, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-23.pdf.

Anonymous:1899:BV

[9592] Anonymous. The biograph in the Vatican. Scientific American, 80(2):24–25, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
24.pdf.

Anonymous:1899:GRS

[9593] Anonymous. Gunboats and railways in the Soudan Campaign. Scientific American, 80(2):25–26, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
25.pdf.

Anonymous:1899:DSS

[9594] Anonymous. Damage by Spanish shells on the cruiser “Brooklyn”. Scientific American, 80(2):27, January 14, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n2/pdf/scientificamerican01141899-27.pdf.

Anonymous:1899:BNBa

[9595] Anonymous. Books: New books, etc. Scientific American, 80 (2):28, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
28a.pdf.

Anonymous:1899:RPIb

[9596] Anonymous. Recently patented inventions. Scientific American, 80(2):28, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n2/pdf/scientificamerican01141899-
28.pdf.

Anonymous:1899:BPb

[9597] Anonymous. Business and personal. Scientific American, 80(2):29, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-29.pdf.

Anonymous:1899:Ia

[9598] Anonymous. To inventors. Scientific American, 80(2):29, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-29b.pdf.

Anonymous:1899:NQb

[9599] Anonymous. Notes and queries. Scientific American, 80(2):29, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-29a.pdf.

Anonymous:1899:IIb

[9600] Anonymous. Index of inventions. Scientific American, 80(2):29–30, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-29c.pdf.

Anonymous:1899:PNPb

[9601] Anonymous. Patent notes: Patents. Scientific American, 80(2):30, January 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n2/pdf/scientificamerican01141899-30.pdf.

Anonymous:1899:DRS

[9602] Anonymous. The diamond reef shoal in New York Harbor. Scientific American, 80(3):34, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
34a.pdf.

Anonymous:1899:GNA

[9603] Anonymous. A greatly needed amendment to the patent laws. Scientific American, 80(3):34, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
34b.pdf.

Anonymous:1899:SGE

[9604] Anonymous. Shall the Government engage in commercial enterprises? Scientific American, 80(3):34, January 21, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n3/pdf/scientificamerican01211899-34.pdf.

Anonymous:1899:DN

[9605] Anonymous. The diseases of nations. Scientific American, 80 (3):34–35, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
34c.pdf.

Anonymous:1899:DRO

[9606] Anonymous. Diligence required from the owners of trade-marks. Scientific American, 80(3):35, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
35.pdf.

Anonymous:1899:EW

[9607] Anonymous. Edible weeds. Scientific American, 80(3):35, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n3/pdf/scientificamerican01211899-35b.pdf.

Anonymous:1899:RLR

[9608] Anonymous. Repairing the lightning rod of St. Peter’s at Rome. Scientific American, 80(3):35, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
35a.pdf.

Sheldon:1899:BSI

[9609] S. L. Sheldon. The beet sugar industry in New York state. Scientific American, 80(3):36, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
36.pdf.

Anonymous:1899:SGW

[9610] Anonymous. Siege guns at the Washington arsenal. Scientific American, 80(3):37, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
37.pdf.

Thurston:1899:RFS

[9611] Robert H. Thurston. Robert Fulton and the steamboat. Scientific American, 80(3):37–38, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
37a.pdf.

Anonymous:1899:Cc

[9612] Anonymous. Correspondence. Scientific American, 80(3):39, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n3/pdf/scientificamerican01211899-39.pdf.

Anonymous:1899:RPC

[9613] Anonymous. The royal palms of Cuba. Scientific American, 80 (3):40–41, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
40.pdf.

Anonymous:1899:ASE

[9614] Anonymous. Air ship experiments. Scientific American, 80(3): 41–43, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
41.pdf.

Anonymous:1899:SCA

[9615] Anonymous. Some curious automata. Scientific American, 80 (3):43, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
43a.pdf.

White:1899:SAS

[9616] John Hellyer White. Simplified apparatus for spectroscopic photography. Scientific American, 80(3):43, January 21, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n3/pdf/scientificamerican01211899-43.pdf.

Anonymous:1899:RPIc

[9617] Anonymous. Recently patented inventions. Scientific American, 80(3):44, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
44.pdf.

Anonymous:1899:BNBb

[9618] Anonymous. Books: New books, etc. Scientific American, 80 (3):44–45, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n3/pdf/scientificamerican01211899-
44a.pdf.

Anonymous:1899:BPc

[9619] Anonymous. Business and personal. Scientific American, 80(3):45, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n3/pdf/scientificamerican01211899-45.pdf.

Anonymous:1899:Ib

[9620] Anonymous. To inventors. Scientific American, 80(3):45, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n3/pdf/scientificamerican01211899-45b.pdf.

Anonymous:1899:NQc

[9621] Anonymous. Notes and queries. Scientific American, 80(3):45, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n3/pdf/scientificamerican01211899-45a.pdf.

Anonymous:1899:IIc

[9622] Anonymous. Index of inventions. Scientific American, 80(3):45–46, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n3/pdf/scientificamerican01211899-45c.pdf.

Anonymous:1899:PNPc

[9623] Anonymous. Patent notes: Patents. Scientific American, 80(3):47, January 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n3/pdf/scientificamerican01211899-47.pdf.

Anonymous:1899:ACT

[9624] Anonymous. The art critic and the tall building. Scientific American, 80(4):50, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
50.pdf.

Anonymous:1899:BPO

[9625] Anonymous. The business of the Patent Office in 1898. Scientific American, 80(4):50, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
50b.pdf.

Anonymous:1899:CATa

[9626] Anonymous. Compressed air traction in New York City. Scientific American, 80(4):50, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
50a.pdf.

Anonymous:1899:BER

[9627] Anonymous. Boston Elevated Road. Scientific American, 80(4):51, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-51f.pdf.

Anonymous:1899:CESa

[9628] Anonymous. Civil engineers in session. Scientific American, 80 (4):51, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
51a.pdf.

Anonymous:1899:DLC

[9629] Anonymous. Death of the Librarian of Congress. Scientific American, 80(4):51, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
51e.pdf.

Anonymous:1899:GSB

[9630] Anonymous. The great strength of bears. Scientific American, 80(4):51, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
51c.pdf.

Anonymous:1899:PBP

[9631] Anonymous. The pirates and brigands of the Philippines. Scientific American, 80(4):51, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
51b.pdf.

Anonymous:1899:PXR

[9632] Anonymous. Printing by X rays. Scientific American, 80(4):51, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-51.pdf.

Anonymous:1899:PZG

[9633] Anonymous. The progress of the Zoological Garden. Scientific American, 80(4):51, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
51d.pdf.

Anonymous:1899:IIS

[9634] Anonymous. An improvement in ice-skates. Scientific American, 80(4):52, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
52.pdf.

Farrington:1899:GAM

[9635] Oliver C. Farrington. A gigantic artificial Moon. Scientific American, 80(4):52, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
52a.pdf.

Anonymous:1899:CTa

[9636] Anonymous. A combination-tool. Scientific American, 80(4): 52–53, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
52b.pdf.

Anonymous:1899:IWW

[9637] Anonymous. An improved wood-working machine. Scientific American, 80(4):53, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
53.pdf.

Anonymous:1899:ROS

[9638] Anonymous. A rotary ore-stamping mill. Scientific American, 80(4):53, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
53a.pdf.

Anonymous:1899:RRA

[9639] Anonymous. Recognition of Rear-Admiral Dewey. Scientific American, 80(4):54, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
54.pdf.

Anonymous:1899:Cd

[9640] Anonymous. Correspondence. Scientific American, 80(4):55, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-55.pdf.

Anonymous:1899:NWa

[9641] Anonymous. Navies of the world. Scientific American, 80(4):56–59, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-56.pdf.

Anonymous:1899:KMF

[9642] Anonymous. The Keely motor fraud. Scientific American, 80 (4):59–60, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
59.pdf.

Anonymous:1899:RPId

[9643] Anonymous. Recently patented inventions. Scientific American, 80(4):60–61, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
60.pdf.

Anonymous:1899:BNBc

[9644] Anonymous. Books: New books, etc. Scientific American, 80 (4):61, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n4/pdf/scientificamerican01281899-
61.pdf.

Anonymous:1899:BPd

[9645] Anonymous. Business and personal. Scientific American, 80(4):61, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-61a.pdf.

Anonymous:1899:Ic

[9646] Anonymous. To inventors. Scientific American, 80(4):61, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-61c.pdf.

Anonymous:1899:IId

[9647] Anonymous. Index of inventions. Scientific American, 80(4):61, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-61d.pdf.

Anonymous:1899:NQd

[9648] Anonymous. Notes and queries. Scientific American, 80(4):61, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-61b.pdf.

Anonymous:1899:PNPd

[9649] Anonymous. Patent notes: Patents. Scientific American, 80(4):63, January 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n4/pdf/scientificamerican01281899-63.pdf.

Anonymous:1899:CPa

[9650] Anonymous. Copyright of photographs. Scientific American, 80 (5):66, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
66b.pdf.

Anonymous:1899:NPCa

[9651] Anonymous. The new Panama Canal. Scientific American, 80 (5):66, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
66.pdf.

Anonymous:1899:PDE

[9652] Anonymous. Prof. Dewar’s experiment with liquid hydrogen. Scientific American, 80(5):66, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
66a.pdf.

Anonymous:1899:FAC

[9653] Anonymous. The Fourth Annual Cycle and Automobile Exhibition. Scientific American, 80(5):66–67, February 4, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n5/pdf/scientificamerican02041899-66c.pdf.

Anonymous:1899:OIT

[9654] Anonymous. Our import trade for 1898. Scientific American, 80(5):67, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
67d.pdf.

Anonymous:1899:PA

[9655] Anonymous. Philippine architecture. Scientific American, 80 (5):67, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
67b.pdf.

Anonymous:1899:VTC

[9656] Anonymous. The verdict in the tank collapse case. Scientific American, 80(5):67, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
67.pdf.

Anonymous:1899:WSU

[9657] Anonymous. Wine stored underground. Scientific American, 80 (5):67, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
67c.pdf.

Serviss:1899:HF

[9658] Garrett P. Serviss. The heavens in February. Scientific American, 80(5):67, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
67a.pdf.

Anonymous:1899:ISBa

[9659] Anonymous. An improved steam-boiler. Scientific American, 80 (5):68, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
68b.pdf.

Anonymous:1899:NWC

[9660] Anonymous. A new way of constructing drawbridges. Scientific American, 80(5):68, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
68a.pdf.

Anonymous:1899:WFD

[9661] Anonymous. A water-feeding device for grinding-wheels. Scientific American, 80(5):68, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
68.pdf.

Anonymous:1899:HKC

[9662] Anonymous. The Hoadley–Knight compressed air motor. Scientific American, 80(5):69–70, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
69.pdf.

Anonymous:1899:Ce

[9663] Anonymous. Correspondence. Scientific American, 80(5):71, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n5/pdf/scientificamerican02041899-71.pdf.

Anonymous:1899:IKL

[9664] Anonymous. Investigations at the Keely Laboratory. Scientific American, 80(5):72–73, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
72.pdf.

Anonymous:1899:NPCb

[9665] Anonymous. The new Panama Canal. Scientific American, 80 (5):73–75, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
73.pdf.

Holder:1899:GS

[9666] Charles Frederick Holder. Glass sponges. Scientific American, 80(5):75, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
75.pdf.

Howard:1899:IBC

[9667] L. O. Howard. An insect breeding in crude petroleum. Scientific American, 80(5):75–76, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
75a.pdf.

Anonymous:1899:RPIe

[9668] Anonymous. Recently patented inventions. Scientific American, 80(5):76, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
76.pdf.

Anonymous:1899:BPe

[9669] Anonymous. Business and personal. Scientific American, 80 (5):77, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
77.pdf.

Anonymous:1899:IIe

[9670] Anonymous. Index of inventions. Scientific American, 80(5):77, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n5/pdf/scientificamerican02041899-77b.pdf.

Anonymous:1899:NQe

[9671] Anonymous. Notes and queries. Scientific American, 80(5):77, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n5/pdf/scientificamerican02041899-77a.pdf.

Anonymous:1899:PNPe

[9672] Anonymous. Patent notes: Patents. Scientific American, 80 (5):79, February 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n5/pdf/scientificamerican02041899-
79.pdf.

Anonymous:1899:EI

[9673] Anonymous. Explorations in Iceland. Scientific American, 80 (6):82, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
82b.pdf.

Anonymous:1899:OFL

[9674] Anonymous. Our first and last sixteen-inch coast-defense gun. Scientific American, 80(6):82, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
82.pdf.

Anonymous:1899:WSP

[9675] Anonymous. The water supply of the Panama Canal. Scientific American, 80(6):82, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
82a.pdf.

Anonymous:1899:HGM

[9676] Anonymous. Honolulu’s great museum. Scientific American, 80 (6):82–83, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
82c.pdf.

Anonymous:1899:CWa

[9677] Anonymous. Colonies of the world. Scientific American, 80(6): 83, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
83b.pdf.

Anonymous:1899:EIG

[9678] Anonymous. The Eighth International Geological Congress, Paris, 1900. Scientific American, 80(6):83, February 11, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n6/pdf/scientificamerican02111899-83.pdf.

Anonymous:1899:NPPa

[9679] Anonymous. A new power in photography. Scientific American, 80(6):83, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
83a.pdf.

Anonymous:1899:ATa

[9680] Anonymous. An automatic Tellurion. Scientific American, 80(6):84, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://en.wikipedia.org/wiki/Tellurion;
 http://www.nature.com/scientificamerican/journal/v80/n6/pdf/
scientificamerican02111899-84a.pdf.

Anonymous:1899:ICS

[9681] Anonymous. An improvement in cutter-stocks. Scientific American, 80(6):84, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
84b.pdf.

Anonymous:1899:POR

[9682] Anonymous. A pneumatically operated railway-sand-box. Scientific American, 80(6):84, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
84.pdf.

Anonymous:1899:SSa

[9683] Anonymous. Skate-sailing. Scientific American, 80(6):85, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n6/pdf/scientificamerican02111899-85.pdf.

Anonymous:1899:JTB

[9684] Anonymous. Japanese torpedo boat destroyer. Scientific American, 80(6):85–86, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
85a.pdf.

Anonymous:1899:NRB

[9685] Anonymous. A non-refillable bottle. Scientific American, 80(6): 86, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
86.pdf.

Anonymous:1899:Cf

[9686] Anonymous. Correspondence. Scientific American, 80(6):87, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n6/pdf/scientificamerican02111899-87a.pdf.

Anonymous:1899:FOM

[9687] Anonymous. The fortifications of Old Manila. Scientific American, 80(6):87, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
87.pdf.

Anonymous:1899:WSL

[9688] Anonymous. The White Star liner “Oceanic”. Scientific American, 80(6):88–89, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
88.pdf.

Anonymous:1899:NLBa

[9689] Anonymous. Novelties at the late bicycle show. Scientific American, 80(6):89–90, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
89.pdf.

Smith:1899:IFA

[9690] Hugh M. Smith. Interesting fish from the Alabama coast. Scientific American, 80(6):91, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
91.pdf.

Anonymous:1899:PCSa

[9691] Anonymous. The pavement of the Cathedral of Siena. Scientific American, 80(6):91–92, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
91a.pdf.

Anonymous:1899:RPIf

[9692] Anonymous. Recently patented inventions. Scientific American, 80(6):92, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
92.pdf.

Anonymous:1899:BNBd

[9693] Anonymous. Books: New books, etc. Scientific American, 80 (6):93, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
93.pdf.

Anonymous:1899:BPf

[9694] Anonymous. Business and personal. Scientific American, 80(6): 93, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
93a.pdf.

Anonymous:1899:Id

[9695] Anonymous. To inventors. Scientific American, 80(6):93, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n6/pdf/scientificamerican02111899-93c.pdf.

Anonymous:1899:IIf

[9696] Anonymous. Index of inventions. Scientific American, 80(6):93, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n6/pdf/scientificamerican02111899-93d.pdf.

Anonymous:1899:NQf

[9697] Anonymous. Notes and queries. Scientific American, 80(6):93, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n6/pdf/scientificamerican02111899-93b.pdf.

Anonymous:1899:PNPf

[9698] Anonymous. Patent notes: Patents. Scientific American, 80(6): 95, February 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n6/pdf/scientificamerican02111899-
95.pdf.

Anonymous:1899:PNCa

[9699] Anonymous. The Panama and Nicaragua Canals — a comparison. Scientific American, 80(7):98, February 18, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n7/pdf/scientificamerican02181899-98.pdf.

Anonymous:1899:HCP

[9700] Anonymous. The horseless carriage and public health. Scientific American, 80(7):98–99, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
98a.pdf.

Anonymous:1899:FPO

[9701] Anonymous. Fiber plants from our new possessions. Scientific American, 80(7):99, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
99.pdf.

Anonymous:1899:GAS

[9702] Anonymous. Great activity in the steel trade. Scientific American, 80(7):99, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
99c.pdf.

Anonymous:1899:GSW

[9703] Anonymous. Great steel and wire combination. Scientific American, 80(7):99, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
99d.pdf.

Anonymous:1899:Pb

[9704] Anonymous. Pullman. Scientific American, 80(7):99, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n7/pdf/scientificamerican02181899-99b.pdf.

Anonymous:1899:TCS

[9705] Anonymous. Testing of Children’s strength. Scientific American, 80(7):99, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
99a.pdf.

Anonymous:1899:IBM

[9706] Anonymous. An ingenious brake and motor-controller for electric cars. Scientific American, 80(7):100, February 18, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n7/pdf/scientificamerican02181899-100.pdf.

Anonymous:1899:PHS

[9707] Anonymous. A portable horse stall for railway cars. Scientific American, 80(7):100, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
100a.pdf.

Anonymous:1899:NNR

[9708] Anonymous. A new non-refillable bottle. Scientific American, 80 (7):100–101, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
100b.pdf.

Anonymous:1899:ALC

[9709] Anonymous. Automobiles at the late cycle exhibition. Scientific American, 80(7):101–102, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
101.pdf.

Anonymous:1899:Cg

[9710] Anonymous. Correspondence. Scientific American, 80(7):103, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n7/pdf/scientificamerican02181899-103.pdf.

Anonymous:1899:PNCb

[9711] Anonymous. The proposed Nicaragua Canal. Scientific American, 80(7):104–106, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
104.pdf.

Anonymous:1899:RKB

[9712] Anonymous. Repairing the keel of a battleship. Scientific American, 80(7):107–108, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
107.pdf.

Anonymous:1899:RPIg

[9713] Anonymous. Recently patented inventions. Scientific American, 80(7):108, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
108.pdf.

Anonymous:1899:BNBe

[9714] Anonymous. Books: New books, etc. Scientific American, 80 (7):109, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
109.pdf.

Anonymous:1899:BPg

[9715] Anonymous. Business and personal. Scientific American, 80(7): 109, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
109a.pdf.

Anonymous:1899:IIg

[9716] Anonymous. Index of inventions. Scientific American, 80(7):109, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n7/pdf/scientificamerican02181899-109c.pdf.

Anonymous:1899:NQg

[9717] Anonymous. Notes and queries. Scientific American, 80(7):109, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n7/pdf/scientificamerican02181899-109b.pdf.

Anonymous:1899:PNPg

[9718] Anonymous. Patent notes: Patents. Scientific American, 80 (7):111, February 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n7/pdf/scientificamerican02181899-
111.pdf.

Prindle:1899:MTI

[9719] E. J. Prindle. Measuring and testing instruments used in the manufacture of Naval ordnance. Scientific American, 80(8):113, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n8/pdf/scientificamerican02251899-113.pdf.

Anonymous:1899:CTB

[9720] Anonymous. Congress and the Technical Bureaus. Scientific American, 80(8):114, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
114.pdf.

Anonymous:1899:MRP

[9721] Anonymous. The mineral resources of the Philippine Islands. Scientific American, 80(8):114, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
114b.pdf.

Anonymous:1899:NB

[9722] Anonymous. The Naval Bill for 1899. Scientific American, 80 (8):114, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
114a.pdf.

Anonymous:1899:TRSa

[9723] Anonymous. Transportation in the recent snowstorm. Scientific American, 80(8):114, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
114c.pdf.

Anonymous:1899:SWS

[9724] Anonymous. The sinking of the White Star steamer “Germanic”. Scientific American, 80(8):114–115, February 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n8/pdf/scientificamerican02251899-114d.pdf.

Anonymous:1899:WWB

[9725] Anonymous. Work of the Weather Bureau. Scientific American, 80(8):115, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
115a.pdf.

Serviss:1899:HMa

[9726] Garrett P. Serviss. The heavens in March. Scientific American, 80(8):115, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
115.pdf.

Anonymous:1899:AMO

[9727] Anonymous. An automatic mechanism for operating mine-doors. Scientific American, 80(8):116, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
116.pdf.

Anonymous:1899:FEL

[9728] Anonymous. A French electric locomotive. Scientific American, 80(8):116, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
116a.pdf.

Anonymous:1899:SSB

[9729] Anonymous. A simple scaffold-bracket. Scientific American, 80 (8):116, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
116b.pdf.

Dailey:1899:ISM

[9730] Howard B. Dailey. Interesting static motors. Scientific American, 80(8):117, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
117.pdf.

Anonymous:1899:Ch

[9731] Anonymous. Correspondence. Scientific American, 80(8):119, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n8/pdf/scientificamerican02251899-119.pdf.

Anonymous:1899:BMN

[9732] Anonymous. Battleships and monitors now building for the Navy. Scientific American, 80(8):121, February 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n8/pdf/scientificamerican02251899-121.pdf.

Jenkins:1899:TPW

[9733] C. Francis Jenkins. The telltale plummet in the Washington Monument. Scientific American, 80(8):122, February 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n8/pdf/scientificamerican02251899-122.pdf.

Anonymous:1899:RKE

[9734] Anonymous. Recent kite experiments at the Blue Hill Observatory. Scientific American, 80(8):123–124, February 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n8/pdf/scientificamerican02251899-123.pdf.

Anonymous:1899:RPIh

[9735] Anonymous. Recently patented inventions. Scientific American, 80 (8):124–125, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
124.pdf.

Anonymous:1899:BNBf

[9736] Anonymous. Books: New books, etc. Scientific American, 80 (8):125, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
125b.pdf.

Anonymous:1899:BPh

[9737] Anonymous. Business and personal. Scientific American, 80(8): 125, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
125.pdf.

Anonymous:1899:Ie

[9738] Anonymous. To inventors. Scientific American, 80(8):125, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n8/pdf/scientificamerican02251899-125c.pdf.

Anonymous:1899:IIh

[9739] Anonymous. Index of inventions. Scientific American, 80(8):125, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n8/pdf/scientificamerican02251899-125d.pdf.

Anonymous:1899:NQh

[9740] Anonymous. Notes and queries. Scientific American, 80(8):125, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n8/pdf/scientificamerican02251899-125a.pdf.

Anonymous:1899:PNPh

[9741] Anonymous. Patent notes: Patents. Scientific American, 80 (8):127, February 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n8/pdf/scientificamerican02251899-
127.pdf.

Anonymous:1899:ABb

[9742] Anonymous. Air bubbles. Scientific American, 80(9):130, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-130.pdf.

Anonymous:1899:CBS

[9743] Anonymous. Cost of the Boston Subway. Scientific American, 80(9):130, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-130c.pdf.

Anonymous:1899:CMP

[9744] Anonymous. Comparative merits of the Panama and Nicaragua Canals. Scientific American, 80(9):130, March 4, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n9/pdf/scientificamerican03041899-130d.pdf.

Anonymous:1899:ETA

[9745] Anonymous. Electricity as a thawing agent. Scientific American, 80(9):130, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
130b.pdf.

Anonymous:1899:FRF

[9746] Anonymous. Flashless rapid-fire guns. Scientific American, 80(9):130, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-130g.pdf.

Anonymous:1899:FVA

[9747] Anonymous. The fastest vessel afloat. Scientific American, 80(9):130, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-130f.pdf.

Anonymous:1899:LA

[9748] Anonymous. Limitations of the autotruck. Scientific American, 80(9):130, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
130a.pdf.

Anonymous:1899:MVW

[9749] Anonymous. Masonry versus wooden dry docks. Scientific American, 80(9):130, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
130e.pdf.

Anonymous:1899:RTP

[9750] Anonymous. Rails and tie-plates. Scientific American, 80(9):130–131, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-130h.pdf.

Anonymous:1899:CATb

[9751] Anonymous. A curious accident on a trolley car. Scientific American, 80(9):131, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
131c.pdf.

Anonymous:1899:CMT

[9752] Anonymous. The Comstock Mines and their drainage. Scientific American, 80(9):131, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
131a.pdf.

Anonymous:1899:NMS

[9753] Anonymous. New machine shop for the New York Navy Yard. Scientific American, 80(9):131, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
131b.pdf.

Anonymous:1899:SCTa

[9754] Anonymous. Suggestions concerning trade with Russia. Scientific American, 80(9):131, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
131.pdf.

Anonymous:1899:SSc

[9755] Anonymous. A strange society. Scientific American, 80(9):131, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-131d.pdf.

Anonymous:1899:BW

[9756] Anonymous. The building of a watch. Scientific American, 80 (9):132–133, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
132.pdf.

Anonymous:1899:PPM

[9757] Anonymous. A percolator package for making coffee. Scientific American, 80(9):134, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
134.pdf.

Anonymous:1899:SPC

[9758] Anonymous. A simple pipe cutting and threading tool. Scientific American, 80(9):134, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
134a.pdf.

Anonymous:1899:VBD

[9759] Anonymous. A variable bicycle driving-gear. Scientific American, 80(9):134–135, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
134b.pdf.

Anonymous:1899:RSV

[9760] Anonymous. Raising sunken vessels. Scientific American, 80(9):136, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-136.pdf.

Anonymous:1899:BBN

[9761] Anonymous. The banana as the basis of a new industry. Scientific American, 80(9):137, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
137.pdf.

Anonymous:1899:RFN

[9762] Anonymous. A remarkable fire in New York. Scientific American, 80(9):137–138, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
137a.pdf.

Anonymous:1899:NREa

[9763] Anonymous. New railway enterprises in Paris. Scientific American, 80(9):139–140, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
139.pdf.

Anonymous:1899:BPi

[9764] Anonymous. Business and personal. Scientific American, 80(9):140, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-140a.pdf.

Anonymous:1899:NQi

[9765] Anonymous. Notes and queries. Scientific American, 80(9):140, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-140b.pdf.

Anonymous:1899:RPIi

[9766] Anonymous. Recently patented inventions. Scientific American, 80(9):140, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n9/pdf/scientificamerican03041899-
140.pdf.

Anonymous:1899:IIi

[9767] Anonymous. Index of inventions. Scientific American, 80(9):141, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-141.pdf.

Anonymous:1899:PNPi

[9768] Anonymous. Patent notes: Patents. Scientific American, 80(9):143, March 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n9/pdf/scientificamerican03041899-143.pdf.

Anonymous:1899:ACP

[9769] Anonymous. An American Canal at Panama? Scientific American, 80(10):146, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
146b.pdf.

Anonymous:1899:ARCa

[9770] Anonymous. Annual report of the Commissioner of Patents. Scientific American, 80(10):146, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
146.pdf.

Anonymous:1899:GAA

[9771] Anonymous. The Government as an advertising agency. Scientific American, 80(10):146, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
146a.pdf.

Anonymous:1899:SNA

[9772] Anonymous. The Senate and the Navy Appropriations. Scientific American, 80(10):146, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
146c.pdf.

Adams:1899:CEL

[9773] Altan D. Adams. Cost of electric light. Scientific American, 80(10):147, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-147b.pdf.

Anonymous:1899:NVH

[9774] Anonymous. Nutritive values of human food. Scientific American, 80(10):147, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
147.pdf.

Anonymous:1899:OCS

[9775] Anonymous. Our consular service in relation to American manufactures. Scientific American, 80(10):147, March 11, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n10/pdf/scientificamerican03111899-147a.pdf.

Anonymous:1899:WWW

[9776] Anonymous. Wax works at Westminster Abbey. Scientific American, 80(10):148, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
148.pdf.

Anonymous:1899:NDAa

[9777] Anonymous. A new double-acting pump. Scientific American, 80(10):149, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
149.pdf.

Anonymous:1899:NMCa

[9778] Anonymous. A novel method of conveying freight. Scientific American, 80(10):149, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
149b.pdf.

Anonymous:1899:PBH

[9779] Anonymous. A pneumatic baggage handler. Scientific American, 80(10):149, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
149a.pdf.

Anonymous:1899:Ma

[9780] Anonymous. The Mareorama. Scientific American, 80(10):150, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-150.pdf.

Anonymous:1899:NL

[9781] Anonymous. The Nernst light. Scientific American, 80(10):150–151, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-150a.pdf.

Anonymous:1899:NWb

[9782] Anonymous. Navies of the world. Scientific American, 80(10):152–155, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-152.pdf.

Anonymous:1899:CI

[9783] Anonymous. A Convict’s ingenuity. Scientific American, 80(10):155, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-155.pdf.

Anonymous:1899:BNBg

[9784] Anonymous. Books: New books, etc. Scientific American, 80(10):156, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-156a.pdf.

Anonymous:1899:BPj

[9785] Anonymous. Business and personal. Scientific American, 80(10):156, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-156b.pdf.

Anonymous:1899:RPIj

[9786] Anonymous. Recently patented inventions. Scientific American, 80(10):156, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n10/pdf/scientificamerican03111899-
156.pdf.

Anonymous:1899:NQj

[9787] Anonymous. Notes and queries. Scientific American, 80(10):156–157, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-156c.pdf.

Anonymous:1899:IIj

[9788] Anonymous. Index of inventions. Scientific American, 80(10):157, March 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n10/pdf/scientificamerican03111899-157.pdf.

Walker:1899:ALE

[9789] Herbert T. Walker. Ancient locomotive engines. Scientific American, 80(11):161, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
161.pdf.

Anonymous:1899:PCSb

[9790] Anonymous. To prevent crowding on street cars. Scientific American, 80(11):162, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
162a.pdf.

Anonymous:1899:SNCa

[9791] Anonymous. Selection of the new Canal Commission. Scientific American, 80(11):162, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
162.pdf.

Anonymous:1899:FPS

[9792] Anonymous. The farmer and the patent system. Scientific American, 80(11):162–163, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
162b.pdf.

Anonymous:1899:DJK

[9793] Anonymous. Death of John Kruesi. Scientific American, 80(11):163, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-163.pdf.

Anonymous:1899:DLD

[9794] Anonymous. Death of Lieut. Dashiell. Scientific American, 80(11):163, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-163a.pdf.

Anonymous:1899:NBC

[9795] Anonymous. Naval Bureau Chiefs. Scientific American, 80(11):163, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-163c.pdf.

Anonymous:1899:SE

[9796] Anonymous. Sportsmen’s exhibition. Scientific American, 80(11):163, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-163b.pdf.

Anonymous:1899:JMM

[9797] Anonymous. Japanese magic mirrors. Scientific American, 80(11):164, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-164a.pdf.

Anonymous:1899:TIG

[9798] Anonymous. Thirteen-inch gun for the “Kearsarge”. Scientific American, 80(11):164, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
164.pdf.

Anonymous:1899:APP

[9799] Anonymous. An adjustable paper-punch. Scientific American, 80(11):165, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
165a.pdf.

Anonymous:1899:EAM

[9800] Anonymous. An electrical alarm-matting. Scientific American, 80(11):165, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
165.pdf.

Anonymous:1899:IP

[9801] Anonymous. An improvement in picks. Scientific American, 80(11):165, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-165c.pdf.

Anonymous:1899:WHS

[9802] Anonymous. A water-heater for steam-boilers. Scientific American, 80(11):165, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
165b.pdf.

Anonymous:1899:AWLa

[9803] Anonymous. An automatic window-locking device. Scientific American, 80(11):166, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
166a.pdf.

Anonymous:1899:Cj

[9804] Anonymous. Correspondence. Scientific American, 80(11):166, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-166.pdf.

Brown:1899:GGW

[9805] Enos Brown. Grape growing and wine making in California. Scientific American, 80(11):168–170, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
168.pdf.

Anonymous:1899:GWM

[9806] Anonymous. George W. Melville, Chief of the Bureau of Steam Engineering. Scientific American, 80(11):170–171, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n11/pdf/scientificamerican03181899-170.pdf.

Anonymous:1899:NAC

[9807] Anonymous. The new America Cup Defender. Scientific American, 80(11):171, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
171.pdf.

Anonymous:1899:BNBh

[9808] Anonymous. Books: New books, etc. Scientific American, 80(11):172, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-172a.pdf.

Anonymous:1899:RPIk

[9809] Anonymous. Recently patented inventions. Scientific American, 80(11):172, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n11/pdf/scientificamerican03181899-
172.pdf.

Anonymous:1899:BPk

[9810] Anonymous. Business and personal. Scientific American, 80(11):173, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-173.pdf.

Anonymous:1899:IIk

[9811] Anonymous. Index of inventions. Scientific American, 80(11):173, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-173b.pdf.

Anonymous:1899:NQk

[9812] Anonymous. Notes and queries. Scientific American, 80(11):173, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-173a.pdf.

Anonymous:1899:PNPj

[9813] Anonymous. Patent notes: Patents. Scientific American, 80(11):175, March 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n11/pdf/scientificamerican03181899-175.pdf.

Anonymous:1899:CPU

[9814] Anonymous. Coal production in the United States. Scientific American, 80(12):178, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
178b.pdf.

Anonymous:1899:FPTa

[9815] Anonymous. Fire protection of tall buildings. Scientific American, 80(12):178, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
178a.pdf.

Anonymous:1899:PNS

[9816] Anonymous. Prof. Newcomb’s successor. Scientific American, 80(12):178, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
178c.pdf.

Anonymous:1899:SAP

[9817] Anonymous. The Senate and the armor plate question. Scientific American, 80(12):178, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
178.pdf.

Anonymous:1899:CPR

[9818] Anonymous. Characteristics of porto rico. Scientific American, 80(12):179, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
179a.pdf.

Anonymous:1899:CTS

[9819] Anonymous. Carriage timber supply. Scientific American, 80(12):179, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-179.pdf.

Anonymous:1899:AEL

[9820] Anonymous. Antarctic explorers land. Scientific American, 80 (12):179–180, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
179b.pdf.

Anonymous:1899:CMD

[9821] Anonymous. A curious means of disposing of the dead. Scientific American, 80(12):180, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
180.pdf.

Baltimore:1899:DTI

[9822] J. M. Baltimore. Discovery of two Indian mummies. Scientific American, 80(12):180, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
180a.pdf.

Anonymous:1899:DAM

[9823] Anonymous. A drop-actuating mechanism for seed-planters. Scientific American, 80(12):181, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
181b.pdf.

Anonymous:1899:EFS

[9824] Anonymous. An efficient fastener for stock-cars. Scientific American, 80(12):181, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
181c.pdf.

Anonymous:1899:IW

[9825] Anonymous. An improvement in wrenches. Scientific American, 80(12):181, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
181.pdf.

Anonymous:1899:SMG

[9826] Anonymous. A sewing-machine gage. Scientific American, 80(12):181, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-181a.pdf.

Anonymous:1899:SNG

[9827] Anonymous. Some notable German arch bridges. Scientific American, 80(12):182, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
182.pdf.

Anonymous:1899:Ck

[9828] Anonymous. Correspondence. Scientific American, 80(12):183, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-183.pdf.

Anonymous:1899:ECSa

[9829] Anonymous. The Electric Cab Service of New York City. Scientific American, 80(12):184–185, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
184.pdf.

Anonymous:1899:CCA

[9830] Anonymous. A curious case of abnormal development in a rattlesnake. Scientific American, 80(12):185, March 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n12/pdf/scientificamerican03251899-185.pdf.

Anonymous:1899:IEG

[9831] Anonymous. An interesting explosion of gas. Scientific American, 80(12):185, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
185a.pdf.

Hovey:1899:ACA

[9832] E. O. Hovey. Antique chain armor in the Caucasus Mountains. Scientific American, 80(12):186, March 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n12/pdf/scientificamerican03251899-186.pdf.

Anonymous:1899:GTPa

[9833] Anonymous. The great telescope at the Paris Exposition of 1900. Scientific American, 80(12):187, March 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n12/pdf/scientificamerican03251899-187.pdf.

Anonymous:1899:BNBi

[9834] Anonymous. Books: New books, etc. Scientific American, 80(12):188, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-188a.pdf.

Anonymous:1899:BPl

[9835] Anonymous. Business and personal. Scientific American, 80(12):188, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-188b.pdf.

Anonymous:1899:RPIl

[9836] Anonymous. Recently patented inventions. Scientific American, 80(12):188, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n12/pdf/scientificamerican03251899-
188.pdf.

Anonymous:1899:NQl

[9837] Anonymous. Notes and queries. Scientific American, 80(12):188–189, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-188c.pdf.

Anonymous:1899:IIl

[9838] Anonymous. Index of inventions. Scientific American, 80(12):189, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-189.pdf.

Anonymous:1899:PNPk

[9839] Anonymous. Patent notes: Patents. Scientific American, 80(12):191, March 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n12/pdf/scientificamerican03251899-191.pdf.

Anonymous:1899:ACS

[9840] Anonymous. Admiral Colomb on the Spanish War. Scientific American, 80(13):194, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
194.pdf.

Anonymous:1899:GCRa

[9841] Anonymous. The Great Central Railway of England. Scientific American, 80(13):194, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
194a.pdf.

Anonymous:1899:GVP

[9842] Anonymous. A gross violation of principle. Scientific American, 80(13):194, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
194c.pdf.

Anonymous:1899:NYTa

[9843] Anonymous. New York Truant School. Scientific American, 80(13):194, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-194b.pdf.

Anonymous:1899:GNL

[9844] Anonymous. Geographical names as legitimate trade marks. Scientific American, 80(13):194–195, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
194d.pdf.

Anonymous:1899:TAG

[9845] Anonymous. Testing acetylene generators. Scientific American, 80(13):195, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
195.pdf.

Anonymous:1899:TWJ

[9846] Anonymous. Trick wood joining. Scientific American, 80(13):196, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-196.pdf.

Holder:1899:IM

[9847] C. F. Holder. Insect migrations. Scientific American, 80(13):196, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-196a.pdf.

Anonymous:1899:Mb

[9848] Anonymous. The multiphone. Scientific American, 80(13):197, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-197a.pdf.

Anonymous:1899:SPH

[9849] Anonymous. A simple pipe-hanger. Scientific American, 80(13):197, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-197.pdf.

Anonymous:1899:MFL

[9850] Anonymous. A machine for filtering liquids. Scientific American, 80(13):197–198, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
197b.pdf.

Anonymous:1899:ISC

[9851] Anonymous. An improvement in smoke-consumers. Scientific American, 80(13):198, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
198.pdf.

Anonymous:1899:ITB

[9852] Anonymous. The improved turrets of the battleship “Texas”. Scientific American, 80(13):199, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
199.pdf.

Hoferer:1899:QCH

[9853] M. J. Hoferer. Quadruple compound harmonic motion. Scientific American, 80(13):200–201, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
200.pdf.

Anonymous:1899:LPC

[9854] Anonymous. The late Prof. O. C. Marsh. Scientific American, 80(13):201–202, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
201.pdf.

Anonymous:1899:T

[9855] Anonymous. Telephotography. Scientific American, 80(13):202–203, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-202.pdf.

Anonymous:1899:Cl

[9856] Anonymous. Correspondence. Scientific American, 80(13):204, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-204.pdf.

Anonymous:1899:RPIm

[9857] Anonymous. Recently patented inventions. Scientific American, 80(13):204, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
204a.pdf.

Anonymous:1899:BNBj

[9858] Anonymous. Books: New books, etc. Scientific American, 80 (13):204–205, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n13/pdf/scientificamerican04011899-
204b.pdf.

Anonymous:1899:BPm

[9859] Anonymous. Business and personal. Scientific American, 80(13):205, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-205.pdf.

Anonymous:1899:NQm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-205a.pdf.

Anonymous:1899:If

[9861] Anonymous. To inventors. Scientific American, 80(13):206, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-206.pdf.

Anonymous:1899:IIm

[9862] Anonymous. Index of inventions. Scientific American, 80(13):206, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-206a.pdf.

Anonymous:1899:PNPl

[9863] Anonymous. Patent notes: Patents. Scientific American, 80(13):209, April 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n13/pdf/scientificamerican04011899-209.pdf.

Anonymous:1899:FST

[9864] Anonymous. French submarine torpedo boats. Scientific American, 80(14):211, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
211.pdf.

Anonymous:1899:PBN

[9865] Anonymous. Proposal to build the New York rapid transit tunnel. Scientific American, 80(14):212, April 8, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n14/pdf/scientificamerican04081899-212a.pdf.

Anonymous:1899:RET

[9866] Anonymous. The recent explosion of a ten-inch Army gun. Scientific American, 80(14):212, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
212.pdf.

Anonymous:1899:USN

[9867] Anonymous. United States Navy and Armstrong guns at the proving grounds. Scientific American, 80(14):212, April 8, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n14/pdf/scientificamerican04081899-212b.pdf.

Anonymous:1899:ASb

[9868] Anonymous. The Audubon Society. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-213b.pdf.

Anonymous:1899:CLI

[9869] Anonymous. Congress of life insurance doctors in Belgium. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
213f.pdf.

Anonymous:1899:DGF

[9870] Anonymous. Death of General Flagler. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-213a.pdf.

Anonymous:1899:EHP

[9871] Anonymous. Electricity at high pressures. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
213.pdf.

Anonymous:1899:FE

[9872] Anonymous. Foreign exhibitions of 1899. Scientific American, 80 (14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
213e.pdf.

Anonymous:1899:SNCb

[9873] Anonymous. Spain’s new cruiser. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-213h.pdf.

Anonymous:1899:SNK

[9874] Anonymous. Some new kite experiments at Bayonne. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
213c.pdf.

Anonymous:1899:TWT

[9875] Anonymous. Tests of wireless telegraphy. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
213d.pdf.

Anonymous:1899:UTC

[9876] Anonymous. An under-trolley car Burns up. Scientific American, 80(14):213, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
213g.pdf.

Anonymous:1899:ILG

[9877] Anonymous. An improved lens-grinding machine. Scientific American, 80(14):214, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
214.pdf.

Anonymous:1899:NPH

[9878] Anonymous. A new power-hammer. Scientific American, 80(14):214, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-214b.pdf.

Anonymous:1899:PFL

[9879] Anonymous. Powerful freight locomotive for the Pennsylvania Railroad. Scientific American, 80(14):214, April 8, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n14/pdf/scientificamerican04081899-214a.pdf.

Anonymous:1899:CQ

[9880] Anonymous. Carrara and its quarries. Scientific American, 80 (14):215–216, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
215.pdf.

Anonymous:1899:NI

[9881] Anonymous. A novel insulator. Scientific American, 80(14):216, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-216.pdf.

Anonymous:1899:SVR

[9882] Anonymous. A simple ventilator for railway-cars. Scientific American, 80(14):216, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
216a.pdf.

Anonymous:1899:Cm

[9883] Anonymous. Correspondence. Scientific American, 80(14):217, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-217.pdf.

Maxim:1899:CRE

[9884] Hudson Maxim. Cause of the recent explosion of the ten-inch gun at Sandy Hook explained. Scientific American, 80(14):218–219, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-218.pdf.

Dodge:1899:GPE

[9885] Charles Richards Dodge. A glimpse of the Paris Exposition of 1900. Scientific American, 80(14):219–220, April 8, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n14/pdf/scientificamerican04081899-219.pdf.

Anonymous:1899:HH

[9886] Anonymous. The hornet and its home. Scientific American, 80 (14):220–221, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
220.pdf.

Anonymous:1899:BNBk

[9887] Anonymous. Books: New books, etc. Scientific American, 80(14):222, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-222a.pdf.

Anonymous:1899:RPIn

[9888] Anonymous. Recently patented inventions. Scientific American, 80(14):222, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n14/pdf/scientificamerican04081899-
222.pdf.

Anonymous:1899:BPn

[9889] Anonymous. Business and personal. Scientific American, 80(14):223, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-223.pdf.

Anonymous:1899:Ig

[9890] Anonymous. To inventors. Scientific American, 80(14):223, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-223b.pdf.

Anonymous:1899:IIn

[9891] Anonymous. Index of inventions. Scientific American, 80(14):223, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-223c.pdf.

Anonymous:1899:NQn

[9892] Anonymous. Notes and queries. Scientific American, 80(14):223, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-223a.pdf.

Anonymous:1899:PNPm

[9893] Anonymous. Patent notes: Patents. Scientific American, 80(14):225, April 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n14/pdf/scientificamerican04081899-225.pdf.

Anonymous:1899:CAa

[9894] Anonymous. A critic answered. Scientific American, 80(15):228, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-228d.pdf.

Anonymous:1899:FPTb

[9895] Anonymous. Fire protection of tall buildings. Scientific American, 80(15):228, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
228a.pdf.

Anonymous:1899:HSF

[9896] Anonymous. High speed on French railways. Scientific American, 80(15):228, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
228b.pdf.

Anonymous:1899:LBC

[9897] Anonymous. Latest battleships and cruisers for the British Navy. Scientific American, 80(15):228, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
228c.pdf.

Anonymous:1899:ORP

[9898] Anonymous. Our relation to the peace Congress. Scientific American, 80(15):228, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
228.pdf.

Anonymous:1899:ONC

[9899] Anonymous. Our Naval constructors to be educated abroad. Scientific American, 80(15):228–229, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
228e.pdf.

Anonymous:1899:HBN

[9900] Anonymous. The honey bee not a native of America. Scientific American, 80(15):229, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
229b.pdf.

Anonymous:1899:NSW

[9901] Anonymous. New system of wireless telegraphy. Scientific American, 80(15):229, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
229.pdf.

Anonymous:1899:WEC

[9902] Anonymous. Woehnelt’s electrostatic current breaker. Scientific American, 80(15):229, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
229a.pdf.

Anonymous:1899:BAGa

[9903] Anonymous. The Bundy acetylene gas lamp. Scientific American, 80(15):230, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
230a.pdf.

Anonymous:1899:NHS

[9904] Anonymous. A new hay-stacker. Scientific American, 80(15):230, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-230.pdf.

Meade:1899:CFA

[9905] Richard K. Meade. A cheap form of air-bath. Scientific American, 80(15):230, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
230b.pdf.

Anonymous:1899:ICB

[9906] Anonymous. An improved chicken-brooder. Scientific American, 80(15):231, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
231.pdf.

Anonymous:1899:SVW

[9907] Anonymous. A simple vehicle wheel. Scientific American, 80(15):231, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-231a.pdf.

Anonymous:1899:TUS

[9908] Anonymous. Turrets of United States Naval Vessels. Scientific American, 80(15):231–232, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
231b.pdf.

Anonymous:1899:Cn

[9909] Anonymous. Correspondence. Scientific American, 80(15):232, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-232.pdf.

Sheffield:1899:OWP

[9910] W. M. Sheffield. Opening the White Pass and Yukon Railway. Scientific American, 80(15):233, April 15, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n15/pdf/scientificamerican04151899-233.pdf.

Anonymous:1899:BBT

[9911] Anonymous. The Beach Broadway Tunnel. Scientific American, 80(15):234, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
234.pdf.

Anonymous:1899:PTD

[9912] Anonymous. Pneumatic tube delivery system at the Waldorf–Astoria Hotel. Scientific American, 80(15):235–236, April 15, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n15/pdf/scientificamerican04151899-235.pdf.

Anonymous:1899:RAC

[9913] Anonymous. Rear-Admiral Charles O’Neil. Scientific American, 80(15):236, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
236.pdf.

Wade:1899:NTL

[9914] Herbert T. Wade. The new testing locomotive at Columbia University. Scientific American, 80(15):237, April 15, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n15/pdf/scientificamerican04151899-237.pdf.

Anonymous:1899:BNBl

[9915] Anonymous. Books: New books, etc. Scientific American, 80(15):238, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-238a.pdf.

Anonymous:1899:BPp

[9916] Anonymous. Business and personal. Scientific American, 80(15):238, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-238b.pdf.

Anonymous:1899:Ih

[9917] Anonymous. To inventors. Scientific American, 80(15):238, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-238d.pdf.

Anonymous:1899:IIo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-238e.pdf.

Anonymous:1899:NQo

[9919] Anonymous. Notes and queries. Scientific American, 80(15):238, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-238c.pdf.

Anonymous:1899:RPIo

[9920] Anonymous. Recently patented inventions. Scientific American, 80(15):238, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n15/pdf/scientificamerican04151899-
238.pdf.

Anonymous:1899:PNPn

[9921] Anonymous. Patent notes: Patents. Scientific American, 80(15):240, April 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n15/pdf/scientificamerican04151899-240.pdf.

Anonymous:1899:EWH

[9922] Anonymous. Echoes of the Windsor Hotel Fire. Scientific American, 80(16):244, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
244a.pdf.

Anonymous:1899:NNW

[9923] Anonymous. Naming the new warships. Scientific American, 80 (16):244, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
244b.pdf.

Anonymous:1899:TEW

[9924] Anonymous. The thin end of the wedge. Scientific American, 80(16):244, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
244.pdf.

Anonymous:1899:RTB

[9925] Anonymous. A record torpedo boat trial. Scientific American, 80(16):244–245, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
244c.pdf.

Anonymous:1899:NSS

[9926] Anonymous. The new satellite of Saturn. Scientific American, 80(16):245, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
245a.pdf.

Anonymous:1899:PZP

[9927] Anonymous. Progress of the Zoological Park. Scientific American, 80(16):245, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
245b.pdf.

Mokton:1899:LAN

[9928] Henry Mokton. Liquid air as a new source of power — another engineering fallacy. Scientific American, 80(16):245, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-245.pdf.

Anonymous:1899:BNC

[9929] Anonymous. A boiler of novel construction. Scientific American, 80(16):246, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
246.pdf.

Anonymous:1899:IFA

[9930] Anonymous. An improved fire-alarm. Scientific American, 80(16):246, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-246b.pdf.

Anonymous:1899:SBSa

[9931] Anonymous. A Ship’s boat supporting and launching device. Scientific American, 80(16):246, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
246a.pdf.

Anonymous:1899:CWb

[9932] Anonymous. Cardinal Wolsey’s hat. Scientific American, 80(16):247, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-247.pdf.

Anonymous:1899:MBa

[9933] Anonymous. Military ballooning. Scientific American, 80(16):247, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-247a.pdf.

Anonymous:1899:CSDa

[9934] Anonymous. A cap for sliding doors. Scientific American, 80 (16):247–248, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
247b.pdf.

Anonymous:1899:CDB

[9935] Anonymous. A convertible dusting-brush. Scientific American, 80(16):248, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
248.pdf.

Anonymous:1899:NFFa

[9936] Anonymous. A new form of fire-escape. Scientific American, 80(16):248, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-248a.pdf.

Anonymous:1899:Co

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-249.pdf.

Anonymous:1899:NWd

[9938] Anonymous. Navies of the world. Scientific American, 80(16):250–252, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-250.pdf.

Holder:1899:SPC

[9939] C. F. Holder. Some Pacific cephalopods. Scientific American, 80(16):253, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
253.pdf.

Anonymous:1899:RPIp

[9940] Anonymous. Recently patented inventions. Scientific American, 80(16):254, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
254.pdf.

Anonymous:1899:BNBm

[9941] Anonymous. Books: New books, etc. Scientific American, 80 (16):254–255, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n16/pdf/scientificamerican04221899-
254a.pdf.

Anonymous:1899:BPq

[9942] Anonymous. Business and personal. Scientific American, 80(16):255, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-255.pdf.

Anonymous:1899:Ij

[9943] Anonymous. To inventors. Scientific American, 80(16):255, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-255b.pdf.

Anonymous:1899:IIp

[9944] Anonymous. Index of inventions. Scientific American, 80(16):255, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-255c.pdf.

Anonymous:1899:NQp

[9945] Anonymous. Notes and queries. Scientific American, 80(16):255, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-255a.pdf.

Anonymous:1899:PNPo

[9946] Anonymous. Patent notes: Patents. Scientific American, 80(16):256, April 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n16/pdf/scientificamerican04221899-256.pdf.

Anonymous:1899:ABCa

[9947] Anonymous. That Athbara bridge contract. Scientific American, 80(17):260, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
260.pdf.

Anonymous:1899:OTN

[9948] Anonymous. Official test of the new Army wire gun. Scientific American, 80(17):260, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
260c.pdf.

Anonymous:1899:SWT

[9949] Anonymous. Sine-wave telegraphy. Scientific American, 80(17):260, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-260b.pdf.

Anonymous:1899:WTAa

[9950] Anonymous. Wireless telegraphy across the English channel. Scientific American, 80(17):260, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
260a.pdf.

Anonymous:1899:WOC

[9951] Anonymous. What our consuls are doing for American trade. Scientific American, 80(17):260–261, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
260d.pdf.

Anonymous:1899:AMT

[9952] Anonymous. American machine tools in England. Scientific American, 80(17):261, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
261.pdf.

Anonymous:1899:PEF

[9953] Anonymous. A Parisian egg forger. Scientific American, 80(17):261, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-261c.pdf.

Benjamin:1899:NAS

[9954] Marcus Benjamin. National Academy of Sciences. Scientific American, 80(17):261, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
261b.pdf.

Serviss:1899:HMb

[9955] Garrett P. Serviss. The heavens in May. Scientific American, 80(17):261, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
261a.pdf.

Anonymous:1899:MER

[9956] Anonymous. The Marvin electric rock drill. Scientific American, 80(17):262–263, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
262.pdf.

Anonymous:1899:CPb

[9957] Anonymous. A convertible plow. Scientific American, 80(17):263, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-263.pdf.

Anonymous:1899:IBS

[9958] Anonymous. An improved band-saw-setting machine. Scientific American, 80(17):263, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
263a.pdf.

Anonymous:1899:CMM

[9959] Anonymous. A cheap method of making a calorimeter. Scientific American, 80(17):263–264, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
263b.pdf.

Anonymous:1899:ASF

[9960] Anonymous. An automatic Spring fish-hook. Scientific American, 80(17):264, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
264.pdf.

Anonymous:1899:ILS

[9961] Anonymous. An improved lemon-squeezer. Scientific American, 80(17):264, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
264a.pdf.

Anonymous:1899:Cr

[9962] Anonymous. Correspondence. Scientific American, 80(17):265, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-265.pdf.

Parsons:1899:SBS

[9963] F. W. Parsons. Sutro Baths of San Francisco. Scientific American, 80(17):266, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
266.pdf.

Anonymous:1899:MKJa

[9964] Anonymous. Manufacture of Krag–Jørgensen rifles at the Springfield Armory — I. Scientific American, 80(17):267–268, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-267.pdf.

Anonymous:1899:CSLa

[9965] Anonymous. The Californian sea lion. Scientific American, 80 (17):268–269, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
268.pdf.

Anonymous:1899:RPIq

[9966] Anonymous. Recently patented inventions. Scientific American, 80(17):269, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n17/pdf/scientificamerican04291899-
269.pdf.

Anonymous:1899:BNBn

[9967] Anonymous. Books: New books, etc. Scientific American, 80(17):270, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-270b.pdf.

Anonymous:1899:BPr

[9968] Anonymous. Business and personal. Scientific American, 80(17):270, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-270.pdf.

Anonymous:1899:Ik

[9969] Anonymous. To inventors. Scientific American, 80(17):270, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-270c.pdf.

Anonymous:1899:IIq

[9970] Anonymous. Index of inventions. Scientific American, 80(17):270, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-270d.pdf.

Anonymous:1899:NQq

[9971] Anonymous. Notes and queries. Scientific American, 80(17):270, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-270a.pdf.

Anonymous:1899:PNPp

[9972] Anonymous. Patent notes: Patents. Scientific American, 80(17):272, April 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n17/pdf/scientificamerican04291899-272.pdf.

Anonymous:1899:CMU

[9973] Anonymous. Crystallization of metal under stress. Scientific American, 80(18):276, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
276.pdf.

Anonymous:1899:PCH

[9974] Anonymous. Prof. Chandler honored. Scientific American, 80(18):276, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-276c.pdf.

Anonymous:1899:UFM

[9975] Anonymous. The uses of Florida moss. Scientific American, 80(18):276, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-276b.pdf.

Anonymous:1899:YC

[9976] Anonymous. The yacht “Columbia”. Scientific American, 80(18):276, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-276a.pdf.

Anonymous:1899:DTR

[9977] Anonymous. The durability of typewriter ribbons. Scientific American, 80(18):277, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
277c.pdf.

Anonymous:1899:MAS

[9978] Anonymous. Meeting of the American Society of Mechanical Engineers. Scientific American, 80(18):277, May 6, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n18/pdf/scientificamerican05061899-277a.pdf.

Anonymous:1899:Oa

[9979] Anonymous. Ozotype. Scientific American, 80(18):277, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-277b.pdf.

Anonymous:1899:RTS

[9980] Anonymous. The Rowland telegraph system. Scientific American, 80(18):277, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
277e.pdf.

Anonymous:1899:TFD

[9981] Anonymous. The Tugrin fog dispeller. Scientific American, 80(18):277, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-277d.pdf.

Anonymous:1899:TM

[9982] Anonymous. Topographical maps. Scientific American, 80(18):277, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-277.pdf.

Anonymous:1899:DMR

[9983] Anonymous. A deep-mine or relay pump. Scientific American, 80(18):278, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
278b.pdf.

Anonymous:1899:IFCa

[9984] Anonymous. An improvement in furnace construction. Scientific American, 80(18):278, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
278a.pdf.

Anonymous:1899:NBTb

[9985] Anonymous. A novel boiler-tube cutter and expander. Scientific American, 80(18):278, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
278.pdf.

Anonymous:1899:QOJ

[9986] Anonymous. A quaint old Japanese clock. Scientific American, 80(18):278, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
278c.pdf.

Anonymous:1899:NSR

[9987] Anonymous. A new shade-roller attachment. Scientific American, 80(18):279, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
279.pdf.

Anonymous:1899:RWM

[9988] Anonymous. A rapid way of marking out cloth patterns. Scientific American, 80(18):279, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
279a.pdf.

Anonymous:1899:DRD

[9989] Anonymous. A device for removing dents from gun-barrels. Scientific American, 80(18):279–280, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
279b.pdf.

Anonymous:1899:KC

[9990] Anonymous. A knockdown cask. Scientific American, 80(18):280, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-280.pdf.

Anonymous:1899:Cs

[9991] Anonymous. Correspondence. Scientific American, 80(18):281, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-281.pdf.

Anonymous:1899:IGE

[9992] Anonymous. The 10-inch gun explosion at Sandy Hook. Scientific American, 80(18):281, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
281a.pdf.

Anonymous:1899:CMa

[9993] Anonymous. The class microscope. Scientific American, 80(18):282, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-282.pdf.

Marion:1899:NAI

[9994] H. Marion. The Natal Academy as it is. Scientific American, 80(18):283–284, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
283.pdf.

Anonymous:1899:SI

[9995] Anonymous. The Samoan Islands. Scientific American, 80(18):285, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-285.pdf.

Anonymous:1899:BNBo

[9996] Anonymous. Books: New books, etc. Scientific American, 80(18):286, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-286c.pdf.

Anonymous:1899:BPs

[9997] Anonymous. Business and personal. Scientific American, 80(18):286, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-286a.pdf.

Anonymous:1899:IIr

[9998] Anonymous. Index of inventions. Scientific American, 80(18):286, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-286d.pdf.

Anonymous:1899:NQr

[9999] Anonymous. Notes and queries. Scientific American, 80(18):286, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-286b.pdf.

Anonymous:1899:RPIr

[10000] Anonymous. Recently patented inventions. Scientific American, 80(18):286, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n18/pdf/scientificamerican05061899-
286.pdf.

Anonymous:1899:PNPq

[10001] Anonymous. Patent notes: Patents. Scientific American, 80(18):289, May 6, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n18/pdf/scientificamerican05061899-289.pdf.

Anonymous:1899:CBa

[10002] Anonymous. The Charm of the bicycle. Scientific American, 80(19):292, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-292.pdf.

Anonymous:1899:FCW

[10003] Anonymous. The fastest cruiser in the world. Scientific American, 80(19):292, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
292c.pdf.

Anonymous:1899:OSC

[10004] Anonymous. Our sea-coast defenses. Scientific American, 80(19):292, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-292b.pdf.

Anonymous:1899:RSE

[10005] Anonymous. Records of speed and endurance on the bicycle. Scientific American, 80(19):292, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
292a.pdf.

Anonymous:1899:AP

[10006] Anonymous. Automobilism in Paris. Scientific American, 80 (19):293–294, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
293.pdf.

Anonymous:1899:WEM

[10007] Anonymous. Woods’ electric motor vehicles. Scientific American, 80(19):294, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
294.pdf.

Anonymous:1899:EMV

[10008] Anonymous. Electric motor vehicles. Scientific American, 80(19):295, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-295.pdf.

Anonymous:1899:SEF

[10009] Anonymous. Some early forms of the automobile. Scientific American, 80(19):296, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
296.pdf.

Anonymous:1899:MBCa

[10010] Anonymous. A mechanical bicycle-chain cleaner. Scientific American, 80(19):297, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
297a.pdf.

Anonymous:1899:NMM

[10011] Anonymous. New method of mounting cyclometers. Scientific American, 80(19):297, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
297.pdf.

Anonymous:1899:NSPa

[10012] Anonymous. A novel stirrup-pedal for bicycles. Scientific American, 80(19):297, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
297b.pdf.

Anonymous:1899:SBSb

[10013] Anonymous. A simple bicycle-support. Scientific American, 80(19):297, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-297c.pdf.

Anonymous:1899:DFTa

[10014] Anonymous. The daring feats of a trick-rider. Scientific American, 80(19):298, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
298.pdf.

Anonymous:1899:HB

[10015] Anonymous. The history of the bicycle. Scientific American, 80 (19):298–299, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
298a.pdf.

Anonymous:1899:ATR

[10016] Anonymous. The art of trick riding. Scientific American, 80(19):299, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-299.pdf.

Anonymous:1899:EFLa

[10017] Anonymous. Early forms of letters patent. Scientific American, 80(19):300, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
300.pdf.

Anonymous:1899:WMC

[10018] Anonymous. Winton motor carriages. Scientific American, 80 (19):300–301, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
300a.pdf.

Anonymous:1899:SRR

[10019] Anonymous. Some recent researches on the Great Nebula in Orion. Scientific American, 80(19):302–303, May 13, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n19/pdf/scientificamerican05131899-302.pdf.

Prindle:1899:RBB

[10020] E. J. Prindle. Riding a bicycle backward. Scientific American, 80(19):303–304, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
303.pdf.

Anonymous:1899:CMC

[10021] Anonymous. Columbia motor carriages. Scientific American, 80 (19):304–305, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
304.pdf.

Anonymous:1899:ETI

[10022] Anonymous. Explosion of a ten-inch gun at Sandy Hook. Scientific American, 80(19):305–306, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
305.pdf.

Anonymous:1899:DMC

[10023] Anonymous. The Decauville motor-carriage. Scientific American, 80(19):306, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
306.pdf.

Anonymous:1899:DBT

[10024] Anonymous. The De Dion–Bouton tricycle. Scientific American, 80(19):307–308, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
307.pdf.

Anonymous:1899:CPAa

[10025] Anonymous. A combination pleasure automobile delivery vehicle. Scientific American, 80(19):308, May 13, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n19/pdf/scientificamerican05131899-308.pdf.

Anonymous:1899:LTB

[10026] Anonymous. Laboratory tests of the bicycle. Scientific American, 80(19):309, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
309.pdf.

Anonymous:1899:BGW

[10027] Anonymous. Bicycle gear — what it is and what it does. Scientific American, 80(19):309–310, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
309a.pdf.

Anonymous:1899:TCB

[10028] Anonymous. The Trebert coaster and brake. Scientific American, 80(19):310, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
310.pdf.

Anonymous:1899:VA

[10029] Anonymous. The Victor automobile. Scientific American, 80(19):310, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-310a.pdf.

Anonymous:1899:ATb

[10030] Anonymous. An amphibious tricycle. Scientific American, 80(19):312, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-312.pdf.

Anonymous:1899:RPIs

[10031] Anonymous. Recently patented inventions. Scientific American, 80(19):312, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n19/pdf/scientificamerican05131899-
312a.pdf.

Anonymous:1899:BNBp

[10032] Anonymous. Books: New books, etc. Scientific American, 80(19):315, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-315.pdf.

Anonymous:1899:BPt

[10033] Anonymous. Business and personal. Scientific American, 80(19):315, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-315a.pdf.

Anonymous:1899:NQs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-315b.pdf.

Anonymous:1899:IIt

[10035] Anonymous. Index of inventions. Scientific American, 80(19):319, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-319.pdf.

Anonymous:1899:PNPr

[10036] Anonymous. Patent notes: Patents. Scientific American, 80(19):321, May 13, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n19/pdf/scientificamerican05131899-321.pdf.

Anonymous:1899:EE

[10037] Anonymous. The electrical exhibition. Scientific American, 80(20):324, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-324a.pdf.

Anonymous:1899:MS

[10038] Anonymous. “Magazine Science”. Scientific American, 80(20):324, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-324.pdf.

Anonymous:1899:LSW

[10039] Anonymous. Loss of speed in warships. Scientific American, 80 (20):324–325, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
324b.pdf.

Anonymous:1899:DPBa

[10040] Anonymous. Death of Prof. Buechner. Scientific American, 80(20):325, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-325c.pdf.

Anonymous:1899:GST

[10041] Anonymous. The Gatemann shell tested. Scientific American, 80 (20):325, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
325b.pdf.

Anonymous:1899:PCa

[10042] Anonymous. Potassium chloroplatinite. Scientific American, 80 (20):325, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
325.pdf.

Scallan:1899:PUD

[10043] R. Scallan. The proposed universal day. Scientific American, 80(20):325, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-325a.pdf.

Anonymous:1899:ASA

[10044] Anonymous. An automatic safety appliance for railroads. Scientific American, 80(20):326, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
326b.pdf.

Anonymous:1899:NHG

[10045] Anonymous. A new hay or grain rack. Scientific American, 80(20):326, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-326.pdf.

Anonymous:1899:NLM

[10046] Anonymous. The new locomotives for the Midland Railway. Scientific American, 80(20):326, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
326a.pdf.

Anonymous:1899:CAP

[10047] Anonymous. A coupling for the air-pipes of railway-cars. Scientific American, 80(20):327, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
327a.pdf.

Anonymous:1899:IPI

[10048] Anonymous. An improved pruning implement. Scientific American, 80(20):327, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
327c.pdf.

Anonymous:1899:PHH

[10049] Anonymous. A Pitman-head for high-speed machines. Scientific American, 80(20):327, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
327.pdf.

Anonymous:1899:WTBa

[10050] Anonymous. A wrench for tire-bolts. Scientific American, 80(20):327, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-327b.pdf.

Anonymous:1899:Cu

[10051] Anonymous. Correspondence. Scientific American, 80(20):328–329, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-328.pdf.

Anonymous:1899:BIa

[10052] Anonymous. A bedstead for invalids. Scientific American, 80(20):329, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-329.pdf.

Anonymous:1899:MKJb

[10053] Anonymous. Manufacture of Krag–Jørgensen rifles at the Springfield Armory — II. Scientific American, 80(20):330–331, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-330.pdf.

Anonymous:1899:PMA

[10054] Anonymous. The physiology of man on the alps. Scientific American, 80(20):331–332, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
331.pdf.

Anonymous:1899:AGH

[10055] Anonymous. An abnormal growth of hair on a horse. Scientific American, 80(20):333, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
333a.pdf.

Anonymous:1899:IRM

[10056] Anonymous. An ingenious reversing mechanism for machine tools. Scientific American, 80(20):333, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
333.pdf.

Anonymous:1899:BNBq

[10057] Anonymous. Books: New books, etc. Scientific American, 80(20):334, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-334a.pdf.

Anonymous:1899:RPIt

[10058] Anonymous. Recently patented inventions. Scientific American, 80(20):334, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n20/pdf/scientificamerican05201899-
334.pdf.

Anonymous:1899:BPu

[10059] Anonymous. Business and personal. Scientific American, 80(20):335, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-335.pdf.

Anonymous:1899:Il

[10060] Anonymous. To inventors. Scientific American, 80(20):335, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-335b.pdf.

Anonymous:1899:IIu

[10061] Anonymous. Index of inventions. Scientific American, 80(20):335, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-335c.pdf.

Anonymous:1899:NQt

[10062] Anonymous. Notes and queries. Scientific American, 80(20):335, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-335a.pdf.

Anonymous:1899:PNPs

[10063] Anonymous. Patent notes: Patents. Scientific American, 80(20):337, May 20, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n20/pdf/scientificamerican05201899-337.pdf.

Anonymous:1899:FFE

[10064] Anonymous. The forty-foot entrance to New York Harbor. Scientific American, 80(21):340, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
340a.pdf.

Anonymous:1899:HOA

[10065] Anonymous. How our art manufactures may affect our exports. Scientific American, 80(21):340, May 27, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n21/pdf/scientificamerican05271899-340c.pdf.

Anonymous:1899:QGE

[10066] Anonymous. A question of Governmental economics. Scientific American, 80(21):340, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
340b.pdf.

Anonymous:1899:SPR

[10067] Anonymous. Is smokeless powder reliable? Scientific American, 80(21):340, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
340.pdf.

Anonymous:1899:MEP

[10068] Anonymous. The mechanical engineers and the Patent Office. Scientific American, 80(21):340–341, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
340d.pdf.

Anonymous:1899:BIb

[10069] Anonymous. The bicycle industry. Scientific American, 80(21):341, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-341a.pdf.

Anonymous:1899:WT

[10070] Anonymous. Wireless telegraphy. Scientific American, 80(21):341, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-341.pdf.

Serviss:1899:HJa

[10071] Garrett P. Serviss. The heavens in June. Scientific American, 80(21):341, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
341b.pdf.

Anonymous:1899:CCb

[10072] Anonymous. The chromo-camera. Scientific American, 80(21):342, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-342.pdf.

Anonymous:1899:NGG

[10073] Anonymous. A new gas-generating machine. Scientific American, 80(21):342–343, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
342a.pdf.

Anonymous:1899:NIC

[10074] Anonymous. Niagara as an industrial center. Scientific American, 80(21):343–344, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
343.pdf.

Anonymous:1899:ASc

[10075] Anonymous. An automatic siphon. Scientific American, 80(21):344, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-344a.pdf.

Anonymous:1899:NPF

[10076] Anonymous. A novel portable folding boat. Scientific American, 80(21):344, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
344.pdf.

Anonymous:1899:Cx

[10077] Anonymous. Correspondence. Scientific American, 80(21):345, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-345.pdf.

Anonymous:1899:OLBa

[10078] Anonymous. The old line-of-battle ship “Pennsylvania”. Scientific American, 80(21):346–347, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
346.pdf.

Anonymous:1899:REE

[10079] Anonymous. The radiophone at the electrical exhibition. Scientific American, 80(21):347, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
347.pdf.

Holder:1899:AC

[10080] C. F. Holder. Animal communities. Scientific American, 80 (21):347–348, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
347a.pdf.

Anonymous:1899:EBa

[10081] Anonymous. The efficiency of the bicycle. Scientific American, 80(21):348–349, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
348.pdf.

Anonymous:1899:ICF

[10082] Anonymous. An improvement in chemical fire-extinguishers. Scientific American, 80(21):349, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
349.pdf.

Anonymous:1899:BNBr

[10083] Anonymous. Books: New books, etc. Scientific American, 80(21):350, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-350c.pdf.

Anonymous:1899:BPv

[10084] Anonymous. Business and personal. Scientific American, 80(21):350, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-350a.pdf.

Anonymous:1899:IIv

[10085] Anonymous. Index of inventions. Scientific American, 80(21):350, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-350d.pdf.

Anonymous:1899:NQu

[10086] Anonymous. Notes and queries. Scientific American, 80(21):350, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-350b.pdf.

Anonymous:1899:RPIu

[10087] Anonymous. Recently patented inventions. Scientific American, 80(21):350, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n21/pdf/scientificamerican05271899-
350.pdf.

Anonymous:1899:PNPt

[10088] Anonymous. Patent notes: Patents. Scientific American, 80(21):353, May 27, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n21/pdf/scientificamerican05271899-353.pdf.

Anonymous:1899:ERT

[10089] Anonymous. The East River Tunnel. Scientific American, 80(22):356, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-356b.pdf.

Anonymous:1899:GAU

[10090] Anonymous. Gruson armor in the United States. Scientific American, 80(22):356, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
356a.pdf.

Anonymous:1899:GP

[10091] Anonymous. Gold in the Philippines. Scientific American, 80(22):356, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-356d.pdf.

Anonymous:1899:LAH

[10092] Anonymous. Liquid air in high explosives. Scientific American, 80(22):356, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
356.pdf.

Anonymous:1899:PAE

[10093] Anonymous. The Pan-American Exposition. Scientific American, 80(22):356, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
356c.pdf.

Anonymous:1899:CPL

[10094] Anonymous. Commercial production of liquid air. Scientific American, 80(22):357, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
357d.pdf.

Anonymous:1899:PPD

[10095] Anonymous. Photography of paintings and drawings. Scientific American, 80(22):357, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
357b.pdf.

Anonymous:1899:PPP

[10096] Anonymous. The plant products of the Philippine Islands. Scientific American, 80(22):357, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
357a.pdf.

Anonymous:1899:RMa

[10097] Anonymous. The “Reina Mercedes”. Scientific American, 80(22):357, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-357c.pdf.

Mgahie:1899:SSP

[10098] Frederick H. M’gahie. Some smokeless powder considerations. Scientific American, 80(22):357, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
357.pdf.

Anonymous:1899:ISJ

[10099] Anonymous. An iceberg at St. John’s, Newfoundland. Scientific American, 80(22):358, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
358a.pdf.

Anonymous:1899:PVE

[10100] Anonymous. Pocket voltmeter and an electric soldering iron. Scientific American, 80(22):358, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
358.pdf.

Anonymous:1899:CTc

[10101] Anonymous. A Chinese typewriter. Scientific American, 80(22):359, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-359a.pdf.

Anonymous:1899:FWP

[10102] Anonymous. A feed-water purifier and skimmer for boilers. Scientific American, 80(22):359, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
359b.pdf.

Anonymous:1899:POC

[10103] Anonymous. A powerful open circuit battery cell. Scientific American, 80(22):359, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
359.pdf.

Anonymous:1899:MSS

[10104] Anonymous. The measurement of shutter speed. Scientific American, 80(22):360, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
360a.pdf.

Anonymous:1899:PCN

[10105] Anonymous. Photography in the colors of nature. Scientific American, 80(22):360, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
360.pdf.

Anonymous:1899:Cy

[10106] Anonymous. Correspondence. Scientific American, 80(22):361, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-361.pdf.

Anonymous:1899:NWe

[10107] Anonymous. Navies of the world. Scientific American, 80(22):362–364, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-362.pdf.

Holder:1899:BG

[10108] Charles Frederick Holder. The bird giants. Scientific American, 80(22):364–365, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
364.pdf.

Anonymous:1899:BPw

[10109] Anonymous. Business and personal. Scientific American, 80(22):366, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-366a.pdf.

Anonymous:1899:Im

[10110] Anonymous. To inventors. Scientific American, 80(22):366, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-366c.pdf.

Anonymous:1899:NQv

[10111] Anonymous. Notes and queries. Scientific American, 80(22):366, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-366b.pdf.

Anonymous:1899:RPIv

[10112] Anonymous. Recently patented inventions. Scientific American, 80(22):366, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n22/pdf/scientificamerican06031899-
366.pdf.

Anonymous:1899:IIw

[10113] Anonymous. Index of inventions. Scientific American, 80(22):366–369, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-366d.pdf.

Anonymous:1899:PNPu

[10114] Anonymous. Patent notes: Patents. Scientific American, 80(22):369, June 3, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n22/pdf/scientificamerican06031899-369.pdf.

Anonymous:1899:LAS

[10115] Anonymous. Liquid air “Surplusage”. Scientific American, 80(23):372, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-372.pdf.

Anonymous:1899:RNC

[10116] Anonymous. Report of the Nicaragua Canal Commission. Scientific American, 80(23):372, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
372b.pdf.

Anonymous:1899:RW

[10117] Anonymous. The reason why. Scientific American, 80(23):372, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-372a.pdf.

Anonymous:1899:BNA

[10118] Anonymous. Brasseys Naval Annual. Scientific American, 80 (23):372–373, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
372c.pdf.

Anonymous:1899:ATC

[10119] Anonymous. The automobile trip from Cleveland to New York. Scientific American, 80(23):373, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
373.pdf.

Anonymous:1899:BRE

[10120] Anonymous. The birth-rate in Europe. Scientific American, 80(23):373, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-373b.pdf.

Anonymous:1899:PR

[10121] Anonymous. Photography in relief. Scientific American, 80(23):373, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-373a.pdf.

Anonymous:1899:IHP

[10122] Anonymous. An improvised hydraulic press. Scientific American, 80(23):374, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
374a.pdf.

Anonymous:1899:SCFa

[10123] Anonymous. A simple cabinet for filing papers. Scientific American, 80(23):374, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
374.pdf.

Anonymous:1899:NFSa

[10124] Anonymous. A new form of stereopticon. Scientific American, 80(23):374–375, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
374b.pdf.

Anonymous:1899:BFW

[10125] Anonymous. The Brickill feed water case. Scientific American, 80(23):375, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
375.pdf.

Maxim:1899:PEW

[10126] Hudson Maxim. Philosophy of explosive wave action. Scientific American, 80(23):376, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
376.pdf.

Anonymous:1899:Cz

[10127] Anonymous. Correspondence. Scientific American, 80(23):377, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-377.pdf.

Anonymous:1899:MHG

[10128] Anonymous. The manufacture of high grade linen ledger paper. Scientific American, 80(23):378–380, June 10, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n23/pdf/scientificamerican06101899-378.pdf.

Anonymous:1899:WER

[10129] Anonymous. Waverly electric runabout and Columbia electric emergency wagon. Scientific American, 80(23):381, June 10, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n23/pdf/scientificamerican06101899-381.pdf.

Anonymous:1899:BNBs

[10130] Anonymous. Books: New books, etc. Scientific American, 80(23):382, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-382c.pdf.

Anonymous:1899:BPx

[10131] Anonymous. Business and personal. Scientific American, 80(23):382, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-382a.pdf.

Anonymous:1899:IIx

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-382d.pdf.

Anonymous:1899:NQw

[10133] Anonymous. Notes and queries. Scientific American, 80(23):382, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-382b.pdf.

Anonymous:1899:RPIw

[10134] Anonymous. Recently patented inventions. Scientific American, 80(23):382, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n23/pdf/scientificamerican06101899-
382.pdf.

Anonymous:1899:PNPv

[10135] Anonymous. Patent notes: Patents. Scientific American, 80(23):385, June 10, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n23/pdf/scientificamerican06101899-385.pdf.

Anonymous:1899:CVW

[10136] Anonymous. Commercial value of wireless telegraphy. Scientific American, 80(24):338, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
388c.pdf.

Anonymous:1899:NCC

[10137] Anonymous. The new Canal Commission. Scientific American, 80(24):388, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
388b.pdf.

Anonymous:1899:PN

[10138] Anonymous. A peril to the Navy. Scientific American, 80(24):388, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n24/pdf/scientificamerican06171899-388.pdf.

Anonymous:1899:WPR

[10139] Anonymous. What poor roads cost our farmers. Scientific American, 80(24):388, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
388a.pdf.

Anonymous:1899:MBE

[10140] Anonymous. Municipal bathing establishments. Scientific American, 80(24):388–389, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
388d.pdf.

Anonymous:1899:COI

[10141] Anonymous. California olive industry. Scientific American, 80(24):389, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n24/pdf/scientificamerican06171899-389.pdf.

Anonymous:1899:DFTb

[10142] Anonymous. The death of Frank Thomson. Scientific American, 80(24):389, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
389b.pdf.

Anonymous:1899:TCO

[10143] Anonymous. Trade with Cuba and our new possessions. Scientific American, 80(24):389, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
389c.pdf.

Anonymous:1899:ACF

[10144] Anonymous. An air-cushion fluid-pressure regulator. Scientific American, 80(24):390, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
390b.pdf.

Anonymous:1899:IAG

[10145] Anonymous. An improved acetylene generator. Scientific American, 80(24):390, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
390a.pdf.

Anonymous:1899:SFG

[10146] Anonymous. A simple fastener for gates. Scientific American, 80(24):390, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
390.pdf.

Anonymous:1899:RBD

[10147] Anonymous. A remarkable bridge disaster. Scientific American, 80(24):391–392, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
391.pdf.

Anonymous:1899:Cbb

[10148] Anonymous. Correspondence. Scientific American, 80(24):393, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n24/pdf/scientificamerican06171899-393.pdf.

Williams:1899:LSM

[10149] Francis H. Williams. A large static machine for X-ray work. Scientific American, 80(24):394–395, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
394.pdf.

Anonymous:1899:CCS

[10150] Anonymous. The Cup Challenger “Shamrock”. Scientific American, 80(24):396, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
396.pdf.

Anonymous:1899:BNG

[10151] Anonymous. The bridges of Niagara Gorge. Scientific American, 80(24):396–397, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
396a.pdf.

Anonymous:1899:BPy

[10152] Anonymous. Business and personal. Scientific American, 80(24):398, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n24/pdf/scientificamerican06171899-398a.pdf.

Anonymous:1899:IIy

[10153] Anonymous. Index of inventions. Scientific American, 80(24):398, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n24/pdf/scientificamerican06171899-398c.pdf.

Anonymous:1899:NQx

[10154] Anonymous. Notes and queries. Scientific American, 80(24):398, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n24/pdf/scientificamerican06171899-398b.pdf.

Anonymous:1899:RPIx

[10155] Anonymous. Recently patented inventions. Scientific American, 80(24):398, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
398.pdf.

Anonymous:1899:WCP

[10156] Anonymous. The World’s coal production. Scientific American, 80(24):399, June 17, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n24/pdf/scientificamerican06171899-
389a.pdf.

Anonymous:1899:PNPw

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n24/pdf/scientificamerican06171899-401.pdf.

Anonymous:1899:CESb

[10158] Anonymous. Comparison of electric and stove heating on street cars. Scientific American, 80(25):404, June 24, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n25/pdf/scientificamerican06241899-404a.pdf.

Anonymous:1899:ISN

[10159] Anonymous. An invention sorely needed. Scientific American, 80(25):404, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
404b.pdf.

Anonymous:1899:PFB

[10160] Anonymous. Petroleum fuel better than coaling stations. Scientific American, 80(25):404, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
404c.pdf.

Anonymous:1899:SYC

[10161] Anonymous. Secrecy in yacht construction. Scientific American, 80(25):404, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
404.pdf.

ScD:1899:LAF

[10162] Henry Morton Ll.D Sc.D. The liquid air fallacy. Scientific American, 80(25):404–405, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
404d.pdf.

Serviss:1899:HJb

[10163] Garrett P. Serviss. The heavens in July. Scientific American, 80(25):405, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
405.pdf.

Anonymous:1899:ABH

[10164] Anonymous. An antiseptic broom for household use. Scientific American, 80(25):406, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
406.pdf.

Anonymous:1899:NTP

[10165] Anonymous. A new two-part bicycle-crank. Scientific American, 80(25):406, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
406b.pdf.

Anonymous:1899:RMb

[10166] Anonymous. The “Reina Mercedes”. Scientific American, 80(25):406, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n25/pdf/scientificamerican06241899-406a.pdf.

Anonymous:1899:SMCa

[10167] Anonymous. Steam motor cars for branch lines. Scientific American, 80(25):407, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
407.pdf.

James:1899:SDMa

[10168] George Wharton James. The snake dance of the Mokis — I. Scientific American, 80(25):408–409, June 24, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v80/
n25/pdf/scientificamerican06241899-408.pdf.

Anonymous:1899:LCY

[10169] Anonymous. The lines and construction of the yacht “Columbia”. Scientific American, 80(25):410–411, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
410.pdf.

Holder:1899:SLR

[10170] C. F. Holder. Sea lion rookeries of Southern California. Scientific American, 80(25):411–412, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
411.pdf.

Anonymous:1899:RPIy

[10171] Anonymous. Recently patented inventions. Scientific American, 80(25):412–413, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v80/n25/pdf/scientificamerican06241899-
412.pdf.

Anonymous:1899:BNBt

[10172] Anonymous. Books: New books, etc. Scientific American, 80(25):413, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n25/pdf/scientificamerican06241899-413b.pdf.

Anonymous:1899:BPz

[10173] Anonymous. Business and personal. Scientific American, 80(25):413, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n25/pdf/scientificamerican06241899-413.pdf.

Anonymous:1899:NQy

[10174] Anonymous. Notes and queries. Scientific American, 80(25):413, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n25/pdf/scientificamerican06241899-413a.pdf.

Anonymous:1899:IIz

[10175] Anonymous. Index of inventions. Scientific American, 80(25):413–414, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n25/pdf/scientificamerican06241899-413c.pdf.

Anonymous:1899:In

[10176] Anonymous. Index. Scientific American, 80(25):415–416, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n25/pdf/scientificamerican06241899-415.pdf.

Anonymous:1899:PNPx

[10177] Anonymous. Patent notes: Patents. Scientific American, 80(25):417, June 24, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v80/n25/pdf/scientificamerican06241899-417.pdf.

Anonymous:1899:FA

[10178] Anonymous. A false alarm. Scientific American, 81(1):4, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-4b.pdf.

Anonymous:1899:PQ

[10179] Anonymous. A pertinent question. Scientific American, 81(1):4, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-4.pdf.

Anonymous:1899:SAR

[10180] Anonymous. Speed and air resistance in cycling. Scientific American, 81(1):4, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
4a.pdf.

Anonymous:1899:SSd

[10181] Anonymous. Science and spiritualism. Scientific American, 81(1):4, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-4c.pdf.

Anonymous:1899:TN

[10182] Anonymous. The transplantation of nerves. Scientific American, 81(1):4–5, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
4d.pdf.

Anonymous:1899:APB

[10183] Anonymous. Armor plate for bank vaults. Scientific American, 81(1): 5, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-5d.pdf.

Anonymous:1899:MAA

[10184] Anonymous. Meeting of the American Association for the Advancement of Science. Scientific American, 81(1):5, July 1, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n1/pdf/scientificamerican07011899-5c.pdf.

Anonymous:1899:MTP

[10185] Anonymous. Method of transferring photographic films. Scientific American, 81(1):5, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
5a.pdf.

Anonymous:1899:PCSc

[10186] Anonymous. Prevention of collisions at sea. Scientific American, 81(1): 5, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-5.pdf.

Carden:1899:WTB

[10187] Lieut. G. L. Carden. Water-tube boilers for our warships. Scientific American, 81(1):5, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
5b.pdf.

Anonymous:1899:ABCb

[10188] Anonymous. Automobile Brougham for a city physician. Scientific American, 81(1):6, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
6a.pdf.

Anonymous:1899:IBG

[10189] Anonymous. An ingenious bicycle gear. Scientific American, 81(1):6, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-6b.pdf.

Anonymous:1899:IHG

[10190] Anonymous. An improved hanger for grindstones. Scientific American, 81(1):6, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
6.pdf.

Anonymous:1899:MTB

[10191] Anonymous. The Mossberg tire-bell. Scientific American, 81(1):7, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-7.pdf.

Gratacap:1899:M

[10192] L. P. Gratacap. The mosasaurs. Scientific American, 81(1):7, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-7a.pdf.

Hale:1899:TCB

[10193] W. H. Hale. The theory and construction of ball bearings. Scientific American, 81(1):8–9, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
8.pdf.

Anonymous:1899:TPM

[10194] Anonymous. Typical perpetual motion fraud. Scientific American, 81(1):9, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
9.pdf.

Anonymous:1899:PER

[10195] Anonymous. The proposed East River Tunnel. Scientific American, 81(1):10–11, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n1/pdf/scientificamerican07011899-
10.pdf.

Anonymous:1899:DLF

[10196] Anonymous. “Deutschland” — the latest and fastest of the transatlantic liners. Scientific American, 81(1):11–12, July 1, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n1/pdf/scientificamerican07011899-11.pdf.

Anonymous:1899:RPIz

[10197] Anonymous. Recently patented inventions. Scientific American, 81(1): 12, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-12.pdf.

Anonymous:1899:BNBu

[10198] Anonymous. Books: New books, etc. Scientific American, 81(1):13, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-13b.pdf.

Anonymous:1899:BPba

[10199] Anonymous. Business and personal. Scientific American, 81(1):13, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-13.pdf.

Anonymous:1899:NQz

[10200] Anonymous. Notes and queries. Scientific American, 81(1):13, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-13a.pdf.

Anonymous:1899:IIba

[10201] Anonymous. Index of inventions. Scientific American, 81(1):13–15, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-13c.pdf.

Anonymous:1899:PNPy

[10202] Anonymous. Patent notes: Patents. Scientific American, 81(1):15, July 1, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n1/pdf/scientificamerican07011899-15.pdf.

Anonymous:1899:AOR

[10203] Anonymous. American ordnance for Russia. Scientific American, 81(2): 18, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-18a.pdf.

Anonymous:1899:CNP

[10204] Anonymous. The Canadian Niagara Power Company. Scientific American, 81(2):18, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n2/pdf/scientificamerican07081899-
18c.pdf.

Anonymous:1899:JIP

[10205] Anonymous. Japan and the International Patent Union. Scientific American, 81(2):18, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n2/pdf/scientificamerican07081899-
18.pdf.

Anonymous:1899:LF

[10206] Anonymous. The largest flag. Scientific American, 81(2):18, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-18g.pdf.

Anonymous:1899:PCb

[10207] Anonymous. The passing of the centerboard. Scientific American, 81(2): 18, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-18e.pdf.

Anonymous:1899:QTS

[10208] Anonymous. Questionable tunnel schemes. Scientific American, 81(2): 18, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-18d.pdf.

Anonymous:1899:RMW

[10209] Anonymous. Railway mileage of the world. Scientific American, 81(2): 18, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-18f.pdf.

Anonymous:1899:TRSb

[10210] Anonymous. The third rail system on the New York elevated railroads. Scientific American, 81(2):18, July 8, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n2/pdf/scientificamerican07081899-18b.pdf.

Anonymous:1899:AI

[10211] Anonymous. The automobile industry. Scientific American, 81(2):19, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-19a.pdf.

Anonymous:1899:MDP

[10212] Anonymous. A million dollars for a patent. Scientific American, 81(2): 19, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-19d.pdf.

Anonymous:1899:SFC

[10213] Anonymous. Smyrna figs in California. Scientific American, 81(2):19, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-19b.pdf.

Anonymous:1899:WFC

[10214] Anonymous. Work of the fish commission. Scientific American, 81(2): 19, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-19c.pdf.

Murray-Aaron:1899:HB

[10215] Eugene Murray-Aaron. Honey-birds. Scientific American, 81(2):19, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-19.pdf.

Anonymous:1899:PFT

[10216] Anonymous. Painting the flagstaffs of tall buildings. Scientific American, 81(2):20, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n2/pdf/scientificamerican07081899-
20.pdf.

Anonymous:1899:NMI

[10217] Anonymous. A new method of instructing and examining trainmen. Scientific American, 81(2):20–21, July 8, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n2/pdf/scientificamerican07081899-20a.pdf.

Hovey:1899:MSE

[10218] E. O. Hovey. Models of some of the extinct lizards of Western North America. Scientific American, 81(2):21–23, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-21.pdf.

Anonymous:1899:CLS

[10219] Anonymous. Complete loss of the steamship “Paris”. Scientific American, 81(2):24, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n2/pdf/scientificamerican07081899-
24.pdf.

Anonymous:1899:USE

[10220] Anonymous. The United States Experimental Model Basin. Scientific American, 81(2):25–26, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n2/pdf/scientificamerican07081899-
25.pdf.

Anonymous:1899:RGS

[10221] Anonymous. A remarkable geyser in Southern California. Scientific American, 81(2):27, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n2/pdf/scientificamerican07081899-
27a.pdf.

Mindeleff:1899:NIG

[10222] Cosmos Mindeleff. Navaho Indian gamblers. Scientific American, 81(2): 27, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-27.pdf.

Anonymous:1899:RPIba

[10223] Anonymous. Recently patented inventions. Scientific American, 81(2): 28, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-28.pdf.

Anonymous:1899:BPbb

[10224] Anonymous. Business and personal. Scientific American, 81(2):29, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-29.pdf.

Anonymous:1899:IIbb

[10225] Anonymous. Index of inventions. Scientific American, 81(2):29, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-29b.pdf.

Anonymous:1899:NQba

[10226] Anonymous. Notes and queries. Scientific American, 81(2):29, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-29a.pdf.

Anonymous:1899:PNPz

[10227] Anonymous. Patent notes: Patents. Scientific American, 81(2):31, July 8, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n2/pdf/scientificamerican07081899-31.pdf.

Anonymous:1899:ERS

[10228] Anonymous. Extreme range of sixteen-inch gun. Scientific American, 81(3):34, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
34a.pdf.

Anonymous:1899:MBR

[10229] Anonymous. Murphy’s bicycle ride a hint to the railroads. Scientific American, 81(3):34, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
34.pdf.

Anonymous:1899:SLE

[10230] Anonymous. Schenectady locomotives for England. Scientific American, 81(3):34, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
34b.pdf.

Adams:1899:EH

[10231] Altan D. Adams. Electric heating. Scientific American, 81(3):35, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-35b.pdf.

Anonymous:1899:CAS

[10232] Anonymous. Corrosion of alloys in sea water. Scientific American, 81(3):35, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
35.pdf.

Anonymous:1899:GSS

[10233] Anonymous. Growth in speed and size of ocean liners. Scientific American, 81(3):35, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
35a.pdf.

Peckham:1899:LLA

[10234] W. C. Peckham. The latest liquid air plant. Scientific American, 81(3):36–38, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
36.pdf.

Anonymous:1899:ADS

[10235] Anonymous. An apparatus for darning stockings. Scientific American, 81(3):38, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
38a.pdf.

Anonymous:1899:ISBb

[10236] Anonymous. An improved sash-balance. Scientific American, 81(3):38, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-38b.pdf.

Anonymous:1899:NTR

[10237] Anonymous. A new type of rotary engine. Scientific American, 81(3):38, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-38.pdf.

Anonymous:1899:OLBb

[10238] Anonymous. Our latest battleship, the “Kearsarge”. Scientific American, 81(3):39, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
39a.pdf.

Anonymous:1899:RCC

[10239] Anonymous. The reconstructed cruiser “Chicago”. Scientific American, 81(3):39, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
39.pdf.

Anonymous:1899:LBD

[10240] Anonymous. Life buoy designed by Rear-Admiral Hichborn. Scientific American, 81(3):40, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
40.pdf.

Anonymous:1899:MLT

[10241] Anonymous. A mile in less than a minute on a bicycle. Scientific American, 81(3):41–42, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
41.pdf.

Anonymous:1899:UAS

[10242] Anonymous. Universal automatic screw machine. Scientific American, 81(3):43, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
43.pdf.

Anonymous:1899:BNBv

[10243] Anonymous. Books: New books, etc. Scientific American, 81(3):44, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-44a.pdf.

Anonymous:1899:RPIbb

[10244] Anonymous. Recently patented inventions. Scientific American, 81(3):44, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n3/pdf/scientificamerican07151899-
44.pdf.

Anonymous:1899:BPbc

[10245] Anonymous. Business and personal. Scientific American, 81(3):45, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-45.pdf.

Anonymous:1899:Io

[10246] Anonymous. To inventors. Scientific American, 81(3):45, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-45b.pdf.

Anonymous:1899:IIbc

[10247] Anonymous. Index of inventions. Scientific American, 81(3):45, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-45c.pdf.

Anonymous:1899:NQbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-45a.pdf.

Anonymous:1899:PNPba

[10249] Anonymous. Patent notes: Patents. Scientific American, 81(3):47, July 15, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n3/pdf/scientificamerican07151899-47.pdf.

Anonymous:1899:COC

[10250] Anonymous. Comparative operating cost of horse and electric delivery wagons. Scientific American, 81(4):50, July 22, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n4/pdf/scientificamerican07221899-50a.pdf.

Anonymous:1899:FP

[10251] Anonymous. Floating of the “Paris”. Scientific American, 81(4):50, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-50.pdf.

Anonymous:1899:SSCa

[10252] Anonymous. Spain sells cruisers. Scientific American, 81(4):50, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-50c.pdf.

Anonymous:1899:WOR

[10253] Anonymous. The wage of original research. Scientific American, 81(4):50, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
50b.pdf.

Anonymous:1899:BGB

[10254] Anonymous. Brooms as germ breeders. Scientific American, 81(4):51, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-51b.pdf.

Anonymous:1899:TTP

[10255] Anonymous. Time of transatlantic passage again reduced. Scientific American, 81(4):51, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
51a.pdf.

Weir:1899:ATU

[10256] James Weir, Jr. Animal tool-users, engineers, and mechanics. Scientific American, 81(4):51, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
51.pdf.

Anonymous:1899:EEC

[10257] Anonymous. The Eastman “Electro-Cycle”. Scientific American, 81(4):52, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
52a.pdf.

Anonymous:1899:SMCb

[10258] Anonymous. A simple mold for composition rollers. Scientific American, 81(4):52, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
52.pdf.

Anonymous:1899:STCa

[10259] Anonymous. A simple tool for cutting pipes. Scientific American, 81(4):52, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
52b.pdf.

Anonymous:1899:MTT

[10260] Anonymous. Moving a 500-ton train-shed arch. Scientific American, 81(4):53–54, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
53.pdf.

Anonymous:1899:Cbc

[10261] Anonymous. Correspondence. Scientific American, 81(4):55, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-55.pdf.

Anonymous:1899:NFP

[10262] Anonymous. Niagara Falls Power Plant. Scientific American, 81(4):56–57, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
56.pdf.

Anonymous:1899:AV

[10263] Anonymous. Alessandro Volta. Scientific American, 81(4):57–58, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-57.pdf.

Holder:1899:CWF

[10264] C. F. Holder. A Catalina walking fish. Scientific American, 81(4):59, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-59.pdf.

Anonymous:1899:BNBw

[10265] Anonymous. Books: New books, etc. Scientific American, 81(4):60, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-60a.pdf.

Anonymous:1899:RPIbc

[10266] Anonymous. Recently patented inventions. Scientific American, 81(4):60, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n4/pdf/scientificamerican07221899-
60.pdf.

Anonymous:1899:BPbd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-61.pdf.

Anonymous:1899:Iq

[10268] Anonymous. To inventors. Scientific American, 81(4):61, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-61b.pdf.

Anonymous:1899:NQbc

[10269] Anonymous. Notes and queries. Scientific American, 81(4):61, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-61a.pdf.

Anonymous:1899:IIbd

[10270] Anonymous. Index of inventions. Scientific American, 81(4):61–63, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-61c.pdf.

Anonymous:1899:PNPbb

[10271] Anonymous. Patent notes: Patents. Scientific American, 81(4):63, July 22, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n4/pdf/scientificamerican07221899-63.pdf.

Anonymous:1899:ARCb

[10272] Anonymous. Annual report of the Commissioner of Patents. Scientific American, 81(5):66, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
66a.pdf.

Anonymous:1899:ARP

[10273] Anonymous. That air resistance problem. Scientific American, 81(5):66, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-66.pdf.

Anonymous:1899:EWF

[10274] Anonymous. Exportations of wheat flour. Scientific American, 81(5):66, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-66b.pdf.

Serviss:1899:HA

[10275] Garrett P. Serviss. The heavens in August. Scientific American, 81(5):66–67, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
66c.pdf.

Anonymous:1899:RCP

[10276] Anonymous. Report of the Commissioner of Patents for the fiscal year ending June 30, 1899. Scientific American, 81(5):67, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-67.pdf.

Anonymous:1899:NFSb

[10277] Anonymous. New features in the Scientific American. Scientific American, 81(5):67–68, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
67a.pdf.

Anonymous:1899:HOT

[10278] Anonymous. A huge Overland traction engine. Scientific American, 81(5):68, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
68.pdf.

Anonymous:1899:ISPb

[10279] Anonymous. An improved stove-pipe coupling. Scientific American, 81(5):69, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
69b.pdf.

Anonymous:1899:JHE

[10280] Anonymous. The Johns hydraulic engine — a combined water and air motor. Scientific American, 81(5):69, July 29, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n5/pdf/scientificamerican07291899-69a.pdf.

Anonymous:1899:NDR

[10281] Anonymous. A novel device for raising and lowering smokestacks. Scientific American, 81(5):69, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
69.pdf.

Anonymous:1899:Cbd

[10282] Anonymous. Correspondence. Scientific American, 81(5):70, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-70.pdf.

Anonymous:1899:PEM

[10283] Anonymous. The Paris Exposition Medal. Scientific American, 81(5):70–71, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
70a.pdf.

Anonymous:1899:NWf

[10284] Anonymous. Navies of the world. Scientific American, 81(5):72–75, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-72.pdf.

Anonymous:1899:TAV

[10285] Anonymous. A transcontinental automobile vehicle. Scientific American, 81(5):75, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
75.pdf.

Anonymous:1899:IPB

[10286] Anonymous. An Italian Prince bound for the North Pole. Scientific American, 81(5):75–76, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
75a.pdf.

Anonymous:1899:RPIbd

[10287] Anonymous. Recently patented inventions. Scientific American, 81(5):76–77, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n5/pdf/scientificamerican07291899-
76.pdf.

Anonymous:1899:BNBx

[10288] Anonymous. Books: New books, etc. Scientific American, 81(5):77, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-77b.pdf.

Anonymous:1899:BPbe

[10289] Anonymous. Business and personal. Scientific American, 81(5):77, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-77.pdf.

Anonymous:1899:Ir

[10290] Anonymous. To inventors. Scientific American, 81(5):77, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-77c.pdf.

Anonymous:1899:IIbe

[10291] Anonymous. Index of inventions. Scientific American, 81(5):77, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-77d.pdf.

Anonymous:1899:NQbd

[10292] Anonymous. Notes and queries. Scientific American, 81(5):77, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-77a.pdf.

Anonymous:1899:PNPbc

[10293] Anonymous. Patent notes: Patents. Scientific American, 81(5):79, July 29, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n5/pdf/scientificamerican07291899-79.pdf.

Anonymous:1899:ALI

[10294] Anonymous. Average life of an iron or steel bridge. Scientific American, 81(6):82, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
82.pdf.

Anonymous:1899:AMa

[10295] Anonymous. Automobile motors. Scientific American, 81(6):82, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-82a.pdf.

Anonymous:1899:ARCc

[10296] Anonymous. An American railway in China. Scientific American, 81(6):82, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
82b.pdf.

Anonymous:1899:CNa

[10297] Anonymous. A commercial need. Scientific American, 81(6):82, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-82c.pdf.

Anonymous:1899:FHE

[10298] Anonymous. Firing high explosives. Scientific American, 81(6):82–83, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-82d.pdf.

Anonymous:1899:AAC

[10299] Anonymous. Activity of American consuls. Scientific American, 81(6):83, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
83b.pdf.

Anonymous:1899:BPbf

[10300] Anonymous. British patents in 1898. Scientific American, 81(6):83, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-83.pdf.

Anonymous:1899:LAM

[10301] Anonymous. Liquid air in medicine and surgery. Scientific American, 81(6):83, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
83a.pdf.

Anonymous:1899:PLSa

[10302] Anonymous. A prize for a life-saving appliance. Scientific American, 81(6):83, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
83d.pdf.

Anonymous:1899:SDN

[10303] Anonymous. Steel dredges for New York Harbor. Scientific American, 81(6):83, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
83c.pdf.

Anonymous:1899:IFS

[10304] Anonymous. An improved flue-stopper. Scientific American, 81(6):84, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-84.pdf.

Anonymous:1899:ITW

[10305] Anonymous. An improved turbine water-wheel. Scientific American, 81(6):84, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
84a.pdf.

Anonymous:1899:NWT

[10306] Anonymous. A new wagon-tongue support. Scientific American, 81(6):84, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
84c.pdf.

Sanders:1899:WDE

[10307] H. P. Sanders. Wurdemann’s dividing engine. Scientific American, 81(6):84, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
84b.pdf.

Anonymous:1899:CEH

[10308] Anonymous. Conservation of energy in the human body. Scientific American, 81(6):85–86, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
85.pdf.

Anonymous:1899:Cbe

[10309] Anonymous. Correspondence. Scientific American, 81(6):86, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-86.pdf.

Smith:1899:PBI

[10310] Hugh M. Smith. The pearl-button industry of the Mississippi River. Scientific American, 81(6):86–87, August 5, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n6/pdf/scientificamerican08051899-86a.pdf.

Anonymous:1899:SVF

[10311] Anonymous. The submarine vessel “La France”. Scientific American, 81(6):88, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
88.pdf.

Anonymous:1899:TPS

[10312] Anonymous. Three-pounder semi-automatic gun. Scientific American, 81(6):88, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
88a.pdf.

Anonymous:1899:SIP

[10313] Anonymous. A subterranean industrial plant. Scientific American, 81(6):89–90, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
89.pdf.

Anonymous:1899:EHC

[10314] Anonymous. Electric heating in a Carmelite Hospice. Scientific American, 81(6):91–92, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
91.pdf.

Anonymous:1899:RPIbe

[10315] Anonymous. Recently patented inventions. Scientific American, 81(6):92–93, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n6/pdf/scientificamerican08051899-
92.pdf.

Anonymous:1899:BPbg

[10316] Anonymous. Business and personal. Scientific American, 81(6):93, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-93a.pdf.

Anonymous:1899:IIbf

[10317] Anonymous. Index of inventions. Scientific American, 81(6):93, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-93c.pdf.

Anonymous:1899:NBE

[10318] Anonymous. New books, etc. Scientific American, 81(6):93, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-93.pdf.

Anonymous:1899:NQbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-93b.pdf.

Anonymous:1899:PNPbd

[10320] Anonymous. Patent notes: Patents. Scientific American, 81(6):95, August 5, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n6/pdf/scientificamerican08051899-95.pdf.

Anonymous:1899:FSY

[10321] Anonymous. The factor of safety in yacht construction. Scientific American, 81(7):98, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
98.pdf.

Anonymous:1899:NPE

[10322] Anonymous. A National Park in the East. Scientific American, 81(7):98, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
98b.pdf.

Anonymous:1899:OFT

[10323] Anonymous. Our foreign trade for the last fiscal year. Scientific American, 81(7):98, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
98a.pdf.

Anonymous:1899:PET

[10324] Anonymous. Postage and the export trade. Scientific American, 81(7):98, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
98c.pdf.

Anonymous:1899:PSa

[10325] Anonymous. The production of slate. Scientific American, 81 (7):98–99, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
98d.pdf.

Anonymous:1899:IOT

[10326] Anonymous. Improvements in our tobacco. Scientific American, 81(7):99, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
99a.pdf.

Anonymous:1899:SNb

[10327] Anonymous. Science notes. Scientific American, 81(7):99, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-99b.pdf.

Anonymous:1899:SSW

[10328] Anonymous. Some soda water fountain statistics. Scientific American, 81(7):99, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
99.pdf.

Anonymous:1899:CLD

[10329] Anonymous. The crankshafts of the liner “Deutschland”. Scientific American, 81(7):100, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
100a.pdf.

Anonymous:1899:NMDa

[10330] Anonymous. A new Miner’s drill. Scientific American, 81(7):100, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-100.pdf.

Anonymous:1899:PAP

[10331] Anonymous. The “Polyphone” — attachment for phonographs. Scientific American, 81(7):100, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
100b.pdf.

Anonymous:1899:BK

[10332] Anonymous. The boxing kangaroo. Scientific American, 81(7):101, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-101.pdf.

Anonymous:1899:EBb

[10333] Anonymous. An electric Brougham. Scientific American, 81(7):101, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-101a.pdf.

Anonymous:1899:Cbf

[10334] Anonymous. Correspondence. Scientific American, 81(7):102–103, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-102.pdf.

Jenkins:1899:FG

[10335] C. Francis Jenkins. Fuel gas. Scientific American, 81(7):104–105, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-104.pdf.

Anonymous:1899:TDC

[10336] Anonymous. Tree dwelling in California. Scientific American, 81 (7):105–106, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
105.pdf.

Anonymous:1899:CSC

[10337] Anonymous. “Columbia” and “Shamrock” — a comparison. Scientific American, 81(7):106–108, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
106.pdf.

Anonymous:1899:BPbh

[10338] Anonymous. Business and personal. Scientific American, 81(7):109, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-109a.pdf.

Anonymous:1899:IIbg

[10339] Anonymous. Index of inventions. Scientific American, 81(7):109, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-109c.pdf.

Anonymous:1899:NQbf

[10340] Anonymous. Notes and queries. Scientific American, 81(7):109, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-109b.pdf.

Anonymous:1899:RPIbf

[10341] Anonymous. Recently patented inventions. Scientific American, 81(7):109, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n7/pdf/scientificamerican08121899-
109.pdf.

Anonymous:1899:PNPbe

[10342] Anonymous. Patent notes: Patents. Scientific American, 81(7):111, August 12, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n7/pdf/scientificamerican08121899-111.pdf.

Anonymous:1899:EAL

[10343] Anonymous. English and American locomotives. Scientific American, 81(8):114, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
114a.pdf.

Anonymous:1899:OSP

[10344] Anonymous. Our shipbuilders and the proposed new cruisers. Scientific American, 81(8):114, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
114.pdf.

Anonymous:1899:EGC

[10345] Anonymous. The education of German consuls. Scientific American, 81(8):114–115, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
114b.pdf.

Anonymous:1899:Pd

[10346] Anonymous. Pawnbroking. Scientific American, 81(8):115, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-115.pdf.

Adams:1899:SIP

[10347] Altan D. Adams. Saving in isolated plants. Scientific American, 81(8):115–116, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
115a.pdf.

Anonymous:1899:NAGa

[10348] Anonymous. A new acetylene gas burner. Scientific American, 81(8):116, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
116.pdf.

Anonymous:1899:PLF

[10349] Anonymous. Peculiar lightning flashes. Scientific American, 81 (8):116, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
116a.pdf.

Anonymous:1899:STCb

[10350] Anonymous. The Stratford Trolley Car Disaster. Scientific American, 81(8):117, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
117.pdf.

Anonymous:1899:Cbg

[10351] Anonymous. Correspondence. Scientific American, 81(8):118, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-118.pdf.

Anonymous:1899:FFO

[10352] Anonymous. Faulty features in our proposed sheathed and coppered cruisers. Scientific American, 81(8):118–119, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-118a.pdf.

Stewart:1899:ETG

[10353] J. A. Stewart. Experimental tea growing in the United States. Scientific American, 81(8):120–121, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
120.pdf.

Anonymous:1899:RAD

[10354] Anonymous. The return of Admiral Dewey. Scientific American, 81(8):121–122, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
121.pdf.

Anonymous:1899:EMI

[10355] Anonymous. The English Mark IV. Cordite ammunition. Scientific American, 81(8):122, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
122.pdf.

Benjamin:1899:EO

[10356] Marcis Benjamin. Edward Orton. Scientific American, 81(8): 123–124, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
123.pdf.

Anonymous:1899:RPIbg

[10357] Anonymous. Recently patented inventions. Scientific American, 81 (8):124–125, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
124.pdf.

Anonymous:1899:BNBy

[10358] Anonymous. Books: New books, etc. Scientific American, 81 (8):125, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n8/pdf/scientificamerican08191899-
125.pdf.

Anonymous:1899:BPbi

[10359] Anonymous. Business and personal. Scientific American, 81(8):125, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-125a.pdf.

Anonymous:1899:Is

[10360] Anonymous. To inventors. Scientific American, 81(8):125, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-125c.pdf.

Anonymous:1899:IIbh

[10361] Anonymous. Index of inventions. Scientific American, 81(8):125, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-125d.pdf.

Anonymous:1899:NQbg

[10362] Anonymous. Notes and queries. Scientific American, 81(8):125, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-125b.pdf.

Anonymous:1899:PNPbf

[10363] Anonymous. Patent notes: Patents. Scientific American, 81(8):127, August 19, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n8/pdf/scientificamerican08191899-127.pdf.

Anonymous:1899:ASS

[10364] Anonymous. Accuracy and style in scientific writing. Scientific American, 81(9):130, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
130c.pdf.

Anonymous:1899:NO

[10365] Anonymous. A national opportunity. Scientific American, 81 (9):130, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
130b.pdf.

Anonymous:1899:PNCc

[10366] Anonymous. The problem of the new cruisers. Scientific American, 81(9):130, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
130.pdf.

Anonymous:1899:STDa

[10367] Anonymous. The Stratford Trolley Disaster. Scientific American, 81(9):130, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
130a.pdf.

Anonymous:1899:CSLb

[10368] Anonymous. “Columbia” and “Shamrock” in light weather. Scientific American, 81(9):130–131, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
130d.pdf.

Anonymous:1899:DDD

[10369] Anonymous. Death of Dr. D. G. Brinton. Scientific American, 81(9):131, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
131d.pdf.

Anonymous:1899:DPBb

[10370] Anonymous. Death of Prof. Bunsen. Scientific American, 81(9):131, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-131b.pdf.

Anonymous:1899:JPB

[10371] Anonymous. Japan purchases bridges. Scientific American, 81 (9):131, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
131e.pdf.

Anonymous:1899:JR

[10372] Anonymous. James Richardson. Scientific American, 81(9):131, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-131c.pdf.

Anonymous:1899:WWT

[10373] Anonymous. The Waltham Watch trade-mark. Scientific American, 81(9):131, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
131a.pdf.

Serviss:1899:HS

[10374] Garrett P. Serviss. The heavens in September. Scientific American, 81(9):131, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
131.pdf.

Fawcett:1899:SPS

[10375] Waldon Fawcett. Self-propelling steel canal boats. Scientific American, 81(9):132, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
132.pdf.

Anonymous:1899:MBCb

[10376] Anonymous. A Mexican bean clock. Scientific American, 81(9):133, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-133.pdf.

Anonymous:1899:NFE

[10377] Anonymous. A novel fire-escape. Scientific American, 81(9):133, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-133a.pdf.

Anonymous:1899:Cbh

[10378] Anonymous. Correspondence. Scientific American, 81(9):134–135, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-134.pdf.

Michelsen:1899:TSR

[10379] Henry Michelsen. The Trans-Siberian Railroad. Scientific American, 81(9):136–138, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
136.pdf.

Anonymous:1899:TA

[10380] Anonymous. A token of appreciation. Scientific American, 81 (9):138, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
138.pdf.

Anonymous:1899:BTW

[10381] Anonymous. A Barrow with two wheels. Scientific American, 81(9):139, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
139.pdf.

Anonymous:1899:NEL

[10382] Anonymous. New express locomotive for the Pennsylvania Railroad. Scientific American, 81(9):139–140, August 26, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n9/pdf/scientificamerican08261899-139a.pdf.

Anonymous:1899:BPbj

[10383] Anonymous. Business and personal. Scientific American, 81(9):141, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-141a.pdf.

Anonymous:1899:NQbh

[10384] Anonymous. Notes and queries. Scientific American, 81(9):141, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-141b.pdf.

Anonymous:1899:RPIbh

[10385] Anonymous. Recently patented inventions. Scientific American, 81(9):141, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
141.pdf.

Anonymous:1899:IIbi

[10386] Anonymous. Index of inventions. Scientific American, 81(9): 141–143, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n9/pdf/scientificamerican08261899-
141c.pdf.

Anonymous:1899:PNPbg

[10387] Anonymous. Patent notes: Patents. Scientific American, 81(9):143, August 26, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n9/pdf/scientificamerican08261899-143.pdf.

Anonymous:1899:NSBb

[10388] Anonymous. Nickel steel in boiler construction. Scientific American, 81(10):146, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
146.pdf.

Anonymous:1899:SOV

[10389] Anonymous. The statistics of our vast railroad system. Scientific American, 81(10):146, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
146b.pdf.

Anonymous:1899:TCG

[10390] Anonymous. Trolley cars on grades. Scientific American, 81 (10):146, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
146a.pdf.

Anonymous:1899:WMM

[10391] Anonymous. Weight of mail matter in the United States. Scientific American, 81(10):146–147, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
146c.pdf.

Anonymous:1899:AOO

[10392] Anonymous. The Allegheny Observatory objective. Scientific American, 81(10):147, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
147b.pdf.

Anonymous:1899:CCM

[10393] Anonymous. Copper coins melted up. Scientific American, 81 (10):147, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
147a.pdf.

Hovey:1899:AAA

[10394] Horace C. Hovey. The American Association for the Advancement of Science. Scientific American, 81(10):147, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n10/pdf/scientificamerican09021899-147.pdf.

Anonymous:1899:IMM

[10395] Anonymous. An ingenious mowing-machine-sickle grinder. Scientific American, 81(10):148, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
148b.pdf.

Anonymous:1899:IREa

[10396] Anonymous. An improvement in rotary engines. Scientific American, 81(10):148, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
148.pdf.

Anonymous:1899:NAGb

[10397] Anonymous. A new acetylene gas generator. Scientific American, 81(10):148, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
148a.pdf.

Anonymous:1899:PPC

[10398] Anonymous. Present and proposed cruisers of the United States Navy compared. Scientific American, 81(10):149–150, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n10/pdf/scientificamerican09021899-149.pdf.

Anonymous:1899:IMN

[10399] Anonymous. Ice manufacture on a new system. Scientific American, 81(10):151, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
151.pdf.

Anonymous:1899:RFE

[10400] Anonymous. Removal of a famous engineering landmark. Scientific American, 81(10):152–153, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
152.pdf.

Brown:1899:LEM

[10401] Enos Brown. The last eruption of Mauna Loa. Scientific American, 81(10):153–154, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
153.pdf.

Anonymous:1899:AEV

[10402] Anonymous. Ancient Egyptian versus modern pin locks. Scientific American, 81(10):155, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
155.pdf.

Anonymous:1899:BNBz

[10403] Anonymous. Books: New books, etc. Scientific American, 81 (10):156, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
156a.pdf.

Anonymous:1899:BPbk

[10404] Anonymous. Business and personal. Scientific American, 81 (10):156, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
156b.pdf.

Anonymous:1899:RPIbi

[10405] Anonymous. Recently patented inventions. Scientific American, 81(10):156, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
156.pdf.

Anonymous:1899:NQbi

[10406] Anonymous. Notes and queries. Scientific American, 81(10): 156–157, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
156c.pdf.

Anonymous:1899:It

[10407] Anonymous. To inventors. Scientific American, 81(10):157, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n10/pdf/scientificamerican09021899-157.pdf.

Anonymous:1899:IIbj

[10408] Anonymous. Index of inventions. Scientific American, 81(10): 157, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
157a.pdf.

Anonymous:1899:PNPbh

[10409] Anonymous. Patent notes: Patents. Scientific American, 81 (10):159, September 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n10/pdf/scientificamerican09021899-
159.pdf.

Anonymous:1899:EEU

[10410] Anonymous. Economy of the electric underground trolley. Scientific American, 81(11):162, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162a.pdf.

Anonymous:1899:FCC

[10411] Anonymous. Fall of the Chicago Coliseum. Scientific American, 81(11):162, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162c.pdf.

Anonymous:1899:GCRb

[10412] Anonymous. Generous criticism — and the reverse. Scientific American, 81(11):162, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162.pdf.

Anonymous:1899:PDP

[10413] Anonymous. The prolonged distress in porto rico. Scientific American, 81(11):162, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162d.pdf.

Anonymous:1899:PSTa

[10414] Anonymous. The preliminary speed trial of the “Alabama”. Scientific American, 81(11):162, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162e.pdf.

Anonymous:1899:SET

[10415] Anonymous. In search of an effective trolley car brake. Scientific American, 81(11):162, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162b.pdf.

Anonymous:1899:VW

[10416] Anonymous. Velocity of the wind. Scientific American, 81(11): 162, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162f.pdf.

Anonymous:1899:NPPb

[10417] Anonymous. The new parcels-post to Germany. Scientific American, 81(11):162–163, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
162g.pdf.

Anonymous:1899:SNd

[10418] Anonymous. Science notes. Scientific American, 81(11):163, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n11/pdf/scientificamerican09091899-163a.pdf.

Hovey:1899:SCCa

[10419] Horace C. Hovey. Scientific Congress at Columbus. Scientific American, 81(11):163, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
163.pdf.

Anonymous:1899:AWLb

[10420] Anonymous. The accurate weighing of locomotives. Scientific American, 81(11):164, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
164a.pdf.

Anonymous:1899:SCTb

[10421] Anonymous. A simple chimney top and cowl. Scientific American, 81(11):164, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
164b.pdf.

Anonymous:1899:SMCc

[10422] Anonymous. Simple means of cooling drinking water. Scientific American, 81(11):164, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
164.pdf.

Anonymous:1899:TSM

[10423] Anonymous. A time-stamp for mercantile use. Scientific American, 81(11):165, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
165.pdf.

Lydekker:1899:CN

[10424] R. Lydekker. A contrast in noses. Scientific American, 81(11): 165–166, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
165a.pdf.

James:1899:SDMb

[10425] G. Wharton James. The snake dance of the Mokis — II. Scientific American, 81(11):167, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
167.pdf.

Anonymous:1899:MHL

[10426] Anonymous. The Mungsten high-level arch bridge. Scientific American, 81(11):168–169, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
168.pdf.

Anonymous:1899:PAO

[10427] Anonymous. Proposed armament for our three latest battleships. Scientific American, 81(11):169–170, September 9, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n11/pdf/scientificamerican09091899-169.pdf.

Anonymous:1899:NEEa

[10428] Anonymous. A novel electrical effect. Scientific American, 81 (11):171, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
171a.pdf.

Anonymous:1899:PSb

[10429] Anonymous. Photography of the stomach. Scientific American, 81(11):171, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
171.pdf.

Anonymous:1899:Cbi

[10430] Anonymous. Correspondence. Scientific American, 81(11):172, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n11/pdf/scientificamerican09091899-172.pdf.

Anonymous:1899:RPIbj

[10431] Anonymous. Recently patented inventions. Scientific American, 81 (11):172–173, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
172a.pdf.

Anonymous:1899:BNBba

[10432] Anonymous. Books: New books, etc. Scientific American, 81 (11):173, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
173.pdf.

Anonymous:1899:BPbl

[10433] Anonymous. Business and personal. Scientific American, 81 (11):173, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
173a.pdf.

Anonymous:1899:IIbk

[10434] Anonymous. Index of inventions. Scientific American, 81(11): 173, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
173c.pdf.

Anonymous:1899:NQbj

[10435] Anonymous. Notes and queries. Scientific American, 81(11): 173, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
173b.pdf.

Anonymous:1899:SAC

[10436] Anonymous. The Scientific American cyclopedia of receipts, notes & queries. Scientific American, 81(11):174, September 9, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n11/pdf/scientificamerican09091899-174.pdf.

Anonymous:1899:PNPbi

[10437] Anonymous. Patent notes: Patents. Scientific American, 81 (11):175, September 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n11/pdf/scientificamerican09091899-
175.pdf.

Anonymous:1899:ASd

[10438] Anonymous. Artificial silk. Scientific American, 81(12):178, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n12/pdf/scientificamerican09161899-178e.pdf.

Anonymous:1899:OFB

[10439] Anonymous. Our fastest battleship. Scientific American, 81(12): 178, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
178a.pdf.

Anonymous:1899:RTG

[10440] Anonymous. Rapid transit by the grace of the politicians. Scientific American, 81(12):178, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
178.pdf.

Anonymous:1899:SHC

[10441] Anonymous. The “Shamrock” and her chances. Scientific American, 81(12):178, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
178d.pdf.

Anonymous:1899:TON

[10442] Anonymous. Trade with our newly acquired territories. Scientific American, 81(12):178, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
178c.pdf.

Anonymous:1899:WIC

[10443] Anonymous. Work of the industrial commission. Scientific American, 81(12):178, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
178b.pdf.

Anonymous:1899:ANa

[10444] Anonymous. Automobile news. Scientific American, 81(12): 179, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
179b.pdf.

Hovey:1899:SCCb

[10445] Horack C. Hovey. Scientific Congress at Columbus. Scientific American, 81(12):179, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
179a.pdf.

MGahie:1899:LAE

[10446] Frederick H. M’Gahie. Liquid air as an explosive. Scientific American, 81(12):179, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
179.pdf.

Anonymous:1899:JFF

[10447] Anonymous. Jig-file for finishing dies. Scientific American, 81 (12):180, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
180.pdf.

Holder:1899:LO

[10448] C. F. Holder. A large octopus. Scientific American, 81(12):180, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n12/pdf/scientificamerican09161899-180a.pdf.

Anonymous:1899:MOC

[10449] Anonymous. A mechanically-operated cross-cut saw. Scientific American, 81(12):180–181, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
180b.pdf.

Anonymous:1899:IREb

[10450] Anonymous. Improvements in rotary engines. Scientific American, 81 (12):181–182, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
181.pdf.

Anonymous:1899:Cbj

[10451] Anonymous. Correspondence. Scientific American, 81(12):182– 183, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
182.pdf.

Anonymous:1899:CATc

[10452] Anonymous. Compressed air traction in New York City. Scientific American, 81(12):184–185, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
184.pdf.

Anonymous:1899:FKC

[10453] Anonymous. Fin keel centerboard for one-rater. Scientific American, 81(12):186, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
186.pdf.

Anonymous:1899:DA

[10454] Anonymous. The Dewey arch. Scientific American, 81(12):187– 188, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
187.pdf.

Anonymous:1899:RPIbk

[10455] Anonymous. Recently patented inventions. Scientific American, 81 (12):188–189, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
188.pdf.

Anonymous:1899:BNBbb

[10456] Anonymous. Books: New books, etc. Scientific American, 81 (12):189, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
189.pdf.

Anonymous:1899:BPbm

[10457] Anonymous. Business and personal. Scientific American, 81 (12):189, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
189a.pdf.

Anonymous:1899:IIbl

[10458] Anonymous. Index of inventions. Scientific American, 81(12): 189, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
189c.pdf.

Anonymous:1899:NQbk

[10459] Anonymous. Notes and queries. Scientific American, 81(12): 189, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
189b.pdf.

Anonymous:1899:PNPbj

[10460] Anonymous. Patent notes: Patents. Scientific American, 81 (12):191, September 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n12/pdf/scientificamerican09161899-
191.pdf.

Anonymous:1899:PPF

[10461] Anonymous. Plant and process for fireproofing wood. Scientific American, 81(13):193, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
193.pdf.

Anonymous:1899:DAB

[10462] Anonymous. Danger in the approaches to the Brooklyn Navy Yard. Scientific American, 81(13):194, September 23, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n13/pdf/scientificamerican09231899-194c.pdf.

Anonymous:1899:FPC

[10463] Anonymous. Financial problems of Cuba. Scientific American, 81 (13):194, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
194a.pdf.

Anonymous:1899:MSTa

[10464] Anonymous. Money spent by travelers. Scientific American, 81 (13):194, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
194b.pdf.

Anonymous:1899:NWU

[10465] Anonymous. New ways of using the big corn crop. Scientific American, 81(13):194, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
194d.pdf.

Anonymous:1899:Ob

[10466] Anonymous. The “Oceanic”. Scientific American, 81(13):194, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n13/pdf/scientificamerican09231899-194.pdf.

Anonymous:1899:WTR

[10467] Anonymous. Wireless telegraphy to report the yacht races. Scientific American, 81(13):194–195, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
194e.pdf.

Anonymous:1899:GAFa

[10468] Anonymous. Government aid in forestry. Scientific American, 81 (13):195, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
195c.pdf.

Anonymous:1899:NLP

[10469] Anonymous. News of Lieut. Peary. Scientific American, 81 (13):195, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
195a.pdf.

Anonymous:1899:OP

[10470] Anonymous. Observations of Polaris. Scientific American, 81 (13):195, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
195b.pdf.

Anonymous:1899:SCFb

[10471] Anonymous. Some calendarial facts about the Twentieth Century. Scientific American, 81(13):195, September 23, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n13/pdf/scientificamerican09231899-195.pdf.

Bennett:1899:SFF

[10472] Herbert I. Bennett. A San Francisco fire fighter. Scientific American, 81(13):196, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
196.pdf.

Anonymous:1899:GRB

[10473] Anonymous. The grant roller-bearing. Scientific American, 81 (13):197, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
197.pdf.

Anonymous:1899:IVT

[10474] Anonymous. An improved vehicle-tire. Scientific American, 81 (13):197, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
197a.pdf.

Anonymous:1899:RIR

[10475] Anonymous. A recently invented rotary engine. Scientific American, 81(13):197, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
197b.pdf.

Anonymous:1899:Cbk

[10476] Anonymous. Correspondence. Scientific American, 81(13):198– 199, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
198.pdf.

Anonymous:1899:PPR

[10477] Anonymous. In the path of the Porto Rican Hurricane. Scientific American, 81(13):200–201, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
200.pdf.

Anonymous:1899:GEO

[10478] Anonymous. The “Great Eastern” and the “Oceanic” — a comparison. Scientific American, 81(13):201–202, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n13/pdf/scientificamerican09231899-201.pdf.

Anonymous:1899:NMCb

[10479] Anonymous. A novel menu card. Scientific American, 81(13): 202, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
202.pdf.

Anonymous:1899:SDa

[10480] Anonymous. The “Shamrock” disabled. Scientific American, 81 (13):203–204, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
203.pdf.

Anonymous:1899:BPbn

[10481] Anonymous. Business and personal. Scientific American, 81 (13):205, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
205a.pdf.

Anonymous:1899:IIbm

[10482] Anonymous. Index of inventions. Scientific American, 81(13): 205, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
205c.pdf.

Anonymous:1899:NQbl

[10483] Anonymous. Notes and queries. Scientific American, 81(13): 205, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
205b.pdf.

Anonymous:1899:RPIbl

[10484] Anonymous. Recently patented inventions. Scientific American, 81 (13):205, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
205.pdf.

Anonymous:1899:PNPbk

[10485] Anonymous. Patent notes: Patents. Scientific American, 81 (13):207, September 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n13/pdf/scientificamerican09231899-
207.pdf.

Anonymous:1899:AMb

[10486] Anonymous. The arms and the man. Scientific American, 81 (14):210, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
210.pdf.

Anonymous:1899:EVR

[10487] Anonymous. The Euphrates Valley Railroad. Scientific American, 81(14):210, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
210b.pdf.

Anonymous:1899:MP

[10488] Anonymous. The Mauser pistol. Scientific American, 81(14): 210, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
210c.pdf.

Anonymous:1899:NEEb

[10489] Anonymous. The national export exposition. Scientific American, 81(14):210, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
210a.pdf.

Anonymous:1899:ASM

[10490] Anonymous. Arrival of Signor Marconi. Scientific American, 81 (14):211, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
211c.pdf.

Anonymous:1899:ELC

[10491] Anonymous. Electric lighting at the covent garden theater. Scientific American, 81(14):211, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
211a.pdf.

Anonymous:1899:IMA

[10492] Anonymous. An improvement in the manufacture of arsenate of lead. Scientific American, 81(14):211, September 30, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n14/pdf/scientificamerican09301899-211b.pdf.

Serviss:1899:HO

[10493] Garrett P. Serviss. The heavens in October. Scientific American, 81(14):211, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
211.pdf.

Anonymous:1899:NEEc

[10494] Anonymous. The National Export Exposition, Philadelphia. Scientific American, 81(14):212–213, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
212.pdf.

Anonymous:1899:CEM

[10495] Anonymous. The Chapman electro-mobile. Scientific American, 81 (14):213, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
213a.pdf.

Anonymous:1899:RBB

[10496] Anonymous. A roller-bearing for bicycles. Scientific American, 81 (14):213, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
213.pdf.

Anonymous:1899:REI

[10497] Anonymous. A rotary engine with improved abutments. Scientific American, 81(14):213–214, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
213b.pdf.

Anonymous:1899:HSI

[10498] Anonymous. The Hayne suspended ink-well. Scientific American, 81 (14):214–215, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
214.pdf.

Anonymous:1899:NWg

[10499] Anonymous. Navies of the world. Scientific American, 81(14): 216–219, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
216.pdf.

Anonymous:1899:BNBbc

[10500] Anonymous. Books: New books, etc. Scientific American, 81 (14):220, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
220a.pdf.

Anonymous:1899:BPbo

[10501] Anonymous. Business and personal. Scientific American, 81 (14):220, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
220b.pdf.

Anonymous:1899:RPIbm

[10502] Anonymous. Recently patented inventions. Scientific American, 81 (14):220, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
220.pdf.

Anonymous:1899:NQbm

[10503] Anonymous. Notes and queries. Scientific American, 81(14): 220–221, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
220c.pdf.

Anonymous:1899:IIbn

[10504] Anonymous. Index of inventions. Scientific American, 81(14): 221, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
221.pdf.

Anonymous:1899:PNPbl

[10505] Anonymous. Patent notes: Patents. Scientific American, 81 (14):223, September 30, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n14/pdf/scientificamerican09301899-
223.pdf.

Anonymous:1899:HRI

[10506] Anonymous. History repeats itself. Scientific American, 81(15):226, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-226b.pdf.

Anonymous:1899:IHa

[10507] Anonymous. The improvement of Havana. Scientific American, 81(15):226, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
226a.pdf.

Anonymous:1899:KAC

[10508] Anonymous. Krupp Armor and Congress. Scientific American, 81(15):226, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
226.pdf.

Anonymous:1899:LCU

[10509] Anonymous. Lake Chelan upheaval. Scientific American, 81(15):226, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-226d.pdf.

Anonymous:1899:PE

[10510] Anonymous. A photographic exhibition. Scientific American, 81 (15):226, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
226e.pdf.

Anonymous:1899:TMD

[10511] Anonymous. Trade mark decision by the Imperial German Patent Office. Scientific American, 81(15):226, October 7, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n15/pdf/scientificamerican10071899-226c.pdf.

Anonymous:1899:CPF

[10512] Anonymous. Compressed peat fuel. Scientific American, 81(15):227, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-227b.pdf.

Anonymous:1899:NSPb

[10513] Anonymous. The new smokeless powder factory. Scientific American, 81(15):227, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
227a.pdf.

Anonymous:1899:SNe

[10514] Anonymous. Science notes. Scientific American, 81(15):227, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-227c.pdf.

Hassack:1899:MPP

[10515] Paul Hassack. The manufacture of paper pulp, paper and cardboard from peat. Scientific American, 81(15):227, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-227.pdf.

Anonymous:1899:AAG

[10516] Anonymous. An automatic acetylene-generator. Scientific American, 81(15):228, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
228a.pdf.

Anonymous:1899:HMC

[10517] Anonymous. A home-made Cold box and water cooler. Scientific American, 81(15):228, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
228.pdf.

Gates:1899:EES

[10518] Elmer Gates. Electrographs of the electro-static current made without a camera. Scientific American, 81(15):228–229, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-228b.pdf.

Anonymous:1899:FYIa

[10519] Anonymous. Fifty years of International Yacht Racing. Scientific American, 81(15):230, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
230.pdf.

Anonymous:1899:NCBb

[10520] Anonymous. A new chainless bicycle. Scientific American, 81(15):231, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-231a.pdf.

Dewar:1899:SH

[10521] James Dewar. Solid hydrogen. Scientific American, 81(15):231, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-231.pdf.

Anonymous:1899:CGC

[10522] Anonymous. A courthouse going to court. Scientific American, 81(15):232, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
232.pdf.

Anonymous:1899:MSM

[10523] Anonymous. Modern stage mechanism. Scientific American, 81 (15):232–233, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
232a.pdf.

Holder:1899:WSI

[10524] C. F. Holder. The wind-swept island of San Nicolas. Scientific American, 81(15):233–235, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
233.pdf.

Anonymous:1899:IBB

[10525] Anonymous. The illumination of the Brooklyn Bridge. Scientific American, 81(15):235–236, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
235.pdf.

Anonymous:1899:RPIbn

[10526] Anonymous. Recently patented inventions. Scientific American, 81 (15):236–237, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n15/pdf/scientificamerican10071899-
236.pdf.

Anonymous:1899:BNBbd

[10527] Anonymous. Books: New books, etc. Scientific American, 81(15):237, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-237c.pdf.

Anonymous:1899:BPbp

[10528] Anonymous. Business and personal. Scientific American, 81(15):237, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-237a.pdf.

Anonymous:1899:Iu

[10529] Anonymous. To inventors. Scientific American, 81(15):237, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-237d.pdf.

Anonymous:1899:NQbn

[10530] Anonymous. Notes and queries. Scientific American, 81(15):237, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-237b.pdf.

Anonymous:1899:SR

[10531] Anonymous. Something to remember. Scientific American, 81(15):237, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-237.pdf.

Anonymous:1899:IIbo

[10532] Anonymous. Index of inventions. Scientific American, 81(15):237–239, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-237e.pdf.

Anonymous:1899:PNPbm

[10533] Anonymous. Patent notes: Patents. Scientific American, 81(15):239, October 7, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n15/pdf/scientificamerican10071899-239.pdf.

Anonymous:1899:CSDb

[10534] Anonymous. “Columbia” and “Shamrock” in dry dock. Scientific American, 81(16):242, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
242a.pdf.

Anonymous:1899:MT

[10535] Anonymous. Marconi telegraphy. Scientific American, 81(16):242, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n16/pdf/scientificamerican10141899-242.pdf.

Anonymous:1899:PCP

[10536] Anonymous. Proposed cycle path across the Brooklyn Bridge. Scientific American, 81(16):242, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
242b.pdf.

Anonymous:1899:AIT

[10537] Anonymous. Amateur investigations with a testing tank. Scientific American, 81(16):242–243, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
242c.pdf.

Anonymous:1899:EFF

[10538] Anonymous. Exchanging fish fry with Europe. Scientific American, 81(16):243, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
243a.pdf.

Anonymous:1899:MIA

[10539] Anonymous. The mining interests of Africa. Scientific American, 81(16):243, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
243b.pdf.

Anonymous:1899:R

[10540] Anonymous. Railroads in 1898. Scientific American, 81(16):243, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n16/pdf/scientificamerican10141899-243c.pdf.

Anonymous:1899:SDb

[10541] Anonymous. Statistical duplication. Scientific American, 81(16): 243, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
243.pdf.

Anonymous:1899:SWR

[10542] Anonymous. Soda water to relieve hunger. Scientific American, 81 (16):243–244, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
243d.pdf.

Anonymous:1899:NMH

[10543] Anonymous. A novel matcher-head for planing mills. Scientific American, 81(16):244, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
244a.pdf.

Anonymous:1899:NREb

[10544] Anonymous. A new Riker electric vehicle. Scientific American, 81(16):244, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
244.pdf.

Crafts:1899:DAG

[10545] B. A. Crafts. Ditching across the Great Divides in the Rocky Mountains. Scientific American, 81(16):245–247, October 14, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n16/pdf/scientificamerican10141899-245.pdf.

Anonymous:1899:FYIb

[10546] Anonymous. Fifty years of International Yacht Racing. Scientific American, 81(16):248–250, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
248.pdf.

Anonymous:1899:BC

[10547] Anonymous. The battle chariot. Scientific American, 81(16):251, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n16/pdf/scientificamerican10141899-251a.pdf.

Caulk:1899:MO

[10548] William B. Caulk. The Magician’s omelette. Scientific American, 81(16):251, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
251.pdf.

Anonymous:1899:BPbq

[10549] Anonymous. Business and personal. Scientific American, 81 (16):252, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
252a.pdf.

Anonymous:1899:NQbo

[10550] Anonymous. Notes and queries. Scientific American, 81(16):252, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n16/pdf/scientificamerican10141899-252b.pdf.

Anonymous:1899:RPIbo

[10551] Anonymous. Recently patented inventions. Scientific American, 81(16):252, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
252.pdf.

Anonymous:1899:BNBbe

[10552] Anonymous. Books: New books, etc. Scientific American, 81 (16):252–253, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
252c.pdf.

Anonymous:1899:Iv

[10553] Anonymous. To inventors. Scientific American, 81(16):253, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n16/pdf/scientificamerican10141899-253.pdf.

Anonymous:1899:IIbp

[10554] Anonymous. Index of inventions. Scientific American, 81(16): 253–255, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
253a.pdf.

Anonymous:1899:PNPbn

[10555] Anonymous. Patent notes: Patents. Scientific American, 81 (16):255, October 14, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n16/pdf/scientificamerican10141899-
255.pdf.

Anonymous:1899:DCC

[10556] Anonymous. The “Denver” class of cruisers. Scientific American, 81(17):258, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
258b.pdf.

Anonymous:1899:HYR

[10557] Anonymous. Hysteria and yacht racing. Scientific American, 81 (17):258, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
258.pdf.

Anonymous:1899:LAP

[10558] Anonymous. Liquid air promoting. Scientific American, 81(17):258, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n17/pdf/scientificamerican10211899-258a.pdf.

Anonymous:1899:TR

[10559] Anonymous. A timely rebuke. Scientific American, 81(17):258, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n17/pdf/scientificamerican10211899-258c.pdf.

Anonymous:1899:GAFb

[10560] Anonymous. Government advertising in France. Scientific American, 81(17):258–259, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
258d.pdf.

Anonymous:1899:ASP

[10561] Anonymous. American steel for a Prince’s palace. Scientific American, 81(17):259, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
259c.pdf.

Anonymous:1899:BG

[10562] Anonymous. The Boers and the giraffe. Scientific American, 81 (17):259, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
259a.pdf.

Anonymous:1899:LPT

[10563] Anonymous. A long photographic telescope. Scientific American, 81(17):259, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
259d.pdf.

Anonymous:1899:VCI

[10564] Anonymous. Vitrified claypipe instead of iron for gas mains. Scientific American, 81(17):259, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
259b.pdf.

Field:1899:DES

[10565] Archelaus G. Field. The defective eyesight of school children. Scientific American, 81(17):259, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
259.pdf.

Anonymous:1899:BRF

[10566] Anonymous. A boiler with a removable fire box. Scientific American, 81(17):260, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
260b.pdf.

Anonymous:1899:CHS

[10567] Anonymous. A compact high speed engine. Scientific American, 81(17):260, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
260a.pdf.

Anonymous:1899:CWS

[10568] Anonymous. A combined walking and sulky plow. Scientific American, 81(17):260, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
260c.pdf.

Anonymous:1899:IFCb

[10569] Anonymous. An improvement in friction clutch-pulleys. Scientific American, 81(17):260, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
260.pdf.

Anonymous:1899:PSTb

[10570] Anonymous. The Philadelphia Subway and Tunnel. Scientific American, 81(17):261–262, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
261.pdf.

Anonymous:1899:Cbl

[10571] Anonymous. Correspondence. Scientific American, 81(17):262–263, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n17/pdf/scientificamerican10211899-262.pdf.

Anonymous:1899:PSBa

[10572] Anonymous. Preliminary skirmishing between “Shamrock” and “Columbia”. Scientific American, 81(17):264, October 21, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n17/pdf/scientificamerican10211899-264.pdf.

Sails:1899:BS

[10573] Walter Burnham Sails. Boats and sails. Scientific American, 81 (17):265–266, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
265.pdf.

Anonymous:1899:XHD

[10574] Anonymous. Xiphopages, or human doubles. Scientific American, 81(17):266, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
266.pdf.

Lumsden:1899:TES

[10575] G. E. Lumsden. Total eclipses of the Sun. Scientific American, 81 (17):267–268, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
267.pdf.

Anonymous:1899:RPIbp

[10576] Anonymous. Recently patented inventions. Scientific American, 81 (17):268–269, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
268.pdf.

Anonymous:1899:BNBbf

[10577] Anonymous. Books: New books, etc. Scientific American, 81 (17):269, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
269b.pdf.

Anonymous:1899:BPbr

[10578] Anonymous. Business and personal. Scientific American, 81 (17):269, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
269.pdf.

Anonymous:1899:Ix

[10579] Anonymous. To inventors. Scientific American, 81(17):269, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n17/pdf/scientificamerican10211899-269c.pdf.

Anonymous:1899:NQbp

[10580] Anonymous. Notes and queries. Scientific American, 81(17):269, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n17/pdf/scientificamerican10211899-269a.pdf.

Anonymous:1899:IIbq 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
269d.pdf.

Anonymous:1899:PNPbo

[10582] Anonymous. Patent notes: Patents. Scientific American, 81 (17):271, October 21, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n17/pdf/scientificamerican10211899-
271.pdf.

Anonymous:1899:ASDa

[10583] Anonymous. Another successful defense of the “America” Cup. Scientific American, 81(18):274, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
274.pdf.

Anonymous:1899:ET

[10584] Anonymous. An engineering triumph. Scientific American, 81 (18):274, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
274a.pdf.

Anonymous:1899:RAH

[10585] Anonymous. Rear-Admiral Hichborn on the needs of the Navy. Scientific American, 81(18):274–275, October 28, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n18/pdf/scientificamerican10281899-274b.pdf.

Anonymous:1899:DAC

[10586] Anonymous. Death of Admiral Colomb. Scientific American, 81 (18):275, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
275a.pdf.

Anonymous:1899:NZV

[10587] Anonymous. A New Zealand Vesuvius. Scientific American, 81 (18):275, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
275b.pdf.

Serviss:1899:HN

[10588] Garrett P. Serviss. The heavens in November. Scientific American, 81(18):275, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
275.pdf.

Anonymous:1899:FSH

[10589] Anonymous. A four-spindle horizontal car-boring machine. Scientific American, 81(18):276, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
276a.pdf.

Anonymous:1899:NGC

[10590] Anonymous. A new gang or cultivator plow. Scientific American, 81(18):276, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
276.pdf.

Anonymous:1899:MBM

[10591] Anonymous. A mechanical butter moulder and cutter. Scientific American, 81(18):276–277, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
276b.pdf.

Anonymous:1899:PAG

[10592] Anonymous. A portable acetylene gas lamp. Scientific American, 81(18):277, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
277a.pdf.

Anonymous:1899:RAa

[10593] Anonymous. A racing automobile. Scientific American, 81(18):277, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n18/pdf/scientificamerican10281899-277.pdf.

Anonymous:1899:Cbm

[10594] Anonymous. Correspondence. Scientific American, 81(18):278, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n18/pdf/scientificamerican10281899-278.pdf.

Anonymous:1899:CSJ

[10595] Anonymous. Control of the San Juan River at Greytown. Scientific American, 81(18):278, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
278a.pdf.

Anonymous:1899:WTY

[10596] Anonymous. Wireless telegraphy at the yacht races. Scientific American, 81(18):279, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
279.pdf.

Pickering:1899:NM

[10597] E. C. Pickering. The November meteors of 1899. Scientific American, 81(18):279, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
279a.pdf.

Holder:1899:SHP

[10598] Charles F. Holder. Some harbors and piers of Southern California. Scientific American, 81(18):280, October 28, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n18/pdf/scientificamerican10281899-280.pdf.

Anonymous:1899:DCA

[10599] Anonymous. Destruction of the city acetylene gas plant at Wabash, Ind. Scientific American, 81(18):281–282, October 28, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n18/pdf/scientificamerican10281899-281.pdf.

Anonymous:1899:GC

[10600] Anonymous. A giant cactus. Scientific American, 81(18):283, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n18/pdf/scientificamerican10281899-283.pdf.

Anonymous:1899:BNBbg

[10601] Anonymous. Books: New books, etc. Scientific American, 81 (18):284, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
284a.pdf.

Anonymous:1899:RPIbq

[10602] Anonymous. Recently patented inventions. Scientific American, 81(18):284, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
284.pdf.

Anonymous:1899:BPbs

[10603] Anonymous. Business and personal. Scientific American, 81 (18):285, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
285.pdf.

Anonymous:1899:Iy

[10604] Anonymous. To inventors. Scientific American, 81(18):285, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n18/pdf/scientificamerican10281899-285b.pdf.

Anonymous:1899:NQbq

[10605] Anonymous. Notes and queries. Scientific American, 81(18):285, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n18/pdf/scientificamerican10281899-285a.pdf.

Anonymous:1899:IIbr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
285c.pdf.

Anonymous:1899:PNPbp

[10607] Anonymous. Patent notes: Patents. Scientific American, 81 (18):287, October 28, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n18/pdf/scientificamerican10281899-
287.pdf.

Anonymous:1899:ANb

[10608] Anonymous. Aerial navigation. Scientific American, 81(19): 290, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
290b.pdf.

Anonymous:1899:CNY

[10609] Anonymous. Comparison of New York cable, electric, and horse car lines. Scientific American, 81(19):290, November 4, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n19/pdf/scientificamerican11041899-290.pdf.

Anonymous:1899:WTAb

[10610] Anonymous. Wireless telegraphy in the Army and Navy. Scientific American, 81(19):290, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
290a.pdf.

Anonymous:1899:SNO

[10611] Anonymous. Stability of the “New Orleans” and the “Albany”. Scientific American, 81(19):290–291, November 4, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n19/pdf/scientificamerican11041899-290c.pdf.

Anonymous:1899:DDT

[10612] Anonymous. Double-deck turrets. Scientific American, 81(19): 291, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
291.pdf.

Anonymous:1899:DPE

[10613] Anonymous. Death of Prof Edward Orton. Scientific American, 81(19):291, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
291b.pdf.

Anonymous:1899:NYC

[10614] Anonymous. New York’s coast defenses. Scientific American, 81 (19):291, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
291c.pdf.

Anonymous:1899:RAE

[10615] Anonymous. Recent applications of electro-metallurgy. Scientific American, 81(19):291, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
291e.pdf.

Anonymous:1899:STP

[10616] Anonymous. A seed testing plant. Scientific American, 81(19): 291, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
291d.pdf.

Moore:1899:IMA

[10617] H. F. Moore. Introducing marine animals into great salt lake. Scientific American, 81(19):291, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
291a.pdf.

Anonymous:1899:IRD

[10618] Anonymous. An improvement in rotary drum-driers. Scientific American, 81(19):292, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
292a.pdf.

Anonymous:1899:UCC

[10619] Anonymous. An underground conduit-crossing for electrical conductors. Scientific American, 81(19):292, November 4, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n19/pdf/scientificamerican11041899-292.pdf.

Stfwart:1899:WEB

[10620] J. A. Stfwart. Waterworks expansion in Boston. Scientific American, 81(19):292–294, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
292b.pdf.

Anonymous:1899:Cbn

[10621] Anonymous. Correspondence. Scientific American, 81(19):294– 295, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
294.pdf.

Fawcett:1899:YGL

[10622] Waldon Fawcett. Yachting on the Great Lakes. Scientific American, 81(19):296–297, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
296.pdf.

Anonymous:1899:RDR

[10623] Anonymous. Recent discoveries in the Roman forum. Scientific American, 81(19):297, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
297.pdf.

Anonymous:1899:GTPb

[10624] Anonymous. The great telescope of the Paris Exposition of 1900. Scientific American, 81(19):298–299, November 4, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n19/pdf/scientificamerican11041899-298.pdf.

Anonymous:1899:NDAb

[10625] Anonymous. The “New Departure” automatic coaster. Scientific American, 81(19):299, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
299.pdf.

Anonymous:1899:ORJ

[10626] Anonymous. Oil-retaining journal. Scientific American, 81(19): 299, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
299a.pdf.

Anonymous:1899:BNBbh

[10627] Anonymous. Books: New books, etc. Scientific American, 81 (19):300, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
300a.pdf.

Anonymous:1899:RPIbr

[10628] Anonymous. Recently patented inventions. Scientific American, 81(19):300, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
300.pdf.

Anonymous:1899:BPbt

[10629] Anonymous. Business and personal. Scientific American, 81 (19):301, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
301.pdf.

Anonymous:1899:Iz

[10630] Anonymous. To inventors. Scientific American, 81(19):301, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n19/pdf/scientificamerican11041899-301b.pdf.

Anonymous:1899:NQbr

[10631] Anonymous. Notes and queries. Scientific American, 81(19): 301, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
301a.pdf.

Anonymous:1899:IIbs

[10632] Anonymous. Index of inventions. Scientific American, 81(19): 301–303, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
301c.pdf.

Anonymous:1899:PNPbq

[10633] Anonymous. Patent notes: Patents. Scientific American, 81 (19):303, November 4, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n19/pdf/scientificamerican11041899-
303.pdf.

Anonymous:1899:BOA

[10634] Anonymous. Bureau of Ordnance on the armor question. Scientific American, 81(20):306, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
306b.pdf.

Anonymous:1899:CDTa

[10635] Anonymous. Commissioner Duell on trade marks. Scientific American, 81(20):306, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
306a.pdf.

Anonymous:1899:EPP

[10636] Anonymous. The electric power problem of Niagara. Scientific American, 81(20):306, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
306.pdf.

Anonymous:1899:AEC

[10637] Anonymous. American equipment for the City of Glasgow. Scientific American, 81(20):306–307, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
306c.pdf.

Anonymous:1899:CDTb

[10638] Anonymous. Commissioner Duell on trade marks. Scientific American, 81(20):307, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
307.pdf.

Anonymous:1899:NGP

[10639] Anonymous. The new Government Printing Office. Scientific American, 81(20):307, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
307b.pdf.

Anonymous:1899:NMDb

[10640] Anonymous. New method of detecting gold. Scientific American, 81(20):307, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
307d.pdf.

Anonymous:1899:OCN

[10641] Anonymous. Our California national parks. Scientific American, 81(20):307, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
307c.pdf.

Anonymous:1899:SNY

[10642] Anonymous. Searchlights for the New York fire department. Scientific American, 81(20):307, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
307a.pdf.

Anonymous:1899:UCS

[10643] Anonymous. Uses for corn stalks. Scientific American, 81(20): 307, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
307e.pdf.

Fawcett:1899:PLC

[10644] Waldon Fawcett. Problems in lake commerce. Scientific American, 81(20):308–309, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
308.pdf.

Anonymous:1899:AHB

[10645] Anonymous. An automatic hoop and basket strip cutting machine. Scientific American, 81(20):309, November 11, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n20/pdf/scientificamerican11111899-309.pdf.

Hopkins:1899:CMB

[10646] N. Monroe Hopkins. Cases of “mysterious” busting. Scientific American, 81(20):309–310, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
309a.pdf.

Anonymous:1899:NFH

[10647] Anonymous. Niagara Falls Hydraulic Power Plant. Scientific American, 81(20):310–311, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
310.pdf.

Anonymous:1899:ONF

[10648] Anonymous. Our new Fleet of torpedo boat destroyers. Scientific American, 81(20):312, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
312.pdf.

Anonymous:1899:AMW

[10649] Anonymous. Acetylene motor wagons and carriages. Scientific American, 81(20):313–314, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
313.pdf.

Anonymous:1899:DGI

[10650] Anonymous. The death of a great inventor. Scientific American, 81(20):314, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
314.pdf.

Anonymous:1899:FWA

[10651] Anonymous. A feed-water apparatus for boilers. Scientific American, 81(20):315, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
315.pdf.

Anonymous:1899:SDM

[10652] Anonymous. A simple driving mechanism for gins. Scientific American, 81(20):315, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
315a.pdf.

Anonymous:1899:MNP

[10653] Anonymous. The “Maestro” a new playing attachment for organs. Scientific American, 81(20):315–316, November 11, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n20/pdf/scientificamerican11111899-315b.pdf.

Anonymous:1899:RPIbs

[10654] Anonymous. Recently patented inventions. Scientific American, 81(20):316, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
316.pdf.

Anonymous:1899:BNBbi

[10655] Anonymous. Books: New books, etc. Scientific American, 81 (20):317, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
317b.pdf.

Anonymous:1899:BPbu

[10656] Anonymous. Business and personal. Scientific American, 81 (20):317, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
317.pdf.

Anonymous:1899:IIbt

[10657] Anonymous. Index of inventions. Scientific American, 81(20): 317, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
317c.pdf.

Anonymous:1899:NQbs

[10658] Anonymous. Notes and queries. Scientific American, 81(20): 317, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
317a.pdf.

Anonymous:1899:PNPbr

[10659] Anonymous. Patent notes: Patents. Scientific American, 81 (20):319, November 11, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n20/pdf/scientificamerican11111899-
319.pdf.

Anonymous:1899:BPCa

[10660] Anonymous. Bids for the proposed cruisers. Scientific American, 81(21):322, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
322b.pdf.

Anonymous:1899:MPS

[10661] Anonymous. The mammoth power stations of New York City. Scientific American, 81(21):322, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
322.pdf.

Anonymous:1899:NTM

[10662] Anonymous. Naval tests of Marconi telegraphy. Scientific American, 81(21):322, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
322c.pdf.

Anonymous:1899:SP

[10663] Anonymous. Soil parasites. Scientific American, 81(21):322, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n21/pdf/scientificamerican11181899-322a.pdf.

Anonymous:1899:ONY

[10664] Anonymous. Opening of the New York Zoological Park. Scientific American, 81(21):322–323, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
322d.pdf.

Anonymous:1899:CBP

[10665] Anonymous. Coloring bromide prints. Scientific American, 81 (21):323, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
323c.pdf.

Anonymous:1899:IEP

[10666] Anonymous. Interesting experiments with photographic plates. Scientific American, 81(21):323, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
323b.pdf.

Anonymous:1899:MSTb

[10667] Anonymous. Masses smaller than atoms. Scientific American, 81 (21):323, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
323a.pdf.

Anonymous:1899:PHA

[10668] Anonymous. Production of hydrogen with the aid of magnesium. Scientific American, 81(21):323, November 18, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n21/pdf/scientificamerican11181899-323d.pdf.

Anonymous:1899:RBS

[10669] Anonymous. Report of the Bureau of Steam Engineering. Scientific American, 81(21):323, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
323.pdf.

DeKubicki:1899:SFT

[10670] C. De Kubicki. Some French types of automobiles. Scientific American, 81(21):324, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
324.pdf.

Anonymous:1899:PLN

[10671] Anonymous. The pea louse, a new and important economic species of the genus Nectarophora. Scientific American, 81(21):325–326, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n21/pdf/scientificamerican11181899-325.pdf.

Anonymous:1899:Cbo

[10672] Anonymous. Correspondence. Scientific American, 81(21):327, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n21/pdf/scientificamerican11181899-327.pdf.

Anonymous:1899:SDT

[10673] Anonymous. A simple device for tightening fence-wires. Scientific American, 81(21):327, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
327a.pdf.

Hatcher:1899:EP

[10674] J. B. Hatcher. Explorations in Patagonia. Scientific American, 81 (21):328–329, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
328.pdf.

Anonymous:1899:PSTc

[10675] Anonymous. The Parsons steam turbine. Scientific American, 81 (21):329–331, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
329.pdf.

Anonymous:1899:BAGb

[10676] Anonymous. The “Becklight” acetylene-gas generator. Scientific American, 81(21):331, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
331a.pdf.

Anonymous:1899:BNBbj

[10677] Anonymous. Books: New books, etc. Scientific American, 81 (21):331, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
332a.pdf.

Anonymous:1899:BPbv

[10678] Anonymous. Business and personal. Scientific American, 81 (21):331, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
332b.pdf.

Anonymous:1899:NQbt

[10679] Anonymous. Notes and queries. Scientific American, 81(21): 331, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
332c.pdf.

Anonymous:1899:RPIbt

[10680] Anonymous. Recently patented inventions. Scientific American, 81(21):331, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
332.pdf.

Anonymous:1899:TWI

[10681] Anonymous. A traction-wheel of improved form. Scientific American, 81(21):331, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
331.pdf.

Anonymous:1899:IIbu

[10682] Anonymous. Index of inventions. Scientific American, 81(21): 333, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
333.pdf.

Anonymous:1899:PNPbs

[10683] Anonymous. Patent notes: Patents. Scientific American, 81 (21):335, November 18, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n21/pdf/scientificamerican11181899-
335.pdf.

Anonymous:1899:BKC

[10684] Anonymous. Bids for 16 1/2-knot cruisers accepted. Scientific American, 81(22):338, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
338a.pdf.

Anonymous:1899:IIGa

[10685] Anonymous. Industrial invasion of Great Britain. Scientific American, 81(22):338, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
338.pdf.

Anonymous:1899:MD

[10686] Anonymous. The meteor display. Scientific American, 81(22): 338, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
338b.pdf.

Anonymous:1899:RAb

[10687] Anonymous. Railroads in Asia. Scientific American, 81(22): 338, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
338c.pdf.

Anonymous:1899:RBO

[10688] Anonymous. Report of the Bureau of Ordnance. Scientific American, 81(22):338, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
338d.pdf.

Anonymous:1899:WTBb

[10689] Anonymous. Wireless telegraphy between balloons. Scientific American, 81(22):338, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
338e.pdf.

Anonymous:1899:DST

[10690] Anonymous. A discussion of the superimposed turret for warships. Scientific American, 81(22):339, November 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n22/pdf/scientificamerican11251899-339a.pdf.

Hale:1899:MNA

[10691] William H. Hale. Meeting of the National Academy of Sciences. Scientific American, 81(22):339, November 25, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n22/pdf/scientificamerican11251899-339.pdf.

Anonymous:1899:NLBb

[10692] Anonymous. New letter-box indicator. Scientific American, 81 (22):340, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
340.pdf.

Heap:1899:MRO

[10693] D. P. Heap. Methods of revolving the optical apparatus for lighthouses. Scientific American, 81(22):340–341, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n22/pdf/scientificamerican11251899-340a.pdf.

Anonymous:1899:PPE

[10694] Anonymous. Progress of the Paris Exposition of 1900. Scientific American, 81(22):342–343, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
342.pdf.

Anonymous:1899:AAP

[10695] Anonymous. American armor plate for a Russian battleship. Scientific American, 81(22):344, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
344.pdf.

Anonymous:1899:GM

[10696] Anonymous. The gypsy moth. Scientific American, 81(22):345, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n22/pdf/scientificamerican11251899-345.pdf.

Anonymous:1899:LSM

[10697] Anonymous. Long-scale measuring instruments. Scientific American, 81(22):345–346, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
345a.pdf.

Benjamin:1899:HYC

[10698] Marcus Benjamin. Hamilton Young Castner. Scientific American, 81(22):346, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
346.pdf.

Holder:1899:DC

[10699] Chas. Fredk. Holder. “Don Coyote”. Scientific American, 81(22): 346–347, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
346a.pdf.

Anonymous:1899:IBI

[10700] Anonymous. An improved branding instrument. Scientific American, 81(22):347, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
347.pdf.

Anonymous:1899:SSV

[10701] Anonymous. A simple slide-valve for steam-engines. Scientific American, 81(22):347, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
347a.pdf.

Anonymous:1899:BPbw

[10702] Anonymous. Business and personal. Scientific American, 81 (22):348, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
348a.pdf.

Anonymous:1899:Iba

[10703] Anonymous. To inventors. Scientific American, 81(22):348, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n22/pdf/scientificamerican11251899-348c.pdf.

Anonymous:1899:IIbv

[10704] Anonymous. Index of inventions. Scientific American, 81(22): 348, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
348d.pdf.

Anonymous:1899:NQbu

[10705] Anonymous. Notes and queries. Scientific American, 81(22): 348, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
348b.pdf.

Anonymous:1899:RPIbu

[10706] Anonymous. Recently patented inventions. Scientific American, 81(22):348, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
348.pdf.

Anonymous:1899:PNPbt

[10707] Anonymous. Patent notes: Patents. Scientific American, 81 (22):351, November 25, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n22/pdf/scientificamerican11251899-
351.pdf.

Anonymous:1899:CNC

[10708] Anonymous. Congress and the Nicaragua Canal. Scientific American, 81(23):354, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
354.pdf.

Anonymous:1899:CNE

[10709] Anonymous. Cables of the new East River Bridge. Scientific American, 81(23):354, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
354d.pdf.

Anonymous:1899:EFA

[10710] Anonymous. The explosive features of acetylene. Scientific American, 81(23):354, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
354b.pdf.

Anonymous:1899:GE

[10711] Anonymous. “A gross error”. Scientific American, 81(23):354, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n23/pdf/scientificamerican12021899-354a.pdf.

Anonymous:1899:YCE

[10712] Anonymous. The yacht “Columbia” in European waters. Scientific American, 81(23):354, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
354c.pdf.

Anonymous:1899:NNP

[10713] Anonymous. The new Naval programme. Scientific American, 81 (23):354–355, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
354e.pdf.

Anonymous:1899:AFP

[10714] Anonymous. American fruits at the Paris Exposition. Scientific American, 81(23):355, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
355c.pdf.

Anonymous:1899:CCU

[10715] Anonymous. Compressed cork and its uses. Scientific American, 81(23):355, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
355b.pdf.

Anonymous:1899:CWM

[10716] Anonymous. Can white men live in the tropics? Scientific American, 81(23):355, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
355d.pdf.

Anonymous:1899:DSW

[10717] Anonymous. Death of Sir William Dawson. Scientific American, 81(23):355, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
355e.pdf.

Anonymous:1899:MRF

[10718] Anonymous. The monument to Robert Fulton. Scientific American, 81(23):355, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
355.pdf.

Serviss:1899:HD

[10719] Garrett P. Serviss. The heavens in December. Scientific American, 81(23):355, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
355a.pdf.

Kobbe:1899:TBM

[10720] Gustav Kobbé. The tidal “bore” at Moncton. Scientific American, 81(23):356, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
356.pdf.

Anonymous:1899:DGE

[10721] Anonymous. A destructive gas explosion. Scientific American, 81(23):357, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
357a.pdf.

Anonymous:1899:EFLb

[10722] Anonymous. An electric flash-light device. Scientific American, 81(23):357, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
357.pdf.

Mels:1899:FSA

[10723] Edgar Mels. The future of South Africa — II. Scientific American, 81(23):358–359, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
358.pdf.

Anonymous:1899:SEP

[10724] Anonymous. Simple experiments in physics and chemistry. Scientific American, 81(23):360, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
360.pdf.

Anonymous:1899:LCK

[10725] Anonymous. The Lecornu cellular kite. Scientific American, 81 (23):360–361, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
360a.pdf.

Anonymous:1899:NYB

[10726] Anonymous. The New York botanical garden. Scientific American, 81(23):361–362, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
361.pdf.

Holder:1899:PLF

[10727] C. F. Holder. Photographing living fishes at Santa Catalina Zoological Station. Scientific American, 81(23):362–363, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n23/pdf/scientificamerican12021899-362.pdf.

Anonymous:1899:Cbq

[10728] Anonymous. Correspondence. Scientific American, 81(23):363, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n23/pdf/scientificamerican12021899-363a.pdf.

Anonymous:1899:IFCc

[10729] Anonymous. An improvement in flue-cutters. Scientific American, 81(23):363, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
363.pdf.

Anonymous:1899:BPbx

[10730] Anonymous. Business and personal. Scientific American, 81 (23):364, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
364a.pdf.

Anonymous:1899:IIbw

[10731] Anonymous. Index of inventions. Scientific American, 81(23):364, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n23/pdf/scientificamerican12021899-364c.pdf.

Anonymous:1899:NQbv

[10732] Anonymous. Notes and queries. Scientific American, 81(23):364, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n23/pdf/scientificamerican12021899-364b.pdf.

Anonymous:1899:RPIbv

[10733] Anonymous. Recently patented inventions. Scientific American, 81(23):364, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
364.pdf.

Anonymous:1899:PNPbu

[10734] Anonymous. Patent notes: Patents. Scientific American, 81 (23):367, December 2, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n23/pdf/scientificamerican12021899-
367.pdf.

Anonymous:1899:AAW

[10735] Anonymous. Airship or aeroplane — which. Scientific American, 81(24):370, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
370.pdf.

Anonymous:1899:DE

[10736] Anonymous. The Dalmeny Experiments. Scientific American, 81 (24):370, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
370a.pdf.

Anonymous:1899:SOSa

[10737] Anonymous. The secrets of our success in the steel trade. Scientific American, 81(24):370–371, December 9, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n24/pdf/scientificamerican12091899-370b.pdf.

Anonymous:1899:ABA

[10738] Anonymous. The automobile in Buenos Ayres. Scientific American, 81(24):371, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
371c.pdf.

Anonymous:1899:DIA

[10739] Anonymous. The danger from the importation of animals. Scientific American, 81(24):371, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
371.pdf.

Anonymous:1899:IPT

[10740] Anonymous. Importance of patents and trade marks in Germany. Scientific American, 81(24):371, December 9, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n24/pdf/scientificamerican12091899-371b.pdf.

Anonymous:1899:WTT

[10741] Anonymous. Wireless telegraphy tests in Switzerland. Scientific American, 81(24):371, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
371a.pdf.

Anonymous:1899:ECSb

[10742] Anonymous. End of the Creusot strike. Scientific American, 81 (24):371–372, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
371d.pdf.

Anonymous:1899:SSL

[10743] Anonymous. A safety switch-locking mechanism for railways. Scientific American, 81(24):372, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
372a.pdf.

MichaudDSc:1899:XRD

[10744] Gustave Michaud, D.Sc. An X-ray delusion. Scientific American, 81(24):372, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
372.pdf.

Jenkins:1899:RS

[10745] C. Francis Jenkins. Railway signals. Scientific American, 81(24): 373–374, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
373.pdf.

Anonymous:1899:ECb

[10746] Anonymous. Education by correspondence. Scientific American, 81 (24):374–375, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
374.pdf.

Fawcett:1899:SBF

[10747] Waldon Fawcett. The submarine boat and its future. Scientific American, 81(24):376–377, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
376.pdf.

Anonymous:1899:NFA

[10748] Anonymous. The new French armor-clad “Henri Quatre”. Scientific American, 81(24):377, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
377.pdf.

Anonymous:1899:Cbr

[10749] Anonymous. Correspondence. Scientific American, 81(24):378, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n24/pdf/scientificamerican12091899-378.pdf.

Holder:1899:CC

[10750] Chas. Fredk. Holder. The crime of a century. Scientific American, 81(24):378–379, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
378a.pdf.

Anonymous:1899:RPIbw

[10751] Anonymous. Recently patented inventions. Scientific American, 81(24):380, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
380.pdf.

Anonymous:1899:BNBbk

[10752] Anonymous. Books: New books, etc. Scientific American, 81 (24):381, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
381b.pdf.

Anonymous:1899:BPby

[10753] Anonymous. Business and personal. Scientific American, 81 (24):381, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
381.pdf.

Anonymous:1899:IIbx

[10754] Anonymous. Index of inventions. Scientific American, 81(24):381, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n24/pdf/scientificamerican12091899-381c.pdf.

Anonymous:1899:NQbw

[10755] Anonymous. Notes and queries. Scientific American, 81(24):381, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n24/pdf/scientificamerican12091899-381a.pdf.

Anonymous:1899:PNPbv

[10756] Anonymous. Patent notes: Patents. Scientific American, 81 (24):383, December 9, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n24/pdf/scientificamerican12091899-
383.pdf.

Anonymous:1899:ONS

[10757] Anonymous. Opening of the new St. Lawrence Canal Locks. Scientific American, 81(25):386, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
386a.pdf.

Anonymous:1899:RRB

[10758] Anonymous. Relation of rainfall to boiler scale. Scientific American, 81(25):386, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
386.pdf.

Anonymous:1899:WFT

[10759] Anonymous. “The wish is father to the thought”. Scientific American, 81(25):386, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
386b.pdf.

Anonymous:1899:TVB

[10760] Anonymous. Tunnels versus bridges for rapid transit. Scientific American, 81(25):386–387, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
386c.pdf.

Anonymous:1899:AC

[10761] Anonymous. The automobile cup. Scientific American, 81(25): 387, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
387b.pdf.

Anonymous:1899:AEP

[10762] Anonymous. Ancient Egyptian porcelain. Scientific American, 81 (25):387, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
387c.pdf.

Anonymous:1899:ASG

[10763] Anonymous. “Advance Sheets” of the German Consular Reports. Scientific American, 81(25):387, December 16, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n25/pdf/scientificamerican12161899-387e.pdf.

Anonymous:1899:NYTb

[10764] Anonymous. New York’s trade in wild animals. Scientific American, 81(25):387, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
387f.pdf.

Anonymous:1899:OEI

[10765] Anonymous. Our exports of iron and steel. Scientific American, 81(25):387, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
387.pdf.

Anonymous:1899:RSA

[10766] Anonymous. Railway safety appliances. Scientific American, 81 (25):387, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
387d.pdf.

Anonymous:1899:TPO

[10767] Anonymous. Tropical products of our new possessions. Scientific American, 81(25):387, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
387a.pdf.

Anonymous:1899:HLM

[10768] Anonymous. Hauling logs by motor trucks. Scientific American, 81(25):388, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
388b.pdf.

Anonymous:1899:NGL

[10769] Anonymous. The Naval guns at Ladysmith. Scientific American, 81(25):388, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
388a.pdf.

Anonymous:1899:TRE

[10770] Anonymous. A tool for reloading empty cartridge shells. Scientific American, 81(25):388, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
388.pdf.

Anonymous:1899:EFC

[10771] Anonymous. The engines of the first-class battleship “Wisconsin”. Scientific American, 81(25):389, December 16, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n25/pdf/scientificamerican12161899-389.pdf.

Anonymous:1899:Cbt

[10772] Anonymous. Correspondence. Scientific American, 81(25):390– 391, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
390.pdf.

Anonymous:1899:THP

[10773] Anonymous. The Trojan horse at the Paris opera house. Scientific American, 81(25):392, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
392.pdf.

Anonymous:1899:NSV

[10774] Anonymous. A new safety valve. Scientific American, 81(25): 393, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
393.pdf.

Holder:1899:CVA

[10775] Charles F. Holder. A California Verde antique quarry. Scientific American, 81(25):393–394, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
393a.pdf.

Anonymous:1899:NFFb

[10776] Anonymous. A new form of fire-escape and lowering apparatus. Scientific American, 81(25):394, December 16, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n25/pdf/scientificamerican12161899-394a.pdf.

Anonymous:1899:NRC

[10777] Anonymous. A novelty in railroad collisions. Scientific American, 81(25):394, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
394.pdf.

Anonymous:1899:HSC

[10778] Anonymous. How sculptors change the scale of their statues. Scientific American, 81(25):395, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
395.pdf.

Anonymous:1899:STDb

[10779] Anonymous. A switch-throwing device for cars. Scientific American, 81(25):395, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
395a.pdf.

Anonymous:1899:BNBbl

[10780] Anonymous. Books: New books, etc. Scientific American, 81 (25):396, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
396a.pdf.

Anonymous:1899:BPbz

[10781] Anonymous. Business and personal. Scientific American, 81 (25):396, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
396b.pdf.

Anonymous:1899:NQbx (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
396c.pdf.

Anonymous:1899:RPIbx

[10783] Anonymous. Recently patented inventions. Scientific American, 81(25):396, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
396.pdf.

Anonymous:1899:IIby

[10784] Anonymous. Index of inventions. Scientific American, 81(25): 397, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
397.pdf.

Anonymous:1899:PNPbw

[10785] Anonymous. Patent notes: Patents. Scientific American, 81 (25):399, December 16, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n25/pdf/scientificamerican12161899-
399.pdf.

Anonymous:1899:ABd

[10786] Anonymous. The American bridge. Scientific American, 81(26): 402, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
402a.pdf.

Anonymous:1899:ECSc

[10787] Anonymous. The end of the century steam engine. Scientific American, 81(26):402, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
402.pdf.

Anonymous:1899:PCR

[10788] Anonymous. Pre-Columbian remains in Massachusetts. Scientific American, 81(26):402–403, December 23, 1899. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
402b.pdf.

Anonymous:1899:PLSb

[10789] Anonymous. The Pollok life-saving prize. Scientific American, 81 (26):403, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
403a.pdf.

Anonymous:1899:PTU

[10790] Anonymous. The proposed tunnel under the Straits of Gibraltar. Scientific American, 81(26):403, December 23, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n26/pdf/scientificamerican12231899-403c.pdf.

Anonymous:1899:RWB

[10791] Anonymous. Report of the Weather Bureau. Scientific American, 81(26):403, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
403b.pdf.

Anonymous:1899:RWD

[10792] Anonymous. The recent Welsbach Decision. Scientific American, 81(26):403, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
403.pdf.

Anonymous:1899:BPBba

[10793] Anonymous. A back-pedaling brake for chain and chain-less bicycles. Scientific American, 81(26):404, December 23, 1899. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v81/
n26/pdf/scientificamerican12231899-404.pdf.

Anonymous:1899:IPD

[10794] Anonymous. An insect-powder duster for plants. Scientific American, 81(26):404, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
404a.pdf.

Anonymous:1899:EHW

[10795] Anonymous. An electric hose wagon. Scientific American, 81(26): 404–405, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
404b.pdf.

Anonymous:1899:LAE

[10796] Anonymous. The late Alaska Earthquake. Scientific American, 81 (26):405–406, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
405.pdf.

Anonymous:1899:RWC

[10797] Anonymous. Railways of the world compared. Scientific American, 81(26):406–407, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
406.pdf.

Reynolds:1899:JH

[10798] Cuyler Reynolds. Joseph Henry. Scientific American, 81(26): 408–410, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
408.pdf.

Anonymous:1899:NREc

[10799] Anonymous. A new rotary engine. Scientific American, 81(26): 411, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
411.pdf.

Anonymous:1899:SOSb

[10800] Anonymous. Some ostrich stories. Scientific American, 81(26): 411, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
411a.pdf.

Anonymous:1899:BPcb

[10801] Anonymous. Business and personal. Scientific American, 81 (26):412, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
412b.pdf.

Anonymous:1899:KW

[10802] Anonymous. Key West. Scientific American, 81(26):412, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v81/n26/pdf/scientificamerican12231899-412a.pdf.

Anonymous:1899:NQby

[10803] Anonymous. Notes and queries. Scientific American, 81(26): 412, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
412c.pdf.

Anonymous:1899:RPIby

[10804] Anonymous. Recently patented inventions. Scientific American, 81(26):412, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
412.pdf.

Anonymous:1899:IIbz

[10805] Anonymous. Index of inventions. Scientific American, 81(26): 413–415, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
413.pdf.

Anonymous:1899:PNPbx

[10806] Anonymous. Patent notes: Patents. Scientific American, 81 (26):415, December 23, 1899. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/ REFERENCES 1803

scientificamerican/journal/v81/n26/pdf/scientificamerican12231899-
415.pdf.